A Bibliography of Publications in Fishery Bulletin

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

12 October 2024 Version 1.12

Title word cross-reference

14 [4618, 2493, 2570, 5159]. α [2344, 2441]. b [3939]. μ [3851]. ω [3032]. t [5193].

-Tocopherol [2344, 2441].

0-Age [2993]. 0.202-mm [5260]. 0.333-mm [5260]. 0ctober [1993].

1 [294, 2294, 3754, 437]. 10-foot [2056]. 101-150 [1705]. 12S [4392]. 12S/16S [4392]. 15 [294]. 151-200 [1708]. 16S [4392, 4433]. 1872 [654]. 1880 [138]. 1882 [211, 317]. 1883 [422, 556, 251, 271, 403, 437, 296, 248, 231]. 1884 [631, 403, 662, 597, 855, 654]. 1885 [626, 676, 589, 605, 637, 701, 738, 755, 772, 789]. 1886 [781, 814, 823, 830, 841, 846, 769, 783, 853]. 1887 [855, 879, 893, 908, 916]. 1888 [1012]. 1892 [1012]. 1896 [1128]. 1897 [1128]. 1898 [1158]. 1899 [1170]. 1960s [3728]. 1963 [1960]. 1964 [1993]. 1969 [2094, 2260]. 1970 [2178]. 1975 [2763]. 1979 [3015]. 1981 [3057]. 1990 [3766, 3788]. 1991 [3965]. 19th [2313].

2 [4173, 3750]. 20-micron [2359]. 201-250 [1711]. 20th [2313]. 21 [3872]. 226 [3872]. 24-hour [3143]. 251-300 [1714]. 26^{o}00’N [3473].

3-meter [2306]. 3/ND [4392]. 301-350 [1717]. 351-400 [1719].

4 [4392]. 401-450 [1701]. 4500-year [4906]. 451-500 [1703].

501-545 [1706].

6- [2056].

7 [533].

A-bomb [4618]. A. [3388, 3594, 3248, 3740, 2247, 4017]. Aaröe [662]. Abaco [827]. Abalone [3489, 2126, 3527, 3693, 3510, 4117, 4801, 4167, 4877, 3970]. Abalones [2341, 3510]. Abashiri [4623]. Abbott [453]. abei [3670]. Abernathy [2384]. Ability [2286, 260, 5256]. Abnormal [3572, 307]. Abnormalities [2078]. abnormalties [2699]. Aboard [3205]. above [3473, 324]. abrogation [631]. Abrolhos [4374]. absence [5284]. absorbing [1471]. absorption [185]. Abundance [2044, 2191, 3007, 3078, 3759, 2568, 2417, 3444, 3447, 4864, 4847, 3464, 4936, 5202, 3065, 4807, 3385, 4390, 2847, 2530, 3420, 3441, 3185, 3315, 3280, 4994, 3549, 1859, 4619, 2128, 3787, 3352, 2010, 4620, 2316, 4980, 3378, 3570, 2860, 2546, 2571, 2574, 5315, 2763, 1861, 2238, 2256, 5311, 3125, 3407, 3409, 4713, 2958, 2355, 3359, 2718, 3243, 3246, 4589, 1993, 2169, 4155, 2429, 3597, 3630, 2663, 2603, 4825, 2260, 2059, 2262, 3076, 623, 3445, 4793, 2313, 2141, 3657, 3682, 5192, 3166, 3268, 2250, 2443, 2447, 1848, 3495, 2095, 3275, 2083, 5002, 1759, 1767, 1836, 3797, 4100, 2656, 5119, 3923]. abundance [1662, 4848, 1988, 5053, 906, 3711, 4669, 4460, 412, 723, 4896, 5170, 5136, 3542, 4404, 3882, 3880, 5107, 4728, 3744, 4480, 4863, 143, 4149, 3679, 3921, 3983, 4908, 1731, 2702, 4366, 4167, 3200, 4960, 4995, 4464, 4607, 3951, 3178, 4975, 4113, 5367, 3851, 5056, 3829, 4066, 5327, 1734, 1779, 1845, 3794, 4233, 4820, 5138, 4096, 2593, 4395, 5175, 1668, 4906, 4973, 5042, 4154, 4917, 4568, 4548, 2538, 3875, 3901, 4058, 5374, 3903, 1987, 3918, 2274, 3730, 3161, 3286, 5230, 4957, 4742, 4598, 4201, 5033, 1785, 4037, 5008, 1218, 1860, 4263, 1666, 3805, 3718, 1808, 1838, 5000]. abundance [4662, 1760, 1971, 3233, 5234, 5318, 1964, 4492, 4302, 5251, 3836, 5134, 2002, 1843, 4986, 4378, 4752, 4228, 5169]. abundances [4028, 3937, 3869, 5040]. abundant [4987, 4800, 122]. Acanthaster [2268]. acanthias [139, 5199, 2579, 5337, 4455, 4569, 5227, 5126, 4969]. Acanthistius [5151].

Acanthochromis [4736]. Acanthocybium [1969, 5147]. Acanthopagrus [4375]. Acanthurus [4135]. acanthus [1874]. Acartia [2761]. Accelerated [1801]. acceleration [5371]. accelerometry [5297]. acclimatiaztion [1586]. Acclimation [2372, 2353, 2544]. Acclimatization [504, 1395, 505, 730, 386]. acclimatize [1094]. acclimatized [640]. accommodates [4474]. accomplish [472]. accomplished [159]. Account [62, 150, 134, 1052, 946, 4819, 941, 973, 974, 5261, 1384, 952, 265, 154]. Accumulation [1879, 3443, 5081, 1796]. Accuracy [4978, 4246, 4005, 4393, 4735, 4751]. acellatus [3866, 3698, 3607, 3909]. achievements [1379]. achieving [5299]. Achirus [2078]. Achoerodus [4920]. Acid [2058, 5067, 2764, 3934, 5006, 4931]. acidification [5339]. Acids [2413, 5174]. Acipenser [2413, 5207, 684, 2983, 4057, 2959, 5305, 4994, 5258, 5351, 5315, 5306, 3914, 4180, 4376, 4320]. Acknowledgment [4437, 4519]. Acoustic [2914, 2310, 2430, 5135, 2780, 4990, 4149, 4968, 5059, 5371, 5056, 5161, 5368, 4156, 5055, 5316, 4238, 5040]. acoustic-trawl [5059]. acoustically [3772]. acoustically-tagged [3772]. acoustics [5206]. Acquisition [2273, 2507]. acronotus [4473]. Across [2391, 5044, 4934, 5279, 4736, 3226]. Act [210, 700, 875, 425, 865, 453]. Actideus [2290]. actinians [1200]. Actiniaria [2518]. Action [2375, 4049, 259]. active [4888]. Activities [2914, 2901, 2944, 3962, 3595, 5142, 3877, 3220, 2697]. Activity [2342, 2423, 2509, 2148, 2199, 2060, 4862, 2180, 2427, 3628, 4123, 4761, 4995, 4064, 3384, 2278, 4420, 3291, 4915, 3718]. aculeatus [1823, 2499]. Acute [3343, 2357, 2900, 4780]. acutifrons [2655]. acutorostrata [3973]. acutus [4946, 2659, 2797]. Adaptation [1648, 1482]. Adaptations [2223, 3602, 4432]. adapted [4816]. Adaptive [2383, 5254, 4582, 1492]. Addenda [1268, 1542]. addition [4456]. Additional [2014, 2347, 2067, 2799, 14, 2348, 4549, 3345]. Additions [2008]. additive [5124]. Address [1049, 1023]. addressed [162, 271]. addresses [248, 264]. adductor [1911]. adetii [4025]. adina [3433]. Adjacent [2805, 3617, 2438, 4784, 2300, 3910, 1502, 2653, 1970, 4022, 4973, 4548, 1904, 4512, 4065, 3905, 3767, 3955]. Adjustment [4419]. adjustments [3345]. Administrators [2328]. Adriatic [2356]. adspersus [4123, 2222, 2536, 2427, 2727, 3067, 3743]. Adult [3694, 2455, 3017, 2437, 2819, 3058, 3461, 3591, 2384, 2118, 5118, 4784, 4441, 5079, 2489, 5249, 1524, 5290, 4273, 2810, 5367, 1966, 1847, 4162, 5374, 3814, 3784, 4352, 4886, 2398, 5273, 4793, 5211, 2889, 3889, 3268, 2043, 1986, 5130, 4912, 3720, 1873, 3456]. adulthood [4334]. Adults [2516, 3260, 3310, 4068, 4041]. Advances [4926]. Advantages [4890]. advection [3485]. adventure [371]. adverse [1619]. Advice [2236, 3651, 3751, 3749, 3951]. advisability [271]. Aegean [3716]. aeglefinus [4388, 3093, 3435, 5073, 3825, 4866, 3810, 5041, 4237, 4269, 1652]. aequispina [3283, 2969, 3302]. aequispinus [5158, 5330, 4205]. Aerial [2749, 3280, 3959, 3921, 2996, 2719, 4091, 2482, 3679, 3844, 3100, 3994, 3861, 5318, 3447]. aestivalis [3224, 2923, 3740]. Aetideidae [1776]. Aetobatus [5054]. affect [5091, 4607, 4725]. Affecting [3216, 3588, 3403, 3076, 3657, 1918, 4748, 1657, 4828, 653, 2557, 516, 1729]. affects [2849]. affinis [2403, 2099, 1472, 3390]. Affinities [2890, 3669]. affinity [5130]. affinus [2891]. Africa [2979, 3569, 5005, 3593, 2187, 4505, 3924, 4612, 4299, 4837, 3995, 4224]. African [2657, 4077, 4525]. After [2852, 3381, 4719, 5016, 3779, 5080, 4749, 3846, 4130, 5235, 5077, 1801, 5069, 3986, 4556, 4951]. again [3363]. against [287, 1422, 4897]. Agassiz [5075]. agassizi [3229]. Age [3708, 3724, 5255, 5301, 3853, 3629, 3035, 3062, 3480, 3501, 3644, 3856, 3183, 4032, 5005, 3091, 3358, 3729, 4412, 5120, 3414, 4642, 5106, 3593, 3486, 5115, 4523, 3979, 4040, 4194, 4461, 2377, 4849, 4213, 2986, 4141, 3044, 3399, 5155, 4463, 4159, 2776, 3648, 3864, 4711, 4230, 4253, 4998, 5278, 1458, 4124, 2279, 3352, 4232, 4454, 4482, 1644, 2993, 4287, 4370, 4151, 3573, 4197, 5325, 1688, 2120, 3042, 3173, 3257, 4168, 4905, 1827, 4694, 3559, 3654, 5353, 5154, 4372, 3405, 3356, 4608, 2271, 2956, 3760, 4170, 2975, 3987, 2220, 5159, 4446, 4898, 3604, 4332, 2958, 3475, 3119]. Age [5219, 3649, 3936, 4512, 4549, 4139, 5201, 3096, 4612, 3235, 3304, 3548, 4985, 4448, 3687, 5084, 1669, 4405, 5052, 3678, 4237, 3614, 2511, 3071, 3296, 4336, 4177, 5160, 4112, 3146, 3375, 3742, 4633, 3826, 3633, 4879, 2560, 4591, 3680, 4556, 4901, 4011, 4320, 2435, 3592, 4573, 4699, 4435, 4534, 3642, 4355, 4240, 4241, 4083, 5282, 2893, 4700, 4969, 3102, 2896, 3571, 3223, 3417, 2428, 4571, 5076, 2537, 4451, 5204, 3995, 4224, 4471, 4425, 5004, 4656, 1789, 3134, 4145, 3953, 5334, 1725, 3837, 5016, 3895, 3898, 1636, 4138, 4863, 5152, 4734, 5061, 5279, 4286, 4075]. age [2784, 4816, 3704, 4526, 3825, 3577, 5013, 3168, 5216, 3739, 4006, 5337, 1867, 4565, 1857, 1842, 3872, 1861, 5195, 4777, 3849, 3807, 4198, 5345, 4200, 3896, 5370, 4292, 3930, 5077, 4431, 4537, 4510, 1693, 1672, 4486, 3903, 5295, 1820, 4399, 4631, 4104, 4359, 4488, 3301, 4005, 4698, 5034, 5111, 4577, 4375, 3083, 5330, 4263, 4915, 4680, 1943, 5064, 3671, 3928, 4266, 5205, 3828, 4814, 4267, 3966, 3443, 4636, 4667, 4746, 1564, 1610, 4378, 4703, 4323, 4271, 5369]. Age- [4901, 4286, 4271]. Age-0 [3642, 3953, 3704, 3903, 4915, 3671, 3928, 4266, 5205, 4378, 4703]. age-1 [3704]. age-based [5013, 5295, 4746]. age-composition [4777]. age-dependence [5061]. age-dependent [5013]. Age-Determination [3654]. Age-Frequency [3035]. age-reading [4698]. Age-specific [3414, 4898, 4656]. age-structured [5345, 4263]. Age-validation [4556]. Age/Size [3102]. Ageing [3405, 3933, 3514, 3693, 3947, 4232, 3499, 3925]. agents [2610, 1959]. Ages [1904, 3025, 3815, 4162, 4467, 5335]. aggregate [3003]. Aggregating [3615, 5166, 4611, 5185]. Aggregation [2730, 3506, 2630, 4357, 3803, 4526, 4649]. Aggregations [2276, 4719, 4245]. Aging [4286, 2658, 4813, 5112, 5346, 5122, 5181, 3367, 3215, 4735, 4449, 5356]. Agnatha [3968]. ago [265, 22]. Agonidae [3817, 4115]. Agonostomus [1784]. agreement [4459]. Agriculture [271]. Aid [2075, 298]. Aided [2976]. Ain [880]. Air [2441, 3344, 3932, 1514]. Airborne [2141, 4330, 3805]. Aires [4161]. Alabama [28, 2804, 954, 4975, 5191, 4448, 3107, 1099, 233]. alabamae [5191]. alalunga [4829, 3800, 4250, 3831, 2529, 3401, 3849, 2496, 2593, 2554, 2565, 2909, 1987, 4061, 4142, 3265, 3565, 2567, 4457, 2005, 2083]. Alamagan [3322]. Alamitos [4268]. Alan [3626]. aland [7]. alascanus [4978, 4687, 4480, 4417]. Alaska [448, 4783, 3228, 3708, 2417, 2516, 3756, 4084, 1662, 2270, 2245, 5348, 3273, 3871, 4706, 3874, 3977, 1998, 4708, 2631, 2474, 3953, 4964, 5372, 2869, 3080, 5110, 5280, 2506, 5136, 1641, 2380, 4404, 3956, 2105, 5010, 4480, 1332, 4827, 4977, 1311, 2894, 4304, 1524, 1590, 1572, 1823, 3611, 4582, 1852, 2626, 2646, 2709, 2775, 2886, 1945, 2677, 2134, 4892, 5114, 2712, 3851, 2666, 3254, 4130, 1629, 2114, 3380, 3579, 3893, 5195, 3586, 5138, 4153, 4445, 5100, 2910, 4898, 4954, 5168, 4293, 4823, 5042, 4095, 1007, 3359, 4208, 2550, 1158, 1220, 4511, 3436, 3722, 1762, 1830, 4105, 3814, 4033]. Alaska [1480, 2581, 2259, 3143, 2197, 2822, 2856, 2927, 4742, 2286, 3306, 4950, 2865, 1578, 1583, 1593, 1605, 1612, 1631, 1634, 1627, 1302, 4297, 2262, 4360, 1603, 1623, 1624, 1640, 1804, 3149, 1162, 5137, 4489, 3410, 5330, 3445, 4263, 4401, 2911, 3075, 5051, 4853, 4901, 4340, 4874, 4976, 5347, 4341, 1943, 4434, 939, 963, 4814, 3949, 1607, 4665, 4969, 4838, 2428, 5071, 3574, 5361, 5025, 4378, 5021, 4386, 4551, 4752, 4248, 4654, 4664, 2500, 4102, 5229]. Alaskan [4442, 3787, 4835, 4980, 2708, 4818, 4174, 3901, 4398, 4902, 4951, 4017, 4407]. albacares [4558, 5212, 3397, 4391, 4605, 2383, 3262, 3463, 4367, 4245, 4595, 5373, 2281, 2682, 4839, 2276, 3484, 3265, 4495, 1894, 3663, 4007, 4338, 3841, 2431, 4011, 4767, 2896, 2920, 4116, 4817, 5083]. Albacore [2092, 2112, 2525, 3401, 2496, 2554, 2909, 2229, 3565, 2083, 4829, 3800, 4250, 3831, 2529, 1887, 3681, 2596, 3849, 2593, 2565, 1820, 1818, 1877, 1987, 1982, 2388, 4061, 4142, 3265, 2567, 4457, 2005]. Albatross [1109, 422, 562, 1292, 1251, 1295, 1158, 1291, 626, 597, 1012, 1289, 600, 1242, 832, 879, 939, 963, 1297]. albatrossi [2361]. Albemarle [997, 5102, 75]. albidus [2684, 4693, 4791, 4738, 4743, 2551, 217]. albigutta [3976, 3479]. albiguttus [1492]. Albula [3048, 4040, 4109, 4195, 3978, 1974]. alcohol [4749]. alcyonaria [1217]. Aleknagik [2262]. Alepisaurus [4495]. Alepocephalidae [2616]. Alepocephaloidea [3670]. Aleutian [4936, 5138, 4172, 4823, 5330, 5359, 4305, 4306]. aleutianus [5288, 4732, 4233, 4881, 4401]. aleuticus [2438]. Alevins [2852]. Alewife [2923, 4216, 2906, 3025, 3215]. alewives [46, 543, 110, 1855]. alfonsino [3911]. alfredi [5300]. alga [2754]. algae [1750, 1502, 1388, 1729, 1250]. Algal [2380, 3491]. Algarve [4900]. Algeria [907]. Algiers [382]. algorithm [4710]. alimentary [1452, 1471]. alkaline [466, 734]. Alkalinity [1555]. Alleghany [941]. allelochemics [2291]. Allen [3335]. allied [1246, 1426]. Allies [2888]. Alligator [3466]. Allocation [3356, 5047]. Allocyttus [3451]. Allometric [2183, 4859, 3942, 4923]. Allometry [3821, 3997, 3897]. Allozyme [4605, 4042, 3347, 4055, 4530, 4035, 4116, 4365]. allozymes [4732, 4125]. Along [2128, 3041, 2367, 2211, 2418, 3205, 2724, 2923, 2836, 2603, 1782, 1813, 4100, 4798, 4453, 4708, 3437, 5095, 4111, 2314, 3748, 3775, 5027, 4418, 5125, 2947, 4446, 5198, 4714, 2958, 4430, 4155, 5030, 2251, 3106, 4872, 4805, 5343, 4514, 793, 5139, 4380]. Alopias [5152, 2940, 5331, 4170, 2444, 4639]. Alosa [4516, 2401, 2460, 2437, 3224, 5122, 1817, 3060, 4216, 2906, 3025, 3215, 5191, 37, 2724, 2923, 5030, 5108, 89, 3459, 1855]. alosiodes [1828]. alosoides [2228]. Alpena [785, 892, 213]. already [458]. Altamaha [5315]. Altered [2082, 2443, 3064]. alternating [4406]. alternative [5276, 4353, 3828, 4969, 3970, 3998]. Alternatives [2793]. altimetry [4222]. altivelis [4687, 4978, 4417, 4483, 3034]. altus [1876]. alutus [4936, 2566, 5138, 3653, 3395]. Always [2811]. amaena [3900]. amber [30]. amber-fish [30]. amberjack [4196, 5371, 4241, 4667]. ambient [4584, 5127]. Amblyraja [4699, 4733]. Amchitka [2270, 2380]. ameba [1580]. America [2317, 1437, 1056, 5153, 2615, 258, 913, 1008, 1233, 194, 1078, 1565, 1574, 1958, 5027, 2598, 167, 2596, 3077, 2947, 914, 1689, 931, 5030, 3316, 1328, 3395, 351, 1676, 1661, 1678, 3889, 3631, 436]. American [3527, 2111, 4329, 4850, 2366, 1052, 1402, 361, 347, 415, 4798, 3474, 3505, 1389, 3374, 4516, 874, 3765, 5375, 898, 4772, 2401, 2460, 2437, 5253, 4479, 4580, 4135, 2183, 2309, 896, 903, 1235, 1619, 5122, 2949, 3140, 4111, 2390, 2563, 663, 2779, 2841, 1414, 2277, 145, 1690, 1765, 779, 4933, 481, 17, 70, 3168, 367, 1036, 1103, 1430, 1817, 829, 3225, 3060, 571, 940, 1857, 4866, 4184, 2339, 2211, 2418, 2593, 3656, 184, 2173, 86, 360, 232, 4834, 2718, 2385, 2501, 298, 2724, 1063, 1081, 4000, 1987, 2711, 5317, 5030, 1653, 2070, 732, 4488, 868, 640, 629]. American [683, 1841, 3047, 1788, 1805, 1882, 197, 4469, 3459, 1692, 1124, 1039, 1071, 3400, 3291, 1035, 1812, 609, 1395, 3256, 5286, 2897, 2422, 848, 147, 1407, 332, 704]. americana [2998, 3500, 2315]. americanus [4860, 3724, 4100, 5301, 3474, 3505, 3590, 4182, 4070, 3765, 2764, 3603, 4772, 5106, 5253, 4479, 4580, 3157, 4990, 2949, 3141, 3172, 3140, 4111, 2779, 2841, 2461, 1778, 3104, 3789, 4184, 2339, 2211, 2418, 2528, 2652, 2881, 2580, 3656, 2220, 5159, 3384, 2718, 3594, 2711, 3281, 2115, 3047, 4469, 2381, 677, 3400, 3891, 4300, 5356, 5205, 3256, 5286, 2897, 3839]. Amherst [122]. Amiurus [217]. Ammodytes [5187, 3388, 3890, 2718, 3594, 5128, 2711, 3281, 2224]. Ammonia [2815, 2357, 2865, 2852, 2653]. ammoniacal [687]. Ammonium [2037]. Among [3184, 2942, 2952, 2076, 2632, 3008, 3642, 4052, 4002, 3733, 4750, 4048, 2490, 4645, 2049, 4823, 4867, 4025, 5374, 1882, 4375, 4955, 84, 4703, 5259]. Amount [2297, 344, 2186, 4803, 313]. Ampelisca [3229]. amphibians [1586, 1563]. Amphiodon [1828]. Amphipod [2491]. amphipoda [1300]. Amphipods [3126, 3229, 3239, 3307]. Amphiprion [2943]. amplified [4002, 4125, 4815]. ampullatus [4416]. Amsterdam [829]. amusement [881]. Anadromous [3621, 4927, 5281, 5372, 1694, 3412, 2906, 1979, 3153, 1855, 4599]. Anadromy [1805, 3778]. Anaerobic [2313]. anal [4534, 3966]. anal-fin [4534, 3966]. analis [4719, 5028, 3974, 4955]. Analog [1950]. analoga [2643]. Analogue [2158]. analyis [3746]. Analyses [2240, 3570, 3050, 3425, 5303, 4501, 4576, 3168, 4178, 3799, 3103, 5277, 5286]. Analysis [3032, 2892, 3834, 4935, 2633, 4121, 3038, 3404, 3555, 3567, 3573, 2580, 2343, 4398, 5026, 2453, 2560, 2679, 3696, 2683, 319, 3625, 2239, 3027, 1795, 2396, 4558, 4081, 4452, 3972, 569, 4249, 5255, 4120, 4934, 4552, 4190, 4002, 4021, 4211, 5067, 5271, 4106, 5050, 5113, 3840, 3935, 4999, 5091, 4929, 4787, 3915, 3943, 3679, 4392, 4442, 4231, 3748, 4992, 2479, 5288, 4544, 1977, 4232, 4991, 1817, 2834, 4289, 4418, 3906, 5154, 4291, 4085, 4785, 4766, 4373, 4445, 5193, 4055, 1948, 5117, 5276, 4396, 4317, 4335, 4985, 4400, 3423, 3942, 4768, 5354, 2991, 3265, 3774, 4872, 3841, 3778, 4822]. analysis [4318, 4298, 4013, 4842, 4457, 4592, 4969, 4815, 5247, 4718, 4272, 4045, 4599, 4471]. analytic [5014]. Analytical [3715, 3564, 2502, 4261, 4599]. Analyze [2815, 3504]. Analyzed [3032, 3105]. Analyzing [2793, 3071, 5208]. Anarhichas [5132]. Anarrhichthys [3006]. Anatomical [3303]. Anatomy [1922, 1431, 1452, 2781, 1953, 2762, 119, 1490]. Anchoa [4028, 2645, 4626, 4260, 3511]. anchoita [5231, 5231]. Anchor [3673, 4131, 4276]. anchoveta [4368, 4637]. Anchovia [1463]. anchovies [4002]. Anchovy [3691, 3694, 2469, 3536, 3346, 3517, 2673, 3324, 2356, 3521, 2153, 2449, 2521, 2759, 2907, 3187, 3196, 2493, 2885, 2840, 2397, 2062, 3188, 3266, 2132, 3310, 2733, 3119, 2164, 2833, 3525, 3235, 2870, 2476, 2575, 3600, 2175, 2644, 3589, 2858, 3538, 2439, 3511, 2338, 2756, 2809, 3523, 3414, 2124, 4028, 2645, 2147, 2842, 3919, 4626, 4330, 4294, 2936, 2818, 3304, 3305, 3796, 2861, 2260, 4260, 2587, 3531, 4209, 3716, 3369, 2348, 2902, 2943, 5242]. ancient [1807]. Ancylopsetta [1482]. Andrew [2429, 2548, 2534]. Andrews [1050]. angel [5237, 5328]. angle [4953]. angler [4832]. Anglerfish [5093, 3270, 2925, 5197, 4422]. Anglerfishes [2812, 2800, 2829]. anglers [4052, 4462]. angling [1485, 1362, 3303]. Angola [1676]. Anguilla [88, 5140, 290, 122, 4000, 3090, 4488, 3291]. angulata [175]. angulatus [4015]. angustimana [2775]. angustirostris [2926, 3994]. animal [1545, 1747, 1686]. Animals [2177, 2198, 2561, 1546, 1042, 946, 1684, 4980, 4184, 86, 1726, 1024, 1309, 256, 1338, 3245, 1094, 1790, 4271]. Ann [2625]. Annelids [1748, 2635, 1205, 1297]. Annotated [1810, 1531, 1584, 1841, 1908, 1695, 1781]. Announcement [4056, 4277, 4362, 2081]. Annual [2191, 4143, 4189, 4640, 3868, 422, 3727, 1614, 2562, 3646, 2823, 2825, 3668, 4818, 1940, 4176, 4397, 3820, 3226, 3306, 3123, 3628, 4249, 2317, 2772, 4311, 4843, 4846, 3842, 4563, 206, 4085, 4068, 2947, 2251, 3106, 4354, 4891, 4678, 5021]. annular [4000]. annulus [3996]. anomalies [2678]. anomalous [4956]. Anomura [3868, 3144, 3798, 2201, 2216, 2688, 2533]. anomuran [1216]. anonychus [4682]. Anoplopoma [3195, 5058, 3276, 3331, 3635, 2517, 4232, 4417, 3872, 3804, 4169, 4088, 4898, 4631, 5160, 3445, 4263, 5211, 5051, 3061]. anoxic [2245]. answer [287, 303]. Answers [129, 733, 299]. Antarctic [5133, 3677, 3004, 3422, 3701, 3614]. antarctica [4018]. antarcticus [4225]. antennarius [2060, 2339]. antennatus [3887]. Antennular [2932]. anthonyi [2060, 3665]. Anthozoa [1451, 3290]. antibodies [4668]. Antillean [2052]. Antilles [2541, 2395]. Antipredation [2750]. antiquorum [103]. antisera [1968]. Anton [540]. Antrim [949]. any [425, 453]. anything [256]. Aorta [3397]. Apalachicola [5351, 2467, 3855]. Apeltes [2]. apex [4957, 3891]. Aphanopus [2762]. Apo [4784]. Apodichthys [4796]. Apostle [1903, 2265]. Apparatus [1393, 2049, 302, 262, 463, 282]. Apparent [3609, 2214, 3683, 1668, 2711, 2141, 2447, 2083, 3853, 1987, 1964]. apparently [538, 452]. appear [388, 533]. Appearance [139, 3953, 206, 4, 307, 230, 284, 585]. Appendage [3410, 4855]. appliances [1091, 1109]. Application [2167, 3483, 4617, 4792, 5059, 4693, 3266, 2999, 3319, 3261, 4222, 3653, 4078, 2514, 3778, 3625, 2253, 2439, 4081, 4281, 1052, 4623, 4560, 2997, 3163, 4919, 4889, 4945, 4544, 3825, 4483, 4373, 4444, 4234, 4982, 4223, 4758, 4550, 4043, 1459, 1761, 4926, 3998, 4307]. Applications [4582, 3774]. Applied [2507, 4145, 5157, 4861, 4663]. applying [5131]. appointed [264, 329]. Appraisal [3513]. Approach [3431, 3319, 2745, 3751, 4348, 4958, 4769, 5340, 639, 4712, 4674, 4692, 3072, 3206, 4550, 5014, 3774, 4891, 3720, 3998, 4146]. approaches [4723, 4053, 3982]. Approaching [2980]. appropriated [313]. Approximate [2238]. Approximation [2354]. Approximations [3944]. April [2178, 792, 877]. Aprion [3552, 4354]. Apristurus [3470]. Aptocyclus [5252]. Aquaria [979, 440, 2409, 1403]. Aquarium [1031, 929, 1043, 848, 1025, 829, 266]. aquatic [1657, 999, 1518, 1338, 218]. aquiculture [1418]. aquosus [3606, 3980]. Arabian [2887, 4757, 5364]. Aragon [888]. araneus [2992]. Arc [3556]. archaeology [4906]. Archeological [4073]. arches [4537]. Archipelago [5054, 1320, 4172, 4612, 4034, 1245, 1307, 1334, 5295, 4076, 2812, 2829, 3322, 3636, 3471, 3175, 3176, 1932]. archival [5305, 4945, 4691, 4693, 4583, 4584, 4621, 4684, 4531, 4851]. archived [3941]. Archosargus [3644, 2645, 3548, 3633, 4086, 5204]. Arctic [2332, 3149]. Arctica [2978, 2960, 3123, 3146, 3255, 2707]. arctifrons [2244]. Arcto [3132]. arcuata [2922]. Area [2417, 2524, 2352, 2265, 2942, 2462, 3243, 3961, 2303, 1947, 2373, 3381, 4707, 2853, 1903, 4231, 4821, 4930, 4620, 3212, 4288, 4316, 4448, 2552, 5316, 3802, 1423, 4636, 4305, 4306, 4380]. Areas [2082, 4226, 2655, 3336, 2704, 3244, 4677, 3295, 5055, 4572, 1893, 3964, 5268, 3907, 4865, 5224, 4986, 4840]. arel [3732]. Arenaeus [4460]. arenaria [3032, 4243, 2625, 2997, 3163, 905, 2544, 3318, 3260, 1184, 2322, 3191, 721, 1787, 1905]. arenarius [4916, 4917, 2935]. arenarla [3449]. Arenas [194]. Arendal [596]. areolatus [4888]. Argentina [4327, 4160, 4329, 3886, 3832, 4331, 4567, 1395]. Argentine [4309, 5133, 5151, 4725, 5088, 5183, 5221, 3583, 4702, 5231]. Argentinian [4797]. Argentinine [2008]. argentinus [4309, 5133, 4247]. Argonauta [3003]. Argopecten [3790, 3490]. arguinensis [5239]. Argulus [1310]. argus [4384, 1517, 2877, 4952, 3773, 4929, 3427, 3241, 3645, 3835, 3344]. Argyripnus [2247]. Argyrosomus [4077]. Argyrozona [4707]. ariakensis [4604]. arid [4460, 4943]. Ariidae [3296]. ARIMA [4824]. Aristeidae [3887]. Aristeus [3887]. Arithmetic [3151]. Arizona [567]. Arkansas [974, 1072, 1100]. Arm [5168]. armatus [3963]. armhook [4682]. armorhead [3454, 3827, 3764, 3571]. Armory [185]. Armstrong [531]. arrange [250]. Arrangement [464]. Arrival [3273, 243, 41]. arrowtooth [4553, 3248, 3938, 4901, 4017, 4102]. artedi [1644, 1597]. artesian [319]. artesian-well [319]. Arthropods [1751]. article [1139, 351]. articles [570, 856, 409]. Artificial [2293, 714, 3773, 2475, 64, 440, 708, 2075, 2015, 2289, 90, 1511, 162, 797, 753, 625, 258, 287, 539, 457, 1036, 1606, 959, 5175, 1346, 1435, 271, 4973, 3974, 4448, 1108, 398, 1135, 1107, 5064, 23, 75, 919, 220, 175]. artificially [684, 1027, 236]. artisanal [5333, 5328, 5164]. Ascelichthys [2981]. Ascension [3645, 3835]. Ascent [3958]. ascertain [370]. ascertained [4531]. ashes [87]. Asian [2111, 4398, 4715]. asking [631]. aspe [4176]. Aspects [3377, 3097, 2225, 2301, 2484, 5110, 3470, 3312, 3219, 2571, 2574, 3135, 4030, 4219, 2650, 2446, 2864, 2685, 3550, 3311, 2731, 3082, 2602, 2546, 2682, 1362, 3221, 4323]. asper [2438, 4118]. aspera [5355, 3109, 4397, 5241]. aspersa [390]. aspersus [442]. aspidura [5282]. assayed [4335]. assays [4887]. Assemblage [3007, 3126, 3246, 4322, 3259, 4121, 4975, 5012, 5062]. Assemblages [3309, 2823, 3131, 3232, 4084, 5170, 2825, 5017, 5223, 4316, 4803, 4695, 4973, 4611, 4374, 4872, 4894, 4759, 4773, 5224, 3894, 4922, 4051]. assess [4829, 4459, 4983, 4735, 5011, 4107]. assessed [3949]. Assessing [3590, 3941, 4968, 4466, 5172, 3275, 4541, 4604, 5318]. Assessment [3394, 5349, 3201, 3280, 4930, 2462, 3401, 4943, 2430, 2721, 3651, 2736, 3319, 4597, 3066, 4420, 3625, 2253, 3371, 4164, 5303, 4686, 4873, 4106, 4602, 4808, 4769, 5061, 4121, 4673, 4757, 4507, 5013, 3919, 4979, 4565, 5195, 5345, 5186, 3971, 3072, 4234, 4547, 3750, 4529, 4101, 4223, 4758, 5089, 3499, 3754, 4078, 3838, 4902, 4986, 5147]. assessmentand [4744]. Assessments [3249, 4829, 5214, 4366, 4215, 5014]. Assignment [3924]. Assimilation [2181, 2931, 2068, 1029]. Associated [2466, 3126, 2478, 3615, 2262, 3553, 3639, 3448, 4883, 3977, 3711, 4191, 4903, 5153, 4026, 4470, 3771, 5176, 5223, 5262, 2448, 5166, 4973, 4855, 3136, 4294, 3974, 4611, 5269, 3865, 5156, 5248, 3843, 2699, 4204, 1760, 1951, 5145, 5062, 4667, 5185, 4654, 4840]. Associates [3689]. Association [898, 4391, 1958, 4293, 3723, 3003, 2593, 4383, 298, 704]. Associations [3650, 2705, 3428, 3079, 4833, 4916, 4624, 3160, 4480, 5086, 3491, 5137, 3747, 4206, 4356]. assumed [4829]. assumption [3838]. assumptions [3721, 4897]. Asterias [1654, 2011]. asteriscus [3898]. Asteroidea [2268]. Asymmetric [4992]. Asymmetry [2239]. at-sea [5322, 4246]. At-vessel [5269]. Atheresthes [4553, 3248, 3938, 4901, 3287, 4017, 4102]. Atherinid [2254]. Atherinidae [2662, 3085]. Atka [3996, 4869, 4868, 4172, 4095, 4818, 4663, 4305, 4306, 4664]. Atlantic [3198, 2296, 2684, 4120, 3629, 3062, 2892, 5310, 4191, 4641, 3372, 3160, 3806, 5237, 4994, 3596, 2314, 5249, 4854, 3650, 4731, 3465, 3352, 3300, 3113, 5337, 4528, 2364, 2203, 2367, 2946, 5331, 2477, 2214, 4738, 4173, 2343, 3812, 2796, 4589, 2960, 3936, 4486, 5274, 5272, 5302, 4139, 4104, 3090, 5156, 4122, 3209, 3114, 4591, 2824, 2013, 2806, 5357, 3018, 3079, 3371, 3545, 4228, 4137, 4655, 3563, 3975, 2413, 4242, 2371, 3379, 3947, 4343, 1782, 1784, 1813, 2346, 4425, 2915, 4829, 4249, 2778, 2807, 2950, 3853, 3910, 4120, 2900, 3062, 2092, 2862, 2654, 2349, 1822, 1853, 4601]. Atlantic [4453, 4686, 4346, 3093, 3435, 4032, 5352, 3402, 3091, 2903, 4477, 4024, 4211, 2983, 4963, 4106, 5050, 5174, 1886, 1876, 5079, 2130, 1725, 5133, 4478, 4830, 3840, 2859, 1840, 2722, 3001, 2637, 852, 3437, 3486, 5115, 2951, 3201, 4523, 3935, 2615, 4946, 3488, 3715, 5379, 3609, 2729, 3880, 2028, 5304, 2240, 4994, 2878, 2931, 3044, 3054, 3081, 3424, 5313, 2489, 4900, 5099, 5132, 2007, 2232, 5152, 2899, 2776, 4749, 1880, 3969, 4012, 3795, 3888, 5258, 5075, 5249, 4046, 4254, 4763, 4524, 4968, 4764, 1778, 2021, 1600, 4196, 4606, 2623, 2784, 3748, 3775, 2966, 4428, 2010, 4544, 3577]. Atlantic [4048, 4214, 6, 3954, 3129, 1896, 1574, 1638, 4942, 4693, 3455, 2571, 5200, 4947, 4288, 2592, 2944, 2199, 3199, 1842, 2029, 2163, 2994, 1075, 2367, 4791, 4866, 3957, 5353, 3383, 5373, 5154, 5215, 4645, 5006, 5311, 3029, 4291, 2436, 2690, 4098, 3760, 2272, 2281, 2321, 2103, 3059, 4546, 4200, 5186, 2975, 3810, 5041, 4966, 5159, 4235, 3581, 5104, 4554, 5306, 4093, 1851, 2129, 5263, 4154, 2867, 4199, 4074, 290, 4174, 1664, 5148, 5339, 2209, 4236, 4548, 1868, 4588, 37, 2792, 2796, 3661, 4144, 4984, 5097, 4677, 4136, 4295, 2539, 4317, 2724, 2923, 4895, 2093, 4139, 3594, 2711]. Atlantic [3703, 3651, 4513, 4886, 3282, 2482, 3045, 3295, 4768, 2798, 4296, 3583, 2085, 2308, 3320, 3786, 3301, 3678, 2205, 4614, 4789, 2334, 2244, 5309, 2929, 4122, 3712, 4419, 4178, 3826, 5335, 4063, 4180, 4376, 1661, 1678, 3290, 4805, 2264, 2208, 3094, 1863, 4013, 1002, 1076, 1089, 3802, 3389, 3745, 5343, 1946, 1884, 4810, 166, 221, 3682, 4679, 3255, 5178, 4320, 2015, 2444, 3740, 192, 1812, 3828, 2797, 2893, 3102, 3256, 5247, 5172, 3873, 4514, 2488, 2458, 2731, 5092, 2707, 2537, 3758, 3931, 2095, 3419, 3497, 4921, 2734, 2789, 5369]. atlantica [2089, 2762]. atlanticum [2762]. atlanticus [4425, 5116, 3876, 3979, 3791, 4125, 3958, 4119]. Atlanto [4081]. Atlas [1865, 1997]. Atmospheric [1048, 2404, 1989]. Atoll [2254, 2632, 3249, 4007]. Atolls [2728]. Atractoscion [4883, 5118, 3759, 4141, 3533]. atrarius [960]. Atresia [3187, 4818]. atretic [4008]. atricauda [4764]. Atrina [1431]. attachment [2519, 4416]. Attacks [3694]. attained [458]. attempt [684, 931]. attempts [1094]. attenuata [4640, 4640, 3234, 2811, 3837, 2838, 3821, 4761, 4367, 3222, 3274, 4259, 2473, 2590, 3879, 4097]. attractants [3241]. Attracting [2289]. Attraction [2015, 5349]. attributable [1963]. Attributes [2917, 5125]. attrita [2616]. attrition [4558]. atun [4525]. atypical [5189]. Atypopenaeus [3516]. audax [4089, 2485]. August [3788, 1993, 437, 855, 296, 654, 589, 814]. Augusta [1531]. Auke [2516, 3306]. auratus [3783, 3580, 599, 4221]. aurea [4327, 4331]. auricilla [5295, 3471]. auriculatus [5175, 4219]. aurita [3808]. aurolineatum [2958, 3209]. aurorubens [4060, 2808, 4287, 2577, 3843, 4137]. Australia [4018, 4190, 4799, 4036, 4090, 4251, 4347, 5087, 4212, 735, 3735, 3809, 3981, 4673, 481, 820, 4151, 3985, 4006, 4197, 4987, 3407, 4714, 3933, 3002, 4025, 4549, 3005, 4298, 4377, 4402, 4616, 4700, 386, 3970, 4603]. Australian [4503, 2895, 4657, 3929, 3542, 4674, 4482, 3791, 3544, 4570, 4246, 4298, 3485, 4701]. australis [3422]. Austria [1414]. Austrian [1407]. authorities [60]. authority [329]. autochthonous [3016]. Autolysis [2045]. Automated [5298, 3671, 4710]. automatic [281]. autonomous [5172]. autoradiography [3393]. Autoregressive [3589]. autospasy [2760]. autotrawl [4754]. Autumn [4166, 681, 161]. Auxis [4038, 1689, 1671]. Availability [2240, 3627, 4989, 5050, 5174, 5249, 2282, 4217, 5350, 5246, 4432]. available [4748, 2186]. average [4750]. Avi [4476]. Avoid [2069, 454]. Avoidance [2962, 5349, 2557, 3736]. awarded [68]. Awards [3626]. awoara [4815]. Ayres [2051, 4083, 2989, 4407]. ayresi [3037]. ayresii [5134]. ayu [4483]. Azores [4588]. aztecus [2468, 2000, 2843, 2772, 4643, 2181, 2032, 3057, 5009, 2283, 2000, 3156, 3339, 2685, 3230]. azurea [3534].

B. [3277, 3129, 3320, 3786, 5335, 4681]. B.C. [2438]. bachei [2307]. Back [3514, 3556, 3688, 4774, 1606, 4399, 4844, 2, 4532]. Back-Calculated [3688]. Back-Calculating [3514]. back-calculation [4774, 4399, 4844, 4532]. backscatter [4149, 5229]. Bacteria [1750, 1791, 89]. bacteria-free [1791]. Bacterial [2786, 3016, 1516]. bacteriology [3147]. bacterium [1684, 1259]. Bahama [2541, 4091, 2395]. Bahamas [3354, 1784, 4357, 3978, 3974, 5162, 3852, 4041]. Bahamian [3048]. Bahia [3645, 3835]. Bahja [3868]. Baily [2039]. bairdi [3932, 2249, 2886, 2559, 2492, 4174, 5082, 2599, 5069, 5137, 4265, 4340, 4902]. Bairdiella [1463, 2040, 2372, 2353, 2803]. Bait [3198, 905, 5196, 5284, 276, 1871, 4227, 3761, 722, 1833]. Bait-Caught [3198]. baited [5001, 5086]. Baitfish [3672]. Baja [3868, 2016, 2542, 1832, 4324, 3702, 4022, 2159, 3087, 4177, 2359, 3534]. Balaena [2875, 4223, 3197, 595]. Balaenoptera [3277, 3973, 2711, 3776, 3598, 3507]. balao [4587, 4587]. Baleen [3630]. Balenoptera [2681]. Balistes [5120, 5176, 3173, 327, 5368, 413]. Ballarat [730]. ballyhoo [4587]. Baltic [575, 579, 901, 298]. baltica [3355]. Banana [2719, 5121]. band [5272, 3106, 3226]. band-pair [5272]. Banded [2803, 3139, 3152, 2327]. Bank [4281, 4542, 5132, 4960, 4995, 4091, 4529, 4374, 3547, 3772, 3758, 905, 4009, 4388, 3093, 723, 3890, 4930, 3825, 3829, 4645, 4010, 4494, 2975, 3810, 5041, 2220, 3232, 4513, 3281, 1645, 1785, 1769, 1696, 3802, 1800, 4269, 1652, 5071, 3419]. banks [852, 4765, 542, 641, 685, 928, 626, 1699]. barba [3296]. Barbara [2423, 2509, 2823, 2825, 2638, 3617, 4316]. barcoding [5165, 5343]. Bare [1823, 1762, 1830]. Barff [574]. barge [3088]. Barkley [4310, 3489, 3148]. barnacle [1223, 4907, 3723]. Barnacles [1427, 2432]. Barotrauma [4026]. Barotrauma-associated [4026]. barracudina [2089]. barramundi [4379]. barred [4030, 4714, 4268]. Barrier [4799, 4348, 3864, 3981, 4736, 3407, 3875, 4025, 4402, 2318]. Barrrier [4628]. bars [1126]. bartramii [4034, 4958, 5155]. Based [2044, 2669, 3435, 3467, 3280, 2629, 2909, 2975, 3137, 3587, 3639, 3271, 5255, 4623, 4346, 4328, 4959, 4617, 4349, 4849, 4787, 3140, 4671, 4286, 5243, 4816, 4314, 4790, 5013, 3951, 4723, 3919, 5262, 5046, 5377, 3100, 5341, 5215, 4766, 5140, 5166, 5218, 5159, 5275, 5368, 991, 4175, 4529, 4295, 5295, 4352, 5128, 1987, 4513, 4399, 4768, 3982, 4925, 4550, 5033, 3925, 4806, 4318, 3799, 4891, 3905, 4203, 3667, 4810, 4011, 4861, 4902, 4951, 5356, 5045, 5323, 5171, 4746, 3830, 4471]. baseline [5101]. basic [4652]. Basin [1764, 636, 1078, 115, 1340, 2298, 1009, 1011, 1070, 1097, 1150, 1073, 1144, 1087, 1086, 3556, 3990]. basins [941, 973, 974]. Basis [2281, 4559, 3771, 1492]. basket [4913]. Bass [2654, 5102, 5103, 2988, 5099, 2279, 2193, 3327, 2493, 3511, 3975, 4856, 5044, 5296, 118, 2656, 4516, 5189, 4686, 2781, 3336, 456, 5222, 2965, 3440, 896, 3064, 5151, 3664, 4945, 116, 367, 4443, 3367, 2192, 5353, 2495, 4628, 3342, 4566, 4030, 1243, 4941, 5236, 3040, 1664, 338, 8, 18, 4886, 1038, 1647, 5230, 3946, 3778, 3925, 4822, 1137, 3799, 1185, 230, 131, 5203, 427, 59, 4701, 4268, 4181, 4574, 3332, 3308, 960, 75, 400, 592, 911, 5308, 606, 332, 436, 455]. bassensis [3929]. Basses [3582, 4887]. bastard [4612]. Batch [2917, 2759, 3498, 5061, 5003, 4161, 4331, 4818, 4397, 4007]. bated [543]. Bathochordaeus [2738]. bathybius [2890]. Bathylagidae [3036]. Bathymetric [3155, 2013]. Bathymetry [2663, 2665, 4118]. Bathynectes [2448]. Bathypterois [4413]. Bathyraja [4892]. Batrachus [754]. Battery [835, 670]. Bay [3228, 3007, 4052, 3708, 936, 3037, 3765, 2474, 4057, 4643, 2082, 3388, 4781, 3044, 1821, 3578, 4069, 2445, 2623, 2784, 2626, 2646, 2709, 2775, 1626, 2860, 3641, 3254, 1970, 3559, 2635, 3050, 3666, 3624, 4395, 2217, 2347, 2467, 5100, 1318, 3131, 50, 66, 3310, 2921, 3211, 4227, 3006, 3766, 3722, 4696, 2429, 2836, 2259, 2288, 3528, 1659, 5084, 2197, 2322, 4344, 4598, 4632, 2548, 3306, 2552, 1099, 3375, 4376, 2973, 1118, 1775, 1450, 1477, 3194, 250, 1933, 963, 2717, 1995, 2020, 2215, 3221, 4240, 2534, 3511, 4268, 4745, 3574, 5204, 429, 4747, 4028, 1608, 3787, 2645, 4626, 4260, 3490]. bay [4793, 4452, 4441, 3853, 3910, 3068, 3336, 5372, 3793, 13, 161, 2660, 1613, 4756, 3878, 4769, 4604, 4028, 1878, 3318, 1588, 2304, 3101, 3662, 4620, 1582, 1720, 1772, 4168, 4418, 93, 4565, 4626, 4371, 5006, 4876, 3059, 4445, 3987, 52, 4972, 1975, 2178, 4548, 3156, 713, 2864, 1904, 3973, 4512, 4895, 8, 895, 4597, 4924, 3723, 3220, 1593, 1612, 4260, 3191, 4715, 3946, 12, 470, 5291, 3778, 3925, 4822, 3598, 955, 2612, 3328, 3587, 4793, 3642, 207, 5234, 4342, 4574, 4575, 934, 334, 370, 527, 4922, 4654, 3743]. bay-wide [4793]. Bayes [4870]. Bayesian [5334, 4804, 5193, 5370, 5097, 5261, 4400, 5242, 4861, 3751]. Bays [3065, 2827, 3622, 3741, 3728, 4965, 5230, 4918]. be [410, 4182, 342, 68, 566, 358, 310, 293]. Beach [3543, 2504, 2836, 4460, 4617, 4436, 5234]. beached [2432]. beaches [2714]. Beagle [3832]. Beak [2982, 5155]. beaked [5092]. beaks [4454]. beam [2853, 889, 4803, 5229]. beam-trawl [889]. beam-trawling [889]. Bean [241, 3465]. Bear [359, 1503]. bearing [1592, 1036, 1482, 1225, 314, 952, 1621, 919]. Beaufort [1488, 1438, 1498, 1493, 4113, 1615, 1639, 1502, 1313, 4488, 1474, 3197]. becoming [773]. Bed [2558, 1936]. Bedford [327, 122]. beds [534, 719, 4227, 1099, 1115]. been [264]. Beeson [2159]. beetle [1522]. beetles [624, 1530]. before [3779, 453]. beginning [654]. begun [298]. Behavior [2466, 3474, 2342, 2631, 2216, 2966, 3422, 2358, 2153, 2452, 3584, 2921, 2718, 2542, 2777, 2180, 2427, 2586, 3776, 2572, 2160, 3618, 4300, 5149, 3344, 2415, 2734, 2789, 3197, 3314, 2303, 5187, 2838, 4871, 2602, 3785, 2876, 4367, 2555, 2147, 3789, 4290, 5297, 4586, 1482, 4383, 4823, 5261, 5317, 2107, 4432, 4971, 4531, 5359, 5217, 4915, 5000, 4912, 5232, 4986]. Behavior-dependent [5149]. Behavioral [3492, 5290, 4416, 2339, 4788, 3047, 3507, 3755, 2689, 2700, 2087]. Behaviors [3612, 5349, 5079, 4722]. Behr [152]. being [555, 640]. Belgium [184, 866, 259, 315]. Belize [4722, 3906]. belliceps [2079]. Belone [65]. belonging [1186, 428]. Beloniformes [4570]. Belostomidæ [801]. below [3332]. Belt [662]. beluga [4404, 4371]. benefit [4784, 4832]. benefits [4992, 4264]. Bengal [3723]. Benguela [3600, 4505, 4525]. Bennet [5213]. Bennett [2120]. Bentheogennema [2421]. Benthic [3229, 2836, 3084, 2020, 4765, 3993, 4911, 2597, 2997, 2489, 4284, 4939, 1891, 3307, 4184, 4200, 4469, 3843, 4841, 5217]. benthic-phase [4184]. benthid [3239]. benthopelagic [4218]. Benzene [2493, 2531, 2495, 2570]. Benzo [3612, 3085]. Bering [2316, 3481, 3999, 3377, 2726, 5053, 2875, 4896, 5010, 5355, 3253, 1442, 3109, 3635, 3704, 3373, 4369, 4770, 4892, 2104, 3561, 2799, 4118, 4176, 4397, 4932, 5141, 3087, 3084, 4352, 1427, 1012, 1634, 3905, 3867, 5359, 3302, 4015, 4853, 5268, 3033, 2428, 3287, 3955]. Berlin [17]. Bermuda [4521]. bermudezi [2477]. Berry [3999, 3354, 3148]. Berryteuthis [3999, 4682]. Bertalanffy [3032, 3035, 2782, 3804, 5193, 5047, 3338, 2485, 4379]. Bertalantfy [3706]. beryllina [1513]. Beryx [4998, 3911]. Best [900]. Better [3613]. Between [2346, 2618, 3513, 2503, 2082, 2331, 2134, 2148, 2364, 3654, 3401, 2339, 3624, 3029, 2397, 2355, 3386, 3674, 3439, 3610, 2836, 3425, 2055, 2098, 2155, 2170, 2237, 3033, 2491, 2419, 2982, 4916, 3003, 904, 4346, 4521, 4020, 5372, 3898, 3142, 1990, 5249, 4763, 4730, 4854, 4944, 3200, 3825, 2370, 2994, 4859, 2813, 4820, 3908, 2682, 2906, 5157, 2947, 4803, 1837, 1897, 3153, 3117, 1122, 3989, 3818, 4236, 2662, 4334, 5029, 4176, 4886, 5354, 3723, 3920, 4950, 3406, 3707, 5043, 1676, 5260, 4433, 5268, 3233, 4666, 3808, 5025, 4922, 3970, 4654, 4380]. between-haul [3818]. Between-Reader [3654]. between-sex [4176]. Bevan’ [2257]. Bias [3335, 3654, 2238, 3530, 4133, 5246, 3069, 4871, 4980, 4585, 5140, 3031]. Biases [3234]. Bibliography [1540, 1547, 1553, 1709, 1888, 1810, 1841, 1908, 1695, 1781]. bidens [5092]. biennial [5274]. Bigeye [2940, 3615, 5078, 4558, 1876, 5365, 5152, 4821, 4245, 1792, 5373, 4170, 4531, 2444, 4435, 4639]. Bight [3596, 3563, 3975, 4557, 3759, 4120, 5312, 4686, 2230, 4561, 5079, 3372, 3160, 3806, 5170, 4503, 3280, 5099, 2696, 3650, 3465, 2549, 3352, 4048, 2307, 3113, 2763, 2946, 2885, 2477, 3334, 4030, 4171, 4913, 3812, 2796, 4294, 2960, 4139, 3703, 4104, 2244, 5156, 3209, 3114, 3162, 2824, 3745, 3891, 5178, 4321, 3256, 5172, 3018, 3079, 3371, 4817, 3628, 4137, 5231]. Bights [4122]. Bikini [2816].

Biliary [1231]. bilinearis [4346, 3124, 3081, 2314, 3029, 796]. bilineata [4407]. bill [210]. Billfish [3407, 3939, 4231, 4992]. billfishermen [3545]. Billfishes [2892, 4524, 4942, 4509]. Bimini [3733]. Bioaccumulation [1938]. Biochemical [3913, 3603, 3109, 3582, 3674, 2055, 3039, 4421, 3480, 4103, 3712]. Biodiversity [4770, 5094]. Bioeconomic [2277, 2287, 2627]. bioelectrical [5050, 4999]. Bioenergetic [2580]. Bioenergetics [3691, 2988, 3526, 4781, 4692]. biogenic [5138]. Biogeography [3567, 3454]. biologic [1488]. Biological [2298, 1998, 1478, 5346, 3354, 4315, 4800, 3822, 1017, 2250, 3221, 4918, 1585, 2371, 1182, 3381, 4873, 1874, 1613, 988, 1978, 4314, 4819, 4987, 1526, 4096, 1113, 4924, 881, 1119]. biologist [2171]. Biology [2014, 3912, 4018, 2950, 2998, 3011, 2527, 3438, 4561, 3470, 3312, 1515, 3321, 2487, 4721, 2410, 2696, 4757, 2808, 2940, 2566, 2193, 3193, 2196, 3373, 4369, 2543, 2712, 2421, 2456, 4218, 2921, 2446, 3274, 3045, 3732, 2551, 3074, 1661, 1678, 2612, 3328, 1994, 3510, 3018, 2365, 2604, 2731, 3082, 4017, 4783, 274, 2048, 3377, 4324, 3480, 4158, 4707, 4389, 3177, 3372, 5093, 5110, 4135, 4670, 3886, 4652, 2738, 3465, 2549, 3178, 4482, 4151, 4528, 3832, 1844, 3859, 4333, 4115, 4741, 3637, 1841, 1908, 1974, 1603, 1667, 1695, 1687, 4634, 3302, 4015, 4949, 2717, 3767, 3909, 1888, 4323]. biology [5123, 4603]. Bioluminescence [2561]. Biomass [2562, 2692, 4974, 3116, 2661, 2870, 2175, 4873, 4283, 4935, 2540, 3151, 4730, 4673, 4712, 4564, 5059, 4544, 3919, 4626, 3829, 5046, 5140, 5166, 3750, 5097, 2818, 3070, 3754, 4550, 5309, 5033, 3799, 4238, 3120, 5171, 4119, 4380]. biomass-based [3919]. biomass-per-recruit [4544]. Biomasses [3522]. biomedical [4775, 5317]. Biometric [1676]. biopsy [4416, 3755]. Biotic [1729, 5375]. biphenyls [3055]. Birds [1296, 1740, 681, 86]. birostris [5318]. birth [3845]. Birthdate [4081, 3688]. Biscayne [2973, 2717]. bites [3762]. biting [2943]. Bivalve [3260, 2350, 4623, 1973]. Bivalvia [2746, 2340]. Black [2770, 5093, 3451, 5099, 2192, 1569, 2560, 3511, 3975, 118, 5189, 4656, 4686, 4910, 4579, 4145, 4194, 5222, 143, 3813, 2924, 367, 4168, 4418, 5353, 5022, 1243, 3815, 2454, 5236, 4162, 3643, 1038, 5230, 4735, 5206, 1137, 3799, 1185, 373, 427, 702, 3308, 3393, 5308, 332, 436, 455]. blackcheek [4240]. blackfin [5116]. blackfish [354, 183, 172]. blackfish- [183]. blackgill [4680]. blacklip [4167, 4877, 3970]. blacknose [4473]. Blacksmith [2872, 3136]. blackspotted [5288]. blacktail [4764]. blacktip [5048, 3817, 3559, 5323, 5232, 3995]. Bladder [3171, 2521, 2781, 2358, 1766, 1514, 1228, 5308]. bladders [756]. Bladderwort [439, 430]. blainville [5037]. blainvillei [4670]. blank [437, 855, 296]. blastodicks [771]. bleeding [4775, 5317]. Blenny [1238]. blind [5363]. Bloch [4025, 1659, 4379]. Block [556, 373]. Blood [3606, 3337, 2235, 2264, 1855, 1920, 1619, 5114, 1966, 65, 1397, 2227, 1308, 4578]. blood-vascular [1308]. blooded [1684]. bloodworm [2642]. blooms [2392]. Blue [3097, 3418, 3283, 3555, 3601, 3254, 2144, 2129, 2739, 3678, 2572, 2485, 2713, 3350, 3993, 3016, 4018, 5303, 1497, 5266, 4920, 5194, 2124, 4748, 54, 133, 3943, 5144, 5278, 4472, 4992, 243, 3462, 4565, 5327, 4621, 4608, 4993, 202, 2157, 4738, 2795, 3323, 4743, 5008, 5291, 3039, 4319, 2703, 4591, 3075, 773, 4679, 5000, 4949, 1994, 1999, 5234, 5247, 5357, 4342, 2681, 3855, 40, 4603]. blue-eye [4018]. Blueback [2923, 1866]. blueblack [3224]. Bluefin [2950, 3596, 3599, 3706, 3709, 3038, 2343, 3537, 2160, 4655, 2778, 2807, 4477, 4963, 4478, 4311, 4073, 4423, 4503, 4710, 4148, 4852, 4428, 4545, 4583, 4584, 2272, 3971, 4091, 4317, 5269, 3159, 2043]. Bluefish [3629, 2715, 3439, 4557, 4249, 4129, 4282, 4283, 4760, 3895, 692, 3775, 4429, 3782, 3812, 254, 486, 4844, 4805, 4915, 5203, 4809, 3828, 4727, 4921]. bluefun [1851]. Bluegill [2342]. Blueline [3074]. bluntnose [4537]. board [267, 464]. boat [4658, 1668, 5323]. boats [981, 5129]. Boca [1995]. bocaccio [4831, 4787, 4786, 4982, 4490]. bocourti [4433, 2897]. bodies [4603]. Body [3343, 2407, 3702, 2383, 2557, 2858, 2744, 4347, 1470, 5007, 4780, 2894, 4004, 4979, 4216, 1673, 4185, 3306, 2929, 4931, 4923, 1378, 3830]. Body-Length [3702]. body-otolith [2894]. bogs [31]. boilers [562]. Bojanus [212]. Bolinas [3167]. bollmani [950]. Bomb [5031, 4618, 4905, 4694, 5159, 4985]. Bomolochid [2854]. bonaci [4579, 4194, 5206]. bonasus [4691, 3328]. Bonavista [3622]. bond [3771]. bondi [5075]. bone [3293, 4727]. bonefish [3048, 4040, 4109, 4195, 3978, 1974]. bones [3580, 4663]. bongo [5066, 4858, 4209]. Bonitos [2389]. Bonnaterre [1894]. Bonne [3221]. bonnetheads [5326]. Bonneville [1847, 2031]. Bony [1701, 1702, 1703, 1705, 1706, 1708, 1711, 1714, 1717, 1719, 1539, 1157]. Boonea [3452]. Boothbay [2621, 2133, 2765]. Bootstrap [3445]. bootstrapping [3818]. Bordeaux [774]. borealijaponica [3856]. borealis [3228, 3708, 2628, 4928, 4732, 2709, 2110, 4233, 2652, 2765, 3295, 3776, 4401, 3625]. Boreogavus [2332]. borer [1529]. born [3353]. Boschma [3283]. bosci [2210]. Boston [52]. BotCam [5001]. both [4920, 3681, 372]. Bothid [2022, 2244]. Bothidae [3028, 2022, 2295, 2285, 2961]. Bothus [2285]. bottlenecks [4181]. Bottlenose [2749, 3464, 3593, 2944, 4485, 2855, 2734, 4878, 4159, 4416, 3766, 4965, 3723, 4187, 4782]. bottlenosed [2719]. bottles [1811]. Bottom [2202, 3125, 3249, 3553, 3400, 4911, 5196, 5284, 4527, 4417, 5046, 3794, 4685, 2906, 5125, 4803, 5350, 4588, 4513, 5033, 5156, 5241, 1423, 5149, 5131]. bottomfish [5001, 3955]. bottoms [1587]. Boulder [1305]. boundary [5209]. Bower [574]. Bowhead [2875, 3197, 4223]. Bowley [5213]. Box [3497, 4426, 3206, 2866]. boxes [1396]. Brachiopoda [1742]. Brachydanio [2426]. Brachyrhyncha [2200]. Brachyura [3575, 1213, 1290, 4433, 2737, 2743, 3121, 3497]. Brachyuran [3682]. Brackish [2374, 4623, 2184]. Branch [110, 53]. Branchiostegal [2964, 2985]. Brandt [409]. Brasil [3296]. brasiliensis [2637, 2086, 2468, 4778, 4587]. Brauer [2154]. Brazil [189, 4670, 3648, 4473, 4653, 5378, 4429, 2655, 4374, 3705, 3808, 4571, 5123]. Brazilian [4327, 5170, 4331, 4633, 4384, 5231]. BRD [4090]. Breach [3902]. breakdown [4078]. Bream [149, 4998, 4668, 2645]. Breast [2441]. bred [3940]. breeders [461]. Breeding [1027, 1698, 1252, 163, 176, 408, 686, 3465, 3994, 1384, 69, 217, 746, 791, 134, 3221, 510, 1522]. Brenham [115]. Bretagne [397]. Breton [4834]. Breuoortia [3129]. breve [2030]. breviceps [3683, 615]. brevipinna [4709]. brevirostris [3948]. brevirostrum [5207, 5315]. brevis [1791, 3686]. brevispinis [4744]. Brevoortia [4327, 2912, 3379, 3947, 4343, 4850, 5098, 5214, 4963, 3713, 2729, 2240, 2878, 2931, 3044, 3054, 2489, 2776, 4749, 3969, 3795, 3888, 3171, 2393, 2029, 5209, 2103, 2879, 4966, 4331, 4235, 3581, 2539, 3278, 2658, 3320, 3786, 3920, 5146, 5335, 3389, 4681, 3417, 3873, 4514, 2095, 4146]. Briarosaccus [3283]. Bridle [4450]. Brief [356, 377, 379, 496, 453, 1148, 929]. bright [4910, 5235]. Brine [2050, 3656, 3432]. Bringing [209]. Bristol [2197, 2474, 5372, 2660, 2304, 4445, 1593, 4715, 963, 3574]. Britain [1355, 889, 686, 1045]. British [569, 3011, 3251, 4310, 4919, 3489, 3457, 3460, 3763, 3006, 887, 932, 3877, 3820, 671, 826, 3480, 5049, 943, 4073, 1078, 4153, 4982, 1044, 4203, 4744]. Broad [456, 4872]. broad-scale [4872]. brocki [2247]. Brood [2516, 2256, 2716, 4812]. brooding [5283, 4868]. Broods [2287, 2627]. broodyear [4398]. Brook [376, 2605, 559, 904, 779, 1382, 151, 1032, 2826]. Brooks [1852]. brought [333, 263, 157, 234]. Broward [4071]. Brown [2843, 2032, 2283, 914, 3156, 3339, 1955, 2772, 4643, 3057, 5009, 2000, 4219, 2685, 2617, 3230, 723]. Brownell [531]. Brownlee [2001, 1986, 2019, 2023, 2026]. Browns [1699, 4645]. brunneus [3470]. Brunswick [3207, 3092, 2648]. Bryde [3277]. Bryozoa [1742, 1436]. bubble [2412, 1180, 2425]. bubblemorph [3520]. buccaneeri [2196]. Bucksport [479]. budegassa [5093]. Budget [2046, 3890, 4004]. Budgets [3054]. Buenos [4161]. buffalo [242, 617]. buffalo-fish [617]. bugs [801]. Buhi [1239]. Building [185]. bulk [4558]. Bulkheading [2443]. Bulletin [4059]. bullnose [5164]. bumper [3760]. buoyancy [2555]. buoys [5166]. Bureau [1379, 271, 1411]. burial [4257]. burkenroadi [2421]. Burnt [3430]. Burrowing [5187, 1500, 4258]. Burst [2902]. burti [3649, 3406]. Burying [3584]. business [1369, 261]. Busycon [4862]. Busycotypus [4862, 5084]. Butcher [1898]. butcheri [4375]. Butte [1657]. Butter [2820, 2832]. Butterfish [3649, 2874, 1324, 4800, 3406, 4122]. butterflyfish [2511]. button [1167]. Buzzard [895]. Buzzards [5084]. By-Products [2177]. Bycatch [4600, 4524, 3624, 3411, 4468, 5057, 4036, 4126, 4090, 4251, 4347, 4657, 2693, 5299, 5153, 4426, 4193, 4581, 5305, 5284, 4797, 4231, 5096, 4909, 4987, 4870, 4800, 4227, 4677, 5026, 4353, 4027, 4614, 4110, 4802, 4491, 4976, 4533, 5145, 4539, 4146]. bycatch-reducing [4090].

C [3964, 4015, 2992, 3105, 2493, 2570, 5159]. C-Benzene [2493]. C. [4916, 1907, 3575, 3882, 2249, 3457, 2060, 4917, 4174, 4984, 3732, 5069, 4433, 4265, 4902, 2961, 3250]. C14 [1898]. Ca [3726, 4599]. Cabezon [3498]. cables [4970]. Cadmium [2356, 2536, 2194, 2480]. caerulea [1835, 1977, 3521, 2139, 2152, 1786, 1837, 1897, 1668, 1766]. Cal. [364].

Calamus [3553, 3018]. Calanoid [1858, 1983, 2290, 1776, 2436]. Calanoida [2261, 2280, 4066, 1793]. Calappidae [3497]. calcarifer [4379]. Calcification [3386]. Calcified [3555, 4449]. calcium [4599]. CalCOFI [5107]. Calculated [3688]. Calculating [2564, 3514]. calculation [3151, 4774, 4429, 4399, 4844, 4532]. calculi [1231]. calculus [4623]. Caledonia [3624, 3911]. calf [4761]. Calhoun [949]. calibrated [5229]. Calibration [3137, 4959]. Calico [2059, 2558, 896, 3790]. California [3759, 3868, 4324, 5312, 4624, 3280, 3702, 4625, 2885, 3334, 4030, 4913, 4704, 4294, 2542, 4177, 4962, 2141, 3688, 4067, 5118, 2167, 4280, 1832, 2977, 3007, 3078, 3167, 4052, 4833, 2926, 3127, 2497, 2568, 2303, 2268, 3444, 3447, 3923, 4864, 3122, 62, 48, 4847, 1947, 2016, 3309, 2597, 3009, 3014, 2423, 2527, 4364, 3536, 4561, 44, 4390, 5334, 4057, 2432, 333, 5101, 2562, 2054, 3126, 3184, 3470, 3277, 5036, 4141, 5107, 5305, 2509, 2823, 2825, 3142, 3360, 3361, 2689, 4787, 241, 3746, 1311, 2553, 2487, 3921, 3983, 4652, 4284, 3700, 4001, 4325, 3719, 4625]. California [2445, 965, 3582, 85, 1117, 1179, 3190, 5017, 4464, 5105, 3178, 2462, 3638, 2791, 2952, 2076, 2490, 3462, 4415, 1366, 2742, 2860, 4274, 506, 5285, 5325, 1085, 1088, 4152, 3050, 4595, 3588, 3666, 2339, 4608, 4820, 4993, 3125, 4022, 3019, 5292, 2341, 4785, 5125, 4444, 4943, 2638, 2774, 2919, 3135, 3174, 3205, 3334, 3617, 4171, 4219, 4803, 3619, 3668, 3994, 4713, 1318, 4484, 2297, 2434, 4446, 2090, 5328, 5370, 4676, 403, 787, 3310, 263, 4867, 2662, 2159, 3246, 3189, 3522, 2542, 3413, 2071, 4696, 3685, 3562, 2836, 3087, 2725, 3627, 2582, 3305, 3796, 4035, 3984, 266, 234]. California [3066, 3757, 1764, 1768, 5024, 515, 4112, 2324, 2476, 3964, 2155, 3022, 2572, 5366, 5358, 2378, 3602, 3585, 3214, 2359, 398, 1341, 1450, 1477, 3805, 3534, 654, 2882, 3510, 3845, 3233, 34, 59, 4765, 4767, 767, 962, 2713, 23, 4268, 3347, 2989, 3519, 5264, 3247, 5038, 4745, 5076, 3258, 4535, 11, 3085, 4051, 2695, 2526, 5329, 4067, 4755, 5040, 4365]. Californian [2901]. californianus [4067, 4625, 2746, 3619, 3668, 4713, 5366, 5358, 4745, 4067]. Californias [2313]. californica [5328]. californicus [5305, 3700, 3666, 4943, 4446, 3685, 4268]. californiensis [2761, 3534, 4365]. call [16]. called [1495, 546]. Callianassidae [3695, 4258]. Callinectes [3016, 4460, 4565, 5327, 2795, 2680, 5008, 4433, 4319, 2824, 4679, 5000, 1994, 1999, 5234, 4342, 3855, 2329, 2897, 3350]. calling [2602]. Callorhinus [2284, 3273, 3492, 4792, 2708, 3316, 3284, 3867, 2889, 2882, 3353]. Callorhynchidae [3886]. Callorhynchus [3886]. callosus [3283]. callyodon [3520]. Caloric [2436, 2218, 2495]. calves [4640, 2303, 4761]. calving [4159]. Calyptopis [2207]. Cambarus [1576]. camera [3937, 5001, 5086, 5217, 5172, 4986]. cameras [5217]. campechanus [5271, 5113, 2773, 4350, 4671, 3000, 4448, 5267, 5111, 4762, 5343, 5064, 4266, 4650, 4451]. Campeche [4281]. camtschatica [2316, 2286, 1873]. camtschaticus [2922, 4445, 5100, 5069, 4875, 4683]. Can [4182, 489, 4725, 256, 3478, 4618, 1567, 675]. Cana [4743]. Canada [3221, 840, 1079, 3207, 1065, 4371, 5006, 3131, 3220, 3148, 3092, 23, 67]. canadensis [5375]. Canadian [840, 874, 3091, 384, 3242, 662, 2748, 4406]. canadum [4452, 3913, 4389, 5095, 5226, 3768, 4014, 4253, 5060, 5287]. Canal [459, 2614]. canaliculatum [4862]. canaliculatus [5084]. Canals [2298, 2217, 2347, 2300]. Canarian [4612]. Canary [2560, 4218]. Canaveral [3922]. Cancer [3457, 3963, 3009, 3014, 3068, 3575, 2451, 2060, 2339, 2504, 2652, 2765, 2470, 2947, 3340, 2850, 2932, 2251, 3047, 3665, 3410, 4679, 3166, 4662, 2714, 5250]. Cancridae [3575]. canicula [5037]. canis [4523, 1426, 4037]. Cannibalism [4012, 5053, 2842]. canning [690, 276, 598]. Canopy [2380]. canyon [4490, 4356, 3243, 5156]. capabilities [5148]. Capacity [2417, 2745, 4929]. Cape [4299, 2625, 2540, 2704, 4111, 2692, 4690, 2192, 3924, 4834, 2355, 2792, 3973, 4895, 3226, 2224, 12, 3922, 3290, 2237, 552, 3021, 2747, 2831]. Capelin [3513, 4768]. capensis [4798, 3924]. Capital [3431]. Capitella [3382]. caprices [327]. caprinus [2990]. capriscus [5120, 5176, 3173, 5368]. Capt. [514]. Captive [3187, 3137, 3643, 3926, 4183, 4428, 2532, 4839, 5300, 3940, 2889, 3845]. captivity [1420, 5275, 2585, 1413, 1375]. Capture [114, 4259, 2569, 4683, 5228, 4961, 3762, 5352, 13, 111, 4653, 97, 4003, 4981, 4029, 1076, 615]. Captured [3601, 3205, 2235, 5039, 4977, 2968, 3303, 3761, 5019, 3640, 5323]. captures [135, 3404, 2227]. Car [269, 76]. cara [2630]. Carangid [2296, 4722]. Carangidae [3760, 4788, 2299, 4837]. carangoid [30]. Caranx [1822, 4788, 4351, 2299, 4837]. carbo [2762]. carbohydrate [1948]. Carbon [2050, 2628, 2770, 2741, 2419, 4618, 5208, 3482]. carbunculus [4252, 3680]. carcass [4358]. Carcasses [3271, 5320, 4193, 5251]. Carcharhincus [3995]. Carcharhinid [2066]. Carcharhinidae [2921]. Carcharhinus [5031, 5048, 4552, 2804, 4328, 4938, 4244, 4562, 4709, 4617, 4781, 4780, 4473, 3559, 5377, 2921, 3211, 3936, 5272, 4633, 4377, 4013, 3792, 5220, 5323, 5232]. carcharias [4358, 5360, 3683, 2955, 4224]. Carcharodon [4358, 5360, 3683, 2955, 4224]. Carcino [3340]. Carcinoma [1455]. Carcinonemertes [3602]. Carcinus [5375, 2003, 2002]. Cardium [1610]. Care [352, 881]. Careproctus [4205]. Caretta [5039, 5065, 4410, 4601, 3385, 4653, 3337, 2867, 4697, 4678, 4615, 4515]. Caribbean [5028, 3773, 5075, 4721, 3695, 4890, 4955, 5318, 2604]. carica [4862]. Caridea [2922, 2515, 2204]. Caridean [3312, 3322]. Carlo [4552]. carnivores [2368]. carnivorous [402]. Carolina [5039, 984, 810, 4963, 5094, 4903, 1577, 456, 3224, 988, 3902, 4462, 2621, 3959, 3781, 1438, 1991, 2808, 3577, 3026, 3737, 3496, 292, 506, 478, 2735, 941, 2192, 5294, 5060, 2103, 3059, 1313, 2510, 4610, 4568, 4739, 4485, 2446, 110, 2792, 4886, 2680, 3226, 4344, 4488, 1474, 4029, 3788, 3074, 3139, 4572, 4661, 5043, 3479, 2558, 3633, 4879, 3790, 131, 997, 615, 5145, 253, 3223, 3873, 4514, 2747, 2831, 2403, 2095, 180, 711, 911, 4323]. Carolinas [1930]. Carolinian [2897]. carolinus [3237, 4465, 4517, 4940]. Carotenoid [2114]. Carp [363, 133, 441, 503, 745, 577, 533, 394, 100, 703, 811, 128, 739, 288, 38, 28, 872, 450, 618, 906, 454, 561, 308, 189, 208, 588, 149, 762, 601, 555, 54, 460, 21, 138, 366, 199, 815, 850, 524, 243, 404, 651, 140, 304, 190, 202, 316, 126, 557, 673, 104, 193, 115, 146, 272, 553, 16, 538, 191, 549, 514, 624, 378, 452, 299, 270, 237, 565, 544, 613, 280, 466, 734, 428, 368, 187, 179, 218, 568, 260, 27, 203, 414, 32, 123, 40, 606, 337, 798, 220, 742]. Carp-culture [577, 703]. Carpentaria [3933, 3005]. carpenter [4707]. carpio [38]. carponotatus [4628]. Carretta [4555]. Carrying [2417, 313]. Cartilage [3293]. Cascadia [3556]. Case [2277, 3494, 2677, 3586, 2302, 2976, 4829, 4527, 5011, 5162]. Casella [551]. Castelnau [4025]. castelnaui [3493]. Castrating [45]. cat [1256, 1252, 1946]. cat-fish [1252]. cat-fishes [1256]. Catalina [3126, 2490, 2572, 2713]. catalogue [1443, 1444, 1491, 164, 1242]. Catch [5039, 2892, 748, 2420, 782, 854, 894, 3543, 343, 2236, 148, 816, 3681, 3906, 2418, 2765, 324, 3116, 3475, 2093, 3415, 3484, 3676, 3639, 1964, 5329, 4499, 5065, 5228, 3374, 2859, 5016, 4426, 5365, 5196, 4863, 1834, 448, 5249, 5129, 1977, 4607, 4483, 5377, 5245, 1668, 168, 673, 4740, 331, 219, 4555, 4758, 3406, 3944, 1769, 4891, 2400, 233, 544, 3092, 5162, 4340, 4875, 4593, 5259, 1795, 4045, 4782, 742]. Catch-at-Age [3475, 4863]. catch-bycatch [4426]. catch-free [4758]. catch-length [4045]. Catch-per-effort [3681]. Catchability [4281, 4016, 3518, 3261, 3047, 4209, 3972, 4563, 4982, 4272]. Catches [4900, 2279, 2528, 2388, 3425, 2548, 5081, 4126, 4211, 4411, 2772, 5133, 3317, 2621, 4821, 4607, 4527, 4215, 4085, 4724, 5035, 2947, 2465, 1797, 2835, 2251, 3865, 5210, 3539, 3716, 4406, 5078, 1721, 4767, 4635, 3955, 4840]. Catching [543, 160, 2652, 511, 345, 210, 528, 549, 4419, 310, 4342]. categorical [4146]. catfish [924, 273, 129, 405, 348, 571, 3908, 184, 567, 1569, 3296, 217, 5291, 630, 24, 848]. Catostomus [1579]. catshark [3236]. Caudal [2143]. caudata [2024]. caufornianus [3127]. Caught [3198, 3143, 2165, 3079, 5212, 4117, 3762, 4617, 5194, 4349, 4849, 4462, 3440, 4748, 99, 3276, 344, 4693, 503, 4987, 459, 4981, 4456, 4614, 12, 4879, 3034, 4533, 5164, 3238, 865, 3247, 5232, 3545, 4228]. Caulolatilus [3074]. caurinus [4381, 5175]. Cause [2086, 3478, 1180, 254, 235, 2500]. caused [667, 4257, 518]. Causes [1414, 1731, 2174, 933, 2255, 2257, 1401, 5087, 178, 4775, 3218, 1035]. causing [1804]. cavalla [4884, 627, 3437, 3486, 4138, 3321, 4366, 4506, 3042, 3854, 4107, 3592, 3699, 3098, 4782]. Cayuga [1824]. Cedar [4179]. Cell [2227, 2235, 2264, 3538, 5089]. cells [1378, 1440]. Census [2749, 3186, 4846, 4167]. Censuses [3624, 2926, 3066]. Central [2096, 851, 5153, 3981, 1565, 3588, 3558, 2100, 3188, 2771, 2499, 3591, 2169, 2113, 2141, 2382, 4833, 3912, 1839, 2089, 2303, 2687, 4357, 5310, 5342, 4348, 4720, 4964, 3360, 2553, 4764, 3866, 4835, 5017, 3954, 4368, 5367, 1887, 4909, 1734, 1792, 1779, 1883, 4608, 5125, 3174, 4713, 4446, 1802, 1969, 3899, 5148, 3859, 3875, 1904, 3974, 3916, 5276, 4162, 5023, 4025, 1820, 1877, 3282, 1983, 4697, 4555, 5052, 3757, 1732, 1794, 3964, 3922, 4179, 4063, 2485, 3852, 1808, 1838, 3233, 4241, 4207, 5264, 5259, 982, 1068, 4051, 2526, 1783, 4067, 2977, 3502, 2219, 3309, 3017]. Central [2433, 2206, 1233, 194, 3650, 2391, 1958, 2470, 2919, 1689, 2603, 185, 1676, 979, 3214, 1140, 3688]. central-northern [4067]. Centrolophidae [4018]. Centropages [4066]. Centropomus [4355, 4131]. Centropristas [4686]. Centropristis [3975, 5189, 5222, 5099, 2192, 5353, 5236, 5230, 3799, 3511, 4887, 3308, 5308]. Centroscymnus [2227, 2235]. Centuries [2313]. Century [441]. cepedianum [1937, 1844, 3357]. Cephalopholis [4521, 4952]. Cephalopod [3504, 2982, 4830]. Cephalopoda [1453, 5133, 4682, 4247, 4755, 3999, 4309, 4034, 3985, 4006, 4197, 3875, 4267, 3554, 3686]. Cephalopods [2941, 2674, 1437, 3954]. Cephalorhynchus [4658]. Cephalurus [4324]. cephalus [1813, 4324, 3368, 4289, 4044, 4082, 2650, 4610, 4568, 4739]. Ceratioid [2812, 2829, 2800]. Ceriantharia [3290, 2518]. cerium [3884]. Certain [2484, 3381, 1190, 174, 921, 920, 965, 1473, 1075, 968, 1489, 1881, 1218]. Certificate [4929]. Cesium137 [1879]. Cestoda [3300, 3790]. Cestode [3397, 1324]. cetacea [657, 4670, 2376, 250, 4571]. Cetacean [4409, 3263, 4907]. Cetaceans [2052, 2108, 2954, 2086, 3916, 3923, 4864, 5202, 5107, 3921, 4619, 5105, 3192, 4589, 3723]. cf [2107]. cf. [2145]. Chaceon [4899, 5178]. Chaetodipterus [3880, 3577]. chaetognath [2368]. Chaetognatha [3650]. chaetognaths [1736]. Chairman [1049, 1023]. chalcogramma [2726, 3871, 5053, 3977, 3572, 3953, 3080, 4672, 3253, 3611, 3704, 4795, 3111, 3373, 3608, 2712, 4859, 3380, 3579, 3893, 4737, 2883, 3901, 3722, 4530, 4275, 4456, 2741, 4489, 4341, 4434, 3949, 4663, 4665, 4378, 4703, 5021, 3526, 4664]. chalcogrammus [5089]. Challenger [3876]. chamaeleonticeps [3465, 3200, 3496, 3041, 1306, 5077, 1308]. chamaelonticeps [3102]. Champlain [431]. Change [3343, 4705, 4920, 5007, 4780, 5058, 5238, 4725, 4915, 5213]. Changes [4911, 2764, 2628, 3336, 5016, 2540, 3876, 2770, 2692, 3652, 4127, 3058, 2521, 2283, 2297, 2442, 1482, 3930, 5077, 4199, 4855, 2093, 3630, 4185, 2055, 2741, 5309, 2929, 4931, 4489, 1831, 77, 3049, 2858, 2756, 2002, 4100, 3853, 3913, 3381, 3163, 5136, 4542, 4689, 4828, 1586, 4545, 1648, 5294, 5175, 1177, 3987, 5321, 5272, 4101, 3161, 4474, 5267, 1397, 4204, 1317, 4650, 4267, 4918, 4654]. changing [4622, 5312]. Channel [2926, 3205, 3832, 5039, 1952, 571, 4316, 2384]. channeled [4862, 5084]. Channels [3385]. Chapman [4133, 2039]. Char [3149, 3446]. Character [514, 487, 370]. Characteristics [2413, 2839, 1685, 2105, 2036, 2022, 2752, 1966, 2592, 3645, 3278, 3162, 3194, 2647, 3379, 3912, 3756, 2928, 5346, 988, 1878, 1978, 3917, 4991, 1885, 4785, 4800, 3973, 3722, 1768, 4177, 197, 3598, 4744, 5184, 5357]. Characterization [5333, 2273, 4063, 5116, 2922, 4965]. Characterizing [5359]. Characters [2915, 2232, 2721, 3663, 4916, 1630, 1574, 2370, 3927, 2209, 5270, 4400]. Charles [169, 182, 531, 12]. Charleston [5039, 2621, 457]. Charlotte [2566, 3092]. charon [2738]. charter [5129, 5323]. charter-boat-based [5323]. Chase [531]. Chatham [1758, 4322]. Cheasapeake [4376]. Cheboygan [1464]. Check [214, 1320]. Check-list [214]. Checking [2866]. Cheilodactylidae [3907]. Cheilopogon [4303]. Cheilotrema [2924].

Chelonia [2267, 3962, 4471]. Chemical [1915, 2628, 2503, 2770, 2036, 2038, 2291, 2024, 2030, 4221, 2741, 2212, 3044, 2001, 1188, 4094, 1526, 1177]. chemicals [1635]. Chemistry [2839, 3606, 3337, 3344, 1744, 4882, 4972]. Chemosensory [2932]. cheni [5127]. Chesapeake [675, 4452, 4441, 3853, 3910, 2654, 3336, 3793, 13, 1613, 3878, 4604, 4028, 4781, 3318, 3101, 2696, 3662, 2549, 1582, 4168, 4418, 4565, 4626, 2477, 4876, 3987, 1975, 2864, 1659, 2244, 4260, 3946, 12, 5291, 3778, 3925, 4822, 3328, 3642, 207, 4240, 4342, 4574, 4575, 334, 370, 527]. Chesapeake- [675]. chief [926]. chiefly [1011, 60, 1653]. Chignik [1612]. Chile [4368, 3409, 3743]. Chilean [4637, 3063]. Chili [682]. Chilian [682]. chilipepper [3988]. Chimaeriformes [3541]. Chimæroids [1538, 1700]. China [4372, 577, 578, 724, 703, 1353, 4422]. Chincoteague [2921, 3211, 470]. chinensis [3801, 3985, 5341, 4157]. Chinese [418]. Chinook [4560, 4559, 2219, 5348, 2520, 4919, 5101, 2412, 2404, 5167, 2234, 2819, 3133, 3408, 3578, 3945, 4069, 3719, 3917, 5290, 1315, 2810, 2755, 2131, 2629, 2252, 4274, 2826, 5233, 2097, 1962, 3588, 4373, 4566, 4484, 2004, 2031, 3181, 4396, 5143, 3784, 4033, 4035, 3398, 1509, 1561, 1596, 1666, 2406, 3271, 2384, 1986, 4128, 3341, 3512, 3770, 2627, 2918, 3990, 2006]. Chionocoetes [2599]. Chionoecetes [3932, 2249, 2886, 2559, 2492, 4174, 5082, 5141, 3710, 5069, 5137, 3986, 4015, 4265, 4340, 4902, 3221, 3622, 3673, 3870, 3311]. chips [3762]. chloride [4311]. Chlorinated [2584, 2382, 2683]. Chlorophthalmidae [4413]. Chlorophyceae [1250]. Chlorophyll [2397, 2037, 2195, 3299]. Chloroscombrus [3760]. chlorostictus [4490]. choice [5284]. choices [5109]. Chondrichthyes [4324]. Chone [2116]. chouicha [62, 48]. Chowan [3224]. Christmas [2033, 2609]. Chromis [2872, 3136]. Chromosomal [3265]. Chronological [676]. chronologies [5344]. chronology [2573]. Chrysophrys [3580]. chrysops [4686, 3128, 4226]. chrysoptera [3095]. chrysoura [2803]. chrysura [1463]. chrysurus [3760, 3940]. chub [3796, 4636]. Chukchi [2332, 4084]. Chum [2417, 3011, 3528, 3480, 4054, 4893, 4934, 4977, 1936, 4950, 5169]. chuss [2036, 2038, 2226, 2305]. Cichlasoma [4463]. cichlid [4463]. Cichlidae [4399]. Ciega [1995]. Ciguatera [2816]. Ciguatoxic [2816]. ciliates [3239]. ciliatus [4648]. cimbrius [2722, 4213]. cinerea [1628, 2024]. circle [4992, 4693, 5043, 5232, 5180]. circulars [437, 855, 296]. Circulation [1549, 1811, 1972]. circulatory [1920]. Circuli [3632, 4689, 3671]. cirratum [5228, 4183]. Cirripedia [1202, 2432, 1328, 1327]. cisactica [595]. cisco [1644]. Citharichthys [2700, 3028, 4984, 2665, 2244, 3031, 3964, 2961]. citulina [2284]. City [1015]. Clackamas [800, 843]. Cladistic [3567]. Clam [2625, 3449, 2374, 3101, 2259, 2322, 3191, 2783, 4243, 1807, 2292, 2544, 1806, 6, 1183, 1184, 2814, 2576, 3070, 3226, 1803, 721, 658, 1787, 2707, 1564, 1621]. Clams [3318, 3365, 905, 1996, 2791, 713, 3106, 3375, 658, 166, 221, 483, 689, 1905]. Clarence [1641]. Clarification [2109]. clarify [5285]. clarkei [2433]. clarki [3637, 3637]. Class [2539, 4459, 1772, 1897, 4074, 4696, 3984]. classes [4507].

Classification [2925, 4192, 994, 4064, 5071]. Classifying [2999, 4459, 1817]. clathratus [4030]. clause [631]. clavata [4588]. Cleaning [2179, 2216, 2076]. cleanup [2683]. Clear [1085]. cleavage [1929, 107, 771]. cleavage-figures [107]. Clemente [3619]. Cleve [2257]. Cleveland [336, 335]. Climate [5008, 5146, 5097, 3796, 5248, 1865, 4918, 4654, 5169]. Climate-related [5008, 5146]. Climatic [1790, 2987, 4954]. cline [5087]. Clinocottus [3746]. Clinton [980]. Clipperton [4007]. closing [2835]. closure [4911, 5080, 4939]. closures [4231, 4821]. Club [693]. Clupea [2900, 4032, 3001, 4130, 3199, 2570, 3383, 2045, 5006, 4098, 2975, 3130, 4074, 3514, 4696, 4513, 3306, 2641, 4360, 1603, 1623, 1624, 3712, 2264, 2208, 3802, 2015, 4341, 2531, 4434, 2893, 2491, 3574, 5025]. Clupeid [2854, 3493]. Clupeidae [4850, 5098, 2878, 2931, 3054, 3859, 4681, 3082]. Clupeids [3605]. Clupeiformes [4850]. clupeiformis [43, 55]. Clupen [1640]. clupeoids [872]. cluster [4582]. clutch [3986]. Cnidoglanis [5016, 3908]. Co [5096, 4350, 4956, 3732, 4535]. Co-occurrence [5096, 4535]. co-occurring [4350, 4956, 3732]. Coal [87, 1529]. Coast [2094, 2654, 2206, 3593, 2210, 2187, 178, 2128, 3457, 4418, 1251, 2715, 2418, 2830, 1113, 2655, 1141, 4155, 3415, 4613, 3398, 3076, 57, 1126, 1153, 807, 2518, 793, 4089, 966, 3961, 4242, 1133, 1782, 1784, 1813, 4845, 4324, 984, 1822, 3013, 5005, 1807, 3634, 2693, 4282, 4412, 5094, 4669, 5333, 4617, 1041, 5153, 878, 139, 927, 3437, 4946, 5222, 5095, 1978, 4889, 1458, 5225, 3748, 3775, 92, 5096, 1574, 1638, 1170, 5027, 5046, 1439, 795, 2596, 961, 3077, 4637, 2947, 4714, 4695, 5240, 5037, 4917, 4199, 1689, 2821, 1664, 5097, 4156, 4549, 2725, 4886, 2274]. coast [5030, 1595, 2251, 1328, 944, 4177, 5043, 952, 4633, 1692, 1676, 3922, 4179, 4879, 3034, 4805, 1006, 1076, 5343, 3889, 3792, 3740, 1812, 3828, 4593, 132, 2043, 4301, 413, 4880, 4514, 2989, 3247, 4571, 2604, 1527, 2707, 3995, 4224, 4380, 5369, 5123, 4603, 2296, 3629, 2749, 4453, 3213, 3543, 1176, 341, 981, 3486, 339, 899, 1008, 1150, 802, 1880, 194, 639, 849, 6, 2993, 167, 3041, 2525, 1075, 2367, 2211, 2528, 3527, 3695, 2470, 2958, 37, 2724, 2923, 2603, 658, 808, 2155, 3826, 2035, 1080, 2141, 221, 689, 328, 900, 1107, 192, 3341, 1028, 3631, 3688]. Coastal [2122, 2942, 3666, 2956, 2438, 3015, 3252, 2552, 2970, 2326, 2534, 3519, 3495, 2289, 5065, 2317, 3910, 3122, 4190, 4389, 5087, 2474, 3953, 5281, 2869, 4193, 4073, 4658, 5199, 3057, 5284, 5355, 4159, 5009, 3055, 4001, 3823, 4945, 4473, 4214, 3462, 4151, 5209, 2593, 3019, 4171, 4161, 4972, 3899, 5374, 3637, 5230, 3788, 3083, 4421, 1341, 1989, 3230, 2648, 1995, 3485, 4207, 4575, 5025, 2111, 3627]. Coasts [2615, 4798, 4364, 4709, 1600, 1765, 1972, 4726, 4355]. cobalt [2596]. cobalt-60 [2596]. Cobb [3822]. Cobble [3125]. Cobble-Bottom [3125]. Cobia [5095, 4452, 3913, 4389, 5226, 3768, 4014, 4253, 5060, 5287]. cockfish [3886]. cockle [1610]. Cod [3228, 3062, 3093, 3435, 3081, 4111, 105, 3405, 3029, 3561, 50, 66, 2883, 3973, 4895, 2332, 3132, 323, 257, 2869, 15, 341, 529, 723, 905, 4896, 258, 339, 822, 5313, 1588, 4730, 206, 4947, 4288, 4988, 2666, 710, 3804, 4866, 4645, 5311, 4200, 3810, 5041, 82, 322, 215, 330, 420, 509, 697, 4906, 475, 4236, 4136, 435, 240, 3774, 4698, 3710, 902, 808, 4419, 306, 1618, 3905, 5244, 300, 226, 4853, 702, 807, 5179, 3758, 291, 1888, 4654, 2224, 2237, 3021]. cod-fishery [15]. Cod-fishing [50, 808]. cod-hatching [710]. cod-nets [15]. coded [3731, 5293, 4131]. codends [4036, 4126, 4210, 4251]. Codfish [12, 13, 537, 870, 695, 774, 917, 183, 238, 499]. codfisheries [132]. Codfishing [109]. Codium [1957, 2351]. cœca [37]. coefficient [4016]. coefficients [3972]. Coelenterata [3290]. coelenterates [1755]. coelestis [4882]. coelolepis [2227, 2235]. coeruleoalba [3861]. Coherence [3094]. Coho [2506, 2242, 2105, 3884, 3408, 3632, 4689, 4442, 2952, 2826, 5233, 2256, 2442, 3088, 3446, 2557, 3442, 3550, 3398, 2425, 2753, 4203, 2648, 4592, 3233, 4043, 2415, 3046, 2287, 2786]. Cohort [4028, 3984, 3946, 5356]. cohort-specific [3946]. Cohorts [3642, 4249]. coioides [5190]. Cojoined [3456]. Cold [3455, 475, 2661, 4087, 3932, 1684, 407, 5114, 3443, 793, 566]. cold-blooded [1684]. cold-core [4087]. cold-water [3443]. Colima [4726]. collapse [4673, 3802, 3870]. collapse/recovery [3802]. collapsible [4875]. collateral [4937]. collect [173]. Collected [3420, 3184, 3596, 3628, 1284, 945, 1202, 3711, 3466, 1292, 3715, 1233, 3704, 4526, 2775, 1077, 1531, 1584, 1571, 169, 182, 1245, 1305, 1178, 1295, 1100, 1015, 1319, 1291, 4344, 1289, 1242, 1297, 4773, 2428, 1203]. collecting [216, 2188, 463, 696, 464, 1091]. collection [1609, 204, 958, 1169, 1000, 1162, 997, 1068]. Collections [2044, 2112, 62, 1216, 1262, 991]. colliei [2120, 2846]. Cololabis [2269, 2282, 5177, 3515]. Colombia [3695]. Colonies [162]. colonorum [4701]. colony [56]. Color [1274, 1558, 2145, 2376, 1482, 69]. Colorado [46, 1576, 1070, 4325, 973, 1305, 1326, 567]. coloration [721]. colors [1492]. Columbia [2219, 3546, 1931, 2106, 700, 569, 3011, 3251, 3480, 4310, 5049, 2687, 3037, 2928, 1685, 4919, 1655, 2959, 4073, 2325, 2009, 2819, 3489, 1106, 1150, 1073, 3917, 1429, 2131, 521, 349, 3457, 3460, 709, 1962, 2256, 4153, 3658, 1866, 3763, 3153, 3218, 4982, 3117, 3914, 1069, 2031, 3006, 2327, 2538, 887, 932, 3877, 3820, 3784, 4885, 671, 826, 1509, 1596, 1660, 1666, 4203, 357, 806, 1160, 4744, 4592, 1829, 3341, 5277, 2415, 3046, 2287, 2627, 2716, 2918, 5062, 5134, 2006]. Columbian [1025]. Column [2964, 2985, 3958]. combating [1420, 1375]. comber [4764]. Combined [1907, 2350]. combining [5056]. Comision [560]. Commander [1566]. commanding [170]. commensal [4907, 3256]. Comment [2351]. Comments [2092, 2499, 2145, 2925, 2890, 2722, 3036, 4774, 3791, 2676, 4333, 4344, 1976, 4701, 2257]. Commercial [2615, 2487, 1611, 2528, 3695, 3205, 3411, 1122, 1364, 3002, 3264, 1933, 2987, 2716, 3711, 4878, 3741, 3876, 4794, 3437, 4896, 5194, 4604, 2309, 4563, 5249, 4889, 4690, 1828, 3846, 3404, 4417, 4987, 1256, 4791, 1885, 1113, 4246, 1595, 1653, 2991, 5126, 4419, 1152, 3640, 4976, 5149, 1621, 4365]. commercial-type [3437]. Commercially [3561, 5344, 3440, 1990, 4167, 5195, 1973, 4865]. commerson [4714]. commersonii [1579]. Commission [125, 1195, 1355, 313, 90, 267, 562, 62, 48, 945, 1128, 150, 189, 21, 1025, 138, 159, 206, 140, 190, 953, 961, 193, 768, 263, 1158, 628, 170, 1012, 1289, 437, 855, 979, 131, 296, 231, 676, 238, 134, 879, 939, 963, 1091, 499, 187, 291]. Commissioners [214, 264, 329, 248]. Commissions [1122, 313]. Committee [545, 1023, 53, 1016]. Common [3280, 3703, 2363, 1712, 5255, 4799, 175, 1602, 3882, 5284, 3967, 88, 4454, 2720, 1184, 5331, 4487, 197, 658, 4355, 4131, 1650]. Communities [2836, 3684, 1747, 5142, 4939, 4835, 4611, 4297]. Community [2380, 3279, 3701, 3164, 4911, 4845, 4997, 4830, 3160, 5080, 1891, 4511, 4885]. companies [448]. Company [470]. Comparative [3542, 3575, 2187, 1953, 3555, 680, 3524, 1792, 2339, 2223, 5037, 2883, 3591, 3562, 2409, 2018, 3016, 4032, 4997, 4991, 3192, 1563, 1387]. Compared [2628, 625, 197, 4662]. Comparing [4785]. Comparison [2915, 4848, 3358, 3467, 3713, 3837, 3317, 2181, 2082, 2183, 5122, 3141, 5279, 4750, 3168, 4723, 2320, 2112, 2525, 3624, 3288, 3617, 2910, 3433, 1948, 3647, 5350, 2795, 5276, 2226, 3087, 2427, 3249, 5033, 1692, 1699, 5335, 3531, 4009, 5019, 5347, 3497, 4897, 3797, 2945, 4283, 3466, 3934, 4709, 4772, 3803, 5106, 5115, 5066, 4581, 4968, 4167, 4393, 4394, 3100, 4502, 5182, 3279, 5345, 3215, 3896, 3989, 1868, 2835, 4467, 3786, 3920, 5043, 65, 4971, 1676, 5260, 3925, 4449, 4125, 4406, 3885, 5064, 3828, 4385, 5264, 5180, 4922, 5250]. Comparisons [3779, 3521, 2904, 3908, 3522, 3722, 2548, 3480, 4350, 1899, 3333, 3129, 4154, 4681, 4746]. Compartmentalized [2578]. compatibility [3338]. compelling [5175]. Compensatory [3651]. Competing [2411]. Competition [2499, 4989, 2701, 4904, 4811, 4715, 4434]. complementary [5056]. Complete [2915, 1259]. Complex [3661, 4881, 4513, 3107, 4837]. complexes [3791]. complexity [4835]. Component [2929, 2531]. Components [3032]. composite [4126, 4251]. Composition [3007, 2992, 3603, 3185, 3315, 2279, 3650, 2763, 2163, 530, 2058, 2211, 2418, 2504, 2765, 3356, 3409, 2424, 3703, 3235, 2736, 2603, 3639, 3625, 3495, 2083, 2887, 1915, 2684, 4310, 4893, 1988, 3160, 5170, 4349, 4952, 3044, 1188, 3200, 4607, 4975, 4482, 2282, 5216, 4151, 4004, 4777, 4502, 3807, 5006, 5321, 2538, 1987, 3918, 5052, 5055, 4931, 4163, 5192, 3120, 3571, 4574, 5139, 5259]. Compositions [3341, 4987, 3155, 4375]. Comprehensive [3430]. Compression [5267]. comprising [2280, 185]. Computer [2633, 2396, 2668, 1950, 3671]. Concentrate [2036, 2213, 2045]. concentrates [2319]. Concentration [2356, 3500, 3299, 2588, 3948, 2127]. Concentrations [2404, 2037, 1680, 2865, 2498]. concentric [4454]. Conception [3622]. Concerning [397, 408, 1590, 70, 727, 202, 1055, 316, 82, 5047, 39, 720, 152, 1514]. Conch [3394, 3584, 3618, 3696, 3733, 4644, 3824, 3718, 3885, 3852, 4041, 4206]. Conchs [3354]. concolor [3245]. Concurrent [4358, 2442, 3542, 4689, 4957]. Condensed [2212]. Condition [728, 3949, 3258, 3950, 1528, 4936, 502, 4624, 5050, 4057, 4953, 4852, 3827, 1837, 4101, 5089, 1603, 512, 4716, 1153, 2849, 484]. conditional [4225]. conditioned [2242]. conditionof [5082]. Conditions [2016, 3646, 2688, 2217, 2347, 2180, 2586, 2200, 2670, 2743, 3154, 3452, 3523, 4477, 1619, 1657, 2533, 5262, 3367, 381, 4993, 3288, 5191, 550, 4374, 1982, 4885, 1598, 5242, 1345, 2737, 767, 5038, 4599]. conducted [1072, 1012, 185]. conductivity [4216]. coney [4521]. Conference [410]. Confidence [3947, 3210, 5192]. configuration [4802]. configured [4406]. confinement [333, 3241, 307]. confinements [479]. confinnation [3248]. confirmation [5036, 4618, 5272]. conflict [4315]. confront [461]. congeneric [4576, 3287]. conger [1224, 88, 88]. conger-eel [88]. Congress [1139]. Congridae [2379]. conical [4875]. Connecticut [1834, 782, 854, 894, 964, 1731, 3342, 5339, 1797, 3357, 1598, 3268, 3297]. connecting [1219]. Connection [143]. connective [1188]. connectivity [5028, 4955]. Consequence [3674]. Consequences [2802, 4351, 2793, 3365, 5087, 3299]. consequent [1141]. Conservation [3690, 4229, 4992, 1595]. Conserving [4646]. Consideration [2901, 2151, 5096, 1326, 1129]. Considerations [2844, 5362]. considered [1011, 4832, 1357, 49]. consistency [3226]. Consistent [3953]. conspecifics [2242]. constant [4508, 2425, 3838, 5131]. constant-F [4508]. constellatus [4337]. constituents [1222, 1529]. constrained [5316]. constraints [3915]. construction [1052, 316, 414]. consumed [4283, 3253, 4663, 4665, 4664]. Consumption [2843, 2452, 3313, 3284, 3325, 4360, 5366, 4052, 3963, 4963, 4781, 1617, 130, 5068, 4443, 3820, 4487, 5358, 173, 4049, 4551]. contact [4970, 5131]. contagious [1722]. container [2849]. contaminant [3796]. contaminants [4094]. contamination [2695]. contemporary [4398]. Content [2258, 4452, 1581, 2764, 1949, 2495, 5006, 4216, 1394, 4768]. contenting [298]. Contents [526, 3081, 3596, 2112, 2114, 3211, 2276, 3325, 3934, 3440, 4014, 2591, 2856]. Continental [2618, 3470, 2314, 3263, 2013, 3150, 3402, 4282, 4478, 3160, 3941, 4619, 4505, 4750, 4066, 4736, 5125, 4695, 3916, 5314, 5156, 4810, 3747, 4321, 4809, 4773, 5224, 2403, 5090, 2138, 3420, 2508, 2224, 2237, 2518, 3419]. continents [1765]. Continuous [2223, 4444, 1396]. contracts [5322]. Contrast [3610]. Contrasting [5042, 3525, 3070, 5211]. Contrasts [3521]. contributing [5371, 4375]. Contribution [2092, 2654, 3372, 2445, 2661, 1603, 2753, 2911, 4015, 2415, 2287, 2627, 2716, 2006, 4884, 120, 992, 972, 103, 2717, 3046]. Contributions [2048, 1834, 1625, 1667, 1687, 1527, 3308, 1104, 2366, 26]. Control [2024, 2405, 2809, 1874, 1592, 1959, 4318, 5330]. Controlled [2011, 2180]. controlling [1654, 1598, 2845]. convention [631]. conventional [4948, 4971]. Convergence [3502]. Conversion [3313, 4933, 2184, 3338]. Conversions [3142]. Convict [1764, 1768]. cooccurring [2508]. cook [561, 366, 2886, 4297]. Cooked [3403]. cookiecutter [3363]. Cooking [2441, 762, 815, 844, 237, 187, 798]. Coonstripe [2459]. Coos [3578, 4069]. Coosa [233]. Copepod [3441, 3260, 2655, 1476, 3851]. Copepoda [1793, 2261, 2280, 2761, 4066, 4407]. Copepodids [3260]. copepodites [4066]. Copepods [3126, 2027, 2854, 3051, 2064, 1776, 1858, 1967, 3050, 2436, 1983, 2290, 1225, 1181]. Copper [2356, 2641, 1856, 2861, 1634]. Copulation [2811]. Copulation-Induced [2811]. Coquimbo [3743]. Coral [2335, 3407, 3079, 4499, 4164, 4799, 3803, 4135, 3819, 3864, 4930, 3303, 4402, 4765, 5171, 5361, 4594, 4307, 4908]. coral-reef [4499, 4594]. coralgrouper [3981, 4888]. Corallimorpharia [2518]. corals [542, 1197]. Corbicula [4623]. corbis [1610]. cordiceps [972]. Core [3204, 2661, 4087]. Coregonid [1589]. Coregonus [9, 43, 55, 2550]. cores [3943]. coriacea [4847, 5352, 4653]. corn [146]. cornutus [4984, 2961]. Coronulidae [3723]. corrected [4980, 248]. Correcting [4085, 4175, 3539]. correction [3715, 4749, 4663, 3482]. Correlates [2720]. Correlating [5138]. Correlation [2346, 2420, 3737, 2370, 2947]. Correlations [3513, 1470]. Correspondence [174]. correspondent [123]. corrugata [3510]. coruscans [3552]. Coryphaena [3926, 5379, 3882, 5240, 4292, 5148, 4487, 3096, 3282, 2085, 2837, 4336, 4879, 3250, 5185]. Coryphaenidae [3882, 2196, 2837]. cost [4691, 4235, 3774]. Costa [3503, 5153, 2938, 1892]. Costia [1406]. Costs [5322, 2479]. Cottbus [208]. Cottidae [4928, 3366, 3746, 3008, 2981, 3203, 4771, 2888]. cottoid [3520]. Cottus [2438, 2557]. couesi [2911]. coulteri [1945, 2221]. Counter [3132]. Countershading [2561]. Counties [949]. counting [1376]. countries [889, 1014]. Counts [3994, 3663, 2235, 2264, 1864, 4124, 3933, 2227]. County [643, 4141, 1088, 1305, 1090, 980, 3666, 4071, 244, 1464, 1764, 3092]. Coupling [4627]. Courland [395]. course [3736]. courses [1104]. Courtship [4722, 2586, 5206]. covariance [3679]. Covariation [3014, 3970, 3009, 4820]. Cove [2417]. coverage [3842]. cowcod [4561]. Cowlahls [3450]. Cownose [3328, 5334, 4691]. cows [47]. Cozumel [945]. Crab [2622, 2420, 3014, 794, 2643, 2696, 2316, 2969, 3457, 2595, 2339, 2504, 2765, 3561, 2470, 2501, 2194, 2932, 2200, 3047, 3487, 2743, 3428, 3665, 2911, 3622, 2118, 3371, 3350, 3377, 3963, 3016, 3009, 3150, 5375, 1497, 4769, 2451, 2553, 2549, 2886, 2559, 2492, 4565, 5327, 4899, 4445, 5100, 5158, 2947, 3832, 3340, 2795, 5082, 5141, 5317, 2850, 2251, 2599, 4597, 2286, 3734, 3710, 2003, 5137, 3986, 2760, 2737, 5330, 971, 955, 3075, 3302, 4205, 4265, 4340, 5178, 4875, 4902, 5000, 1994, 4662, 4436, 3870, 3233, 1230, 5234, 5145, 3720, 1873, 4683, 2480, 2002, 3855, 2714, 2897, 3099]. crab [5250, 4045, 4603]. Crabs [3068, 2216, 3283, 3254, 2060, 2652, 2589, 3682, 3166, 3673, 3311, 2329, 3497, 3868, 3402, 3932, 4460, 1223, 4933, 4620, 2104, 4775, 346, 528, 4174, 4614, 3107, 5069, 5008, 4433, 4319, 3410, 4793, 4015, 4679, 3640, 4951, 3103, 1999, 3221, 4342, 4578]. Craig [1032]. crameri [4553, 2706]. Crangon [2775]. crangonid [2775]. Crangonidae [3193]. cranial [3717, 3879]. crashes [3881]. crassidens [2801]. crassipes [2060]. Crassostrea [4604, 4924, 2498, 3452]. Crater [2086, 3363]. crawfish [516, 1104]. Crayfishes [1576]. creature [321]. Creek [2516, 48, 2506, 2677, 1764, 2865, 573, 4759, 511, 512]. Creeks [3392]. creel [5171]. creole [4521]. creole-fish [4521]. creosote [1529]. Crepuscular [2901]. crest [4890]. crevalle [4351, 4837]. crevalles [1822, 30]. cribrarius [4460]. crimson [4564, 5135]. crinacea [771]. Crisfield [971]. Cristine [376]. Criteria [3183, 2397, 2326, 4252]. critical [1909]. critique [1597]. Croaker [2010, 2040, 3455, 3429, 3135, 3604, 5097, 3389, 2989, 2537, 3853, 3910, 4120, 4968, 2924, 1638, 5154, 4291, 3059, 4567, 4548, 2792, 4139, 4296]. Croaking [884]. croicensis [2634]. Croix [3208]. cromis [4145, 3813, 4168, 4418, 5022, 3815, 4162]. Crop [3618]. Croppies [255]. Cross [4695, 1968, 4756]. Cross-reactive [1968]. Cross-shelf [4695, 4756]. crossotus [2961]. crossroads [4679]. Crowding [2170, 2400]. crude [1642]. cruise [211, 5107, 2668, 626, 597]. Cruises [2367, 3425]. Crustacea [2077, 1843, 2805, 2915, 1751, 3868, 3569, 2232, 2678, 2769, 3182, 3364, 3948, 2201, 2216, 2333, 2410, 2688, 2533, 2922, 3193, 2813, 2972, 2721, 2864, 4258, 2079, 4630, 3295, 1973, 2731, 2315]. crustacean [3355, 5157]. Crustaceans [3597, 3495, 1930, 3419, 3432, 2440, 1493, 5195, 2477]. Crustecdysone [2060]. Cruz [2582]. crysoleucas [2407]. crystal [3948]. Ctenolabrus [442]. Ctenophore [3662, 2368]. ctenophores [1755]. Cuba [1261, 645]. Cuban [532, 2737, 2743]. cubana [275]. Cubiceps [4096]. Cucumaria [4776]. cucumber [4776]. cucumberHolothuria [5239]. Cull [2528]. cultivated [898, 680]. Cultivating [399, 1373]. cultivation [1381, 347, 26, 255, 753, 155, 346, 1309, 813, 434]. cultural [1049, 576, 1505, 1388, 1050, 1411, 711]. Culture [898, 563, 650, 2449, 3695, 2101, 984, 983, 851, 1056, 618, 502, 1026, 1372, 72, 287, 620, 990, 1004, 1240, 416, 899, 63, 96, 1592, 273, 1054, 64, 564, 95, 1606, 1520, 302, 1098, 1183, 1485, 1136, 311, 2814, 1243, 577, 271, 1361, 458, 1344, 1368, 1466, 666, 4258, 845, 1063, 732, 1019, 857, 683, 1020, 236, 358, 512, 952, 847, 578, 1869, 724, 418, 164, 472, 703, 1116, 1142, 811, 1872, 259, 315, 1499, 1530, 203, 1104, 590, 128, 664, 739, 220, 447]. Cultured [3060, 3314, 2594, 4082, 5135, 2845]. cultures [1791, 1898]. culturists [1421, 3]. Cumberland [1000]. cuneata [2374]. Cunner [2536, 2427, 2727, 2222, 442]. cunners [3067]. cunnet [4123]. Cured [2075]. curema [1782, 4289, 4609]. curing [655]. curious [870]. Current [4499, 2054, 2541, 2090, 2888, 3600, 1947, 1757, 4724, 4752, 2497, 2568, 4864, 4624, 3536, 5101, 2562, 5036, 5017, 2462, 2395, 5285, 4096, 4867, 3522, 3562, 3305, 5038, 3258, 2526, 5040]. Current-Geographic [3522]. Currents [2872, 1652]. curvature [2905]. Curve [2904, 2782, 1977, 5179]. Curves [2485, 3908, 3818, 1935, 4385]. cutlassfishes [4372]. cutlassfishTrichiurus [5254]. cutthroat [1810, 3637]. cuttlefish [4285]. Cuv [627]. Cuvier [1778, 1969, 1868, 3592, 5194, 4358, 4295]. cuvieri [3358]. Cybium [61]. Cycle [2191, 3009, 2625, 3498, 2978, 2510, 3705, 2603, 2322, 3191, 3314, 2317, 3910, 4656, 3727, 4285, 2613, 2874, 2924, 3028, 4095, 5274, 4981, 3375, 4733, 5021, 4422]. Cycles [3190, 3181, 3628, 2505, 2746, 3212, 1846, 4587]. Cyclic [3014]. Cycling [2185, 2917, 2156]. Cyclopoid [1967]. Cyclopsetta [2022]. Cyclopteridae [2466]. Cyclopterus [5319, 3491]. Cyclostomata [3631]. cygnus [3216, 3472, 4855, 3002]. Cylcostomes [1713]. Cymothiodae [2604].

Cymothoa [2604]. Cymothoidae [3031]. Cynoglossidae [4280, 3661, 4144, 3732, 4962, 4240]. Cynoglossus [3732, 3381, 4432, 3345]. Cynoscion [4916, 3438, 4576, 3488, 1470, 2986, 3416, 2619, 2487, 5074, 1600, 3748, 3896, 3987, 4161, 4917, 2446, 4101, 1659, 4359, 4598, 4632, 4726, 4661, 3114, 3162, 2935, 5203, 3642, 3671, 5130, 4836, 2488, 1459]. Cyprinidae [3086]. Cyprinodon [2034, 968]. Cyprinodontidae [3212]. cyprinoid [1084, 975]. Cyprinus [38]. cytochrome [3939]. cytometry [5089].

D [507, 2257]. D.C. [185, 979]. dace [1384]. Dactylopsaron [4596]. Daily [3435, 4123, 2872, 2455, 2914, 2634, 3081, 2827, 5265, 4113, 3514, 2278, 4101, 4104, 3071, 3680, 2435, 2916, 3493, 2893, 2896, 3481, 3393, 4129, 4963, 3134, 4028, 4454, 4444, 1225, 3159, 2511, 3499, 3792, 3250, 4921, 4119]. Dakota [462]. dalei [4414]. Dall [3516, 3576, 2974, 2882, 3659]. dalli [3576, 2882, 3659]. Dam [1931, 2810, 2106, 3332, 1601, 1847, 1954, 2031]. damage [4855]. Damariscotta [2906, 3323]. dams [239]. damselfish [4736, 4882]. damselflies [1499]. Dana [3575, 2280]. danae [2680]. dangers [4229, 220]. Daphnia [2472]. darkblotched [4553]. Darkness [2148]. dart [2807, 5287, 4276]. dart-tag [5287]. darter [1577, 1531, 980, 1253, 1446]. darting [4416]. Dasyatis [2199]. Dasyrhynchus [3397]. Data [3483, 3477, 2892, 3280, 2692, 3706, 3521, 3186, 3613, 3570, 3354, 2273, 2507, 2067, 3475, 2482, 3313, 2569, 3625, 2253, 3347, 3631, 4558, 4829, 5057, 4250, 3711, 4959, 4919, 2540, 4804, 3157, 5194, 4581, 4710, 5107, 3744, 3842, 4863, 5363, 4149, 4652, 5326, 5074, 4826, 4075, 5336, 3333, 4545, 5371, 3118, 4987, 3804, 5331, 5245, 4445, 5166, 4917, 4660, 4175, 4529, 5261, 4530, 4513, 3730, 4031, 4027, 5287, 3982, 4467, 5011, 4532, 3905, 4377, 4340, 4861, 5347, 4824, 5177, 3828, 4593, 4700, 3770, 3345, 4717, 5256, 3998, 4897, 4146]. data-poor [5011]. Data-processing [2253]. date [4795, 4696]. Dates [230, 264]. dating [5031]. Dauphin [2804]. Day [2928, 2509, 3917, 4995, 2548, 3079, 4652, 3104, 2388]. Daylight [2148]. days [1116]. daytime [4493]. DDD [2297]. DDE [2297]. DDT [2054, 2068, 2297, 2434, 2266]. deacon [5303]. Dead [795, 411, 538, 452]. deani [3631]. death [531]. deaths [3766]. decadactylus [4998]. decadal [5248]. decade [159, 77]. decades [5113, 5309]. Decapod [3495, 1930, 3419, 3432, 1493]. Decapoda [2915, 3569, 2232, 3182, 3364, 3948, 2922, 2972, 4258, 4630, 2992, 4414, 3575, 2769, 3170, 3144, 3193, 3798, 3516, 2318, 3924, 2785, 3695, 2448, 3295, 2070, 2053, 2200, 2680, 2515, 3076, 2204, 3887, 2824, 3870, 2731, 2329, 3099, 3456, 4603]. Decapterus [3352]. December [631, 422, 1128, 789, 769]. Decision [2507, 2793, 4175]. Decline [4708, 4852, 1683, 3359, 1816, 2255, 2257, 3983, 1804, 3233, 4492]. Declines [4171, 4957]. decomposition [5320]. decompression [4910]. Decrease [3025, 154, 744, 256, 438, 933, 1035]. Decreased [3612]. decreasing [2772, 3989]. deduced [1811]. Deductions [1514]. Deep [2433, 2124, 2696, 3547, 2200, 3319, 2929, 2911, 1109, 2195, 2731, 1311, 2549, 4930, 4218, 4555, 2991, 1020, 3887, 3471, 1567, 3747, 4765, 1663, 4322, 4356].

Deep-Reef [3547]. Deep-Sea [2433, 2696, 2731, 2929, 2911, 2549, 4930, 2991, 4322]. deep-trap-net [1663]. deep-water [1020, 3887, 4356]. Deepsea [2890, 4899, 5178]. Deepstar [2124]. Deepwater [3312, 3362, 3176, 3311, 4833, 3834, 4564, 3073, 4554, 3647, 5295, 5308]. Defenses [2077]. define [4772]. Defines [3538]. Defining [4861]. Definition [1840, 4832, 2176]. Definitions [1673]. Deflation [3171, 3888]. deformity [1963]. Degeneration [4828, 2208, 3969, 1414]. degree [1561]. Dekay [2110]. Del [2324]. Delaware [2983, 2401, 2460, 4769, 4620, 816, 2635, 2948, 4548, 713, 8, 628, 4597, 4924, 955, 2612, 4793, 5234, 3836]. delay [4036, 1931, 4053]. Delayed [2051, 5250]. deleterious [5112]. delineating [4834]. Delineation [3041, 3907]. Delphinapterus [4404, 2072, 4371]. Delphinid [2145]. Delphinidae [2376, 3723, 3952, 4571]. Delphinus [3280, 3967, 2376, 2720, 3703]. delphis [3280, 3967, 2376, 2720, 3703]. delta [4819]. Demand [2503, 4832, 4692, 2834, 5162]. Demersal [3391, 4084, 2942, 4585, 3701, 3232, 2663, 4833, 3466, 3160, 5080, 4505, 4995, 5096, 5046, 4226, 4511, 5012, 4806, 4321]. Democratic [2267]. Demographic [4052, 3566, 3935, 4991, 4565, 5190, 4013, 4552, 5005, 3811, 4671, 4298]. Demographics [5015, 5039, 5108]. demography [5102, 5087, 5295, 4614, 4594]. demonstrate [4621]. Demonstration [3564]. Dendraster [2274]. denizens [1130]. Denmark [794, 587]. Densities [2872, 3281, 4585]. Density [2839, 3169, 3472, 3376, 2332, 3459, 3680, 3389, 1551, 4864, 4542, 4977, 4968, 2547, 5182, 4973, 3814, 4185, 4597, 1397, 4009, 4340, 4874, 1317, 5145, 5172, 4636]. density-dependence [4636]. Density-Dependent [3169, 3459]. Density-independent [3459]. Dental [3137]. dentatus [3563, 5002, 3950, 5079, 4150, 4627, 4876, 3903, 2180, 4449, 2237, 2612, 5205]. dentex [4383]. Dentition [1890]. denying [40]. deoxyribonucleic [2764]. Department [840]. dependence [5061, 4636]. Dependent [3169, 4021, 2640, 5013, 4988, 5331, 4082, 5292, 3304, 4467, 5012, 3459, 3639, 5149, 4271]. Depletion [359, 5354, 3754, 1126, 427]. depositing [336, 335]. deposition [5265, 4454, 5373, 4647, 3130, 5272, 3106, 3226]. depot [94]. Depredation [4782, 3955]. Depredations [534, 719]. depressor [2133]. Depth [3762, 3185, 4417, 2946, 2548, 2846, 3707, 5137, 2624, 3389, 5178, 2809, 4100, 4390, 5015, 4423, 3473, 3948, 5371, 2742, 4584, 1024, 1225, 4555, 5012, 5316, 1567, 5131, 1992, 5185]. depth-constrained [5316]. depth-dependent [5012]. depths [4226, 5211]. Depuration [2493, 2570]. derelict [5145]. derived [3842, 4787, 4730, 60, 4428, 5171]. Dermasterias [2098]. Dermochelys [4847, 5352, 4653]. describe [4877, 4859, 4629, 4379]. describing [1673]. Description [1767, 3502, 2349, 3950, 3017, 2433, 4928, 213, 2773, 3126, 2102, 636, 1221, 1009, 2619, 3364, 3453, 241, 1057, 2279, 1179, 850, 321, 2459, 2626, 2646, 2709, 2886, 2969, 3170, 3129, 2337, 2393, 2559, 1320, 2192, 2421, 4019, 4198, 4659, 2577, 4869, 2103, 3670, 1802, 1253, 1208, 1672, 2070, 2800, 1446, 1031, 2606, 3988, 1238, 34, 2315, 3497, 2361, 3197, 15, 3880, 5324, 2899, 950, 1600, 1117, 2463, 1531, 1305, 980, 1259, 4547, 4352, 5128, 1768, 4337, 4837, 4915, 1109, 3753]. Descriptions [2017, 2231, 1123, 2201, 2249, 1571, 2757, 1194, 1186, 93, 1258, 3661, 2812, 2803, 4681, 945, 4038, 1233, 1584, 2758, 3516, 1262, 167, 1169, 1251, 1330, 3904, 2829, 1946, 3968]. Desert [2034, 227]. Desiccation [3344, 2760]. Design [2202, 3378, 2844, 2326, 5307, 5059, 3807, 4264, 4490]. designated [5224]. designs [4483]. desirability [720]. desire [590]. Desmarest [3604]. Desmodema [2606]. Desolventizing [2213]. Desor [1654, 2011]. desotoi [5351]. destroyed [778, 619]. destroying [801, 624]. Destruction [439, 369, 518, 1255, 10, 485, 460]. destructive [558, 404]. detail [4269]. detect [3744]. detected [1951]. Detecting [3679, 3655, 3797]. Detection [3391, 1651, 2850, 3796, 3132, 3736]. Determination [4893, 3183, 2986, 2120, 3654, 3588, 3849, 3405, 2560, 3966, 2537, 1789, 4978, 5278, 4124, 5337, 1857, 5195, 4431, 4537, 1672, 1820, 3548, 4631, 2511, 4633, 1777, 4616, 4320, 1459, 3995, 4224]. Determine [2016, 2503, 4643, 5194, 680]. Determined [3477, 2692, 3573, 2624, 2893, 4498, 4893, 4934, 5049, 3195, 4500, 4477, 4390, 2540, 4687, 4994, 5155, 4149, 4075, 4526, 4393, 4394, 4454, 4583, 4584, 3794, 4233, 5154, 3815, 4917, 4162, 3096, 4985, 1803, 5335, 4822, 5150, 4805, 4340, 4534, 3238, 3250, 4815, 4269, 4836, 5071, 4746, 3061, 3816]. Determining [1725, 2564, 1867, 3257, 1842, 2114, 3431, 2745, 5004, 3934, 1949, 4792, 3468, 5216, 4926]. deterrent [4938]. developed [5098, 4110, 107]. Developing [4602, 1177]. Development [1809, 2915, 5116, 2296, 2767, 2622, 3418, 3817, 2292, 5307, 1886, 1876, 3572, 3517, 3020, 2441, 4796, 5167, 3036, 5313, 3366, 2776, 3056, 2688, 3368, 3128, 2022, 3193, 3564, 2462, 1615, 2110, 2595, 3060, 2362, 2972, 1845, 2785, 2788, 2881, 2908, 2126, 2771, 2999, 2978, 3500, 3533, 2883, 2981, 3270, 2353, 2299, 1868, 2721, 2159, 2154, 2345, 5127, 3544, 3685, 3705, 2308, 2837, 2964, 2985, 2848, 1875, 2334, 2706, 4940, 65, 61, 2204, 2743, 5330, 3534, 1579, 2710, 2439, 2961, 4086, 2989, 3275, 1378, 1440, 3085, 4655, 4280, 2723, 1782, 1813, 1912, 4070, 3913, 3871, 5310, 3929, 4414]. development [1127, 1655, 1907, 1996, 3768, 3882, 3860, 4350, 1224, 3746, 3813, 4312, 2461, 4001, 1778, 1825, 2533, 3111, 1125, 1036, 1103, 1513, 1639, 1638, 1637, 120, 3516, 3789, 3576, 3791, 2813, 1463, 1479, 1494, 4869, 2676, 5158, 5157, 4234, 4714, 4839, 3203, 281, 4739, 4447, 4333, 4570, 4630, 1854, 4275, 3293, 3976, 3984, 1801, 14, 103, 2, 89, 185, 217, 657, 754, 4962, 2737, 1342, 4204, 3330, 1995, 4701, 3110, 3958, 3743]. Developmental [3848, 2781, 2032, 2078, 2115, 2357, 2266, 3480, 1862, 5300, 5034]. device [2591, 1672, 1376]. Devices [3615, 1403, 5039, 4090, 3803, 4504, 1421, 5166, 2795, 4611, 5026, 3788, 4110, 4491, 5185]. Devil [462]. Devilfish [3544]. Dexistes [4741]. DGGE [4335]. dhufish [4482]. diaconus [5303]. Diagnostic [2232]. Diagramma [4757]. diameter [4347]. diamond [1606]. diamond-back [1606]. diamondback [1511]. diaphragm [358]. Diaphus [5324]. Diatoms [2137, 2489]. Dibothria [956]. Dibothrium [972]. dibranchiata [2642]. Dichelaspis [1223]. Dichelopandalus [3194]. Diego [4141, 3666]. Diel [4845, 2342, 3874, 2530, 2671, 3185, 2860, 2521, 2946, 3658, 5022, 1979, 3698, 2424, 3674, 3181, 3143, 2572, 2624, 3291, 4238, 5234, 4639, 4883, 2702, 3810, 2846, 4971, 5130, 2526]. Diel-Depth [2946]. Diet [2293, 3228, 4160, 3858, 4952, 2267, 4443, 3029, 3668, 3436, 5314, 3703, 4697, 3316, 4650, 2941, 4682, 4574, 5139, 4248, 4654, 4860, 3505, 4936, 4989, 4501, 4963, 5375, 4478, 3224, 5257, 4026, 4625, 4121, 5336, 4792, 3954, 4502, 4256, 3782, 5289, 4469, 4178, 5343, 4402, 4814, 5173, 4921, 5090, 3287]. Dietary [2344, 4069, 2058, 4956, 5358, 3158, 4213, 5355, 4121, 5205]. Diets [2622, 2851, 2181, 3656, 4867, 4588, 4838, 3519, 5067, 4830, 3757, 2686, 3655]. differ [4385]. Difference [2906, 3386, 3292, 4623, 4053]. Differences [4478, 3517, 3898, 2952, 3199, 2155, 3808, 4798, 5312, 5372, 3825, 4175, 5029, 5111, 3655, 5234]. Different [3590, 2181, 2643, 4245, 2011, 2655, 3154, 2786, 2664, 4750, 4987, 4215, 1680, 4429, 550, 2388, 5188, 2686, 4572, 2425]. Differential [3710, 2235, 2264, 3642, 4268, 4407, 5137, 2227]. Differentiating [3777]. Differentiation [2413, 3237, 3383, 4645, 4778, 4739, 4457]. difficulties [461]. diffusion [3982]. Diffusivity [3506]. Digenea [1874, 2231]. digested [4727]. digestibility [4099]. Digestion [3351, 1917, 743, 4663]. Digestive [2222, 1563, 3554, 4247, 4275, 1335]. Digestive-gland [3554]. digging [2791]. Digital [2253]. digitization [3671]. dilute [2394]. Dimensions [3180, 2982, 3412, 1673]. Dimethylamine [2815]. diminutive [4984]. dimorphicum [4596]. dimorphism [2959, 3576, 3833]. dinoflagellate [3285, 2392, 4319]. Dinoflagellates [2575, 3307, 2587]. Diodon [2676]. diodontid [1967]. Diodontidae [2676]. Dioxide [2050, 2628, 2770, 2741]. Diplodus [2179]. diploproa [2597, 2945]. Dipturus [5341]. Direct [4311, 5129, 4290, 3815, 4162, 1757, 5195, 3971, 2818, 5179]. directed [4686]. directing [1434]. Directions [216, 316, 696, 15]. directive [1467]. Disadvantage [871]. disappearance [178, 11]. Discard [5308, 4147, 5129, 4226, 4614, 5051, 4902, 4951]. discarded [4685, 4027, 5019, 5250]. discards [4686, 4494, 4740, 4027]. Discerning [4146]. Discharge [3432, 4325, 2746, 698]. Discovery [2648]. Discrete [3292, 2624, 5157, 3712]. Discrete-time [3292]. Discriminant [2660, 4249]. Discriminating [5354, 4540]. discrimination [4190, 4002, 5288, 4125]. Discussion [410, 2688, 2154, 3853, 2533, 1656, 3987]. Disease [3453, 84, 2786, 2412, 1516, 1180, 653, 516, 1382, 1722, 2425, 1401]. Diseases [2615, 1166, 1973, 2699]. Disequilibria [3667, 3872]. disk [107]. dispar [4630]. Dispersal [2318, 2727, 3005, 4376, 4351, 5148, 1681, 2897]. Dispersion [3586, 3915]. dispersive [4599]. Displacement [2419, 2994, 3482]. displaces [4747]. Disposition [731]. disrupt [4296]. dissociated [1440]. Dissolved [2050, 2404, 1385, 1723, 2588, 1394]. Dissostichus [4614]. distance [4357, 4919]. distances [1681]. Distant [2123]. Distant-Water [2123]. distillates [1529]. distinct [3854]. Distinction [4346, 5244, 3248]. distinctness [3866]. Distinguish [3621, 613]. distinguished [4842]. Distinguishing [1574, 1889, 4897]. distributed [892]. Distribution [5044, 1759, 1836, 3078, 2300, 3999, 2568, 2767, 3502, 3122, 2202, 2016, 2373, 2524, 2895, 3464, 3213, 3150, 3381, 4669, 2847, 2530, 2484, 2138, 2228, 3420, 3157, 3441, 3312, 3546, 3280, 5237, 2019, 3578, 3945, 3204, 3207, 4313, 2128, 3664, 2022, 2376, 3352, 2391, 2316, 3178, 2131, 2196, 3851, 2592, 2199, 2763, 5327, 2367, 2715, 2946, 3380, 4899, 2493, 3794, 4233, 2765, 2256, 3125, 2593, 2768, 3407, 2000, 2771, 3698, 5020, 2958, 2499, 2616, 2129, 193, 269, 833, 2355, 3119, 3243, 3362, 3875, 3980, 2159, 2345, 3246, 4984, 2305, 3901, 1956, 2923, 2429, 3597, 2661, 1987, 2274]. Distribution [2259, 3281, 2545, 2663, 3528, 3637, 1507, 2800, 3425, 4614, 2332, 4201, 2073, 2224, 2059, 3076, 3964, 2855, 3034, 1860, 4716, 3585, 4996, 2237, 3587, 4634, 3389, 3302, 1884, 5072, 3682, 3166, 3534, 3120, 3618, 1808, 2250, 2018, 3485, 3132, 4667, 3371, 2198, 2095, 2315, 4535, 4228, 2083, 2674, 704, 1767, 1809, 1832, 3167, 3759, 4100, 4845, 1955, 4847, 1988, 3874, 3977, 4624, 4807, 1886, 1876, 5102, 4390, 4460, 79, 785, 3806, 1941, 3895, 3965, 5257, 1643, 4423, 5199, 3473, 3768, 3882, 3860, 3880, 4141, 5107, 5091, 5010, 3036, 4994, 1011, 1859, 1562, 1588, 2280, 4908, 2702, 4619]. distribution [138, 3948, 4673, 4995, 5216, 506, 1972, 3460, 5262, 4626, 3893, 4371, 5331, 2813, 2570, 4737, 1845, 5138, 5311, 4022, 3911, 1771, 1979, 5193, 3810, 4713, 4484, 5032, 1802, 1969, 768, 803, 935, 1985, 1965, 4548, 1868, 4294, 1993, 3877, 4677, 5302, 4512, 4895, 4033, 5030, 2471, 1232, 4825, 4555, 4742, 4344, 4598, 5052, 1598, 2846, 4950, 2260, 5309, 3155, 1340, 5137, 4339, 2748, 4489, 3802, 4679, 5178, 4874, 3740, 3852, 1790, 1971, 2797, 4241, 5277, 1652, 5172, 3223, 2428, 5038, 4134, 3836, 1992, 3758, 4422, 2526, 3958, 4017]. Distributional [3205, 1491]. Distributions [3566, 2633, 3579, 2137, 3688, 2396, 5040, 5365, 3811, 5340, 4750, 6, 2742, 3192, 4417, 4777, 4176, 4696, 4466, 4632, 4341, 5264]. District [448, 700, 357, 806, 1160]. Disturbance [2836, 4911, 5168, 4239]. ditropis [4774]. Diurnal [2621, 3171, 3386, 4789]. Dive [4555, 4662]. Dive-depth [4555]. diver [4646, 5276]. diver-harvesting [4646]. divergence [4048, 2813, 4433]. divergens [4304]. divers [5228, 4393, 4394]. Diverse [2663, 2665]. Diversity [4856, 2860, 3597, 4960, 4048, 4174, 5374, 4080, 4874, 5234, 4815, 4181]. Dividing [4919]. Diving [3660, 4586]. Dixon [1564, 1621]. DNA [4916, 4054, 4310, 4002, 4617, 4576, 4889, 4442, 4506, 4048, 4942, 3383, 5341, 4291, 5165, 4766, 4509, 5270, 4396, 4530, 4080, 4203, 4125, 5343, 4842, 4457, 4815, 4887, 4181, 4116, 4836]. DNA-based [4617]. DNA-specific [4887]. Do [4787, 4062, 4507, 3613, 338, 331, 4296, 4670, 256, 3734, 394, 4385, 4551]. Documentation [3123]. documented [4820]. Documenting [4193]. Does [4652, 4607, 4784]. dofleini [4961, 5110]. Dog [139, 160]. Dog-fish [139, 160]. dogfish [485, 4523, 5257, 5199, 1348, 2579, 5337, 1944, 4455, 4569, 5227, 1426, 1467, 4037, 5126, 4969]. Dollar [3568, 2274]. Dolly [5372]. Dolphin [2980, 3231, 3689, 2749, 3464, 3280, 3549, 4149, 3222, 3376, 3378, 3570, 2944, 2768, 2830, 2777, 2601, 3690, 3425, 2855, 3049, 3217, 2734, 2789, 4640, 3926, 4878, 3837, 2838, 3882, 3744, 3821, 3842, 4147, 4761, 3967, 3869, 3844, 2720, 2659, 2719, 2971, 4292, 4485, 3899, 2821, 3766, 3096, 3282, 2590, 3879, 3861, 2085, 2837, 3723, 3289, 3883, 3052, 2882, 2797, 3833, 3850, 3298, 4571, 2367, 2224]. dolphinfish [5379, 5240, 5148, 4487, 4336, 5043, 4879, 5185]. Dolphins [3234, 2811, 3593, 3186, 2196, 2996, 3240, 2049, 3137, 3274, 3703, 4160, 3797, 3711, 4946, 4658, 3717, 3771, 4159, 4367, 4965, 4259, 3865, 4097, 3736, 4099, 3250, 3952, 4187, 4782]. dome [5179, 1892]. dome-shaped [5179]. Domestic [2910]. domestication [1083]. dominant [3829, 4867, 1933]. dominated [2380, 3340]. Dominican [4743]. door [5131]. d’Oreille [289]. Dormant [2761]. Dorosoma [1937, 1844, 3357]. dorothy [4771]. Dorsal [3397, 5337, 4221, 3301]. dorsal-fin [5337]. dorsalis [5312]. Dosidicus [5243, 4704]. Double [3027, 4088]. double-tagged [4088]. Double-Tagging [3027]. Douglas [1464]. Dover [4364, 2937, 3721, 4417, 4905, 3725, 2584, 2545, 2665, 3551]. Dr. [241]. drag [4691]. Dragonflies [1499]. Drainage [2918]. drainages [5191]. drawn [479]. dredge [4677, 4342]. Dredged [2018, 1251]. Dredging [2312, 170, 2443, 4747, 1170, 4646, 4257, 832, 1995, 3720]. Dresden [410]. Dried [2061]. Drift [3347, 3800, 4034, 1811, 3305, 4029, 5092, 3090]. driftnets [3831]. drill [1874, 1628]. Drills [2024]. drivers [5170, 5344]. driving [5145]. droebachiensis [4602, 4712]. Dromiidae [2785]. drop [2636]. Droughts [2667]. drug [1777]. Drum [3698, 3607, 2803, 3139, 3152, 326, 4948, 3501, 4103, 4145, 2519, 3898, 4462, 3813, 3785, 3866, 5367, 2930, 4168, 4418, 4914, 5022, 3815, 4162, 4474, 5052, 4262, 4318, 3928, 4049, 3909, 4276]. drum-fish [326]. drums [1638]. Dry [2419, 80, 2994, 3482, 4955]. Dryfoos [2330]. Drying [2213, 774, 340]. Dual [2512, 5306, 5242]. dual-season [5306]. dual-time [5242]. dubius [3388, 3890, 3594]. Duboscquodinida [3307]. Due [2858, 90, 4829, 4911, 906, 4980, 968, 4419, 1135, 4753]. dumerili [4196, 5371, 4241, 4667]. dumping [3834]. dumpsite [3834]. Dunaliella [1898]. Dunbar [1057]. Dunes [634]. Dungeness [3963, 2420, 3009, 3014, 2451, 2553, 2470, 2947, 3340, 2850, 2932, 2251, 3428, 3410, 3166, 4662, 3233, 2714, 5250]. duorarum [1941, 1981, 4756, 1862, 3026, 3737, 1993, 4047]. Duration [2437, 4761, 2545, 2694, 4050, 5003, 4868, 5104]. During [3269, 3536, 2099, 2352, 3420, 2314, 2445, 3368, 2148, 2956, 2580, 2771, 2011, 2556, 2178, 269, 2031, 3436, 3143, 3442, 2053, 3252, 2641, 2313, 2756, 3574, 3993, 5039, 3999, 4309, 1839, 936, 4070, 3913, 1528, 2764, 5170, 3965, 4391, 364, 3057, 839, 3956, 4761, 2041, 5099, 448, 3408, 4652, 4285, 954, 835, 834, 1469, 3787, 886, 344, 4367, 958, 3404, 4368, 680, 4775, 3851, 941, 973, 4371, 1086, 4956, 3288, 961, 3025, 3994, 4713, 4234, 4862, 3782, 586, 833, 837, 4199, 4071, 974, 991, 1072, 5191, 1015, 2550, 4589, 626, 597, 3703, 4886, 818, 4789, 5160, 2953, 5108, 107]. during [470, 4469, 3776, 5366, 5358, 4931, 1960, 77, 765, 867, 909, 1567, 773, 3297, 4300, 939, 963, 865, 4682, 5277, 4302, 4636, 934, 4986, 4270, 4051, 2561, 5040]. Dusky [2789, 4160, 4964, 5115, 5280, 3899, 3936, 4648, 4377, 3833]. dwarf [4726]. Dye [3605]. Dynamic [2273, 2408, 4577, 3749, 5097, 3847, 4049]. Dynamics [4558, 2092, 2401, 2730, 3609, 2990, 2857, 1942, 3692, 4395, 2470, 2866, 2935, 3152, 3568, 2537, 4498, 4622, 4705, 3292, 4018, 3853, 5214, 4561, 4562, 4709, 4212, 5015, 5379, 4148, 4092, 5367, 5294, 4876, 4234, 5198, 4676, 4071, 5141, 3960, 4222, 4661, 4157, 5291, 3764, 4875, 3873].

E. [2922, 2132, 3631]. each [1284, 941, 973, 974]. eagle [5054, 227]. ear [1766, 1416]. earlier [771]. Early [3563, 3975, 1782, 2296, 2995, 3474, 2349, 3513, 3517, 2032, 3420, 3388, 1862, 3424, 4827, 3700, 1825, 5151, 3582, 3611, 2922, 3144, 3222, 2110, 2571, 2574, 2790, 3789, 2294, 2577, 3656, 2126, 3270, 2798, 219, 3320, 2641, 1509, 2870, 3802, 3745, 2710, 3493, 4702, 3110, 4303, 5163, 3574, 3085, 2005, 5255, 3505, 5348, 4070, 3950, 3977, 3817, 4928, 4796, 4749, 1625, 2461, 4606, 3463, 4795, 3170, 3129, 2546, 4988, 3516, 3904, 4184, 5154, 3692, 4993, 4788, 1866, 4839, 3725, 3782, 330, 3914, 4334, 4115, 3722, 3643, 1653, 4104, 3786, 4789, 2861, 3824, 2766, 4063, 5268, 1116, 207, 5277, 53]. early [2065]. Early-Juvenile [3611, 3656, 4749]. early-life [5348]. Early-life-history [3700, 3725, 3722, 3786]. ears [1426]. East [4418, 3041, 1839, 4282, 4412, 1481, 3866, 3954, 5276, 4162, 4317, 3922, 4355, 4514, 3995, 4224, 3593, 2318, 2715, 3766, 3556, 4422]. east-central [3866, 5276, 4162, 3922]. Eastern [3704, 3558, 3561, 2601, 3690, 3753, 3620, 4655, 1809, 5212, 4640, 3377, 2726, 2862, 4190, 1988, 2373, 5053, 3381, 3711, 4191, 3729, 4212, 4642, 4935, 4038, 5365, 1131, 4605, 3717, 5010, 1675, 3771, 3821, 4147, 1576, 4111, 3869, 4149, 2702, 2738, 4764, 4774, 4821, 4930, 4367, 4369, 4770, 4892, 4287, 4370, 2104, 2546, 5360, 4907, 5046, 4371, 2321, 2971, 5240, 2799, 4823, 1082, 3040, 2327, 4118, 4176, 4397, 4932, 5141, 4374, 3814, 4487, 5201, 4027, 4259, 2471, 2590, 3879, 3861, 4825, 4924, 3865, 3883, 4097, 952, 3390, 4007, 4420, 4531, 4851, 3905, 5330, 3867, 5359, 1076]. eastern [3302, 4015, 5072, 5268, 5078, 4665, 3808, 3966, 4269, 5173, 4302, 3850, 3952, 3298, 1527, 5025, 527, 3287, 5169, 2044, 2191, 3231, 3477, 3385, 2847, 2530, 2231, 3182, 2143, 3349, 2705, 3038, 3222, 3376, 3378, 3570, 2571, 2574, 2269, 2214, 2768, 3240, 3274, 2145, 2473, 2601, 2800, 2332, 2059, 470, 2160, 2431, 4299, 3049, 3589, 2710, 2037, 2336, 3217, 2518, 2151, 2982, 3481]. eastern-central [4764]. Eastport [920, 598]. Eastropac [2191, 2044]. Eastropic [1839]. eat [432, 222, 331, 394]. eaten [3052]. eater [488]. eating [47, 430, 533, 374]. Eberbach [5335]. ecdysis [2480]. Echeneibothrium [3790]. Echinarachnius [3568]. echinata [1965, 1963]. Echinoderm [1745]. Echinodermata [2268]. echinoderms [1214, 1248]. echo [5071]. echogram [4897]. echosounder [5166, 5229]. Eckhardt [303]. eclipse [1960]. Ecological [4719, 2046, 2122, 2210, 1665, 1518, 5125, 5161, 3365, 3249, 2558, 2378, 4315, 1464, 4872]. Ecology [2466, 2524, 2225, 2301, 2484, 3004, 3611, 2252, 2993, 2360, 2464, 3695, 5364, 2324, 3290, 2828, 3684, 3686, 2522, 2734, 2789, 2373, 4989, 4963, 4413, 3900, 3768, 5225, 4369, 4627, 5292, 5186, 4484, 2650, 5037, 1856, 3153, 5289, 2864, 4144, 2817, 3103, 3369]. econometric [3423]. Economic [3483, 5202, 2479, 2910, 5121, 1345, 1006, 3154, 2371, 2363, 1515, 1061, 3040, 1476, 1490, 1522]. economical [159, 1129]. economy [4627]. Ecosystem [5036, 3094, 2672, 2979, 5040, 4864, 5101, 4505, 4066, 4867, 3206, 5126, 5045, 4525]. ecosystem-based [5045]. Ecosystems [2328, 2135]. Ecotypic [4823]. Ectoparasitic [3452]. edeni [3277]. edge [641, 3747]. Edible [627, 221, 645, 368, 1223, 563, 57, 197, 237, 187]. Edinburgh [102, 68, 142]. Edisto [536]. Editor [2161]. education [2142]. educational [1399, 1410, 1398]. edulis [120, 175, 3097, 2386, 1515, 2746, 200, 2754, 3627, 3323]. Edw [3107]. Edwards [3569, 2207, 2115]. edwardsii [2740, 4657]. Eel [3281, 1224, 1235, 60, 321, 3168, 88, 3908, 5140, 532, 3006, 4000, 591, 746, 1087]. eel-like [321]. Eelgrass [3474, 3490, 3350]. eelpots [665]. Eelpout [2611]. Eels [665, 825, 122, 543, 520, 708, 780, 4488, 604, 3291, 510]. Effect [2298, 2622, 3343, 5102, 1723, 4426, 2694, 2588, 2038, 2213, 2009, 3811, 4734, 1952, 3111, 407, 2994, 4858, 2058, 2045, 2126, 1680, 2106, 2353, 3340, 687, 1936, 2984, 1797, 5063, 707, 1881, 2025, 2697, 3005, 2170, 2400, 3673, 2647, 2266, 3217, 1652, 5038, 2794, 3452, 5169, 4599, 3167, 228, 2744, 2764, 355, 4760, 4580, 4404, 276, 1665, 4564, 3608, 4979, 4754, 5003, 1770, 4566, 1394, 1447, 2754, 5242, 1047, 1135, 4635, 5256, 4752, 2096, 2113]. Effective [4970, 5028, 4762]. Effectiveness [2853, 2652, 4110, 1681]. Effects [5296, 2839, 5142, 5348, 4747, 3397, 3831, 4910, 4036, 4347, 3603, 4878, 4522, 3513, 4896, 5320, 2472, 2344, 2441, 1996, 2877, 4808, 2404, 2965, 2185, 2407, 2234, 2819, 3549, 3842, 2328, 3133, 3492, 4977, 3869, 2243, 2779, 3472, 3285, 3700, 4001, 1642, 4939, 3331, 4960, 4692, 1356, 5105, 2131, 3427, 3186, 2746, 2930, 1632, 3524, 2863, 3241, 2761, 2312, 1360, 2495, 2814, 2418, 3403, 5209, 4724, 5100, 2350, 4429, 5221, 3581, 3500, 1856, 2051, 2372, 2353, 5339, 4227, 3339, 5082, 2655, 3181, 1830, 4118, 5295, 5317, 4981, 1791, 2861, 3047, 5034, 2457, 2135, 2204, 5210, 2760, 1397, 4180, 3400]. Effects [1787, 4915, 2531, 5078, 2916, 3726, 3344, 2118, 4578, 4636, 2905, 2480, 4718, 4378, 3085, 3830, 4622, 2945, 4938, 4210, 4211, 5214, 2693, 3134, 1907, 678, 4780, 3773, 3785, 3465, 4050, 4819, 4909, 4169, 4866, 5018, 4315, 4257, 5125, 5035, 5088, 725, 5168, 5097, 4295, 1762, 4611, 4885, 1854, 5246, 3796, 4142, 5008, 2406, 5241, 4615, 4041, 5000, 1317, 1995, 3720]. Efficacy [5284, 4613, 5000]. Efficiencies [3543, 2133]. Efficiency [2046, 2931, 2025, 3707, 4491, 4875, 4499, 4961, 2306, 4959, 3070, 4419, 3338, 4265, 4450, 4841, 4342]. Efficient [2640, 3499]. Effluent [2352]. Effluents [2503, 2564, 2502]. Effort [2892, 3697, 2093, 3415, 2482, 3425, 3676, 2447, 2535, 5129, 3681, 4215, 3906, 5267, 2176, 4419, 173, 1964]. Efforts [416, 672, 90]. Egg [936, 3513, 3517, 3646, 2553, 2842, 3579, 3188, 2004, 2053, 2848, 3976, 2334, 3602, 1759, 3779, 4542, 1224, 3140, 4692, 3608, 2930, 4094, 5319, 2676, 1177, 4637, 4444, 4567, 5088, 4647, 3340, 4855, 3859, 4597, 51, 107, 89, 2760, 1114, 5231]. egg-fungus [89]. egg-size [4637]. Egg-to-migrant [2004]. Eggs [1773, 1929, 2063, 2839, 2349, 2618, 2524, 2619, 2314, 3366, 3662, 3582, 2337, 2393, 2761, 3237, 2367, 3266, 3500, 2299, 3685, 2820, 2832, 2890, 2766, 2237, 3531, 2989, 1767, 1836, 1832, 410, 125, 684, 451, 391, 227, 2292, 308, 333, 785, 3715, 3878, 726, 1907, 4028, 5010, 389, 539, 1859, 1588, 2451, 731, 4001, 3664, 3129, 1513, 4668, 478, 3893, 1408, 1845, 5165, 4913, 419, 268, 566, 768, 876, 1901, 263, 4294, 707, 818, 947, 4768, 4742, 234, 1875, 4360, 307, 236, 314, 3986, 94, 3511, 642, 1652, 568, 730, 3958, 682, 1390, 108, 80, 704]. egress [5233]. Ehrenberg [4285]. Ehu [3680]. Eigenmann [2159]. Eight [2982, 555]. Eighteen [1382, 299]. Elasmobranch [5153, 3147, 5333, 5328, 5037, 5026]. Elasmobranchii [5219]. elasmobranchs [4987, 4533, 1335]. elassodon [4742]. electra [2471]. Electric [1594, 422, 2352, 611]. Electrical [4953, 2323, 2025, 2326, 4216, 4171, 1881]. electroencephalographic [2242]. electron [3159, 3907, 4599]. electron-probe [3907]. Electrophoretic [3463, 2117, 2320, 2492, 3103]. electropositive [4938]. Electrotaxis [4184]. eleginoides [4614]. Eleginus [5244]. Elemental [2992, 4672, 4882]. Elements [2156, 4190]. Elephant [1657, 2926, 3363, 3994]. Elevated [2383, 678, 4971]. Eleven [3702, 4348]. Ellobiopsidae [2361]. elongatus [5363, 3641, 3348, 3763, 4431, 5045, 3816]. Elops [1825, 4430]. elsewhere [1150]. Elton [2182, 2171, 2166]. elucidated [4445]. elvers [2403]. emaciated [2936]. emarginata [2622, 2595]. Embassichthys [2890]. embayment [3080, 4800]. embayments [5137]. Embedding [5242]. Embiotocidae [2613, 4812]. embiotocoid [696]. Embryo [2776, 5158, 4538, 65, 185]. embryogenesis [4931]. Embryological [1854]. embryology [1463, 1472, 1479, 1494, 71, 960]. embryonated [94]. Embryonic [3871, 2776, 3056, 2353, 3612, 1519, 1996, 4312, 3111, 4869, 4234, 4333, 5300, 5034]. Embryos [3572, 2988, 2115, 2967, 2457, 3727, 2292, 2340, 696, 2458]. emerge [2728]. Emergence [3612, 2852, 3466]. Emerging [2442, 4874]. Emerita [2643]. Emigration [3569, 2026, 4660]. emperor [4757, 5184]. emphaeus [4158]. Emphasis [2605, 3126, 2196, 3403, 3954, 3153, 2260]. Empirical [3118, 4870]. employed [1032, 1393, 300, 1091]. employment [943]. empusa [2721, 2864, 3155]. enable [5303]. enables [4673]. Enchelyopus [2722, 4213]. Enclosure [3327]. encompass [4507]. encounter [4768]. Encountered [3455, 5305]. encounters [5069]. Encrasicholina [3346, 3517]. encrasicholus [2356]. encrasicolus [4002, 5242, 3589, 3716]. Encroachment [2106]. end [5116, 4484]. endangered [4801, 4132, 5315, 5297, 4373, 5298, 5143, 5188]. endemic [5007]. ending [1158, 317]. endocrine [2087]. endoparasitic [4834]. endure [260, 675]. enemies [851, 491, 1166]. enerasicolus [4209]. Energetic [3492, 2756]. Energetics [4103, 3771, 3524, 2907, 2431, 4049, 2338, 3821, 3369, 2905]. Energy [2992, 2046, 3054, 3284, 2929, 2567, 3511, 4089, 4977, 3890, 4692, 4004, 5047, 2836, 5034, 5224]. Enewetak [2632, 2816]. engaged [448]. engine [562]. England [1176, 341, 639, 2978, 3102, 631, 898, 4478, 3259, 139, 987, 4946, 5313, 1221, 7, 441, 35, 70, 367, 1300, 1170, 2607, 1439, 5311, 961, 2351, 3971, 500, 19, 20, 2305, 5227, 3415, 5314, 1816, 1618, 3245, 2578, 1260, 3267, 1790, 5205, 23, 84, 3758, 589, 605, 637, 701, 738, 755, 761, 772, 781, 789, 792, 814, 823, 830, 841, 846, 860, 769, 783, 853, 5085]. English [655, 3862, 663, 4094, 3244, 360, 3476, 2953, 2970, 4065, 4204, 245, 4861, 1951].

Engraulidae [2196, 2435]. Engraulis [2497, 4002, 3691, 3694, 2469, 3414, 3536, 2673, 3324, 2356, 3521, 4368, 2147, 2153, 2449, 2521, 2759, 2842, 2907, 3187, 3196, 3919, 2493, 2885, 2840, 2139, 2152, 2062, 4637, 3188, 3266, 2132, 3310, 2733, 3119, 4294, 1766, 2833, 2936, 3525, 2818, 3235, 3304, 3305, 3796, 2861, 2260, 2870, 5242, 2476, 3022, 2575, 2587, 2175, 3531, 4209, 3589, 3716, 2858, 3369, 3538, 2348, 2338, 2902, 2943, 2756, 2809, 3523, 5231]. Enhance [3653]. Enhanced [2394]. enhancement [4627, 4044, 5121, 3233]. Enhancing [4178]. Enhydra [4708, 2631]. Enoploteuthis [3017]. ENORMSEP [2396]. enough [4504]. ensifer [4337, 5163]. Entanglement [3492, 4187, 4067, 3831, 4050]. entering [4273, 4488]. Enteroctopus [4961, 5110]. Enteropneusta [1742]. enters [1957]. entire [5300]. entitles [210]. entomelas [3013, 2908, 4003, 3838, 3275, 4270]. Entosphenus [5134]. entrance [4690, 4767]. entropy [5177]. enumeration [3671]. Environment [3032, 2901, 2297, 3537, 2262, 2514, 2165, 3016, 5312, 5167, 1938, 1030, 4098, 1936, 1338, 2087, 1787, 4051]. Environmental [1918, 5240, 5191, 4965, 2655, 2101, 2239, 3848, 5348, 4624, 4211, 3134, 5133, 5170, 1643, 5262, 4820, 5138, 4547, 2680, 5242, 4872, 5038]. Environmentally [3892, 4623]. Environments [2643, 3525, 2020]. Enzyme [2075]. enzymes [1563]. Eopsetta [3892, 4526, 5292, 2545]. ephippiatus [2247]. Ephippidae [3880, 3577]. Epibenthic [2467, 2515]. epibiotic [2451]. Epibulia [2121]. epidemic [946]. Epinephelidae [5115, 5007]. Epinephelinae [5032, 3675]. Epinephelini [5032]. Epinephelus [4281, 4357, 3729, 5115, 4332, 5190, 4758, 4080, 3742, 4815, 4323]. epipelagic [4830, 5017, 2742]. episodic [3776]. Epizoites [2448]. Epizootic [3318]. eponym [3535]. Eptatretus [2899, 4220, 3631]. Equation [3032, 3035, 4053, 3944, 3338]. equations [4844, 3482]. Equatorial [2391, 1839, 1858, 4531, 4851, 1687, 1783, 2169]. Equilibrium [2512, 2354, 2223, 2997]. Equipment [2041, 1478]. equipped [4684]. equiselis [3882, 2837, 3250]. equivalent [3720]. era [3714]. erectus [4600]. Erie [459, 1937, 656, 583, 1625, 1192, 1688, 1021, 1232, 1250, 1142, 1934]. Erimacrus [3377]. Erosion [3184, 2584]. errans [3340]. Errata [4496]. Erratum [3626, 4023, 4072, 4108, 4439]. error [4999, 4232, 4527, 4175]. errors [4980, 4979, 4631]. ERTS [2294]. ERTS-1 [2294]. erumei [5346]. erythrinus [5255]. Erythrocyte [2208]. erythrocytic [2748, 3165]. Escambia [954]. Escape [2652, 2107, 3344, 3831, 4210, 4504]. Escapement [3494, 4690, 2091, 3271]. escapements [1864, 1998]. Eschrichtius [2069, 2542, 3087, 3066, 2639]. escolar [4218]. Esox [253, 5265]. especially [775]. Espiritu [4595]. essential [4643]. Establish [2326]. Establishing [5086, 457]. Establishment [596, 1119, 2918, 4911, 163, 176, 4719, 576, 888, 880, 1379]. Estimate [2031, 3659, 4873, 3741, 4933, 4064, 3118, 4693, 4870, 3807, 4216, 5166, 4161, 4597, 4727, 4119]. Estimated [3800, 4034, 2875, 4147, 2653, 3119, 2833, 4467, 2920, 5107, 5305, 3869, 5326, 4016, 4510, 4740, 4696, 4466, 4377, 4793, 5040]. Estimates [4250, 3374, 4601, 4794, 2046, 5007, 4470, 3549, 3744, 4079, 4152, 3586, 2496, 4088, 3530, 3325, 3587, 3680, 3568, 2938, 1838, 3341, 4746, 3961, 2740, 4640, 4281, 5004, 5202, 1998, 4801, 4106, 1909, 5346, 3542, 4404, 4710, 4728, 3842, 4230, 4730, 4673, 4832, 4826, 4427, 4816, 4553, 5059, 4393, 4394, 4607, 3739, 4170, 5159, 5368, 3936, 5063, 2581, 5246, 4359, 4005, 5033, 4298, 4009, 5192, 3792, 4573, 4699, 4663, 4700, 1964, 3952]. Estimating [4403, 5119, 1864, 3711, 3385, 3467, 4781, 4304, 4712, 4428, 4545, 2629, 3613, 3376, 3804, 3356, 4068, 2971, 4660, 3475, 3991, 3730, 3313, 4031, 4768, 4456, 3678, 5055, 4662, 2037, 3493, 4117, 5057, 4848, 5050, 5174, 4846, 3898, 4252, 5340, 4863, 4167, 4472, 2651, 4723, 4777, 4483, 4429, 3647, 4353, 2818, 3918, 3286, 4027, 3206, 3982, 4550, 1785, 4806, 2679, 5000, 4342, 1761, 3998]. Estimation [3444, 3447, 3035, 3644, 2997, 4581, 3081, 2776, 2074, 3706, 3709, 2272, 3857, 1726, 3116, 2866, 1693, 5239, 2482, 2670, 4261, 1686, 4263, 2644, 5051, 3421, 3850, 2714, 4225, 4310, 5334, 3864, 4734, 2957, 3844, 4691, 4585, 4502, 4444, 5370, 3818, 4399, 3738, 4238, 4923, 3828, 3571, 4717, 3353, 4840]. estimator [5218, 5316]. estimators [4133]. Estuaky [2893]. Estuaries [3563, 2206, 3455, 2156, 4462, 3057, 4941, 3812, 4351, 4430, 4610, 4568, 4739, 4155, 3903, 4811, 4375, 4065, 4918, 4514]. Estuarine [3269, 2185, 3246, 2970, 2295, 2218, 2534, 4557, 4856, 3963, 4100, 5299, 4997, 5016, 1938, 2489, 5080, 4121, 2636, 4945, 4150, 3244, 3908, 5198, 2650, 4139, 3920, 4950, 4572, 1893, 4802, 4666, 5130, 4571]. Estuary [3569, 2983, 2608, 3023, 2125, 1891, 3237, 3008, 2103, 3436, 3166, 3448, 2831, 2095, 3975, 4327, 2245, 3037, 4129, 5079, 3086, 2325, 3781, 5290, 2379, 4675, 4082, 3059, 4943, 4331, 4567, 5104, 3117, 3340, 2538, 3784, 3323, 4488, 3296, 5012, 3476, 4262, 4037, 3291, 5203, 2018, 4701, 4759, 3873, 5062, 5134, 2998, 2755, 2761, 2360, 3658, 4484, 4679]. etc [267, 889, 456, 239, 30, 475, 597, 486, 236, 1218, 317, 280, 427, 370]. Eteline [3552, 5086]. Etelis [4252, 3552, 4721, 3680]. Ethanol [3582]. Ethanol-preserved [3582]. ethylcarbazole [1948]. Etiology [3430]. Etropus [2244, 2766, 2961]. Etrumeus [3713, 2546]. Eualus [2922]. Eubalaena [3422, 4895, 3021]. Eucalanus [2261]. Euchaetidae [1776]. Euchone [2116]. Eulamia [1850]. Eumetopias [4706, 4730, 4464, 5114, 4586, 3359, 3989, 2927, 4825, 5024, 5366, 5358, 4665, 4838, 4551, 4664]. Eumicrotremus [5336]. Euphausia [2527, 2068, 2694, 3422, 2362, 2813, 2035]. Euphausiacea [2333, 2410, 2207, 2362, 2813]. Euphausiid [2962, 3419, 2068, 4022]. Euphausiids [3522, 2526]. Eupleura [2024]. Eurasian [4842]. Eureka [4284]. Europe [851, 1056, 913, 940, 197, 630]. European [4002, 4720, 889, 1004, 4524, 60, 4008, 5124, 5242, 4209]. Euryhaline [3686]. Euterpina [2655]. Euthynnus [3643, 2099, 2891, 2454, 1689, 3390, 1670]. Eutrophic [2252]. evacuation [3351]. evaluate [4870, 5350]. Evaluating [5057, 5189, 2660, 5373, 4613, 5316, 4806, 4873, 3825, 1768, 5250]. Evaluation [4961, 3590, 4560, 2202, 5050, 5174, 5101, 5066, 3937, 4183, 5371, 2954, 5056, 4216, 3769, 5287, 3653, 3445, 2384, 2255, 2289, 3275, 3350, 4276, 4075, 4472, 3721, 5215, 1959, 3040, 5052, 4261, 4532, 5000, 3928, 1898, 3907, 2257, 3393].

Evaluations [4090]. Evasive [4367]. even [4398]. even-broodyear [4398]. Evening [2148]. Event [3331, 4051]. Events [3504, 5057, 4692, 4993, 4956, 5078, 4767]. Everglades [4598, 4632, 4463]. Evermann [4881]. evermanni [3248, 3287, 4017]. Evidence [2179, 3388, 3584, 3854, 4684, 2751, 4932, 4904, 4811, 3162, 5312, 4719, 4073, 4871, 3276, 4255, 5114, 2370, 2492, 4096, 5175, 5128, 4035, 3159, 3712, 2348, 5130, 4514, 5179, 4703, 4515]. evident [5295]. eviscerated [3306]. evolans [3237, 4465, 4517]. Evolution [3939, 2154, 2291]. Evolutionary [4916, 4850, 4790]. evolving [4649]. Evorthodus [1630]. Ex [3267]. Ex-Vessel [3267]. Examination [2281, 4629, 3861, 622, 4856, 872, 3140, 3608, 680, 4382, 4019, 4198, 2682, 2867, 4530, 3988, 365, 977]. examine [3138]. examined [973, 974]. examining [5375, 4933]. Example [2298, 2263, 2439, 5140, 4547, 5330, 4356]. examples [3897]. excellent [533]. excentricus [2274]. exception [4594]. exceptional [4920]. excess [2697]. Exchange [2383, 419, 279, 588, 1730]. excisa [2604]. excluded [239]. excluder [5039, 4504, 3788]. Exclusive [5202]. excrescences [968]. Excretion [2931, 2132, 2653, 4049]. excretory [1923]. Executive [545]. Exercise [2055]. exhibit [1424, 1399, 1410, 1398, 325]. exhibited [1528]. exhibiting [5061, 4750]. Exhibition [102, 1417, 938, 17, 302, 68, 142, 325, 545]. exilis [2665]. eximia [2813]. exist [381]. existence [2348, 40]. Exocoetidae [4303]. expansion [5236, 5259, 1380]. expectations [4897]. expedition [1839, 1195, 1170]. experiements [3830]. Experience [244, 479, 1382]. Experiment [2294, 3261, 3322, 4311, 1572, 4079, 3754]. Experimental [3606, 2380, 3489, 3472, 1592, 1637, 1799, 2411, 2387, 3696, 3885, 5286, 4522, 1609, 4674, 4101, 4614, 1961, 1898]. experimentally [4319]. experimentation [4315]. Experimentel [917]. Experiments [2386, 858, 273, 1524, 3709, 3327, 1520, 519, 1346, 104, 3313, 1596, 4802, 2326, 1172, 3027, 675, 1037, 4794, 4391, 1758, 2041, 1606, 1726, 4982, 281, 298, 5354, 1578, 1583, 1605, 1627, 952, 1686, 1607, 3061, 548]. expert [4982]. Explanation [554, 716, 737, 970, 976, 1001, 1005, 1060, 1067, 1101, 1111, 1138, 1156, 1168, 1234, 1257, 1298, 1316, 1329, 1333, 1365, 1432]. Explanations [951, 2370]. explicit [5015, 5215]. exploitable [4712, 3750, 3754]. Exploitation [2912, 2240, 2135, 2170, 2411, 2387, 5054, 5164, 4250, 5189, 5332, 3381, 3876, 3779, 3089, 2169, 5295, 1799, 1961]. Exploited [2263, 3116, 3636, 2431, 4991, 4174, 5219, 5239]. Exploration [3371, 717, 5283, 1109]. Explorations [939, 1735, 1079, 954, 941, 973, 1087, 974, 1072]. Exploring [5375, 4779, 207]. exponential [3151]. Export [283]. exportation [268]. exported [364]. Exports [839, 380, 317]. Exposed [3572, 2900, 3932, 1757, 4803]. Exposition [251, 2176, 929, 1025]. Exposure [3612, 3344, 4910, 2965, 3241, 4862, 5235, 4951]. Exposures [2852]. expression [3338]. extended [1151, 4432, 4051]. Extension [1075, 5340]. extent [875, 1153, 1587]. External [3690, 3576]. Extractable [3426, 4565].

Extracted [2038]. extraction [2683, 4578]. Extracts [457]. Extraordinary [91, 4302]. extrinsic [5344]. Extrusion [5260, 3140]. Exxon [4130]. eye [4018, 4275]. Eyes [2854, 1426].

F [4508]. F. [5007, 4778]. faber [3880, 3577]. fabricii [2922]. Factor [3627, 5050, 4101, 1951]. Factors [1552, 5053, 3216, 2099, 1694, 4675, 1626, 3588, 3403, 2845, 3076, 1804, 3508, 3657, 2015, 1760, 2987, 5145, 1918, 5348, 4748, 5091, 4828, 4933, 5371, 5283, 5294, 5138, 1147, 2557, 4375, 4663]. facts [727]. FAD [4611]. FADs [4611]. failure [1359]. Fairport [1478, 1499, 1530]. falcatus [4461]. falciformis [4552]. Falkland [4685]. Fall [3578, 2131, 2332, 2627, 3999, 3756, 3923, 3480, 2412, 4069, 1972, 2006]. fall- [4069]. Falling [649]. falls [289]. false [3832, 2801]. Fam [1630]. families [1776, 1186]. Family [2014, 3727, 2102, 3768, 2143, 2117, 2881, 2908, 3008, 3243, 2159, 2154, 2345, 3610, 2800, 2706, 2743, 2731, 1988, 1886, 4192, 4796, 4026, 3882, 3880, 1793, 1638, 4019, 4198, 4659, 3208, 3988, 2737, 4681, 1671, 4535]. Fantail [3685]. Farallon [4464, 2873]. Farallones [4484]. Farfantepenaeus [4643, 4756, 5009, 4778]. farming [1143]. fasciatus [2803, 3139, 4339, 3152]. fast [1469]. faster [4062]. Fat [2407, 3190, 1471, 1469, 2495, 5006, 3288, 3306]. fat-absorbing [1471]. Fathoms [2235, 2227]. fattening [1495, 546, 733]. Fatty [2413, 2058, 5067, 5006, 4931]. Fauna [3677, 2478, 1443, 1134, 999, 1307, 1165, 1339, 1423]. Faunal [2424]. faustinum [3089]. Fear [2747, 2831]. Feasibility [2910, 3827, 3971, 1114]. Features [2554, 2442, 2111, 2250, 4860, 921, 3576, 2565, 1975, 5206]. February [2178, 533, 841, 916]. fecal [4501, 4813]. feces [2856]. fecundation [162, 175, 797]. fecundities [4397]. Fecundity [3013, 2625, 3068, 5281, 3346, 3201, 4687, 3646, 2917, 2729, 2949, 3368, 1852, 2607, 2759, 3721, 4094, 3498, 1786, 4647, 2510, 3242, 4610, 2796, 3295, 3705, 1788, 3665, 4656, 4165, 4843, 4060, 4846, 3809, 3779, 4542, 5304, 3813, 5061, 4764, 4675, 5319, 5292, 4161, 4331, 5088, 3218, 4818, 2792, 3304, 5034, 5160, 4007, 4575, 4926, 4422, 4119, 4102]. Fed [2441, 3656]. federal [5052]. Fedorov [5336]. fedorovi [5336]. feed [338]. Feeding [2417, 4989, 3065, 3603, 4963, 2631, 4720, 4413, 2222, 2608, 2437, 2671, 2963, 4195, 3224, 3004, 2931, 3054, 2937, 5176, 2216, 2623, 3611, 4150, 2966, 3422, 3577, 2254, 2335, 2993, 2414, 2153, 2452, 3380, 2360, 2464, 2885, 4153, 2397, 2580, 1770, 2638, 2011, 2499, 3153, 2051, 3899, 2921, 3136, 2550, 3674, 3181, 146, 2140, 3392, 3087, 3084, 2180, 2427, 1083, 2665, 3143, 4726, 3553, 2889, 1905, 2817, 2828, 4434, 2916, 5205, 3369, 4309, 2497, 4798, 3150, 4938, 4129, 4478, 3713, 1523, 2878, 5132, 2573, 5338, 4121, 1648, 2555, 4274, 2147, 4584, 3367, 3050, 1518, 4494, 1680, 4235, 5037]. feeding [1392, 2278, 2711, 4101, 2003, 185, 3209, 3843, 3471, 4716, 1396, 4299, 5203, 4300, 3332, 2639, 3067, 5139, 5021]. feeding-deterrent [4938]. Felder [3433]. Female [3513, 2696, 2907, 3274, 4783, 4123, 5271, 4964, 3172, 3660, 2549, 4526, 4064, 3721, 4455, 5327, 4628, 4333, 4741, 5227, 5274, 5160, 4661, 2889, 4853, 4901, 5252]. females [2775, 1866, 3986]. fennerl [3371]. Fenneropenaeus [5121]. ferox [4495]. ferruginea [2607, 2863, 3058, 4382, 3043, 3415, 1816, 2578, 2691, 2624, 5149, 5085]. fertile [1084]. Fertilization [2096, 2100, 2353, 2113, 3986, 2498, 1929]. fertilizations [539]. fertilized [236]. fertilizing [1762, 1830]. fevers [1051]. few [4592]. fewer [4555]. Fiber [460]. fibre [2664]. Fiddler [2194, 2118, 2480]. fidelity [4317, 5188]. Field [3449, 3201, 3613, 3171, 2397, 4000, 3452, 2945, 5073, 4581, 2700, 3468, 2814, 4445, 5186, 5354, 4300]. fields [601, 4434]. Fifteen [3029, 1194]. fifty [388, 22]. fifty-six [22]. figures [107]. filamentosus [4026, 3937, 4564, 4076, 5135, 4354, 3071, 5316, 3176]. file [327]. file-fish [327]. filefish [1165]. Filefishes [1853]. Fillets [3403, 4421]. Filling [2443]. films [2382]. Filter [2878, 2931, 3054, 2140, 4235, 358]. Filter-Feeding [2931, 3054, 2878]. Filtering [1803, 2306, 4053]. filth [1051]. Filtration [2133]. fimbria [3195, 5058, 3276, 3331, 3635, 2517, 4232, 4417, 3872, 3804, 4169, 4088, 4898, 4631, 5160, 3445, 4263, 5211, 5051, 3061]. fimbriata [2022]. Fin [3184, 2964, 2266, 3929, 3262, 5337, 2584, 4431, 2711, 2837, 2985, 3598, 4534, 3966, 3507, 2794]. fin-ray [4431]. financial [2479]. Finback [3277]. findings [4539]. Fine [5297, 5220, 5298]. Fine-scale [5297, 5220, 5298]. finescale [4681]. finetooth [2804, 4562]. Finfish [3029, 3628, 3055, 5321]. Finfishes [2692, 2540]. Finger [1457, 1506]. fingerling [1881]. fingerlings [1757]. finite [4623]. Finland [271, 1081]. Finned [2999, 4024, 2876]. Fins [2964, 2837, 2985, 4421]. finwhale [878]. Fire [2727]. First [5207, 3435, 41, 2738, 3570, 2762, 3294, 2397, 3656, 5186, 4, 3442, 2471, 2585, 4495, 284, 5361, 3993, 3996, 55, 159, 5336, 4415, 680, 4010, 5236, 3357, 472]. first-year [4415, 5236]. Fish [3391, 2044, 2191, 3007, 2913, 1052, 125, 267, 562, 2050, 2177, 62, 48, 945, 3309, 2872, 2481, 1128, 2815, 2433, 2222, 2206, 3201, 5170, 189, 72, 517, 620, 2185, 2036, 2038, 2213, 899, 2823, 1195, 2776, 2243, 21, 2143, 3204, 1025, 3105, 3702, 788, 159, 4075, 3662, 4674, 2513, 2517, 2579, 812, 2857, 5223, 206, 2478, 2629, 140, 95, 1582, 2254, 3615, 2860, 3524, 5068, 302, 2042, 190, 3257, 953, 2310, 2058, 3586, 2045, 3624, 961, 2064, 1178, 2467, 2088, 4316, 2583, 3131, 3116, 2106, 4973, 193, 269, 1050, 3701, 1122, 263, 3339, 1158, 3506, 3189, 845, 827, 584]. Fish [3232, 2408, 3313, 628, 1019, 2025, 3614, 2569, 3249, 170, 1012, 1289, 5156, 437, 855, 979, 847, 2326, 578, 3478, 724, 4894, 2411, 131, 214, 313, 296, 231, 676, 2212, 2313, 238, 773, 3639, 3747, 2247, 939, 963, 1091, 499, 1411, 4759, 2987, 187, 4635, 3519, 730, 291, 793, 259, 315, 3605, 664, 326, 4784, 4493, 1832, 4622, 660, 729, 558, 90, 700, 1037, 3292, 569, 426, 4164, 410, 347, 5081, 490, 3675, 4084, 305, 216, 983, 117, 1049, 3762, 2664, 1056, 4521, 4210, 4873, 3147, 4283, 340, 5349, 3934, 10, 898, 1027, 4830, 1796, 375, 150, 213]. fish [871, 838, 785, 4643, 1456, 2722, 3259, 139, 211, 5015, 4541, 4953, 4999, 318, 76, 409, 287, 714, 1393, 1400, 576, 636, 4871, 1359, 5344, 801, 4710, 839, 5196, 411, 534, 2825, 1221, 4900, 5080, 1030, 1134, 790, 345, 663, 63, 5061, 439, 946, 3387, 4012, 1417, 1421, 96, 138, 160, 719, 4121, 2957, 4427, 2636, 30, 1180, 639, 2547, 1188, 1222, 2610, 3073, 3787, 4835, 4428, 3, 155, 5017, 4750, 4930, 130, 4393, 4394, 148, 1125, 4050, 47, 2742, 2555, 744, 25, 519, 2834, 4004, 829, 1098, 286, 348, 1307, 861, 5046, 1526, 1360, 1485, 1505, 4226, 1408]. fish [709, 210, 795, 1472, 3807, 1051, 4395, 1177, 5165, 5125, 1055, 1129, 5161, 1979, 5166, 5218, 462, 4218, 4803, 3780, 1306, 327, 4927, 1837, 271, 809, 360, 4695, 1388, 4, 937, 3647, 2465, 3117, 281, 458, 768, 5350, 4660, 1594, 431, 488, 4227, 430, 666, 617, 1672, 4511, 3426, 4058, 550, 756, 844, 224, 4611, 369, 4374, 1063, 1081, 540, 124, 331, 597, 796, 31, 947, 743, 4399, 3918, 4983, 353, 747, 3942, 4185, 4768, 254, 5246, 732, 4981, 3796, 868, 720, 640, 629, 1791, 5012, 2552, 4813, 4297, 4698, 1376, 4751, 65, 307, 352, 696, 1984, 3138, 5260]. fish [1696, 705, 4806, 623, 518, 359, 1799, 1950, 1961, 1308, 374, 2699, 230, 418, 164, 173, 317, 356, 377, 379, 496, 357, 373, 472, 632, 635, 647, 806, 749, 804, 819, 862, 891, 1094, 1252, 1567, 2319, 1339, 1363, 3539, 3655, 5217, 585, 1035, 4009, 4810, 645, 5019, 1043, 134, 654, 1116, 4842, 94, 154, 4650, 4592, 4593, 4114, 4814, 2233, 34, 5234, 1193, 760, 282, 3421, 1403, 283, 380, 4773, 5224, 511, 289, 805, 45, 2905, 5062, 1992, 5256, 386, 1462, 5185, 1761, 1620, 1201, 4594, 1310, 1522, 2732, 4752, 4922, 711, 911, 1374, 590, 770, 1390]. fish [108, 220, 80, 447, 423, 704, 3207, 628, 170, 248, 867, 909, 207, 879, 429, 370, 401]. fish-bladders [756]. fish-breeding [134]. Fish-cultural [1050, 1411, 1049, 576, 1505, 1388, 711]. Fish-Culture [898, 72, 620, 899, 1019, 847, 578, 724, 664, 983, 63, 666, 1063, 732, 1116, 590, 447]. fish-destroying [801]. fish-eater [488]. fish-eating [47, 430, 374]. fish-food [439]. Fish-guiding [2025]. Fish-habitat [3747]. fish-hatching [281, 654, 282]. fish-hawks [230]. fish-hook [375]. fish-oil [317, 283, 380]. fish-ova [124]. fish-ponds [31]. fish-trap [636]. fish-traps [709]. fish-way [540, 34, 289]. fished [3981, 4041, 334]. Fisheries [2802, 125, 2275, 874, 3014, 2730, 1478, 480, 2331, 102, 508, 1524, 1572, 3638, 3354, 1355, 2123, 3401, 68, 142, 325, 545, 2149, 2948, 2501, 4264, 3252, 1260, 427, 3033, 1933, 3510, 3098, 3102, 2415, 2287, 2627, 4558, 4622, 660, 249, 840, 1052, 2535, 228, 490, 4686, 3009, 706, 2693, 4283, 444, 522, 523, 5333, 252, 4617, 1041, 4884, 465, 1349, 1601, 136, 529, 655, 921, 905, 943, 987, 995, 875, 4193, 1655, 751, 794, 828, 3900, 4581, 457, 448, 4652, 4730, 1347, 4992, 1034, 820, 70, 385, 541, 659, 864, 4544, 92, 443, 764, 1125, 1720, 901, 262]. fisheries [384, 1065, 4152, 1360, 1357, 3906, 105, 1149, 5215, 295, 1045, 4085, 2682, 725, 42, 83, 106, 19, 127, 181, 232, 301, 420, 509, 467, 474, 493, 513, 697, 593, 594, 4906, 679, 1007, 1069, 1113, 4740, 1364, 1158, 887, 932, 435, 602, 740, 757, 895, 1343, 2581, 397, 382, 476, 671, 826, 5011, 5364, 158, 3774, 1593, 1612, 1631, 1634, 1345, 2127, 1645, 1893, 923, 1017, 763, 907, 5126, 5210, 777, 77, 264, 300, 356, 377, 379, 496, 632, 635, 647, 749, 804, 819, 862, 891, 989, 1002, 1014, 1066, 1039, 1059, 1080, 1089, 4421, 5162, 1367, 1047]. fisheries [5051, 278, 622, 661, 633, 824, 821, 1351, 4976, 5268, 4824, 1046, 910, 609, 144, 5045, 1371, 775, 786, 1781, 3046, 1154, 766, 5323, 4745, 589, 605, 637, 701, 738, 755, 761, 772, 781, 789, 792, 814, 823, 830, 841, 846, 860, 769, 783, 853, 877, 885, 893, 908, 916, 925, 1201, 1028, 22, 153, 365, 429, 370, 649, 736, 53, 1353, 5250, 4307, 220, 840, 2150, 313, 1379, 1411]. fisheries-independent [4824]. Fisherman [1122, 656]. Fishermen [3216, 2525, 2160, 631, 125, 905, 2309, 580]. Fishery [2977, 1139, 2892, 2654, 2730, 2503, 2564, 2487, 2779, 2246, 2277, 2279, 17, 3038, 3652, 2131, 3496, 869, 3431, 3202, 251, 2273, 2507, 3135, 3334, 3645, 78, 81, 82, 4676, 3411, 2866, 110, 298, 3522, 3278, 2093, 2405, 2171, 2061, 2753, 1860, 2745, 2578, 3639, 3589, 2938, 2439, 2976, 2447, 2716, 3217, 3595, 2422, 704, 4784, 3379, 1133, 4117, 631, 2366, 4640, 5322, 5065, 4281, 4911, 5057, 4961, 4600, 3374, 5189, 4036, 4126, 4090, 4657, 5299, 2230, 4191, 257, 5307, 5113, 3834, 4772, 5153, 1127, 1174, 878, 15, 174, 161, 341, 799, 889, 964, 2502, 277]. fishery [5194, 5095, 2877, 3819, 4349, 4849, 4470, 839, 5305, 4994, 276, 920, 5313, 4929, 5080, 4863, 186, 309, 4543, 4730, 4524, 4797, 35, 4712, 4472, 1828, 3333, 4821, 92, 5096, 4214, 1146, 1683, 4693, 406, 3241, 4417, 2598, 614, 2871, 4483, 3681, 5331, 5377, 4653, 4685, 4315, 4494, 2971, 2576, 3072, 1029, 5328, 36, 5, 4071, 3040, 4588, 5276, 4677, 4246, 4353, 39, 1987, 1982, 3769, 3822, 4467, 5011, 4614, 2991, 1012, 4468, 5033, 1603, 2142, 5043, 4802, 4047, 1831, 196, 969, 971, 955, 1095, 1118, 1167, 1164, 1173, 245, 882, 3805, 265, 4491, 1161, 4875, 4533]. fishery [4041, 2683, 817, 1046, 5164, 3870, 767, 962, 612, 1964, 1663, 3952, 3298, 5213, 5092, 1621, 5232, 5180, 5259, 4539, 1795, 4755, 4840, 5308, 4782, 4059]. Fishery-dependent [3639, 5331, 4467]. fishery-derived [4730]. fishery-independent [5065, 4772, 4994, 5276, 4467, 5033]. Fishery-induced [3652]. Fishes [2802, 3078, 1737, 2296, 2614, 2524, 3183, 2423, 2455, 2179, 4903, 2608, 2657, 2530, 2301, 2484, 2671, 2851, 2963, 2008, 2854, 3051, 3184, 2122, 3004, 596, 2917, 2509, 1322, 2827, 3324, 3494, 1144, 1442, 1337, 2383, 1179, 2942, 2391, 1565, 2076, 2148, 2335, 2490, 2632, 2901, 3613, 2414, 2358, 3113, 2086, 1331, 1334, 2360, 3125, 2217, 2347, 2771, 2638, 3205, 3617, 3131, 2223, 2556, 3012, 2868, 3518, 3246, 3597, 2663, 3547, 3319, 2548, 2816, 2558, 1412, 2326, 3428, 2973, 1112, 2141, 3657, 3534, 2295, 2806, 2858, 2534, 2239, 3347, 2747, 2831, 3079, 2289, 2494, 2887, 4499, 4833]. fishes [1270, 1539, 1535, 1701, 1702, 1703, 1705, 1706, 1708, 1711, 1714, 1717, 1719, 1710, 1048, 2497, 1032, 1419, 1402, 5081, 5119, 1879, 936, 945, 983, 1947, 4021, 1389, 1646, 949, 1886, 44, 3803, 1602, 1694, 2704, 3160, 1516, 1580, 1677, 678, 3819, 3473, 3057, 2689, 1008, 1261, 3937, 1011, 1070, 1092, 1097, 1106, 1123, 1204, 1219, 1233, 1332, 1481, 1491, 1424, 1625, 2702, 4001, 5129, 1455, 194, 950, 954, 1078, 1422, 2505, 5176, 30, 3115, 4945, 1117, 994, 1473, 4995, 1356, 1586, 204, 407, 958, 1077, 1157, 1241, 4113, 1531, 1584, 1571, 1615, 1967, 74]. fishes [1846, 1420, 1391, 1383, 1399, 2186, 2762, 1958, 1194, 1186, 1262, 2370, 93, 169, 182, 167, 417, 638, 941, 973, 1085, 1169, 1258, 1251, 1245, 1320, 1330, 1425, 1305, 1075, 1090, 1256, 1247, 1485, 1505, 530, 1563, 4684, 980, 1000, 2323, 1589, 1479, 1494, 5070, 2751, 2690, 3279, 1770, 86, 4257, 4724, 975, 1171, 1313, 1409, 1410, 5020, 1394, 20, 1482, 3728, 3153, 876, 1418, 974, 991, 1082, 1072, 1100, 1844, 2209, 1398, 2538, 1015, 1868, 4294, 4890, 1083, 1246, 1299, 1434, 1448, 2288, 1489, 1507, 1659, 1362, 486, 157, 4957, 2837, 3293, 4344, 1902, 3499, 1384, 1464, 1908, 2260]. fishes [3052, 3303, 4572, 1162, 1340, 65, 71, 49, 185, 394, 4178, 3022, 4063, 1397, 3083, 4894, 1218, 1172, 985, 997, 1006, 1095, 1160, 1239, 1242, 1341, 1450, 1477, 3738, 1413, 4321, 3120, 1317, 3503, 1514, 1228, 1395, 4865, 4268, 4880, 3443, 1679, 5171, 1476, 4926, 977, 982, 1068, 5090, 1375, 5040, 682, 1407]. Fishing [3483, 3697, 641, 63, 145, 2566, 130, 2629, 3564, 3225, 3205, 2910, 2482, 3261, 1044, 759, 3249, 3322, 3707, 3693, 3267, 335, 2447, 294, 966, 395, 4911, 2535, 831, 5352, 4411, 4878, 4020, 10, 4958, 5016, 852, 943, 981, 4896, 364, 4808, 1131, 583, 4929, 859, 4900, 4563, 498, 5129, 4854, 4564, 3465, 4939, 1356, 4367, 1366, 536, 4987, 1325, 1021, 4870, 5100, 5109, 4800, 50, 52, 215, 425, 110, 5082, 4175, 492, 827, 5143, 195, 3769, 3423, 4981, 5267, 4888, 2176, 808, 5316, 4163, 4402, 487, 4340, 5177, 611, 939, 963, 552, 1964, 926, 4187, 4718, 4067, 453]. fishing-banks [852]. fishing-ground [859]. fishing-site [5109]. fishponds [1530]. fishway [2906, 235]. Fishways [1415, 397]. Fishy [2441]. Fistulicola [3300]. Fit [2633]. Fitting [2354, 322]. Five [2296, 3385, 3190, 2320, 3405, 3008, 2456, 2812, 3487, 147, 798, 4478, 4994, 1233, 4967, 3104, 2508, 1479, 2321, 3728, 5023, 2829, 1239, 1946, 4603]. five-day-old [3104]. fixation [201]. fixed [5245, 3714]. fixed-station [5245]. fjord [4493, 4989, 3901, 5012, 5025]. fjords [4848]. Fla. [1571, 157]. flag [4840]. flag-associated [4840]. flagellate [1570]. flat [1807, 5020]. Flatfish [2022, 2078, 4971, 2285, 5346, 1439, 4811, 4970, 4450, 4841]. Flatfishes [2993, 2663, 2244, 2710, 4728, 2508, 4984]. flathead [3739, 4742, 5364, 999]. flatworms [1736]. flavescens [4329, 1821, 4216, 4842]. Flavidus [2908, 3727, 4796, 2788, 3288, 3960, 3772, 3889, 4238, 4270]. flavipinnis [5295]. flavolimbatus [4332]. flavolineatum [3208]. Flavor [2344, 2441]. Flavors [2042]. Fleet [3564, 4411, 3842, 2479, 4228]. Fleming [88]. flesh [870, 917, 1324, 1409]. Flexible [5193]. Flicking [2932]. flies [143]. floating [4383]. Floats [546, 753]. Flödevig [1393]. floods [91]. floor [2728]. flora [1444, 3016, 1423]. Florida [3419, 966, 4242, 4498, 2512, 4164, 2618, 4412, 3385, 3466, 3733, 5028, 4884, 1127, 927, 1941, 1981, 3979, 4040, 4109, 4194, 4195, 4461, 1517, 4756, 4644, 741, 1131, 1862, 1978, 4929, 1134, 3399, 63, 5258, 5351, 4967, 3866, 4506, 482, 2010, 4930, 958, 1077, 1125, 1366, 3241, 2735, 2944, 2996, 417, 1970, 3559, 2719, 2217, 2347, 4257, 2467, 4091, 5032, 4430, 4587, 4071, 2821, 2830, 1113, 4227, 5276, 1993, 1141, 3607, 4162, 4155, 2801, 2429, 39, 486, 4405, 4758, 4598, 4632, 2548, 2368, 4940, 2059, 4107, 923, 1115, 4384, 3922, 4179, 4047, 4955, 2973, 4163, 1126, 1152, 1173, 661]. Florida [824, 821, 4491, 1153, 1933, 1994, 1999, 2717, 1995, 2020, 2215, 3767, 4355, 2941, 2043, 4666, 2534, 1848, 4187, 5232, 3855, 793, 1143, 5204, 4515, 982, 4782]. Flotsam [2705, 2432, 1958]. Flounder [3563, 3603, 5079, 3606, 2022, 2863, 3058, 3043, 2881, 2580, 2220, 3415, 2180, 2457, 2578, 2381, 2237, 2612, 2624, 5085, 5002, 4100, 2900, 4182, 4070, 3950, 4453, 3381, 1519, 2764, 4990, 4150, 4553, 2607, 3104, 3789, 4382, 4627, 1492, 5003, 4876, 5117, 3903, 3976, 4432, 3248, 3938, 1816, 3479, 2766, 4449, 2691, 4901, 3891, 4300, 5205, 5149, 3345, 3839, 3830, 4102]. Flounders [1690, 2295, 1482, 4017, 3743]. Flow [4210, 2336, 5089, 1980, 3990]. Flow-related [4210]. flows [4128, 3855]. fluctuating [4885]. Fluctuations [2977, 1662, 2562, 1821, 1720, 1772, 2467, 3630, 1623, 1624, 1666, 278, 3972, 1731, 4507, 5105, 1760]. Fluffy [3219]. Fluorochrome [3605]. fluorometric [3934]. Flushed [2831]. fluviatilis [4842]. flying [4958, 4830, 5155, 597]. flying-fish [597]. flyingfish [4303]. focus [3963, 4921]. Focusing [2583]. foetens [5075, 4909]. follicle [3969, 3145]. follicles [4828, 5003]. Following [2342, 3455, 3612, 5352, 5377, 4185]. fontinalis [152]. Food [4089, 4329, 3198, 3963, 2497, 3251, 2219, 2903, 3124, 2423, 3158, 2179, 1523, 3080, 5266, 4946, 3185, 2028, 2187, 644, 2934, 2573, 2623, 2784, 3331, 3635, 3352, 1835, 2010, 2196, 2134, 2452, 2158, 1887, 3429, 2120, 2029, 445, 2508, 2369, 3043, 3077, 2638, 2919, 3619, 2049, 2499, 3500, 2718, 37, 3491, 3189, 2427, 4487, 3627, 3313, 3284, 3015, 2822, 2816, 1732, 3325, 5024, 3487, 2575, 3209, 3214, 2295, 3792, 3448, 2882, 1829, 3132, 5092, 1934, 3287, 3481, 2887, 1917, 228, 906, 1042, 5050, 5174, 1027, 2869, 918, 3895, 409, 1806, 4026, 4141, 1348, 1428, 9, 43, 55, 439]. food [4166, 199, 524, 2645, 2666, 2370, 348, 571, 1326, 530, 1792, 4217, 4494, 1409, 20, 937, 1392, 2184, 553, 1387, 743, 1489, 1507, 1062, 1102, 157, 1352, 351, 69, 49, 185, 705, 2587, 3034, 1961, 1095, 1239, 1161, 1579, 2233, 1459, 1374]. food-fish [348]. food-fishes [530, 20, 157, 1095, 1239]. Food-Seeking [3429]. Foods [1419, 2474, 3057, 2726, 2856, 3083]. foot [2056]. forage [1883, 4814]. Foraging [2872, 4282, 5065, 5355, 3785, 3660, 4965, 4629, 4037, 3291]. foraminifera [1206]. forbesi [1654, 2011]. forced [4615]. forceps [2477]. forces [2371]. Forecast [2866, 3863]. Forecasting [4047]. Foregut [2112]. Foreign [740, 3595]. foreigner [732]. Forest [2509, 3125, 3214]. forestomach [5293]. Forests [3309, 2423, 2823, 2638, 2825]. forgotten [5245]. fork [3142]. Form [3461, 2601, 4279, 4308, 4326, 4345, 4361, 4387, 4408, 4424, 4440, 4458, 4475, 4497, 4518, 5372, 2590, 358, 957, 4765]. formalin [4668, 4142]. formalin-preserved [4668]. format [3994]. Formation [3581, 3386, 3181, 2967, 4343, 4348, 1486, 4696, 4003, 3138, 4616]. formosus [3516]. forms [1784, 1052, 437, 855, 296, 464, 1670]. Forsskål [4757]. Fort [2228]. FORTRAN [2633, 2396]. Fortune [3388]. Fossil [2270, 2143]. Foul [1051]. Fouling [2475, 2451, 4611, 1587]. found [2270, 870, 5293, 1221, 241, 4245, 167, 941, 940, 973, 1320, 1112, 645, 192, 4759, 1310, 1499, 1530]. Four [2017, 3017, 3361, 1179, 3555, 2438, 3591, 3541, 313, 3680, 4242, 4961, 555, 2280, 4001, 4427, 2201, 1117, 4960, 4987, 4866, 1802, 4110, 2, 4810, 3682, 4616, 3103, 2295]. four-pound [555]. four-spined [2]. fourbeard [2722, 4213]. Fourteen [2851, 1615]. Fowler [5075, 1974, 2285]. FPC [2213]. Fraction [2359]. Fractions [3572]. fragile [4816]. fragment [4766]. Fragments [3605]. Frailes [5054]. Framework [3564, 4547]. France [361, 347, 559, 415, 502, 620, 576, 990, 745, 504, 279, 382, 640, 629, 683, 880, 763, 907, 642, 505]. franciscanas [4670]. franciscanus [4522, 4618, 2582, 4590]. Francisco [4057, 364, 4484, 3310, 4696, 3191, 3587, 580]. franciscorum [2775]. Fraser [3011, 4560, 4559, 3370, 4396, 3052, 1645]. Free [1736, 2061, 4758, 1791, 1181]. Free-living [1736]. free-swimming [1181]. Freedom [861]. Freeze [2061]. Freeze-Dried [2061]. freezing [838, 1728, 5320, 4142, 4951, 27]. freminvillei [5164]. French [3208, 1237]. frequencies [1693, 4295, 3808]. Frequency [3566, 3035, 3346, 2633, 2917, 3329, 3498, 3581, 2396, 3467, 3157, 4060, 3809, 3779, 3811, 4787, 3915, 3813, 4826, 2651, 4675, 4584, 4161, 4331, 4714, 4317, 4156, 4466, 4007, 4338, 4717, 4575, 3998]. Fresh [2437, 1496, 3582, 320, 1374, 1784, 1528, 355, 1614, 1523, 1510, 409, 828, 1516, 1677, 686, 1261, 1619, 790, 1520, 473, 1088, 1075, 1090, 1247, 1148, 1149, 1346, 1435, 1309, 2184, 1031, 1166, 1454, 910, 1476]. Fresh-water [1496, 320, 1374, 1784, 1528, 355, 1614, 1523, 1510, 828, 1516, 1677, 1261, 1619, 1520, 1088, 1075, 1090, 1247, 1148, 1149, 1346, 1435, 1309, 1166, 1454, 910, 1476]. Freshwater [2413, 5348, 2489, 4325, 5290, 2186, 3357]. fried [4842]. frigate [4038, 1671]. Fringed [3448]. frog [618]. frogs [222]. frondosa [4776]. Front [576, 5203]. frontal [2465]. frost [796, 585]. frost-fish [796, 585]. frostfish [230]. Frozen [3656, 501, 565, 934]. frozen-herring [934]. Fry [2417, 2516, 2839, 2442, 866, 62, 1255, 3884, 1406, 481, 687, 833, 272, 2409, 2557, 201, 1396]. Fuca [3251]. fucensis [3520]. fucicola [4746]. fully [5307]. Fulton [5050, 224]. fulva [4521]. function [4249, 4410, 4910, 4132, 1592, 1471, 4859, 5193, 5235, 5047, 1299, 4185, 4259]. Functional [2608, 2213, 5355, 3461, 2828, 3482, 5187, 3141, 2817]. Functions [2033, 1478, 5013, 37, 1416, 4590, 1379]. Fundulus [3212, 2967, 2794]. Fundy [5006, 3131, 3220, 3068, 934]. fungi [1750]. fungus [1420, 89, 1413, 642, 1375]. Funka [3722]. fur-farming [1143]. Fur-seal [1566]. fur-seals [250]. furcatus [5291]. furcifer [4521]. furnieri [4567, 3604]. furnish [125]. furnished [138]. Further [1511, 1506, 507, 1639, 2464, 3478, 2406]. furunculosis [1777]. fushery [4747]. future [1054, 1064, 512, 1040, 1154, 484]. Fyke [1014]. Fyke-net [1014].

g [3453]. G. [2210]. gadid [2722, 5244]. Gadidae [3806, 2883, 2226, 2305, 389]. gadids [2894]. Gadiformes [5338, 3243]. gadoid [3485]. Gadus [3228, 3062, 3093, 3435, 2869, 4896, 3081, 5313, 4730, 4947, 4288, 4988, 2666, 3804, 4866, 4645, 3405, 5311, 3029, 3561, 4200, 3810, 5041, 4906, 2883, 4236, 4136, 5089, 4698, 3710, 4419, 3905, 4853, 3132, 5179, 3758, 1888, 4654]. gag [3372, 4165, 4199]. Gaimard [3735, 3809]. Gains [2275]. gairdneri [3412, 2819, 2826, 2385, 3386, 2357, 2250, 2786]. Galápagos [4220]. Galathea [2688, 2533]. galatheid [2477]. Galatheidae [2805, 3868, 2688, 2533, 3556]. Galena [2259]. Galeocerdo [3358, 5194, 4358, 4295]. Galeorhinus [3648, 4588, 3705, 4225]. galeus [3648, 4588, 3705, 4225]. Gallbladder [2594]. Galveston [1955, 4643, 1626, 3156]. Gamba [3364]. Gambusia [1472]. game [1362]. gamete [2498]. Gammarideans [3229]. gangions [4406]. gar [65]. garfish [4570]. Gas [2404, 2084, 2425, 3639, 3657, 2781, 2412, 4787, 1180, 698, 1447, 1766, 154, 1228]. Gas-bubble [2425, 1180]. Gases [1385, 1989]. gaspereau [495]. Gasterosteus [1823, 2499]. gastric [3351]. gastrointestinal [2867]. gastrophysus [3270]. gastropods [4124]. gate [850]. gathering [1148]. Gear [2779, 3601, 3707, 3371, 4552, 5352, 5299, 4871, 5196, 5284, 4854, 4621, 4791, 5035, 4429, 5269, 5246, 4344, 4802, 4841, 4406, 3345, 4187, 4067, 4137]. gear-specific [4137]. Gel [2583]. Gemma [1807]. Gempylidae [4525, 4218]. gen [2231]. gene [4392, 4731]. genealogies [4293]. Genera [2915, 2757, 3497, 2722, 1806, 1123, 2563, 1690, 2922, 2758, 3516, 940, 1258, 2888]. General [1545, 1546, 1922, 1016, 3126, 1198, 2263, 2360, 1023, 4271, 2494, 4547, 3942, 1109]. generalization [1422]. Generalized [2074, 2354, 3908, 5124, 3897]. generating [4171]. generation [162, 797]. generative [200]. Generic [2890, 1939]. generosa [6]. genes [4335]. Genesee [99]. genesis [65]. Genetic [5303, 3733, 3866, 3109, 4732, 3521, 4214, 5341, 4645, 4055, 5148, 5029, 4335, 3248, 3398, 3841, 3636, 3175, 3176, 4592, 4815, 3341, 3512, 3770, 4836, 5025, 4365, 5098, 4829, 3480, 4453, 4919, 5101, 3878, 3719, 3748, 4172, 4966, 4174, 5374, 4398, 4530, 4400, 4638, 4035, 4857, 3712, 4318, 4125, 4043, 5277]. Genetically [5150]. Genetics [3582, 5095, 5262, 5287, 3039]. Geneva [946]. genitalia [657]. genomic [4942]. Genotypes [2786]. Genus [2017, 2767, 3540, 2433, 2066, 2231, 2102, 2611, 3182, 3364, 3453, 2261, 1953, 2333, 2410, 3582, 3527, 3433, 2109, 3413, 2070, 2925, 2308, 2311, 3757, 2606, 2488, 2329, 3497, 3631, 4850, 5098, 3158, 1577, 1643, 3036, 1576, 5363, 2390, 2007, 2563, 2678, 2769, 2899, 1793, 2280, 1458, 1690, 1301, 1967, 3924, 4785, 4234, 290, 4144, 4352, 5128, 2290, 4202, 4205, 1946, 3103, 1287, 1671, 3753, 1310, 1490, 3968]. Genyonemus [2505, 3429, 3135, 3022, 2989]. geoduck [6]. Geographic [3717, 3719, 5243, 4481, 4381, 2367, 4172, 2159, 3522, 3690, 3663, 3114, 4853, 3833, 4703, 3853, 4559, 1941, 1643, 1011, 3879, 4297]. Geographical [4798, 3606, 2601, 1357, 2590, 164]. Geographically [3162]. geography [4118]. geolocation [4806]. Geological [53]. Geology [1738]. geometric [5247]. geometry [5241, 4753]. Geomorphology [2013]. geophysical [5206]. George [3216, 3929, 3472, 4151, 531, 4510]. Georges [4010, 2975, 4513, 4388, 3093, 3890, 3825, 3829, 4645, 4494, 3810, 5041, 2220, 3232, 3281, 1785, 1769, 1699, 1696, 3802, 1800, 4269, 1652, 3758, 3419]. Georgetown [324, 365]. Georgia [3701, 3820, 2122, 2210, 4462, 3064, 950, 3200, 140, 3496, 1531, 5315, 4695, 1064, 3918, 2295, 2976, 32]. Geostrophic [2336, 1972]. German [28, 776, 308, 149, 138, 17, 190, 381, 104, 193, 298, 743, 378, 452, 299, 270, 237, 187, 27, 123, 704]. Germany [288, 21, 243, 959, 1098, 861, 516, 147, 739, 436, 455]. germinal [107]. germo [1887]. Geryon [2696, 2549, 2200, 3121, 3371]. Geryonidae [3121]. Gestation [5227, 5300, 842]. Ghent [24]. Ghost [3695, 4340, 4411, 4900, 5100, 4258, 3769]. ghost-fishing [4411, 4900]. Giant [3126, 2324, 4961, 5110, 2738, 4381, 4788, 3971, 4091, 3159, 658, 969, 3330, 5318]. gibboides [2362]. gibbus [3790]. Gibraltar [4191]. giganteus [2695]. gigas [3394, 3733, 4644, 5243, 1648, 3354, 3584, 4704, 3824, 2498, 3618, 3718, 3885, 3852, 4041, 4206]. Gilbert [4928, 3036, 5144, 169, 182, 4913, 4303]. Gill [15, 2049, 4352, 3143, 1935, 2548, 3663, 2973, 2534, 2933, 257, 205, 341, 529, 667, 3827, 2659, 372, 50, 66, 109, 322, 215, 330, 420, 509, 697, 2465, 3384, 3031, 217, 5203, 1876]. gill-net [257, 341, 529, 215, 420, 509, 697]. Gill-nets [15, 205, 667, 50, 66, 109, 322, 330]. Gill-Raker [3663]. gilled [1484]. Gillichthys [2185]. Gillnet [5048, 3800, 4193, 4152, 5092]. gillnets [4939]. Gillnetting [2246, 4050]. Gills [1928, 1914, 3016, 1223, 1592, 721]. Ginglymostoma [5228, 4183]. ginsburgi [2210]. Girard [1977, 3244, 2062, 1897, 3330]. Girella [3534]. gizzard [1937, 1844, 3357]. glacial [4493, 4848, 5168]. glacialis [4895, 3021]. gladifer [3675]. gladius [5243, 4191, 2914, 4349, 4849, 4313, 4606, 4042, 3954, 4064, 3300, 2964, 2985, 4406, 4534, 3767, 2941, 3966, 3247]. Gland [2199, 3554]. glands [657]. Glandular [2261]. glassy [4830]. glauca [5266, 5194, 4748, 4729, 5278, 2144, 2157, 2739, 2572, 4591, 2713, 5357]. Glaucosoma [4482]. Gld. [6]. gliding [4915]. Global [4202, 4116, 4907, 5248]. Globicephala [2735, 3703]. globicipitis [4907, 3723]. glochida [1476]. Gloucester [249, 631, 465, 480, 206, 109, 4, 78, 81, 82, 42, 83, 36, 106, 127, 181, 215, 301, 330, 420, 509, 467, 474, 493, 513, 697, 593, 594, 238, 499, 877, 885, 893, 908, 916, 925]. Glugea [2748]. glutinosa [4093, 4173]. Glycera [2642]. Glycimeris [6]. Glycogen [1486, 3458, 2936]. Glyptocephalus [4783, 3381, 2543, 2545, 2665, 4432, 3345]. Gmelin [327, 212]. Gobiesocidae [2723]. Gobiesox [2723]. Gobiid [2185]. Gobiidae [2210, 1630]. Gobiosoma [2210]. Gold [3082, 1938, 352, 40]. gold-fish [352]. goldband [4549]. Golden [2407, 3283, 2969, 3371, 5334, 599, 5158, 5330, 3302, 4205]. goldeneye [5295]. Goldeye [2228, 1828]. Goldfish [1361]. goldflag [5295]. goldlined [4675, 5374]. goldorfe [288]. goliath [4758]. Gonad [3190, 2040, 3565, 3628, 4064, 4739, 3247]. Gonad-Somatic [3628]. Gonadal [3646, 2978, 4204, 5252, 3288]. gonadectomy [1801]. gonads [3306]. Gonatidae [4682, 3999]. goniurus [2626]. Gonostoma [2762]. Gonville [620]. Gonyaulax [2575, 2587]. good [859]. Goode [3465]. goodei [2398, 3988]. Goodness [2633]. Goose [2432, 2810]. goosefish [4860, 3724, 5301, 5356]. gorbuscha [2417, 2516, 2839, 5049, 1636, 1641, 2677, 4334, 4398, 2865, 2852, 4715, 1681]. gouldii [4920]. governing [1552]. government [5322, 38, 275, 174, 259]. governor [453]. Grace [835, 670, 307, 319]. gracilis [5244]. gradient [3089, 5198]. gradients [4100, 4430, 4872]. grading [4246]. Grampus [912, 2768, 961]. Grand [542, 641, 685, 928, 626]. Grande [4670, 1011]. grandis [3351]. granulosa [3832]. Grass [2204, 2215, 4198, 4219, 2394]. Grasses [2267]. Gratifying [149]. Gravel [2516, 2839]. gravimetric [4687]. gravity [1514]. Gray [2069, 3173, 2542, 3087, 3066, 4498, 2303, 4412, 5120, 4711, 4967, 5176, 5368, 1659, 4857, 4918, 767, 2639, 3918, 3833]. Grayling [1819]. Grays [3166]. Grazing [2489]. grease [2502]. Great [433, 3864, 3981, 1355, 4628, 852, 4358, 557, 4985, 2096, 936, 4799, 874, 227, 4348, 1354, 889, 4503, 686, 1350, 1219, 1189, 1425, 4736, 1589, 372, 1045, 2100, 3407, 251, 4091, 1029, 2499, 3875, 1319, 4025, 2113, 1018, 1358, 77, 4402, 427]. Greater [2359, 5299, 4196, 5371, 4241, 4667]. Grebe [1819]. Greece [4008, 5238]. Green [2267, 5375, 4602, 5305, 1617, 4712, 1484, 1141, 3962, 2003, 69, 721, 2002, 3998, 4272, 4471, 1720, 1904, 1507, 1775]. green-gilled [1484]. Greenland [4500, 3091, 4843, 4816, 5359, 230, 4912, 3481]. greenling [4360]. greenspotted [4490]. Greenwood [1584]. Gremaz [880]. Grenadier [3418, 3943]. grid [4090]. Griffin [2871]. grilse [274, 67]. griseus [4498, 4412, 4711, 4967, 2768, 4537, 4857, 4918]. groenlandicus [2646]. grooved [4015]. groper [4920]. gross [3338]. groun [211]. Ground [2993, 700, 4423, 859, 2288, 3920, 487]. Groundfish [2258, 2138, 2910, 5322, 2903, 5376, 5313, 4730, 5096, 4585, 5046, 5350, 4529, 4613, 1785, 4976, 3907, 3897]. groundfishes [3961, 5182]. Grounds [1941, 2970, 4047, 966, 4478, 4958, 2460, 699, 912, 1981, 988, 4795, 5378, 5236, 4317, 470, 1153, 939, 963]. group [3842, 1296, 994, 3857, 4200, 3973, 3944, 1943]. grouper [4281, 4357, 4579, 4165, 5115, 4194, 5007, 4952, 5223, 5294, 4332, 5032, 3647, 5190, 4758, 4080, 5206, 373, 4815, 4323]. groupers [5340, 4163]. Groupings [2163]. Groups [2562, 2673, 3613, 4107]. grow [4062]. Growing [2992, 5377]. Growth [3708, 3032, 3629, 28, 3566, 3505, 3035, 2219, 3644, 3926, 3569, 2622, 3540, 3691, 3093, 3435, 3358, 2928, 4103, 4024, 2455, 3216, 3536, 4919, 3572, 4709, 3467, 2046, 3593, 3486, 3312, 2377, 3735, 5226, 2988, 2189, 3399, 3648, 3967, 3632, 3472, 5326, 4254, 3101, 2643, 138, 56, 199, 4774, 3368, 3352, 3706, 2010, 140, 3555, 3222, 3000, 3327, 3524, 2863, 2449, 2452, 4289, 3573, 190, 3173, 3199, 2904, 3380, 2097, 2168, 2782, 3666, 2504, 2271, 2956, 2909, 3656, 2214, 2126, 2341, 2144, 2157, 4217, 1771, 4445, 2350, 2975, 2919, 3645, 2220, 3581, 3604, 2958, 3500, 5306, 4954]. Growth [1851, 3728, 4568, 2733, 3514, 3518, 3386, 3674, 2184, 2960, 3607, 3649, 4295, 4258, 3643, 3627, 3235, 2259, 3313, 5362, 3687, 2473, 2601, 4632, 3678, 3614, 2686, 1561, 3940, 3123, 2953, 2970, 2476, 5273, 3633, 4065, 4377, 2485, 2035, 3680, 280, 3568, 3166, 2435, 3592, 3699, 3642, 1999, 2916, 3622, 3493, 2893, 3102, 2896, 3250, 187, 3121, 2920, 3417, 3450, 4207, 1873, 2756, 3574, 1564, 2783, 3082, 3452, 3839, 3688, 4323, 3314, 2494, 4783, 4498, 1782, 1813, 2926, 3999, 3724, 5255, 4623, 5004, 3853, 3501, 4158, 4070, 3856, 4410, 2945, 3195, 4032, 5005, 3091, 906, 4706]. growth [4624, 3163, 3729, 4412, 5120, 1789, 3134, 5334, 4132, 1614, 3713, 4602, 4642, 4964, 5016, 5106, 4541, 5115, 4523, 4920, 3979, 4040, 4194, 4461, 3898, 1907, 1996, 686, 4849, 4138, 5344, 4213, 4141, 1903, 3044, 5058, 5155, 4827, 4463, 5152, 4159, 3864, 5265, 4671, 4711, 3945, 4689, 3781, 5279, 3783, 4230, 4312, 4253, 1899, 4046, 4763, 4998, 5278, 4285, 4183, 4286, 4124, 4877, 3333, 4795, 3577, 4507, 4545, 3168, 4368, 4482, 1644, 1772, 4287, 4370, 2104, 2555, 2282, 4988, 3739, 4151, 4381, 3985, 4197, 5325, 3789, 1688, 2826, 3042, 4168, 5238, 4859, 1827, 3559, 2162, 3804, 4169]. growth [4736, 2495, 4628, 5154, 5209, 4372, 3849, 3692, 4019, 4198, 4608, 4659, 4010, 3908, 3760, 5124, 1726, 4170, 5193, 4172, 3987, 4429, 4446, 4332, 5370, 4292, 3218, 4517, 5236, 4510, 4548, 4447, 5219, 1693, 4334, 1904, 3901, 4136, 3936, 4486, 1830, 4155, 4512, 2754, 4549, 4118, 4139, 5047, 5239, 4000, 5201, 3096, 4612, 4101, 2685, 4448, 1669, 4104, 3106, 4405, 4354, 5246, 4474, 4467, 4237, 4432, 3159, 2511, 3499, 3920, 3296, 2845, 1729, 4940, 4260, 4336, 4177, 4112, 4590, 4262, 3375, 4037, 4715, 5108, 3946, 658, 573, 5137, 5111, 3742, 3988, 4633, 4375, 3826, 3138]. growth [4879, 4180, 4376, 3114, 4716, 3338, 4996, 4591, 5211, 3738, 1787, 4853, 4011, 4680, 4320, 1579, 1905, 4573, 4699, 4435, 4534, 4266, 1475, 4355, 4240, 4114, 3369, 4267, 4241, 3485, 4083, 5282, 4700, 4969, 3966, 4385, 4717, 5254, 3223, 5251, 4651, 4571, 4667, 4746, 4817, 1610, 1621, 4718, 4451, 5204, 3995, 4224, 4270, 3970, 4271, 4379, 3998, 5169, 3393, 5369, 4471, 5123, 2131]. Growth-Rate [3644, 4545]. Growth-Ring [3573]. growth-transition [4602]. Grunion [2051, 2266, 3085]. grunions [2662]. grunt [4612]. Grunts [3208]. Guam [4865, 5171]. guano [569, 1349, 160]. Guatemala [5333]. Guiana [1044]. guianensis [4571]. guidance [1976]. Guidelines [2243]. guiding [2025]. Guinea [2920, 2073]. guitarfish [4083]. Gulf [3629, 3871, 3358, 3697, 3441, 3315, 3424, 2487, 3105, 4619, 3349, 2478, 2571, 2574, 2712, 3173, 4008, 5238, 3380, 3579, 3893, 2456, 3698, 2868, 3722, 2073, 1893, 5111, 5248, 3508, 3585, 3389, 3657, 3417, 4783, 4280, 2912, 2887, 5098, 1549, 1552, 1535, 1560, 1710, 1740, 1735, 1743, 1741, 1744, 2296, 2268, 4600, 4409, 2892, 3501, 3913, 1646, 4087, 4453, 4251, 4347, 3874, 3977, 5214, 5271, 4579, 3147, 3729, 4807, 4020, 4244, 4562, 4709, 5334, 4903, 3713, 4772, 4964, 2540, 3317, 383, 494, 699, 995, 4165, 4479, 4541, 5110, 5280, 1939, 3935, 2615, 3277]. Gulf [5376, 1811, 4990, 3609, 4213, 2986, 2611, 3473, 3768, 3882, 3860, 3880, 3956, 5196, 5304, 5091, 4480, 3399, 4671, 1776, 1793, 3204, 2934, 5279, 5351, 4014, 4253, 4763, 2692, 4325, 4166, 4625, 5129, 1600, 1611, 4366, 4196, 4852, 3115, 2128, 4757, 2479, 4939, 3611, 2942, 4582, 5223, 2040, 4048, 3178, 4975, 3129, 5371, 5367, 3000, 4287, 4370, 5114, 3171, 2546, 4288, 4289, 5325, 3225, 4909, 2364, 3854, 2203, 2659, 3041, 1827, 1970, 5327, 2871, 4899, 3294, 2323, 4595, 5209, 1885, 4096, 4010, 3558, 4098, 3760, 2879, 2975, 4484, 5032, 4704, 5328, 5370, 3433, 4093, 4173, 5042, 1975]. Gulf [5236, 4834, 5350, 3537, 5368, 2718, 5191, 4965, 3933, 2662, 4208, 4511, 3916, 3649, 3278, 4155, 4512, 4317, 2658, 3413, 3814, 3597, 5269, 4353, 2661, 4513, 2290, 2685, 4448, 3281, 3630, 3143, 3423, 1669, 535, 254, 2822, 2856, 2927, 4742, 3320, 3786, 3976, 4344, 3676, 5052, 5267, 2453, 5364, 3920, 3406, 3616, 4080, 3155, 4112, 4107, 2630, 2551, 3005, 156, 3479, 4762, 4962, 5008, 5146, 4039, 3776, 5335, 623, 4339, 4489, 3445, 4996, 3389, 4401, 2911, 3152, 4853, 4901, 278, 487, 622, 1989, 5347, 2806, 4573, 4699, 4733, 5064, 1933, 4266, 3503, 1119, 1971, 3625]. Gulf [3154, 3949, 4241, 4681, 4767, 962, 4969, 3808, 3417, 805, 5071, 4134, 2488, 5361, 5025, 793, 3605, 3909, 4378, 4451, 5021, 153, 982, 4248, 4654, 5329, 4067, 4102, 4146, 4365, 5229]. Gulland [3626]. gully [5005, 4299]. gummy [4225]. gunnel [4796]. gunteri [4681]. Günther [3743]. guttata [3334]. guttatus [928, 5374, 3742]. guttulatus [4846]. Gymnapistes [3544]. Gymnocanthus [4369]. Gymnodinium [1791, 3285, 2392, 2030, 2575, 2587, 2233]. gymnorhinus [4984, 2961]. Gymnothorax [5369]. Gyre [5324, 2382, 3965, 2771].

H [241, 169, 182, 531, 2900, 2992]. H. [3569, 4587, 4570, 3107, 3510]. Habit [2179, 3429]. Habitat [2669, 2423, 4505, 4909, 4256, 2638, 3591, 3156, 2868, 4105, 3814, 2727, 3490, 3696, 4809, 4814, 3345, 3079, 5090, 4356, 4860, 5044, 4833, 5142, 5312, 4847, 5187, 4958, 4643, 5199, 3773, 4480, 4787, 3941, 5249, 4945, 4930, 5367, 4585, 5360, 5056, 5331, 5297, 4646, 4257, 5125, 5161, 4943, 4803, 5104, 5350, 4965, 5086, 4529, 3918, 3557, 4076, 5135, 3226, 4201, 5033, 5137, 3824, 4531, 3083, 3747, 4206, 4650, 4765, 4268, 4880, 5361]. habitat-based [4529, 5033]. habitat-related [5350]. Habitat-specific [3814]. habitat-use [4643]. Habitats [3391, 3231, 3590, 3263, 3666, 3392, 3164, 2747, 3495, 4557, 3993, 4409, 5115, 3869, 5290, 4150, 4482, 4585, 4044, 4430, 4587, 5023, 5374, 3883, 5156, 4321, 4266, 4759, 4535]. Habits [3198, 2417, 3065, 2179, 2222, 2608, 388, 2963, 3185, 2623, 3331, 3635, 3352, 1036, 2120, 1256, 3043, 2919, 3619, 2718, 796, 2427, 2665, 3143, 1501, 3015, 2816, 3487, 3553, 3214, 2295, 3481, 4089, 4329, 3963, 408, 3158, 4282, 4720, 135, 5266, 4195, 4946, 4213, 4141, 5355, 2573, 4166, 1422, 5176, 4150, 1377, 3577, 6, 1103, 599, 1513, 1391, 4274, 571, 1485, 1472, 4153, 1518, 3077, 1243, 3899, 3491, 3136, 4487, 4629, 4598, 4726, 1384, 5024, 2003, 4037, 1698, 2, 217, 746, 3471, 1678, 3034, 1166, 1252, 1579, 3792, 1808, 4434, 250, 5205, 5092, 998, 3287]. Haddock [3093, 3435, 5073, 4388, 1840, 35, 3825, 1867, 4866, 3810, 5041, 82, 4237, 1769, 1699, 1696, 1800, 807, 4269, 1652]. Hadropenaeus [2563]. Haeckel [2121]. haemastoma [1874]. Haemulidae [3208, 4612, 3209, 3095]. Haemulon [2958, 3208, 3209]. Haff [395, 126]. hagfish [4093, 4173, 4220]. Hagfishes [3631, 2899, 3968]. hair [3377]. Hake [2995, 3124, 2377, 2036, 2038, 3081, 2314, 2045, 3029, 2067, 5088, 3583, 3325, 2551, 2253, 1773, 3848, 4346, 4010, 5124, 4725, 5183, 5221, 3242, 4156, 5329, 4720]. Hakes [3567, 2226, 2305]. Halcyon [886]. Halibut [3700, 3666, 3685, 3481, 1133, 4540, 4500, 3091, 4910, 4794, 136, 412, 4843, 5346, 5305, 5249, 4816, 2513, 4290, 105, 5035, 4943, 4446, 5235, 82, 5359, 609, 1951, 4268, 4912]. halibut-fisheries [609]. Haliotis [4801, 3489, 4167, 4877, 2126, 3527, 3693, 3510, 3970]. Haliporoides [2563]. Halland [775]. halleri [5370]. Halobates [2847]. hamatus [4066]. hammerhead [3434, 5225, 4985]. Hampshire [376, 650, 97, 154]. Hancock [3764, 3571]. hand [905]. hand-line [905]. Handline [3249, 2991]. Handling [2858, 4470, 3846]. Hanford [2959, 3546]. Hansen [2972]. haplotype [3808]. hapuku [4230]. harak [5184]. Harbor [3167, 3444, 3447, 936, 3065, 3207, 3823, 2642, 321, 206, 2791, 2765, 475, 566, 1598, 3166, 2284, 4848, 5067, 4390, 4193, 5293, 5136, 3679, 3983, 4166, 1498, 4944, 4980, 5068, 3100, 5168, 3820, 5314, 4629, 3161, 2822, 2856, 4638, 4201, 3220, 4029, 4813, 3245, 3092, 4239, 336, 335, 499]. Harbour [3207]. Hard [3101, 3553, 2814, 1820, 3106, 3226, 1803, 3375, 5156, 5330]. hard-bottom [5156]. hard-to-age [5330]. hardbottom [4257]. Hardy [3006]. Harengula [2337, 2574, 4405]. harengus [3199, 2570, 2045, 3130, 3306, 3712, 2264, 2208, 2015, 2893, 2015, 2900, 4032, 3001, 3199, 3383, 5006, 4098, 2975, 4074, 4513, 2641, 3712, 2264, 2208, 3802, 2531, 2893, 2491]. Harmonic [2033]. Harness [2519]. Harvest [2910, 2215, 5142, 5332, 4522, 4135, 4304, 3846, 4510, 4071, 669, 5052, 5330, 2714, 2006]. Harvestable [2652]. Harvestable-sized [2652]. Harvested [3415, 4283, 3734]. Harvesting [3638, 3365, 2326, 2745, 4747, 4315, 4646]. Harvests [3341]. Haswell [3516]. hatch [401]. hatchdate [4632]. hatched [206, 2775, 3798, 566, 238, 499, 291]. Hatcheries [2131, 2256, 715, 248, 910, 1403, 2006]. Hatchery [3269, 507, 2952, 3060, 3461, 3653, 185, 2753, 2287, 2627, 2716, 4243, 213, 1393, 446, 4543, 1057, 4273, 349, 720, 4950, 4940, 3885, 582, 4276, 1380]. Hatchery-Cultured [3060]. Hatchery-Reared [2753, 2716, 4243, 4940, 3885, 4276]. Hatchetfish [2767]. Hatching [3603, 610, 596, 822, 354, 663, 4795, 2761, 2580, 2353, 2385, 902, 642, 4034, 227, 150, 625, 714, 726, 1400, 1058, 835, 481, 482, 670, 2930, 710, 5018, 281, 5191, 4236, 947, 747, 628, 2760, 1396, 607, 654, 319, 282, 180, 203, 986]. Hatteras [2792, 2540, 2704, 2692, 2355, 2224, 3290, 612]. Haul [5168, 326, 3167, 3818, 3161]. Haul-out [5168, 3161]. haulback [4036]. Haute [576]. Havana [157]. Haverstraw [2623, 2784]. Havre [835, 670, 307, 319]. Hawaii [4558, 3926, 2225, 2301, 2484, 5257, 3312, 4349, 4849, 4952, 4543, 2335, 2728, 1330, 4044, 2556, 2650, 2698, 3557, 2145, 2991, 1860, 2247, 3082]. Hawaii-based [4349, 4849]. Hawaiian [5202, 4542, 4470, 3752, 4222, 1270, 831, 1453, 4034, 2671, 2851, 2963, 3346, 3517, 1292, 4026, 4542, 4688, 5007, 1675, 3937, 3552, 1296, 1293, 2148, 1194, 1186, 1262, 1258, 1724, 4082, 2454, 1295, 2424, 393, 3750, 1693, 3137, 2777, 1291, 1820, 3769, 4076, 2107, 1327, 3071, 3754, 1290, 1289, 2155, 1692, 1865, 2165, 3636, 3039, 3175, 3176, 1831, 2012, 2703, 1242, 2285, 2435, 1297, 1287, 1964, 2447, 3250, 2522, 1795, 2005, 2248, 2674, 4471]. Hawaiian-Caught [2165]. hawaiiensis [4197]. Hawk [615, 628, 170, 867, 909, 207, 429, 370, 401]. hawks [230]. Head [2908, 2033, 2609, 1106, 3332]. headed [2471]. Health [3452]. Hearing [1246]. Heart [3384, 2697]. Heat [2383]. Heated [2520, 2036]. Heavily [2710]. Heavy [2284, 5057, 5016, 3375, 2707]. hebraicum [4482]. Hecate [3862]. Heceta [4995, 4529, 3547, 3772]. Hector [3418, 4658]. hectori [4658]. Height [3538]. Held [2628, 2770, 2741, 68, 881]. helix [64, 390]. Helminth [2231, 3115]. Helminths [2197, 4834]. helvomaculatus [2706]. Hematodinium [4319]. hematology [2965, 4319]. Hemilepidotus [3203]. Hemiramphidae [4570]. Hemiramphus [4587]. Hemitripterus [2461]. hemocyanin [5317]. hemocyanins [3103]. Hemolymph [2843, 3344]. Henry [514]. Hepomadus [2232]. Heptacarpus [2922]. her [1109]. Herbst [3924]. herbstii [3107, 3099]. herding [4970, 4971, 5241]. heredity [1787]. Hereford [3237]. Heritable [2412]. Herklotsichthys [3493, 3082]. Hermaphrodite [2365, 5189, 3799]. hermaphroditic [4920]. hermit [1230]. Hermosilla [3534]. Herr [152]. Herring [3001, 2128, 2125, 2571, 3199, 3383, 2045, 2975, 3514, 2923, 2641, 2334, 2075, 2264, 2208, 1260, 2938, 2015, 2531, 2893, 2976, 2491, 3574, 3082, 294, 700, 2900, 4032, 581, 522, 523, 388, 3713, 161, 501, 655, 3224, 239, 1906, 644, 276, 539, 489, 4062, 587, 343, 344, 385, 541, 659, 864, 449, 2546, 4130, 2570, 295, 5006, 4098, 82, 232, 3130, 4074, 4696, 4513, 353, 4405, 2205, 3306, 4360, 1603, 1623, 1624, 1640, 3712, 5273, 777, 230, 1775, 3092, 3802, 245, 4341, 3740, 4434, 144, 1597, 934, 5025, 370]. herring-nets [449]. hertwigi [2748]. Heterocarpus [3312, 3053, 3362, 3750, 3322, 3754]. heteroclitus [3212, 2967, 2794]. heterogeneity [1727]. Heteroptera [2847]. Heterostichus [3330]. hexagrammid [1902]. Hexagrammidae [4172, 4360]. Hexanchus [4537]. Hickory [5108]. hierarchical [5261]. High [1798, 2246, 3263, 5377, 2430, 2836, 1935, 2933, 3831, 355, 1895, 4787, 1347, 1586, 3089, 4226, 3218, 4529, 1343, 5354, 4078, 4806, 1351]. high- [5354]. high-breakdown [4078]. High-energy [2836]. high-frequency [4787]. high-gradient [3089]. high-resolution [4529]. High-Seas [2246, 2430, 2933, 3831]. High-speed [1798]. High-Use [3263]. highlands [1644]. Highly [3636, 4174, 5219, 4031]. Hillsborough [2020]. hind [4952, 3742]. Hinnites [2695]. hints [237]. Hiodon [2228]. Hippidae [2643]. Hippocampus [4600, 4846, 103]. Hippoglossoides [4866, 4834, 4742, 4500, 3091, 4843, 4816, 5359, 4912, 3481]. Hippoglossus [4540, 4910, 4794, 5249, 2513, 4290, 5035, 5235, 1951]. Hippolytid [2646, 3620]. Hippolytidae [2922, 3170, 3193]. hippos [4351, 4837]. Hippospongia [5362]. hippurus [3926, 5379, 3882, 4292, 5148, 4487, 3096, 3282, 2837, 4336, 4879, 3250, 5185]. Hirudinea [2687, 1319]. Histiophorus [305]. Histochemical [2936, 3375]. Histological [2833, 2647, 3913, 1492, 3238, 3949]. Histology [3565, 3145, 1452, 3554]. historic [3114]. Historical [1735, 3521, 2667, 53, 2460, 3987, 5143, 2679, 4861, 3850, 4752]. histories [1625, 1391, 4030, 4219, 1653, 1384, 1527, 1530]. History [2466, 3563, 2802, 2995, 2895, 3513, 2608, 1026, 3023, 2506, 3388, 2105, 2263, 3219, 2265, 2307, 2110, 2571, 2574, 2790, 3504, 3135, 3334, 281, 3413, 2641, 2324, 3162, 2911, 3194, 654, 3684, 3311, 3493, 2976, 3574, 3975, 1559, 1511, 5312, 3950, 5303, 1937, 4364, 3163, 4562, 4928, 5281, 5372, 375, 1497, 5106, 838, 1510, 15, 3372, 5015, 4779, 1655, 1517, 3086, 242, 1906, 1227, 5340, 3700, 1654, 1073, 5151, 60, 4606, 1823, 4077, 4525, 1377, 1608, 4127, 1430, 1615, 1639, 2546, 302, 3089, 1944, 1184, 3904, 5294, 1148, 3692, 4993, 5140, 972, 4546, 5159, 3725, 3782, 3914]. history [1187, 3980, 4984, 3722, 2725, 3236, 1595, 1647, 3320, 3786, 5364, 2861, 1509, 1249, 1618, 1342, 1006, 1094, 1775, 3745, 1850, 4744, 5019, 3330, 1107, 1475, 5184, 4701, 1597, 4702, 1848, 4303, 3308, 1650, 1824, 1490, 1522, 2005, 2065, 3958, 5369]. hitherto [1071]. Hogfish [3020, 5150]. hoki [4068]. Hokkaido [3722, 4157, 869]. holbrooki [2179]. holding [2041, 5250]. Hole [390, 1182, 1443, 1444, 1190, 1226, 1248, 1562, 1473, 1304, 1451, 710, 1178, 1171, 1191, 1436, 1112, 1181, 1310]. Holland [191, 1260]. Holocephali [3886]. Holstein [959]. holtbyrnia [3670]. Homaridae [2053]. Homarinus [3924]. Homarus [3474, 3505, 3590, 3765, 4772, 5253, 4479, 4580, 2949, 3141, 3172, 3140, 4111, 2779, 2841, 4184, 3924, 2339, 2211, 2418, 2528, 2652, 3656, 3384, 2115, 3047, 4469, 677, 3400, 3256, 5286, 2897]. Home [5099, 3763, 5316]. Homing [2234, 3370, 3550, 2617, 2415, 3046, 1636, 2242, 3763, 3088, 2406]. Homogeneity [2633]. Hood [791]. Hook [3707, 375, 4871, 4992, 387, 5026, 533, 549, 4904, 5043, 5180, 556]. hook-and-line [4871, 5026]. hooked [3762]. hooking [4290]. hooks [543, 4992, 4693, 5232]. Hoplostethus [3876, 3791, 4125, 3958, 4119]. Horizontal [4477, 3601, 3615, 5200, 4503, 4803, 5023, 3784]. Hormonal [2381]. Hormonal-Induced [2381]. Hormone [2040, 2199, 2060, 2144, 2157]. Hormone- [2144]. Hormone-Induced [2040]. Hormuz [5121]. horned [1384]. horse [4008, 103]. horseshoe [3150, 3402, 4769, 4933, 4620, 4775, 5317, 4597, 4793, 4578]. Hosa [3740]. hosei [3052]. Host [2027, 3051, 3602, 1770, 2604, 5134]. host-parasite [2604]. hosts [3031, 1454]. Hot [2352]. hound [5219]. hour [3143]. hours [555]. house [464]. hubbsi [4725, 5088, 5183, 5221, 3583]. Hudson [2623, 2784, 2998, 2654, 748, 4129, 784, 1120, 3440, 1834, 4443, 4371, 4217, 679, 897, 1797, 191, 240, 100, 4679, 4320, 1760]. human [5094]. humans [4283]. Hume [507]. Hump [4719, 5161]. Humpback [2966, 3281, 3756, 4669, 2711, 4301, 3755, 4134, 3931, 4752]. Humpy [2626]. hundred [411, 299, 265, 527]. Hunter [4090]. Huron [1663, 1821, 1597]. Hutchinson [4515]. hyaline [4003]. hyaline-zone [4003]. Hyaloteuthis [4830]. Hyas [3710, 2992]. Hybrid [3327, 904, 2492, 188]. Hybridization [4521, 4576, 2662]. hybrids [1084, 2385]. hydra [1255]. Hydractinia [1965, 1963]. hydrated [5003]. hydraulic [1995]. Hydroacoustic [2253, 2976, 4009]. hydroacoustic-survey [4009]. Hydrocarbon [3432, 2683]. Hydrocarbons [2386, 2932, 2375, 2584, 2850, 2382]. hydrodynamic [5209]. Hydroelectric [2063, 3332]. Hydrographic [2367, 600, 5248, 832]. Hydrography [3310, 4344]. hydroid [1963]. Hydroids [1291, 1438, 1191]. Hydrolagus [2120, 2846]. Hydrological [1878, 1978, 2217, 2347, 5008, 5146]. Hydrophilus [1522]. Hydrostatic [2223]. Hydrothermal [3432]. Hydrous [1522]. hygiene [917]. Hymenopenaeus [2048, 2563]. Hyperiid [2491]. Hyperoche [2491]. Hyperoglyphe [4018]. Hyperoodon [4416]. Hyperostosic [3580]. Hypoconcha [2785]. Hypomesus [4819]. Hyporhamphus [4570]. Hyporthodus [5007, 5294]. hypotheses [3726]. Hypothesis [2934, 3478, 4382]. Hypoxia [3252, 5296, 4180]. hypoxic [5376]. hypsinotus [2459].

I. [2066]. Iberian [4081, 4828]. Ice [2628, 2770, 2038, 2741, 5168]. Iceland [412, 828, 5319, 602, 910, 609]. Icelinus [4928, 3366]. Ichthyologia [275]. ichthyology [267]. ichthyoneuston [2621]. Ichthyoplankton [3213, 5310, 4997, 5066, 3315, 2946, 3409, 4885, 4845, 3122, 3874, 3956, 4975, 5165, 4444, 4344, 3836]. Icichthys [2555]. icistia [2040, 2372, 2353]. Ictalurus [571, 5291]. Icy [1758]. Idaho [1503, 1093, 1106]. Identification [3806, 2660, 5181, 4392, 3582, 3193, 4668, 2583, 2109, 2866, 3180, 3557, 4339, 3311, 4887, 1955, 4054, 4560, 4388, 5303, 4501, 4617, 3840, 5101, 4804, 3488, 3878, 3463, 4731, 4214, 5208, 4785, 4738, 4554, 4509, 4538, 4638, 4337, 4178, 4063, 4203, 4421, 3907, 4043]. identified [3756, 241, 3776, 3598, 5163]. identify [4958, 3412, 5344, 4889]. Identifying [3616, 2987, 5247, 3827, 3927, 5378]. Idiosepius [3504, 3801]. Idotea [3355]. Idus [599]. if [124]. IGF1 [5174]. ignobilis [4788]. ignoring [4229]. II [1210, 3447, 4103, 3869, 3921, 2517, 2886, 1779, 2207, 2100, 1612, 1769, 1678]. III [2096, 2579, 1634]. Iliamna [2134]. illecebrosus [4211, 4923]. Illex [4309, 4211, 5133, 4923, 4247]. Illumina [5116]. illustrated [398]. Illustrations [716, 737, 951, 976, 1001, 1005, 1060, 1067, 1101, 1111, 1138, 1156, 1168, 1365, 554, 970, 1212, 1210, 1234, 1257, 1275, 1286, 1294, 1298, 1316, 1329, 1333, 1432]. illustrative [426]. imagery [4529]. Imaging [5066]. imbricata [2098]. immature [4410, 4586, 1669, 1933, 4515]. immediate [5194, 1306]. immersion [3138]. Immigration [2614]. immunochemical [1889]. Immunodiffusion [3237]. Immunologic [4063]. Impact [2352, 4543, 3225, 3651, 4745, 4012, 3072, 4238]. impacts [4973, 3070]. Impairment [2932, 5376, 4947, 4902, 4951]. impedance [5050, 4999]. Imperial [271]. impingement [4171]. implementation [5307]. implements [831, 10, 174]. Implication [3586]. Implications [3169, 4835, 2857, 3354, 3202, 3180, 3690, 3252, 3154, 4552, 4328, 5073, 5197, 4541, 5346, 5095, 2877, 4026, 4761, 5061, 4730, 3785, 4092, 4585, 4215, 4737, 1942, 5161, 4895, 4925, 4744, 4490, 4267, 5045]. Implied [3524]. Importance [3695, 2090, 4880, 1151, 4830, 2894]. Important [3561, 2887, 5065, 1990, 4167, 5195, 4494, 71, 1973, 235]. importation [682]. importing [865]. imports [317]. imposed [4691]. imprecision [5140]. impregnate [684]. impregnated [124]. impressa [3452]. imprinted [2617]. Imprinting [3550]. impropriety [335]. improve [5004, 4728, 5313, 5059, 5288, 5089, 5343, 4663]. Improved [2815, 3605, 943, 1421, 3272, 957, 1363]. improvement [1396, 1039, 1104]. Improvements [1193, 1045]. Improving [4501, 4816, 4394, 4251, 664]. Inauguration [501]. Incertae [2361]. Incidence [3172, 2759, 2528, 2748]. Incidental [4653, 3240, 4740, 3092, 3217, 3298, 3595, 4640, 5352, 2971, 3952]. incidentally [3545, 4228]. incipiency [261, 421]. incisor [1868]. incisus [4612]. inclosed [489]. inclosures [479]. includes [5246]. Including [3343, 2203, 5346, 4780, 5276]. inclusive [150, 230, 313, 300, 357, 134]. incognitus [5365]. Income [2581]. Incorporating [2535, 4863, 3706, 5218, 4558, 5261, 4269]. Incorporation [3818, 4221]. incorrect [5346]. Increase [2175, 489, 744, 256, 1141, 3289]. Increased [2472, 3284, 4719, 1127]. Increases [4957, 2694, 4918]. increasing [4126, 4210, 2479, 3930, 3989, 4419, 1367]. Increment [3257, 3181, 2967, 4343, 4124, 3739, 3367, 5373, 4101, 3138]. Increments [3435, 2455, 2975, 3581, 3514, 3071, 3680, 2435, 3493, 2893, 2896, 3195, 3134, 4311, 5265, 4877, 4882, 4101, 4354, 3159, 3499, 3138, 4616, 3250, 3393]. Incubation [2053, 2384, 308]. Incubators [2516, 2839]. independent [5065, 4772, 4994, 5276, 4467, 5033, 3459, 4824]. indeterminate [5061]. Index [98, 177, 247, 525, 674, 873, 930, 948, 967, 978, 996, 1010, 1022, 1074, 1096, 1110, 1155, 1163, 1175, 1209, 1229, 1254, 1270, 1272, 1285, 1288, 1303, 1321, 1323, 1336, 1386, 1433, 1441, 1445, 1460, 1468, 1483, 1487, 1504, 1508, 1512, 1525, 1534, 1568, 1575, 1591, 1599, 1604, 1616, 1633, 1649, 1658, 1697, 1682, 1712, 1715, 1733, 1749, 1774, 1780, 1814, 1826, 1849, 1870, 1900, 1921, 2190, 2416, 2523, 2600, 2787, 2880, 2939, 3010, 3112, 3179, 3227, 3326, 3396, 3469, 3560, 3623, 3992, 4059, 4143, 4189, 4278, 4363, 4438, 4520, 2420, 5133, 4770, 4951]. indexing [3937, 4255]. India [3723]. Indian [859, 5346, 2719, 1100, 146, 4354, 3732, 4468, 5210, 4011, 5078, 4666, 5083]. Indiana [48, 941, 1087, 131]. Indicate [3566, 4799]. Indicated [2240, 2554, 2435, 2896, 2565, 4405]. indicating [5363, 5274]. indication [248]. indications [2936]. indicative [1475]. Indicator [2239, 3848, 4688, 3458, 4894, 5213]. Indicators [3677, 3628, 4103, 5094, 2856, 3245]. Indices [4366, 2929, 3445, 3628, 3797, 5073, 3951, 4064, 3827, 4954, 4983, 3286, 2251, 5033, 2127, 3949, 3247]. indicus [3569]. Indirect [4553, 4698, 49]. indistinguishable [3770]. Individual [4070, 3513, 3582, 3706, 3425, 4651, 4870, 1529, 4717]. individually [3756, 3776, 3598]. individually-identified [3776]. individuals [4243, 3734]. indivisa [4407]. Indo [2854, 3516]. Indo-West [2854, 3516]. Indonesian [2667]. Induced [2811, 2040, 2891, 2062, 2381, 2691, 3892, 3652, 2532, 5241, 3238, 2458]. Inducements [125]. Induction [3416]. industries [1496, 920, 869, 952, 1312, 1046]. Industry [2512, 3267, 2371, 1349, 4426, 58, 1131, 903, 1592, 468, 598, 1021, 1064, 4110, 1167, 1237, 1566, 1013, 484, 911, 1121]. industry-developed [4110]. Infauna [2514]. infected [4319]. Infection [3397, 3254, 3403]. inference [4959, 4803]. inferences [4021, 5302]. Inferred [5143, 2160, 3347, 4916, 4948, 4034, 4576, 4732, 5331, 4055, 4293, 5117, 5270, 2205, 4955, 4457, 4116]. Inferring [3566, 3811]. Infestation [3596]. infesting [1037, 994, 972]. Inflation [3171, 2521, 2781, 3795]. Influence [2568, 5081, 2872, 2605, 5196, 3204, 2810, 3739, 4215, 4993, 4082, 2341, 3461, 1866, 1954, 2194, 1426, 3969, 92, 4094, 4882, 5294, 1024, 1434, 1467, 3946, 5146, 4163, 398, 5251, 3855]. Influenced [3432]. influences [5065, 1048, 3985, 1729, 5173]. Influencing [3508, 2015, 2987, 5053, 1694, 4675, 1626]. influx [3776]. Information [3431, 202, 919, 4830]. ingested [4178]. ingestion [2751, 3761]. inhabit [1160]. inhabiting [5136, 946, 1085, 5175, 4612, 982]. inhibiting [2610]. Inhibition [2426]. Inhibitors [2984]. Inhibitory [2754]. Initial [2875, 2051, 4296]. Initiation [3603]. injected [4114]. injecting [3387]. Injection [2441]. injured [924]. Injuries [901, 667, 2877]. injurious [228, 725]. injury [3410]. Inland [4716, 5136, 621, 752, 1135]. Inlet [3902, 2886, 3460, 3237, 3877, 2727, 4950, 4297, 4113]. Inlets [3385]. inner [1766, 4809, 5090]. inodorus [4077]. inquilinus [2466]. Inquiries [70]. inquiry [13, 55, 881]. Insecta [2847]. insecticides [2458]. insects [1518]. insensibility [27]. Inshore [2779, 2076, 2397, 5060, 2450, 3079, 2474, 3959, 5099, 4481, 4909, 2368, 4110, 4923]. Insights [5277, 5067, 3939, 5262, 5297]. Inspection [357, 806]. instability [3848]. instable [5074]. Installation [2063]. installment [363]. Instantaneous [3266, 4225]. Instar [3311]. instinct [1636]. Institution [396]. instruction [267, 881]. Instructions [391]. insular [3937]. insularum [2556]. Intake [2243]. intakes [1979]. integrated [5293, 4863]. Integrating [5004, 5165, 4547]. Integration [4529, 3509, 3499]. Integrity [3477]. Integumental [2261]. intensities [4020]. Intensity [2561, 4760, 1656, 5127]. Intensive [3525, 3261, 3322, 4564, 1423]. Intensively [2831]. interacting [3292]. Interaction [3401, 1961, 2371, 4096, 2682, 2794]. Interactions [2380, 4854, 5013, 2490, 2339, 2556, 4886, 2408, 3047, 5268, 3033, 5094, 4730, 3153, 4950]. Interactive [3134]. Interannual [3869, 4415, 3561, 3409, 3522, 3903, 3757, 3883, 3688, 5319, 5248, 4636, 5062]. intercept [4819]. interchange [1980]. Interdecadal [5058]. interest [953]. interested [164]. interests [276]. Interference [2815]. intergeneric [2888]. intermediate [2664]. intermedius [5075]. Intermingling [4500]. Internal [3478, 2077, 1404, 4388, 2130, 3106, 4131]. International [1347, 1130, 251, 325, 545, 1343, 1351, 102, 1355, 68]. internationally [2363]. Interoceanic [2298]. Interpretation [3180, 4678]. Interrelation [2973]. Interrelationships [3117, 2558]. interruptus [2339, 2378, 3602]. Intersex [2678]. interspecies [4063]. Interspecific [3650, 4154, 2135]. interstitialis [4332]. Intertidal [2183, 1763, 2475, 3067, 2650, 3491, 2714]. Intertropical [2336]. interval [5063]. intervals [5192]. Intra [5319, 5130]. Intra- [5319]. intra-estuarine [5130]. Intragravel [1980]. intraovarian [842]. Intrapopulation [4309, 5074]. intrarelationships [4192]. intraspecific [4779]. introduce [931]. introduced [4952, 360, 514, 883]. introducing [271]. Introduction [1541, 1543, 1557, 1704, 7, 713, 3082, 11, 1402, 942, 15, 663, 111, 6, 184, 330, 393, 270, 59, 24, 2351]. Invasion [2298]. invasive [5375, 5279, 5291]. Inverse [3738, 4877, 2370]. inverse-logistic [4877]. Inversions [2146]. Invertebrate [3126, 2467, 2836, 946, 999, 5224]. Invertebrates [3478, 2997, 4765]. investigate [3778]. Investigation [4576, 1227, 2117, 1598, 3523, 966, 984, 572, 1151, 852, 1563, 961, 4966, 3766, 3282, 1145, 1302, 1095, 278]. Investigations [1590, 5216, 2302, 1620, 912, 2700, 993, 1030, 1093, 1106, 1150, 1198, 457, 965, 1073, 1086, 1050, 1072, 1220, 1012, 185, 250, 963]. Investing [3154].

Investment [3431, 3801, 3423]. involves [3944]. Iowa [991, 1082, 666, 1499, 1530]. Ipswich [50, 66]. Ireland [1045]. iridescens [2247]. irideua [152]. irideus [85]. iris [3693]. Irish [1034, 390, 3203]. Iron [574]. irradians [3490]. irroratus [3068, 2504, 2652, 2765, 4679]. Isaacs [2306, 2056, 2835]. isenbeckii [3377]. ISIIS [5066]. isinglass [727]. Isistius [2086]. Island [3127, 2270, 717, 3273, 2804, 3097, 3449, 2052, 4919, 3126, 3606, 2380, 2105, 3864, 1931, 835, 1823, 521, 2490, 3457, 2666, 3254, 3407, 5100, 1656, 3619, 3668, 2106, 4609, 2438, 3763, 5321, 5339, 3978, 4105, 2427, 2727, 2955, 3146, 5137, 2572, 3148, 3400, 5164, 2713, 4752, 4515, 4922, 911, 869, 5162, 4840, 4784, 556, 2138, 3044, 3408, 2779, 2841, 1972, 795, 2499, 4972, 3701, 662, 827, 3322, 3487, 1162, 2033, 2609, 373, 3291, 595, 527, 3353]. islandica [2978, 2960, 3123, 3146, 3255, 2707]. Islands [5202, 4464, 5068, 3354, 5138, 3205, 4823, 1480, 4201, 1331, 1334, 4596, 1105, 2541, 2873, 729, 2926, 1453, 4936, 1292, 4542, 4470, 1903, 1322, 1296, 1293, 2265, 2254, 2632, 2395, 1194, 1186, 1262, 1258, 105, 4685, 4218, 2454, 1295, 3208, 4220, 3672, 3053, 3750, 1291, 3752, 4222, 3754, 2816, 1290, 1289, 4469, 2155, 1865, 1691, 2703, 5330, 5359, 1239, 1242, 1566, 2285, 2435, 1297, 1287, 4305, 4306, 1932]. Isle [1015, 777]. Isles [97]. isodon [2804, 4562]. Isoelectric [2583]. isolated [3454]. Isolation [2463, 2774, 3605, 3770]. isolepis [2820, 2832]. isopleths [4508]. isopod [3355, 2079, 1230]. isopoda [1199, 2079, 3031, 2604]. Isopods [1289, 1302]. Isopsetta [2820, 2832]. Isotope [3105, 5074]. isotopes [5208, 4735, 4107]. isotopic [4967]. issues [4010]. istiophorid [1778, 4509]. Istiophoridae [2892, 3939, 4942, 3407]. Istiophorus [2862, 4642, 4935, 1778, 2021, 4738, 3301, 4996]. Isurus [4729, 5200, 3294, 4333, 5302, 5076, 4921]. itajara [3729, 4758]. Italian [747, 580]. Italy [1389, 225, 1345]. Iterative [2875]. ITOs [3714]. IV [2633, 3877, 3881, 1631, 2396]. Ives [2032, 2000, 3339]. Ivory [3543].

J [4693, 5043, 5232, 5180]. J-hooks [5232]. Jack [3502, 2132, 2055, 2647, 1767, 5312, 1822, 1859, 4383, 2849, 3238, 5247]. jacket [327]. Jacks [3684, 4351, 4837]. jaguana [2337, 2574, 4405]. Jalisco [4726]. Jamaica [942, 3906]. James [5207, 363, 2374, 5087, 4371, 515]. January [271, 3766, 597, 676, 637, 823, 908]. Japan [4623, 490, 751, 3823, 4795, 4004, 869, 1251, 1357, 5341, 1361, 1309, 3722, 5128, 1427, 4157, 1066, 1238, 1312, 4436, 4207, 4380]. Japanese [3542, 1240, 2246, 2483, 1251, 2123, 5003, 5140, 4716, 3509, 4207, 4380]. japonica [4623, 5140, 2783]. japonicus [2618, 2790, 2532, 2140, 3796, 2697, 5254, 4636, 5038]. jar [281]. Jasus [2740, 4657, 4212, 4690, 4507]. Java [5078]. Jeboehlkia [3675]. Jenkins [2866]. Jerome [512, 573]. Jersey [1797, 453, 3563, 3975, 4856, 936, 3150, 3834, 1834, 25, 3237, 425, 4037, 1006, 4809, 552, 413, 750, 4514, 5090, 5308, 453]. jet [4114]. jet-injected [4114]. jettied [3962]. jew [373]. jew-fish [373]. jewfish [3729]. Joaquin [3086, 1340]. jobfish [4564, 5295, 5135]. John [608, 741, 3626, 2928, 5087, 3917]. johnii [5087]. Johns [5258, 1994, 1999, 1848, 401]. Johnston [3249]. Jolla [2432, 2307, 2178]. Jonah [2765]. Jordan [5144, 4913, 4881]. jordani [2202, 3892, 1009, 3652, 3846, 4255, 4526, 4960, 3692, 5292, 2545, 3687, 4005, 4550, 2515, 3076]. Joseph [569]. Josie [211]. Journal [470]. Jr. [352]. Juan [3251, 5068, 4201]. jubatus [4706, 4730, 4464, 5114, 4586, 3359, 3989, 2927, 4825, 5024, 5366, 5358, 4665, 4838, 4551, 4664]. Judd [3229]. judgment [4982]. July [3057, 626, 294, 403, 1012, 1960, 605, 830]. jumbo [5243, 4704]. June [3965, 294, 2094, 3057, 403, 1158, 626, 3015, 248, 231, 317, 701, 846, 893]. jurisdiction [425, 453]. Jutland [926]. Juvenile [3502, 2096, 3474, 3590, 2219, 2520, 4624, 3536, 2179, 2608, 2773, 3546, 2404, 2019, 2084, 3492, 3578, 3472, 1630, 2623, 3611, 2755, 2252, 2134, 3666, 2023, 2577, 3656, 2341, 2499, 3339, 3436, 3181, 5374, 3528, 3015, 2332, 2262, 2970, 2402, 3428, 1691, 2026, 3214, 3166, 3696, 2916, 3311, 1670, 4773, 3350, 3688, 2083, 3947, 5296, 3963, 5048, 3505, 4243, 2597, 2945, 3195, 4936, 4989, 2903, 4477, 4938, 3634, 3731, 3817, 1886, 2474, 4244, 4643, 4311, 3895, 3412, 3898, 4796, 5167, 4781, 4780, 4728, 1675, 3773, 3937, 4977, 4734, 4749, 3945, 4069, 5351, 4014, 4254, 3785, 4967, 3917, 3463, 3333]. juvenile [4795, 4507, 3000, 5262, 5154, 4256, 4019, 4198, 4659, 4820, 4153, 3658, 4785, 4724, 4445, 1979, 4943, 3810, 4484, 5198, 2650, 5306, 3153, 3117, 3088, 4568, 5321, 2795, 3006, 4447, 3875, 3491, 5127, 4139, 4105, 2079, 3643, 4101, 3557, 4076, 5135, 3548, 3637, 4104, 5230, 4221, 4344, 4359, 4632, 4237, 3757, 3962, 2686, 3940, 4844, 3476, 4811, 4572, 4037, 3988, 5008, 4039, 3824, 3138, 4180, 3034, 4377, 1905, 3885, 3230, 2648, 4650, 1999, 4186, 4592, 4114, 4681, 4268, 5277, 4131, 4651, 5062, 2639, 5076, 3758, 5021, 4276, 3830, 3393]. Juveniles [2657, 2715, 2788, 2908, 3119, 2345, 3653, 2967, 2706, 3534, 3803, 3806, 5249, 3129, 4788, 4778, 4696, 4041, 3121].

K. [3403, 1868]. Kachemak [2626, 2646, 2709, 2775]. Kadiak [448, 1162]. Kalamazoo [949]. Kalman [4053]. Kamchatka [1442, 3248, 4017]. kamtschatkana [3489]. Karluk [1852, 2257, 1662, 2105, 1899, 1590, 1823, 1629, 4954, 3446, 1804, 2255]. kate [555]. Katsuwonus [4403, 5212, 3198, 2669, 3477, 2373, 2895, 2744, 1675, 2898, 2383, 3329, 2968, 3202, 3265, 4420, 4851, 4186, 1964, 2447, 2896, 1670, 2920, 2151, 2065]. Kawakawa [3390]. Kay [37]. keeping [938]. Keitel [395]. Kelp [3309, 2423, 3126, 2380, 2509, 2823, 3125, 2638, 2324, 3214, 2825, 4835, 4030]. kelpfish [3330]. kelts [5263]. Kemp [3858, 4678]. kempi [2867]. kempii [3858, 4132, 4678]. Kemps [4132]. Kennebec [890]. kenojei [5270]. Kensington [663]. Kentucky [1000, 1491, 980, 977]. Keokuk [1601, 1602]. kept [110, 307]. keta [2417, 3011, 3480, 4054, 4893, 4934, 4977, 3528, 4950, 5169]. Key [1535, 1710, 3356, 2070, 3620, 2982, 1939, 5338, 3516, 3807, 4766, 1571, 157]. Keys [4164, 3733, 4040, 4109, 4195, 4644, 417, 4179, 4163, 1533, 1707, 3825, 4577, 5028]. Khor [2267]. Kidd [2306, 2056, 2835]. Kidney [2786]. kill [4640, 3952]. killed [788, 2659, 836, 793]. Killer [2069, 2072, 3955, 4293, 4823, 5042, 2801, 3507]. killifish [2327]. Kilohertz [2780]. Kimberley [4549]. Kinds [2044, 2191, 410, 94]. Kinematics [3024]. King [3014, 3486, 3321, 3283, 2316, 2969, 3254, 2911, 3592, 3699, 3098, 3009, 4884, 3437, 4138, 4366, 4506, 1452, 1471, 1469, 1465, 2104, 3042, 3854, 4445, 5100, 5158, 3832, 2286, 4107, 5069, 3034, 5330, 955, 3075, 3302, 4205, 3640, 4875, 3669, 1873, 4683, 4782, 3929, 4151, 4510]. king-of-the-salmon [3034]. Kingfish [5106, 627, 648, 3157, 4286]. kingfishers [491, 230]. kingfishes [1638]. Kiribati [3859]. kishi [4414]. kisutch [3009, 3014, 2242, 3884, 3408, 4689, 4442, 2952, 2826, 5233, 2256, 2442, 3088, 2557, 3442, 3550, 3398, 2425, 2753, 4203, 4592, 4043, 3046, 2786]. Kitty [615]. Klamath [3770, 1144]. Klip [729, 353]. Klip-fish [729, 353]. klsutch [2506]. knife [870]. knobbed [4862]. know [1033]. Knowledge [2888, 120, 992]. Known [2238, 3675, 1092, 1123, 1219, 1491, 1160]. kob [4077]. Kodiak [2105, 3254, 5100, 1823, 2666, 4105, 2286, 5137, 4340, 4752]. Kogia [3683, 615]. Koko [2033, 2609]. Kona [2335]. Krill [3422]. Kröyer [3708, 2116, 2089]. Kudoa [3403]. Kure [2728]. Kuroshio [2653, 4636]. Kurren [395]. Kuwait [4272]. Kvichak [1861, 2197]. Kyphosid [3534]. Kyphosus [1868].

L [170, 4039, 120]. L. [4655, 3394, 4243, 3001, 5365, 575, 3823, 4042, 3300, 3260, 4008, 2975, 4093, 4610, 4739, 3491, 4025, 3323, 4597, 3220, 4789, 2003, 2208, 3092, 4534, 3966, 2002, 4407]. Labeling [3137, 2375]. Labial [1926]. Laboratory [2915, 2622, 3355, 2386, 3201, 2032, 2034, 2688, 2857, 3333, 2459, 2969, 3171, 2078, 2110, 3327, 2452, 2162, 2168, 2972, 2339, 2126, 2064, 3581, 1959, 2981, 2140, 2180, 2586, 2200, 2476, 2743, 3538, 3526, 2945, 4414, 2700, 5265, 2533, 3170, 3798, 4819, 3516, 2690, 4447, 3643, 4630, 3976, 2686, 2737, 4300, 4086, 3121, 3830, 5250, 1050]. laboratory-hatched [3798]. laboratory-holding [5250]. Laboratory-Reared [3581, 2981, 2622, 2476, 5265, 4819, 4447, 3976, 4086]. Labrador [3765, 3381, 148, 633]. labrid [4920]. Labridae [4441, 4123, 3020, 1186, 2586, 5150, 2365, 3067]. Lacépède [3864, 1887]. lachne [5362]. Lachnolaimus [3020, 5150]. Lack [3712, 4433]. lacking [4903, 3262]. Lactating [3284]. lady [4679]. ladyfish [4430]. laevigatus [3312, 3750, 3322, 3754]. Lagenodelphis [3052]. Lagenorhynchus [4160, 4946, 2659, 3899, 2882, 2797, 3833, 2789]. Lago [2066]. Lagodon [2010, 2464, 4155, 4512, 2817, 2828]. Lagoon [3167, 3866, 2632, 3296, 4666, 5079, 2728, 4609, 1856, 3407, 3875, 5184]. Lagoons [3543]. Lake [376, 227, 1821, 111, 2134, 1085, 1629, 1819, 2906, 462, 2499, 4954, 1762, 1830, 1507, 1464, 1768, 1775, 1503, 1824, 361, 559, 415, 904, 1581, 1789, 1906, 55, 1683, 725, 1698, 1567, 1597, 4623, 2096, 1937, 1789, 656, 576, 1906, 1903, 411, 1625, 2573, 946, 1823, 2265, 1683, 1192, 1688, 638, 1325, 1021, 2100, 972, 3040, 431, 1253, 1904, 2113, 1232, 2262, 518, 989, 1239, 1250, 1461, 427, 1142, 1597, 1843, 1992, 1934]. Lakes [1457, 1506, 1828, 1852, 1326, 359, 1385, 1003, 1030, 1644, 953, 1526, 1485, 752, 1500, 1489, 1501, 1764, 1768, 1775, 1339, 874, 227, 1354, 1350, 1219, 946, 1189, 1425, 1589, 372, 1029, 1319, 1018, 1358, 77, 1663]. lalandi [4286]. lalandii [4690, 4507]. lamellibranchiate [992]. Lamna [4312, 4729, 4774, 4528, 4486, 5274]. lamnid [4538]. lamnoids [4333]. Lampanyctus [5342]. Lampetra [3037, 5134]. Lampfish [2780]. lamprey [3037, 1770, 1854, 604, 3165, 3268, 5251, 1824]. lampreys [1536, 137, 3297, 1140, 5134]. Lampris [5365, 928]. Lampsilis [1528]. Lance [3388, 2718, 5187, 3890, 2711, 2224]. lancelets [1756]. lanceolatus [2803]. Lances [3594, 5128]. lancetfish [4495]. land [361, 111, 5377, 3100, 2708, 114]. land-based [5377, 3100]. land-locked [361, 111, 114]. landed [3374, 5019]. Landing [1696, 865]. Landings [3026, 3633, 3264, 2987, 4325, 4933, 3737, 1831]. landlocked [350, 408, 362, 650, 459, 419]. lane [4857]. Lanternfish [2102, 2154, 5310]. Lanternfishes [2225, 3349, 3494, 2154]. lanthanum [3884]. lapilli [5265]. lapillus [3898]. Lapland [878]. Large [326, 3094, 3684, 5004, 4311, 5101, 497, 4073, 4504, 2188, 3387, 4939, 4079, 4214, 2814, 3994, 4800, 5350, 2835, 2955, 5012, 5126, 1137, 4402, 4574, 4594, 2695, 455, 742]. large-format [3994]. large-mouthed [1137, 455]. large-scale [4311, 4079]. larger [1500]. Larimus [2803, 3139, 3152]. larva [2723, 3675]. Larvacea [2738]. Larvae [2293, 2044, 2191, 2992, 2568, 2349, 2618, 2622, 2524, 3694, 2206, 3001, 2988, 2619, 2314, 3366, 3324, 3582, 2316, 2626, 2646, 2709, 2752, 2969, 3193, 3613, 3171, 3455, 2337, 2393, 3327, 3524, 2863, 2449, 2521, 2318, 2192, 2367, 2715, 3586, 4596, 2788, 2908, 3008, 2397, 2580, 3407, 2139, 2064, 3266, 2350, 3698, 3581, 3500, 3533, 3520, 2883, 2051, 2372, 3310, 3537, 2733, 2299, 3514, 2721, 2345, 2164, 2833, 3685, 2545, 2967, 2073, 3616, 2244, 2706, 2820, 2888, 2890, 3005, 3076, 2237, 2824, 3534, 2710, 2647, 3511, 2491, 1679, 2989, 2809, 1767, 1773, 1832, 4499, 2497, 2900, 5342]. larvae [3414, 3934, 3803, 3806, 4541, 1806, 1907, 1996, 3542, 3880, 4028, 5324, 1859, 4012, 4908, 4889, 2201, 3664, 4150, 3787, 3144, 3129, 1513, 3104, 3516, 3777, 3367, 3893, 2340, 1845, 5018, 4788, 2152, 3059, 5157, 5032, 2884, 3203, 1802, 3930, 1901, 4294, 1993, 2754, 4890, 2936, 3323, 4222, 4456, 4742, 3305, 1598, 4432, 1875, 2224, 2552, 2832, 4260, 3946, 2766, 5260, 4339, 4716, 4996, 3539, 2498, 3485, 1652, 3095, 5163, 2902, 2943, 5038, 4817]. Larval [5002, 4280, 2191, 2723, 1784, 1813, 2915, 1912, 3195, 3093, 3397, 2455, 3418, 3929, 3536, 4414, 3020, 2773, 3420, 3965, 3768, 3882, 3860, 4350, 3204, 2216, 2688, 2128, 2125, 3128, 3331, 3611, 3635, 2022, 2533, 2752, 4368, 3613, 2153, 3429, 2595, 3199, 3257, 3791, 3380, 3579, 3586, 2362, 2813, 2972, 2785, 2577, 2881, 2956, 2690, 2126, 2341, 3530, 2771, 2470, 2975, 2883, 2981, 2109, 2353, 4465, 3119, 2721, 3518, 2302, 3189, 3544, 2539, 4570, 3525, 4630, 2288, 2200, 2848, 4344, 4598, 2803, 2334, 4962, 2204, 2737, 2743, 2575, 2378, 3478, 3428, 2624, 3802, 3389, 2858, 3538, 3132, 2893, 4701]. Larval [2961, 2338, 3347, 3417, 3515, 3519, 2756, 3523, 3526, 3743, 1759, 4021, 1528, 5310, 4032, 1519, 4103, 3817, 1886, 3713, 1939, 4796, 3473, 3746, 4734, 4749, 5265, 4672, 3781, 3785, 4121, 4366, 3115, 4313, 4124, 3276, 4750, 4113, 3608, 2930, 1637, 2645, 2147, 3516, 5262, 1463, 1479, 1494, 5154, 5209, 3927, 4019, 4198, 4659, 4096, 4395, 3760, 2676, 4217, 4724, 956, 5186, 3810, 4738, 4695, 1969, 3130, 5339, 4236, 2538, 4447, 2864, 3426, 3901, 1956, 4139, 4374, 5063, 3643, 4352, 3548, 3942, 4185, 4104, 4275, 5089, 3976, 4359, 4632, 4237, 4432, 4005, 4550, 3920, 2260, 2398, 3940, 4940]. larval [3988, 2394, 3022, 3138, 2587, 2691, 4209, 3738, 3718, 4041, 1838, 2849, 3369, 3949, 2233, 4681, 4049, 4887, 3223, 3873, 4514, 4207, 5251, 5361, 4051]. larvel [3330]. Lasker [3535, 3532]. last [4840]. Late [3611, 4920, 5324, 5262, 5233, 4066, 3643, 1342]. Late-Larval [3611, 3643]. late-stage [4066]. latent [4459]. Lateral [2346, 1299]. Lateral-Line [2346, 1299]. Lates [4379]. latifrons [3817, 3670]. latitude [3473, 4710]. latitudes [4806]. Latitudinal [3524, 4637, 3845, 5087, 5306]. Latreille [3835]. Latris [4700]. lavage [2591]. law [372]. Lawrence [3040, 383, 494, 4479, 3424, 4763, 3225, 4834, 4339]. Laws [621, 1192, 1029]. Lay [2120]. Layer [2780, 5209, 2583]. Layers [2397, 3137]. Laysan [1296]. Lead [2356, 2480]. leader [5210]. leading [281, 5286]. learn [124]. learned [2557]. Least [2354, 2525, 4133]. Least-Squares [2354, 4133]. Leather [327, 2098, 38]. Leather-jacket [327]. Leatherback [3761, 4847, 5352, 4653]. Leatherjacket [2179]. Lebbeus [2646, 2922]. Lebranche [5378]. Lee [3978]. Leech [564]. leeches [2687, 64, 1208]. Leeward [1296]. Legal [2652, 2371, 4510]. legal-size [4510]. Legal-sized [2652]. legislation [1359, 397, 1047]. Legislature [210]. Leidy [972]. leidyi [3662]. Leiostomus [4734, 3781, 3455, 2464, 3059, 3392, 2848, 3389, 3223]. Lemaitre [4258]. Length [2778, 3467, 4749, 4729, 3702, 4526, 3283, 2120, 2163, 3957, 4812, 3356, 2321, 5035, 5328, 2599, 3479, 3633, 3799, 3667, 3625, 2083, 2684, 5004, 5016, 4843, 3157, 1470, 4688, 3360, 3915, 1990, 4826, 4075, 4394, 3825, 3168, 2651, 3807, 5182, 2690, 3857, 3025, 5218, 4510, 4295, 5201, 4399, 3942, 4185, 4456, 5267, 3754, 4751, 4577, 4901, 5356, 3828, 4717, 4636, 5264, 2428, 4753, 5131, 4746, 3897, 3998, 5329, 4045]. length- [4746]. length-age [3168]. length-at-age [5004, 4510, 4636]. length-at-maturity [4901]. Length-based [3799, 3667, 5218]. Length-frequency [3467, 3157, 3915, 4826, 4717, 3998]. length-mode [5356]. Length-specific [4812]. Length-width-weight [2599]. Lengths [2128, 2656, 3142, 4393]. Leon [453]. Leone [2073]. Leopard [2620, 3587, 3981, 4556]. leopardus [4799, 3864, 3981, 4402]. Lepidochelys [3858, 4132, 2867, 4555, 4678]. Lepidophthalmus [3695, 4258]. Lepidopsetta [5355, 5018, 4932, 4352, 4407]. Lepomis [2342]. Leptinogaster [3260]. Leptocephali [2311, 1235, 2379, 3090]. leptocephalus [1235]. leptocerus [3194]. Leptocottus [3963]. Lernaea [1490]. Lernaeopodidae [4407]. lesions [2594]. Leslie [4314, 3261]. Lesser [2052]. lessoniana [3573, 4267]. LeSueur [1597, 4377, 3291]. lethostigma [4453, 5117, 3976]. lethrinid [4354]. Lethrinus [4757, 4354, 5184]. leucas [4404, 2072, 4371]. Leucichthys [1644, 1597]. Leuciscus [1305]. leucopsarus [4493]. Leucoraja [4573]. leucosteus [3553, 3018]. Leuresthes [2051, 2662, 2266, 3085]. Leuroglossus [4493, 3036]. Levant [2298]. Level [3513, 3116, 2155, 3752, 3863, 4957, 1961]. Levels [2236, 2091, 3255, 4242, 4021, 5174, 5253, 4462, 4174, 5317, 4891]. Leven [883, 419]. levis [4561]. lewini [3434]. liberated [1190]. Libinia [2622, 2595]. lida [3732]. lidar [4330]. Lieut. [170]. Life [2466, 3563, 2802, 2995, 396, 2895, 1937, 3449, 4562, 3513, 2608, 5372, 1497, 3023, 2506, 3388, 3086, 2105, 1906, 3424, 2263, 3219, 1823, 4525, 2265, 2358, 2110, 2571, 2574, 3089, 2790, 3504, 1944, 3266, 4546, 5041, 3135, 3334, 198, 2725, 2545, 2798, 2641, 2870, 3162, 1775, 2911, 4744, 3330, 3194, 3684, 3311, 3493, 5184, 1597, 3574, 1824, 1522, 3314, 5369, 3975, 5255, 5312, 5348, 4070, 3950, 5303, 3037, 3977, 4364, 3163, 4928, 5281, 5106, 480, 3372, 5015, 4779, 1657, 5340, 1625, 508, 4285, 3700, 5151, 60, 4606, 4077, 4127, 1615, 1639, 520, 1846, 1391, 2546, 4988, 1184, 3904]. life [5294, 3692, 4993, 4022, 5140, 972, 4030, 4219, 5159, 3725, 3782, 3914, 3980, 4115, 4984, 3722, 1480, 3236, 1647, 1653, 3320, 3786, 5364, 4789, 1384, 2861, 1618, 4063, 464, 3745, 5019, 5268, 1475, 4701, 4702, 1848, 4303, 1527, 3308, 1650, 1490, 1530, 2005, 2065, 770]. life-history [5015, 60]. life-needs [770]. life-saving [464]. Life-stage [5041]. Life-Stage-Specific [3266]. lifespan [4859]. lift [4691]. Ligament [1919, 1940]. light [4910, 3803, 5280, 4724, 1024, 5235, 492, 5127, 4806, 464, 611, 2233]. light-aggregation [3803]. light-based [4806]. light-house [464]. light-trap [4724]. lighting [422]. Lights [2015]. lignorum [1529]. like [321, 2144, 2157, 1908, 1679]. Likelihood [2782, 4623, 4604, 4717, 4225]. Limanda [5355, 3109, 2607, 2863, 3058, 4382, 3043, 4397, 1816, 2578, 2691, 2624, 5241, 5149, 5085]. limbatus [5048, 3559, 5323, 5232, 3995]. limitation [294, 4622, 276]. Limitations [3770, 1643, 3942, 1142]. Limited [5098, 3431, 104]. Limiting [3627]. Limits [3210, 2898, 248, 3970]. Limnological [1532, 1629, 1762, 1764, 1457, 1506]. limnology [2001]. Limnoria [1529]. Limulus [3150, 3402, 4769, 4933, 4620, 4775, 5317, 4597, 4793, 4578]. Line [2346, 3549, 3376, 5202, 905, 4658, 4871, 4564, 5026, 1299]. line-transect [5202, 4658]. Linear [2693, 2047, 2745, 3744, 5316]. Linear-programming [2047]. lineata [4700]. lineatus [3643, 4135, 2505, 2078, 3135, 2454, 1689, 338, 8, 3022, 2989, 75]. lined [4600]. Lines [3123]. ling [435]. Lingcod [3641, 3348, 3763, 4431, 5045, 3816]. Link [2331, 3233]. Linkages [3267, 5374]. linked [5200]. Linn [1650]. Linn. [1653]. Linnaeus [3003, 5266, 3648, 3967, 5075, 1825, 1823, 3954, 2192, 4256, 1868, 3095, 1815]. Linne [3123, 3146, 3255]. Linseed [2344]. liolepis [3685]. Lion [3359, 2927, 4329, 4706, 4625, 3989, 4825, 5358, 4745]. lionfish [5279]. Lions [2977, 3127, 3619, 3668, 4730, 4464, 5114, 4586, 4713, 5024, 5366, 3845, 4665, 4838, 4551, 4067, 4664]. Liopsetta [2900, 2881]. Liparidae [4641]. Liparis [2466, 3520]. Lipid [2061]. lipids [2319]. lipofuscin [3426, 3443]. lipofuscins [4565]. Liquid [2177]. Lironeca [3031]. List [1, 112, 1212, 1210, 1275, 1286, 1294, 1284, 1219, 194, 169, 182, 1085, 1251, 1305, 1015, 157, 1160, 2018, 179, 1092, 1531, 1584, 74, 1320, 68, 1100, 855, 214, 296, 248, 264, 329, 676, 1499]. lists [437, 977, 982]. literature [392]. Lithodes [3283, 2969, 5158, 5330, 2911, 3302, 4205]. lithodid [3144]. Lithodidae [3144, 3798, 3832]. Litopenaeus [2390, 5198, 4651]. Little [2810, 433, 1123, 662, 3149, 328]. little-known [1123]. Littleneck [2259]. Littoral [3007, 2657, 5115, 1939, 2689]. litulon [4422]. Live [25, 705, 117, 4748, 4618, 489, 2188, 1377, 1383, 279, 1418, 4227, 4456, 1172, 1567, 616, 1833]. live-bait [4227]. lived [4267]. Liver [2055, 2936]. livers [323, 306]. lives [555]. Living [1807, 3260, 2561, 1736, 2440, 931, 881, 423]. liza [5208, 5378]. Lizard [3864, 3407]. lizardfish [5075, 4909]. Lizardfishes [2346]. Llmanda [3415]. Lobate [3662]. Lobotes [3860]. Lobotidae [3860]. Lobster [3474, 3216, 2779, 2277, 2318, 2339, 2211, 2418, 2652, 3645, 3002, 2115, 3047, 3636, 2378, 3175, 3602, 3344, 2740, 2366, 4657, 4212, 4772, 5253, 1174, 4479, 4580, 1517, 726, 2877, 3779, 4542, 4470, 3773, 4929, 2949, 3140, 4111, 1370, 2841, 4889, 4690, 4507, 1036, 1103, 1146, 1236, 1430, 4215, 3924, 1033, 1344, 4855, 3384, 863, 3769, 3752, 3863, 4222, 5364, 857, 4469, 4384, 677, 3256, 5286, 2897]. lobster-handling [4470]. lobster-trap [4470]. Lobsters [3590, 1370, 3472, 3427, 3225, 2528, 3656, 2053, 3400, 2731, 3556, 3505, 3765, 1190, 5253, 714, 610, 822, 4688, 2309, 3141, 3172, 3241, 4184, 3835, 438, 944, 527, 675, 5123]. Local [4460, 2584, 4896, 5162]. localities [155, 3165, 877, 885, 893, 908, 916, 925]. Locality [2012, 1571]. Localized [2582]. Location [3673, 582, 2994, 4003]. Locations [2910, 2415, 3660, 4041, 3046]. Loch [883, 419]. loci [4576, 5341, 4955]. locked [361, 111, 114]. Lockington [3178]. lockingtoni [2555]. locomotion [4300]. locomotor [2555]. log [4147, 5193]. log-skew- [5193]. logbook [5194, 4483]. Logbooks [2160]. Loggerhead [3385, 3337, 5039, 5065, 4410, 4601, 4653, 2867, 4697, 4555, 4678, 4615, 4515]. logistic [4877]. Loire [576]. Loliginid [3573, 3801, 3985]. Loliginidae [3985, 3875, 4267, 3554, 3686]. Loligo [4024, 4211, 4652, 2876, 3333, 4481, 2452, 2649, 3801, 3985, 4647, 2999, 4676, 4039, 3148, 4923, 3314, 4755]. Lolliguncula [3686]. London [251, 545, 302, 325]. Long [3474, 4357, 3731, 3453, 3570, 2467, 2536, 2999, 2727, 3232, 5188, 3063, 2852, 4024, 5028, 4846, 2965, 3902, 2876, 4464, 4545, 616, 717, 3097, 3449, 2138, 3606, 1656, 5321, 5339, 2427, 2955, 3146, 3400, 595, 527, 4922]. Long-distance [4357]. Long-Finned [2999, 4024, 2876]. long-snouted [4846]. Long-Term [2467, 3063, 3474, 3731, 3453, 2536, 3232, 5188, 2852, 5028, 2965, 3902, 4464, 4545]. Longbill [3545]. Longevity [2986, 2104, 3301, 3762, 4920, 4723, 3118, 4680, 4240]. longfin [3028, 4481, 5200, 3294, 3203, 4923]. longhorn [4494, 3110]. longimanus [5220]. longirostris [2811, 3717, 4367, 2821, 2830, 3053, 3137, 2777, 2145, 2107, 2601, 3690, 65, 3850]. Longitudinal [4483]. Longline [4748, 3707, 2422, 5057, 3762, 4794, 5194, 4349, 4849, 5196, 5284, 4797, 4854, 4621, 4791, 4653, 2682, 5035, 5350, 4588, 5026, 5269, 1987, 4981, 4555, 4904, 5210, 1831, 3761, 5051, 4406, 4635, 4228, 5259, 4539, 3955]. Longline-caught [4748, 3762, 4981]. longlines [4231]. longnose [4913, 4495]. Longspine [2990, 4687]. Lontra [5375]. Lookout [2192, 3226, 607, 765, 572, 2237]. Loop [4096]. Lophiid [3270]. Lophiiform [2925]. Lophiiformes [3724]. Lophius [4860, 3724, 5301, 5093, 5197, 3270, 5356, 4422]. Lophobranchiates [103]. Lopholatilus [3200, 3496, 3041, 1306, 5077, 1308, 3102]. Lopholithodes [3798]. Lopkolatilus [3465]. lord [3203]. Loss [2437, 2386, 480, 508, 2246, 2053, 3788, 2068, 1931, 4088, 3720]. losses [2366]. lost [4635]. Louis [2330]. Louisiana [62, 5142, 3644, 2772, 3315, 4711, 3813, 5009, 1642, 3252, 4110, 3639, 4451, 1121]. Louisianaa [3420]. Low [2063, 2898, 4174, 4156, 5313, 5354, 3332]. Low-frequency [4156]. low-head [3332]. Low-oxygen [2898]. low-quality [5354]. Low-Temperature [2063]. lower [4452, 4441, 3878, 204, 2886, 4675, 3914, 2327, 1659, 4297, 985, 5234, 5062, 4575, 370, 3131]. lowlands [950]. lucetia [1809]. Lucia [3569]. lucius [5265]. Luis [1088]. lumpfish [5319, 3491]. lumpsucker [5336, 5252]. lumpus [5319, 3491]. Lunar [2698, 4883, 3212, 4981, 3406]. lunulatus [5289]. lupus [5132]. lurida [1637]. luteola [1528]. lutherii [2754]. Lutianidae [1611]. lutjanid [4354, 4063]. Lutjanidae [2017, 4026, 4721, 3933, 3843, 2604]. lutjanids [4991]. Lutjaninae [3840]. Lutjanus [4498, 5271, 4412, 4719, 5113, 4348, 5087, 5028, 2773, 3735, 3809, 4350, 4671, 4711, 4967, 3000, 4628, 3974, 4025, 4448, 5267, 4857, 4177, 5111, 4762, 4955, 5343, 5064, 4266, 4650, 4918, 4451]. lutris [4708, 2631]. Luzon [1331, 1239]. Lyopsetta [2665]. lyricus [1630].

M [531, 3107, 3317, 3851]. M. [1907, 4289, 4701]. Maatjes [2075]. Maatjes-Cured [2075]. maccoyii [4311, 4423, 4503, 4148, 3599, 3706, 3709, 4545]. macdonaldi [2487, 2159, 4112]. MacGilchrist [3053]. Mackenzie [1078]. Mackerel [3502, 2349, 2637, 3486, 3399, 3321, 2790, 285, 2132, 2796, 122, 2140, 3651, 2055, 3592, 3699, 2647, 3098, 294, 1767, 3996, 4190, 778, 3793, 2859, 4884, 41, 135, 205, 469, 485, 556, 383, 494, 655, 927, 912, 3715, 4138, 644, 354, 1675, 1859, 4012, 4366, 4506, 1034, 3042, 3854, 4008, 1360, 1845, 4869, 4868, 2532, 4172, 4714, 87, 82, 101, 5, 168, 229, 3930, 4095, 4818, 421, 1132, 4104, 219, 3796, 5309, 2697, 4107, 61, 3826, 1661, 1678, 4716, 196, 300, 1164, 3669, 246, 2849, 3238, 4663, 786, 865, 5247, 4636, 5038, 4305, 4306, 4664, 4782]. Mackerels [2049, 3177, 913, 4202, 1671]. mackini [3453]. macleayi [4937]. Macquaria [4701]. macro [5224]. macro-invertebrate [5224]. Macrobenthic [2312]. Macrobrachium [3089, 2701]. macrocephala [2170]. macrocephalus [3228, 5016, 2869, 4896, 4730, 4988, 2666, 3804, 3405, 3908, 3561, 4906, 2883, 4698, 3905, 4853, 5179, 4654]. macrochirus [2342]. Macrocrustaceans [2443]. Macrocystis [3309, 3126, 2324]. macroinvertebrate [4960]. Macroinvertebrates [2217, 2347, 2558, 3279]. macrolepidotus [3086]. macrophyte [3120]. macropterus [1689, 1693, 3907, 1670]. macrorhynchus [2735]. Macroscopic [2198]. Macrouridae [1251, 3243]. Macrozooplankton [2450, 1970, 3562]. Macrura [1213, 1290]. Macruronus [3418, 3943, 4068, 3485]. Mactra [4157]. macularius [2034]. maculatum [61]. maculatus [3793, 3437, 3451, 3399, 2271, 4104, 3826]. Made [2044, 569, 62, 211, 993, 954, 958, 1262, 941, 1087, 961, 4803, 5275, 974, 991, 370, 453]. maderensis [2616]. maenas [5375, 2003, 2002]. Magdalen [4469]. Magdalena [3868]. magellanicus [2466, 3458, 4544, 4979, 1940, 4677, 2929, 5172]. magister [3999, 3963, 3009, 3014, 2451, 3457, 2470, 2947, 3340, 2850, 2932, 2251, 3410, 3166, 4662, 2714, 5250]. magnets [5026]. Magnitude [4677]. Mahadeva [4962]. mahsena [4354]. main [3750, 3754]. Maine [4010, 2975, 4513, 1549, 1552, 1535, 1560, 1710, 479, 1646, 1807, 4602, 4772, 2206, 2540, 1174, 1957, 3023, 3895, 1811, 4990, 4213, 2309, 276, 2692, 4166, 4852, 4712, 2128, 2125, 2642, 4939, 2942, 1891, 321, 4288, 2659, 1827, 1090, 1247, 1485, 4776, 2211, 2418, 2528, 2504, 2765, 3558, 3008, 4098, 4395, 2906, 419, 4093, 4173, 5263, 5236, 5350, 2718, 3597, 3281, 3630, 1669, 3323, 2630, 3776, 3165, 969, 3194, 1989, 4573, 4699, 4733, 3625, 807, 2893, 23, 4134, 11]. mainly [90]. mainstem [4128]. maintenance [1039]. Majidae [2992, 2886, 2559, 2595, 3870]. Major [2631, 1998, 3664, 4668, 3260, 2361]. Majuro [2254]. Makaim [3555]. Makaira [4472, 4992, 4621, 4738, 4743, 3039, 2703, 2485, 4949]. Makalra [3678]. Makers [2328]. making [4700]. Mako [5076, 3294, 4333, 5302, 4921]. makos [5200]. Malacca [590]. Malaclemmys [1301]. Malaysian [2701]. Maldives [4403]. Male [2060, 3622, 3673, 3336, 4670, 387, 88, 2708, 5275, 5141, 780, 2599, 2286, 591, 615, 1873, 67]. males [88]. malformations [2458]. maliger [4694, 5034, 4931]. Mallotus [3513, 4768]. malma [5372, 5372]. Malmgren [2116, 303]. Malta [649, 736]. mammal [5094, 4152]. Mammalia [4670]. Mammals [1743, 3595, 4669, 3253, 4854, 2591, 1563]. mammary [657]. mammoth [226]. Man [777, 4803, 836]. man-made [4803]. Management [2802, 3749, 2119, 4328, 2331, 4929, 2277, 3496, 3354, 5215, 2273, 2507, 2793, 2047, 3510, 3751, 2976, 337, 4120, 2693, 5197, 5095, 2877, 4092, 3951, 4482, 4418, 1942, 4895, 3881, 4925, 929, 4401, 4744, 5045, 123]. managing [3819, 4613]. manalus [2719]. Manar [3030]. manatee [2719]. Manatees [2996]. mandatory [5039]. mandtii [3798]. Mangrove [3448]. Mangrove-Fringed [3448]. Manila [2783]. Manipulation [3472]. manmade [4441]. manned [4233, 5070, 5182]. manner [187]. manta [5300, 5318]. mantis [3155]. Mantle [1924, 1528]. manufacture [698]. manufactured [524, 553]. manufacturer [244]. many [4027, 84, 218]. Maori [4657]. maorum [4657]. Mapping [2462, 4801, 4284]. Maps [2013]. Mar [2324]. maracaiboensis [4433]. Marble [2291]. March [2178, 925, 3788, 855, 860]. Marennes [1484]. Margarita [5164, 4609]. Margin [3419]. marginal [3739, 4616]. marginatus [5115, 3779, 4542, 4688, 2680, 3752, 4222, 3636, 3175]. Mariana [3053, 3471, 5295, 3322, 3556]. Marianas [3362, 3319]. Mariculture [3354, 3527]. marina [3490, 3350]. Marine [2413, 2802, 1741, 2568, 162, 1226, 2179, 2657, 3424, 2328, 2475, 3524, 1502, 3260, 2158, 2256, 4044, 2064, 2297, 2149, 2948, 1050, 2302, 4611, 4033, 3918, 2408, 2375, 2135, 3922, 4179, 2012, 2141, 3347, 3595, 3574, 2198, 1930, 3432, 4784, 1443, 1444, 1879, 5348, 3037, 1042, 2997, 4707, 5094, 4669, 1796, 5101, 2704, 1393, 4827, 4977, 1899, 4254, 2894, 3253, 1057, 4854, 3787, 2591, 4930, 4527, 2186, 4152, 1075, 3904, 3050, 4153, 1409, 4786, 4927, 279, 2867, 4660, 3997, 2209, 4867, 1309, 256, 4334, 3426, 5374, 2079, 4264, 3643, 2288, 1659, 1102, 4031, 471, 4789, 3296, 1730]. marine [4572, 4715, 5126, 1529, 1973, 1960, 464, 357, 806, 4401, 4490, 4616, 4041, 1790, 4114, 5045, 4865, 5277, 4926]. marine-stage [4827]. marinus [2230, 1827, 1669, 842, 3165, 3268, 3297, 5251]. Maritimes [2748]. Mark [3271, 4114, 4311, 5261, 3106, 4793, 4861]. mark-recapture [4311, 5261, 3106, 4793, 4861]. mark-recaptured [3925]. Marked [2629, 2415, 4459, 1789, 4232, 1857, 4000, 4698, 4844, 3046, 5076]. marker [4243, 4804]. markers [4916, 5098, 5116, 4889, 4565, 5003, 4509, 4208, 4396, 5244, 4836, 4515]. Market [3314, 838, 943, 4652, 4647, 4676, 756, 353, 3148, 4755]. Markets [3509, 74, 157]. Marking [3138, 4794, 2041, 3884, 1726, 4221, 1596, 1686, 3928, 2486, 4276]. Markov [2793]. marks [4284, 1940, 4678]. Marlin [3555, 3601, 2129, 3678, 2485, 4089, 2684, 4472, 4992, 4693, 4621, 4791, 4738, 4743, 3039, 2703, 4949]. marmoratus [3498, 3544]. Marone [3664, 3925]. Marquesan [1956]. Marquesas [1932]. Marsh [3156, 3164, 2443, 218, 2747, 4651]. Marshall [2254, 2632, 2816]. marshes [4548]. Marshland [2352]. Martin [409]. Martyn [1610]. Maryland [5222, 210, 5230, 1446, 307, 971, 985, 1013, 2987, 5180]. Mary’s [401]. masculinization [5376]. Masking [2042]. masou [2385]. Mass [2634, 3884, 47, 2830, 3503, 3868, 2245, 4780, 3460, 2821, 2801, 2582, 465, 467, 474, 493, 513]. Mass. [390, 327, 215, 330, 420, 509, 697, 593, 594, 265, 877, 885, 893, 908, 916, 925]. Massachusetts [249, 2618, 2625, 5281, 161, 4111, 5132, 1370, 1588, 56, 1473, 5377, 50, 66, 109, 78, 81, 82, 42, 83, 127, 181, 301, 3973, 3357, 3161, 5084, 3598, 1869, 2748, 2198, 3839]. masses [3989]. Mastigonema [2231]. Matagorda [3766]. Matarese [4407]. match [4382]. match-mismatch [4382]. matching [4710]. mate [2286, 2299]. mated [3986]. Material [2548, 4866, 5210]. materials [3437]. Maternal [5088, 5034]. mating [4868, 5141, 2585]. Matrix [3386, 4602, 4314]. Matsubara [4881]. Mattaponi [4516]. Matter [2054, 579, 154]. Maturation [4109, 3809, 3552, 2040, 2532, 3672, 3314, 4920, 4060, 5007, 3862, 4046, 1896, 2968, 4628, 5275, 4258, 3304, 4853, 4355, 5254, 3839, 5252]. mature [2230, 88, 5141, 2599]. maturing [1866]. Maturity [3910, 3062, 4656, 3068, 4964, 3361, 3368, 2211, 2238, 2256, 2999, 2792, 3938, 3146, 3659, 4102, 3091, 5073, 3837, 5016, 4843, 4920, 1636, 4252, 4349, 4688, 3141, 4729, 1899, 4764, 1458, 4774, 4526, 4481, 2720, 3721, 1966, 4006, 4008, 2879, 5328, 5077, 4095, 4199, 1904, 4118, 1818, 1877, 5084, 1561, 4261, 5160, 3986, 3826, 3075, 4901, 4083, 1951, 1564, 4137]. Maumee [1086]. Maurolicus [4004]. Maxima [2195]. maximizing [2498]. Maximum [2397, 4136, 2938, 4717, 4623, 3987, 1567, 5177]. maximus [3020, 188, 5150]. May [552, 637, 761, 885, 3965, 2260]. Mayan [4463]. mayflies [1500]. Mazatlan [169]. McCloud [44, 333, 653, 834, 33, 134, 424, 654]. McLeod [2039]. Md. [835, 670, 470, 319]. Me. [227, 920, 598]. meadi [3236]. Meadows [3696, 3279, 3120, 4206]. meal [2834]. mean [4117, 2317, 3151, 2947, 4175, 4031]. means [4002, 4311, 4124, 4905, 87, 3499, 2897, 664]. measure [5002, 4953, 4947]. measured [4171, 1668, 4676]. Measurement [4979, 2535, 2853, 1651, 4527, 4599]. Measurements [4831, 4999, 2088, 2780, 2235, 2218, 2858, 2317, 1470, 3857, 4156, 4456, 5089, 4751, 2227, 4727, 4897]. Measures [3915, 4705, 4459, 4873]. Measuring [2202, 3564, 5213, 2547, 1901, 1898, 1390]. Meat [2441, 2576]. Mecca [2243]. Mechanical [3365]. Mechanism [2154, 2750, 1648, 1770]. Mechanisms [2750, 4820, 1976, 3824, 5286]. mechanistic [4021]. Medialuna [3534]. Median [3062]. medicinal [64]. mediocris [5108]. medirostris [5305]. Mediterranean [4655, 4720, 4413, 5093, 5197, 5338, 5124, 4292, 4740, 3887, 3589, 3966]. mediterraneus [4413]. Medusæ [1304, 1295, 584]. medusarum [2491]. Megafauna [3558]. megafaunal [4911]. Megalopa [3433, 2559]. Megalopae [3457, 3433, 3428]. Megalopal [3575]. Megalops [4425, 3979, 445, 2962]. megalopsis [5365]. megalopterus [5005, 4299]. Megaptera [4669, 2966, 2711, 3281, 4301, 3755, 4134, 3931, 4752]. Mekong [5363]. melanin [2376]. Melanocetidae [2800]. melanochir [4570]. Melanogrammus [4388, 3093, 3435, 5073, 3825, 4866, 3810, 5041, 4237, 4269, 1652]. melanops [4656, 4910, 2770, 2788, 2908, 4735, 3393]. melanostictus [5288, 4881, 4207, 4380]. melanostomus [4680]. melon [2471]. melon-headed [2471]. membranes [1505]. Memoranda [350, 479]. Memorandum [90, 710, 458, 484]. Memorial [3626, 271]. Memoriam [3030, 2182, 2039, 2330]. Mendota [411, 946, 946]. Menhaden [2912, 2240, 2878, 2931, 3054, 2776, 3171, 2393, 2163, 2294, 2367, 2103, 3581, 2539, 3278, 2093, 2658, 2453, 3389, 3417, 2095, 294, 4327, 3379, 3947, 4343, 5098, 228, 261, 5214, 4963, 4106, 2130, 3713, 342, 277, 58, 2729, 3044, 2489, 4749, 3969, 3795, 3888, 186, 309, 498, 343, 1356, 92, 3129, 1896, 1842, 2029, 1360, 5209, 2879, 4966, 4331, 4235, 338, 1062, 244, 3320, 3786, 3920, 1841, 1875, 1908, 5146, 5335, 1095, 585, 4681, 310, 3873, 4514, 293, 4146]. menhaden-like [1908]. menhadens [4850, 1908]. Menidia [2951, 3201, 1513, 5186, 314, 2458]. Menippe [3433, 2760, 2737, 2743]. Menippidae [2737, 2743]. mentella [4554, 4339]. Menticirrhus [5106, 3157]. Mercenaria [3101, 2340, 2814, 3106, 3226, 3375, 1907, 1803, 2760]. Mercury [4242, 2258, 2243, 2356, 2513, 2517, 2579, 2241, 2703, 2118, 2965, 2480]. merguiensis [5121]. Meristic [2209, 2226, 1902, 2370]. Merlucciidae [3485]. Merluccius [1773, 3848, 2995, 4346, 3124, 4720, 2377, 3081, 2314, 3567, 2203, 2045, 3403, 3029, 5124, 3077, 4725, 5088, 5183, 5221, 3242, 2984, 3522, 4156, 3583, 3325, 2551, 5329]. Merlucius [796]. Merrimac [604]. Mesh [2139, 2152, 4036, 4126, 4210, 4251, 4347, 342, 4690, 372, 1935, 5260, 518]. meshed [778]. meshing [4617]. Meso [3392]. Meso-Oligohaline [3392]. Mesocosm [3574]. Mesopelagic [2530, 2484, 2671, 2851, 2414, 3670, 2424, 3052, 4493, 2702, 4004]. Mesopenaeus [2563]. mesophotic [5223]. Mesoplankton [3656]. Mesoplodon [5092]. mesoscale [4703]. mesozooplankton [3562]. meta [5271, 5113, 5014]. meta-analysis [5271, 5113]. meta-analytic [5014]. Metabolic [3343, 3455, 4235, 5296, 2744, 4780, 4691]. Metabolism [4004, 2655, 2118]. Metabolite [2426]. Metabolites [2297, 2434]. Metal [3255, 2807, 2695]. metals [2284, 4938, 4427, 3375, 2707]. metamorphic [3964]. metamorphic-stage [3964]. Metamorphosis [2293, 2580, 3725, 1912, 1825, 2029, 2953]. Metanauplius [2207]. Metapenaeopsis [4414, 3516]. Metapenaeus [4937]. metapopulation [4229]. Metazoan [2269, 4208]. meteorological [5146]. meteorology [1741]. meter [2306]. metering [2853]. Method [2875, 2815, 3385, 2503, 2660, 2776, 3844, 2440, 2904, 2114, 2430, 528, 2999, 3313, 549, 1768, 3445, 173, 3667, 1761, 5228, 4623, 5112, 1949, 3840, 1026, 2502, 4804, 4953, 4846, 409, 5181, 3387, 1417, 1654, 2547, 2610, 4255, 3468, 3825, 4064, 1391, 1383, 3335, 5195, 4870, 1656, 3780, 3896, 2867, 4431, 1901, 1344, 1368, 5276, 4353, 1373, 2818, 3918, 4223, 4698, 1785, 1984, 418, 1363, 3539, 3738, 2498, 221, 1898, 4663, 1597, 4342, 568, 1459, 2732, 3393, 4119, 1375, 1390, 5229]. Methodological [3264]. Methodologies [3275, 3849]. methodology [3100, 4532]. Methods [1032, 460, 2108, 1420, 3376, 2782, 3405, 101, 756, 2833, 1145, 1384, 2560, 3605, 5004, 5057, 3122, 831, 4848, 4747, 4709, 655, 5346, 1393, 990, 1004, 4581, 1424, 4968, 4286, 4774, 4075, 4167, 4816, 4792, 302, 4987, 1408, 4776, 3279, 3215, 4429, 1948, 1132, 947, 4400, 4597, 4638, 1889, 4751, 1020, 4063, 881, 4125, 1142, 5356, 3928, 1109, 3828, 3238, 5163, 4926, 5250]. methylmercury [2794]. Mexican [2780]. Mexicana [560]. mexicanus [2780]. Mexico [4089, 1740, 1735, 1743, 1741, 2296, 4281, 3868, 3629, 4600, 4409, 2892, 3501, 3913, 4087, 4453, 3358, 5214, 5271, 4579, 3147, 3729, 4807, 3697, 4020, 4244, 4562, 4709, 5334, 3317, 699, 4165, 4541, 1939, 3935, 2615, 5376, 3441, 3609, 2986, 2611, 3315, 3473, 3768, 3882, 3860, 3880, 5196, 5304, 5091, 1233, 3399, 4671, 1776, 1793, 2934, 5279, 4014, 4253, 3105, 4619, 3702, 4625, 4889, 3948, 5129, 3349, 1611, 4366, 4196, 3115, 5223, 2478, 4048, 4975, 5371, 5367, 3000, 4287, 4370, 2546, 2571, 2574, 4289, 5325, 4909, 2364, 3173, 3854, 169, 1169, 2203, 3041, 1970]. Mexico [5327, 2871, 4899, 3294, 2323, 1885, 4096, 3760, 2456, 3645, 3835, 3698, 5032, 5240, 4704, 5328, 5370, 4778, 5368, 5191, 4965, 2868, 3916, 4512, 4317, 2542, 5269, 4353, 2290, 2685, 4448, 535, 254, 3676, 5052, 5267, 4726, 3616, 4080, 4177, 3155, 4107, 2551, 156, 5111, 4762, 4962, 5008, 5146, 5248, 4039, 623, 3508, 4996, 3389, 3152, 3657, 278, 487, 622, 2806, 5064, 1933, 4266, 1119, 1971, 3154, 4241, 3808, 5318, 4301, 3417, 805, 2488, 793, 3605, 3909, 4451, 153, 982, 1068, 4067, 4324]. Michael [3765]. Michican [1720]. Michigan [949, 1789, 213, 2573, 1683, 263, 1904, 1019, 1464, 518, 1775, 1663, 1843, 1992]. micro [1806, 1680]. micro-organisms [1806, 1680]. microanalysis [3907]. microbial [2610]. Microcarnivorous [2423]. microcell [3453]. microchemical [4822]. Microchemistry [3621, 5378, 3778]. Microcopepods [2568]. Microgadus [2623, 2784, 2883, 2798]. microlepis [3372, 4165, 4199]. micron [2359]. Micronekton [1988, 2424, 2450]. Micronesia [4865]. Micropanope [2915]. Microplankton [3656]. Micropogon [2010]. Micropogonias [3853, 3910, 4120, 4968, 3455, 5154, 4291, 3059, 4567, 3604, 4548, 2792, 5097, 4139, 4296, 3389, 2537]. micropredators [4206]. microprobe [3726, 4599]. microps [3074, 3743]. micropyle [51]. Microsatellite [4310, 4914, 4857, 4916, 5098, 5116, 4934, 4506, 5341, 4396, 4530, 4955, 4203, 5244, 4457, 4836]. microsatellites [4560, 4893, 5049]. microscope [3159, 5335]. microscopic [197]. Microsporidans [3229]. microsporidian [2748]. Microstomus [4364, 2937, 3721, 4417, 4905, 3725, 2584, 2545, 2665, 3551, 2244, 2766]. Microstructure [3180, 3678, 5255, 5155, 5154, 5209, 4010, 4609, 4292, 4405, 4011]. Microstructures [3504, 4498, 4399]. Mid [3465, 121, 2289, 4249, 3748, 3775, 2370, 5104, 3153, 3218, 4957, 2085, 3931, 3563, 3975, 4686, 5079, 5099, 4048, 2080, 3703, 3745, 5178, 3256, 5172, 2707]. Mid-Atlantic [3465, 4249, 3748, 3775, 5104, 2085, 3931, 3563, 3975, 4686, 5079, 5099, 4048, 3703, 3745, 5178, 3256, 5172, 2707]. mid-Columbia [3153, 3218]. Mid-Ocean [121]. Mid-Subarctic [2080]. mid-trophic [4957]. Mid-water [2289, 2370]. Middle [3806, 3650, 4122, 5156, 1089, 3894, 2371, 4120, 2859, 3160, 3488, 3596, 2314, 2367, 2946, 2796, 2960, 4139, 5156, 3114, 2824, 2013, 3102]. midges [1518]. Midgut [3538]. midshipman [3134]. midshipmen [267]. midsummer [1464]. Midwater [2056, 3494, 1883, 2561, 2306, 5170, 2751, 2835, 4186, 4986]. Midway [2728]. Might [4555]. Migrant [2303, 4655, 2004, 4317]. Migrating [2851, 2430, 2780, 1694, 5173]. Migration [2530, 3370, 2010, 4583, 2023, 2129, 2424, 4569, 1985, 3006, 3436, 2724, 3005, 2151, 2597, 2951, 5036, 3057, 575, 3945, 2702, 1429, 1469, 681, 5285, 5233, 861, 4169, 4737, 2593, 5175, 3025, 1954, 4162, 3357, 2725, 1509, 2515, 4822, 2567, 4805, 1218, 4238, 4861, 3297, 1943, 3221, 5277, 4639, 2526]. Migrations [1941, 2240, 2304, 3105, 2173, 2343, 3442, 1618, 1678, 1048, 4500, 1641, 3133, 2596, 1771, 1024, 3810, 2584, 5141, 1225, 1976, 1812]. Migratory [5079, 3917, 3583, 5003, 3658, 4941, 4886]. migratory-nucleus [5003]. Miguel [3127]. Mikrocytos [3453]. milberti [1850]. miles [293]. Milford [1598]. Milk [2177]. mill [1665, 1632]. Miller [551]. million [227]. Milne [3569, 2207, 2115]. Milne-Edwards [3569, 2207, 2115]. Milner [2039]. mineral [151]. minimal [4970, 4682]. minimally [5112]. minimize [4231]. Minimum [2529, 2281]. minisatellite [4054]. Minister [162]. minke [3973]. Minnesota [1828]. minnow [968]. minnows [336, 335]. minor [4190, 437, 855, 296]. Minute [531, 355, 119]. mirabilia [4639]. mirabilis [2185]. Mirounga [2926, 3994]. mirror [140]. miscellaneous [1748]. misidentification [4980]. mismatch [4382]. Miss. [1584]. missing [3996, 5340, 5261]. Mississippi [1099, 1601, 1602, 3157, 5009, 1532, 204, 1893, 1167, 1585, 123]. Missouri [1009, 915, 974]. Mitch [796]. Mitchill [1821, 1426, 3826]. mitchilli [1463, 4028, 2645, 4626, 4260, 3511]. mitigate [4937]. Mitochondrial [4889, 4731, 4048, 3383, 4291, 4080, 4916, 4002, 4576, 4392, 4442, 5341, 4785, 4766, 4509, 5270, 4530, 4335, 4125, 4116]. mitsukurii [3912]. Mixed [2276, 5212, 2693, 4884, 4245]. Mixed-Species [2276]. mixture [4777, 4027]. Mixtures [2736, 4804, 4400, 1462]. mm [5260]. Mnemiopsis [3662]. Mo. [1092]. Mobile [1099]. Mobula [5300, 3413, 5318]. Mock [2610]. modal [2656]. Mode [2358, 2375, 3262, 520, 201, 658, 5356]. Model [2913, 2520, 2119, 2074, 2354, 2304, 2277, 3401, 2580, 3261, 2670, 2431, 2578, 3589, 2439, 2494, 4558, 3749, 3292, 4623, 5057, 5189, 4328, 4643, 4604, 4781, 4863, 4832, 4877, 4819, 3919, 2871, 5215, 5157, 5166, 3072, 4547, 4982, 5261, 3942, 4027, 3206, 4474, 4758, 3305, 3734, 5242, 4263, 2679, 4861, 5347, 3847, 4269, 4594, 3897, 4379, 3958]. Modeling [4212, 3819, 5199, 5152, 5124, 4566, 3266, 4330, 3881, 4590, 5291, 3523, 5005, 4958, 4769, 4627, 4700]. Modelling [3911]. Models [2167, 3566, 3183, 3169, 2640, 3706, 2857, 2238, 2866, 2793, 1976, 2151, 4459, 3811, 4286, 4428, 5013, 3113, 4979, 4777, 4502, 4536, 5345, 5124, 5140, 5041, 5109, 4547, 5097, 4105, 5246, 5011, 4078, 1950, 4824, 5177, 4969, 4746, 4271]. moderately [1567]. Modes [2756, 4050]. Modification [3934, 4819, 4126]. modifications [5299]. Modified [4937, 2628, 2770, 2741]. modifiers [3134]. mokarran [4985]. Molecular [4638, 4337, 4850, 4804, 3939, 4966, 5128, 4887, 5163]. Mollusca [2997, 3163, 3449, 2746, 4006, 2340, 1318, 1973]. mollusk [1207, 57]. Mollusks [1753, 3260, 2018, 2020, 992, 1100, 1963]. Molt [3622, 3673, 3140, 3986]. Molting [2060, 3172]. Monacanthidae [1853]. Monadnock [376]. money [313]. Monica [4052, 3050]. monitor [5256]. Monitoring [3299, 3378, 3570, 4373, 2141, 3275, 5322, 5307, 5165, 2971, 5298, 5001, 5135, 3805, 5172]. Mono [1764]. monoclonal [4668]. monofilament [4406]. monogenean [3827]. Monognathus [2311]. Monophyly [4192]. monopterygius [3996, 4869, 4868, 4172, 4095, 4818, 4663, 4305, 4306, 4664]. Monster [688, 471]. Montagu [3464]. Montague [2631]. montagui [3194]. Montana [2228, 318, 1003, 999, 289, 2221]. Montauk [556]. Monte [4552]. Monterey [2303, 2445, 2791, 1318, 2836, 1450, 4490, 4745]. Monthly [2155, 3716]. Months [3213, 3435, 1382]. monticola [1784]. Montpellier [642]. moratorium [4808, 5052]. moray [5369]. mordax [2497, 2900, 3691, 3694, 2469, 3414, 3536, 5281, 2673, 3324, 2573, 3521, 2147, 2153, 2449, 2521, 2759, 2842, 2907, 3187, 3196, 3919, 2493, 2885, 2840, 2139, 2152, 2062, 3188, 3266, 2132, 3310, 2733, 3119, 4294, 1766, 2833, 2936, 3525, 2818, 3235, 3304, 3305, 3796, 2861, 2260, 2870, 2476, 3022, 2575, 2587, 2748, 2175, 2858, 3369, 3538, 2348, 2338, 2902, 2943, 2756, 2809, 3523]. morhua [3062, 3093, 3435, 3081, 5313, 4947, 4288, 4866, 4645, 5311, 3029, 4200, 3810, 5041, 4236, 4136, 3710, 4419, 3132, 3758, 1888]. Morid [2143]. Moridae [2143]. moringa [5369]. morio [4281, 4080]. Morning [2148]. Morone [4856, 5044, 5296, 2656, 2998, 2654, 4516, 2781, 5102, 5103, 3336, 2965, 3440, 2988, 4945, 3327, 4443, 3367, 2495, 2493, 3342, 4217, 4566, 3500, 4941, 4886, 3946, 3778, 4822, 5203, 4181, 4574, 3332]. morpholine [2617]. morpholine-imprinted [2617]. Morphological [2232, 3276, 2752, 3193, 4173, 4447, 2647, 4860, 4916, 2922, 2813, 4785, 5270, 5029, 5128, 3238]. Morphology [1916, 5342, 2608, 2389, 3177, 3262, 2752, 3060, 1969, 3690, 3565, 1815, 3480, 4799, 5187, 3717, 5363, 5243, 3300, 992, 1472, 4352, 3879, 3940, 103, 2849, 3970]. Morphometric [4249, 3878, 4252, 3521, 4778, 2226, 3395, 3663, 5288, 1894, 3712]. morphometrically [5141]. morphometrics [4388, 4346, 4671, 5027, 5178, 5247]. Morphometries [4093]. Morphometry [4039]. Morro [2860]. Mortalities [2900, 2451, 2629, 2425, 2553, 3404]. Mortality [2933, 3708, 2740, 3394, 3566, 3467, 2437, 2046, 1757, 1981, 3312, 2986, 3472, 3709, 3524, 3173, 3586, 3666, 2238, 183, 3240, 3266, 3530, 2343, 5368, 2031, 3518, 3607, 3459, 3693, 4319, 4298, 3531, 3587, 33, 3699, 3102, 3217, 3515, 5085, 3993, 3853, 4158, 5005, 3831, 4937, 4657, 4412, 5120, 5113, 4541, 4843, 5346, 4040, 5194, 3735, 4470, 5226, 4028, 3842, 4863, 4463, 4711, 1899, 4553, 5013, 4723, 5371, 3118, 4287, 4370, 2282, 4775, 3460, 3042, 4168, 4152, 4736, 4628, 2708, 3760, 4217, 2971, 5218, 4446, 4647, 4800, 5263, 3997, 1901, 4133, 4155, 4512, 4549, 5143, 5269, 2582, 254, 486]. mortality [3881, 4632, 4649, 1881, 4614, 3734, 4260, 4177, 4262, 5069, 4996, 3738, 5051, 4902, 4951, 3503, 1800, 4240, 3238, 4700, 3421, 3298, 4717, 805, 5323, 4187, 1621, 5232, 3931, 5250, 5308]. mosquito [1412, 3866]. mosquitoes [619]. moss [555, 143]. mossambica [2400, 2387]. Most [2525, 947, 4700]. mostly [347, 411]. mother [4761]. mothers [5188]. Mothership [2246]. motion [5241, 2905]. Mount [227]. mountain [942, 333, 85, 603, 263, 787]. mouth [12, 2766, 5149]. mouthed [3382, 1137, 455]. movable [540]. Movement [2980, 3048, 4479, 4990, 2019, 3709, 4288, 3641, 3348, 5263, 3297, 3669, 3061, 5118, 4948, 4357, 5102, 5226, 4111, 4169, 1979, 4898, 3938, 4851, 5211, 5130, 4651, 5131, 5085]. Movements [46, 2872, 3065, 4129, 469, 3609, 1903, 3408, 2841, 3207, 2990, 2128, 343, 2254, 3601, 3615, 2944, 3666, 2554, 2768, 3645, 3131, 168, 229, 285, 3108, 3649, 2427, 3772, 4743, 4531, 2855, 2935, 2624, 3152, 3889, 4490, 246, 4269, 3816, 4856, 5044, 5352, 4477, 1190, 135, 4503, 3064, 5099, 2810, 1622, 5200, 1192, 3342, 2565, 5298, 4154, 5023, 2278, 3784, 1426, 1434, 2205, 2846, 3220, 5316, 4806, 3905, 5203, 879, 5220, 4666, 1843]. moving [5217]. Mr. [569]. mtDNA [5270, 4035, 4433, 4318, 3808, 3605, 4515]. much [4784, 4551]. mucosa [2750]. mucous [4942]. Mud [3474, 685, 3107, 3103, 3099]. muddy [454, 260]. muelleri [4004]. Mugil [1782, 1813, 3368, 4289, 4044, 5208, 5378, 2650, 4609, 4610, 4568, 4739]. Mugilidae [4289]. Mugli [4082]. Müller [390]. Mullet [3368, 1782, 1784, 1813, 942, 473, 4289, 4044, 4082, 5378, 2650, 4609, 4610, 4568, 4739, 393, 476, 919]. Mullidae [1886]. multibeam [4801, 4529, 5347, 5071]. Multidecadal [5271]. multidens [4549]. Multidirectional [4666]. multifilament [4406]. Multiple [3183, 3201, 4924, 4925, 2745, 3271, 4181, 5044, 4501, 4794, 4846, 4944, 5283, 5216, 5003, 5157, 4955, 4615, 4717]. multiple-batch [5003]. multiplex [4857]. multipunctatus [2154]. Multiscale [4833, 5091]. Multispecies [3638, 3261, 4164, 5081, 4728, 4992, 4607, 2991]. Multistage [2857]. multivariate [5106]. Multivessel [2191]. multiyear [4863, 5261]. Mummichog [2967, 2794]. Municipal [3184, 2695]. Munida [2477]. Munidopsis [2805, 2477]. Munroe [4962]. munroi [4190]. murdochi [3170]. murphyi [3502]. muscarum [2723]. Muscle [2984, 2055, 1911, 2957, 3458, 5089]. muscles [2697]. muscular [1469]. musculature [1465]. musculus [2681]. museum [1424, 1417, 267]. musk [488]. musk-rat [488]. Muskingum [141]. muskrat [460]. muskrats [432]. Mussel [3097, 2386, 3627, 1496, 1515, 2754, 3323, 1167, 1476]. mussels [1528, 1614, 1523, 1496, 1510, 1619, 1348, 1428, 155, 1520, 1346, 1435, 1166, 1454]. Mustelus [4523, 5289, 5219, 1426, 4037, 4225]. mutton [4719, 5028, 3974, 4955]. Mutual [4283]. Mya [3032, 4243, 2625, 2997, 3163, 3449, 905, 2544, 3318, 3260, 1184, 2322, 3191, 721, 1787, 1905]. Mycteroperca [4579, 3372, 4165, 4194, 4332, 4199, 5206]. Myctophid [2391]. Myctophidae [5310, 2225, 2102, 3349, 3494, 2154, 2302]. myctophids [5173]. mydas [2267, 3962, 4471]. mykiss [3674, 3612, 3637, 5277]. Myliobatis [5164]. Myopsida [4755]. myotomes [2664]. Myoxocephalus [4494, 3110]. Myripristis [3900]. Mysid [2046, 2361]. Mysidacea [2315]. Mysids [3126]. mysis [1939]. mystes [5325]. mysticetus [2875, 4223, 3197]. Mystinus [2908, 3993, 5303, 5144, 3462, 4608, 4993]. Mytilidae [3097]. Mytilus [3097, 2386, 1515, 2746, 2754, 3627, 3323]. Myxine [4093, 4173, 3968]. Myxinidae [2899, 4220, 3631, 3968]. Myxosporeans [3403]. Myxosporidia [1488, 994].

N [2992, 2231, 3453, 2057, 3602, 2361, 3105, 2483, 1948]. N-ethylcarbazole [1948]. N-nitrosamines [2483]. N.C [1498]. N.C. [1493, 1615, 1639, 1502, 612, 400]. Naknek [1945, 1943]. names [1270, 1445, 1712, 1715, 248, 264]. Nanaimo [2755]. Nannoplankton [2445, 2090]. nannus [4726]. Nanomia [2630]. Nantucket [4645, 265]. Naobranchia [4407]. Naobranchiidae [4407]. Naples [225]. narinari [5054]. Narragansett [3044, 4972, 4922]. narrow [4714]. narrow-barred [4714]. narrownose [5219]. Nassau [4357]. nasus [4312, 4729, 4528, 4486, 5274]. natal [4273, 4941]. Nation [898]. National [1139, 267, 999, 953, 1819, 956, 3918, 1362, 4598, 4632, 517, 372]. Native [1270]. Natural [3993, 2516, 625, 1510, 4843, 2082, 1654, 1608, 3709, 2629, 1430, 2307, 2211, 3617, 3413, 1595, 2324, 1249, 1342, 3490, 1850, 1800, 2443, 4441, 1511, 4158, 3831, 1027, 1517, 988, 242, 4787, 539, 1073, 4553, 4723, 5218, 4803, 1187, 1338, 5156, 4469, 1006, 1135, 1107, 1454, 3421, 1374, 4356]. natural-history [1006]. Naturalistic [3474]. Naturalization [1407]. naturally [4652]. Nature [3689, 2061, 1721, 1587, 4254, 1495, 4354]. nauplii [3851, 3441]. Nauplius [2207]. Naval [267]. navigable [759]. ND [4392]. ND-3 [4392]. ND-3/ND-4 [4392]. ND-4 [4392]. ND3 [4785]. ND3/ND4 [4785]. ND4 [4785]. Neah [3641, 250]. Near [3269, 2202, 2225, 2484, 3184, 2010, 2944, 3617, 1955, 1202, 4034, 581, 2301, 412, 4908, 835, 4960, 2490, 2666, 3559, 105, 2765, 3407, 2031, 662, 2071, 2582, 3226, 4488, 3306, 5156, 12, 470, 4007, 2572, 777, 3657, 645, 609, 2713, 5224, 1499, 4752, 2695]. Near-Bottom [2202]. near-reef [4908]. near-surface [2071]. nearly [5300]. Nearshore [3078, 2901, 3573, 4257, 3189, 4051, 3993, 3983, 4835, 5367, 3985, 5175, 4171, 3022, 4065, 2382]. nearsurface [4493]. Nebraska [1576, 1082]. nebulifer [4030, 4268]. nebulosus [3438, 2619, 5074, 4757, 4359, 4598, 4632, 4661, 4836]. necatrix [1406]. necessarily [3734]. Necker [4542]. Necropsy [4539]. necrosis [2748, 3165]. Nectobrachia [4407]. Need [372, 5363, 5245, 397]. needed [349]. Needlefishes [2027]. needs [484, 770]. negative [3770]. Nehalem [4960]. Nehu [3346, 3517, 2435]. Nekton [3164, 2979, 5142, 3941, 4505, 4867, 2835]. nektonic [2651]. nelsoni [2550]. Nemadactylus [3907]. Nematistius [5201]. Nematoda [2231]. nematode [4760]. nematodes [1736]. Nematoscelis [2333, 2410, 2962]. Nemertea [3602]. Nemerteans [1292, 1736, 1211]. nemertes [3340]. nemoptera [1974]. Nene [367]. Neomysis [2315]. neon [4958, 4882]. Neonatal [2177, 4706]. neonate [4244]. Neoscopelidae [2433]. Neosho [1092]. Neothunnus [1689, 1693, 1676, 1670]. Nephropidae [3924]. Nereis [3023]. Nereocystis [3309]. Neritic [3315, 3909]. nerka [1662, 4310, 2474, 2660, 4827, 4273, 3370, 2134, 1866, 2499, 4954, 1830, 2197, 2262, 4715, 1943, 2716]. nerve [4862]. nervous [1925, 1157]. nests [1685]. Net [2172, 2306, 3492, 2445, 2414, 2112, 2359, 4265, 257, 341, 529, 5066, 583, 2621, 2636, 321, 215, 420, 509, 697, 1797, 3423, 1935, 4456, 4029, 2686, 1014, 1663, 5131]. net-caught [4456]. Netarts [3065, 3528]. Netplankton [2090]. Nets [2139, 3143, 2548, 2534, 2962, 2933, 1052, 778, 3803, 2838, 15, 205, 529, 449, 667, 4858, 2659, 2751, 2152, 50, 66, 109, 322, 330, 2465, 758, 4890, 5260, 518, 1014, 4209, 3539, 5203, 334]. Netuma [3296]. neural [4537]. Neuse [1097]. neuston [2621]. Neustonic [3956, 3428]. Nevada [16]. Newfoundland [3221, 3381, 4123, 3388, 839, 2949, 3141, 130, 148, 3622, 3870, 934]. Newly [2783, 2968, 5137]. Newport [3007, 2360, 3059]. news [372, 78]. niche [4121, 5358]. Nichols [3502]. Nickel [2356]. Nicolas [3668]. Nicoya [3503].

Night [2509, 3079, 4652, 4995, 2728, 421, 4633]. night-seining [421]. nigricans [3678, 4472, 4992, 3555, 4621, 4738, 4743, 3039, 2703, 2485, 3534, 4949]. nilotica [2916]. nine [1571, 3968]. Ninespine [2265]. nineteen [93]. Niño [4169, 5078, 3269, 3299, 3331, 4692, 3288, 4735, 2667, 3063, 4767, 4302, 4270, 4051]. Niño [?]SouthernTorres-Orozco:2006:VYT. niphonius [4716]. Nitrogen [2441, 2931, 3054, 2009, 2084, 2132, 2979, 2425, 4049]. Nitrogenous [2178]. nitrosamines [2483]. niveatus [5294, 4323]. No [4871, 4666]. no-take [4666]. nobilis [4883, 5118, 3759, 4141, 3533]. Nocturnal [2901, 4436, 4789]. nodifrons [2737, 2743]. noise [4824]. Nomeidae [4096]. Non [3621, 4273, 4941]. Non-anadromous [3621]. non-natal [4273, 4941]. nonanadromous [3412]. Noncoevolved [3246]. Nonequilibrium [2670]. nonextractive [5001]. Nonindigenous [2918]. nonlethal [5050, 5181, 2591]. nonparametric [3818]. Nonrandomness [3549]. nonsalmonid [3117]. nonspawning [4068]. Nordmore [4090]. Norfolk [155, 3243, 5156]. Normal [2283, 1619, 1617, 3288, 1801]. North [2892, 3224, 2143, 1438, 2966, 3038, 3193, 3985, 2192, 3681, 2768, 2103, 1313, 110, 4589, 2482, 3316, 3226, 2800, 4344, 4488, 3395, 1474, 3398, 2244, 2160, 2710, 615, 2747, 2831, 2731, 2095, 3497, 911, 1853, 4250, 1886, 1876, 5367, 4909, 1851, 1868, 2792, 3916, 1820, 1983, 5052, 1012, 4241, 3855, 53, 4850, 3912, 2684, 2089, 2317, 4961, 1437, 3856, 4601, 3454, 5310, 1056, 3831, 4477, 4191, 4963, 5375, 5094, 4903, 4641, 2722, 1577, 5110, 5365, 2615, 456, 3735, 3809, 5237, 1008, 1619, 3959, 5155, 3967, 3781, 5278, 4774, 2201, 4606, 3109, 1991, 4731, 4042, 2808, 2752, 3144, 3026]. North [3737, 1574, 4693, 4197, 5337, 5027, 4528, 941, 2974, 5046, 4791, 3401, 5331, 3957, 5294, 2596, 2477, 2436, 2496, 2593, 2909, 4868, 3527, 3059, 3077, 2771, 2947, 4738, 5159, 2510, 4173, 4154, 4569, 2446, 2385, 2501, 2111, 4984, 3936, 4295, 4486, 5274, 5272, 5302, 4895, 3594, 4352, 4648, 5128, 4886, 4697, 5030, 2680, 1328, 4825, 4555, 4957, 3090, 4029, 3788, 4872, 3074, 3139, 4572, 1788, 1805, 1882, 5043, 351, 3479, 2558, 3633, 4879, 1661, 1678, 3790, 2485, 4591, 131, 4013, 997, 3889, 2444, 2250, 702, 3845, 2797, 5357, 5145, 253, 3223, 3873, 4514, 2403, 5092, 2382, 2361, 3631]. North [3545, 4228, 5259, 4539, 180, 711, 4323, 5169]. north-central [5367, 3916, 4241]. northcentral [4541, 5008, 3698]. Northeast [2395, 5247, 3595, 4540, 4989, 3036, 5058, 4672, 2541, 4066, 4169, 1982, 2888, 5126, 4634, 3792, 4266, 4593, 4682, 4880, 5224, 3540, 2513, 2517, 2579, 3263, 2883]. northeastern [5322, 5301, 3158, 2693, 4244, 4562, 3434, 4165, 4687, 4843, 5346, 5199, 4710, 4900, 5144, 4253, 4978, 4481, 4473, 1644, 2282, 1887, 4226, 1087, 4170, 4588, 3980, 4633, 4436, 2681, 3061, 3575, 2478, 2320, 2364, 3173, 2598, 2421, 3043, 2456, 3670, 2057, 3450]. Northern [3228, 3708, 3273, 3691, 3694, 2469, 3536, 3489, 3492, 3864, 2673, 3324, 2709, 3521, 2110, 2449, 2521, 2759, 2907, 3187, 3196, 2493, 2885, 2271, 2840, 2062, 3188, 3266, 3310, 3359, 2733, 3119, 2833, 3525, 3235, 3316, 3284, 2053, 2115, 3796, 2870, 2476, 2575, 2175, 2644, 3625, 2858, 3538, 2439, 2809, 3523, 2284, 2926, 4409, 3501, 3913, 3874, 4364, 5214, 5271, 4807, 3414, 4669, 4390, 4522, 4928, 5372, 5280, 5266, 4073, 3473, 3768, 3882, 3860, 3880, 5196, 5304, 5091, 3036, 3360, 5355, 5265, 4671, 2553, 5279, 4014, 2894, 5129, 3823, 3890, 3660, 4852, 5225, 3115, 4792, 4795, 3144]. northern [4975, 5371, 3462, 4415, 4416, 4274, 2842, 3919, 5353, 4608, 4820, 4153, 5018, 3363, 3760, 3994, 4713, 3108, 5042, 4465, 5077, 5190, 5368, 5191, 4867, 4965, 4294, 4932, 2936, 4246, 5269, 2818, 2685, 3304, 3295, 4448, 3305, 5267, 3984, 4201, 2861, 2260, 5024, 4762, 5146, 4039, 1890, 5273, 2587, 4996, 4298, 3867, 1006, 1341, 2889, 4923, 2882, 5064, 3369, 2348, 5318, 4702, 218, 750, 2902, 2943, 5361, 3909, 4451, 1068, 3551, 3353, 5329, 4067, 4407, 3629, 4624, 751, 3441, 3315, 3424, 2487, 3105, 2356, 3190, 2475, 3193, 3409, 3205, 2868, 3053, 2661, 3585, 3389, 3657, 3417]. Northumberland [244]. Northville [150, 785, 892]. northward [3765, 4514]. Northwest [3091, 4956, 3348, 4612, 1133, 4346, 2903, 4024, 4211, 4523, 4006, 2994, 2690, 2272, 2711, 5366, 5358, 3290, 4921, 3062, 4958, 4994, 5249, 3300, 3029, 4554, 3651, 3295, 3712, 4178, 3094, 4810, 3682, 2806, 3512, 2979]. Northwestern [4542, 1526, 3450, 3402, 4020, 1840, 5093, 1939, 1079, 1003, 3000, 5200, 5124, 4974, 4677, 4549, 3155, 1884, 4616, 1971, 3358, 3697, 4470, 2364, 2343, 3752, 4222, 2073, 3255, 3515]. Norton [4404]. Norway [581, 714, 610, 726, 822, 1393, 764, 614, 665, 691, 720, 902, 132, 283, 380]. Norwegian [15, 757, 3132]. notatus [3134]. Note [2161, 4087, 355, 485, 927, 3458, 387, 1513, 928, 1494, 1394, 683, 212, 378, 591, 689, 510, 62, 4652, 2957, 2541, 290, 532, 4295, 5047, 3148]. Notemigonus [2407]. Notes [1133, 274, 660, 249, 347, 4324, 490, 48, 717, 945, 3037, 618, 44, 444, 522, 523, 3434, 375, 465, 838, 1223, 136, 161, 135, 655, 852, 799, 918, 964, 339, 364, 1058, 1131, 242, 920, 1481, 2143, 186, 309, 498, 2461, 950, 1078, 60, 137, 1034, 1117, 1179, 1077, 651, 744, 1098, 638, 1088, 1169, 1245, 1330, 1325, 1090, 1247, 1439, 980, 1000, 1472, 1479, 2908, 1409, 42, 83, 36, 106, 5, 19, 20, 127, 181, 215, 301, 420, 509, 467, 474, 493, 513, 697, 593, 594, 2883, 475, 1082, 666, 626, 597, 895, 39, 3236, 747, 535, 486]. Notes [2750, 2816, 158, 438, 857, 516, 1162, 69, 2, 2743, 1218, 237, 547, 632, 635, 647, 749, 804, 819, 862, 891, 677, 971, 957, 955, 985, 1076, 1080, 1071, 1095, 1173, 1239, 245, 882, 824, 427, 1454, 1046, 910, 1287, 2961, 1671, 386, 769, 783, 853, 934, 3419, 548, 2723, 570, 856, 1182, 1511, 1955, 1646, 1528, 5342, 223, 889, 943, 1106, 1233, 2738, 1778, 954, 70, 3178, 1571, 1639, 2762, 1425, 2813, 81, 82, 4162, 1063, 2680, 476, 4354, 2888, 236, 2737, 1124, 1396, 356, 377, 379, 496, 1006, 1014, 1107, 877, 885, 893, 908]. notes [916, 925, 599]. notheastern [4084]. nothus [4916, 2986, 4917]. notialis [4778]. Notice [975, 1165, 615, 71, 119, 217, 754]. Notolabrus [4746]. Nototheniops [3614]. Nototodarus [4197]. Nova [3068, 905, 2138, 2355, 1669, 3290, 3740]. novaeangliae [4669, 2966, 3281, 4301, 3755, 4134, 3931, 4752]. novaengliae [2711]. novaezelandiae [3418, 3943, 4068, 3485]. novemaculeata [4701]. November [1128, 556, 138, 662, 772, 853]. Nuclear [4509, 4145, 4576, 3412, 4335, 5089, 107]. Nuclei [3180]. nucleic [5174, 3934]. nucleotide [5101]. nucleus [5003]. nudifrons [3614]. Number [389, 2261, 4428, 2651, 680, 2947, 256, 3304]. Numbers [2346, 4371, 4748, 2188, 1696, 3353]. Numerical [3499]. Nuremberg [588]. nurse [5228, 4183]. Nurseries [2747, 2460]. Nursery [2352, 2993, 4430, 4076, 2970, 4266, 5355, 4254, 4795, 4892, 3244, 4044, 5378, 4913, 5236, 5135, 2288, 3920, 4811, 4572, 5268, 4321, 3907, 4651, 4703]. Nushagak [2091]. Nutrient [3012, 2037, 2146, 4927]. Nutrient-Poor [2037]. Nutrients [2178, 5251]. Nutrition [1486, 1495]. Nutritional [2036, 4234, 3960, 2575, 2101, 3538, 3258, 4094, 2587, 2849]. Nutritive [2038, 2212, 2887, 426]. nylon [4406].

O [3872]. O. [3009, 3014, 175, 3408, 200, 5233, 1680, 1830, 3637, 3398, 4715]. O/Ra [3872]. O/Ra-226 [3872]. Oahu [2033, 2609]. Oak [2103, 2095]. obesus [4558, 4821, 4245, 5373, 4531, 3103, 4435, 5078]. Obispo [1088]. objections [453]. Objective [3183, 1962, 2999, 4121, 4673]. objects [4383]. oblique [3715]. obliquidens [2882]. obscurus [4160, 3899, 3936, 4377, 3833, 2789]. Observation [2444, 1864, 5129, 1422, 5300, 5361]. observational [4824]. Observations [2191, 1037, 983, 776, 2895, 1456, 1675, 2487, 1778, 2478, 2791, 2555, 3171, 2659, 1184, 2162, 2168, 2464, 2339, 2126, 3695, 1055, 3537, 2821, 2830, 3750, 3877, 780, 3547, 2798, 2085, 2955, 1794, 185, 3633, 979, 1260, 3328, 2247, 3686, 1674, 3021, 3574, 986, 1506, 529, 5222, 2124, 4990, 1532, 2876, 4428, 88, 1637, 4290, 5275, 991, 5276, 3822, 4027, 1232, 170, 1801, 2630, 14, 65, 89, 4300, 3103, 3507, 917]. Observed [2914, 1839, 2317, 936, 4503, 1106, 194, 638, 5182, 974, 3676, 1043]. Observer [4976, 3217, 3744, 3842, 4863, 4980]. Observer-reported [4976]. observing [1391]. Obstructions [289, 1135]. obtain [4652, 2732]. Obtained [2112, 410, 433, 4751, 71, 977]. obtaining [658]. occidentalis [4407]. Occupancy [3207]. occupied [832]. Occupying [2868]. Occur [2355, 1219]. Occurrence [294, 2089, 2804, 2230, 4244, 3437, 2325, 2314, 3318, 2483, 2379, 473, 3457, 1970, 2477, 2788, 2908, 3835, 327, 3973, 3357, 2706, 5248, 2855, 3585, 3790, 373, 1933, 121, 2961, 3256, 413, 4386, 4493, 1782, 4441, 5207, 305, 1955, 5342, 3793, 542, 927, 3880, 5258, 5096, 5105, 5283, 928, 3791, 4907, 4595, 3059, 975, 3203, 2327, 1484, 2278, 1598, 4097, 4384, 3165, 4436, 5224, 2403, 4535]. Occurrences [2192]. Occurring [2805, 2497, 4350, 4956, 968, 2792, 3732]. Ocean [2167, 1899, 2304, 2131, 2629, 2238, 2554, 2960, 3442, 5030, 3090, 3123, 3146, 2753, 3600, 3255, 121, 4557, 1784, 3917, 1188, 4255, 4960, 4993, 2708, 2565, 5339, 1561, 4715, 1047, 2707, 5212, 3198, 3912, 4829, 2684, 2089, 3477, 3062, 2862, 2892, 1988, 2373, 3454, 3540, 4936, 5310, 5342, 5352, 3158, 3711, 4191, 4958, 4830, 4641, 2530, 2225, 2301, 4523, 4687, 4843, 5365, 5346, 5379, 3717, 2028, 5237, 4994, 5058, 3771, 3821, 4147, 859, 4900, 5155, 5152, 3967, 4672, 4312, 5249, 2702, 4978, 5278, 4524, 2738, 4854, 4196, 3109, 4731, 1858, 4042, 2566, 4079, 3038, 4821, 3144]. Ocean [3193, 4367, 3300, 3376, 5200, 2282, 5337, 4528, 2320, 2364, 4907, 2203, 5331, 5353, 5373, 2477, 5138, 4868, 2214, 4738, 5240, 4173, 2883, 2129, 2343, 3670, 1969, 4569, 5289, 2209, 4589, 4144, 4677, 5023, 3936, 4486, 5274, 5272, 5302, 4487, 4352, 4648, 5201, 1818, 1982, 1983, 4027, 4259, 2800, 4354, 3583, 3653, 2308, 3865, 2057, 3395, 4061, 1012, 5309, 4468, 4872, 5210, 3390, 3663, 4007, 4420, 4851, 3841, 2160, 2431, 2485, 4591, 166, 4011, 2806, 1808, 4435, 4949, 2250, 702, 5078, 2037, 2146, 2336, 5173, 4302, 2195, 5357, 5254, 3450, 3515, 2488, 3061, 3620, 2151, 3556, 3545]. Ocean [4228, 4921, 5259, 5169, 2083, 2065, 5147, 5083]. ocean-type [3917]. Oceania [1320]. Oceanic [4274, 3586, 2724, 3583, 4822, 3684, 2137, 3856, 4601, 4036, 4126, 1675, 1221, 3387, 4150, 4596, 3782, 4954, 4906, 4697, 2835, 2552, 1976, 1691, 1884, 5220]. oceanic-stage [4601]. Oceanographic [1559, 2554, 1982, 3428, 5173, 4477, 2565, 4956, 4374, 4885, 1865, 5259]. Oceanography [1839, 2229]. Oceanside [3078]. ocellata [2519, 2930, 4573]. ocellate [5270]. ocellatus [4948, 3501, 4103, 3731, 3898, 4462, 3785, 5367, 4914, 3006, 4474, 5052, 4262, 4318, 4679, 3928, 4049, 4276]. ochotensis [2922]. October [2260, 654, 755, 783]. octodecemspinosus [4494, 3110]. Octomitus [1570]. octonemus [3609]. octopod [3003]. octopus [4961, 4657, 5110, 4900, 4454, 597, 4657, 4454]. oculatus [4721]. Oculina [4930]. Ocyurus [3940]. Odobenus [4304, 3084]. Odontaspis [3056]. Odontocete [2954]. Odontopyxis [3817]. odorant [4862]. Oegopsida [3017]. Off [2528, 2545, 2558, 3639, 3595, 5118, 4089, 4327, 4403, 966, 1832, 2977, 3078, 4160, 4329, 2805, 1784, 4833, 4845, 4018, 2317, 3122, 2094, 4847, 1822, 2016, 2524, 3309, 2597, 3013, 3213, 3150, 2804, 3358, 2423, 2527, 3634, 4364, 2693, 5094, 4669, 4709, 4903, 5333, 4617, 3834, 4602, 4642, 4935, 2859, 4884, 336, 5153, 556, 927, 5266, 3420, 4946, 3470, 5257, 4060, 5222, 3315, 4141, 5107, 2187, 4358, 2509, 2823, 2825, 1221, 2621, 4787, 4159, 3981, 1311, 3408, 3945, 4689, 2841, 3983, 4284, 4998, 2937, 3948, 4673, 5176, 5225, 4313, 3200, 3331, 4930, 4481, 4795, 4473, 5105]. off [4368, 4975, 3496, 3462, 2742, 2543, 1170, 2762, 4274, 1972, 3985, 4197, 4585, 5360, 3457, 5262, 2598, 4987, 1251, 1439, 2715, 4226, 4776, 5294, 4653, 2596, 4820, 4993, 3043, 4685, 2788, 2840, 2908, 3125, 4956, 961, 3077, 4637, 4444, 4974, 3409, 4546, 4218, 2470, 2638, 2774, 2919, 3135, 3174, 2297, 2434, 4161, 4331, 4332, 4695, 5240, 1668, 5037, 2799, 4823, 3763, 1975, 4982, 4917, 3537, 2655, 3980, 2159, 3189, 5276, 5097, 5227, 4156, 4549, 4374, 3415, 2071, 2427, 2725, 5314, 4886, 2274, 4885, 4448, 2229, 2450, 2663, 2665, 3637, 3705, 3316, 1669, 4758, 3796, 2955, 4743, 5011]. off [3395, 4201, 4726, 4029, 3757, 2073, 2260, 2398, 2552, 2706, 2820, 2832, 2870, 3938, 5024, 4336, 4177, 2059, 3146, 2324, 3074, 3076, 1816, 5043, 4469, 3964, 4633, 1618, 4879, 3553, 3034, 3428, 4894, 4065, 4263, 2560, 4299, 2359, 2035, 373, 3531, 5211, 2313, 1884, 2141, 3889, 5019, 335, 3534, 4340, 3792, 3740, 2882, 5078, 4809, 4592, 4593, 4765, 5318, 4301, 3341, 4880, 4492, 4702, 3223, 4207, 2989, 3519, 5264, 3247, 3894, 5076, 2403, 3308, 2518, 2198, 5180, 4451, 4386, 4551, 5085, 3688, 649, 5090, 4323, 4272, 3551, 4051, 5308, 5369, 5123]. offal [569]. offered [125]. Official [102]. Offshore [3394, 2951, 2478, 1965, 2923, 2450, 2053, 2473, 2551, 4845, 3501, 2373, 5115, 1613, 4111, 3967, 4908, 1958, 5327, 4169, 3244, 4786, 2590, 5052, 3920]. Offsore [3222]. Ogeechee [503]. oglinum [2571, 2205, 2334]. Ohio [336, 477, 165, 74, 975, 335]. Oil [3572, 2502, 2441, 2328, 2058, 3617, 3639, 3657, 323, 261, 4787, 160, 1642, 209, 4130, 4316, 4803, 244, 306, 317, 283, 380, 1462]. Oils [2344, 2042]. Okamejei [5270]. Okanogan [1954]. Okhotsk [4771]. old [56, 3104, 220, 521]. Older [3613]. Olfactory [2242]. Olga [3254]. Oligochaeta [1319]. Oligocottus [3219]. Oligohaline [3392]. oligomerus [4962]. Oligoplites [2179]. olivacea [4555]. olivaceus [5003]. Olive [2774, 2919, 3174, 4555]. Olympia [1637]. Oman [5190]. omanensis [2066]. Ommastrephes [4034, 4958, 5155]. ommastrephid [5072]. Ommastrephidae [5133, 4247, 4309, 4034, 4830, 4197]. Onaga [3552]. once [3986]. Oncorhynchus [2417, 2516, 2839, 1662, 3011, 3251, 3480, 4054, 4310, 4560, 4559, 4893, 4934, 5049, 62, 48, 2219, 5348, 2520, 4989, 3009, 3014, 4919, 2474, 5101, 2660, 2412, 2506, 1643, 1636, 1641, 5167, 2242, 2819, 3884, 3133, 4827, 241, 4977, 3408, 3945, 4069, 4689, 4442, 1458, 3917, 5290, 4273, 3370, 2810, 2755, 2677, 2629, 2252, 2952, 2134, 4274, 2826, 5233, 1962, 3588, 2256, 4373, 4566, 4484, 1866, 2004, 2499, 4954, 2442, 3088, 2385, 4334, 3674, 3181, 4396, 4398, 5143, 3784, 4033, 3612, 2557, 3442, 3528, 3637, 2197, 3550, 4035, 4950, 3398, 2865, 2852, 2262, 2425, 4715, 2753, 5366, 1681, 1666]. Oncorhynchus [4203, 1943, 3271, 4592, 4128, 3341, 3512, 4043, 5277, 3046, 2716, 2918, 3990, 2786, 2006, 5169]. One [2757, 269, 227, 388, 241, 3262, 2758, 4913, 299, 265, 527]. onitis [4441, 5181, 4972, 2278, 2427, 2586, 2727, 4335, 3745, 4575]. Onofre [3078, 3122, 3125, 3519]. Onset [4383]. Onshore [4908]. Onslow [3059, 4344]. Ontario [3040, 989, 11, 67]. ontogenesis [2764]. Ontogenetic [5338, 4988, 5006, 3782, 4469, 3824, 4267, 4100, 5312, 4819, 5289]. ontogenic [4474]. Ontogeny [2620, 4788, 5272, 2515]. Onychoteuthis [3856]. Oocyte [3862, 3368, 3780, 4818]. oocytes [3727, 4866, 5003]. opah [5365]. opakapaka [5316]. opalescens [4652, 2452, 2649, 4647, 4676, 3148, 3314, 4755]. opaque [4348]. Open [2970, 1784, 4903, 5062]. Opening [456, 5313, 2835]. opening-closing [2835]. openings [4504]. operated [5070, 5182]. Operating [3154]. Operational [2481]. operations [150, 596, 835, 670, 1158, 628, 4981, 470, 607, 1379, 134, 654, 879, 592]. Opha [928]. Ophiodon [3641, 3348, 3763, 4431, 5045, 3816]. opilio [2492, 2249, 4174, 5082, 5141, 3710, 5069, 3986, 4265, 4902, 3221, 3622, 3673, 3870]. opinion [92, 254, 732, 303]. Opinions [720, 544]. Opisthobranchiate [1318]. Opisthonema [2571, 2205, 2334, 2938]. Opossum [2315]. Opportunistic [4494]. Opsanus [2602, 1377]. Opsopæodus [950]. Optical [2458]. Optimal [4117, 2119, 3807, 2405, 3736, 3523, 4958, 4627, 4264]. Optimization [4919, 2793]. optimizing [2668]. Optimum [2363, 3356, 2273, 2233, 1998]. orange [3876, 4062, 3791, 5190, 4125, 3958, 4119]. orange-spotted [5190]. orbital [4639]. orca [2069, 2072, 4293, 4823, 5042, 3507, 3955]. Orcinus [2069, 2072, 4293, 4823, 5042, 3507, 3955]. Order [2311, 451]. ore [1856]. Oreek [1985]. Oregon [3638, 2470, 2603, 690, 2805, 4845, 3444, 3447, 3269, 3013, 3213, 3037, 3634, 4364, 3065, 2258, 399, 507, 446, 3408, 3578, 3945, 4069, 4689, 2937, 3719, 3331, 4960, 4995, 2952, 2993, 2543, 2762, 5262, 2761, 603, 2508, 3244, 2788, 2840, 2908, 2956, 5292, 1668, 2799, 4529, 4629, 2274, 4885, 2229, 2288, 2450, 2545, 2663, 2665, 3528, 3547, 3637, 3772, 2399, 3015, 2260, 2398, 2552, 2706, 2820, 2832, 2870, 5024, 4811, 3076, 2753, 3034, 3428, 2560, 2035, 791, 1341, 1339, 5211, 3747, 939, 4592, 582, 3894, 4535, 5250, 3551]. oregonenisis [1829]. oregonensis [3575, 3457]. Oregoniinae [2249, 2559]. Oren [531]. Oreo [3451, 4055]. Oreosomatidae [4055]. orfe [7]. organ [1241, 212]. Organic [1581, 579, 3386, 3255, 1723, 1222]. Organism [2030, 2233]. Organisms [2312, 2375, 2218, 3132, 355, 1806, 2651, 1883, 1680, 1960, 1091]. Organization [1369, 3164, 1478, 1342, 296, 231, 676, 1379]. Organochlorine [2806]. Organs [1927, 2261, 1917, 88, 200, 1472, 1299, 1426, 591]. orientalis [4710, 4583, 4584, 5364]. orientallis [3850]. Orientation [2099, 1694]. Oriented [2254]. Origin [4035, 4515, 2862, 4624, 1042, 58, 4033, 1722]. original [4178, 4923, 4727, 4379]. Origins [1895]. Orland [227]. Orr [4407]. Orthonopias [3746]. Orthopristis [3095]. Orwell [2826]. Oscar [2182, 2171, 2166]. oscillating [5238]. Oscillation [5133, 4169, 2667, 3063, 5078, 4767]. Oslofjord [4455]. Osmerus [2900, 5281, 2573, 2748]. Osmolality [2843]. osphradium [4862]. ostenfeldi [5324]. Osteological [2308]. Osteology [2925, 1306]. Ostrea [1635, 1648, 1637, 120, 212]. Oswego [643, 114]. ot [3706]. Otaria [4329]. Other [2484, 3260, 3229, 3029, 2178, 2793, 2663, 2750, 2061, 2858, 3082, 1048, 3374, 4669, 3837, 5106, 889, 456, 1470, 409, 801, 1235, 3937, 5340, 491, 155, 1571, 3851, 1425, 3647, 752, 1480, 1387, 1373, 1426, 1595, 170, 65, 1618, 5244, 357, 547, 806, 1014, 1094, 328, 4406, 4186, 3845, 4268, 877, 885, 893, 908, 916, 925, 5040]. Otoconia [3541]. Otolith [3435, 3451, 3702, 5027, 3257, 3621, 2975, 4292, 3514, 3674, 3181, 3180, 3678, 3071, 3947, 4498, 3996, 5255, 4388, 4799, 3195, 4346, 3412, 5344, 3360, 3943, 4734, 3783, 2894, 4816, 5288, 3789, 3367, 4859, 4882, 5373, 5154, 5209, 3927, 4010, 5378, 3215, 3896, 4609, 4972, 5117, 4236, 3933, 4399, 4532, 3778, 3925, 4822, 4616, 4011, 3726, 5247]. otolith-based [4816, 3925]. Otoliths [3644, 2455, 2143, 3386, 2967, 2402, 3680, 2916, 3493, 2893, 4343, 4190, 4459, 3134, 4348, 4311, 4672, 3739, 5208, 3130, 4000, 3096, 4101, 4003, 4354, 4359, 3301, 3159, 2511, 4813, 4698, 4844, 3138, 3907, 4663, 3250, 4270, 3393, 4599]. Otter [2631, 2380, 3031, 4686, 4708, 4939, 3272, 4491]. Otter-dominated [2380]. otters [5375, 2791]. Our [1047, 120, 992, 256, 1500]. outer [4619, 5125, 5156, 5025]. outer-coastal [5025]. Outfall [3184, 3136, 2695]. Outlet [1893]. Outline [1410]. outplanting [3885]. output [5327]. ova [481, 2968, 419, 124, 14, 65, 89]. Ovalipes [4679, 4436]. Ovarian [5376, 2917, 3813, 3368, 4008, 3910, 3913, 4656, 4866]. Ovaries [3058, 5073, 4760, 1505, 4142]. Ovary [3187]. overexploited [5190]. overfished [4508, 4786, 5256]. overfishing [5077, 4983]. Overlap [3029, 3158, 5355, 4069, 4121, 3757, 5126, 5358, 5205, 5139, 5090, 3287]. oversampled [3715]. Overseeding [2677]. overview [3725]. Overwinter [5360]. overwintering [4343, 4443, 4517]. oviducts [1505]. ovigerous [2775]. Ovulation [2811, 2426, 2381, 2968]. oxide [2394]. Oxycline [2592]. Oxygen [2669, 2843, 2503, 1617, 4049, 2588, 3948, 2898, 5208, 1449, 4735, 2425]. oxygen-nitrogen [2425]. oxygeneios [4230]. Oxyrhyncha [2249, 2886]. oxyrhynchus [2413, 2983]. oxyrinchus [5258, 5351, 5306, 4994, 5258, 4729, 5306, 4333, 5302, 4180, 4376, 4320, 5076, 4921]. oxytetracycline [4183, 5272, 5076]. Oyashio [4636]. Oyster [2024, 470, 2371, 1918, 1917, 984, 5142, 572, 175, 502, 1949, 964, 1874, 1806, 636, 988, 990, 1004, 1240, 534, 1628, 2602, 1592, 719, 468, 200, 1635, 1648, 1637, 120, 1136, 4315, 4646, 1656, 1147, 845, 1064, 461, 3881, 4924, 4925, 1598, 1099, 1020, 1115, 351, 69, 201, 197, 119, 463, 512, 1126, 957, 1367, 1013, 1153, 484, 548, 1121]. oyster-bars [1126]. oyster-beds [1115]. oyster-culture [984, 990, 1004, 1136]. oyster-grounds [988, 1153]. Oysters [2183, 3453, 1136, 3452, 162, 797, 753, 1723, 1996, 4604, 1609, 1617, 1642, 1665, 965, 680, 1632, 1626, 1622, 1656, 1680, 1447, 1486, 1495, 1484, 298, 662, 1392, 1598, 1352, 236, 546, 358, 733, 573, 1869, 483, 616]. Ozark [787].

p’ [2266, 2266]. p’-DDT [2266]. P. [2181, 4030, 4465, 4517, 5295, 4246, 3976, 3471, 4842, 3103, 4365]. Pa. [583]. Pachaly [76]. Pachygrapsus [2060]. Pacific [3912, 2089, 3477, 3454, 3540, 5342, 4958, 2530, 5365, 5155, 3967, 5278, 3109, 4042, 3038, 3144, 3193, 3376, 2318, 2320, 2364, 2080, 4868, 3527, 2771, 2883, 3670, 4569, 4352, 4648, 2800, 2057, 4061, 4872, 2160, 2431, 2485, 2250, 702, 2037, 2146, 2336, 3450, 3515, 2151, 3556, 5259, 5169, 2083, 3961, 1759, 1773, 1809, 1836, 2044, 2191, 2977, 3848, 5212, 3228, 4640, 3231, 1839, 2995, 3502, 4961, 3251, 4934, 4250, 3856, 5187, 1988, 2373, 4540, 4936, 4989, 3831, 4910, 3158, 3634, 3711, 3017, 2433, 3414, 5333, 2847, 4794, 2225, 2301, 5153, 2869, 4038, 981, 4896, 5110, 4687]. Pacific [4843, 5365, 2854, 2231, 339, 3299, 2377, 1643, 3575, 3819, 5036, 4605, 4710, 3717, 899, 3036, 2084, 5058, 1675, 3771, 3821, 4147, 1008, 1150, 5324, 802, 3182, 1727, 3869, 4149, 2143, 4672, 5144, 4312, 4543, 4730, 2702, 4978, 2738, 4889, 4304, 4673, 194, 1458, 2874, 3145, 4774, 2201, 849, 2705, 1858, 4944, 2566, 209, 2513, 2517, 2579, 4079, 1835, 4821, 2463, 2594, 2391, 1977, 2752, 4367, 3521, 6, 3555, 3222, 3601, 3378, 3570, 5114, 2269, 2282, 2790, 3329, 4988, 4381, 4583, 4584, 1887, 3516, 5285, 5027, 2666, 4130, 167, 2974, 4290, 4907, 5046, 2162, 2168, 3804, 4169, 1734]. Pacific [1792, 1779, 1883, 3401, 3681, 2362, 2570, 2045, 1845, 2421, 3403, 3849, 3405, 5182, 5138, 4596, 4153, 2496, 2593, 2909, 2768, 4956, 2532, 3077, 3561, 2971, 3240, 4444, 4974, 5035, 1786, 1897, 5235, 5240, 2884, 5328, 4906, 3242, 2883, 3348, 1802, 1969, 3930, 4982, 3130, 3997, 1689, 5289, 2984, 3859, 3514, 3522, 2169, 3436, 5023, 3274, 4156, 2079, 2140, 4696, 4487, 4629, 5128, 5201, 1820, 1818, 1877, 1982, 1983, 4697, 2409, 2665, 3143, 5030, 2070, 4027, 4259, 2145, 2471, 2473, 2601, 2590, 3690, 3879, 3861, 1935, 1328, 4825, 3484, 4555, 4957, 3653, 3796, 3865, 4638, 3306, 3395, 944]. Pacific [1012, 3565, 3883, 1732, 3325, 2641, 2888, 4813, 4698, 4097, 4177, 4360, 1603, 1794, 1976, 658, 808, 1676, 1667, 3390, 3663, 4007, 4420, 4531, 4851, 3841, 5366, 3039, 1695, 1687, 3905, 3148, 1080, 1094, 3049, 3531, 3680, 4634, 5072, 4853, 166, 221, 689, 2938, 4341, 328, 900, 1107, 1808, 1838, 2882, 2531, 4434, 2710, 4435, 4949, 4239, 5177, 2648, 4186, 3845, 4814, 2253, 4682, 4301, 1951, 3512, 5173, 4302, 2195, 2491, 2681, 3952, 2287, 2627, 3217, 3298, 3753, 5254, 3450, 3515, 4207, 5038, 5179, 5134, 3061, 3574, 1564, 1610, 1621, 3620, 1761, 5025, 2382, 3556, 2361, 4386]. Pacific [3631, 2982, 4539, 4248, 4654, 2065, 1783, 5329, 4840, 4380, 5147, 5040]. pacifica [2527, 2068, 2694, 2035]. pacificum [2231]. pacificus [4364, 2937, 5114, 3721, 4417, 4905, 3725, 2584, 2545, 2665, 4248, 3551]. Packages [3467]. packed [555]. packing [849, 808, 1363, 900]. page [3529, 687]. Pagellus [5255]. Pages [1263, 1264, 1265, 1266, 1267, 1269, 1271, 1273, 1276, 1277, 1278, 1279, 1280, 1281, 1282, 1283, 1701, 1702, 1703, 1705, 1706, 1708, 1711, 1714, 1717, 1719]. Pagrus [4808, 4871, 3783, 5176, 4127, 4668, 4370, 4256, 2510, 4221, 4467, 4492]. paint [2486]. painted [4757]. pair [5272]. paired [5116, 4959]. paired-end [5116]. paired-tow [4959]. Palaemonetes [2204]. Palatability [2058]. Palinuridea [2318]. Pallas [4648]. pallashii [1623]. Pallasi [2641, 2531, 2491, 4130, 2570, 2045, 3130, 4696, 3306, 4360, 4341, 4434]. pallasii [3514, 1603, 1624, 1640, 3574, 5025]. pallial [4862]. palmensis [3516]. Palps [1926, 721]. pamphlet [676]. Pamunkey [4516]. Panama [2300, 93, 182, 2604, 4817]. Panamic [5282]. Panay [1331]. Pandalid [2263, 2752, 2320]. Pandalidae [3193, 4630, 3076]. Pandalopsis [4630]. Pandalus [3228, 3708, 2202, 2628, 3652, 3846, 4255, 4960, 2459, 2626, 2709, 2752, 3460, 2585, 2686, 2845, 2515, 3076, 3194, 3625]. pandora [5255]. Pandulus [3295]. panel [4126]. panels [4036, 4210, 4857]. Panguitch [359]. paniformis [3403]. Panopeus [3107, 3103, 3099]. Pantosteus [1009]. pantropical [4640, 4259, 3879, 4097]. Panulirus [3216, 1517, 2877, 3779, 4542, 4688, 3773, 4929, 3472, 4889, 3427, 3241, 2339, 3645, 3835, 4855, 3002, 3752, 4222, 4384, 3636, 2378, 3175, 3602, 3344]. papers [1139, 102]. Papua [2920]. Paralabrax [3582, 4030, 4268]. paralarvae [5072, 4302, 4247]. Paralarval [3686, 4034, 3554]. Paralepis [2089, 2762]. Paralichthyidae [4984, 3685, 3976]. Paralichthys [1492, 3563, 5002, 3950, 4453, 5079, 5305, 3700, 1690, 4150, 4627, 3666, 5003, 4876, 4943, 4446, 5117, 3903, 3685, 2180, 3976, 3479, 4449, 2237, 2612, 5205, 4268, 3743]. Paralichtyhys [1482]. Paralithodes [3777, 3283, 2316, 3254, 4445, 5100, 2286, 5069, 3075, 4875, 1873, 4683].

Paralomis [3832]. Paralytic [2500]. Parameter [3032, 2074, 2670, 5133, 5340]. Parameterizing [4777]. Parameters [2219, 3467, 2564, 2640, 3706, 2418, 3751, 3493, 3381, 4364, 3163, 5271, 3837, 3821, 5279, 4909, 3804, 5140, 4429, 5159, 5302, 4549, 5239, 2590, 4467, 1881, 5019, 5282, 3998]. Parana [4571]. Paranthias [4521]. Parasite [2027, 3051, 3677, 3596, 2361, 3827, 972, 1324, 3031, 442, 2604]. Parasites [2615, 2231, 2269, 2120, 3229, 3239, 3307, 3202, 1171, 1313, 3174, 2973, 4540, 1677, 695, 994, 1473, 5285, 975, 1055, 1129, 1178, 1409, 4554, 4208, 3256, 4247, 1404, 1310, 1476]. Parasitic [1739, 2854, 1223, 1570, 1580, 1967, 4319, 1043, 1230]. Parasitism [4205, 1528, 4213, 3115, 4407]. Parasitized [3283, 2984]. Pared [5164]. parent [1378]. Parental [3116, 2353]. Paris [848]. Park [999, 953, 1819, 4598, 517, 956, 4632]. parma [3568]. Parmaturus [4324, 3470]. parmifera [4892]. Parophrys [3244, 2956, 3476, 2953, 2970, 4861, 1951]. Parorchis [1874]. Parrotfish [2634]. part [680, 1779, 1087, 370, 527, 4067, 1210, 3923, 4103, 3921, 4093, 4173, 3750, 3754, 1593, 1612, 1631, 1634, 1769, 1696, 4063, 1678, 4049]. Partial [2177, 5233]. participation [545]. Particle [3132]. Particular [3403]. Particularly [2109, 1083, 4247]. Particulate [2054, 2140]. partition [3099]. partitioning [4811, 3345, 4919]. parts [1820]. parturition [4656, 4812]. Party [3617]. pass [3962]. Passage [2106, 1986, 1952]. Passamaquoddy [3131]. passed [631]. Passenger [3205]. passing [1931]. Passive [2838, 5293, 4990, 4968, 5161]. Past [269, 1054, 1021, 1064, 1040, 1154]. patachonicus [5151]. Patagonia [4160, 4329, 4702]. Patagonian [4309, 3886, 5088, 5183]. patch [3974, 4163]. patches [2645]. Patchiness [3258, 3733, 3930]. patchy [4700, 4146]. patent [101]. path [2905]. Pathfinder [3471]. pathogenic [1684, 2463]. Pathology [2973]. Pathways [3448, 4750]. Patinopecten [4381]. Patos [3296]. patronus [2912, 5098, 5214, 3713, 3129, 3171, 5209, 2879, 3278, 2658, 3320, 3786, 3920, 5146, 5335, 3389, 4681, 3417, 4146]. Pattern [2261, 3583, 3355, 2376, 5238, 4725, 1482, 3903]. Patterns [3563, 2802, 3078, 3566, 2342, 2914, 3467, 2671, 3201, 2660, 2228, 3101, 2168, 4736, 3008, 3205, 3208, 2154, 3246, 2174, 3525, 2145, 5206, 3508, 5211, 2916, 5130, 3347, 5118, 3167, 4441, 4948, 4798, 4123, 5102, 5106, 4165, 3048, 5365, 4779, 4990, 4671, 4254, 4797, 3660, 4092, 3787, 5017, 4995, 4481, 4988, 4381, 4583, 5285, 4066, 2162, 3810, 4172, 5020, 4898, 4292, 5168, 4293, 5042, 4569, 4511, 5230, 4296, 4122, 3824, 4851, 4805, 3802, 4874, 3669, 4918, 4146]. patula [1564, 1621]. pauciradiatus [4096]. paucispinis [4831, 4787, 4786, 4982, 4490]. paucus [5200, 3294]. Paul [3273, 3353]. Paulo [2655]. Pavlof [3228, 3708, 4654]. Pavlova [2754]. Pb [3872]. Pb-21 [3872]. PCR [4002, 3840]. PCR-amplified [4002]. PCR-RFLP [3840]. peacock [4952]. pealei [4024, 4211, 2876, 3333, 2999, 4039]. pealeii [4481, 4923]. pearl [706, 820, 1149, 1167, 1367, 962, 766]. pearl-oyster [1367]. Pearls [766, 1148, 1149, 1367]. pearly [1166]. Peat [31]. Peat-bogs [31]. Peck [2228]. Peconic [2827]. pectinata [5297, 5298, 5188]. Pectoral [2266, 3262]. pectoralis [5201]. peculiarities [54, 1153]. Pelagic [3078, 2524, 3454, 4191, 2046, 2462, 2421, 2956, 2721, 2545, 2890, 2326, 2313, 2141, 3534, 4912, 2289, 3688, 2674, 4558, 1832, 3868, 3003, 5057, 1947, 3762, 4903, 3803, 3806, 5194, 4349, 4849, 5344, 4729, 2738, 4854, 4231, 2547, 4525, 4150, 2555, 2762, 3192, 4621, 4791, 4653, 4019, 4198, 4659, 1024, 4200, 4740, 5023, 2725, 5269, 4981, 4344, 3796, 4029, 3757, 4337, 3988, 3964, 5248, 5210, 3764, 3805, 4406, 1808, 1091, 3571, 4635, 4667, 5092, 4228, 5040]. pelagica [4830]. pelagicus [4603]. pelamis [4403, 5212, 3198, 2669, 3477, 2373, 2895, 2744, 1675, 2898, 2383, 3329, 2968, 3202, 3265, 4420, 4851, 4186, 1964, 2447, 2896, 1670, 2920, 2151, 2065]. Pelecypoda [2997, 3097, 3163, 3449]. Pella [2675, 2670]. PEN [3165]. Penaeid [2615, 4325, 2421, 3695, 2429, 3005, 3508, 2101, 3448, 2041, 2769, 1990, 3516, 1942, 2468, 4258]. Penaeidae [2014, 3569, 4414, 1939, 2232, 2678, 3948, 3516, 1653, 2070, 3456]. Penaeidea [2522]. penaeids [4365]. penaeoid [3965]. Penaeoidea [3516, 3182, 3364]. Penaeopsis [2678, 2769]. Penaeus [3569, 2843, 2772, 2181, 2032, 1941, 1981, 3902, 3057, 1862, 2390, 2007, 3948, 3026, 3737, 2468, 2283, 1771, 2000, 3156, 3339, 1993, 2685, 1653, 4047, 3230, 1971, 1650, 4272, 1815, 4365]. Penaeuse [4246]. Pend [289]. penicillata [3178]. Peninsula [1524, 1572, 4778, 4708, 3004, 1593]. Peninsular [2996, 4506]. Penning [451, 519]. Penobscot [479, 163, 718, 650, 4395, 5263, 1118, 3194]. penpoint [4796]. pens [2686]. Pensacola [2010, 661, 824]. Pentacerotidae [3827]. Pentapodus [4800]. People [2267]. Peponocephala [2471]. Peprilus [2874, 3649, 3406, 4122]. Per-recruit [5147]. Perca [1821, 4216, 4842]. Percent [3069, 3842]. Percentage [2833]. perception [4980]. Perch [2998, 2566, 3500, 3653, 2803, 3395, 4936, 896, 411, 1821, 884, 1688, 5138, 4216, 1501, 51, 4842, 4701, 986]. Percichthyidae [4701]. Perciformes [5115, 4812, 4596, 4291, 4701]. Percoidei [3929]. Percophidae [4596]. percula [2943]. perezi [4319]. Performance [2056, 3612, 3625, 4705, 5296, 4929, 4992, 4754, 4330, 4981, 5043, 2902, 5180]. Period [308, 3570, 2956, 2580, 3143, 5039, 4309, 1236, 5077, 5227, 4432, 5108]. periodical [1024]. periodicities [4883]. Periodicity [4343, 4348, 3106, 3155]. Periods [2556, 4129, 4545, 4868, 2465, 4465]. peripheral [1157]. periphery [5324]. Peritoneal [1505]. permanent [5026]. permit [4461]. perpetuity [3714]. persistence [4790, 5245]. Persistent [4066]. personatus [5187]. personnel [231]. persons [164, 264]. Perspective [3084]. perspectives [4513]. Peru [2094, 1349, 5266, 3145, 5225, 3899, 3531, 4177]. Peruvian [3502, 3531]. pest [4258]. pesticides [1996]. Petition [692]. Petrale [2545, 3892, 4526, 5292]. Petroleum [2386, 2932, 2375, 2531, 2850]. Petromyzon [3165, 3268, 3297, 5251]. Petromyzontidæ [137]. Pettaquamscutt [3487]. pfluegeri [3545]. pH [2045]. pharaonis [4285]. Phase [3386, 4953, 4184, 3782, 3406]. phases [1618]. Phenotypic [2952].

Pheromonal [2426]. Pheromone [2060]. Philippine [1239, 1322, 1334, 2812, 2829, 1670, 1679, 1671, 1674]. Philometra [4760]. Phoca [2284, 4848, 5067, 3065, 4944, 5068, 5168, 3820, 4629, 3161, 2822, 2856, 4638, 4813, 3245, 4239]. Phocoena [3444, 3447, 4390, 4193, 5293, 5136, 3983, 4166, 3207, 3823, 4980, 3100, 5314, 4201, 3220, 3585, 3092]. Phocoenoides [3576, 2882, 3659]. Phoenix [1691]. Pholidae [4796]. Phoronida [1742]. phosphate [1728]. Phosphoglucomutase [2468]. Phosphorus [1730]. Photic [2901]. Photobehavior [2442]. photogrammetry [3861]. photographed [2000]. Photographic [2547, 3558, 1997, 3971]. photographs [3994]. photography [3844, 1338, 1172]. Photololigo [4006, 3875]. Photoperiod [2185, 2407, 2916, 2945]. Photophore [2154]. Photosensitive [2674]. Photosynthetic [2195]. phrictus [2799]. Phycis [3806]. Phyllodoce [2750]. phyllosoma [4889]. Phylogenetic [2154, 4055]. phylogenetically [3520]. Phylogeny [3567]. Phylogeographic [4442]. phylogeography [4202]. Physalus [3277, 2711, 3598, 3507]. Physical [2628, 2770, 2001, 4098, 2030, 2741, 2165, 988, 1073, 4096, 961, 5082, 1598, 4051]. Physics [1744]. Physiological [1315, 4484, 3141, 1665, 4615, 1317, 5232]. physiology [1431, 1335]. Phytoplankton [2082, 2189, 2575, 2359, 2037, 1546, 1730, 1803, 2587, 1898]. Piankatank [116]. picarel [5238]. Picker [2423, 73]. Picker-Type [2423]. Pickerel [253]. pictum [4757]. picturatus [5247]. Piedra [888]. Pig [2058]. pigfish [1475, 3095]. Pigmentation [2145]. Pigmented [2710]. Pigments [2901, 3443]. pigmy [694, 615]. Pike [253, 5265]. pilchard [1727, 1668, 4380]. pilchards [644]. pilchardus [4081, 4828, 2356]. Pilot [3703, 2735]. Pimelometopon [2365]. Pinealectomy [2407]. Pinfish [2179, 2010, 2464, 4155, 4512]. Pinguipedidae [4702]. pinguis [2254]. Pink [3228, 2063, 2417, 2516, 2839, 2202, 2628, 1758, 3652, 2677, 3026, 2114, 2852, 3510, 5049, 1941, 1981, 4756, 1636, 1641, 4026, 1862, 3937, 2613, 1763, 3846, 3737, 1966, 4812, 4778, 1901, 1936, 1985, 4334, 1993, 4398, 4076, 2865, 4715, 4047, 1681]. pinniger [2706]. Pinniped [4502, 4501]. pinnipeds [2873, 3019, 3845, 4663]. Pintsch [540]. pipes [122]. piscatorius [5197]. Pisces [2466, 2017, 3724, 2767, 5310, 4707, 4413, 3020, 3806, 2210, 2611, 3860, 2878, 2931, 3054, 3366, 3746, 4908, 3349, 3494, 5151, 4525, 3128, 2193, 3577, 2196, 3212, 4668, 2110, 2757, 2758, 4289, 3429, 3791, 4096, 3407, 4788, 2456, 4429, 2510, 2616, 3670, 2372, 2353, 2355, 2921, 4447, 3243, 5219, 2345, 4115, 3877, 2226, 2305, 3685, 2586, 4399, 2308, 3976, 2057, 2073, 3616, 2888, 2890, 2606, 2620, 3209, 2817, 2828, 2285, 2435, 2710, 4086, 3753, 2365, 2989, 3095, 4303, 3085]. Piscicultura [560]. piscicultural [888, 880]. Pisciculture [500]. piscine [2748, 3165]. Piscivorous [668, 4129, 4178, 5090]. piscivorus [2687]. pistilliger [4369]. pits [5223]. Pitt [34]. Pituitary [2199]. place [791]. placed [367]. places [926]. Placetron [3144]. Placopecten [2466, 3458, 4544, 4979, 1940, 4677, 2929, 5172]. plagiusa [4240]. plagusia [3661]. Plaice [2293, 5010, 4866, 188, 4834, 2428]. plainfin [3134]. Plan [68, 1354, 1350, 159, 1399, 1398, 1358]. planipes [3868]. Planktivore [2878, 2931, 3054, 2943]. Planktivorous [2872]. Planktology [2399]. Plankton [2094, 2445, 2632, 2139, 3189, 2164, 1798, 1544, 1545, 3505, 4087, 1796, 3715, 1562, 2886, 2152, 1018, 1145, 5260, 1250, 3539, 1620]. Planktonic [3843, 2672, 1546, 2440, 2368, 1729, 1971, 1843]. planktonic-stage [1971]. plans [4508]. Plant [1747, 3072, 298, 402, 374]. planted [403, 866]. Planting [390, 752, 233, 90, 227, 656, 871, 965, 481, 165, 897, 662, 747, 141, 890]. Plants [450, 1546, 668, 1232, 1031, 179, 218]. plasma [5174, 4204]. plastic [2807, 5254]. Plata [4327, 4331, 4567]. Plateau [3876]. Plates [1, 112, 113, 312, 922, 1274]. Platessa [188, 2293]. platessoides [4866, 4834]. Platforms [2478, 3617, 3657, 4787, 4316, 4803]. Platycephalus [3739]. platyceras [3460, 2845]. platyceros [2585, 2686]. Platyhelminthes [3790]. platypterus [2862, 4642, 4935, 2021, 4738, 3301, 4996]. platypus [3283, 3254, 3075]. Platytroctids [3670]. plausible [4861]. plea [1422]. pleas [1125]. Plecoglossus [4483]. Plectropomus [4799, 3864, 3981, 4888, 4402]. plei [4039]. pleopod [4688, 3140]. Pleoticus [2563]. Pleurobrachia [2307]. Pleurogrammus [3996, 4869, 4868, 4172, 4095, 4818, 4663, 4305, 4306, 4664]. Pleuroncodes [3868]. Pleuronectes [2293, 4100, 4182, 4070, 5010, 3862, 3789, 4094, 4118, 4176, 4065, 4204, 3891, 3839]. Pleuronectid [2956]. Pleuronectidae [4352, 4192, 2881, 2820, 2832, 2890]. Pleuronectids [3382, 4105]. Pleuronectiformes [4280, 4192, 3661, 4144, 4984, 3732, 4962, 2295, 2710, 4240]. plicatus [3300]. plotosid [5016]. plumbeus [5031, 4328, 4938, 4244, 4781, 4780, 5377, 2921, 3211, 5272, 4013, 3792]. Plume [3662, 4885]. plus [3944]. Plymouth [56]. poacher [3817]. Poachers [558]. Poatomus [3775]. pocket [101]. Podothecus [4115]. Poecilostomatoid [3260]. Poey [275, 4579]. Pogonias [4145, 3813, 4168, 4418, 5022, 3815, 4162]. Pogonichthys [3086]. poikilothermal [1563]. point [4314, 4925, 1681]. point-based [4925]. points [4873, 4924, 4925, 119]. Poisoned [156]. poisoning [2500]. Poisonous [760, 788]. polaris [2922]. Policies [3638]. Policy [2328, 2793]. politus [3646, 2917, 3185, 2505, 3022]. polkadot [5341]. Pollachius [4237]. Pollock [3572, 3611, 3373, 2712, 3380, 3579, 3586, 52, 2883, 5089, 2741, 3526, 2726, 3871, 5053, 3977, 3953, 3080, 4672, 3253, 3704, 4795, 3111, 3608, 4859, 3893, 4737, 3901, 3722, 4530, 4275, 4456, 4237, 4489, 4341, 4434, 3949, 4663, 4665, 4986, 4378, 4703, 5021, 4664]. Pollock-fishing [52]. Pollutant [2564, 2158]. pollutants [2683]. Pollution [1746, 2699, 1651, 1642, 1632, 698, 1585]. Pollution-associated [2699]. polyacanthus [4736]. Polyacrylamide [2583]. Polychaeta [2116]. Polychaete [2750]. Polychaetous [2635, 1297, 1205]. Polychelidae [2731]. polychlorinated [3055]. Polydactylus [3609, 4543, 2698]. polyedra [2575, 2587]. polyethylene [2394]. Polyhaline [3392]. polylphemus [4793]. polymer [2394]. Polymodal [2396]. polymorphic [5341, 4125, 4815]. polymorphism [5101, 2468, 4043]. Polynomial [2660]. polyphasic [4132]. polyphemus [3150, 3402, 4769, 4933, 4620, 4775, 5317, 4597, 4578]. Polyprion [4230, 5159]. Polyspecific [3689]. polyspinis [5280, 5361]. Polystream [2024]. polyxystra [5355, 5018, 4932, 4407]. Pomacentrus [4882]. Pomadasys [4612]. Pomatomidae [4429]. Pomatomus [4557, 4249, 3629, 4129, 4282, 4283, 4760, 3895, 2715, 4429, 3782, 3812, 3439, 4844, 4805, 4915, 5203, 4809, 3828, 4727, 4921]. Pomolobus [46]. pompano [3882, 4940]. Pompanos [1880, 30]. Pond [3327, 26, 851, 273, 8, 236, 32, 220]. pondfish [1499, 1530]. ponds [558, 38, 255, 872, 450, 149, 56, 728, 850, 651, 25, 316, 4972, 825, 4258, 31, 1038, 358, 3230, 428, 179, 218, 1620, 414, 606]. Pontellina [2280]. Pontoporia [4670]. Pool [1179, 3749, 1117, 1162, 3847]. Poor [2037, 5011, 180]. poorly [3675]. pop [5305, 4472, 4945, 4691, 4693, 4621, 4684]. pop-up [5305, 4472, 4945, 4691, 4693, 4621, 4684]. popular [351, 881]. Population [2167, 3961, 3379, 2977, 2926, 3394, 3756, 3011, 4934, 5049, 2092, 2875, 2527, 2625, 5028, 4135, 5095, 3609, 2240, 1727, 3981, 3472, 2263, 2990, 3210, 4196, 4506, 3109, 2566, 2857, 3652, 4464, 5367, 2211, 2418, 5198, 3475, 5270, 3278, 4396, 3066, 5108, 4762, 4157, 3693, 3598, 3175, 3176, 2431, 2935, 4955, 4203, 3667, 3764, 3152, 3568, 3510, 4301, 3512, 3371, 2537, 3912, 3292, 3853, 4864, 4560, 4559, 4893, 4453, 5189, 3163, 4561, 4801, 4106, 4919, 4562, 1909, 3336, 4772, 3876, 5365, 4769, 4542, 4644, 1821, 5155, 3719, 4366, 3827, 5027, 4909, 5341, 1942, 3807, 4291, 4085, 4566]. population [5100, 4966, 1837, 4093, 4173, 4136, 5274, 4025, 4549, 5143, 3106, 4825, 3881, 2205, 3289, 3734, 1729, 4080, 4177, 3887, 3841, 4318, 5150, 4263, 1950, 2679, 4874, 5000, 4918, 4269, 4492, 3850, 3720, 4181, 4116, 4323, 5169, 5123]. Populations [2094, 3402, 3449, 2605, 2857, 2860, 2312, 2467, 2774, 2438, 3313, 2569, 3636, 2411, 2387, 3049, 3033, 1759, 1918, 3480, 4500, 1807, 4708, 3733, 5372, 1981, 4604, 3141, 2451, 2553, 4930, 2813, 1885, 3908, 5218, 4171, 4173, 5042, 3647, 4074, 4174, 3989, 5270, 4133, 4398, 4335, 3881, 4924, 4925, 1686, 3922, 4179, 4955, 1799, 1961, 1943, 3770, 5247, 4181, 4746, 1761, 5025, 5169, 4045]. porbeagle [4312, 4528, 4486]. porbeagles [5274]. Porcellanid [2216]. Porcellanidae [2201]. porcupinefishes [2676]. Pores [2261]. PORGS [4919]. Porgy [2990, 2510, 3553, 3018, 4808, 4871, 5176, 4127, 4370, 4256, 4467, 4492]. Porichthys [3134]. Porpoise [3444, 3447, 3207, 3823, 3576, 2276, 2473, 712, 3659, 552, 4390, 5293, 5136, 3679, 3983, 4980, 2974, 3100, 2107, 4201, 4029, 284, 3092, 2882, 612, 1287]. Porpoise-fishing [552]. Porpoises [2231, 2145, 4193, 4166, 5314, 3220, 464, 2486]. Port [3006, 2253]. Portable [2636]. Portions [2018]. Porto [1197, 1202, 1196, 1214, 1211, 1207, 1200, 1204, 1195, 1206, 1217, 1208, 1199, 1213, 1205, 1201, 1203]. Portugal [4900, 5037, 5239, 5011]. Portuguese [175, 197, 2227, 2235]. Portunidae [2448, 2680, 4433, 2824, 2329, 4603]. Portunus [4603]. POSEIDON [4222]. Position [2154, 487]. Positional [2099]. possessing [4639]. Possibilities [1127, 1526, 1141]. Possible [2568, 3204, 2239, 2809, 1380, 2862, 3948, 4231, 3262, 2370, 4973, 4375, 3233, 2500]. possibly [1722]. Post [3746, 4046, 1939, 2538]. post-larval [2538]. Post-smolt [4046]. Post-yolksac [3746]. postcapture [2751]. Posthatch [2494]. Postlarva [2721]. postlarvae [4756, 3902, 3516]. Postlarval [2360, 2464, 1689, 2831, 4343, 1955, 2188, 3920]. Postmortem [2283]. Postnatal [3222]. Postovulatory [3145, 3969, 4828, 5003]. Postrelease [5352, 5305, 5023, 4472, 4945, 4992, 5371, 4693, 5269, 5323, 5232]. postsettlement [4296, 4262]. postsmolt [5050, 5174, 4763]. Postsmolts [3424, 4254, 4789]. postspawning [1855]. poststratification [4728]. Postulated [2897]. pot [4961, 5082, 4614, 4340]. potamogetons [1466]. Potential [2092, 3677, 3662, 2571, 2574, 4775, 2580, 4786, 3687, 2750, 3865, 3665, 4989, 4210, 3834, 5280, 3819, 4012, 3458, 4821, 2546, 4947, 4184, 3927, 4820, 4044, 5378, 5221, 1866, 4208, 5364, 4950, 4262, 3824, 4401, 4434, 5177, 5083]. Potomac [210, 700, 388, 477, 344, 385, 541, 659, 864, 324, 91, 567, 698, 586, 837, 217, 5273, 985, 334]. pots [4411, 4340, 4875, 5000, 4662, 5145]. Pound [583, 555, 372, 1466, 1797]. pound-net [1797]. pounder [1825]. pounds [1696]. Power [3744, 3072, 4394, 4373, 4171, 4175, 5300]. Practical [3625, 4053, 1368, 5362, 1145, 1363]. practice [1505]. practices [5073, 1411]. præstabilis [37]. Pranesus [2254, 2556]. Prawn [3364, 4036, 4126, 4090, 4210, 4347, 4800, 5121, 4246, 2701, 158, 2585, 2686, 2845, 4533, 4272]. prawn-trawl [4036, 4126, 4090, 4210]. prawns [4126, 4251, 4937, 3460, 4246, 3998]. Pre [2169]. Pre-exploitation [2169]. Precision [2673, 4826, 3530, 4751, 3844, 4816, 4393, 3721, 4466, 3726]. Preconstruction [1893]. Predation [2873, 3228, 3590, 3382, 3324, 3662, 2064, 3561, 3533, 3446, 3339, 2079, 3149, 4021, 3414, 5167, 4012, 4415, 5114, 4684, 3019, 3683, 4629, 2409, 2557, 5291, 4810, 2943, 4752, 3830, 4921]. Predator [2556, 2098, 3602, 3891, 2491, 4525, 3340, 5045]. predators [3468, 4415, 2814, 5104, 4957, 4810]. Predatory [3694, 2491, 5196, 4823]. predictability [3991]. Predicting [2656, 5177, 3720, 5133, 1656, 5229]. Prediction [5036, 1769, 4951, 3069, 4157, 5347]. predictions [3140]. predominant [3025]. Preface [1053, 1754, 1910, 1215]. prefer [905]. preference [4584, 5045]. preferences [3468, 4803, 5240]. Prehatch [2494]. prejuvenile [4343, 4681]. Preliminary [4453, 4719, 2240, 946, 1977, 4382, 5195, 4161, 4537, 4735, 71, 119, 217, 754, 2101, 4457, 1650, 1093, 999, 4255, 3608, 3282, 963]. Premigratory [2097]. Preparation [2045, 762, 4734, 4816]. Preparations [2075]. prepare [648]. prepared [1968]. preparing [838, 1424, 1417, 756, 1193]. prerecruit [4033]. Presence [5274, 3965, 4317, 3507]. Present [2460, 767, 990, 1054, 1021, 1064, 512, 1154, 484]. Preservation [2050, 2858, 722, 790, 4749, 210, 3780, 758, 4185, 1161]. preservative [4866]. preservatives [529]. preserve [256]. preserved [3582, 2651, 4668]. Preserving [809, 216, 409, 286, 1367]. presettlement [4890]. prespawning [1855]. Pressure [3001, 2425]. pressures [355]. Prevalence [3397, 4760, 4907, 3340]. prevention [1406]. previous [1072, 3723]. previously [3770]. Prey [3127, 2631, 2196, 3310, 3630, 2927, 2750, 4097, 2098, 3867, 5045, 2713, 4551, 4089, 4989, 4283, 4952, 4149, 4944, 3276, 3468, 1356, 4200, 2556, 3820, 4886, 4638, 4432, 3710, 4178, 5358, 4099, 3245, 3891, 4434, 2491, 4134]. preying [402]. Prezoeae [2249]. Pribilof [1909, 1480]. Price [3267]. Prices [288]. Primary [5293, 2445, 5285, 2080, 2090, 2113, 2747, 4305, 5345, 49]. primers [4914]. primitive [5346]. primolecta [1898]. Prince [2631, 2259, 4130, 4293, 1634, 4360, 1623, 4341, 4434]. Principal [3032, 2704, 2752, 1973]. principle [4790, 1418]. principles [929]. Prionace [5266, 5194, 4748, 4729, 5278, 2144, 2157, 2739, 2572, 4591, 2713, 5357]. Prionotus [3237, 4154, 4465, 4517, 2620]. prior [5290, 4526, 4790, 4611, 2701]. Pristipomoides [4026, 4252, 3937, 4564, 4549, 5295, 4076, 5135, 4354, 3071, 5316, 3471, 3176]. Pristis [5297, 5298, 5188]. private [5322]. prizes [68]. probabilities [4601, 4777]. probability [4185, 4891, 4683, 3830]. probability-based [4891]. Probable [2677, 2086, 3218, 4313, 3852]. probably [2659]. probatocephalus [3644, 3548, 3633, 4086, 5204]. probe [3907]. Problem [2238, 1370, 1183, 1147, 1412]. Problems [3475, 3616, 3264, 2644, 1592, 1061, 1482, 1145]. procedure [4426, 2683]. procedures [4416, 4261, 3755]. Proceedings [1139]. Process [2857, 286, 3654, 340, 4775, 809, 5300, 5317, 4649, 1367]. Processed [2036]. Processes [3566, 3583, 3811, 4296]. Processing [2512, 2253]. Processor [2910]. procurement [174]. procuring [250]. Prodelphinus [1287]. produced [4393, 4394, 1586]. Producers [2090]. producing [2438]. Production [2516, 2520, 2605, 3169, 2082, 1588, 2934, 2074, 2354, 2252, 2761, 2826, 2080, 3461, 3188, 272, 2113, 3528, 2368, 2670, 3568, 2037, 3515, 4883, 3134, 1354, 4769, 4604, 1350, 2640, 1609, 4692, 2653, 1688, 1526, 4444, 5022, 5161, 3072, 4567, 5198, 4855, 3859, 4334, 1830, 4076, 3306, 4550, 1358, 398, 2498, 4041, 1898, 1620]. Productivity [3007, 2275, 2445, 3564, 2091, 4983, 5213]. Products [2177, 2114, 2061, 364, 839, 687, 698, 357, 806, 712, 1161, 2683, 220]. productus [1773, 3848, 2995, 2377, 3575, 2045, 3403, 3077, 3242, 2984, 3522, 4156, 3325, 4083, 5329]. Professor [275, 303]. profile [4093, 4173, 5190, 4931]. profiles [5118]. profit [95]. Profitability [3154]. Progeny [2677, 3621, 1985]. Progesterone [3659]. Program [2633, 4929, 2273, 2141, 2396, 5307, 3805]. Programming [2745, 2693, 2047]. programs [4772, 3542, 5165]. Progress [96]. progression [5356]. prohibit [692, 425, 453]. prohibited [293]. project [1893, 2243]. projection [1672]. Projections [3210]. projector [5335]. Prolactin [2144, 2157]. Prolactin-like [2144, 2157]. prometheus [4218]. Promethichthys [4218]. promoting [1354, 1350, 1358]. Propagating [617, 149, 263]. Propagation [693, 643, 517, 1405, 1038, 90, 118, 1511, 753, 656, 1510, 133, 457, 70, 1036, 959, 348, 1346, 1435, 586, 837, 1132, 18, 708, 1108, 398, 1107, 23, 75, 400, 919]. proper [881]. Properties [2770, 2038, 2213, 4536, 2030, 2212, 1972, 1855, 4593, 1968]. Propertires [1462]. property [2363, 480, 508]. proportion [4068]. Proportions [2297]. Proposed [294, 2063, 276, 348, 709, 268, 393, 2154, 259, 2151, 682, 942, 2502, 794, 102, 210, 184, 8, 4223, 196, 221, 278, 24]. Proposition [648]. propriety [336]. Prosomal [4933]. Prosomal-width-to-weight [4933]. Prosopium [1945, 2221]. prospective [4961]. prospects [512]. Prospectus [592]. protandric [4355]. Protease [2984]. protect [700]. protected [4707, 4930, 4865]. Protecting [719, 568]. Protection [29, 1125, 1146, 1405, 1130, 1141, 1362, 1166, 280]. protective [4617, 1359]. Protein [2036, 2213, 2045, 2488, 2764, 2061, 4063, 2319]. Proteins [2144, 2157, 1459]. Proteolysis [2984]. Proteolytic [2075]. Protista [2361]. protocol [4959, 4373]. Protogynous [2365, 4705, 5189, 4873, 4579, 4165, 5015, 4920, 4871, 3799]. protophytes [49]. Protothaca [2259]. Protozoa [1752, 1192, 1226, 49]. protozoan [1677, 994, 975, 1043]. protozoean [1939]. provide [700, 210]. Providence [521]. Providing [3951]. province [4161, 887, 932, 845, 671, 826, 11]. Provincetown [47]. provisional [1320, 4384]. prowfish [4634]. Proximate [2887, 3672]. proximity [5349]. proximus [2883]. Psednos [4641]. Psettodes [5346]. pseudaharengus [3740]. Pseudaristeus [3364]. Pseudocaranx [4383]. Pseudocyttus [3451]. pseudoharengus [4216, 2906, 3025, 3215, 2923, 1855]. Pseudopentaceros [3454, 3827, 3764, 3571]. Pseudopercis [4702]. Pseudophyllidea [3300]. Pseudopleuronectes [2764, 3603, 4990, 2881, 2580, 2220, 2381, 4300, 5205]. Pseudopleuronectus [3104]. Pseudopriacanthus [1876]. Pseudorca [2801]. Psorosperms [968, 975]. Psychrolutes [2799]. Pterois [5279]. Ptilichthys [2398]. Ptychocheilus [1829, 3351]. public [313]. Publications [3626, 437, 855, 296, 676]. published [214, 329]. Puerto [3394, 4336, 3742]. Puerulus [2378]. Puffer [2271]. puffers [2717]. Puget [2475, 3382, 4158, 5080, 812, 2463, 4094, 5233, 3498, 3108, 3591, 2589, 2846, 2845, 5012, 1645, 4204, 171, 1046, 2714]. pugettensis [2079]. pugilator [2194, 2118, 2480]. pulchrum [2365]. pulex [2472]. Pulley [5223]. pulp [1665, 1632]. pulsating [1757]. pulsed [3811]. pumilio [2193]. pump [2071]. pumping [1031]. punctata [3929, 4151, 4510]. punctatus [3352, 4436]. punctipinnis [2872, 3136]. punctuatus [571]. Pungitius [2265]. Punta [194, 4743]. pup [3989]. Pupfish [2034]. pupping [5227]. pups [4706, 3353]. Purchase [275]. Purple [2098, 4746]. purpose [658]. Purposes [2793, 1417, 881, 938]. purpuratus [2098]. purpurea [3346, 3517]. purpureus [2435]. Purse [3543, 2730, 2164, 2093, 3425, 4640, 3803, 2838, 4761, 4524, 4821, 2971, 87, 232, 425, 4468, 3805, 3952, 453]. Purse-Seine [2093, 4640, 4761, 4524, 4821, 232, 4468, 3805, 3952]. purse-seines [87]. put [191]. putnami [4100, 2900, 2881]. Putting [606]. Pycnogonida [1751]. pygmaeus [3504, 3801]. Pygmy [2193, 1945, 2221, 3817, 4659, 3683].

Pyloric [37]. Pyramidellid [3452]. pyrene [3612, 3085]. pyrifera [3126, 2324].

quadracus [2]. quadrimaculatus [3082]. quadriseriatus [3366]. quadrituberculatus [5010]. Quahog [2960, 3255, 2707]. Quahogs [3123, 3146]. Qualitative [3012, 3147]. qualities [627, 149, 445, 237, 368, 187]. Quality [3403, 2973, 3696, 4943, 5088, 5354, 4703]. Quantification [4691, 5069]. quantify [5167]. Quantifying [4563, 5074, 5217, 5014, 2498, 4921]. Quantitative [3147, 3280, 1586, 2307, 3012, 4616, 4674, 3951, 5157]. Quantity [3251]. Queen [3394, 2566, 3354, 3584, 3618, 3696, 3852, 4041, 3733, 4644, 4721, 3824, 3718, 3885, 4206]. Queenfish [3646, 2917, 3185]. Queensland [3985]. queenslandicus [4190]. Queries [352]. quernus [5007]. Questions [786, 408, 129, 733, 299, 152]. quillback [4694, 5034, 4931]. quillfish [2398]. Quinnat [1249, 1107, 504, 505]. quinquedens [2696, 2549, 4899, 2200, 5178, 3121]. quinquelineatus [4025]. quota [5313, 3714]. Quoy [3735, 3809].

R [2257]. R. [507, 303]. Ra-226 [3872]. rabbits [1968]. race [664]. Races [1640, 1817, 40]. Rachycentridae [3768]. Rachycentron [4452, 3913, 4389, 5095, 5226, 3768, 4014, 4253, 5060, 5287]. radar [4787]. radiata [3710, 4699, 4733]. Radical [1406]. Radicals [2061]. Radio [3220]. radioactive [1796, 1938, 1803]. Radiocarbon [4145, 5031, 4905, 4694, 5159, 4985]. radiologic [2867]. Radiometric [4425, 3943, 4680, 4711]. radiometry [3933]. Rafinesque [3056, 1828, 571, 4333]. Raft [1869]. rafting [725]. Raia [771]. Rainbow [915, 3386, 3674, 3612, 2357, 2402, 2900, 5281, 3412, 2573, 653, 2184, 1387, 1373, 234, 1801, 1108, 2748]. Rainfall [3063]. raise [672]. Raised [2452, 3538]. Raising [601, 734, 151, 362, 258, 87, 1114]. Raja [4913, 4588, 3710]. Rajidae [2355, 4874]. Rakaia [4128]. Raker [2049, 3663]. rakes [3070]. Randall [3575]. Random [2074, 4649, 4125, 4815]. randomized [5307]. Randomly [3210]. range [5099, 1075, 1135]. Rangeley [1485]. ranges [3763, 5316]. Rangia [2374]. RAPD [4605, 4815]. Rapid [906, 1949, 2968, 4910, 3218, 1984]. Rappahannock [239]. rare [2121, 5057]. Raritan [3375, 4679]. rastrefliger [4198]. rastrelliger [4219]. rat [488]. Rate [3032, 3644, 3632, 2643, 3613, 2035, 3587, 3344, 2744, 3163, 3741, 4843, 1903, 3864, 4553, 3111, 4507, 4545, 4482, 3335, 1726, 4235, 4660, 4101, 4768, 4259, 3406, 658, 573, 3138, 2889, 1800, 4593, 5329]. Rates [3234, 3590, 3343, 2605, 2240, 2878, 3210, 2898, 3709, 3187, 2045, 2496, 2909, 3266, 2011, 3339, 3514, 2115, 3484, 3881, 3289, 3531, 3639, 3688, 4403, 4052, 5039, 5065, 5255, 2807, 3926, 4250, 3831, 2997, 4878, 4020, 5016, 4580, 1981, 4781, 4780, 5196, 4863, 4827, 4671, 5326, 1899, 5129, 4079, 3168, 3118, 4443, 4988, 4008, 4226, 4870, 5373, 4685, 2272, 4171, 4133, 4136, 5272, 5143, 4487, 5269, 4474, 3920, 1803, 5069, 4375, 4822, 4377, 4861, 4266, 3351, 4385, 4651, 4746, 4225, 3958].

Ratfish [2120, 2325, 2846]. Rathbun [4414, 4630, 1994, 1999, 3311]. Ratio [2211, 3146, 3098, 1728, 4165, 4349, 5007, 4199, 1904, 3304, 3268, 3247, 5131, 5329, 4137]. Ration [3481, 4129, 4963, 3792, 4921]. Rationales [2501]. Rations [2177, 3081]. Ratios [3105, 2510, 2906, 3155, 2425, 5178, 3726, 4599]. raven [2461]. Ray [3328, 5334, 4431, 4588, 5054, 5164]. Rays [3413, 2964, 2266, 1537, 4691, 4685, 5300, 2985, 1474, 4534]. razor [1564, 1621]. Re [4538, 4840, 3880, 4673, 4507]. re-assessment [4673, 4507]. re-description [3880]. Re-estimation [4840]. Re-identification [4538]. Reach [3546, 1954]. reaches [4675]. reactance [4999]. Reaction [3186]. Reactions [5070, 4416, 1225, 4239, 3755]. reactive [1968]. read [5335]. Reader [3654]. Reading [3514, 4698]. realized [4846]. Reappearance [291, 1159]. Reappraisal [3681]. Reared [2032, 2619, 2459, 2626, 2646, 2709, 2969, 2078, 2110, 2168, 2972, 3581, 2981, 2200, 2753, 2716, 4243, 2622, 1027, 4414, 333, 5265, 138, 3170, 4819, 3516, 3367, 2162, 2690, 4447, 2184, 4630, 3976, 234, 4940, 2476, 3885, 4086, 4276]. Rearing [2293, 2034, 85, 1871, 236, 466, 2849, 1032, 79, 686, 3333, 8, 1083, 1396, 2691, 568]. reasonable [2581]. reasoning [863]. Reasons [277, 2809, 905, 1359]. rebuilding [5096, 4508, 4786, 5256]. Recalculated [4921]. Recapture [3508, 3271, 5031, 4311, 4580, 1903, 3140, 5326, 3804, 5261, 3106, 3905, 4377, 4793, 4861]. recaptured [4232, 3301, 4698, 3925, 4556]. receipt [481, 815]. receipts [798]. received [396, 187]. receives [1560]. Recently [3490, 4141]. recipes [844]. Reciprocal [2662]. recolonization [4674]. Recommendations [3496, 2255, 2257]. recommended [779, 813]. Reconnaissance [482, 429, 1003]. reconnoissance [953, 1464, 1099, 1080]. Reconstruction [4923, 4178]. reconstructions [4850]. Record [2143, 832, 2121, 5222, 3294, 2454, 2471, 2585, 4495, 3390]. Recorded [2313, 1075, 4862, 23]. recorders [5167]. recording [1391, 1384, 4986]. Records [2687, 2398, 2361, 2738, 2762, 2799, 110, 2680, 3723, 2897]. recovered [4813, 4663]. Recoveries [2629, 4805]. Recovery [5235, 4513, 5102, 5103, 4939, 4536, 4373, 4660, 4136, 3802, 4340]. Recreational [2892, 4052, 4617, 2859, 4543, 4832, 4472, 4992, 4693, 4483, 5109, 4800, 4593, 5323, 4187, 5232, 5180, 3545, 5308]. Recrudescence [3368]. Recruit [3113, 2281, 4705, 4120, 4935, 4544, 4418, 5140, 3799, 5147]. Recruitment [2912, 2625, 2374, 2605, 3467, 5351, 2857, 3427, 3455, 4626, 3588, 3812, 3208, 3978, 3002, 3439, 2259, 4296, 4649, 4488, 3751, 3098, 4514, 5002, 4081, 3454, 3892, 4541, 3811, 4787, 3781, 4012, 4255, 4790, 3335, 4151, 4382, 5154, 5209, 4098, 4217, 5041, 4171, 4725, 4609, 4568, 3991, 4058, 3903, 3752, 3863, 5242, 4262, 3946, 5146, 4206, 3873]. recruits [4747, 3827, 5124, 4611]. recurrent [2801]. recurring [5261]. Red [2048, 2773, 2036, 2038, 4618, 4671, 2696, 3698, 2510, 2030, 3607, 2200, 2055, 3510, 4281, 3868, 4948, 1662, 4798, 1864, 3501, 4103, 5271, 5113, 2519, 4522, 699, 799, 3898, 4808, 4871, 4462, 4350, 5155, 5279, 4998, 3785, 1524, 1590, 3866, 5176, 4075, 2549, 4127, 5223, 1852, 5367, 1574, 4668, 2930, 3000, 4370, 4914, 1861, 4899, 4256, 917, 4445, 5100, 1959, 1830, 4448, 2582, 4474, 4467, 5052, 5267, 4080, 2697, 4177, 4590, 4262, 5069, 5111, 3742, 4762, 4318, 5343, 5178, 4875, 5064, 3928, 4266, 4650, 4049, 4492, 1154, 4683, 3909, 4451, 4276, 1828]. red-salmon [1590]. red-snapper [799, 1154]. Red-Tide [2030, 1959]. Redds [2442, 2865]. Redescription [2116, 2066, 2881, 3095, 5144, 4648, 4881]. redescriptions [5075]. redfish [2230, 1827, 4554, 1595, 842, 4339, 1863]. reduce [4787, 4730, 4821, 4802]. reduced [5256]. Reducing [3411, 5131, 4126, 4090, 4251, 5284, 5308]. Reduction [4270, 4036, 5299, 5026, 4110, 5069, 4491, 4119]. Reductions [3572]. Reef [4799, 2872, 3466, 2509, 2823, 3981, 2148, 3113, 4628, 3407, 3875, 4025, 3547, 3471, 4499, 4164, 5081, 5119, 5142, 3803, 5015, 4135, 3819, 2825, 5099, 4908, 5129, 5176, 5223, 4393, 4394, 3906, 5175, 4646, 4913, 5020, 5321, 5300, 4890, 3918, 4448, 2511, 3303, 3843, 4894, 4402, 3747, 4865, 5171, 4594, 4307, 4348, 3864, 4736]. reef-associated [5176, 3843]. Reef-Fish [2823, 2825, 4393]. Reefs [2335, 3617, 3553, 3981, 4315, 4803, 4973, 3763, 3974, 4025, 4163, 5064]. reentry [3466]. Reevaluation [2350, 2447, 3345]. Reeves [211]. Reexamination [3412, 3879]. Referees [4140]. Reference [3677, 2246, 3373, 2267, 984, 4873, 1011, 1030, 4314, 1189, 1192, 1526, 1485, 1518, 1482, 1659, 4924, 4925, 1095, 1250, 1091, 4765, 1663, 1585]. References [2014, 302, 547]. referring [5332]. refine [4801]. Reflex [4947, 4902, 4951]. Reforms [1045]. Refrigerated [2050, 2628, 2770, 2741]. Refuge [3350, 4356]. refuging [3067]. refuse [154]. regalis [4576, 3488, 1470, 3416, 3748, 3896, 3987, 2446, 4101, 1659, 3114, 3162, 5203, 3642, 3671, 5130, 1459]. regard [392, 271]. regarding [92, 5161, 2406, 303]. Regeneration [2979, 378, 2480, 2794]. regime [4770, 4906, 4715]. Regimes [2407, 3181, 5306, 4954, 5008, 5146, 5248]. Region [2749, 1248, 999, 1304, 1171, 2975, 1191, 2663, 2665, 997, 1181, 1310, 5065, 3853, 2859, 3488, 1488, 1562, 4392, 1451, 1493, 3516, 4168, 941, 4010, 4785, 3987, 787, 1319, 4374, 1436, 4318, 3148, 3758, 4378, 2526, 2566, 2635, 2080, 3600, 3102, 2920]. Regional [5017, 2199, 3235, 4260, 5111, 5021, 5312, 5270, 5231]. Regions [2111, 3680, 2300, 1641, 4967, 4980, 1502, 941, 1425, 4511]. Regression [3183, 5109, 4078, 3482]. Regrowth [2215]. regularis [4570]. regulate [210, 372]. Regulation [2281, 2405, 2561]. Regulations [2501, 4878, 1347, 1343, 1351]. Regulatory [2243]. regurgitation [4026]. rehabilitate [4644]. Rehabilitation [2255, 2257]. reinforce [4296]. Reinhardtius [4500, 3091, 4843, 4816, 5359, 4912, 3481]. relate [5283]. Related [2016, 3584, 2793, 2644, 1809, 4210, 1690, 2922, 1518, 2000, 5350, 5321, 1844, 4058, 5362, 4003, 5008, 5146, 5248, 3726, 4767, 1888]. relating [102, 965, 299, 865]. Relation [2063, 2084, 1990, 2358, 2397, 2554, 1837, 2465, 3439, 3281, 3630, 2548, 2055, 3707, 2262, 2170, 4477, 5102, 5103, 79, 4827, 4325, 3948, 1061, 2029, 4859, 2565, 4217, 1897, 4235, 4954, 4383, 1113, 4236, 1466, 4334, 5127, 5141, 4374, 1766, 1982, 4076, 4432, 1145, 4715, 1017, 1412, 4128, 3351, 2002, 1499, 1522, 1530, 4380]. Relations [3409, 1988, 3634, 3462, 2690, 2321, 1306, 1122, 1476, 1490]. Relationship [3032, 4020, 3360, 3368, 2265, 3026, 2134, 2120, 4820, 3674, 3425, 2098, 2491, 2684, 3157, 5249, 4763, 3200, 3704, 4255, 2994, 5201, 3942, 3406]. Relationships [2657, 2503, 2027, 3051, 2183, 3702, 3650, 2942, 3283, 2335, 2632, 3701, 3002, 2836, 2925, 1882, 2135, 1339, 2713, 2419, 2982, 2778, 3141, 3939, 2894, 4933, 4944, 4790, 2340, 3957, 3857, 4803, 4055, 4867, 2599, 4924, 2888, 3479, 3034, 4923, 3891, 2604, 3990, 3897]. Relative [2654, 1806, 4688, 3315, 2056, 2316, 3613, 2090, 2355, 3310, 2718, 3522, 3070, 2575, 3805, 3618, 1621, 479, 5119, 162, 4959, 4884, 174, 129, 5367, 5327, 2593, 2538, 4175, 2701, 4957, 4572, 5316, 2587, 3083, 1971, 786, 4228, 5123]. Relatives [2389]. relaxation [5267]. Release [5113, 3461, 5228, 4748, 3133, 4627, 5377, 4082]. Released [2415, 3601, 4621, 4791, 5263, 4376, 3046]. Releases [3653, 4543]. releasing [2041]. relevance [4482]. reliability [4459]. Remains [2198, 4178, 4727]. Remarkable [565, 4641]. Remarks [2767, 3602, 67, 1793, 11]. remedies [1359]. remembrance [3532]. reminiscences [1116]. remote [4676]. remotely [5070, 5182, 5177]. Removal [131, 709, 1388, 4611, 289]. Removing [4824]. repeat [3224]. repeatedly [2286]. Repel [2072]. Replenishment [3587]. Replies [152, 408, 352]. Reply [2257]. Report [966, 840, 646, 569, 562, 872, 936, 942, 1016, 572, 162, 797, 502, 227, 785, 211, 987, 596, 1004, 21, 954, 965, 835, 849, 481, 670, 886, 385, 541, 659, 864, 958, 1170, 506, 478, 1262, 941, 973, 1087, 1086, 961, 897, 698, 586, 768, 803, 837, 935, 1007, 974, 991, 1072, 662, 931, 888, 224, 628, 1099, 307, 512, 607, 765, 867, 909, 989, 997, 1002, 1043, 250, 702, 207, 879, 962, 595, 334, 365, 370, 401, 977, 982, 1068, 400, 53, 123, 436, 422, 949, 995, 993, 1093, 1106, 1198, 1233, 457, 507, 946, 999, 1073, 521]. report [1192, 142, 1089, 963, 1650, 1158]. reported [2772, 4976]. reporting [3741, 4462, 4079]. reports [214, 329]. Representative [2917, 2114]. reproducing [937]. Reproduction [3765, 3097, 4579, 2374, 3593, 3609, 3361, 3552, 4312, 2990, 3056, 3352, 2010, 4473, 1638, 4455, 2271, 2840, 2919, 3604, 4609, 5236, 3607, 3649, 2473, 2601, 2739, 266, 3139, 3162, 2935, 4299, 3152, 3592, 2537, 3082, 1927, 3724, 4018, 2684, 3355, 3729, 3434, 5106, 4194, 4461, 5257, 3967, 5336, 3577, 680, 4287, 4370, 4197, 1435, 4974, 3780, 1936, 3960, 5047, 4879, 1135, 5054, 5164, 5369]. Reproductive [4783, 2950, 3234, 4158, 4707, 3438, 4389, 5214, 4165, 5093, 5197, 5280, 4060, 3312, 4670, 2185, 3900, 3886, 4148, 3321, 2487, 4285, 2613, 2696, 2808, 2549, 3373, 4130, 5292, 3832, 2510, 2446, 3859, 4333, 3274, 4741, 3610, 3045, 3705, 2590, 2322, 3074, 4661, 3665, 3328, 4949, 3767, 5357, 2365, 4575, 3909, 3551, 4422, 5083, 4603, 3727, 5271, 4057, 5281, 3837, 3372, 5110, 4109, 5376, 3809, 3465, 2876, 1236, 4482, 2746, 3212, 4064, 4151, 3801, 4528, 5327, 1472, 4234, 5221, 4714, 1866, 4093, 4095, 4115, 5274, 5302, 3375, 4204, 3302, 3718, 3852, 4300, 4733, 4323, 5123]. reproductively [4888]. reptiles [1563]. Republic [2267, 1565, 4743]. Republica [560]. repulsive [4938]. require [4551]. Requirements [2669, 3012, 2101, 1449, 2233]. requiring [289]. rerio [2426]. resampling [4261]. Research [2692, 3378, 3570, 2528, 3883, 2150, 5113, 2540, 5320, 4563, 1061, 3951, 3643, 1017, 881, 4307].

Researches [200]. Reservation [1524, 1572]. reserve [4784, 4041, 4719, 5161]. Reserves [2407, 2929, 3819, 4264, 5034, 4401, 4490]. Reservoir [2228, 2001, 2252, 1986, 1657, 2928, 2019, 3917, 2023, 2106, 2026]. Residence [2437, 3578, 5104, 3528, 4245, 2701]. Residency [5233, 5099, 5285, 5042]. resident [5142, 4293]. residual [2319]. residuals [4536]. Residue [2503]. Residues [2054, 2806, 2068, 710]. Resilience [3259]. Resistance [2469, 2589, 2786, 5375, 4999, 2412]. resistant [1777]. resistivities [1881]. Resolution [631, 4529, 5242, 5343]. resolving [4010, 4315]. resonance [4338]. resorption [4828, 1469]. Resource [3362, 3249, 2363, 4811]. Resources [3319, 4989, 3834, 1127, 4613, 1379, 3805]. respecting [921, 1072, 512, 250]. Respiration [2878, 2898, 2452, 2809, 1914, 4180]. Respiratory [2087]. Responding [2980]. Response [2566, 2350, 2932, 5349, 2323, 5306, 5267, 4971, 5248, 5217, 4918, 5323, 4134, 5232, 1833]. Responses [3932, 2615, 3455, 3232, 3651, 1960, 2828, 2943, 5050, 5174, 2242, 2689, 2700, 4971, 2087, 2817]. Responsiveness [3694]. resting [5274]. Restocking [604]. restoration [452, 1126]. restraining [2440, 5131]. Restricted [4919, 342, 2087, 5316, 1135, 1423, 310]. restricting [277]. Restriction [3605, 4392, 4785, 4766]. restrictive [397]. Restructuring [2675]. Result [165, 141]. resulting [4604]. Results [656, 912, 3902, 1079, 2294, 890, 3805, 3605, 90, 5189, 433, 149, 1195, 96, 4286, 159, 330, 897, 458, 3863, 71, 3922, 4179, 173, 1094, 1142, 23]. Resumed [3674]. resurrection [5075, 4648, 4881, 1612]. Resuscitation [538, 565]. retaken [403]. retardation [14, 89]. Retention [2125, 2139, 3673, 2831, 1879, 3731, 4858, 2152, 4209, 4114, 4131]. Retention/Survival [3673]. retia [4639]. reticulata [3017]. Retrospective [4106, 4164]. Return [238, 499, 2384, 4545, 5250]. Returned [3216]. returning [5192]. Returns [2240, 2819, 2238, 2909, 2453, 3168, 4295, 2581, 2205]. Reuben [3535, 3532]. revealed [5098, 5067, 4506, 4609, 4851, 1597, 4365]. reveals [4592]. reversal [4355]. Review [2433, 2102, 2328, 5061, 1600, 3038, 1606, 2123, 2088, 3012, 2241, 2077, 4837, 2212, 3968, 1437, 1056, 913, 3144, 167, 940, 1251, 1425, 4907, 1593, 1612, 1631, 1634, 2699, 1094, 1946, 192, 3421, 3990]. Reviewers [4188, 4437, 4519]. Revised [3952, 4192]. Revision [2017, 2008, 4038, 2769, 3364, 2757, 2758, 4352, 5363, 1301]. Revisiting [5312]. Rex [2545, 2665, 4783, 2543]. RFLP [3840, 4335]. Rhabdosargus [4675]. rhessodon [2723]. rhina [4913]. Rhincodon [4807, 5275, 5318]. Rhine [1405]. Rhinobatos [4083]. Rhinolithodes [3144]. Rhinoptera [5334, 4691, 3328]. Rhizocephalan [3283, 3254]. Rhizoprionodon [3935, 5304, 4048, 4546, 3045, 4298]. Rhode [3044, 3487, 2779, 2841, 795, 4972, 3291]. rhodorus [2981]. rhomboidalis [2645]. rhomboides [2010, 2464, 4155, 4512, 2817, 2828]. Rhomboplites [4060, 2808, 4287, 2577, 3843, 4137]. Rhombus [188]. rhotheus [2557]. rhynchoteuthion [4247]. rhythms [4915]. Ribbonfish [2606]. ribonucleic [2764]. ribosomal [4887]. Rica [5153, 1892, 2938, 3503]. Rican [1197, 1199]. rice [601]. richardi [2284]. richardii [4944]. richardsi [4629, 2822, 4638]. Richardson [2298, 4546, 3045, 1829]. richness [5119, 4874]. Rico [3394, 1202, 1196, 1214, 1211, 1207, 1200, 1204, 1195, 1206, 1217, 1208, 1213, 4336, 3742, 1205, 1201, 1203]. ridge [2270, 5223]. ridges [4880]. ridley [3858, 4132, 2867, 4555, 4678]. Right [3422, 3021, 3374, 4895, 595]. rigid [773]. rigida [1431]. Rikuzen [4741]. rikuzenius [4741]. Riley [4719, 5161]. Rim [4934]. Rimicola [2723]. Ring [3573, 3130, 3933, 4616]. ringens [4368, 4637, 2132, 3531]. Rings [3204, 2661, 4087, 4454, 4000, 3096, 2511]. Rio [4670, 1011, 4567, 4327, 4331]. Ripening [2075]. Rise [4322, 3556]. risk [4021, 3834, 3774]. Risso [2768]. ritteriana [2121]. River [479, 5207, 1662, 2998, 3011, 46, 2687, 2374, 44, 2608, 643, 3224, 741, 2105, 3064, 2234, 1097, 1150, 5258, 1731, 1665, 1073, 1144, 1590, 2784, 99, 653, 1891, 2755, 116, 2826, 1000, 1087, 1086, 2360, 3658, 2103, 2004, 4954, 5263, 567, 3357, 4629, 2197, 3323, 3487, 1804, 1893, 3149, 230, 233, 398, 791, 985, 3164, 424, 654, 154, 1994, 1999, 34, 2893, 4666, 4128, 2918, 1848, 289, 2747, 2831, 2095, 3990, 3037, 1998, 5375, 5167, 3089, 941, 973, 725, 3340, 974, 5191, 4570, 5273, 3740, 5134, 210, 800, 700, 843, 3480, 4560, 4559, 2219, 363, 748, 3037, 2928]. River [4090, 2983, 4129, 1685, 5102, 5103, 608, 388, 784, 1120, 2401, 2460, 333, 1601, 1602, 371, 1655, 2959, 517, 3546, 3440, 239, 477, 2325, 2009, 2819, 1106, 5351, 1899, 1931, 5009, 4325, 1078, 1521, 3917, 2623, 834, 1429, 165, 3370, 495, 344, 385, 541, 659, 864, 1945, 2131, 367, 74, 349, 503, 4443, 536, 5315, 816, 1861, 709, 1962, 3342, 2256, 3059, 4217, 4373, 4566, 1866, 1954, 2106, 91, 897, 3153, 3218, 3117, 3914, 403, 698, 586, 837, 1069, 3446, 3040, 2031, 2327, 2538, 4396, 122, 3784, 195, 4885, 191, 628, 933, 1596, 1634, 1660, 604, 515, 1645, 5273]. River [100, 1666, 2406, 141, 890, 1118, 1167, 4320, 3268, 3297, 33, 134, 1943, 1760, 4592, 1829, 775, 4759, 3341, 5277, 2415, 3046, 2287, 2627, 2716, 2918, 3836, 5062, 5134, 1585, 3855, 22, 334, 2006, 53]. river-dominated [3340]. Rivers [776, 4516, 2279, 482, 2719, 5348, 456, 1030, 241, 457, 3664, 1169, 184, 1797, 866, 1509, 5205, 84, 67, 401, 977, 982, 1068, 1834, 954, 3770]. RNA [5089]. Roanoke [5102, 2654, 5103, 403, 911]. Robert [2330]. Robinson [514]. Robson [4133]. robust [2936, 4078]. Robustness [3667]. robustus [2048, 2069, 2542, 3087, 3066, 2639]. Roccus [1664, 338, 8, 1647, 75]. Rocinela [2079]. Rock [3216, 3068, 3182, 3472, 2339, 2504, 2765, 3002, 3665, 2740, 4657, 4212, 5355, 4690, 4507, 5018, 4855, 5321, 4932, 4679, 2428, 911, 2695, 4407, 1931, 2106]. Rockfish [810, 2770, 3190, 3638, 2844, 2774, 2919, 2159, 3687, 3653, 3264, 2560, 3894, 3275, 3993, 4798, 4158, 5303, 4656, 2597, 3013, 3727, 4910, 4964, 5280, 5144, 4392, 4092, 4553, 4582, 5059, 5288, 4692, 4732, 3462, 5262, 5056, 4694, 3794, 4016, 4233, 3692, 3927, 4019, 4198, 4608, 4659, 4993, 5175, 4956, 3288, 3174, 4171, 4219, 4786, 3108, 3877, 5127, 708, 3960, 3814, 4881, 3772, 1647, 3822, 4003, 4735, 4550, 3757, 5034, 4872, 3838, 4931, 4401, 3889, 4238, 4744, 4680, 5163, 5361, 4986, 4270, 3393]. Rockfishes [3361, 2788, 2908, 3591, 2706, 3214, 3688, 3158, 3360, 5283, 4820, 4785, 4766, 4982, 4208, 4648, 3984, 4337, 5055, 4490, 5264, 4535, 4051, 4356]. rockling [2722, 4213]. rocklings [2722]. rockpools [4674]. Rocks [3474]. Rocky [2504, 3597, 4913, 3763, 4009, 1954]. rodent [460]. rodents [491]. Rodrigues [4258]. roe [2483, 691, 353]. Role [2267, 3429, 2708, 2154, 2302, 3120, 2664, 3954, 3244, 4401, 4305, 4306, 5308]. roles [2701]. rookeries [3994, 3989]. Roosen [809]. roosterfish [5201]. rosaceus [2613, 4812]. rosefish [1669]. rosenbergii [2701]. Rosenstihl [352]. rosmarus [4304, 3084]. rostrata [2688, 2533, 122, 4000, 4488, 3291]. rostratus [3330]. Rosylip [2981]. Rotaria [1189]. rotational [4544]. Rotatoria [1244]. roudi [4218]. rougheye [5288, 4732, 4233, 4881, 4401]. roughleyi [3453]. roughy [3876, 4062, 3791, 4125, 3958, 4119]. Rouleina [2616]. Round [3352, 3713, 2546, 5370]. Rounsefell [2257]. routes [3763]. routine [4780]. ROVs [4652]. Royal [2048, 829]. rRNA [4392]. ruberrimus [3814]. rubescens [2248]. rubra [4167, 4877, 3970]. rudder [1868]. Rudolphi [3300]. rufescens [3510]. rule [5087, 4175]. rules [4705, 5330]. Run [2091, 2031, 4069, 387, 4566, 2550, 5192, 3268, 3297]. runcatus [4187]. Runs [2401, 2793, 2255, 2257, 1590, 1797, 1660, 2918]. Russia [4893]. Russian [1105]. Rustless [574]. RV [2367, 2224].

S [125, 945, 1025, 4978, 138, 961, 768, 2145, 2107, 628, 1289, 979, 676, 939, 963, 1091, 187, 3105]. S. [5303, 3437, 5280, 4687, 5075, 5144, 5288, 4732, 2922, 4367, 4417, 4233, 2908, 5175, 4219, 5029, 4339, 4401, 4490, 4270]. S.C. [457, 365]. s.l [3099]. s.l. [3107]. Saanich [3460, 3877]. Sabellidae [2116]. sabina [2199]. Sable [2704]. Sablefish [3331, 3635, 3445, 3509, 3195, 5058, 3276, 4232, 4417, 3872, 3804, 4169, 4088, 4898, 4631, 5160, 4263, 5211, 5051, 3061, 2517]. sabulosa [2785]. Sac [2863, 2372, 2809, 4028, 1875]. Saccopharyngiformes [2311]. sachi [4115]. Sacramento [3086, 4373, 4566, 933, 1509, 1340, 398, 3351]. Safflower [2344]. saffron [5244]. Sagamichthys [3670]. sagax [3521, 2497, 3414, 5036, 4673, 3145, 5285, 5027, 4444, 4974, 3796, 3531, 4386, 5040]. Saginaw [1821, 1772]. sagitta [3898]. sagittae [5265, 4816, 4405, 3581, 2435, 2896]. Sahara [2524]. said [700]. saida [2332]. sail [305]. sail-fish [305]. Sailfish [2021, 2862, 4642, 4935, 1778, 4738, 3301, 4996]. Saint [576, 608, 741, 195, 573]. Saint-Front [576]. Saipan [5184]. saira [2269, 2282, 5177, 3450]. sairn [3515]. salar [5050, 5174, 3424, 575, 4254, 4763, 5263, 4789]. Sälbling [56]. sale [224]. Salinities [2033, 2609]. Salinity [1554, 2544, 3650, 2761, 2350, 2372, 2353, 2194, 2457, 2204, 2118, 1551, 4100, 1907, 2588, 2930, 1648, 5198, 4430, 5203, 1317]. Salish [5067]. Salmo [5050, 5174, 3412, 3424, 2819, 575, 4254, 4763, 85, 2826, 504, 5263, 2385, 3386, 4789, 2357, 2617, 152, 2250, 505, 2786]. Salmon [800, 690, 2063, 2417, 2516, 2839, 843, 2096, 3269, 3011, 3251, 2219, 2520, 2119, 3014, 784, 5101, 2660, 2506, 2404, 2105, 715, 924, 2019, 3424, 2009, 2084, 2234, 2819, 1150, 802, 507, 3578, 3632, 2246, 2304, 99, 834, 3370, 2755, 2677, 2131, 2629, 2252, 2952, 2134, 2136, 2097, 3588, 2023, 2238, 2256, 3461, 2273, 2430, 2507, 2499, 2438, 2442, 2091, 2793, 2031, 2111, 3436, 3181, 3143, 3442, 3528, 2197, 3015, 3550, 3398, 2852, 1578, 1583, 1605, 1627, 2262, 2047, 3149, 2753, 2026, 1118, 3271, 2384, 1607, 775, 3341, 3512, 2255, 2257, 2415, 2287, 2627, 2716, 2918, 2786, 2933, 274, 163, 176]. salmon [350, 479, 408, 718, 646, 569, 361, 362, 1662, 3480, 4054, 4310, 4560, 4559, 4893, 4934, 5049, 62, 48, 1864, 776, 5348, 4989, 3009, 451, 3634, 1998, 1685, 5050, 5174, 4919, 2474, 1120, 1895, 643, 1757, 2412, 339, 1643, 1636, 1641, 686, 858, 5167, 2242, 1758, 3884, 1093, 1106, 3133, 4827, 241, 4977, 575, 446, 448, 3408, 3945, 4069, 4689, 1899, 1931, 4046, 4254, 4763, 2483, 3719, 4442, 1458, 1521, 1524, 1590, 1572, 3917, 5290, 4774, 111, 849, 1315, 1429, 1452, 1471, 1469, 1465, 4273, 387, 1763, 2463, 2594, 1852, 2810, 650, 1405, 349, 4274, 1966, 1847, 2826]. salmon [603, 5233, 1861, 1962, 4153, 4373, 4566, 459, 4484, 419, 1866, 1954, 2004, 4954, 897, 5263, 3088, 1007, 1069, 3997, 3446, 1901, 1936, 1985, 114, 4334, 1158, 1220, 1997, 1830, 4396, 4398, 2079, 5143, 3784, 4033, 2581, 397, 2409, 2557, 1935, 1881, 4035, 266, 683, 4789, 933, 4950, 2865, 1509, 1561, 1596, 1593, 1612, 1631, 1634, 1660, 1302, 515, 1645, 1804, 1976, 1722, 2425, 4715, 1249, 107, 808, 5366, 3034, 1681, 1666, 2406, 4203, 398, 791, 1076, 33, 328, 424, 900, 1107, 1943, 2648, 4592, 3233, 1986, 23, 4128, 3770, 642, 4043, 3046, 582, 84, 750, 1404, 289, 4745]. salmon [3990, 11, 67, 2006, 5169, 682, 643, 2826]. salmon-breeding [163, 176]. salmon-fisheries [397]. salmon-fry [62]. salmon-packing [808]. Salmon-producing [2438]. Salmon-tagging [1578, 1583, 1605, 1627, 1607]. Salmonid [5293, 3180, 4624, 1580, 5261, 4638]. Salmonidae [1396, 735, 1373, 1788, 1805, 1882, 938, 147]. Salmonids [3546, 3621, 4229, 1952, 3658, 3117, 4629, 5192, 5062, 4599]. salmonis [1570, 1580]. salmonoid [1032, 1419, 1406, 1455, 1425, 1505]. Salmonoids [1387, 752, 1401, 1104]. salps [3003]. Salt [3156, 355, 905, 4426, 714, 720, 1031]. salt-box [4426]. salt-water [355, 714, 720]. saltatrix [4557, 4249, 3629, 4129, 4282, 4283, 4760, 3895, 3775, 2715, 4429, 3782, 3812, 3439, 4844, 4805, 4915, 5203, 4809, 3828, 4727, 4921]. Salting [691]. salts [2480]. Salvador [1565]. Salvelinus [5372, 152, 56]. Salvo [56]. same [3172, 1571, 1072]. Samoa [4135, 1320, 5295, 1987, 2422]. sample [569, 1864, 4826, 5216, 4890]. sampled [2853, 5170]. Sampler [2202, 1798, 5066, 4444, 3272, 2515]. Samplers [2172]. Samples [3467, 4501, 4042, 4526, 4510, 2936, 4466, 4813, 4842]. Sampling [2056, 3570, 2554, 2164, 2514, 3264, 2624, 4186, 2672, 4499, 4117, 3122, 5073, 5307, 3803, 3542, 3902, 2591, 4582, 5056, 4184, 3807, 2565, 3279, 3658, 4445, 5109, 2795, 2071, 5063, 2835, 4597, 3031, 5260, 5178]. San [3086, 1340, 3078, 3127, 3122, 4057, 364, 4141, 5068, 1088, 3666, 3125, 3619, 3668, 4484, 3310, 4696, 4201, 3191, 3587, 3519]. sanctuaries [4660]. Sanctuary [3918]. Sand [3388, 2643, 3056, 2718, 3594, 3281, 2935, 3568, 4916, 5187, 3890, 4030, 4917, 5128, 2711, 2274, 2224, 4268, 4880]. Sandbar [2921, 3211, 5031, 4328, 4938, 4244, 4781, 4780, 5377, 5272, 4013, 1850, 3792]. Sanddab [2665, 2700, 3028, 3031]. sanddabs [3964]. sandfish [2884, 4814]. Sandfishes [3610]. sandperch [4702]. Sandworm [3023]. sandy [4436, 5234, 556]. Sanibel [1941, 1981]. Sanitary [521, 1592]. Santa [4052, 2423, 3126, 2509, 2823, 2825, 2490, 3050, 2638, 3617, 4316, 2582, 2572, 2713]. Santo [4595]. São [2655]. sapidissima [4516, 2401, 2460, 2437, 5122, 1817, 3060, 2724, 5030, 89, 3459]. sapidus [3016, 4565, 5327, 2795, 5008, 4319, 4679, 5000, 1994, 1999, 5234, 4342, 3855, 3350]. Saprolegnia [1037]. Sarasota [2944]. sarba [4675]. Sarcoma [3318]. Sarda [2389]. Sardillops [3414]. Sardina [4081, 4828, 2356, 2662]. Sardine [2356, 3521, 2337, 2574, 2168, 3531, 1759, 1836, 4081, 3414, 5036, 903, 4828, 4673, 3145, 1835, 1977, 5285, 5027, 598, 2162, 4444, 4974, 1786, 1837, 1897, 3796, 1237, 3808, 4207, 1761, 4386, 5040]. Sardinella [3808]. sardines [178]. Sardini [2389]. Sardinops [2497, 5036, 4673, 3145, 1835, 1977, 3521, 5285, 5027, 2139, 2152, 4444, 4974, 1786, 1837, 1897, 1668, 1766, 3796, 3531, 4207, 4386, 4380, 5040]. Sargasso [2661]. Sargassum [4903, 5248, 4667]. SARIMA [5011]. Saronikos [4008, 5238]. Sars [3023]. Sashin [1985, 2865, 2506, 2677]. Satellite [3537, 3676, 5305, 4472, 4945, 4691, 4693, 5200, 4621, 4684, 4676, 4222]. satellite-linked [5200]. satellites [4710]. saturation [5253, 5286]. saturnum [2924]. saugers [335]. Saurida [2298]. Sauries [2757, 2758]. saurus [2179, 1825, 4430]. Saury [2269, 3450, 3515, 2282, 5177]. Savannah [3064]. Saving [396, 464]. Savoy [190]. sawfish [5297, 5298, 5188]. saxatilis [4856, 5044, 5296, 2656, 2654, 4516, 2781, 5102, 5103, 3336, 2965, 3440, 2988, 3664, 4945, 3327, 4443, 3367, 2495, 2493, 3342, 4566, 4941, 1664, 4886, 1647, 3946, 3778, 3925, 4822, 5203, 4181, 4574, 3332]. saxicola [4019]. Saxony [304]. Say [1628]. Scad [3352]. Scale [2660, 3632, 2168, 2111, 2137, 4311, 4541, 4689, 4079, 2162, 2554, 2565, 5297, 3215, 3896, 5298, 4872, 3925, 3671, 5220, 1597, 4840]. scale-based [3925]. Scaled [2337, 2574, 4405]. Scales [2346, 2174, 4594, 5044, 4182, 1789, 1725, 1456, 1222, 4944, 1867, 1857, 1842, 4215, 1962, 3815, 1672, 1997, 378, 5335, 1475, 5130, 1597]. Scallop [2466, 3225, 2929, 2059, 2558, 4747, 3458, 1608, 4544, 4381, 4494, 1940, 4677, 3790, 969, 3490, 2695]. scalloped [3434]. scallops [4979, 5172]. Scanning [3159]. Scarcity [183, 172, 328, 807, 383, 494, 67]. Scaridae [1186]. Scarus [2634]. scat [4638]. scats [4663]. scattering [4831, 1116]. Scavenger [3332, 1522]. scavengers [4494]. scavenging [4358, 4684]. scenarios [4627]. Schaefer [2039]. Schedules [2238, 1899]. Schizamoeba [1580]. Schleswig [959]. Schleswig-Holstein [959]. schmidti [4493, 3036]. schmitti [5219]. schomburgkii [3929]. Schoodic [176, 646, 718]. School [3549, 3596, 3648, 2673, 2649, 3705, 2644, 4190, 4937, 3929, 4147, 3844, 4428, 4259, 3861, 4225]. Schooling [2099, 2730, 3196, 2556, 2920]. schoolingbehavior [5127]. Schools [2913, 2980, 3689, 3477, 2481, 3376, 2885, 285, 2160, 5212, 2547, 4428, 2742, 3865]. schooner [912, 961]. Sciaenaps [3898, 3866, 3909]. sciaenid [2505, 5011, 3503, 2608]. Sciaenidae [3646, 3185, 2924, 1638, 3429, 4291, 3135, 2372, 2353, 2446, 2848, 2803, 4112, 3753, 2989, 1527]. Sciaenids [3420, 5106, 3878, 1595, 5139]. Sciaenops [4948, 3501, 4103, 3731, 2519, 4462, 3785, 5367, 2930, 4914, 3698, 3607, 4474, 5052, 4262, 4318, 3928, 4049, 4276]. Science [2331]. Scientific [2236, 1715, 1151, 1195, 159, 1061, 5014, 71]. scientists [2142]. scitulus [2620]. Scomber [2349, 2618, 2859, 3715, 4012, 2790, 2532, 2796, 2140, 3796, 5309, 2697, 4202, 1661, 1678, 4636, 5038]. Scomberesocidae [2757, 2758]. Scomberjaponicus [3930]. Scomberomorus [4190, 3793, 4884, 2637, 3177, 627, 3437, 3486, 4138, 1675, 3399, 3321, 4366, 4506, 3042, 3854, 4714, 4104, 4107, 3826, 4716, 3592, 3699, 3098, 4782]. Scombrid [3051, 2383]. Scombridae [4908, 2196, 3616, 2389, 3177, 4038, 4447, 2308, 3565, 2073]. scombrids [5004, 2321]. Scombrinæ [913]. Scombroid [2223, 3293]. Scombroidei [3939]. scombrus [2349, 2618, 2859, 3715, 4012, 2796, 5309, 1661]. Scopelengys [2433]. Scopelopsis [2154]. Scophthalmus [3606, 3980]. Scorpaena [5325, 3334]. Scorpaenichthys [3498]. Scorpaenidae [3727, 4687, 3190, 4692, 2117, 4019, 4198, 2788, 2908, 2844, 3174, 2159, 2345, 5029, 3544, 2706, 3988, 4339]. Scorpaenidaea [2057]. Scorpaeniformes [5029, 4771, 3817, 5144, 4648, 4881]. Scorpionfish [3334, 5325]. Scotch [660, 444, 522, 523, 295, 435, 144]. Scotia [3068, 905, 2138, 2355, 1669, 3290, 3740]. Scotian [3062, 4237]. Scotland [38, 371, 858, 715, 1057, 262, 68, 419, 809, 1260]. Scotopic [2901]. screen [1594]. scrombrus [1678]. sculentus [4246]. Sculpin [3366, 3219, 2981, 3963, 4928, 4369, 4494, 2799, 2409, 2557, 4771, 3110]. Sculpins [3008, 2438, 2888, 3746]. sculptipes [2915]. Scup [3128, 4686, 4226]. Scyliorhinid [3470]. Scyliorhinidae [4324, 1946]. Scyliorhinus [5037, 3236]. Scyllaridae [5123]. Scyllarides [4688]. Sea [2466, 2977, 3127, 4084, 3216, 2433, 3697, 2631, 2380, 2189, 5099, 1348, 2696, 1991, 2193, 2267, 3225, 2192, 3407, 3619, 3668, 3337, 198, 3359, 2733, 1015, 2174, 4890, 2833, 2200, 2927, 3676, 907, 2155, 2033, 2609, 3511, 2731, 3975, 4329, 1737, 5322, 5039, 5065, 5057, 4410, 4601, 5189, 4686, 4706, 3858, 4708, 4020, 4522, 4132, 4602, 1041, 1728, 5320, 5222, 3299, 1400, 4404, 4710, 4618, 3959, 539, 1428, 1515, 3408, 4730, 2461, 4625, 5151, 3458, 4712, 2547, 3276, 2549, 4544, 4930, 4464, 764, 5105, 3404, 1638, 2791, 4668, 440, 5114, 2645, 519, 4381, 4979, 1130, 4776, 2291, 5353, 4653]. sea [1770, 4586, 4713, 5236, 3989, 1940, 4677, 492, 4570, 4246, 5239, 669, 4697, 2582, 1659, 382, 5230, 1854, 4825, 3752, 3863, 4555, 688, 2991, 5024, 2929, 4590, 103, 763, 5366, 5358, 3799, 3165, 1076, 4678, 2911, 4615, 3268, 3297, 1423, 1109, 3845, 3238, 4765, 4665, 4838, 4887, 5172, 4635, 4636, 5251, 4745, 3308, 4322, 1824, 2382, 960, 4551, 4228, 4515, 4539, 4067, 5308, 4664, 108, 434, 3999, 3377, 2726, 5053, 2875, 5067, 4720, 5028, 5093, 4896, 5010, 5355, 4908, 3253, 5338, 1442, 3109, 3635, 3704, 2316, 3373, 4369, 4770, 4892, 2104, 4004, 4372, 5124, 3561, 2799]. Sea [4740, 4118, 4176, 4397, 4932, 5141, 3087, 3084, 4352, 2661, 4771, 1427, 2332, 1012, 3905, 4716, 3867, 5359, 3302, 4015, 4853, 3716, 5268, 3033, 5318, 2428, 3197, 3287, 3481, 3955, 392]. sea-caught [3276, 3238]. sea-fish [1400, 519]. sea-fisheries [382]. sea-fishing [492]. sea-herring [539]. sea-horse [103]. sea-run [3268, 3297]. ‘Sea-serpent’ [392]. sea-serpents [688]. Sea-stars [1991]. Sea-Surface [2174, 2033, 2609, 3676, 5105, 2382]. sea-surface-temperature [4710]. sea-water [440]. Seabass [3533, 4883, 5118, 3759, 4141]. Seabed [5071]. Seabird [3689, 4152]. Seabirds [3269, 2933, 4797, 4870]. seabream [4707, 4675]. seafloor [4970, 5069, 5347, 4897]. seafood [4052]. Seagrass [3618, 3696, 4747, 3279, 4227, 3070, 3120, 4206, 4759]. seahorse [4846]. seahorses [4600]. Seal [3167, 3033, 2284, 2926, 1909, 2708, 4629, 2822, 2856, 4638, 4813, 1890, 230, 882, 1040, 1105, 1566, 2882, 121, 1371, 3353]. Seals [3273, 3065, 3492, 3316, 3284, 4848, 5067, 3660, 4944, 4792, 901, 5068, 3363, 3994, 5168, 3820, 1480, 3161, 3245, 3867, 2889, 4239, 250, 1371]. SEAMAP [5260]. Seamount [4595, 3912, 4245, 3827, 3822, 3764]. seamounts [3454, 3911, 3571]. seaperch [4812]. Search [2482, 556, 5097, 595]. Searching [3169, 3425]. Searobin [2620, 4154, 4465]. Searobins [2456, 4517]. Seas [2246, 1935, 2933, 3831, 1895, 1347, 2653, 2430, 1343, 1351, 1670, 4422]. Season [2352, 2445, 3199, 3656, 269, 3439, 5012, 4783, 785, 135, 3172, 4159, 4998, 4764, 835, 834, 670, 886, 4368, 2994, 4628, 4082, 3994, 5306, 586, 768, 837, 4610, 628, 5267, 4488, 5160, 547, 607, 765, 867, 909, 900, 4300, 865, 180, 592]. Season- [5012]. Seasonal [4883, 5118, 3007, 4441, 4100, 1955, 3517, 3606, 5107, 4111, 1562, 4689, 5258, 3948, 2505, 2874, 2924, 3028, 3368, 3190, 4975, 2860, 3058, 3801, 3985, 5262, 5331, 2593, 3498, 5175, 3668, 4567, 3131, 2978, 4510, 3119, 4548, 2305, 5141, 4374, 2727, 2450, 2603, 2929, 4037, 4715, 3905, 4810, 3589, 3740, 2250, 1971, 5062, 5134, 1843, 1992, 3495, 5231, 4500, 4123, 4478, 3902, 5036, 3880, 5099, 4944, 5017, 4464, 4737, 4085, 4637, 5022, 3835, 4695, 4917, 4074, 4965, 2278, 4474, 2846, 3938, 4853, 4915, 5130, 4422]. Seasonality [3315, 3473, 3361, 3672, 2868, 3323, 3718, 4109, 4060, 3809, 5007, 4010, 5227, 4061, 4266, 4575]. seasonally [4943]. Seasons [2011, 5170, 802]. Seatrout [3438, 2986, 2619, 5074, 2935, 4916, 4917, 4359, 4598, 4632, 4661, 4836]. Seatrouts [2488]. seaturtles [4471]. Seaward [3436, 3133, 1509]. Seawater [2628, 2770, 2054, 2097, 2741, 2292]. seaweed [1312]. seaweeds [1314]. Sebago [1404]. Sebastes [3993, 4798, 4158, 5303, 4656, 2597, 2945, 3013, 3540, 4936, 3727, 4910, 3158, 2230, 4561, 4964, 2770, 5280, 4831, 3142, 4787, 5144, 4392, 3190, 2566, 4553, 5288, 4692, 4732, 3462, 4415, 5283, 5262, 1827, 4694, 4233, 3692, 3927, 4019, 4198, 4608, 4659, 4820, 4993, 5138, 2788, 2908, 5175, 3288, 4785, 4766, 2774, 2919, 3174, 4219, 4786, 4234, 4554, 3108, 4982, 2159, 3877, 5029, 5127, 3960, 3814, 4648, 4881, 3772, 3687, 4003, 1669, 4735, 3653, 3395, 4005, 4550, 3757, 2706, 4337, 5034, 4872, 842, 3988, 3838, 4339, 4931, 3214, 4401, 3889, 4238, 4744, 4490, 4680, 5163, 5264, 5361, 3275]. Sebastes [3688, 4270, 3393]. Sebastidae [4535, 5144, 4659]. Sebastolobus [4687, 4480, 4978, 4417, 2345]. Second [2191, 1572, 206, 2454, 2150, 3390, 2121, 2585, 3986]. secondary [4090, 5198, 4334]. Secretary [1016]. sectatrix [1868]. Section [1545, 1546, 1049, 17]. sectioning [3739]. sector [3562]. secure [451, 21]. securing [875]. Sediment [2313, 1938, 3200]. Sediments [2434, 2663, 2665, 2198, 5125, 4773]. sedis [2361]. seed [1609]. Seeding [2839]. Seeking [3429]. seen [302]. Seepage [3432]. segmentation [107]. Segregation [652, 5212]. Sehastes [2057]. sei [3776]. Seine [2730, 2164, 2093, 3425, 4640, 3803, 2838, 4761, 4524, 4821, 2971, 232, 4468, 3805, 3952]. Seines [3543, 342, 277, 87, 425, 310, 453]. seining [692, 421, 293]. Selected [2294, 3630, 2101, 3657, 3628, 2689, 2621, 3055, 4766, 4208]. Selection [2136, 3156, 2727, 4952, 3276, 4256, 3818, 4932, 1935, 3710, 4531, 3867, 3539, 5134, 414]. Selective [2387, 4622]. selectivities [4417]. Selectivity [2172, 2779, 2937, 2885, 2418, 2049, 2534, 5048, 4961, 3800, 4552, 4989, 4251, 4794, 4871, 5013, 4429, 5246, 3838, 5149, 3345, 5179]. selenium [2703]. Self [937, 73, 4604]. self-picker [73]. Self-reproducing [937]. self-sustaining [4604]. semifasciata [3587, 4556, 4702]. Semilunar [3212]. semisulcatus [4246, 4272]. Senate [271]. Sending [630]. Sennet [2110]. sense [1241, 1426, 1448, 1467]. sensed [5177]. senses [1246]. sensilla [2261]. sensing [4676]. Sensitivity [3163, 3001, 4832, 5140, 2641, 1777]. Sensory [2099, 2770, 1635, 3533, 2741]. sent [731, 140, 190, 566, 436, 455]. Separate [2402]. separated [1139]. Separation [3594, 4249, 4426, 4761, 4077, 3780]. Sepia [4285]. sepioid [3801]. Sepioteuthis [3573, 4267]. September [2763, 211, 1170, 2763, 879, 738, 781]. sequences [4731, 5341, 5270, 4433, 4318]. sequencing [5116]. Sergestes [2112, 2972, 2725]. sergestid [2522]. Sergestidae [2972, 2522]. Serial [2917]. Series [3540, 4906, 4547, 4824]. Seriola [5312, 4286, 4196, 5371, 4241, 4667]. Seriphus [3646, 2917, 3185, 2505, 3022]. Serological [1863]. serologically [1951]. serpent’ [392]. serpents [688]. Serraniculus [2193]. serranid [3675, 5190, 4402]. Serranidae [3675, 940, 3372, 4165, 4764, 5151, 2193, 5032]. serranids [4521]. serranoides [2774, 2919, 3174]. Serranus [4764, 960]. Serum [3659, 1951, 2732]. service [264, 464, 396, 2150]. serving [248]. sessile [3723]. session [210]. sessional [1369]. setiferus [2772, 2181, 1771, 5198, 1653, 4651, 1650, 1815]. Seto [4716]. sets [4147, 4761, 4050, 4555]. Sette [2182, 2171, 2166]. setting [1626, 1637, 1656, 1598, 4891]. Settled [2783, 4141, 5137]. Settlement [4321, 4206, 3759, 4498, 4499, 2997, 4465, 4296]. settlement-stage [4499]. Settling [2126]. seup [230]. Seven [2355, 3137, 2710, 2690, 4316, 4055]. Seventeenth [441]. Several [2622, 2114, 2357, 544, 448, 3165, 5244]. Severe [4492]. sewage [3834]. Sex [4705, 2684, 3185, 2120, 5238, 2060, 4628, 2211, 2510, 3146, 3098, 4622, 4182, 4935, 4165, 4920, 4349, 5007, 4871, 5199, 4978, 1630, 4899, 2906, 5035, 4199, 1904, 5302, 4176, 3304, 4003, 3155, 613, 5178, 3268, 4355, 3247, 5329, 4137]. Sex-change [4705]. sex-changing [4622]. Sex-Specific [3185, 4628, 4935]. sexes [652]. sexfilis [4543, 2698]. Sexual [3062, 4349, 1896, 3576, 2999, 5275, 4739, 1818, 1877, 3705, 3146, 3659, 3314, 2303, 3091, 3837, 4252, 4688, 88, 4008, 2879, 4095, 5084, 4261, 197, 3075, 4083, 3833, 4137]. sexuality [175]. Shad [2401, 2437, 205, 741, 239, 446, 292, 536, 3060, 2173, 566, 2724, 195, 3459, 171, 22, 203, 700, 125, 4516, 363, 748, 1937, 608, 1725, 782, 854, 894, 388, 2460, 456, 5122, 457, 1834, 1731, 145, 835, 165, 482, 670, 886, 70, 344, 385, 541, 659, 864, 1817, 478, 816, 1857, 97, 324, 679, 687, 586, 833, 803, 837, 935, 5191, 1844, 110, 37, 1797, 3357, 39, 707, 5030, 628, 14, 89, 307, 230, 141, 233, 547, 607, 765, 867, 890, 909, 1135, 1760, 1812, 319, 1848, 365, 429, 370, 401, 180, 998, 1380, 53, 5108]. shad-fisheries [365, 429]. Shad-fishing [536]. shad-hatchery [1380]. shad-hatching [835, 670, 607, 319]. shade [1482]. shades [1492]. Shallow [2860, 3666, 3164, 3684, 3519, 2747, 3122, 3763, 5210]. Shallow-water [2860]. shank [4693]. Shape [2652, 4338, 5117, 4267, 5247]. shaped [5179]. Shapes [2644]. Shark [3689, 3358, 3648, 3056, 2086, 2144, 3045, 3705, 2739, 2572, 2235, 3587, 2713, 5031, 5228, 4552, 5005, 2804, 4328, 4562, 4709, 5266, 3935, 4748, 4312, 5278, 5225, 4473, 4214, 5114, 4528, 4684, 3559, 5331, 3019, 3363, 2157, 4170, 4546, 3683, 5328, 5275, 4538, 4588, 602, 3936, 4295, 4486, 4985, 4633, 2227, 4298, 4591, 4299, 4013, 4421, 4556, 1850, 2444, 3792, 5318, 5357, 4639, 5180, 3995, 4224, 4225, 4248]. shark-fisheries [602]. Sharks [1537, 1716, 2066, 3470, 2921, 3211, 3024, 4242, 2873, 5228, 5048, 4938, 4191, 4807, 4244, 4617, 4779, 5194, 2124, 4781, 4780, 5237, 5284, 5304, 5091, 4358, 5340, 4729, 4183, 4774, 4731, 4048, 4214, 5360, 5377, 3957, 4537, 4740, 5276, 5023, 5272, 2955, 1474, 4377, 464, 585, 1946, 1808, 5220, 5323, 4322, 5232]. Sharpnose [3045, 3935, 5304, 4048, 4546, 4298]. Shedding [2807, 3599, 2496, 4403, 4079, 2272, 4225]. Sheepscot [3023, 2125, 1891, 2893]. Sheepshead [3644, 3633, 3548, 4086, 5204]. sheepswool [5362]. Shelf [2618, 3735, 3809, 3204, 1993, 3682, 2518, 4498, 5053, 3150, 3402, 3158, 4282, 4478, 3160, 4756, 3937, 4284, 4619, 4505, 4750, 4006, 4066, 4736, 5125, 3835, 4695, 4511, 5314, 4344, 5156, 5126, 4810, 3747, 4321, 4809, 4773, 5224, 3894, 4535, 5090, 4309, 3062, 2138, 3420, 2314, 3263, 2508, 2978, 4834, 4237, 2224, 2237, 2013]. Shelikof [3871, 3851, 3380, 3579, 3893, 3586, 3722]. Shell [3449, 3318, 3101, 1963, 2322, 3191, 1916, 4243, 2625, 2544, 477, 1431, 155, 1622, 346, 1940, 645, 1905]. shell-fish [645]. Shellfish [2513, 2517, 2579, 2987, 2747, 1796, 4315, 1449, 2699, 2500]. shells [339, 390, 3106, 3226]. shelter [2701]. shelters [3773]. Shepherd [3625]. Shetland [729]. shift [4770, 3782, 5289, 4715, 3824, 4134]. Shifts [5080, 5312, 5338, 4906, 4469, 4650]. Shiner [2407]. Ship [2980, 3923, 4390, 1587, 3444]. Shipboard [3537, 3100]. Shipment [730, 62, 48, 1193]. shipping [5039]. shipworms [1124, 1342]. shipwreck [5156]. shirred [425, 453]. Shoals [97, 3392, 4645]. Shock [2342, 2900, 3104]. Shore [2490, 3066, 470, 3347, 4111, 1489, 1242, 5234, 5323, 5171, 293]. shore-based [5323, 5171]. shoreface [4880]. Shoreward [3478]. Short [2589, 2852, 4131, 5050, 5174, 1876, 4426, 2694, 968, 4267]. Short- [2852]. short-lived [4267]. Short-Term [2589, 4131, 5050, 5174, 4426]. Shortbelly [3687, 3692]. Shortfin [5076, 4309, 5133, 4333, 5302, 4923, 4921]. shortnose [5207, 5315]. shortraker [4732, 4233, 4401]. Shortspine [4480, 4687]. should [342, 310, 293]. shovelnose [4083]. showing [313]. shown [2071]. Shrimp [2167, 3228, 2014, 2512, 2048, 3708, 2202, 2843, 2628, 3697, 2046, 2032, 3312, 3182, 2934, 2263, 3105, 3652, 2459, 2626, 2646, 2709, 2752, 3026, 2320, 2114, 2421, 3624, 3656, 3695, 2283, 3156, 3339, 3362, 3322, 3005, 2204, 3508, 2101, 3625, 3154, 3620, 2315, 4600, 1955, 5299, 2772, 4020, 4414, 4643, 5153, 3317, 3437, 4426, 1941, 1981, 2068, 4756, 4581, 3057, 1862, 2041, 2678, 2769, 1990, 2188, 5009, 4325, 3948, 2479, 3846, 4255, 4960, 2775, 3404, 3737, 2598, 2871, 1885, 1942, 1771, 2000, 5198, 475, 4778, 4227, 1993, 4258, 2725, 4353, 4630, 2685, 3295, 1653, 3423, 3788, 3155]. shrimp [2394, 3887, 4047, 4491, 4533, 3230, 1971, 4651, 1650, 3456, 1374, 1815]. shrimpers [3741]. Shrimps [2615, 3126, 3695, 2429, 2070, 3448, 3965, 339, 2390, 2007, 2563, 2468, 2522]. Shrinkage [818, 3949]. shrip [158]. Shumagin [105]. Sibling [2305, 2280]. sic [1464]. sicklefin [5289]. Sicyonia [3182, 3178]. Side [2099]. Side-to-Side [2099]. sided [4946, 2659, 2882]. Sidescan [3391, 4284, 4340]. Sidescan-sonar [4284]. sidestriped [4630]. sieboldii [4252]. sierra [1675, 2073, 1675]. Siganus [5374]. sighted [2974]. Sightings [3425, 3711, 3869, 2681]. signals [2291, 3802]. signatures [5067, 4672, 4967]. signatus [4633]. Significance [1528, 2261, 2383, 3459, 2756, 2979, 1505]. significant [3520]. silenus [4634]. Siliqua [1564, 1621]. silky [4552]. Sillaginidae [3929]. Sillaginodes [3929, 4151, 4510]. Sillago [3929]. Siltation [3427]. Siluriformes [3296]. siluroid [167]. Silver [3014, 3124, 2986, 3081, 2314, 2356, 3029, 2803, 1782, 4916, 4346, 3009, 896, 4077, 4917, 1881, 65, 3233]. silvergray [4744]. Silverside [3201, 2556, 2951, 5186, 2458]. silversides [1513, 314]. Similar [2750, 1484]. Similarity [3069]. similis [2112, 2972, 2725]. simillimus [2874]. Simple [3566, 5299, 2277, 2793, 3313, 3749, 3840, 3811, 3072, 124, 3774, 2679, 3847, 2732, 3897]. Simplification [2569]. simplified [2683]. simulate [4222]. Simulated [3467, 2673, 2645, 2507, 2442, 4808, 4581, 4075, 4532]. simulating [3292]. Simulation [3401, 2408, 3090, 2578, 2439, 5057, 5340, 4547, 5354, 3774, 4718]. simulations [2693, 4053, 4613]. Simulator [2263]. simultaneously [3387]. Sinaloa [3948]. Single [2012, 3163, 5101]. sinuensis [3695, 4258]. sinus [3585]. siphonophore [2121, 2630]. siphonophores [2300]. Site [3525, 4441, 3953, 4392, 4892, 4785, 5109, 4317, 3161, 5188, 5130]. Site-Intensive [3525]. site-utilization [4441]. Sites [2136, 4034, 4832, 4960, 4294, 4354, 5206, 5178, 414]. Situ [5066, 2626, 2646, 2709, 2326, 3132, 2270, 2876]. Six [3435, 3672, 4130, 4590, 3928, 3099, 22]. sixgill [4537]. sixty [1625]. sixty-two [1625]. Size [2913, 2568, 3566, 3062, 3251, 2862, 4021, 2875, 3343, 3603, 3068, 3513, 3517, 4423, 3646, 3185, 3324, 3702, 3368, 2163, 2211, 2418, 2504, 2281, 3533, 3974, 3436, 2429, 2852, 2262, 3146, 2359, 2035, 3049, 2644, 3098, 3102, 2982, 4558, 4622, 4117, 4961, 1955, 2744, 4251, 4347, 2772, 5028, 342, 5115, 3779, 4349, 4542, 4688, 4871, 3044, 3811, 3842, 3773, 3141, 3133, 5061, 2894, 4286, 3844, 3276, 4481, 4507, 4607, 4482, 3608, 4006, 4979, 5319, 4899, 4736, 372, 4812, 4082, 5292, 1771, 4637, 3987, 1837, 1897, 4906, 3930, 4154, 4517, 5077, 4510, 4199, 4236, 1693, 4611, 4317]. size [4246, 1987, 3730, 2557, 2701, 5084, 4259, 4488, 2205, 3289, 4261, 3155, 5034, 5055, 4762, 4157, 1686, 3034, 4163, 3075, 3539, 4810, 4923, 3268, 5064, 4650, 2849, 4267, 4663, 4301, 3850, 3345, 5134, 4378, 3830, 3970, 5231]. size- [5061, 4082]. size-based [4286, 4810]. Size-dependent [4021, 5292]. Size-Frequency [3566, 3811, 4317]. size-maturity [3141]. size-selective [4622]. size-selectivity [4251]. Size-Specific [3533, 3324, 4558]. sized [2652]. Sizes [3253, 3475, 4665, 4664, 5007, 4764, 4826, 5216, 4178, 1696, 2679, 4727]. Skagit [2322]. skate [4892, 5341, 4913, 5270, 3710, 4976, 4573, 4699, 4733]. Skates [2355, 4874, 5268]. skeletal [4501, 1465, 4678]. skeletochronology [4471]. skeleton [1185]. Skeletons [2143, 250, 595]. Sketch [163, 176, 1735]. skew [5193]. skewed [4031]. skin [442]. Skipjack [3198, 2669, 3477, 2895, 2028, 2187, 2898, 2383, 3329, 2592, 3202, 2165, 2447, 2896, 2920, 2151, 1932, 4403, 5212, 2373, 2744, 1675, 2968, 2454, 3643, 3265, 4420, 4851, 1691, 1860, 4186, 1964, 1795, 2065, 1833]. Skipped [5183, 5221, 5160]. skull [3833]. sleeper [5114, 4248]. Slender [2665]. slimy [260]. Slipper [5123, 4688]. Slope [4197, 4413, 3941, 4284, 4092, 4511, 3916, 3471, 5211, 3894, 2403, 4535, 3470, 2518]. slough [685]. slow [4920]. sludge [3834]. smack [211]. Small [3382, 4541, 4658, 2108, 2136, 2954, 3504, 2554, 2565, 3436, 2663, 2137, 5044, 778, 871, 5344, 2591, 5105, 2651, 3851, 3801, 5104, 4778, 115, 369, 5188, 2766, 518, 5162, 4491, 5203, 2648, 4276, 2732, 4840]. Small-boat [4658]. small-meshed [778]. Small-mouthed [3382]. Small-Scale [2137, 4541, 2554, 4840]. small-tooth [5188]. Smaller [3522, 1104]. smalleye [5365]. smalltooth [5297, 5298]. smaris [5238]. smell [1448, 1467]. smelt [2900, 5281, 2573, 70, 4819, 2748]. smelts [1573]. Smith [2048, 2200]. smithi [3129, 2393]. Smithsonian [396]. smoked [627]. Smoking [587, 648]. Smolt [3461, 4046, 4763]. Smolts [2819, 2238, 3149, 1861]. Smooth [3451, 2881, 4100, 2900, 4523, 5225, 5219, 4037, 5252]. smooth-hound [5219]. smoothhound [5289]. smoothing [4859]. smoothtongue [3036]. snail [3452]. snailfish [4205]. snails [563]. snake [511, 517, 2234, 2406, 3770, 3990]. Snakehead [5071]. Snakes [345, 404, 331]. Snapper [5087, 2773, 2808, 2577, 3071, 3176, 4498, 5271, 4412, 4719, 5113, 5028, 699, 799, 4060, 4026, 4350, 3937, 4671, 4711, 3580, 4967, 4721, 4075, 4564, 3000, 4287, 3933, 3974, 4549, 4076, 4448, 4221, 5267, 4857, 3940, 4177, 5111, 4762, 3843, 4955, 5343, 5064, 4266, 4650, 4918, 1154, 4451, 4137].

Snappers [3552, 3840, 4252, 1611, 1574, 5086, 5295, 3557, 3471, 2604]. snoek [4525]. snook [4355, 4131]. Snout [2959]. snouted [4846]. Snow [3622, 3673, 2886, 2559, 2492, 4174, 5082, 5141, 2599, 5069, 3986, 4265, 4902, 3221, 3870]. snowy [5294, 4323]. snyderi [3219]. so-called [1495, 546]. soak [2535, 5081, 4635]. Society [829, 730, 53]. Socioeconomic [3483]. Sockeye [2096, 2660, 2304, 3370, 2134, 2136, 2499, 2091, 2111, 2197, 2262, 3149, 2255, 2257, 2716, 4310, 1998, 2474, 1895, 4827, 1899, 4273, 1954, 4954, 3446, 1997, 1804, 4715, 1943]. Soft [2625, 3449, 3318, 2322, 2929, 3191, 2514, 454, 2544, 477, 346, 1787, 1905]. Soft-body [2929]. Soft-Shell [3449, 2322, 3191, 2625, 2544, 477, 346, 1905]. softshell [4243]. software [4919, 2668]. solandri [1969, 5147]. solar [1960]. sold [74]. soldierfish [3900]. Sole [2937, 3109, 3685, 2545, 2665, 2820, 2890, 2970, 4783, 4364, 3892, 5363, 3862, 5355, 4526, 2543, 3721, 4417, 4094, 4905, 3244, 5018, 5292, 3725, 2584, 4741, 4176, 4397, 4932, 4742, 2832, 3476, 2953, 4065, 4204, 5241, 3640, 4861, 1951, 2428, 3551, 4407]. solenocerid [2563]. soles [931]. solidissima [2292, 5215, 5339, 2707]. Solomon [3672]. Soluble [3572]. Solubles [2177, 2212]. Solution [2047]. Solutions [2640, 4862, 2394]. Solvent [2038]. Solvent-Extracted [2038]. solving [3944]. Somatic [4410, 3646, 3674, 3628, 3783, 4285, 4507, 4859]. Some [1052, 2116, 1519, 223, 2530, 2225, 2301, 529, 905, 751, 3004, 1359, 2105, 1438, 2022, 1817, 2269, 3113, 3229, 2312, 727, 2418, 2844, 1177, 330, 2030, 2442, 1975, 3475, 1147, 844, 1063, 3286, 476, 461, 1018, 1124, 2973, 2141, 821, 1116, 2250, 1995, 2218, 3493, 23, 2731, 3497, 770, 2887, 90, 62, 3158, 44, 801, 1899, 2702, 2894, 111, 3178, 88, 1639, 2742, 1256, 1247, 530, 2323, 2436, 1518, 2682, 458, 2209, 1653, 1659, 185, 119, 307, 4375, 452, 4099, 248, 1035, 4593, 3256, 4926]. sometimes [432, 537, 222]. Somniosus [5114, 4248]. Sonar [3391, 2462, 4801, 4284, 4529, 4340, 5347]. Sonic [1847]. sordidus [2665, 3964]. sorenseni [4801, 2126]. sorting [4937]. Sotalia [4571]. soul [4118]. Sound [4310, 3382, 2631, 5102, 4404, 3489, 5080, 2566, 2475, 2463, 4094, 4130, 5233, 3498, 3108, 4293, 1434, 2589, 2259, 2846, 2845, 5012, 1099, 4360, 1623, 3148, 4204, 4341, 4434, 2714, 4883, 5022, 5161, 717, 4158, 3097, 3449, 3606, 812, 1656, 3591, 5321, 5339, 1634, 1645, 3400, 171, 1046, 4922, 75]. sounder [5071]. soundings [832]. Sounds [2069, 3277, 2072]. source [4942, 2596, 49]. Sources [1560, 2715, 3415, 4631, 4999, 1102, 3852]. South [4117, 5039, 3502, 3629, 4036, 4126, 4251, 4719, 4617, 4212, 174, 3902, 2621, 3113, 2598, 5294, 5161, 3812, 4299, 3970, 1782, 1784, 1813, 457, 1880, 4835, 172, 4587, 1877, 1002, 4329, 2296, 4798, 984, 5342, 5005, 4657, 3438, 3929, 810, 2657, 4884, 3372, 2615, 3979, 988, 4462, 859, 5324, 663, 145, 4077, 4525, 2808, 3352, 4464, 3577, 3496, 3329, 292, 506, 478, 2735, 3924, 4372, 3849, 5060, 4610, 4568, 4739, 4485, 3701, 4570, 5239, 4104, 4614, 4061, 3565, 2368, 3962, 3074, 4122, 4661, 3209, 2973, 4355, 3018, 3079, 3371, 3995, 4224, 2734, 2789, 4323, 2083, 4137]. south-central [4835]. southeast [5005, 4709, 3437, 4750, 4596, 4257, 4695, 4177, 4773, 3931, 5348, 5136, 3764, 4665, 4838]. Southeastern [2506, 2677, 5065, 3756, 1822, 4410, 5120, 3953, 2869, 3080, 5170, 4060, 5095, 1641, 4463, 4998, 5176, 4313, 3787, 4127, 3042, 4153, 4546, 4199, 5289, 4033, 5287, 3106, 3306, 2865, 4107, 1624, 1640, 3826, 5150, 4894, 3410, 1339, 5343, 5019, 3669, 4814, 4700, 4492, 2428, 1992, 3308, 5139, 4386, 3955, 2500, 5369, 5123, 5231, 4977, 3321, 2316, 2592, 2958, 3537, 3436, 1631, 3553, 3699, 2146, 3098, 3495]. Southern [3569, 5106, 3593, 3650, 3582, 3422, 3599, 3706, 3709, 5105, 3237, 4169, 5117, 3630, 3705, 578, 5078, 2797, 3758, 4856, 4018, 62, 4453, 2597, 3381, 4657, 4389, 4579, 4561, 5079, 5334, 4311, 912, 1874, 3157, 4670, 4423, 4503, 4990, 1221, 4148, 3746, 3948, 4967, 4505, 965, 4757, 4545, 958, 4050, 2742, 1300, 1170, 2607, 3791, 1000, 4653, 5311, 5378, 4785, 961, 4171, 4219, 3832, 3994, 4429, 4484, 4446, 4778, 4834, 5082, 4589, 5227, 4570, 2071, 3562, 2725, 5314, 4771, 4758, 3976, 3796, 5069, 5137, 3776, 3022, 1618, 3245, 1164, 4809, 5205, 3485, 3907, 4765, 3808, 3102, 2989]. southern [3247, 4571, 4134, 5076, 4535, 2695, 4407, 3759, 5312, 2527, 1517, 3184, 3470, 3280, 5107, 3648, 2885, 2341, 2774, 3334, 4030, 4913, 2297, 2434, 2978, 2159, 4294, 3189, 2542, 2667, 3063, 2476, 2855, 3600, 3602, 2578, 2141, 3510, 4767, 3347]. Southside [693]. Southwest [3394, 5133, 4312, 1945, 4157, 5210, 5147]. Southwestern [3583, 1998, 5372, 4479, 5115, 5293, 4673, 4151, 4987, 915, 4377, 3198, 3068, 3497]. Sowerby [5092]. Sp [3451, 3602, 3003, 46, 2231, 3453, 241, 3703, 2057, 2224, 2361]. Space [2174, 2186, 2000, 5097]. Spacing [2099, 3632, 4689, 2547]. spadefish [3880, 3577]. Spain [888]. spangled [4757]. Spanish [275, 2524, 3793, 2637, 3177, 3437, 354, 3399, 4714, 4104, 61, 3826, 4716, 3808]. Sparid [2657]. Sparidae [4707, 4871, 3783, 3580, 3128, 4675, 4668, 2510, 4221, 4375, 3553, 2817, 2828, 4086]. sparse [5261]. Spat [2783, 463]. Spatial [3078, 3793, 4884, 3312, 3546, 4138, 5304, 3773, 4797, 4625, 4944, 4092, 5032, 4972, 4917, 4511, 4895, 3597, 5230, 3550, 4950, 4262, 4341, 3758, 5044, 4498, 5102, 4541, 3860, 4671, 4712, 4066, 4215, 5245, 5035, 5240, 3930, 4530, 4742, 4122, 4489, 4269, 4594]. spatially [5015, 5215]. Spatio [3887]. Spatio-temporal [3887]. Spatiotemporal [5365, 5379, 5126]. Spawn [2884, 4652, 5273]. spawned [4309, 4351, 4809]. spawner [4846, 5003, 4058, 5251]. Spawners [2677, 1985]. Spawning [4327, 2516, 4516, 2625, 581, 333, 3346, 3517, 2634, 3020, 3201, 3388, 2917, 2986, 2827, 3781, 4764, 2040, 3373, 2136, 2759, 2907, 3329, 1724, 3498, 2062, 2879, 1656, 2067, 4331, 4839, 4517, 4587, 3914, 3310, 3672, 115, 2796, 4317, 3439, 2586, 3525, 4061, 2870, 3191, 4007, 3600, 2175, 2531, 3271, 253, 3873, 3628, 455, 4883, 4655, 4783, 4498, 4705, 1782, 1813, 700, 3910, 28, 4357, 4873, 1685, 4719, 3336, 4935, 2460, 2704, 3224, 4060, 3878, 3809, 4423, 3779, 5007, 3172, 3416, 3813, 5061, 4998, 138, 5290, 2505, 2874, 2924, 3028, 3458, 4313, 4722, 1469, 1763, 4255, 4526]. spawning [4368, 4675, 1896, 1626, 1637, 3721, 4626, 2891, 2968, 5327, 5294, 4628, 5003, 4010, 5060, 324, 2532, 3025, 5140, 1771, 5161, 1897, 4161, 4725, 5183, 5221, 4714, 5306, 3242, 2698, 5236, 4610, 1936, 4074, 2550, 2792, 4696, 1818, 1877, 2818, 254, 4598, 3796, 4743, 1598, 3920, 2585, 5309, 3938, 3155, 5160, 4122, 5206, 4420, 3826, 3799, 1218, 3148, 5359, 2691, 3297, 4300, 1943, 865, 4128, 187, 1674, 4575, 998, 1932, 2005, 1783, 4380, 337]. spawning-biomass-per-recruit [5140]. spawning-date [4696]. spawning-ground [700]. spawning-per-recruit [4705]. spawning-stock [4626, 1897]. spawnings [3304]. spear [4163]. spearfish [3545]. spearfishes [1794]. special [5096, 302, 1189, 1192, 1326, 1485, 1518, 1482, 1659, 2260, 1095, 1250, 1091, 4765, 1663, 1585]. Species [2805, 2017, 2116, 3017, 2433, 2631, 2847, 2657, 2671, 2851, 2637, 3160, 2231, 3677, 3451, 2102, 2611, 3361, 3364, 3453, 3702, 3650, 1179, 3190, 2757, 2320, 4987, 4866, 2421, 3008, 2341, 3695, 2456, 3561, 2583, 2355, 3672, 2538, 3661, 3594, 3610, 4881, 2408, 2589, 2070, 2276, 2603, 2812, 2800, 2311, 3047, 3616, 2262, 3487, 2606, 2575, 2745, 3639, 2295, 2247, 2285, 2018, 2713, 3079, 3628, 3497, 3631, 2982, 4493, 4242, 1809, 1284, 5119, 1955, 945, 1822, 5310, 5342, 2903, 3977, 4624, 2693, 4348, 4641, 3840, 1223, 5170, 5376, 5344, 4952, 2588, 4350, 5196, 3142, 2621, 1092]. species [1123, 1219, 1233, 1491, 2899, 1625, 5061, 1793, 2280, 1990, 5075, 5144, 3055, 4889, 5338, 1778, 950, 1600, 4722, 1117, 4731, 4750, 5288, 5096, 4482, 64, 1584, 1574, 1571, 4980, 2758, 5216, 829, 1194, 1186, 1262, 3777, 93, 167, 940, 1088, 1169, 1258, 1251, 1320, 1330, 3791, 1305, 3829, 4907, 1439, 3957, 2508, 1479, 2690, 5175, 4956, 4373, 956, 5166, 4786, 4055, 4800, 5037, 1802, 5350, 4220, 290, 5321, 5001, 2385, 3859, 3933, 5086, 1253, 3875, 5023, 4352, 5128, 2829, 4771, 2085, 1889, 5055, 5156, 4063, 2587, 4099, 4894, 4163, 5244, 547, 4837, 1071, 1238, 1461, 4205]. species [1946, 4810, 4616, 4874, 4406, 3120, 4902, 3103, 192, 1933, 5234, 4666, 3753, 2604, 4639, 3099, 1888, 3968, 3897, 3287, 682]. species-habitat [5086]. Species-specific [4866]. Specific [3566, 3185, 3582, 3266, 3533, 3162, 4558, 4656, 3414, 4935, 5010, 3324, 4866, 4628, 4812, 4898, 4509, 3814, 3248, 3946, 1514, 4049, 5277, 4887, 4137]. Specifying [3461]. Specimens [2413, 396, 279, 2398, 250, 3808]. speckled [2700, 405]. spectra [4831]. Speculations [2417]. speculator [3739]. Speed [2980, 1798, 2588, 2529, 3285, 4858, 4724, 4235, 5127]. Speeds [2878, 2664]. spengleri [2717]. sperm [4622, 1929, 694, 3683, 615]. sperm-whale [694, 615]. Spermatogenesis [3565]. Spermatophores [2390]. spermatozoa [1681]. Sphoeroides [2717, 2271, 2717]. Sphyraena [2110]. Sphyraenidae [2110]. Sphyrna [3434, 5326, 5225, 4985]. Spicara [5238]. Spider [2622, 2595, 3311]. spiderfish [4413]. spill [4130]. Spillover [4865, 4784]. spilopterus [2961]. Spin [2375]. Spin-labeling [2375]. spined [2]. spinefoot [5374]. Spines [2908, 5337, 3777, 4221, 3301, 3966]. Spinner [2830, 3137, 2777, 2145, 2601, 3690, 4709, 3717, 4367, 2821, 2107, 3850]. spinning [2]. Spiny [3427, 2339, 3645, 3002, 3863, 3636, 2378, 3175, 3602, 3344, 1517, 2877, 3779, 4542, 4688, 5199, 3773, 4929, 4889, 3241, 5337, 1944, 4455, 3835, 4569, 5227, 3752, 4222, 5126, 4384, 4969, 2579]. Spirontocaris [2922, 3170]. Spiruroidea [2231]. Spisula [2292, 5215, 5339, 2707]. splendens [3285, 2392, 3911, 2575, 2587, 2233]. split [5229]. split-beam [5229]. splittail [3086]. Spokane [328]. Sponge [1372, 3079, 1364, 1368, 5362, 3303, 923, 1173]. Sponge-Coral [3079, 3303].

Sponges [1755, 1498, 1364, 320, 1152, 1114, 1203, 1378, 1440]. spontaneous [3785, 2968]. Sporozoön [1473]. Sport [1366, 2753, 2230]. sportfish [4666]. Sportfishing [3601, 3617]. Spot [3455, 2464, 3392, 2848, 3389, 3082, 4734, 3781, 1638, 3059, 5270, 2585, 2686, 2845, 3223]. Spotfin [2022, 4962]. Spottail [2179]. Spotted [3234, 3438, 2619, 5074, 3222, 3274, 2473, 4640, 4190, 5005, 3837, 2838, 3821, 4761, 4367, 4778, 5190, 4259, 2590, 3879, 4359, 4598, 4632, 4097, 4661, 4299, 5054, 4836, 5369]. Spp [3382, 1880, 3049, 4989, 4460, 4617, 4889, 4392, 4415, 5283, 5262, 4372, 3927, 4820, 4785, 4766, 4217, 5240, 4154, 3877, 2276, 3710, 3031, 4872, 5366, 2938, 4186, 1971, 4887, 5264, 3688]. Spratt [569]. Spring [3127, 2726, 3213, 3093, 3420, 5132, 3578, 681, 3026, 2252, 3342, 475, 566, 3703, 2918, 5224, 3836, 581, 2859, 79, 41, 912, 5170, 3878, 3956, 3064, 3133, 4069, 3921, 3787, 344, 3851, 1972, 2004, 1975, 4176, 1164, 207, 3221, 4682, 4128, 3990, 334, 370, 400, 5040]. spring-run [4069]. spring-spawning [3878]. Squaloid [2086]. Squalus [3912, 139, 5199, 2579, 5337, 4455, 5037, 4569, 5227, 5126, 4969]. Squamation [2964, 2985]. squammosus [4688]. Squamscot [154]. squamulosus [5257]. square [4036, 4126, 4210, 4251, 4175, 4875]. square-mesh [4036, 4126, 4210, 4251]. Squares [2354, 4133]. squaretail [4888]. squarfish [1829]. Squat [3556]. Squatina [5328]. Squatinidae [5237]. squawfish [3351]. Squeteague [1227, 1231, 1416, 1475]. Squid [3017, 2452, 2649, 3573, 2999, 3686, 3314, 3999, 4309, 3856, 4024, 5133, 4958, 4830, 497, 918, 5155, 4652, 5243, 2876, 3333, 4481, 3801, 3985, 4197, 4685, 4647, 4704, 4676, 4802, 3148, 3761, 585, 4810, 4923, 4267, 4682, 4302, 4247, 4755]. squid-trawl [4802]. Squids [2692, 4211, 2540, 3554]. Squilla [2721, 2864, 3155]. Squillidae [2721]. Sr [3726, 4599]. Sr/Ca [4599]. SRLCA [3625]. St [4251, 4347, 2052, 383, 494, 4479, 3424, 4763, 3225, 4834, 4339, 3273]. St. [3569, 3765, 5258, 3624, 3208, 1050, 3040, 2429, 2548, 512, 3742, 1994, 1999, 2534, 1848, 401, 3353]. Stability [3600, 2266]. Stabilization [1728]. Stabilizes [3116]. Stable [3105, 5208, 4735, 4924, 4925, 4107]. Stage [5010, 3266, 3433, 2378, 3947, 4499, 5348, 4601, 4328, 4827, 3333, 3170, 4066, 5041, 5302, 3964, 1971, 2249, 2775, 2886, 3798]. stage-based [4328]. Stage-I [3798]. Stage-specific [5010]. Stages [3536, 2032, 3575, 3700, 2216, 2207, 2999, 2302, 2200, 2798, 2357, 5255, 4070, 3950, 1519, 3977, 4928, 1939, 1862, 5324, 2922, 3144, 4988, 1966, 4022, 5157, 3725, 4115, 3304, 1854, 2861, 771, 4469, 4063, 5268, 4303]. staghorn [3963]. Staging [5073]. stagnant [260]. staminea [2259]. Stand [2324]. Standard [3343, 4780, 2744, 3142, 4456]. Standing [2445, 2134, 2100, 2424, 2450, 3618, 3230, 3562]. standpoint [1357]. Star [2803, 2098, 534, 719]. star-fish [534, 719]. Starfish [2011, 1654, 1187]. starfishes [1293, 1311]. Starks [4158]. stars [1991]. Starvation [3581, 3674, 2647, 3238, 2764, 4977, 3276]. Starvation-induced [3238]. Starved [3694]. Starving [2992, 2833, 3950]. State [2358, 4404, 1106, 210, 4093, 1072, 425, 5097, 4133, 453, 2687, 1720, 2149, 1122, 2655, 313, 428]. state-space [5097]. stated [164]. Statement [448]. States [90, 267, 562, 396, 62, 48, 1128, 150, 987, 189, 21, 159, 3465, 206, 521, 140, 1244, 190, 953, 3041, 3263, 325, 2958, 193, 263, 1158, 170, 437, 855, 131, 313, 296, 264, 231, 1002, 1094, 1164, 1379, 238, 134, 499, 291, 214, 248, 329, 1076, 1089, 5162, 1782, 1784, 1813, 2296, 5322, 5065, 410, 422, 5301, 942, 1822, 4389, 2693, 5120, 4709, 252, 213, 1496, 15, 918, 921, 995, 3437, 3486, 875, 4060, 5095, 5199, 5237, 1079, 1123, 1880, 3321, 4998, 3055, 1600, 5176, 30, 4313, 728, 4750, 4127, 4481, 6, 1638, 302, 2834, 5360, 1189, 3042, 2598]. States [4418, 1355, 1526, 1075, 2715, 4226, 1148, 1149, 3043, 5345, 4974, 4546, 545, 4695, 4199, 3537, 393, 37, 3980, 4156, 4983, 5287, 2251, 3106, 759, 158, 944, 3788, 952, 3826, 3553, 5150, 4894, 1166, 173, 214, 270, 248, 264, 329, 300, 883, 1014, 1059, 1080, 1165, 1314, 5343, 3509, 5019, 1107, 1116, 3699, 3669, 192, 3098, 4880, 4492, 4773, 218, 3595, 4514, 1527, 3308, 3495, 2707, 2518, 3931, 5139, 5369, 455]. Station [2352, 670, 979, 1284, 150, 225, 835, 710, 5245, 1113, 319, 1119, 1478, 653, 424]. stationary [4724]. Stations [892, 1284, 785, 457, 4171, 170, 134, 654, 832, 2382, 573]. Statistical [4959, 2844, 1593, 1612, 1631, 1634, 2239, 995, 4394, 3825, 1817, 4373, 4330, 3286, 1089, 4593]. Statistics [252, 921, 887, 932, 671, 826, 2667, 808, 763, 317, 1059, 875, 751, 802, 4712, 173, 1795]. Statolith [3504, 3573, 3985]. statoliths [4124, 4267]. Status [2940, 3496, 2203, 3904, 2743, 3538, 5002, 1793, 4464, 4064, 4094, 3681, 2737, 5184, 3551]. steady [4133]. steady-state [4133]. Steam [2352, 425, 453]. Steam-electric [2352]. Steamer [422, 562, 1158, 628, 1012, 1289, 765, 867, 832, 1109, 429, 401, 572, 1292, 886, 1251, 1295, 1291, 170, 607, 909, 1242, 207, 879, 939, 963, 1297]. steamers [228, 267, 943, 277, 92, 196]. Steelhead [2404, 2009, 2084, 2234, 2819, 2402, 2250, 2786, 2826, 2409, 3637, 2406, 5277]. steepness [4790]. Steindachner [3743]. steindachneri [5334, 2248]. Stellar [3989]. stellatus [684]. Steller [4706, 4730, 4464, 5114, 4586, 2927, 4825, 5024, 5366, 5358, 4665, 4838, 4551, 4664]. Stellifer [2803]. Stellwagen [5132]. Stenella [4640, 3234, 2811, 3837, 2838, 2231, 3717, 3821, 4761, 4367, 3222, 2821, 2830, 3137, 3274, 2777, 4259, 2145, 2107, 2276, 2473, 2601, 2590, 3690, 3879, 3861, 4097, 3049, 3850]. stenellae [2231]. Stenobrachius [4493]. stenolepis [4540, 4910, 4794, 2513, 4290, 5035, 5235, 1951]. Stenotomus [4686, 2990, 3128, 4226]. Stephidae [1776]. stepwise [5340]. stereo [4393, 4394, 5086, 5217, 4986]. stereo-video [4393, 4394, 5086]. stereological [4687]. Sternoptyx [2767]. steroids [4204]. Stettiner [126]. stickle [2]. stickle-back [2]. Stickleback [2265, 2499, 1823]. stigmaeus [3964, 2700]. Stilomysis [2361]. Stimulating [2199]. Stimulation [2426, 1635]. stimulations [1968]. Stimuli [3672, 1225]. Stingray [2199, 5370, 5282]. Stizostedion [2928, 3153, 3218]. Stochastic [2913, 3035, 2273, 5340]. Stock [4190, 4002, 2875, 4873, 3488, 3677, 2640, 2354, 3775, 2462, 3373, 2134, 2282, 4508, 4169, 3431, 2100, 5208, 3188, 3116, 2424, 3002, 3180, 3651, 2736, 3261, 2670, 3587, 3751, 2253, 3341, 5002, 4622, 4705, 3999, 4829, 4054, 4310, 4893, 4799, 5303, 5332, 5214, 4106, 5102, 5103, 4212, 4602, 3840, 5101, 4804, 4576, 4808, 5061, 4427, 4712, 3748, 4077, 4255, 4790, 5013, 3951, 2607, 4979, 4626, 4627, 3681, 4215, 5353, 4044, 5345, 5159, 1897, 5088, 5183, 4554, 4547, 5077, 5117, 5097, 5121, 3282, 3730, 4400, 5287, 3861, 4223, 4758, 5052, 3499, 5014, 4078, 4157, 3838, 2400, 4203, 4125]. stock [4744, 3230, 3669, 4457, 3907, 4665, 5184, 4043, 5277, 4664, 5147]. stock-enhancement [4044]. Stock-production [2640]. Stock-rebuilding [4508]. Stock-Recruitment [3751, 4255, 4790]. stock-specific [5277]. Stocking [318, 126, 3978, 297]. Stocks [2654, 2660, 2445, 2566, 2952, 3041, 2450, 3653, 2786, 4705, 4164, 4388, 4540, 4346, 3837, 4884, 1840, 4046, 4092, 5096, 4508, 3854, 4834, 4660, 3562, 4983, 3796, 3712, 5330, 3885, 3852, 4593, 3421, 5256]. Stockton [470]. Stolephorus [192, 2196, 2435]. Stomach [4452, 3440, 3081, 3596, 4014, 3211, 2276, 2856, 3325, 2591, 4768, 4495]. Stomachless [2222]. Stomachs [2112]. Stomatopod [3495]. stomatopoda [1196, 2721, 2864]. stomias [4553, 3938, 4901, 3287, 4017, 4102]. Stomiatoid [2301, 2963, 2247]. Stomiatoidei [2767]. stomius [3248]. Stone [2743, 5325, 3734, 2760, 2737]. Stonecrab [3433]. stony [1197]. Storage [2038, 1469, 1486]. Storer [3043, 3415]. Storm [3189]. stoutii [3631]. straight [4693]. straight-shank [4693]. Strait [3871, 2631, 1758, 3851, 3380, 3579, 3893, 3586, 3722, 3251, 4191, 1758, 3862, 5121, 3820, 2976]. Straits [4091, 1641, 5032, 3767, 2941]. Stranded [3245, 4193]. Stranding [694, 2735, 2830, 3931, 4878, 4020, 2821, 2801]. Strandings [5094, 3697, 3868, 4485]. strange [464]. Strategies [2910, 4120, 4730, 5338, 4121, 4418]. strategy [5197, 4644, 5215, 5139]. stratification [3022]. stray [240]. Straying [4273]. Stream [2520, 608, 1120, 1651, 2409, 2648, 4903, 3204, 1975, 3537, 2661, 4344]. Streambed [2677, 1980]. streamer [4580]. streamer-tag [4580]. Streams [2136, 2438, 318, 1003, 297, 953, 980, 1500, 4932, 1341, 1450, 1477, 328]. Strength [2088, 2539, 1772, 3984]. strengthen [634]. Stress [2536, 2239, 5323, 3848, 5232]. striata [3511, 3975, 5189, 4686, 5222, 5099, 2192, 5353, 5236, 5230, 3799, 3511, 3308, 5308]. striatus [4357, 4161]. strigeoidea [1846]. strikes [3363]. Strings [4919]. Striped [2654, 5102, 5103, 2634, 2988, 2279, 3368, 116, 3327, 2493, 2485, 4089, 4856, 5044, 1813, 5296, 2656, 4516, 2781, 3336, 2965, 3440, 3064, 3664, 4945, 4443, 4289, 3367, 2495, 3342, 4044, 4082, 4566, 4161, 2650, 4383, 4941, 4465, 4610, 4568, 4739, 1664, 338, 8, 18, 4886, 1647, 3861, 3946, 3778, 3925, 4822, 5203, 59, 4700, 4181, 4574, 3332, 75, 400, 592, 911]. stripetail [4019]. stripey [4628]. Stromateidae [2874]. Stromateoid [2358, 2555]. strombid [4124]. Strombus [3394, 3733, 4644, 3354, 3584, 3824, 3618, 3718, 3885, 3852, 4041, 4206]. Strong [5249, 5356]. Strongylocentrotus [4522, 4602, 4618, 4712, 2582, 4590, 2098]. strontium [4311, 4221, 4599]. strontium-to-calcium [4599]. Structural [2512, 4828, 5125]. Structure [2240, 3648, 3109, 3652, 3373, 2358, 2649, 4071, 3189, 1416, 3693, 3175, 3176, 2035, 2893, 3102, 3512, 3371, 5098, 4117, 1916, 3999, 4309, 4829, 4559, 4893, 4934, 5049, 4799, 4453, 4997, 5028, 3336, 4772, 4229, 4576, 5155, 3981, 4671, 3719, 4427, 4196, 4506, 3748, 3775, 5027, 4169, 5341, 4291, 5035, 4966, 4154, 5148, 5117, 5270, 3974, 4396, 4025, 5374, 4398, 4335, 3282, 5287, 3861, 2837, 4035, 3754, 4080, 2, 4762, 3887, 3841, 4318, 5150, 4203, 1943, 1475, 4116, 4323, 5123, 4365]. structured [5345, 4263]. Structures [2481, 3555, 3639, 2289, 4441, 4501, 5122, 4427, 4786, 4449]. Studied [2524]. Studies [2050, 2046, 1602, 1874, 3023, 1677, 2380, 1628, 1435, 2064, 2347, 1664, 1696, 2378, 2101, 4500, 3355, 1998, 5320, 4026, 1665, 2636, 1315, 1629, 1526, 1563, 1962, 3279, 3780, 4704, 1392, 1018, 4857, 3922, 4179, 1863, 2406, 5179]. Study [3474, 2730, 2027, 3051, 2122, 2210, 3575, 2187, 2232, 3453, 1931, 2277, 3494, 3586, 2339, 2223, 2556, 3591, 2302, 2140, 1820, 3525, 2569, 1769, 2514, 3587, 3639, 3526, 4081, 3016, 1457, 4032, 4411, 1909, 1613, 1488, 4728, 1592, 1609, 60, 4945, 1586, 1103, 1326, 1792, 4776, 1518, 3763, 4537, 5011, 1764, 2515, 1893, 1894, 1340, 71, 4793, 1423, 4457]. Studying [2375, 2610, 1384]. stumps [2270]. Sturgeon [2413, 2983, 5207, 5112, 4057, 2959, 1058, 5305, 4994, 5258, 5351, 406, 5315, 959, 5306, 3914, 952, 4180, 4376, 4320]. sturgeons [952]. Style [3610]. stylirostris [4365]. Subadult [3591, 5233, 3342, 4039, 3332]. Subarctic [2530, 2702, 2080]. subfamily [3840, 2559]. subgenus [2390]. submarine [4490, 4356]. submerged [2270, 4980]. submergences [4615]. Submersible [3547, 3249, 3794, 4016, 4233, 5070, 5182, 3750, 3877, 4529, 2630, 4009, 5264]. submersible-survey [4009]. submersibles [3941]. subnormal [1188]. suboptimal [3367]. Subpopulation [2870, 2348]. Subsampler [2651]. subsamples [4607]. Subscription [4279, 4308, 4326, 4345, 4361, 4387, 4408, 4424, 4440, 4458, 4475, 4497, 4518]. subsequent [79, 111]. subsidy [4494]. subspecies [4038, 4384]. substances [1723]. substantiating [2348]. Substitute [2177, 4832]. Substrata [2475]. Substrate [2514, 4835]. Substrates [3474, 4897]. Substratum [3339]. subsurface [2382]. Subtidal [2183, 2475, 3597, 2836]. subtilis [2468]. Subtropical [3063, 4541, 4800, 4957, 4262, 5254, 3502]. subtropically [4351]. subyearling [2826]. Success [3483, 362, 257, 341, 726, 2126, 325, 263, 447, 4251, 4541, 4130, 4234, 3796, 2760]. Successful [118, 2525, 2397, 4652, 297, 232, 5191, 931]. sucker [1009, 1579]. suckers [273]. suckleyi [2922]. sudden [4673]. Sueur [1937, 1644, 217, 1579]. Suez [2300, 2614]. Suffolk [155]. Suggested [2564, 1359, 4826, 1045]. Suggestions [943, 3, 166, 6, 39, 1166]. suggests [3970]. Suitability [2130, 5367]. suitable [4958, 5249, 155, 218]. Sul [4670]. sulfate [1665]. sulfonamide [1777]. sulfonamide-resistant [1777]. Sullivan [2642]. Suminoe [4604]. Summary [1195, 2752, 2294, 269, 110, 1012, 711, 2471, 3723]. Summer [3563, 3127, 3756, 3213, 3977, 5079, 2530, 2869, 2634, 2134, 2666, 2946, 2771, 2031, 2180, 3442, 3083, 2237, 2612, 5002, 3999, 2726, 3923, 936, 3950, 5170, 2702, 1524, 1572, 4150, 973, 1251, 4627, 4371, 1087, 1086, 3342, 4876, 961, 4713, 4941, 1015, 4589, 4176, 3903, 1507, 470, 4449, 1567, 963, 4809, 5205, 4128, 3990, 3830]. summer-spawned [4809]. Summer-winter [2634]. Summering [3197]. summers [3408]. Sumner [1641]. sun [1188]. sun-fish [1188]. sunlight [707]. sunray [4157]. sunspot [2947]. superba [3422]. superbus [2448]. superciliosus [5152, 2940, 4170, 2444, 4639]. Superintendents [248]. Superior [1906, 1903, 2265, 638]. Supersaturation [2009, 2084, 1989]. Supplement [2441]. Supplemental [1646, 2315]. supply [4756, 1356, 2370, 1360, 104, 5063, 1031, 1623, 1624]. Support [3478]. Supports [2964, 2837, 2985]. Supposed [305, 734]. Suprabenthic [3429]. surf [2292, 2576, 2868, 2707]. Surface [3502, 1811, 4942, 2350, 2174, 3707, 2033, 2609, 2336, 2920, 2639, 2317, 2597, 3299, 4710, 5105, 2682, 3272, 2071, 3676, 5173, 4636, 2382]. surface-migrating [5173]. Surface-Schooling [2920]. surface-to-benthic [2597]. surfclam [5215, 5339, 4157]. surfperch [2613]. surgeonfish [4135]. surinamensis [3860]. Surplus [3169, 5047]. Surplus-Production [3169]. surrounding [1493, 4666]. surveillance [5253]. Survey [3483, 2191, 1947, 3385, 4730, 2692, 3055, 3186, 3378, 3558, 3249, 2816, 3625, 1545, 1546, 5081, 4848, 5202, 4211, 4959, 4390, 4772, 2540, 5059, 4527, 5046, 4483, 4754, 3100, 5070, 5182, 2719, 2668, 4330, 5161, 4445, 4091, 5350, 5276, 4175, 4513, 4466, 4265, 4450, 4841, 5241, 4009, 4683, 5131, 5179, 2707, 1585]. Survey- [4730]. Surveyed [2141]. Surveys [3444, 3447, 2673, 2996, 2844, 2976, 5119, 3923, 4658, 4728, 4994, 3959, 4149, 3679, 3921, 3951, 4585, 4016, 5245, 4444, 4982, 3286, 3754, 3883, 2224, 4904, 5033, 5055, 5318, 5264, 5171, 5040]. Survival [2568, 2839, 3474, 2622, 2240, 2234, 2677, 3613, 3327, 3584, 4791, 2256, 4685, 2397, 2580, 3656, 2350, 3500, 3088, 3119, 3525, 2852, 2457, 3076, 3640, 3673, 4128, 2118, 3574, 2783, 3350, 3523, 3505, 5348, 4601, 4426, 2694, 1996, 5305, 3773, 4827, 1859, 5351, 4046, 4001, 4472, 4945, 4992, 2930, 4693, 2645, 4947, 2340, 4621, 5018, 4082, 2004, 4954, 1985, 2385, 4334, 5023, 4981, 4432, 2845, 5242, 3946, 2394, 4180, 4065, 2406, 5211, 1652, 3990]. Survive [3613]. survivorship [5377]. Susan [2253]. susceptibility [4983]. Susquehanna [22, 53]. Sustainability [4533, 4747]. sustainable [4906]. sustaining [4604]. sustenance [5009]. Suwannee [4759]. swallow [537]. swash [4436]. Sweden [251, 725, 232, 811, 775, 813]. sweeps [4971]. sweetlips [4757]. swept [3961]. swept-area [3961]. Swiftsure [1645]. Swim [2358, 3171, 2521, 1228, 5308]. Swim-bladder [2358, 1228]. Swimbladder [3795, 3888, 4338]. swimmer [4603]. Swimming [2878, 2216, 2147, 2153, 4584, 2223, 2338, 3024, 2756, 2809, 5185, 2329, 2664, 4460, 2588, 2529, 4761, 3285, 3262, 4235, 5127, 4433, 4915, 4436, 2902, 2905, 1181, 2897]. swimmingspeed [2697]. Swiss [779]. Switch [2638]. Switch-Feeding [2638]. switching [4646]. Switzerland [731, 268]. Swordfish [2914, 2964, 2941, 4191, 4349, 4849, 4313, 4606, 4042, 3954, 4064, 3300, 4942, 4740, 2985, 3774, 3761, 4406, 4534, 3767, 3966, 3247, 836]. swordspine [5163]. Swpe [3319]. Symbiont [3602]. Symbiosis [2076]. Symbiotic [3602]. symmetricus [1767, 2497, 1859, 2132, 2055, 2647, 2849, 3238]. Sympatric [3621, 2438, 5042, 4074, 3710, 3743]. Symphurus [4280, 3661, 4144, 4962, 4240]. Symphysanodon [2017]. synagris [4857]. Synaphobranchus [2403]. Synaptas [1176]. Syndinidae [3307]. Synergistic [2655, 2118]. Synergy [5259]. Synodontidae [5075, 2346]. Synodus [5075, 4909]. Synopsis [1846, 570, 856, 1072]. synoptic [4856]. synthesis [4992]. synthetic [2292, 2394]. System [2568, 2222, 2564, 2105, 2675, 3533, 2429, 2197, 2025, 2253, 4856, 1920, 1923, 1925, 5079, 5222, 2188, 3465, 4393, 4394, 1945, 1157, 3462, 2668, 4330, 5001, 5086, 3562, 1308, 1943, 5172, 5038, 5066, 2054, 2090]. Systematic [2203, 1844, 2302, 2514, 2018, 2672, 1445, 1488]. Systematics [4850, 2767, 2389, 2280, 2676, 4144, 5128, 2800, 2816, 1974, 3177, 3036, 1953]. Systems [2061, 2326, 1998, 4754, 4986].

T [241, 4276]. T-anchor [4276]. T. [4558, 4245, 5373]. Taaningichthys [2102]. Table [149]. Tables [1551, 426, 3449, 313]. tabs [5293, 5285, 4305]. Tag [4462, 2240, 3140, 3599, 2496, 2909, 3673, 4403, 5031, 3731, 3741, 4580, 5326, 4945, 4079, 4545, 3168, 4416, 4693, 3804, 4536, 2272, 4088, 4660, 4295, 5287, 3301, 2205, 4805, 3905, 4377, 4225, 5250]. Tag-recapture [3140, 5031, 5326, 3804, 3905, 4377]. tag-recaptured [3301]. tag-recovery [4536, 4660]. Tag-reporting [4462, 3741, 4079]. tag-return [4545, 5250]. tag-shedding [4403, 4079]. Tagged [3641, 2343, 2453, 3508, 3699, 2309, 1903, 4111, 3408, 2841, 2104, 4288, 4169, 5373, 4088, 3108, 3348, 3772, 3889, 5076, 5085]. Tagging [3477, 2108, 3706, 3709, 2954, 4704, 3587, 3027, 4558, 4948, 4250, 4500, 4580, 1758, 4863, 3387, 1524, 1572, 3073, 4079, 4776, 4295, 5135, 3982, 1578, 1583, 1605, 1627, 1984, 3922, 4179, 4861, 1607, 3061]. tags [2807, 2130, 5293, 4710, 5305, 4472, 4691, 5371, 5200, 4583, 4584, 4621, 4684, 4834, 4531, 4851, 4131, 4276]. Tahiti [706]. Tahoe [1325, 1461]. tail [1990, 3777]. tailed [3908]. tainha [5208]. Taiwan [3434, 4642, 4935, 4170, 4534]. Take [3595, 533, 4666]. Taken [2484, 3617, 205, 321, 4607, 3994, 626, 597, 1987, 1659, 515, 4468, 1095, 424, 5092, 4539]. taking [391]. Taku [4950]. talismani [3397]. Tampa [1878, 1970, 3559, 2217, 2347, 4227, 4512, 1933, 2215, 5204]. Tangipahoa [495]. tango [230]. Tank [1872, 423]. tanks [825]. Tanner [48, 3932, 4174, 5082, 170, 5069, 5137, 4015, 4265, 3640, 4340, 4902]. tanneri [4015]. tannerl [3311]. Tapes [2783]. tar [1447, 1529]. Taractes [2248]. Taractichthys [2248]. tardigrades [1736]. Target [2088, 5096, 5217]. targeted [5344]. targeting [5129]. targets [5305, 5055]. tarpon [4425, 3979]. Tarpum [445]. Tasmania [4018, 4877]. Tasmanian [3418]. taste [454, 1241]. tau [2602, 1377, 754]. Taunton [230]. taurus [3056]. Tautog [2427, 2586, 2727, 4441, 5181, 4972, 2278, 4335, 3745, 4575]. Tautoga [4441, 5181, 4972, 2278, 2427, 2586, 2727, 4335, 3745, 4575]. Tautogolabrus [4123, 2222, 2536, 2427, 2727, 3067]. Taxa [2831, 5340]. taxidermy [267]. Taxonomic [2562, 2261, 2940, 5343]. Taxonomy [1913, 2410, 2616, 2488]. Tay [371]. taylori [4298]. Technique [3590, 3257, 4937, 3937, 4734, 3073, 3367]. Techniques [2954, 2350, 2793, 2375, 4801, 5050, 3934, 4617, 4687, 3647, 3286]. technology [4945, 4693]. TEDs [5039, 3788]. Teleconnections [2364]. telemetric [3763]. Telemetry [2914, 2310, 4948, 4477, 4391, 4503, 4990, 5368, 5316, 5203, 3816, 4307]. Teleost [2568, 2087, 4616, 1968]. Teleostean [2335, 1615, 1479, 1494]. Teleostei [4352, 3675, 4928, 4641, 3939, 5075, 3567, 4055, 5029, 3544, 4648, 4881, 4771, 3553, 3843, 4837, 3485]. Teleosts [2302, 1639, 3904]. telson [3777]. temperate [4674, 5175, 4154, 4351, 3485, 3907]. temperate-water [5175]. Temperature [1550, 2063, 2669, 3343, 3603, 2138, 2472, 2185, 2189, 3969, 3785, 3700, 3650, 3026, 1622, 3524, 2863, 2761, 2340, 2173, 2126, 2341, 2350, 3500, 2372, 2353, 5321, 2174, 3181, 2194, 2457, 2204, 2118, 5118, 1552, 1551, 5296, 2945, 1907, 678, 2407, 4710, 4780, 3285, 4001, 3948, 3111, 5105, 3737, 2930, 4988, 4584, 4215, 4737, 5003, 5018, 4217, 4235, 1954, 4334, 1854, 170, 1768, 3946, 4180, 5178, 2233, 3351, 4636, 2002, 675]. temperature-dependent [4988]. Temperature-related [5321]. Temperature-Salinity [3650]. Temperature-time [2340]. Temperatures [2383, 3455, 2045, 2011, 2115, 3676, 2033, 2609, 1586, 3384, 4789, 2697, 4951]. Temporal [3078, 4498, 5065, 5136, 4542, 2602, 5294, 4742, 4297, 4122, 4817, 4137, 3793, 4884, 4541, 4138, 3860, 4797, 4625, 4481, 4066, 4215, 5006, 5032, 4972, 4511, 4895, 4530, 5230, 4950, 4260, 4262, 3887, 4269, 3758]. temporarily [2245]. Ten [408, 1806, 1825, 317]. ten-pounder [1825]. Tenakee [2500]. tench [288, 561, 813]. Tennessee [1481, 1491, 954, 941, 1000, 511]. tenuis [4010, 2051, 2662, 2226, 2305, 2266, 3085]. Teredinidae [1342]. teres [3713, 2546]. Term [2467, 2589, 3063, 3474, 5332, 3731, 5050, 5174, 5028, 4426, 2965, 3902, 3453, 4464, 4545, 2536, 3232, 5188, 2852, 4131]. terms [248, 264]. terraenovae [3935, 5304, 4048, 4546, 3045]. terrapin [1511, 477, 1606]. Territory [448, 1100, 939]. Test [4476, 4317, 124, 3106, 4385]. Testicular [4655]. testing [4145, 5313, 4547]. testis [1801]. Testosterone [3659]. tests [3797, 3375]. testudineus [2717]. testudinum [2215]. Tethering [3590]. Tetrabrachium [2925]. Tetracycline [3137, 3387, 4000, 4844, 3138]. tetracycline-marked [4000, 4844]. tetraodontid [1967]. Tetraodontiformes [2676]. Tetrapturus [4089, 2684, 4693, 4791, 4738, 4743, 2485, 3545]. Teuthoidea [4830]. Texas [2749, 1955, 46, 3438, 2772, 3741, 4643, 4426, 2082, 3057, 993, 1011, 4159, 1626, 190, 1136, 3728, 4917, 3766, 115, 1595, 3788, 3962, 2855, 3083, 3230, 3154, 4918, 128]. Texture [3403]. Thais [1874]. Thalassia [2215]. Thalassinidea [3695]. Thalassomyces [2361]. Thanks [545]. Tharybidae [1793]. thazard [1689]. thec [2401]. Their [2802, 2568, 3677, 3081, 2566, 2677, 2267, 2192, 2045, 3656, 3131, 3537, 2501, 2174, 3442, 2262, 2558, 2165, 2982, 3993, 4281, 1052, 5342, 451, 5102, 5103, 1725, 79, 2389, 15, 4423, 3860, 3880, 1011, 1219, 5099, 663, 2894, 4427, 4944, 4092, 1036, 4482, 2196, 2901, 1846, 2653, 4979, 1192, 1867, 1842, 1505, 5378, 4913, 4234, 4484, 1482, 1361, 1985, 4629, 4886, 1982, 3070, 1232, 4789, 1145, 1338, 1384, 1340, 3986, 5008, 5146, 1412, 1166, 1126, 264, 1153, 4701, 5145, 1154, 1621, 1476, 4051]. thelyca [2390]. them [388, 3387, 111, 1356, 320, 237, 221, 278, 982]. Thenus [5364]. Theoretical [2730, 2369, 2682, 5362]. Theory [2507, 2405, 3430]. Theraga [2712]. Theragra [2726, 3871, 5053, 3977, 3572, 3953, 3080, 4672, 3253, 3611, 3704, 4795, 3111, 3373, 3608, 4859, 3380, 3579, 3893, 4737, 2883, 3901, 3722, 4530, 4275, 4456, 2741, 4489, 4341, 4434, 3949, 4663, 4665, 4378, 4703, 5021, 3526, 4664]. thereto [287]. Thermal [2342, 2469, 2352, 2084, 2689, 2700, 2589, 2900, 2694, 2746, 3104, 5306, 3136, 5023, 1768, 3958]. thermally [4459]. thermocline [1567]. thermometer [551]. thermoregulation [5290]. Thermoregulatory [2342]. these [298]. Thickness [2742]. Thin [2583]. Thin-layer [2583]. thing [656]. things [881]. thirteen [3840]. Thirty [3361]. Thirty-Four [3361]. Thomas [3030, 3742]. thompsoni [2285]. Thoracica [2432]. thornback [4588]. thorny [3710, 4699, 4733]. thornyhead [4687, 4480]. thornyheads [4978, 4417]. those [1189, 941, 1563, 1484, 65, 1091, 877, 885, 893, 908, 916, 925, 1499, 1530, 4922]. Thread [2571, 2334, 2938, 2205, 314]. thread-bearing [314]. threaded [4369]. Threadfin [3609, 4543, 2698]. threaten [220]. Three [2116, 3474, 2423, 2605, 2657, 3677, 2341, 3561, 2638, 4220, 2311, 3049, 1503, 3534, 2831, 2151, 3631, 2497, 3336, 1990, 4729, 4075, 3115, 1584, 4980, 4415, 1966, 4987, 167, 3791, 3849, 4956, 4973, 3859, 3933, 4155, 4344, 3796, 3022, 4125, 2604]. Threespine [2499, 1823]. Thresher [2940, 5152, 5331, 4170, 2444, 4639]. Threshold [2063, 3196, 2323]. throughout [4126, 5091, 173]. thumbprint [5184]. Thunberg [4757]. Thunnidae [1671]. Thunnini [4908]. Thunnus [4655, 4558, 5212, 5116, 4829, 2778, 2807, 3800, 4250, 3397, 3831, 4477, 4963, 4478, 4311, 4073, 4391, 4423, 4503, 4605, 4710, 2529, 3596, 4148, 1953, 4852, 2383, 3262, 3463, 3599, 3706, 3709, 3038, 4428, 4821, 4545, 4367, 4245, 4583, 4584, 1887, 3401, 4595, 5373, 3849, 2496, 2593, 2554, 2565, 2909, 2281, 2682, 3971, 4091, 4839, 2343, 2109, 3537, 4447, 4317, 5269, 1987, 2276, 3484, 2308, 3159, 4061, 4142, 3265, 3565, 3616, 4495, 1894, 3663, 4007, 4338, 4531, 3841, 2431, 2567, 4011, 4435, 5078, 4457, 4186, 2043, 4767, 2896, 2920, 4116, 4817, 2005, 2083, 5083]. thynnus [4710, 4428, 4655, 2778, 2807, 4477, 4963, 4478, 4073, 3596, 4852, 3038, 4428, 3971, 4091, 2343, 3537, 4447, 4317, 5269, 3159, 2043]. Thyroid [2199, 1455]. Thyrsites [4525]. thyrsitis [3403]. Thysanopoda [2207]. tiburo [5326]. Tickfaw [62]. Tidal [3392, 2831, 1807, 3212, 4932, 5205, 4759, 4651, 3836]. tidal-creek [4759]. Tide [1179, 2030, 1117, 1959, 1162]. tide-pool [1117, 1162]. tidepool [5020]. tidepools [3259]. Tiger [3358, 3056, 5194, 4358, 4295, 4246, 3998, 4272]. tilapia [1871, 1872, 2170, 2400, 2387, 2916]. tilapias [4399]. Tile [3397, 211, 1306, 1308]. tile-fish [211, 1306, 1308]. Tilefish [3496, 3041, 3074, 3102, 1159, 3465, 3200, 3647, 5077]. Tilesius [4648, 1873]. Tillamook [3065]. Time [2622, 3540, 3517, 3169, 2827, 2640, 639, 2548, 1943, 3292, 2535, 5081, 3762, 3441, 3133, 3781, 4231, 4821, 4417, 4508, 2340, 1656, 5104, 4906, 4547, 4981, 5242, 4007, 4489, 4824, 4635, 3873, 4271]. time-area [4231, 4821]. Time-dependent [2640, 4271]. time-varying [4417]. Timely [2976]. Times [3578, 3528, 2704, 4245, 4226, 4136, 2388, 3351]. Timing [2597, 3762, 4868, 4003, 3459, 4348, 4580, 4675, 4993, 5022, 5034, 4616, 3845]. Tippecanoe [1253]. tissue [1949, 4285, 1188, 1469, 1378, 1440]. Tissues [2058]. Title [3529]. toadfish [2602, 1377, 754]. Tobago [4332]. Tocopherol [2344, 2441]. together [149, 1106, 3704, 236]. Togiak [3149]. Tolerance [2469, 2084, 2898, 3104, 2372, 2353, 1586]. toluene [2570]. Tomakomai [4157]. Tomales [1477]. Tomcod [2623, 2883, 2798, 2784, 2623, 2784, 2798]. Tomlinson [2675, 2670]. Tomtate [2958]. tonging [4646]. tongs [957]. tonguefish [4962, 4240]. Tonguefishes [3661, 4144, 3732]. Tonnage [300]. tons [411]. took [680]. Tool [3391, 5000]. tools [5056, 5070]. tooth [5188]. toothfish [4614]. top [2906, 5045]. tope [4588]. TOPEX [4222]. TOPEX-POSEIDON [4222]. Topical [570, 856]. Topography [1548, 3192]. Torpedoes [1718]. torptate [3209]. torrent [2409, 2557]. Tortugas [1981, 4955, 4719, 1941, 3965, 1245, 1307, 5161, 1993, 4047]. Total [344, 4831, 3142, 3360, 1990, 4216, 2425]. total-body [4216]. Totoaba [2487, 4112, 4112]. tourist [5162]. tow [4959]. Towed [2172, 2962, 3803, 2071]. Towed-Net [2172]. towed-pump [2071]. towers [1864]. towing [4858]. Townsville [3981, 3985]. tows [3715]. Toxic [2615, 2531]. Toxicities [1529]. Toxicity [2194, 2357, 4183]. Toxins [2030]. Trace [4427, 3255, 2695, 4190]. tracer [1938]. Trachichthyidae [3791]. Trachinotus [4461, 1880, 4940]. Trachipteridae [2606]. Trachipterus [3034]. Trachurus [1767, 2497, 3502, 1859, 4008, 2132, 2055, 2647, 2849, 3238, 5247]. Trachypenaeus [2070]. tracked [5200]. Tracking [4710, 2554, 4391, 1847, 2565, 3220, 5217, 5203]. tracks [2668]. tract [1452, 1471, 2867, 4275, 1335, 4247]. Tracy [5168]. trade [501, 208, 339, 173, 934, 1201]. traditional [5066, 3896]. Traitors [2417]. Traits [3513, 5362, 3786, 4718]. trajectories [4787]. Trajectory [3982, 4207]. Trajectory-based [3982]. Transect [3280, 3549, 2673, 3376, 5202, 4658, 4529, 3918]. transects [3961, 4662]. Transfer [477, 4558, 38]. Transferrin [4043, 2786]. Transferring [567]. transformation [5324]. transforming [2398]. transient [5042, 4572]. transiting [5349]. Transition [2593, 4200, 4997, 4602]. Transitions [4860, 2148, 5157]. Translocation [4644]. transmentanus [4057]. transmission [682, 80]. transmitter [2519]. transmontanus [2959, 3914]. transpacificus [4819]. transparency [1556]. Transplant [3489]. Transplanted [3490, 3765, 4035, 3816]. Transplanting [675, 944, 658, 166, 221, 527]. transponder [5293]. Transport [1914, 3420, 2395, 2539, 3478, 3965, 4756, 2541, 4750, 4932, 4222, 4514]. Transportation [117, 2234, 2819, 478, 483, 1952, 104, 1418, 818, 2406, 1363, 773, 423]. transported [3088, 4207]. Transporting [189, 76, 1383, 1408]. Trap [2366, 2779, 2418, 2535, 5081, 636, 4470, 4929, 4724, 5276, 4340, 1663, 5286, 334]. trap-nets [334]. trap-video [5276]. Trapping [491, 495, 3754, 460]. Traps [2652, 3466, 5253, 4900, 4690, 709]. trauma [3303]. Trawl [3385, 3638, 2844, 3079, 4600, 4686, 4036, 4126, 4090, 4210, 2306, 5299, 5153, 889, 5170, 4581, 5305, 5313, 4284, 4968, 5059, 4607, 5096, 4527, 4417, 4585, 5046, 4754, 4016, 2751, 4685, 4445, 4982, 5350, 4175, 4353, 4513, 2835, 3286, 4466, 3031, 5033, 4971, 4802, 4265, 4450, 4841, 5241, 4533, 1721, 5149, 5264, 4683, 4753, 5131, 5179, 5040]. Trawl-Caught [3079]. trawl-survey [4445, 4466]. trawlability [5347, 5071, 5229]. trawlable [4585, 4535]. trawled [4937]. Trawling [3400, 889, 4960, 3404, 4226, 1883, 3794, 5125, 4800, 4227, 3303, 4419]. Trawlnet [3624]. Trawls [2056, 3494, 3961, 4347, 2853, 3317, 3437, 2133, 4939, 5069, 4491, 3640, 4186, 4986]. trays [478]. Treatment [551, 2369, 1777, 1413, 108, 1388]. treatments [4866]. treaty [631]. tree [2270]. Trematoda [1874, 2231, 1846]. Trends [3234, 2772, 2138, 3570, 3561, 2667, 5171, 3797, 5065, 3711, 3744, 3679, 4968, 4464, 3739, 1885, 4965, 1790]. Trepang [817]. trevalla [4018]. trevally [4788]. triacanthus [4122]. triacis [3746]. Triakis [5005, 4299, 3587, 4556]. trials [5026]. triangle [4015]. triangularis [1522]. tributaries [517, 1878, 1978, 1000, 5273, 429]. tributary [1625, 2826, 1450, 1477, 4128, 982]. Trichechus [2719]. Trichiurus [4372]. Trichodon [2884, 4814]. Trichodontidae [3610]. tricuspidata [2207]. tridens [2752]. tridentatus [5134]. Triggerfish [3173, 5120, 5176, 5368]. Triglidae [2456, 4517, 2620]. Triglops [4771]. Trimethylamine [2815, 2957]. Trinidad [3604, 4332, 3592]. trip [21, 535, 207, 595, 370, 401]. Triphoturus [2780]. tripletail [3860]. triploid [4604]. tripod [2636]. trispinosa [3817]. Trocadero [266, 848]. Troll [2131, 2525, 5043, 4782]. Trophic [5212, 3634, 2657, 2894, 5225, 2265, 3704, 2942, 2490, 2632, 3462, 3008, 3701, 5289, 2408, 2620, 3164, 3900, 4876, 4867, 4957, 4305, 4306]. trophodynamics [4997, 4012]. Tropical [2044, 2191, 3231, 2016, 3343, 2847, 2231, 2705, 3222, 3376, 3378, 3570, 2592, 3504, 3573, 3240, 3274, 2473, 2601, 3049, 2037, 2146, 3493, 3217, 2982, 5212, 4640, 1988, 2373, 3711, 4460, 4830, 5346, 3299, 3900, 3717, 3771, 3821, 4147, 3869, 4564, 4079, 4991, 3089, 3801, 3985, 4197, 4907, 4736, 2321, 2971, 5020, 4609, 3859, 3933, 3875, 3643, 4027, 4259, 2590, 4354, 2085, 2511, 3499, 3883, 4097, 5072, 4533, 4186, 4267, 5173, 4302, 2681, 3952, 3298, 2897, 167, 2203]. Trotline [3411]. trouble [642]. Trout [2605, 2404, 2009, 2084, 2234, 311, 2023, 3386, 3674, 3612, 2357, 2402, 2026, 2250, 2786, 361, 559, 415, 376, 904, 1255, 4799, 1789, 333, 693, 1810, 3412, 1570, 399, 416, 715, 143, 3864, 273, 56, 1054, 779, 85, 728, 653, 491, 524, 1683, 297, 506, 603, 1325, 381, 972, 419, 914, 915, 403, 787, 263, 2184, 619, 553, 1083, 1387, 1373, 2409, 3637, 1659, 234, 1801, 1382, 1698, 2617, 1108, 2406, 883, 672, 1777, 1461, 652, 151]. trout-culture [1054]. trout-fishing [1325]. trouts [1414, 1638, 1326]. trumpeter [4700]. truncatus [2749, 3464, 4878, 3593, 4159, 2944, 2719, 4485, 3766, 4965, 3723, 2855, 2734, 4782]. trutta [2617]. truttae [1259]. tryannus [3129]. tschawytscha [1666]. tshawytscha [4560, 4559, 2219, 5348, 2520, 3009, 3014, 4919, 5101, 2412, 5167, 2819, 3133, 3408, 3945, 4069, 3917, 5290, 2810, 2755, 2629, 2252, 4274, 2826, 5233, 1962, 3588, 4373, 4566, 4484, 2004, 3181, 4396, 5143, 3784, 4033, 4035, 3398, 3271, 4128, 3341, 3512, 2918, 3990, 2006]. Tuamotu [1932]. Tucca [1967]. Tuccidae [1967]. tumor [1569]. Tuna [3198, 3689, 2950, 2669, 3477, 2895, 3397, 2099, 2730, 2344, 2441, 3542, 3596, 2898, 2383, 2705, 3599, 3706, 3709, 3038, 3615, 3329, 2592, 2214, 2281, 3202, 2343, 2109, 3537, 3672, 2276, 3425, 3484, 3676, 3565, 2073, 3616, 3663, 2160, 2165, 2431, 5078, 2447, 2896, 3217, 2920, 3430, 2151, 5083, 4655, 4403, 4558, 5212, 5116, 4640, 2778, 2807, 1988, 2373, 2744, 4477, 3711, 4963, 4478, 4311, 2838, 4073, 4391, 4423, 4503, 4605, 4710, 3744, 3842, 4147, 4761, 4148, 3869, 4908, 4852, 3262, 3463, 4079, 4428, 4821, 4545, 4367, 1871, 4245, 4583, 4584, 1887, 1724, 2891, 2968, 1792, 4595, 5373]. tuna [2272, 2682, 2971, 3971, 4091, 4839, 1851, 4740, 1693, 4317, 5269, 4027, 3865, 3159, 3265, 1732, 1889, 4468, 4495, 5043, 1692, 4007, 4338, 4420, 4531, 4851, 3841, 1831, 4011, 1838, 4435, 4186, 2043, 4767, 1964, 1781, 1679, 1674, 3952, 3298, 4116, 4817, 1932, 2065, 1783, 4524]. tuna-like [1679]. tuna-vessel [3744]. Tunas [2016, 3343, 2028, 2187, 2196, 2049, 2169, 2241, 2308, 1781, 4038, 2588, 3771, 1953, 1673, 1956, 3643, 1676, 1667, 1695, 1687, 1884, 4186]. tune [1802]. tuned [4078]. tunicate [2738]. Tunicates [1756]. turbid [2000]. Turbidity [3339, 4724]. turbine [1979]. turbot [1381]. Turkey [2344]. Turkeys [2441]. Tursiops [2749, 3464, 4878, 3593, 4159, 2944, 2719, 4485, 3766, 4965, 3723, 2855, 4187, 2734, 4782]. Turtle [3385, 3697, 4504, 2267, 3337, 2215, 5039, 5057, 3858, 4020, 4132, 5320, 3404, 4152, 1141, 4677, 3788, 3922, 4179, 3761]. turtles [1737, 5039, 5065, 4847, 4410, 4601, 5352, 3959, 1301, 4653, 2867, 4697, 4555, 3962, 4678, 4615, 280, 4228, 4515, 4539]. twelve [1494]. twenty [1169, 4917]. twice [3986]. Twilight [2556]. twin [3456, 1326]. twine [4347]. twinning [4074]. Two [3449, 3467, 3470, 2231, 411, 3552, 3364, 3453, 2643, 3613, 2757, 829, 2392, 1084, 2030, 2655, 3661, 3594, 3610, 3047, 2415, 4493, 5004, 4521, 4719, 3741, 3934, 4772, 3837, 4038, 4252, 2588, 3141, 3746, 2899, 1625, 1793, 4046, 4167, 4042, 4792, 2463, 2379, 4545, 4723, 1574, 1571, 2758, 3192, 3777, 1169, 4152, 3050, 5070, 2477, 2468, 4536, 3908, 3279, 956, 5037, 1948, 5042, 3647, 4430, 3875, 1208, 4294, 4984, 4677, 3070, 2085, 3732, 3984, 2368, 4751, 3471, 1071, 4205, 4402, 4810, 4902, 3443, 3046, 293, 3287, 3743, 4365]. Tychoplanktonic [2642]. Type [3251, 2423, 3437, 5196, 3842, 4734, 3917]. Types [2910, 3591, 2235, 4961, 2664, 943, 5284, 4973, 4344, 2227, 5347].

Typhlachirus [5363]. typical [4594]. typus [4807, 5275, 5318]. tyrannus [3379, 3947, 4343, 4963, 2729, 2240, 2878, 2931, 3044, 3054, 2489, 2776, 4749, 3969, 3795, 3888, 2029, 2103, 4966, 4235, 3581, 2539, 3320, 3786, 5335, 3873, 4514, 2095].

U [125]. U. [945, 1025, 138, 961, 768, 2226, 2305, 628, 1289, 979, 676, 939, 963, 1091, 187]. U.S. [3629, 2892, 2275, 4410, 4346, 5202, 4282, 3806, 5194, 3842, 4619, 4854, 4231, 5096, 4214, 3113, 4066, 5046, 3681, 3812, 4589, 5097, 4353, 1012, 5111, 5126, 5248, 3792, 4593, 5224, 4228, 4840]. U.S.A. [3107]. Uca [2194, 2118, 2480]. Uku [3552]. ultra [5313]. ultra-low-opening [5313]. Ultracentrifugation [3605]. Ultrasonic [2554, 4307, 4948, 4477, 2519, 4391, 4503, 2565, 3816]. ultrasound [5300]. Ultrastructure [3451]. Umaira [2267]. Umbrina [3753]. Umpqua [4629]. unarmored [2392]. unassisted [4761]. unassociated [5185]. Uncertain [3751]. uncertainty [4829, 5218, 5014]. unconsolidated [5125, 4773]. Uncoupling [3783]. uncultivated [680]. undecimalis [4355, 4131]. Undergraduate [2142]. Undersea [3192]. Undersize [3216]. understand [5253]. Underwater [5253, 2486, 5228, 5349, 5222, 4846, 5167, 4167, 5070, 4751, 5172]. Undesirable [2042]. Undisturbed [3684]. undosquamis [2298]. Undredged [2018]. undue [1422]. undulatus [3853, 3910, 4120, 4968, 2010, 3455, 5154, 4291, 3059, 4548, 2792, 5097, 4139, 4296, 3389, 2537]. uneatus [3429]. unequal [4753]. Unexplained [323, 306]. Unexploited [3319, 3974]. unfished [3981, 4041]. Ungava [4371]. unialgal [1791]. unicellular [1546, 1750]. unidentified [1778]. unilateral [1801]. unique [4243, 1684]. Unit [3415]. United [3321, 521, 1355, 2715, 3263, 3043, 2958, 3537, 3553, 3699, 3098, 3495, 2518, 1782, 1784, 1813, 90, 2296, 5322, 5065, 410, 267, 422, 562, 5301, 62, 48, 942, 1822, 1128, 4389, 2693, 5120, 4709, 252, 150, 213, 1496, 15, 918, 921, 3437, 3486, 875, 189, 4060, 5095, 5199, 5237, 1079, 1123, 1880, 21, 4998, 3055, 1600, 5176, 30, 159, 4313, 728, 3465, 4750, 4127, 206, 4481, 6, 140, 1638, 302, 2834, 5360, 1189, 190, 3042, 2598, 4418, 953, 3041, 1526, 1075, 4226, 1148, 1149, 5345, 4974, 4546, 325, 545, 4695, 193, 4199, 263, 393, 37, 3980, 1158, 4156, 4983]. United [5287, 2251, 3106, 759, 170, 158, 944, 3788, 952, 3826, 437, 855, 5150, 4894, 1166, 173, 131, 214, 313, 270, 296, 248, 264, 329, 231, 300, 883, 1014, 1059, 1080, 1165, 1314, 1379, 5343, 238, 3509, 5019, 134, 1107, 1116, 3669, 192, 499, 4880, 4492, 4773, 218, 3595, 4514, 1527, 3308, 2707, 291, 3931, 5139, 5369, 455]. Unites [1244, 1164, 396]. unobserved [4640, 5069]. unrecognized [1071]. Unreported [5162]. Unselective [2387]. Unstructured [2158, 2369]. unsuitable [10]. untrawlable [4585, 5056, 5055, 4986]. Unusual [723, 2327]. unwanted [4126, 4210]. Updated [5302]. Upland [2217, 2347]. Upon [3280, 3204, 2135, 966, 660, 228, 717, 983, 949, 227, 355, 211, 276, 3140, 993, 1093, 1106, 946, 965, 1073, 481, 385, 541, 659, 864, 1356, 92, 958, 1036, 744, 1360, 1087, 1086, 961, 531, 1394, 897, 687, 991, 597, 707, 952, 1397, 248, 356, 377, 379, 496, 398, 544, 632, 635, 647, 749, 804, 819, 862, 891, 1047, 1135, 1317, 963, 962, 548, 675, 982, 400, 123, 919]. Upper [1758, 4822, 1601, 1009, 4325, 1532, 4511, 3916, 4112, 4163, 3894, 1585, 3007, 3470]. uproots [4747]. upstream [2810, 1954]. Uptake [1796, 2386, 2068, 2493, 2570]. Upwelling [2094, 2016, 2524, 2420, 2445, 2251, 2399, 2359, 1892, 2317, 3462, 2552, 3964, 2979]. urban [2043]. Urchin [2098, 4602, 4712, 2582, 4590]. urchins [4522, 4618]. Urea [2178]. Urobatis [5370]. Urochordata [2738]. urophthalmus [4463]. Urophycis [3806, 2036, 2038, 4010, 2226, 2305]. Urosalpinx [1628, 2024]. Urotrygon [5282]. ursinnus [2284]. ursinus [3273, 3492, 4792, 2708, 3316, 3284, 3867, 2889, 2882, 3353]. Uruguay [4327, 4331, 4567]. USA [38, 3419]. Use [4557, 1533, 1707, 4829, 2050, 5332, 4251, 5222, 3677, 4945, 2462, 4064, 3827, 3504, 5285, 3621, 3263, 4621, 2871, 3401, 4444, 4554, 4941, 3591, 4834, 492, 5011, 4107, 2402, 5203, 3659, 5318, 4986, 4379, 4860, 5044, 5116, 5039, 4054, 4388, 4799, 4540, 4459, 4801, 4643, 15, 918, 905, 5346, 4999, 3412, 277, 4990, 3941, 4652, 4505, 4427, 4366, 387, 4821, 3118, 5337, 5360, 5331, 4184, 3927, 4010, 5297, 4085, 5378, 5161, 5193, 4786, 5109, 4537, 5086, 1672, 4317, 4629, 4735, 4201, 4844, 2127, 196, 611, 4809, 4268, 1403, 5172, 5361, 4276, 5229]. Useable [2576]. used [5228, 4182, 4873, 4772, 4426, 5320, 1938, 4933, 92, 3241, 5345, 1673, 2515, 358, 4125, 4491, 4875]. useful [4731, 256]. usefulness [298]. uses [4623, 1817, 1031]. Using [2293, 3035, 3644, 3183, 3467, 5194, 5344, 4728, 3105, 3494, 4877, 5288, 3601, 3570, 3872, 3237, 4870, 3405, 3356, 2350, 4982, 2866, 2793, 5272, 4983, 5089, 3445, 5343, 3605, 4727, 4249, 3590, 3711, 3466, 4846, 4581, 5305, 4781, 3941, 3679, 4075, 4712, 3706, 4428, 4668, 3777, 4565, 4694, 3804, 5215, 3849, 3807, 101, 4660, 3818, 3933, 4396, 4890, 4000, 4530, 5287, 4456, 3301, 2686, 4698, 4590, 5316, 4532, 5335, 5244, 5217, 5347, 4824, 4186, 1898, 3966, 4887, 5247, 4119, 4146]. Utah [973, 1137, 359, 1503]. Uticularia [439]. utility [1400, 4231, 1113, 1132, 5276]. Utilization [2755, 2863, 3476, 3511, 4441, 1517, 5313, 155, 3918, 3245, 1314]. utilized [2363, 3805]. Utilizing [3164, 2511]. Utricularia [430].

Va. [244]. Valdez [4130]. Valid [2285, 5075]. Validated [4486, 5112, 4064, 3896]. Validation [4846, 5337, 3367, 4399, 4354, 4359, 4844, 3925, 4678, 5356, 4425, 5301, 5005, 4145, 4348, 4311, 3140, 3864, 4711, 4232, 3872, 4905, 4694, 5353, 3760, 5159, 3815, 4162, 5272, 4549, 4000, 4985, 3301, 4237, 4698, 4556, 4680, 4700, 5076, 3393]. Validity [1789, 1857, 4431]. Valley [204]. valuable [57, 94, 4865]. valuation [5121]. Value [3483, 3584, 2575, 3490, 2887, 426, 875, 1806, 1428, 1515, 571, 1387, 4572, 2587, 1367, 1459]. values [4089, 5340, 2436]. Valves [2114]. Vancouver [3457, 2499, 3148, 4919, 3408, 1972, 2438, 3763]. Vaquita [3585]. Varden [5372]. variabiliity [4098]. variabilis [4964, 5280, 4648]. Variability [3234, 3062, 4756, 2823, 2261, 2074, 3521, 3829, 3654, 3668, 4714, 3180, 2071, 2661, 3752, 3445, 3716, 4239, 3079, 1833, 4052, 4884, 4541, 4028, 5304, 2825, 3869, 5074, 4042, 4835, 4975, 4113, 4382, 5006, 4176, 3984, 3883, 5012, 4260, 4262, 5248, 1721, 4717, 4836, 5131, 5259]. Variabilty [4025]. Variable [3506, 3261]. Variables [2655, 4820]. variance [3335, 4031, 4762]. Variation [3032, 4948, 4364, 2722, 2634, 4967, 3706, 3521, 3608, 2860, 4006, 4737, 5378, 3409, 3119, 3415, 3690, 2308, 3395, 3398, 4375, 3663, 2514, 3636, 3175, 3176, 1567, 2828, 4767, 2336, 4753, 2315, 3688, 4845, 3480, 4054, 4310, 4070, 3892, 4138, 4605, 3717, 4563, 3719, 4625, 4442, 4464, 4732, 4415, 3801, 5319, 4637, 4172, 4055, 4695, 4173, 4972, 4823, 3818, 4374, 3903, 4530, 3879, 1902, 3757, 4297, 1340, 4203, 2817, 3845, 3833, 5254, 4116, 4636, 5062, 4817, 5025, 4594, 4703, 5021, 5090, 4272, 4137, 5231, 4365]. Variations [3606, 3646, 3185, 3300, 2163, 5046, 3337, 3674, 3235, 2450, 3143, 2399, 3063, 2012, 3847, 323, 2621, 5243, 4956, 3835, 5270, 306, 3114, 4853]. variegates [968]. variegatus [2057]. Varieties [304, 40]. Various [2183, 2404, 2045, 2115, 2575, 2587, 3098, 1052, 3384, 1225, 214, 313, 248, 264, 329]. Vary [3210]. Varying [3467, 4417, 1681]. vascular [1308]. vectors [4927]. Vegetable [695, 1462]. Vegetated [3156]. vehicle [5070, 5182, 5217, 5172]. vehicles [5349, 5182]. Velocity [2395, 2025, 4499, 1952, 3982, 2541]. velvet [4414, 5257]. Vendean [178]. Venezuela [4609, 4433, 5054, 5164]. Venice [845]. Vent [2652]. ventilatory [3384]. Venting [2779, 5308]. ventless [5286]. ventricosus [5252]. Venus [1803]. verdict [287]. verified [5207]. verify [5055]. Vermilion [2808, 2577, 4060, 4287, 3843, 4137]. verreauxi [4212]. Version [3261]. Versus [3392, 3249, 3490, 4020, 2509, 3151, 3632, 4285, 3702]. Vertebra [3648]. Vertebrae [2346, 5337]. Vertebral [2964, 5272, 2985]. vertebrata [402]. vertebrates [1563, 65]. Vertical [1832, 2202, 3213, 2530, 2484, 2228, 5257, 4503, 2702, 3601, 3615, 3460, 3125, 4022, 2771, 3810, 3698, 2424, 4294, 3189, 3784, 3005, 4851, 3022, 2674, 2526, 4845, 4477, 3874, 5010, 3844, 4750, 5200, 3893, 2813, 1979, 5023, 2725, 2515, 4238, 5220, 1843, 4639, 3958]. Vertically [2851]. very [3122]. Vesicles [2674]. Vessel [2692, 3186, 3378, 3617, 2910, 3267, 2540, 921, 875, 3744, 3842, 3869, 5215, 5168, 5269, 3883, 5241]. vessel-based [5215]. Vessels [3570, 3205, 3154, 395, 3711, 943, 981, 4391, 4870, 425, 3423, 300, 453]. vetula [413]. vetulus [3862, 4094, 3244, 2956, 3476, 2953, 2970, 4065, 4204, 4861, 1951]. via [2405]. viability [4566]. vibrios [2463]. Vibriosis [2786]. vicinity [1443, 1444, 1092, 1498, 1473, 1531, 1584, 1318, 895, 1112, 1160, 250, 911]. Victoria [569, 386]. video [3961, 5119, 5253, 5222, 3937, 4393, 4394, 5086, 5276, 3918]. videotapes [3941]. Vietnam [5374]. view [457, 17, 482, 4980, 680]. villosus [3513, 4768]. Vincent [4251, 4347, 3624, 2052]. vinciguerria [1809]. virens [3023, 4237]. virescens [3552, 4354]. Virgin [3208, 4564]. Virginia [5207, 4516, 3464, 3358, 2374, 2608, 4781, 3596, 1665, 1337, 2279, 3664, 116, 941, 2576, 2921, 3211, 1484, 4139, 3903, 4886, 1659, 5273, 3164, 4240, 4575, 1028]. virginica [1907, 1635, 1680, 4924, 212, 3452]. virtual [4106, 4085]. virus [1722]. visceral [3288]. visit [926]. visitors [558]. Viskan [775]. Visual [1556, 2901, 3196, 3624, 4910, 4846, 5235, 4751]. Visualization [3605]. visually [4393, 4394]. Vital [3210]. Vitality [452, 355, 4947]. vitellins [1968]. vitreum [2928, 3153, 3218, 2928, 3153, 3218]. vitro [4099]. vitta [4800]. vittus [3735, 3809]. vitulina [4629, 2822, 4638, 3245, 4848, 5067, 3065, 4944, 5068, 5168, 3820, 3161, 2856, 4813]. vitulinarichardsi [4239]. Viviparous [1008, 1472, 4234, 3960]. volatile [2483]. volitans [5279]. Volume [4059, 3190, 2419, 2994, 4338, 3482]. volumes [3288]. Volumetric [1392]. Voluntary [2878]. vulgaris [88, 4454, 188, 3031, 2204]. Vulnerability [3324, 3533, 5189, 3414, 4983, 2409, 4888]. vulpes [3048, 4040, 4109, 4195, 3978, 5029]. vulpinus [5331].

W [531]. wahoo [1969, 5043, 5147]. Walbaum [1662, 3091, 3664, 1962, 1647]. Wales [4036, 4212, 3970, 4117, 4126, 4617, 174, 84]. Walleye [3611, 3373, 2712, 3380, 3579, 3586, 2883, 5089, 2741, 3526, 2726, 3871, 5053, 2928, 3977, 3953, 3080, 4672, 3253, 3704, 4795, 3111, 1772, 3608, 4859, 3893, 4737, 3153, 3218, 3901, 3722, 4530, 4275, 4456, 4489, 4341, 4434, 3949, 4663, 4665, 4378, 4703, 5021, 4664]. walleyes [1934]. Walrus [3084, 4304]. Wanapum [2106]. wants [320]. warlike [968]. Warm [3204, 1684, 4922]. warm-blooded [1684]. warm-water [4922]. warming [5311, 5236]. warp [4753, 5131]. Wash [70]. Wash. [250]. Washington [2687, 185, 979, 631, 38, 3444, 3447, 5049, 3382, 3634, 4364, 4669, 4522, 2959, 4073, 5293, 5080, 3408, 3945, 4689, 3331, 2566, 2475, 2463, 3638, 2952, 5068, 1972, 3641, 4585, 5262, 4094, 5233, 1090, 2840, 3498, 5292, 2004, 1668, 3108, 3006, 2274, 2589, 3637, 2322, 2846, 4201, 3398, 2845, 5012, 2820, 2832, 2870, 3938, 3476, 4811, 5366, 5358, 4065, 4204, 3166, 2882, 939, 4592, 2253, 3341, 3894, 2714]. Waste [2503, 2564, 2502, 687, 698]. wastes [1665]. Wastewater [3184]. watched [388]. Water [2839, 2872, 3603, 2374, 3572, 2437, 2472, 2069, 1657, 3204, 2123, 2173, 2011, 2548, 2025, 624, 2973, 2037, 1530, 1784, 1746, 1914, 4623, 2245, 1581, 1528, 433, 1949, 4903, 355, 1614, 1523, 79, 1496, 1510, 1728, 828, 714, 1516, 1677, 686, 2242, 555, 1261, 1619, 1311, 1952, 788, 850, 440, 2860, 1520, 473, 1972, 4584, 2370, 1088, 1075, 1090, 1247, 710, 1148, 1149, 4216, 5175, 1346, 1435, 2000, 1394, 104, 1447, 1309, 2184, 122, 320, 1881, 720, 1031, 1020, 2425, 156, 3887, 1397, 1166, 1094, 466, 734, 1317, 1454, 319, 910, 1980, 3443, 616, 179, 511, 260, 2289, 151, 1476]. water [1522, 4922, 1374, 1104, 4356, 3958, 4603]. Water-beetles [624]. water-gas [1447]. water-pipes [122]. Water-Soluble [3572]. Waterproofing [449]. Waters [2805, 3644, 2618, 3418, 2671, 2851, 2963, 2138, 3315, 2956, 2910, 3604, 2424, 4095, 3607, 3627, 2545, 3252, 2012, 2359, 3021, 3519, 3432, 2674, 4160, 1560, 1744, 2656, 3910, 3923, 3122, 984, 3501, 4190, 1385, 4410, 2895, 4346, 4389, 3929, 3858, 5087, 2474, 5133, 5281, 3434, 4642, 4935, 2869, 1957, 5197, 3048, 5266, 3420, 1613, 3979, 4040, 4109, 4194, 4195, 4461, 4073, 5257, 5222, 5136, 5199, 3886, 1862, 5237, 1675, 2949, 3959, 1106, 1219, 4977, 3967, 663, 446, 489, 1625, 3921, 3983, 4908, 5009, 4619, 145, 3055, 1642, 3823, 1690, 779, 1858, 4077, 4473, 5367, 1720, 2543]. waters [1958, 4151, 940, 3791, 1724, 5327, 4169, 4776, 210, 381, 4153, 2593, 4022, 3019, 5292, 961, 4170, 2351, 3971, 621, 4161, 360, 1388, 3242, 1802, 4818, 3899, 425, 1693, 3491, 1904, 1956, 4512, 5374, 4139, 4570, 1820, 747, 3637, 4405, 4344, 5052, 759, 3732, 2368, 3788, 2260, 2398, 4110, 5248, 4384, 1692, 3039, 4065, 3905, 1076, 1160, 5211, 1035, 1884, 4744, 1989, 3592, 4534, 3767, 3485, 4666, 1964, 3250, 4301, 1197, 5145, 1679, 4207, 5062, 4575, 2382, 3909, 4922, 3955, 4272, 2005, 2248, 4017, 4407, 453, 1407]. waterway [2043]. waterways [1130]. Wave [2836, 5241]. wave-induced [5241]. wavelength [4599]. Waves [3478]. way [540, 34, 289]. Weakfish [2446, 3162, 3642, 4576, 3488, 1470, 3416, 3748, 3896, 3987, 4161, 4101, 4726, 3114, 5203, 3671, 5130]. weakfishes [1600, 1638]. weather [5248]. weather-related [5248]. Web [2934, 3189]. Webbing [2548]. Webs [2158, 2369]. weekly [3972]. Weight [2437, 3702, 3846, 3283, 2120, 2419, 2684, 2778, 3157, 1470, 4933, 4094, 2994, 3957, 2690, 2321, 3857, 550, 5201, 3942, 2599, 3306, 3479, 3633, 2428, 3482, 3897]. Weights [424, 5109]. Weirs [369, 3092]. Weldon [400]. well [319]. West [4498, 1571, 4197, 157, 4613, 57, 966, 4324, 1978, 5096, 5027, 5046, 2596, 3077, 2947, 4714, 2821, 4156, 2251, 1692, 4179, 3889, 1153, 4355, 2043, 4603, 2082, 3420, 2854, 2187, 1337, 4482, 3457, 2525, 2719, 2830, 1126, 4837, 3516]. west-central [4179]. Western [3216, 852, 3472, 2128, 2203, 3594, 3999, 2684, 2778, 2317, 1437, 1853, 1937, 4477, 3977, 1886, 1876, 4413, 4641, 3840, 1577, 4920, 4756, 5379, 3956, 5237, 1079, 1261, 1003, 5324, 5075, 5278, 5338, 4196, 4774, 4606, 4731, 4939, 4079, 4693, 5337, 4528, 941, 4791, 5331, 3957, 2477, 4738, 4800, 4292, 1851, 4154, 1082, 4855, 5148, 2209, 1868, 4144, 4984, 3936, 4295, 4486, 5274, 5272, 4398, 3282, 1669, 1232, 4742, 5309, 3887, 4063, 1687, 4013, 1946, 821, 4011, 2444, 4573, 4699, 4733, 4435, 4949, 4186, 2797, 3949, 2221, 5134, 5092, 5361, 4378, 5021, 3545, 4228, 4840, 4664]. western [5083, 2346, 2915, 2950, 2892, 2016, 2637, 2007, 2232, 2899, 2966, 2712, 2996, 2910, 298, 3002, 3661, 2482, 3316, 3484, 3395, 2244, 3508, 2731, 3419]. westonii [4384]. Wet [2419, 3482]. Wetherall [3667]. Whale [2875, 2072, 2966, 2542, 3087, 3281, 3066, 4807, 371, 4358, 209, 694, 2735, 614, 3683, 5275, 2801, 2711, 2471, 300, 882, 265, 767, 615, 595, 5318, 4301, 3507, 2681, 2639, 3955]. whalebone [3374, 317]. whaler [4617]. Whales [2069, 3277, 3677, 2072, 3422, 29, 3703, 3630, 3021, 3197, 3756, 2303, 3374, 223, 4669, 4404, 4416, 4371, 4293, 4823, 5042, 3973, 4895, 4223, 3776, 3598, 464, 3755, 4134, 5092, 3931, 4752]. whaling [808]. wheeleri [3454, 3827, 3764, 3571]. wheels [2853]. whelk [5084]. whelks [4862]. where [320]. whether [4385]. Which [2016, 3116, 2355, 631, 1560, 680, 381, 461, 3734, 1567, 218]. While [3656]. White [2998, 2072, 3429, 2974, 3019, 3135, 3500, 3533, 2129, 2055, 2989, 4883, 5118, 3759, 2684, 1955, 4801, 2772, 4057, 5197, 4946, 2959, 4141, 4358, 2390, 4693, 4289, 5360, 2659, 4791, 4010, 3683, 4738, 5198, 4609, 3914, 5127, 2955, 4743, 2697, 51, 1579, 2882, 4651, 4224, 1815, 2103, 2095]. white-sided [4946, 2659, 2882]. Whitebone [3553, 3018]. Whitefish [559, 892, 2221, 415, 391, 227, 79, 336, 625, 656, 1026, 1354, 517, 1350, 1903, 9, 43, 55, 731, 481, 1945, 667, 268, 752, 2550, 1904, 818, 1358, 335, 1142, 1663, 730]. whitefish-culture [1026]. whitefishes [1425, 1071, 1503]. Whitemouth [3604, 4567]. whitesided [2797]. whitetip [5220, 4303]. Whiting [2977, 3403, 2984, 3522, 3929, 4151, 3077, 4510, 796]. Whitings [1765, 1638]. who [388, 320, 164, 264]. Whole [2050, 2036, 2038, 4841, 4190, 4346, 4216]. whole-body [4216]. Whole-gear [4841]. wholesomeness [1352]. wickhami [3602]. wide [4793]. Widow [3275, 3013, 4003, 3838, 4270]. Width [3581, 3071, 4933, 4101, 2599, 5131]. widths [4882]. Wilbert [2039]. Wild [2605, 2952, 3187, 2753, 5332, 2451, 2553, 5135, 4950, 3885, 4592, 4181]. will [616]. Willapa [3006]. Willard [101]. William [2631, 4130, 4293, 2259, 1634, 4360, 1623, 4341, 4434, 352]. Williams [3260, 3433]. williamsi [4280]. Willow [110]. Wilson [3060]. wilsoni [4659]. Wind [3523, 5224, 2918]. Windowpane [3606]. windowpanne [3980]. wing [1431]. wing-shell [1431]. Winter [3603, 3420, 3441, 3064, 3026, 2881, 2580, 2220, 4029, 2457, 2381, 4309, 4100, 4182, 4070, 1519, 2764, 4884, 2634, 139, 341, 469, 2951, 4990, 3921, 35, 206, 4368, 3737, 3104, 3789, 4812, 4373, 4566, 4886, 3161, 3891, 4300, 4573, 5205, 934, 3839, 3830]. winter-run [4566]. winter-spawned [4309]. Winter-time [3441]. Winyaw [429]. wire [3731, 5293, 4131]. Wis. [1507]. Wisconsin [1385, 411, 946, 1644, 1489, 1501]. witch [3381, 4432, 3345]. Within [3662, 3199, 3334, 2721, 3425, 4249, 5202, 5079, 4994, 4121, 2125, 2547, 5161, 3701, 425, 4867, 3226, 3964, 5172, 293, 453]. within-estuary [5079]. Within-Season [3199]. Without [2441, 3212]. wolf [3006]. wolffish [5132]. Womens [5100]. wood [1529, 710, 390]. Wooden [423]. Woodhull [111]. Woods [1182, 1443, 1444, 1190, 1226, 1248, 1562, 1473, 1304, 1451, 1178, 1171, 1191, 1436, 1112, 1181, 1310]. Work [1355, 718, 646, 572, 1128, 159, 886, 292, 478, 1192, 512, 600, 398, 765, 867, 909, 711, 1380]. working [562]. works [687, 154]. world [468, 353, 2471, 1908, 1781, 3968, 1025]. worms [1748, 1739]. wosnessenskii [3144]. Wounds [2086, 3363]. wrasse [4746]. wreckfish [5159]. Wyoming [1003, 999, 953, 1819, 53].

x [2492]. xaniurus [4324, 3470]. Xanthid [2915]. Xanthidae [2915, 2737, 2743, 3099]. xanthostigma [3028]. xanthurus [4734, 3781, 3455, 2464, 3059, 3392, 2848, 3389, 3223]. Xeneretmus [3817]. Xenobalanus [4907, 3723]. xenopterus [4303]. Xiphias [4191, 2914, 4349, 4849, 4313, 4606, 4042, 3954, 4064, 3300, 2964, 2985, 4406, 4534, 3767, 2941, 3966, 3247]. xiphiid [4509]. Xiphiidae [3939, 4942]. Xiphoid [2223]. Xiphosura [1751]. Xystreurys [3685].

yaito [1670]. Yakima [2004]. Yaquina [3037, 2288, 2552]. Year [3570, 2011, 4074, 2539, 3439, 4696, 2399, 4557, 4860, 3993, 3759, 2928, 4129, 3437, 364, 448, 1772, 4415, 680, 5360, 4010, 3288, 1897, 3782, 4906, 3591, 5236, 4548, 1158, 887, 4155, 3304, 4003, 3984, 763, 4434, 4667, 4378, 1934]. Year-Class [2539, 4074, 4696, 1772, 1897, 3984]. Year-to-Year [2399]. yearling [4376]. Yearly [3161, 3953, 3159]. Years [2150, 3381, 4719, 388, 4618, 4130, 5275, 4917, 317, 4556, 265, 23, 22, 4840]. yelk [185]. Yellow [3665, 4842, 3929, 1821, 1688, 4216, 1501, 1252, 4815, 986, 4422]. Yellowchin [3366]. yellowedge [4332, 3647]. yelloweye [3814]. Yellowfin [3397, 2028, 2187, 2383, 3109, 3615, 2393, 2214, 2281, 2276, 3484, 3676, 3663, 2431, 2896, 2920, 1783, 5083, 4558, 5212, 4391, 4605, 5355, 3262, 3463, 4367, 3129, 4245, 1724, 1792, 4595, 5373, 2682, 4839, 1693, 4118, 4176, 4397, 3265, 1732, 1875, 4495, 1894, 5043, 1692, 1676, 4007, 4338, 3841, 5241, 4011, 4767, 4116, 2428, 4817]. yellowmouth [4332]. Yellowstone [1819, 999, 953, 972, 956]. Yellowtail [2863, 3058, 3043, 3415, 2578, 2560, 2624, 5085, 5312, 3727, 4286, 2607, 4382, 3288, 3960, 3772, 3940, 1816, 2691, 3889, 4238, 5149, 4270]. Yemen [2267]. Yield [2167, 4120, 4544, 2131, 2571, 2574, 3113, 4418, 2045, 2281, 3362, 3687, 3319, 2059, 2170, 3587, 2938, 2363, 323, 2546, 2871, 2708, 5140, 5374, 306, 3799, 1367]. Yield- [4544]. Yield-per-recruit [4120, 4418, 5140]. Yields [2092, 2576]. Yolk [2863, 2372, 2809, 4028, 1875]. Yolk-Sac [2863, 2372, 2809, 4028, 1875]. yolksac [3746]. York [4557, 2230, 1834, 2763, 2149, 3162, 3891, 4321, 3628, 2656, 1457, 1506, 3858, 2608, 41, 643, 2827, 1665, 2784, 111, 99, 209, 2826, 566, 114, 224, 1797, 3439, 2427, 2727, 2955, 3146, 1698, 3164, 1140, 1824]. Young [1822, 778, 2619, 3632, 2021, 3613, 2504, 2103, 3670, 3591, 3208, 2921, 3211, 619, 3439, 2427, 2727, 2308, 3678, 2244, 750, 2095, 2151, 4557, 4860, 3759, 1032, 1419, 46, 2928, 227, 4129, 10, 1027, 656, 13, 3437, 3036, 1227, 9, 43, 55, 206, 4583, 4584, 5360, 1867, 2029, 3244, 3782, 937, 687, 4548, 4155, 272, 747, 1659, 4488, 1596, 1376, 4337, 394, 658, 705, 238, 4434, 499, 3238, 1401, 4667, 568, 291, 1934]. Young-of-the-Year [3439, 3591, 4860, 3759, 2928, 4129, 3437, 5360, 4548, 4155, 4434, 4667, 1934]. young-of-year [4557, 3782]. YOY [4443]. Yucatan [945, 4778]. Yukon [3480, 1521].

Z [170]. zacentrus [2908]. zachirus [4783, 2543, 2545, 2665]. Zalembius [2613, 4812]. Zalophotrema [2231]. Zalophus [3127, 4625, 3619, 3668, 4713, 5366, 5358, 4745, 4067]. Zameus [5257]. Zaprora [4634]. zapus [3203]. Zealand [2740, 1402, 3876, 72, 241, 4230, 4729, 3580, 443, 4050, 4068, 3541, 4035, 868, 3693, 847, 3714, 4421, 4128, 4322]. Zealanders [86]. Zebrafish [2426]. zero [5057, 4027]. zero-heavy [5057]. zibethicus [460]. zinc [1796, 2794]. Zoanthidea [2518]. Zoarcidae [2611]. Zoeae [2249, 2459, 2589, 2775, 2886, 2922, 3170, 3798]. zoeal [2922, 3144]. Zonation [2836]. zonatus [5029, 3471]. Zone [2593, 2868, 3597, 4997, 4348, 5376, 4003, 4436, 5234, 4666, 3502, 5202, 2399, 3015]. Zones [3555, 3159, 4402, 5177]. Zoogeography [2333, 2676]. Zoological [829, 225]. Zooplankters [2490, 2728]. Zooplankton [2417, 2872, 2562, 2134, 2763, 1734, 1779, 3310, 3522, 2661, 2603, 3094, 2012, 2419, 3258, 4989, 3466, 2653, 3851, 2994, 3829, 4858, 2071, 3482, 1545, 2100]. Zostera [3490, 3350]. zygaena [5225].

References

Anonymous:1881:LPa

[1] Anonymous. List of plates. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/list-plates;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb1%20plates.
pdf.

Ryder:1881:NDS

[2] John A. Ryder. Notes on the development, spinning habits, and structure of the four-spined stickle-back (Apeltes quadracus). Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/notes-
development-spinning-habits-and-structure-four-spined-stickle-
back-apeltes-quadracus;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb1.8.pdf.

Harding:1881:SFC

[3] Garrick M. Harding. Suggestions to fish culturists. Fishery Bulletin, 1(1): ??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/suggestions-
fish-culturists;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb1.16.pdf.

Martin:1881:FAF

[4] S. J. Martin. First appearance of fish at Gloucester, 1881. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/first-appearance-fish-gloucester-1881;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb1.24.pdf.

Martin:1881:NMF

[5] S. J. Martin. Notes on the mackerel fishery of 1881. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/notes-

mackerel-fishery-1881;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb1.32.pdf.

Hemphill:1881:HDG

[6] Henry Hemphill. On the habits and distributions of the geoduck, a clam of the Pacific (Glycimeris generosa, Gld.), with suggestions as to its introduction into the Atlantic Coast of the United States. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/habits-
and-distributions-geoduck-clam-pacific-glycimeris-generosa-
gld-suggestions-its;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb1.40.pdf.

Field:1881:IAO

[7] The Field. Introduction of the aland, or orfe, into England. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
introduction-aland-or-orfe-england;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb1.48.pdf.

Norny:1881:PPR

[8] E. R. Norny. A proposed pond for rearing striped bass (Roccus lineatus) in Delaware Bay. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/proposed-pond-rearing-striped-bass-roccus-
lineatus-delaware-bay;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb1.56.pdf.

Forbes:1881:FYWa

[9] S. A. Forbes. On the food of young whitefish (Coregonus). Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/food-young-whitefish-coregonus;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb1.64.pdf.

Chadwick:1881:DYF

[10] B. P. Chadwick. The destruction of young fish by unsuitable fishing implements. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/destruction-young-fish-unsuitable-fishing-
implements;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb1.72.pdf.

Wilmot:1881:ICS

[11] Samuel Wilmot. Introduction of California salmon into Ontario, with remarks on the disappearance of Maine salmon from that Province. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/introduction-california-salmon-ontario-remarks-
disappearance-maine-salmon-province;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb1.80.pdf.

Ryder:1881:CCN

[12] Francis W. Ryder. Codfish caught near Cape Charles, mouth of Chesapeake Bay, in 1834. Fishery Bulletin, 1(1):??, 1881. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/codfish-caught-near-cape-charles-
mouth-chesapeake-bay-183;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb1.88.pdf.

Collins:1881:ICY

[13] J. W. Collins. An inquiry as to the capture of young codfish in Chesapeake Bay. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/inquiry-capture-young-codfish-chesapeake-bay;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb1.96.pdf.

Ryder:1881:AOR

[14] John A. Ryder. Additional observations on the retardation of the development of the ova of the shad. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/additional-
observations-retardation-development-ova-shad;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb1.
104.pdf.

Collins:1881:GNC

[15] J. W. Collins. Gill-nets in the cod-fishery; a description of the Norwegian cod-nets, with directions for their use, and a history of their introduction into the United States. Fishery Bulletin, 1 (1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/gill-
nets-cod-fishery-description-norwegian-cod-nets-directions-
their-use-and-history-their;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb1.1.pdf.

Pasco:1881:CCN

[16] I. D. Pasco. A call for carp from Nevada. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/call-carp-nevada;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb1.9.pdf.

Haack:1881:GVA

[17] ?. Haack. A German view of the American section in the Berlin Fishery Exhibition. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/german-view-american-section-berlin-fishery-
exhibition;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb1.17.pdf.

Norny:1881:PSB

[18] E. R. Norny. On the propagation of the striped bass. Fishery Bulletin, 1 (1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/propagation-
striped-bass;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb1.25.pdf.

Martin:1881:NNEa

[19] S. J. Martin. Notes on New England fisheries. Fishery Bulletin, 1 (1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/notes-new-
england-fisheries;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb1.33.pdf.

Martin:1881:NNEb

[20] S. J. Martin. Notes on New England food-fishes. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/notes-
new-england-food-fishes;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb1.41.pdf.

Finsch:1881:RTG

[21] O. Finsch. Report on a trip in Germany to secure carp for the United States Fish Commission. Fishery Bulletin, 1(1):??, 1881. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/report-trip-germany-secure-carp-
united-states-fish-commission;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb1.49.pdf.

Willis:1881:SFS

[22] H. Willis. Shad fisheries of the Susquehanna River fifty-six years ago. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/shad-fisheries-susquehanna-river-fifty-six-years-ago;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb1.57.pdf.

USFC:1881:SRA

[23] United States Fish Commission. Some results of the artificial propagation of Maine and California salmon in New England and Canada, recorded in the years 1879 and 1880. Fishery Bulletin, 1 (1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/some-
results-artificial-propagation-maine-and-california-salmon-
new-england-and-canada;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb1.65.pdf.

Wilson:1881:PIC

[24] Thomas Wilson. The proposed introduction of catfish into Ghent. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/proposed-introduction-catfish-ghent;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb1.
73.pdf.

Hughes:1881:LPF

[25] Smith E. Hughes. Live ponds for fish in New Jersey. Fishery Bulletin, 1 (1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/live-ponds-
fish-new-jersey;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb1.81.pdf.

Banditten:1881:RCP

[26] Eben Banditten. Recent contributions to pond cultivation. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/recent-contributions-pond-cultivation;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb1.89.pdf.

Wigg:1881:IGC

[27] George Wigg. On the insensibility of the German carp to freezing. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
insensibility-german-carp-freezing;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb1.97.pdf.

Barnes:1881:GSG

[28] A. G. Barnes. Growth and spawning of German carp in Alabama. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/growth-and-spawning-german-carp-alabama;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb1.105.pdf.

Jacobson:1881:PW

[29] Herman (Translator) Jacobson. Protection of whales. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/protection-whales;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb.2.pdf.

Goode:1881:CFU

[30] G. Brown Goode. The carangoid fishes of the United States — pompanos, crevallés, amber-fish, etc. Fishery Bulletin, 1(1):??, 1881. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/carangoid-fishes-united-states-
pompanos-crevall%C3%A9s-amber-fish-e;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb1.10.pdf.

OF:1881:PBF

[31] Oesterreichisch-Ungarische Fischerei-Zeitung. Peat-bogs as fish-ponds. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/peat-bogs-fish-ponds;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb1.18.pdf.

Wright:1881:GCP

[32] Abel A. Wright. A Georgia carp pond. Fishery Bulletin, 1(1): ??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/georgia-
carp-pond;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb1.26.pdf.

Stone:1881:MMR

[33] Livingston Stone. Mortality of McCloud River salmon in 1881. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/mortality-mccloud-river-salmon-1881;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb1.
34.pdf.

Throckmorton:1881:DFW

[34] S. R. Throckmorton. Description of the fish-way in Pitt River, California. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/description-fish-way-pitt-river-
california;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb1.42.pdf.

Goode:1881:WHF

[35] G. Brown Goode and Capt. J. W. Collins. The winter haddock fishery of New England. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/winter-haddock-fishery-new-england;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb1.50.pdf.

Martin:1881:NGFa

[36] S. J. Martin. Notes on the Gloucester fishery. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/notes-
gloucester-fishery;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb1.58.pdf.

Mordecai:1881:FSA

[37] E. R. Mordecai. Food of the shad of the Atlantic Coast of the United States (Alosa præstabilis De Kay), and the functions of the pyloric cœca. Fishery Bulletin, 1(1):277–282, 1881. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/food-shad-atlantic-coast-united-
states-alosa-pr%C3%A6stabilis-de-kay-and-functions-pyloric-
c%C3%A6ca;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb1.66.pdf.

Armistead:1881:TLC

[38] A. Wilson Armistead. A transfer of leather carp (Cyprinus carpio) from the government ponds at Washington, USA, to Scotland. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
transfer-leather-carp-cyprinus-carpio-government-ponds-washington-
usa-scotland;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb1.74.pdf.

Osborn:1881:NSC

[39] J. H. Osborn. Notes and suggestions concerning the Florida shad fishery. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/notes-and-suggestions-concerning-
florida-shad-fishery;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb1.82.pdf.

Zentz:1881:RVC

[40] F. Zentz. On the races or varieties of carp, denying the existence of blue carp and gold carp. Fishery Bulletin, 1(1):??, 1881. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/races-or-varieties-carp-denying-
existence-blue-carp-and-gold-carp;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb1.90.pdf.

Collins:1881:FAM

[41] J. W. Collins. First arrival of mackerel in New York in the spring of 1882. Fishery Bulletin, 1(1):??, 1881. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/first-arrival-mackerel-new-york-
spring-1882;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb1.98.pdf.

Martin:1881:NFGa

[42] S. J. Martin. Notes on the fisheries of Gloucester, Massachusetts. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/notes-
fisheries-gloucester-massachusetts-0;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb1.106.pdf.

Forbes:1881:FYWb

[43] S. A. Forbes. The food of young whitefish (Coregonus clupeiformis). Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
food-young-whitefish-coregonus-clupeiformis;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb1.
3.pdf.

Campbell:1881:NMR

[44] J. B. Campbell. Notes on McCloud River, California, and some of its fishes. Fishery Bulletin, 1(1):??, 1881. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/notes-mccloud-river-california-
and-some-its-fishes;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb1.11.pdf.

Weddige:1881:CF

[45] Herr. Weddige. Castrating fish. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/castrating-fish;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb1.
19.pdf.

Bean:1881:MYA

[46] Tarleton H. Bean. Movements of young alewives (?) (Pomolobus, sp.) in Colorado River, Texas. Fishery Bulletin, 1(1):??, 1881. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/movements-young-alewives-
pomolobus-sp-colorado-river-texas;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb1.27.pdf.

Hinckley:1881:FEC

[47] Isaac Hinckley. The fish-eating cows of provincetown, mass. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/fish-
eating-cows-provincetown-mass;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb1.35.pdf.

Bean:1881:NSU

[48] Tarleton H. Bean. Notes on a shipment, by the United States Fish Commission, of California salmon (Oncorhynchus chouicha) to Tanner’s Creek, Indiana, in 1876. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic).  URL https://spo.nmfs.noaa.gov/content/notes-
shipment-united-states-fish-commission-california-salmon-oncorhynchus-

chouicha;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/ffb1.43.pdf.

Ryder:1881:PPC

[49] John A. Ryder. The protozoa and protophytes considered as the primary or indirect source of the food of fishes. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/protozoa-and-protophytes-
considered-primary-or-indirect-source-food-fishes;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb1.51.pdf.

Martin:1881:CFG

[50] S. J. Martin. Cod-fishing with gill-nets in Ipswich Bay, Massachusetts. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/cod-fishing-gill-nets-ipswich-bay-massachusetts;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb1.59.pdf.

Ryder:1881:MEW

[51] John A. Ryder. The micropyle of the egg of the white perch. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/micropyle-egg-white-perch;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb1.67.pdf.

Martin:1881:PFB

[52] S. J. Martin. Pollock-fishing in Boston Bay. Fishery Bulletin, 1 (1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/pollock-
fishing-boston-bay;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb1.75.pdf.

Wright:1881:RCW

[53] Harrison Wright. Report of a committee of the Wyoming Historical and Geological Society on the early shad fisheries of the north branch of the Susquehanna River. Fishery Bulletin, 1(1):??, 1881. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/report-committee-wyoming-
historical-and-geological-society-early-shad-fisheries-north;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb1.83.pdf.

Eckhardt:1881:PBC

[54] ?. Eckhardt. The peculiarities of blue carp. Fishery Bulletin, 1(1): ??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/peculiarities-
blue-carp;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb1.91.pdf.

Forbes:1881:IFF

[55] S. A. Forbes. An inquiry into the first food of young lake whitefish (Coregonus clupeiformis). Fishery Bulletin, 1(1):??, 1881. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/inquiry-first-food-
young-lake-whitefish-coregonus-clupeiformis;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb1.
99.pdf.

Gilbert:1881:GSS

[56] W. L. Gilbert. Growth of the Sälbling (Salvo salvelinus) in the old colony trout ponds at Plymouth, Massachusetts. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/growth-s%C3%A4lbling-salvo-salvelinus-old-
colony-trout-ponds-plymouth-massachusetts;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb1.107.
pdf.

Ryder:1881:VEM

[57] John A. Ryder. A valuable edible mollusk of the west coast. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
valuable-edible-mollusk-west-coast;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb1.4.pdf.

Deblois:1881:OMI

[58] E. T. Deblois. The origin of the menhaden industry. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/origin-
menhaden-industry;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb1.12.pdf.

Throckmorton:1881:ISB

[59] S. R. Throckmorton. The introduction of striped bass into Cal-

ifornia.  Fishery Bulletin, 1(1):??, 1881.  CODEN FSYBAY.

ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/introduction-striped-bass-california;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb1.20.pdf.

Goode:1881:NLH

[60] G. Brown Goode. Notes on the life-history of the eel, chiefly derived from a study of recent European authorities. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/notes-life-history-
eel-chiefly-derived-study-recent-european-authorities;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb1.28.pdf.

Ryder:1881:DSM

[61] John A. Ryder. Development of the Spanish mackerel (Cybium maculatum). Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/development-spanish-mackerel-cybium-maculatum;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb1.36.pdf.

Bean:1881:ASU

[62] Tarleton H. Bean. Account of a shipment, by the United States Fish Commission, of California salmon-fry (Oncorhynchus chouicha) to southern Louisiana, with a note on some collections made at Tickfaw. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/account-shipment-united-states-fish-commission-
california-salmon-fry-oncorhynchus-chouicha;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb1.
44.pdf.

Fisher:1881:FFC

[63] F. B. Fisher. Fishing and fish-culture in Florida. Fishery Bulletin, 1(1): ??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/fishing-and-
fish-culture-florida;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb1.52.pdf.

Hessel:1881:ACM

[64] Rudolph Hessel. Artificial culture of medicinal leeches and species of helix. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY.

ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/artificial-culture-medicinal-leeches-
and-species-helix;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb1.60.pdf.

Ryder:1881:DSG

[65] John A. Ryder. Development of the silver gar (Belone longirostris), with observations on the genesis of the blood in embryo fishes, and a comparison of fish ova with those of other vertebrates. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/development-silver-
gar-belone-longirostris-observations-genesis-blood-embryo-fishes-
and;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb1.68.pdf.

Martin:1881:CGNa

[66] S. J. Martin. Cod gill-nets in Ipswich Bay, Massachusetts. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/cod-gill-nets-ipswich-bay-massachusetts;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb1.76.pdf.

Wilmot:1881:RSM

[67] Samuel Wilmot. Remarks on the scarcity of male and grilse salmon in the rivers of Ontario, Canada. Fishery Bulletin, 1(1):??, 1881. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/remarks-scarcity-male-and-grilse-
salmon-rivers-ontario-canada;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb1.84.pdf.

Leonard:1881:PIF

[68] J. A. Leonard. Plan of the International Fisheries Exhibition to be held at Edinburgh, Scotland, and a list of prizes to be awarded. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/plan-
international-fisheries-exhibition-be-held-edinburgh-scotland-
and-list-prizes-be;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb1.92.pdf.

Ryder:1881:NBF

[69] John A. Ryder. Notes on the breeding, food, and green color of the oyster. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/notes-breeding-food-and-green-color-oyster;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb1.100.pdf.

Harding:1881:ICP

[70] Charles W. Harding. Inquiries concerning the propagation of American smelt and shad, and notes on the fisheries of the Wash of England. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/inquiries-concerning-propagation-american-
smelt-and-shad-and-notes-fisheries-wash-england;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb1.108.pdf.

Ryder:1881:PNM

[71] John A. Ryder. Preliminary notice of the more important scientific results obtained from a study of the embryology of fishes. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
preliminary-notice-more-important-scientific-results-obtained-
study-embryology-fishes;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb1.5.pdf.

Creighton:1881:FCN

[72] R. J. Creighton. Fish-culture in New Zealand. Fishery Bulletin, 1(1): ??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/fish-culture-
new-zealand;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb1.13.pdf.

Clark:1881:SP

[73] Frank N. Clark. The self-picker. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/self-picker;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb1.21.pdf.

Hobbs:1881:LOR

[74] Orlando Hobbs. A list of Ohio River fishes sold in the markets. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/list-ohio-river-fishes-sold-markets;   https://spo.

nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb1.
29.pdf.

Worth:1881:APS

[75] S. G. Worth. The artificial propagation of the striped bass (Roccus lineatus) in Albemarle Sound. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/artificial-
propagation-striped-bass-roccus-lineatus-albemarle-sound;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb1.37.pdf.

DF:1881:PCT

[76] Deutsche Fisherei-Zeitung. Pachaly’s car for transporting fish. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/pachalys-car-transporting-fish;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb1.45.pdf.

Smiley:1881:CFG

[77] Chas W. Smiley. Changes in the fisheries of the Great Lakes during the decade 1870–1880. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/changes-fisheries-great-lakes-during-
decade-1870-1880;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb1.53.pdf.

Martin:1881:FNGa

[78] S. J. Martin. Fishery news from Gloucester, Massachusetts. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fishery-news-gloucester-massachusetts;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb1.61.pdf.

Clark:1881:RWS

[79] Frank N. Clark. On the rearing of whitefish in spring water, and its relation to their subsequent distribution. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/rearing-whitefish-spring-
water-and-its-relation-their-subsequent-distribution;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb1.69.pdf.

vondemBorne:1881:DTF

[80] Max von dem Borne. Dry transmission of fish eggs. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/dry-transmission-fish-eggs;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb1.77.pdf.

Martin:1881:FNGb

[81] S. J. Martin. Fishery notes from Gloucester, Massachusetts. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishery-notes-gloucester-massachusetts;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb1.85.pdf.

Martin:1881:FNGc

[82] S. J. Martin. Fishery notes from Gloucester, Massachusetts, concerning cod, haddock, herring, halibut, and mackerel. Fishery Bulletin, 1(1): ??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/fishery-notes-
gloucester-massachusetts-concerning-cod-haddock-herring-halibut-
and-mackerel;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb1.93.pdf.

Martin:1881:NFGb

[83] S. J. Martin. Notes on the fisheries of Gloucester, Massachusetts. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/notes-fisheries-gloucester-massachusetts;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb1.101.pdf.

Walpole:1881:DAS

[84] S. Walpole and T. H. Huxley. Disease among the salmon of many rivers of England and Wales. Fishery Bulletin, 1(1):??, 1881. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/disease-among-salmon-
many-rivers-england-and-wales;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb1.109.pdf.

Green:1881:RCM

[85] Seth Green. Rearing of California mountain trout (Salmo irideus). Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
rearing-california-mountain-trout-salmo-irideus;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb1.6.pdf.

Leslie:1881:ABA

[86] Frank Leslie. American birds, animals, and fishes for New Zealanders. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/american-birds-animals-and-fishes-new-zealanders;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb1.14.pdf.

Martin:1881:CAM

[87] S. J. Martin. Coal ashes as a means of raising mackerel in purseseines. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/coal-ashes-means-raising-mackerel-purse-seines;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb1.22.pdf.

Hermes:1881:MMS

[88] Otto Hermes. On the mature males sexual organs of the congereel (Conger vulgaris), with some observations on the male of the common eel (Anguilla vulgaris, Fleming). Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/mature-males-
sexual-organs-conger-eel-conger-vulgaris-some-observations-
male-common-eel;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb1.30.pdf.

Ryder:1881:RDO

[89] John A. Ryder. On the retardation of the development of the ova of the shad (Alosa sapidissima), with observations on the egg-fungus and bacteria. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/retardation-development-ova-shad-alosa-sapidissima-
observations-egg-fungus-and-bacteria;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb1.38.pdf.

Anonymous:1881:MSR

[90] Anonymous. Memorandum of some results of the artificial propagation and planting of fish, due mainly to the efforts of the United States Fish Commission. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/memorandum-some-results-artificial-propagation-
and-planting-fish-due-mainly-efforts-united;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb1.
46.pdf.

Marquette:1881:EFP

[91] George R. Marquette. Extraordinary floods in the Potomac River. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
extraordinary-floods-potomac-river;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb1.54.pdf.

Hawkins:1881:ORI

[92] J. W. Hawkins. An opinion regarding the influence upon the coast fisheries of the steamers used in the menhaden fishery. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/opinion-
regarding-influence-upon-coast-fisheries-steamers-used-menhaden-
fishery;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb1.62.pdf.

Jordan:1881:DNN

[93] David S. Jordan and Charles H. Gilbert. Descriptions of nineteen new species of fishes from the Bay of Panama. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/descriptions-
nineteen-new-species-fishes-bay-panama;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb1.70.pdf.

Swetitsch:1881:DEE

[94] Joseph Swetitsch. A depot for embryonated eggs of all the valuable kinds of fish. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/depot-embryonated-eggs-all-valuable-kinds-fish;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb1.78.pdf.

Hiester:1881:FCP

[95] Charles E. Hiester. Fish culture for profit. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/fish-culture-profit;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb1.86.pdf.

GDN:1881:PRF

[96] Galveston Daily News. Progress and results of fish culture. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
progress-and-results-fish-culture;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb1.94.pdf.

Laighton:1881:CSI

[97] Cedric Laighton. The capture of shad at Isles of Shoals, New Hampshire. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/capture-shad-isles-shoals-new-hampshire;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb1.102.pdf.

Anonymous:1881:I

[98] Anonymous. Index. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/index-39;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb1.index_.pdf.

Green:1881:SCG

[99] Seth Green. Salmon caught in Genesee River, New York. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/salmon-caught-genesee-river-new-york;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb1.7.pdf.

Shears:1881:CHR

[100] E. E. Shears. Carp in the Hudson River. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/carp-hudson-
river;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb1.15.pdf.

Martin:1881:MUW

[101] S. J. Martin. Methods of using Willard’s patent pocket for mackerel. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY.

ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/methods-using-willards-patent-pocket-
mackerel;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb1.23.pdf.

DS:1881:OPR

[102] Department of State. Official papers relating to the proposed International Fisheries Exhibition at Edinburgh in 1882. Fishery Bulletin, 1 (1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/official-
papers-relating-proposed-international-fisheries-exhibition-
edinburgh-1882;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb1.31.pdf.

Ryder:1881:CDM

[103] John A. Ryder. A contribution to the development and morphology of the Lophobranchiates (Hippocampus antiquorum, the sea-horse). Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
contribution-development-and-morphology-lophobranchiates-hippocampus-
antiquorum-sea-horse;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb1.39.pdf.

McDonald:1881:ETG

[104] Marshall McDonald. Experiments in the transportation of the German carp in a limited supply of water. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/experiments-
transportation-german-carp-limited-supply-water;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb1.47.pdf.

Krause:1881:CHF

[105] ?. Krause. Cod and halibut fisheries near the Shumagin Islands. Fishery Bulletin, 1(1):259–260, 1881. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
cod-and-halibut-fisheries-near-shumagin-islands;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb1.55.pdf.

Martin:1881:NGFb

[106] S. J. Martin. Notes on the Gloucester fisheries. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/notes-
gloucester-fisheries;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb1.63.pdf.

Ryder:1881:NCF

[107] John A. Ryder. On the nuclear cleavage-figures developed during the segmentation of the germinal disk of the egg of the salmon. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/nuclear-cleavage-figures-developed-during-
segmentation-germinal-disk-egg-salmon;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb1.71.pdf.

vonBehr:1881:TFE

[108] ?. von Behr. Treatment of fish eggs at sea. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
treatment-fish-eggs-sea;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb1.79.pdf.

Martin:1881:CGNb

[109] S. J. Martin. Codfishing with gill-nets from Gloucester, Massachusetts. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/codfishing-gill-nets-gloucester-
massachusetts;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb1.87.pdf.

Milner:1881:SFR

[110] J. W. Milner. Summary of fishing records, or shad and alewives, kept at Willow Branch Fishery, North Carolina, from 1835 to 1874. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
summary-fishing-records-or-shad-and-alewives-kept-willow-branch-
fishery-north-carolina-1835;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb1.95.pdf.

Green:1881:ILL

[111] Seth Green. The introduction of land-locked salmon into Woodhull Lake, New York, and the subsequent capture of some of them. Fishery Bulletin, 1(1):??, 1881. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
introduction-land-locked-salmon-woodhull-lake-new-york-and-

subsequent-capture-some-them;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb1.103.pdf.

Anonymous:1881:LPb

[112] Anonymous. List of plates. Fishery Bulletin, 1(1):??, 1881. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/list-plates-0;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
plates.pdf.

Anonymous:1882:P

[113] Anonymous. Plates. Fishery Bulletin, 2(1):??, 1882. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/plates;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb2%20plates.
pdf.

Morgan:1882:CLL

[114] George Morgan. Capture of land-locked salmon at Oswego, New York. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/capture-land-locked-salmon-oswego-new-york;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb2.8.pdf.

Morriss:1882:SCS

[115] T. W. Morriss. Spawning of carp, in a small basin, at Brenham, Texas. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/spawning-carp-small-basin-brenham-texas;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb2.16.pdf.

Healy:1882:SBP

[116] R. Healy. Striped bass in Piankatank River, Virginia. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
striped-bass-piankatank-river-virginia;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb2.24.pdf.

Berichte:1882:TLF

[117] Amtliche Berichte. Transportation of live fish. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/transportation-live-
fish;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb2.32.pdf.

Arnold:1882:SPB

[118] Isaac Arnold, Jr. Successful propagation of black bass. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
successful-propagation-black-bass;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb2.40.pdf.

Ryder:1882:PNS

[119] John A. Ryder. Preliminary notice of some points in the minute anatomy of the oyster. Fishery Bulletin, 2(1):??, 1882. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/preliminary-notice-
some-points-minute-anatomy-oyster;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb2.48.pdf.

Horst:1882:COK

[120] R Horst. A contribution to our knowledge of the development of the oyster (Ostrea edulis L.). Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/contribution-our-knowledge-development-
oyster-ostrea-edulis-l;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb2.56.pdf.

Tanner:1882:OSM

[121] Z. L. Tanner. Occurrence of fur seal in mid-ocean. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
occurrence-fur-seal-mid-ocean;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb2.64.pdf.

Nye:1882:EAR

[122] Willard Nye, Jr. Eels (Anguilla rostrata) in New Bedford water-pipes — mackerel abundant in Amherst River. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/eels-anguilla-
rostrata-new-bedford-water-pipes-mackerel-abundant-amherst-
river;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb2.72.pdf.

Yost:1882:RUM

[123] John Yost. Report upon the management of German carp by a Mississippi correspondent. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/report-upon-management-german-carp-mississippi-
correspondent;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb2.80.pdf.

Nussbaum:1882:STL

[124] M. Nussbaum. A simple test to learn if fish-ova are impregnated. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
simple-test-learn-if-fish-ova-are-impregnated;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb2.
88.pdf.

Baird:1882:IOF

[125] Spencer F. Baird. Inducements offered fishermen to furnish shad eggs for the u. s. commission of fish and fisheries. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/inducements-offered-
fishermen-furnish-shad-eggs-u-s-commission-fish-and-fisheries;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb2.96.pdf.

MZ:1882:SSH

[126] Magdeburische Zeitung. Stocking the Stettiner Haff with carp. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
stocking-stettiner-haff-carp;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb2.1.pdf.

Martin:1882:NFGa

[127] S. J. Martin. Notes on the fisheries of Gloucester, Massachusetts. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/notes-
fisheries-gloucester-massachusetts-1;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb2.9.pdf.

Yoakum:1882:CCT

[128] F. L. Yoakum. Carp culture in Texas. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/carp-culture-
texas;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb2.17.pdf.

Hiester:1882:AQR

[129] Charles E. Hiester. Answers to questions relative to catfish. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
answers-questions-relative-catfish;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb2.25.pdf.

Harvey:1882:FCF

[130] M. Harvey. Fishing and consumption of fish in Newfoundland. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/fishing-and-consumption-fish-newfoundland;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb2.33.pdf.

Smiley:1882:RBI

[131] Chas W. Smiley. Removal of bass from Indiana to North Carolina by the United States Fish Commission. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/removal-bass-indiana-
north-carolina-united-states-fish-commission;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb2.
41.pdf.

Tidende:1882:CCN

[132] Fiskeri Tidende. The codfisheries on the coast of Norway. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/codfisheries-coast-norway;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb2.49.pdf.

Eckhardt:1882:CPB

[133] George Eckhardt. Carp propagation and blue carp. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/carp-propagation-and-blue-carp;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb2.57.pdf.

Stone:1882:AOM

[134] Livingston Stone. Account of operations at the McCloud River fishbreeding stations of the United States Fish Commission, from 1872 to 1882, inclusive. Fishery Bulletin, 2(1):??, 1882. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/account-operations-mccloud-river-
fish-breeding-stations-united-states-fish-commission-1872;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb2.65.pdf.

Collins:1882:NMH

[135] J. W. Collins. Notes on the movements, habits, and captures of mackerel for the season of 1882. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic).  URL https://spo.nmfs.noaa.gov/content/notes-
movements-habits-and-captures-mackerel-season-1882;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb2.73.pdf.

Collins:1882:NHFa

[136] J. W. Collins. Notes on the halibut fisheries of 1881–1882. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
notes-halibut-fisheries-1881-1882;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb2.81.pdf.

Goode:1882:NLP

[137] G. Brown Goode. Notes on the lampreys — petromyzontidæ. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/notes-
lampreys-petromyzontid%C3%A6;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb2.89.pdf.

Gaines:1882:GSD

[138] Kemp Gaines. Growth, spawning, and distribution of German carp reared from 20 fish furnished by the U. S. Fish Commission, November 29, 1880. Fishery Bulletin, 2(1):??, 1882. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/growth-spawning-and-distribution-
german-carp-reared-20-fish-furnished-u-s-fish-commission;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb2.97.pdf.

Collins:1882:ADF

[139] J. W. Collins. Appearance of dog-fish (Squalus acanthias) on the New England coast in winter. Fishery Bulletin, 2(1):??, 1882. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/appearance-dog-fish-
squalus-acanthias-new-england-coast-winter;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb2.2.pdf.

Heyser:1882:GMC

[140] E. Heyser. Growth of mirror carp sent to Georgia by the United States Fish Commission. Fishery Bulletin, 2(1):??, 1882. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/growth-mirror-carp-sent-georgia-
united-states-fish-commissio;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb2.10.pdf.

Smiley:1882:RPS

[141] Chas W. Smiley. Result of planting shad in the Muskingum River. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/result-planting-shad-muskingum-river;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb2.18.pdf.

Leonard:1882:REF

[142] J. Leonard. A report on the Edinburgh Fisheries Exhibition. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/report-edinburgh-fisheries-exhibition;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb2.26.pdf.

Farmer:1882:CAM

[143] Scientific Farmer. Connection of abundance of moss and of black flies with abundance of trout. Fishery Bulletin, 2(1):??, 1882. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/connection-abundance-
moss-and-black-flies-abundance-trout;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb2.34.pdf.

Tidende:1882:SHF

[144] Fiskeri Tidende. The Scotch herring fisheries in 1882. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
scotch-herring-fisheries-1882;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb2.42.pdf.

GT:1882:FSS

[145] Gloucester Telegraph. Fishing for shad in South American waters. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishing-shad-south-american-waters;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb2.50.pdf.

Nicklas:1882:FCI

[146] Carl Nicklas. Feeding carp with Indian corn. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/feeding-carp-indian-
corn;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb2.58.pdf.

vonBehr:1882:FAS

[147] ?. von Behr. Five American salmonidæ in Germany. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/five-american-salmonid%C3%A6-germany;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb2.66.pdf.

Hayward:1882:CFN

[148] Jos S. Hayward. Catch of fish in Newfoundland and Labrador in 1881. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/catch-fish-newfoundland-and-labrador-1881;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb2.74.pdf.

Davis:1882:GRP

[149] H. B. Davis. Gratifying results of propagating German carp — bream and carp in ponds together — table qualities of carp. Fishery Bulletin, 2(1): ??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/gratifying-
results-propagating-german-carp-bream-and-carp-ponds-together-
table-qualities;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb2.82.pdf.

Clark:1882:AON

[150] Frank N. Clark. Account of operations at the northville fish hatching station of the United States Fish Commission, from 1874 to 1882 inclusive. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/account-operations-northville-fish-hatching-station-
united-states-fish-commission-1874-1882;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb2.90.pdf.

Wilson:1882:RBT

[151] Charles A. Wilson. Raising brook trout in mineral water. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
raising-brook-trout-mineral-water;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb2.98.pdf.

Stone:1882:RQH

[152] Livingston Stone. Replies to questions of Herr von Behr concerning Salvelinus fontinalis and Salmo irideua. Fishery Bulletin, 2(1): ??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/replies-
questions-herr-von-behr-concerning-salvelinus-fontinalis-and-
salmo-irideua;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb2.3.pdf.

Wood:1882:FGM

[153] M. L. Wood. The fisheries of the Gulf of Mexico. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fisheries-gulf-mexico;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb2.11.pdf.

Swett:1882:DFS

[154] S. B. Swett. Decrease of fish in the Squamscot River, New Hampshire, on account of refuse matter from gas works. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
decrease-fish-squamscot-river-new-hampshire-account-refuse-
matter-gas-works;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb2.19.pdf.

Harding:1882:ULN

[155] Charles W. Harding. The utilization of localities in Norfolk and Suffolk suitable for the cultivation of mussels and other shell fish. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
utilization-localities-norfolk-and-suffolk-suitable-cultivation-
mussels-and-other-shell;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb2.27.pdf.

ST:1882:PWG

[156] Sunland Tribune. Poisoned water in the Gulf of Mexico. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/poisoned-water-gulf-mexico;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb2.35.pdf.

Poey:1882:LFF

[157] Felipe Poey. List of food-fishes brought from Key West, Fla., into the markets of Havana. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/list-food-fishes-brought-
key-west-fla-markets-havana;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb2.43.pdf.

Rathbun:1882:NSP

[158] Richard Rathbun. Notes on the shrip and prawn fisheries of the United States. Fishery Bulletin, 2(1):??, 1882. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/notes-shrip-and-prawn-fisheries-
united-states;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb2.51.pdf.

Goode:1882:FDU

[159] G. Brown Goode. The first decade of the United States Fish Commission: its plan of work and accomplished results, scientific and economical. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/first-decade-united-states-fish-commission-its-
plan-work-and-accomplished-results-0;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb2.59.pdf.

Gallup:1882:CDF

[160] George Gallup. Catching dog-fish for oil and guano. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/catching-dog-fish-oil-and-guano;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb2.
67.pdf.

Collins:1882:NHFb

[161] J. W. Collins. Notes on the herring fishery of Massachusetts Bay in the autumn of 1882. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/notes-herring-fishery-massachusetts-bay-
autumn-1882;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb2.75.pdf.

Bouchon-Brandely:1882:RRG

[162] G. Bouchon-Brandely. Report relative to the generation and artificial fecundation of oysters, addressed to the Minister of the Marine and the Colonies. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/report-relative-generation-and-artificial-
fecundation-oysters-addressed-minister-marine-and.

Atkins:1882:SPS

[163] Charles G. Atkins. Sketch of the Penobscot salmon-breeding establishment. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/sketch-penobscot-salmon-breeding-establishment;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb2.91.pdf.

Smiley:1882:GCP

[164] Chas W. Smiley. A geographical catalogue of persons who have stated that they are interested in fish culture. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/geographical-
catalogue-persons-who-have-stated-they-are-interested-fish-
culture;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb2.99.pdf.

Griffith:1882:RPS

[165] William Griffith. Result of planting shad in the Ohio River. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/result-planting-shad-ohio-river;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb2.
4.pdf.

Stearns:1882:STC

[166] R. E. C. Stearns. Suggestions for transplanting clams from the Pacific Ocean to the Atlantic. Fishery Bulletin, 2(1):??, 1882. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/suggestions-transplanting-clams-
pacific-ocean-atlantic;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb2.12.pdf.

Jordan:1882:RSF

[167] David S. Jordan and Charles H. Gilbert. A review of the siluroid fishes found on the Pacific Coast of Tropical America, with descriptions of three new species. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/review-siluroid-fishes-found-pacific-
coast-tropical-america-descriptions-three-new-species;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb2.20.pdf.

Martin:1882:MCM

[168] S. J. Martin. Movements and catch of mackerel. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/movements-and-catch-
mackerel;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb2.28.pdf.

Jordan:1882:LFCa

[169] David S. Jordan and Charles H. Gilbert. List of fishes collected at Mazatlan, Mexico, by Charles H. Gilbert. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/list-fishes-
collected-mazatlan-mexico-charles-h-gilbert;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb2.
36.pdf.

Rathbun:1882:DSU

[170] Richard Rathbun. Dredging stations of the United States Fish Commission steamer Fish Hawk, Lieut. Z. L. Tanner, commanding, for 1880, 1881, and 1882, with temperature and other observations. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/dredging-stations-unites-states-fish-
commission-steamer-fish-hawk-lieut-z-l-tanner;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb2.
44.pdf.

Swan:1882:SPS

[171] James G. Swan. Shad in Puget Sound. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/shad-puget-
sound;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb2.52.pdf.

Leslie:1882:SBS

[172] Chas C. Leslie. Scarcity of the blackfish in the south. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/scarcity-
blackfish-south;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb2.60.pdf.

Smiley:1882:MRE

[173] Chas W. Smiley. Method and results of an effort to collect statistics of the fish trade and consumption of fish throughout the United States. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
method-and-results-effort-collect-statistics-fish-trade-and-
consumption-fish-throughout;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb2.68.pdf.

Collins:1882:CRP

[174] J. W. Collins. Correspondence relative to the procurement of certain fishery implements for the government of New South Wales. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
correspondence-relative-procurement-certain-fishery-implements-
government-new-south-wales;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb2.76.pdf.

Bouchon-Brandely:1882:SCO

[175] G. Bouchon-Brandely. On the sexuality of the common oyster (O. edulis), and that of the Portuguese oyster (O. angulata). Artificial fecundation of the Portuguese oyster. Fishery Bulletin, 2(1):??, 1882. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/sexuality-common-oyster-o-edulis-
and-portuguese-oyster-o-angulata-artificial-fecundation;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb2.84.pdf.

Atkins:1882:SSS

[176] Charles G. Atkins. Sketch of the schoodic salmon-breeding establishment. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/sketch-schoodic-salmon-breeding-
establishment;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb2.92.pdf.

Anonymous:1882:I

[177] Anonymous. Index. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/index-40;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb2.index_.pdf.

Gifford:1882:DSV

[178] George Gifford. The disappearance of sardines from the vendean coast, and its causes. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/disappearance-sardines-vendean-coast-
and-its-causes;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb2.5.pdf.

Ward:1882:LWP

[179] Lester F. Ward. List of water plants for carp ponds. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/list-water-plants-carp-ponds-0;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb2.13_0.
pdf.

Worth:1882:PSS

[180] S. G. Worth. A poor season for shad hatching in North Carolina. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/poor-season-shad-hatching-north-carolina;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb2.21.pdf.

Martin:1882:NFGb

[181] S. J. Martin. Notes on the fisheries of Gloucester, Massachusetts. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/notes-
fisheries-gloucester-massachusetts-2;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb2.29.pdf.

Jordan:1882:LFCb

[182] David S. Jordan and Charles H. Gilbert. List of fishes collected at Panama by Charles H. Gilbert. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/list-fishes-
collected-panama-charles-h-gilbert;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb2.37.pdf.

Leslie:1882:SBM

[183] Chas C. Leslie. Scarcity of blackfish- mortality of codfish. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/scarcity-blackfish-mortality-codfish;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb2.45.pdf.

Lefebvre:1882:PIA

[184] Alfred Lefebvre. The proposed introduction of American catfish into the rivers of Belgium. Fishery Bulletin, 2(1):??, 1882. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/proposed-introduction-american-
catfish-rivers-belgium;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb2.53.pdf.

Ryder:1882:OAY

[185] John A. Ryder. Observations on the absorption of the yelk, the food, feeding, and development of embryo fishes, comprising some investigations conducted at the Central Hatchery, Armory Building, Washington, D.C., in 1882. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/observations-

absorption-yelk-food-feeding-and-development-embryo-fishes-
comprising-some;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb2.61.pdf.

Friedlaender:1882:NMF

[186] Oscar O. Friedlaender. Notes on the menhaden fishery. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/notes-
menhaden-fishery;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb2.69.pdf.

VanAntwerp:1882:GSE

[187] W. Van Antwerp. Growth, spawning, edible qualities, and manner of cooking German carp received from the U. S. Fish Commission in 1880. Fishery Bulletin, 2(1):??, 1882. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/growth-spawning-edible-qualities-
and-manner-cooking-german-carp-received-u-s-fish;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb2.77.pdf.

Krause:1882:HPP

[188] K. E. H. Krause. A hybrid plaice — Platessa vulgaris with Rhombus maximus. Fishery Bulletin, 2(1):??, 1882. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/hybrid-plaice-platessa-vulgaris-
rhombus-maximus;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb2.85.pdf.

Couchman:1882:TCU

[189] J. W. Couchman. Transporting carp from the United States Fish Commission to Brazil. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/transporting-carp-united-
states-fish-commission-brazil;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb2.93.pdf.

Johnson:1882:GGC

[190] Samuel Johnson. Growth of German carp sent to Savoy, Texas, by the United States Fish Commission. Fishery Bulletin, 2(1): ??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/growth-
german-carp-sent-savoy-texas-united-states-fish-commission;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb2.6.pdf.

Phillips:1882:HCP

[191] Barnet Phillips. Holland carp put into the Hudson River about 1830. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/holland-carp-put-hudson-river-about-1830;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb2.14.pdf.

Swain:1882:RSS

[192] Joseph Swain. A review of the species of stolephorus found on the Atlantic Coast of the United States. Fishery Bulletin, 2(1): ??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/review-
species-stolephorus-found-atlantic-coast-united-states;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb2.22.pdf.

McDonald:1882:DGC

[193] Marshall McDonald. Distribution of German carp by the United States Fish Commission. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/distribution-german-carp-
united-states-fish-commission;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb2.30.pdf.

Gilbert:1882:LFO

[194] Charles H. Gilbert. List of fishes observed at Punta Arenas, on the Pacific Coast of Central America. Fishery Bulletin, 2 (1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/list-
fishes-observed-punta-arenas-pacific-coast-central-america;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb2.38.pdf.

Osborn:1882:SFS

[195] John H. Osborn. Shad fishing on the Saint John’s River. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/shad-fishing-saint-johns-river;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb2.46.pdf.

Smiley:1882:PUS

[196] Chas W. Smiley. The proposed use of steamers in the mackerel fishery. Fishery Bulletin, 2(1):??, 1882. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/proposed-use-steamers-mackerel-
fishery;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb2.54.pdf.

Ryder:1882:MSC

[197] John A. Ryder. The microscopic sexual characteristics of the American, Portuguese, and common edible oyster of Europe compared. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
microscopic-sexual-characteristics-american-portuguese-and-
common-edible-oyster-europe;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb2.62.pdf.

Martens:1882:LS

[198] J. B. Martens. Life in the sea. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/life-sea;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb2.70.pdf.

Gillespie:1882:GFC

[199] Samuel Gillespie. Growth and food of carp. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/growth-and-food-carp;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb2.78.pdf.

Hoek:1882:RGO

[200] P. P. C. Hoek. Researches on the generative organs of the oyster (O. edulis). Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/researches-generative-organs-oyster-o-edulis;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb2.86.pdf.

Ryder:1882:MFF

[201] John A. Ryder. On the mode of fixation of the fry of the oyster. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/mode-
fixation-fry-oyster;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb2.94.pdf.

Leonhardt:1882:ICB

[202] G. Leonhardt. Information concerning the blue carp. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/information-concerning-blue-carp;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb2.7.pdf.

Worth:1882:SHC

[203] S. G. Worth. Shad hatching and carp culture. Fishery Bulletin, 2(1): ??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/shad-hatching-
and-carp-culture;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb2.15.pdf.

Hay:1882:CFL

[204] O. P. Hay. On a collection of fishes from the lower Mississippi Valley. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/collection-fishes-lower-mississippi-valley;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb2.23.pdf.

Collins:1882:STM

[205] J. W. Collins. Shad taken in mackerel gill-nets. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/shad-taken-mackerel-
gill-nets;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb2.31.pdf.

Haskell:1882:SAA

[206] E. H. Haskell. Second annual appearance of young cod hatched by the United States Fish Commission in Gloucester harbor in the winter of 1879–1880. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/second-annual-appearance-
young-cod-hatched-united-states-fish-commission-gloucester-
harbor;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb2.39.pdf.

Tanner:1882:RET

[207] Z. L. Tanner. Report of an exploring trip of the steamer Fish Hawk, in Chesapeake Bay in the early spring of 1882. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/report-exploring-trip-steamer-fish-hawk-chesapeake-
bay-early-spring-1882;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb2.47.pdf.

DF:1882:CCT

[208] Deutsche Fisherei-Zeitung. The Cottbus carp trade. Fishery Bulletin, 2 (1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/cottbus-carp-
trade;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb2.55.pdf.

Habershaw:1882:BWO

[209] Frederick Habershaw. Bringing whale oil from the Pacific to New York. Fishery Bulletin, 2(1):215, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/bringing-whale-oil-pacific-new-york;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb2.
63.pdf.

Knight:1882:BPM

[210] William Knight. A bill proposed to the Maryland Legislature at the session of 1878, and entitles ‘An Act to regulate the catching and to provide for the preservation of fish in the waters of the state and of the Potomac River’. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/bill-proposed-
maryland-legislature-session-1878-and-entitles-act-regulate-
catching-and;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb2.71.pdf.

Collins:1882:RUC

[211] J. W. Collins. Report upon a cruise made to the tile-fish groun in the smack Josie Reeves, September, 1882. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/report-upon-cruise-
made-tile-fish-groun-smack-josie-reeves-september-1882;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb2.79.pdf.

Ryder:1882:NOB

[212] John A. Ryder. Note on the organ of Bojanus in Ostrea virginica Gmelin. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/note-organ-bojanus-ostrea-virginica-gmelin;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb2.87.pdf.

Clark:1882:DUS

[213] Frank N. Clark. Description of the United States fish hatchery at Alpena, Michigan. Fishery Bulletin, 2(1):??, 1882. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/description-united-states-fish-hatchery-alpena-
michigan;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb2.95.pdf.

Smiley:1883:CLP

[214] Chas W. Smiley. Check-list of the published reports of the fish commissioners of the various states of the United States. Fishery Bulletin, 3(1): ??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/check-list-
published-reports-fish-commissioners-various-states-united-
states;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb3.3.pdf.

Martin:1883:NCG

[215] S. J. Martin. Notes on the cod gill-net fishing at Gloucester, Mass., 1882–1883. Fishery Bulletin, 3(1):??, 1883. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/notes-cod-gill-net-
fishing-gloucester-mass-1882-1883;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb3.11.pdf.

Bean:1883:DCP

[216] Tarleton H. Bean. Directions for collecting and preserving fish. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/directions-collecting-and-preserving-fish;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb3.19.pdf.

Ryder:1883:PND

[217] John A. Ryder. Preliminary notice of the development and breeding habits of the Potomac catfish, Amiurus albidus (Le Sueur) gill. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
preliminary-notice-development-and-breeding-habits-potomac-
catfish-amiurus-albidus-le-sueur;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb3.27.pdf.

Ward:1883:MAP

[218] Lester F. Ward. Marsh and aquatic plants of the northern United States, many of which are suitable for carp ponds. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
marsh-and-aquatic-plants-northern-united-states-many-which-
are-suitable-carp-ponds;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb3.35.pdf.

Phillips:1883:ECM

[219] Barnet Phillips. Early catch of mackerel. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/early-catch-mackerel;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb3.43.pdf.

vonTreskow-Weissagk:1883:WDT

[220] ?. von Treskow-Weissagk. What dangers threaten the products of the old pond carp fisheries from artificial fish culture. Fishery Bulletin, 3(1):303– 304, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/what-dangers-
threaten-products-old-pond-carp-fisheries-artificial-fish-culture;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb3.51.pdf.

Stearns:1883:ECP

[221] R. E. C. Stearns. Edible clams of the Pacific Coast, and a proposed method of transplanting them to the Atlantic Coast. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/edible-clams-pacific-coast-and-proposed-method-
transplanting-them-atlantic-coast;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb3.59.pdf.

Downman:1883:WFS

[222] R. H. Downman. What frogs sometimes eat. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/what-frogs-sometimes-
eat;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb3.67.pdf.

Brown:1883:SNW

[223] James Temple Brown. Some notes on whales. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/some-notes-whales;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb3.75.pdf.

NYMIJ:1883:RSF

[224] New York Market Index and Journal. Report for 1879 and 1880 of the sale of fish in Fulton Market, New York. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/report-1879-and-1880-
sale-fish-fulton-market-new-york;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb3.83.pdf.

Dohrn:1883:ZSN

[225] Anton Dohrn. The zoological station at Naples, Italy. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/zoological-station-naples-italy;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb3.
91.pdf.

Smith:1883:MC

[226] Everett Smith. A mammoth cod. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/mammoth-cod;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb3.99.pdf.

Buck:1883:RUH

[227] H. H. Buck. Report upon hatching one million whitefish eggs from the Great Lakes, at Orland, Me., and upon planting the young in Eagle Lake, Mount Desert. Fishery Bulletin, 3(1):??, 1883. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/report-upon-hatching-one-million-
whitefish-eggs-great-lakes-orland-me-and-upon-planting;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb3.108.pdf.

Babbitt:1883:IEM

[228] John O. Babbitt. The injurious effect of menhaden steamers upon the food fisheries. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/injurious-effect-menhaden-steamers-upon-food-
fisheries;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb3.116.pdf.

Martin:1883:MM

[229] S. J. Martin. Movements of mackerel. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/movements-mackerel;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb3.124.pdf.

Slade:1883:DAH

[230] Elisha Slade. Dates of the appearance of herring, shad, bass, tango, seup, frostfish, fish-hawks, kingfishers, and Greenland seal in the taunton river, from 1871 to 1883, inclusive. Fishery Bulletin, 3(1):??, 1883. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/dates-appearance-herring-
shad-bass-tango-seup-frostfish-fish-hawks-kingfishers-and;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb3.132.pdf.

Smiley:1883:OPU

[231] Chas W. Smiley. The organization and personnel of the United States Fish Commission, June 1, 1883. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/organization-
and-personnel-united-states-fish-commission-june-1-1883;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb3.4.pdf.

Martin:1883:APS

[232] S. J. Martin. The American purse-seine successful in the herring fisheries of Sweden. Fishery Bulletin, 3(1):??, 1883. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/american-purse-seine-successful-
herring-fisheries-sweden;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb3.12.pdf.

Smiley:1883:PCS

[233] Chas W. Smiley. Planting and catch of shad in Coosa River, Alabama. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/planting-and-catch-shad-coosa-river-alabama;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb3.20.pdf.

Redmond:1883:RTR

[234] Roland Redmond. On rainbow trout reared from eggs brought from California. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/rainbow-trout-reared-eggs-brought-california;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb3.28.pdf.

Snell:1883:IFC

[235] Merwin P. Snell. An important fishway cause. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
important-fishway-cause;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb3.36.pdf.

Ryder:1883:ROA

[236] John A. Ryder. Rearing oysters from artificially fertilized eggs, together with notes on pond culture, etc. Fishery Bulletin, 3(1): ??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/rearing-
oysters-artificially-fertilized-eggs-together-notes-pond-culture-
e;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb3.44.pdf.

Smiley:1883:NEQ

[237] Chas W. Smiley. Notes on the edible qualities of German carp, and hints about cooking them. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/notes-edible-qualities-german-carp-and-
hints-about-cooking-them;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb3.52_0.pdf.

Spinney:1883:RGY

[238] Benjamin H. Spinney. Return to Gloucester of young codfish hatched by United States Fish Commission in 1879. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
return-gloucester-young-codfish-hatched-united-states-fish-
commission-1879;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb3.60.pdf.

Downman:1883:SHE

[239] R. H. Downman. Shad, herring, etc., excluded from the Rappahannock River by dams. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/shad-herring-e-excluded-rappahannock-
river-dams;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb3.68.pdf.

Phillips:1883:SCH

[240] Barnet Phillips. A stray cod up the Hudson. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/stray-cod-hudson;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb3.76.pdf.

Farr:1883:DCS

[241] S. C. Farr. Description of a California salmon (Oncorhynchus sp.) found in one of the rivers of New Zealand and identified by Dr. T. H. Bean. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
description-california-salmon-oncorhynchus-sp-found-one-rivers-
new-zealand-and-identified;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb3.84.pdf.

Dresser:1883:NNH

[242] Henry Dresser. Notes on the natural history of the buffalo. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
notes-natural-history-buffalo;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb3.92.pdf.

Hessel:1883:ABC

[243] Rud Hessel. Arrival of blue carp from Germany. Fishery Bulletin, 3(1): ??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/arrival-blue-
carp-germany;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb3.100.pdf.

Polk:1883:EMO

[244] William R. Polk. Experience of a menhaden oil manufacturer in Northumberland County, Va. Fishery Bulletin, 3(1):??, 1883. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/experience-menhaden-
oil-manufacturer-northumberland-county-va;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb3.109.
pdf.

Southwell:1883:NEH

[245] Thomas Southwell. Notes on the English herring fishery of 1882. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/notes-
english-herring-fishery-1882;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb3.117.pdf.

Tanner:1883:MM

[246] Z. L. Tanner. Movements of mackerel. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/movements-mackerel-0;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb3.125.pdf.

Anonymous:1883:I

[247] Anonymous. Index. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/index-41;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb3.index_.pdf.

Smiley:1883:LNAa

[248] Chas W. Smiley. A list of the names and addresses of the Fish Commissioners of the various states of the United States, and of Superintendents of hatcheries, corrected to June 1, 1883, with some indication of the limits of the terms upon with they are now serving. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
list-names-and-addresses-fish-commissioners-various-states-
united-states-and;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb3.5.pdf.

Anonymous:1883:NFG

[249] Anonymous. Notes on the fisheries of Gloucester, Massachusetts. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/notes-
fisheries-gloucester-massachusetts-3;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb3.13.pdf.

Swan:1883:RIN

[250] James G. Swan. Report of investigations at Neah Bay, Wash., respecting the habits of fur-seals of that vicinity, and to arrange for procuring specimens of skeletons of Cetacea. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/report-investigations-neah-bay-wash-respecting-
habits-fur-seals-vicinity-and-arrange;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb3.21.pdf.

Ljungman:1883:SGI

[251] Axel Vilhelm Ljungman. Sweden at the Great International Fishery Exposition at London, 1883. Fishery Bulletin, 3(1):??, 1883. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/sweden-great-international-fishery-
exposition-london-1883;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb3.29.pdf.

Census:1883:SFU

[252] Tenth Census. Statistics of the fisheries of the United States in 1880. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/statistics-fisheries-united-states-1880;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb3.37.pdf.

Walke:1883:SEP

[253] E. H. Walke. Spawning of esox (pike or pickerel) in North Carolina. Fishery Bulletin, 3(1):295, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/spawning-esox-pike-or-pickerel-north-
carolina;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb3.45.pdf.

Pierce:1883:SBO

[254] H. D. Pierce. The spawning of bluefish — an opinion of the cause of mortality of fish in the Gulf of Mexico. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
spawning-bluefish-opinion-cause-mortality-fish-gulf-mexico;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb3.53.pdf.

Bartlett:1883:CCP

[255] S. P. Bartlett. The cultivation of croppies in ponds. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/cultivation-croppies-ponds;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb3.61.pdf.

Mobius:1883:CWD

[256] Karl Möbius. Can we do anything to decrease, preserve, or increase the number of our useful marine animals? Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/can-we-do-anything-
decrease-preserve-or-increase-number-our-useful-marine-animals;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb3.69.pdf.

CAA:1883:SCG

[257] Cape Ann Advertiser. Success of the cod gill-net fishery. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/success-
cod-gill-net-fishery;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb3.77.pdf.

DFZ:1883:ARC

[258] Deutsche Fischerei Zeitung. The artificial raising of cod in America. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/artificial-raising-cod-america;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb3.85.pdf.

Willequet:1883:FCBa

[259] E. Willequet. Fish culture in Belgium — proposed government action. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fish-culture-belgium-proposed-government-action;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb3.93.pdf.

White:1883:ACE

[260] C. M. White. Ability of carp to endure muddy, slimy, and stagnant water. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/ability-carp-endure-muddy-slimy-and-stagnant-water;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb3.101.pdf.

Barker:1883:IMO

[261] E. L. Barker. The incipiency of the menhaden oil business. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/incipiency-menhaden-oil-business;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb3.110.pdf.

JohnstonSons:1883:SFA

[262] Jos. Johnston & Sons. Scotland fisheries and apparatus. Fishery Bulletin, 3(1):466, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/scotland-
fisheries-and-apparatus;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb3.118.pdf.

Metcalf:1883:SPC

[263] Martin Metcalf. Success in propagating California mountain trout from eggs brought to Michigan by the United States Fish Commission. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
success-propagating-california-mountain-trout-eggs-brought-
michigan-united-states-fish;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb3.126.pdf.

Smiley:1883:LNAb

[264] Chas W. Smiley. A list of the names of all persons who have been appointed commissioners of fisheries of the various states of the United States, with their addresses, and the dates of their terms of service. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/list-names-all-persons-who-have-been-appointed-
commissioners-fisheries-various-states;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb3.6.pdf.

StJohn:1883:AWF

[265] J. Hector St. John. An account of the whale fishery of Nantucket, Mass., one hundred years ago. Fishery Bulletin, 3(1):??, 1883. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/account-whale-fishery-nantucket-
mass-one-hundred-years-ago;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb3.14.pdf.

Raveret-Wattel:1883:RCS

[266] C. Raveret-Wattel and ?. Bartet. Reproduction of California salmon in the Aquarium of Trocadero. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/reproduction-california-salmon-
aquarium-trocadero;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb3.22.pdf.

Baird:1883:INM

[267] Spencer F. Baird. The instruction of naval midshipmen in taxidermy, ichthyology, etc., at the United States National Museum and on board the steamers of the United States Fish Commission. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/instruction-naval-midshipmen-taxidermy-ichthyology-
e-united-states-national-museum-and;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb3.30.pdf.

Mason:1883:PEW

[268] Frank H. Mason. Proposed exportation of whitefish eggs to Switzerland. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/proposed-exportation-whitefish-eggs-switzerland;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb3.38.pdf.

McDonald:1883:SDF

[269] Marshall McDonald. Summary of distribution of fish by one car during the past season. Fishery Bulletin, 3(1):295, 1883. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/summary-distribution-fish-car-no-

1-during-past-season;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb3.46.pdf.

Smiley:1883:GCI

[270] Chas W. Smiley. The German carp, and its introduction in the United States. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/german-carp-and-its-introduction-united-states;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb3.54.pdf.

Malmgren:1883:MAB

[271] A. J. Malmgren. Memorial addressed to the Bureau of Agriculture of the Imperial Senate for Finland, January 20, 1883, in regard to the advisability of introducing artificial fish culture in Finland. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/memorial-addressed-bureau-agriculture-imperial-
senate-finland-january-20-1883-regard;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb3.62.pdf.

Nicklas:1883:PYF

[272] Carl Nicklas. Production of young fry of the carp. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/production-young-fry-carp;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb3.70.pdf.

Gardner:1883:EPC

[273] A. P. Gardner. Experiments in the pond culture of trout, suckers, and catfish. Fishery Bulletin, 3(1):??, 1883. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/experiments-pond-culture-trout-
suckers-and-catfish;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb3.78.pdf.

Anderson:1883:NBS

[274] John Anderson. Notes on the biology of the salmon and grilse. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/notes-biology-salmon-and-grilse;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb3.
86.pdf.

Baird:1883:PPP

[275] Spencer F. Baird. Purchase of Professor Poey’s Ichthyologia cubana by the Spanish government. Fishery Bulletin, 3(1):??, 1883. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/purchase-professor-poeys-
ichthyologia-cubana-spanish-government;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb3.94.pdf.

Earll:1883:PLH

[276] R. E. Earll. Proposed limitation of the herring fishery in Maine and its effect upon bait and canning interests. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/proposed-limitation-
herring-fishery-maine-and-its-effect-upon-bait-and-canning-
interests;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb3.102.pdf.

DHomergue:1883:RRU

[277] Louis C. D’Homergue. Reasons for restricting the use of steamers and seines in the menhaden fishery. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/reasons-restricting-
use-steamers-and-seines-menhaden-fishery;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb3.111.
pdf.

Stearns:1883:FFG

[278] Silas Stearns. Fluctuations in the fisheries of the Gulf of Mexico and the proposed investigation of them. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/fluctuations-
fisheries-gulf-mexico-and-proposed-investigation-them;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb3.119.pdf.

Mather:1883:ELM

[279] Fred Mather. Exchange of live marine specimens with France. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/exchange-live-marine-specimens-france;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb3.127.pdf.

Stabler:1883:GPC

[280] Edward Stabler. Growth and protection of carp from turtles, etc. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
growth-and-protection-carp-turtles-etc;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb3.7.pdf.

McDonald:1883:HEL

[281] Marshall McDonald. History of the experiments leading to the development of the automatic fish-hatching jar. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/history-experiments-
leading-development-automatic-fish-hatching-jar;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb3.15.pdf.

VS:1883:NFH

[282] von La Valette St. George. A new fish-hatching apparatus. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/new-
fish-hatching-apparatus;   https://spo.nmfs.noaa.gov/content/
new-fish-hatching-apparatus-0;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb3.106.pdf;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb3.23.pdf.

Wallem:1883:EFO

[283] Fredrik M. Wallem. Export of fish-oil from Norway, 1878–1882. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/export-fish-oil-norway-1878-1882;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb3.31.pdf.

Smiley:1883:FAP

[284] Chas W. Smiley. First appearance of porpoise. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/first-
appearance-porpoise;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb3.39.pdf.

Martin:1883:MSM

[285] S. J. Martin. The movements of schools of mackerel. Fishery Bulletin, 3(1):295–396, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/movements-schools-mackerel;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb3.47.pdf.

Jennings:1883:PPF

[286] Ralph S. Jennings. Process of preserving fish. Fishery Bulletin, 3 (1):336, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/process-
preserving-fish;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb3.55.pdf.

DFZ:1883:VAA

[287] Deutsche Fischerei Zeitung. A verdict against artificial fish culture, and an answer thereto. Fishery Bulletin, 3(1):??, 1883. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/verdict-agains-artificial-fish-
culture-and-answer-thereto;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb3.63.pdf.

Anonymous:1883:PCT

[288] Anonymous. Prices of carp, tench, and goldorfe in Germany. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
prices-carp-tench-and-goldorfe-germany;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb3.71.pdf.

Waterbury:1883:OSP

[289] E. B. Waterbury et al. Obstructions to salmon in Pend d’Oreille River, Montana, requiring a fish-way or removal of falls. Fishery Bulletin, 3(1): ??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/obstructions-
salmon-pend-doreille-river-montana-requiring-fish-way-or-removal-
falls;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb3.79.pdf.

Meek:1883:NAS

[290] Seth E. Meek. A note on the Atlantic species of the genus Anguilla. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY.

ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/note-atlantic-species-genus-anguilla;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb3.87.pdf.

Wilcox:1883:RYC

[291] W. A. Wilcox. Reappearance of young cod hatched by the United States Fish Commission. Fishery Bulletin, 3(1):??, 1883. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/reappearance-young-cod-hatched-
united-states-fish-commission;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb3.95.pdf.

Huske:1883:SWS

[292] C. J. Huske. Shad work in South Carolina. Fishery Bulletin, 3(1): ??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/shad-work-
south-carolina;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb3.103.pdf.

Wharton:1883:WMS

[293] Joseph Wharton. Why menhaden seining should not be prohibited within two miles of the shore. Fishery Bulletin, 3(1): ??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/why-
menhaden-seining-should-not-be-prohibited-within-two-miles-
shore;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb3.113.pdf.

Abbe:1883:PLM

[294] W. A. Abbe. Proposed limitation of menhaden fishing to July 1, and of mackerel fishing to June 15 — occurrence of herring. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/proposed-limitation-menhaden-fishing-july-
1-and-mackerel-fishing-june-15-occurrence-herring;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb3.120.pdf.

LDT:1883:SHF

[295] London Daily Telegraph. The Scotch herring fisheries. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/scotch-
herring-fisheries;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb3.128.pdf.

Smiley:1883:LBF

[296] Chas W. Smiley and Chas W. Scudder. A list of the blank forms, circulars, and minor publications of the United States Fish Commission, from its organization, in 1871, to August 1, 1883. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/list-
blank-forms-circulars-and-minor-publications-united-states-
fish-commission-its;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb3.8.pdf.

Holberton:1883:SSS

[297] Wakeman Holberton. The successful stocking of streams with trout. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
successful-stocking-streams-trout;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb3.16.pdf.

Mobius:1883:EBP

[298] Karl Möbius. On experiments, begun in 1880, to plant American oysters in the Western Baltic, and the usefulness of contenting these experiments with the aid of the German Fishery Association. Fishery Bulletin, 3(1): ??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/experiments-
begun-1880-plant-american-oysters-western-baltic-and-usefulness-
contenting;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb3.24.pdf.

Smiley:1883:AOH

[299] Chas W. Smiley. Answers to one hundred and eighteen questions relating to German carp. Fishery Bulletin, 3(1):??, 1883. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/answers-one-hundred-
and-eighteen-questions-relating-german-carp;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb3.
32.pdf.

Smiley:1883:TVU

[300] Chas W. Smiley. Tonnage of vessels of the United States employed in the whale, cod, and mackerel fisheries, from 1860 to 1882, inclusive. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
tonnage-vessels-united-states-employed-whale-cod-and-mackerel-
fisheries-1860-1882-inclusive;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb3.40.pdf.

Martin:1883:NFG

[301] S. J. Martin. Notes on the fisheries of Gloucester, Massachusetts. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/notes-
fisheries-gloucester-massachusetts-4;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb3.48.pdf.

Hubrecht:1883:FCS

[302] A. A. W. Hubrecht. Fish culture as seen at the London Exhibition, with special references to its history, apparatus, and methods in the United States. Fishery Bulletin, 3(1):??, 1883. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/fish-culture-seen-london-
exhibition-special-references-its-history-apparatus-and-methods;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb3.56.pdf.

vonPolenz:1883:ORR

[303] Chamberlain von Polenz. An opinion regarding R. Eckhardt’s answer to Professor Malmgren. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/opinion-regarding-r-eckhardts-answer-
professor-malmgren;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb3.64.pdf.

Hunger:1883:VCS

[304] Oscar Hunger. Varieties of carp in Saxony. Fishery Bulletin, 3(1): ??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/varieties-
carp-saxony;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb3.72.pdf.

Barnes:1883:SOS

[305] William M. Barnes. Supposed occurrence of sail-fish (Histiophorus). Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/supposed-occurrence-sail-fish-histiophorus;   https:

//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb3.80.pdf.

SchieffelinCo:1883:UVY

[306] W. H. Schieffelin & Co. Unexplained variations in the yield of oil from cod livers. Fishery Bulletin, 3(1):431, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/unexplained-variations-yield-oil-cod-
livers;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb3.88.pdf.

Ryder:1883:RAA

[307] John A. Ryder. Report on the abnormal appearance of some shad eggs from a fish kept in confinement at Havre de Grace, Maryland. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/report-abnormal-appearance-some-shad-eggs-fish-
kept-confinement-havre-de-grace-maryland;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb3.96.pdf.

Cary:1883:PIE

[308] H. H. Cary. Period of incubation of eggs of German carp. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
period-incubation-eggs-german-carp;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb3.104.pdf.

Friedlaender:1883:NMF

[309] Oscar O. Friedlaender. Notes on the menhaden fishery. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/notes-
menhaden-fishery-0;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb3.112.pdf.

Vail:1883:WCM

[310] David F. Vail. Why the catching of menhaden with seines, et., should be restricted. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/why-catching-menhaden-seines-et-should-be-
restricted;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb3.121.pdf.

Koss:1883:TTC

[311] R. A. Koss. Trout and trout culture. Fishery Bulletin, 3(1): ??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/trout-and-
trout-culture;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb3.129.pdf.

Anonymous:1883:P

[312] Anonymous. Plates. Fishery Bulletin, 3(1):??, 1883. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/plates-0;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb3%20plates.pdf.

Smiley:1883:FTS

[313] Chas W. Smiley. Four tables, showing the amount of public money appropriated for carrying on the United States Commission of Fish and Fisheries, and the various State Commissions, from 1865 to 1882, inclusive. Fishery Bulletin, 3(1):??, 1883. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/four-tables-showing-
amount-public-money-appropriated-carrying-united-states-commission-
fish;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb3.9.pdf.

Ryder:1883:TBE

[314] John A. Ryder. On the thread-bearing eggs of the silversides (Menidia). Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/thread-bearing-eggs-silversides-menidia;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb3.17.pdf.

Willequet:1883:FCBb

[315] E. Willequet. Fish culture in Belgium. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/fish-culture-belgium;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb3.25.pdf.

MFCR:1883:DCC

[316] Maryland Fish Commission Report. Directions concerning the construction of carp ponds. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY.

ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/directions-concerning-construction-carp-
ponds;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb3.33.pdf.

Smiley:1883:SIE

[317] Chas W. Smiley. Statistics of the imports and exports of fish, fish-oil, whalebone, etc., for ten years ending June 30, 1882. Fishery Bulletin, 3 (1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/statistics-
imports-and-exports-fish-fish-oil-whalebone-e-ten-years-ending-
june-30-1882;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb3.41.pdf.

Crosby:1883:SSM

[318] John Schuyler Crosby. Stocking the streams of Montana with fish. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
stocking-streams-montana-fish;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb3.49.pdf.

Taylor:1883:AAW

[319] Fred W. Taylor. Analysis of artesian-well water from the shadhatching station at Havre de Grace, Md. Fishery Bulletin, 3(1): ??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/analysis-
artesian-well-water-shad-hatching-station-havre-de-grace-md;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb3.57.pdf.

Potts:1883:FWS

[320] Edward Potts. Fresh-water sponges: what, where, when, and who wants them. Fishery Bulletin, 3(1):??, 1883. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/fresh-water-sponges-what-where-
when-and-who-wants-them;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb3.65.pdf.

Hanna:1883:DEL

[321] S. W. Hanna. Description of an eel-like creature taken in a net at New Harbor, Maine, in 1880. Fishery Bulletin, 3(1):??, 1883. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/description-eel-creature-taken-

net-new-harbor-maine-1880;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb3.73.pdf.

Martin:1883:FCG

[322] S. J. Martin. Fitting out with cod gill-nets. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/fitting-out-cod-gill-
nets;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb3.81.pdf.

Atwood:1883:UVY

[323] N. E. Atwood. Unexplained variations in the yield of oil from cod livers. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/unexplained-variations-yield-oil-cod-
livers-0;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb3.89.pdf.

Lennurt:1883:CSS

[324] J. T. Lennurt. Catch of spawning shad in the Potomac, above Georgetown. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/catch-spawning-shad-potomac-above-georgetown;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb3.97.pdf.

Lowell:1883:SUS

[325] James Russell Lowell. Success of the United States Exhibit at the London International Fisheries Exhibition. Fishery Bulletin, 3(1): ??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/success-
united-states-exhibit-london-international-fisheries-exhibition;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb3.105.pdf.

ACR:1883:LHD

[326] Atlantic City Review. Large haul of drum-fish. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/large-haul-drum-fish;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb3.114.pdf.

Luce:1883:OBC

[327] Thomas R. Luce. Occurrence of Balistes caprices Gmelin (leather-jacket or file-fish) at New Bedford, Mass. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/occurrence-balistes-
caprices-gmelin-leather-jacket-or-file-fish-new-bedford-mass;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb3.122.pdf.

Stone:1883:SSL

[328] Livingston Stone. Scarcity of salmon in the Little Spokane and other streams on Pacific Coast. Fishery Bulletin, 3(1):??, 1883. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/scarcity-salmon-little-spokane-
and-other-streams-pacific-coast;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb3.130.pdf.

Smiley:1883:LPR

[329] Chas W. Smiley. A list of the published reports of the commissioners appointed by authority of various states of the United States. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/list-published-reports-commissioners-appointed-
authority-various-states-united-states;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb3.2.pdf.

Martin:1883:SER

[330] S. J. Martin. Some of the early results of the introduction of cod gillnets at Gloucester, Mass., 1880–1881. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/some-early-results-
introduction-cod-gill-nets-gloucester-mass-1880-1881;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb3.10.pdf.

Nye:1883:DSC

[331] Willard Nye, Jr. Do snakes catch and eat fish? Fishery Bulletin, 3(1):196, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/do-snakes-catch-and-
eat-fish;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb3.18.pdf.

vondemBorne:1883:ABB

[332] Max von dem Borne. The American black bass. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/american-black-bass;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb3.26.pdf.

Clark:1883:SCM

[333] Frank N. Clark. Spawning of California mountain trout, reared in confinement, from eggs brought from McCloud River. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/spawning-california-mountain-trout-reared-
confinement-eggs-brought-mccloud-river;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb3.34.pdf.

Wood:1883:RTN

[334] W. M. Wood. Report on the trap-nets fished in the spring of 1883 in the Potomac River and Chesapeake Bay. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/report-trap-
nets-fished-spring-1883-potomac-river-and-chesapeake-bay;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb3.42.pdf.

Sterling:1883:IDW

[335] E. Sterling. On the impropriety of depositing whitefish minnows off the Harbor of Cleveland, Ohio — fishing for saugers. Fishery Bulletin, 3(1): ??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/impropriety-
depositing-whitefish-minnows-harbor-cleveland-ohio-fishing-
saugers;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb3.50.pdf.

Clark:1883:PDW

[336] Frank N. Clark and Seymour Bower. On the propriety of depositing whitefish minnows off the Harbor of Cleveland, Ohio. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/propriety-
depositing-whitefish-minnows-harbor-cleveland-ohio;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb3.58.pdf.

vonBehr:1883:MSC

[337] ?. von Behr. Management of spawning carp. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/management-spawning-
carp;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb3.66.pdf.

Mosher:1883:DSB

[338] Gideon Mosher. Do striped bass (Roccus lineatus) feed on menhaden? Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/do-striped-bass-roccus-lineatus-feed-
menhaden;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb3.74.pdf.

Dall:1883:NPC

[339] William H. Dall. Notes on the Pacific Coast trade in shells, shrimps cod and salmon. Fishery Bulletin, 3(1):??, 1883. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/notes-pacific-coast-trade-shells-
shrimps-cod-and-salmon;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb3.82.pdf.

CAA:1883:NPD

[340] Cape Ann Advertiser. A new process of drying fish. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
new-process-drying-fish;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb3.90.pdf.

Collins:1883:SGN

[341] J. W. Collins. Success of the gill-net cod fishery on the New England Coast, winter of 1882–1883. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/success-gill-net-cod-
fishery-new-england-coast-winter-1882-1883;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb3.98.pdf.

Church:1883:WSM

[342] Daniel T. Church. Why the size of mesh in menhaden seines should not be restricted. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY.

ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/why-size-mesh-menhaden-seines-should-
not-be-restricted;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb3.115.pdf.

Grant:1883:MMC

[343] Isaac H. Grant. Movements of menhaden — catch of herring. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/movements-menhaden-catch-herring;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb3.123.pdf.

Harris:1883:TAS

[344] Gwynn Harris. Total amount of shad and herring caught on the Potomac River during the spring of 1882. Fishery Bulletin, 3(1):??, 1883. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/total-amount-
shad-and-herring-caught-potomac-river-during-spring-1882;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb3.131.pdf.

FS:1884:SCF

[345] Forest and Stream. Snakes catching fish. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/snakes-catching-fish;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb4.129.pdf.

Leslie:1884:CSS

[346] Charles C. Leslie. On the cultivation of soft-shell crabs. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
cultivation-soft-shell-crabs;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb4.137.pdf.

BSd:1884:NCF

[347] Bulletin de la Société d’acclimatation. Notes on the cultivation of fish — mostly American — in France. Fishery Bulletin, 4 (1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/notes-
cultivation-fish-mostly-american-france;   https://spo.nmfs.

noaa.gov/sites/default/files/pdf-content/fish-bull/fb4.145.
pdf.

Jordan:1884:PPC

[348] David S. Jordan. Proposed propagation of catfish as a food-fish. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
proposed-propagation-catfish-food-fish;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb4.153.
pdf.

Holden:1884:CRS

[349] E. C. Holden. The Columbia River salmon — a hatchery needed. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/columbia-river-salmon-hatchery-needed;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.161.pdf.

Atkins:1884:MLS

[350] Charles G. Atkins. Memoranda on landlocked salmon. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/memoranda-
landlocked-salmon;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb4.170.pdf.

Rumpff:1884:OPA

[351] Carl Rumpff. The oyster as a popular article of food in North America. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/oyster-popular-article-food-north-america;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb4.178.pdf.

Ryder:1884:CGF

[352] John A. Ryder. Care of gold-fish. — queries of William Rosenstihl, Jr., with replies. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/care-gold-fish-queries-william-rosenstihl-
jr-replies;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.186.pdf.

Paulsen:1884:WMK

[353] Paul George Paulsen. The world’s market for klip-fish, roe, and herring. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/worlds-market-klip-fish-roe-and-herring;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.194.pdf.

Earll:1884:HBS

[354] R. E. Earll. Hatching blackfish and Spanish mackerel. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/hatching-blackfish-and-spanish-mackerel;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.202.pdf.

Certes:1884:NUE

[355] A. Certes. Note upon the effect of high pressures on the vitality of minute fresh-water and salt-water organisms. Fishery Bulletin, 4 (1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/note-upon-
effect-high-pressures-vitality-minute-fresh-water-and-salt-
water-organisms;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb4.210.pdf.

Smiley:1884:BNUa

[356] Chas W. Smiley. Brief notes upon fish and fisheries. Fishery Bulletin, 4(1):456, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/brief-notes-upon-fish-and-fisheries-1;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.218.pdf.

Smiley:1884:IFO

[357] Chas W. Smiley. Inspection of fish and other marine products in the District of Columbia, 1870 to 1883, inclusive. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/inspection-
fish-and-other-marine-products-district-columbia-1870-1883-
inclusive;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.1.pdf.

Ryder:1884:NFF

[358] John A. Ryder. On a new form of filter or diaphragm to be used in the culture of oysters in ponds. Fishery Bulletin, 4(1): ??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/new-form-
filter-or-diaphragm-be-used-culture-oysters-ponds;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.9.pdf.

Siler:1884:DFP

[359] Andrew L. Siler. Depletion of fish in Panguitch and Bear Lakes, Utah. Fishery Bulletin, 4(1):51, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/depletion-fish-panquitch-and-bear-lakes-utah;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb4.17.pdf.

Manley:1884:AFI

[360] J. J. Manley. American fish introduced in English waters. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/american-fish-introduced-english-waters;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.25.pdf.

BSd:1884:ALL

[361] Bulletin de la Société d’acclimatation. American land-locked salmon and lake trout in France. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/american-land-locked-salmon-and-lake-
trout-france;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.33.pdf.

Barber:1884:SRL

[362] C. H. Barber. Success in raising landlocked salmon. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
success-raising-landlocked-salmon;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb4.41.pdf.

Blackford:1884:CIS

[363] E. G. Blackford. Carp in an installment of shad from James River. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
carp-installment-shad-james-river;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb4.49.pdf.

Dall:1884:NFP

[364] W. H. Dall. Notes on fishing products exported from San Francisco, Cal., during the year 1883. Fishery Bulletin, 4(1):??, 1884. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/notes-fishing-products-exported-
san-francisco-cal-during-year-1883;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb4.57.pdf.

Wood:1884:RES

[365] W. M. Wood. Report of an examination of the shad-fisheries in Georgetown, S.C. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/report-examination-shad-fisheries-georgetown-
s-c;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb4.65.pdf.

Gerber:1884:HCC

[366] C. Gerber, Jr. How to cook carp. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/how-cook-carp;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb4.73.pdf.

Hensman:1884:ABB

[367] John T. Hensman. American black bass placed in the River Nene, England. Fishery Bulletin, 4(1):??, 1884. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/american-black-bass-placed-river-
nene-england;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.81.pdf.

Thompson:1884:EQC

[368] Edward Thompson. Edible qualities of carp. Fishery Bulletin, 4(1): ??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/edible-
qualities-carp;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.89.pdf.

Nickerson:1884:DSF

[369] Seth Nickerson. Destruction of small fish in weirs. Fishery Bulletin, 4(1): ??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/destruction-
small-fish-weirs;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb4.97.pdf.

Wood:1884:RTM

[370] W. M. Wood. Report of a trip made by the Fish Hawk to the lower part of Chesapeake Bay, to ascertain the character of the fisheries for shad, herring, etc., in the spring of 1884. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/report-trip-
made-fish-hawk-lower-part-chesapeake-bay-ascertain-character-
fisheries-shad;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.105.pdf.

Collins:1884:AWR

[371] J. W. Collins. An adventure with a whale in the River Tay, Scotland. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/adventure-whale-river-tay-scotland;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb4.113.pdf.

L:1884:NNL

[372] Laflin & Co. Need of a national law to regulate the size of mesh of both pound and gill news on the Great Lakes. Fishery Bulletin, 4(1):223–224, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/need-national-law-regulate-size-mesh-both-
pound-and-gill-news-great-lakes;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb4.121.pdf.

Smiley:1884:OBG

[373] Chas W. Smiley. Occurrence of black grouper or jew-fish off Block Island. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/occurrence-black-grouper-or-jew-fish-
block-island;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.130.pdf.

Simms:1884:FEP

[374] G. E. Simms, Jr. A fish-eating plant. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/fish-eating-
plant;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb4.138.pdf.

Christensen:1884:NHF

[375] R. Christensen. Notes on the history of the fish-hook. Fishery Bulletin, 4 (1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/notes-history-
fish-hook;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.146.pdf.

Bean:1884:BTM

[376] Tarleton H. Bean. Brook trout from Monadnock Lake and Cristine Lake, New Hampshire. Fishery Bulletin, 4(1):??, 1884. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/brook-trout-monadnock-lake-and-
cristine-lake-new-hampshire;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb4.154.pdf.

Smiley:1884:BNUb

[377] Chas W. Smiley. Brief notes upon fish and the fisheries. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/brief-
notes-upon-fish-and-fisheries;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb4.162.pdf.

Ryder:1884:NRS

[378] John A. Ryder. Note on the regeneration of the scales of the German carp. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/note-regeneration-scales-german-carp;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb4.171.pdf.

Smiley:1884:BNUc

[379] Chas W. Smiley. Brief notes upon fish and fisheries. Fishery Bulletin, 4(1):359–368, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/brief-notes-upon-fish-and-fisheries-0;   https:

//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.179.pdf.

Wallem:1884:EFO

[380] Fredrik Wallem. Exports of fish-oil from Norway, 1878–1882. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
exports-fish-oil-norway-1878-1882;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb4.187.pdf.

Kunkel:1884:CUW

[381] ?. Kunkel. On the conditions under which trout exist in the German waters. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/conditions-under-which-trout-exist-german-waters;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb4.195.pdf.

Peixotto:1884:SFF

[382] Benjamin F. Peixotto. The sea-fisheries of France and Algiers. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/sea-fisheries-france-and-algiers;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.203.pdf.

Collins:1884:SMGa

[383] J. W. Collins. On the scarcity of mackerel in the Gulf of St. Lawrence. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/scarcity-mackerel-gulf-st-lawrence-0;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb4.211.pdf.

Joncas:1884:CF

[384] L. Z. Joncas. The Canadian fisheries. Fishery Bulletin, 4(1): ??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/canadian-
fisheries;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.219.pdf.

Harris:1884:RUSa

[385] Gwynn Harris. Report upon the shad and herring fisheries of the Potomac River for 1883. Fishery Bulletin, 4(1):??, 1884. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/report-upon-shad-and-
herring-fisheries-potomac-river-1883;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb4.2.pdf.

Whitcombe:1884:NAF

[386] W. P. Whitcombe. Notes on the acclimatization of fish in Victoria, Australia. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/notes-acclimatization-fish-victoria-australia;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb4.10.pdf.

Habersham:1884:NUM

[387] William Ney A. Habersham. Note on the use of the male salmon hook, and the run of 1883. Fishery Bulletin, 4(1):52, 1884. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/note-use-male-salmon-
hook-and-run-1883;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb4.18.pdf.

Chapman:1884:HSH

[388] Pearson Chapman, Sr. Habits of the shad and herring, as they appear in the Potomac River to one who has watched them for fifty years. Fishery Bulletin, 4(1):61–64, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
habits-shad-and-herring-they-appear-potomac-river-one-who-has-
watched-them-fifty-years;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb4.26.pdf.

Dunn:1884:NEG

[389] Matthias Dunn. Number of eggs in the Gadidæ. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
number-eggs-gadid%C3%A6;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb4.34.pdf.

Koons:1884:PIS

[390] B. F. Koons. Planting Irish shells — Helix aspersa Müller — at Wood’s Hole, Mass. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/planting-irish-shells-helix-aspersa-m%C3%BCller-
woods-holl-mass;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb4.42.pdf.

Bower:1884:ITW

[391] Seymour Bower. Instructions for taking whitefish eggs. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
instructions-taking-whitefish-eggs;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb4.50.pdf.

Garman:1884:RSS

[392] Samuel Garman. In regard to the ‘Sea-serpent’ of literature. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
regard-sea-serpent-literature;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb4.58.pdf.

Miller:1884:PIH

[393] John F. Miller. Proposed introduction of Hawaiian mullet into the United States. Fishery Bulletin, 4(1):??, 1884. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/proposed-introduction-hawaiian-
mullet-united-states;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb4.66.pdf.

Ryder:1884:CDE

[394] John A. Ryder. Carp do eat young fishes. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/carp-do-eat-young-fishes;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb4.74.pdf.

Ancher:1884:KKF

[395] Ernst Ancher. The kurren and keitel (fishing vessels) of the courland haff. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/kurren-and-keitel-fishing-vessels-courland-haff;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.82.pdf.

Baird:1884:SRS

[396] S. F. Baird. On the specimens received by the Smithsonian Institution from the Unites States Life Saving Service. Fishery Bulletin, 4 (1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/specimens-
received-smithsonian-institution-unites-states-life-saving-
service;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb4.90.pdf.

PF:1884:CSF

[397] La Petite France. Concerning the salmon-fisheries of Bretagne, France, and the need of fishways and restrictive legislation. Fishery Bulletin, 4(1): ??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/concerning-
salmon-fisheries-bretagne-france-and-need-fishways-and-restrictive-
legislation;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.98.pdf.

Smiley:1884:IAP

[398] Chas W. Smiley. The influence of artificial propagation upon production illustrated by the salmon work of the Sacramento River, California. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/influence-artificial-propagation-
upon-production-illustrated-salmon-work-sacramento-river;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb4.106.pdf.

Dowell:1884:CTO

[399] B. F. Dowell. Cultivating trout in Oregon. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/cultivating-
trout-oregon;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.114.pdf.

Worth:1884:RUP

[400] S. G. Worth. Report upon the propagation of striped bass at Weldon, N.C., in the spring of 1884. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/report-upon-
propagation-striped-bass-weldon-n-c-spring-1884;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.122.pdf.

Wood:1884:RTS

[401] W. M. Wood. Report of a trip by the steamer Fish Hawk to the St. Mary’s and St. John’s rivers to hatch shad. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/report-trip-steamer-
fish-hawk-st-marys-and-st-johns-rivers-hatch-shad;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.131.pdf.

Moseley:1884:CPP

[402] H. N. Moseley. A carnivorous plant preying on vertebrata. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/carnivorous-plant-preying-vertebrata;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.139.pdf.

McDonald:1884:CTP

[403] Marshall McDonald. California trout planted in Roanoke River in July, 1883, retaken in June, 1884. Fishery Bulletin, 4(1):??, 1884. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/california-trout-planted-
roanoke-river-july-1883-retaken-june-1884;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb4.147.
pdf.

Hessel:1884:SDC

[404] Rud Hessel. Snakes destructive to carp. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/snakes-destructive-
carp;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb4.155.pdf.

Jones:1884:SC

[405] J. F. Jones. The speckled catfish. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/speckled-catfish;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb4.163.pdf.

Hovey:1884:SF

[406] H. C. Hovey. The sturgeon fishery. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/sturgeon-fishery;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb4.172.pdf.

Heath:1884:ECF

[407] Neil Heath. Effect of cold on fishes. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/effect-cold-fishes;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb4.180.pdf.

Atkins:1884:TQC

[408] Charle G. Atkins. Ten questions concerning the habits and breeding of landlocked salmon, with replies. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/ten-questions-concerning-habits-
and-breeding-landlocked-salmon-replies;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb4.188.
pdf.

DF:1884:MBM

[409] Deutsche Fischerei-Zeitung. Martin Brandt’s method of preserving fresh fish and other articles of food. Fishery Bulletin, 4(1):395– 397, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/martin-
brandts-method-preserving-fresh-fish-and-other-articles-food;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb4.196.pdf.

BF:1884:DDC

[410] Bayerische Fischerei-Zeitung. Discussion at the Dresden Conference in 1883, of the kinds of fish eggs to be obtained from the United States. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/discussion-dresden-conference-1883-kinds-fish-eggs-
be-obtained-united-states;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb4.204.pdf.

Dunning:1884:THT

[411] Philo Dunning. Two hundred tons of dead fish, mostly perch, at Lake Mendota, Wisconsin. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/two-hundred-tons-dead-fish-mostly-perch-

lake-mendota-wisconsin;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb4.212.pdf.

Collins:1884:AHN

[412] J. W. Collins. On the abundance of halibut near Iceland. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/abundance-halibut-near-iceland;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb4.220.
pdf.

VanMater:1884:OBV

[413] J. H. Van Mater. Occurrence of Balistes vetula on the New Jersey coast. Fishery Bulletin, 4(1):??, 1884. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/occurrence-balistes-vetula-new-
jersey-coast;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.3.pdf.

Worth:1884:SSC

[414] S. G. Worth. The selection of sites and the construction of carp ponds. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/selection-sites-and-construction-
carp-ponds;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.11.pdf.

Banmeyer:1884:ALT

[415] ?. Banmeyer. American lake trout and whitefish in France. Fishery Bulletin, 4(1):52, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/american-lake-trout-and-whitefish-france;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.19.pdf.

Dowell:1884:ETC

[416] B. F. Dowell. Efforts in trout culture. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/efforts-trout-
culture;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb4.27.pdf.

Jordan:1884:FFK

[417] David S. Jordan. The fishes of the Florida keys. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/fishes-florida-keys;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb4.35.pdf.

Smart:1884:CMF

[418] ?. Smart, ?. Goldsmith, and ?. Johnson. A Chinese method of fish culture. Fishery Bulletin, 4(1):88, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/chinese-method-fish-culture;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.43.pdf.

Maitland:1884:ELS

[419] James G. Maitland. Exchange of landlocked salmon eggs from Maine, for Loch Leven trout ova from Scotland. Fishery Bulletin, 4(1): ??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/exchange-
landlocked-salmon-eggs-maine-loch-leven-trout-ova-scotland;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb4.51.pdf.

Martin:1884:NCGa

[420] S. J. Martin. Notes on the cod gill-net fisheries of Gloucester, Mass., 1883–1884. Fishery Bulletin, 4(1):??, 1884. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/notes-cod-gill-net-
fisheries-gloucester-mass-1883-1884;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb4.59.pdf.

Merchant:1884:INS

[421] George Merchant, Jr. The incipiency of night-seining for mackerel. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
incipiency-night-seining-mackerel;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb4.67.pdf.

Baird:1884:ARE

[422] G. W. Baird. Annual report on the electric lighting of the United States steamer Albatross, December 31, 1883. Fishery Bulletin, 4(1):

??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/annual-report-
electric-lighting-united-states-steamer-albatross-december-
31-1883;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb4.75.pdf.

vondemBorne:1884:WTT

[423] Max von dem Borne. Wooden tank for the transportation of living fish. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/wooden-tank-transportation-living-fish;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb4.
83.pdf.

Stone:1884:WST

[424] Livingston Stone. Weights of salmon taken at McCloud River Station in 1880. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/weights-salmon-taken-mccloud-river-station-1880;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.91.pdf.

Miller:1884:APF

[425] W. B. Miller. An act to prohibit fishing by steam vessels with shirred or purse seines in any of the waters within the jurisdiction of the state of New Jersey. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/act-prohibit-fishing-
steam-vessels-shirred-or-purse-seines-any-waters-within-jurisdiction;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb4.99.pdf.

Atwater:1884:TIN

[426] W. O. Atwater. Tables illustrative of the nutritive value of fish. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/tables-illustrative-nutritive-value-fish;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.107.pdf.

Sterling:1884:NGL

[427] E. Sterling. Notes on the Great Lake fisheries, depletion of black bass, etc. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY.

ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/notes-great-lake-fisheries-depletion-
black-bass-e;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.115.pdf.

TFR:1884:CPB

[428] Texas Farm and Ranch. The carp ponds belonging to the State of Texas. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/carp-ponds-belonging-state-texas;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb4.123.pdf.

Wood:1884:RSF

[429] W. M. Wood. Reconnaissance of the shad-fisheries of winyaw bay and its tributaries by the steamer Fish Hawk. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/reconnaissance-shad-
fisheries-winyaw-bay-and-its-tributaries-steamer-fish-hawk;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb4.132.pdf.

Moseley:1884:FEU

[430] H. N. Moseley. The fish-eating utricularia, or bladderwort. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/fish-
eating-utricularia-or-bladderwort;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb4.140.pdf.

Merriam:1884:FLC

[431] C. Hart Merriam. The fish of Lake Champlain. Fishery Bulletin, 4 (1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fish-lake-
champlain;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.148.pdf.

Carpenter:1884:WMS

[432] Charles Carpenter. What muskrats sometimes eat. Fishery Bulletin, 4 (1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/what-muskrats-
sometimes-eat;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.156.pdf.

Brumme:1884:GRO

[433] Brumme. Great results obtained with little water. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/great-
results-obtained-little-water;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb4.165.pdf.

vonScherzer:1884:CS

[434] Karl von Scherzer. The cultivation of the sea. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
cultivation-sea;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb4.173.pdf.

NF:1884:SCL

[435] Norsk Fiskeritidende. The Scotch cod and ling fisheries. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
scotch-cod-and-ling-fisheries;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb4.181.pdf.

vondemBorne:1884:RBB

[436] Max von dem Borne. Report on black bass sent from America to Germany in 1883. Fishery Bulletin, 4(1):??, 1884. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/report-black-bass-sent-america-
germany-1883;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.189.pdf.

Scudder:1884:LBF

[437] Charles W. Scudder. A lists of the blank forms, circulars, and minor publications of the United States Fish Commission, from August 1, 1883, to August 1, 1884. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic).  URL https://spo.nmfs.noaa.gov/content/lists-
blank-forms-circulars-and-minor-publications-united-states-
fish-commission-august-1;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb4.197.pdf.

Rathbun:1884:NDL

[438] Richard Rathbun. Notes on the decrease of lobsters. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
notes-decrease-lobsters;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb4.205.pdf.

Forbes:1884:DFF

[439] S. A. Forbes. Destruction of fish-food by the bladderwort (Uticularia). Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/destruction-fish-food-bladderwort-uticularia;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb4.213.pdf.

Hoffman:1884:ASW

[440] R. E. Hoffman. Artificial sea-water for aquaria. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/artificial-sea-water-
aquaria;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb4.221.pdf.

Fuller:1884:CES

[441] Thomas Fuller. Carp in England in the Seventeenth Century. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/carp-england-seventeenth-century;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.4.pdf.

Ryder:1884:SPC

[442] John A. Ryder. On a skin parasite of the Cunner (Ctenolabrus aspersus). Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/skin-parasite-cunner-ctenolabrus-aspersus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb4.12.pdf.

Hector:1884:FNZ

[443] James Hector. The fisheries of New Zealand. Fishery Bulletin, 4(1): ??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/fisheries-new-
zealand;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb4.20.pdf.

Carr:1884:NSF

[444] T. F. Robertson Carr. Notes on the Scotch fisheries. Fishery Bulletin, 4 (1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/notes-scotch-
fisheries;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.28.pdf.

Kenworthy:1884:FQT

[445] C. J. Kenworthy. Food qualities of Tarpum (Megalops). Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/food-qualities-tarpum-megalops;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb4.36.pdf.

Finely:1884:SOW

[446] Chas I. Finély. Shad in Oregon waters. A new salmon hatchery. Fishery Bulletin, 4(1):88, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/shad-
oregon-waters-new-salmon-hatchery;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb4.44.pdf.

vondemBorne:1884:SFC

[447] Max von dem Borne. Success of fish-culture. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/success-fish-culture;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb4.52.pdf.

Fisher:1884:SCS

[448] Wm. J. Fisher. Statement of the catch of the several companies engaged in the salmon fisheries in Kadiak District, Alaska Territory, during the year 1883. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/statement-catch-several-companies-engaged-
salmon-fisheries-kadiak-district-alaska-territory;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.60.pdf.

Hearder:1884:WHN

[449] William Hearder. Waterproofing for herring-nets. Fishery Bulletin, 4(1): ??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/waterproofing-

herring-nets;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.68.pdf.

Brakeley:1884:PCP

[450] John H. Brakeley. Plants for carp ponds. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/plants-carp-
ponds;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb4.76.pdf.

Bottemanne:1884:PSO

[451] C. J. Bottemanne. Penning of salmon in order to secure their eggs. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
penning-salmon-order-secure-their-eggs;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb4.84.pdf.

Scudder:1884:VGC

[452] Charles W. Scudder. Vitality of German carp, and restoration of some apparently dead. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/vitality-german-carp-and-restoration-some-
apparently-dead;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb4.92.pdf.

dHomergue:1884:BOM

[453] Louis C. d’Homergue. Brief of the objections made before Leon Abbott, governor of New Jersey, to the ‘An act to prohibit fishing by steam vessels with shirred or purse seines in any of the waters within the jurisdiction of the state of New Jersey’. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/brief-objections-made-leon-abbott-governor-new-
jersey-act-prohibit-fishing-steam-vessels;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb4.100.
pdf.

Butcher:1884:HAS

[454] E. Z. Butcher. How to avoid a soft or muddy taste of carp. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/how-
avoid-soft-or-muddy-taste-carp;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb4.108.pdf.

vondemBorne:1884:SGL

[455] Max von dem Borne. Spawning in Germany of the large-mouthed black bass sent from the United States in 1882. Fishery Bulletin, 4 (1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/spawning-
germany-large-mouthed-black-bass-sent-united-states-1882;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb4.116.pdf.

Coxe:1884:OBO

[456] Frank Coxe. Opening the broad and other rivers of North Carolina to shad, bass, etc. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/opening-broad-and-other-rivers-north-
carolina-shad-bass-e;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb4.124.pdf.

Ferguson:1884:ERI

[457] T. B. Ferguson. Extracts from a report of investigations of the shad fisheries and rivers south of Charleston, S.C., with a view to establishing stations for artificial propagation. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/extracts-report-investigations-
shad-fisheries-and-rivers-south-charleston-s-c-view;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.133.pdf.

McDonald:1884:MSR

[458] Marshall McDonald. Memorandum of some results of fish culture already attained. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/memorandum-some-results-fish-culture-
already-attained;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb4.141.pdf.

Loomis:1884:LSC

[459] Watts T. Loomis. A landlocked salmon caught in erie canal. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/landlocked-salmon-caught-erie-canal;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb4.
149.pdf.

Elliott:1884:DCM

[460] Henry W. Elliott. The destruction of carp by the muskrat (Fiber zibethicus) — methods of trapping the rodent. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
destruction-carp-muskrat-fiber-zibethicus-methods-trapping-
rodent;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb4.157.pdf.

Pierce:1884:SDW

[461] H. H. Pierce. Some of the difficulties which confront oyster breeders. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/some-difficulties-which-confront-
oyster-breeders;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb4.166.pdf.

Lord:1884:FDL

[462] Henry W. Lord. The fish of Devil’s Lake, Dakota. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fish-devils-lake-dakota;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb4.174.pdf.

Ryder:1884:ACO

[463] John A. Ryder. On apparatus for collecting oyster spat. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/apparatus-collecting-oyster-spat;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.182.pdf.

Smiley:1884:ALS

[464] Chas W. Smiley. Arrangement with the life-saving service and the light-house board for collecting whales, porpoises, sharks, and strange forms of marine life. Fishery Bulletin, 4(1):??, 1884. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/arrangement-life-saving-service-
and-light-house-board-collecting-whales-porpoises-sharks;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb4.190.pdf.

Clark:1884:NFG

[465] A. Howard Clark. Notes on the fisheries of Gloucester, Mass. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
notes-fisheries-gloucester-mass-1;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb4.198.pdf.

Stover:1884:RCA

[466] E. S. Stover. Rearing carp in alkaline water. Fishery Bulletin, 4(1): ??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/rearing-carp-
alkaline-water;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.206.pdf.

Martin:1884:NFGa

[467] S. J. Martin. Notes on the fisheries of Gloucester, Mass. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
notes-fisheries-gloucester-mass-3;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb4.214.pdf.

Goode:1884:OIW

[468] G. Brown Goode. The oyster industry of the world. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/oyster-industry-world;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb4.222.pdf.

Collins:1884:MMW

[469] J. W. Collins. Movements of mackerel in winter. Fishery Bulletin, 4 (1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/movements-
mackerel-winter;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb4.5.pdf.

Ryder:1884:JOG

[470] John A. Ryder. Journal of operations on the grounds of the Eastern Shore Oyster Company, on Chincoteague Bay, near Stockton, Md., during the summer of 1883. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/journal-operations-
grounds-eastern-shore-oyster-company-chincoteague-bay-near-

stockton-md;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.13.pdf.

Proctor:1884:MM

[471] Richard A. Proctor. A marine monster. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/marine-monster;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.21.pdf.

Smiley:1884:WFC

[472] Chas W. Smiley. What fish culture has first to accomplish. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/what-
fish-culture-has-first-accomplish;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb4.29.pdf.

Hoxie:1884:OMF

[473] Walter Hoxie. Occurrence of mullet in fresh water. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
occurrence-mullet-fresh-water;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb4.37.pdf.

Martin:1884:NFGb

[474] S. J. Martin. Notes on the fisheries of Gloucester, Mass. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/notes-fisheries-gloucester-mass;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb4.
45.pdf.

Mather:1884:NCS

[475] Fred Mather. Notes on cod, shrimp, etc., at cold spring harbor. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/notes-cod-shrimp-e-cold-spring-harbor;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.53.pdf.

Phillips:1884:SNM

[476] Barnet Phillips. Some notes on the mullet fisheries. Fishery Bulletin, 4 (1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/some-notes-
mullet-fisheries;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb4.61.pdf.

Dukehard:1884:TSS

[477] J. P. Dukehard. Transfer of soft-shell terrapin from the Ohio to the Potomac River. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/transfer-soft-shell-terrapin-ohio-potomac-
river;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb4.69.pdf.

Huske:1884:RSW

[478] C. J. Huske. Report on the shad work in South Carolina in 1883 — transportation of shad eggs on trays. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/report-shad-work-south-carolina-
1883-transportation-shad-eggs-trays;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb4.77.pdf.

Atkins:1884:MRI

[479] Charles G. Atkins. Memoranda relative to inclosures for the confinements of salmon, drawn from experience at Bucksport, Penobscot River, Maine. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/memoranda-relative-inclosures-confinements-salmon-drawn-
experience-bucksport-penobscot;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb4.85.pdf.

Collins:1884:LLP

[480] J. W. Collins. Loss of life and property in the Gloucester fisheries. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/loss-
life-and-property-gloucester-fisheries;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb4.93.pdf.

Greenfield:1884:RUR

[481] Alfred Greenfield. Report upon the receipt and hatching of American whitefish ova and planting of the fry in Australia. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/report-upon-receipt-and-hatching-american-
whitefish-ova-and-planting-fry-australia;   https://spo.nmfs.

noaa.gov/sites/default/files/pdf-content/fish-bull/fb4.101.
pdf.

Hamlen:1884:RFR

[482] Wm Hamlen. Reconnaissance of Florida rivers, with a view of shad hatching. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/reconnaissance-florida-rivers-view-shad-hatching;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb4.109.pdf.

Stearns:1884:TCO

[483] R. E. C. Stearns. Transportation of clams and oysters. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/transportation-clams-and-oysters;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.117.pdf.

Winslow:1884:MPC

[484] Francis Winslow. Memorandum of the present condition and future needs of the oyster industry. Fishery Bulletin, 4(1):??, 1884. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/memorandum-present-condition-and-
future-needs-oyster-industry;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb4.125.pdf.

Collins:1884:NDM

[485] J. W. Collins. Note on the destruction of mackerel by dogfish. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
note-destruction-mackerel-dogfish;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb4.134.pdf.

Pierce:1884:NBM

[486] H. D. Pierce. Notes on the bluefish, mortality of Florida fishes, etc. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/notes-bluefish-mortality-florida-fishes-e;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb4.142.pdf.

Stearns:1884:PCF

[487] Silas Stearns. On the position and character of the fishing ground of the Gulf of Mexico. Fishery Bulletin, 4(1):??, 1884. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/position-and-character-fishing-
ground-gulf-mexico;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb4.150.pdf.

Merriam:1884:MRF

[488] C. Hart Merriam. The musk-rat as a fish-eater. Fishery Bulletin, 4(1): ??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/musk-rat-fish-
eater;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb4.158.pdf.

Finn:1884:CHL

[489] W. Finn. Can herring live and increase in inclosed waters? Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/can-
herring-live-and-increase-inclosed-waters;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb4.167.
pdf.

Baldwin:1884:NFF

[490] Charles H. Baldwin. Notes on the fish and fisheries of Japan. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/notes-fish-and-fisheries-japan;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb4.175.
pdf.

Haack:1884:TKR

[491] ?. Haack. Trapping kingfishers, rodents, and other enemies of trout. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/trapping-kingfishers-rodents-and-other-enemies-
trout;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb4.183.pdf.

NF:1884:ULS

[492] Norsk Fiskeritidende. Use of light in sea-fishing. Fishery Bulletin, 4(1): ??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/use-light-sea-
fishing;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb4.191.pdf.

Martin:1884:NFGc

[493] S. J. Martin. Notes on the fisheries of Gloucester, Mass. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
notes-fisheries-gloucester-mass-2;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb4.199.pdf.

Collins:1884:SMGb

[494] J. W. Collins. On the scarcity of mackerel in the Gulf of St. Lawrence. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/scarcity-mackerel-gulf-st-lawrence;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb4.207.pdf.

Harrell:1884:TGT

[495] J. Dock Harrell. Trapping gaspereau in Tangipahoa River. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/trapping-gaspereau-tangipahoa-river;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb4.
215.pdf.

Smiley:1884:BNUd

[496] Chas W. Smiley. Brief notes upon fish and fisheries. Fishery Bulletin, 4(1):469–470, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/brief-notes-upon-fish-and-fisheries-2;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.223.pdf.

Collins:1884:LS

[497] J. W. Collins. A large squid. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/large-squid;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb4.6.pdf.

Friedlaender:1884:NMF

[498] Oscar O. Friedlaender. Notes on the menhaden fishing of 1883. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/notes-
menhaden-fishing-1883;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb4.14.pdf.

Tarr:1884:RGH

[499] R. S. Tarr. Return to Gloucester Harbor of the young codfish hatched by the United States Fish Commission. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/return-gloucester-harbor-young-codfish-hatched-
united-states-fish-commission;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb4.22.pdf.

Manley:1884:PE

[500] J. J. Manley. Pisciculture in England. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/pisciculture-england;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb4.30.pdf.

Collins:1884:IFH

[501] J. W. Collins. Inauguration of the frozen herring trade. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
inauguration-frozen-herring-trade;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb4.38.pdf.

Brocchi:1884:RCO

[502] P. Brocchi. Report on the condition of oyster culture in France in 1881. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/report-condition-oyster-culture-
france-1881;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.46.pdf.

Hudson:1884:CCO

[503] George A. Hudson. Carp caught on Ogeechee River. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/carp-caught-ogeechee-river;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb4.54.pdf.

Maslieurat-Lagemard:1884:ASQ

[504] ?. Maslieurat-Lagémard. Acclimatization of Salmo quinnat in France. Fishery Bulletin, 4(1):144, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/acclimatization-salmo-quinnat-france-0;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb4.
70.pdf.

Valery-Mayot:1884:ASQ

[505] ?. Valéry-Mayot. Acclimatization of Salmo quinnat in France. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/acclimatization-salmo-quinnat-france;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.62.pdf.

Huske:1884:RCT

[506] C. J. Huske. Report on California trout distribution in South Carolina in 1883. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/report-california-trout-distribution-
south-carolina-1883;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb4.78.pdf.

Finely:1884:FRR

[507] Chas I. Finély. Further report of R. D. Hume’s Salmon Hatchery, Oregon. Fishery Bulletin, 4(1):??, 1884. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/further-report-r-d-humes-salmon-
hatchery-oregon;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb4.86.pdf.

Forbes:1884:LLP

[508] R. B. Forbes. Loss of life and property in the fisheries. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/loss-life-and-property-fisheries;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.94.pdf.

Martin:1884:NCGb

[509] S. J. Martin. Notes on the cod gill-net fisheries of Gloucester, Mass., 1883–1884. Fishery Bulletin, 4(1):??, 1884. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/notes-cod-gill-net-
fisheries-gloucester-mass-1883-1884-0;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb4.102.pdf.

Thompson:1884:NBE

[510] Edward H. Thompson. Note on the breeding of eels. Fishery Bulletin, 4 (1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/note-breeding-
eels;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb4.110.pdf.

Warner:1884:CFC

[511] J. S. Warner. Catching fish in a creek in Tennessee by a water snake. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/catching-fish-creek-tennessee-water-snake;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb4.118.pdf.

Ryder:1884:RRP

[512] John A. Ryder. Report respecting the present condition and future prospects at St. Jerome Creek for the work of oyster culture. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/report-respecting-present-condition-and-future-
prospects-st-jerome-creek-work-oyster;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb4.126.pdf.

Martin:1884:NFGd

[513] S. J. Martin. Notes on the fisheries of Gloucester, Mass. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
notes-fisheries-gloucester-mass-0;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb4.135.pdf.

Redding:1884:CCI

[514] Joseph D. Redding. Character of the carp introduced by Capt. Henry Robinson about 1830. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY.

ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/character-carp-introduced-capt-henry-
robinson-about-1830;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb4.143.pdf.

Robinson:1884:CST

[515] W. Russell Robinson. A California salmon taken in James River. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/california-salmon-taken-james-river;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb4.
151.pdf.

Raveret-Wattel:1884:NDA

[516] C. Raveret-Wattel. Notes on a disease affecting crawfish in Germany. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/notes-disease-affecting-crawfish-germany;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.159.pdf.

Curtis:1884:FNP

[517] J. E. Curtis. Fish in the national park and tributaries of Snake River — propagation of whitefish. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/fish-national-park-
and-tributaries-snake-river-propagation-whitefish;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.168.pdf.

Sheley:1884:DFC

[518] G. A. Sheley. Destruction of fish caused by nets of small mesh in Lake Michigan. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/destruction-fish-caused-nets-small-mesh-lake-michigan;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb4.176.pdf.

Hughes:1884:EPS

[519] Smith E. Hughes. Experiments in penning sea-fish. Fishery Bulletin, 4 (1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/experiments-

penning-sea-fish;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb4.184.pdf.

Hinkelmann:1884:MLE

[520] ?. Hinkelmann. The mode of life of eels. Fishery Bulletin, 4(1): 389–390, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/mode-
life-eels;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.192.pdf.

Herndon:1884:SRO

[521] C. G. Herndon. Sanitary report on Old Providence Island, United States of Columbia. Fishery Bulletin, 4(1):??, 1884. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/sanitary-report-old-providence-
island-united-states-columbia;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb4.200.pdf.

Carr:1884:NSHa

[522] T. F. Robertson Carr. Notes on the Scotch herring fisheries. Fishery Bulletin, 4(1):60, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/notes-scotch-herring-fisheries;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb4.24.pdf.

Carr:1884:NSHb

[523] T. F. Robertson Carr. Notes on the Scotch herring fisheries. Fishery Bulletin, 4(1):431, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/notes-scotch-herring-fisheries-0;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.208.pdf.

Harz:1884:MFT

[524] C. O. Harz. On manufactured food for trout and carp. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/manufactured-food-trout-and-carp;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.216.pdf.

Anonymous:1884:I

[525] Anonymous. Index. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/index-42;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb4.index_.pdf.

Anonymous:1884:C

[526] Anonymous. Contents. Fishery Bulletin, 4(1):??, 1884. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/contents-487;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb4.contents.
pdf.

Wood:1884:TOH

[527] W. M. Wood. The transplanting of one hundred lobsters from the eastern part of Long Island to Chesapeake Bay. Fishery Bulletin, 4(1): ??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/transplanting-
one-hundred-lobsters-eastern-part-long-island-chesapeake-bay;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb4.7.pdf.

M:1884:MCC

[528] McMenamin & Co. Method of catching crabs. Fishery Bulletin, 4 (1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/method-
catching-crabs;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.15.pdf.

Collins:1884:SOC

[529] J. W. Collins. Some observations on the cod gill-net fisheries and preservatives for nets. Fishery Bulletin, 4(1):??, 1884. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/some-observations-cod-gill-net-
fisheries-and-preservatives-nets;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb4.23.pdf.

Kensington:1884:CSF

[530] E. T. Kensington. Composition of some of the food-fishes. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
composition-some-food-fishes;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb4.31.pdf.

MBFC:1884:MUD

[531] Michigan Board of Fish Commissioners. Minute upon the death of Oren M. Chase, George W. Armstrong, and Charles H. Brownell. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
minute-upon-death-oren-m-chase-george-w-armstrong-and-charles-
h-brownell;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.39.pdf.

Meek:1884:NCE

[532] Seth E. Meek. A note on the Cuban eel. Fishery Bulletin, 4 (1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/note-
cuban-eel;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.47.pdf.

Palmer:1884:CAF

[533] B. D. Palmer. Carp appear February 7, take the hook, and are excellent eating. Fishery Bulletin, 4(1):124, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/carp-appear-february-7-take-hook-and-
are-excellent-eating;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb4.55.pdf.

ER:1884:DOB

[534] Evening Register. Depredations to oyster beds by star-fish. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
depredations-oyster-beds-star-fish;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb4.63.pdf.

Phillips:1884:NTG

[535] Barnet Phillips. Notes of a trip in the Gulf of Mexico. Fishery Bulletin, 4(1):144, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/notes-
trip-gulf-mexico;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb4.71.pdf.

Huske:1884:SFE

[536] C. J. Huske. Shad-fishing on the Edisto River. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/shad-fishing-edisto-
river;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb4.79.pdf.

Collins:1884:WCS

[537] J. W. Collins. What codfish sometimes swallow. Fishery Bulletin, 4(1): ??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/what-codfish-
sometimes-swallow;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb4.87.pdf.

Peirce:1884:RAD

[538] Milton P. Peirce. Resuscitation of apparently dead carp. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
resuscitation-apparently-dead-carp;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb4.95.pdf.

Ewart:1884:NAF

[539] J. Cossar Ewart. On the natural and artificial fertilizations of seaherring eggs. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/natural-and-artificial-fertilizations-sea-
herring-eggs;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.103.pdf.

Nowicki:1884:APM

[540] M. Nowicki. Anton Pintsch’s movable fish-way. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/anton-
pintschs-movable-fish-way;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb4.111.pdf.

Harris:1884:RUSb

[541] Gwynn Harris. Report upon the shad and herring fisheries of the Potomac River for 1884. Fishery Bulletin, 4(1):??, 1884. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/report-upon-shad-and-
herring-fisheries-potomac-river-1884;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb4.119.pdf.

Collins:1884:OCG

[542] J. W. Collins. On the occurrence of corals in the Grand Banks. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
occurrence-corals-grand-banks;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb4.127.pdf.

Crittenden:1884:CAH

[543] A. R. Crittenden. Catching alewives with hooks bated with eels. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
catching-alewives-hooks-bated-eels;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb4.136.pdf.

Smiley:1884:SOU

[544] Chas W. Smiley. Several opinions upon how to catch carp. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/several-opinions-upon-how-catch-carp;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.144.pdf.

Lowell:1884:TEC

[545] James Russell Lowell. Thanks of the Executive Committee of the London International Fisheries Exhibition for the participation by the United States. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/thanks-executive-committee-london-international-
fisheries-exhibition-participation-united;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb4.152.
pdf.

Ryder:1884:FCF

[546] John A. Ryder. Floats for the so-called fattening of oysters. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
floats-so-called-fattening-oysters;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb4.160.pdf.

Smiley:1884:NSS

[547] Chas W. Smiley. Notes on the shad season of 1884, with references to other species. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/notes-shad-season-1884-references-other-species;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb4.169.pdf.

Winslow:1884:NUO

[548] Francis Winslow. Notes upon oyster experiments in 1883. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/notes-
upon-oyster-experiments-1883;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb4.177.pdf.

Quattlebaum:1884:MCC

[549] Paul Quattlebaum. Method of catching carp with a hook. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/method-catching-carp-hook;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb4.185.pdf.

NF:1884:WFD

[550] Norsk Fiskeritidende. The weight of fish in different conditions. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/weight-fish-different-conditions;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.193.pdf.

Casella:1884:TCM

[551] Louis P. Casella. Treatment of the Casella–Miller thermometer. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/treatment-casella-miller-thermometer;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.201.pdf.

True:1884:PFC

[552] Frederick W. True. Porpoise-fishing at Cape May, New Jersey. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/porpoise-fishing-cape-may-new-jersey;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.209.pdf.

Nicklas:1884:MFT

[553] Carl Nicklas. On manufactured food for trout and carp. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
manufactured-food-trout-and-carp-0;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb4.217.pdf.

Anonymous:1884:EI

[554] Anonymous. Explanation of illustrations. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/explanation-
illustrations;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.plates.pdf.

Duke:1884:FPC

[555] R. T. W. Duke. A four-pound carp lives eight hours out of water by being packed in kate moss. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/four-pound-
carp-lives-eight-hours-out-water-being-packed-kate-moss;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.8.pdf.

Collins:1884:SMB

[556] J. W. Collins. A search for mackerel off Block Island, Montauk, and Sandy Hook, in November, 1883. Fishery Bulletin, 4(1): ??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/search-
mackerel-block-island-montauk-and-sandy-hook-november-1883;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb4.16.pdf.

MacKrill:1884:GC

[557] Alfred MacKrill. A great carp. Fishery Bulletin, 4(1):75, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/great-carp;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb4.32.pdf.

Annin:1884:PDV

[558] James Annin, Jr. Poachers or destructive visitors of fish ponds. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
poachers-or-destructive-visitors-fish-ponds;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb4.
40.pdf.

BSd:1884:WLT

[559] Bulletin de la Société d’acclimatation. Whitefish, lake trout, and brook trout in France. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/whitefish-lake-trout-and-brook-trout-
france;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb4.48.pdf.

LaMotte:1884:CPR

[560] Alfred V. La Motte. Comision de Piscicultura de la Republica Mexicana. Fishery Bulletin, 4(1):124, 1884. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/comision-de-piscicultura-de-la-
republica-mexicana;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb4.56.pdf.

CJ:1884:HCC

[561] Chambers’ Journal. How to cook carp and tench. Fishery Bulletin, 4(1): ??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/how-cook-carp-
and-tench;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.64.pdf.

Baird:1884:RWB

[562] G. W. Baird. Report on the working of the boilers and engine of the United States Fish Commission steamer Albatross. Fishery Bulletin, 4 (1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/report-
working-boilers-and-engine-united-states-fish-commission-steamer-
albatross;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.72.pdf.

Hessel:1884:CES

[563] Rud. Hessel. Culture of edible snails. Fishery Bulletin, 4(1): ??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/culture-
edible-snails;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb4.80.pdf.

Hessel:1884:LC

[564] Rud Hessel. Leech culture. Fishery Bulletin, 4(1):??, 1884. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/leech-culture;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.88.pdf.

Smiley:1884:RRF

[565] Chas W. Smiley. Remarkable resuscitation of frozen carp. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/remarkable-resuscitation-frozen-carp;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb4.96.pdf.

McDonald:1884:SES

[566] Marshall McDonald. Shad eggs sent to Cold Spring Harbor, New York, to be hatched. Fishery Bulletin, 4(1):??, 1884. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/shad-eggs-sent-cold-
spring-harbor-new-york-be-hatched;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb4.104.pdf.

McDonald:1884:TCP

[567] Marshall McDonald. Transferring catfish from the Potomac to the Colorado River, Arizona. Fishery Bulletin, 4(1):??, 1884. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/transferring-catfish-
potomac-colorado-river-arizona;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb4.112.pdf.

Wheeler:1884:NMP

[568] L. T. Wheeler. A new method of protecting the eggs of carp and rearing the young. Fishery Bulletin, 4(1):??, 1884. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/new-method-protecting-eggs-carp-
and-rearing-young;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb4.120.pdf.

Atwater:1884:RAS

[569] W. O. Atwater. Report of analysis of a sample of fish guano made from salmon offal, by Mr. Joseph Spratt, of Victoria, British Columbia. Fishery Bulletin, 4(1):??, 1884. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
report-analysis-sample-fish-guano-made-salmon-offal-mr-joseph-
spratt-victoria-british;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb4.128.pdf.

Anonymous:1885:TSA

[570] Anonymous. Topical synopsis of articles and notes. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/topical-synopsis-articles-and-notes;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb5%
20contents.pdf.

Jordan:1885:HVF

[571] David S. Jordan. The habits and the value for food of the American channel catfish (Ictalurus punctuatus, Rafinesque). Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/habits-and-value-food-american-channel-catfish-
ictalurus-punctuatus-rafinesque;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb5.8.pdf.

Blackford:1885:RWO

[572] Eugene G. Blackford. Report of the work in an oyster investigation with the steamer Lookout. Fishery Bulletin, 5(1):??, 1885. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/report-work-oyster-investigation-
steamer-lookout;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb5.16.pdf.

Ryder:1885:RGO

[573] John A. Ryder. The rate of growth of oysters at Saint Jerome Creek Stations. Fishery Bulletin, 5(1):??, 1885. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/rate-growth-oysters-saint-jerome-
creek-stations;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb5.24.pdf.

Kidder:1885:BBR

[574] J. H. Kidder. The Bower–Barff rustless iron. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/bower-barff-rustless-

iron;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb5.32.pdf.

Fiedler:1885:MSS

[575] Judge Fiedler. The migration of salmon (Salmo salar, L.) in the Baltic. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/migration-salmon-salmo-salar-l-baltic;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.40.pdf.

DeCausans:1885:FCE

[576] Viscount De Causans. The fish-cultural establishment of Lake Saint-Front, Haute Loire, France. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/fish-cultural-establishment-lake-
saint-front-haute-loire-france;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb5.48.pdf.

MacGowan:1885:CCC

[577] D. J. MacGowan. Carp-culture in China. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/carp-culture-
china;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb5.56.pdf.

Seymour:1885:FCS

[578] Charles Seymour. Fish-culture in southern China. Fishery Bulletin, 5(1): ??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/fish-culture-
southern-china;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb5.64.pdf.

Hensen:1885:OMB

[579] V. Hensen. Organic matter in the Baltic. Fishery Bulletin, 5(1): ??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/organic-
matter-baltic;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb5.72.pdf.

SFC:1885:IFS

[580] San Francisco Chronicle. Italian fishermen in San Francisco. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/italian-fishermen-san-francisco;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb5.
80.pdf.

Buch:1885:SSH

[581] S. A. Buch. Spawning of spring herring near Norway. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/spawning-spring-herring-near-norway;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb5.
88.pdf.

Walker:1885:LSH

[582] C. H. Walker. Location of a salmon hatchery in Oregon. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/location-salmon-hatchery-oregon;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb5.
96.pdf.

Dunlap:1885:PNF

[583] M. E. Dunlap. Pound and net fishing at Erie, Pa. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/pound-and-
net-fishing-erie-pa.

Oesterbol:1885:FM

[584] A. Oesterbol. Fish and medusæ. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/fish-and-medus%C3%A6;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb5.112.pdf.

Southwick:1885:ASS

[585] J. M. K. Southwick. The appearance of sharks, squid, menhaden, and frost-fish. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/appearance-sharks-squid-menhaden-and-
frost-fish;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb5.120.pdf.

McDonald:1885:RSP

[586] Marshall McDonald. Report of shad propagation on the Potomac River during the season of 1885. Fishery Bulletin, 5(1): ??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/report-
shad-propagation-potomac-river-during-season-1885;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.128.pdf.

Garde:1885:SHD

[587] Aug Garde. Smoking herring in Denmark. Fishery Bulletin, 5(1): ??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/smoking-
herring-denmark;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb5.136.pdf.

DF:1885:NCE

[588] Deutsche Fischerei-Zeitung. Nuremberg carp exchange. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
nuremberg-carp-exchange;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb5.144.pdf.

Wilcox:1885:NEFa

[589] W. A. Wilcox. New England fisheries in August, 1885. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/new-
england-fisheries-august-1885;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb5.152.pdf.

Wray:1885:DFC

[590] Leonard Wray. A desire for fish-culture in Malacca. Fishery Bulletin, 5 (1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/desire-fish-
culture-malacca;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb5.160.pdf.

Ryder:1885:NMO

[591] John A. Ryder. Note on the male organs of the eel. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/note-
male-organs-eel;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb5.1_0.pdf.

Worth:1885:PSB

[592] S. G. Worth. Prospectus of striped bass operations for the season of 1885. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/prospectus-striped-bass-operations-season-1885;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb5.9.pdf.

Martin:1885:NFGa

[593] S. J. Martin. Notes on the fisheries of Gloucester, Mass. Fishery Bulletin, 5(1):205–208, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/notes-fisheries-glocester-mass;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb5.49.pdf.

Martin:1885:NFGb

[594] S. J. Martin. Notes on the fisheries of Gloucester, Mass. Fishery Bulletin, 5(1):57–58, 1885. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
notes-fisheries-gloucester-mass-4;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb5.17.pdf.

True:1885:RTL

[595] F. W. True. Report of a trip to Long Island in search of skeletons of the right whale (Balæna cisactica). Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/report-trip-long-island-search-
skeletons-right-whale-bal%C3%A6na-cisactica;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb5.
25.pdf.

Dannevig:1885:ROH

[596] G. M. Dannevig. Report of operations at the hatching establishment for fishes, arendal, 1884. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/report-operations-hatching-establishment-
fishes-arendal-1884;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb5.33.pdf.

Nye:1885:NUO

[597] Willard Nye, Jr. Notes upon the octopus, flying-fish, etc., taken during the Albatross cruise in January, 1884. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/notes-upon-octopus-
flying-fish-c-taken-during-albatross-cruise-january-1884;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb5.41.pdf.

Kemp:1885:SCI

[598] N. H. Kemp et al. The sardine canning industry of Eastport, Me. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/sardine-canning-industry-eastport-me;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.57.pdf.

Hessel:1885:NHG

[599] Rud Hessel. Notes on the habits of the golden die (Idus auratus). Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/notes-habits-golden-die-idus-auratus;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.65.pdf.

Schroeder:1885:HWA

[600] Seaton Schroeder. Hydrographic work of the Albatross in 1884. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/hydrographic-work-albatross-1884;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.73.pdf.

Donaldson:1885:RCR

[601] B. J. Donaldson. Raising carp in rice fields. Fishery Bulletin, 5(1): ??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/raising-carp-
rice-fields;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb5.81.pdf.

NF:1885:ISF

[602] Norsk Fiskeritidende. The Iceland shark-fisheries. Fishery Bulletin, 5(1): ??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/iceland-shark-
fisheries;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb5.89.pdf.

Jordan:1885:MST

[603] David S. Jordan. The mountain or salmon trout of Oregon. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/mountain-or-salmon-trout-oregon;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb5.
97.pdf.

Riddle:1885:RMR

[604] George W. Riddle. Restocking the Merrimac River with lamprey eels. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/restocking-merrimac-river-lamprey-eels;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb5.
105.pdf.

Wilcox:1885:NEFb

[605] W. A. Wilcox. New England fisheries in July, 1885. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/new-england-fisheries-july-1885;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb5.
113.pdf.

vomdemBorne:1885:PBC

[606] Max vom dem Borne. Putting bass into carp ponds. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/putting-
bass-carp-ponds;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb5.121.pdf.

Smith:1885:RSH

[607] James A. Smith. Report of shad-hatching operations by the steamer lookout, season of 1885. Fishery Bulletin, 5(1):??, 1885. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/report-shad-hatching-
operations-steamer-lookout-season-1885;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb5.129.
pdf.

Cary:1885:SJR

[608] H. H. Cary. The Saint John’s River as a shad stream. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/saint-
johns-river-shad-stream;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb5.137.pdf.

Thorsteinson:1885:AHF

[609] A. Thorsteinson. American halibut-fisheries near Iceland. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/american-halibut-fisheries-near-iceland;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.145.pdf.

Dannevig:1885:HLN

[610] G. M. Dannevig. Hatching lobsters in Norway. Fishery Bulletin, 5 (1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/hatching-
lobsters-norway;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb5.153.pdf.

Tanner:1885:UEL

[611] Z. L. Tanner. The use of the electric light in fishing. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/use-electric-light-fishing;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb5.161.
pdf.

True:1885:PFH

[612] F. W. True. The porpoise fishery of Hatteras, N.C. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
porpoise-fishery-hatterns-n-c;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb5.2.pdf.

Smiley:1885:HDS

[613] Chas W. Smiley. How to distinguish the sex of a carp. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/how-
distinguish-sex-carp;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb5.10.pdf.

Juel:1885:WFN

[614] Niels Juel. The whale fishery of Norway. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/whale-fishery-norway;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb5.18.pdf.

True:1885:NCM

[615] F. W. True. Notice of the capture of a male pigmy sperm-whale (Kogia breviceps), at Kitty Hawk, North Carolina. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/notice-capture-male-
pigmy-sperm-whale-kogia-breviceps-kitty-hawk-north-carolina;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb5.26.pdf.

Verrill:1885:HLW

[616] A. E. Verrill. How long will oysters live out of water. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/how-long-will-oysters-live-out-water;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.34.pdf.

Mosher:1885:PBF

[617] A. A. Mosher. Propagating buffalo-fish. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/propagating-buffalo-fish;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb5.42.pdf.

Brakeley:1885:NCF

[618] John H. Brakeley. Notes on carp and frog culture. Fishery Bulletin, 5(1): ??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/notes-carp-
and-frog-culture;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb5.50.pdf.

Murray:1885:YTD

[619] C. H. Murray. Young trout destroyed by mosquitoes. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/young-trout-destroyed-mosquitoes;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.58.pdf.

DHalloy:1885:FCG

[620] Leon D’Halloy. Fish-culture at Gonville, France. Fishery Bulletin, 5(1): ??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/fish-culture-
gonville-france;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb5.66.pdf.

Lyman:1885:LIW

[621] Theodore Lyman. Laws of inland waters. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/laws-inland-waters;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb5.74.pdf.

Stearns:1885:EFG

[622] Silas Stearns. Examination of the fisheries in the Gulf of Mexico. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
examination-fisheries-gulf-mexico;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb5.82.pdf.

Sellars:1885:AFG

[623] L. H. Sellars. Abundance of fish in the Gulf of Mexico. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/abundance-fish-gulf-mexico;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb5.90.pdf.

Riley:1885:WBD

[624] C. V. Riley. Water-beetles destroying carp. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/water-beetles-
destroying-carp;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb5.98.pdf.

Clark:1885:NCA

[625] Frank N. Clark. Natural as compared with artificial hatching of whitefish. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/natural-compared-artificial-hatching-whitefish;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb5.106.pdf.

Nye:1885:NTD

[626] Willard Nye, Jr. Notes taken during cruise of the Albatross to Grand Banks in June and July, 1885. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/notes-taken-during-
cruise-albatross-grand-banks-june-and-july-1885;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.114.pdf.

Collins:1885:EQS

[627] J. W. Collins. Edible qualities of smoked kingfish (Scomberomorus cavalla, Cuv.). Fishery Bulletin, 5(1):??, 1885. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/edible-qualities-smoked-kingfish-
scomberomorus-cavalla-cuv;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb5.122_0.pdf.

Piepmeyer:1885:ROU

[628] L. W. Piepmeyer. Report of the operations of the U. S. Fish Commission steamer Fish Hawk, in shad hatching, on the Delaware River, season of 1885. Fishery Bulletin, 5(1):??, 1885. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/report-operations-u-
s-fish-commission-steamer-fish-hawk-shad-hatching-delaware-
river-season;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb5.130.pdf.

Raveret-Wattel:1885:AFF

[629] C. Raveret-Wattel. American fish in France. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/american-fish-france;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb5.138.pdf.

Smiley:1885:SCE

[630] Chas W. Smiley. Sending catfish to Europe. Fishery Bulletin, 5 (1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/sending-

catfish-europe;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb5.146.pdf.

Anonymous:1885:RAA

[631] Anonymous. Resolution asking for the abrogation of the fishery clause of the Washington treaty, passed at a convention of New England fishermen, which met at Gloucester, December 27, 1884. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/resolution-
asking-abrogation-fishery-clause-washington-treaty-passed-convention-
new-england;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb5.154.pdf.

Smiley:1885:NUFa

[632] Chas W. Smiley. Notes upon fish and the fisheries. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/notes-upon-fish-and-fisheries-1;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb5.
162.pdf.

Stearns:1885:LF

[633] W. A. Stearns. The Labrador fisheries. Fishery Bulletin, 5(1): ??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/labrador-
fisheries;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb5.3.pdf.

Boeck:1885:HSD

[634] G. Boeck. How to strengthen the dunes. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/how-strengthen-dunes;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb5.11.pdf.

Smiley:1885:NUFb

[635] Chas W. Smiley. Notes upon the fish and the fisheries. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/notes-
upon-fish-and-fisheries;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb5.19.pdf.

DePoincy:1885:DFT

[636] J. W. De Poincy. Description of a fish-trap and oyster basin. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
description-fish-trap-and-oyster-basin;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb5.27.pdf.

Wilcox:1885:NEFc

[637] W. A. Wilcox. New England fisheries from January to May, 1885. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
new-england-fisheries-january-may-1885;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb5.35.pdf.

Jordan:1885:NFO

[638] David S. Jordan. Notes on fishes observed in Lake Superior. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/notes-
fishes-observed-lake-superior;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb5.43.pdf.

Gould:1885:TAF

[639] N. E. Gould. Time of approach of fish to the New England Coast. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/time-approach-fish-new-england-coast;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.51.pdf.

Raveret-Wattel:1885:AFB

[640] C. Raveret-Wattel. American fish being acclimatized in France. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/american-fish-being-acclimatized-france;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.59.pdf.

Collins:1885:FEG

[641] J. W. Collins. Fishing on an edge of the Grand Banks. Fishery Bulletin, 5 (1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/fishing-edge-
grand-banks;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb5.67.pdf.

Valery-Mayet:1885:HSE

[642] ?. Valéry-Mayet. Hatching salmon eggs at Montpellier, France, and trouble with fungus. Fishery Bulletin, 5(1):??, 1885. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/hatching-salmon-eggs-montpellier-
france-and-trouble-fungus;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb5.75.pdf.

Collins:1885:PSS

[643] John D. Collins. Propagation of salmon in Salmon River, Oswego County, New York. Fishery Bulletin, 5(1):??, 1885. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/propagation-salmon-salmon-river-
oswego-county-new-york;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb5.83.pdf.

Dunn:1885:FMP

[644] Matthias Dunn. Food of mackerel, pilchards, and herring. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/food-mackerel-pilchards-and-herring;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb5.
91.pdf.

Stearns:1885:ESF

[645] Robt. E. C. Stearns. Edible shell-fish found near Cuba. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
edible-shell-fish-found-near-cuba;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb5.99.pdf.

Atkins:1885:RSS

[646] Chas G. Atkins. Report on the schoodic salmon work of 1884–1885. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/report-schoodic-salmon-work-1884-1885;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.107.pdf.

Smiley:1885:NUFc

[647] Chas W. Smiley. Notes upon fish and the fisheries. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/notes-upon-fish-and-fisheries-0;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb5.
115.pdf.

Collins:1885:PPK

[648] J. W. Collins. Proposition to prepare kingfish by smoking. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/proposition-prepare-kingfish-smoking;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.123.pdf.

Worthington:1885:FFM

[649] John Worthington. Falling off in the fisheries of Malta. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/falling-
fisheries-malta;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb5.131.pdf.

Hodge:1885:CLP

[650] E. B. Hodge. Culture of landlocked and Penobscot salmon in New Hampshire. Fishery Bulletin, 5(1):??, 1885. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/culture-landlocked-and-penobscot-
salmon-new-hampshire;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb5.139.pdf.

Hessel:1885:NCP

[651] Rud Hessel. Notes from the carp ponds. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/notes-carp-
ponds;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb5.147.pdf.

Stone:1885:SST

[652] Livingston Stone. Segregation of the sexes of trout. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
segregation-sexes-trout;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb5.155.pdf.

Green:1885:DAR

[653] Loren W. Green. On a disease affecting the rainbow trout at Mc-Cloud River Station. Fishery Bulletin, 5(1):??, 1885. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/disease-affecting-rainbow-trout-
mccloud-river-station;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb5.164.pdf.

Stone:1885:HOF

[654] Livingston Stone. History of operations at the fish-hatching stations on the McCloud River, California, from the beginning, August, 1872, to October, 1884. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/history-operations-fish-hatching-stations-
mccloud-river-california-beginning-august-1872;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb5.
4.pdf.

Collins:1885:NEH

[655] J. W. Collins. Notes on the English herring and mackerel fisheries, and the methods of curing. Fishery Bulletin, 5(1):??, 1885. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/notes-english-herring-and-mackerel-
fisheries-and-methods-curing;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb5.12.pdf.

Clark:1885:RPY

[656] Frank N. Clark. Results of planting young whitefish in Lake Erie. what the fisherman thing of whitefish propagation. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/results-
planting-young-whitefish-lake-erie-what-fisherman-thing-whitefish-
propagation;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb5.20.pdf.

Ryder:1885:DMG

[657] John A. Ryder. On the development of the mammary glands and genitalia of the cetacea. Fishery Bulletin, 5(1):??, 1885. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/development-mammary-glands-and-
genitalia-cetacea;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb5.28.pdf.

Ryder:1885:RGC

[658] John A. Ryder. On the rate of growth of the common clam, and on a mode of obtaining the young of the giant clams of the Pacific Coast for the purpose of transplanting. Fishery Bulletin, 5(1):??, 1885. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/rate-growth-common-clam-and-mode-
obtaining-young-giant-clams-pacific-coast-purpose;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.36.pdf.

Harris:1885:RUS

[659] Gwynn Harris. Report upon the shad and herring fisheries of the Potomac River for 1885. Fishery Bulletin, 5(1):??, 1885. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/report-upon-shad-and-
herring-fisheries-potomac-river-1885;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb5.44.pdf.

Anderson:1885:NUS

[660] John Anderson. Notes upon the Scotch fish and fisheries. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/notes-upon-scotch-fish-and-fisheries;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.52.pdf.

Stearns:1885:FPF

[661] Silas Stearns. The fisheries of Pensacola, Florida. Fishery Bulletin, 5 (1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fisheries-
pensacola-florida;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb5.60.pdf.

Mobius:1885:RPC

[662] Karl Möbius. Report on planting Canadian oysters near the Island of Aaröe, in the Little Belt, November 6, 1884. Fishery Bulletin, 5 (1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/report-
planting-canadian-oysters-near-island-aar%C3%B6e-little-belt-
november-6-1884;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb5.68.pdf.

Ffennell:1885:HAF

[663] Henry Ffennell. Hatching American fish at South Kensington, and their introduction to English waters. Fishery Bulletin, 5(1): ??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/hatching-
american-fish-south-kensington-and-their-introduction-english-
waters;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb5.76.pdf.

vonK:1885:FCM

[664] Baron von K. Fish-culture as a means of improving the race of fish. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fish-culture-means-improving-race-fish;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb5.84.pdf.

NF:1885:EEN

[665] Norsk Fiskeritidende. Eels and eelpots in Norway. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
eels-and-eelpots-norway;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb5.92.pdf.

Mosher:1885:NFC

[666] A. A. Mosher. Notes on fish-culture in Iowa. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
notes-fish-culture-iowa;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb5.100.pdf.

Howell:1885:ICG

[667] D. Y. Howell. Injuries caused by gill-nets to whitefish. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/injuries-caused-gill-nets-whitefish;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb5.
108.pdf.

Halperine:1885:PP

[668] E. Halpérine. Piscivorous plants. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/piscivorous-plants;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb5.116.pdf.

PP:1885:HS

[669] Philadelphia Press. The harvest of the sea. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/harvest-sea;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb5.124.pdf.

Hamlen:1885:RSH

[670] William Hamlen. Report of shad-hatching operations at Battery Station, Havre de Grace, Md., season of 1885. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/report-shad-hatching-
operations-battery-station-havre-de-grace-md-season-1885;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb5.132.pdf.

Pittendrigh:1885:SFP

[671] George Pittendrigh. Statistics of the fisheries of the Province of British Columbia for 1884. Fishery Bulletin, 5(1):??, 1885. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/statistics-fisheries-
province-british-columbia-1884;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb5.140.pdf.

Smith:1885:ERT

[672] S. B. Smith. Efforts to raise trout. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/efforts-raise-trout;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb5.148.pdf.

Martin:1885:HCC

[673] Robert A. Martin. How to catch carp. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/how-catch-carp;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.156.pdf.

Anonymous:1885:I

[674] Anonymous. Index. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/index-43;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb5.index_.pdf.

Wood:1885:TLC

[675] W. M. Wood. Transplanting lobsters to the chesapeake- experiments upon the temperature they can endure. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/transplanting-
lobsters-chesapeake-experiments-upon-temperature-they-can-endure;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb5.5.pdf.

Smiley:1885:CLP

[676] Chas W. Smiley and Chas W. Scudder. Chronological list of pamphlet publications of the U. S. Fish Commission from its organization in 1871 to January, 1885. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/chronological-list-
pamphlet-publications-u-s-fish-commission-its-organization-
1871-january;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb5.13.pdf.

Smith:1885:NLH

[677] A. C. Smith. Notes on the lobster (Homarus americanus). Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/notes-lobster-homarus-americanus;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.21.pdf.

Day:1885:EET

[678] Francis Day. The effects of an elevated temperature on fishes. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/effects-elevated-temperature-fishes;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb5.
29.pdf.

Mather:1885:SFH

[679] Fred Mather. The shad fisheries of the Hudson. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/shad-
fisheries-hudson;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb5.37.pdf.

Hoek:1885:CEC

[680] P. P. C. Hoek. Comparative examination of cultivated and uncultivated oysters, with the view to determine the number which, during the first year, took part in reproduction. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/comparative-
examination-cultivated-and-uncultivated-oysters-view-determine-
number-which;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb5.45.pdf.

Harvie-Brown:1885:MBS

[681] J. A. Harvie-Brown. On the migration of birds in the spring and autumn of 1884. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/migration-birds-spring-and-autumn-1884;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.53.pdf.

delaCCerda:1885:PTS

[682] Juan de la C.Cerda. Proposed transmission of salmon eggs to chili and importation of Chilian species of fishes. Fishery Bulletin, 5(1): ??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/proposed-
transmission-salmon-eggs-chili-and-importation-chilian-species-
fishes;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb5.61.pdf.

Raveret-Wattel:1885:NCA

[683] C. Raveret-Wattel. Note on the culture of American salmon in France. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/note-culture-american-salmon-france;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb5.
69.pdf.

Borodin:1885:AIA

[684] N. Borodin. An attempt to impregnate artificially the eggs of Acipenser stellatus. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/attempt-impregnate-artificially-eggs-acipenser-
stellatus;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb5.77.pdf.

Johnson:1885:MSG

[685] George A. Johnson. Mud slough on the Grand Banks. Fishery Bulletin, 5 (1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/mud-slough-
grand-banks;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb5.85.pdf.

Day:1885:RGB

[686] Francis Day. On the rearing, growth, and breeding of salmon in fresh water in Great Britain. Fishery Bulletin, 5(1):??, 1885. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/rearing-growth-and-
breeding-salmon-fresh-water-great-britain;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb5.93.pdf.

McDonald:1885:EWP

[687] Marshall McDonald. Effect of waste products from Page’s ammoniacal works upon young shad fry. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/effect-waste-
products-pages-ammoniacal-works-upon-young-shad-fry;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.101.pdf.

Proctor:1885:MSS

[688] Richard A. Proctor. Monster sea-serpents. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/monster-sea-serpents;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb5.109.pdf.

Stearns:1885:NCP

[689] Robt E. Stearns. Note on the clams of the Pacific Coast. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/note-
clams-pacific-coast;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb5.117.pdf.

Adams:1885:SCO

[690] Emma H. Adams. Salmon canning in Oregon. Fishery Bulletin, 5 (1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/salmon-
canning-oregon;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb5.125.pdf.

NF:1885:SRN

[691] Norsk Fiskeritidende. Salting roe in Norway. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/salting-roe-norway;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb5.133.pdf.

Dickinson:1885:PPS

[692] Fred Dickinson. Petition to prohibit the seining of bluefish. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
petition-prohibit-seining-bluefish;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb5.141.pdf.

Clarke:1885:PTS

[693] H. B. Clarke. Propagation of trout by the southside club. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/propagation-trout-southside-club;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.149.pdf.

Hobbs:1885:SPS

[694] James R. Hobbs. Stranding of a pigmy sperm-whale. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/stranding-
pigmy-sperm-whale;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb5.157.pdf.

Farlow:1885:VPC

[695] W. G. Farlow. Vegetable parasites of codfish. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/vegetable-parasites-
codfish;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb6.1.pdf.

Ryder:1885:DCE

[696] John A. Ryder. Directions for collecting embiotocoid fish embryos. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/directions-collecting-embiotocoid-fish-embryos;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb5.6.pdf.

Martin:1885:NCG

[697] S. J. Martin. Notes on the cod gill-net fisheries of Gloucester, Mass., 1884–1885. Fishery Bulletin, 5(1):??, 1885. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/notes-cod-gill-net-
fisheries-gloucester-mass-1884-1885;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb5.14.pdf.

McDonald:1885:RPP

[698] Marshall McDonald. Report on the pollution of the Potomac River by the discharge of waste products from gas manufacture. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/report-pollution-potomac-river-discharge-waste-
products-gas-manufacture;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb5.22.pdf.

Collins:1885:RSG

[699] J. W. Collins. The red snapper grounds in the Gulf of Mexico. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/red-snapper-grounds-gulf-mexico;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb5.
30.pdf.

Anonymous:1885:APF

[700] Anonymous. An act to protect the fish in the Potomac River in the District of Columbia, and to provide a spawning-ground for shad and herring in the said Potomac River. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/act-protect-fish-potomac-river-

district-columbia-and-provide-spawning-ground-shad-and;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.38.pdf.

Wilcox:1885:NEFd

[701] W. A. Wilcox. New England fisheries in June, 1885. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/new-england-fisheries-june-1885;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb5.
46.pdf.

Swan:1885:RBC

[702] James G. Swan. Report on black cod of the North Pacific Ocean. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/report-black-cod-north-pacific-ocean;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.54.pdf.

Smithers:1885:CCC

[703] E. J. Smithers. Carp-culture in China. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/carp-culture-china-0;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb5.62.pdf.

vondemBorne:1885:DAF

[704] Max von dem Borne. Distribution of American fish and fish eggs by the German Fishery Association. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/distribution-american-fish-and-
fish-eggs-german-fishery-association;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb5.70.pdf.

Schwaab:1885:LFY

[705] ?. Schwaab. Live food for young fish. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/live-food-young-fish;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb5.78.pdf.

Bouchon-Brandely:1885:PFT

[706] G. Bouchon-Brandely. The pearl fisheries of Tahiti. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/pearl-
fisheries-tahiti;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb5.86.pdf.

Page:1885:ESU

[707] W. F. Page. Effect of sunlight upon shad eggs. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/effect-sunlight-upon-
shad-eggs;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb5.94.pdf.

Norny:1885:APR

[708] E. R. Norny. Artificial propagation of rockfish and eels. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/artificial-propagation-rockfish-and-eels;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.102.pdf.

Kinney:1885:PRF

[709] Aug C. Kinney. Proposed removal of fish-traps from Columbia River. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/proposed-removal-fish-traps-columbia-river;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.110.pdf.

Kidder:1885:MWR

[710] J. H. Kidder. Memorandum on water residues from cod-hatching station at Wood’s Hole. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/memorandum-water-residues-cod-hatching-station-
woods-holl;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb5.118.pdf.

Worth:1885:SFC

[711] S. G. Worth. Summary of fish-cultural work in North Carolina. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/summary-fish-cultural-work-north-carolina;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb5.126.pdf.

Sparks:1885:PP

[712] George L. Sparks. Porpoise products. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/porpoise-products;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb5.134.pdf.

Moore:1885:ICD

[713] C. B. Moore. Introduction of clams into Delaware Bay. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/introduction-clams-delaware-bay;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb5.
142.pdf.

Dannevig:1885:AHS

[714] G. M. Dannevig. Artificial hatching of salt-water fish and lobsters in Norway. Fishery Bulletin, 5(1):??, 1885. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/artificial-hatching-salt-water-
fish-and-lobsters-norway;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb5.150.pdf.

Duncan:1885:STH

[715] J. Barker Duncan. Salmon and trout hatcheries in Scotland. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/salmon-and-trout-hatcheries-scotland;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.158.pdf.

Anonymous:1885:EI

[716] Anonymous. Explanation of illustrations. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/explanation-
illustrations-0;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb5%20plates.pdf.

Bean:1885:NUE

[717] Tarleton H. Bean. Notes upon an exploration on Long Island Sound. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/notes-upon-exploration-long-island-sound;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.7.pdf.

Atkins:1885:PSS

[718] Chas G. Atkins. The Penobscot and Schoodic salmon work of 1884– 1885. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/penobscot-and-schoodic-salmon-work-1884-1885;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb5.15.pdf.

Garman:1885:POB

[719] Samuel Garman. Protecting the oyster beds from star-fish depredations. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/protecting-oyster-beds-star-fish-
depredations;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb5.23.pdf.

Rasch:1885:OCD

[720] H. Rasch, A. Landmark, and G. O. Sars. Opinions concerning the desirability of a hatchery for salt-water fish in Norway. Fishery Bulletin, 5(1):146–150, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/opinions-concerning-desirability-hatchery-salt-
water-fish-norway;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb5.31.pdf.

Ryder:1885:GCG

[721] John A. Ryder. On the green coloration of the gills and palps of the clam (Mya arenaria). Fishery Bulletin, 5(1):??, 1885. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/green-coloration-gills-and-palps-
clam-mya-arenaria;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb5.39.pdf.

Waage:1885:PB

[722] P. Waage. Preservation of bait. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/preservation-bait;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb5.47.pdf.

Collins:1885:UAC

[723] J. W. Collins. Unusual abundance of cod on Brown’s Bank. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
unusual-abundance-cod-browns-bank;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb5.55.pdf.

Shephard:1885:FCC

[724] Isaac Shephard. Fish-culture in China. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/fish-culture-
china;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb5.63.pdf.

Malmgren:1885:IER

[725] A. J. Malmgren. The injurious effects of rafting on the lake and river fisheries of Sweden. Fishery Bulletin, 5(1):??, 1885. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/injurious-effects-rafting-lake-
and-river-fisheries-sweden;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb5.71.pdf.

Dannevig:1885:SHL

[726] G. M. Dannevig. Success in hatching lobster eggs in Norway. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/success-hatching-lobster-eggs-norway;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.79.pdf.

Knudsen:1885:SFC

[727] Henrik Knudsen. Some facts concerning isinglass. Fishery Bulletin, 5(1): ??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/some-facts-
concerning-isinglass;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb5.87.pdf.

Green:1885:CUS

[728] Loren W. Green. Condition of the United States trout ponds. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/condition-united-states-trout-ponds;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb5.
95.pdf.

Annaniassen:1885:KFS

[729] Abr. Annaniassen.  Klip-fish at the Shetland Islands.  Fish-

ery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/klip-fish-shetland-islands;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb5.103.
pdf.

Whitcombe:1885:SWE

[730] W. P. Whitcombe. Shipment of whitefish eggs to the Ballarat Fish Acclimatization Society. Fishery Bulletin, 5(1):329, 1885. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/shipment-whitefish-eggs-ballarat-
fish-acclimatization-society;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb5.111.pdf.

Frey:1885:DWE

[731] E. Frey. Disposition of whitefish eggs sent to Switzerland. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
disposition-whitefish-eggs-sent-switzerland;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb5.
119.pdf.

Playfair:1885:FOA

[732] Lyon Playfair. A foreigner’s opinion of American fish-culture. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/foreigners-opinion-american-fish-culture;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.127.pdf.

Ryder:1885:AQA

[733] John A. Ryder. Answers to questions about fattening oysters. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/answers-questions-about-fattening-oysters;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb5.135.pdf.

Stover:1885:RCS

[734] E. S. Stover. Raising carp in a supposed alkaline water. Fishery Bulletin, 5(1):??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/raising-carp-supposed-alkaline-water;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb5.143.pdf.

Dannevig:1885:SA

[735] G. M. Dannevig. Salmonidæ in Australia. Fishery Bulletin, 5(1): ??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/salmonid%
C3%A6-australia;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb5.151.pdf.

Worthington:1885:FM

[736] John Worthington. The fisheries of Malta. Fishery Bulletin, 5(1): ??, 1885. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fisheries-
malta;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb5.159.pdf.

Anonymous:1886:EI

[737] Anonymous. Explanation of illustrations. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/explanation-
illustrations-1;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb6%20plates.pdf.

Wilcox:1886:NEFa

[738] W. A. Wilcox. New England fisheries in September, 1885. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/new-england-fisheries-september-1885;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb6.9.pdf.

vonStabrowski:1886:CCG

[739] Xaver von Stabrowski. Carp culture in Germany. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/carp-
culture-germany;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb6.17.pdf.

NF:1886:FF

[740] Norsk Fiskeritidende. Foreign fisheries in 1885. Fishery Bulletin, 6 (1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/foreign-
fisheries-1885;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb6.25.pdf.

Dempsey:1886:SSJ

[741] W. Dempsey. Shad in the Saint John’s River, Florida. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/shad-saint-johns-river-florida;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb6.33.pdf.

vondemBorne:1886:LCC

[742] Max von dem Borne. A large catch of carp. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/large-catch-carp;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb6.41.pdf.

Pancritius:1886:FDG

[743] P. Pancritius. On the food and digestion of German fish. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/food-and-digestion-german-fish;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb6.49.pdf.

Hubner:1886:NUI

[744] A. Hübner. Notes upon the increase and decrease of fish. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/notes-upon-increase-and-decrease-fish;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb6.57.pdf.

MP:1886:CF

[745] Moniteur de la Pisciculture. Carp in France. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/carp-france;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb6.65.pdf.

Sawyer:1886:BHE

[746] J. N. Sawyer. The breeding habits of the eel. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/breeding-habits-eel;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb6.73.pdf.

Pavesi:1886:NHP

[747] Pietro Pavesi. Notes on hatching and planting young fish in Italian waters. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/notes-hatching-and-planting-young-fish-italian-
waters;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb6.81.pdf.

Blackford:1886:CSH

[748] E. G. Blackford. Catch of shad in the Hudson River for 1885. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/catch-
shad-hudson-river-1885;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb6.89.pdf.

Smiley:1886:NUFa

[749] Chas W. Smiley. Notes upon fish and the fisheries. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/notes-upon-fish-and-fisheries-3;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb6.
97.pdf.

Ward:1886:YSN

[750] F. M. Ward. Young salmon in northern New Jersey. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/young-salmon-northern-new-jersey;   https:

//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb6.105.pdf.

Cutter:1886:SSF

[751] John C. Cutter. Some statistics of the fisheries of Northern Japan. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/some-statistics-fisheries-northern-japan;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb6.113.pdf.

Metcalf:1886:PWO

[752] Martin Metcalf. Planting whitefish and other salmonoids in inland lakes. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/planting-whitefish-and-other-salmonoids-inland-
lakes;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb6.121.pdf.

Brooks:1886:APC

[753] W. K. Brooks. On the artificial propagation and cultivation of oysters in floats. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/artificial-propagation-and-cultivation-oysters-
floats;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb6.129.pdf.

Ryder:1886:PND

[754] John A. Ryder. Preliminary notice of the development of the toadfish (Batrachus tau). Fishery Bulletin, 6(1):??, 1886. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/preliminary-notice-development-
toadfish-batrachus-tau;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb6.2.pdf.

Wilcox:1886:NEFb

[755] W. A. Wilcox. New England fisheries in October, 1885. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/new-
england-fisheries-october-1885;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb6.10.pdf.

NF:1886:MPF

[756] Norsk Fiskeritidende. Methods of preparing fish-bladders for market. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/methods-preparing-fish-bladders-market;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb6.
18.pdf.

NF:1886:NF

[757] Norsk Fiskeritidende. The Norwegian fisheries in 1885. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/norwegian-
fisheries-1885;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb6.26.pdf.

NF:1886:PN

[758] Norsk Fiskeritidende. The preservation of nets. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/preservation-nets;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb6.34.pdf.

RUSHR:1886:FNW

[759] Report to the United States House of Representatives. Fishing in the navigable waters of the United States. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishing-navigable-waters-united-states;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb6.42.pdf.

Tybring:1886:PF

[760] Oscar Tybring. Poisonous fish. Fishery Bulletin, 6(1):??, 1886. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/poisonous-fish;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb6.50.pdf.

Wilcox:1886:NEFc

[761] W. A. Wilcox. New England fisheries in may, 1886. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/new-england-fisheries-may-1886;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb6.58.pdf.

Dolge:1886:PCC

[762] Alfred Dolge. The preparation and cooking of carp. Fishery Bulletin, 6 (1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/preparation-
and-cooking-carp;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb6.66.pdf.

SPM:1886:SSF

[763] Statistique des Pêches Maritimes. Statistics of the sea fisheries of France for the year 1884. Fishery Bulletin, 6(1):??, 1886. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/statistics-sea-fisheries-france-
year-1884;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb6.74.pdf.

Heincke:1886:SFN

[764] Friedrich Heincke. The sea fisheries of Norway. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/sea-fisheries-norway;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb6.82.pdf.

Smith:1886:RSWa

[765] James A. Smith. Report on the shad work of the steamer lookout during the season of 1886. Fishery Bulletin, 6(1):??, 1886. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/report-shad-work-steamer-lookout-
during-season-1886;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb6.90.pdf.

Weber:1886:PPF

[766] M. Weber. Pearls and pearl fisheries. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/pearls-and-
pearl-fisheries;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb6.98.pdf.

Townsend:1886:PCC

[767] Charles H. Townsend. Present conditions of the California gray whale fishery. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/present-conditions-california-gray-whale-fishery;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb6.106.pdf.

McDonald:1886:RDF

[768] Marshall McDonald. Report on distribution of fish and eggs by the U. S. Fish Commission for the season of 1885–1886. Fishery Bulletin, 6 (1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/report-
distribution-fish-and-eggs-u-s-fish-commission-season-1885-
1886;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb6.114.pdf.

Wilcox:1886:NNEa

[769] W. A. Wilcox. Notes on the New England fisheries in December, 1886. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/notes-new-england-fisheries-december-1886;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb6.122.pdf.

Zacharias:1886:SLN

[770] Otto Zacharias. Some of the life-needs of fish. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/some-life-needs-fish;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb6.130.pdf.

Ryder:1886:ESC

[771] John A. Ryder. On the earlier stages of cleavage of the blastodicks of Raia crinacea. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/earlier-stages-cleavage-blastodicks-raia-crinacea;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb6.3.pdf.

Wilcox:1886:NEFd

[772] W. A. Wilcox. New England fisheries in November, 1885. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/new-england-fisheries-november-1885;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb6.
11.pdf.

Stange:1886:FBB

[773] J. Stange. Fish becoming blue and rigid during transportation. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fish-becoming-blue-and-rigid-during-transportation;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb6.19.pdf.

Gundersen:1886:DCB

[774] H. Gundersen. Drying codfish at Bordeaux. Fishery Bulletin, 6(1):84–86, 1886. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
drying-codfish-bordeaux;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb6.27.pdf.

Trybom:1886:SFH

[775] Filip Trybom. Salmon fisheries of Halland, Sweden, especially in the River Viskan. Fishery Bulletin, 6(1):??, 1886. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/salmon-fisheries-halland-sweden-
especially-river-viskan;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb6.35.pdf.

Benecke:1886:OSG

[776] Berthold Benecke. Observations on salmon in German rivers. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
observations-salmon-german-rivers;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb6.43.pdf.

Simonsen:1886:HFN

[777] Carl Simonsen. The herring fisheries near the isle of man. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/herring-fisheries-near-isle-man;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb6.
51.pdf.

Chadwick:1886:YMD

[778] B. P. Chadwick. Young mackerel destroyed by small-meshed nets. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/young-

mackerel-destroyed-small-meshed-nets;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb6.59.pdf.

Goll:1886:ABT

[779] Hermann Goll. The American brook trout recommended for Swiss waters. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/american-brook-trout-recommended-
swiss-waters;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb6.67.pdf.

Pavesi:1886:OME

[780] Pietro Pavesi. Observations on male eels. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/observations-
male-eels;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb6.75.pdf.

Wilcox:1886:NEFe

[781] W. A. Wilcox. New England fisheries in September, 1886. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/new-england-fisheries-september-1886;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb6.83.pdf.

Chalker:1886:CSCa

[782] Robert B. Chalker. Catch of shad in Connecticut for 1886. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/catch-
shad-connecticut-1886;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb6.91.pdf.

Wilcox:1886:NNEb

[783] W. A. Wilcox. Notes on the New England fisheries in October, 1886. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/notes-new-england-fisheries-october-1886;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb6.99.pdf.

Cheney:1886:SHR

[784] A. N. Cheney. Salmon in the Hudson River. Fishery Bulletin, 6(1): ??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/salmon-hudson-
river;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb6.107.pdf.

Clark:1886:RDF

[785] Frank N. Clark. Report on distribution of fish and eggs from northville and alpena stations for the season of 1885–1886. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/report-distribution-
fish-and-eggs-northville-and-alpena-stations-season-1885-1886;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb6.115.pdf.

UC:1886:QRM

[786] U.S.Fish Commission. Questions relative to mackerel and the mackerel fisheries. Fishery Bulletin, 6(1):??, 1886. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/questions-relative-mackerel-and-
mackerel-fisheries;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb6.123.pdf.

McDonald:1886:CTO

[787] Marshall McDonald. California trout for the Ozark Mountain region. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/california-trout-ozark-mountain-region;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb6.
131.pdf.

Glennan:1886:FKP

[788] A. H. Glennan. Fish killed by poisonous water. Fishery Bulletin, 6(1): ??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/fish-killed-
poisonous-water;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb6.4.pdf.

Wilcox:1886:NEFf

[789] W. A. Wilcox. New England fisheries in December, 1885. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/new-england-fisheries-december-1885;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb6.
12.pdf.

Ewart:1886:PFF

[790] J. Cossar Ewart. The preservation of fresh fish. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
preservation-fresh-fish;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb6.20.pdf.

Smith:1886:HRO

[791] E. L. Smith. Hood River, Oregon, as a place for salmon breeding. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/hood-river-oregon-place-salmon-breeding;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb6.28.pdf.

Wilcox:1886:NEFg

[792] W. A. Wilcox. New England fisheries in April, 1886. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/new-england-fisheries-april-1886;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb6.36.pdf.

Willcox:1886:FKC

[793] Joseph Willcox. Fish killed by cold along the Gulf of Mexico and coast of Florida. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fish-killed-cold-along-gulf-mexico-and-coast-florida;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb6.44.pdf.

DF:1886:CFP

[794] Danish Fiskeritidende. Crab fisheries proposed in Denmark. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/crab-fisheries-proposed-denmark;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb6.
52.pdf.

Knowles:1886:DFC

[795] Herbert M. Knowles. Dead fish on the coast of Rhode Island. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/dead-
fish-coast-rhode-island;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb6.60.pdf.

Nye:1886:HWF

[796] Willard Nye, Jr. Habits of whiting or frost-fish (Merlucius bilinearis Mitch.). Fishery Bulletin, 6(1):208, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/habits-whiting-or-frost-fish-merlucius-bilinearis-
mitch;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb6.68.pdf.

Bouchon-Brandely:1886:RAF

[797] G. Bouchon-Brandely. Report on the artificial fecundation and generation of oysters. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/report-artificial-fecundation-and-generation-
oysters;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb6.76.pdf.

vonBehr:1886:FRC

[798] E. von Behr. Five receipts for cooking carp. Fishery Bulletin, 6(1): ??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/five-receipts-
cooking-carp;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb6.84.pdf.

Collins:1886:NRS

[799] J. W. Collins. Notes on the red-snapper fishery. Fishery Bulletin, 6 (1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/notes-red-
snapper-fishery;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb6.92.pdf.

Abernethy:1886:SCR

[800] A. S. Abernethy. Salmon in the Clackamas River. Fishery Bulletin, 6 (1):332, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/salmon-
clackamas-river-0;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb6.100.pdf.

Dimmock:1886:BSO

[801] George Dimmock. Belostomidæ and some other fish-destroying bugs. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/belostomid%C3%A6-and-some-other-fish-destroying-bugs;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb6.108.pdf.

FS:1886:SSP

[802] Field and Stone. Salmon statistics on the Pacific Coast for the seasons of 1884–1886. Fishery Bulletin, 6(1):??, 1886. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/salmon-statistics-pacific-coast-
seasons-1884-1886;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb6.116.pdf.

McDonald:1886:RSD

[803] Marshall McDonald. Report on shad distribution for 1886. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
report-shad-distribution-1886;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb6.124.pdf.

Smiley:1886:NUFb

[804] Chas W. Smiley. Notes upon fish and the fisheries. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/notes-upon-fish-and-fisheries-5;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb6.
132.pdf.

Webb:1886:MFG

[805] John G. Webb. The mortality of fish in the Gulf of Mexico. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/mortality-fish-gulf-mexico;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb6.5.pdf.

Smiley:1886:IFO

[806] Chas W. Smiley. Inspection of fish and other marine products in the District of Columbia. Fishery Bulletin, 6(1):??, 1886. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/inspection-fish-and-other-marine-
products-district-columbia;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb6.13.pdf.

Tibbetts:1886:SCH

[807] N. V. Tibbetts. Scarcity of cod and haddock on the coast of Maine. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/scarcity-cod-and-haddock-coast-maine;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb6.21.pdf.

SFC:1886:SWC

[808] San Francisco Chronicle. Statistics of whaling, cod-fishing, and salmonpacking on the Pacific Coast. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/statistics-whaling-cod-fishing-
and-salmon-packing-pacific-coast;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb6.29.pdf.

Malmros:1886:PFS

[809] Oscar Malmros. Preserving fish in Scotland by the Roosen process. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/preserving-fish-scotland-roosen-process;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb6.37.pdf.

Burns:1886:RSC

[810] Frank Burns. Rockfish in South Carolina. Fishery Bulletin, 6(1): ??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/rockfish-
south-carolina;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb6.45.pdf.

Trybom:1886:CCS

[811] Filip Trybom. Carp culture in Sweden. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/carp-culture-sweden;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb6.53.pdf.

Hammond:1886:FPS

[812] J. P. Hammond. Fish in Puget Sound. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/fish-puget-
sound;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb6.61.pdf.

Trybom:1886:TRC

[813] Filip Trybom. The tench recommended for cultivation in Sweden. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/tench-recommended-cultivation-sweden;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb6.69.pdf.

Wilcox:1886:NEFh

[814] W. A. Wilcox. New England fisheries in August, 1886. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/new-
england-fisheries-august-1886;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb6.77.pdf.

Gross:1886:ARC

[815] Julius Gross. Another receipt for cooking carp. Fishery Bulletin, 6 (1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/another-
receipt-cooking-carp;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb6.85.pdf.

Jann:1886:CSD

[816] P. F. Jann. Catch of shad in the Delaware River for 1885–1886. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/catch-
shad-delaware-river-1885-1886;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb6.93.pdf.

Swan:1886:TF

[817] James G. Swan. The trepang fishery. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/trepang-fishery;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb6.101.pdf.

Page:1886:SWE

[818] W. F. Page. Shrinkage of whitefish eggs during transportation. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/shrinkage-whitefish-eggs-during-transportation;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb6.109.pdf.

Smiley:1886:NUFc

[819] Chas W. Smiley. Notes upon fish and the fisheries. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/notes-upon-fish-and-fisheries-4;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb6.
117.pdf.

Griffin:1886:PFA

[820] G. W. Griffin. The pearl fisheries of Australia. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/pearl-
fisheries-australia;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb6.125.pdf.

Stearns:1886:SFW

[821] Silas Stearns. Some of the fisheries of western Florida. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/some-fisheries-western-florida;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb6.133.
pdf.

Dannevig:1886:HLC

[822] G. M. Dannevig. Hatching lobsters and cod in Norway. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/hatching-
lobsters-and-cod-norway;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb6.6.pdf.

Wilcox:1886:NEFi

[823] W. A. Wilcox. New England fisheries in January, 1886. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/new-

england-fisheries-january-1886;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb6.14.pdf.

Stearns:1886:NFP

[824] Silas Stearns. Notes on the fisheries of Pensacola, Florida. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
notes-fisheries-pensacola-florida;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb6.22.pdf.

NF:1886:ETP

[825] Norsk Fiskeritidende. Eels in tanks and ponds. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/eels-tanks-and-ponds;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb6.30.pdf.

Pittendrigh:1886:SFP

[826] George Pittendrigh. Statistics of the fisheries of the Province of British Columbia for 1885. Fishery Bulletin, 6(1):??, 1886. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/statistics-fisheries-
province-british-columbia-1885;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb6.38.pdf.

Nye:1886:FFA

[827] Willard Nye, Jr. Fish and fishing at Abaco Island. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fish-and-fishing-abaco-island;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb6.46.pdf.

DF:1886:IFW

[828] Danish Fiskeritidende. The Iceland fresh-water fisheries. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
iceland-fresh-water-fisheries;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb6.54.pdf.

JFN:1886:TSA

[829] Journal of Fisheries in the Netherlands. Two species of American fish in the Aquarium of the Royal Zoological Society at Amsterdam. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/two-
species-american-fish-aquarium-royal-zoological-society-amsterdam;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb6.62.pdf.

Wilcox:1886:NEFj

[830] W. A. Wilcox. New England fisheries in July, 1886. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/new-england-fisheries-july-1886;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb6.
70.pdf.

Beckley:1886:HFI

[831] Emma Metcalf Beckley. Hawaiian fishing implements and methods of fishing. Fishery Bulletin, 6(1):??, 1886. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/hawaiian-fishing-implements-and-
methods-fishing;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb6.78.pdf.

Tanner:1886:RHS

[832] Z. L. Tanner. Record of hydrographic soundings and dredging stations occupied by the steamer Albatross in 1886. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/record-hydrographic-
soundings-and-dredging-stations-occupied-steamer-albatross-
1886;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb6.86.pdf.

McDonald:1886:DSF

[833] Marshall McDonald. Distribution of shad fry during 1886. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
distribution-shad-fry-during-1886;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb6.94.pdf.

Green:1886:SMR

[834] Loren W. Green. Salmon in the McCloud River during the season of 1886. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/salmon-mccloud-river-during-season-1886;   https://

spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb6.102.pdf.

Grabill:1886:ROS

[835] L. R. Grabill. Report of operations at the shad-hatching station on Battery Island, near Havre de Grace, Md., during the season of 1886. Fishery Bulletin, 6(1):??, 1886. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/report-operations-shad-hatching-
station-battery-island-near-havre-de-grace-md-during-season;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb6.110.pdf.

Wilcox:1886:MKS

[836] W. A. Wilcox. A man killed by a swordfish. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/man-killed-swordfish;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb6.118.pdf.

McDonald:1886:RSP

[837] Marshall McDonald. Report on shad propagation on the Potomac River during the season of 1886. Fishery Bulletin, 6(1): ??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/report-
shad-propagation-potomac-river-during-season-1886;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb6.126.pdf.

Clark:1886:NHP

[838] A. Howard Clark. Notes on the history of preparing fish for market by freezing. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/notes-history-preparing-fish-market-freezing;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb6.134.pdf.

Donnelly:1886:EFF

[839] W. J. S. Donnelly. Exports of fish and fishery products from Newfoundland during 1884. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/exports-fish-and-fishery-

products-newfoundland-during-1884;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb6.7.pdf.

Anonymous:1886:FCR

[840] Anonymous. The fisheries of Canada in 1884: Report of the Canadian Department of Fisheries. Fishery Bulletin, 6(1):51–56, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/report-
canadian-department-fisheries;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb6.15.pdf.

Wilcox:1886:NEFk

[841] W. A. Wilcox. New England fisheries in February, 1886. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/new-england-fisheries-february-1886;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb6.
23.pdf.

Ryder:1886:IGR

[842] John A. Ryder. On the intraovarian gestation of the redfish (Sebastes marinus). Fishery Bulletin, 6(1):??, 1886. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/intraovarian-gestation-redfish-
sebastes-marinus;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb6.31.pdf.

Barin:1886:SCR

[843] L. T. Barin. Salmon in the Clackamas River. Fishery Bulletin, 6(1): 111–112, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/salmon-
clackamas-river;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb6.39.pdf.

NF:1886:SRC

[844] Norsk Fiskeritidende. Some recipes for cooking fish. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/some-recipes-cooking-fish;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb6.47.pdf.

Ninni:1886:FOC

[845] Alexander P. Ninni. Fish and oyster culture in the Province of Venice. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fish-and-oyster-culture-province-venice;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb6.55.pdf.

Wilcox:1886:NEFl

[846] W. A. Wilcox. New England fisheries in June, 1886. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/new-england-fisheries-june-1886;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb6.
63.pdf.

Seed:1886:FCN

[847] William Seed. Fish-culture in New Zealand. Fishery Bulletin, 6(1): ??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/fish-culture-
new-zealand-0;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb6.71.pdf.

deBellesme:1886:ACT

[848] Jousset de Bellesme. American catfish in the Trocadéro Aquarium of Paris. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/american-catfish-trocad%C3%A9ro-aquarium-paris;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb6.79.pdf.

Green:1886:RPS

[849] Loren W. Green. Report on the packing of salmon on the Pacific Coast from 1883 to 1886. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/report-packing-salmon-pacific-coast-1883-
1886;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb6.87.pdf.

Gross:1886:DCP

[850] Julius Gross. Description of carp ponds and water gate. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/description-carp-ponds-and-water-gate;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb6.95.pdf.

Benecke:1886:EPC

[851] Berthold Benecke. The enemies of pond culture in central Europe. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/enemies-pond-culture-central-europe;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb6.
103.pdf.

Collins:1886:NIG

[852] J. W. Collins. Notes on an investigation of the great fishing-banks of the western Atlantic. Fishery Bulletin, 6(1):??, 1886. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/notes-investigation-
great-fishing-banks-western-atlantic;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb6.111.pdf.

Wilcox:1886:NNEc

[853] W. A. Wilcox. Notes on the New England fisheries in November, 1886. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/notes-new-england-fisheries-november-1886;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb6.119.pdf.

Chalker:1886:CSCb

[854] Robert B. Chalker. Catch of shad in Connecticut for 1885. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/catch-
shad-connecticut-1885;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb6.127.pdf.

Scudder:1886:LBF

[855] Charles W. Scudder. A list of blank forms, circulars, and minor publications of the United States Fish Commission from August 1, 1884, to March 1, 1887. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/list-blank-forms-circulars-and-minor-
publications-united-states-fish-commission-august-1;   https:

//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb6.135.pdf.

Anonymous:1886:TSA

[856] Anonymous. Topical synopsis of the articles and notes. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/topical-
synopsis-articles-and-notes-0;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb6%20contents.pdf.

Rathbun:1886:NLC

[857] Richard Rathbun. Notes on lobster culture. Fishery Bulletin, 6 (1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/notes-
lobster-culture;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb6.8.pdf.

Day:1886:ESS

[858] Francis Day. Experiments with salmon in Scotland. Fishery Bulletin, 6 (1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/experiments-
salmon-scotland;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb6.16.pdf.

Endicott:1886:GFG

[859] Charles Endicott. A good fishing-ground in the South Indian Ocean. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/good-fishing-ground-south-indian-ocean;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb6.
24.pdf.

Wilcox:1886:NEFm

[860] W. A. Wilcox. New England fisheries in March, 1886. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/new-england-fisheries-march-1886;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb6.32.pdf.

Keller:1886:FMF

[861] H. Keller. Freedom of migration for fish in Germany. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/freedom-migration-fish-germany;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb6.40.pdf.

Smiley:1886:NUFd

[862] Chas W. Smiley. Notes upon fish and the fisheries. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/notes-upon-fish-and-fisheries-2;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb6.
48.pdf.

Nye:1886:RL

[863] Willard Nye, Jr. A reasoning lobster. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/reasoning-lobster;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb6.56.pdf.

Harris:1886:RUS

[864] Gwynn Harris. Report upon the shad and herring fisheries of the Potomac River for 1886. Fishery Bulletin, 6(1):??, 1886. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/report-upon-shad-and-
herring-fisheries-potomac-river-1886;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb6.64.pdf.

UHR:1886:ARI

[865] U.S.Senate and House of Representatives. An act relating to the importing and landing of mackerel caught during the spawning season. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/act-relating-importing-and-landing-mackerel-caught-
during-spawning-season;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb6.72.pdf.

RBCF:1886:FPR

[866] Report of the Belgian Commission on Fish-Culture. Fry planted in rivers of Belgium in 1886. Fishery Bulletin, 6(1):??, 1886. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/fry-planted-rivers-belgium-1886;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb6.80.pdf.

Smith:1886:RSWb

[867] James A. Smith. Report on the shad work of the steamer Fish Hawk during the season of 1886. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/rep-ort-shad-work-steamer-fish-
hawk-during-season-1886;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb6.88.pdf.

RWWAS:1886:AFN

[868] Report of the Wellington and Wairarapa Acclimatization Society. American fish in New Zealand. Fishery Bulletin, 6(1):304, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/report-wellington-and-wairarapa-
acclimatization-society;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb6.96.pdf.

Ito:1886:FII

[869] K. Ito. Fishery industries of the island of Hokkaido, Japan. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fishery-industries-island-hokkaido-japan;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb6.104.pdf.

Collins:1886:CKF

[870] J. W. Collins. A curious knife found in the flesh of a codfish. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
curious-knife-found-flesh-codfish;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb6.112.pdf.

Clark:1886:DPS

[871] Frank N. Clark. Disadvantage of planting small fish. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/disadvantage-planting-small-fish;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb6.120.pdf.

Bean:1886:REC

[872] Tarleton H. Bean. Report on examination of clupeoids from carp ponds. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/report-examination-clupeoids-carp-ponds;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb6.128.pdf.

Anonymous:1886:I

[873] Anonymous. Index. Fishery Bulletin, 6(1):??, 1886. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/index-44;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb6.index_.pdf.

Bissell:1887:CAF

[874] John H. Bissell. The Canadian and American fisheries of the Great Lakes. Fishery Bulletin, 7(1):??, 1887. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/canadian-and-american-fisheries-
great-lakes;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb7.2.pdf.

Congress:1887:ASS

[875] U. S. Congress. An act for securing statistics of the extent and value of the vessel fisheries of the United States. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/act-securing-
statistics-extent-and-value-vessel-fisheries-united-states;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb7.10.pdf.

McIntosh:1887:EF

[876] ?. McIntosh. The eggs of fishes. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/eggs-fishes;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb7.18.pdf.

Wilcox:1887:FGMa

[877] W. A. Wilcox. The fisheries of Gloucester, Mass., in April, 1887, with notes on those of other localities. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/fisheries-gloucester-mass-april-
1887-notes-those-other-localities;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb7.26.pdf.

Cocks:1887:FFL

[878] Alfred Heneage Cocks. The finwhale fishery on the Lapland coast in 1886. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/finwhale-fishery-lapland-coast-1886;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb7.
35.pdf.

Tanner:1887:RMO

[879] Z. L. Tanner. Report of the movements and operations of the Fish Commission steamer Albatross from September 15 to 20, 1887. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
report-movements-and-operations-fish-commission-steamer-albatross-
september-15-20-1887;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb7.43.pdf.

Raveret-Wattel:1887:PEG

[880] C. Raveret-Wattel. The piscicultural establishment at Gremaz (Ain), France. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/piscicultural-establishment-gremaz-ain-france;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb7.51.pdf.

Seal:1887:IPM

[881] William P. Seal. An inquiry as to the proper methods for the care of living things held for popular amusement and instruction, or for purposes of biological research. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/inquiry-proper-methods-care-living-things-
held-popular-amusement-and-instruction-or;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb7.59.pdf.

Southwell:1887:NSW

[882] Thomas Southwell. Notes on the seal and whale fishery of 1880. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
notes-seal-and-whale-fishery-1880;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb7.3.pdf.

Smiley:1887:LLT

[883] Charles W. Smiley. Loch leven trout introduced in the United States. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/loch-leven-trout-introduced-united-states;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb7.11.pdf.

Hamilton:1887:CP

[884] W. R. Hamilton. Croaking of the perch. Fishery Bulletin, 7(1): ??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/croaking-
perch;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb7.19.pdf.

Wilcox:1887:FGMb

[885] W. A. Wilcox. The fisheries of Gloucester, Mass., in may, 1887, with notes on those of other localities. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/fisheries-gloucester-
mass-may-1887-notes-those-other-localities;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb7.27.pdf.

Hamlen:1887:RSW

[886] William Hamlen. Report on the shad work of the steamer halcyon during the season of 1887. Fishery Bulletin, 7(1):??, 1887. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/report-shad-work-steamer-halcyon-
during-season-1887;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb7.36.pdf.

Mowat:1887:SFPa

[887] Thomas Mowat. Statistics of the fisheries of the Province of British Columbia for the year 1887. Fishery Bulletin, 7(1):??, 1887. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/statistics-fisheries-
province-british-columbia-year-1887;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb7.44.pdf.

Muntadas:1887:RPE

[888] F. Muntadas. Report on the piscicultural establishment of Piedra, Aragon, Spain. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/report-piscicultural-establishment-piedra-
aragon-spain;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb7.52.pdf.

Collins:1887:BTF

[889] J. W. Collins. The beam-trawl fishery of Great Britain, with notes on beam-trawling in other European countries, etc. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
beam-trawl-fishery-great-britain-notes-beam-trawling-other-
european-countries-etc;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb7.60.pdf.

Smith:1887:RPS

[890] Everett Smith. Results of planting shad in the Kennebec River. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/results-planting-shad-kennebec-river;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb7.4.pdf.

Smiley:1887:NUF

[891] Chas W. Smiley. Notes upon fish and the fisheries. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/notes-upon-fish-and-fisheries-6;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb7.
12.pdf.

Clark:1887:WDN

[892] Frank N. Clark. Whitefish distributed from northville and alpena stations. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/whitefish-distributed-northville-and-
alpena-stations;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb7.20.pdf.

Wilcox:1887:FGMc

[893] W. A. Wilcox. The fisheries of Gloucester, Mass., in June, 1887, with notes on those of other localities. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/fisheries-gloucester-mass-june-

1887-notes-those-other-localities;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb7.28.pdf.

Chalker:1887:CSC

[894] Robert B. Chalker. Catch of shad in Connecticut for 1887. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/catch-
shad-connecticut-1887;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb7.37.pdf.

Nye:1887:NFB

[895] Willard Nye, Jr. Notes on the fisheries of Buzzard’s Bay and vicinity. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/notes-fisheries-buzzards-bay-and-vicinity;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb7.45.pdf.

Duclos:1887:ASP

[896] Gilbert Duclos. American silver perch or calico bass. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/american-silver-perch-or-calico-bass;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb7.53.pdf.

Mather:1887:RUR

[897] Fred Mather. Report upon the results of salmon planting in the Hudson River. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/report-upon-results-salmon-planting-hudson-river;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb7.61.pdf.

Chambers:1887:AFC

[898] W. Oldham Chambers. American fish cultivated by the Nation Fish-Culture Association of England. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/american-fish-cultivated-nation-
fish-culture-association-england;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb7.5.pdf.

Dunn:1887:FCP

[899] Horace D. Dunn. Fish-culture on the Pacific Coast. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fish-culture-pacific-coast;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb7.13.pdf.

Stone:1887:BSP

[900] Livingston Stone. Best season for packing salmon on the Pacific Coast. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/best-season-packing-salmon-pacific-coast;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb7.21.pdf.

Hinckelmann:1887:IFB

[901] ?. Hinckelmann. Injuries to the fisheries in the Baltic by seals. Fishery Bulletin, 7(1):81–83, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/injuries-fisheries-baltic-seals;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb7.
29.pdf.

Rognerud:1887:HCN

[902] Carl Rognerud. Hatching cod in Norway. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/hatching-cod-norway;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb7.38.pdf.

Earll:1887:ASI

[903] R. Edward Earll and Hugh M. Smith. The American sardine industry in 1880. Fishery Bulletin, 7(1):??, 1887. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/american-sardine-industry-1880;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb7.46.pdf.

Bean:1887:HBL

[904] Tarleton H. Bean. A hybrid between the lake trout and brook trout. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/hybrid-between-lake-trout-and-brook-trout;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb7.54.pdf.

Collins:1887:SRW

[905] J. W. Collins. Some reasons why the fishermen of Nova Scotia prefer to use salt clams (Mya arenaria) for bait in the bank handline cod fisheries. Fishery Bulletin, 7(1):??, 1887. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/some-reasons-why-fishermen-nova-
scotia-prefer-use-salt-clams-mya-arenaria-bait-bank-hand;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb7.62.pdf.

Brakeley:1887:RGC

[906] John H. Brakeley. Rapid growth of carp due to abundance of food. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/rapid-growth-carp-due-abundance-food;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb7.6.pdf.

SPM:1887:SFF

[907] Statistique des Pêches Maritimes. Sea fisheries of France and Algeria in 1885. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/sea-fisheries-france-and-algeria-1885;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb7.14.pdf.

Wilcox:1887:FGMd

[908] W. A. Wilcox. The fisheries of Gloucester, Mass., in January, 1887, with notes on those of other localities. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/fisheries-gloucester-
mass-january-1887-notes-those-other-localities;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb7.
22.pdf.

Smith:1887:RSW

[909] James A. Smith. Report on the shad work of the steamer Fish Hawk during the season of 1887. Fishery Bulletin, 7(1):??, 1887. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/report-shad-work-steamer-
fish-hawk-during-season-1887;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb7.30.pdf.

Thorarinsson:1887:NHF

[910] Bened. S. Thorarinsson. Notes on the hatcheries and fresh-water fisheries of Iceland. Fishery Bulletin, 7(1):??, 1887. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/notes-hatcheries-and-fresh-water-
fisheries-iceland;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb7.39.pdf.

Worth:1887:SBR

[911] S. G. Worth. The striped bass or rock fish industry of Roanoke Island, North Carolina, and vicinity. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/striped-bass-or-rock-
fish-industry-roanoke-island-north-carolina-and-vicinity;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb7.47.pdf.

Collins:1887:RIS

[912] D. E. Collins, T. H. Bean, and Richard Rathbun. Results of investigations by the schooner Grampus, on the southern mackerel grounds in the spring of 1887. Fishery Bulletin, 7(1):??, 1887. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/results-investigations-schooner-
grampus-southern-mackerel-grounds-spring-1887;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb7.
55.pdf.

Dresslar:1887:RMS

[913] Fletcher B. Dresslar and Bert Fesler. A review of the mackerels (Scombrinæ) of America and Europe. Fishery Bulletin, 7(1): ??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/review-
mackerels-scombrin%C3%A6-america-and-europe;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb7.
63.pdf.

Mather:1887:BTA

[914] Fred Mather. Brown trout in America. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/brown-trout-america;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb7.7.pdf.

Maynard:1887:RTS

[915] H. J. Maynard. Rainbow trout in the southwestern Missouri. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/rainbow-trout-southwestern-missouri;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb7.
15.pdf.

Wilcox:1887:FGMe

[916] W. A. Wilcox. The fisheries of Gloucester, Mass., in February, 1887, with notes on those of other localities. Fishery Bulletin, 7 (1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fisheries-
gloucester-mass-february-1887-notes-those-other-localities;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb7.23.pdf.

Layet:1887:EHO

[917] Alexandre Layet. Experimentel hygiene. Observations on the red flesh of the codfish. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/experimentel-hygiene-observations-red-flesh-
codfish;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb7.31.pdf.

Collins:1887:NUS

[918] J. W. Collins. Notes on the use of squid for food in the United States. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/notes-use-squid-food-united-states;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb7.40.pdf.

deCRavenel:1887:IBU

[919] W. de C.Ravenel. Information bearing upon the artificial propagation of mullet. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/information-bearing-upon-
artificial-propagation-mullet;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb7.48.pdf.

Earll:1887:NCF

[920] R. Edward Earll. Notes on certain fishery industries of Eastport, Me., in 1886. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/notes-certain-fishery-industries-eastport-
me-1886;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb7.56.pdf.

Collins:1887:SRC

[921] J. W. Collins. Statistics respecting certain features of the vessel fisheries of the United States. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/statistics-respecting-certain-features-
vessel-fisheries-united-states;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb7.64.pdf.

Anonymous:1887:P

[922] Anonymous. Plates. Fishery Bulletin, 7(1):??, 1887. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/plates-1;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb7%20plates.pdf.

Ruge:1887:SFF

[923] J. G. Ruge. The sponge fisheries of Florida. Fishery Bulletin, 7 (1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/sponge-
fisheries-florida;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb7.8.pdf.

Dunn:1887:SIC

[924] Horace D. Dunn. Salmon not injured by catfish. Fishery Bulletin, 7(1): ??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/salmon-not-
injured-catfish;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb7.16.pdf.

Wilcox:1887:FGMf

[925] W. A. Wilcox. The fisheries of Gloucester, Mass., in march, 1887, with notes on those of other localities. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/fisheries-gloucester-mass-march-
1887-notes-those-other-localities;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb7.24.pdf.

VonOterendorp:1887:VCF

[926] C. G. Von Oterendorp. A visit to the chief fishing places of Jutland. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/visit-chief-fishing-places-jutland;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb7.33.pdf.

Collins:1887:NOM

[927] J. W. Collins. Note on the occurrence of mackerel off the coast of Florida. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/note-occurrence-mackerel-coast-florida;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb7.
41.pdf.

Jordan:1887:NOO

[928] David S. Jordan. Note on the occurrence of the opha (Lampris guttatus) on the Grand Banks. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/note-occurrence-opha-
lampris-guttatus-grand-banks;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb7.49.pdf.

Seal:1887:ABE

[929] William P. Seal. The aquarium. A brief exposition of its principles and management. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/aquarium-brief-exposition-its-principles-and-
management;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb7.57.pdf.

Anonymous:1887:I

[930] Anonymous. Index. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/index-45;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb7.index_.pdf.

Moore:1887:RSA

[931] Thomas J. Moore. Report on a successful attempt to introduce living soles to America. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/report-successful-attempt-introduce-living-
soles-america;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb7.1.pdf.

Mowat:1887:SFPb

[932] Thomas Mowat. Statistics of the fisheries of the Province of British Columbia for 1886. Fishery Bulletin, 7(1):??, 1887. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/statistics-fisheries-
province-british-columbia-1886;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb7.9.pdf.

Redding:1887:CDS

[933] Joseph D. Redding. Causes of the decrease of salmon in the Sacramento River. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/causes-decrease-salmon-sacramento-river;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb7.17.pdf.

Wilcox:1887:NFH

[934] W. A. Wilcox. Notes on the frozen-herring trade with Newfoundland and the Bay of Fundy during the winter of 1886–1887. Fishery Bulletin, 7(1): ??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/notes-frozen-
herring-trade-newfoundland-and-bay-fundy-during-winter-1886-
1887;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb7.25.pdf.

McDonald:1887:RSD

[935] Marshall McDonald. Report of shad distribution for 1887. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
report-shad-distribution-1887;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb7.34.pdf.

Bean:1887:RFO

[936] Tarleton H. Bean. Report on the fishes observed in Great Egg Harbor Bay, New Jersey, during the summer of 1887. Fishery Bulletin, 7(1): ??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/report-fishes-
observed-great-egg-harbor-bay-new-jersey-during-summer-1887;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb7.42.pdf.

Mason:1887:SRF

[937] Frank Mason II. Self-reproducing food for young fish. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/self-reproducing-food-young-fish;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb7.50.pdf.

Seal:1887:KSP

[938] William P. Seal. The keeping of salmonidæ for purposes of exhibition. Fishery Bulletin, 7(1):??, 1887. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/keeping-salmonid%C3%A6-purposes-
exhibition;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb7.58.pdf.

Tanner:1888:EFG

[939] Z. L. Tanner et al. Explorations of the fishing grounds of Alaska, Washington Territory, and Oregon, during 1888, by the U. S. Fish Commission steamer Albatross. Fishery Bulletin, 8(1):1–95, 1888. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/explorations-fishing-
grounds-alaska-washington-territory-and-oregon-during-1888-
u-s-fish;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb8.1.pdf.

Jordan:1888:RGS

[940] David Starr Jordan and Carl H. Eigenmann. A review of the genera and species of serranidæ found in the waters of American and Europe. Fishery Bulletin, 8(1):??, 1888. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/review-genera-and-species-serranid%C3%A6-found-
waters-american-and-europe;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb8.9.pdf.

Jordan:1888:REM

[941] David Starr Jordan. Report of explorations made during 1888 in the Alleghany region of Virginia, North Carolina, and Tennessee, and in western Indiana, with an account of the fishes found in each of the river basins of those regions. Fishery Bulletin, 8(1):97, 173, 1888. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/report-explorations-made-during-1888-alleghany-
region-virginia-north-carolina-and-tennessee;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb8.
2.pdf.

Bean:1888:RPI

[942] Tarleton H. Bean. Report on the proposed introduction of the Jamaica mountain mullet into the United States. Fishery Bulletin, 8 (1):??, 1888. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/report-
proposed-introduction-jamaica-mountain-mullet-united-states;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb8.10.pdf.

Collins:1888:SEI

[943] J. W. Collins. Suggestions for the employment of improved types of vessels in the the market fisheries, with notes on British fishing steamers. Fishery Bulletin, 8(1):??, 1888. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/suggestions-employment-improved-
types-vessels-market-fisheries-notes-british-fishing;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb8.3.pdf.

Rathbun:1888:TLP

[944] Richard Rathbun. The transplanting of lobsters to the Pacific coast of the United States. Fishery Bulletin, 8(1):??, 1888. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/transplanting-lobsters-pacific-coast-
united-states;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb8.11.pdf.

Bean:1888:NFC

[945] Tarleton H. Bean. Notes on fishes collected at Cozumel, Yucatan, by the U. S. Fish Commission, with descriptions of new species. Fishery Bulletin, 8(1):??, 1888. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
notes-fishes-collected-cozumel-yucatan-u-s-fish-commission-
descriptions-new-species;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb8.4.pdf.

Forbes:1888:PRU

[946] S. A. Forbes. Preliminary report upon the invertebrate animals inhabiting Lakes Geneva and Mendota, Wisconsin, with an account of the fish epidemic in Lake Mendota in 1884. Fishery Bulletin, 8(1):??, 1888. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/preliminary-report-
upon-invertebrate-animals-inhabiting-lakes-geneva-and-mendota-
wisconsin;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb8.12.pdf.

Page:1888:MRM

[947] William F. Page. The most recent methods of hatching fish eggs. Fishery Bulletin, 8(1):??, 1888. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/most-
recent-methods-hatching-fish-eggs;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb8.5.pdf.

Anonymous:1888:I

[948] Anonymous. Index. Fishery Bulletin, 8(1):??, 1888. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/index-46;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb8.index_.pdf.

Bollman:1888:RUF

[949] Charles H. Bollman. A report upon the fishes of Kalamazoo, Calhoun, and Antrim Counties, Michigan. Fishery Bulletin, 8(1): ??, 1888. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/report-
upon-fishes-kalamazoo-calhoun-and-antrim-counties-michigan;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb8.6.pdf.

Gilbert:1888:NFL

[950] Chas H. Gilbert. Notes on the fishes from the lowlands of Georgia, with a description of a new species (Opsopæodus bollmani). Fishery Bulletin, 8(1):??, 1888. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/notes-fishes-lowlands-georgia-description-
new-species-opsop%C3%A6odus-bollmani;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb8.7.pdf.

Anonymous:1888:EI

[951] Anonymous. Explanations of illustrations. Fishery Bulletin, 8(1):??, 1888. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/explanations-
illustrations;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb8%20plates.pdf.

Ryder:1888:SSI

[952] John A. Ryder. The sturgeons and sturgeon industries of the eastern coast of the United States, with an account of experiments bearing upon sturgeon culture. Fishery Bulletin, 8(1):??, 1888. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/sturgeons-and-
sturgeon-industries-eastern-coast-united-states-account-experiments-
bearing;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb8.8.pdf.

Jordan:1889:RSL

[953] David Starr Jordan. A reconnoissance of the streams and lakes of the Yellowstone National Park, Wyoming, in the interest of the United States Fish Commission. Fishery Bulletin, 9(1):??, 1889. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/reconnoissance-streams-and-lakes-
yellowstone-nation-park-wyoming-interest-united-states;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb9.2.pdf.

Gilbert:1889:REM

[954] Charles H. Gilbert. Report of explorations made in Alabama during 1889, with notes on the fishes of the Tennessee, Alabama, and Escambia Rivers. Fishery Bulletin, 9(1):??, 1889. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/report-explorations-made-alabama-during-1889-notes-
fishes-tennessee-alabama-and-escambia;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb9.10.pdf.

Smith:1889:NKC

[955] Hugh M. Smith. Notes on the king crab fishery of Delaware Bay. Fishery Bulletin, 9(1):??, 1889. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/notes-king-crab-fishery-delaware-bay;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb9.19.pdf.

Linton:1889:TSL

[956] Edwin Linton. On two species of larval Dibothria from the Yellowstone National Park. Fishery Bulletin, 9(1):??, 1889. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/two-species-larval-
dibothria-yellowstone-nation-park;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb9.3.pdf.

Smith:1889:NIF

[957] Hugh M. Smith. Notes on an improved form of oyster tongs. Fishery Bulletin, 9(1):??, 1889. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/notes-improved-form-oyster-tongs;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb9.11.pdf.

Henshall:1889:RUC

[958] James A. Henshall. Report upon a collection of fishes made in southern Florida during 1889. Fishery Bulletin, 9(1):??, 1889. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/report-upon-collection-fishes-made-southern-
florida-during-1889;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb9.20.pdf.

Jacobson:1889:APS

[959] Herman (translated by) Jacobson. The artificial propagation of sturgeon in Schleswig-Holstein, Germany. Fishery Bulletin, 9(1):??, 1889. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/artificial-
propagation-sturgeon-schleswig-holstein-germany;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb9.4.pdf.

Wilson:1889:ESB

[960] Henry V. Wilson. The embryology of the sea bass (Serranus atrarius). Fishery Bulletin, 9(1):??, 1889. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/embryology-sea-bass-serranus-atrarius;   https:

//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb9.13.pdf.

Libbey:1889:RUP

[961] William Libbey, Jr. Report upon a physical investigation of the waters off the southern coast of New England, made during the summer of 1889, by the U. S. Fish Commission schooner Grampus. Fishery Bulletin, 9(1):??, 1889. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/report-upon-physical-investigation-waters-southern-
coast-new-england-made-during-summer;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb9.21.pdf.

Townsend:1889:RUP

[962] Charles H. Townsend. Report upon the pearl fishery of the Gulf of California. Fishery Bulletin, 9(1):??, 1889. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/report-upon-pearl-fishery-gulf-california;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb9.5.pdf.

Tanner:1889:FGB

[963] Z. L. Tanner. The fishing grounds of Bristol Bay, Alaska: a preliminary report upon the investigations of the U. S. Fish Commission steamer Albatross during the summer of 1890. Fishery Bulletin, 9(1):??, 1889. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishing-grounds-bristol-bay-alaska-preliminary-report-upon-
investigations-u-s-fish;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb9.14.pdf.

Collins:1889:NOF

[964] J. W. Collins. Notes on the oyster fishery of Connecticut. Fishery Bulletin, 9(1):??, 1889. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/notes-oyster-fishery-connecticut;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb9.22.pdf.

Gilbert:1889:RUC

[965] Charles H. Gilbert. Report upon certain investigations relating to the planting of oysters in southern California. Fishery Bulletin, 9(1):??, 1889. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/report-upon-certain-
investigations-relating-planting-oysters-southern-california;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb9.6.pdf.

Adams:1889:RUI

[966] A. C. Adams and W. C. Kendall. Report upon an investigation of the fishing grounds off the west coast of Florida. Fishery Bulletin, 9(1):??, 1889. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/report-upon-
investigation-fishing-grounds-west-coast-florida;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb9.15.pdf.

Anonymous:1889:I

[967] Anonymous. Index. Fishery Bulletin, 9(1):??, 1889. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/index-47;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb9.index_.pdf.

Linton:1889:CWE

[968] Edwin Linton. On certain warlike excrescences occurring on the short minnow, Cyprinodon variegates, due to Psorosperms. Fishery Bulletin, 9(1):??, 1889. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/certain-
warlike-excrescences-occurring-short-minnow-cyprinodon-variegates-
due-psorosperms;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb9.7.pdf.

Smith:1889:GSF

[969] Hugh M. Smith. The giant scallop fishery of Maine. Fishery Bulletin, 9 (1):??, 1889. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/giant-scallop-
fishery-maine;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb9.16.pdf.

Anonymous:1889:EI

[970] Anonymous. Explanation of illustrations. Fishery Bulletin, 9(1):??, 1889. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/explanation-
illustrations-2;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb9%20plates.pdf.

Smith:1889:NCF

[971] Hugh M. Smith. Notes on the crab fishery in Crisfield, Maryland. Fishery Bulletin, 9(1):??, 1889. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/notes-crab-fishery-crisfield-maryland;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb9.8.pdf.

Linton:1889:CLH

[972] Edwin Linton. A contribution to the life history of Dibothrium cordiceps Leidy, a parasite infesting the trout of Yellowstone Lake. Fishery Bulletin, 9(1):??, 1889. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/contribution-life-history-dibothrium-cordiceps-
leidy-parasite-infesting-trout-yellowstone;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb9.17.pdf.

Jordan:1889:REC

[973] David Starr Jordan. Report of explorations in Colorado and Utah during the summer of 1889, with an account of the fishes found in each of the river basins examined. Fishery Bulletin, 9(1):??, 1889. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/report-explorations-colorado-and-
utah-during-summer-1889-account-fishes-found-each-river;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb9.1.pdf.

Meek:1889:REM

[974] Seth Eugene Meek. Report of explorations made in Missouri and Arkansas during 1889, with an account of the fishes observed in each of the river basins examined. Fishery Bulletin, 9(1):??, 1889. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/report-explorations-made-
missouri-and-arkansas-during-1889-account-fishes-observed-each;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb9.9.pdf.

Linton:1889:NOP

[975] Edwin Linton. Notice of the occurrence of protozoan parasites (Psorosperms) on cyprinoid fishes in Ohio. Fishery Bulletin, 9(1): ??, 1889. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/notice-
occurrence-protozoan-parasites-psorosperms-cyprinoid-fishes-
ohio;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb9.18.pdf.

Anonymous:1890:EI

[976] Anonymous. Explanation of illustrations. Fishery Bulletin, 10(1):??, 1890. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/explanation-
illustrations-3;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb10%20plates.pdf.

Woolman:1890:RER

[977] Albert J. Woolman. Report of an examination of the rivers of Kentucky, with lists of the fishes obtained. Fishery Bulletin, 10(1): ??, 1890. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/report-
examination-rivers-kentucky-lists-fishes-obtained;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb10.8.pdf.

Anonymous:1890:I

[978] Anonymous. Index. Fishery Bulletin, 10(1):??, 1890. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-48;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb10.index_.pdf.

Seal:1890:OAU

[979] William P. Seal. Observations on the aquaria of the U. S. Fish Commission at Central Station, Washington, D.C. Fishery Bulletin, 10(1): ??, 1890. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/observations-
aquaria-u-s-fish-commission-central-station-washington-d-c;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb10.1.pdf.

Kirsch:1890:NSF

[980] Philip H. Kirsch. Notes on the streams and fishes of Clinton County, Kentucky, with a description of a new darter. Fishery Bulletin, 10 (1):??, 1890. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/notes-
streams-and-fishes-clinton-county-kentucky-description-new-

darter;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb10.9.pdf.

Collins:1890:FVB

[981] J. W. Collins. The fishing vessels and boats of the Pacific Coast. Fishery Bulletin, 10(1):??, 1890. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fishing-vessels-and-boats-pacific-coast;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb10.2.pdf.

Woolman:1890:RUR

[982] Albert J. Woolman. Report upon the rivers of central Florida tributary to the Gulf of Mexico, with lists of fishes inhabiting them. Fishery Bulletin, 10(1):??, 1890. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/report-
upon-rivers-central-florida-tributary-gulf-mexico-lists-fishes-
inhabiting-them;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb10.10.pdf.

Bean:1890:OUF

[983] Tarleton H. Bean. Observations upon fishes and fish-culture. Fishery Bulletin, 10(1):??, 1890. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/observations-upon-fishes-and-fish-culture;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb10.3.pdf.

Battle:1890:ICW

[984] John D. Battle. An investigation of the coast waters of South Carolina with reference to oyster-culture. Fishery Bulletin, 10(1):??, 1890. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/investigation-coast-
waters-south-carolina-reference-oyster-culture;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb10.
11.pdf.

Smith:1890:NFL

[985] Hugh M. Smith. Notes on fishes from the lower Potomac River, Maryland. Fishery Bulletin, 10(1):??, 1890. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/notes-fishes-lower-potomac-river-maryland;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb10.4.pdf.

Worth:1890:OHY

[986] S. G. Worth. Observations on the hatching of the yellow perch. Fishery Bulletin, 10(1):??, 1890. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
observations-hatching-yellow-perch;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb10.12.pdf.

Collins:1890:RFN

[987] J. W. Collins and Hugh M. Smith. Report on the fisheries of the New England states. Fishery Bulletin, 10(1):??, 1890. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/report-fisheries-new-
england-states;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb10.5.pdf.

Dean:1890:PBC

[988] Bashford Dean. The physical and biological characteristics of the natural oyster-grounds of South Carolina. Fishery Bulletin, 10(1):??, 1890. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/physical-and-
biological-characteristics-natural-oyster-grounds-south-carolina;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb10.13.pdf.

Smith:1890:RFL

[989] Hugh M. Smith. Report on the fisheries of Lake Ontario. Fishery Bulletin, 10(1):??, 1890. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
report-fisheries-lake-ontario;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb10.6.pdf.

Dean:1890:PMO

[990] Bashford Dean. The present methods of oyster-culture in France. Fishery Bulletin, 10(1):??, 1890. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/present-methods-oyster-culture-france;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb10.14.pdf.

Meek:1890:RUF

[991] Seth Eugene Meek. Report upon the fishes of Iowa, based upon observations and collections made during 1889, 1890, and 1891. Fishery Bulletin, 10(1):??, 1890. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/report-
upon-fishes-iowa-based-upon-observations-and-collections-made-
during-1889-1890-and;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb10.7.pdf.

Kellogg:1890:COK

[992] James L. Kellogg. A contribution to our knowledge of the morphology of lamellibranchiate mollusks. Fishery Bulletin, 10(1):??, 1890. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/contribution-
our-knowledge-morphology-lamellibranchiate-mollusks;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb10.15.pdf.

Evermann:1891:RUI

[993] Barton W. Evermann. A report upon investigations made in Texas in 1891. Fishery Bulletin, 11(1):??, 1891. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/report-upon-investigations-made-
texas-1891;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb11.2.pdf.

Gurley:1891:CMG

[994] R. R. Gurley. On the classification of the Myxosporidia, a group of protozoan parasites infesting fishes. Fishery Bulletin, 11(1):??, 1891. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/classification-myxosporidia-group-
protozoan-parasites-infesting-fishes;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb11.10.pdf.

Collins:1891:SRF

[995] J. W. Collins and Hugh M. Smith. A statistical report on the fisheries of the Gulf States. Fishery Bulletin, 11(1):??, 1891. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/statistical-report-
fisheries-gulf-states;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb11.3.pdf.

Anonymous:1891:I

[996] Anonymous. Index. Fishery Bulletin, 11(1):??, 1891. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-49;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb11.index_.pdf.

Smith:1891:RCF

[997] Hugh M. Smith. Report on a collection of fishes from the albemarle region of North Carolina. Fishery Bulletin, 11(1):??, 1891. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/report-collection-fishes-
albemarle-region-north-carolina;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb11.4.pdf.

Worth:1891:SHS

[998] S. G. Worth. The spawning habits of the shad. Fishery Bulletin, 11(1):??, 1891. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/spawning-habits-shad;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb11.5.pdf.

Forbes:1891:PRA

[999] S. A. Forbes. A preliminary report on the aquatic invertebrate fauna of the Yellowstone National Park, Wyoming, and the Flathead Region of Montana. Fishery Bulletin, 11(1):??, 1891. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/preliminary-report-
aquatic-invertebrate-fauna-yellowstone-national-park-wyoming-
and;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb11.6.pdf.

Kirsch:1891:NCF

[1000] Philip H. Kirsch. Notes on a collection of fishes from the southern tributaries of the Cumberland River in Kentucky and Tennessee. Fishery Bulletin, 11(1):??, 1891. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/notes-collection-fishes-southern-
tributaries-cumberland-river-kentucky-and-tennessee;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb11.7.pdf.

Anonymous:1891:EI

[1001] Anonymous. Explanation of illustrations. Fishery Bulletin, 11(1):??, 1891. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/explanation-
illustrations-4;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb11%20plates.pdf.

Smith:1891:RFS

[1002] Hugh M. Smith. Report on the fisheries of the south Atlantic states. Fishery Bulletin, 11(1):??, 1891. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
report-fisheries-south-atlantic-states;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb11.8.pdf.

Evermann:1891:RSL

[1003] Barton W. Evermann. A reconnaissance of the streams and lakes of western Montana and northwestern Wyoming. Fishery Bulletin, 11(1):??, 1891. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
reconnaissance-streams-and-lakes-western-montana-and-northwestern-
wyoming;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb11.1.pdf.

Dean:1891:REM

[1004] Bashford Dean. Report on the European methods of oyster-culture. Fishery Bulletin, 11(1):??, 1891. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
report-european-methods-oyster-culture;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb11.9.pdf.

Anonymous:1892:EI

[1005] Anonymous. Explanation of illustrations. Fishery Bulletin, 12(1):??, 1892. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/explanation-
illustrations-5;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb12%20plates.pdf.

Smith:1892:ENH

[1006] Hugh M. Smith. Economic and natural-history notes on fishes of the northern coast of New Jersey. Fishery Bulletin, 12(1): ??, 1892. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/economic-
and-natural-history-notes-fishes-northern-coast-new-jersey;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb12.8.pdf.

McDonald:1892:RSF

[1007] Marshall McDonald. Report on the salmon fisheries of Alaska. Fishery Bulletin, 12(1):??, 1892. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/report-salmon-fisheries-alaska;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb12.1.pdf.

Eigenmann:1892:VFP

[1008] Carl H. Eigenmann. On the viviparous fishes of the Pacific Coast of North America. Fishery Bulletin, 12(1):??, 1892. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/viviparous-fishes-pacific-coast-
north-america;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb12.9.pdf.

Evermann:1892:DNS

[1009] Barton W. Evermann. Description of a new sucker (Pantosteus jordani) from the upper Missouri Basin. Fishery Bulletin, 12(1):??, 1892. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/description-
new-sucker-pantosteus-jordani-upper-missouri-basin;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb12.2.pdf.

Anonymous:1892:I

[1010] Anonymous. Index. Fishery Bulletin, 12(1):??, 1892. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-50;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb12.index_.pdf.

Evermann:1892:FTR

[1011] Barton W. Evermann and William C. Kendall. The fishes of Texas and the Rio Grande Basin, considered chiefly with reference to their geographic distribution. Fishery Bulletin, 12(1):??, 1892. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/fishes-texas-and-rio-
grande-basin-considered-chiefly-reference-their-geographic;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb12.3.pdf.

Rathbun:1892:SFI

[1012] Richard Rathbun. Summary of fishery investigations conducted in the north Pacific Ocean and Bering Sea from July 1, 1888, to July 1, 1892, by the U.S. Fish Commission steamer Albatross. Fishery Bulletin, 12(1):??, 1892. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/summary-fishery-investigations-conducted-north-
pacific-ocean-and-bering-sea-july-1-1888;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb12.4.pdf.

Stevenson:1892:OIM

[1013] Charles H. Stevenson. The oyster industry of Maryland. Fishery Bulletin, 12(1):??, 1892. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/oyster-
industry-maryland;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb12.5.pdf.

Smith:1892:FNF

[1014] Hugh M. Smith. The fyke nets and fyke-net fisheries of the United States, with notes on the fyke nets of other countries. Fishery Bulletin, 12(1):??, 1892. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fyke-nets-
and-fyke-net-fisheries-united-states-notes-fyke-nets-other-
countries;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb12.6.pdf.

Moore:1892:LFC

[1015] H. F. Moore. List of fishes collected at sea isle city during the summer of 1892. Fishery Bulletin, 12(1):??, 1892. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/list-fishes-collected-sea-isle-
city-during-summer-1892;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb12.7.pdf.

Bean:1893:RSG

[1016] Tarleton H. Bean. Report as Secretary of the General Committee. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/report-secretary-general-committee;

https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb13.1.pdf.

Ryder:1893:BRR

[1017] John A. Ryder. Biological research in relation to the fisheries. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
biological-research-relation-fisheries;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb13.9.pdf.

Reighard:1893:SPS

[1018] J. E. Reighard. Some plankton studies in the Great Lakes. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
some-plankton-studies-great-lakes;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb13.17.pdf.

Post:1893:FCM

[1019] Hoyt Post. Fish-culture in Michigan. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/fish-culture-
michigan;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb13.25.pdf.

Rowe:1893:MDW

[1020] Henry C. Rowe. The methods of deep-water oyster culture. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
methods-deep-water-oyster-culture;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb13.33.pdf.

Keyes:1893:FIL

[1021] C. M. Keyes. The fishing industry of Lake Erie, past and present. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishing-industry-lake-erie-past-and-present;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb13.
41.pdf.

Anonymous:1893:I

URL https://spo.nmfs.noaa.gov/content/index-51;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb13.index_.pdf.

McDonald:1893:ACG

[1023] Marshall McDonald. Address as chairman of the general committee. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/address-chairman-general-committee;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb13.2.pdf.

Loeb:1893:ILP

[1024] Jacques Loeb. The influence of light on the periodical depth migrations of pelagic animals. Fishery Bulletin, 13(1):??, 1893. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/influence-light-periodical-depth-
migrations-pelagic-animals;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb13.10.pdf.

Forbes:1893:AUF

[1025] S. A. Forbes. The aquarium of the U. S. Fish Commission at the World’s Columbian Exposition. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/aquarium-u-s-
fish-commission-worlds-columbian-exposition;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb13.
18.pdf.

Clark:1893:HMW

[1026] F. N. Clark. History and method of whitefish-culture. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/history-and-method-whitefish-culture;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb13.26.pdf.

Cheney:1893:BNF

[1027] A. N. Cheney. Breeding natural food artificially for young fish artificially reared. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/breeding-natural-food-artificially-young-fish-artificially-

reared;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb13.34.pdf.

Wilkins:1893:FVC

[1028] J. T. Wilkins. The fisheries of the Virginia Coast. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fisheries-
virginia-coast;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb13.42.pdf.

MacCallum:1893:AFL

[1029] G. A. MacCallum. The assimilation of the fishery laws of the Great Lakes. Fishery Bulletin, 13(1):??, 1893. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/assimilation-fishery-laws-great-
lakes;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb13.3.pdf.

Evermann:1893:IRL

[1030] Barton W. Evermann. The investigations of rivers and lakes with reference to the fish environment. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/investigations-rivers-and-lakes-
reference-fish-environment;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb13.11.pdf.

Reeves:1893:DFS

[1031] I. S. K. Reeves. Description of the fresh and salt water supply and pumping plants uses for the aquarium. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
description-fresh-and-salt-water-supply-and-pumping-plants-
uses-aquarium;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb13.19.pdf.

Atkins:1893:MEC

[1032] Charles G. Atkins. Methods employed at Craig Brook in rearing young salmonoid fishes. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/methods-employed-craig-brook-rearing-
young-salmonoid-fishes;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb13.27.pdf.

Mather:1893:WWK

[1033] Fred Mather. What we know about the lobster. Fishery Bulletin, 13(1): ??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/what-we-know-
about-lobster;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb13.35.pdf.

Green:1893:NIM

[1034] W. S. Green. Notes on the Irish mackerel fisheries. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/notes-irish-mackerel-fisheries;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb13.43.
pdf.

Spangler:1893:DFA

[1035] A. M. Spangler. The decrease of fish in American waters, and some of the causes. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/decrease-fish-american-waters-and-some-causes;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb13.4.pdf.

Herrick:1893:HDA

[1036] F. H. Herrick. Habits and development of the American lobster, and their bearing upon its artificial propagation. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/habits-and-
development-american-lobster-and-their-bearing-upon-its-artificial-
propagation;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb13.12.pdf.

Anonymous:1893:OES

[1037] Anonymous. Observations and experiments on saprolegnia infesting fish. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/observations-and-experiments-saprolegnia-infesting-
fish;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb13.20.pdf.

Page:1893:PBB

[1038] William F. Page. Propagation of black bass in ponds. Fishery Bulletin, 13 (1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/propagation-
black-bass-ponds;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb13.28.pdf.

Smith:1893:MIA

[1039] Hugh M. Smith. The maintenance and improvement of the American fisheries. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/maintenance-and-improvement-american-fisheries;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb13.36.pdf.

Stanley-Brown:1893:PFS

[1040] J. Stanley-Brown. The past and future of the fur seal. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/past-and-
future-fur-seal;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb13.44.pdf.

Church:1893:SCF

[1041] Daniel T. Church. The sea and coast fisheries. Fishery Bulletin, 13(1): ??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/sea-and-coast-
fisheries;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb13.5.pdf.

Brooks:1893:OFM

[1042] W. K. Brooks. The origin of the food of marine animals. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/origin-food-marine-animals;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb13.13.
pdf.

Stiles:1893:RPP

[1043] C. W. Stiles. Report on a parasitic protozoan observed on fish in the aquarium. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/report-parasitic-protozoan-observed-fish-aquarium;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb13.21.pdf.

Quelch:1893:FBG

[1044] J. J. Quelch. Fishing in British Guiana. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/fishing-british-guiana;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb13.29.pdf.

Lawrence-Hamilton:1893:RIS

[1045] J. Lawrence-Hamilton. Reforms and improvements suggested for the fisheries of Great Britain and Ireland. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/reforms-and-
improvements-suggested-fisheries-great-britain-and-ireland;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb13.37.pdf.

Swan:1893:NFF

[1046] James G. Swan. Notes on the fisheries and fishery industries of Puget Sound. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/notes-fisheries-and-fishery-industries-puget-sound;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb13.45.pdf.

Southwick:1893:OOF

[1047] J. M. K. Southwick. Our ocean fisheries, and the effect of legislation upon the fisheries. Fishery Bulletin, 13(1):??, 1893. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/our-ocean-fisheries-and-effect-
legislation-upon-fisheries;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb13.6.pdf.

Armistead:1893:AOI

[1048] J. J. Armistead. Atmospheric and other influences on the migrations of fishes. Fishery Bulletin, 13(1):??, 1893. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/atmospheric-and-other-influences-
migrations-fishes;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb13.14.pdf.

Blackford:1893:ACF

[1049] Eugene G. Blackford. Address as chairman of the fish-cultural section. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/address-chairman-fish-cultural-section;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb13.
22.pdf.

McIntosh:1893:FCI

[1050] W. C. McIntosh. Fish-cultural investigations at the St. Andrews Marine Laboratory. Fishery Bulletin, 13(1):??, 1893. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/fish-cultural-investigations-st-
andrews-marine-laboratory;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb13.30.pdf.

Lawrence-Hamilton:1893:FFF

[1051] J. Lawrence-Hamilton. Foul fish and filth fevers. Fishery Bulletin, 13(1): ??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/foul-fish-and-
filth-fevers;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb13.38.pdf.

Augur:1893:FNS

[1052] C. H. Augur. Fish nets: Some account of their construction and the application of their various forms in the American fisheries. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/fish-
nets-some-account-their-construction-and-application-their-
various-forms-american;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb13.46.pdf.

Anonymous:1893:P

[1053] Anonymous. Preface. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/preface-5;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb13%
20preface.pdf.

Gilbert:1893:PPF

[1054] W. L. Gilbert. The past, present, and future of trout-culture. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/past-present-and-future-trout-culture;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb13.7.pdf.

Linton:1893:OCF

[1055] Edwin Linton. Observations concerning fish parasites. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
observations-concerning-fish-parasites;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb13.15.
pdf.

Borodine:1893:RFC

[1056] Nicolas Borodine. A review of fish culture in Europe and North America. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/review-fish-culture-europe-and-north-america;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb13.23.pdf.

Fulton:1893:DMH

[1057] T. Wemyss Fulton. Description of the marine hatchery at Dunbar, Scotland. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/description-marine-hatchery-dunbar-scotland;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb13.31.pdf.

Dean:1893:NHS

[1058] Bashford Dean. Notes on the hatching of sturgeon. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
notes-hatching-sturgeon;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb13.39.pdf.

Smith:1893:SFU

[1059] Hugh M. Smith. Statistics of the fisheries of the United States. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
statistics-fisheries-united-states;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb13.47.pdf.

Anonymous:1893:EI

[1060] Anonymous. Explanation of illustrations. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/explanation-

illustrations-6;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb13%20plates.pdf.

Goode:1893:RSR

[1061] George Brown Goode. The relation of scientific research to economic problems. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/relation-scientific-research-economic-problems;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb13.8.pdf.

Peck:1893:FM

[1062] James I. Peck. On the food of the menhaden. Fishery Bulletin, 13(1): ??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/food-menhaden;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb13.16.pdf.

Nordqvist:1893:SNA

[1063] Oscar Nordqvist. Some notes about American fish-culture. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/some-
notes-about-american-fish-culture;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb13.24.pdf.

Oemler:1893:PPF

[1064] A. Oemler. The past, present, and future of the oyster industry of Georgia. Fishery Bulletin, 13(1):??, 1893. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/past-present-and-future-oyster-
industry-georgia;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb13.32.pdf.

Joncas:1893:FC

[1065] L. Z. Joncas. The fisheries of Canada. Fishery Bulletin, 13(1):??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/fisheries-canada;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb13.40.pdf.

Smith:1893:FJ

[1066] Hugh M. Smith. The fisheries of Japan. Fishery Bulletin, 13(1): ??, 1893. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fisheries-
japan;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb13.48.pdf.

Anonymous:1894:EI

[1067] Anonymous. Explanation of illustrations. Fishery Bulletin, 14(1):??, 1894. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/explanation-
illustrations-7;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb14%20plates.pdf.

Woolman:1894:RCF

[1068] Albert J. Woolman. Report on a collection of fishes from the rivers of central and northern Mexico. Fishery Bulletin, 14(1):??, 1894. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/report-collection-fishes-rivers-
central-and-northern-mexico;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb14.8.pdf.

McDonald:1894:SFC

[1069] Marshall McDonald. The salmon fisheries of the Columbia River. Fishery Bulletin, 14(1):??, 1894. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/salmon-fisheries-columbia-river;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb14.
16.pdf.

Evermann:1894:FCB

[1070] B. W. Evermann and Cloudsley Rutter. The fishes of the Colorado Basin. Fishery Bulletin, 14(1):??, 1894. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fishes-colorado-basin;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb14.24.pdf.

Smith:1894:NTH

[1071] Hugh M. Smith. Notes on two hitherto unrecognized species of American whitefishes. Fishery Bulletin, 14(1):??, 1894. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/notes-two-hitherto-unrecognized-
species-american-whitefishes;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb14.1.pdf.

Meek:1894:RIR

[1072] Seth Eugene Meek. Report of investigations respecting the fishes of Arkansas, conducted during 1891, 1892, and 1893, with a synopsis of previous explorations in the same state. Fishery Bulletin, 14(1): ??, 1894. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/report-
investigations-respecting-fishes-arkansas-conducted-during-
1891-1892-and-1893;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb14.9.pdf.

Gilbert:1894:RUP

[1073] C. H. Gilbert and B. W. Evermann. A report upon physical and natural history investigations in the Columbia River Basin. Fishery Bulletin, 14 (1):??, 1894. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/report-upon-
physical-and-natural-history-investigations-columbia-river-
basin;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb14.17.pdf.

Anonymous:1894:I

[1074] Anonymous. Index. Fishery Bulletin, 14(1):??, 1894. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-52;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb14.index_.pdf.

Kendall:1894:ERR

[1075] W. C. Kendall and Hugh M. Smith. Extension of the recorded range of certain marine and fresh-water fishes of the Atlantic Coast of the United States. Fishery Bulletin, 14(1):??, 1894. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/extension-recorded-range-certain-
marine-and-fresh-water-fishes-atlantic-coast-united-states;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb14.2.pdf.

Smith:1894:NCA

[1076] Hugh M. Smith. Notes on the capture of Atlantic salmon at sea and in the coast waters of the eastern states. Fishery Bulletin, 14(1):??, 1894. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
notes-capture-atlantic-salmon-sea-and-coast-waters-eastern-

states;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb14.10.pdf.

Henshall:1894:NFC

[1077] James A. Henshall. Notes on fishes collected in Florida in 1892. Fishery Bulletin, 14(1):??, 1894. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/notes-fishes-collected-florida-1892;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb14.
18.pdf.

Gilbert:1894:NFB

[1078] Charles H. Gilbert. Notes on fishes from the basin of the Mackenzie River in British America. Fishery Bulletin, 14(1):??, 1894. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/notes-fishes-basin-
mackenzie-river-british-america;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb14.3.pdf.

Eigenmann:1894:REW

[1079] Carl H. Eigenmann. Results of explorations in western Canada and the northwestern United States. Fishery Bulletin, 14(1):??, 1894. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/results-
explorations-western-canada-and-northwestern-united-states;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb14.11.pdf.

Smith:1894:NRF

[1080] Hugh M. Smith. Notes on a reconnoissance of the fisheries of the Pacific Coast of the United States in 1894. Fishery Bulletin, 14(1):??, 1894. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
notes-reconnoissance-fisheries-pacific-coast-united-states-
1894;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb14.19.pdf.

Nordqvist:1894:AFF

[1081] Oscar Nordqvist. An American fish in Finland. Fishery Bulletin, 14(1): ??, 1894. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/american-fish-
finland;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb14.4.pdf.

Meek:1894:NFW

[1082] Seth Eugene Meek. Notes on the fishes of western Iowa and eastern Nebraska. Fishery Bulletin, 14(1):??, 1894. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/notes-fishes-western-iowa-and-eastern-nebraska;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb14.12.pdf.

Page:1894:FRF

[1083] William F. Page. Feeding and rearing fishes, particularly trout, under domestication. Fishery Bulletin, 14(1):??, 1894. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/feeding-and-rearing-
fishes-particularly-trout-under-domestication;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb14.
20.pdf.

Knauthe:1894:TFC

[1084] Karl Knauthe.  Two fertile cyprinoid hybrids.  Fishery Bul-

letin, 14(1):??, 1894. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
two-fertile-cyprinoid-hybrids;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb14.5.pdf.

Jordan:1894:LFI

[1085] D. S. Jordan and C. H. Gilbert. List of the fishes inhabiting Clear Lake, California. Fishery Bulletin, 14(1):??, 1894. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/list-fishes-inhabiting-clear-lake-california;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb14.13.pdf.

Kirsch:1894:RUI

[1086] Philip H. Kirsch. Report upon investigations in the Maumee River Basin during the summer of 1893. Fishery Bulletin, 14(1):??, 1894. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/report-upon-
investigations-maumee-river-basin-during-summer-1893;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb14.21.pdf.

Kirsch:1894:RUE

[1087] Philip H. Kirsch. Report upon explorations made in Eel River Basin in the northeastern part of Indiana in the summer of 1892. Fishery Bulletin, 14(1):??, 1894. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
report-upon-explorations-made-eel-river-basin-northeastern-
part-indiana-summer-1892;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb14.6.pdf.

Jordan:1894:NFW

[1088] David Starr Jordan. Notes on the fresh-water species of San Luis Obispo County, California. Fishery Bulletin, 14(1):??, 1894. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/notes-fresh-water-species-san-luis-obispo-
county-california;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb14.14.pdf.

Smith:1894:SRF

[1089] Hugh M. Smith. A statistical report on the fisheries of the middle Atlantic states. Fishery Bulletin, 14(1):??, 1894. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/statistical-report-
fisheries-middle-atlantic-states;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb14.22.pdf.

Kendall:1894:NFW

[1090] William C. Kendall. Notes on the fresh-water fishes of Washington County, Maine. Fishery Bulletin, 14(1):??, 1894. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/notes-fresh-water-fishes-
washington-county-maine;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb14.7.pdf.

Tanner:1894:ACP

[1091] Z. L. Tanner. On the appliances for collecting pelagic organisms, with special reference to those employed by the U. S. Fish Commission. Fishery Bulletin, 14(1):??, 1894. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
appliances-collecting-pelagic-organisms-special-reference-those-
employed-u-s-fish;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb14.15.pdf.

Evermann:1894:LSF

[1092] B. W. Evermann and W. C. Kendall. A list of the species of fishes known from the vicinity of Neosho, Mo. Fishery Bulletin, 14(1):??, 1894. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/list-species-fishes-
known-vicinity-neosho-mo;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb14.23.pdf.

Evermann:1895:PRU

[1093] Barton W. Evermann. A preliminary report upon salmon investigations in Idaho in 1894. Fishery Bulletin, 15(1):??, 1895. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/preliminary-report-upon-salmon-
investigations-idaho-1894;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb15.2.pdf.

Smith:1895:RHR

[1094] Hugh M. Smith. A review of the history and results of the attempts to acclimatize fish and other water animals in the Pacific states. Fishery Bulletin, 15(1):??, 1895. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
review-history-and-results-attempts-acclimatize-fish-and-other-
water-animals-pacific-states;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb15.10.pdf.

Smith:1895:NIM

[1095] Hugh M. Smith. Notes on an investigation of the menhaden fishery in 1894, with special reference to the food-fishes taken. Fishery Bulletin, 15(1):??, 1895. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/notes-
investigation-menhaden-fishery-1894-special-reference-food-
fishes-taken;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb15.3.pdf.

Anonymous:1895:I

[1096] Anonymous. Index. Fishery Bulletin, 15(1):??, 1895. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-53;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb15.index_.pdf.

Evermann:1895:FNR

[1097] Barton W. Evermann and Ulysses O. Cox. The fishes of the Neuse River Basin. Fishery Bulletin, 15(1):??, 1895. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/fishes-neuse-river-basin;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb15.4.pdf.

Jaffe:1895:NFC

[1098] S. Jaffé. Notes on fish culture in Germany. Fishery Bulletin, 15(1):??, 1895. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/notes-fish-
culture-germany;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb15.5.pdf.

Ritter:1895:RRO

[1099] Homer P. Ritter. Report on a reconnoissance of the oyster beds of Mobile Bay and Mississippi Sound, Alabama. Fishery Bulletin, 15 (1):??, 1895. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/report-
reconnoissance-oyster-beds-mobile-bay-and-mississippi-sound-
alabama;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb15.6.pdf.

Meek:1895:LFM

[1100] Seth Eugene Meek. A list of fishes and mollusks collected in Arkansas and Indian Territory in 1894. Fishery Bulletin, 15(1):??, 1895. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/list-fishes-and-mollusks-
collected-arkansas-and-indian-territory-1894;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb15.
7.pdf.

Anonymous:1895:EI

[1101] Anonymous. Explanation of illustrations. Fishery Bulletin, 15(1):??, 1895. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/explanation-
illustrations-8;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb15%20plates.pdf.

Peck:1895:SMF

[1102] James I. Peck. The sources of marine food. Fishery Bulletin, 15(1):??, 1895. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/sources-marine-food;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb15.8.pdf.

Herrick:1895:ALS

[1103] Francis Hobart Herrick. The American lobster; a study of its habits and development. Fishery Bulletin, 15(1):??, 1895. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/american-lobster-study-its-habits-and-development;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb15.1.pdf.

Wozelka-Iglau:1895:CTI

[1104] Karl Wozelka-Iglau. Contributions toward the improvement of the culture of salmonoids and crawfish in smaller water courses. Fishery Bulletin, 15(1):??, 1895. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
contributions-toward-improvement-culture-salmonoids-and-crawfish-
smaller-water-courses;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb15.9.pdf.

Stejneger:1896:RSI

[1105] Leonhard Stejneger. The Russian fur seal islands. Fishery Bulletin, 16(1): ??, 1896. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/russian-fur-
seal-islands;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb16.1.pdf.

Evermann:1896:RUS

[1106] Barton W. Evermann. A report upon salmon investigations in the head waters of the Columbia River, in the state of Idaho, in 1895, together with notes upon the fishes observed in that state in 1894 and 1895. Fishery Bulletin, 16(1):??, 1896. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/report-upon-salmon-
investigations-head-waters-columbia-river-state-idaho-1895-
together;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb16.2.pdf.

Stone:1896:APS

[1107] Livingston Stone. The artificial propagation of salmon on the Pacific Coast of the United States, with notes on the natural history of the quinnat salmon. Fishery Bulletin, 16(1):??, 1896. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/artificial-propagation-salmon-pacific-coast-
united-states-notes-natural-history-quinnat;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb16.
3.pdf.

Seagle:1896:APR

[1108] George A. Seagle. The artificial propagation of the rainbow trout. Fishery Bulletin, 16(1):??, 1896. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/artificial-propagation-rainbow-trout;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb16.4.pdf.

Tanner:1896:DSE

[1109] Z. L. Tanner. Deep sea exploration: a general description of the steamer albatross, her appliances and methods. Fishery Bulletin, 16(1):??, 1896. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
deep-sea-exploration-general-description-steamer-albatross-
her-appliances-and-methods;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb16.5.pdf.

Anonymous:1896:I

[1110] Anonymous. Index. Fishery Bulletin, 16(1):??, 1896. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-54;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb16.index_.pdf.

Anonymous:1896:EI

[1111] Anonymous. Explanation of illustrations. Fishery Bulletin, 16(1):??, 1896. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/explanation-
illustrations-9;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb16%20plates.pdf.

Smith:1897:FFV

[1112] Hugh M. Smith. Fishes found in the vicinity of Woods Hole. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/fishes-found-vicinity-woods-hole;   https:

//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb17.3.pdf.

Meek:1897:UBS

[1113] Seth E. Meek. The utility of a biological station on the Florida coast in its relation to the commercial fisheries. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/utility-biological-
station-florida-coast-its-relation-commercial-fisheries;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb17.11.pdf.

Wilson:1897:FRS

[1114] H. V. Wilson. On the feasibility of raising sponges from the egg. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/feasibility-raising-sponges-egg;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb17.
19.pdf.

Ruge:1897:OOB

[1115] John G. Ruge. The oyster and oyster-beds of Florida. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/oyster-
and-oyster-beds-florida;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb17.27.pdf.

Stone:1897:SSR

[1116] Livingston Stone. Some scattering reminiscences of the early days of fish-culture in the United States. Fishery Bulletin, 17 (1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/some-
scattering-reminiscences-early-days-fish-culture-united-states;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb17.35.pdf.

Greenley:1897:NTP

[1117] Arthur White Greenley. Notes on the tide-pool fishes of California, with a description of four new species. Fishery Bulletin, 17 (1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/notes-
tide-pool-fishes-california-description-four-new-species;

https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb19.2.pdf.

Smith:1897:SFP

[1118] Hugh M. Smith. Salmon fishery of Penobscot Bay and River in 1895 and 1896. Fishery Bulletin, 17(1):??, 1897. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/salmon-fishery-penobscot-bay-and-
river-1895-and-1896;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb17.4.pdf.

Taylor:1897:EBS

[1119] W. Edgar Taylor. Establishment of a biological station on the Gulf of Mexico. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/establishment-biological-station-gulf-mexico;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb17.12.pdf.

Cheney:1897:HRS

[1120] A. Nelson Cheney. The Hudson River as a salmon stream. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/hudson-river-salmon-stream;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb17.20.
pdf.

Zacharie:1897:LOI

[1121] F. C. Zacharie. The Louisiana oyster industry. Fishery Bulletin, 17 (1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/louisiana-
oyster-industry;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb17.28.pdf.

Meehan:1897:RBS

[1122] W. E. Meehan. The relations between State Fish Commissions and commercial fisherman. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/relations-between-state-fish-commissions-
and-commercial-fisherman;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb17.36.pdf.

Evermann:1897:DNL

[1123] Barton W. Evermann and William C. Kendall. Descriptions of new or little-known genera and species of fishes from the United States. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
descriptions-new-or-little-known-genera-and-species-fishes-
united-states;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb17.5.pdf.

Sigerfoos:1897:SNA

[1124] Charles P. Sigerfoos. Some notes on American shipworms. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
some-notes-american-shipworms;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb17.13.pdf.

Henshall:1897:PDP

[1125] James A. Henshall. A pleas for the development and protection of Florida fish and fisheries. Fishery Bulletin, 17(1):??, 1897. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/pleas-development-and-protection-
florida-fish-and-fisheries;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb17.21.pdf.

Smeltz:1897:OBW

[1126] H. A. Smeltz. The oyster-bars of the West Coast of Florida; their depletion and restoration. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/oyster-bars-west-coast-florida-their-
depletion-and-restoration;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb17.29.pdf.

Cobb:1897:PID

[1127] John N. Cobb. Possibilities for an increased development of Florida’s fishery resources. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/possibilities-increased-development-floridas-
fishery-resources;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb17.37.pdf.

Brice:1897:WUS

[1128] John J. Brice. The work of the United States Fish Commission from December 1, 1896, to November 3, 1897. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/work-united-states-fish-commission-december-1-1896-
november-3-1897;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb17.6.pdf.

Linton:1897:ECF

[1129] Edwin Linton. An economical consideration of fish parasites. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/economical-consideration-fish-parasites;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb17.14.pdf.

James:1897:IPD

[1130] Bushrod W. James. International protection for denizens of the sea and waterways. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/international-protection-denizens-sea-and-
waterways;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb17.22.pdf.

Detwiler:1897:NFI

[1131] John Y. Detwiler. Notes on the fishing industry of eastern Florida. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
notes-fishing-industry-eastern-florida;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb17.30.
pdf.

Moore:1897:UMM

[1132] J. Percy Moore. The utility and methods of mackerel propagation. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/utility-and-methods-mackerel-propagation;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb17.38.pdf.

Alexander:1897:NHF

[1133] A. B. Alexander. Notes on the halibut fishery of the northwest coast in 1896. Fishery Bulletin, 17(1):??, 1897. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/notes-halibut-fishery-northwest-

coast-1896;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb17.7.pdf.

Evermann:1897:FFF

[1134] Barton W. Evermann. The fish fauna of Florida. Fishery Bulletin, 17(1): ??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/fish-fauna-
florida;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb17.15.pdf.

Stevenson:1897:RIR

[1135] Charles H. Stevenson. The restricted inland range of shad due to artificial obstructions and its effect upon natural reproduction. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
restricted-inland-range-shad-due-artificial-obstructions-and-
its-effect-upon-natural;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb17.23.pdf.

Kibbe:1897:OOC

[1136] I. P. Kibbe. Oysters and oyster-culture in Texas. Fishery Bulletin, 17(1): ??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/oysters-and-
oyster-culture-texas;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb17.31.pdf.

Sharp:1897:LMB

[1137] John Sharp. The large-mouthed black bass in Utah. Fishery Bulletin, 17 (1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/large-mouthed-
black-bass-utah;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb17.39.pdf.

Anonymous:1897:EI

[1138] Anonymous. Explanation of illustrations. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/explanation-
illustrations-10;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb17%20plates.pdf.

Anonymous:1897:PPN

[1139] Anonymous. Proceedings and papers of the national fishery congress, separated by article. Fishery Bulletin, 17(1):??, 1897. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/proceedings-and-papers-national-
fishery-congress-separated-article;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb17.8.pdf.

Surface:1897:LCN

[1140] H. A. Surface. The lampreys of Central New York. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/lampreys-central-new-york;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb17.16.pdf.

Munroe:1897:GTP

[1141] Ralph M. Munroe. The green turtle, and the possibilities of its protection and consequent increase on the Florida coast. Fishery Bulletin, 17(1): ??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/green-turtle-
and-possibilities-its-protection-and-consequent-increase-florida-
coast;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb17.24.pdf.

Stranahan:1897:MLR

[1142] J. J. Stranahan. The methods, limitations, and results of whitefish culture in Lake Erie. Fishery Bulletin, 17(1):??, 1897. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/methods-limitations-and-results-
whitefish-culture-lake-erie;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb17.32.pdf.

Willson:1897:FF

[1143] J. M. Willson, Jr. Florida fur-farming. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/florida-fur-farming;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb17.40.pdf.

Gilbert:1897:FKR

[1144] Charles H. Gilbert. Fishes of the Klamath River Basin. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fishes-klamath-river-basin;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb17.1.pdf.

Reighard:1897:MPI

[1145] Jacob Reighard. Methods of plankton investigation in their relation to practical problems. Fishery Bulletin, 17(1):??, 1897. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/methods-plankton-investigation-
their-relation-practical-problems;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb17.9.pdf.

Herrick:1897:PLF

[1146] Francis H. Herrick. The protection of the lobster fishery. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/protection-lobster-fishery;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb17.17.
pdf.

Moore:1897:SFO

[1147] H. F. Moore. Some factors in the oyster problem. Fishery Bulletin, 17(1): ??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/some-factors-
oyster-problem;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb17.25.pdf.

Kunz:1897:BHG

[1148] George F. Kunz. A brief history of the gathering of freshwater pearls in the United States. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/brief-history-
gathering-fresh-water-pearls-united-states;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb17.33.
pdf.

Kunz:1897:FWP

[1149] George F. Kunz. The fresh-water pearls and pearl fisheries of the United States. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fresh-water-pearls-and-pearl-fisheries-united-states;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb17.41.pdf.

Evermann:1897:SIC

[1150] Barton Warren Evermann and Seth Eugene Meek. Salmon investigations in the Columbia River Basin and elsewhere on the Pacific Coast in 1896. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/salmon-investigations-columbia-river-
basin-and-elsewhere-pacific-coast-1896;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb17.2.pdf.

Bumpus:1897:IES

[1151] H. C. Bumpus. The importance of extended scientific investigation. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/importance-extended-scientific-investigation;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb17.10.pdf.

Smith:1897:FCS

[1152] Hugh M. Smith. The Florida commercial sponges. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/florida-commercial-sponges;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb17.18.
pdf.

Swift:1897:OGW

[1153] Franklin Swift. The oyster-grounds of the west Florida coast; their extent, condition, and peculiarities. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/oyster-grounds-west-
florida-coast-their-extent-condition-and-peculiarities;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb17.26.pdf.

Warren:1897:RSF

[1154] Andrew F. Warren. The red-snapper fisheries; their past, present, and future. Fishery Bulletin, 17(1):??, 1897. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/red-snapper-fisheries-their-past-
present-and-future;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb17.34.pdf.

Anonymous:1897:I

URL https://spo.nmfs.noaa.gov/content/index-55;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb17.index_.pdf.

Anonymous:1898:EI

[1156] Anonymous. Explanation of illustrations. Fishery Bulletin, 18(1):??, 1898. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/explanation-
illustrations-11;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb18%20plates.pdf.

Herrick:1898:PNS

[1157] C. Judson Herrick. The peripheral nervous system of the bony fishes. Fishery Bulletin, 18(1):??, 1898. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/peripheral-nervous-system-bony-fishes;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb18.8.pdf.

Moser:1898:SSF

[1158] Jefferson F. Moser. The salmon and salmon fisheries of Alaska. Report of the operations of the United States Fish Commission steamer Albatross for the year ending June 30, 1898. Fishery Bulletin, 18(1):??, 1898. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/salmon-and-salmon-fisheries-alaska-report-
operations-united-states-fish-commission-steamer;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb18.1_0.pdf.

Bumpus:1898:RT

[1159] Hermon C. Bumpus. The reappearance of the tilefish. Fishery Bulletin, 18(1):321–333, 1898. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/reappearance-tilefish;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb18.9.pdf.

Smith:1898:LFK

[1160] Hugh M. Smith and Barton A. Bean. List of fishes known to inhabit the waters of the District of Columbia and vicinity. Fishery Bulletin, 18(1):??, 1898. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/list-
fishes-known-inhabit-waters-district-columbia-and-vicinity;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb18.2.pdf.

Stevenson:1898:PFP

[1161] Charles H. Stevenson. The preservation of fishery products for food. Fishery Bulletin, 18(1):??, 1898. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
preservation-fishery-products-food;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb18.10.pdf.

Rutter:1898:NCT

[1162] Cloudsley Rutter. Notes on a collection of tide-pool fishes from Kadiak Island in Alaska. Fishery Bulletin, 18(1):189–192, 1898. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/notes-collection-tide-pool-fishes-
kadiak-island-alaska;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb18.3.pdf.

Anonymous:1898:I

[1163] Anonymous. Index. Fishery Bulletin, 18(1):??, 1898. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-56;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb18.index_.pdf.

Smith:1898:SSM

[1164] Hugh M. Smith. The southern spring mackerel fishery of the unites states. Fishery Bulletin, 18(1):??, 1898. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/southern-spring-mackerel-fishery-unites-states;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb18.4.pdf.

Smith:1898:NFN

[1165] Hugh M. Smith. Notice of a filefish new to the fauna of the United States. Fishery Bulletin, 18(1):??, 1898. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/notice-filefish-new-fauna-united-states;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb18.5.pdf.

Simpson:1898:PFW

[1166] Charles T. Simpson. The pearly fresh-water mussels of the United States; their habits, enemies, and diseases, with suggestions for their protection. Fishery Bulletin, 18(1):??, 1898. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/pearly-fresh-water-mussels-united-states-their-
habits-enemies-and-diseases-suggestions;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb18.6.pdf.

Smith:1898:MFP

[1167] Hugh M. Smith. The mussel fishery and pearl button industry of the Mississippi River. Fishery Bulletin, 18(1):??, 1898. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/mussel-fishery-and-pearl-
button-industry-mississippi-river;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb18.7.pdf.

Anonymous:1899:EI

[1168] Anonymous. Explanation of illustrations. Fishery Bulletin, 19(1):??, 1899. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/explanation-
illustrations-12;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb19%20plates.pdf.

Jordan:1899:NCF

[1169] David Starr Jordan and John O. Snyder. Notes on a collection of fishes from the rivers of Mexico, with descriptions of twenty two new species. Fishery Bulletin, 19(1):??, 1899. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/notes-collection-fishes-rivers-mexico-descriptions-
twenty-two-new-species;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb19.9.pdf.

Howe:1899:RDE

[1170] Freeland Howe. Report of a dredging expedition off the southern coast of New England, September, 1899. Fishery Bulletin, 19(1): ??, 1899. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/report-
dredging-expedition-southern-coast-new-england-september-1899;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb19.17.pdf.

Linton:1899:PFW

[1171] Edwin Linton. Parasites of fishes of the Woods Hole region. Fishery Bulletin, 19(1):??, 1899. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
parasites-fishes-woods-hole-region;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb19.25.pdf.

Shufeldt:1899:EPL

[1172] R. W. Shufeldt. Experiments in photography of live fishes. Fishery Bulletin, 19(1):??, 1899. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/experiments-photography-live-fishes;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb19.
1.pdf.

Smith:1899:NFS

[1173] Hugh M. Smith. Notes on the Florida sponge fishery in 1899. Fishery Bulletin, 19(1):??, 1899. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
notes-florida-sponge-fishery-1899;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb19.10.pdf.

Cobb:1899:LFM

[1174] John N. Cobb. The lobster fishery of Maine. Fishery Bulletin, 19 (1):??, 1899. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/lobster-
fishery-maine;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb19.18.pdf.

Anonymous:1899:I

[1175] Anonymous. Index. Fishery Bulletin, 19(1):??, 1899. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-57;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb19.Index_.pdf.

Clark:1899:SNE

[1176] Hubert Lyman Clark. The synaptas of the New England Coast. Fishery Bulletin, 19(1):??, 1899. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/synaptas-new-england-coast;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb19.3.pdf.

Levene:1899:SCC

[1177] P. A. Levene. Some chemical changes in the developing fish egg. Fishery Bulletin, 19(1):??, 1899. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/some-chemical-changes-developing-fish-egg;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb19.11.pdf.

Linton:1899:FPC

[1178] Edwin Linton. Fish parasites collected at Woods Hole in 1898. Fishery Bulletin, 19(1):??, 1899. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fish-parasites-collected-woods-hole-1898;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb19.19.pdf.

Greenly:1899:NTP

[1179] Arthur White Greenly. Notes on the tide: Pool fishes of California, with a description of four new species. Fishery Bulletin, 19(1):??, 1899. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/notes-tide%
C2%B7pool-fishes-california-description-four-new-species;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/greenley.pdf.

Gorham:1899:GBD

[1180] F. P. Gorham. The gas-bubble disease of fish and its cause. Fishery Bulletin, 19(1):??, 1899. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/gas-
bubble-disease-fish-and-its-cause;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb19.4.pdf.

Wheeler:1899:FSC

[1181] William Morton Wheeler. The free-swimming copepods of the Woods Hole region. Fishery Bulletin, 19(1):??, 1899. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/free-swimming-copepods-woods-hole-region;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb19.12.pdf.

Anonymous:1899:WHB

[1182] Anonymous. Woods Hole biological notes. Fishery Bulletin, 19(1):??, 1899. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/woods-hole-
biological-notes;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb19.20.pdf.

Kellogg:1899:CPC

[1183] James L. Kellogg. The clam problem and clam culture. Fishery Bulletin, 19(1):??, 1899. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
clam-problem-and-clam-culture;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb19.5.pdf.

Kellogg:1899:OLH

[1184] James L. Kellogg. Observations on the life history of the common clam, Mya arenaria. Fishery Bulletin, 19(1):??, 1899. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/observations-life-history-common-clam-mya-
arenaria;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb19.13.pdf.

Shufeldt:1899:SBB

[1185] R. W. Shufeldt. The skeleton of the black bass. Fishery Bulletin, 19(1):??, 1899. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/skeleton-black-bass;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb19.21.pdf.

Jenkins:1899:DNS

[1186] Oliver P. Jenkins. Descriptions of new species of fishes from the Hawaiian Islands, belonging to the families of Labridae and Scaridae. Fishery Bulletin, 19(1):??, 1899. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
descriptions-new-species-fishes-hawaiian-islands-belonging-
families-labridae-and-scaridae;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb19.6.pdf.

Mead:1899:NHS

[1187] A. D. Mead. The natural history of the starfish. Fishery Bulletin, 19 (1):??, 1899. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/natural-
history-starfish;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb19.14.pdf.

Green:1899:CCS

[1188] Erik H. Green. The chemical composition of the subnormal connective tissue of the ocean sun-fish. Fishery Bulletin, 19(1):??, 1899. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/chemical-composition-subnormal-
connective-tissue-ocean-sun-fish;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb19.22.pdf.

Jennings:1899:RUS

[1189] H. S. Jennings. Rotaria of the United States with special reference to those of the Great Lakes. Fishery Bulletin, 19(1):??, 1899. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/rotaria-
unites-states-special-reference-those-great-lakes;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb19.7.pdf.

Bumpus:1899:MCL

[1190] Hermon C. Bumpus. On the movements of certain lobsters liberated at Woods Hole. Fishery Bulletin, 19(1):??, 1899. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/movements-certain-lobsters-liberated-
woods-hole;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb19.15.pdf.

Nutting:1899:HWH

[1191] C. C. Nutting. The hydroids of the Woods Hole region. Fishery Bulletin, 19(1):??, 1899. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/hydroids-woods-hole-region;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb19.23.
pdf.

Jennings:1899:RWP

[1192] H. S. Jennings. A report of work on the protozoa of Lake Erie, with special reference to the laws of their movements. Fishery Bulletin, 19(1):??, 1899. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/report-work-protozoa-
lake-erie-special-reference-laws-their-movements;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb19.8.pdf.

Tower:1899:IPF

[1193] Ralph W. Tower. Improvements in preparing fish for shipment. Fishery Bulletin, 19(1):??, 1899. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/improvements-preparing-fish-shipment;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb19.16.pdf.

Jenkins:1899:DFN

[1194] Oliver P. Jenkins. Descriptions of fifteen new species of fishes from the Hawaiian Islands. Fishery Bulletin, 19(1):??, 1899. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/descriptions-fifteen-new-species-
fishes-hawaiian-islands;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb19.24.pdf.

Evermann:1900:SSR

[1195] Barton Warren Evermann. Summary of the scientific results of the fish commission expedition to Porto Rico. Fishery Bulletin, 20(1): ??, 1900. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/summary-
scientific-results-fish-commission-expedition-porto-rico;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb20.2.pdf.

Bigelow:1900:SPR

[1196] Robert Payne Bigelow. The stomatopoda of Porto Rico. Fishery Bulletin, 20(1):??, 1900. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
stomatopoda-porto-rico;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb20.10.pdf.

Vaughan:1900:SCP

[1197] T. Wayland Vaughan. The stony corals of the porto rican waters. Fishery Bulletin, 20(1):??, 1900. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/stony-corals-porto-rican-waters;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb20.
18.pdf.

Evermann:1900:GRI

[1198] Barton Warren Evermann. General report on the investigations. Fishery Bulletin, 20(1):??, 1900. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
general-report-investigations;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb20.3.pdf.

Moore:1900:PRI

[1199] H. F. Moore. The Porto Rican isopoda. Fishery Bulletin, 20(1):??, 1900. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/porto-rican-isopoda;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb20.11.pdf.

Duerden:1900:APR

[1200] J. E. Duerden. The actinians of Porto Rico. Fishery Bulletin, 20(1): ??, 1900. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/duerden-j-e;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb20.19.pdf.

Wilcox:1900:FFT

[1201] William A. Wilcox. The fisheries and fish trade of Porto Rico. Fishery Bulletin, 20(1):??, 1900. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fisheries-and-fish-trade-porto-rico;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb20.
4.pdf.

Bigelow:1900:CCN

[1202] Maurice A. Bigelow. The Cirripedia collected near Porto Rico. Fishery Bulletin, 20(1):??, 1900. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/cirripedia-collected-near-porto-rico;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb20.12.pdf.

Wilson:1900:SCP

[1203] H. V. Wilson. The sponges collected in Porto Rico in 1899. Fishery Bulletin, 20(1):??, 1900. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
sponges-collected-porto-rico-1899;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb20.20.pdf.

Evermann:1900:FPR

[1204] Barton Warren Evermann and Millard Caleb Marsh. The fishes of Porto Rico. Fishery Bulletin, 20(1):??, 1900. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fishes-porto-rico;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb20.5.pdf.

Treadwell:1900:PAP

[1205] A. L. Treadwell. The polychætous annelids of Porto Rico. Fishery Bulletin, 20(1):??, 1900. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/polych%C3%A6tous-annelids-porto-rico;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb20.13.pdf.

Flint:1900:FPR

[1206] James M. Flint. The foraminifera of Porto Rico. Fishery Bulletin, 20(1):??, 1900. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
foraminifera-porto-rico;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb20.21.pdf.

Dall:1900:MPR

[1207] William Healy Dall and Charles Torrey Simpson. The mollusk of Porto Rico. Fishery Bulletin, 20(1):??, 1900. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/mollusk-porto-rico;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb20.6.pdf.

Moore:1900:DTN

[1208] J. Percy Moore. Description of two new leeches from Porto Rico. Fishery Bulletin, 20(1):??, 1900. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
description-two-new-leeches-porto-rico;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb20.14.
pdf.

Anonymous:1900:I

[1209] Anonymous. Index. Fishery Bulletin, 20(1):??, 1900. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-58;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb20.index_.pdf.

Anonymous:1900:LIP

[1210] Anonymous. List of illustrations: Part II. Fishery Bulletin, 20(1):??, 1900. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/list-illustrations-
part-ii;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb20.7.pdf.

Coe:1900:NPR

[1211] Wesley R. Coe. The nemerteans of Porto Rico. Fishery Bulletin, 20(1): ??, 1900. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/nemerteans-
porto-rico;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb20.15.pdf.

Anonymous:1900:LI

[1212] Anonymous. List of illustrations. Fishery Bulletin, 20(1):??, 1900. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/list-illustrations;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb20%20plates.pdf.

Rathbun:1900:BMP

[1213] Mary J. Rathbun. The Brachyura and Macrura of Porto Rico. Fishery Bulletin, 20(1):??, 1900. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/brachyura-and-macrura-porto-rico;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb20.8.pdf.

Clark:1900:EPR

[1214] Hubert Lyman Clark. The echinoderms of Porto Rico. Fishery Bulletin, 20(1):??, 1900. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
echinoderms-porto-rico;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb20.16.pdf.

Bowers:1900:P

[1215] George Bowers. Preface. Fishery Bulletin, 20(1):??, 1900. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/preface-6;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb20.
1.pdf.

Benedict:1900:AC

[1216] James E. Benedict. The anomuran collections. Fishery Bulletin, 20(1):??, 1900. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/anomuran-collections;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb20.9.pdf.

Hargitt:1900:APR

[1217] Charles W. Hargitt and Charles G. Rogers. The alcyonaria of Porto Rico. Fishery Bulletin, 20(1):??, 1900. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/alcyonaria-porto-rico;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb20.17.pdf.

Sherwood:1901:NMS

[1218] George H. Sherwood and Vinal N. Edwards. Notes on the migration, spawning, abundance, etc., of certain fishes in 1900. Fishery Bulletin, 21(1):??, 1901. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
notes-migration-spawning-abundance-etc-certain-fishes-1900;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb21.2.pdf.

Evermann:1901:LSF

[1219] Barton Warren Evermann. List of species of fishes known to occur in the Great Lakes or their connecting waters. Fishery Bulletin, 21(1):??, 1901. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/list-species-fishes-known-occur-great-lakes-or-
their-connecting-waters;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb21.10.pdf.

Moser:1901:ASI

[1220] Jefferson F. Moser. Alaska salmon investigations in 1900 and 1901. Fishery Bulletin, 21(1):??, 1901. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
alaska-salmon-investigations-1900-and-1901;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb21.18.
pdf.

Eigenmann:1901:DNO

[1221] Carl H. Eigenmann. Description of a new oceanic fish found off southern New England. Fishery Bulletin, 21(1):??, 1901. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/description-new-oceanic-fish-found-southern-
new-england;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb21.3.pdf.

Green:1901:OCS

[1222] Erik H. Green and Ralph W. Tower. The organic constituents of the scales of fish. Fishery Bulletin, 21(1):??, 1901. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/organic-constituents-scales-fish;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb21.11.pdf.

Coker:1901:NSB

[1223] Robert E. Coker. Notes on a species of barnacle (Dichelaspis) parasitic on the gills of edible crabs. Fishery Bulletin, 21(1):??, 1901. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/notes-species-
barnacle-dichelaspis-parasitic-gills-edible-crabs;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb21.19.pdf.

Eigenmann:1901:EDC

[1224] Carl H. Eigenmann. The egg and development of the conger eel. Fishery Bulletin, 21(1):??, 1901. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/egg-and-development-conger-eel;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb21.4.pdf.

Parker:1901:RCV

[1225] G. H. Parker. The reactions of copepods to various stimuli, and the bearing of this on daily depth migrations. Fishery Bulletin, 21(1):??, 1901. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
reactions-copepods-various-stimuli-and-bearing-daily-depth-
migrations;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb21.12.pdf.

Calkins:1901:MPW

[1226] Gary N. Calkins. Marine protozoa from Woods Hole. Fishery Bulletin, 21(1):??, 1901. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/marine-protozoa-woods-hole;   https://spo.nmfs.

noaa.gov/sites/default/files/pdf-content/fish-bull/fb21.20.
pdf.

Eigenmann:1901:IHY

[1227] Carl H. Eigenmann. Investigation into the history of the young squeteague. Fishery Bulletin, 21(1):??, 1901. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/investigation-history-young-squeteague;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb21.5.pdf.

Tower:1901:GSB

[1228] Ralph W. Tower. The gas in the swim-bladder of fishes. Fishery Bulletin, 21(1):??, 1901. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
gas-swim-bladder-fishes;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb21.13.pdf.

Anonymous:1901:I

[1229] Anonymous. Index. Fishery Bulletin, 21(1):??, 1901. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-59;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb21.index_.pdf.

Thompson:1901:NIP

[1230] Millett T. Thompson. A new isopod parasitic on the hermit crab. Fishery Bulletin, 21(1):??, 1901. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/new-isopod-parasitic-hermit-crab;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb21.6.pdf.

Tower:1901:BCS

[1231] Ralph W. Tower. Biliary calculi in the squeteague. Fishery Bulletin, 21(1):??, 1901. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/biliary-calculi-squeteague;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb21.14.
pdf.

Pieters:1901:PWL

[1232] A. J. Pieters. The plants of western Lake Erie, with observations on their distribution. Fishery Bulletin, 21(1):??, 1901. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/plants-western-lake-erie-
observations-their-distribution;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb21.7.pdf.

Evermann:1901:RFC

[1233] Barton Warren Evermann and E. L. Goldsborough. A report on fishes collected in Mexico and Central America, with notes and descriptions of five new species. Fishery Bulletin, 21(1):??, 1901. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/report-fishes-collected-mexico-
and-central-america-notes-and-descriptions-five-new-species;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb21.15.pdf.

Anonymous:1901:EI

[1234] Anonymous. Explanation of illustrations. Fishery Bulletin, 21(1):??, 1901. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/explanation-
illustrations-13;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb21%20plates.pdf.

Eigenmann:1901:LAE

[1235] Carl H. Eigenmann and C. H. Kennedy. The leptocephalus of the American eel and other American leptocephali. Fishery Bulletin, 21(1):??, 1901. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
leptocephalus-american-eel-and-other-american-leptocephali;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb21.8.pdf.

Herrick:1901:RPL

[1236] Francis H. Herrick. The reproductive period in the lobster. Fishery Bulletin, 21(1):??, 1901. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
reproductive-period-lobster;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb21.16.pdf.

Smith:1901:FSI

[1237] Hugh M. Smith. The French sardine industry. Fishery Bulletin, 21(1):??, 1901. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
french-sardine-industry;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb21.1.pdf.

Smith:1901:DNS

[1238] Hugh M. Smith. Description of a new species of blenny from Japan. Fishery Bulletin, 21(1):??, 1901. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/description-new-species-blenny-japan;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb21.9.pdf.

Smith:1901:NFF

[1239] Hugh M. Smith. Notes on the five food-fishes of Lake Buhi, Luzon, Philippine Islands. Fishery Bulletin, 21(1):??, 1901. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/notes-five-food-fishes-lake-buhi-
luzon-philippine-islands;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb21.17.pdf.

Dean:1902:JOC

[1240] Bashford Dean. Japanese oyster culture. Fishery Bulletin, 22(1): ??, 1902. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/japanese-
oyster-culture;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb22.2.pdf.

Herrick:1902:OST

[1241] C. Judson Herrick. The organ and sense of taste in fishes. Fishery Bulletin, 22(1):??, 1902. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/organ-
and-sense-taste-fishes;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb22.10.pdf.

Snyder:1902:CSF

[1242] John Otterbein Snyder. A catalogue of the shore fishes collected by the steamer Albatross about the Hawaiian Islands in 1902. Fishery Bulletin, 22(1):??, 1902. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/catalogue-shore-fishes-collected-steamer-albatross-
about-hawaiian-islands-1902;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb22.18.pdf.

Lydell:1902:HCB

[1243] Dwight Lydell. The habits and culture of the black bass. Fishery Bulletin, 22(1):??, 1902. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
habits-and-culture-black-bass;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb22.3.pdf.

Jennings:1902:RUS

[1244] H. S. Jennings. Rotatoria of the unites states. Fishery Bulletin, 22(1):??, 1902. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
rotatoria-unites-states;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb22.11.pdf.

Jordan:1902:NFC

[1245] David Starr Jordan. Notes on fishes collected in the Tortugas Archipelago. Fishery Bulletin, 22(1):??, 1902. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/notes-fishes-collected-tortugas-
archipelago;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb22.19.pdf.

Parker:1902:HAS

[1246] G. H. Parker. Hearing and allied senses in fishes. Fishery Bulletin, 22(1): ??, 1902. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/hearing-and-
allied-senses-fishes;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb22.4.pdf.

Kendall:1902:NSF

[1247] William Converse Kendall. Notes on some fresh-water fishes from Maine. Fishery Bulletin, 22(1):??, 1902. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/notes-some-fresh-water-fishes-maine;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb22.
12.pdf.

Clark:1902:EWH

[1248] Hubert Lyman Clark. The echinoderms of the Woods Hole region. Fishery Bulletin, 22(1):??, 1902. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
echinoderms-woods-hole-region;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb22.20.pdf.

Rutter:1902:NHQ

[1249] Cloudsley Rutter. Natural history of the quinnat salmon. Fishery Bulletin, 22(1):??, 1902. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/natural-history-quinnat-salmon;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb22.5.pdf.

Snow:1902:PAL

[1250] Julia W. Snow. The plankton algae of Lake Erie, with special reference to the Chlorophyceae. Fishery Bulletin, 22(1):??, 1902. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/plankton-algae-lake-erie-special-
reference-chlorophyceae;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb22.13.pdf.

Jordan:1902:LFD

[1251] David Starr Jordan and Edwin Chapin Starks. List of fishes dredged by the steamer Albatross off the coast of Japan in the summer of 1900, with descriptions of new species and a review of the Japanese Macrouridae. Fishery Bulletin, 22(1):??, 1902. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/list-fishes-dredged-
steamer-albatross-coast-japan-summer-1900-descriptions-new-
species-and;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb22.21.pdf.

Smith:1902:BHY

[1252] Hugh M. Smith and L. G. Harron. Breeding habits of the yellow cat-fish. Fishery Bulletin, 22(1):??, 1902. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/breeding-habits-yellow-cat-fish;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb22.6.pdf.

Moenkhaus:1902:DNS

[1253] William J. Moenkhaus. Description of a new species of darter from Tippecanoe Lake. Fishery Bulletin, 22(1):??, 1902. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/description-new-species-darter-
tippecanoe-lake;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb22.14.pdf.

Anonymous:1902:I

[1254] Anonymous. Index. Fishery Bulletin, 22(1):??, 1902. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-60;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb22.index_.pdf.

Beardsley:1902:DTF

[1255] A. E. Beardsley. The destruction of trout fry by hydra. Fishery Bulletin, 22(1):??, 1902. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
destruction-trout-fry-hydra;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb22.7.pdf.

Kendall:1902:HSC

[1256] William Converse Kendall. Habits of some of the commercial cat-fishes. Fishery Bulletin, 22(1):??, 1902. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/habits-some-commercial-cat-fishes;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb22.15.
pdf.

Anonymous:1902:EI

[1257] Anonymous. Explanation of illustrations. Fishery Bulletin, 22(1):??, 1902. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/explanation-
illustrations-14;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb22%20plates.pdf.

Jordan:1902:DNG

[1258] David Starr Jordan and Barton Warren Evermann. Descriptions of new genera and species of fishes from the Hawaiian Islands. Fishery Bulletin, 22(1):??, 1902. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
descriptions-new-genera-and-species-fishes-hawaiian-islands;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb22.8.pdf.

Marsh:1902:MCD

[1259] M. C. Marsh. A more complete description of bacterium truttae. Fishery Bulletin, 22(1):??, 1902. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
more-complete-description-bacterium-truttae;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb22.
16.pdf.

Smith:1902:OHF

[1260] Hugh M. Smith. Observations on the herring fisheries of England, Scotland, and Holland. Fishery Bulletin, 22(1):??, 1902. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/observations-herring-
fisheries-england-scotland-and-holland;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb22.1.pdf.

Eigenmann:1902:FWF

[1261] Carl H. Eigenmann. The fresh-water fishes of western Cuba. Fishery Bulletin, 22(1):??, 1902. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fresh-water-fishes-western-cuba;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb22.
9.pdf.

Jenkins:1902:RCF

[1262] Oliver P. Jenkins. Report on collections of fishes made in the Hawaiian Islands, with descriptions of new species. Fishery Bulletin, 22(1):??, 1902. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/report-collections-
fishes-made-hawaiian-islands-descriptions-new-species;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb22.17.pdf.

Anonymous:1903:PTa

[1263] Anonymous. Pages 51 through 100. Fishery Bulletin, 23(1):??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/pages-51-through-100;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb23(1).2.pdf.

Anonymous:1903:PTb

[1264] Anonymous. Pages 451 through 500. Fishery Bulletin, 23(1): ??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/pages-
451-through-500;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb23(1).10.pdf.

Anonymous:1903:PTc

[1265] Anonymous. Pages 101 through 150. Fishery Bulletin, 23(1): ??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/pages-
101-through-150;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb23(1).3.pdf.

Anonymous:1903:PTd

[1266] Anonymous. Pages 501 through 533. Fishery Bulletin, 23(1): ??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/pages-
501-through-533;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb23(1).11.pdf.

Anonymous:1903:PTe

[1267] Anonymous. Pages 151 through 200. Fishery Bulletin, 23(1): ??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/pages-
151-through-200;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb23(1).4.pdf.

Anonymous:1903:A

[1268] Anonymous. Addenda. Fishery Bulletin, 23(1):??, 1903. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/addenda;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb23(1).12.pdf.

Anonymous:1903:PTf

[1269] Anonymous. Pages 201 through 250. Fishery Bulletin, 23(1): ??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/pages-
201-through-250;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb23(1).5.pdf.

Anonymous:1903:INN

[1270] Anonymous. Index to native names of Hawaiian fishes. Fishery Bulletin, 23(1):??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/index-
native-names-hawaiian-fishes;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb23(1).13.pdf.

Anonymous:1903:PTg

[1271] Anonymous. Pages 251 through 300. Fishery Bulletin, 23(1): ??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/pages-
251-through-300;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb23(1).6.pdf.

Anonymous:1903:Ia

[1272] Anonymous. Index. Fishery Bulletin, 23(1):??, 1903. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-61;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb23(1).13_0.pdf.

Anonymous:1903:PTh

[1273] Anonymous. Pages 301 through 350. Fishery Bulletin, 23(1): ??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/pages-
301-through-350;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb23(1).7.pdf.

Anonymous:1903:CP

[1274] Anonymous. Color plates. Fishery Bulletin, 23(1):??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/color-plates;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb23(1).
colorplates.pdf.

Anonymous:1903:LIa

[1275] Anonymous. List of illustrations. Fishery Bulletin, 23(1):??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/list-illustrations-0;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb23(1)%20plates.pdf.

Anonymous:1903:PTi

[1276] Anonymous. Pages 351 through 400. Fishery Bulletin, 23(1): ??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/pages-
351-through-400;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb23(1).8.pdf.

Anonymous:1903:PTj

[1277] Anonymous. Pages 1 through 50. Fishery Bulletin, 23(1):??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/pages-1-through-50;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb23(1).1.pdf.

Anonymous:1903:PTk

[1278] Anonymous. Pages 401 through 450. Fishery Bulletin, 23(1): ??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/pages-
401-through-450;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb23(1).9.pdf.

Anonymous:1903:PTl

[1279] Anonymous. Pages 577 through 600. Fishery Bulletin, 23(2): ??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/pages-
577-through-600;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb23(2).1.pdf.

Anonymous:1903:PTm

[1280] Anonymous. Pages 601 through 650. Fishery Bulletin, 23(2): ??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/pages-
601-through-650;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb23(2).2.pdf.

Anonymous:1903:PTn

[1281] Anonymous. Pages 651 through 699. Fishery Bulletin, 23(2): ??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/pages-
651-through-699;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb23(2).3.pdf.

Anonymous:1903:PTo

[1282] Anonymous. Pages 717 through 750. Fishery Bulletin, 23(2): ??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/pages-
717-through-750;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb23(2).4.pdf.

Anonymous:1903:PTp

[1283] Anonymous. Pages 751 through 765. Fishery Bulletin, 23(2): ??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/pages-
751-through-765;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb23(2).5.pdf.

Anonymous:1903:LSS

[1284] Anonymous. List of stations and of species collected at each station. Fishery Bulletin, 23(2):??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
list-stations-and-species-collected-each-station;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb23(2).6.pdf.

Anonymous:1903:Ib

[1285] Anonymous. Index. Fishery Bulletin, 23(2):??, 1903. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-62;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb23(2).index_.pdf.

Anonymous:1903:LIb

[1286] Anonymous. List of illustrations. Fishery Bulletin, 23(2):??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/list-illustrations-1;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb23(2)%20plates.pdf.

True:1903:NPG

[1287] Frederick W. True. Notes on a porpoise of the genus Prodelphinus from the Hawaiian Islands. Fishery Bulletin, 23(3):??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/notes-porpoise-genus-
prodelphinus-hawaiian-islands;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb23(3).2.pdf.

Anonymous:1903:Ic

[1288] Anonymous. Index. Fishery Bulletin, 23(3):??, 1903. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-63;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb23(3).index_.pdf.

Richardson:1903:ICH

[1289] Harriet Richardson. Isopods collected at the Hawaiian Islands by the U. S. Fish Commission steamer Albatross. Fishery Bulletin, 23 (3):??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/isopods-
collected-hawaiian-islands-u-s-fish-commission-steamer-albatross;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb23(3).3.pdf.

Rathbun:1903:BMH

[1290] Mary J. Rathbun. The Brachyura and Macrura of the Hawaiian Islands. Fishery Bulletin, 23(3):??, 1903. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/brachyura-and-macrura-hawaiian-
islands;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb23(3).4.pdf.

Nutting:1903:HHI

[1291] C. C. Nutting. Hydroids of the Hawaiian Islands collected by the steamer Albatross in 1902. Fishery Bulletin, 23(3):??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/hydroids-hawaiian-islands-collected-steamer-
albatross-1902;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb23(3).5.pdf.

Coe:1903:NHI

[1292] Wesley R. Coe. Nemerteans of the Hawaiian Islands collected by the steamer Albatross in 1902. Fishery Bulletin, 23(3):??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/nemerteans-
hawaiian-islands-collected-steamer-albatross-1902;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb23(3).7.pdf.

Fisher:1903:SHI

[1293] Walter K. Fisher. The starfishes of the Hawaiian Islands. Fishery Bulletin, 23(3):??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
starfishes-hawaiian-islands;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb23(3).8.pdf.

Anonymous:1903:LIc

[1294] Anonymous. List of illustrations. Fishery Bulletin, 23(3):??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/list-illustrations-2;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb23(3)%20plates.pdf.

Mayer:1903:MHI

[1295] Alfred Goldsborough Mayer. Medusæ of the Hawaiian Islands collected by the steamer Albatross in 1902. Fishery Bulletin, 23(3):??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/medus%C3%A6-
hawaiian-islands-collected-steamer-albatross-1902;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb23(3).9.pdf.

Fisher:1903:BLL

[1296] Walter K. Fisher. Birds of Laysan and the Leeward Islands, Hawaiian group. Fishery Bulletin, 23(3):??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/birds-laysan-and-leeward-islands-hawaiian-
group;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb23(3).1.pdf.

Treadwell:1903:PAH

[1297] Aaron L. Treadwell. Polychætous annelids of the Hawaiian Islands collected by the steamer Albatross in 1902. Fishery Bulletin, 23(3): ??, 1903. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/polych%C3%
A6tous-annelids-hawaiian-islands-collected-steamer-albatross-
1902;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb23(3).10.pdf.

Anonymous:1904:EI

[1298] Anonymous. Explanation of illustrations. Fishery Bulletin, 24(1):??, 1904. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/explanation-
illustrations-15;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb24%20plates.pdf.

Parker:1904:FLL

[1299] G. H. Parker. The function of the lateral-line organs in fishes. Fishery Bulletin, 24(1):??, 1904. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/function-lateral-line-organs-fishes;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb24.
8.pdf.

Holmes:1904:ASN

[1300] S. J. Holmes. The amphipoda of southern New England. Fishery Bulletin, 24(1):??, 1904. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/amphipoda-southern-new-england;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb24.16.
pdf.

Hay:1904:RMG

[1301] William Perry Hay. A revision of malaclemmys, a genus of turtles. Fishery Bulletin, 24(1):??, 1904. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/revision-malaclemmys-genus-turtles;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb24.1.pdf.

Richardson:1904:IAS

[1302] Harriet Richardson. Isopods from the Alaska salmon investigation. Fishery Bulletin, 24(1):??, 1904. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/isopods-alaska-salmon-investigation;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb24.
9.pdf.

Anonymous:1904:I

[1303] Anonymous. Index. Fishery Bulletin, 24(1):??, 1904. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-64;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb24.index_.pdf.

Hargitt:1904:MWH

[1304] Charles W. Hargitt. The medusæ of the Woods Hole region. Fishery Bulletin, 24(1):??, 1904. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/medusae-woods-hole-region;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb24.2.pdf.

Juday:1904:LFC

[1305] Chancey Juday. List of the fishes collected in Boulder County, Colorado, with a description of a new species of Leuciscus. Fishery Bulletin, 24(1): ??, 1904. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/list-fishes-
collected-boulder-county-colorado-description-new-species-leuciscus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb24.10.pdf.

Lucas:1904:OIR

[1306] Frederic A. Lucas. The osteology and immediate relations of the tile-fish, Lopholatilus chamaeleonticeps. Fishery Bulletin, 24(1):??, 1904. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/osteology-and-
immediate-relations-tile-fish-lopholatilus-chamaeleonticeps;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb24.3.pdf.

Jordan:1904:FFT

[1307] David Starr Jordan and Joseph C. Thompson. The fish fauna of the Tortugas Archipelago. Fishery Bulletin, 24(1):??, 1904. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/fish-fauna-tortugas-archipelago;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb24.11.pdf.

Silvester:1904:BVS

[1308] C. F. Silvester. The blood-vascular system of the tile-fish, Lopholatilus chamaeleonticeps. Fishery Bulletin, 24(1):??, 1904. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/blood-vascular-system-tile-fish-
lopholatilus-chamaeleonticeps;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb24.4.pdf.

Mitsukuri:1904:CMF

[1309] K. Mitsukuri. The cultivation of marine and fresh-water animals in Japan. Fishery Bulletin, 24(1):??, 1904. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/cultivation-marine-and-fresh-water-animals-japan;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb24.12.pdf.

Wilson:1904:FPG

[1310] Charles B. Wilson. The fish parasites of the genus Argulus found in the Woods Hole region. Fishery Bulletin, 24(1):??, 1904. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/fish-parasites-genus-
argulus-found-woods-hole-region;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb24.5.pdf.

Fisher:1904:NSD

[1311] Walter K. Fisher. New starfishes from deep water off California and Alaska. Fishery Bulletin, 24(1):??, 1904. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/new-starfishes-deep-water-california-
and-alaska;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb24.13.pdf.

Smith:1904:SIJ

[1312] Hugh M. Smith. The seaweed industries of Japan. Fishery Bulletin, 24 (1):??, 1904. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/seaweed-
industries-japan;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb24.6.pdf.

Linton:1904:PFB

[1313] Edwin Linton. Parasites of fishes of Beaufort, North Carolina. Fishery Bulletin, 24(1):??, 1904. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/parasites-fishes-beaufort-north-carolina;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb24.14.pdf.

Smith:1904:USU

[1314] Hugh M. Smith. The utilization of seaweeds in the United States. Fishery Bulletin, 24(1):??, 1904. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
utilization-seaweeds-united-states;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb24.7.pdf.

Greene:1904:PSC

[1315] Charles Wilson Greene. Physiological studies of the Chinook salmon. Fishery Bulletin, 24(1):??, 1904. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/physiological-studies-chinook-salmon;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb24.15.pdf.

Anonymous:1905:EI

[1316] Anonymous. Explanation of illustrations. Fishery Bulletin, 25(1):??, 1905. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/explanation-
illustrations-16;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb25%20plates.pdf.

Sumner:1905:PEU

[1317] Francis B. Sumner. The physiological effects upon fishes of changes in the density and salinity of water. Fishery Bulletin, 25(1):??, 1905. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/physiological-
effects-upon-fishes-changes-density-and-salinity-water;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb25.2.pdf.

MacFarland:1905:OMM

[1318] F. M. MacFarland. Opisthobranchiate Mollusca from Monterey Bay, California, and vicinity. Fishery Bulletin, 25(1):111–151 + 20 plates, 1905. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
opisthobranchiate-mollusca-monterey-bay-california-and-vicinity;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb25.3.pdf.

Moore:1905:HOC

[1319] J. Percy Moore. Hirudinea and Oligochaeta collected in the Great Lakes region. Fishery Bulletin, 25(1):??, 1905. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/hirudinea-and-oligochaeta-collected-great-

lakes-region;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb25.4.pdf.

Jordan:1905:FSD

[1320] David Starr Jordan and Alvin Seale. The fishes of Samoa: Description of the species found in the archipelago, with a provisional check list of the fishes of Oceania. Fishery Bulletin, 25(1):??, 1905. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/fishes-samoa-description-species-
found-archipelago-provisional-check-list-fishes-oceania;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb25.5.pdf.

Anonymous:1905:I

[1321] Anonymous. Index. Fishery Bulletin, 25(1):??, 1905. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-65;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb25.index_.pdf.

Evermann:1906:FPI

[1322] Barton Warren Evermann and Alvin Seale. Fishes of the Philippine Islands. Fishery Bulletin, 26(1):??, 1906. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/fishes-philippine-islands;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb26.2.pdf.

Anonymous:1906:I

[1323] Anonymous. Index. Fishery Bulletin, 26(1):??, 1906. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-66;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb26.index_.pdf.

Linton:1906:CPF

[1324] Edwin Linton. A cestode parasite in the flesh of the butterfish. Fishery Bulletin, 26(1):??, 1906. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
cestode-parasite-flesh-butterfish;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb26.3.pdf.

Juday:1906:NLT

[1325] Chancey Juday. Notes on Lake Tahoe, its trout and trout-fishing. Fishery Bulletin, 26(1):??, 1906. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/notes-lake-tahoe-its-trout-and-trout-fishing;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb26.4.pdf.

Juday:1906:STL

[1326] Chancey Juday. A study of Twin Lakes, Colorado, with special consideration of the food of the trouts. Fishery Bulletin, 26(1):??, 1906. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/study-twin-
lakes-colorado-special-consideration-food-trouts;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb26.5.pdf.

Pilsbry:1906:HC

[1327] Henry A. Pilsbry. Hawaiian Cirripedia. Fishery Bulletin, 26(1):??, 1906. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/hawaiian-cirripedia;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb26.6.pdf.

Pilsbry:1906:CPC

[1328] Henry A. Pilsbry. Cirripedia from the Pacific coast of North America. Fishery Bulletin, 26(1):??, 1906. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/cirripedia-pacific-coast-north-america;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb26.
7.pdf.

Anonymous:1906:EI

[1329] Anonymous. Explanation of illustrations. Fishery Bulletin, 26(1):??, 1906. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/explanation-
illustrations-17;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb26%20plates.pdf.

Jordan:1906:NFH

[1330] David Starr Jordan and John Otterbein Snyder. Notes on fishes of Hawaii, with descriptions of new species. Fishery Bulletin, 26 (1):??, 1906. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/notes-
fishes-hawaii-descriptions-new-species;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb26.8.pdf.

Jordan:1906:FIL

[1331] David Starr Jordan and Alvin Seale. Fishes of the islands of luzon and panay. Fishery Bulletin, 26(1):??, 1906. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/fishes-islands-luzon-and-panay;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb26.1.pdf.

Evermann:1906:FA

[1332] Barton Warren Evermann and Edmund Lee Goldsborough. The fishes of Alaska. Fishery Bulletin, 26(1):??, 1906. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/evermann-barton-warren-and-edmund-lee-goldsborough;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb26.9.pdf.

Anonymous:1907:EI

[1333] Anonymous. Explanation of illustrations. Fishery Bulletin, 27(1):??, 1907. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/explanation-
illustrations-18;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb27%20plates.pdf.

Jordan:1907:FIP

[1334] David Starr Jordan and Robert Earl Richardson. Fishes from islands of the Philippine Archipelago. Fishery Bulletin, 27(1):??, 1907. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishes-
islands-philippine-archipelago;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb27.8.pdf.

Sullivan:1907:PDT

[1335] Michael X. Sullivan. The physiology of the digestive tract of elasmobranchs. Fishery Bulletin, 27(1):??, 1907. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/physiology-digestive-tract-elasmobranchs;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb27.1.pdf.

Anonymous:1907:I

[1336] Anonymous. Index. Fishery Bulletin, 27(1):??, 1907. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-67;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb27.index_.pdf.

Goldsborough:1907:FWV

[1337] Edmund Lee Goldsborough and H. Walton Clark. Fishes of West Virginia. Fishery Bulletin, 27(1):??, 1907. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/fishes-west-virginia;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb27.2.pdf.

Reighard:1907:PAA

[1338] Jacob Reighard. The photography of aquatic animals in their natural environment. Fishery Bulletin, 27(1):??, 1907. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/photography-aquatic-
animals-their-natural-environment;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb27.3.pdf.

Snyder:1907:RFF

[1339] John Otterbein Snyder. Relationships of the fish fauna of the lakes of southeastern Oregon. Fishery Bulletin, 27(1):??, 1907. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/relationships-fish-fauna-lakes-
southeastern-oregon;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb27.4.pdf.

Rutter:1907:FSS

[1340] Cloudsley Rutter. The fishes of the Sacramento–San Joaquin basin, with a study of their distribution and variation. Fishery Bulletin, 27(1):??, 1907. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/fishes-sacramento-san-joaquin-basin-study-
their-distribution-and-variation;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb27.5.pdf.

Snyder:1907:FCS

[1341] John Otterbein Snyder. The fishes of the coastal streams of Oregon and northern California. Fishery Bulletin, 27(1):??, 1907. CODEN FSYBAY.

ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/fishes-coastal-streams-oregon-and-northern-
california;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb27.6.pdf.

Sigerfoos:1907:NHO

[1342] Charles P. Sigerfoos. Natural history, organization, and late development of the teredinidae, or shipworms. Fishery Bulletin, 27(1): ??, 1907. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/natural-
history-organization-and-late-development-teredinidae-or-shipworms;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb27.7.pdf.

Olsen:1908:IRF

[1343] O. T. Olsen. International regulations of the fisheries on the high seas. Fishery Bulletin, 28(1):79–90, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/international-regulations-fisheries-high-seas;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb28(1).2.pdf.

Mead:1908:MLC

[1344] A. D. Mead. A method of lobster culture. Fishery Bulletin, 28(1): ??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/method-
lobster-culture;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb28(1).10.pdf.

Rossati:1908:ECF

[1345] Guido Rossati. Economic conditions of the fisheries in Italy. Fishery Bulletin, 28(1):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/economic-conditions-fisheries-italy;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb28(
1).18.pdf.

Lefevre:1908:EAP

[1346] George Lefevre and W. C. Curtis. Experiments in the artificial propagation of fresh-water mussels. Fishery Bulletin, 28(1):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/experiments-
artificial-propagation-fresh-water-mussels;   https://spo.nmfs.

noaa.gov/sites/default/files/pdf-content/fish-bull/fb28(1).
27.pdf.

Fryer:1908:IRF

[1347] Charles Edward Fryer. International regulations of the fisheries on the high seas. Fishery Bulletin, 28(1):93–102, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/international-regulations-fisheries-high-seas-
0;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb28(1).3.pdf.

Field:1908:SMD

[1348] Irving A. Field. Sea mussels and dogfish as food. Fishery Bulletin, 28(1): ??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/sea-mussels-
and-dogfish-food;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb28(1).11.pdf.

Coker:1908:FGI

[1349] Robert E. Coker. The fisheries and the guano industry of Peru. Fishery Bulletin, 28(1):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fisheries-and-guano-industry-peru;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb28(1).19.pdf.

Downing:1908:PPW

[1350] S. W. Downing. A plan for promoting the whitefish production of the Great Lakes. Fishery Bulletin, 28(1):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/plan-promoting-whitefish-production-great-
lakes;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb28(1).28.pdf.

Stevenson:1908:IRF

[1351] Charles Hugh Stevenson. International regulations of the fisheries on the high seas. Fishery Bulletin, 28(1):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/international-regulations-fisheries-high-seas-
1;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb28(1).4.pdf.

Rowe:1908:WOF

[1352] Henry C. Rowe. The wholesomeness of oysters as food. Fishery Bulletin, 28(1):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/wholesomeness-oysters-food;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb28(1).
12.pdf.

Yen:1908:FC

[1353] Wei-Ching W. Yen. The fisheries of China. Fishery Bulletin, 28 (1):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fisheries-
china;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb28(1).20.pdf.

Clark:1908:PPW

[1354] Frank N. Clark. A plan for promoting the whitefish production of the Great Lakes. Fishery Bulletin, 28(1):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/plan-promoting-whitefish-production-great-
lakes-0;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb28(1).29.pdf.

Jordan:1908:WIF

[1355] David Starr Jordan. Work of the International Fisheries Commission of Great Britain and the United States. Fishery Bulletin, 28(1):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/work-international-
fisheries-commission-great-britain-and-united-states;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb28(1).5.pdf.

Hathaway:1908:EMF

[1356] Walter E. Hathaway. Effects of menhaden fishing upon the supply of menhaden and of the fishes that prey upon them. Fishery Bulletin, 28 (1):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/effects-
menhaden-fishing-upon-supply-menhaden-and-fishes-prey-upon-
them;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb28(1).13.pdf.

Kitahara:1908:FJC

[1357] T. Kitahara. The fisheries of Japan considered from a geographical standpoint. Fishery Bulletin, 28(1):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fisheries-japan-considered-geographical-standpoint;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb28(1).21.pdf.

Reighard:1908:PPW

[1358] Paul Reighard. A plan for promoting the whitefish production of the Great Lakes. Fishery Bulletin, 28(1):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/plan-promoting-whitefish-production-great-
lakes-1;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb28(1).30.pdf.

Dennis:1908:SRF

[1359] Oregon Milton Dennis. Some reasons for failure of fish protective legislation and some suggested remedies. Fishery Bulletin, 28(1):189–192, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/some-reasons-failure-
fish-protective-legislation-and-some-suggested-remedies;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb28(1).6.pdf.

Kendall:1908:EMM

[1360] W. C. Kendall. Effects of the menhaden and mackerel fisheries upon the fish supply. Fishery Bulletin, 28(1):??, 1908. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/effects-menhaden-and-
mackerel-fisheries-upon-fish-supply;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb28(1).14.pdf.

Matsubara:1908:GTC

[1361] Shinnosuke Matsubara. Goldfish and their culture in Japan. Fishery Bulletin, 28(1):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/goldfish-and-their-culture-japan;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb28(1).22.pdf.

Perce:1908:NAA

[1362] H. Wheeler Perce. National aspects of angling and the protection of game fishes. Fishery Bulletin, 28(1):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/national-aspects-angling-
and-protection-game-fishes;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb28(1).7.pdf.

Solling:1908:IPM

[1363] A. Sölling. An improved and practical method of packing fish for transportation. Fishery Bulletin, 28(1):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/improved-and-practical-method-packing-fish-
transportation;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb28(1).15.pdf.

Moore:1908:CSS

[1364] H. F. Moore. Commercial sponges and the sponge fisheries. Fishery Bulletin, 28(1):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/commercial-sponges-and-sponge-fisheries;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb28(1).23.pdf.

Anonymous:1908:EI

[1365] Anonymous. Explanation of illustrations. Fishery Bulletin, 28(1):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/explanation-
illustrations-19;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb28(1)%20plates.pdf.

Holder:1908:SFC

[1366] Charles F. Holder. Sport fishing in California and Florida. Fishery Bulletin, 28(1):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/sport-fishing-california-and-florida;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb28(1).8.pdf.

Solomon:1908:PPP

[1367] John I. Solomon. A process for preserving the pearl-oyster fisheries and for increasing the value of the yield of pearls. Fishery Bulletin, 28(1):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/process-preserving-pearl-oyster-fisheries-
and-increasing-value-yield-pearls;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb28(1).16.pdf.

Moore:1908:PMS

[1368] H. F. Moore. A practical method of sponge culture. Fishery Bulletin, 28(1):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/practical-method-sponge-culture;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb28(
1).25.pdf.

Anonymous:1908:OSB

[1369] Anonymous. Organization and sessional business. Fishery Bulletin, 28 (1):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/organization-
and-sessional-business;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb28(1).1.pdf.

Field:1908:LLP

[1370] George W. Field. Lobsters and the lobster problem in Massachusetts. Fishery Bulletin, 28(1):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/lobsters-and-lobster-problem-massachusetts;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb28(1).9.pdf.

Townsend:1908:SSF

[1371] Charles H. Townsend. Fur seals and the seal fisheries. Fishery Bulletin, 28 (1):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/fur-seals-and-
seal-fisheries;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb28(1).17.pdf.

Cotte:1908:SC

[1372] Jules Cotte. Sponge culture. Fishery Bulletin, 28(1):??, 1908. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/sponge-culture;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb28(1).26.pdf.

Paige:1908:MCR

[1373] Charles L. Paige. A method of cultivating rainbow trout and other salmonidae. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/method-cultivating-rainbow-trout-and-
other-salmonidae;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb28(2).3.pdf.

Worth:1908:FWS

[1374] S. G. Worth. Fresh-water shrimp, a natural fish food. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fresh-water-shrimp-natural-fish-food;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb28(2).11.pdf.

Zirzow:1908:NMC

[1375] Paul Zirzow. A new method of combating fungus on fishes in captivity. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/zirzow-paul;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb28(2).19.pdf.

Robinson:1908:DCY

[1376] Robert K. Robinson. A device for counting young fish. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/device-counting-young-fish;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb28(2).
27.pdf.

Gudger:1908:HLH

[1377] E. W. Gudger. The habits and live history of the toadfish (Opsanus tau). Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/habits-and-live-history-toadfish-opsanus-tau;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb28(2).35.pdf.

Wilson:1908:DST

[1378] H. V. Wilson. Development of sponges from tissue cells outside the body of the parent. Fishery Bulletin, 28(2):??, 1908. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/development-sponges-tissue-cells-
outside-body-parent;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb28(2).43.pdf.

Smith:1908:USB

[1379] Hugh M. Smith. The United States Bureau of Fisheries: Its establishment, functions, organization, resources, operations, and achievements. Fishery Bulletin, 28(2):??, 1908. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/united-states-bureau-fisheries-
its-establishment-functions-organization-resources;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb28(2).51.pdf.

Worth:1908:PES

[1380] S. G. Worth. Possible expansion of shad-hatchery work. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/possible-expansion-shad-hatchery-work;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb28(2).4.pdf.

Anthony:1908:CT

[1381] R. Anthony. The cultivation of the turbot. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/cultivation-turbot;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb28(2).12.pdf.

Rosenberg:1908:EME

[1382] Albert Rosenberg. Eighteen months experience with a disease of brook trout. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/eighteen-months-experience-disease-brook-trout;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb28(2).20.pdf.

Holder:1908:MTL

[1383] Charles F. Holder. A method of transporting live fishes. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/method-transporting-live-fishes;   https://spo.

nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb28(
2).28.pdf.

Reighard:1908:MSH

[1384] Jacob Reighard. Methods of studying the habits of fishes and recording their life histories; with an account of the breeding habits of the horned dace. Fishery Bulletin, 28(2):??, 1908. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/methods-studying-habits-fishes-
and-recording-their-life-histories-account-breeding-habits;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb28(2).36.pdf.

Birge:1908:GDW

[1385] E. A. Birge. Gases dissolved in the waters of Wisconsin lakes. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
gases-dissolved-waters-wisconsin-lakes;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb28(2).
44.pdf.

Anonymous:1908:I

[1386] Anonymous. Index. Fishery Bulletin, 28(2):??, 1908. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-68;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb28(2).index_.pdf.

Paige:1908:CVF

[1387] Charles L. Paige. The comparative value of food for rainbow trout and other salmonoids. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/comparative-value-food-rainbow-trout-
and-other-salmonoids;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb28(2).5.pdf.

Marsh:1908:TFC

[1388] M. C. Marsh and R. K. Robinson. The treatment of fish-cultural waters for the removal of algae. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/treatment-fish-cultural-waters-
removal-algae;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb28(2).13.pdf.

Besana:1908:AFI

[1389] Giuseppe Besana. American fishes in Italy. Fishery Bulletin, 28(2): ??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/american-
fishes-italy;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb28(2).21.pdf.

vonBayer:1908:MMF

[1390] H. von Bayer. A method of measuring fish eggs. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/method-measuring-fish-eggs;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb28(2).
29.pdf.

Holder:1908:MOH

[1391] Charles F. Holder. A method of observing the habits and recording the life histories of fishes. Fishery Bulletin, 28(2):??, 1908. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/method-observing-habits-and-recording-
life-histories-fishes;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb28(2).37.pdf.

Moore:1908:VSF

[1392] H. F. Moore. Volumetric studies of the food and feeding of oysters. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
volumetric-studies-food-and-feeding-oysters;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb28(
2).45.pdf.

Dannevig:1908:AME

[1393] G. M. Dannevig. Apparatus and methods employed at the marine fish hatchery at Flödevig, Norway. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/apparatus-and-
methods-employed-marine-fish-hatchery-fl%C3%B6devig-norway;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb28(2).6.pdf.

Marsh:1908:NDC

[1394] M. C. Marsh. Note on the dissolved content of water in its effect upon fishes. Fishery Bulletin, 28(2):??, 1908. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/note-dissolved-content-water-its-
effect-upon-fishes;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb28(2).14.pdf.

Tulian:1908:AAF

[1395] E. A. Tulian. Acclimatization of American fishes in Argentina. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/acclimatization-american-fishes-argentina;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb28(2).22.pdf.

Simms:1908:IHR

[1396] G. E. Simms. An improvement in hatching and rearing boxes; with notes on continuous feeding of the fry of salmonidae. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/improvement-hatching-
and-rearing-boxes-notes-continuous-feeding-fry-salmonidae;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb28(2).30.pdf.

Scott:1908:ECD

[1397] George G. Scott. Effects of changes in the density of water upon the blood of fishes. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/effects-changes-density-water-upon-blood-fishes;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb28(2).38.pdf.

Miner:1908:PEE

[1398] Roy W. Miner. A plan for an educational exhibit of fishes. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
plan-educational-exhibit-fishes-0;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb28(2).47.pdf.

Holder:1908:PEE

[1399] Charles F. Holder. A plan for an educational exhibit of fishes. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/plan-educational-exhibit-fishes;   https://spo.

nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb28(
2).46.pdf.

Dannevig:1908:USF

[1400] G. M. Dannevig. The utility of sea-fish hatching. Fishery Bulletin, 28(2): ??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/utility-sea-
fish-hatching;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb28(2).7.pdf.

Vincent:1908:CDY

[1401] Eugene Vincent. Causes of disease in young salmonoids. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/causes-disease-young-salmonoids;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb28(
2).15.pdf.

Ayson:1908:IAF

[1402] L. F. Ayson. The introduction of American fishes into New Zealand. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/introduction-american-fishes-new-zealand;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb28(2).23.pdf.

Vincent:1908:DUF

[1403] Eugene Vincent. Devices for use in fish hatcheries and aquaria. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/devices-use-fish-hatcheries-and-aquaria;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb28(2).31.pdf.

Ward:1908:IPS

[1404] Henry B. Ward. Internal parasites of the sebago salmon. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/internal-
parasites-sebago-salmon;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb28(2).39.pdf.

Hoek:1908:PPR

[1405] P. P. C. Hoek. Propagation and protection of the Rhine salmon. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/propagation-and-protection-rhine-salmon;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb28(2).8.pdf.

Franke:1908:RPC

[1406] Johann Franke. Radical prevention of costia necatrix in salmonoid fry. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/radical-prevention-costia-necatrix-salmonoid-fry;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb28(2).16.pdf.

vonPirko:1908:NAF

[1407] Franz von Pirko. Naturalization of American fishes in Austrian waters. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/naturalization-american-fishes-austrian-waters;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb28(2).24.pdf.

Kincaid:1908:NMT

[1408] W. S. Kincaid. New methods of transporting eggs and fish. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
new-methods-transporting-eggs-and-fish;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb28(2).
32.pdf.

Linton:1908:NFP

[1409] Edwin Linton. Notes on the flesh parasites of marine food fishes. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/notes-flesh-parasites-marine-food-fishes;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb28(2).40.pdf.

Lucas:1908:OEE

[1410] F. A. Lucas. Outline for an educational exhibit of fishes. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/outline-
educational-exhibit-fishes;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb28(2).48.pdf.

Titcomb:1908:FCP

[1411] John W. Titcomb. Fish-cultural practices in the Bureau of Fisheries. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fish-cultural-practices-bureau-fisheries;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb28(2).1.pdf.

Seal:1908:FTR

[1412] William P. Seal. Fishes in their relation to the mosquito problem. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishes-their-relation-mosquito-problem;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb28(2).
9.pdf.

Spencer:1908:TFF

[1413] L. B. Spencer. Treatment of fungus on fishes in captivity. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
treatment-fungus-fishes-captivity;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb28(2).17.pdf.

Franke:1908:CDA

[1414] Johann Franke. Causes of degeneration of American trouts in Austria. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/causes-degeneration-american-trouts-austria;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb28(2).25.pdf.

vonBayer:1908:F

[1415] H. von Bayer. Fishways. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/fishways.

Parker:1908:SFE

[1416] G. H. Parker. Structure and functions of the ear of the squeteague. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
structure-and-functions-ear-squeteague;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb28(2).
41.pdf.

Franklin:1908:MPF

[1417] Dwight Franklin. A method of preparing fish for museum and exhibition purposes. Fishery Bulletin, 28(2):??, 1908. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/method-preparing-fish-museum-and-
exhibition-purposes;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb28(2).49.pdf.

Mead:1908:NPA

[1418] A. D. Mead. New principle of aquiculture and transportation of live fishes. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/new-principle-aquiculture-and-transportation-live-
fishes;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb28(2).2.pdf.

Atkins:1908:FYS

[1419] Charles G. Atkins. Foods for young salmonoid fishes. Fishery Bulletin, 28 (2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/foods-young-
salmonoid-fishes;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb28(2).10.pdf.

Holder:1908:MCF

[1420] Charles F. Holder. Methods of combating fungus on fishes in captivity. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/methods-combating-fungus-fishes-captivity;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb28(2).18.pdf.

Fuller:1908:NID

[1421] Alfred E. Fuller. New and improved devices for fish culturists. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/new-and-improved-devices-fish-culturists;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb28(2).26.pdf.

Gill:1908:POH

[1422] Theodore Gill. A plea for observation of the habits of fishes and against undue generalization. Fishery Bulletin, 28(2):??, 1908. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/plea-observation-habits-fishes-
and-against-undue-generalization;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb28(2).34.pdf.

Sumner:1908:ISF

[1423] Francis B. Sumner. An intensive study of the fauna and flora of a restricted area of sea bottom. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic).  URL https://spo.nmfs.noaa.gov/content/intensive-
study-fauna-and-flora-restricted-area-sea-bottom;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb28(2).42.pdf.

Figgins:1908:NMP

[1424] J. D. Figgins. New methods of preparing fishes for museum exhibit. Fishery Bulletin, 28(2):??, 1908. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
new-methods-preparing-fishes-museum-exhibit;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb28(
2).50.pdf.

Jordan:1909:RSF

[1425] David Starr Jordan and Barton Warren Evermann. A review of the salmonoid fishes of the Great Lakes, with notes on the whitefishes of other regions. Fishery Bulletin, 29(1):??, 1909. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/review-salmonoid-fishes-great-lakes-notes-
whitefishes-other-regions;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb29.1.pdf.

Parker:1909:IEE

[1426] G. H. Parker. Influence of the eyes, ears, and other allied sense organs on the movements of the dogfish, Mustelus canis (Mitchill). Fishery Bulletin, 29(1):??, 1909. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/influence-eyes-ears-and-other-allied-
sense-organs-movements-dogfish-mustelus-canis-mitchill;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb29.2.pdf.

Pilsbry:1909:BJB

[1427] Henry A. Pilsbry. Barnacles of Japan and Bering Sea. Fishery Bulletin, 29(1):??, 1909. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/barnacles-japan-and-bering-sea;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb29.3.pdf.

Field:1909:FVS

[1428] Irving A. Field. The food value of sea mussels. Fishery Bulletin, 29(1): ??, 1909. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/food-value-
sea-mussels;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb29.4.pdf.

Greene:1909:MSC

[1429] Charles W. Greene. The migration of salmon in the Columbia River. Fishery Bulletin, 29(1):??, 1909. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/migration-salmon-columbia-river;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb29.5.pdf.

Herrick:1909:NHA

[1430] Francis Hobart Herrick. Natural history of the American lobster. Fishery Bulletin, 29(1):??, 1909. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/natural-history-american-lobster;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb29.6.pdf.

Grave:1909:APW

[1431] Benjamin H. Grave. Anatomy and physiology of the wing-shell atrina rigida. Fishery Bulletin, 29(1):??, 1909. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/anatomy-and-physiology-wing-shell-atrina-rigida;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb29.7.pdf.

Anonymous:1909:EI

[1432] Anonymous. Explanation of illustrations. Fishery Bulletin, 29(1):??, 1909. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/explanation-

illustrations-20;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb29.plates.pdf.

Anonymous:1909:I

[1433] Anonymous. Index. Fishery Bulletin, 29(1):??, 1909. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-69;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb29.index_.pdf.

Parker:1911:SDI

[1434] G. H. Parker. Sound as a directing influence in the movements of fishes. Fishery Bulletin, 30(1):??, 1911. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/sound-directing-influence-movements-fishes;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb30.3.pdf.

Lefevre:1911:SRA

[1435] George Lefevre and Winterton C. Curtis. Studies on the reproduction and artificial propagation of fresh-water mussels. Fishery Bulletin, 30(1):109–201 + 12 plates, 1911. CODEN FSY-

BAY.  ISSN 0090-0656 (print), 1937-4518 (electronic).  URL

https://spo.nmfs.noaa.gov/content/studies-reproduction-and-
artificial-propagation-fresh-water-mussels;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb30.4.pdf.

Osburn:1911:BWH

[1436] Raymond C. Osburn. The Bryozoa of the Woods Hole region. Fishery Bulletin, 30(1):??, 1911. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/bryozoa-woods-hole-region;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb30.5.pdf.

Berry:1911:RCW

[1437] S. Stillman Berry. A review of the cephalopods of western North America. Fishery Bulletin, 30(1):??, 1911. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/review-cephalopods-western-north-america;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb30.6.pdf.

Fraser:1911:SHB

[1438] C. McLean Fraser. Some hydroids of Beaufort, North Carolina. Fishery Bulletin, 30(1):??, 1911. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/some-hydroids-beaufort-north-carolina;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb30.7.pdf.

Kendall:1911:NNS

[1439] William C. Kendall. Notes on a new species of flatfish from off the coast of New England. Fishery Bulletin, 30(1):??, 1911. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/notes-new-species-flatfish-coast-new-
england;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb30.8.pdf.

Wilson:1911:DSD

[1440] H. V. Wilson. Development of sponges from dissociated tissue cells. Fishery Bulletin, 30(1):??, 1911. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/development-sponges-dissociated-tissue-cells;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb30.1.pdf.

Anonymous:1911:I

[1441] Anonymous. Index. Fishery Bulletin, 30(1):??, 1911. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-70;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb30.index_.pdf.

Gilbert:1911:FBS

[1442] Charles Henry Gilbert and Charles Victor Burke. Fishes from Bering Sea and Kamchatka. Fishery Bulletin, 30(1):??, 1911. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/fishes-bering-sea-and-kamchatka;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb30.2.pdf.

Anonymous:1911:CMFa

[1443] Anonymous. A catalogue of the marine fauna of Woods Hole and vicinity. Fishery Bulletin, 31(2):??, 1911. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/catalogue-marine-fauna-woods-hole-and-vicinity;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb31(2).1.pdf.

Anonymous:1911:CMFb

[1444] Anonymous. A catalogue of the marine flora of Woods Hole and vicinity. Fishery Bulletin, 31(2):??, 1911. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/catalogue-marine-flora-woods-hole-and-vicinity;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb31(2).2.pdf.

Anonymous:1911:ISN

[1445] Anonymous. Index to systematic names. Fishery Bulletin, 31(2):??, 1911. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/index-systematic-names;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb31(2).3.pdf.

Radcliffe:1912:DND

[1446] Lewis Radcliffe and William W. Welsh. Description of a new darter from Maryland. Fishery Bulletin, 32(1):??, 1912. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/description-new-darter-maryland;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb32.3.pdf.

Mitchell:1912:EWG

[1447] Philip H. Mitchell. The effect of water-gas tar on oysters. Fishery Bulletin, 32(1):??, 1912. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
effect-water-gas-tar-oysters;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb32.11.pdf.

Parker:1912:SSF

[1448] G. H. Parker and R. E. Sheldon. The sense of smell in fishes. Fishery Bulletin, 32(1):??, 1912. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/sense-
smell-fishes;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb32.4.pdf.

Mitchell:1912:ORS

[1449] Philip H. Mitchell. The oxygen requirements of shellfish. Fishery Bulletin, 32(1):??, 1912. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
oxygen-requirements-shellfish;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb32.12.pdf.

Snyder:1912:FST

[1450] John Otterbein Snyder. The fishes of the streams tributary to Monterey Bay, California. Fishery Bulletin, 32(1):??, 1912. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/fishes-streams-tributary-
monterey-bay-california;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb32.5.pdf.

Hargitt:1912:AWH

[1451] Charles W. Hargitt. The Anthozoa of the Woods Hole region. Fishery Bulletin, 32(1):??, 1912. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/anthozoa-woods-hole-region;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb32.13.
pdf.

Greene:1912:AHA

[1452] Charles W. Greene. Anatomy and histology of the alimentary tract of the king salmon. Fishery Bulletin, 32(1):??, 1912. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/anatomy-and-histology-alimentary-
tract-king-salmon;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb32.6.pdf.

Berry:1912:CHI

[1453] S. Stillman Berry. The cephalopoda of the Hawaiian Islands. Fishery Bulletin, 32(1):??, 1912. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
cephalopoda-hawaiian-islands;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb32.14.pdf.

Surber:1912:NNH

[1454] Thaddeus Surber. Notes on the natural hosts of fresh-water mussels. Fishery Bulletin, 32(1):??, 1912. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/notes-natural-hosts-fresh-water-mussels;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb32.7.pdf.

Gaylord:1912:CTS

[1455] Harvey R. Gaylord and Millard C. Marsh. Carcinoma of the thyroid in the salmonoid fishes. Fishery Bulletin, 32(1):??, 1912. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/carcinoma-thyroid-salmonoid-fishes;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb32.15.pdf.

Cockerell:1912:OFS

[1456] T. D. A. Cockerell. Observations on fish scales. Fishery Bulletin, 32(1): ??, 1912. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/observations-
fish-scales;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb32.8.pdf.

Birge:1912:LSF

[1457] Edward A. Birge and Chancey Juday. A limnological study of the Finger Lakes of New York. Fishery Bulletin, 32(1):??, 1912. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/limnological-study-finger-lakes-
new-york;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb32.16.pdf.

Gilbert:1912:AMP

[1458] Charles H. Gilbert. Age at maturity of the Pacific coast salmon of the genus Oncorhynchus. Fishery Bulletin, 32(1):??, 1912. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/age-maturity-pacific-
coast-salmon-genus-oncorhynchus;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb32.1.pdf.

White:1912:NMD

[1459] George F. White and Adrian Thomas. A new method for the determination of the food value of proteins, with application to Cynoscion regalis. Fishery Bulletin, 32(1):??, 1912. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/new-method-determination-
food-value-proteins-application-cynoscion-regalis;   https:

//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb32.9.pdf.

Anonymous:1912:I

[1460] Anonymous. Index. Fishery Bulletin, 32(1):??, 1912. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-71;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb32.index_.pdf.

Snyder:1912:NST

[1461] John Otterbein Snyder. A new species of trout from Lake Tahoe. Fishery Bulletin, 32(1):??, 1912. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
new-species-trout-lake-tahoe;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb32.2.pdf.

White:1912:PFV

[1462] George F. White and Adrian Thomas. Propertires of fish and vegetable oil mixtures. Fishery Bulletin, 32(1):??, 1912. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/propertires-fish-and-vegetable-
oil-mixtures;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb32.10.pdf.

Kuntz:1913:ELD

[1463] Albert Kuntz. The embryology and larval development of bairdiella chrysura and anchovia mitchilli. Fishery Bulletin, 33(1):??, 1913. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/embryology-and-larval-development-
bairdiella-chrysura-and-anchovia-mitchilli;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb33.1.pdf.

Reighard:1913:ERS

[1464] Jacob Reighard. An ecological reconnoissance [sic] of the fishes of Douglas Lake, Cheboygan County, Michigan, in midsummer. Fishery Bulletin, 33(1):219–249, 1913. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
ecological-reconnoissance-fishes-douglas-lake-cheboygan-county-
michigan-midsummer;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb33.9.pdf.

Greene:1913:SMK

[1465] Charles Wilson Greene and Carl Hartley Greene. The skeletal musculature of the king salmon. Fishery Bulletin, 33(1):??, 1913. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/skeletal-musculature-king-salmon;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb33.2.pdf.

Moore:1913:PRP

[1466] Emmeline Moore. The potamogetons in relation to pound culture. Fishery Bulletin, 33(1):??, 1913. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/potamogetons-relation-pound-culture;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb33.
10.pdf.

Parker:1913:DIS

[1467] G. H. Parker. The directive influence of the sense of smell in the dogfish. Fishery Bulletin, 33(1):??, 1913. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/directive-influence-sense-smell-dogfish;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb33.3.pdf.

Anonymous:1913:I

[1468] Anonymous. Index. Fishery Bulletin, 33(1):??, 1913. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-72;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb33.index_.pdf.

Greene:1913:SFM

[1469] Charles W. Greene. The storage of fat in the muscular tissue of the king salmon and its resorption during the fast of the spawning migration. Fishery Bulletin, 33(1):??, 1913. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/storage-fat-muscular-
tissue-king-salmon-and-its-resorption-during-fast-spawning-
migration;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb33.4.pdf.

Crozier:1913:CWL

[1470] William J. Crozier and Selig Hecht. Correlations of weight, length, and other body measurements in the weakfish, Cynoscion regalis. Fishery Bulletin, 33(1):??, 1913. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/correlations-weight-length-and-other-body-
measurements-weakfish-cynoscion-regalis;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb33.5.pdf.

Greene:1913:FAF

[1471] Charles W. Greene. The fat-absorbing function of the alimentary tract of the king salmon. Fishery Bulletin, 33(1):??, 1913. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/fat-absorbing-function-alimentary-tract-
king-salmon;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb33.6.pdf.

Kuntz:1913:NHM

[1472] Albert Kuntz. Notes on the habits, morphology of the reproductive organs, and embryology of the viviparous fish Gambusia affinis. Fishery Bulletin, 33(1):??, 1913. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
notes-habits-morphology-reproductive-organs-and-embryology-
viviparous-fish-gambusia-affinis;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb33.7.pdf.

Hahn:1913:SPC

[1473] C. W. Hahn. Sporozoön parasites of certain fishes in the vicinity of Woods Hole, Massachusetts. Fishery Bulletin, 33(1):??, 1913. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/sporozo%C3%B6n-parasites-
certain-fishes-vicinity-woods-hole-massachusetts;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb33.8.pdf.

Radcliffe:1914:SRB

[1474] Lewis Radcliffe. The sharks and rays of Beaufort, North Carolina. Fishery Bulletin, 34(1):??, 1914. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/sharks-and-rays-beaufort-north-carolina;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb34.3.pdf.

Taylor:1914:SGS

[1475] Harden F. Taylor. The structure and growth of the scales of the squeteague and the pigfish as indicative of life history. Fishery Bulletin, 34(1):??, 1914. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/structure-and-growth-scales-squeteague-and-pigfish-
indicative-life-history;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb34.4.pdf.

Wilson:1914:CPF

[1476] Charles Branch Wilson. Copepod parasites of fresh-water fishes and their economic relations to mussel glochida. Fishery Bulletin, 34(1): ??, 1914. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/copepod-
parasites-fresh-water-fishes-and-their-economic-relations-mussel-
glochida;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb34.5.pdf.

Snyder:1914:FST

[1477] John Otterbein Snyder. The fishes of the streams tributary to Tomales Bay, California. Fishery Bulletin, 34(1):??, 1914. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/fishes-streams-tributary-tomales-bay-
california;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb34.6.pdf.

Coker:1914:FFB

[1478] Robert E. Coker. The Fairport Fisheries Biological Station: Its equipment, organization, and functions. Fishery Bulletin, 34(1):??, 1914. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/fairport-fisheries-
biological-station-its-equipment-organization-and-functions;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb34.7.pdf.

Kuntz:1914:NEL

[1479] Albert Kuntz. Notes on the embryology and larval development of five species of teleostean fishes. Fishery Bulletin, 34(1):409–429, 1914. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/notes-embryology-and-
larval-development-five-species-teleostean-fishes;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb34.8.pdf.

Osgood:1914:SOL

[1480] Wilfred H. Osgood, Edward A. Preble, and George H. Parker. The fur seals and other life of the Pribilof Islands, Alaska, in 1914. Fishery Bulletin, 34(1):1–214, 1914. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/fur-seals-and-other-life-pribilof-islands-
alaska-1914;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb34.1.pdf.

Evermann:1914:NFE

[1481] Barton Warren Evermann and Samuel F. Hildebrand. Notes on the fishes of east Tennessee. Fishery Bulletin, 34(1):??, 1914. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/notes-fishes-east-tennessee;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb34.9.pdf.

Mast:1914:CSC

[1482] S. O. Mast. Changes in shade, color, and pattern in fishes, and their bearing on the problems of adaptation and behavior, with special reference to the flounders paralichtyhys and ancylopsetta. Fishery Bulletin, 34(1):??, 1914. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/changes-shade-color-and-pattern-
fishes-and-their-bearing-problems-adaptation-and-behavior;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb34.2.pdf.

Anonymous:1914:I

[1483] Anonymous. Index. Fishery Bulletin, 34(1):??, 1914. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-73;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb34.index_.pdf.

Mitchell:1915:OVG

[1484] Philip H. Mitchell and Raymond L. Barney. The occurrence in Virginia of green-gilled oysters similar to those of Marennes. Fishery Bulletin, 35(1):??, 1915–1916. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
occurrence-virginia-green-gilled-oysters-similar-those-marennes;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb35.4.pdf.

Kendall:1915:RLM

[1485] William Converse Kendall. The Rangeley lakes, Maine; with special reference to the habits of the fishes, fish culture, and angling. Fishery Bulletin, 35(1):??, 1915–1916. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
rangeley-lakes-maine-special-reference-habits-fishes-fish-culture-
and-angling;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb35.12.pdf.

Mitchell:1915:NOG

[1486] Philip H. Mitchell. Nutrition of oysters: Glycogen formation and storage. Fishery Bulletin, 35(1):??, 1915–1916. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/nutrition-oysters-glycogen-formation-
and-storage;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb35.5.pdf.

Anonymous:1915:I

[1487] Anonymous. Index. Fishery Bulletin, 35(1):??, 1915–1916. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-74;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb35.index_.pdf.

Davis:1915:MBR

[1488] H. S. Davis. The Myxosporidia of the Beaufort region, a systematic and biologic study. Fishery Bulletin, 35(1):??, 1915– 1916. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/myxosporidia-
beaufort-region-systematic-and-biologic-study;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb35.
7.pdf.

Pearse:1915:FSF

[1489] A. S. Pearse. The food of the shore fishes of certain Wisconsin lakes. Fishery Bulletin, 35(1):??, 1915–1916. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/food-shore-fishes-certain-wisconsin-lakes;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb35.8.pdf.

Wilson:1915:ERA

[1490] Charles Branch Wilson. The economic relations, anatomy and life history of the genus Lernaea. Fishery Bulletin, 35(1):??, 1915–1916. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/wilson-charles-branch;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb35.6_0.pdf.

Evermann:1915:FKT

[1491] Barton Warren Evermann. The fishes of Kentucky and Tennessee: a distributional catalogue of the known species. Fishery Bulletin, 35(1): ??, 1915–1916. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishes-
kentucky-and-tennessee-distributional-catalogue-known-species;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb35.9.pdf.

Kuntz:1915:HBA

[1492] Albert Kuntz. The histological basis of adaptive shades and colors in the flounder paralichthys albiguttus. Fishery Bulletin, 35(1):??, 1915– 1916. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/histological-basis-
adaptive-shades-and-colors-flounder-paralichthys-albiguttus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb35.1.pdf.

Hay:1915:DCB

[1493] W. P. Hay and C. A. Shore. The decapod crustaceans of Beaufort, N.C., and the surrounding region. Fishery Bulletin, 35(1):??, 1915–1916. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/decapod-
crustaceans-beaufort-n-c-and-surrounding-region;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb35.10.pdf.

Kuntz:1915:NEL

[1494] Albert Kuntz and Lewis Radcliffe. Note on the embryology and larval development of twelve teleostean fishes. Fishery Bulletin, 35(1): ??, 1915–1916. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/note-
embryology-and-larval-development-twelve-teleostean-fishes;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb35.3.pdf.

Mitchell:1915:NON

[1495] Philip H. Mitchell. Nutrition of oysters: The nature of the so-called ‘fattening’ of oysters. Fishery Bulletin, 35(1):??, 1915–1916. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/nutrition-oysters-nature-
so-called-fattening-oysters;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb35.11.pdf.

Coker:1917:FWM

[1496] Robert E. Coker. Fresh-water mussels and mussel industries of the United States. Fishery Bulletin, 36(1):??, 1917–1918. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/fresh-water-mussels-and-mussel-industries-
united-states;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb36.2.pdf.

Churchill:1917:LHB

[1497] E. P. Churchill, Jr. Life history of the blue crab. Fishery Bulletin, 36 (1):??, 1917–1918. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/life-
history-blue-crab;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb36.3.pdf.

George:1917:SBC

[1498] W. C. George and H. V. Wilson. Sponges of Beaufort (N.C.) harbor and vicinity. Fishery Bulletin, 36(1):??, 1917–1918. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/sponges-beaufort-n-c-harbor-and-
vicinity;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb36.4.pdf.

Wilson:1917:DDR

[1499] Charles Branch Wilson. Dragonflies and damselflies in relation to pondfish culture, with a list of those found near Fairport, Iowa. Fishery Bulletin, 36(1):??, 1917–1918. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
dragonflies-and-damselflies-relation-pondfish-culture-list-
those-found-near-fairport-iowa;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb36.5.pdf.

Needham:1917:BMO

[1500] James G. Needham. Burrowing mayflies of our larger lakes and streams. Fishery Bulletin, 36(1):??, 1917–1918. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/burrowing-mayflies-our-larger-lakes-and-streams;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb36.6.pdf.

Pearse:1917:HYP

[1501] A. S. Pearse and Henrietta Achtenberg. Habits of yellow perch in Wisconsin lakes. Fishery Bulletin, 36(1):??, 1917–1918. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/habits-yellow-perch-
wisconsin-lakes;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb36.7.pdf.

Hoyt:1917:MAB

[1502] W. D. Hoyt. Marine algae of Beaufort, N.C., and adjacent regions. Fishery Bulletin, 36(1):??, 1917–1918. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/marine-algae-beaufort-n-c-and-adjacent-regions;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb36.8.pdf.

Snyder:1917:TNW

[1503] John Otterbein Snyder. Three new whitefishes from Bear Lake, Idaho and Utah. Fishery Bulletin, 36(1):??, 1917–1918. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/three-new-whitefishes-bear-lake-
idaho-and-utah;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb36.1.pdf.

Anonymous:1917:I

[1504] Anonymous. Index. Fishery Bulletin, 36(1):??, 1917–1918. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-75;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb36.index_.pdf.

Kendall:1919:PMO

[1505] William Converse Kendall. Peritoneal membranes, ovaries, and oviducts of salmonoid fishes and their significance in fish-cultural practice. Fishery Bulletin, 37(1):??, 1919–1920. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/peritoneal-membranes-ovaries-and-oviducts-
salmonoid-fishes-and-their-significance-fish;   https://spo.

nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb37.
3.pdf.

Birge:1919:FLO

[1506] Edward A. Birge and Chancey Juday. Further limnological observations on the Finger Lakes of New York. Fishery Bulletin, 37(1):??, 1919–1920. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/further-
limnological-observations-finger-lakes-new-york;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb37.4.pdf.

Pearse:1919:DFF

[1507] A. S. Pearse. Distribution and food of the fishes of Green Lake, Wis., in summer. Fishery Bulletin, 37(1):??, 1919–1920. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/distribution-and-food-fishes-green-lake-wis-
summer;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb37.5.pdf.

Anonymous:1919:I

[1508] Anonymous. Index. Fishery Bulletin, 37(1):??, 1919–1920. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-76;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb37.index_.pdf.

Rich:1919:EHS

[1509] Willis H. Rich. Early history and seaward migration of Chinook salmon in the Columbia and Sacramento rivers. Fishery Bulletin, 37(1):??, 1919–1920. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/early-history-and-seaward-migration-chinook-salmon-
columbia-and-sacramento-rivers;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb37.1.pdf.

Coker:1919:NHP

[1510] R. E. Coker, A. F. Shira, H. W. Clark, and A. D. Howard. Natural history and propagation of fresh-water mussels. Fishery Bulletin, 37(1): ??, 1919–1920. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/natural-
history-and-propagation-fresh-water-mussels;   https://spo.

nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb37.
2.pdf.

Barney:1921:FNN

[1511] R. L. Barney. Further notes on the natural history and artificial propagation of the diamondback terrapin. Fishery Bulletin, 38(1):??, 1921– 1922. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/further-notes-
natural-history-and-artificial-propagation-diamondback-terrapin;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb38.3.pdf.

Anonymous:1921:I

[1512] Anonymous. Index. Fishery Bulletin, 38(1):??, 1921–1922. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-77;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb38.index_.pdf.

Hildebrand:1921:NHD

[1513] Samuel F. Hildebrand. Note on habits and development of eggs and larvae of the silversides Menidia menidia and Menidia beryllina. Fishery Bulletin, 38(1):??, 1921–1922. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/note-habits-and-development-eggs-and-
larvae-silversides-menidia-menidia-and-menidia;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb38.
4.pdf.

Taylor:1921:DCA

[1514] Harden F. Taylor. Deductions concerning the air bladder and the specific gravity of fishes. Fishery Bulletin, 38(1):??, 1921– 1922. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/deductions-
concerning-air-bladder-and-specific-gravity-fishes;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb38.5.pdf.

Field:1921:BEV

[1515] Irving A. Field. Biology and economic value of the sea mussel Mytilus edulis. Fishery Bulletin, 38(1):??, 1921–1922. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/biology-and-economic-value-sea-

mussel-mytilus-edulis;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb38.6.pdf.

Davis:1921:NBD

[1516] H. S. Davis. A new bacterial disease of fresh-water fishes. Fishery Bulletin, 38(1):??, 1921–1922. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/new-bacterial-disease-fresh-water-fishes;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb38.7.pdf.

Crawford:1921:SLP

[1517] D. R. Crawford and W. J. J. De Smidt. The spiny lobster, Panulirus argus, of Southern Florida: Its natural history and utilization. Fishery Bulletin, 38(1):281–310, 1921–1922. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/spiny-lobster-panulirus-argus-southern-florida-its-
natural-history-and-utilization;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb38.8.pdf.

Leathers:1921:ESA

[1518] Adelbert L. Leathers. Ecological study of aquatic midges and some related insects with special reference to feeding habits. Fishery Bulletin, 38(1):??, 1921–1922. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
ecological-study-aquatic-midges-and-some-related-insects-special-
reference-feeding-habits;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb38.1.pdf.

Breder:1921:SEL

[1519] C. M. Breder, Jr. Some embryonic and larval stages of the winter flounder. Fishery Bulletin, 38(1):??, 1921–1922. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/some-embryonic-and-larval-stages-
winter-flounder;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb38.9.pdf.

Howard:1921:ECF

[1520] Arthur Day Howard. Experiments in the culture of fresh-water mussels. Fishery Bulletin, 38(1):??, 1921–1922. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/experiments-culture-fresh-water-mussels;

https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb38.2.pdf.

Gilbert:1921:SYR

[1521] Charles H. Gilbert. The salmon of the Yukon River. Fishery Bulletin, 38 (1):??, 1921–1922. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/salmon-
yukon-river;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb38.10.pdf.

Wilson:1923:LHS

[1522] Charles Branch Wilson. Life history of the scavenger water beetle Hydrous (Hydrophilus) triangularis, and its economic relation to fish breeding. Fishery Bulletin, 39(1):??, 1923–1924. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/life-history-scavenger-water-beetle-hydrous-
hydrophilus-triangularis-and-its-economic;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb39.2.pdf.

Churchill:1923:FFF

[1523] E. P. Churchill, Jr. and Sara I. Lewis. Food and feeding in fresh-water mussels. Fishery Bulletin, 39(1):??, 1923–1924. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/food-and-feeding-fresh-water-mussels;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb39.10.pdf.

Gilbert:1923:ETA

[1524] Charles H. Gilbert. Experiments in tagging adult red salmon, Alaska Peninsula Fisheries Reservation, summer of 1922. Fishery Bulletin, 39(1):??, 1923–1924. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
experiments-tagging-adult-red-salmon-alaska-peninsula-fisheries-
reservation-summer-1922;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb39.3.pdf.

Anonymous:1923:I

[1525] Anonymous. Index. Fishery Bulletin, 39(1):??, 1923–1924. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-78;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb39.index_.pdf.

Kemmerer:1923:NLU

[1526] George Kemmerer, J. F. Bovard, and W. R. Boorman. Northwestern lakes of the United States: biological and chemical studies with reference to possibilities in production of fish. Fishery Bulletin, 39(1):51–140, 1923–1924. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/northwestern-
lakes-united-states-biological-and-chemical-studies-reference-
possibilities;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb39.4.pdf.

Welsh:1923:CLH

[1527] William W. Welsh and C. M. Breder, Jr. Contributions to life histories of Sciaenidae of the eastern United States coast. Fishery Bulletin, 39(1):??, 1923–1924. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/contributions-
life-histories-sciaenidae-eastern-united-states-coast;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb39.5.pdf.

Blystad:1923:SLM

[1528] Chester N. Blystad. Significance of larval mantle of fresh-water mussels during parasitism, with notes on a new mantle condition exhibited by lampsilis luteola. Fishery Bulletin, 39(1):??, 1923–1924. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/significance-larval-mantle-fresh-
water-mussels-during-parasitism-notes-new-mantle-condition;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb39.6.pdf.

Shackell:1923:TCT

[1529] L. F. Shackell. Toxicities of coal tar creosote, creosote distillates, and individual constituents for the marine wood borer Limnoria lignorum. Fishery Bulletin, 39(1):??, 1923–1924. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/toxicities-coal-tar-
creosote-creosote-distillates-and-individual-constituents-marine-
wood;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb39.7.pdf.

Wilson:1923:WBR

[1530] Charles Branch Wilson. Water beetles in relation to pondfish culture, with life histories of those found in fishponds at Fairport, Iowa. Fishery Bulletin, 39(1):??, 1923–1924. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/water-beetles-relation-pondfish-
culture-life-histories-those-found-fishponds-fairport-iowa;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb39.8.pdf.

Hildebrand:1923:ALF

[1531] Samuel F. Hildebrand. Annotated list of fishes collected in vicinity of Augusta, Georgia, with description of a new darter. Fishery Bulletin, 39(1): ??, 1923–1924. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/annotated-
list-fishes-collected-vicinity-augusta-georgia-description-
new-darter;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb39.1.pdf.

Galtsoff:1923:LOU

[1532] P. S. Galtsoff. Limnological observations in the upper Mississippi, 1921. Fishery Bulletin, 39(1):??, 1923–1924. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/limnological-observations-upper-
mississippi-1921;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb39.9.pdf.

Anonymous:1924:UK

[1533] Anonymous. Use of the keys. Fishery Bulletin, 40(1):??, 1924. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/use-keys;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb40(1).2.pdf.

Anonymous:1924:Ia

[1534] Anonymous. Index. Fishery Bulletin, 40(1):??, 1924. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-79;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb40(1).index_.pdf.

Anonymous:1924:KGM

[1535] Anonymous. Key to Gulf of Maine fishes. Fishery Bulletin, 40(1): ??, 1924. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/key-gulf-

maine-fishes;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb40(1).3.pdf.

Anonymous:1924:L

[1536] Anonymous. The lampreys. Fishery Bulletin, 40(1):??, 1924. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/lampreys;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb40(1).4.pdf.

Anonymous:1924:SR

[1537] Anonymous. Sharks and rays. Fishery Bulletin, 40(1):??, 1924. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/sharks-and-rays;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb40(
1).5.pdf.

Anonymous:1924:Ca

[1538] Anonymous. Chimæroids. Fishery Bulletin, 40(1):??, 1924. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/chim%C3%A6roids;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb40(
1).6.pdf.

Anonymous:1924:BF

[1539] Anonymous. The bony fishes. Fishery Bulletin, 40(1):??, 1924. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/bony-fishes;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb40(1).
7.pdf.

Anonymous:1924:Ba

[1540] Anonymous. Bibliography. Fishery Bulletin, 40(1):??, 1924. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/bibliography;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb40(1).
8.pdf.

Anonymous:1924:Ib

[1541] Anonymous. Introduction. Fishery Bulletin, 40(1):??, 1924. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/introduction;   https://spo.nmfs.

noaa.gov/sites/default/files/pdf-content/fish-bull/fb40(1).
1.pdf.

Anonymous:1924:Aa

[1542] Anonymous. Addenda. Fishery Bulletin, 40(1):??, 1924. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/addenda-0;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb40(
1).9.pdf.

Anonymous:1924:Ic

[1543] Anonymous. Introduction. Fishery Bulletin, 40(2):??, 1924. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/introduction-0;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb40(2).1.pdf.

Anonymous:1924:P

[1544] Anonymous. The plankton. Fishery Bulletin, 40(2):??, 1924. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/plankton-0;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb40(2).
2.pdf.

Anonymous:1924:SGSa

[1545] Anonymous. Section 1: General survey of the animal plankton (zoöplankton). Fishery Bulletin, 40(2):??, 1924. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/section-1-general-survey-animal-
plankton-zo%C3%B6plankton;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb40(2).3.pdf.

Anonymous:1924:SGSb

[1546] Anonymous. Section 2: General survey of the planktonic plants (phytoplankton) and unicellular animals. Fishery Bulletin, 40(2):??, 1924. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/section-2-general-survey-
planktonic-plants-phytoplankton-and-unicellular-animals;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb40(2).4.pdf.

Anonymous:1924:Bb

[1547] Anonymous. Bibliography. Fishery Bulletin, 40(2):??, 1924. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/bibliography-0;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb40(2).5.pdf.

Anonymous:1924:Ta

[1548] Anonymous. Topography. Fishery Bulletin, 40(3):??, 1924. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/topography;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb40(3).
3.pdf.

Anonymous:1924:CGM

[1549] Anonymous. Circulation in the Gulf of Maine. Fishery Bulletin, 40(3): ??, 1924. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/circulation-
gulf-maine;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb40(3).11.pdf.

Anonymous:1924:Tb

[1550] Anonymous. Temperature. Fishery Bulletin, 40(3):??, 1924. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/temperature;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb40(3).
4.pdf.

Anonymous:1924:TTS

[1551] Anonymous. Tables of temperature, salinity, and density. Fishery Bulletin, 40(3):??, 1924. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/tables-temperature-salinity-and-density;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb40(3).12.pdf.

Anonymous:1924:FGT

[1552] Anonymous. Factors governing the temperature of the Gulf of Maine. Fishery Bulletin, 40(3):??, 1924. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/factors-governing-temperature-gulf-maine;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb40(3).5.pdf.

Anonymous:1924:Bc

[1553] Anonymous. Bibliography. Fishery Bulletin, 40(3):??, 1924. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/bibliography-1;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb40(3).13.pdf.

Anonymous:1924:S

[1554] Anonymous. Salinity. Fishery Bulletin, 40(3):??, 1924. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/salinity;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb40(3).6.pdf.

Anonymous:1924:Ab

[1555] Anonymous. Alkalinity. Fishery Bulletin, 40(3):??, 1924. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/alkalinity;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb40(3).
7.pdf.

Anonymous:1924:VT

[1556] Anonymous. Visual transparency. Fishery Bulletin, 40(3):??, 1924. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/visual-transparency;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb40(3).8.pdf.

Anonymous:1924:Id

[1557] Anonymous. Introduction. Fishery Bulletin, 40(3):??, 1924. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/introduction-1;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb40(3).1.pdf.

Anonymous:1924:Cb

[1558] Anonymous. Color. Fishery Bulletin, 40(3):??, 1924. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/color;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb40(3).9.pdf.

Anonymous:1924:OH

[1559] Anonymous. Oceanographic history. Fishery Bulletin, 40(3):??, 1924. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/oceanographic-
history;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb40(3).2.pdf.

Anonymous:1924:SWG

[1560] Anonymous. Sources from which the Gulf of Maine receives its waters. Fishery Bulletin, 40(3):??, 1924. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/sources-which-gulf-maine-receives-its-waters;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb40(3).10.pdf.

Rich:1925:GDM

[1561] Willis H. Rich. Growth and degree of maturity of Chinook salmon in the ocean. Fishery Bulletin, 41(1):??, 1925. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/growth-and-degree-maturity-chinook-salmon-
ocean;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb41.3.pdf.

Fish:1925:SDP

[1562] Charles J. Fish. Seasonal distribution of the plankton of the Woods Hole region. Fishery Bulletin, 41(1):??, 1925. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/seasonal-distribution-plankton-woods-hole-
region;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb41.4.pdf.

Kenyon:1925:DEP

[1563] Walter A. Kenyon. Digestive enzymes in poikilothermal vertebrates. An investigation of enzymes in fishes, with comparative studies on those of amphibians, reptiles, and mammals. Fishery Bulletin, 41(1):??, 1925. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
digestive-enzymes-poikilothermal-vertebrates-investigation-
enzymes-fishes-comparative;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb41.5.pdf.

Weymouth:1925:GAM

[1564] F. W. Weymouth, H. C. McMilllin, and H. B. Holmes. Growth and age at maturity of the Pacific razor clam, Siliqua patula (Dixon). Fishery Bulletin, 41(1):201–236, 1925. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
growth-and-age-maturity-pacific-razor-clam-siliquia-patula-
dixon;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb41.6.pdf.

Hildebrand:1925:FRS

[1565] Samuel F. Hildebrand. Fishes of the Republic of El Salvador, Central America. Fishery Bulletin, 41(1):??, 1925. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/fishes-republic-el-salvador-central-america;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb41.7.pdf.

Stejneger:1925:SIC

[1566] Leonhard Stejneger. Fur-seal industry of the Commander Islands, 1897 to 1922. Fishery Bulletin, 41(1):??, 1925. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fur-seal-industry-commander-islands-1897-1922;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb41.8.pdf.

Smith:1925:VMD

[1567] Frank Smith. Variation in the maximum depth at which fish can live during summer in a moderately deep lake with a thermocline. Fishery Bulletin, 41(1):??, 1925. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/variation-maximum-depth-which-fish-can-live-during-
summer-moderately-deep-lake-thermocline;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb41.1.pdf.

Anonymous:1925:I

[1568] Anonymous. Index. Fishery Bulletin, 41(1):??, 1925. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-80;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb41.index_.pdf.

Osburn:1925:BTC

[1569] Raymond C. Osburn. Black tumor of the catfish. Fishery Bulletin, 41(1): ??, 1925. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/black-tumor-
catfish;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb41.2.pdf.

Davis:1927:OSP

[1570] H. S. Davis. Octomitus salmonis, a parasitic flagellate of trout. Fishery Bulletin, 42(1):??, 1927. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
octomitus-salmonis-parasitic-flagellate-trout;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb42.
2.pdf.

Hildebrand:1927:DTN

[1571] Samuel F. Hildebrand and Isaac Ginsburg. Descriptions of two new species of fishes from Key West, Fla., with notes on nine other species collected in the same locality. Fishery Bulletin, 42(1):??, 1927. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/descriptions-two-new-
species-fishes-key-west-fla-notes-nine-other-species-collected-
same;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb42.10.pdf.

Gilbert:1927:SET

[1572] Charles H. Gilbert and Willis H. Rich. Second experiment in tagging salmon in the Alaska Peninsula Fisheries Reservation, summer of 1923. Fishery Bulletin, 42(1):??, 1927. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/second-experiment-tagging-salmon-
alaska-peninsula-fisheries-reservation-summer-1923;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb42.3.pdf.

Kendall:1927:S

[1573] William Converse Kendall. The smelts. Fishery Bulletin, 42(1):??, 1927. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/kendall-william-converse;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb42.11.pdf.

Hildebrand:1927:DCT

[1574] Samuel F. Hildebrand and Isaac Ginsburg. Distinguishing characters of two species of red snappers of the Atlantic coast of North America. Fishery Bulletin, 42(1):??, 1927. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/distinguishing-characters-two-species-red-snappers-
atlantic-coast-north-america;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb42.4.pdf.

Anonymous:1927:Ia

[1575] Anonymous. Index. Fishery Bulletin, 42(1):??, 1927. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-81;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb42.index_.pdf.

Engle:1927:CGC

[1576] Earl Theron Engle. Crayfishes of the genus Cambarus in Nebraska and eastern Colorado. Fishery Bulletin, 42(1):??, 1927. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/crayfishes-genus-cambarus-nebraska-and-
eastern-colorado;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb42.5.pdf.

Coker:1927:NGD

[1577] R. E. Coker. New genus of darter from western North Carolina. Fishery Bulletin, 42(1):??, 1927. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/new-genus-darter-western-north-carolina;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb42.6.pdf.

Rich:1927:STEa

[1578] Willis H. Rich. Salmon-tagging experiments in Alaska, 1924 and 1925. Fishery Bulletin, 42(1):??, 1927. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/salmon-tagging-experiments-alaska-1924-and-1925;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb42.7.pdf.

Stewart:1927:DGF

[1579] Norman Hamilton Stewart. Development, growth, and food habits of the white sucker, Catostomus commersonii Le Sueur. Fishery Bulletin, 42(1):

??, 1927. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/development-
growth-and-food-habits-white-sucker-catostomus-commersonii-
le-sueur;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb42.8.pdf.

Davis:1927:SSN

[1580] H. S. Davis. Schizamoeba salmonis, a new ameba parasitic in salmonid fishes. Fishery Bulletin, 42(1):??, 1927. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/schizamoeba-salmonis-new-ameba-
parasitic-salmonid-fishes;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb42.1.pdf.

Birge:1927:OCL

[1581] Edward A. Birge and Chancey Juday. Organic content of lake water. Fishery Bulletin, 42(1):??, 1927. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/organic-content-lake-water;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb42.9.pdf.

Hildebrand:1927:FCB

[1582] Samuel F. Hildebrand and William C. Schroeder. Fish of the Chesapeake Bay. Fishery Bulletin, 43(1):1–366, 1927. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/fish-chesapeake-bay;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
hildebrand.pdf.

Rich:1927:STEb

[1583] Willis H. Rich and Arnie J. Suomela. Salmon-tagging experiments in Alaska, 1926. Fishery Bulletin, 43(2):71–104, 1927. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/salmon-tagging-experiments-alaska-1926;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/rich.pdf.

Hildebrand:1927:ALF

[1584] Samuel F. Hildebrand and Irving L. Towers. Annotated list of fishes collected in the vicinity of Greenwood, Miss., with descriptions of three new species. Fishery Bulletin, 43(2):??, 1927. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/annotated-list-fishes-collected-

vicinity-greenwood-miss-descriptions-three-new-species;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb43(2).3.pdf.

Wiebe:1927:BSU

[1585] A. H. Wiebe. Biological survey of the upper Mississippi River with special reference to pollution. Fishery Bulletin, 43(2):??, 1927. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/biological-
survey-upper-mississippi-river-special-reference-pollution;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/wiebe.pdf.

Hathaway:1927:QSC

[1586] Edward S. Hathaway. Quantitative study of the changes produced by acclimatiaztion in the tolerance of high temperatures by fishes and amphibians. Fishery Bulletin, 43(2):??, 1927. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/quantitative-study-
changes-produced-acclimatiaztion-tolerance-high-temperatures-
fishes-and;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/hathaway.pdf.

Visscher:1927:NEF

[1587] J. Paul Visscher. Nature and extent of fouling of ships’ bottoms. Fishery Bulletin, 43(2):??, 1927. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
nature-and-extent-fouling-ships-bottoms.

Fish:1927:PDC

[1588] Charles J. Fish. Production and distribution of cod eggs in Massachusetts Bay in 1924 and 1925. Fishery Bulletin, 43(2):??, 1927. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/production-and-distribution-cod-
eggs-massachusetts-bay-1924-and-1925;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb43(2).7.pdf.

Koelz:1927:CFG

[1589] Walter Koelz. Coregonid fishes of the Great Lakes. Fishery Bulletin, 43(2):??, 1927. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/coregonid-
fishes-great-lakes;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb43(2).8.pdf.

Gilbert:1927:ICR

[1590] Charles H. Gilbert and Willis H. Rich. Investigations concerning the red-salmon runs to the Karluk River, Alaska. Fishery Bulletin, 43(2):??, 1927. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/investigations-concerning-red-salmon-runs-
karluk-river-alaska;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb43(2).1.pdf.

Anonymous:1927:Ib

[1591] Anonymous. Index. Fishery Bulletin, 43(2):??, 1927. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-82;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb43(2).index_.pdf.

Galtsoff:1928:ESF

[1592] Paul S. Galtsoff. Experimental study of the function of the oyster gills and its bearing on the problems of oyster culture and sanitary control of the oyster industry. Fishery Bulletin, 44(1):??, 1928. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/experimental-study-
function-oyster-gills-and-its-bearing-problems-oyster-culture-
and;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb44.1.pdf.

Rich:1928:SRA

[1593] Willis H. Rich and Edward M. Ball. Statistical review of the Alaska salmon fisheries. Part I: Bristol Bay and the Alaska Peninsula. Fishery Bulletin, 44(1):??, 1928. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
statistical-review-alaska-salmon-fisheries-part-i-bristol-bay-
and-alaska-peninsula;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb44.2.pdf.

McMillan:1928:EFS

[1594] F. O. McMillan. Electric fish screen. Fishery Bulletin, 44(1):??, 1928. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/electric-fish-screen;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb44.3.pdf.

Pearson:1928:NHC

[1595] John C. Pearson. Natural history and conservation of the redfish and other commercial sciaenids on the Texas coast. Fishery Bulletin, 44(1):??, 1928. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/natural-history-and-conservation-redfish-and-other-
commercial-sciaenids-texas-coast;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb44.4.pdf.

Rich:1928:EMY

[1596] Willis H. Rich and Harlan B. Holmes. Experiments in marking young Chinook salmon on the Columbia River, 1916 to 1927. Fishery Bulletin, 44(1):??, 1928. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
experiments-marking-young-chinook-salmon-columbia-river-1916-
1927;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb44.5.pdf.

VanOosten:1928:LHL

[1597] John Van Oosten. Life history of the lake herring (Leucichthys artedi LeSueur) of Lake Huron as revealed by its scales, with a critique of the scale method. Fishery Bulletin, 44(1):??, 1928. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/life-history-lake-herring-leucichthys-artedi-
lesueur-lake-huron-revealed-its-scales;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb44.6.pdf.

Prytherch:1928:IPC

[1598] Herbert F. Prytherch. Investigation of the physical conditions controlling spawning of oysters and the occurrence, distribution, and setting of oyster larvae in Milford Harbor, Connecticut. Fishery Bulletin, 44(1): ??, 1928. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/investigation-
physical-conditions-controlling-spawning-oysters-and-occurrence-
distribution;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb44.7.pdf.

Anonymous:1928:Ia

[1599] Anonymous. Index. Fishery Bulletin, 44(1):??, 1928. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-83;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb44.index_.pdf.

Ginsburg:1928:RWC

[1600] Isaac Ginsburg. Review of the weakfishes (Cynoscion) of the Atlantic and Gulf coasts of the United States, with a description of a new species. Fishery Bulletin, 45(1):??, 1928. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/review-weakfishes-cynoscion-atlantic-
and-gulf-coasts-united-states-description-new-species;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb45.3.pdf.

Coker:1928:KDF

[1601] Robert E. Coker. Keokuk Dam and the fisheries of the upper Mississippi River. Fishery Bulletin, 45(1):??, 1928. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/keokuk-dam-and-fisheries-
upper-mississippi-river;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb45.4.pdf.

Coker:1928:SCF

[1602] Robert E. Coker. Studies of common fishes of the Mississippi River at Keokuk. Fishery Bulletin, 45(1):??, 1928. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/studies-common-fishes-mississippi-river-keokuk;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb45.5.pdf.

Rounsefell:1928:CBP

[1603] George A. Rounsefell. Contribution to the biology of the Pacific herring, Clupea pallasii, and the condition of the fishery in Alaska. Fishery Bulletin, 45(1):??, 1928. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/contribution-biology-pacific-herring-
clupea-pallasii-and-condition-fishery-alaska;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb45.
6.pdf.

Anonymous:1928:Ib

[1604] Anonymous. Index. Fishery Bulletin, 45(1):??, 1928. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-84;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb45.index_.pdf.

Rich:1928:STE

[1605] Willis H. Rich and Frederick G. Morton. Salmon-tagging experiments in Alaska 1927 and 1928. Fishery Bulletin, 45(1):??, 1928. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/salmon-
tagging-experiments-alaska-1927-and-1928;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb45.1.pdf.

Hildebrand:1928:REA

[1606] Samuel F. Hildebrand. Review of experiments on artificial culture of diamond-back terrapin. Fishery Bulletin, 45(1):??, 1928. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/hildebrand-samuel-f-
review-experiments-artificial-culture-diamond-back-terrapin;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb45.2.pdf.

Thompson:1929:STE

[1607] Seton H. Thompson. Salmon-tagging experiments in Alaska, 1929. Fishery Bulletin, 46(1):??, 1929. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
salmon-tagging-experiments-alaska-1929;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb46.3.pdf.

Gutsell:1929:NHB

[1608] James S. Gutsell. Natural history of the bay scallop. Fishery Bulletin, 46(1):??, 1929. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/natural-
history-bay-scallop;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb46.11.pdf.

Galtsoff:1929:ESP

[1609] P. S. Galtsoff, H. F. Prytherch, and H. C. McMillin. An experimental study in production and collection of seed oysters. Fishery Bulletin, 46(1):197–263, 1929. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
experimental-study-production-and-collection-seed-oysters;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb46.4.pdf.

Weymouth:1929:AGP

[1610] F. W. Weymouth and Seton H. Thompson. The age and growth of the Pacific cockle (Cardium corbis Martyn). Fishery Bulletin, 46 (1):??, 1929. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/age-and-
growth-pacific-cockle-cardium-corbis-martyn;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb46.
12.pdf.

Ginsburg:1929:CSL

[1611] Isaac Ginsburg. Commercial snappers (Lutianidae) of Gulf of Mexico. Fishery Bulletin, 46(1):??, 1929. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/commercial-snappers-lutianidae-gulf-mexico;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb46.5.pdf.

Rich:1929:SRA

[1612] Willis H. Rich and Edward M. Ball. Statistical review of the Alaska salmon fisheries. Part II: Chignik to Resurrection Bay. Fishery Bulletin, 46(1):??, 1929. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
statistical-review-alaska-salmon-fisheries-part-ii-chignik-
resurrection-bay;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb46.13.pdf.

Cowles:1929:BSO

[1613] R. P. Cowles. A biological study of the offshore waters of Chesapeake Bay. Fishery Bulletin, 46(1):??, 1929. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/biological-study-offshore-waters-
chesapeake-bay;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb46.6.pdf.

Chamberlain:1929:AGF

[1614] Thomas K. Chamberlain. Annual growth of fresh-water mussels. Fishery Bulletin, 46(1):??, 1929. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
annual-growth-fresh-water-mussels;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb46.14.pdf.

Hildebrand:1929:DLH

[1615] Samuel F. Hildebrand and Louella E. Cable. Development and life history of fourteen teleostean fishes at Beaufort, N.C. Fishery Bulletin, 46(1):??, 1929. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/development-and-life-
history-fourteen-teleostean-fishes-beaufort-n-c;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb46.7.pdf.

Anonymous:1929:I

[1616] Anonymous. Index. Fishery Bulletin, 46(1):??, 1929. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-85;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb46.index_.pdf.

Galtsoff:1929:OCN

[1617] Paul S. Galtsoff and Dorothy V. Whipple. Oxygen consumption of normal and green oysters. Fishery Bulletin, 46(1):??, 1929. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/oxygen-consumption-
normal-and-green-oysters;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb46.8.pdf.

Schroeder:1929:MOP

[1618] William C. Schroeder. Migrations and other phases in the life history of the cod off southern New England. Fishery Bulletin, 46(1):??, 1929. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/migrations-and-other-
phases-life-history-cod-southern-new-england;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb46.
1.pdf.

Ellis:1929:BNA

[1619] M. M. Ellis, Amanda D. Merrick, and Marion D. Ellis. The blood of North American fresh-water mussels under normal and adverse conditions. Fishery Bulletin, 46(1):509–542, 1929. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/blood-north-american-fresh-water-
mussels-under-normal-and-adverse-conditions;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb46.
9.pdf.

Wiebe:1929:IPP

[1620] A. H. Wiebe. Investigations on plankton production in fish ponds. Fishery Bulletin, 46(1):??, 1929. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
investigations-plankton-production-fish-ponds;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb46.
2.pdf.

Weymouth:1929:RGM

[1621] F. W. Weymouth and H. C. McMillin. Relative growth and mortality of the Pacific razor clam (Siliqua patula Dixon) and their bearing on the commercial fishery. Fishery Bulletin, 46(1):543–567, 1929. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/relative-growth-and-
mortality-pacific-razor-clam-siliqua-patula-dixon-and-their-
bearing;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb46.10.pdf.

Hopkins:1931:TSM

[1622] A. E. Hopkins. Temperature and the shell movements of oysters. Fishery Bulletin, 47(1):??, 1931–1933. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/temperature-and-shell-movements-oysters;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb47.1.pdf.

Rounsefell:1931:FSHa

[1623] George A. Rounsefell and Edwin H. Dahlgren. Fluctuations in the supply of herring, Clupea pallashii, in Prince William Sound, Alaska. Fishery Bulletin, 47(1):??, 1931–1933. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fluctuations-supply-herring-clupea-pallashii-prince-
william-sound-alaska;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb47.9.pdf.

Rounsefell:1931:FSHb

[1624] George A. Rounsefell. Fluctuations in the supply of herring (Clupea pallasii) in southeastern Alaska. Fishery Bulletin, 47(1):??, 1931–1933. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/fluctuations-supply-
herring-clupea-pallasii-southeastern-alaska;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb47.
2.pdf.

Fish:1931:CEL

[1625] Marie Poland Fish. Contributions to the early life histories of sixty-two species of fishes from Lake Erie and its tributary waters. Fishery Bulletin, 47(1):??, 1931–1933. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
contributions-early-life-histories-sixty-two-species-fishes-
lake-erie-and-its-tributary;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb47.10.pdf.

Hopkins:1931:FIS

[1626] A. E. Hopkins. Factors influencing the spawning and setting of oysters in Galveston Bay, Texas. Fishery Bulletin, 47(1):

??, 1931–1933.  CODEN FSYBAY.  ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
factors-influencing-spawning-and-setting-oysters-galveston-
bay-texas;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb47.3.pdf.

Rich:1931:STE

[1627] Willis H. Rich. Salmon-tagging experiments in Alaska, 1930. Fishery Bulletin, 47(1):??, 1931–1933. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
salmon-tagging-experiments-alaska-1930;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb47.11.
pdf.

Federighi:1931:SOD

[1628] Henry Federighi. Studies on the oyster drill (Urosalpinx cinerea, say). Fishery Bulletin, 47(1):??, 1931–1933. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/studies-oyster-drill-urosalpinx-
cinerea-say;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb47.4.pdf.

Juday:1931:LSK

[1629] C. Juday, Willish H. Rich, G. I. Kemmerer, and Albert Mann. Limnological studies of Karluk Lake, Alaska, 1926–1930. Fishery Bulletin, 47(1):407–436, 1931–1933. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
limnological-studies-karluk-lake-alaska-1926-1930;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb47.12.pdf.

Ginsburg:1931:JSC

[1630] Isaac Ginsburg. Juvenile and sex characters of Evorthodus lyricus (fam. Gobiidae). Fishery Bulletin, 47(1):??, 1931–1933. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/juvenile-and-sex-characters-evorthodus-
lyricus-fam-gobiidae;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb47.5.pdf.

Rich:1931:SRAa

[1631] Willis H. Rich and Edward M. Ball. Statistical review of the Alaska salmon fisheries. Part IV: Southeastern Alaska. Fishery Bulletin, 47(1):??, 1931–1933. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/statistical-review-alaska-salmon-fisheries-part-iv-
southeastern-alaska;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb47.13.pdf.

Hopkins:1931:EPM

[1632] A. E. Hopkins, Paul S. Galtsoff, and H. C. McMillin. Effects of pulp mill pollution on oysters. Fishery Bulletin, 47(1):125–186, 1931– 1933. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/effects-pulp-
mill-pollution-oysters;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb47.6.pdf.

Anonymous:1931:I

[1633] Anonymous. Index. Fishery Bulletin, 47(1):??, 1931–1933. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-86;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb47.index_.pdf.

Rich:1931:SRAb

[1634] Willis H. Rich and Edward M. Ball. Statistical review of the Alaska salmon fisheries. Part III: Prince William Sound, Copper River, and Bering River. Fishery Bulletin, 47(1):??, 1931–1933. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/rich-willis-h-and-edward-m-ball;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb47.7.pdf.

Hopkins:1931:SSO

[1635] A. E. Hopkins. Sensory stimulation of the oyster, Ostrea virginica, by chemicals. Fishery Bulletin, 47(1):??, 1931–1933. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/sensory-stimulation-
oyster-ostrea-virginica-chemicals;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb47.8.pdf.

Davidson:1933:HIA

[1636] Frederick Davidson. The homing instinct and age at maturity of pink salmon (Oncorhynchus gorbuscha). Fishery Bulletin, 48(1):??, 1933– 1936. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/homing-instinct-and-
age-maturity-pink-salmon-oncorhynchus-gorbuscha;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb48.2.pdf.

Hopkins:1933:EOS

[1637] A. E. Hopkins. Experimental observations on spawning, larval development, and setting in the Olympia oyster, Ostrea lurida. Fishery Bulletin, 48(1):??, 1933–1936. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/experimental-observations-spawning-larval-
development-and-setting-olympia-oyster-ostrea;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb48.
10.pdf.

Hildebrand:1933:RDW

[1638] Samuel F. Hildebrand and Louella E. Cable. Reproduction and development of whitings or kingfishes, drums, spot, croaker, and weakfishes or sea trouts, family Sciaenidae, or the Atlantic coast of the United States. Fishery Bulletin, 48(1):??, 1933–1936. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/reproduction-and-development-whitings-or-kingfishes-
drums-spot-croaker-and-weakfishes-or;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb48.3.pdf.

Hildebrand:1933:FND

[1639] Samuel F. Hildebrand and Louella E. Cable. Further notes on the development and life history of some teleosts at Beaufort, N.C. Fishery Bulletin, 48(1):??, 1933–1936. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
further-notes-development-and-life-history-some-teleosts-beaufort-
n-c;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb48.11.pdf.

Rounsefell:1933:RHC

[1640] George A. Rounsefell and Edwin H. Dahlgren. Races of herring, Clupen pallasii, in southeastern Alaska. Fishery Bulletin, 48(1): ??, 1933–1936. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/races-
herring-clupen-pallasii-southeastern-alaska;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb48.
4.pdf.

Davidson:1933:MPS

[1641] Frederick A. Davidson and Leroy S. Christey. The migrations of pink salmon (Oncorhynchus gorbuscha) in the Clarence and Sumner Straits regions of southeastern Alaska. Fishery Bulletin, 48 (1):??, 1933–1936. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
migrations-pink-salmon-oncorhynchus-gorbuscha-clarence-and-
sumner-straits-regions;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb48.12.pdf.

Galtsoff:1933:ECO

[1642] Paul S. Galtsoff, Herbert F. Prytherch, Robert O. Smith, and Vera Kochring. Effects of crude oil pollution on oysters in Louisiana waters. Fishery Bulletin, 48(1):??, 1933–1936. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/effects-crude-oil-pollution-oysters-louisiana-waters;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb48.5.pdf.

Davidson:1933:GDE

[1643] Frederick A. Davidson and Samuel J. Hutchinson. The geographic distribution and environmental limitations of the Pacific salmon (genus Oncorhynchus). Fishery Bulletin, 48(1):??, 1933– 1936. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/geographic-
distribution-and-environmental-limitations-pacific-salmon-genus-
oncorhynchus;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb48.13.pdf.

Hile:1933:AGC

[1644] Ralph Hile. Age and growth of the cisco, Leucichthys artedi (Le Sueur), in the lakes of the northeastern highlands, Wisconsin. Fishery Bulletin, 48(1):??, 1933–1936. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/age-and-growth-cisco-leucichthys-
artedi-le-sueur-lakes-northeastern-highlands-wisconsin;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb48.6.pdf.

Rounsefell:1933:SSF

[1645] George A. Rounsefell and George B. Kelez. The salmon and salmon fisheries of Swiftsure Bank, Puget Sound, and the Fraser River. Fishery Bulletin, 48(1):??, 1933–1936. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
salmon-and-salmon-fisheries-swiftsure-bank-puget-sound-and-
fraser-river;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb48.14.pdf.

Bigelow:1933:SNF

[1646] Henry B. Bigelow and William C. Schroeder. Supplemental notes on the fishes of the Gulf of Maine. Fishery Bulletin, 48(1):??, 1933– 1936. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/supplemental-
notes-fishes-gulf-maine;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb48.7.pdf.

Pearson:1933:LHS

[1647] John C. Pearson. The life history of the striped bass, or rockfish, Roccus saxatilis (Walbaum). Fishery Bulletin, 48(1):??, 1933– 1936. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/life-history-
striped-bass-or-rockfish-roccus-saxatilis-walbaum;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb48.15.pdf.

Hopkins:1933:AFM

[1648] A. E. Hopkins. Adaptation of the feeding mechanism of the oyster, Ostrea gigas, to changes in salinity. Fishery Bulletin, 48(1):??, 1933– 1936. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/adaptation-feeding-
mechanism-oyster-ostrea-gigas-changes-salinity;   https://spo.

nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb48.
8.pdf.

Anonymous:1933:I

[1649] Anonymous. Index. Fishery Bulletin, 48(1):??, 1933–1936. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-87;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb48.index_.pdf.

Weymouth:1933:PRL

[1650] F. W. Weymouth, Milton J. Lindner, and W. W. Anderson. Preliminary report on the life history of the common shrimp, Penaeus setiferus (Linn.). Fishery Bulletin, 48(1):1–26, 1933–1936. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/preliminary-report-
life-history-common-shrimp-penaeus-setiferus-linn;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb48.1.pdf.

Ellis:1933:DMS

[1651] M. M. Ellis. Detection and measurement of stream pollution. Fishery Bulletin, 48(1):??, 1933–1936. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/detection-and-measurement-stream-pollution;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb48.9.pdf.

Walford:1938:ECD

[1652] Lionel A. Walford. Effect of currents on distribution and survival of the eggs and larvae of the haddock (Melanogrammus aeglefinus) on Georges Bank. Fishery Bulletin, 49(1):??, 1938. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/effect-currents-distribution-and-survival-eggs-
and-larvae-haddock-melanogrammus-aeglefinus;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb49.
1.pdf.

Pearson:1938:ELH

[1653] John C. Pearson. The early life histories of some American Penaeidae, chiefly the commercial shrimp, Penaeus setiferus (Linn.). Fishery Bulletin, 49(1):1–73, 1938. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/early-

life-histories-some-american-penaeidae-chiefly-commercial-shrimp-
penaeus-setiferus;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb49.2.pdf.

Galtsoff:1938:NHM

[1654] Paul S. Galtsoff and Victor L. Loosanoff. Natural history and method of controlling the starfish (Asterias forbesi, Desor). Fishery Bulletin, 49 (1):??, 1938. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/natural-
history-and-method-controlling-starfish-asterias-forbesi-desor;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb49.3.pdf.

Craig:1938:HDF

[1655] Joseph A. Craig and Robert L. Hacker. The history and development of the fisheries of the Columbia River. Fishery Bulletin, 49(1):??, 1938. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
history-and-development-fisheries-columbia-river;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb49.4.pdf.

Loosanoff:1938:SSO

[1656] Victor L. Loosanoff and James B. Engle. Spawning and setting of oysters in Long Island Sound in 1937, and discussion of the method for predicting the intensity and time of oyster setting. Fishery Bulletin, 49(1):??, 1938. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/spawning-and-setting-oysters-long-island-sound-
1937-and-discussion-method-predicting;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb49.5.pdf.

Ellis:1938:WCA

[1657] M. M. Ellis. Water conditions affecting aquatic life in elephant butte reservoir. Fishery Bulletin, 49(1):??, 1938. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/water-conditions-affecting-aquatic-life-elephant-
butte-reservoir;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb49.6.pdf.

Anonymous:1938:I

URL https://spo.nmfs.noaa.gov/content/index-88;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb49.index_.pdf.

Pearson:1941:YSM

[1659] John C. Pearson. The young of some marine fishes taken in lower Chesapeake Bay, Virginia, with special reference to the gray sea trout Cynoscion regalis (Bloch). Fishery Bulletin, 50(1):??, 1941–1946. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/young-some-marine-fishes-
taken-lower-chesapeake-bay-virginia-special-reference-gray-
sea;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb50.2.pdf.

Rich:1941:SRC

[1660] Willis H. Rich. The salmon runs of the Columbia River in 1938. Fishery Bulletin, 50(1):??, 1941–1946. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/salmon-runs-columbia-river-1938;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb50.3.pdf.

Sette:1941:BAM

[1661] Oscar Elton Sette. Biology of the Atlantic mackerel (Scomber scombrus) of North America. Fishery Bulletin, 50(1):??, 1941– 1946. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/biology-
atlantic-mackerel-scomber-scombrus-north-america;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb50.4.pdf.

Barnaby:1941:FAR

[1662] Joseph T. Barnaby. Fluctuations in abundance of red salmon, Oncorhynchus nerka (Walbaum), of the Karluk River, Alaska. Fishery Bulletin, 50(1):??, 1941–1946. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fluctuations-abundance-red-salmon-oncorhynchus-
nerka-walbaum-karluk-river-alaska;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb50.5.pdf.

VanOosten:1941:WFL

[1663] John Van Oosten, Ralph Hile, and Frank W. Jobes. The whitefish fishery of Lakes Huron and Michigan with special reference to the deep-trap-net fishery. Fishery Bulletin, 50(1):297–394, 1941– 1946. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic).  URL https://spo.nmfs.noaa.gov/content/whitefish-
fishery-lakes-huron-and-michigan-special-reference-deep-trap-
net-fishery;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb50.6.pdf.

Merriman:1941:SSB

[1664] Daniel Merriman. Studies on the striped bass (Roccus saxatilis) of the Atlantic coast. Fishery Bulletin, 50(1):??, 1941–1946. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/studies-striped-bass-
roccus-saxatilis-atlantic-coast;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb50.1.pdf.

Galtsoff:1952:EPS

[1665] Paul S. Galtsoff, Walter A. Chipman, Jr., James B. Engle, and Howard N. Calderwood. Ecological and physiological studies of the effect of sulfate pulp mill wastes on oysters in the York River, Virginia. Fishery Bulletin, 51(1):??, 1952. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/ecological-and-physiological-studies-
effect-sulfate-pulp-mill-wastes-oysters-york-river;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb51.3.pdf.

Silliman:1952:FAC

[1666] Ralph P. Silliman. Fluctuations in abundance of Columbia River Chinook salmon (Oncorhynchus tschawytscha) 1935–45. Fishery Bulletin, 51(1): ??, 1952. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/fluctuations-
abundance-columbia-river-chinook-salmon-oncorhynchus-tschawytscha-
1935-45;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb51.11.pdf.

Schaefer:1952:CBP

[1667] Milner B. Schaefer and John C. Marr. Contributions to the biology of the Pacific tunas. Fishery Bulletin, 51(1):??, 1952. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/contributions-biology-pacific-tunas;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb51.4.pdf.

Marr:1952:AAP

[1668] John C. Marr. Apparent abundance of the pilchard (Sardinops caerulea) off Oregon and Washington, 1935–43, as measured by the catch per boat. Fishery Bulletin, 51(1):??, 1952. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/apparent-abundance-pilchard-sardinops-caerulea-
oregon-and-washington-1935-43-measured-catch;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb51.
12.pdf.

Perlmutter:1952:AGI

[1669] Alfred Perlmutter and George M. Clarke. Age and growth of immature rosefish (Sebastes marinus) in the Gulf of Maine off western Nova Scotia. Fishery Bulletin, 51(1):??, 1952. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/age-and-growth-immature-rosefish-sebastes-
marinus-gulf-maine-western-nova-scotia;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb51.5.pdf.

Wade:1952:JFN

[1670] Charles B. Wade. Juvenile forms of Neothunnus macropterus, Katsuwonus pelamis, and Euthynnus yaito from Philippine seas. Fishery Bulletin, 51(1):??, 1952. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
juvenile-forms-neothunnus-macropterus-katsuwonus-pelamis-and-
euthynnus-yaito-philippine;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb51.13.pdf.

Wade:1952:NPF

[1671] Charles B. Wade. Notes on the Philippine frigate mackerels, family Thunnidae, genus Auxis. Fishery Bulletin, 51(1):??, 1952. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/notes-philippine-frigate-
mackerels-family-thunnidae-genus-auxis;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb51.6.pdf.

Mosher:1952:DPD

[1672] Kenneth H. Mosher. Description of a projection device for use in age determination from fish scales. Fishery Bulletin, 51(1):??, 1952. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/description-
projection-device-use-age-determination-fish-scales;   https:

//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb51.14.pdf.

Marr:1952:DBD

[1673] John C. Marr and Milner B. Schaefer. Definitions of body dimensions used in describing tunas. Fishery Bulletin, 51(1):??, 1952. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/definitions-body-dimensions-used-
describing-tunas;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb51.7.pdf.

Wade:1952:OSP

[1674] Charles B. Wade. Observations on the spawning of Philippine tuna. Fishery Bulletin, 51(1):??, 1952. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/observations-spawning-philippine-tuna;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb51.15.pdf.

Eckles:1952:OJO

[1675] Howard H. Eckles. Observations on juvenile oceanic skipjack (Katsuwonus pelamis) from Hawaiian waters and sierra mackerel (Scomberomorus sierra) from the eastern Pacific. Fishery Bulletin, 51(1):245–250, 1952. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/observations-
juvenile-oceanic-skipjack-katsuwonus-pelamis-hawaiian-waters-
and-sierra;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb51.8.pdf.

Schaefer:1952:BCB

[1676] Milner B. Schaefer and Lionel A. Walford. Biometric comparison between yellowfin tunas (Neothunnus) of Angola and of the Pacific coast of Central America. Fishery Bulletin, 51(1):??, 1952. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/biometric-comparison-
between-yellowfin-tunas-neothunnus-angola-and-pacific-coast-
central;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb51.16.pdf.

Davis:1952:SPP

[1677] H. S. Davis. Studies of the protozoan parasites of fresh-water fishes. Fishery Bulletin, 51(1):??, 1952. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/studies-protozoan-parasites-fresh-water-fishes;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb51.1.pdf.

Sette:1952:BAM

[1678] Oscar Elton Sette. Biology of the Atlantic mackerel (Scomber scrombrus) of North America. Part II — migrations and habits. Fishery Bulletin, 51 (1):??, 1952. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/biology-
atlantic-mackerel-scomber-scrombrus-north-america-part-ii-migrations-
and-habits;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb51.9.pdf.

Wade:1952:LTT

[1679] Charles B. Wade. Larvae of tuna and tuna-like fishes from Philippine waters. Fishery Bulletin, 51(1):??, 1952. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/larvae-tuna-and-tuna-fishes-philippine-waters;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb51.17.pdf.

Loosanoff:1952:EDC

[1680] V. L. Loosanoff and James B. Engle. Effect of different concentrations of micro-organisms on the feeding of oysters (O. virginica). Fishery Bulletin, 51(1):??, 1952. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/effect-different-concentrations-
micro-organisms-feeding-oysters-o-virginica;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb51.
2.pdf.

Shuman:1952:ESP

[1681] Richard F. Shuman. On the effectiveness of spermatozoa of the pink salmon (Oncorhynchus gorbuscha) as varying distances from point of dispersal. Fishery Bulletin, 51(1):??, 1952. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/effectiveness-spermatozoa-pink-
salmon-oncorhynchus-gorbuscha-varying-distances-point;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb51.10.pdf.

Anonymous:1952:I

[1682] Anonymous. Index. Fishery Bulletin, 51(1):??, 1952. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-89;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb51.index_.pdf.

Hile:1951:DLT

[1683] Ralph Hile, Paul H. Eschmeyer, and George F. Lunger. Decline of the lake trout fishery in Lake Michigan. Fishery Bulletin, 52(1):77–95, 1951–1953. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/decline-lake-trout-fishery-lake-michigan;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb52.3.pdf.

Griffin:1951:UBP

[1684] Philip J. Griffin and Stanislas F. Snieszko. A unique bacterium pathogenic for warm-blooded and cold-blooded animals. Fishery Bulletin, 52(1):??, 1951–1953. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
unique-bacterium-pathogenic-warm-blooded-and-cold-blooded-animals;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb52.11.pdf.

Burner:1951:CSN

[1685] Clifford J. Burner. Characteristics of spawning nests of Columbia River salmon. Fishery Bulletin, 52(1):??, 1951–1953. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/characteristics-spawning-nests-
columbia-river-salmon;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb52.4.pdf.

Schaefer:1951:ESA

[1686] Milner B. Schaefer. Estimation of size of animal populations by marking experiments. Fishery Bulletin, 52(1):??, 1951–1953. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/estimation-size-animal-populations-marking-
experiments;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb52.12.pdf.

Shimada:1951:CBT

[1687] Bell M. Shimada. Contributions to the biology of tunas from the western equatorial Pacific. Fishery Bulletin, 52(1):??, 1951– 1953. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/contributions-
biology-tunas-western-equatorial-pacific;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb52.5.pdf.

Jobes:1951:AGP

[1688] Frank W. Jobes. Age, growth, and production of yellow perch in Lake Erie. Fishery Bulletin, 52(1):??, 1951–1953. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/age-growth-and-production-yellow-
perch-lake-erie;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb52.13.pdf.

Mead:1951:PNM

[1689] Giles W. Mead. Postlarval Neothunnus macropterus, Auxis thazard, and Euthynnus lineatus from the Pacific coast of Central America. Fishery Bulletin, 52(1):??, 1951–1953. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/postlarval-neothunnus-macropterus-auxis-thazard-and-
euthynnus-lineatus-pacific-coast;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb52.6.pdf.

Ginsburg:1951:FGP

[1690] Isaac Ginsburg. Flounders of the genus Paralichthys and related genera in American waters. Fishery Bulletin, 52(1):??, 1951–1953. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/flounders-genus-paralichthys-and-
related-genera-american-waters;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb52.14.pdf.

Shimada:1951:JOS

[1691] Bell M. Shimada. Juvenile oceanic skipjack from the Phoenix Islands. Fishery Bulletin, 52(1):??, 1951–1953. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/juvenile-oceanic-skipjack-phoenix-islands;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb52.7.pdf.

Schaefer:1951:CYT

[1692] Milner B. Schaefer. Comparison of yellowfin tuna of Hawaiian waters and of the American west coast. Fishery Bulletin, 52(1):??, 1951– 1953. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/comparison-yellowfin-
tuna-hawaiian-waters-and-american-west-coast;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb52.
15.pdf.

Moore:1951:EAG

[1693] Harvey L. Moore. Estimation of age and growth of yellowfin tuna (Neothunnus macropterus) in Hawaiian waters by size frequencies. Fishery Bulletin, 52(1):??, 1951–1953. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/estimation-age-and-growth-yellowfin-tuna-neothunnus-
macropterus-hawaiian-waters-size;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb52.8.pdf.

Collins:1951:FIO

[1694] Gerald B. Collins. Factors influencing the orientation of migrating anadromous fishes. Fishery Bulletin, 52(1):??, 1951–1953. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/factors-influencing-orientation-
migrating-anadromous-fishes;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb52.16.pdf.

Shimada:1951:ABB

[1695] Bell M. Shimada. An annotated bibliography on the biology of Pacific tunas. Fishery Bulletin, 52(1):??, 1951–1953. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/annotated-bibliography-biology-
pacific-tunas;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb52.1.pdf.

Schuck:1951:SGB

[1696] Howard A. Schuck. Studies of Georges Bank haddock; Part I: Landing by pounds, numbers, and sizes of fish. Fishery Bulletin, 52 (1):??, 1951–1953. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
studies-georges-bank-haddock-part-i-landing-pounds-numbers-
and-sizes-fish;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb52.9.pdf.

Anonymous:1951:I

[1697] Anonymous. Index. Fishery Bulletin, 52(1):??, 1951–1953. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-90;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb52.index_.pdf.

Royce:1951:BHL

[1698] William F. Royce. Breeding habits of lake trout in New York. Fishery Bulletin, 52(1):??, 1951–1953. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/breeding-habits-lake-trout-new-york;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb52.
2.pdf.

Schuck:1951:CHG

[1699] Howard A. Schuck and Edgar L. Arnold, Jr. Comparison of haddock from Georges and Browns Banks. Fishery Bulletin, 52 (1):??, 1951–1953. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
comparison-haddock-georges-and-browns-banks;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb52.
10.pdf.

Anonymous:1953:Ca

[1700] Anonymous. Chimæroids. Fishery Bulletin, 53(1):??, 1953. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/chim%C3%A6roids-0;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb53.7.pdf.

Anonymous:1953:BFPa

[1701] Anonymous. Bony fishes: pages 401-450. Fishery Bulletin, 53(1):??, 1953. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/bony-fishes-
pages-401-450;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb53.15.pdf.

Anonymous:1953:BFPb

[1702] Anonymous. Bony fishes: pages 80 to 100. Fishery Bulletin, 53(1):??, 1953. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/bony-fishes-pages-80-
100;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb53.8.pdf.

Anonymous:1953:BFPc

[1703] Anonymous. Bony fishes: pages 451-500. Fishery Bulletin, 53(1):??, 1953. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/bony-fishes-
pages-451-500;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb53.16.pdf.

Anonymous:1953:Ia

[1704] Anonymous. Introduction. Fishery Bulletin, 53(1):??, 1953. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/introduction-2;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb53.1.pdf.

Anonymous:1953:BFPd

[1705] Anonymous. Bony fishes: pages 101-150. Fishery Bulletin, 53(1):??, 1953. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/bony-fishes-
pages-101-150;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb53.9.pdf.

Anonymous:1953:BFPe

[1706] Anonymous. Bony fishes: pages 501-545. Fishery Bulletin, 53(1):??, 1953. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/bony-fishes-
pages-501-545;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb53.17.pdf.

Anonymous:1953:UK

[1707] Anonymous. Use of the keys. Fishery Bulletin, 53(1):??, 1953. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/use-keys-0;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb53.2.pdf.

Anonymous:1953:BFPf

[1708] Anonymous. Bony fishes: pages 151-200. Fishery Bulletin, 53(1):??, 1953. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/bony-fishes-

pages-151-200;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb53.10.pdf.

Anonymous:1953:B

[1709] Anonymous. Bibliography. Fishery Bulletin, 53(1):??, 1953. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/bibliography-2;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb53.18.pdf.

Anonymous:1953:KGM

[1710] Anonymous. Key to Gulf of Maine fishes. Fishery Bulletin,

53(1):??, 1953.  CODEN FSYBAY.  ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
key-gulf-maine-fishes-0;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb53.3.pdf.

Anonymous:1953:BFPg

[1711] Anonymous. Bony fishes: pages 201-250. Fishery Bulletin, 53(1):??, 1953. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/bony-fishes-
pages-201-250;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb53.11.pdf.

Anonymous:1953:ICN

[1712] Anonymous. Index to common names. Fishery Bulletin, 53(1):??, 1953. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/index-common-
names;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb53.19.pdf.

Anonymous:1953:Cb

[1713] Anonymous. Cylcostomes. Fishery Bulletin, 53(1):??, 1953. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/cylcostomes;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb53.4.pdf.

Anonymous:1953:BFPh

[1714] Anonymous. Bony fishes: pages 251-300. Fishery Bulletin, 53(1):??, 1953. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/bony-fishes-
pages-251-300;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb53.12.pdf.

Anonymous:1953:ISN

[1715] Anonymous. Index to scientific names. Fishery Bulletin, 53(1):??, 1953. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/index-scientific-names;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb53.20.pdf.

Anonymous:1953:S

[1716] Anonymous. Sharks. Fishery Bulletin, 53(1):??, 1953. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/sharks;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb53.5.pdf.

Anonymous:1953:BFPi

[1717] Anonymous. Bony fishes: pages 301-350. Fishery Bulletin, 53(1):??, 1953. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/bony-fishes-
pages-301-350;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb53.13.pdf.

Anonymous:1953:T

[1718] Anonymous. Torpedoes. Fishery Bulletin, 53(1):??, 1953. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/torpedoes;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb53.
6.pdf.

Anonymous:1953:BFPj

[1719] Anonymous. Bony fishes: pages 351-400. Fishery Bulletin, 53(1):??, 1953. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/bony-fishes-
pages-351-400;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb53.14.pdf.

Hile:1953:FFS

[1720] Ralph Hile, George F. Lunger, and Howard J. Buettner. Fluctuations in the fisheries of State of Michican waters of Green Bay. Fishery Bulletin, 54(1):1–34, 1953–1954. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fluctuactions-fisheries-state-michican-waters-green-bay;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb54.1.pdf.

Taylor:1953:NVT

[1721] Clyde C. Taylor. Nature of variability in trawl catches. Fishery Bulletin, 54(1):??, 1953–1954. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/nature-variability-trawl-catches;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb54.9.pdf.

Rucker:1953:CDS

[1722] R. R. Rucker, W. J. Whipple, J. R. Parvin, and C. A. Evans. A contagious disease of salmon possibly of virus origin. Fishery Bulletin, 54(1):??, 1953–1954. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
contagious-disease-salmon-possibly-virus-origin;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb54.2.pdf.

Collier:1953:EDO

[1723] Albert Collier, S. M. Ray, A. W. Magnitzky, and Joe O. Bell. Effect of dissolved organic substances on oysters. Fishery Bulletin, 54(1):??, 1953–1954. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
effect-dissolved-organic-substances-oysters;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb54.
10.pdf.

June:1953:SYT

[1724] Fred C. June. Spawning of yellowfin tuna in Hawaiian waters. Fishery Bulletin, 54(1):??, 1953–1954. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/spawning-yellowfin-tuna-hawaiian-waters;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb54.3.pdf.

Cating:1953:DAA

[1725] James P. Cating. Determining age of Atlantic shad from their scales. Fishery Bulletin, 54(1):??, 1953–1954. CODEN FSY-

BAY.  ISSN 0090-0656 (print), 1937-4518 (electronic).  URL

https://spo.nmfs.noaa.gov/content/determining-age-atlantic-
shad-their-scales;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb54.11.pdf.

Lindner:1953:EGR

[1726] Milton J. Lindner. Estimation of growth rate in animals by marking experiments. Fishery Bulletin, 54(1):??, 1953–1954. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/estimation-growth-rate-animals-
marking-experiments;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb54.4.pdf.

Felin:1953:PHP

[1727] Frances E. Felin. Population heterogeneity in the Pacific pilchard. Fishery Bulletin, 54(1):??, 1953–1954. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/population-heterogeneity-pacific-pilchard;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb54.12.pdf.

Collier:1953:SPR

[1728] Albert W. Collier and Kenneth T. Marvin. Stabilization of the phosphate ratio of sea water by freezing. Fishery Bulletin, 54 (1):??, 1953–1954. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
stabilization-phosphate-ratio-sea-water-freezing;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb54.5.pdf.

Rice:1953:BIA

[1729] Theodore R. Rice. Biotic influences affecting population growth of planktonic algae. Fishery Bulletin, 54(1):??, 1953–1954. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/biotic-influences-affecting-population-
growth-planktonic-algae;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb54.13.pdf.

Rice:1953:PEM

[1730] Theodore R. Rice. Phosphorus exchange in marine phytoplankton. Fishery Bulletin, 54(1):??, 1953–1954. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/phosphorus-exchange-marine-phytoplankton;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb54.6.pdf.

Fredin:1953:CFA

[1731] Reynold A. Fredin. Causes of fluctuations in abundance of Connecticut river shad. Fishery Bulletin, 54(1):??, 1953–1954. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/causes-fluctuations-
abundance-connecticut-river-shad;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb54.14.pdf.

Reintjes:1953:FYT

[1732] John W. Reintjes and Joseph E. King. Food of yellowfin tuna in the central Pacific. Fishery Bulletin, 54(1):??, 1953–1954. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/food-yellowfin-tuna-
central-pacific;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb54.7.pdf.

Anonymous:1953:Ib

[1733] Anonymous. Index. Fishery Bulletin, 54(1):??, 1953–1954. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-91;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb54.index_.pdf.

King:1953:ZAC

[1734] Joseph King and Joan Demond. Zooplankton abundance in the central Pacific. Fishery Bulletin, 54(1):??, 1953–1954. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/zooplankton-abundance-central-pacific;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb54.8.pdf.

Anonymous:1954:HSE

[1735] Anonymous. Historical sketch of the explorations in the Gulf of Mexico. Fishery Bulletin, 55(1):??, 1954. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/historical-sketch-explorations-
gulf-mexico;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb55.2.pdf.

Anonymous:1954:FLF

[1736] Anonymous. Free-living flatworms, nemerteans, nematodes, tardigrades, and chaetognaths. Fishery Bulletin, 55(1):??, 1954. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/free-living-flatworms-nemerteans-
nematodes-tardigrades-and-chaetognaths;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb55.10.
pdf.

Anonymous:1954:FST

[1737] Anonymous. Fishes and sea turtles. Fishery Bulletin, 55(1): ??, 1954. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishes-
and-sea-turtles;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb55.18.pdf.

Anonymous:1954:G

[1738] Anonymous. Geology. Fishery Bulletin, 55(1):??, 1954. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/geology;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb55.3.pdf.

Anonymous:1954:PWa

[1739] Anonymous. Parasitic worms. Fishery Bulletin, 55(1):??, 1954. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/parasitic-worms;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb55.
11.pdf.

Anonymous:1954:BGM

[1740] Anonymous. The birds of the Gulf of Mexico. Fishery Bulletin, 55(1):??, 1954. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/birds-gulf-mexico;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb55.19.pdf.

Anonymous:1954:MMG

[1741] Anonymous. Marine meteorology of the Gulf of Mexico. Fishery Bulletin, 55(1):??, 1954. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/marine-
meteorology-gulf-mexico;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb55.4.pdf.

Anonymous:1954:BBP

[1742] Anonymous. Bryozoa, Brachiopoda, Phoronida, and Enteropneusta. Fishery Bulletin, 55(1):??, 1954. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/bryozoa-brachiopoda-phoronida-and-enteropneusta;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb55.12.pdf.

Anonymous:1954:MGM

[1743] Anonymous. Mammals of the Gulf of Mexico. Fishery Bulletin, 55(1):??, 1954. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/mammals-gulf-mexico;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb55.20.pdf.

Anonymous:1954:PCG

[1744] Anonymous. Physics and chemistry of Gulf waters. Fishery Bulletin, 55(1):??, 1954. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
physics-and-chemistry-gulf-waters;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb55.5.pdf.

Anonymous:1954:E

[1745] Anonymous. Echinoderm. Fishery Bulletin, 55(1):??, 1954. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/echinoderm;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb55.13.
pdf.

Anonymous:1954:PWb

[1746] Anonymous. Pollution of water. Fishery Bulletin, 55(1):??, 1954. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/pollution-water;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb55.21.pdf.

Anonymous:1954:PAC

[1747] Anonymous. Plant and animal communities. Fishery Bulletin, 55 (1):??, 1954. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/plant-and-
animal-communities;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb55.6.pdf.

Anonymous:1954:AMW

[1748] Anonymous. Annelids and miscellaneous worms. Fishery Bulletin, 55 (1):??, 1954. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/annelids-
and-miscellaneous-worms;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb55.14.pdf.

Anonymous:1954:Ia

[1749] Anonymous. Index. Fishery Bulletin, 55(1):??, 1954. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-92;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb55.index_.pdf.

Anonymous:1954:BFU

[1750] Anonymous.  Bacteria, fungi, and unicellular algae.  Fishery

Bulletin, 55(1):??, 1954. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/bacteria-fungi-and-unicellular-algae;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb55.7.pdf.

Anonymous:1954:AXP

[1751] Anonymous. Arthropods: Xiphosura, Pycnogonida, and Crustacea. Fishery Bulletin, 55(1):??, 1954. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/arthropods-xiphosura-pycnogonida-and-crustacea;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb55.15.pdf.

Anonymous:1954:Pa

[1752] Anonymous. Protozoa. Fishery Bulletin, 55(1):??, 1954. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/protozoa;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb55.8.pdf.

Anonymous:1954:M

[1753] Anonymous. Mollusks. Fishery Bulletin, 55(1):??, 1954. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/mollusks;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb55.16.pdf.

Anonymous:1954:Pb

[1754] Anonymous. Preface. Fishery Bulletin, 55(1):??, 1954. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/preface-7;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb55.
1.pdf.

Anonymous:1954:SCC

[1755] Anonymous. Sponges, coelenterates, ctenophores. Fishery Bulletin, 55(1):??, 1954. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
sponges-coelenterates-ctenophores;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb55.9.pdf.

Anonymous:1954:TL

[1756] Anonymous. Tunicates and lancelets. Fishery Bulletin, 55(1):??, 1954. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/tunicates-and-
lancelets;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb55.17.pdf.

Collins:1954:MSF

[1757] Gerald B. Collins, Charles D. Volz, and Parker S. Trefethen. Mortality of salmon fingerlings exposed to pulsating direct current. Fishery Bulletin, 56(1):61–81, 1954–1956. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/mortality-salmon-fingerlings-exposed-pulsating-
direct-current;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb56.3.pdf.

Elling:1954:PST

[1758] Carl H. Elling and Paul T. Macy. Pink salmon tagging experiments in Icy Strait and Upper Chatham Strait, 1950. Fishery Bulletin, 56(1):??, 1954–1956. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/pink-salmon-tagging-experiments-icy-strait-and-
upper-chatham-strait-1950;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb56.11.pdf.

Ahlstrom:1954:DAE

[1759] Elbert H. Ahlstrom. Distribution and abundance of egg and larval populations of the Pacific sardine. Fishery Bulletin, 56(1):??, 1954– 1956. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/distribution-
and-abundance-egg-and-larval-populations-pacific-sardine;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fb56.4.pdf.

Talbot:1954:FAF

[1760] Gerald B. Talbot. Factors associated with fluctuations in abundance of Hudson River shad. Fishery Bulletin, 56(1):??, 1954–1956. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/factors-associated-fluctuations-
abundance-hudson-river-shad;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb56.12.pdf.

Widrig:1954:MEF

[1761] T. M. Widrig. Method of estimating fish populations, with application to Pacific sardine. Fishery Bulletin, 56(1):??, 1954– 1956. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/method-
estimating-fish-populations-application-pacific-sardine;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb56.5.pdf.

Nelson:1954:LEF

[1762] Philip R. Nelson and W. T. Edmondson. Limnological effects of fertilizing Bare Lake, Alaska. Fishery Bulletin, 56(1):??, 1954–1956. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/limnological-effects-fertilizing-
bare-lake-alaska;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb56.13.pdf.

Hanavan:1954:ISP

[1763] Mitchell G. Hanavan and Bernard Einar Skud. Intertidal spawning of pink salmon. Fishery Bulletin, 56(1):??, 1954–1956. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/intertidal-spawning-pink-salmon;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb56.6.pdf.

Reimers:1954:LSL

[1764] Norman Reimers, John A. Maciolek, and Edwin P. Pister. Limnological study of the lakes in Convict Creek Basin, Mono County, California. Fishery Bulletin, 56(1):??, 1954–1956. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/limnological-study-lakes-convict-
creek-basin-mono-county-california;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb56.14.pdf.

Ginsburg:1954:WCA

[1765] Isaac Ginsburg. Whitings on the coasts of the American continents. Fishery Bulletin, 56(1):??, 1954–1956. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/whitings-coasts-american-continents;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb56.
7.pdf.

OConnell:1954:GBR

[1766] Charles P. O’Connell. The gas bladder and its relation to the inner ear in Sardinops caerulea and Engraulis mordax. Fishery Bulletin, 56(1):??, 1954–1956. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/gas-bladder-
and-its-relation-inner-ear-sardinops-caerulea-and-engraulis-
mordax;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb56.15.pdf.

Ahlstrom:1954:DEL

[1767] Elbert H. Ahlstrom and Orville P. Ball. Description of eggs and larvae of jack mackerel (Trachurus symmetricus) and distribution and abundance of larvae in 1950 and 1951. Fishery Bulletin, 56(1):??, 1954–1956. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/description-eggs-and-larvae-jack-mackerel-
trachurus-symmetricus-and-distribution-and;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fb56.8.pdf.

Reimers:1954:MET

[1768] Norman Reimers and Bobby D. Combs. Method of evaluating temperature in lakes, with description of thermal characteristics of Convict Lake, California. Fishery Bulletin, 56(1):??, 1954– 1956. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/method-
evaluating-temperature-lakes-description-thermal-characteristics-
convict-lake;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb56.16.pdf.

Royce:1954:SGB

[1769] William F. Royce and Howard A. Schuck. Study of Georges Bank haddock. Part II: Prediction of the catch. Fishery Bulletin, 56(1):??, 1954– 1956. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/study-georges-bank-
haddock-part-ii-prediction-catch;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb56.1.pdf.

Lennon:1954:FMS

[1770] Robert E. Lennon. Feeding mechanism of the sea lamprey and its effect on host fishes. Fishery Bulletin, 56(1):??, 1954–1956. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/feeding-mechanism-sea-lamprey-and-its-
effect-host-fishes;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb56.9.pdf.

Lindner:1954:GMS

[1771] Milton J. Lindner and William W. Anderson. Growth, migrations, spawning and size distribution of shrimp Penaeus setiferus. Fishery Bulletin, 56(1):??, 1954–1956. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
growth-migrations-spawning-and-size-distribution-shrimp-penaeus-
setiferus;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fb56.17.pdf.

Hile:1954:FGY

[1772] Ralph Hile. Fluctuations in growth and year-class strength of the walleye in Saginaw Bay. Fishery Bulletin, 56(1):??, 1954–1956. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/fluctuations-growth-
and-year-class-strength-walleye-saginaw-bay;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb56.
2.pdf.

Ahlstrom:1954:ELP

[1773] Elbert H. Ahlstrom and Robert C. Counts. Eggs and larvae of the Pacific hake, Merluccius productus. Fishery Bulletin, 56 (1):??, 1954–1956. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
eggs-and-larvae-pacific-hake-merluccius-productus;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb56.10.pdf.

Anonymous:1954:Ib

[1774] Anonymous. Index. Fishery Bulletin, 56(1):??, 1954–1956. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-93;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb56.index_.pdf.

Smith:1956:LHL

[1775] Stanford H. Smith. Life history of lakes herring of Green Bay, Lake Michigan. Fishery Bulletin, 57(1):??, 1956–1958. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/life-history-lakes-herring-green-bay-lake-
michigan;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb57.3.pdf.

Fleminger:1956:NCC

[1776] Abraham Fleminger. New calanoid copepods of the families aetideidae, euchaetidae, and stephidae from the Gulf of Mexico. Fishery Bulletin, 57(1):??, 1956–1958. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/new-calanoid-copepods-families-aetideidae-
euchaetidae-and-stephidae-gulf-mexico;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb57.11.pdf.

Snieszko:1956:TSR

[1777] S. F. Snieszko and G. L. Bullock. Treatment of sulfonamide-resistant furunculosis in trout and determination of drug sensitivity. Fishery Bulletin, 57(1):??, 1956–1958. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/treatment-sulfonamide-resistant-furunculosis-trout-
and-determination-drug-sensitivity;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb57.19.pdf.

Gehringer:1956:ODA

[1778] Jack W. Gehringer. Observations on the development of the Atlantic sailfish Istiophorus americanus (Cuvier), with notes on an unidentified species of istiophorid. Fishery Bulletin, 57(1):??, 1956–1958. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/observations-development-
atlantic-sailfish-istiophorus-americanus-cuvier-notes-unidentified;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb57.4.pdf.

King:1956:ZAC

[1779] Joseph E. King and Thomas S. Hida. Zooplankton abundance in the central Pacific, part II. Fishery Bulletin, 57(1):??, 1956– 1958. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/zooplankton-
abundance-central-pacific-part-ii;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fb57.12.pdf.

Anonymous:1956:I

[1780] Anonymous. Index. Fishery Bulletin, 57(1):??, 1956–1958. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-94;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb57.index_.pdf.

VanCampen:1956:TTF

[1781] Wilvan G. Van Campen and Earl E. Hoven. Tunas and tuna fisheries of the world: an annotated bibliography, 1930–53. Fishery Bulletin, 57(1):??, 1956–1958. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
tunas-and-tuna-fisheries-world-annotated-bibliography-1930-
53;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb57.5.pdf.

Anderson:1956:EDS

[1782] William W. Anderson. Early development, spawning, growth, and occurrence of the silver mullet (Mugil curema) along the south Atlantic coast of the United States. Fishery Bulletin, 57(1):??, 1956–1958. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/early-
development-spawning-growth-and-occurrence-silver-mullet-mugil-
curema-along-south;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb57.13.pdf.

Yuen:1956:YTS

[1783] Heeny S. H. Yuen and Fred C. June. Yellowfin tuna spawning in the central equatorial Pacific. Fishery Bulletin, 57(1):??, 1956–1958. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/yellowfin-tuna-spawning-central-
equatorial-pacific;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb57.6.pdf.

Anderson:1956:LFF

[1784] William W. Anderson. Larval forms of the fresh-water mullet (Agonostomus monticola) from the open ocean off the Bahamas and south Atlantic coast of the United States. Fishery Bulletin, 57 (1):??, 1956–1958. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
larval-forms-fresh-water-mullet-agonostomus-monticola-open-
ocean-bahamas-and-south-atlantic;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fb57.14.pdf.

Rounsefell:1956:MEA

[1785] George A. Rounsefell. A method of estimating abundance of groundfish on Georges Bank. Fishery Bulletin, 57(1):??, 1956– 1958. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/method-
estimating-abundance-groundfish-georges-bank;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb57.
7.pdf.

MacGregor:1956:FPS

[1786] John S. MacGregor. Fecundity of the Pacific sardine (Sardinops caerulea). Fishery Bulletin, 57(1):??, 1956–1958. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/fecundity-pacific-sardine-sardinops-caerulea;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb57.15.pdf.

Spear:1956:EEH

[1787] Harlan S. Spear and John B. Glude. Effects of environment and heredity on growth of the soft clam (Mya arenaria). Fishery Bulletin, 57(1): ??, 1956–1958. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/effects-
environment-and-heredity-growth-soft-clam-mya-arenaria;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb57.8.pdf.

Rounsefell:1956:FNA

[1788] George A. Rounsefell. Fecundity of North American salmonidae. Fishery Bulletin, 57(1):??, 1956–1958. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fecundity-north-american-salmonidae;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fb57.
16.pdf.

Cable:1956:VAD

[1789] Louella E. Cable. Validity of age determination from scales, and growth of marked Lake Michigan lake trout. Fishery Bulletin, 57(1):??, 1956– 1958. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/validity-age-
determination-scales-and-growth-marked-lake-michigan-lake-trout;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb57.1.pdf.

Taylor:1956:CTD

[1790] Clyde C. Taylor, Henry B. Bigelow, and Herbert W. Graham. Climatic trends and the distribution of marine animals in New England. Fishery Bulletin, 57(1):293–345, 1956–1958. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/climatic-trends-and-distribution-
marine-animals-new-england;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fb57.9.pdf.

Ray:1956:EUB

[1791] Sammy M. Ray and William B. Wilson. Effects of unialgal and bacteriafree cultures of gymnodinium brevis on fish. Fishery Bulletin, 57(1): ??, 1956–1958. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/effects-
unialgal-and-bacteria-free-cultures-gymnodinium-brevis-fish;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fb57.17.pdf.

King:1956:CSF

[1792] Joseph E. King and Isaac I. Ikehara. Comparative study of food of bigeye and yellowfin tuna in the central Pacific. Fishery Bulletin, 57(1):??, 1956–1958. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/comparative-
study-food-bigeye-and-yellowfin-tuna-central-pacific;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fb57.2.pdf.

Fleminger:1956:NGT

[1793] Abraham Fleminger. New genus and two new species of Tharybidae (Copepoda calanoida) from the Gulf of Mexico with remarks on the status of the family. Fishery Bulletin, 57(1):??, 1956– 1958. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/new-genus-and-

two-new-species-tharybidae-copepoda-calanoida-gulf-mexico-remarks-
status;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fb57.10.pdf.

Royce:1956:OSC

[1794] William F. Royce. Observations on the spearfishes of the central Pacific. Fishery Bulletin, 57(1):??, 1956–1958. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/observations-spearfishes-
central-pacific;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fb57.18.pdf.

Yamashita:1958:ACS

[1795] Daniel T. Yamashita. Analysis of catch statistics of the Hawaiian skipjack fishery. Fishery Bulletin, 58(1):??, 1958. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/analysis-catch-statistics-hawaiian-skipjack-fishery;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/yamashita_0.pdf.

Chipman:1958:UAR

[1796] Walter A. Chipman, Theodore R. Rice, and Thomas J. Price. Uptake and accumulation of radioactive zinc by marine plankton, fish, and shellfish. Fishery Bulletin, 58(1):279–292, 1958. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/uptake-and-accumulation-radioactive-zinc-marine-
plankton-fish-and-shellfish;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/chipman_0.pdf.

Nichols:1958:ENJ

[1797] Paul R. Nichols. Effect of New Jersey–New York pound-net catches on shad runs of Hudson and Connecticut rivers. Fishery Bulletin, 58 (1):??, 1958. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/effect-
new-jersey-new-york-pound-net-catches-shad-runs-hudson-and-
connecticut-rivers;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/nichols_0.pdf.

Ahlstrom:1958:HSP

[1798] Elbert H. Ahlstrom, John D. Isaacs, James R. Thrailkill, and Lewis W. Kidd. High-speed plankton sampler. Fishery Bulletin, 58(1):??, 1958. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/high-speed-plankton-
sampler;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/ahlstrom_1.pdf.

Silliman:1958:EEF

[1799] Ralph P. Silliman, James S. Gutsell, Clarence E. Dunbar, and Saufley B. Friddle. Experimental exploitation of fish populations. Fishery Bulletin, 58(1):??, 1958. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
experimental-exploitation-fish-populations;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/silliman_
0.pdf.

Taylor:1958:NMR

[1800] Clyde C. Taylor. Natural mortality rate of Georges Bank haddock. Fishery Bulletin, 58(1):??, 1958. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
natural-mortality-rate-georges-bank-haddock;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/taylor_
1.pdf.

Robertson:1958:ADT

[1801] O. H. Robertson. Accelerated development of testis after unilateral gonadectomy, with observations on normal testis of rainbow trout. Fishery Bulletin, 58(1):??, 1958. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/accelerated-development-testis-after-
unilateral-gonadectomy-observations-normal-testis;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
robertson_0.pdf.

Matsumoto:1958:DDL

[1802] Walter M. Matsumoto. Description and distribution of larvae of four species of tune in central Pacific waters. Fishery Bulletin, 58(1):??, 1958. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/description-
and-distribution-larvae-four-species-tune-central-pacific-waters;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/matsumoto_0.pdf.

Rice:1958:FRH

[1803] Theordore R. Rice and Rebecca J. Smith. Filtering rates of the hard clam (Venus mercenaria) determined with radioactive phytoplankton. Fishery Bulletin, 58(1):??, 1958. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/filtering-rates-hard-clam-venus-
mercenaria-determined-radioactive-phytoplankton;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
rice_0.pdf.

Rounsefell:1958:FCD

[1804] George A. Rounsefell. Factors causing decline in sockeye salmon of Karluk River, Alaska. Fishery Bulletin, 58(1):??, 1958. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/factors-causing-decline-sockeye-
salmon-karluk-river-alaska;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/rounsefell_1.pdf.

Rounsefell:1958:ANA

[1805] George A. Rounsefell. Anadromy in North American salmonidae. Fishery Bulletin, 58(1):??, 1958. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
anadromy-north-american-salmonidae;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/rounsefell1_0.
pdf.

Davis:1958:RVT

[1806] Harry C. Davis and Robert R. Guillard. Relative value of ten genera of micro-organisms as food for oyster and clam larvae. Fishery Bulletin, 58(1):??, 1958. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/relative-value-ten-genera-micro-organisms-food-
oyster-and-clam-larvae;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/davis_1.pdf.

Bradley:1958:LAP

[1807] W. H. Bradley and Peter Cooke. Living and ancient populations of the clam Gemma gemma in a Maine coast tidal flat. Fishery Bulletin, 58 (1):??, 1958. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/living-
and-ancient-populations-clam-gemma-gemma-maine-coast-tidal-
flat;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/bradley_0.pdf.

Strasburg:1958:DAH

[1808] Donald W. Strasburg. Distribution, abundance, and habits of pelagic sharks in the central Pacific Ocean. Fishery Bulletin, 58(1):??, 1958. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/distribution-
abundance-and-habits-pelagic-sharks-central-pacific-ocean;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/strasburg_1.pdf.

Ahlstrom:1958:DDV

[1809] Elbert H. Ahlstrom and Robert C. Counts. Development and distribution of vinciguerria lucetia and related species in the eastern Pacific. Fishery Bulletin, 58(1):??, 1958. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/development-and-distribution-vinciguerria-
lucetia-and-related-species-eastern-pacific;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/ahlstrom1_
1.pdf.

Cope:1958:ABC

[1810] Oliver B. Cope. Annotated bibliography on the cutthroat trout. Fishery Bulletin, 58(1):??, 1958. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
annotated-bibliography-cutthroat-trout;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/cope_0.pdf.

Day:1958:SCG

[1811] C. Godfrey Day. Surface circulation in the Gulf of Maine as deduced from drift bottles. Fishery Bulletin, 58(1):??, 1958. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/surface-circulation-gulf-maine-
deduced-drift-bottles;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/day_0.pdf.

Talbot:1958:ACM

[1812] Gerald B. Talbot and James E. Sykes. Atlantic coast migrations of American shad. Fishery Bulletin, 58(1):??, 1958. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/atlantic-coast-migrations-american-shad;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/talbot_0.pdf.

Anderson:1958:LDG

[1813] William W. Anderson. Larval development, growth, and spawning of striped mullet (Mugil cephalus) along the south Atlantic coast of the United States. Fishery Bulletin, 58(1):??, 1958. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/larval-development-
growth-and-spawning-striped-mullet-mugil-cephalus-along-south-
atlantic;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/anderson_0.pdf.

Anonymous:1959:Ia

[1814] Anonymous. Index. Fishery Bulletin, 59(1):??, 1959. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-99;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
index_4.pdf.

Young:1959:MWS

[1815] Joseph H. Young. Morphology of the white shrimp, Penaeus setiferus, (Linnaeus 1758). Fishery Bulletin, 59(1):??, 1959. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/morphology-white-shrimp-penaeus-
setiferus-linnaeus-1758;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/young.pdf.

Royce:1959:DYF

[1816] William F. Royce, Raymond J. Buller, and Ernest D. Premetz. Decline of the yellowtail flounder (Limanda ferruginea) off New England. Fishery Bulletin, 59(1):169–267, 1959. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/decline-yellowtail-flounder-limanda-ferruginea-new-
england;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/royce.pdf.

Hill:1959:SUS

[1817] Donald R. Hill. Some uses of statistical analysis in classifying races of American shad (Alosa sapidissima). Fishery Bulletin, 59 (1):??, 1959. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/some-uses-
statistical-analysis-classifying-races-american-shad-alosa-
sapidissima;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/hill.pdf.

Otsu:1959:SMS

[1818] Tamio Otsu and Richard N. Uchida. Sexual maturity and spawning of albacore in the Pacific Ocean. Fishery Bulletin, 59(1):??, 1959. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/sexual-maturity-and-
spawning-albacore-pacific-ocean;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/otsu_0.pdf.

Kruse:1959:GGL

[1819] Thomas E. Kruse. Grayling of Grebe Lake, Yellowstone National Park, Wyoming. Fishery Bulletin, 59(1):??, 1959. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/grayling-grebe-lake-
yellowstone-national-park-wyoming;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/kruse.pdf.

Otsu:1959:SAD

[1820] Tamio Otsu and Richard N. Uchida. Study of age determination by hard parts of albacore from central north Pacific and Hawaiian waters. Fishery Bulletin, 59(1):??, 1959. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
study-age-determination-hard-parts-albacore-central-north-pacific-
and-hawaiian-waters;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/otsu1.pdf.

El-Zarka:1959:FPY

[1821] Salah El-Din El-Zarka. Fluctuations in the population of yellow perch, Perca flavescens (Mitchill), in Saginaw Bay, Lake Huron. Fishery Bulletin, 59(1):??, 1959. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fluctuations-population-yellow-perch-perca-flavescens-mitchill-
saginaw-bay-lake-huron;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/elzarka.pdf.

Berry:1959:YJC

[1822] Frederick H. Berry. Young jack crevalles (Caranx species) off the southeastern Atlantic coast of the United States. Fishery Bulletin, 59(1):??, 1959. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/young-jack-crevalles-
caranx-species-southeastern-atlantic-coast-united-states;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/berry_0.pdf.

Greenbank:1959:LHT

[1823] John Greenbank and Philip R. Nelson. Life history of the threespine stickleback, Gasterosteus aculeatus Linnaeus, in Karluk Lake and Bare Lake, Kodiak Island, Alaska. Fishery Bulletin, 59(1):??, 1959. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/life-history-
threespine-stickleback-gasterosteus-aculeatus-linnaeus-karluk-
lake-and-bare;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/greenbank.pdf.

Wigley:1959:LHS

[1824] Roland L. Wigley. Life history of the sea lamprey of Cayuga Lake, New York. Fishery Bulletin, 59(1):??, 1959. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/life-history-sea-lamprey-cayuga-lake-
new-york;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/wigley.pdf.

Gehringer:1959:EDM

[1825] Jack W. Gehringer. Early development and metamorphosis of the ten-pounder, Elops saurus Linnaeus. Fishery Bulletin, 59(1): ??, 1959. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/early-
development-and-metamorphosis-ten-pounder-elops-saurus-linnaeus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/gehringer.pdf.

Anonymous:1959:Ib

[1826] Anonymous. Index. Fishery Bulletin, 60(1):??, 1959–1960. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-96;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
index_0.pdf.

Kelly:1959:AGR

[1827] George F. Kelly and Robert S. Wolf. Age and growth of the redfish (Sebastes marinus) in the Gulf of Maine. Fishery Bulletin, 60(1):??, 1959–1960. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/age-and-growth-redfish-sebastes-marinus-gulf-maine;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/kelly.pdf.

Grosslein:1959:GAA

[1828] Marvin D. Grosslein and Llyod L. Smith, Jr. The goldeye, Amphiodon alosiodes (Rafinesque), in the commercial fishery of the Red Lakes, Minnesota. Fishery Bulletin, 60(1):??, 1959– 1960. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/goldeye-
amphiodon-alosiodes-rafinesque-commercial-fishery-red-lakes-
minnesota;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/grosslein.pdf.

Thompson:1959:FSP

[1829] Richard B. Thompson. Food of the squarfish, Ptychocheilus oregonenisis (Richardson) of the Columbia River. Fishery Bulletin, 60(1): ??, 1959–1960. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/food-
squarfish-ptychocheilus-oregonenisis-richardson-columbia-river;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/thompson.pdf.

Nelson:1959:EFB

[1830] Philip R. Nelson. Effects of fertilizing Bare Lake, Alaska, on growth and production of red salmon (O. nerka). Fishery Bulletin, 60 (1):??, 1959–1960. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
effects-fertilizing-bare-lake-alaska-growth-and-production-
red-salmon-o-nerka;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/nelson.pdf.

Shomura:1959:CTL

[1831] Richard S. Shomura. Changes in tuna landings of the Hawaiian longline fishery, 1948–1956. Fishery Bulletin, 60(1):??, 1959– 1960. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/changes-
tuna-landings-hawaiian-longline-fishery-1948-1956;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
shomura.pdf.

Ahlstrom:1959:VDP

[1832] Elbert H. Ahlstrom. Vertical distribution of pelagic fish eggs and larvae off California and Baja California. Fishery Bulletin, 60(1):??, 1959–1960. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/vertical-
distribution-pelagic-fish-eggs-and-larvae-california-and-baja-
california;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/ahlstrom.pdf.

Yuen:1959:VSR

[1833] Heeny S. H. Yuen. Variability of skipjack response to live bait. Fishery Bulletin, 60(1):??, 1959–1960. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/variability-skipjack-response-live-bait;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
yuen.pdf.

Fischler:1959:CHC

[1834] Kenneth J. Fischler. Contributions of Hudson and Connecticut Rivers to New York–New Jersey shad catch of 1956. Fishery Bulletin, 60(1):??, 1959–1960. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/contributions-
hudson-and-connecticut-rivers-new-york-new-jersey-shad-catch-
1956;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fischler.pdf.

Hand:1959:FPS

[1835] Cadet H. Hand and Leo Berner, Jr. Food of the Pacific sardine (Sardinops caerulea). Fishery Bulletin, 60(1):??, 1959–1960. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/food-pacific-sardine-sardinops-
caerulea;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/hand.pdf.

Ahlstrom:1959:DAE

[1836] Elbert H. Ahlstrom. Distribution and abundance of eggs of the Pacific sardine 1952–1956. Fishery Bulletin, 60(1):??, 1959–1960. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/distribution-and-abundance-eggs-
pacific-sardine-1952-1956;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/ahlstrom1.pdf.

MacGregor:1959:RBF

[1837] John S. MacGregor. Relation between fish condition and population size in the sardine (Sardinops caerulea). Fishery Bulletin, 60(1):??, 1959–1960. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/relation-between-fish-condition-and-population-
size-sardine-sardinops-caerulea;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/macgregor.pdf.

Strasburg:1959:ELT

[1838] Donald W. Strasburg. Estimates of larval tuna abundance in the central Pacific. Fishery Bulletin, 60(1):??, 1959–1960. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/estimates-larval-tuna-abundance-
central-pacific;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/strasburg.pdf.

Austin:1959:OEC

[1839] Thomas S. Austin. Oceanography of the east central equatorial Pacific as observed during expedition eastropic. Fishery Bulletin, 60(1):??, 1959–1960. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/oceanography-
east-central-equatorial-pacific-observed-during-expedition-
eastropic;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/austin.pdf.

Clark:1959:DHS

[1840] John R. Clark and Vadim D. Vladykov. Definition of haddock stocks of the northwestern Atlantic. Fishery Bulletin, 60(1):??, 1959–1960. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/definition-haddock-
stocks-northwestern-atlantic;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/clark_0.pdf.

Reintjes:1959:ABB

[1841] John W. Reintjes, James Y. Christmas, Jr., and Richard A. Collins. Annotated bibliography on biology of American menhaden. Fishery Bulletin, 60(1):??, 1959–1960. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
annotated-bibliography-biology-american-menhaden;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
reintjes_0.pdf.

June:1959:DAA

[1842] Fred C. June and Charles M. Roithmayr. Determining age of Atlantic menhaden from their scales. Fishery Bulletin, 60(1):??, 1959–1960. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/determining-age-atlantic-
menhaden-their-scales;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/june.pdf.

Wells:1959:SAV

[1843] LaRue Wells. Seasonal abundance and vertical movements of planktonic crustacea in Lake Michigan. Fishery Bulletin, 60(1):??, 1959– 1960. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/seasonal-
abundance-and-vertical-movements-planktonic-crustacea-lake-
michigan;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/wells.pdf.

Miller:1959:SBG

[1844] Robert Rush Miller. Systematic and biology of the gizzard shad (Dorosoma cepedianum) and related fishes. Fishery Bulletin, 60(1):??, 1959– 1960. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/systematic-
and-biology-gizzard-shad-dorosoma-cepedianum-and-related-fishes;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/miller.pdf.

Kramer:1959:DEL

[1845] David Kramer. Development of eggs and larvae of Pacific mackerel and distribution and abundance of larvae 1952–1956. Fishery Bulletin, 60(1):??, 1959–1960. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
development-eggs-and-larvae-pacific-mackerel-and-distribution-
and-abundance-larvae-1952;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/kramer.pdf.

Hoffman:1959:SST

[1846] Glenn L. Hoffman. Synopsis of strigeoidea (Trematoda) of fishes and their life cycles. Fishery Bulletin, 60(1):??, 1959–1960. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/synopsis-strigeoidea-
trematoda-fishes-and-their-life-cycles;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/hoffman.
pdf.

Johnson:1959:STA

[1847] James H. Johnson. Sonic tracking of adult salmon at Bonneville Dam, 1957. Fishery Bulletin, 60(1):??, 1959–1960. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/sonic-tracking-adult-
salmon-bonneville-dam-1957;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/johnson.pdf.

Walburg:1959:ALH

[1848] Charles H. Walburg. Abundance and life history of shad, St. Johns River, Florida. Fishery Bulletin, 60(1):??, 1959–1960. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/abundance-and-life-history-shad-
st-johns-river-florida;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/walburg.pdf.

Anonymous:1960:I

[1849] Anonymous. Index. Fishery Bulletin, 61(1):??, 1960–1962. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-98;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
index_2.pdf.

Springer:1960:NHS

[1850] Stewart Springer. Natural history of the sandbar shark, Eulamia milberti. Fishery Bulletin, 61(1):??, 1960–1962. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/natural-history-sandbar-shark-eulamia-milberti;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/springer.pdf.

Mather:1960:GBT

[1851] Frank J. Mather III and Howard A. Schuck. Growth of bluefun tuna of the western north Atlantic. Fishery Bulletin, 61(1):??, 1960–1962. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/growth-bluefun-tuna-
western-north-atlantic;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/mather.pdf.

Hartman:1960:FRS

[1852] Wilbur L. Hartman and Charles Y. Conkle. Fecundity of red salmon at Brooks and Karluk Lakes, Alaska. Fishery Bulletin, 61(1):??, 1960– 1962. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/fecundity-red-salmon-
brooks-and-karluk-lakes-alaska;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/hartman.pdf.

Berry:1960:FMW

[1853] Frederick H. Berry and Louis E. Voegle. Filefishes (Monacanthidae) of the western north Atlantic. Fishery Bulletin, 61(1):

??, 1960–1962.  CODEN FSYBAY.  ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
filefishes-monacanthidae-western-north-atlantic;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
berry.pdf.

Piavis:1960:ESS

[1854] George W. Piavis. Embryological stages in the sea lamprey and effects of temperature on development. Fishery Bulletin, 61(1):??, 1960–1962. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/embryological-stages-sea-
lamprey-and-effects-temperature-developent;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/piavis.pdf.

Sindermann:1960:BPP

[1855] Carl J. Sindermann and Donald F. Mairs. Blood properties of prespawning and postspawning anadromous alewives (Alosa pseudoharengus). Fishery Bulletin, 61(1):??, 1960–1962. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/blood-properties-prespawning-and-
postspawning-anadromous-alewives-alosa-pseudoharengus;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
sindermann.pdf.

Marvin:1960:ECO

[1856] Kenneth T. Marvin, Larence M. Lansford, and Ray S. Wheeler. Effects of copper ore on the ecology of a lagoon. Fishery Bulletin, 61 (1):153–160, 1960–1962. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
effects-copper-ore-ecology-lagoon;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/marvin.pdf.

Judy:1960:VAD

[1857] Mayo H. Judy. Validity of age determination from scales of marked American shad. Fishery Bulletin, 61(1):??, 1960–1962. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/validity-age-determination-scales-marked-
american-shad;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/judy.pdf.

Grice:1960:CCE

[1858] George D. Grice. Calanoid copepods from equatorial waters of the Pacific Ocean. Fishery Bulletin, 61(1):??, 1960–1962. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/calanoid-copepods-equatorial-waters-pacific-
ocean;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/grice.pdf.

Farris:1960:ADE

[1859] David A. Farris. Abundance and distribution of eggs and larvae and survival of larvae of jack mackerel (Trachurus symmetricus). Fishery Bulletin, 61(1):??, 1960–1962. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/abundance-and-distribution-eggs-and-larvae-and-
survival-larvae-jack-mackerel-trachurus;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/farris.pdf.

Shippen:1960:DAS

[1860] Herbert H. Shippen. Distribution and abundance of skipjack in the Hawaii fishery, 1952–53. Fishery Bulletin, 61(1):??, 1960– 1962. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/distribution-
and-abundance-skipjack-hawaii-fishery-1952-53;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/shippen.
pdf.

Kerns:1960:AAK

[1861] Orra E. Kerns, Jr. Abundance and age of Kvichak River red salmon smolts. Fishery Bulletin, 61(1):??, 1960–1962. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/abundance-and-age-kvichak-river-
red-salmon-smolts;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/kerns.pdf.

Dobkin:1960:EDS

[1862] Sheldon Dobkin. Early developmental stages of pink shrimp, Penaeus duorarum, from Florida waters. Fishery Bulletin, 61(1):??, 1960–1962. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/early-developmental-
stages-pink-shrimp-penaeus-duorarum-florida-waters;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
dobkin.pdf.

Sindermann:1960:SSA

[1863] Carl J. Sindermann. Serological studies of Atlantic redfish. Fishery Bulletin, 61(1):??, 1960–1962. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/serological-studies-atlantic-redfish;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
sindermann1.pdf.

Becker:1960:ERS

[1864] Clarence Dale Becker. Estimating red salmon escapements by sample counts from observation towers. Fishery Bulletin, 61(1):??, 1960– 1962. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/estimating-
red-salmon-escapements-sample-counts-observation-towers;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
becker.pdf.

Seckel:1960:AOC

[1865] Gunter R. Seckel. Atlas of the oceanographic climate of the Hawaiian Islands. Fishery Bulletin, 61(1):??, 1960–1962. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/atlas-oceanographic-climate-hawaiian-
islands;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/seckel.pdf.

Major:1960:IEM

[1866] Richard L. Major and Donovan R. Craddock. Influence of early maturing females on reproductive potential of Columbia River blueback salmon (Oncorhynchus nerka). Fishery Bulletin, 61(1):??, 1960–1962. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/influence-early-maturing-
females-reproductive-potential-columbia-river-blueback-salmon;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/major.pdf.

Jensen:1960:DAY

[1867] Albert C. Jensen and John P. Wise. Determining age of young haddock from their scales. Fishery Bulletin, 61(1):??, 1960–1962. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/determining-age-young-haddock-their-
scales;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/jensen.pdf.

Moore:1960:DDC

[1868] Donald Moore. Development, distribution, and comparison of rudder fishes Kyphosus sectatrix (Linnaeus) and K. incisor (Cuvier) in the western north Atlantic. Fishery Bulletin, 61(1):??, 1960–1962. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/development-distribution-
and-comparison-rudder-fishes-kyphosus-sectatrix-linnaeus-and-
k;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/moore.pdf.

Shaw:1960:RCO

[1869] William N. Shaw. Raft culture of oysters in Massachusetts. Fishery Bulletin, 61(1):??, 1960–1962. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
raft-culture-oysters-massachusetts;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/shaw.pdf.

Anonymous:1962:I

[1870] Anonymous. Index. Fishery Bulletin, 62(1):??, 1962–1963. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-95;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
index.pdf.

Hida:1962:RTT

[1871] Thomas S. Hida, Joseph R. Harada, and Joseph E. King. Rearing tilapia for tuna bait. Fishery Bulletin, 62(1):1–20, 1962–1963. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/rearing-tilapia-tuna-
bait;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/hida.pdf.

Uchida:1962:TCT

[1872] Richard N. Uchida and Joseph E. King. Tank culture of tilapia. Fishery Bulletin, 62(1):??, 1962–1963. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/tank-culture-tilapia;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/uchida.pdf.

Weber:1962:GAM

[1873] Douglas D. Weber and Takashi Miyahara. Growth of the adult male king crab Paralithodes camtschatica (Tilesius). Fishery Bulletin, 62(1):??, 1962–1963. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/growth-adult-
male-king-crab-paralithodes-camtschatica-tilesius;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
weber.pdf.

Cooley:1962:SPA

[1874] Nelson R. Cooley. Studies on Parorchis acanthus (Trematoda: Digenea) as a biological control for the southern oyster drill, Thais haemastoma. Fishery Bulletin, 62(1):??, 1962–1963. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/studies-parorchis-acanthus-trematoda-
digenea-biological-control-southern-oyster-drill-thais;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
cooley.pdf.

Reintjes:1962:DEY

[1875] John W. Reintjes. Development of eggs and yolk-sac larvae of yellowfin menhaden. Fishery Bulletin, 62(1):??, 1962–1963. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/development-eggs-and-yolk-sac-larvae-
yellowfin-menhaden;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/reintjes.pdf.

Caldwell:1962:DDS

[1876] David K. Caldwell. Development and distribution of the short bigeye Pseudopriacanthus altus (Gill) in the western north Atlantic. Fishery Bulletin, 62(1):??, 1962–1963. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/development-and-distribution-short-bigeye-
pseudopriacanthus-altus-gill-western-north;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/caldwell.
pdf.

Otsu:1962:SMS

[1877] Tamio Otsu and Richard J. Hansen. Sexual maturity and spawning of the albacore in the central south Pacific. Fishery Bulletin, 62(1):??, 1962– 1963. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/sexual-maturity-and-
spawning-albacore-central-south-pacific;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/otsu.pdf.

Dragovich:1962:HCT

[1878] Alexander Dragovich and Billie Z. May. Hydrological characteristics of Tampa Bay tributaries. Fishery Bulletin, 62(1):??, 1962– 1963. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/hydrological-
characteristics-tampa-bay-tributaries;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/dragovich.pdf.

Baptist:1962:ARC

[1879] John P. Baptist and Thomas J. Price. Accumulation and retention of cesium137 by marine fishes. Fishery Bulletin, 62(1):??, 1962– 1963. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/accumulation-
and-retention-cesium137-marine-fishes;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/baptist.pdf.

Fields:1962:PTS

[1880] Hugh M. Fields. Pompanos (Trachinotus spp.) of south Atlantic Coast of the United States. Fishery Bulletin, 62(1):??, 1962–1963. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/pompanos-trachinotus-
spp-south-atlantic-coast-united-states;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fields.pdf.

Pugh:1962:ECE

[1881] John R. Pugh. Effect of certain electrical parameters and water resistivities on mortality of fingerling silver salmon. Fishery Bulletin, 62(1):??, 1962–1963. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/effect-certain-electrical-parameters-and-water-
resistivities-mortality-fingerling-silver;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/pugh.pdf.

Rounsefell:1962:RAN

[1882] George A. Rounsefell. Relationships among North American salmonidae. Fishery Bulletin, 62(1):??, 1962–1963. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/relationships-among-north-american-salmonidae;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/rounsefell.pdf.

King:1962:MTF

[1883] Joseph E. King and Robert T. B. Iverson. Midwater trawling for forage organisms in the central Pacific. Fishery Bulletin, 62(1):??, 1962–1963. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/midwater-trawling-forage-organisms-central-pacific;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/king.pdf.

Squire:1962:DTO

[1884] James L. Squire, Jr. Distribution of tunas in oceanic waters off the northwestern Atlantic. Fishery Bulletin, 62(1):??, 1962–1963. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/distribution-tunas-
oceanic-waters-northwestern-atlantic;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/squire.pdf.

Kutkuhn:1962:GMC

[1885] Joseph H. Kutkuhn. Gulf of Mexico commercial shrimp populations — trends and characteristics 1956–59. Fishery Bulletin, 62(1):??, 1962– 1963. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/gulf-mexico-
commercial-shrimp-populations-trends-and-characteristics-1956-
59;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/kutkuhn.pdf.

Caldwell:1962:DDL

[1886] Melba C. Caldwell. Development and distribution of larval and juvenile fishes of the family Mullidae of the western north Atlantic. Fishery Bulletin, 62(1):??, 1962–1963. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
development-and-distribution-larval-and-juvenile-fishes-family-
mullidae-western-north;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/caldwell1.pdf.

Iversen:1962:FAT

[1887] Robert T. B. Iversen. Food of the albacore tuna, Thunnus germo (Lacépède), in the central and northeastern Pacific. Fishery Bulletin, 62(1):??, 1962–1963. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/food-
albacore-tuna-thunnus-germo-lac%C3%A9p%C3%A8de-central-and-
northeastern-pacific;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/iversen.pdf.

Wise:1962:BBC

[1888] John P. Wise. Bibliography on the biology of the cod, Gadus morhua and related species. Fishery Bulletin, 62(1):??, 1962–1963. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/bibliography-biology-cod-gadus-
morhua-and-related-species;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/wise.pdf.

Ridgway:1963:DTS

[1889] George J. Ridgway. Distinguishing tuna species by immunochemical methods. Fishery Bulletin, 63(1):??, 1963. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/distinguishing-tuna-species-immunochemical-
methods;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/ridgway.pdf.

Scheffer:1964:DNS

[1890] Victor B. Scheffer and Bertram S. Kraus. Dentition of the northern fur seal. Fishery Bulletin, 63(2):??, 1964. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/dentition-northern-fur-seal;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
scheffer.pdf.

Hanks:1964:BCS

[1891] Robert W. Hanks. A benthic community in the Sheepscot River Estuary, Maine. Fishery Bulletin, 63(2):??, 1964. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/benthic-community-sheepscot-river-estuary-
maine;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/hanks.pdf.

Wyrtki:1964:UCR

[1892] Klaus Wyrtki. Upwelling in the Costa Rica Dome. Fishery Bulletin, 63(2):??, 1964. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/upwelling-
costa-rica-dome;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/wyrtki.pdf.

Rounsefell:1964:PSF

[1893] George A. Rounsefell. Preconstruction study of the fisheries of the estuarine areas by the Mississippi River–Gulf Outlet project.

Fishery Bulletin, 63(2):??, 1964. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/preconstruction-study-fisheries-estuarine-
areas-mississippi-river-gulf-outlet-project;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/rounsefell_
2.pdf.

Royce:1964:MSY

[1894] William F. Royce. A morphometric study of yellowfin Thunnus albacares (Bonnaterre). Fishery Bulletin, 63(2):??, 1964. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/morphometric-study-
yellowfin-thunnus-albacares-bonnaterre;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/royce%20(
1).pdf.

Cleaver:1964:OHS

[1895] Fred Cleaver. Origins of high seas sockeye salmon. Fishery Bulletin, 63 (2):??, 1964. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/origins-high-
seas-sockeye-salmon;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/cleaver.pdf.

Higham:1964:SMS

[1896] Joseph R. Higham and William R. Nicholson. Sexual maturation and spawning of Atlantic menhaden. Fishery Bulletin,

63(2):??, 1964.  CODEN FSYBAY.  ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
sexual-maturation-and-spawning-atlantic-menhaden;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
higham.pdf.

MacGregor:1964:RBS

[1897] John S. MacGregor. The relation between spawning-stock size and yearclass size for the Pacific sardine Sardinops caerulea (Girard). Fishery Bulletin, 63(2):??, 1964. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
relation-between-spawning-stock-size-and-year-class-size-pacific-
sardine-sardinops-caerulea;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/macgregor_0.pdf.

Thomas:1964:EEC

[1898] William H. Thomas. An experimental evaluation of the C14 method for measuring phytoplankton production using cultures of Dunaliella primolecta Butcher. Fishery Bulletin, 63(2):??, 1964. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/experimental-evaluation-c14-method-measuring-
phytoplankton-production-using-cultures;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/thomas.pdf.

Fredin:1963:OMM

[1899] R. A. Fredin. Ocean mortality and maturity schedules of Karluk River sockeye salmon and some comparisons of marine growth and mortality rates. Fishery Bulletin, 63(3):??, 1963–1964. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/ocean-mortality-and-
maturity-schedules-karluk-river-sockeye-salmon-and-some-comparisons;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/fredin.pdf.

Anonymous:1963:I

[1900] Anonymous. Index. Fishery Bulletin, 63(3):??, 1963–1964. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/index-100;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/index%
20(1).pdf.

McNeil:1963:MMM

[1901] William J. McNeil. A method of measuring mortality of pink salmon eggs and larvae. Fishery Bulletin, 63(3):??, 1963–1964. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/method-measuring-mortality-pink-
salmon-eggs-and-larvae;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/ncneil.pdf.

Quast:1963:MVH

[1902] J. C. Quast. Meristic variation in the hexagrammid fishes. Fishery Bulletin, 63(3):??, 1963–1964. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/meristic-variation-hexagrammid-fishes;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
quast.pdf.

Dryer:1963:MGR

[1903] William R. Dryer. Movements, growth, and rate of recapture of whitefish tagged in the Apostle Islands area of Lake Superior. Fishery Bulletin, 63 (3):??, 1963–1964. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/movements-
growth-and-rate-recapture-whitefish-tagged-apostle-islands-
area-lake-superior;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/dryer1.pdf.

Mraz:1963:AGS

[1904] Donald Mraz. Ages, growth, sex ratio, and maturity of the whitefish in central Green Bay and adjacent waters of Lake Michigan. Fishery Bulletin, 63(3):??, 1963–1964. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
ages-growth-sex-ratio-and-maturity-whitefish-central-green-
bay-and-adjacent-waters-lake;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/mraz.pdf.

Stickney:1963:FGJ

[1905] Alden P. Stickney. Feeding and growth of juvenile soft-shell clams, Mya arenaria. Fishery Bulletin, 63(3):??, 1963–1964. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/feeding-and-growth-
juvenile-soft-shell-clams-mya-arenaria;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/stickney.
pdf.

Dryer:1963:LHL

[1906] William R. Dryer and Joseph Beil. Life history of lake herring in Lake Superior. Fishery Bulletin, 63(3):??, 1963–1964. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/life-history-lake-herring-lake-superior;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/dryer%20(1).pdf.

Davis:1963:CET

[1907] Harry C. Davis and Anthony Calabrese. Combined effects of temperature and salinity on development of eggs and growth of larvae of M. mercenaria, and C. virginica. Fishery Bulletin, 63 (3):??, 1963–1964. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
combined-effects-temperature-and-salinity-development-eggs-

and-growth-larvae-m-mercenaria;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/davis_2.pdf.

Reintjes:1963:ABB

[1908] John W. Reintjes. Annotated bibliography on biology of menhadens and menhaden-like fishes of the world. Fishery Bulletin, 63(3):??, 1963–1964. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/annotated-
bibliography-biology-menhadens-and-menhaden-fishes-world;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/reintjes_1.pdf.

Chapman:1963:CSP

[1909] D. G. Chapman. A critical study of Pribilof fur seal population estimates. Fishery Bulletin, 63(3):??, 1963–1964. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/critical-study-pribilof-fur-seal-
population-estimates;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/chapman.pdf.

Anonymous:1964:P

[1910] Anonymous. Preface. Fishery Bulletin, 64(1):??, 1964. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/preface-8;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/preface.
pdf.

Anonymous:1964:AM

[1911] Anonymous. The adductor muscle. Fishery Bulletin, 64(1):??, 1964. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/adductor-muscle;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
ch8.pdf.

Anonymous:1964:LDM

[1912] Anonymous. Larval development and metamorphosis. Fishery Bulletin, 64(1):??, 1964. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/larval-
development-and-metamorphosis;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/ch16.pdf.

Anonymous:1964:T

[1913] Anonymous. Taxonomy. Fishery Bulletin, 64(1):??, 1964. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/taxonomy;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
ch1.pdf.

Anonymous:1964:TWG

[1914] Anonymous. Transport of water by the gills and respiration. Fishery Bulletin, 64(1):??, 1964. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/transport-water-gills-and-respiration;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
ch9.pdf.

Anonymous:1964:CC

[1915] Anonymous. Chemical composition. Fishery Bulletin, 64(1):??, 1964. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/chemical-composition;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/ch17.pdf.

Anonymous:1964:MSS

[1916] Anonymous. Morphology and structure of shell. Fishery Bulletin, 64(1): ??, 1964. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/morphology-
and-structure-shell;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/ch2.pdf.

Anonymous:1964:ODF

[1917] Anonymous. The organs of digestion and food of the oyster. Fishery Bulletin, 64(1):??, 1964. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/organs-digestion-and-food-oyster;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
ch10.pdf.

Anonymous:1964:EFA

[1918] Anonymous. Environmental factors affecting oyster populations. Fishery Bulletin, 64(1):??, 1964. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/environmental-factors-affecting-oyster-populations;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/ch18.pdf.

Anonymous:1964:L

[1919] Anonymous. The ligament. Fishery Bulletin, 64(1):??, 1964. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/ligament;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
ch3.pdf.

Anonymous:1964:CSB

[1920] Anonymous. The circulatory system and blood. Fishery Bulletin, 64(1): ??, 1964. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/circulatory-
system-and-blood;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/ch11.pdf.

Anonymous:1964:I

[1921] Anonymous. Index. Fishery Bulletin, 64(1):??, 1964. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/index-101;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/index%
20(2).pdf.

Anonymous:1964:GA

[1922] Anonymous. General anatomy. Fishery Bulletin, 64(1):??, 1964. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/general-anatomy;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/ch4.
pdf.

Anonymous:1964:ES

[1923] Anonymous. The excretory system. Fishery Bulletin, 64(1):??, 1964. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/excretory-system;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/ch12.pdf.

Anonymous:1964:M

[1924] Anonymous. The mantle. Fishery Bulletin, 64(1):??, 1964. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/mantle;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/ch5.pdf.

Anonymous:1964:NS

[1925] Anonymous. The nervous system. Fishery Bulletin, 64(1):??, 1964. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/nervous-system;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
ch13.pdf.

Anonymous:1964:LP

[1926] Anonymous. The labial palps. Fishery Bulletin, 64(1):??, 1964. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/labial-palps;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/ch6.pdf.

Anonymous:1964:OR

[1927] Anonymous. Organs of reproduction. Fishery Bulletin, 64(1):??, 1964. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/organs-reproduction;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/ch14.pdf.

Anonymous:1964:G

[1928] Anonymous. The gills. Fishery Bulletin, 64(1):??, 1964. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/gills;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/ch7.pdf.

Anonymous:1964:ESF

[1929] Anonymous. Eggs, sperm, fertilization, and cleavage. Fishery Bulletin, 64(1):??, 1964. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/eggs-sperm-fertilization-and-cleavage;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
ch15.pdf.

Williams:1965:MDC

[1930] Austin B. Williams. Marine decapod crustaceans of the Carolinas. Fishery Bulletin, 65(1):??, 1965. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/marine-decapod-crustaceans-carolinas;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/williams.pdf.

French:1965:SLD

[1931] Robert R. French and Roy L. Wahle. Study of loss and delay of salmon passing Rock Island Dam, Columbia River, 1954–56. Fishery Bulletin, 65 (2):??, 1965. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/study-loss-
and-delay-salmon-passing-rock-island-dam-columbia-river-1954-
56;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/french.pdf.

Yoshida:1965:STS

[1932] Howard O. Yoshida. Skipjack tuna spawning in the Marquesas Islands and Tuamotu Archipelago. Fishery Bulletin, 65(2):??, 1965. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/skipjack-tuna-
spawning-marquesas-islands-and-tuamotu-archipelago;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
yoshida.pdf.

Sykes:1965:OTB

[1933] James E. Sykes and John H. Finucane. Occurrence in Tampa Bay, Florida, of immature species dominant in Gulf of Mexico commercial fisheries. Fishery Bulletin, 65(2):??, 1965. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/occurrence-tampa-bay-floridia-immature-species-
dominant-gulf-mexico-commercial-fisheries;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/sykes.pdf.

Wolfert:1965:FYY

[1934] David R. Wolfert. Food of young-of-the-year walleyes in Lake Erie. Fishery Bulletin, 65(2):??, 1965. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
food-young-year-walleyes-lake-erie;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/wolfert.pdf.

Peterson:1965:GNM

[1935] Alvin E. Peterson. Gill net mesh selection curves for Pacific salmon on the high seas. Fishery Bulletin, 65(2):??, 1965. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/gill-net-mesh-selection-curves-
pacific-salmon-high-seas;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/peterson.pdf.

McNeil:1965:ESB

[1936] William J. McNeil. Effect of the spawning bed environment on reproduction of pink and chum salmon. Fishery Bulletin, 65(2):??, 1965. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/effect-spawning-bed-environment-
reproduction-pink-and-chum-salmon;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/mcneil.pdf.

Bodola:1965:LHG

[1937] Anthony Bodola. Life history of the gizzard shad, Dorosoma cepedianum (Le Sueur), in western Lake Erie. Fishery Bulletin, 65(2):??, 1965. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/life-history-gizzard-
shad-dorosoma-cepedianum-le-sueur-western-lake-erie;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
bodola.pdf.

Duke:1965:BRG

[1938] Thomas W. Duke, John P. Baptist, and Donald E. Hoss. Bioaccumulation of radioactive gold used as a sediment tracer in the estuarine environment. Fishery Bulletin, 65(2):427–436, 1965. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/bioaccumulation-radioactive-gold-
used-sediment-tracer-estuarine-environment;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/duke.pdf.

Cook:1965:GKP

[1939] Harry L. Cook. A generic key to the protozoean mysis and post larval stages of the littoral Penaeidae of the northwestern Gulf of Mexico. Fishery Bulletin, 65(2):??, 1965. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/generic-key-protozoean-mysis-and-
post-larval-stages-littoral-penaeidae-northwestern-gulf;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
cook.pdf.

Merrill:1965:AMS

[1940] Arthur S. Merrill, Julius A. Posgay, and Fred E. Nichy. Annual marks on shell and ligament of sea scallop Placopecten magellanicus. Fishery Bulletin, 65(2):299–311, 1965. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/annual-marks-shell-and-ligament-

sea-scallop-placopecten-magellanicus;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/merrill.pdf.

Costello:1965:MGD

[1941] T. J. Costello and Donald M. Allen. Migrations and geographic distribution of pink shrimp, Penaeus duorarum, of the Tortugas and Sanibel Grounds; Florida. Fishery Bulletin, 65(2):??, 1965. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/migrations-
and-geographic-distribution-pink-shrimp-penaeus-duorarum-tortugas-
and-sanibel;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/costello.pdf.

Kutkuhn:1965:DPS

[1942] Joseph H. Kutkuhn. Dynamics of a penaeid shrimp population and management implications. Fishery Bulletin, 65(2):??, 1965. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/dynamics-
penaeid-shrimp-population-and-management-implications;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
kutkuhn_0.pdf.

Straty:1965:TMA

[1943] Richard R. Straty. Time of migration and age group structure of sockeye salmon (Oncorhynchus nerka) spawning populations in the Naknek River system, Alaska. Fishery Bulletin, 65(2):??, 1965. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/time-migration-and-
age-group-structure-sockeye-salmon-oncorhynchus-nerka-spawning;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/straty.pdf.

Jensen:1965:LHS

[1944] Albert C. Jensen. Life history of the spiny dogfish. Fishery Bulletin, 65(3):??, 1965. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/life-history-spiny-dogfish;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/jensen_0.
pdf.

Heard:1965:PWP

[1945] William R. Heard and Wilbur L. Hartman. Pygmy whitefish Prosopium coulteri in the Naknek River system of southwest Alaska. Fishery Bulletin, 65(3):??, 1965. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/pygmy-whitefish-prosopium-coulteri-naknek-
river-system-southwest-alaska;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/heard.pdf.

Springer:1965:RWA

[1946] Stewart Springer. A review of western Atlantic cat sharks, Scyliorhinidae, with descriptions of a new genus and five new species. Fishery Bulletin, 65(3):??, 1965. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/review-
western-atlantic-cat-sharks-scyliorhinidae-descriptions-new-
genus-and-five-new;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/springer_0.pdf.

Berry:1965:SPF

[1947] Frederick H. Berry and Herbert C. Perkins. Survey of pelagic fishes of the California current area. Fishery Bulletin, 65(3):??, 1965. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/survey-
pelagic-fishes-california-current-area;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/berry%20(
1).pdf.

Marvin:1965:CTM

[1948] Kenneth T. Marvin and Raphael R. Proctor, Jr. Comparison of two methods of N-ethylcarbazole carbohydrate analysis. Fishery Bulletin, 65 (3):??, 1965. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/comparison-
two-methods-n-ethylcarbazole-carbohydrate-analysis;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
marvin_0.pdf.

Carver:1965:RMD

[1949] Thomas C. Carver, Jr. Rapid method for determining water content in oyster tissue. Fishery Bulletin, 65(3):??, 1965. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/rapid-method-determining-water-
content-oyster-tissue;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/carver.pdf.

Silliman:1967:ACM

[1950] Ralph P. Silliman. Analog computer models of fish population. Fishery Bulletin, 66(1):??, 1967. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
analog-computer-models-fish-population;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/silliman_
1.pdf.

Utter:1967:SDS

[1951] Fred M. Utter and George J. Ridgway. A serologically detected serum factor associated with maturity in English sole Parophrys vetulus, and Pacific halibut, Hippoglossus stenolepis. Fishery Bulletin, 66(1):??, 1967. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/serologically-
detected-serum-factor-associated-maturity-english-sole-parophrys-
vetulus-and;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/utter.pdf.

Gauley:1967:EWV

[1952] Joseph R. Gauley. Effect of water velocity on passage of salmonids in a transportation channel. Fishery Bulletin, 66(1):??, 1967. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/effect-water-velocity-passage-salmonids-
transportation-channel;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/gauley_0.pdf.

Gibbs:1967:CAS

[1953] Robert H. Gibbs, Jr. and Bruce B. Collette. Comparative anatomy and systematics of the tunas, genus Thunnus. Fishery Bulletin, 66(1):??, 1967. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/comparative-
anatomy-and-systematics-tunas-genus-thunnus;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/gibbs.
pdf.

Major:1967:IRR

[1954] Richard L. Major and James L. Mighell. Influence of Rocky Reach Dam and the temperature of the Okanogan River on the upstream migration of sockeye salmon. Fishery Bulletin, 66(1): ??, 1967. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/influence-
rocky-reach-dam-and-temperature-okanogan-river-upstream-migration-

sockeye-salmon;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/major_0.pdf.

Baxter:1967:SOS

[1955] Kenneth N. Baxter and William C. Renfro. Seasonal occurrence and size distribution of postlarval brown and white shrimp near Galveston, Texas, with notes on species identification. Fishery Bulletin, 66(1):??, 1967. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/seasonal-occurrence-and-size-distribution-
postlarval-brown-and-white-shrimp-near-galveston;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
baxter.pdf.

Nakamura:1967:DLT

[1956] Eugene L. Nakamura and Walter Matsumoto. Distribution of larval tunas in marquesan waters. Fishery Bulletin, 66(1):??, 1967. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/distribution-
larval-tunas-marquesan-waters;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/nakamura.pdf.

Coffin:1967:CEM

[1957] Gareth W. Coffin and Alden P. Stickney. Codium enters Maine waters. Fishery Bulletin, 66(1):??, 1967. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/codium-enters-maine-waters;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/coffin.pdf.

Hunter:1967:AFF

[1958] John R. Hunter and Charles T. Mitchell. Association of fishes with flotsam in the offshore waters of Central America. Fishery Bulletin, 66(1):??, 1967. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
association-fishes-flotsam-offshore-waters-central-america;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/hunter.pdf.

Marvin:1967:LER

[1959] Kenneth T. Marvin and Raphael R. Proctor, Jr. Laboratory evaluation of red-tide control agents. Fishery Bulletin, 66(1):??, 1967. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/laboratory-evaluation-red-tide-

control-agents;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/marvin_1.pdf.

Skud:1967:RMO

[1960] Bernard E. Skud. Responses of marine organisms during the solar eclipse of July 1963. Fishery Bulletin, 66(2):??, 1967. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/responses-marine-organisms-during-solar-eclipse-
july-1963;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/skud_0.pdf.

Silliman:1967:IFL

[1961] Ralph P. Silliman. Interaction of food level and exploitation in experimental fish populations. Fishery Bulletin, 66(2):??, 1967. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/interaction-food-level-and-exploitation-
experimental-fish-populations;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/silliman%20(1).pdf.

Koo:1967:OSS

[1962] Ted S. Y. Koo. Objective studies of scales of Columbia River Chinook salmon, Oncorhynchus tshawytscha (Walbaum). Fishery Bulletin, 66 (2):??, 1967. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/objective-
studies-scales-columbia-river-chinook-salmon-oncorhynchus-tshawytscha-
walbaum;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/koo.pdf.

Merrill:1967:SDM

[1963] Arthur S. Merrill. Shell deformity of mollusks attributable to the hydroid, Hydractinia echinata. Fishery Bulletin, 66(2):??, 1967. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/shell-deformity-mollusks-
attributable-hydroid-hydractinia-echinata;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/merrill_
0.pdf.

Uchida:1967:CEF

[1964] Richard N. Uchida. Catch and estimates of fishing effort and apparent abundance in the fishery for skipjack tuna (Katsuwonus pelamis) in Hawaiian waters, 1952–62. Fishery Bulletin, 66(2):??, 1967. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/catch-and-estimates-
fishing-effort-and-apparent-abundance-fishery-skipjack-tuna-
katsuwonus;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/uchida_0.pdf.

Merrill:1967:ODH

[1965] Arthur S. Merrill. Offshore distribution of Hydractinia echinata. Fishery Bulletin, 66(2):??, 1967. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
offshore-distribution-hydractinia-echinata;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/merrill1.
pdf.

Hutton:1967:CBA

[1966] Kenneth E. Hutton. Characteristics of the blood of adult pink salmon at three stages of maturity. Fishery Bulletin, 66(2):??, 1967. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/characteristics-blood-adult-pink-
salmon-three-stages-maturity;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/hutton.pdf.

Ho:1967:CCG

[1967] Ju shey Ho. Cyclopoid copepods of the genus Tucca (Tuccidae), parasitic on diodontid and tetraodontid fishes. Fishery Bulletin, 66(2):??, 1967. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/cyclopoid-copepods-genus-
tucca-tuccidae-parasitic-diodontid-and-tetraodontid-fishes;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/ho.pdf.

Utter:1967:CRP

[1968] Fred M. Utter and George J. Ridgway. Cross-reactive properties of antisera prepared in rabbits by stimulations with teleost vitellins. Fishery Bulletin, 66(2):??, 1967. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/cross-
reactive-properties-antisera-prepared-rabbits-stimulations-
teleost-vitellins;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/utter%20(1).pdf.

Matsumoto:1967:MDL

[1969] Walter M. Matsumoto. Morphology and distribution of larval wahoo Acanthocybium solandri (Cuvier) in the central Pacific Ocean. Fishery Bulletin, 66(2):??, 1967. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
morphology-and-distribution-larval-wahoo-acanthocybium-solandri-
cuvier-central-pacific;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/matsumoto_2.pdf.

Kelly:1967:OMT

[1970] John A. Kelly, Jr. and Alexander Dragovich. Occurrence of macrozooplankton in Tampa Bay, Florida, and the adjacent Gulf of Mexico. Fishery Bulletin, 66(2):??, 1967. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
occurrence-macrozooplankton-tampa-bay-florida-and-adjacent-
gulf-mexico;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/kelly_0.pdf.

Temple:1967:SDR

[1971] Robert F. Temple and Clarence C. Fischer. Seasonal distribution and relative abundance of planktonic-stage shrimp (Penaeus spp.) in the northwestern Gulf of Mexico, 1961. Fishery Bulletin, 66(2): ??, 1967. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/seasonal-
distribution-and-relative-abundance-planktonic-stage-shrimp-
penaeus-spp;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/temple.pdf.

Ingraham:1967:GCD

[1972] W. James Ingraham, Jr. The geostrophic circulation and distribution of water properties off the coasts of Vancouver Island and Washington, spring and fall 1963. Fishery Bulletin, 66(2):??, 1967. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/geostrophic-
circulation-and-distribution-water-properties-coasts-vancouver-
island-and;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/ingraham.pdf.

Sindermann:1967:PDC

[1973] Carl J. Sindermann and Aaron Rosenfield. Principal diseases of commercially important marine bivalve Mollusca and Crustacea. Fishery Bulletin, 66(2):??, 1967. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/principal-diseases-commercially-important-
marine-bivalve-mollusca-and-crustacea;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/sindermann_1.
pdf.

Rivas:1967:SBB

[1974] Luis R. Rivas and Stanley M. Warlen. Systematics and biology of the bonefish, Albula nemoptera (Fowler). Fishery Bulletin, 66(2):??, 1967. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
systematics-and-biology-bonefish-albula-nemoptera-fowler;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/rivas.pdf.

Mazeika:1967:SFG

[1975] P. A. Mazeika. Some features of the Gulf Stream off Chesapeake Bay in the spring of 1963. Fishery Bulletin, 66(2):??, 1967. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/some-features-gulf-stream-chesapeake-
bay-spring-1963;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/mazeika.pdf.

Royce:1967:MOM

[1976] William F. Royce, Lynwood S. Smith, and Allan C. Hartt. Models of oceanic migrations of Pacific salmon and comments on guidance mechanisms. Fishery Bulletin, 66(3):441–462, 1967. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/models-oceanic-migrations-pacific-salmon-and-
comments-guidance-mechanisms;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/royce_0.pdf.

Hayashi:1967:PAC

[1977] Sigeiti Hayashi. Preliminary analysis of the catch curve of the Pacific sardine, Sardinops caerulea Girard. Fishery Bulletin, 66(3):??, 1967. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/preliminary-analysis-
catch-curve-pacific-sardine-sardinops-caerulea-girard;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
hayashi.pdf.

Dragovich:1967:HBC

[1978] Alexander Dragovich, John A. Kelly, Jr., and H. Grant Goodell. Hydrological and biological characteristics of Florida’s west coast tributaries. Fishery Bulletin, 66(3):??, 1967. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/hydrological-and-biological-characteristics-floridas-
west-coast-tributaries;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/dragovich_0.pdf.

Long:1967:DMV

[1979] Clifford W. Long. Diel movement and vertical distribution of juvenile anadromous fish in turbine intakes. Fishery Bulletin, 66(3):??, 1967. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/diel-movement-
and-vertical-distribution-juvenile-anadromous-fish-turbine-
intakes;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/long.pdf.

Vaux:1967:IFI

[1980] Walter G. Vaux. Intragravel flow and interchange of water in a streambed. Fishery Bulletin, 66(3):??, 1967. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/intragravel-flow-and-interchange-
water-streambed;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/vaux.pdf.

Costello:1967:MRP

[1981] T. J. Costello and Donald M. Allen. Mortality rates in populations of pink shrimp, Penaeus duorarum, on the Sanibel and Tortugas grounds, Florida. Fishery Bulletin, 66(3):491–502, 1967. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/mortality-rates-populations-pink-
shrimp-penaeus-duorarum-sanibel-and-tortugas-grounds;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
costello_0.pdf.

Owen:1967:OCN

[1982] Robert W. Owen, Jr. Oceanographic conditions in the northeast Pacific Ocean and their relation to the albacore fishery. Fishery Bulletin, 66(3):??, 1967. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
oceanographic-conditions-northeast-pacific-ocean-and-their-
relation-albacore-fishery;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/owen.pdf.

Park:1967:CCC

[1983] Tai Soo Park. Calanoid copepods from the central north Pacific Ocean. Fishery Bulletin, 66(3):??, 1967. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/calanoid-copepods-central-north-
pacific-ocean;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/park.pdf.

Sakuda:1967:RMT

[1984] Henry M. Sakuda. A rapid method of tagging fish. Fishery Bulletin, 66(3):??, 1967. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/rapid-method-tagging-fish;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/sakuda.pdf.

McNeil:1967:MDP

[1985] William J. McNeil. Migration and distribution of pink salmon spawners in sashin oreek in 1965, and survival of their progeny. Fishery Bulletin, 66 (3):??, 1967. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/migration-and-
distribution-pink-salmon-spawners-sashin-oreek-1965-and-survival-
their;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/mcneil_0.pdf.

Trefethen:1968:PAC

[1986] Parker S. Trefethen and Doyle F. Sutherland. Passage of adult Chinook salmon through brownlee reservoir, 1960–62. Fishery Bulletin, 67(1): ??, 1968. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/passage-adult-
chinook-salmon-through-brownlee-reservoir-1960-62;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
trefethen.pdf.

Otsu:1968:DAA

[1987] Tamio Otsu and Ray F. Sumida. Distribution, apparent abundance, and size composition of albacore (Thunnus alalunga) taken in the longline fishery based in American Samoa, 1954–65. Fishery Bulletin, 67(1): ??, 1968. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/distribution-
apparent-abundance-and-size-composition-albacore-thunnus-alalunga-
taken;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/otsu_1.pdf.

Blackburn:1968:MET

[1988] Maurice Blackburn. Micronekton of the eastern tropical Pacific Ocean: family composition, distribution, abundance, and relations to tuna. Fishery Bulletin, 67(1):??, 1968. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/micronekton-eastern-tropical-pacific-
ocean-family-composition-distribution-abundance-and;   https:

//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
blackburn.pdf.

Stickney:1968:SAG

[1989] Alden P. Stickney. Supersaturation of atmospheric gases in the coastal waters of the Gulf of Maine. Fishery Bulletin, 67(1):??, 1968. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/supersaturation-atmospheric-gases-
coastal-waters-gulf-maine;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/stickney_0.pdf.

Fontaine:1968:RBT

[1990] Clark T. Fontaine and Richard A. Neal. Relation between tail length and total length for three commercially important species of penaeid shrimp. Fishery Bulletin, 67(1):??, 1968. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/relation-between-tail-length-and-
total-length-three-commercially-important-species-penaeid;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/fontaine.pdf.

Gray:1968:SSN

[1991] I. E. Gray, Maureen E. Downey, and M. J. Cerame-Vivas. Seastars of North Carolina. Fishery Bulletin, 67(1):127–163, 1968. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/sea-stars-north-carolina;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/gray.pdf.

Wells:1968:SDD

[1992] LaRue Wells. Seasonal depth distribution of fish in southeastern Lake Michigan. Fishery Bulletin, 67(1):??, 1968. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/seasonal-depth-distribution-fish-
southeastern-lake-michigan;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/wells_1.pdf.

Munro:1968:ADL

[1993] J. L. Munro, A. C. Jones, and D. Dimitriou. Abundance and distribution of the larvae of the pink shrimp (Penaeus duorarum) on the Tortugas Shelf of Florida, August 1962–0ctober 1964. Fishery Bulletin, 67(1):165– 181, 1968. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/abundance-and-

distribution-larvae-pink-shrimp-penaeus-duorarum-tortugas-shelf-
florida;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/munro.pdf.

Tagatz:1968:BBC

[1994] Marlin E. Tagatz. Biology of the blue crab, Callinectes sapidus Rathbun, in the St. Johns River, Florida. Fishery Bulletin, 67(1):??, 1968. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/biology-blue-
crab-callinectes-sapidus-rathbun-st-johns-river-florida;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
tagatz.pdf.

Taylor:1968:SEH

[1995] John L. Taylor and Carol H. Saloman. Some effects of hydraulic dredging and coastal development in Boca Ciega Bay, Florida. Fishery Bulletin, 67 (2):??, 1968. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/some-effects-
hydraulic-dredging-and-coastal-development-boca-ciega-bay-florida;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/taylor_2.pdf.

Davis:1968:EPE

[1996] Harry C. Davis and Herbert Hidu. Effects of pesticides on embryonic development of clams and oysters and on survival and growth of the larvae. Fishery Bulletin, 67(2):??, 1968. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/effects-pesticides-embryonic-development-clams-and-
oysters-and-survival-and-growth-larvae;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/davis_3.
pdf.

Mosher:1968:PAS

[1997] Kenneth H. Mosher. Photographic atlas of sockeye salmon scales. Fishery Bulletin, 67(2):??, 1968. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/photographic-atlas-sockeye-salmon-scales;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
mosher.pdf.

Burgner:1968:BSE

[1998] Robert L. Burgner, Charles J. DiCostanzo, Robert J. Ellis, Georgie Y. Harry, Jr., Wilbur L. Hartman, Orra E. Kerns, Jr., Ole A. Mathisen, and William R. Royce. Biological studies and estimates of optimum escapements of sockeye salmon in the major river systems in southwestern Alaska. Fishery Bulletin, 67(2):??, 1968. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/biological-studies-and-estimates-
optimum-escapements-sockeye-salmon-major-river-systems;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
burgner.pdf.

Tagatz:1968:GJB

[1999] Marlin E. Tagatz. Growth of juvenile blue crabs, Callinectes sapidus Rathbun, in the St. Johns River, Florida. Fishery Bulletin, 67(2):??, 1968. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/growth-juvenile-blue-
crabs-callinectes-sapidus-rathbun-st-johns-river-florida;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/tagatz%20(1).pdf.

Lindner:1968:DBS

[2000] Milton J. Lindner and James S. Bailey. Distribution of brown shrimp (Penaeus aztecus aztecus Ives) as related to turbid water photographed from space. Fishery Bulletin, 67(2):??, 1968. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/distribution-brown-shrimp-penaeus-aztecus-
aztecus-ives-related-turbid-water-photographed;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/lindner.
pdf.

Ebel:1968:PCL

[2001] Wesley J. Ebel and Charles H. Koski. Physical and chemical limnology of brownlee reservoir, 1962–64. Fishery Bulletin,

67(2):??, 1968.  CODEN FSYBAY.  ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
physical-and-chemical-limnology-brownlee-reservoir-1962-64;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/ebel.pdf.

Welch:1968:CAG

[2002] Walter R. Welch. Changes in abundance of the green crab, Carcinus maenas (L.), in relation to recent temperature changes. Fishery Bulletin, 67(2):??, 1968. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/changes-abundance-green-crab-carcinus-maenas-l-

relation-recent-temperature-changes;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/welch.pdf.

Ropes:1968:FHG

[2003] John W. Ropes. The feeding habits of the green crab, Carcinus maenas (L.). Fishery Bulletin, 67(2):??, 1968. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/feeding-habits-green-crab-carcinus-maenas-
l;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/ropes_0.pdf.

Major:1968:EMS

[2004] Richard L. Major and James L. Mighell. Egg-to-migrant survival of spring Chinook salmon (Oncorhynchus tshawytscha) in the Yakima River, Washington. Fishery Bulletin, 67(2):??, 1968. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/egg-migrant-survival-
spring-chinook-salmon-oncorhynchus-tshawytscha-yakima-river-
washington;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/major%20(1).pdf.

Yoshida:1968:ELH

[2005] Howard O. Yoshida. Early life history and spawning of the albacore, Thunnus alalunga, in Hawaiian waters. Fishery Bulletin, 67(2):??, 1968. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/early-life-history-
and-spawning-albacore-thunnus-alalunga-hawaiian-waters;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
yoshida_0.pdf.

Worlund:1968:CCR

[2006] Donald D. Worlund, Roy J. Wahle, and Paul D. Zimmer. Contribution of Columbia River hatcheries to harvest of fall Chinook salmon (Oncorhynchus tshawytscha). Fishery Bulletin, 67(2):361–391, 1968. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/contribution-
columbia-river-hatcheries-harvest-fall-chinook-salmon-oncorhynchus-
tshawytscha;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/worlund.pdf.

Farfante:1968:WAS

[2007] Isabel Pérez Farfante. Western Atlantic shrimps of the genus Penaeus. Fishery Bulletin, 67(3):??, 1968. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/western-atlantic-shrimps-genus-penaeus;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/perez.
pdf.

Cohen:1970:ARA

[2008] Daniel M. Cohen and Samuel P. Atsaides. Additions to a revision of Argentinine fishes. Fishery Bulletin, 68(1):??, 1970. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/additions-revision-argentinine-
fishes;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1970/681/cohen.pdf.

Ebel:1970:SNC

[2009] Wesley J. Ebel. Supersaturation of nitrogen in the Columbia River and its effect on salmon and steelhead trout. Fishery Bulletin, 68(1):??, 1970. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/supersaturation-nitrogen-
columbia-river-and-its-effect-salmon-and-steelhead-trout;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1970/
681/ebel.pdf.

Hansen:1970:FGM

[2010] David J. Hansen. Food, growth, migration, reproduction, and abundance of pinfish, Lagodon rhomboides, and Atlantic croaker, Micropogon undulatus, near Pensacola, Florida, 1963–65. Fishery Bulletin, 68(1):??, 1970. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/food-growth-
migration-reproduction-and-abundance-pinfish-lagodon-rhomboides-
and-atlantic;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1970/681/hansen.pdf.

MacKenzie:1970:FRS

[2011] Clyde L. MacKenzie, Jr. Feeding rates of starfish, Asterias forbesi (Desor), at controlled water temperatures and during different seasons of the year. Fishery Bulletin, 68(1):??, 1970. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/feeding-rates-starfish-asterias-
forbesi-desor-controlled-water-temperatures-and-during;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1970/681/
mackenzie.pdf.

Shomura:1970:VMZ

[2012] Richard S. Shomura and Eugene L. Nakamura. Variations in marine zooplankton from a single locality in Hawaiian waters. Fishery Bulletin, 68 (1):??, 1970. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/variations-
marine-zooplankton-single-locality-hawaiian-waters;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1970/681/
shomura.pdf.

Stearns:1970:BMG

[2013] Franklin Stearns. Bathymetric maps and geomorphology of the Middle Atlantic Continental Shelf. Fishery Bulletin, 68(1):??, 1970. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/bathymetric-maps-and-
geomorphology-middle-atlantic-continental-shelf;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1970/681/
stearns.pdf.

Allen:1970:ARB

[2014] Donald M. Allen and T. J. Costello. Additional references on the biology of shrimp, family Penaeidae. Fishery Bulletin, 68(1):??, 1970. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/additional-
references-biology-shrimp-family-penaeidae;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1970/681/allen.pdf.

Stickney:1970:FIA

[2015] Alden P. Stickney. Factors influencing the attraction of Atlantic herring, Clupea harengus Harengus, to artificial lights. Fishery Bulletin, 68(1):??, 1970. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/factors-influencing-
attraction-atlantic-herring-clupea-harengus-harengus-artificial-
lights;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1970/681/stickney.pdf.

Blackburn:1970:CRU

[2016] Maurice Blackburn. Conditions related to upwelling which determine distribution of tropical tunas off Western Baja California. Fishery Bulletin, 68(1):??, 1970. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
conditions-related-upwelling-which-determine-distribution-tropical-
tunas-western-baja;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1970/681/blackburn.pdf.

Anderson:1970:RGS

[2017] William D. Anderson, Jr. Revision of the genus Symphysanodon (Pisces: Lutjanidae) with descriptions of four new species. Fishery Bulletin, 68(2):??, 1970. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/revision-genus-symphysanodon-pisces-lutjanidae-
descriptions-four-new-species;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1970/682/anderson.pdf.

Sykes:1970:CDM

[2018] James E. Sykes and John R. Hall. Comparative distribution of mollusks in dredged and undredged portions of an estuary, with a systematic list of species. Fishery Bulletin, 68(2):??, 1970. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/comparative-distribution-
mollusks-dredged-and-undredged-portions-estuary-systematic-
list;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1970/682/sykes.pdf.

Durkin:1970:DMJ

[2019] Joseph T. Durkin, Donn L. Park, and Robert F. Raleigh. Distribution and movement of juvenile salmon in Brownlee Reservoir, 1962–65. Fishery Bulletin, 68(2):??, 1970. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/distribution-and-movement-juvenile-salmon-brownlee-
reservoir-1962-65;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1970/682/durkin.pdf.

Taylor:1970:MBE

[2020] John L. Taylor, John R. Hall, and Carl H. Saloman. Mollusks and benthic environments in Hillsborough Bay, Florida. Fishery Bulletin, 68(2):??, 1970. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
mollusks-and-benthic-environments-hillsborough-bay-florida;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1970/682/taylor.pdf.

Gehringer:1970:YAS

[2021] Jack W. Gehringer. Young of the Atlantic sailfish, Istiophorus platypterus. Fishery Bulletin, 68(2):??, 1970. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/young-atlantic-sailfish-istiophorus-platypterus;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1970/682/gehringer.pdf.

Gutherz:1970:CSL

[2022] Elmer J. Gutherz. Characteristics of some larval bothid flatfish, and development and distribution of larval spotfin flounder, Cyclopsetta fimbriata (Bothidae). Fishery Bulletin, 68(2):??, 1970. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/characteristics-some-
larval-bothid-flatfish-and-development-and-distribution-larval-
spotfin;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1970/682/gutherz.pdf.

Krcma:1970:MJS

[2023] Richard F. Krcma and Robert F. Raleigh. Migration of juvenile salmon and trout into Brownlee Reservoir, 1962–65. Fishery Bulletin, 68(2):??, 1970. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/migration-juvenile-
salmon-and-trout-brownlee-reservoir-1962-65;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1970/682/krcma.
pdf.

MacKenzie:1970:COD

[2024] Clyde L. MacKenzie, Jr. Control of oyster drills, Eupleura caudata and Urosalpinx cinerea, with the chemical polystream. Fishery Bulletin, 68 (2):??, 1970. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/control-
oyster-drills-eupleura-caudata-and-urosalpinx-cinerea-chemical-
polystream;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1970/682/mackenzie.pdf.

Pugh:1970:EWV

[2025] John R. Pugh, Gerald E. Monan, and Jim R. Smith. Effect of water velocity on the fish-guiding efficiency of an electrical guiding system. Fishery Bulletin, 68(2):??, 1970. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
effect-water-velocity-fish-guiding-efficiency-electrical-guiding-
system;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1970/682/pugh.pdf.

Sims:1970:EJS

[2026] Carl W. Sims. Emigration of juvenile salmon and trout from Brownlee Reservoir, 1963–65. Fishery Bulletin, 68(2):??, 1970. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/emigration-juvenile-
salmon-and-trout-brownlee-reservoir-1963-65;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1970/682/sims.
pdf.

Cressey:1970:CNS

[2027] Roger F. Cressey and Bruce B. Collette. Copepods and needlefishes: a study in host–parasite relationships. Fishery Bulletin, 68(3):??, 1970. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/copepods-and-
needlefishes-study-host-parasite-relationships;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1970/683/cressey.
pdf.

Dragovich:1970:FSY

[2028] Alexander Dragovich. The food of skipjack and yellowfin tunas in the Atlantic Ocean. Fishery Bulletin, 68(3):??, 1970. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/food-skipjack-and-yellowfin-tunas-atlantic-
ocean;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1970/683/dragovich.pdf.

June:1970:FYA

[2029] Fred C. June and Frank T. Carlson. Food of young Atlantic menhaden, Brevoortia tyrannus, in relation to metamorphosis. Fishery Bulletin, 68 (3):??, 1970. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/food-young-
atlantic-menhaden-brevoortia-tyrannus-relation-metamorphosis;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1970/683/june.pdf.

Martin:1970:SCP

[2030] Dean F. Martin and Ashim B. Chatterjee. Some chemical and physical properties of two toxins from the red-tide organism, Gymnodinium breve. Fishery Bulletin, 68(3):??, 1970. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/some-chemical-and-physical-properties-two-
toxins-red-tide-organism-gymnodinium-breve;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1970/683/martin.pdf.

Merrell:1970:EMC

[2031] Theodore R. Merrell, Jr., Melvin D. Collins, and Joseph W. Greenough. An estimate of mortality of Chinook salmon in the Columbia River near Bonneville Dam during the summer run of 1955. Fishery Bulletin, 68 (3):??, 1970. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/estimate-
mortality-chinook-salmon-columbia-river-near-bonneville-dam-
during-summer-run-1955;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1970/683/merrell.pdf.

Cook:1971:EDS

[2032] Harry L. Cook and M. Alice Murphy. Early developmental stages of the brown shrimp, Penaeus aztecus Ives, reared in the laboratory. Fishery Bulletin, 69(1):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/early-developmental-stages-brown-
shrimp-penaeus-aztecus-ives-reared-laboratory;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1971/691/cook.
pdf.

Seckel:1971:HFS

[2033] Günther R. Seckel and Marian Y. Y. Yong. Harmonic functions for sea-surface temperatures and salinities, Koko Head, Oahu, 1956–69, and sea-surface temperatures, Christmas Island, 1954–69. Fishery Bulletin, 69(1):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
harmonic-functions-sea-surface-temperatures-and-salinities-
koko-head-oahu-1956-69-and-sea;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1971/691/seckel.pdf.

Crear:1971:LRD

[2034] David Crear and Irwin Haydock. Laboratory rearing of the desert pupfish, Cyprinodon macularius. Fishery Bulletin, 69(1):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/laboratory-rearing-
desert-pupfish-cyprinodon-macularius;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1971/691/crear.pdf.

Smiles:1971:SSG

[2035] Michael C. Smiles, Jr. and William G. Pearcy. Size structure and growth rate of Euphausia pacifica off the Oregon Coast. Fishery Bulletin, 69(1):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/size-
structure-and-growth-rate-euphausia-pacifica-oregon-coast;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1971/
691/smiles.pdf.

Dubrow:1971:CNC

[2036] David L. Dubrow and Bruce R. Stillings. Chemical and nutritional characteristics of fish protein concentrate processed from heated whole red hake, Urophycis chuss. Fishery Bulletin, 69(1):??, 1971. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/chemical-and-nutritional-
characteristics-fish-protein-concentrate-processed-heated-whole;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1971/691/dubrow.pdf.

Thomas:1971:EPP

[2037] William H. Thomas and Robert W. Owen, Jr. Estimating phytoplankton production from ammonium and chlorophyll concentrations in nutrientpoor water of the Eastern Tropical Pacific Ocean. Fishery Bulletin, 69(1): ??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/estimating-
phytoplankton-production-ammonium-and-chlorophyll-concentrations-
nutrient-poor;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1971/691/thomas.pdf.

Dubrow:1971:EIS

[2038] David L. Dubrow, Norman L. Brown, E. R. Pariser, Harry Miller, Jr., V. D. Sidwell, and Mary E. Ambrose. Effect of ice storage on the chemical and nutritive properties of solvent-extracted whole fish — red hake, Urophycis chuss. Fishery Bulletin, 69(1):??, 1971. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/effect-ice-storage-chemical-and-
nutritive-properties-solvent-extracted-whole-fish-red-hake;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1971/691/dubrow1.pdf.

Roedel:1971:MWM

[2039] Philip M. Roedel. In memoriam — Wilbert McLeod Chapman and Milner Baily Schaefer. Fishery Bulletin, 69(1):??, 1971. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/memoriam-wilbert-mcleod-chapman-
and-milner-baily-schaefer;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/inmemoriam.pdf.

Haydock:1971:GMH

[2040] Irwin Haydock. Gonad maturation and hormone-induced spawning of the Gulf croaker, Bairdiella icistia. Fishery Bulletin, 69(1):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/gonad-maturation-and-
hormone-induced-spawning-gulf-croaker-bairdiella-icistia;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1971/
691/haydock.pdf.

Emiliani:1971:EHR

[2041] Dennis A. Emiliani. Equipment for holding and releasing penaeid shrimp during marking experiments. Fishery Bulletin, 69(1):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/equipment-holding-and-releasing-
penaeid-shrimp-during-marking-experiments;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/emiliani.
pdf.

Jellinek:1971:MUF

[2042] Gisela Jellinek and Maurice E. Stansby. Masking undesirable flavors in fish oils. Fishery Bulletin, 69(1):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/masking-undesirable-flavors-fish-oils;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1971/691/
jellinek.pdf.

Topp:1971:ABT

[2043] Robert W. Topp and Frank H. Hoff. An adult bluefin tuna, Thunnus thynnus, from a Florida west coast urban waterway. Fishery Bulletin, 69 (1):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/adult-bluefin-
tuna-thunnus-thynnus-florida-west-coast-urban-waterway;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
topp.pdf.

Ahlstrom:1971:KAF

[2044] Elbert H. Ahlstrom. Kinds and abundance of fish larvae in the Eastern Tropical Pacific, based on collections made on EAS-TROPAC I. Fishery Bulletin, 69(1):??, 1971. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/kinds-and-abundance-fish-larvae-
eastern-tropical-pacific-based-collections-made-eastropac-i;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1971/691/ahlstrom.pdf.

Koury:1971:PAR

[2045] Barbara Koury, John Spinelli, and Dave Wieg. Protein autolysis rates at various pH’s and temperatures in hake, Merluccius productus, and Pacific herring, Clupea harengus pallasi, and their effect on yield in the preparation of fish protein concentrate. Fishery Bulletin, 69(1):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/protein-autolysis-rates-various-
phs-and-temperatures-hake-merluccius-productus-and-pacific;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1971/691/koury.pdf.

Clutter:1971:EEP

[2046] Robert I. Clutter and Gail H. Theilacker. Ecological efficiency of a pelagic mysid shrimp; estimates from growth, energy budget, and mortality studies. Fishery Bulletin, 69(1):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/ecological-efficiency-pelagic-mysid-shrimp-estimates-
growth-energy-budget-and-mortality;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1971/691/clutter.pdf.

Rothschild:1971:LPS

[2047] Brian J. Rothschild and James W. Balsiger. A linear-programming solution to salmon management. Fishery Bulletin, 69(1):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/linear-programming-
solution-salmon-management;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1971/691/rothschild.pdf.

Anderson:1971:CBR

[2048] William W. Anderson and Milton J. Lindner. Contributions to the biology of the royal red shrimp, Hymenopenaeus robustus Smith. Fishery Bulletin, 69(2):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
contributions-biology-royal-red-shrimp-hymenopenaeus-robustus-
smith;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1971/692/anderson.pdf.

Magnuson:1971:GRA

[2049] John J. Magnuson and Jean G. Heitz. Gill raker apparatus and food selectivity among mackerels, tunas, and dolphins. Fishery Bulletin, 69(2):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/gill-
raker-apparatus-and-food-selectivity-among-mackerels-tunas-
and-dolphins;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1971/692/magnuson.pdf.

Barnett:1971:SUC

[2050] Harold J. Barnett, Richard W. Nelson, Patrick J. Hunter, Steven Bauer, and Herman Groninger. Studies on the use of carbon dioxide dissolved in refrigerated brine for the preservation of whole fish. Fishery Bulletin, 69 (2):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/studies-use-
carbon-dioxide-dissolved-refrigerated-brine-preservation-whole-
fish;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1971/692/barnett.pdf.

May:1971:EDI

[2051] Robert C. May. Effects of delayed initial feeding on larvae of the grunion, Leuresthes tenuis (Ayres). Fishery Bulletin, 69(2):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/effects-delayed-initial-feeding-
larvae-grunion-leuresthes-tenuis-ayres;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1971/692/may.pdf.

Caldwell:1971:CLA

[2052] David K. Caldwell, Melba C. Caldwell, Warren F. Rathjen, and John R. Sullivan. Cetaceans from the Lesser Antillean Island of St. Vincent. Fishery Bulletin, 69(2):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/cetaceans-lesser-antillean-island-st-vincent;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1971/
692/caldwell.pdf.

Perkins:1971:ELD

[2053] Herbert C. Perkins. Egg loss during incubation from offshore northern lobsters (Decapoda: Homaridae). Fishery Bulletin, 69 (2):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/egg-
loss-during-incubation-offshore-northern-lobsters-decapoda-
homaridae;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1971/692/perkins.pdf.

Cox:1971:DRS

[2054] James L. Cox. DDT residues in seawater and particulate matter in the California Current System. Fishery Bulletin, 69(2):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/ddt-residues-
seawater-and-particulate-matter-california-current-system;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1971/
692/cox.pdf.

Pritchard:1971:RBE

[2055] Austin W. Pritchard, John R. Hunter, and Reuben Lasker. The relation between exercise and biochemical changes in red and white muscle and liver in the jack mackerel, Trachurus symmetricus. Fishery Bulletin, 69(2):379–386, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/relation-between-exercise-and-biochemical-
changes-red-and-white-muscle-and-liver-jack;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1971/692/pritchard.
pdf.

Friedl:1971:RSP

[2056] William A. Friedl. The relative sampling performance of 6- and 10-foot Isaacs–Kidd midwater trawls. Fishery Bulletin, 69(2):427– 432, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/relative-
sampling-performance-6-and-lo-foot-isaacs-kidd-midwater-trawls;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1971/692/friedl.pdf.

Quast:1971:SVS

[2057] Jay C. Quast. Sehastes variegatus, sp. n. from the Northeastern Pacific Ocean (Pisces, Scorpaenidaea). Fishery Bulletin, 69(2):??, 1971. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/sehastes-variegatus-sp-n-
northeastern-pacific-ocean-pisces-scorpaenidae;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1971/692/quast.
pdf.

Kifer:1971:EDF

[2058] Robert R. Kifer, Preston Smith, Jr., and Edgar P. Young. Effect of dietary fish oil on the fatty acid composition and palatability of pig tissues. Fishery Bulletin, 69(2):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/effect-dietary-fish-oil-fatty-acid-composition-and-
palatability-pig-tissues;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1971/692/kifer.pdf.

Roe:1971:CSD

[2059] Richard B. Roe, Robert Cummins, Jr., and Harvey R. Bullis, Jr. Calico scallop distribution, abundance, and yield off Eastern Florida, 1967–68. Fishery Bulletin, 69(2):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/calico-scallop-distribution-abundance-and-
yield-eastern-florida-1967-68;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1971/692/roe.pdf.

Kittredge:1971:SPA

[2060] James S. Kittredge, Michelle Terry, and Francis T. Takahashi. Sex pheromone activity of the molting hormone, crustecdysone, on male crabs (Pachygrapsus crassipes, Cancer antennarius, and C. anthonyi). Fishery Bulletin, 69(2):337–343, 1971. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/sex-pheromone-activity-molting-
hormone-crustecdysone-male-crabs-pachygrapsus-crassipes;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1971/692/
kittredge.pdf.

Roubal:1971:NFR

[2061] William T. Roubal. Nature of free radicals in freeze-dried fishery products and other lipid–protein systems. Fishery Bulletin, 69(2):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/nature-free-radicals-
freeze-dried-fishery-products-and-other-lipid-protein-systems;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1971/692/roubal.pdf.

Leong:1971:ISN

[2062] Roderick Leong. Induced spawning of the northern anchovy, Engraulis mordax Girard. Fishery Bulletin, 69(2):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/induced-spawning-northern-anchovy-engraulis-
mordax-girard;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1971/692/leong.pdf.

Bailey:1971:LTT

[2063] Jack E. Bailey and Dale R. Evans. The low-temperature threshold for pink salmon eggs in relation to a proposed hydroelectric installation. Fishery Bulletin, 69(3):587–593, 1971. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/low-temperature-threshold-pink-
salmon-eggs-relation-proposed-hydroelectric-installation;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1971/
693/bailey.pdf.

Lillelund:1971:LSP

[2064] Kurt Lillelund and Reuben Lasker. Laboratory studies of predation by marine copepods on fish larvae. Fishery Bulletin, 69(3):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/laboratory-studies-
predation-marine-copepods-fish-larvae;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1971/693/lillelund.pdf.

Yoshida:1971:ELH

[2065] Howard O. Yoshida. The early life history of skipjack tuna, Katsuwonus pelamis, in the Pacific Ocean. Fishery Bulletin, 69(3):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/early-life-history-
skipjack-tuna-katsuwonus-pelamis-pacific-ocean;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/yoshida%
20(1).pdf.

Compagno:1971:LNG

[2066] Leonard J. V. Compagno and Stewart Springer. Lago, a new genus of carcharhinid sharks, with a redescription of I. omanensis. Fishery Bulletin, 69(3):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/lago-new-genus-carcharhinid-sharks-redescription-i-
omanensis;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1971/693/compagno.pdf.

MacGregor:1971:ADS

[2067] John S. MacGregor. Additional data on the spawning of the hake. Fishery Bulletin, 69(3):581–585, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/additional-data-spawning-hake;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1971/
693/macgregor.pdf.

Cox:1971:UAL

[2068] James L. Cox. Uptake, assimilation, and loss of DDT residues by Euphausia pacifica, a euphausiid shrimp. Fishery Bulletin, 69(3):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/uptake-assimilation-
and-loss-ddt-residues-euphausia-pacifica-euphausiid-shrimp;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/cox_1.pdf.

Cummings:1971:GWE

[2069] William C. Cummings and Paul O. Thompson. Gray whales, Eschrichtius robustus, avoid the under water sounds of killer whales, Orcinus orca. Fishery Bulletin, 69(3):??, 1971. CODEN FSY-

BAY.  ISSN 0090-0656 (print), 1937-4518 (electronic).  URL

https://spo.nmfs.noaa.gov/content/gray-whales-eschrichtius-
robustus-avoid-under-water-sounds-killer-w-hales-orcinus-orca;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1971/693/cummings.pdf.

PerezFarfante:1971:KAP

[2070] Isabel Perez Farfante. A key to the American Pacific shrimps of the genus Trachypenaeus (Decapoda, Penaeidae), with the description of a new species. Fishery Bulletin, 69(3):??, 1971. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/key-american-pacific-
shrimps-genus-trachypenaeus-decapoda-penaeidae-description-
new-species;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1971/693/perez.pdf.

OConnell:1971:VNS

[2071] Charles P. O’Connell. Variability of near-surface zooplankton off southern California, as shown by towed-pump sampling. Fishery Bulletin, 69 (3):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/variability-
near-surface-zooplankton-southern-california-shown-towed-pump-
sampling;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/oconnell.pdf.

Fish:1971:KWO

[2072] James F. Fish and John S. Vania. Killer whale, Orcinus orca, sounds repel white whales, Delphinapterus leucas. Fishery Bulletin, 69(3):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/killer-whale-orcinus-
orca-sounds-repel-white-whales-delphinapterus-leucas;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1971/693/
fish.pdf.

Richards:1971:DTL

[2073] William J. Richards and David C. Simmons. Distribution of tuna larvae (Pisces, Scombridae) in the Northwestern Gulf of Guinea and off Sierra Leone. Fishery Bulletin, 69(3):??, 1971. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/distribution-tuna-larvae-pisces-
scombridae-northwestern-gulf-guinea-and-sierra-leone;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1971/693/
richards.pdf.

Fox:1971:RVP

[2074] William W. Fox, Jr. Random variability and parameter estimation for the generalized production model. Fishery Bulletin, 69(3):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/random-variability-
and-parameter-estimation-generalized-production-model;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1971/693/
fox.pdf.

Ritskes:1971:ARM

[2075] T. M. Ritskes. Artificial ripening of Maatjes-cured herring with the aid of proteolytic enzyme preparations. Fishery Bulletin, 69(3):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/artificial-ripening-
maatjes-cured-herring-aid-proteolytic-enzyme-preparations;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1971/
693/ritskes.pdf.

Hobson:1971:CSA

[2076] Edmund S. Hobson. Cleaning symbiosis among California inshore fishes. Fishery Bulletin, 69(3):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/cleaning-symbiosis-among-california-inshore-fishes;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1971/693/hobson.pdf.

Sindermann:1971:IDC

[2077] Carl J. Sindermann. Internal defenses of crustacea: a review. Fishery Bulletin, 69(3):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
internal-defenses-crustacea-review;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1971/693/sindermann.pdf.

Houde:1971:DAF

[2078] Edward D. Houde. Developmental abnormalities of the flatfish Achirus lineatus reared in the laboratory. Fishery Bulletin, 69(3):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/developmental-
abnormalities-flatfish-achirus-lineatus-reared-laboratory;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1971/
693/houde.pdf.

Novotny:1971:PJP

[2079] Anthony J. Novotny and Conrad V. W. Mahnken. Predation on juvenile Pacific salmon by a marine isopod Rocinela belliceps pugettensis (Crustacea, Isopoda). Fishery Bulletin, 69(3):??, 1971. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/predation-juvenile-
pacific-salmon-marine-isopod-rocinela-belliceps-pugettensis-
crustacea;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/novotny.pdf.

Larrance:1971:PPM

[2080] Jerry D. Larrance. Primary production in the Mid-Subarctic Pacific Region, 1966–68. Fishery Bulletin, 69(3):??, 1971. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/primary-production-mid-subarctic-
pacific-region-1966-68;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1971/693/larrance.pdf.

Lasker:1971:A

[2081] Reuben Lasker. Announcement. Fishery Bulletin, 69(3):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/annoucement;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
announcement.pdf.

Corliss:1971:CPP

[2082] Jane Corliss and Lee Trent. Comparison of phytoplankton production between natural and altered areas in West Bay, Texas. Fishery Bulletin, 69 (4):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/comparison-
phytoplankton-production-between-natural-and-altered-areas-
west-bay-texas;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1971/694/corliss.pdf.

Yoshida:1971:DAA

[2083] Howard O. Yoshida. Distribution, apparent abundance, and length composition of juvenile albacore, Thunnus alalunga, in the South Pacific Ocean. Fishery Bulletin, 69(4):??, 1971. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/distribution-apparent-abundance-
and-length-composition-juvenile-albacore-thunnus-alalunga;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1971/
694/yoshida.pdf.

Ebel:1971:TTJ

[2084] Wesley J. Ebel, Earl M. Dawley, and Bruce H. Monk. Thermal tolerance of juvenile Pacific salmon and steelhead trout in relation to supersaturation of nitrogen gas. Fishery Bulletin, 69(4):833– 843, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/thermal-
tolerance-juvenile-pacific-salmon-and-steelhead-trout-relation-
supersaturation;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1971/694/ebel.pdf.

Potthoff:1971:OTS

[2085] Thomas Potthoff. Observations on two species of dolphin (Coryphaena) from the tropical mid-Atlantic. Fishery Bulletin, 69(4):??, 1971. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/observations-two-species-
dolphin-coryphaena-tropical-mid-atlantic;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/potthoff.
pdf.

Jones:1971:IBS

[2086] Everet C. Jones. Isistius brasiliensis, a squaloid shark, the probable cause of crater wounds on fishes and cetaceans. Fishery Bulletin, 69(4):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
isistius-brasiliensis-squaloid-shark-probable-cause-crater-
wounds-fishes-and-cetaceans;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1971/694/jones.pdf.

Sage:1971:RBE

[2087] Martin Sage. Respiratory, behavioral, and endocrine responses of a teleost to a restricted environment. Fishery Bulletin, 69(4):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/respiratory-
behavioral-and-endocrine-responses-teleost-restricted-environment;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/sage.pdf.

Love:1971:MFT

[2088] Richard H. Love. Measurements of fish target strength: a review. Fishery Bulletin, 69(4):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/measurements-fish-target-strength-review;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1971/694/
love.pdf.

Bakkala:1971:OBP

[2089] Richard G. Bakkala. Occurrence of the barracudina, Paralepis atlantica Krøyer, in the central North Pacific Ocean. Fishery Bulletin, 69(4):881, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/occurrence-
barracudina-paralepis-atlantica-kr%C3%B8yer-central-north-pacific-
ocean;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/bakkala.pdf.

Malone:1971:RIN

[2090] Thomas C. Malone. The relative importance of nannoplankton and netplankton as primary producers in the California Current System. Fishery Bulletin, 69(4):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
relative-importance-nannoplankton-and-netplankton-primary-producers-
california-current;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1971/694/malone.pdf.

Mathisen:1971:ELP

[2091] Ole A. Mathisen. Escapement levels and productivity of the Nushagak sockeye salmon run from 1908 to 1966. Fishery Bulletin, 69(4):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
escapement-levels-and-productivity-nushagak-sockeye-salmon-
run-1908-1966;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1971/694/mathisen.pdf.

Beardsley:1971:CPD

[2092] Grant L. Beardsley. Contribution to the population dynamics of Atlantic albacore with comments on potential yields. Fishery Bulletin, 69(4):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/contribution-population-dynamics-atlantic-albacore-
comments-potential-yields;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1971/694/beardsley.pdf.

Nicholson:1971:CCE

[2093] William R. Nicholson. Changes in catch and effort in the Atlantic menhaden purse-seine fishery 1940–68. Fishery Bulletin, 69(4):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/changes-catch-
and-effort-atlantic-menhaden-purse-seine-fishery-1940-68;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1971/
694/nicholson.pdf.

Beers:1971:PPU

[2094] John R. Beers, Merritt R. Stevenson, Richard W. Eppley, and Elaine R. Brooks. Plankton populations and upwelling off the coast of Peru, June 1969. Fishery Bulletin, 69(4):??, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/plankton-populations-and-
upwelling-coast-peru-june-1969;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1971/694/beers.pdf.

Wilkens:1971:ADY

[2095] E. Peter H. Wilkens and Robert M. Lewis. Abundance and distribution of young Atlantic menhaden, Brevoortia tyrannus, in the White Oak River Estuary, North Carolina. Fishery Bulletin, 69(4):783–789, 1971. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/abundance-and-distribution-young-atlantic-menhaden-
brevoortia-tyrannus-white-oak-river;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1971/694/wilkens.pdf.

Barraclough:1972:FGC

[2096] W. E. Barraclough and D. Robinson. The fertilization of Great Central Lake. III. Effect on juvenile sockeye salmon. Fishery Bulletin, 70(1):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/fertilization-great-
central-lake-iii-effect-juvenile-sockeye-salmon;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1972/701/
barraclough.pdf.

Kepshire:1972:GPC

[2097] Bernard M. Kepshire, Jr. and William J. McNeil. Growth of premigratory Chinook salmon in seawater. Fishery Bulletin, 70(1):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/growth-premigratory-
chinook-salmon-seawater;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1972/701/kepshire.pdf.

Rosenthal:1972:PPR

[2098] Richard J. Rosenthal and James R. Chess. A predator–prey relationship between the leather star, Dermasterias imbricata, and the purple urchin, Strongylocentrotus purpuratus. Fishery Bulletin, 70(1):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/predator-prey-
relationship-between-leather-star-dermasterias-imbricata-and-
purple-urchin;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1972/701/rosenthal.pdf.

Cahn:1972:SFS

[2099] Phyllis H. Cahn. Sensory factors in the side-to-side spacing and positional orientation of the tuna, Euthynnus affinis, during schooling. Fishery Bulletin, 70(1):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
sensory-factors-side-side-spacing-and-positional-orientation-
tuna-euthynnus-affinis-during;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1972/701/cahn.pdf.

Lebrasseur:1972:FGC

[2100] R. J. Lebrasseur and O. D. Kennedy. The fertilization of Great Central Lake. II. Zooplankton standing stock. Fishery Bulletin, 70(1):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/fertilization-
great-central-lake-ii-zooplankton-standing-stock;   https:/

/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1972/701/
lebrasseur.pdf.

Sick:1972:PSS

[2101] Lowell V. Sick, James W. Andrews, and David B. White. Preliminary studies of selected environmental and nutritional requirements for the culture of penaeid shrimp. Fishery Bulletin, 70 (1):101–109, 1972. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
preliminary-studies-selected-environmental-and-nutritional-
requirements-culture-penaeid;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1972/701/sick.pdf.

Davy:1972:RLG

[2102] Brent Davy. A review of the lanternfish genus Taaningichthys (family Myctophidae) with the description of a new species. Fishery Bulletin, 70(1):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
review-lanternfish-genus-taaningichthys-family-myctophidae-
description-new-species;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1972/701/davy.pdf.

Lewis:1972:DYA

[2103] Robert M. Lewis, E. Peter H. Wilkens, and Herbert R. Gordy. A description of young Atlantic menhaden, Brevoortia tyrannus, in the White Oak River Estuary, North Carolina. Fishery Bulletin, 70(1):115–118, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/description-young-atlantic-menhaden-brevoortia-
tyrannus-white-oak-river-estuary-north;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1972/701/lewis.pdf.

Hoopes:1972:LGT

[2104] David T. Hoopes and John F. Karinen. Longevity and growth of tagged king crabs in the eastern Bering Sea. Fishery Bulletin, 70(1): ??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/longevity-and-
growth-tagged-king-crabs-eastern-bering-sea;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/hoopes.
pdf.

Drucker:1972:SLH

[2105] Benson Drucker. Some life history characteristics of Coho salmon of the Karluk River System, Kodiak Island, Alaska. Fishery Bulletin, 70(1):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
some-life-history-characteristics-coho-salmon-karluk-river-
system-kodiak-island-alaska;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1972/701/drucker.pdf.

Major:1972:EEW

[2106] Richard L. Major and Gerald J. Paulik. Effect of encroachment of Wanapum Dam Reservoir on fish passage over Rock Island Dam, Columbia River. Fishery Bulletin, 70(1):??, 1972. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/effect-encroachment-wanapum-dam-
reservoir-fish-passage-over-rock-island-dam-columbia-river;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1972/701/major.pdf.

Perrin:1972:EBH

[2107] William F. Perrin and John R. Hunter. Escape behavior of the Hawaiian spinner porpoise (Stenella cf. S. longirostris). Fishery Bulletin, 70(1):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/escape-behavior-hawaiian-spinner-porpoise-stenella-
cf-s-longirostris;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/perrin.pdf.

Evans:1972:MTS

[2108] W. E. Evans, J. D. Hall, A. B. Irvine, and J. S. Leatherwood. Methods for tagging small cetaceans. Fishery Bulletin, 70(1):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/methods-tagging-small-cetaceans;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1972/701/evans.pdf.

Matsumoto:1972:CLT

[2109] Walter M. Matsumoto, Elbert H. Ahlstrom, S. Jones, Witold L. Klawe, William J. Richards, and Shoji Ueyanagi. On the clarification of larval tuna identification particularly in the genus Thunnus. Fishery Bulletin, 70(1):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
clarification-larval-tuna-identification-particularly-genus-
thunnus;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1972/701/matsumoto.pdf.

Houde:1972:DEL

[2110] Edward D. Houde. Development and early life history of the northern sennet, Sphyraena borealis Dekay (Pisces: Sphyraenidae) reared in the laboratory. Fishery Bulletin, 70(1):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/development-and-early-life-history-northern-
sennet-sphyraena-borealis-dekay-pisces;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1972/701/houde.pdf.

Mosher:1972:SFS

[2111] Kenneth H. Mosher. Scale features of sockeye salmon from Asian and North American Coastal regions. Fishery Bulletin, 70(1): ??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/scale-
features-sockeye-salmon-asian-and-north-american-coastal-regions;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1972/701/mosher.pdf.

Judkins:1972:CFC

[2112] David C. Judkins and Abraham Fleminger. Comparison of foregut contents of Sergestes similis obtained from net collections and albacore stomachs. Fishery Bulletin, 70(1):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/comparison-foregut-contents-sergestes-similis-obtained-
net-collections-and-albacore;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1972/701/judkins.pdf.

Parsons:1972:FGC

[2113] T. R. Parsons, K. Stephens, and M. Takahashi. The fertilization of Great Central Lake. I. Effect of primary production. Fishery Bulletin, 70(1):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/fertilization-great-central-lake-i-effect-
primary-production;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1972/701/parsons.pdf.

Kelley:1972:MDC

[2114] Carolyn E. Kelley and Anthony W. Harmon. Method of determining carotenoid contents of Alaska pink shrimp and representative valves for several shrimp products. Fishery Bulletin, 70(1):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/method-determining-
carotenoid-contents-alaska-pink-shrimp-and-representative-valves-
several;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1972/701/kelley.pdf.

Perkins:1972:DRV

[2115] Herbert C. Perkins. Developmental rates at various temperatures of embryos of the northern lobster (Homarus americanus Milne-Edwards). Fishery Bulletin, 70(1):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/developmental-rates-various-temperatures-embryos-northern-
lobster-homarus-americanus-milne;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1972/701/perkins.pdf.

Banse:1972:RSS

[2116] Karl Banse. Redescription of some species of Chone Kröyer and Euchone Malmgren, and three new species (Sabellidae, Polychaeta). Fishery Bulletin, 70(2):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/redescription-some-species-chone-
kroyer-and-euchone-malmgren-and-three-new-species;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1972/702/
banse.pdf.

Johnson:1972:EIF

[2117] Allyn G. Johnson, Fred M. Utter, and Harold O. Hodgins. Electrophoretic investigation of the family Scorpaenidae. Fishery Bulletin, 70(2):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
electrophoretic-investigation-family-scorpaenidae;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1972/702/
johnson.pdf.

Vernberg:1972:SET

[2118] Winona B. Vernberg and John Vernberg. The synergistic effects of temperature, salinity, and mercury on survival and metabolism of the adult fiddler crab, Uca pugilator. Fishery Bulletin, 70(2):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/synergistic-effects-temperature-
salinity-and-mercury-survival-and-metabolism-adult-fiddler;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1972/702/vernberg.pdf.

Booth:1972:MOS

[2119] Douglas E. Booth. A model for optimal salmon management. Fishery Bulletin, 70(2):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/model-optimal-salmon-management;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1972/702/booth.
pdf.

Johnson:1972:LWR

[2120] Allyn G. Johnson and Howard F. Horton. Length–weight relationship, food habits, parasites, and sex and age determination of the ratfish, Hydrolagus colliei (Lay and Bennett). Fishery Bulletin, 70(2):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/length-weight-
relationship-food-habits-parasites-and-sex-and-age-determination-
ratfish;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1972/702/johnson1.pdf.

Alvarino:1972:SRR

[2121] A. Alvarino. A second record of a rare siphonophore Epibulia ritteriana Haeckel 1888. Fishery Bulletin, 70(2):??, 1972. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/second-record-rare-siphonophore-
epibulia-ritteriana-haeckel-1888;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/alvarino.pdf.

Dahlberg:1972:ESG

[2122] Michael D. Dahlberg. An ecological study of Georgia coastal fishes. Fishery Bulletin, 70(2):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/ecological-study-georgia-coastal-fishes;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1972/702/
dahlberg.pdf.

Kasahara:1972:JDW

[2123] Hiroshi Kasahara. Japanese distant-water fisheries: a review. Fishery Bulletin, 70(2):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/japanese-distant-water-fisheries-review;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1972/702/
kasahara.pdf.

Davies:1972:DOA

[2124] Isaac E. Davies and Robert P. Bradley. Deep observations of anchovy and blue sharks from deepstar 4000. Fishery Bulletin, 70 (2):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/deep-
observations-anchovy-and-blue-sharks-deepstar-4000;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
davies.pdf.

Graham:1972:RLH

[2125] Joseph J. Graham. Retention of larval herring within the Sheepscot Estuary of Maine. Fishery Bulletin, 70(2):??, 1972. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/retention-larval-herring-within-
sheepscot-estuary-maine;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1972/702/graham.pdf.

Leighton:1972:LOE

[2126] David L. Leighton. Laboratory observations on the early growth of the abalone, Haliotis sorenseni, and the effect of temperature on larval development and settling success. Fishery Bulletin, 70(2):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/laboratory-
observations-early-growth-abalone-haliotis-sorenseni-and-effect-
temperature;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1972/702/leighton.pdf.

Rothschild:1972:UCI

[2127] B. J. Rothschild and D. S. Robson. The use of concentration indices in fisheries. Fishery Bulletin, 70(2):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/use-concentration-indices-fisheries;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
rothschild.pdf.

Graham:1972:ADM

[2128] Joseph J. Graham, Stanley B. Chenoweth, and Clarence W. Davis. Abundance, distribution, movements, and lengths of larval herring along the western coast of the Gulf of Maine. Fishery Bulletin, 70 (2):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/abundance-
distribution-movements-and-lengths-larval-herring-along-western-
coast-gulf-maine;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1972/702/graham1.pdf.

Mather:1972:MDW

[2129] Frank J. Mather III, Albert C. Jones, and Grant L. Beardsley, Jr. Migration and distribution of white marlin and blue marlin in the Atlantic Ocean. Fishery Bulletin, 70(2):??, 1972. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/migration-and-distribution-white-
marlin-and-blue-marlin-atlantic-ocean;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1972/702/mather.pdf.

Carlson:1972:SIT

[2130] Frank T. Carlson and John W. Reintjes. Suitability of internal tags for Atlantic menhaden. Fishery Bulletin, 70(2):??, 1972. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/suitability-internal-
tags-atlantic-menhaden;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/carlson.pdf.

Henry:1972:ODG

[2131] Kenneth A. Henry. Ocean distribution. Growth, and effects of the troll fishery on yield of fall Chinook salmon from Columbia River hatcheries. Fishery Bulletin, 70(2):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/ocean-distribution-growth-and-effects-troll-fishery-
yield-fall-chinook-salmon-columbia;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1972/702/henry.pdf.

McCarthy:1972:NEA

[2132] James J. McCarthy and Terry E. Whitledge. Nitrogen excretion by anchovy (Engraulis mordax and E. ringens) and jack mackerel (Trachurus symmetricus). Fishery Bulletin, 70(2):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/nitrogen-excretion-anchovy-engraulis-
mordax-and-e-ringens-and-jack-mackerel-trachurus;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1972/702/
mccarthy.pdf.

Graham:1972:FEB

[2133] Joseph J. Graham. Filtration efficiencies of Boothbay depressor trawls. Fishery Bulletin, 70(2):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/filtration-efficiencies-boothbay-
depressor-trawls;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/graham2.pdf.

Hoag:1972:RBS

[2134] Stephen H. Hoag. The relationship between the summer food of juvenile sockeye salmon, Oncorhynchus nerka, and the standing stock of zooplankton in Iliamna Lake, Alaska. Fishery Bulletin, 70(2):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/relationship-between-
summer-food-juvenile-sockeye-salmon-oncorhynchus-nerka-and-
standing;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1972/702/hoag.pdf.

Saila:1972:EEU

[2135] Saul B. Saila and James D. Parrish. Exploitation effects upon interspecific relationships in marine ecosystems. Fishery Bulletin, 70(2):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/exploitation-
effects-upon-interspecific-relationships-marine-ecosystems;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1972/702/saila.pdf.

Hoopes:1972:SSS

[2136] David T. Hoopes. Selection of spawning sites by sockeye salmon in small streams. Fishery Bulletin, 70(2):??, 1972. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/selection-spawning-sites-sockeye-
salmon-small-streams;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1972/702/hoopes.pdf.

Venrick:1972:SSD

[2137] E. L. Venrick. Small-scale distributions of oceanic diatoms. Fishery Bulletin, 70(2):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/small-scale-distributions-oceanic-diatoms;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1972/702/venrick.pdf.

Colton:1972:TTD

[2138] John B. Colton, Jr. Temperature trends and the distribution of groundfish in Continental Shelf waters, Nova Scotia to Long Island. Fishery Bulletin, 70(3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
temperature-trends-and-distribution-groundfish-continental-
shelf-waters-nova-scotia-long;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1972/703/colton.pdf.

Lenarz:1972:MRLa

[2139] William H. Lenarz. Mesh retention of larvae of Sardinops caerulea and Engraulis mordax by plankton nets. Fishery Bulletin, 70 (3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/mesh-
retention-larvae-sardinops-caerulea-and-engraulis-mordax-plankton-
nets;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1972/703/lenarz.pdf.

OConnell:1972:LSP

[2140] Charles P. O’Connell and James R. Zweifel. A laboratory study of particulate and filter feeding of the Pacific mackerel, Scomber japonicus. Fishery Bulletin, 70(3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic)q. URL https://

spo.nmfs.noaa.gov/content/laboratory-study-particulate-and-
filter-feeding-pacific-mackerel-scomber-japonicus;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1972/703/
oconnell.pdf.

Squire:1972:AAS

[2141] James L. Squire, Jr. Apparent abundance of some pelagic marine fishes off the Southern and Central California Coast as surveyed by an airborne monitoring program. Fishery Bulletin, 70(3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/apparent-abundance-
some-pelagic-marine-fishes-southern-and-central-california-
coast;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1972/703/squire.pdf.

Royce:1972:UEF

[2142] William F. Royce. Undergraduate education of fishery scientists. Fishery Bulletin, 70(3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
undergraduate-education-fishery-scientists;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/royce%20(
1)_0.pdf.

Fitch:1972:FFM

[2143] John E. Fitch and Lloyd W. Barker. The fish family Moridae in the Eastern North Pacific with notes on morid otoliths, caudal skeletons, and the fossil record. Fishery Bulletin, 70(3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/fish-family-moridae-eastern-north-pacific-
notes-morid-otoliths-caudal-skeletons-and-fossil;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1972/703/
fitch.pdf.

Lewis:1972:GHPa

[2144] U. J. Lewis, R. N. P. Singh, B. K. Seavey, R. Lasker, and Grace E. Pickford. Growth hormone- and prolactin-like proteins of the blue shark (Prionace glauca). Fishery Bulletin, 70(3):933–939, 1972. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/growth-hormone-and-
prolactin-proteins-blue-shark-prionace-glauca;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1972/703/lewis.
pdf.

Perrin:1972:CPS

[2145] William F. Perrin. Color patterns of spinner porpoises (Stenella cf. S. longirostris) of the Eastern Pacific and Hawaii, with comments on delphinid pigmentation. Fishery Bulletin, 70(3):??, 1972. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/color-patterns-spinner-porpoises-
stenella-cf-s-longirostris-eastern-pacific-and-hawaii;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1972/703/
perrin.pdf.

Thomas:1972:NIS

[2146] William H. Thomas. Nutrient inversions in the Southeastern Tropical Pacific Ocean. Fishery Bulletin, 70(3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/nutrient-inversions-southeastern-tropical-
pacific-ocean;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1972/703/thomas.pdf.

Hunter:1972:SFBa

[2147] John R. Hunter. Swimming and feeding behavior of larval anchovy, Engraulis mordax. Fishery Bulletin, 70(3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/swimming-and-feeding-behavior-larval-
anchovy-engraulis-mordax-0;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/hunter%20(1).pdf.

Hobson:1972:AHR

[2148] Edmund S. Hobson. Activity of Hawaiian reef fishes during the evening and morning transitions between daylight and darkness. Fishery Bulletin, 70(3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/activity-
hawaiian-reef-fishes-during-evening-and-morning-transitions-
between-daylight-and;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1972/703/hobson.pdf.

McHugh:1972:MFN

[2149] J. L. McHugh. Marine fisheries of New York State. Fishery Bulletin, 70(3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/marine-fisheries-new-york-state;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1972/703/mchugh.
pdf.

Roedel:1972:RFS

[2150] Philip M. Roedel. Research and the Fisheries Service’s second 100 years. Fishery Bulletin, 70(3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/research-and-fisheries-services-second-100-years;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1972/
703/roedel1.pdf.

Williams:1972:CTP

[2151] F. Williams. Consideration of three proposed models of the migration of young skipjack tuna (Katsuwonus pelamis) into the Eastern Pacific Ocean. Fishery Bulletin, 70(3):??, 1972. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/consideration-three-
proposed-models-migration-young-skipjack-tuna-katsuwonus-pelamis;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1972/703/williams.pdf.

Lenarz:1972:MRLb

[2152] William H. Lenarz. Mesh retention of larvae of Sardinops caerulea and Engraulis mordax: by plankton nets. Fishery Bulletin, 70 (3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/mesh-
retention-larvae-sardinops-caerulea-and-engraulis-mordax-plankton-
nets-0;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/lenarz.pdf.

Hunter:1972:SFBb

[2153] John R. Hunter. Swimming and feeding behavior of larval anchovy Engraulis mordax. Fishery Bulletin, 70(3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/swimming-and-feeding-behavior-larval-
anchovy-engraulis-mordax;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1972/703/hunter.pdf.

Moser:1972:DLS

[2154] H. Geoffrey Moser and Elbert H. Ahlstrom. Development of the lanternfish, Scopelopsis multipunctatus Brauer 1906, with a discussion of its phylogenetic position in the family Myctophidae and its role in a proposed mechanism for the evolution of photophore patterns in lanternfishes. Fishery Bulletin, 70(3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/development-lanternfish-scopelopsis-multipunctatus-
brauer-1906-discussion-its-phylogenetic;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1972/703/moser.pdf.

Saur:1972:MSL

[2155] J. F. T. Saur. Monthly sea level differences between the Hawaiian Islands and the California Coast. Fishery Bulletin, 70(3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/monthly-sea-
level-differences-between-hawaiian-islands-and-california-coast;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1972/703/saur.pdf.

Wolfe:1972:CEE

[2156] D. A. Wolfe and T. R. Rice. Cycling of elements in estuaries. Fishery Bulletin, 70(3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/cycling-elements-estuaries;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1972/703/wolfe.pdf.

Lewis:1972:GHPb

[2157] U. J. Lewis, R. N. P. Singh, B. K. Seavey, R. Lasker, and Grace E. Pickford. Growth hormone and prolactin-like proteins of the blue shark (Prionace glauca). Fishery Bulletin, 70(3):??, 1972. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/growth-hormone-and-
prolactin-proteins-blue-shark-prionace-glauca-0;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
lewis.pdf.

Isaacs:1972:UMF

[2158] John D. Isaacs. Unstructured marine food webs and ‘pollutant analogues’. Fishery Bulletin, 70(3):??, 1972. CODEN FSY-

BAY.  ISSN 0090-0656 (print), 1937-4518 (electronic).  URL

https://spo.nmfs.noaa.gov/content/unstructured-marine-food-
webs-and-pollutant-analogues;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1972/703/isaacs.pdf.

Moser:1972:DGD

[2159] H. Geoffrey Moser. Development and geographic distribution of the rockfish, Sebastes macdonaldi (Eigenmann and Beeson, 1893), family Scorpaenidae, off Southern California and Baja California. Fishery Bulletin, 70(3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
development-and-geographic-distribution-rockfish-sebastes-macdonaldi-
eigenmann-and-beeson;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1972/703/moser1.pdf.

Scott:1972:BBT

[2160] J. Michael Scott and Glenn A. Flittner. Behavior of bluefin tuna schools in the Eastern North Pacific Ocean as inferred from fishermen’s logbooks, 1960–67. Fishery Bulletin, 70(3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/behavior-bluefin-tuna-schools-eastern-north-pacific-
ocean-inferred-fishermens-logbooks-1960;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1972/703/scott.pdf.

Anonymous:1972:EN

[2161] Anonymous. Editor’s note. Fishery Bulletin, 70(3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/editors-note;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/editorsnote.
pdf.

Kimura:1972:OSPa

[2162] Makoto Kimura and Gary T. Sakagawa. Observations on scale patterns and growth of the Pacific sardine reared in the laboratory. Fishery Bulletin, 70(3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/observations-scale-patterns-and-growth-pacific-
sardine-reared-laboratory-0;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/kimura_1.pdf.

June:1972:VSL

[2163] Fred C. June. Variations in size and length composition of Atlantic menhaden groupings. Fishery Bulletin, 70(3):??, 1972. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/variations-size-and-
length-composition-atlantic-menhaden-groupings;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1972/703/june.
pdf.

Murphy:1972:SAL

[2164] Garth I. Murphy and Robert I. Clutter. Sampling anchovy larvae with a plankton purse seine. Fishery Bulletin, 70(3):??, 1972. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/sampling-anchovy-larvae-plankton-
purse-seine;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1972/703/murphy.pdf.

Seckel:1972:HCS

[2165] Gunter R. Seckel. Hawaiian-caught skipjack tuna and their physical environment. Fishery Bulletin, 70(3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/hawaiian-caught-skipjack-
tuna-and-their-physical-environment;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1972/703/seckel.pdf.

Roedel:1972:OES

[2166] Philip M. Roedel. Oscar Elton Sette. Fishery Bulletin, 70(3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/oscar-elton-
sette;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/roedel.pdf.

Abramson:1972:AYM

[2167] Norman J. Abramson and Patrick K. Tomlinson. An application of yield models to a California ocean shrimp population. Fishery Bulletin, 70(3):1021–1041, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/application-yield-models-california-ocean-
shrimp-population;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1972/703/abramson.pdf.

Kimura:1972:OSPb

[2168] Makoto Kimura and Gary T. Sakagawa. Observations on scale patterns and growth of the Pacific sardine reared in the laboratory. Fishery Bulletin, 70(3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/observations-scale-patterns-and-growth-pacific-
sardine-reared-laboratory;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1972/703/kimura.pdf.

Murphy:1972:PEA

[2169] Garth I. Murphy and Richard S. Shomura. Pre-exploitation abundance of tunas in the Equatorial Central Pacific. Fishery Bulletin, 70(3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/pre-
exploitation-abundance-tunas-equatorial-central-pacific;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1972/703/
murphy1.pdf.

Silliman:1972:ECR

[2170] Ralph P. Silliman. Effect of crowding on relation between exploitation and yield in Tilapia macrocephala. Fishery Bulletin, 70(3):693–698, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/effect-crowding-relation-
between-exploitation-and-yield-tilapia-macrocephala;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1972/703/
silliman.pdf.

Powell:1972:OES

[2171] Patricia Powell. Oscar Elton Sette: Fishery biologist. Fishery Bulletin, 70(3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/oscar-
elton-sette-fishery-biologist;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/powell.pdf.

Barkley:1972:STN

[2172] Richard A. Barkley. Selectivity of towed-net samplers. Fishery Bulletin, 70(3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/selectivity-towed-net-samplers;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1972/703/
barkley.pdf.

Leggett:1972:WTM

[2173] William C. Leggett and Richard R. Whitney. Water temperature and the migrations of American shad. Fishery Bulletin, 70 (3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/water-
temperature-and-migrations-american-shad;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1972/703/leggett.pdf.

Namias:1972:SSS

[2174] Jerome Namias. Space scales of sea-surface temperature patterns and their causes. Fishery Bulletin, 70(3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/space-scales-sea-surface-temperature-
patterns-and-their-causes;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1972/703/namias.pdf.

Smith:1972:ISB

[2175] Paul E. Smith. The increase in spawning biomass of northern anchovy, Engraulis mordax. Fishery Bulletin, 70(3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/increase-spawning-biomass-northern-anchovy-
engraulis-mordax;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1972/703/smith.pdf.

Rothschild:1972:EDF

[2176] Brian J. Rothschild. An exposition on the definition of fishing effort. Fishery Bulletin, 70(3):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/exposition-definition-fishing-effort;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/rothschild%20(1).pdf.

Bauersfeld:1972:FSP

[2177] Paul E. Bauersfeld, Jr. and Joseph H. Soares, Jr. Fish solubles as a partial substitute for milk and milk by-products in liquid rations for neonatal animals. Fishery Bulletin, 70(4):1275–1279, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/fish-solubles-partial-substitute-milk-and-
milk-products-liquid-rations-neonatal-animals;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1972/704/bauersfeld.
pdf.

McCarthy:1972:UON

[2178] James J. McCarthy and Daniel Kamykowski. Urea and other nitrogenous nutrients in La Jolla Bay during February, March, and April 1970. Fishery Bulletin, 70(4):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/urea-and-other-nitrogenous-nutrients-
la-jolla-bay-during-february-march-and-april-1970;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1972/704/
mccarthy.pdf.

Carr:1972:FHJ

[2179] William E. S. Carr and Clayton A. Adams. Food habits of juvenile marine fishes: Evidence of the cleaning habit in the leatherjacket, Oligoplites saurus, and the spottail pinfish, Diplodus holbrooki. Fishery Bulletin, 70(4):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
food-habits-juvenile-marine-fishes-evidence-cleaning-habit-
leatherjacket-oligoplites-saurus;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1972/704/carr.pdf.

Olla:1972:AFB

[2180] Bori L. Olla, Carol E. Samet, and Anne L. Studholme. Activity and feeding behavior of the summer flounder (Paralichthys dentatus) under controlled laboratory conditions. Fishery Bulletin, 70(4):1127– 1136, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/activity-
and-feeding-behavior-summer-flounder-paralichthys-dentatus-
under-controlled;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1972/704/olla.pdf.

Condrey:1972:CAD

[2181] Richard E. Condrey, James G. Gosselink, and Harry J. Bennett. Comparison of the assimilation of different diets by Penaeus setiferus and P. aztecus. Fishery Bulletin, 70(4):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/comparison-assimilation-different-diets-penaeus-

setiferus-and-p-aztecus;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1972/704/condrey.pdf.

Lasker:1972:MOE

[2182] Reuben Lasker and Thomas A. Manar. In memoriam: Oscar Elton Sette. Fishery Bulletin, 70(4):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/memoriam-oscar-elton-sette;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/inmemoriam%
20(1).pdf.

Dame:1972:CVA

[2183] Richard F. Dame. Comparison of various allometric relationships in intertidal and subtidal American oysters. Fishery Bulletin, 70(4):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
comparison-various-allometric-relationships-intertidal-and-
subtidal-american-oysters;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1972/704/dame.pdf.

Murai:1972:GFC

[2184] Takeshi Murai and James W. Andrews. Growth and food conversion of rainbow trout reared in brackish and fresh water. Fishery Bulletin, 70(4): ??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/growth-and-
food-conversion-rainbow-trout-reared-brackish-and-fresh-water;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/murai.pdf.

DeVlaming:1972:ETP

[2185] Victor L. De Vlaming. The effects of temperature and photoperiod on reproductive cycling in the estuarine gobiid fish, Gillichthys mirabilis. Fishery Bulletin, 70(4):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
effects-temperature-and-photoperiod-reproductive-cycling-estuarine-
gobiid-fish-gillichthys;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1972/704/devlaming.pdf.

Horn:1972:ASA

[2186] Michael H. Horn. The amount of space available for marine and freshwater fishes. Fishery Bulletin, 70(4):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/amount-space-available-marine-and-freshwater-

fishes;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/horn.pdf.

Dragovich:1972:CSF

[2187] Alexander Dragovich and Thomas Potthoff. Comparative study of food of skipjack and yellowfin tunas off the coast of West Africa. Fishery Bulletin, 70(4):1087–1110, 1972. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/comparative-study-food-skipjack-
and-yellowfin-tunas-coast-west-africa;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1972/704/dragovich.pdf.

Fontaine:1972:SCL

[2188] C. T. Fontaine, S. E. P. Gislason, and W. L. Trent. A system for collecting large numbers of live postlarval shrimp. Fishery Bulletin, 70(4): 1298–1302, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/system-
collecting-large-numbers-live-postlarval-shrimp;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fontaine%20(1).pdf.

Eppley:1972:TPG

[2189] Richard W. Eppley. Temperature and phytoplankton growth in the sea. Fishery Bulletin, 70(4):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/temperature-and-phytoplankton-growth-sea;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1972/704/
eppley.pdf.

Anonymous:1972:I

[2190] Anonymous. Index. Fishery Bulletin, 70(4):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/index-102;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/index_
5.pdf.

Ahlstrom:1972:KAF

[2191] Elbert H. Ahlstrom. Kinds and abundance of fish larvae in the Eastern Tropical Pacific on the second multivessel eastropac survey, and observations on the annual cycle of larval abundance. Fishery Bulletin, 70(4):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/kinds-and-abundance-
fish-larvae-eastern-tropical-pacific-second-multivessel-eastropac;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1972/704/ahlstrom.pdf.

Kendall:1972:DBS

[2192] Arthur W. Kendall, Jr. Description of black sea bass, Centropristis striata (Linnaeus), larvae and their occurrences north of Cape Lookout, North Carolina, in 1966. Fishery Bulletin, 70(4):??, 1972. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/description-
black-sea-bass-centropristis-striata-linnaeus-larvae-and-their-
occurrences;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1972/704/kendall.pdf.

Hastings:1973:BPS

[2193] Robert W. Hastings. Biology of the pygmy sea bass, Serraniculus pumilio (Pisces: Serranidae). Fishery Bulletin, 71(1):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/biology-pygmy-
sea-bass-serraniculus-pumilio-pisces-serranidae;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1973/711/
hastings.pdf.

OHara:1973:ITS

[2194] James O’Hara. The influence of temperature and salinity on the toxicity of cadmium to the fiddler crab, Uca pugilator. Fishery Bulletin, 71(1):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/influence-
temperature-and-salinity-toxicity-cadmium-fiddler-crab-uca-
pugilator;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1973/711/ohara.pdf.

Venrick:1973:DMP

[2195] E. L. Venrick, J. A. McGowan, and A. W. Mantyla. Deep maxima of photosynthetic chlorophyll in the Pacific Ocean. Fishery Bulletin, 71(1):41–52, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/deep-
maxima-photosynthetic-chlorophyll-pacific-ocean;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1973/711/
venrick.pdf.

Hida:1973:FTD

[2196] Thomas S. Hida. Food of tunas and dolphins (Pisces: Scombridae and Coryphaenidae) with emphasis on the distribution and biology of their prey Stolephorus buccaneeri (Engraulidae). Fishery Bulletin, 71(1):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/food-tunas-and-dolphins-pisces-scombridae-
and-coryphaenidae-emphasis-distribution-and;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1973/711/hida.
pdf.

Pennell:1973:HSS

[2197] David A. Pennell, C. Dale Becker, and Nora R. Scofield. Helminths of sockeye salmon (Oncorhynchus nerka) from the Kvichak River System, Bristol Bay, Alaska. Fishery Bulletin, 71(1):??, 1973. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/helminths-sockeye-salmon-
oncorhynchus-nerka-kvichak-river-system-bristol-bay-alaska;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1973/711/pennell.pdf.

Wigley:1973:DMR

[2198] Roland L. Wigley and Frederick Charles Stinton. Distribution of macroscopic remains of recent animals from marine sediments off Massachusetts. Fishery Bulletin, 71(1):??, 1973. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/distribution-macroscopic-remains-
recent-animals-marine-sediments-massachusetts;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1973/711/wigley.
pdf.

Jackson:1973:RDT

[2199] Rodney G. Jackson and Martin Sage. Regional distribution of thyroid stimulating hormone activity in the pituitary gland of the Atlantic stingray, Dasyatis sabina. Fishery Bulletin, 71(1): ??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/regional-
distribution-thyroid-stimulating-hormone-activity-pituitary-
gland-atlantic;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1973/711/jackson.pdf.

Perkins:1973:LSD

[2200] Herbert C. Perkins. The larval stages of the deep sea red crab, Geryon quinquedens Smith, reared under laboratory conditions (Decapoda: Brachyrhyncha). Fishery Bulletin, 71(1):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/larval-stages-deep-sea-red-crab-
geryon-quinquedens-smith-reared-under-laboratory-conditions;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1973/711/perkins.pdf.

Gonor:1973:DLF

[2201] S. L. Gonor and J. J. Gonor. Descriptions of the larvae of four North Pacific Porcellanidae (Crustacea: Anomura). Fishery Bulletin, 71(1):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/descriptions-larvae-
four-north-pacific-porcellanidae-crustacea-anomura;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
gonor.pdf.

Beardsley:1973:DES

[2202] Alan J. Beardsley. Design and evaluation of a sampler for measuring the near-bottom vertical distribution of pink shrimp, Pandalus jordani. Fishery Bulletin, 71(1):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
design-and-evaluation-sampler-measuring-near-bottom-vertical-
distribution-pink-shrimp;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1973/711/beardsley.pdf.

Karnella:1973:SSM

[2203] Charles Karnella. The systematic status of Merluccius in the Tropical Western Atlantic Ocean including the Gulf of Mexico. Fishery Bulletin, 71(1):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
systematic-status-merluccius-tropical-western-atlantic-ocean-
including-gulf-mexico;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1973/711/karnella.pdf.

Sandifer:1973:ETS

[2204] Paul A. Sandifer. Effects of temperature and salinity on larval development of grass shrimp, Palaemonetes vulgaris (Decapoda, Caridea). Fishery Bulletin, 71(1):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/effects-temperature-and-salinity-
larval-development-grass-shrimp-palaemonetes-vulgaris;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1973/711/
sandifer.pdf.

Pristas:1973:ATH

[2205] Paul J. Pristas and Randall P. Cheek. Atlantic thread herring (Opisthonema oglinum) — movements and population size inferred from tag returns. Fishery Bulletin, 71(1):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/atlantic-thread-herring-opisthonema-oglinum-
movements-and-population-size-inferred-tag;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/pristas.
pdf.

Chenoweth:1973:FLE

[2206] Stanley B. Chenoweth. Fish larvae of the estuaries and coast of Central Maine. Fishery Bulletin, 71(1):??, 1973. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/fish-larvae-estuaries-and-coast-
central-maine;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1973/711/chenoweth.pdf.

Knight:1973:NIM

[2207] Margaret D. Knight. The nauplius II, metanauplius, and calyptopis stages of Thysanopoda tricuspidata Milne-Edwards (Euphausiacea). Fishery Bulletin, 71(1):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/nauplius-ii-metanauplius-and-calyptopis-stages-
thysanopoda-tricuspidata-milne-edwards;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1973/711/knight.pdf.

Sherburne:1973:EDA

[2208] Stuart W. Sherburne. Erythrocyte degeneration in the Atlantic herring, Clupea harengus harengus L. Fishery Bulletin, 71(1):??, 1973. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/erythrocyte-degeneration-
atlantic-herring-clupea-harengus-harengus-l;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1973/711/sherburne.
pdf.

Miller:1973:MCS

[2209] Grant L. Miller and Sherrell C. Jorgenson. Meristic characters of some marine fishes of the western Atlantic Ocean. Fishery Bulletin, 71(1):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/meristic-characters-some-marine-fishes-western-

atlantic-ocean;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/miller_0.pdf.

Dahlberg:1973:ESG

[2210] Michael D. Dahlberg and James C. Conyers. An ecological study of Gobiosoma bosci and G. ginsburgi (Pisces, Gobiidae) on the Georgia coast. Fishery Bulletin, 71(1):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/ecological-study-gobiosoma-bosci-and-
g-ginsburgi-pisces-gobiidae-georgia-coast;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1973/711/dahlberg.
pdf.

Krouse:1973:MSR

[2211] Jay S. Krouse. Maturity, sex ratio, and size composition of the natural population of American lobster, Homarus americanus, along the Maine Coast. Fishery Bulletin, 71(1):??, 1973. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/maturity-sex-ratio-
and-size-composition-natural-population-american-lobster-homarus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1973/711/krouse.pdf.

Soares:1973:RCN

[2212] Joseph Soares, Jr., David Miller, Susan Cuppett, and Paul Bauersfeld, Jr. A review of the chemical and nutritive properties of condensed fish solubles. Fishery Bulletin, 71(1):??, 1973. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/review-chemical-and-
nutritive-properties-condensed-fish-solubles;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1973/711/soares.
pdf.

Dubrow:1973:EDD

[2213] David L. Dubrow. Effect of drying and desolventizing on the functional properties of fish protein concentrate (FPC). Fishery Bulletin, 71(1): ??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/effect-drying-
and-desolventizing-functional-properties-fish-protein-concentrate-
fpc;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1973/711/dubrow.pdf.

LeGuen:1973:AGY

[2214] J. C. Le Guen and Gary T. Sakagawa. Apparent growth of yellowfin tuna from the Eastern Atlantic Ocean. Fishery Bulletin, 71 (1):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/apparent-
growth-yellowfin-tuna-eastern-atlantic-ocean;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1973/711/leguen.
pdf.

Taylor:1973:HRT

[2215] John L. Taylor, Carl H. Saloman, and Kenneth W. Prest, Jr. Harvest and regrowth of turtle grass (Thalassia testudinum) in Tampa Bay, Florida. Fishery Bulletin, 71(1):??, 1973. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/harvest-and-regrowth-
turtle-grass-thalassia-testudinum-tampa-bay-florida;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1973/711/
taylor.pdf.

Gonor:1973:FCS

[2216] S. L. Gonor and J. J. Gonor. Feeding, cleaning, and swimming behavior in larval stages of porcellanid crabs (Crustacea: Anomura). Fishery Bulletin, 71(1):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
feeding-cleaning-and-swimming-behavior-larval-stages-porcellanid-
crabs-crustacea-anomura;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1973/711/gonor1.pdf.

Lindall:1973:FMH

[2217] William N. Lindall, Jr., John R. Hall, and Carl H. Saloman. Fishes, macroinvertebrates, and hydrological conditions of upland canals in Tampa Bay, Florida. Fishery Bulletin, 71(1):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/fishes-macroinvertebrates-and-hydrological-
conditions-upland-canals-tampa-bay-florida;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1973/711/lindall.pdf.

Thayer:1973:CMS

[2218] Gordon W. Thayer, William E. Schaaf, Joseph W. Angelovic, and Michael W. Lacroix. Caloric measurements of some estuarine organisms. Fishery Bulletin, 71(1):289–296, 1973. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/caloric-measurements-
some-estuarine-organisms;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1973/711/thayer.pdf.

Becker:1973:FGP

[2219] C. Dale Becker. Food and growth parameters of juvenile Chinook salmon, Oncorhynchus tshawytscha, in Central Columbia River. Fishery Bulletin, 71(2):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/food-
and-growth-parameters-juvenile-chinook-salmon-oncorhynchus-
tshawytscha-central;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1973/712/becker.pdf.

Lux:1973:AGW

[2220] Fred E. Lux. Age and growth of the winter flounder, Pseudopleuronectes americanus, on Georges Bank. Fishery Bulletin, 71(2):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/age-and-growth-winter-flounder-
pseudopleuronectes-americanus-georges-bank;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1973/712/lux.pdf.

Weisel:1973:PWP

[2221] G. F. Weisel, D. A. Hanzel, and R. L. Newell. The pygmy whitefish, Prosopium coulteri, in western Montana. Fishery Bulletin, 71 (2):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/pygmy-
whitefish-prosopium-coulteri-western-montana;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1973/712/weisel.
pdf.

Chao:1973:DSF

[2222] Labbish Ning Chao. Digestive system and feeding habits of the Cunner, Tautogolabrus adspersus, a stomachless fish. Fishery Bulletin, 71(2): 565–586, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/digestive-
system-and-feeding-habits-cunner-tautogolabrus-adspersus-stomachless-
fish;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1973/712/chao.pdf.

Magnuson:1973:CSA

[2223] John J. Magnuson. Comparative study of adaptations for continuous swimming and hydrostatic equilibrium of scombroid and xiphoid fishes. Fishery Bulletin, 71(2):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/comparative-study-adaptations-continuous-swimming-and-
hydrostatic-equilibrium-scombroid-and;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1973/712/magnuson.pdf.

Richards:1973:DSL

[2224] Sarah W. Richards and Arthur W. Kendall, Jr. Distribution of sand lance, Ammodytes sp., larvae on the Continental Shelf from Cape Cod to Cape Hatteras from RV Dolphin surveys in 1966. Fishery Bulletin, 71(2): ??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/distribution-
sand-lance-ammodytes-sp-larvae-continental-shelf-cape-cod-cape-
hatteras-rv;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/richards.pdf.

Clarke:1973:SAE

[2225] Thomas A. Clarke. Some aspects of the ecology of lanternfishes (Myctophidae) in the Pacific Ocean near Hawaii. Fishery Bulletin, 71(2):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/some-aspects-ecology-lanternfishes-myctophidae-
pacific-ocean-near-hawaii;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1973/712/clarke.pdf.

Musick:1973:MMC

[2226] J. A. Musick. A meristic and morphometric comparison of the hakes, Urophycis chuss and U. tenuis (Pisces, Gadidae). Fishery Bulletin, 71(2): ??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/meristic-and-
morphometric-comparison-hakes-urophycis-chuss-and-u-tenuis-
pisces-gadidae;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1973/712/musick.pdf.

Sherburne:1973:CTDa

[2227] Stuart W. Sherburne. Cell types, differential cell counts, and blood cell measurements of a Portuguese shark, Centroscymnus coelolepis, captures at 700 fathoms. Fishery Bulletin, 71(2):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/cell-types-
differential-cell-counts-and-blood-cell-measurements-portuguese-
shark-0;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/sherburne.pdf.

Cooper:1973:VDP

[2228] James L. Cooper. Vertical distribution patterns of goldeye, Hiodon alosoides, in Fort Peck Reservoir, Montana. Fishery Bulletin, 71(2): ??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/vertical-
distribution-patterns-goldeye-hiodon-alosoides-fort-peck-reservoir-
montana;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1973/712/cooper.pdf.

Pearcy:1973:AOO

[2229] William G. Pearcy. Albacore oceanography off Oregon — 1970. Fishery Bulletin, 71(2):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
albacore-oceanography-oregon-1970;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1973/712/pearcy.pdf.

Buchanan:1973:OMR

[2230] Chester C. Buchanan. Occurrence of mature redfish, Sebastes marinus, in sport fishery of the New York Bight. Fishery Bulletin, 71(2): ??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/occurrence-
mature-redfish-sebastes-marinus-sport-fishery-new-york-bight;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/buchanan.pdf.

Dailey:1973:HPP

[2231] Murray D. Dailey and William F. Perrin. Helminth parasites of porpoises of the genus Stenella in the Eastern Tropical Pacific, with descriptions of two new species: Mastigonema stenellae gen. et sp. n. (Nematoda: Spiruroidea) and Zalophotrema pacificum sp. n. (Trematoda: Digenea). Fishery Bulletin, 71(2):??, 1973. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/helminth-parasites-
porpoises-genus-stenella-eastern-tropical-pacific-descriptions-
two-new;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1973/712/dailey.pdf.

Farfante:1973:MSD

[2232] Isabel Perez Farfante. Morphological study of diagnostic characters in Western Atlantic Hepomadus (Crustacea, Decapoda, Penaeidae). Fishery Bulletin, 71(2):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
morphological-study-diagnostic-characters-western-atlantic-
hepomadus-crustacea-decapoda;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1973/712/perez.pdf.

Thomas:1973:OLT

[2233] William H. Thomas, Anne N. Dodson, and Carol A. Linden. Optimum light and temperature requirements for Gymnodinium splendens, a larval fish food organism. Fishery Bulletin, 71(2):599–601, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/optimum-light-
and-temperature-requirements-gymnodinium-splendens-larval-fish-
food-organism;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/thomas_0.pdf.

Ebel:1973:ETS

[2234] Wesley J. Ebel, Donn L. Park, and Richard C. Johnsen. Effects of transportation on survival and homing of Snake River Chinook salmon and steelhead trout. Fishery Bulletin, 71(2):??, 1973. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/effects-transportation-survival-
and-homing-snake-river-chinook-salmon-and-steelhead-trout;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1973/
712/ebel.pdf.

Sherburne:1973:CTDb

[2235] Stuart W. Sherburne. Cell types, differential cell counts, and blood cell measurements of a Portuguese shark, Centroscymnus coelolepis, captured at 700 fathoms. Fishery Bulletin, 71(2): ??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/cell-
types-differential-cell-counts-and-blood-cell-measurements-
portuguese-shark;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1973/712/sherburne.pdf.

Gulland:1973:SAC

[2236] J. A. Gulland and L. K. Boerema. Scientific advice on catch levels. Fishery Bulletin, 71(2):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/scientific-advice-catch-levels;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1973/712/
gulland.pdf.

Smith:1973:DSF

[2237] W. G. Smith. The distribution of summer flounder, Paralichthys dentatus, eggs and larvae on the Continental Shelf between Cape Cod and Cape Lookout, 1965–66. Fishery Bulletin, 71(2):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/distribution-summer-
flounder-paralichthys-dentatus-eggs-and-larvae-continental-
shelf;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1973/712/smith.pdf.

Lander:1973:PBM

[2238] Robert H. Lander. Problem of bias in models to approximate ocean mortality, maturity, and abundance schedules of salmon from known smolts and returns. Fishery Bulletin, 71(2):??, 1973. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/problem-bias-models-approximate-
ocean-mortality-maturity-and-abundance-schedules-salmon;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1973/712/
lander.pdf.

Valentine:1973:AAF

[2239] David W. Valentine, Michael E. Soule, and Paul Samollow. Asymmetry analysis in fishes: a possible statistical indicator of environmental stress. Fishery Bulletin, 71(2):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/asymmetry-analysis-fishes-possible-statistical-
indicator-environmental-stress;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1973/712/valentine.pdf.

Dryfoos:1973:PAA

[2240] Robert L. Dryfoos, Randall P. Cheek, and Richard L. Kroger. Preliminary analyses of Atlantic menhaden, Brevoortia tyrannus, migrations, population structure, survival and exploitation rates, and availability as indicated from tag returns. Fishery Bulletin,

71(3):??, 1973.  CODEN FSYBAY.  ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
preliminary-analyses-atlantic-menhaden-brevoortia-tyrannus-
migrations-population-structure;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1973/713/dryfoos.pdf.

Peterson:1973:MTR

[2241] C. L. Peterson, W. L. Klawe, and G. D. Sharp. Mercury in tunas: a review. Fishery Bulletin, 71(3):??, 1973. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/mercury-tunas-review;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1973/713/
peterson.pdf.

Dizon:1973:OER

[2242] Andrew E. Dizon, Ross M. Horrall, and Arthur D. Hasler. Olfactory electroencephalographic responses of homing Coho salmon, Oncorhynchus kisutch, to water conditioned by conspecifics. Fishery Bulletin, 71(3):893–896, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/olfactory-electroencephalographic-responses-homing-
coho-salmon-oncorhynchus-kisutch-water;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/dizon.pdf.

Finch:1973:ERG

[2243] Roland Finch. Effects of regulatory guidelines on the intake of mercury from fish — the Mecca Project. Fishery Bulletin,

71(3):??, 1973.  CODEN FSYBAY.  ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
effects-regulatory-guidelines-intake-mercury-fish%E2%80%94-
mecca-project;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1973/713/finch.pdf.

Richardson:1973:LYW

[2244] Sally L. Richardson and Edwin B. Joseph. Larvae and young of Western North Atlantic bothid flatfishes Etropus microstomus and Citharichthys arctifrons in the Chesapeake Bight. Fishery Bulletin, 71 (3):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/larvae-
and-young-western-north-atlantic-bothid-flatfishes-etropus-
microstomus-and;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1973/713/richardson1.pdf.

Barr:1973:TAW

[2245] Louis Barr and James R. Knull. A temporarily anoxic water mass in an Alaska estuary. Fishery Bulletin, 71(3):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/temporarily-anoxic-water-mass-alaska-
estuary;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/barr.pdf.

French:1973:LSH

[2246] Robert R. French and J. Richard Dunn. Loss of salmon from high-seas gillnetting with reference to the Japanese salmon mothership fishery. Fishery Bulletin, 71(3):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/loss-salmon-high-seas-gillnetting-reference-japanese-
salmon-mothership-fishery;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1973/713/french.pdf.

Struhsaker:1973:ABN

[2247] Paul Struhsaker. Argyripnus brocki, a new species of stomiatoid fish from Hawaii, with observations on A. ephippiatus and A. iridescens. Fishery Bulletin, 71(3):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
argyripnus-brocki-new-species-stomiatoid-fish-hawaii-observations-
ephippiatus-and;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1973/713/struhsaker.pdf.

Yoshida:1973:TRT

[2248] Howard O. Yoshida. Taractes rubescens and Taractichthys steindachneri from Hawaiian waters. Fishery Bulletin, 71(3):900–902, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/taractes-rubescens-
and-taractichthys-steindachneri-hawaiian-waters;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
yoshida%20(2).pdf.

Haynes:1973:DPS

[2249] Evan Haynes. Descriptions of prezoeae and Stage I zoeae of Chionoecetes bairdi and C. Opilio (Oxyrhyncha, Oregoniinae). Fishery Bulletin, 71(3): ??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/descriptions-
prezoeae-and-stage-i-zoeae-chionoecetes-bairdi-and-c-opilio-
oxyrhyncha;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1973/713/haynes.pdf.

Sutherland:1973:DSA

[2250] Doyle F. Sutherland. Distribution, seasonal abundance, and some biological features of steelhead trout, Salmo gairdneri, in the North Pacific Ocean. Fishery Bulletin, 71(3):??, 1973. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/distribution-seasonal-abundance-

and-some-biological-features-steelhead-trout-salmo;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1973/713/
sutherland.pdf.

Peterson:1973:UIA

[2251] William Thornton Peterson. Upwelling indices and annual catches of Dungeness crab, Cancer magister, along the west coast of the United States. Fishery Bulletin, 71(3):902–??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/upwelling-indices-and-annual-catches-
dungeness-crab-cancer-magister-along-west-coast-united;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
peterson1.pdf.

Higley:1973:EPJ

[2252] Duane L. Higley and Carl E. Bond. Ecology and production of juvenile spring Chinook salmon, Oncorhynchus tshawytscha, in a eutrophic reservoir. Fishery Bulletin, 71(3):??, 1973. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/ecology-and-production-juvenile-
spring-chinook-salmon-oncorhynchus-tshawytscha-eutrophic;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1973/
713/higley.pdf.

Thorne:1973:DHD

[2253] Richard E. Thorne. Digital hydroacoustic data-processing system and its application to Pacific hake stock assessment in Port Susan, Washington. Fishery Bulletin, 71(3):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/digital-hydroacoustic-data-processing-system-and-its-
application-pacific-hake-stock;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1973/713/thorne.pdf.

Hobson:1973:FOM

[2254] Edmund S. Hobson and James R. Chess. Feeding oriented movements of the atherinid fish Pranesus pinguis at Majuro Atoll, Marshall Islands. Fishery Bulletin, 71(3):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/feeding-oriented-movements-atherinid-fish-pranesus-
pinguis-majuro-atoll-marshall-islands;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1973/713/hobson.pdf.

VanCleve:1973:ECD

[2255] Richard Van Cleve and Donald E. Bevan. Evaluation of causes for the decline of the Karluk sockeye salmon runs and recommendations for rehabilitation. Fishery Bulletin, 71(3):??, 1973. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/evaluation-causes-decline-karluk-
sockeye-salmon-runs-and-recommendations-rehabilitation;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1973/713/
vancleve.pdf.

Lander:1973:SMA

[2256] Robert H. Lander and Kenneth A. Henry. Survival, maturity, abundance, and marine distribution of 1965–66 brood Coho salmon, Oncorhynchus kisutch, from Columbia River hatcheries. Fishery Bulletin, 71(3):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/survival-
maturity-abundance-and-marine-distribution-1965-66-brood-coho-
salmon-oncorhynchus;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1973/713/lander.pdf.

VanCleve:1973:RRC

[2257] Richard Van Cleve and Donald E. Bevan. Reply to Rounsefell’s ‘Comments on ‘evaluation of Causes for the Decline of the Karluk Sockeye Salmon Runs and Recommendations for Rehabilitation,’ by R. Van Cleve and D. E. Bevan’. Fishery Bulletin, 71 (3):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/reply-
rounsefells-comments-evaluation-causes-decline-karluk-sockeye-
salmon-runs-and;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1973/713/vancleve1.pdf.

Childs:1973:MCO

[2258] E. A. Childs and J. N. Gaffke. Mercury content of Oregon groundfish. Fishery Bulletin, 71(3):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/mercury-content-oregon-groundfish;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1973/713/childs.pdf.

Paul:1973:GRD

[2259] A. J. Paul and Howard M. Feder. Growth, recruitment, and distribution of the littleneck clam, Protothaca staminea, in Galena Bay, Prince William Sound, Alaska. Fishery Bulletin, 71(3):??, 1973.

CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/growth-recruitment-
and-distribution-littleneck-clam-protothaca-staminea-galena-
bay-prince;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1973/713/paul.pdf.

Richardson:1973:ADL

[2260] Sally L. Richardson. Abundance and distribution of larval fishes in waters off Oregon, May–October 1969, with special emphasis on the northern anchovy, Engraulis mordax. Fishery Bulletin, 71(3):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
abundance-and-distribution-larval-fishes-waters-oregon-may-
october-1969-special-emphasis;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/richardson.pdf.

Fleminger:1973:PNV

[2261] Abraham Fleminger. Pattern, number, variability, and taxonomic significance of integumental organs (sensilla and glandular pores) in the genus Eucalanus (Copepoda, Calanoida). Fishery Bulletin, 71 (4):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/pattern-
number-variability-and-taxonomic-significance-integumental-
organs-sensilla-and;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1973/714/fleminger.pdf.

Rogers:1973:ASJ

[2262] Donald E. Rogers. Abundance and size of juvenile sockeye salmon, Oncorhynchus nerka, and associated species in Lake Aleknagik, Alaska, in relation to their environment. Fishery Bulletin, 71(4):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/abundance-and-size-
juvenile-sockeye-salmon-oncorhynchus-nerka-and-associated-species-
lake;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1973/714/rogers.pdf.

Fox:1973:GLH

[2263] William W. Fox, Jr. A general life history exploited population simulator with pandalid shrimp as an example. Fishery Bulletin, 71(4):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/general-life-history-exploited-population-

simulator-pandalid-shrimp-example;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1973/714/fox.pdf.

Sherburne:1973:DBC

[2264] Stuart W. Sherburne. Differential blood cell counts of Atlantic herring, Clupea harengus harengus. Fishery Bulletin, 71(4):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/differential-blood-cell-counts-
atlantic-herring-clupea-harengus-harengus;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1973/714/sherburne.
pdf.

Griswold:1973:LHT

[2265] Bernard L. Griswold and Lloyd L. Smith, Jr. The life history and trophic relationship of the ninespine stickleback, Pungitius pungitius, in the Apostle Islands area of Lake Superior. Fishery Bulletin, 71(4):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/life-history-and-trophic-
relationship-ninespine-stickleback-pungitius-pungitius-apostle;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1973/714/griswold.pdf.

Valentine:1973:EPP

[2266] David W. Valentine and Michael Soule. Effect of p,p’-DDT on developmental stability of pectoral fin rays in the grunion, Leuresthes tenuis. Fishery Bulletin, 71(4):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/effect-pp-ddt-developmental-stability-pectoral-fin-
rays-grunion-leuresthes-tenuis;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1973/714/valentine.pdf.

Hirth:1973:SGK

[2267] Harold F. Hirth, Lionel G. Klikoff, and K. T. Harper. Sea grasses at Khor Umaira, People’s Democratic Republic of Yemen with reference to their role in the diet of the green turtle, Chelonia mydas. Fishery Bulletin, 71(4):??, 1973. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/sea-grasses-khor-umaira-peoples-
democratic-republic-yemen-reference-their-role-diet-green;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1973/
714/hirth.pdf.

Barham:1973:AEA

[2268] Eric G. Barham, Robert W. Gowdy, and Fay H. Wolfson. Acanthaster (Echinodermata, Asteroidea) in the Gulf of California. Fishery Bulletin, 71(4):927–942, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/acanthaster-echinodermata-asteroidea-gulf-
california;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/barham.pdf.

Hughes:1973:SMP

[2269] Steven E. Hughes. Some metazoan parasites of the Eastern Pacific saury, Cololabis saira. Fishery Bulletin, 71(4):??, 1973. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/some-metazoan-parasites-eastern-
pacific-saury-cololabis-saira;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1973/714/hughes.pdf.

Barr:1973:FTS

[2270] Louis Barr, Robert J. Ellis, and John H. Helle. Fossil tree stumps found in situ on submerged ridge at Amchitka Island, Alaska. Fishery Bulletin, 71(4):1099–1103, 1973. CODEN FSY-

BAY.  ISSN 0090-0656 (print), 1937-4518 (electronic).  URL

https://spo.nmfs.noaa.gov/content/fossil-tree-stumps-found-
situ-submerged-ridge-amchitka-island-alaska;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/barr%
20(1).pdf.

Laroche:1973:AGR

[2271] Joanne L. Laroche and Jackson Davis. Age, growth, and reproduction of the northern puffer, Sphoeroides maculatus. Fishery Bulletin, 71(4): ??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/age-growth-
and-reproduction-northern-puffer-sphoeroides-maculatus;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1973/714/
laroche.pdf.

Lenarz:1973:ERT

[2272] W. H. Lenarz, F. J. Mather III, J. S. Beckett, A. C. Jones, and J. M. Mason, Jr. Estimation of rates of tag shedding by northwest Atlantic bluefin tuna. Fishery Bulletin, 71(4):??, 1973. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/estimation-rates-tag-
shedding-northwest-atlantic-bluefin-tuna;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/lenarz%20(
1).pdf.

Lord:1973:COD

[2273] Gary E. Lord. Characterization of the optimum data acquisition and management of a salmon fishery as a stochastic dynamic program. Fishery Bulletin, 71(4):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/characterization-optimum-data-acquisition-
and-management-salmon-fishery-stochastic-dynamic;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1973/714/
lord.pdf.

Parks:1973:DAS

[2274] Norman B. Parks. Distribution and abundance of the sand dollar, Dendraster excentricus, off the coast of Oregon and Washington. Fishery Bulletin, 71(4):??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/distribution-and-abundance-sand-
dollar-dendraster-excentricus-coast-oregon-and-washington;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/parks.pdf.

Bell:1973:PGU

[2275] Frederick W. Bell and Richard K. Kinoshita. Productivity gains in U.S. fisheries. Fishery Bulletin, 71(4):??, 1973. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/productivity-gains-us-fisheries;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1973/714/bell.pdf.

Perrin:1973:SCP

[2276] W. F. Perrin, R. R. Warner, C. H. Fiscus, and D. B. Holts. Stomach contents of porpoise, Stenella spp., and yellowfin tuna, Thunnus albacares, in mixed-species aggregations. Fishery Bulletin, 71(4): ??, 1973. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/stomach-
contents-porpoise-stenella-spp-and-yellowfin-tuna-thunnus-albacares-
mixed-species;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1973/714/perrin.pdf.

Fullenbaum:1974:SBF

[2277] Richard F. Fullenbaum and Frederick W. Bell. A simple bioeconomic fishery management model: a case study of the American lobster fishery. Fishery Bulletin, 72(1):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/simple-bioeconomic-fishery-management-model-case-study-
american-lobster-fishery;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1974/721/fullenbaum.pdf.

Olla:1974:DAM

[2278] Bori L. Olla, Allen J. Bejda, and A. Dale Martin. Daily activity, movements, feeding, and seasonal occurrence in the tautog, Tautoga onitis. Fishery Bulletin, 72(1):27–35, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/daily-activity-movements-feeding-and-
seasonal-occurrence-tautog-tautoga-onitis;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/olla.pdf.

Grant:1974:ACS

[2279] George C. Grant. The age composition of striped bass catches in Virginia rivers, 1967–1971, and a description of the fishery. Fishery Bulletin, 72(1):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/age-
composition-striped-bass-catches-virginia-rivers-1967-1971-
and-description-fishery;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1974/721/grant.pdf.

Fleminger:1974:SDF

[2280] A. Fleminger and K. Hulsemann. Systematics and distribution of the four sibling species comprising the genus Pontellina dana (Copepoda, Calanoida). Fishery Bulletin, 72(1):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/systematics-and-distribution-four-sibling-
species-comprising-genus-pontellina-dana-copepoda;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fleminger.pdf.

Lenarz:1974:EYP

[2281] W. H. Lenarz, W. W. Fox, Jr., G. T. Sakagawa, and B. J. Rothschild. An examination of the yield per recruit basis for a minimum size regulation for Atlantic yellowfin tuna, Thunnus albacares. Fishery Bulletin, 72(1):??, 1974. CODEN FSYBAY.

ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/examination-yield-recruit-basis-
minimum-size-regulation-atlantic-yellowfin-tuna-thunnus;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1974/721/
lenarz.pdf.

Hughes:1974:SCG

[2282] Steven E. Hughes. Stock composition, growth, mortality, and availability of Pacific saury, Cololabis saira, of the northeastern Pacific Ocean. Fishery Bulletin, 72(1):??, 1974. CODEN FSY-

BAY.  ISSN 0090-0656 (print), 1937-4518 (electronic).  URL

https://spo.nmfs.noaa.gov/content/stock-composition-growth-
mortality-and-availability-pacific-saury-cololabis-saira;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/hughes.pdf.

Lightner:1974:NPC

[2283] Donald V. Lightner. Normal postmortem changes in the brown shrimp, Penaeus aztecus. Fishery Bulletin, 72(1):??, 1974. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/normal-postmortem-changes-brown-
shrimp-penaeus-aztecus;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1974/721/lightner.pdf.

Anas:1974:HMN

[2284] Raymond E. Anas. Heavy metals in the northern fur seal, Callorhinus ursinnus, and harbor seal, Phoca citulina richardi. Fishery Bulletin, 72 (1):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/heavy-metals-
northern-fur-seal-callorhinus-ursinnus-and-harbor-seal-phoca-
citulina-richardi;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/anas.pdf.

Struhsaker:1974:BTF

[2285] Paul Struhsaker and Robert M. Moncrief. Bothus thompsoni (Fowler) 1923, a valid species of flatfish (Pisces: Bothidae) from the Hawaiian Islands. Fishery Bulletin, 72(1):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/bothus-thompsoni-fowler-1923-valid-species-flatfish-
pisces-bothidae-hawaiian-islands;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1974/721/struhsaker.pdf.

Powell:1974:AMK

[2286] Guy C. Powell, Kenneth E. James, and Charles L. Hurd. Ability of male king crab, Paralithodes camtschatica, to mate repeatedly, Kodiak, Alaska, 1973. Fishery Bulletin, 72(1):171–179, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/ability-male-
king-crab-paralithodes-camtschatica-mate-repeatedly-kodiak-
alaska-1973;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/powell%20(1).pdf.

Wahle:1974:BCC

[2287] Roy J. Wahle, Robert R. Vreeland, and Robert H. Lander. Bioeconomic contribution of Columbia River hatchery Coho salmon, 1965 and 1966 broods, to the Pacific salmon fisheries. Fishery Bulletin, 72(1):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/bioeconomic-contribution-
columbia-river-hatchery-coho-salmon-1965-and-1966-broods-pacific;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1974/721/wahle.pdf.

Pearcy:1974:LFY

[2288] William G. Pearcy and Sharon S. Myers. Larval fishes of Yaquina Bay, Oregon: a nursery ground for marine fishes? Fishery Bulletin, 72(1): ??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/larval-fishes-
yaquina-bay-oregon-nursery-ground-marine-fishes;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
pearcy_1.pdf.

Wickham:1974:EMW

[2289] Donald A. Wickham and Gary M. Russell. An evaluation of mid-water artificial structures for attracting coastal pelagic fishes. Fishery Bulletin, 72 (1):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/evaluation-
mid-water-artificial-structures-attracting-coastal-pelagic-
fishes;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1974/721/wickham.pdf.

Park:1974:CCG

[2290] Taisoo Park. Calanoid copepods of the genus Actideus from the Gulf of Mexico. Fishery Bulletin, 72(1):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/calanoid-copepods-genus-actideus-gulf-mexico;

https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/park%20(1).pdf.

Kittredge:1974:CSS

[2291] J. S. Kittredge, Francis T. Takahashi, James Lindsey, and Reuben Lasker. Chemical signals in the sea: Marble allelochemics and evolution. Fishery Bulletin, 72(1):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/chemical-signals-sea-marble-allelochemics-
and-evolution;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/kittredge.pdf.

Cable:1974:DEE

[2292] Wayne D. Cable and Warren S. Landers. Development of eggs and embryos of the surf clam,Spisula solidissima, in synthetic seawater. Fishery Bulletin, 72(1):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
development-eggs-and-embryos-surf-clamspisula-solidissima-synthetic-
seawater;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/cable.pdf.

Adron:1974:RPP

[2293] J. W. Adron, A. Blair, and C. B. Cowey. Rearing of plaice (Pleuronectes platessa) larvae to metamorphosis using an artificial diet. Fishery Bulletin, 72(2):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/rearing-plaice-pleuronectes-platessa-
larvae-metamorphosis-using-artificial-diet;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1974/722/adron.pdf.

Kemmerer:1974:SSE

[2294] Andrew J. Kemmerer, Joseph A. Benigno, Gladys B. Reese, and Frederick C. Minkler. Summary of selected early results from the ERTS-1 menhaden experiment. Fishery Bulletin, 72(2):375– 389, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/summary-
selected-early-results-erts-1-menhaden-experiment;   https:
//spo.nmfs.noaa.gov/content/summary-selected-early-results-
erts-1-menhaden-experiment-0;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/kemmerer.pdf.

Stickney:1974:FHG

[2295] Robert R. Stickney, Gary L. Taylor, and Richard W. Hearda III. Food habits of Georgia estuarine fishes I. Four species of flounders (Pleuronectiformes: Bothidae). Fishery Bulletin, 72(2):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/food-habits-georgia-estuarine-fishes-
i-four-species-flounders-pleuronectiformes-bothidae;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1974/722/
stickney.pdf.

Aprieto:1974:EDF

[2296] Virginia L. Aprieto. Early development of five carangid fishes of the Gulf of Mexico and the South Atlantic Coast of the United States. Fishery Bulletin, 72(2):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/early-development-five-carangid-fishes-gulf-mexico-
and-south-atlantic-coast-united-states;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1974/722/aprieto.pdf.

MacGregor:1974:CAP

[2297] John S. MacGregor. Changes in the amount and proportions of DDT and its metabolites, DDE and DDD, in the marine environment off Southern California, 1949–72. Fishery Bulletin, 72(2): ??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/changes-
amount-and-proportions-ddt-and-its-metabolites-dde-and-ddd-
marine-environment;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1974/722/macgregor.pdf.

Ben-Yami:1974:ISU

[2298] M. Ben-Yami and T. Glaser. The invasion of Saurida undosquamis (Richardson) into the Levant Basin — an example of biological effect of interoceanic canals. Fishery Bulletin, 72(2):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/invasion-saurida-undosquamis-richardson-
levant-basin-example-biological-effect-interoceanic;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1974/722/
benyami.pdf.

Miller:1974:DEL

[2299] John M. Miller and Barbara Y. Sumida. Development of eggs and larvae of Caranx mate (Carangidae). Fishery Bulletin, 72(2):??, 1974.

CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/development-eggs-and-
larvae-caranx-mate-carangidae;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1974/722/miller.pdf.

Alvarino:1974:DSR

[2300] Angeles Alvarino. Distribution of siphonophores in the regions adjacent to the Suez and Panama Canals. Fishery Bulletin, 72(2):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/distribution-
siphonophores-regions-adjacent-suez-and-panama-canals;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
alvarino%20(1).pdf.

Clarke:1974:SAE

[2301] Thomas A. Clarke. Some aspects of the ecology of stomiatoid fishes in the Pacific Ocean near Hawaii. Fishery Bulletin, 72(2):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/some-aspects-ecology-
stomiatoid-fishes-pacific-ocean-near-hawaii;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1974/722/clarke.
pdf.

Moser:1974:RLS

[2302] H. Geoffrey Moser and Elbert H. Ahlstrom. Role of larval stages in systematic investigations of marine teleosts: the Myctophidae, a case study. Fishery Bulletin, 72(2):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/role-larval-stages-systematic-investigations-marine-
teleosts-myctophidae-case-study;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1974/722/moser.pdf.

Baldridge:1974:MGW

[2303] Alan Baldridge. Migrant gray whales with calves and sexual behavior of gray whales in the Monterey area of central California, 1967–73. Fishery Bulletin, 72(2):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
migrant-gray-whales-calves-and-sexual-behavior-gray-whales-
monterey-area-central-california;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/baldridge.pdf.

French:1974:NMO

[2304] Robert R. French and Richard G. Bakkala. A new model of ocean migrations of Bristol Bay sockeye salmon. Fishery Bulletin, 72(2): ??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/new-model-
ocean-migrations-bristol-bay-sockeye-salmon;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1974/722/french.
pdf.

Musick:1974:SDS

[2305] John A. Musick. Seasonal distribution of sibling hakes, Urophycis chuss and U. tenuis (Pisces, Gadidae) in New England. Fishery Bulletin, 72 (2):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/seasonal-
distribution-sibling-hakes-urophycis-chuss-and-u-tenuis-pisces-
gadidae-new-england;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1974/722/musick.pdf.

Brooks:1974:NFE

[2306] A. L. Brooks, C. L. Brown, Jr., and P. H. Scully-Power. Net filtering efficiency of a 3-meter Isaacs–Kidd midwater trawl. Fishery Bulletin, 72(2): ??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/net-filtering-
efficiency-3-meter-isaacs-kidd-midwater-trawl;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/brooks.
pdf.

Hirota:1974:QNH

[2307] Jed Hirota. Quantitative natural history of Pleurobrachia bachei in La Jolla Bight. Fishery Bulletin, 72(2):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/quantitative-natural-history-pleurobrachia-
bachei-la-jolla-bight;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1974/722/hirota.pdf.

Potthoff:1974:ODV

[2308] Thomas Potthoff. Osteological development and variation in young tunas, genus Thunnus (Pisces, Scombridae), from the Atlantic Ocean. Fishery Bulletin, 72(2):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/osteological-development-and-variation-young-tunas-
genus-thunnus-pisces-scombridae-atlantic;   https://spo.nmfs.

noaa.gov/sites/default/files/pdf-content/1974/722/potthoff.
pdf.

Dow:1974:ALT

[2309] Robert L. Dow. American lobsters tagged by Maine commercial fishermen, 1957–59. Fishery Bulletin, 72(2):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/american-lobsters-tagged-maine-commercial-
fishermen-1957-59;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/dow.pdf.

Kanwisher:1974:ATF

[2310] John Kanwisher, Kenneth Lawson, and Gunnar Sundnes. Acoustic telemetry from fish. Fishery Bulletin, 72(2):??, 1974. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/acoustic-telemetry-fish;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1974/722/
kanwisher.pdf.

Raju:1974:TNS

[2311] Solomon N. Raju. Three new species of the genus Monognathus and the leptocephali of the order Saccopharyngiformes. Fishery Bulletin, 72(2):547–562, 1974. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
three-new-species-genus-monognathus-and-leptocephali-order-
saccopharyngiformes;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1974/722/raju.pdf.

Kaplan:1974:SED

[2312] Eugene H. Kaplan, J. R. Welker, and M. Gayle Kraus. Some effects of dredging on populations of macrobenthic organisms. Fishery Bulletin, 72(2):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/some-effects-dredging-populations-macrobenthic-
organisms;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1974/722/kaplan.pdf.

Soutar:1974:APF

[2313] Andrew Soutar and John D. Isaacs. Abundance of pelagic fish during the 19th and 20th Centuries as recorded in anaerobic sediment off the Californias. Fishery Bulletin, 72(2):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/abundance-pelagic-fish-during-19th-and-20th-centuries-

recorded-anaerobic-sediment;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1974/722/soutar.pdf.

Fahay:1974:OSH

[2314] Michael P. Fahay. Occurrence of silver hake, Merluccius bilinearis, eggs and larvae along the Middle Atlantic Continental Shelf during 1966. Fishery Bulletin, 72(3):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
occurrence-silver-hake-merluccius-bilinearis-eggs-and-larvae-
along-middle-atlantic;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1974/723/fahay.pdf.

Williams:1974:DVS

[2315] Austin B. Williams, Thomas E. Bowman, and David M. Damkaer. Distribution, variation, and supplemental description of the opossum shrimp, Neomysis americana (Crustacea: Mysidacea). Fishery Bulletin, 72(3): ??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/distribution-
variation-and-supplemental-description-opossum-shrimp-neomysis-
americana;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1974/723/williams1.pdf.

Haynes:1974:DRA

[2316] Evan B. Haynes. Distribution and relative abundance of larvae of king crab, Paralithodes camtschatica, in the Southeastern Bering Sea, 1969–

70. Fishery Bulletin, 72(3):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/distribution-and-relative-abundance-larvae-king-crab-
paralithodes-camtschatica-southeastern;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1974/723/haynes.pdf.

Bakun:1974:MAC

[2317] Andrew Bakun, Douglas R. McLain, and Frank V. Mayo. The mean annual cycle of coastal upwelling off western North America as observed from surface measurements. Fishery Bulletin, 72(3):843–844, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/mean-annual-
cycle-coastal-upwelling-western-north-america-observed-surface-
measurements;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/bakun.pdf.

Johnson:1974:DLL

[2318] Martin W. Johnson. On the dispersal of lobster larvae into the East Pacific Barrier (Decapoda, Palinuridea). Fishery Bulletin, 72 (3):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/dispersal-
lobster-larvae-east-pacific-barrier-decapoda-palinuridea;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1974/
723/johnson.pdf.

Smith:1974:RLF

[2319] Vega J. Smith, James S. Lin, and Harold S. Olcott. The residual lipids of fish protein concentrates. Fishery Bulletin, 72(3):845–847, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/residual-lipids-fish-
protein-concentrates;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/smith.pdf.

Johnson:1974:ECF

[2320] Allyn G. Johnson, Fred M. Utter, and Harold O. Hodgins. Electrophoretic comparison of five species of pandalid shrimp from the Northeastern Pacific Ocean. Fishery Bulletin, 72(3):??, 1974. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/electrophoretic-comparison-five-
species-pandalid-shrimp-northeastern-pacific-ocean;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1974/723/
johnson1.pdf.

Lenarz:1974:LWR

[2321] William H. Lenarz. Length–weight relations for five eastern tropical Atlantic scombrids. Fishery Bulletin, 72(3):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/length-weight-relations-five-eastern-
tropical-atlantic-scombrids;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/lenarz%20(2).pdf.

Porter:1974:RCS

[2322] Russell G. Porter. Reproductive cycle of the soft-shell clam, Mya arenaria, at Skagit Bay, Washington. Fishery Bulletin, 72(3):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/reproductive-
cycle-soft-shell-clam-mya-arenaria-skagit-bay-washington;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1974/
723/porter.pdf.

Klima:1974:ETR

[2323] Edward F. Klima. Electrical threshold response of some Gulf of Mexico fishes. Fishery Bulletin, 72(3):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/electrical-threshold-response-some-gulf-mexico-
fishes;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/klima.pdf.

Rosenthal:1974:ENH

[2324] Richard J. Rosenthal, William D. Clarke, and Paul K. Dayton. Ecology and natural history of a stand of giant kelp, Macrocystis pyrifera, off Del Mar, California. Fishery Bulletin, 72(3):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/ecology-and-natural-
history-stand-giant-kelp-macrocystis-pyrifera-del-mar-california;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1974/723/rosenthal.pdf.

Durkin:1974:ORC

[2325] Joseph T. Durkin and David A. Misitano. Occurrence of a ratfish in the Columbia River estuary. Fishery Bulletin, 72(3):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/occurrence-ratfish-
columbia-river-estuary;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/durkin.pdf.

Seidel:1974:SEC

[2326] Wilber R. Seidel and Edward F. Klima. In situ experiments with coastal pelagic fishes to establish design criteria for electrical fish harvesting systems. Fishery Bulletin, 72(3):??, 1974. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/situ-experiments-coastal-pelagic-
fishes-establish-design-criteria-electrical-fish;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1974/723/
seidel.pdf.

Misitano:1974:UOE

[2327] David A. Misitano and Carl W. Sims. Unusual occurrence of an eastern banded killifish in the lower Columbia River. Fishery Bulletin, 72(3):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
unusual-occurrence-eastern-banded-killifish-lower-columbia-

river;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/misitano.pdf.

Evans:1974:EOM

[2328] Dale R. Evans and Stanley D. Rice. Effects of oil on marine ecosystems: a review for administrators and policy makers. Fishery Bulletin, 72(3):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/effects-oil-marine-
ecosystems-review-administrators-and-policy-makers;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1974/723/
evans.pdf.

Williams:1974:SCG

[2329] Austin B. Williams. The swimming crabs of the genus Callinectes (Decapoda: Portunidae). Fishery Bulletin, 72(3):??, 1974. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/swimming-crabs-genus-callinectes-
decapoda-portunidae;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1974/723/williams.pdf.

Sherman:1974:MRL

[2330] Kenneth Sherman. In memoriam: Robert Louis Dryfoos, 1939–1974. Fishery Bulletin, 72(3):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/memoriam-robert-louis-dryfoos-1939-1974;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
inmemoriam%20(2).pdf.

Cushing:1974:LBS

[2331] D. H. Cushing. A link between science and management in fisheries. Fishery Bulletin, 72(4):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
link-between-science-and-management-fisheries;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1974/724/cushing.
pdf.

Quast:1974:DDJ

[2332] Jay C. Quast. Density distribution of juvenile Arctic cod, Boreogavus saida, in the Eastern Chukchi Sea in the fall of 1970. Fishery Bulletin, 72(4):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/density-
distribution-juvenile-arctic-cod-boreogavus-saida-eastern-chukchi-

sea-fall-1970;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1974/724/quast.pdf.

Gopalakrishnan:1974:ZGN

[2333] K. Gopalakrishnan. Zoogeography of the genus Nematoscelis (Crustacea, Euphausiacea). Fishery Bulletin, 72(4):??, 1974. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/zoogeography-genus-nematoscelis-
crustacea-euphausiacea;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1974/724/gopalakrishnan.pdf.

Richards:1974:ELD

[2334] William J. Richards, Robert Victor Miller, and Edward D. Houde. Egg and larval development of the Atlantic thread herring, Opisthonema oglinum. Fishery Bulletin, 72(4):1123–1136, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/egg-and-larval-development-atlantic-thread-
herring-opisthonema-oglinum;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1974/724/richards.pdf.

Hobson:1974:FRT

[2335] Edmund S. Hobson. Feeding relationships of teleostean fishes on coral reefs in Kona, Hawaii. Fishery Bulletin, 72(4):??, 1974. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/feeding-relationships-teleostean-
fishes-coral-reefs-kona-hawaii;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1974/724/hobson.pdf.

Tsuchiya:1974:VSG

[2336] Mizuki Tsuchiya. Variation of the surface geostrophic flow in the Eastern Intertropical Pacific Ocean. Fishery Bulletin, 72(4): ??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/variation-
surface-geostrophic-flow-eastern-intertropical-pacific-ocean;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1974/724/tsuchiya.pdf.

Houde:1974:DEL

[2337] Edward D. Houde, William J. Richards, and Vishnu P. Saksena. Description of eggs and larvae of scaled sardine, Harengula jaguana. Fishery Bulletin, 72(4):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
description-eggs-and-larvae-scaled-sardine-harengula-jaguana;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1974/724/houde.pdf.

Vlymen:1974:SEL

[2338] William J. Vlymen III. Swimming energetics of the larval anchovy, Engraulis mordax. Fishery Bulletin, 72(4):??, 1974. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/swimming-energetics-
larval-anchovy-engraulis-mordax;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1974/724/vlymen.pdf.

Krekorian:1974:LSB

[2339] C. O’Neil Krekorian, David C. Sommerville, and Richard F. Ford. Laboratory study of behavioral interactions between the American lobster, Homarus americanus, and the California spiny lobster, Panulirus interruptus, with comparative observations on the rock crab, Cancer antennarius. Fishery Bulletin, 72(4):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/laboratory-study-behavioral-interactions-between-
american-lobster-homarus-americanus-and;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1974/724/krekorian.
pdf.

Kennedy:1974:MMM

[2340] V. S. Kennedy, W. H. Roosenburg, M. Castagna, and J. A. Mihursky. Mercenaria mercenaria (Mollusca: Bivalvia): Temperature-time relationships for survival of embryos and larvae. Fishery Bulletin, 72(4):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/mercenaria-
mercenaria-mollusca-bivalvia-temperature-time-relationships-
survival-embryos-and;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/kennedy.pdf.

Leighton:1974:ITL

[2341] David L. Leighton. The influence of temperature on larval and juvenile growth in three species of Southern California abalones. Fishery Bulletin, 72(4):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/influence-temperature-larval-and-juvenile-
growth-three-species-southern-california-abalones;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1974/724/
leighton.pdf.

Beitinger:1974:TBD

[2342] Thomas L. Beitinger. Thermoregulatory behavior and diel activity patterns of bluegill, Lepomis macrochirus, following thermal shock. Fishery Bulletin, 72(4):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
thermoregulatory-behavior-and-diel-activity-patterns-bluegill-
lepomis-macrochirus-following;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1974/724/beitinger.pdf.

Mather:1974:AMM

[2343] F. J. Mather III, B. J. Rothschild, G. J. Paulik, and W. H. Lenarz. Analysis of migrations and mortality of bluefin tuna, Thunnus thynnus, tagged in the Northwestern Atlantic Ocean. Fishery Bulletin, 72(4):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/analysis-migrations-and-mortality-bluefin-
tuna-thunnus-thynnus-tagged-northwestern-atlantic;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1974/724/
mather.pdf.

Crawford:1974:EDT

[2344] L. Crawford, D. W. Peterson, M. J. Kretsch, A. L. Lilyblade, and H. S. Olcott. The effects of dietary α-tocopherol and tuna, safflower, and linseed oils on the flavor of turkey. Fishery Bulletin, 72(4):??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/effects-dietary-tocopherol-and-
tuna-safflower-and-linseed-oils-flavor-turkey;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1974/724/crawford.
pdf.

Moser:1974:DDL

[2345] H. Geoffrey Moser. Development and distribution of larvae and juveniles of Sebastolobus (Pisces; family Scorpaenidae). Fishery Bulletin, 72(4): ??, 1974. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/development-
and-distribution-larvae-and-juveniles-sebastolobus-pisces-family-
scorpaenidae;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1974/724/moser.pdf.

Anderson:1975:CBN

[2346] William W. Anderson, Jack W. Gehringer, and Frederick H. Berry. The correlation between numbers of vertebrae and lateral-line scales in Western Atlantic lizardfishes (Synodontidae). Fishery Bulletin, 73(1):202–206, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/correlation-between-
numbers-vertebrae-and-lateral-line-scales-western-atlantic-
lizardfishes;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1975/731/anderson1.pdf.

Lindall:1975:ASF

[2347] William N. Lindall, Jr., William A. Fable, Jr., and L. Alan Collins. Additional studies of the fishes, macroinvertebrates, and hydrological conditions of upland canals in Tampa Bay, Florida. Fishery Bulletin, 73(1):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/additional-studies-
fishes-macroinvertebrates-and-hydrological-conditions-upland-
canals;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1975/731/lindall.pdf.

Tillman:1975:AES

[2348] Michael F. Tillman. Additional evidence substantiating existence of northern subpopulation of northern anchovy, Engraulis mordax. Fishery Bulletin, 73(1):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
additional-evidence-substantiating-existence-northern-subpopulation-
northern-anchovy;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/tillman.pdf.

Berrien:1975:DAM

[2349] Peter L. Berrien. A description of Atlantic mackerel, Scomber scombrus, eggs and early larvae. Fishery Bulletin, 73(1):??, 1975. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/description-atlantic-
mackerel-scomber-scombrus-eggs-and-early-larvae;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1975/731/
berrien.pdf.

Lough:1975:RCE

[2350] R. Gregory Lough. A reevaluation of the combined effects of temperature and salinity on survival and growth of bivalve larvae using response surface techniques. Fishery Bulletin, 73(1):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/reevaluation-
combined-effects-temperature-and-salinity-survival-and-growth-
bivalve-larvae;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1975/731/lough.pdf.

Loosanoff:1975:CIC

[2351] Victor L. Loosanoff. Comment. Introduction of codium in New England waters. Fishery Bulletin, 73(1):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/comment-introduction-codium-new-england-waters;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/loosanoff.pdf.

Carr:1975:ITE

[2352] William E. S. Carr and James T. Giesel. Impact of thermal effluent from a steam-electric station on a marshland nursery area during the hot season. Fishery Bulletin, 73(1):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/impact-thermal-effluent-steam-electric-station-
marshland-nursery-area-during-hot-season;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1975/731/carr.pdf.

May:1975:ETS

[2353] Robert C. May. Effects of temperature and salinity on fertilization, embryonic development, and hatching in Bairdiella icistia (Pisces: Sciaenidae), and the effect of parental salinity acclimation on embryonic and larval salinity tolerance. Fishery Bulletin, 73(1):??, 1975. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/effects-temperature-and-salinity-
fertilization-embryonic-development-and-hatching;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1975/731/
may.pdf.

Fox:1975:FGS

[2354] William W. Fox, Jr. Fitting the generalized stock production model by least-squares and equilibrium approximation. Fishery Bulletin, 73(1):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/fitting-generalized-
stock-production-model-least-squares-and-equilibrium-approximation;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1975/731/fox.pdf.

McEachran:1975:DRA

[2355] John D. McEachran and J. A. Musick. Distribution and relative abundance of seven species of skates (Pisces: Rajidae) which occur between Nova Scotia and Cape Hatteras. Fishery Bulletin, 73(1):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/distribution-
and-relative-abundance-seven-species-skates-pisces-rajidae-
which-occur-between;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1975/731/mceachran.pdf.

Gilmartin:1975:CMC

[2356] Malvern Gilmartin and Noelia Revelante. The concentration of mercury, copper, nickel, silver, cadmium, and lead in the Northern Adriatic anchovy, Engraulis encrasicholus, and sardine, Sardina pilchardus. Fishery Bulletin, 73(1):193–201, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/concentration-mercury-copper-nickel-silver-cadmium-
and-lead-northern-adriatic-anchovy;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1975/731/gilmartin.pdf.

Rice:1975:ATA

[2357] Stanley D. Rice and Robert M. Stokes. Acute toxicity of ammonia to several developmental stages of rainbow trout, Salmo gairdneri. Fishery Bulletin, 73(1):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
acute-toxicity-ammonia-several-developmental-stages-rainbow-
trout-salmo-gairdneri;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1975/731/rice.pdf.

Horn:1975:SBS

[2358] Michael H. Horn. Swim-bladder state and structure in relation to behavior and mode of life in stromateoid fishes. Fishery Bulletin, 73(1):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/swim-bladder-state-and-structure-relation-behavior-
and-mode-life-stromateoid-fishes;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1975/731/horn.pdf.

Smayda:1975:NPG

[2359] Theodore J. Smayda. Net phytoplankton and the greater than 20-micron phytoplankton size fraction in upwelling waters off Baja California. Fishery Bulletin, 73(1):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/net-phytoplankton-and-greater-20-
micron-phytoplankton-size-fraction-upwelling-waters-baja;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1975/
731/smayda.pdf.

Kjelson:1975:GFE

[2360] Martin A. Kjelson, David S. Peters, Gordon W. Thayer, and George N. Johnson. The general feeding ecology of postlarval fishes in the Newport River Estuary. Fishery Bulletin, 73(1):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/general-feeding-ecology-postlarval-fishes-
newport-river-estuary;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1975/731/kjelson.pdf.

Wing:1975:NRE

[2361] Bruce L. Wing. New records of Ellobiopsidae (Protista (Incertae sedis)) from the North Pacific with a description of Thalassomyces albatrossi n. sp., a parasite of the mysid Stilomysis major. Fishery Bulletin, 73(1): ??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/new-records-
ellobiopsidae-protista-incertae-sedis-north-pacific-description-
thalassomyces;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1975/731/wing.pdf.

Knight:1975:LDP

[2362] Margaret D. Knight. The larval development of Pacific Euphausia gibboides (Euphausiacea). Fishery Bulletin, 73(1):??, 1975. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/larval-development-
pacific-euphausia-gibboides-euphausiacea;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1975/731/knight.pdf.

Anderson:1975:OEY

[2363] Lee G. Anderson. Optimum economic yield of an internationally utilized common property resource. Fishery Bulletin, 73(1): ??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/optimum-
economic-yield-internationally-utilized-common-property-resource;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/anderson_1.pdf.

Johnson:1975:TBN

[2364] James H. Johnson and Douglas R. McLain. Teleconnections between Northeastern Pacific Ocean and the Gulf of Mexico and Northwestern Atlantic Ocean. Fishery Bulletin, 73(2):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/teleconnections-between-northeastern-

pacific-ocean-and-gulf-mexico-and-northwestern;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1975/732/johnson1.
pdf.

Warner:1975:RBP

[2365] Robert R. Warner. The reproductive biology of the protogynous hermaphrodite Pimelometopon pulchrum (Pisces: Labridae). Fishery Bulletin, 73(2):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/reproductive-biology-protogynous-hermaphrodite-
pimelometopon-pulchrum-pisces-labridae;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1975/732/warner.pdf.

Anonymous:1975:TCL

[2366] Anonymous. Trap contributions to losses in the American lobster fishery. Fishery Bulletin, 73(2):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/trap-contributions-losses-american-lobster-fishery;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/sheldon.pdf.

Kendall:1975:GHD

[2367] Arthur W. Kendall, Jr. and John W. Reintjes. Geographic and hydrographic distribution of Atlantic menhaden eggs and larvae along the Middle Atlantic Coast from RV Dolphin cruises, 1965–66. Fishery Bulletin, 73(2):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
geographic-and-hydrographic-distribution-atlantic-menhaden-
eggs-and-larvae-along-middle;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1975/732/kendall.pdf.

Reeve:1975:PTP

[2368] M. R. Reeve and L. D. Baker. Production of two planktonic carnivores (chaetognath and ctenophore) in South Florida inshore waters. Fishery Bulletin, 73(2):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/production-two-planktonic-carnivores-
chaetognath-and-ctenophore-south-florida-inshore;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
reeve.pdf.

Lange:1975:TTU

[2369] G. D. Lange and A. C. Hurley. A theoretical treatment of unstructured food webs. Fishery Bulletin, 73(2):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/theoretical-treatment-unstructured-food-webs;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1975/732/lange.pdf.

Johnson:1975:ICB

[2370] Robert Karl Johnson and Michael A. Barnett. An inverse correlation between meristic characters and food supply in mid-water fishes: evidence and possible explanations. Fishery Bulletin, 73(2):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/inverse-correlation-
between-meristic-characters-and-food-supply-mid-water-fishes-
evidence;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/johnson_0.pdf.

Agnello:1975:IEB

[2371] Richard J. Agnello and Lawrence P. Donnelley. The interaction of economic, biological, and legal forces in the Middle Atlantic oyster industry. Fishery Bulletin, 73(2):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/interaction-economic-biological-and-legal-
forces-middle-atlantic-oyster-industry;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1975/732/agnello.pdf.

May:1975:EAT

[2372] Robert C. May. Effects of acclimation on the temperature and salinity tolerance of the yolk-sac larvae of Bairdiella icistia (Pisces: Sciaenidae). Fishery Bulletin, 73(2):??, 1975. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/effects-acclimation-temperature-
and-salinity-tolerance-yolk-sac-larvae-bairdiella-icistia;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1975/
732/may.pdf.

Blackburn:1975:DES

[2373] Maurice Blackburn and Francis Williams. Distribution and ecology of skipjack tuna, Katsuwonus pelamis, in an offshore area of the eastern tropical Pacific Ocean. Fishery Bulletin, 73(2):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/distribution-
and-ecology-skipjack-tuna-katsuwonus-pelamis-offshore-area-
eastern-tropical;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/blackburn%20(1).pdf.

Cain:1975:RRB

[2374] Thomas D. Cain. Reproduction and recruitment of the brackish water clam Rangia cuneata in the James River, Virginia. Fishery Bulletin, 73 (2):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/reproduction-
and-recruitment-brackish-water-clam-rangia-cuneata-james-river-
virginia;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1975/732/cain.pdf.

Roubal:1975:SLT

[2375] William T. Roubal and Tracy K. Collier. Spin-labeling techniques for studying mode of action of petroleum hydrocarbons on marine organisms. Fishery Bulletin, 73(2):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/spin-labeling-techniques-
studying-mode-action-petroleum-hydrocarbons-marine-organisms;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1975/732/roubal.pdf.

Gwinn:1975:DMC

[2376] Sharon Gwinn and William F. Perrin. Distribution of melanin in the color pattern of Delphinus delphis (Cetacea; Delphinidae). Fishery Bulletin, 73 (2):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/distribution-
melanin-color-pattern-delphinus-delphis-cetacea-delphinidae;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/gwinn.pdf.

Dark:1975:AGP

[2377] Thomas A. Dark. Age and growth of Pacific hake, Merluccius productus. Fishery Bulletin, 73(2):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/age-and-growth-pacific-hake-merluccius-productus;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1975/
732/dark.pdf.

Serfling:1975:ESP

[2378] Steven A. Serfling and Richard F. Ford. Ecological studies of the puerulus larval stage of the California spiny lobster, Panulirus interruptus. Fishery Bulletin, 73(2):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
ecological-studies-puerulus-larval-stage-california-spiny-lobster-
panulirus-interruptus;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1975/732/serfling.pdf.

Hauser:1975:OTC

[2379] William J. Hauser. Occurrence of two Congridae leptocephali in an estuary. Fishery Bulletin, 73(2):??, 1975. CODEN FSY-

BAY.  ISSN 0090-0656 (print), 1937-4518 (electronic).  URL

https://spo.nmfs.noaa.gov/content/occurrence-two-congridae-
leptocephali-estuary;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/hauser.pdf.

Dayton:1975:ESA

[2380] Paul K. Dayton. Experimental studies of algal canopy interactions in a sea otter-dominated kelp community at Amchitka Island, Alaska. Fishery Bulletin, 73(2):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/experimental-studies-algal-canopy-interactions-
sea-otter-dominated-kelp-community-amchitka;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1975/732/dayton.
pdf.

Smigielski:1975:HIO

[2381] Alphonse S. Smigielski. Hormonal-induced ovulation of the winter flounder, Pseudopleuronectes americanus. Fishery Bulletin, 73(2): ??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/hormonal-
induced-ovulation-winter-flounder-pseudopleuronectes-americanus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1975/732/smigielski.pdf.

Williams:1975:CHS

[2382] P. M. Williams and K. J. Robertson. Chlorinated hydrocarbons in sea-surface films and subsurface waters at nearshore stations and in the North Central Pacific gyre. Fishery Bulletin, 73(2):??, 1975. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/chlorinated-hydrocarbons-
sea-surface-films-and-subsurface-waters-nearshore-stations-
and;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/williams_0.pdf.

Graham:1975:HEY

[2383] Jeffrey B. Graham. Heat exchange in the yellowfin tuna, Thunnus albacares, and skipjack tuna, Katsuwonus pelamis, and the adaptive significance of elevated body temperatures in scombrid fishes. Fishery Bulletin, 73(2):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/heat-exchange-yellowfin-tuna-thunnus-albacares-and-
skipjack-tuna-katsuwonus-pelamis-and;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1975/732/graham.pdf.

Thomas:1975:ERA

[2384] Allan E. Thomas. Evaluation of the return of adult Chinook salmon to the Abernathy Incubation Channel. Fishery Bulletin, 73(2):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/evaluation-
return-adult-chinook-salmon-abernathy-incubation-channel;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1975/
732/thomas.pdf.

Mighell:1975:HSH

[2385] James L. Mighell and James R. Dangel. Hatching survival of hybrids of Oncorhynchus masou with Salmo gairdneri and with North American species of Oncorhynchus. Fishery Bulletin, 73(2): ??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/hatching-
survival-hybrids-oncorhynchus-masou-salmo-gairdneri-and-north-
american-species;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/mighell.pdf.

Clark:1975:ULP

[2386] Robert C. Clark, Jr. and John S. Finley. Uptake and loss of petroleum hydrocarbons by the mussel, Mytilus edulis, in laboratory experiments. Fishery Bulletin, 73(3):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/uptake-and-loss-petroleum-hydrocarbons-mussel-mytilus-
edulis-laboratory-experiments;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1975/733/clark.pdf.

Silliman:1975:SUE

[2387] Ralph P. Silliman. Selective and unselective exploitation of experimental populations of Tilapia mossambica. Fishery Bulletin, 73(3):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/selective-and-
unselective-exploitation-experimental-populations-tilapia-mossambica;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1975/733/silliman.pdf.

Pearcy:1975:CAD

[2388] William G. Pearcy, Daniel A. Panshin, and Donald F. Keene. Catches of albacore at different times of the day. Fishery Bulletin, 73(3):691–693, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/catches-albacore-different-times-day;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
pearcy%20(1).pdf.

Collette:1975:SMB

[2389] Bruce B. Collette and Labbish N. Chao. Systematics and morphology of the bonitos (Sarda) and their relatives (Scombridae, Sardini). Fishery Bulletin, 73(3):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/systematics-and-morphology-bonitos-sarda-and-their-
relatives-scombridae-sardini;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1975/733/collette.pdf.

Farante:1975:STA

[2390] Isabel Perez Farante. Spermatophores and thelyca of the American white shrimps, genus Penaeus, subgenus Litopenaeus. Fishery Bulletin, 73(3):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/spermatophores-and-thelyca-american-white-
shrimps-genus-penaeus-subgenus-litopenaeus;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/perez%20(
1).pdf.

Hartmann:1975:DMF

[2391] A. Rucker Hartmann and Thomas A. Clarke. The distribution of myctophid fishes across the Central Equatorial Pacific. Fishery Bulletin, 73 (3):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/distribution-

myctophid-fishes-across-central-equatorial-pacific;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1975/733/
hartmann.pdf.

Kiefer:1975:TBG

[2392] Dale A. Kiefer and Reuben Lasker. Two blooms of Gymnodinium splendens, an unarmored dinoflagellate. Fishery Bulletin, 73(3):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/two-blooms-
gymnodinium-splendens-unarmored-dinoflagellate;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/kiefer.
pdf.

Houde:1975:DEL

[2393] Edward D. Houde and L. J. Swanson, Jr. Description of eggs and larvae of yellowfin menhaden, Brevoortia smithi. Fishery Bulletin, 73(3):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/description-eggs-and-
larvae-yellowfin-menhaden-brevoortia-smithi;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1975/733/houde.
pdf.

Sandifer:1975:ESL

[2394] Paul A. Sandifer, Walter E. Castro, and Paul B. Zielinski. Enhanced survival of larval grass shrimp in dilute solutions of the synthetic polymer, polyethylene oxide. Fishery Bulletin, 73(3):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/enhanced-survival-larval-grass-
shrimp-dilute-solutions-synthetic-polymer-polyethylene-oxide;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/sandifer.pdf.

Ingham:1975:VTA

[2395] Merton C. Ingham. Velocity and transport of the Antilles Current northeast of the Bahama Islands. Fishery Bulletin, 73(3):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/velocity-and-transport-antilles-
current-northeast-bahama-islands;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1975/733/ingham.pdf. See note

Yong:1975:CPA

[2396] Marian Y. Y. Yong and Robert A. Skillman. A computer program for analysis of polymodal frequency distributions (ENORMSEP), FORTRAN IV. Fishery Bulletin, 73(3):681, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/computer-program-analysis-polymodal-frequency-distributions-
enormsep-fortran-iv;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/yong.pdf.

Lasker:1975:FCS

[2397] Reuben Lasker. Field criteria for survival of anchovy larvae: the relation between inshore chlorophyll maximum layers and successful first feeding. Fishery Bulletin, 73(3):??, 1975. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/field-criteria-survival-anchovy-
larvae-relation-between-inshore-chlorophyll-maximum-layers;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1975/733/lasker.pdf.

Richardson:1975:RLT

[2398] Sally L. Richardson and Douglas A. DeHart. Records of larval, transforming, and adult specimens of the quillfish, Ptilichthys goodei, from waters off Oregon. Fishery Bulletin, 73(3):??, 1975. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/records-larval-transforming-and-
adult-specimens-quillfish-ptilichthys-goodei-waters-oregon;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/richardson%20(1).pdf.

Peterson:1975:YYV

[2399] William T. Peterson and Charles B. Miller. Year-to-year variations in the planktology of the Oregon Upwelling Zone. Fishery Bulletin, 73(3):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/year-year-variations-
planktology-oregon-upwelling-zone;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1975/733/peterson.pdf.

Silliman:1975:ECS

[2400] Ralph P. Silliman. Effect of crowding on stock and catch in Tilapia mossambica. Fishery Bulletin, 73(3):??, 1975. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/effect-crowding-stock-and-catch-

tilapia-mossambica;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/silliman1.pdf.

Chittenden:1975:DAS

[2401] Mark E. Chittenden, Jr. Dynamics of American shad, Alosa sapidissima, runs in thec Delaware River. Fishery Bulletin, 73(3): ??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/dynamics-
american-shad-alosa-sapidissima-runs-delaware-river;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1975/733/
chittenden.pdf.

Rybock:1975:UOS

[2402] J. T. Rybock, H. F. Horton, and J. L. Fessler. Use of otoliths to separate juvenile steelhead trout from juvenile rainbow trout. Fishery Bulletin, 73 (3):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/use-otoliths-
separate-juvenile-steelhead-trout-juvenile-rainbow-trout;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1975/
733/rybock.pdf.

Wenner:1975:OES

[2403] Charles A. Wenner. The occurrence of elvers of synaphobranchus affinis on the continental slope off North Carolina. Fishery Bulletin, 73(3): ??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/occurrence-
elvers-synaphobranchus-affinis-continental-slope-north-carolina;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/wenner.pdf.

Dawley:1975:EVC

[2404] Earl M. Dawley and Wesley J. Ebel. Effects of various concentrations of dissolved atmospheric gas on juvenile Chinook salmon and steelhead trout. Fishery Bulletin, 73(4):??, 1975. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/effects-various-concentrations-
dissolved-atmospheric-gas-juvenile-chinook-salmon-and;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1975/734/
dawley.pdf.

Palm:1975:FRO

[2405] William J. Palm. Fishery regulation via optimal control theory. Fishery Bulletin, 73(4):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/fishery-regulation-optimal-control-theory;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1975/734/palm.pdf.

Slatick:1975:FSR

[2406] Emil Slatick, Donn L. Park, and Wesley J. Ebel. Further studies regarding effects of transportation on survival and homing of Snake River Chinook salmon and steelhead trout. Fishery Bulletin, 73(4): 925–931, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/further-
studies-regarding-effects-transportation-survival-and-homing-
snake-river-chinook;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/slatick.pdf.

DeVlaming:1975:EPT

[2407] Victor L. De Vlaming. Effects of photoperiod–temperature regimes and pinealectomy on body fat reserves in the golden shiner, Notemigonus crysoleucas. Fishery Bulletin, 73(4):??, 1975. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/effects-photoperiod-temperature-
regimes-and-pinealectomy-body-fat-reserves-golden-shiner;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1975/
734/devlaming.pdf.

Parrish:1975:MTI

[2408] James D. Parrish. Marine trophic interactions by dynamic simulation of fish species. Fishery Bulletin, 73(4):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/marine-trophic-interactions-dynamic-simulation-
fish-species;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1975/734/parrish.pdf.

Patten:1975:CVF

[2409] Benjamin J. Patten. Comparative vulnerability of fry of Pacific salmon and steelhead trout to predation by torrent sculpin in stream aquaria. Fishery Bulletin, 73(4):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/comparative-vulnerability-fry-pacific-salmon-and-steelhead-
trout-predation-torrent-sculpin;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/patten.pdf.

Gopalakrishnan:1975:BTG

[2410] K. Gopalakrishnan. Biology and taxonomy of the genus Nematoscelis (Crustacea, Euphausiacea). Fishery Bulletin, 73(4):??, 1975. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/biology-and-taxonomy-
genus-nematoscelis-crustacea-euphausiacea;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1975/734/gopalakrishnan.
pdf.

Silliman:1975:EEC

[2411] Ralph P. Silliman. Experimental exploitation of competing fish populations. Fishery Bulletin, 73(4):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/experimental-exploitation-competing-fish-populations;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1975/734/silliman.pdf.

Cramer:1975:HRG

[2412] S. P. Cramer and J. D. McIntyre. Heritable resistance to gas bubble disease in fall Chinook salmon, Oncorhynchus tshawytscha. Fishery Bulletin, 73(4):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
heritable-resistance-gas-bubble-disease-fall-chinook-salmon-
oncorhynchus-tshawytscha;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/cramer.pdf.

Ackman:1975:DFC

[2413] R. G. Ackman, C. A. Eaton, and B. A. Linke. Differentiation of freshwater characteristics of fatty acids in marine specimens of the Atlantic sturgeon, Acipenser oxyrhynchus. Fishery Bulletin, 73(4):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
differentiation-freshwater-characteristics-fatty-acids-marine-
specimens-atlantic-sturgeon;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1975/734/ackman.pdf.

Hopkins:1975:NFM

[2414] Thomas L. Hopkins and Ronald C. Baird. Net feeding in mesopelagic fishes. Fishery Bulletin, 73(4):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/net-feeding-mesopelagic-fishes;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1975/734/
hopkins.pdf.

Vreeland:1975:HBC

[2415] Robert R. Vreeland, Roy J. Wahle, and Arthur H. Arp. Homing behavior and contribution to Columbia River fisheries of marked Coho salmon released at two locations. Fishery Bulletin, 73(4):??, 1975. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/homing-behavior-and-contribution-
columbia-river-fisheries-marked-coho-salmon-released-two;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1975/
734/vreeland.pdf.

Anonymous:1975:I

[2416] Anonymous. Index. Fishery Bulletin, 73(4):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/index-103;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/index%
20(2)_0.pdf.

Bailey:1975:ZAF

[2417] Jack E. Bailey, Bruce L. Wing, and Chester R. Mattson. Zooplankton abundance and feeding habits of fry of pink salmon, Oncorhynchus gorbuscha, and chum salmon, Oncorhynchus keta, in Traitors Cove, Alaska, with speculations on the carrying capacity of the area. Fishery Bulletin, 73(4):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
zooplankton-abundance-and-feeding-habits-fry-pink-salmon-oncorhynchus-
gorbuscha-and-chum;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1975/734/bailey.pdf.

Krouse:1975:ETS

[2418] Jay S. Krouse and James C. Thomas. Effects of trap selectivity and some population parameters on size composition of the American lobster, Homarus americanus, catch along the Maine coast. Fishery Bulletin, 73 (4):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/effects-trap-
selectivity-and-some-population-parameters-size-composition-
american-lobster;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1975/734/krouse.pdf.

Wiebe:1975:RBZ

[2419] Peter H. Wiebe, Steven Boyd, and James L. Cox. Relationships between zooplankton displacement volume, wet weight, dry weight, and carbon. Fishery Bulletin, 73(4):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/relationships-between-zooplankton-displacement-
volume-wet-weight-dry-weight-and-carbon;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1975/734/wiebe.pdf.

Botsford:1975:CUI

[2420] Louis W. Botsford and Daniel E. Wickham. Correlation of upwelling index and Dungeness crab catch. Fishery Bulletin, 73(4):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/correlation-upwelling-index-and-
dungeness-crab-catch;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1975/734/botsford.pdf.

Krygier:1975:DBN

[2421] Earl E. Krygier and Robert A. Wasmer. Description and biology of a new species of pelagic penaeid shrimp, Bentheogennema burkenroadi, from the Northeastern Pacific. Fishery Bulletin, 73(4):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
description-and-biology-new-species-pelagic-penaeid-shrimp-
bentheogennema-burkenroadi;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1975/734/krygier.pdf.

Yoshida:1975:ASL

[2422] Howard O. Yoshida. The American Samoa longline fishery, 1966–71. Fishery Bulletin, 73(4):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/american-samoa-longline-fishery-1966-71;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1975/734/
yoshida.pdf.

Bray:1975:FAH

[2423] Richard N. Bray and Alfred W. Ebeling. Food, activity, and habitat of three ‘picker-type’ microcarnivorous fishes in the kelp forests off Santa Barbara, California. Fishery Bulletin, 73(4):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/food-activity-
and-habitat-three-picker-type-microcarnivorous-fishes-kelp-
forests-santa;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1975/734/bray.pdf.

Maynard:1975:MMH

[2424] Sherwood D. Maynard, Fletcher V. Riggs, and John F. Walters. Mesopelagic micronekton in Hawaiian waters: Faunal composition, standing stock, and diel vertical migration. Fishery Bulletin, 73(4):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
mesopelagic-micronekton-hawaiian-waters-faunal-composition-
standing-stock-and-diel-vertical;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1975/734/maynard.pdf.

Rucker:1975:GBD

[2425] Robert R. Rucker. Gas-bubble disease: Mortalities of Coho salmon, Oncorhynchus kisutch, in water with constant total gas pressure and different oxygen-nitrogen ratios. Fishery Bulletin, 73(4):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/gas-bubble-
disease-mortalities-coho-salmon-oncorhynchus-kisutch-water-
constant-total-gas;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/rucker.pdf.

Chen:1975:PSM

[2426] Lo-Chai Chen and Robert L. Martinich. Pheromonal stimulation and metabolite inhibition of ovulation in the zebrafish, Brachydanio rerio. Fishery Bulletin, 73(4):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/pheromonal-stimulation-and-metabolite-inhibition-ovulation-
zebrafish-brachydanio-rerio;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1975/734/chen.pdf.

Olla:1975:AMF

[2427] Bori L. Olla, Allen J. Bejda, and A. Dale Martin. Activity, movements, and feeding behavior of the cunner, Tautogolabrus adspersus, and comparison of food habits with young tautog, Tautoga onitis, off Long Island, New York. Fishery Bulletin, 73(4):895–900, 1975. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/activity-movements-and-feeding-
behavior-cunner-tautogolabrus-adspersus-and-comparison-food;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1975/734/olla.pdf.

Weber:1975:ALW

[2428] Douglas D. Weber and Herbert H. Shippen. Age–length–weight and distribution of Alaska plaice, rock sole, and yellowfin sole collected from the southeastern Bering Sea in 1961. Fishery Bulletin, 73 (4):??, 1975. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/age-
length-weight-and-distribution-alaska-plaice-rock-sole-and-
yellowfin-sole-collected;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/weber_0.pdf.

Ogren:1976:DAS

[2429] Larry H. Ogren and Harold A. Brusher. Distribution, abundance, and size of penaeid shrimps in the St. Andrew Bay System, Florida. Fishery Bulletin, 74(1):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/distribution-abundance-and-size-penaeid-shrimps-st-
andrew-bay-system-florida;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1976/741/brusher.pdf.

Lord:1976:AMH

[2430] Gary Lord, William C. Acker, Allan C. Hartt, and Brian J. Rothschild. An acoustic method for the high-seas assessment of migrating salmon. Fishery Bulletin, 74(1):104–111, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/acoustic-method-high-seas-assessment-migrating-
salmon;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1976/741/lord.pdf.

Sharp:1976:EME

[2431] Gary D. Sharp and Robert C. Francis. An energetics model for the exploited yellowfin tuna, Thunnus albacares, population in the Eastern Pacific Ocean. Fishery Bulletin, 74(1):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/energetics-model-exploited-yellowfin-
tuna-thunnus-albacares-population-eastern-pacific;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1976/741/
sharp.pdf.

Cheng:1976:GBC

[2432] Lanna Cheng and Ralph A. Lewin. Goose barnacles (Cirripedia: Thoracica) on flotsam beached at La Jolla, California. Fishery Bulletin, 74(1):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/goose-
barnacles-cirripedia-thoracica-flotsam-beached-la-jolla-california;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/cheng.pdf.

Butler:1976:RDS

[2433] John L. Butler and Elbert H. Ahlstrom. Review of the deep-sea fish genus Scopelengys (Neoscopelidae) with a description of a new species, Scopelengys clarkei, from the Central Pacific. Fishery Bulletin, 74(1):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/review-deep-sea-fish-
genus-scopelengys-neoscopelidae-description-new-species-scopelengys;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1976/741/butler.pdf.

MacGregor:1976:DMS

[2434] John S. MacGregor. DDT and its metabolites in the sediments off Southern California. Fishery Bulletin, 74(1):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/ddt-and-its-metabolites-sediments-southern-
california;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1976/741/macgregor.pdf.

Struhsaker:1976:AGN

[2435] Paul Struhsaker and James H. Uchiyama. Age and growth of the nehu, Stolephorus purpureus (Pisces: Engraulidae), from the Hawaiian Islands as indicated by daily growth increments of Sagittae. Fishery Bulletin, 74(1):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/age-and-growth-nehu-stolephorus-purpureus-
pisces-engraulidae-hawaiian-islands-indicated;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1976/741/struhsaker.
pdf.

Laurence:1976:CVS

[2436] Geoffrey C. Laurence. Caloric values of some North Atlantic calanoid copepods. Fishery Bulletin, 74(1):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/caloric-values-some-north-atlantic-calanoid-
copepods;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/laurence.pdf.

Chittenden:1976:WLM

[2437] Mark E. Chittenden, Jr. Weight loss, mortality, feeding, and duration of residence of adult American shad, Alosa sapidissima, in fresh water. Fishery Bulletin, 74(1):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/weight-loss-mortality-feeding-and-duration-residence-
adult-american-shad-alosa-sapidissima;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1976/741/chittenden.pdf.

Mason:1976:PSS

[2438] J. C. Mason and S. Machidorl. Populations of sympatric sculpins, Cottus aleuticus and Cottus asper, in four adjacent salmon-producing coastal streams on Vancouver Island, B.C. Fishery Bulletin, 74(1):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/populations-
sympatric-sculpins-cottus-aleuticus-and-cottus-asper-four-adjacent-
salmon;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1976/741/mason.pdf.

Tillman:1976:DEA

[2439] Michael F. Tillman and Donald Stadelman. Development and example application of a simulation model of the northern anchovy fishery. Fishery Bulletin, 74(1):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/development-and-example-application-simulation-
model-northern-anchovy-fishery;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1976/741/tillman.pdf.

Haury:1976:MRL

[2440] Loren R. Haury. Method for restraining living planktonic crustaceans. Fishery Bulletin, 74(1):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/method-restraining-living-planktonic-crustaceans;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/haury.pdf.

Crawford:1976:ECA

[2441] L. Crawford and M. J. Kretsch. Effects of cooking in air or in nitrogen on the development of fishy flavor in the breast meat of turkeys fed tuna oil with and without α-tocopherol supplement or injection. Fishery Bulletin, 74(1):89–92, 1976. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/effects-cooking-air-or-nitrogen-
development-fishy-flavor-breast-meat-turkeys-fed-tuna-oil;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1976/
741/crawford.pdf.

Mason:1976:SFC

[2442] J. C. Mason. Some features of Coho salmon, Oncorhynchus kisutch, fry emerging from simulated redds and concurrent changes in photobehavior. Fishery Bulletin, 74(1):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/some-features-coho-salmon-oncorhynchus-kisutch-
fry-emerging-simulated-redds-and-concurrent;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1976/741/mason1.
pdf.

Trent:1976:AMN

[2443] Lee Trent, Edward J. Pullen, and Raphael Proctor. Abundance of macrocrustaceans in a natural marsh and a marsh altered by dredging, bulkheading, and filling. Fishery Bulletin, 74(1):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic).  URL https://spo.nmfs.noaa.gov/content/abundance-
macrocrustaceans-natural-marsh-and-marsh-altered-dredging-bulkheading-
and-filling;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1976/741/trent.pdf.

Stillwell:1976:OBT

[2444] Charles E. Stillwell and John G. Casey. Observation on the bigeye thresher shark, Alopias superciliosus, in the western North Atlantic. Fishery Bulletin, 74(1):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/observation-bigeye-thresher-shark-alopias-
superciliosus-western-north-atlantic;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/stillwell.pdf.

Garrison:1976:CNP

[2445] David L. Garrison. Contribution of the net plankton and nannoplankton to the standing stocks and primary productivity in Monterey Bay, California during the upwelling season. Fishery Bulletin, 74(1):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/contribution-
net-plankton-and-nannoplankton-standing-stocks-and-primary-
productivity;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1976/741/garrison.pdf.

Merriner:1976:ARB

[2446] John V. Merriner. Aspects of the reproductive biology of the weakfish, Cynoscion regalis (Sciaenidae), in North Carolina. Fishery Bulletin, 74 (1):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/aspects-
reproductive-biology-weakfish-cynoscion-regalls-sciaenidae-
north-carolina;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1976/741/merriner.pdf.

Uchida:1976:RFE

[2447] Richard N. Uchida. Reevaluation of fishing effort and apparent abundance in the Hawaiian fishery for skipjack tuna, Katsuwonus pelamis, 1948–70. Fishery Bulletin, 74(1):??, 1976. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/reevaluation-fishing-effort-and-
apparent-abundance-hawaiian-fishery-skipjack-tuna;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1976/741/
uchida.pdf.

Lewis:1976:EAB

[2448] Elizabeth G. Lewis. Epizoites associated with Bathynectes superbus (Decapoda: Portunidae). Fishery Bulletin, 74(1):??, 1976. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/epizoites-associated-bathynectes-
superbus-decapoda-portunidae;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/lewis1.pdf.

Hunter:1976:CGN

[2449] John R. Hunter. Culture and growth of northern anchovy, Engraulis mordax, larvae. Fishery Bulletin, 74(1):??, 1976. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/culture-and-growth-
northern-anchovy-engraulis-mordax-larvae;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1976/741/hunter.pdf.

Pearcy:1976:SIO

[2450] William G. Pearcy. Seasonal and inshore–offshore variations in the standing stocks of micronekton and macrozooplankton off Oregon. Fishery Bulletin, 74(1):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
seasonal-and-inshore-offshore-variations-standing-stocks-micronekton-
and-macrozooplankton;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1976/741/pearcy.pdf.

Fisher:1976:MEF

[2451] William S. Fisher and Daniel W. Wickham. Mortalities and epibiotic fouling of eggs from wild populations of the Dungeness crab, Cancer magister. Fishery Bulletin, 74(1):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/mortalities-and-epibiotic-fouling-eggs-wild-populations-
dungeness-crab-cancer-magister;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fisher.pdf.

Hurley:1976:FBF

[2452] Ann C. Hurley. Feeding behavior, food consumption, growth, and respiration of the squid Loligo opalescens raised in the laboratory. Fishery Bulletin, 74(1):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
feeding-behavior-food-consumption-growth-and-respiration-squid-
loligo-opalescens-raised;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1976/741/hurley.pdf.

Pristas:1976:ART

[2453] Paul J. Pristas, Eldon J. Levi, and Robert L. Dryfoos. Analysis of returns of tagged Gulf menhaden. Fishery Bulletin, 74(1): ??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/analysis-
returns-tagged-gulf-menhaden;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1976/741/pristas.pdf.

Matsumoto:1976:SRB

[2454] Walter M. Matsumoto. Second record of black skipjack, Euthynnus lineatus, from the Hawaiian Islands. Fishery Bulletin, 74(1):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/second-record-black-
skipjack-euthynnus-lineatus-hawaiian-islands;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/matsumoto%
20(1).pdf.

Brothers:1976:DGI

[2455] Edward B. Brothers, Christopher P. Mathews, and Reuben Lasker. Daily growth increments in otoliths from larval and adult fishes. Fishery Bulletin, 74(1):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
daily-growth-increments-otoliths-larval-and-adult-fishes;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1976/
741/brothers.pdf.

Lewis:1976:BFS

[2456] Thomas C. Lewis and Ralph W. Yerger. Biology of five species of searobins (Pisces, Triglidae) from the Northeastern Gulf of Mexico. Fishery Bulletin, 74(1):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
biology-five-species-searobins-pisces-triglidae-northeastern-
gulf-mexico;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1976/741/lewis.pdf.

Rogers:1976:ETS

[2457] Carolyn A. Rogers. Effects of temperature and salinity on the survival of winter flounder embryos. Fishery Bulletin, 74(1):??, 1976. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/effects-temperature-and-salinity-
survival-winter-flounder-embryos;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1976/741/rogers.pdf.

Weis:1976:OMI

[2458] Judith S. Weis and Peddrick Weis. Optical malformations induced by insecticides in embryos of the Atlantic silverside, Menidia menidia. Fishery Bulletin, 74(1):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/optical-malformations-induced-insecticides-embryos-
atlantic-silverside-menidia-menidia;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/weis.pdf.

Haynes:1976:DZC

[2459] Evan Haynes. Description of zoeae of coonstripe shrimp, Pandalus hypsinotus, reared in the laboratory. Fishery Bulletin, 74(2):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/description-
zoeae-coonstripe-shrimp-pandalus-hypsinotus-reared-laboratory;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1976/742/haynes.pdf.

Chittenden:1976:PHS

[2460] Mark E. Chittenden, Jr. Present and historical spawning grounds and nurseries of American shad, Alosa sapidissima, in the Delaware River. Fishery Bulletin, 74(2):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/present-and-historical-spawning-
grounds-and-nurseries-american-shad-alosa-sapidissima;   https:

//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
chittenden.pdf.

Fuiman:1976:NED

[2461] Lee A. Fuiman. Notes on the early development of the sea raven, Hemitripterus americanus. Fishery Bulletin, 74(2):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/notes-early-development-sea-raven-
hemitripterus-americanus;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fuiman.pdf.

Hewitt:1976:DUS

[2462] Roger P. Hewitt, Paul E. Smith, and John C. Brown. Development and use of sonar mapping for pelagic stock assessment in the California Current area. Fishery Bulletin, 74(2):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/development-and-use-sonar-mapping-pelagic-
stock-assessment-california-current-area;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1976/742/hewitt.pdf.

Harrell:1976:IDT

[2463] Lee W. Harrell, Anthony J. Novotny, Michael H. Schiewe, and Harold O. Hodgins. Isolation and description of two vibrios pathogenic to Pacific salmon in Puget Sound, Washington. Fishery Bulletin, 74(2):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/isolation-and-
description-two-vibrios-pathogenic-pacific-salmon-puget-sound-
washington;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/harrell.pdf.

Kjelson:1976:FOF

[2464] Martin A. Kjelson and George N. Johnson. Further observations of the feeding ecology of postlarval pinfish, Lagodon rhomboides, and spot, Leiostomus xanthurus. Fishery Bulletin, 74(2):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/further-observations-
feeding-ecology-postlarval-pinfish-lagodon-rhomboides-and-spot;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1976/742/kjelson.pdf.

May:1976:RFC

[2465] Nelson May, Lee Trent, and Paul J. Pristas. Relation of fish catches in gill nets to frontal periods. Fishery Bulletin, 74(2):449–453, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/relation-fish-catches-gill-nets-
frontal-periods;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/may_1.pdf.

Able:1976:LHE

[2466] K. W. Able and J. A. Musick. Life history, ecology, and behavior of Liparis inquilinus (Pisces: Cyclopteridae) associated with the sea scallop, Placopecten magellanicus. Fishery Bulletin, 74(2):??, 1976. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/life-history-ecology-and-
behavior-liparis-inquilinus-pisces-cyclopteridae-associated-
sea;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1976/742/able.pdf.

Livingston:1976:LTF

[2467] Robert J. Livingston, Gerard J. Kobylinski, Frank G. Lewis III, and Peter F. Sheridan. Long-term fluctuations of epibenthic fish and invertebrate populations in Apalachicola Bay, Florida. Fishery Bulletin, 74 (2):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/long-term-
fluctuations-epibenthic-fish-and-invertebrate-populations-apalachicola-
bay;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1976/742/livingston.pdf.

Lansford:1976:PPT

[2468] Lawrence M. Lansford, Charles W. Caillouet, and Kenneth T. Marvin. Phosphoglucomutase polymorphism in two penaeid shrimps, Penaeus brasiliensis and Penaeus aztecus subtilis. Fishery Bulletin, 74(2):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/phosphoglucomutase-
polymorphism-two-penaeid-shrimps-penaeus-brasiliensis-and-penaeus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/lansford.pdf.

Brewer:1976:TTR

[2469] Gary D. Brewer. Thermal tolerance and resistance of the northern anchovy, Engraulis mordax. Fishery Bulletin, 74(2):??, 1976. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/thermal-tolerance-and-resistance-
northern-anchovy-engraulis-mordax;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1976/742/brewer.pdf.

Lough:1976:LDD

[2470] R. Gregory Lough. Larval dynamics of the Dungeness crab, Cancer magister, off the Central Oregon Coast, 1970–71. Fishery Bulletin, 74(2):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/larval-dynamics-dungeness-crab-cancer-magister-
central-oregon-coast-1970-71;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1976/742/lough.pdf.

Perrin:1976:FRM

[2471] William F. Perrin. First record of the melon-headed whale, Peponocephala electra, in the eastern Pacific, with a summary of world distribution. Fishery Bulletin, 74(2):??, 1976. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/first-record-melon-headed-whale-
peponocephala-electra-eastern-pacific-summary-world;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
perrin1.pdf.

Craddock:1976:EIW

[2472] Donovan R. Craddock. Effects of increased water temperature on Daphnia pulex. Fishery Bulletin, 74(2):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/effects-increased-water-temperature-daphnia-
pulex;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1976/742/craddock.pdf.

Perrin:1976:GRS

[2473] William F. Perrin, James M. Coe, and James R. Zweifel. Growth and reproduction of the spotted porpoise, Stenella attenuata, in the offshore Eastern Tropical Pacific. Fishery Bulletin, 74(2):??, 1976. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/growth-and-reproduction-spotted-
porpoise-stenella-attenuata-offshore-eastern-tropical;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1976/742/
perrin.pdf.

Carlson:1976:FJS

[2474] H. Richard Carlson. Foods of juvenile sockeye salmon, Oncorhynchus nerka, in the inshore coastal waters of Bristol Bay, Alaska, 1966–67. Fishery Bulletin, 74(2):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
foods-juvenile-sockeye-salmon-oncorhynchus-nerka-inshore-coastal-
waters-bristol-bay-alaska;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/carlson%20(1).pdf.

Hanson:1976:SIM

[2475] Charles H. Hanson and Jonathan Bell. Subtidal and intertidal marine fouling on artificial substrata in Northern Puget Sound, Washington. Fishery Bulletin, 74(2):??, 1976. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/subtidal-and-intertidal-marine-
fouling-artificial-substrata-northern-puget-sound-washington;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1976/742/hanson.pdf.

Sakagawa:1976:GLR

[2476] Gary T. Sakagawa and Makoto Kimura. Growth of laboratory-reared northern anchovy, Engraulis mordax, from Southern California. Fishery Bulletin, 74(2):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
growth-laboratory-reared-northern-anchovy-engraulis-mordax-
southern-california;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1976/742/sakagawa.pdf.

Laird:1976:OTG

[2477] Chae E. Laird, Elizabeth G. Lewis, and Paul A. Haefner, Jr. Occurrence of two galatheid crustaceans, Munida forceps and Munidopsis bermudezi, in the Chesapeake Bight of the western North Atlantic Ocean. Fishery Bulletin, 74(2):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/occurrence-two-galatheid-crustaceans-
munida-forceps-and-munidopsis-bermudezi-chesapeake;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
laird.pdf.

Hastings:1976:OFF

[2478] Robert W. Hastings, Larry H. Ogren, and Michael T. Mabry. Observations on the fish fauna associated with offshore platforms in the Northeastern Gulf of Mexico. Fishery Bulletin, 74(2):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/observations-
fish-fauna-associated-offshore-platforms-northeastern-gulf-

mexico;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1976/742/hastings.pdf.

Griffin:1976:EFA

[2479] Wade L. Griffin, Newton J. Wardlaw, and John P. Nichols. Economic and financial analysis of increasing costs in the Gulf shrimp fleet. Fishery Bulletin, 74(2):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
economic-and-financial-analysis-increasing-costs-gulf-shrimp-
fleet;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/griffin.pdf.

Weis:1976:EMC

[2480] Judith S. Weis. Effects of mercury, cadmium, and lead salts on regeneration and ecdysis in the fiddler crab, Uca pugilator. Fishery Bulletin, 74 (2):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/effects-
mercury-cadmium-and-lead-salts-regeneration-and-ecdysis-fiddler-
crab-uca-pugilator;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/weis%20(1).pdf.

Breder:1976:FSO

[2481] Charles M. Breder, Jr. Fish schools as operational structures. Fishery Bulletin, 74(3):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fish-schools-operational-structures;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1976/743/
breder.pdf.

Parrack:1976:EFE

[2482] M. L. Parrack. Estimation of fishing effort in the Western North Atlantic from aerial search data. Fishery Bulletin, 74(3):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/estimation-fishing-
effort-western-north-atlantic-aerial-search-data;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1976/743/
parrack.pdf.

Gadbois:1976:OVN

[2483] D. F. Gadbois, E. M. Ravesi, and R. C. Lundstrom. Occurrence of volatile N-nitrosamines in Japanese salmon roe. Fishery Bulletin, 74(3):683–686, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/occurrence-

volatile-n-nitrosamines-japanese-salmon-roe;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/gadbois.
pdf.

Clarke:1976:VDO

[2484] Thomas A. Clarke and Patricia J. Wagner. Vertical distribution and other aspects of the ecology of certain mesopelagic fishes taken near Hawaii. Fishery Bulletin, 74(3):??, 1976. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/vertical-distribution-and-other-
aspects-ecology-certain-mesopelagic-fishes-taken-near;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1976/743/
clarke.pdf.

Skillman:1976:BGC

[2485] Robert A. Skillman and Marian Y. Y. Yong. Von Bertalanffy growth curves for striped marlin, Tetrapturus audax, and blue marlin, Makaira nigricans, in the central North Pacific Ocean. Fishery Bulletin, 74 (3):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/von-
bertalanffy-growth-curves-striped-marlin-tetrapturus-audax-
and-blue-marlin-makaira;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1976/743/skillman.pdf.

Watkins:1976:UPM

[2486] William A. Watkins and William E. Schevill. Underwater paint marking of porpoises. Fishery Bulletin, 74(3):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/underwater-paint-marking-
porpoises;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/watkins.pdf.

Flanagan:1976:OCF

[2487] Christine A. Flanagan and John R. Hendrickson. Observations on the commercial fishery and reproductive biology of the totoaba, Cynoscion macdonaldi, in the Northern Gulf of California. Fishery Bulletin, 74(3): ??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/observations-
commercial-fishery-and-reproductive-biology-totoaba-cynoscion-
macdonaldi;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1976/743/flanagan.pdf.

Weinstein:1976:PTG

[2488] Michael P. Weinstein and Ralph W. Yerger. Protein taxonomy of the Gulf of Mexico and Atlantic Ocean seatrouts, genus Cynoscion. Fishery Bulletin, 74(3):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/protein-taxonomy-gulf-mexico-and-atlantic-
ocean-seatrouts-genus-cynoscion;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1976/743/weinstein.pdf.

Edgar:1976:GFE

[2489] Robert K. Edgar and James G. Hoff. Grazing of freshwater and estuarine, benthic diatoms by adult Atlantic menhaden, Brevoortia tyrannus. Fishery Bulletin, 74(3):??, 1976. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/grazing-freshwater-and-estuarine-
benthic-diatoms-adult-atlantic-menhaden-brevoortia;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
edgar.pdf.

Hobson:1976:TIA

[2490] Edmund S. Hobson and James R. Chess. Trophic interactions among fishes and zooplankters near shore at Santa Catalina Island, California. Fishery Bulletin, 74(3):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/trophic-interactions-among-fishes-and-zooplankters-
near-shore-santa-catalina-island;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1976/743/hobson.pdf.

VonWesternhagen:1976:PPR

[2491] Hein Von Westernhagen and Harald Rosenthal. Predator–prey relationship between Pacific herring, Clupea harengus Pallasi, larvae and a predatory hyperiid amphipod, Hyperoche medusarum. Fishery Bulletin, 74(3):669–674, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/predator-prey-relationship-between-pacific-herring-
clupea-harengus-pallasi-larvae-and;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1976/743/westernhagen.pdf.

Johnson:1976:EEH

[2492] Allyn G. Johnson. Electrophoretic evidence of hybrid snow crab, Chionoecetes bairdi x opilio. Fishery Bulletin, 74(3):??, 1976. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/electrophoretic-evidence-hybrid-
snow-crab-chionoecetes-bairdi-x-opilio;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/johnson%
20(1).pdf.

Korn:1976:UDD

[2493] Sid Korn, Nina Hirsch, and Jeannette W. Struhsaker. Uptake, distribution, and depuration of ^{14}C-benzene in northern anchovy, Engraulis mordax, and striped bass, Morone saxatilis. Fishery Bulletin, 74(3):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/uptake-distribution-and-depuration-14c-benzene-
northern-anchovy-engraulis-mordax-and;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1976/743/korn.pdf.

Zweifel:1976:PPG

[2494] James R. Zweifel and Reuben Lasker. Prehatch and posthatch growth of fishes — a general model. Fishery Bulletin, 74(3):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/prehatch-and-
posthatch-growth-fishes-general-model;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1976/743/zweifel.pdf.

Korn:1976:EBG

[2495] Sid Korn, Jeannette W. Struhsaker, and Pete Benville, Jr. Effects of benzene on growth, fat content, and caloric content of striped bass, Morone saxatilis. Fishery Bulletin, 74(3):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/effects-benzene-growth-fat-content-and-caloric-
content-striped-bass-morone-saxatilis;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/korn1.pdf.

Laurs:1976:ERT

[2496] R. Michael Laurs, William H. Lenarz, and Robert N. Nishimoto. Estimates of rates of tag shedding by North Pacific albacore, Thunnus alalunga. Fishery Bulletin, 74(3):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/estimates-rates-tag-shedding-north-pacific-
albacore-thunnus-alalunga;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1976/743/laurs.pdf.

Arthur:1976:FFL

[2497] David K. Arthur. Food and feeding of larvae of three fishes occurring in the California Current, Sardinops sagax, Engraulis mordax, and Trachurus symmetricus. Fishery Bulletin, 74(3):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/food-and-feeding-larvae-three-fishes-
occurring-california-current-sardinops-sagax-engraulis;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
arthur.pdf.

Staeger:1976:FMQ

[2498] William H. Staeger and Howard F. Horton. Fertilization method quantifying gamete concentrations and maximizing larvae production in Crassostrea gigas. Fishery Bulletin, 74(3):??, 1976. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/fertilization-method-quantifying-
gamete-concentrations-and-maximizing-larvae-production;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
staeger.pdf.

Manzer:1976:DFF

[2499] J. I. Manzer. Distribution, food, and feeding of the threespine stickleback, Gasterosteus aculeatus, in Great Central Lake, Vancouver Island, with comments on competition for food with juvenile sockeye salmon, Oncorhynchus nerka. Fishery Bulletin, 74(3):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/distribution-
food-and-feeding-threespine-stickleback-gasterosteus-aculeatus-
great-central;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1976/743/manzer.pdf.

Zimmerman:1976:PSP

[2500] Steven T. Zimmerman and Robert S. McMahon. Paralytic shellfish poisoning in Tenakee, southeastern Alaska: a possible cause. Fishery Bulletin, 74(3):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/paralytic-shellfish-poisoning-tenakee-southeastern-
alaska-possible-cause;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/zimmerman_0.pdf.

Miller:1976:NAC

[2501] R. J. Miller. North American crab fisheries: Regulations and their rationales. Fishery Bulletin, 74(3):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/north-american-crab-fisheries-regulations-
and-their-rationales;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1976/743/miller.pdf.

Collins:1976:OGP

[2502] Jeff Collins. Oil and grease: a proposed analytical method for fishery waste effluents. Fishery Bulletin, 74(3):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/oil-and-grease-proposed-analytical-method-
fishery-waste-effluents;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/collins_1.pdf.

Collins:1976:FWE

[2503] Jeff Collins and Richard D. Tenney. Fishery waste effluents: a method to determine relationships between chemical oxygen demand and residue. Fishery Bulletin, 74(4):??, 1976. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/fishery-waste-effluents-method-
determine-relationships-between-chemical-oxygen-demand-and;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1976/744/collins.pdf.

Krouse:1976:SCG

[2504] Jay S. Krouse. Size composition and growth of young rock crab, Cancer irroratus, on a rocky beach in Maine. Fishery Bulletin, 74(4):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/size-composition-and-growth-young-rock-crab-cancer-
irroratus-rocky-beach-maine;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1976/744/krouse1.pdf.

Goldberg:1976:SSC

[2505] Stephen R. Goldberg. Seasonal spawning cycles of the sciaenid fishes Genyonemus lineatus and Seriphus politus. Fishery Bulletin, 74(4): 983–984, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/seasonal-
spawning-cycles-sciaenid-fishes-genyonemus-lineatus-and-seriphus-
politus;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/goldberg.pdf.

Crone:1976:LHC

[2506] Richard A. Crone and Carl E. Bond. Life history of Coho salmon, Oncorhynchus klsutch, in Sashin Creek, Southeastern Alaska. Fishery Bulletin, 74(4):897–923, 1976. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
life-history-coho-salmon-oncorhynchus-klsutch-sashin-creek-
southeastern-alaska;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1976/744/crone.pdf.

Lord:1976:DTA

[2507] Gary E. Lord. Decision theory applied to the simulated data acquisition and management of a salmon fishery. Fishery Bulletin, 74 (4):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/decision-
theory-applied-simulated-data-acquisition-and-management-salmon-
fishery;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1976/744/lord.pdf.

Kravitz:1976:FFS

[2508] Michael J. Kravitz, William G. Pearcy, and M. P. Guin. Food of five species of cooccurring flatfishes on Oregon’s Continental Shelf. Fishery Bulletin, 74(4):984–??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
food-five-species-cooccurring-flatfishes-oregons-continental-
shelf;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/kravitz.pdf.

Ebeling:1976:DVN

[2509] Alfred W. Ebeling and Richard N. Bray. Day versus night activity of reef fishes in a kelp forest off Santa Barbara, California. Fishery Bulletin, 74 (4):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/day-versus-
night-activity-reef-fishes-kelp-forest-santa-barbara-california;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1976/744/ebeling.pdf.

Manooch:1976:RCF

[2510] Charles S. Manooch III. Reproductive cycle, fecundity, and sex ratios of the red porgy, Pagrus pagrus (Pisces: Sparidae) in North Carolina. Fishery Bulletin, 74(4):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/reproductive-cycle-fecundity-and-
sex-ratios-red-porgy-pagrus-pagrus-pisces-sparidae-north;

https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1976/
744/manooch.pdf.

Ralston:1976:ADT

[2511] Stephen Ralston. Age determination of a tropical reef butterflyfish utilizing daily growth rings of otoliths. Fishery Bulletin, 74 (4):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/age-
determination-tropical-reef-butterflyfish-utilizing-daily-growth-
rings-otoliths;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/ralston.pdf.

Alvarez:1976:DSE

[2512] Jose Alvarez, Chris O. Andrew, and Fred J. Prochaska. Dual structural equilibrium in the Florida shrimp processing industry. Fishery Bulletin, 74(4):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/dual-
structural-equilibrium-florida-shrimp-processing-industry;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1976/
744/alvarez.pdf.

Hall:1976:MFSa

[2513] Alice S. Hall, Fuad M. Teeny, Laura G. Lewis, William H. Hardman, and Erich J. Gauglitz, Jr. Mercury in fish and shellfish of the Northeast Pacific. I. Pacific halibut, Hippoglossus stenolepis. Fishery Bulletin, 74(4):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/mercury-fish-and-shellfish-northeast-pacific-i-
pacific-halibut-hippoglossus-stenolepis;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1976/744/hall.pdf.

Scherba:1976:ASS

[2514] Stephen Scherba, Jr. and Vincent F. Gallucci. The application of systematic sampling to a study of infauna variation in a soft substrate environment. Fishery Bulletin, 74(4):??, 1976. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/application-systematic-sampling-
study-infauna-variation-soft-substrate-environment;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1976/744/
scherba.pdf.

Rothlisberg:1976:ESU

[2515] Peter C. Rothlisberg and William G. Pearcy. An epibenthic sampler used to study the ontogeny of vertical migration of Pandalus jordani (Decapoda, Caridea). Fishery Bulletin, 74(4):??, 1976. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/epibenthic-sampler-
used-study-ontogeny-vertical-migration-pandalus-jordani-decapoda-
caridea;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/rothlisberg.pdf.

Bailey:1976:PFA

[2516] Jack E. Bailey, Jerome J. Pella, and Sidney G. Taylor. Production of fry and adults of the 1972 brood of pink salmon, Oncorhynchus gorbuscha, from gravel incubators and natural spawning at Auke Creek, Alaska. Fishery Bulletin, 74(4):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/production-fry-and-adults-1972-brood-pink-
salmon-oncorhynchus-gorbuscha-gravel-incubators;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1976/744/
bailey.pdf.

Hall:1976:MFSb

[2517] Alice S. Hall, Fuad M. Teeny, and Erich J. Gauglitz, Jr. Mercury in fish and shellfish of the Northeast Pacific. II. Sablefish, Anoplopoma fimbria. Fishery Bulletin, 74(4):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/mercury-fish-and-shellfish-northeast-
pacific-ii-sablefish-anoplopoma-fimbria;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1976/744/hall1.pdf.

Widersten:1976:CZC

[2518] Bernt Widersten. Ceriantharia, Zoanthidea, Corallimorpharia, and Actiniaria from the Continental Shelf and Slope off the Eastern Coast of the United States. Fishery Bulletin, 74(4):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/ceriantharia-zoanthidea-corallimorpharia-and-
actiniaria-continental-shelf-and-slope-eastern;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1976/744/widersten.
pdf.

Carr:1976:HAU

[2519] William E. S. Carr and Thomas B. Chaney. Harness for attachment of an ultrasonic transmitter to the red drum, Sciaenops ocellata. Fishery Bulletin, 74(4):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/harness-attachment-ultrasonic-transmitter-red-drum-
sciaenops-ocellata;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/carr.pdf.

Bisson:1976:PJC

[2520] Peter A. Bisson and Gerald E. Davis. Production of juvenile Chinook salmon, Oncorhynchus tshawytscha, in a heated model stream. Fishery Bulletin, 74(4):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
production-juvenile-chinook-salmon-oncorhynchus-tshawytscha-
heated-model-stream;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1976/744/bisson.pdf.

Hunter:1976:DCS

[2521] John R. Hunter and Carol Sanchez. Diel changes in swim bladder inflation of the larvae of the northern anchovy, Engraulis mordax. Fishery Bulletin, 74(4):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
diel-changes-swim-bladder-inflation-larvae-northern-anchovy-
engraulis-mordax;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1976/744/hunter.pdf.

Walters:1976:EHS

[2522] John F. Walters. Ecology of Hawaiian sergestid shrimps (Penaeidea: Sergestidae). Fishery Bulletin, 74(4):799–836, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/ecology-hawaiian-sergestid-shrimps-penaeidea-
sergestidae;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/walters.pdf.

Anonymous:1976:I

[2523] Anonymous. Index. Fishery Bulletin, 74(4):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/index-104;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/index%
20(3).pdf.

Blackburn:1976:DEP

[2524] Maurice Blackburn and Walter Nellen. Distribution and ecology of pelagic fishes studied from eggs and larvae in an upwelling area off Spanish Sahara. Fishery Bulletin, 74(4):??, 1976. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/distribution-and-ecology-pelagic-
fishes-studied-eggs-and-larvae-upwelling-area-spanish;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1976/744/
blackburn.pdf.

Keene:1976:CMS

[2525] Donald F. Keene and William G. Pearcy. Comparison of the most successful and least successful West Coast albacore troll fishermen. Fishery Bulletin, 74(4):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
comparison-most-successful-and-least-successful-west-coast-
albacore-troll-fishermen;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1976/744/keene.pdf.

Youngbluth:1976:VDD

[2526] Marsh J. Youngbluth. Vertical distribution and diel migration of euphausiids in the central region of the California Current. Fishery Bulletin, 74(4):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/vertical-
distribution-and-diel-migration-euphausiids-central-region-
california-current;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/youngbluth.pdf.

Brinton:1976:PBE

[2527] Edward Brinton. Population biology of Euphausia pacifica off Southern California. Fishery Bulletin, 74(4):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/population-biology-euphausia-pacifica-southern-
california;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1976/744/brinton.pdf.

Krouse:1976:ICL

[2528] Jay S. Krouse. Incidence of cull lobsters, Homarus americanus, in commercial and research catches off the Maine Coast. Fishery Bulletin, 74 (4):??, 1976. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/incidence-
cull-lobsters-homarus-americanus-commercial-and-research-catches-
maine-coast;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1976/744/krouse.pdf.

Dotson:1976:MSS

[2529] Ronald C. Dotson. Minimum swimming speed of albacore, Thunnus alalunga. Fishery Bulletin, 74(4):??, 1976. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/minimum-swimming-speed-albacore-
thunnus-alalunga;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/dotson.pdf.

Clark:1977:VDD

[2530] Stephen H. Clark and Bradford E. Brown. Vertical distribution, diel vertical migration, and abundance of some mesopelagic fishes in the Eastern Subarctic Pacific Ocean in summer. Fishery Bulletin, 75(1): 1–21, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/vertical-
distribution-diel-vertical-migration-and-abundance-some-mesopelagic-
fishes-eastern;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1977/751/clark.pdf.

Struhsaker:1977:EBT

[2531] Jeannette W. Struhsaker. Effects of benzene (a toxic component of petroleum) on spawning Pacific herring, Clupea harengus Pallasi. Fishery Bulletin, 75(1):??, 1977. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/effects-benzene-toxic-component-
petroleum-spawning-pacific-herring-clupea-harengus-pallasi;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1977/751/struhsaker.pdf.

Leong:1977:MIS

[2532] Roderick Leong. Maturation and induced spawning of captive Pacific mackerel, Scomber japonicus. Fishery Bulletin, 75(1):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/maturation-and-induced-spawning-
captive-pacific-mackerel-scomber-japonicus;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/leong.pdf.

Haefner:1977:LDG

[2533] Paul A. Haefner, Jr. Larval development of Galathea rostrata under laboratory conditions, with a discussion of larval development in the Galatheidae (Crustacea anomura). Fishery Bulletin, 75(1):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/larval-development-
galathea-rostrata-under-laboratory-conditions-discussion-larval-
0;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1977/751/haefner.pdf.

Trent:1977:SGN

[2534] Lee Trent and Paul J. Pristas. Selectivity of gill nets on estuarine and coastal fishes from St. Andrew Bay, Florida. Fishery Bulletin, 75(1):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/selectivity-gill-nets-estuarine-and-coastal-fishes-
st-andrew-bay-florida;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1977/751/trent.pdf.

Austin:1977:IST

[2535] C. Bruce Austin. Incorporating soak time into measurement of fishing effort in trap fisheries. Fishery Bulletin, 75(1):??, 1977. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/incorporating-soak-
time-measurement-fishing-effort-trap-fisheries;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/austin_
0.pdf.

MacInnes:1977:LTC

[2536] J. R. MacInnes, F. P. Thurberg, R. A. Greig, and E. Gould. Long-term cadmium stress in the cunner, Tautogolabrus adspersus. Fishery Bulletin, 75(1):199–203, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/long-term-cadmium-stress-cunner-
tautogolabrus-adspersus;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1977/751/macinnes.pdf.

White:1977:ADR

[2537] Michael L. White and Mark E. Chittenden, Jr. Age determination, reproduction, and population dynamics of the Atlantic croaker, Micropogonias undulatus. Fishery Bulletin, 75(1):??, 1977. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/age-determination-reproduction-
and-population-dynamics-atlantic-croaker-micropogonias;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1977/751/
white.pdf.

Misitano:1977:SCR

[2538] David A. Misitano. Species composition and relative abundance of larval and post-larval fishes in the Columbia River estuary, 1973. Fishery Bulletin, 75(1):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
species-composition-and-relative-abundance-larval-and-post-
larval-fishes-columbia-river;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/misitano%20(1).pdf.

Nelson:1977:LTY

[2539] Walter R. Nelson, Merton C. Ingham, and William E. Schaaf. Larval transport and year-class strength of Atlantic menhaden, Brevoortia tyrannus. Fishery Bulletin, 75(1):??, 1977. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/larval-transport-and-year-class-
strength-atlantic-menhaden-brevoortia-tyrannus;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1977/751/nelson.
pdf.

Clark:1977:CBF

[2540] Stephen H. Clark and Bradford E. Brown. Changes in biomass of finfishes and squids from the Gulf of Maine to Cape Hatteras, 1963–74, as determined from research vessel survey data. Fishery Bulletin, 75(1):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/changes-biomass-finfishes-and-squids-gulf-
maine-cape-hatteras-1963-74-determined-research-0;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
clark_1.pdf.

Gunn:1977:NVT

[2541] John T. Gunn and Merton C. Ingham. A note on: ‘Velocity and transport of the Antilles Current northeast of the Bahama Islands’. Fishery Bulletin, 75(1):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/note-velocity-and-transport-antilles-
current-northeast-bahama-islands;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/gunn.pdf. See [2395].

Norris:1977:BCG

[2542] Kenneth S. Norris, Robert M. Goodman, Bernardo Villa-Ramirez, and Larry Hobbs. Behavior of California gray whale, Eschrichtius robustus, in Southern Baja California, Mexico. Fishery Bulletin, 75(1):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/behavior-california-gray-
whale-eschrichtius-robustus-southern-baja-california-mexico;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1977/751/norris.pdf.

Hosie:1977:BRS

[2543] Michael J. Hosie and Howard F. Horton. Biology of the rex sole, Glyptocephalus zachirus, in waters off Oregon. Fishery Bulletin, 75(1):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/biology-rex-
sole-glyptocephalus-zachirus-waters-oregon;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/hosie.pdf.

Creaser:1977:SAS

[2544] Edwin P. Creaser, Jr. and David A. Clifford. Salinity acclimation in the soft-shell clam, Mya arenaria. Fishery Bulletin,

75(1):??, 1977.  CODEN FSYBAY.  ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
salinity-acclimation-soft-shell-clam-mya-arenaria;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
creaser.pdf.

Pearcy:1977:DDP

[2545] William G. Pearcy, Michael J. Hosie, and Sally L. Richardson. Distribution and duration of pelagic life of larvae of Dover sole, Microstomus pacificus; rex sole, Glyptocephalus zachirus; and petrale sole, Eopsetta jordani, in waters off Oregon. Fishery Bulletin, 75(1):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/distribution-
and-duration-pelagic-life-larvae-dover-sole-microstomus-pacificus-
rex-sole;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1977/751/pearcy.pdf.

Houde:1977:APYa

[2546] Edward D. Houde. Abundance and potential yield of the round herring, Etrumeus teres, and aspects of its early life history in the eastern Gulf of Mexico. Fishery Bulletin, 75(1):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/abundance-and-potential-yield-round-herring-etrumeus-
teres-and-aspects-its-early-life;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/houde.pdf.

Graves:1977:PMM

[2547] John Graves. Photographic method for measuring spacing and density within pelagic fish schools at sea. Fishery Bulletin, 75(1):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/photographic-method-
measuring-spacing-and-density-within-pelagic-fish-schools-sea;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/graves.pdf.

Pristas:1977:CCF

[2548] Paul J. Pristas and Lee Trent. Comparisons of catches of fishes in gill nets in relation to webbing material, time of day, and water depth in St. Andrew Bay, Florida. Fishery Bulletin, 75(1):??, 1977. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/comparisons-catches-fishes-gill-
nets-relation-webbing-material-time-day-and-water-depth-st;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1977/751/pristas.pdf.

Haefner:1977:RBF

[2549] Paul A. Haefner, Jr. Reproductive biology of the female deep-sea red crab, Geryon quinquedens, from the Chesapeake Bight. Fishery Bulletin, 75(1):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
reproductive-biology-female-deep-sea-red-crab-geryon-quinquedens-
chesapeake-bight-0;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/haefner.pdf.

Morrow:1977:FAW

[2550] James E. Morrow, Eldor W. Schallock, and Glenn E. Bergtold. Feeding by Alaska whitefish, Coregonus nelsoni, during the spawning run. Fishery Bulletin, 75(1):234–235, 1977. CODEN FSY-

BAY.  ISSN 0090-0656 (print), 1937-4518 (electronic).  URL

https://spo.nmfs.noaa.gov/content/feeding-alaska-whitefish-
coregonus-nelsoni-during-spawning-run;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/morrow.pdf.

Rohr:1977:BOH

[2551] Bennie A. Rohr and Elmer J. Gutherz. Biology of offshore hake, Merluccius albidus, in the Gulf of Mexico. Fishery Bulletin, 75(1): ??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/biology-
offshore-hake-merluccius-albidus-gulf-mexico;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1977/751/rohr.
pdf.

Richardson:1977:COF

[2552] Sally L. Richardson and William G. Pearcy. Coastal and oceanic fish larvae in an area of upwelling off Yaquina Bay, Oregon. Fishery Bulletin, 75(1):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
coastal-and-oceanic-fish-larvae-area-upwelling-yaquina-bay-
oregon;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/richardson%20(2).pdf.

Fisher:1977:EMW

[2553] William S. Fisher and Daniel E. Wickham. Egg mortalities in wild populations of the Dungeness crab in central and northern California. Fishery Bulletin, 75(1):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/egg-mortalities-wild-populations-
dungeness-crab-central-and-northern-california;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/fisher%
20(1).pdf.

Laurs:1977:SSMa

[2554] R. Michael Laurs, Heeny S. H. Yuen, and James H. Johnson. Small-scale movements of albacore, Thunnus alalunga, in relation to ocean features as indicated by ultrasonic tracking and oceanographic sampling. Fishery Bulletin, 75(2):??, 1977. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/small-scale-movements-albacore-
thunnus-alalunga-relation-ocean-features-indicated;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1977/752/
laurs.pdf.

Horn:1977:OFG

[2555] Michael H. Horn. Observations on feeding, growth, locomotor behavior, and buoyancy of a pelagic stromateoid fish, Icichthys lockingtoni. Fishery Bulletin, 75(2):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/observations-feeding-growth-locomotor-behavior-and-
buoyancy-pelagic-stromateoid-fish;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/horn%20(1).pdf.

Major:1977:PPI

[2556] Peter F. Major. Predator–prey interactions in schooling fishes during periods of twilight: a study of the silverside Pranesus insularum in Hawaii. Fishery Bulletin, 75(2):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/predator-prey-interactions-schooling-fishes-during-
periods-twilight-study-silverside;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1977/752/major.pdf.

Patten:1977:BSL

[2557] Benjamin G. Patten. Body size and learned avoidance as factors affecting predation on Coho salmon, Oncorhynchus kisutch, fry by torrent sculpin, Cottus rhotheus. Fishery Bulletin, 75(2):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/body-size-and-learned-avoidance-
factors-affecting-predation-coho-salmon-oncorhynchus;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
patten%20(1).pdf.

Schwartz:1977:FMT

[2558] Frank J. Schwartz and Hugh J. Porter. Fishes, macroinvertebrates, and their ecological interrelationships with a calico scallop bed off North Carolina. Fishery Bulletin, 75(2):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fishes-macroinvertebrates-and-their-ecological-
interrelationships-calico-scallop-bed-north;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1977/752/schwartz.
pdf.

Jewett:1977:DMS

[2559] Stephen C. Jewett and Richard E. Haight. Description of megalopa of snow crab, Chionoecetes bairdi (Majidae, subfamily Oregoniinae). Fishery Bulletin, 75(2):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/description-megalopa-snow-crab-chionoecetes-bairdi-
majidae-subfamily-oregoniinae;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/jewett.pdf.

Six:1977:AAD

[2560] Lawrence D. Six and Howard F. Horton. Analysis of age determination methods for yellowtail rockfish, canary rockfish, and black rockfish off Oregon. Fishery Bulletin, 75(2):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/analysis-age-determination-methods-yellowtail-rockfish-
canary-rockfish-and-black-rockfish;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1977/752/six.pdf.

Young:1977:IRB

[2561] Richard Edward Young and Clyde F. E. Roper. Intensity regulation of bioluminescence during countershading in living midwater animals. Fishery Bulletin, 75(2):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
intensity-regulation-bioluminescence-during-countershading-
living-midwater-animals;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1977/752/young.pdf.

Colebrook:1977:AFB

[2562] J. M. Colebrook. Annual fluctuations in biomass of taxonomic groups of zooplankton in the California Current, 1955–59. Fishery Bulletin, 75 (2):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/annual-
fluctuations-biomass-taxonomic-groups-zooplankton-california-
current-1955-59;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1977/752/colebrook.pdf.

Farfante:1977:ASS

[2563] Isabel Perez Farfante. American solenocerid shrimps of the genera Hymenopenaeus, Haliporoides, Pleoticus, Hadropenaeus new genus, and Mesopenaeus new genus. Fishery Bulletin, 75(2):261–346, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/american-solenocerid-
shrimps-genera-hymenopenaeus-haliporoides-pleoticus-hadropenaeus-
new;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/perez%20(2).pdf.

Collins:1977:FWE

[2564] Jeff Collins and Richard D. Tenney. Fishery waste effluents: a suggested system for determining and calculating pollutant parameters. Fishery Bulletin, 75(2):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishery-waste-effluents-suggested-system-determining-and-calculating-
pollutant-parameters;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1977/752/collins.pdf.

Laurs:1977:SSMb

[2565] R. Michael Laurs, Heeny S. H. Yuen, and James H. Johnson. Small scale movements of albacore, Thunnus alalunga, in relation to ocean features as indicated by ultrasonic tracking and oceanographic sampling. Fishery Bulletin, 75(2):347–355, 1977. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/small-scale-movements-albacore-
thunnus-alalunga-relation-ocean-features-indicated-0;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
laurs.pdf.

Gunderson:1977:PBP

[2566] Donald R. Gunderson. Population biology of Pacific Ocean perch, Sebastes alutus, stocks in the Washington–Queen Charlotte Sound Region, and their response to fishing. Fishery Bulletin, 75(2):369–403, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/population-biology-
pacific-ocean-perch-sebastes-alutus-stocks-washington-queen-
charlotte;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1977/752/gunderson.pdf.

Sharp:1977:EMA

[2567] Gary D. Sharp and Ronald C. Dotson. Energy for migration in albacore, Thunnus alalunga. Fishery Bulletin, 75(2):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/energy-migration-albacore-thunnus-alalunga;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/sharp.pdf.

Arthur:1977:DSA

[2568] David K. Arthur. Distribution, size, and abundance of microcopepods in the California Current system and their possible influence on survival of marine teleost larvae. Fishery Bulletin, 75(3):??, 1977. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/distribution-size-and-abundance-
microcopepods-california-current-system-and-their-possible;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1977/753/arthur.pdf.

Rafail:1977:SSF

[2569] Samir Z. Rafail. A simplification for the study of fish populations by capture data. Fishery Bulletin, 75(3):??, 1977. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/simplification-study-
fish-populations-capture-data;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1977/753/rafail.pdf.

Korn:1977:UDD

[2570] Sid Korn, Nina Hirsch, and Jeannette W. Struhsaker. The uptake, distribution, and depuration of ^{14}C benzene and ^{14}C toluene in Pacific herring, Clupea harengus pallasi. Fishery Bulletin, 75(3):633–636, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/uptake-distribution-
and-depuration-14c-benzene-and-14c-toluene-pacific-herring-
clupea;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/korn.pdf.

Houde:1977:APYb

[2571] Edward D. Houde. Abundance and potential yield of the Atlantic thread herring, Opisthonema oglinum, and aspects of its early life history in the Eastern Gulf of Mexico. Fishery Bulletin, 75(3):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/abundance-and-
potential-yield-atlantic-thread-herring-opisthonema-oglinum-
and-aspects-its;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1977/753/houde.pdf.

Sciarrotta:1977:DBB

[2572] Terry C. Sciarrotta and Donald R. Nelson. Diel behavior of the blue shark, Prionace glauca, near Santa Catalina Island, California. Fishery Bulletin, 75(3):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
diel-behavior-blue-shark-prionace-glauca-near-santa-catalina-
island-california;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1977/753/sciarrotta.pdf.

Foltz:1977:FHF

[2573] Jeffrey W. Foltz and Carroll R. Norden. Food habits and feeding chronology of rainbow smelt, Osmerus mordax, in Lake Michigan. Fishery Bulletin, 75(3):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/food-
habits-and-feeding-chronology-rainbow-smelt-osmerus-mordax-
lake-michigan;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/foltz.pdf.

Houde:1977:APYc

[2574] Edward D. Houde. Abundance and potential yield of the scaled sardine, Harengula jaguana, and aspects of its early life history in the Eastern Gulf of Mexico. Fishery Bulletin, 75(3):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/abundance-and-potential-yield-scaled-
sardine-harengula-jaguana-and-aspects-its-early-life;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1977/753/
houde1.pdf.

Scura:1977:VSPa

[2575] Edward D. Scura and Charles W. Jerde. Various species of phytoplankton as food for larval northern anchovy, Engraulis mordax, and relative nutritional value of the dinoflagellates Gymnodinium splendens and Gonyaulax polyedra. Fishery Bulletin, 75(3): ??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/various-
species-phytoplankton-food-larval-northern-anchovy-engraulis-
mordak-and-relative;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1977/753/scura.pdf.

Loesch:1977:UMY

[2576] Joseph G. Loesch. Useable meat yields in the Virginia surf clam fishery. Fishery Bulletin, 75(3):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/useable-meat-yields-virginia-surf-clam-fishery;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/loesch.pdf.

Laroche:1977:DLE

[2577] Wayne A. Laroche. Description of larval and early juvenile vermilion snapper, Rhomboplites aurorubens. Fishery Bulletin, 75(3):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/description-larval-
and-early-juvenile-vermilion-snapper-rhomboplites-aurorubens;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1977/753/laroche.pdf.

Sissenwine:1977:CSM

[2578] Michael P. Sissenwine. A compartmentalized simulation model of the Southern New England yellowtail flounder, Limanda ferruginea, fishery. Fishery Bulletin, 75(3):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/compartmentalized-simulation-model-southern-new-england-
yellowtail-flounder-limanda;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1977/753/sissenwine.pdf.

Hall:1977:MFS

[2579] Alice S. Hall, Fuad M. Teeny, and Erich J. Gauglitz, Jr. Mercury in fish and shellfish of the Northeast Pacific. III. Spiny dogfish, Squalus acanthias. Fishery Bulletin, 75(3):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/mercury-fish-and-shellfish-northeast-
pacific-iii-spiny-dogfish-squalus-acanthias;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/hall.
pdf.

Laurence:1977:BMA

[2580] Geoffrey C. Laurence. A bioenergetic model for the analysis of feeding and survival potential of winter flounder, Pseudopleuronectes americanus, larvae during the period from hatching to metamorphosis. Fishery Bulletin, 75(3):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
bioenergetic-model-analysis-feeding-and-survival-potential-
winter-flounder;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1977/753/laurence.pdf.

Owers:1977:IER

[2581] James E. Owers. Income estimates and reasonable returns in Alaska’s salmon fisheries. Fishery Bulletin, 75(3):??, 1977. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/income-estimates-and-
reasonable-returns-alaskas-salmon-fisheries;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/owers.
pdf.

Pearse:1977:LMM

[2582] John S. Pearse, Daniel P. Costa, Marc B. Yellin, and Catherine R. Agegian. Localized mass mortality of red sea urchin, Strongylocentrotus franciscanus, near Santa Cruz, California. Fishery Bulletin, 75(3):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
localized-mass-mortality-red-sea-urchin-strongylocentrotus-
franciscanus-near-santa-cruz;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/pearse.pdf.

Lundstrom:1977:IFS

[2583] Ronald C. Lundstrom. Identification of fish species by thin-layer polyacrylamide gel isoelectric focusing. Fishery Bulletin, 75(3):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/identification-fish-
species-thin-layer-polyacrylamide-gel-isoelectric-focusing;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1977/753/lundstrom.pdf.

McDermott-Ehrlich:1977:CHD

[2584] D. J. McDermott-Ehrlich, M. J. Sherwood, b T. C. Heesen, D. R. Young, and A. J. Mearns. Chlorinated hydrocarbons in Dover sole, Microstomus pacificus: Local migrations and fin erosion. Fishery Bulletin, 75(3): ??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/chlorinated-
hydrocarbons-dover-sole-microstomus-pacificus-local-migrations-
and-fin-erosion;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/mcdermott.pdf.

Rensel:1977:FRS

[2585] John E. Rensel and Earl F. Prentice. First record of a second mating and spawning of the spot prawn, Pandalus platyceros, in captivity. Fishery Bulletin, 75(3):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/first-record-second-mating-and-spawning-spot-prawn-
pandalus-platyceros-captivity;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/rensel.pdf.

Olla:1977:CSB

[2586] Bori L. Olla and Carol Samet. Courtship and spawning behavior of the tautog, Tautoga onitis (Pisces: Labridae), under laboratory conditions. Fishery Bulletin, 75(3):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/courtship-and-spawning-behavior-tautog-tautoga-
onitis-pisces-labridae-under-laboratory;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1977/753/olla.pdf.

Scura:1977:VSPb

[2587] Edward D. Scura and Charles W. Jerde. Various species of phytoplankton as food for larval northern anchovy, Engraulis mordax, and relative nutritional value of the dinoflagellates Gymnodinium splendens and Gonyaulax polyedra. Fishery Bulletin, 75(3):

??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/various-
species-phytoplankton-food-larval-northern-anchovy-engraulis-
mordax-and-relative;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/scura.pdf.

Dizon:1977:EDO

[2588] Andrew E. Dizon. Effect of dissolved oxygen concentration and salinity on swimming speed of two species of tunas. Fishery Bulletin, 75 (3):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/effect-
dissolved-oxygen-concentration-and-salinity-swimming-speed-
two-species-tunas;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/dizon%20(1).pdf.

Patten:1977:STT

[2589] Benjamin G. Patten. Short-term thermal resistance of zoeae of 10 species of crabs from Puget Sound, Washington. Fishery Bulletin, 75(3):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/short-term-thermal-
resistance-zoeae-10-species-crabs-puget-sound-washington;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1977/
753/patten.pdf.

Perrin:1977:RPO

[2590] William F. Perrin, Ruth B. Miller, and Priscilla A. Sloan. Reproductive parameters of the offshore spotted dolphin, a geographical form of Stenella attenuata, in the eastern tropical Pacific, 1973– 75. Fishery Bulletin, 75(3):629–633, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/reproductive-parameters-offshore-
spotted-dolphin-geographical-form-stenella-attenuata;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
perrin%20(1).pdf.

Hall:1977:NLD

[2591] John D. Hall. A nonlethal lavage device for sampling stomach contents of small marine mammals. Fishery Bulletin, 75(3):??, 1977. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/nonlethal-lavage-device-sampling-
stomach-contents-small-marine-mammals;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/hall1.pdf.

Ingham:1977:OCS

[2592] Merton C. Ingham, Steven K. Cook, and Keith A. Hausknecht. Oxycline characteristics and skipjack tuna distribution in the Southeastern Tropical Atlantic. Fishery Bulletin, 75(4):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/oxycline-characteristics-
and-skipjack-tuna-distribution-southeastern-tropical-atlantic;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1977/754/ingham.pdf.

Laurs:1977:SMN

[2593] R. Michael Laurs and Ronald J. Lynn. Seasonal migration of North Pacific albacore, Thunnus alalunga, into North American coastal waters: Distribution, relative abundance, and association with transition zone waters. Fishery Bulletin, 75(4):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/seasonal-migration-north-pacific-albacore-
thunnus-alalunga-north-american-coastal-waters;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/laurs%
20(1).pdf.

Harrell:1977:GLC

[2594] Lee W. Harrell. Gallbladder lesions in cultured Pacific salmon. Fishery Bulletin, 75(4):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
gallbladder-lesions-cultured-pacific-salmon;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/harrell%
20(1).pdf.

Johns:1977:LDS

[2595] D. Michael Johns and William H. Lang. Larval development of the spider crab, Libinia emarginata (Majidae). Fishery Bulletin, 75(4):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/larval-development-
spider-crab-libinia-emarginata-majidae;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1977/754/johns.pdf.

Krygier:1977:SCM

[2596] Earl E. Krygier and William G. Pearcy. The source of cobalt-60 and migrations of albacore off the west coast of North America. Fishery Bulletin, 75(4):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/source-
cobalt-60-and-migrations-albacore-west-coast-north-america;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/krygier.pdf.

Boehlert:1977:TSB

[2597] George W. Boehlert. Timing of the surface-to-benthic migration in juvenile rockfish, Sebastes diploproa, off southern California. Fishery Bulletin, 75(4):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/timing-surface-benthic-migration-juvenile-rockfish-
sebastes-diploproa-southern-california;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/boehlert.
pdf.

Jones:1977:USS

[2598] Albert C. Jones and Alexander Dragovich. The United States shrimp fishery off Northeastern South America (1972–74). Fishery Bulletin, 75 (4):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/united-states-
shrimp-fishery-northeastern-south-america-1972-74;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1977/754/
jones.pdf.

Phinney:1977:LWW

[2599] Duane E. Phinney. Length-width-weight relationships for mature male snow crab, Chionocoetes bairdi. Fishery Bulletin, 75(4):870–871, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/length-width-weight-
relationships-mature-male-snow-crab-chionocoetes-bairdi;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
phinney.pdf.

Anonymous:1977:I

[2600] Anonymous. Index. Fishery Bulletin, 75(4):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/index-105;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/index%
20(4).pdf.

Perrin:1977:GRE

[2601] William F. Perrin, David B. Holts, and Ruth B. Miller. Growth and reproduction of the eastern spinner dolphin, a geographical form of Stenella longirostris in the Eastern Tropical Pacific. Fishery Bulletin, 75 (4):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/growth-
and-reproduction-eastern-spinner-dolphin-geographical-form-
stenella-longirostris;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1977/754/perrin.pdf.

Fine:1977:TAC

[2602] Michael L. Fine, Howard E. Winn, Linda Joest, and Paul J. Perkins. Temporal aspects of calling behavior in the oyster toadfish, Opsanus tau. Fishery Bulletin, 75(4):871–??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/temporal-aspects-calling-
behavior-oyster-toadfish-opsanus-tau;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/fine.pdf.

Peterson:1977:SCZ

[2603] William T. Peterson and Charles B. Miller. Seasonal cycle of zooplankton abundance and species composition along the Central Oregon Coast. Fishery Bulletin, 75(4):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/seasonal-cycle-zooplankton-abundance-
and-species-composition-along-central-oregon-coast;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1977/754/
peterson.pdf.

Weinstein:1977:BHP

[2604] Michael P. Weinstein and Kenneth L. Heck, Jr. Biology and host-parasite relationships of Cymothoa excisa (Isopoda, Cymothiodae) with three species of snappers (Lutjanidae) on the Caribbean coast of Panama. Fishery Bulletin, 75(4):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
biology-and-host-parasite-relationships-cymothoa-excisa-isopoda-
cymothiodae-three-species;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/weinstein.pdf.

Carline:1977:PTP

[2605] Robert F. Carline. Production by three populations of wild brook trout with emphasis on influence of recruitment rates. Fishery Bulletin, 75(4): ??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/production-
three-populations-wild-brook-trout-emphasis-influence-recruitment-
rates;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1977/754/carline.pdf.

Rosenblatt:1977:RGD

[2606] Richard H. Rosenblatt and John L. Butler. The ribbonfish genus Desmodema, with the description of a new species (Pisces, Trachipteridae). Fishery Bulletin, 75(4):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/ribbonfish-genus-desmodema-description-new-species-
pisces-trachipteridae;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1977/754/rosenblatt.pdf.

Howell:1977:FSN

[2607] W. Huntting Howell and David H. Kesler. Fecundity of the southern New England stock of yellowtail flounder, Limanda ferruginea. Fishery Bulletin, 75(4):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/fecundity-southern-new-england-stock-
yellowtail-flounder-limanda-ferruginea;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/howell.pdf.

Chao:1977:LHF

[2608] Labbish N. Chao and John A. Musick. Life history, feeding habits, and functional morphology of juvenile Sciaenid fishes in the York River Estuary, Virginia. Fishery Bulletin, 75(4):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/life-history-
feeding-habits-and-functional-morphology-juvenile-sciaenid-
fishes-york-river;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1977/754/chao.pdf.

Seckel:1977:KHO

[2609] Gunter R. Seckel and Marian Y. Y. Yong. Koko Head, Oahu, seasurface temperatures and salinities, 1956–73, and Christmas Island sea-surface temperatures, 1954–73. Fishery Bulletin, 75(4):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/koko-head-oahu-sea-
surface-temperatures-and-salinities-1956-73-and-christmas-island-
sea;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1977/754/seckel.pdf.

Green:1977:MFM

[2610] John H. Green and Louis J. Ronsivalli. ‘Mock fish’ method for studying microbial inhibiting agents. Fishery Bulletin, 75(4):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic).  URL https://spo.nmfs.noaa.gov/content/mock-fish-
method-studying-microbial-inhibiting-agents;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/howell_
0.pdf.

Dewitt:1977:NGS

[2611] Hugh H. Dewitt. A new genus and species of eelpout (Pisces, Zoarcidae) from the Gulf of Mexico. Fishery Bulletin, 75(4):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic).  URL https://spo.nmfs.noaa.gov/content/new-genus-
and-species-eelpout-pisces-zoarcidae-gulf-mexico;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1977/754/
dewitt.pdf.

Smith:1977:BSF

[2612] Ronal W. Smith and Franklin C. Daiber. Biology of the summer flounder, Paralichthys dentatus, in Delaware Bay. Fishery Bulletin, 75(4):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
biology-summer-flounder-paralichthys-dentatus-delaware-bay;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1977/754/smith.pdf.

Goldberg:1977:RCP

[2613] Stephen R. Goldberg and William C. Ticknor, Jr. Reproductive cycle of the pink surfperch, Zalembius rosaceus (Embiotocidae). Fishery Bulletin, 75(4):??, 1977. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
reproductive-cycle-pink-surfperch-zalembius-rosaceus-embiotocidae;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/goldberg%20(1).pdf.

Ben-Tuvia:1978:IFT

[2614] Adam Ben-Tuvia. Immigration of fishes through the Suez Canal. Fishery Bulletin, 76(1):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/immigration-fishes-through-suez-canal;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1978/761/
bentuvia.pdf.

Couch:1978:DPT

[2615] John A. Couch. Diseases, parasites, and toxic responses of commercial penaeid shrimps of the Gulf of Mexico and South Atlantic coasts of North America. Fishery Bulletin, 76(1):??, 1978. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/diseases-parasites-
and-toxic-responses-commercial-penaeid-shrimps-gulf-mexico-
and-south;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1978/761/couch.pdf.

Markle:1978:TDR

[2616] Douglas F. Markle. Taxonomy and distribution of Rouleina attrita and Rouleina maderensis (Pisces: Alepocephalidae). Fishery Bulletin, 76(1):79–87, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/taxonomy-and-distribution-rouleina-attrita-and-
rouleina-maderensis-pisces-alepocephalidae;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1978/761/markle.pdf.

Scholz:1978:HMI

[2617] Allan T. Scholz, Jon C. Cooper, Ross M. Horrall, and Arthur D. Hasler. Homing of morpholine-imprinted brown trout, Salmo trutta. Fishery Bulletin, 76(1):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
homing-morpholine-imprinted-brown-trout-salmo-trutta;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
scholz.pdf.

Berrien:1978:ELS

[2618] Peter L. Berrien. Eggs and larvae of Scomber scombrus and Scomber japonicus in continental shelf waters between Massachusetts and Florida. Fishery Bulletin, 76(1):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/eggs-and-larvae-scomber-scombrus-and-
scomber-japonicus-continental-shelf-waters-between;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1978/761/
berrien.pdf.

Fable:1978:DRE

[2619] William A. Fable, Jr., Theodore D. Williams, and C. R. Arnold. Description of reared eggs and young larvae of the spotted seatrout, Cynoscion nebulosus. Fishery Bulletin, 76(1):??, 1978. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/description-reared-
eggs-and-young-larvae-spotted-seatrout-cynoscion-nebulosus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1978/761/fable.pdf.

Ross:1978:TOL

[2620] Stephen T. Ross. Trophic ontogeny of the leopard searobin, Prionotus scitulus (Pisces: Triglidae). Fishery Bulletin, 76(1):??, 1978. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/trophic-ontogeny-leopard-
searobin-prionotus-scitulus-pisces-triglidae;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1978/761/ross.
pdf.

Eldridge:1978:DVC

[2621] Peter J. Eldridge, Frederick H. Berry, and M. Clinton Miller III. Diurnal variations in catches of selected species of ichthyoneuston by the Boothbay neuston net off Charleston, South Carolina. Fishery Bulletin, 76(1):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
diurnal-variations-catches-selected-species-ichthyoneuston-
boothbay-neuston-net-charleston;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/eldridge.pdf.

Bigford:1978:ESD

[2622] Thomas E. Bigford. Effect of several diets on survival, development time, and growth of laboratory-reared spider crab, Libinia emarginata, larvae. Fishery Bulletin, 76(1):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/effect-several-diets-survival-development-time-and-
growth-laboratory-reared-spider-crab;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1978/761/bigford.pdf.

Grabe:1978:FFH

[2623] Stephen A. Grabe. Food and feeding habits of juvenile Atlantic tomcod, Microgadus tomcod, from Haverstraw Bay, Hudson River. Fishery Bulletin, 76(1):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/food-and-feeding-habits-juvenile-atlantic-
tomcod-microgadus-tomcod-haverstraw-bay-hudson;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1978/761/grabe.
pdf.

Smith:1978:DML

[2624] W. G. Smith, J. D. Sibunka, and A. Wells. Diel movements of larval yellowtail flounder, Limanda ferruginea, determined from discrete depth sampling. Fishery Bulletin, 76(1):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/diel-movements-larval-yellowtail-flounder-
limanda-ferruginea-determined-discrete-depth;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1978/761/smith.
pdf.

Brousseau:1978:SCF

[2625] Diane J. Brousseau. Spawning cycle, fecundity, and recruitment in a population of soft-shell clam, Mya arenaria, from Cape Ann, Massachusetts. Fishery Bulletin, 76(1):??, 1978. CODEN FSY-

BAY.  ISSN 0090-0656 (print), 1937-4518 (electronic).  URL

https://spo.nmfs.noaa.gov/content/spawning-cycle-fecundity-
and-recruitment-population-soft-shell-clam-mya-arenaria-cape-
ann;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1978/761/brousseau.pdf.

Haynes:1978:DLHa

[2626] Evan Haynes. Description of larvae of the humpy shrimp, Pandalus goniurus, reared in situ in Kachemak Bay, Alaska. Fishery Bulletin, 76(1): ??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/description-
larvae-humpy-shrimp-pandalus-goniurus-reared-situ-kachemak-
bay-alaska;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1978/761/haynes.pdf.

Wahle:1978:BCC

[2627] Roy J. Wahle and Robert R. Vreeland. Bioeconomic contribution of Columbia River hatchery fall Chinook salmon, 1961 through 1964 broods, to the Pacific salmon fisheries. Fishery Bulletin, 76(1):??, 1978. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/bioeconomic-contribution-
columbia-river-hatchery-fall-chinook-salmon-1961-through-1964;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1978/761/wahle.pdf.

Bullard:1978:PCC

[2628] Fern A. Bullard and Jeff Collins. Physical and chemical changes of pink shrimp, Pandalus borealis, held in carbon dioxide modified refrigerated seawater compared with pink shrimp held in ice. Fishery Bulletin, 76(1): ??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/physical-and-
chemical-changes-pink-shrimp-pandalus-borealis-held-carbon-
dioxide-modified;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1978/761/bullard.pdf.

Henry:1978:ENF

[2629] Kenneth A. Henry. Estimating natural and fishing mortalities of Chinook salmon, Oncorhynchus tshawytscha, in the ocean, based on recoveries of marked fish. Fishery Bulletin, 76(1):??, 1978. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/estimating-natural-and-fishing-
mortalities-chinook-salmon-oncorhynchus-tshawytscha-ocean;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1978/
761/henry.pdf.

Rogers:1978:ASN

[2630] Carolyn A. Rogers, Douglas C. Biggs, and Richard A. Cooper. Aggregation of the siphonophore Nanomia cara in the Gulf of Maine: observations from a submersible. Fishery Bulletin, 76(1):??, 1978. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/aggregation-siphonophore-
nanomia-cara-gulf-maine-observations-submersible;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
rogers.pdf.

Calkins:1978:FBM

[2631] Donald G. Calkins. Feeding behavior and major prey species of the sea otter, Enhydra lutris, in Montague Strait, Prince William Sound, Alaska. Fishery Bulletin, 76(1):??, 1978. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/feeding-behavior-and-major-prey-
species-sea-otter-enhydra-lutris-montague-strait-prince;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1978/761/
calkins.pdf.

Hobson:1978:TRA

[2632] Edmund S. Hobson and James R. Chess. Trophic relationships among fishes and plankton in the lagoon at Enewetak Atoll, Marshall Islands. Fishery Bulletin, 76(1):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/trophic-relationships-among-fishes-and-plankton-lagoon-

enewetak-atoll-marshall-islands;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1978/761/hobson.pdf.

Dahlberg:1978:CPA

[2633] Michael L. Dahlberg. Computer program for analysis of the homogeneity and goodness of fit of frequency distributions, FORTRAN IV. Fishery Bulletin, 76(1):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/computer-program-analysis-homogeneity-and-goodness-
fit-frequency-distributions-fortran-iv;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/dahlberg.
pdf.

Colin:1978:DSW

[2634] Patrick L. Colin. Daily and summer-winter variation in mass spawning of the striped parrotfish, Scarus croicensis. Fishery Bulletin, 76 (1):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/daily-and-
summer-winter-variation-mass-spawning-striped-parrotfish-scarus-
croicensis;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1978/761/colin.pdf.

Kinner:1978:PAD

[2635] Peter Kinner and Don Maurer. Polychaetous annelids of the Delaware Bay Region. Fishery Bulletin, 76(1):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/polychaetous-annelids-delaware-bay-region;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1978/761/kinner.pdf.

Gilmore:1978:PTD

[2636] R. Grant Gilmore, John K. Holt, Robert S. Jones, George R. Kulczycki, Louis G. MacDowell III, and Wayne C. Magley. Portable tripod drop net for estuarine fish studies. Fishery Bulletin, 76(1): ??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/portable-
tripod-drop-net-estuarine-fish-studies;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/gilmore.
pdf.

Collette:1978:SBN

[2637] Bruce B. Collette, Joseph L. Russo, and Luis Alberto Zavala-Camin. Scomberomorus brasiliensis, a new species of Spanish mackerel from the Western Atlantic. Fishery Bulletin, 76(1):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/scomberomorus-
brasiliensis-new-species-spanish-mackerel-western-atlantic;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1978/761/collette.pdf.

Love:1978:FHT

[2638] Milton S. Love and Alfred W. Ebeling. Food and habitat of three switchfeeding fishes in the kelp forests off Santa Barbara, California. Fishery Bulletin, 76(1):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
food-and-habitat-three-switch-feeding-fishes-kelp-forests-santa-
barbara-california;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1978/761/love.pdf.

Wellington:1978:SFJ

[2639] G. M. Wellington and Shane Anderson. Surface feeding by a juvenile gray whale, Eschrichtius robustus. Fishery Bulletin, 76(1): ??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/surface-
feeding-juvenile-gray-whale-eschrichtius-robustus;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
wellington.pdf.

Fletcher:1978:TDS

[2640] R. Ian Fletcher. Time-dependent solutions and efficient parameters for stock-production models. Fishery Bulletin, 76(2):??, 1978. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/time-dependent-solutions-
and-efficient-parameters-stock-production-models;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1978/762/
fletcher.pdf.

Rice:1978:CSP

[2641] D. W. Rice, Jr. and F. L. Harrison. Copper sensitivity of Pacific herring, Clupea harengus Pallasi, during its early life history. Fishery Bulletin, 76(2):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
copper-sensitivity-pacific-herring-clupea-harengus-pallasi-
during-its-early-life-history;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1978/762/rice.pdf.

Graham:1978:TBG

[2642] Joseph J. Graham and Edwin P. Creaser, Jr. Tychoplanktonic bloodworm, Glycera dibranchiata, in Sullivan Harbor, Maine. Fishery Bulletin, 76(2):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/tychoplanktonic-bloodworm-glycera-dibranchiata-
sullivan-harbor-maine;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/graham.pdf.

Fusaro:1978:GRS

[2643] Craig Fusaro. Growth rate of the sand crab, Emerita analoga, (Hippidae) in two different environments. Fishery Bulletin, 76(2):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/growth-rate-
sand-crab-emerita-analoga-hippidae-two-different-environments;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1978/762/fusaro.pdf.

Squire:1978:NAS

[2644] James L. Squire, Jr. Northern anchovy school shapes as related to problems in school size estimation. Fishery Bulletin, 76(2):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/northern-anchovy-school-shapes-
related-problems-school-size-estimation;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1978/762/squire.pdf.

Houde:1978:SFP

[2645] Edward D. Houde and Richard C. Schekter. Simulated food patches and survival of larval bay anchovy, Anchoa mitchilli, and sea bream, Archosargus rhomboidalis. Fishery Bulletin, 76(2):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic).  URL https://spo.nmfs.noaa.gov/content/simulated-
food-patches-and-survival-larval-bay-anchovy-anchoa-mitchilli-
and-sea-bream;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/houde%20(1).pdf.

Haynes:1978:DLHb

[2646] Evan Haynes. Description of larvae of a hippolytid shrimp, Lebbeus groenlandicus, reared in situ in Kachemak Bay, Alaska. Fishery Bulletin, 76 (2):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/description-
larvae-hippolytid-shrimp-lebbeus-groenlandicus-reared-situ-

kachemak-bay-alaska;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1978/762/haynes.pdf.

Theilacker:1978:ESH

[2647] Gail H. Theilacker. Effect of starvation on the histological and morphological characteristics of jack mackerel, Trachurus symmetricus, larvae. Fishery Bulletin, 76(2):??, 1978. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/effect-starvation-histological-
and-morphological-characteristics-jack-mackerel-trachurus;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1978/
762/theilacker.pdf.

Symons:1978:DJP

[2648] Philip E. K. Symons and James D. Martin. Discovery of juvenile Pacific salmon (Coho) in a small coastal stream of New Brunswick. Fishery Bulletin, 76(2):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
discovery-juvenile-pacific-salmon-coho-small-coastal-stream-
new-brunswick;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/symons.pdf.

Hurley:1978:SSS

[2649] Ann C. Hurley. School structure of the squid Loligo opalescens. Fishery Bulletin, 76(2):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/school-structure-squid-loligo-opalescens;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1978/762/
hurley.pdf.

Major:1978:AEI

[2650] Peter F. Major. Aspects of estuarine intertidal ecology of juvenile striped mullet, Mugil cephalus, in Hawaii. Fishery Bulletin, 76(2): ??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/aspects-
estuarine-intertidal-ecology-ofjuvenile-striped-mullet-mugil-
cephalus-hawaii;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/major%20(2).pdf.

Herke:1978:SEN

[2651] William H. Herke. Subsampler for estimating the number and length frequency of small preserved nektonic organisms. Fishery Bulletin, 76(2):

??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/subsampler-
estimating-number-and-length-frequency-small-preserved-nektonic-
organisms;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/herke.pdf.

Krouse:1978:EEV

[2652] Jay S. Krouse. Effectiveness of escape vent shape in traps for catching legal-sized lobster, Homarus americanus, and harvestablesized crabs, Cancer borealis and Cancer irroratus. Fishery Bulletin, 76(2):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
effectiveness-escape-vent-shape-traps-catching-legal-sized-
lobster-homarus-americanus-and;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1978/762/krouse.pdf.

Ikeda:1978:EZP

[2653] Tsutomu Ikeda and Sigeru Motoda. Estimated zooplankton production and their ammonia excretion in the Kuroshio and adjacent seas. Fishery Bulletin, 76(2):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/estimated-zooplankton-production-and-their-ammonia-
excretion-kuroshio-and-adjacent-seas;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/ikeda.pdf.

Berggren:1978:RCH

[2654] Thomas J. Berggren and Joel T. Lieberman. Relative contribution of Hudson, Chesapeake, and Roanoke striped bass, Morone saxatilis, stocks to the Atlantic coast fishery. Fishery Bulletin, 76(2): ??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/relative-
contribution-hudson-chesapeake-and-roanoke-striped-bass-morone-
saxatilis-stocks;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1978/762/berggren.pdf.

Moreira:1978:SEE

[2655] G. Soares Moreira and W. B. Vernberg. Synergistic effects of environmental variables on the metabolism of the copepod Euterpina acutifrons from two different areas off the coast of the State of São Paulo, Brazil. Fishery Bulletin, 76(2):??, 1978. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/synergistic-effects-
environmental-variables-metabolism-copepod-euterpina-acutifrons-

two;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1978/762/moreira.pdf.

Austin:1978:PAS

[2656] Herbert M. Austin and Clarence R. Hickey, Jr. Predicting abundance of striped bass, Morone saxatilis, in New York waters from modal lengths. Fishery Bulletin, 76(2):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/predicting-abundance-striped-bass-morone-saxatilis-
new-york-waters-modal-lengths;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/austin%20(1).pdf.

Christensen:1978:TRJ

[2657] M. S. Christensen. Trophic relationships in juveniles of three species of sparid fishes in the South African marine littoral. Fishery Bulletin, 76 (2):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/trophic-
relationships-juveniles-three-species-sparid-fishes-south-african-
marine-littoral;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1978/762/christensen.pdf.

Nicholson:1978:AGM

[2658] William R. Nicholson and William E. Schaaf. Aging of Gulf menhaden, Brevoortia patronus. Fishery Bulletin, 76(2):??, 1978. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/aging-gulf-menhaden-brevoortia-
patronus;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1978/762/nicholson.pdf.

Katona:1978:OWS

[2659] Steven K. Katona, Salvatore A. Testaverde, and Bradley Barr. Observations on a white-sided dolphin, Lagenorhynchus acutus, probably killed in gill nets in the Gulf of Maine. Fishery Bulletin, 76(2):475–476, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/observations-white-sided-
dolphin-lagenorhynchus-acutus-probably-killed-gill-nets-gulf-
maine;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/katona.pdf.

Cook:1978:ISB

[2660] Rodney C. Cook and Gary E. Lord. Identification of stocks of Bristol Bay sockeye salmon, Oncorhynchus nerka, by evaluating scale patterns with a polynomial discriminant method. Fishery Bulletin, 76(2):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/identification-stocks-bristol-bay-sockeye-
salmon-oncorhynchus-nerka-evaluating-scale;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1978/762/cook.pdf.

Ortner:1978:VZB

[2661] Peter B. Ortner, Peter H. Wiebe, Loren Haury, and Steven Boyd. Variability in zooplankton biomass distribution in the Northern Sargasso Sea: the contribution of Gulf Stream cold core rings. Fishery Bulletin, 76(2): ??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/variability-
zooplankton-biomass-distribution-northern-sargasso-sea-contribution-
gulf-stream;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1978/762/ortner.pdf.

Moffatt:1978:RHB

[2662] Nancy M. Moffatt and Donald A. Thomson. Reciprocal hybridization between the California and Gulf of California grunions, Leuresthes tenuis and Leuresthes sardina (Atherinidae). Fishery Bulletin, 76(2):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/reciprocal-hybridization-
between-california-and-gulf-california-grunions-leuresthes-
tenuis;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/moffatt.pdf.

Pearcy:1978:DAS

[2663] William G. Pearcy. Distribution and abundance of small flatfishes and other demersal fishes in a region of diverse sediments and bathymetry off Oregon. Fishery Bulletin, 76(3):??, 1978. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/distribution-and-abundance-small-
flatfishes-and-other-demersal-fishes-region-diverse;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1978/763/
pearcy.pdf.

Bone:1978:RDF

[2664] Quentin Bone, Joe Kiceniuk, and David R. Jones. On the role of the different fibre types in fish myotomes at intermediate swimming speeds. Fishery Bulletin, 76(3):691–699, 1978. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/role-different-fibre-types-fish-

myotomes-intermediate-swimming-speeds;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/bone.pdf.

Pearcy:1978:FHD

[2665] William G. Pearcy and Danil Hancock. Feeding habits of Dover sole, Microstomus pacificus; rex sole, Glyptocephalus zachirus; slender sole, Lyopsetta exilis; and Pacific sanddab, Citharichthys sordidus, in a region of diverse sediments and bathymetry off Oregon. Fishery Bulletin, 76(3):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/feeding-habits-dover-
sole-microstomus-pacificus-rex-sole-glyptocephalus-zachirus-
slender;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1978/763/pearcy1.pdf.

Jewett:1978:SFP

[2666] Stephen C. Jewett. Summer food of the Pacific cod, Gadus macrocephalus, near Kodiak Island, Alaska. Fishery Bulletin, 76(3):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/summer-food-pacific-
cod-gadus-macrocephalus-near-kodiak-island-alaska;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
jewett%20(1).pdf.

Quinn:1978:HTS

[2667] William H. Quinn, David O. Zopf, Kent S. Short, and Richard T. W. Kuo Yang. Historical trends and statistics of the Southern Oscillation, El Niño, and Indonesian droughts. Fishery Bulletin, 76(3):663–678, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/historical-
trends-and-statistics-southern-oscillation-el-nino-and-indonesian-
droughts;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1978/763/quinn.pdf.

Leming:1978:CSS

[2668] Thomas D. Leming and Hillman J. Holley. A computer software system for optimizing survey cruise tracks. Fishery Bulletin, 76(3): ??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/computer-
software-system-optimizing-survey-cruise-tracks;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
leming.pdf.

Barkley:1978:STK

[2669] Richard A. Barkley, William H. Neill, and Reginald M. Gooding. Skipjack tuna, Katsuwonus pelamis, habitat based on temperature and oxygen requirements. Fishery Bulletin, 76(3):??, 1978. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/skipjack-tuna-katsuwonus-pelamis-
habitat-based-temperature-and-oxygen-requirements;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1978/763/
barkley.pdf.

Rivard:1978:PEP

[2670] D. Rivard and L. J. Bledsoe. Parameter estimation for the Pella–Tomlinson stock production model under nonequilibrium conditions. Fishery Bulletin, 76(3):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/parameter-estimation-pella-tomlinson-
stock-production-model-under-nonequilibrium-conditions;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1978/763/
rivard.pdf.

Clarke:1978:DFP

[2671] Thomas A. Clarke. Diel feeding patterns of 16 species of mesopelagic fishes from Hawaiian waters. Fishery Bulletin, 76(3):??, 1978. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/diel-feeding-patterns-16-
species-mesopelagic-fishes-hawallan-waters;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1978/763/clarke.pdf.

Venrick:1978:SSP

[2672] E. L. Venrick. Systematic sampling in a planktonic ecosystem. Fishery Bulletin, 76(3):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/systematic-sampling-planktonic-ecosystem;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1978/763/
venrick.pdf.

Fiedler:1978:PST

[2673] Paul C. Fiedler. The precision of simulated transect surveys of northern anchovy, Engraulis mordax, school groups. Fishery Bulletin, 76(3):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/precision-simulated-
transect-surveys-northern-anchovy-engraulis-mordax-school-groups;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1978/763/fiedler.pdf.

Young:1978:VDP

[2674] Richard Edward Young. Vertical distribution and photosensitive vesicles of pelagic cephalopods from Hawaiian waters. Fishery Bulletin, 76 (3):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/vertical-
distribution-and-photosensitive-vesicles-pelagic-cephalopods-
hawaiian-waters;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1978/763/young.pdf.

Fletcher:1978:RPT

[2675] R. Ian Fletcher. On the restructuring of the Pella–Tomlinson system. Fishery Bulletin, 76(3):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/restructuring-pella-tomlinson-system;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1978/763/fletcher.pdf.

Leis:1978:SZP

[2676] Jeffrey M. Leis. Systematics and zoogeography of the porcupinefishes (Diodon, Diodontidae, Tetraodontiformes), with comments on egg and larval development. Fishery Bulletin, 76(3):535–567, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/systematics-
and-zoogeography-porcupinefishes-diodon-diodontidae-tetraodontiformes-
comments;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/leis.pdf.

Heard:1978:PCS

[2677] William R. Heard. Probable case of streambed overseeding — 1967 pink salmon, Oncorhynchus gorbuscha, spawners and survival of their progeny in Sashin Creek, Southeastern Alaska. Fishery Bulletin, 76(3): ??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/probable-case-
streambed-overseeding-1967-pink-salmon-oncorhynchus-gorbuscha-
spawners-and;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1978/763/heard.pdf.

Farfante:1978:IAS

[2678] Isabel Perez Farfante. Intersex anomalies in shrimp of the genus Penaeopsis (Crustacea: Penaeidae). Fishery Bulletin, 76(3):??, 1978.

CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/intersex-anomalies-
shrimp-genus-penaeopsis-crustacea-penaeidae;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/perez%
20(3).pdf.

Smith:1978:ASM

[2679] T. D. Smith and T. Polacheck. Analysis of a simple model for estimating historical population sizes. Fishery Bulletin, 76(4):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/analysis-
simple-model-estimating-historical-population-sizes;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
smith%20(1).pdf.

Perschbacher:1978:RRC

[2680] Peter W. Perschbacher and Frank J. Schwartz. Recent records of Callinectes danae and Callinectes marginatus (Decapoda: Portunidae) from North Carolina with environmental notes. Fishery Bulletin, 76 (4):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/recent-
records-callinectes-danae-and-callinectes-marginatus-decapoda-
portunidae-north;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/perschbacher.pdf.

Wade:1978:RSB

[2681] Lawrence S. Wade and Gary L. Friedrichsen. Recent sightings of the blue whale, Balenoptera musculus, in the northeastern tropical Pacific. Fishery Bulletin, 76(4):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
recent-sightings-blue-whale-balenoptera-musculus-northeastern-
tropical-pacific;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/wade.pdf.

Lenarz:1978:TES

[2682] William H. Lenarz and James R. Zweifel. A theoretical examination of some aspects of the interaction between longline and surface fisheries for yellowfin tuna, Thunnus albacares. Fishery Bulletin, 76(4):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/theoretical-
examination-some-aspects-interaction-between-longline-and-surface-
fisheries;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/lenarz%20(3).pdf.

Stout:1978:ACH

[2683] Virginia F. Stout and F. Lee Beezhold. Analysis of chlorinated hydrocarbon pollutants: a simplified extraction and cleanup procedure for fishery products. Fishery Bulletin, 76(4):??, 1978. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/analysis-chlorinated-hydrocarbon-
pollutants-simplified-extraction-and-cleanup-procedure;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
stout.pdf.

Baglin:1978:SCL

[2684] Raymond E. Baglin, Jr. Sex composition, length–weight relationship, and reproduction of the white marlin, Tetrapturus albidus, in the western North Atlantic Ocean. Fishery Bulletin, 76 (4):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/sex-
composition-length-weight-relationship-and-reproduction-ofthe-
white-marlin-tetrapturus;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/baglin.pdf.

Parrack:1978:ABS

[2685] Michael L. Parrack. Aspects of brown shrimp, Penaeus aztecus, growth in the northern Gulf of Mexico. Fishery Bulletin, 76(4):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/aspects-brown-shrimp-
penaeus-aztecus-growth-northern-gulf-mexico;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/parrack.
pdf.

Rensel:1978:GJS

[2686] John E. Rensel and Earl F. Prentice. Growth of juvenile spot prawn, Pandalus platyceros, in the laboratory and in net pens using different diets. Fishery Bulletin, 76(4):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/growth-juvenile-spot-prawn-pandalus-platyceros-
laboratory-and-net-pens-using-different;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/rensel%20(
1).pdf.

Becker:1978:RPL

[2687] C. Dale Becker and Dennis D. Dauble. Records of piscivorus leeches (Hirudinea) from the central Columbia River, Washington State. Fishery Bulletin, 76(4):??, 1978. CODEN FSYBAY.

ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/records-piscivorus-leeches-hirudinea-
central-columbia-river-washington-state;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/becker_0.
pdf.

Gore:1978:LDG

[2688] Robert H. Gore. Larval development of Galathea rostrata under laboratory conditions, with a discussion of larval development in the Galatheidae (Crustacea Anomura). Fishery Bulletin, 76(4):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/larval-development-
galathea-rostrata-under-laboratory-conditions-discussion-larval;
 https://spo.nmfs.noaa.gov/sites/default/files/gore.pdf.

Ehrlich:1978:TBRa

[2689] Karl F. Ehrlich, J. Myron Hood, Gerald Muszynski, and Gerald E. McGowen. Thermal behavioral responses of selected California littoral fishes. Fishery Bulletin, 76(4):??, 1978. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/thermal-behavioral-
responses-selected-california-littoral-fishes;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/ehrlich.
pdf.

Laurence:1978:LLW

[2690] Geoffrey C. Laurence. Larval length–weight relations for seven species of northwest Atlantic fishes reared in the laboratory. Fishery Bulletin, 76 (4):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/larval-length-
weight-relations-seven-species-northwest-atlantic-fishes-reared-
laboratory;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/laurence%20(1).pdf.

Smigielski:1978:ISL

[2691] Alphonse S. Smigielski. Induced spawning and larval rearing of the yellowtail flounder, Limanda ferruginea. Fishery Bulletin, 76(4):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/induced-spawning-and-
larval-rearing-yellowtail-flounder-limanda-ferruginea;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
smigielski.pdf.

Frost:1978:CBF

[2692] Bruce W. Frost and Lawrence E. McCrone. Changes in biomass of finfishes and squids from the Gulf of Maine to Cape Hatteras, 1963–74, as determined from research vessel survey data. Fishery Bulletin, 76(4):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/changes-biomass-finfishes-and-squids-gulf-
maine-cape-hatteras-1963-74-determined-research;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1978/764/
frost.pdf.

Brown:1978:LPS

[2693] B. E. Brown, J. A. Brennan, and J. E. Palmer. Linear programming simulations of the effects of bycatch on the management of mixed species fisheries off the northeastern coast of the United States. Fishery Bulletin, 76(4):851–860, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/linear-programming-simulations-effects-bycatch-
management-mixed-species-fisheries;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/brown_1.pdf.

Craddock:1978:ETI

[2694] Donovan R. Craddock. Effect of thermal increases of short duration on survival of Euphausia pacifica. Fishery Bulletin, 76(4):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/effect-thermal-increases-short-
duration-survival-euphausia-pacifica;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/craddock.pdf.

Young:1978:TMC

[2695] David R. Young and Tsu-Kai Jan. Trace metal contamination of the rock scallop, Hinnites giganteus, near a large southern California municipal outfall. Fishery Bulletin, 76(4):??, 1978. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/trace-metal-contamination-rock-
scallop-hinnites-giganteus-near-large-southern-california;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/young_0.pdf.

Gore:1978:RBF

[2696] Robert H. Gore. Reproductive biology of the female deep-sea red crab, Geryon quinquedens, from the Chesapeake Bight. Fishery Bulletin, 76 (4):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/reproductive-
biology-female-deep-sea-red-crab-geryon-quinquedens-chesapeake-
bight;   https://spo.nmfs.noaa.gov/sites/default/files/gore_
0.pdf.

Roberts:1978:ESE

[2697] John L. Roberts and Jeffrey B. Graham. Effect of swimmingspeed on the excess temperatures and activities of heart and red and white muscles in the mackerel, Scomber japonicus. Fishery Bulletin, 76(4):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/effect-swimmingspeed-
excess-temperatures-and-activities-heart-and-red-and-white-
muscles;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/roberts.pdf.

May:1978:LST

[2698] Robert C. May, Gerald S. Akiyama, and Michael T. Santerre. Lunar spawning of the threadfin, Polydactylus sexfilis, in Hawaii. Fishery Bulletin, 76(4):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/lunar-
spawning-threadfin-polydactylus-sexfilis-hawaii;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
may%20(1).pdf.

Sindermann:1978:PAD

[2699] Carl J. Sindermann. Pollution-associated diseases and abnormalties of fish and shellfish: a review. Fishery Bulletin, 76(4):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/pollution-associated-
diseases-and-abnormalties-fish-and-shellfish-review;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
sindermann%20(1).pdf.

Ehrlich:1978:TBRb

[2700] Karl F. Ehrlich, John S. Stephens, Gerald Muszynski, and J. Myron Hood. Thermal behavioral responses of the speckled sanddab, Citharichthys stigmaeus: laboratory and field investigations. Fishery Bulletin, 76(4):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
thermal-behavioral-responses-speckled-sanddab-citharichthys-
stigmaeus-laboratory-and-field;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/ehrlich1.pdf.

Peebles:1978:RPR

[2701] John B. Peebles. The roles of prior residence and relative size in competition for shelter by the Malaysian prawn, Macrobrachium rosenbergii. Fishery Bulletin, 76(4):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/roles-prior-residence-and-relative-size-competition-
shelter-malaysian-prawn-macrobrachium;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/peebles.pdf.

Frost:1978:VDD

[2702] Bruce W. Frost and Lawrence E. McCrone. Vertical distribution, diel vertical migration, and abundance of some mesopelagic fishes in the eastern subarctic Pacific Ocean in summer. Fishery Bulletin, 76 (4):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/vertical-
distribution-diel-vertical-migration-and-abundance-some-mesopelagic-
fishes-0;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/frost.pdf.

Shultz:1978:MSB

[2703] Cynthia D. Shultz and Bernard M. Ito. Mercury and selenium in blue marlin, Makaira nigricans, from the Hawaiian Islands. Fishery Bulletin, 76(4):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/mercury-and-selenium-blue-marlin-makaira-nigricans-
hawaiian-islands;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/schultz.pdf.

Colton:1978:PSA

[2704] John B. Colton, Jr., Wallace G. Smith, Arthur W. Kendall, Jr., Peter L. Berrien, and Michael P. Fahay. Principal spawning areas and times of marine fishes, Cape Sable to Cape Hatteras. Fishery Bulletin, 76(4):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/principal-spawning-areas-and-times-marine-
fishes-cape-sable-cape-hatteras;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/colton.pdf.

Greenblatt:1979:ATF

[2705] Paul R. Greenblatt. Associations of tuna with flotsam in the Eastern Tropical Pacific. Fishery Bulletin, 77(1):??, 1979. CODEN FSYBAY.

ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/associations-tuna-flotsam-eastern-tropical-
pacific;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1979/771/greenblatt.pdf.

Richardson:1979:DOL

[2706] Sally L. Richardson and Wayne A. Laroche. Development and occurrence of larvae and juveniles of the rockfishes Sebastes crameri, Sebastes pinniger, and Sebastes helvomaculatus (family Scorpaenidae) off Oregon. Fishery Bulletin, 77(1):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/development-and-occurrence-larvae-and-juveniles-rockfishes-
sebastes-crameri-sebastes;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1979/771/richardson.pdf.

Wenzloff:1979:SHM

[2707] D. R. Wenzloff, R. A. Greig, A. S. Merrill, and J. W. Ropes. A survey of heavy metals in the surf clam, Spisula solidissima, and the ocean quahog, Arctica islandica, of the Mid-Atlantic coast of the United States. Fishery Bulletin, 77(1):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/survey-heavy-metals-surf-
clam-spisula-solidissima-and-ocean-quahog-arctica-islandica-
mid;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/wenzloff.pdf.

Lander:1979:RLO

[2708] Robert H. Lander. Role of land and ocean mortality in yield of male Alaskan fur seal, Callorhinus ursinus. Fishery Bulletin, 77(1):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/role-land-and-ocean-
mortality-yield-male-alaskan-fur-seal-callorhinus-ursinus;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/lander.pdf.

Haynes:1979:DLN

[2709] Evan Haynes. Description of larvae of the northern shrimp, Pandalus borealis, reared in situ in Kachemak Bay, Alaska. Fishery Bulletin, 77(1):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/description-larvae-
northern-shrimp-pandalus-borealis-reared-situ-kachemak-bay-
alaska;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1979/771/haynes.pdf.

Sumida:1979:EDS

[2710] B. Y. Sumida, E. H. Ahlstrom, and H. G. Moser. Early development of seven flatfishes of the Eastern North Pacific with heavily pigmented larvae (Pisces, Pleuronectiformes). Fishery Bulletin, 77(1):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/early-
development-seven-flatfishes-eastern-north-pacific-heavily-
pigmented-larvae-pisces;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1979/771/sumida.pdf.

Overholtz:1979:AFF

[2711] William J. Overholtz and John R. Nicolas. Apparent feeding by the fin whale, Balaenoptera physalus, and the humpback whale, Megaptera novaengliae, on the American sand lance, Ammodytes americanus, in the northwest Atlantic. Fishery Bulletin, 77(1): ??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/apparent-
feeding-fin-whale-balaenoptera-physalus-and-humpback-whale-
megaptera-novaengliae;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/overholtz.pdf.

Hughes:1979:BWP

[2712] Steven E. Hughes and George Hirschhorn. Biology of walleye pollock, Theraga chalcogramma, in the Western Gulf of Alaska, 1973–75. Fishery Bulletin, 77(1):263–274, 1979. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/biology-walleye-pollock-theraga-
chalcogramma-western-gulf-alaska-1973-75;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1979/771/hughes.pdf.

Tricas:1979:RBS

[2713] Timothy C. Tricas. Relationships of the blue shark, Prionace glauca, and its prey species near Santa Catalina Island, California. Fishery Bulletin, 77(1):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
relationships-blue-shark-prionace-glauca-and-its-prey-species-
near-santa-catalina-island;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1979/771/tricas.pdf.

Williams:1979:EIH

[2714] John G. Williams. Estimation of intertidal harvest of Dungeness crab, Cancer magister, on Puget Sound, Washington, beaches. Fishery Bulletin, 77(1):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
estimation-intertidal-harvest-dungeness-crab-cancer-magister-
puget-sound-washington-beaches;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/williams%20(1).pdf.

Kendall:1979:SDB

[2715] Arthur W. Kendall, Jr. and Lionel A. Walford. Sources and distribution of bluefish, Pomatomus saltatrix, larvae and juveniles off the East Coast of the United States. Fishery Bulletin, 77(1):??, 1979. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/sources-and-distribution-
bluefish-pomatomus-saltatrix-larvae-and-juveniles-east-coast;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1979/771/kendall.pdf.

Wahle:1979:CBH

[2716] Roy J. Wahle, Reino O. Koski, and Robert Z. Smith. Contribution of 1960–63 brood hatchery-reared sockeye salmon, Oncorhynchus nerka, to the Columbia River commercial fishery. Fishery Bulletin, 77(1):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/contribution-1960-63-brood-hatchery-reared-sockeye-
salmon-oncorhynchus-nerka-columbia-river;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1979/771/wahle.pdf.

Targett:1979:CBP

[2717] Timothy E. Targett. A contribution to the biology of the puffers Sphoeroides testudineus and sphoeroides spengleri from biscayne bay, Florida. Fishery Bulletin, 77(1):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/contribution-biology-puffers-sphoeroides-testudineus-
and-sphoeroides-spengleri-biscayne-bay;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/targett.
pdf.

Meyer:1979:RAB

[2718] Thomas L. Meyer, Richard A. Cooper, and Richard W. Langton. Relative abundance, behavior, and food habits of the American sand lance, Ammodytes americanus, from the Gulf of Maine. Fishery Bulletin, 77(1):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/relative-abundance-
behavior-and-food-habits-american-sand-lance-ammodytes-americanus-

gulf;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1979/771/meyer.pdf.

Leatherwood:1979:ASB

[2719] Stephen Leatherwood. Aerial survey of the bottlenosed dolphin, Tursiops truncatus, and the West Indian manatee, Trichechus manalus, in the Indian and Banana Rivers, Florida. Fishery Bulletin, 77(1):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/aerial-survey-
bottlenosed-dolphin-tursiops-truncatus-and-west-indian-manatee-
trichechus;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/leatherwood.pdf.

Hui:1979:CMC

[2720] Clifford A. Hui. Correlates of maturity in the common dolphin, Delphinus delphis. Fishery Bulletin, 77(1):??, 1979. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/correlates-maturity-
common-dolphin-delphinus-delphis;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/hui.pdf.

Morgan:1979:DPL

[2721] Steven G. Morgan and Anthony J. Provenzano, Jr. Development of pelagic larvae and postlarva of Squilla empusa (Crustacea, Stomatopoda), with an assessment of larval characters within the Squillidae. Fishery Bulletin, 77(1):61–90, 1979. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/development-pelagic-larvae-and-
postlarva-squilla-empusa-crustacea-stomatopoda-assessment;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1979/
771/morgan.pdf.

Cohen:1979:VFR

[2722] Daniel M. Cohen and Joseph L. Russo. Variation in the fourbeard rockling, Enchelyopus cimbrius, a North Atlantic gadid fish, with comments on the genera of rocklings. Fishery Bulletin, 77(1):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/variation-fourbeard-
rockling-enchelyopus-cimbrius-north-atlantic-gadid-fish-comments-
genera;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/cohen.pdf.

Allen:1979:LDG

[2723] Larry G. Allen. Larval development of Gobiesox rhessodon (Gobiesocidae) with notes on the larva of Rimicola muscarum. Fishery Bulletin, 77(1):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/larval-development-gobiesox-rhessodon-gobiesocidae-
notes-larva-rimicola-muscarum;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/allen.pdf.

Neves:1979:OMA

[2724] Richard J. Neves and Linda Depres. The oceanic migration of American shad, Alosa sapidissima, along the Atlantic Coast. Fishery Bulletin, 77 (1):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/oceanic-
migration-american-shad-alosa-sapidissima-along-atlantic-coast;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1979/771/neves.pdf.

Omori:1979:LHV

[2725] Makoto Omori and David Gluck. Life history and vertical migration of the pelagic shrimp Sergestes similis off the southern California coast. Fishery Bulletin, 77(1):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/life-history-and-vertical-migration-pelagic-shrimp-
sergestes-similis-southern-california;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/omori.pdf.

Bailey:1979:SSF

[2726] Kevin Bailey and Jean Dunn. Spring and summer foods of walleye pollock, Theragra chalcogramma, in the eastern Bering Sea. Fishery Bulletin, 77(1):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
spring-and-summer-foods-walleye-pollock-theragra-chalcogramma-
eastern-bering-sea.

Olla:1979:SDH

[2727] Bori L. Olla, Allen J. Bejda, and A. Dale Martin. Seasonal dispersal and habitat selection of cunner, Tautogolabrus adspersus, and young tautog, Tautoga onitis, in Fire Island Inlet, Long Island, New York. Fishery Bulletin, 77(1):??, 1979. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/seasonal-dispersal-and-habitat-
selection-cunner-tautogolabrus-adspersus-and-young-tautog;

https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1979/
771/olla.pdf.

Hobson:1979:ZEL

[2728] Edmund S. Hobson and James R. Chess. Zooplankters that emerge from the lagoon floor at night at Kure and Midway Atolls, Hawaii. Fishery Bulletin, 77(1):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/zooplankters-emerge-lagoon-floor-
night-kure-and-midway-atolls-hawaii;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/hobson.pdf.

Dietrich:1979:FAM

[2729] Charles S. Dietrich, Jr. Fecundity of the Atlantic menhaden, Brevoortia tyrannus. Fishery Bulletin, 77(1):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/fecundity-atlantic-menhaden-brevoortia-tyrannus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/dietrich.pdf.

Clark:1979:AFD

[2730] Colin W. Clark and Marc Mangel. Aggregation and fishery dynamics: a theoretical study of schooling and the purse seine tuna fisheries. Fishery Bulletin, 77(2):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/aggregation-and-fishery-dynamics-
theoretical-study-schooling-and-purse-seine-tuna-fisheries;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1979/772/clark.pdf.

Wenner:1979:SAB

[2731] Elizabeth Lewis Wenner. Some aspects of the biology of deepsea lobsters of the family polychelidae (Crustacea, Decapoda) from the Western North Atlantic. Fishery Bulletin, 77(2):435–444, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/some-aspects-biology-
deep-sea-lobsters-family-polychelidae-crustacea-decapoda-western-
north;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1979/772/wenner.pdf.

Winstead:1979:SMO

[2732] James T. Winstead. A simple method to obtain serum from small fish. Fishery Bulletin, 77(2):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/simple-method-obtain-serum-small-fish;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
winstead.pdf.

Methot:1979:GNA

[2733] Richard D. Methot, Jr. and David Kramer. Growth of northern anchovy, Engraulis mordax, larvae in the sea. Fishery Bulletin, 77(2): ??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/growth-
northern-anchovy-engraulis-mordax-larvae-sea;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1979/772/methot.
pdf.

Wursig:1979:BEB

[2734] Bernd Würsig and Melany Würsig. Behavior and ecology of the bottlenose dolphin, Tursiops truncatus, in the South Atlantic. Fishery Bulletin, 77(2):399–41, 1979. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
behavior-and-ecology-bottlenose-dolphin-tursiops-truncatus-
south-atlantic;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1979/772/wursig.pdf.

Irvine:1979:SPW

[2735] A. Blair Irvine, Michael D. Scott, Randall S. Wells, and James G. Mead. Stranding of the pilot whale, Globicephala macrorhynchus, in Florida and South Carolina. Fishery Bulletin, 77(2):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/stranding-pilot-whale-globicephala-macrorhynchus-
florida-and-south-carolina;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/irvine.pdf.

Pella:1979:ACS

[2736] Jerome J. Pella and Timothy L. Robertson. Assessment of composition of stock mixtures. Fishery Bulletin, 77(2):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/assessment-composition-stock-mixtures;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1979/772/pella.pdf.

Scotto:1979:LDCa

[2737] Liberta E. Scotto. Larval development of the Cuban stone crab, Menippe nodifrons (Brachyura, Xanthidae), under laboratory conditions with notes on the status of the family Menippidae. Fishery Bulletin, 77(2):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/larval-development-
cuban-stone-crab-menippe-nodifrons-brachyura-xanthidae-under-
laboratory;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/scotto.pdf.

Galt:1979:FRG

[2738] Charles P. Galt. First records of a giant pelagic tunicate, Bathochordaeus charon (Urochordata, Larvacea), from the eastern Pacific Ocean, with notes on its biology. Fishery Bulletin, 77(2):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/first-records-giant-pelagic-tunicate-
bathochordaeus-charon-urochordata-larvacea-eastern;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
galt.pdf.

Pratt:1979:RBS

[2739] Harold L. Pratt, Jr. Reproduction in the blue shark, Prionace glauca. Fishery Bulletin, 77(2):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/reproduction-blue-shark-prionace-glauca;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1979/772/
pratt.pdf.

Annala:1979:MEN

[2740] John H. Annala. Mortality estimates for the New Zealand rock lobster, Jasus edwardsii. Fishery Bulletin, 77(2):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/mortality-estimates-new-zealand-rock-
lobster-jasus-edwardsii;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/annala.pdf.

Reppond:1979:WPT

[2741] Kermit D. Reppond, Fern A. Bullard, and Jeff Collins. Walleye pollock, Theragra chalcogramma: Physical, chemical, and sensory changes when held in ice and in carbon dioxide modified refrigerated seawater. Fishery Bulletin, 77(2):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/walleye-pollock-theragra-chalcogramma-physical-
chemical-and-sensory-changes-when-held-ice;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1979/772/reppond.pdf.

Holliday:1979:TDD

[2742] D. V. Holliday and H. L. Larsen. Thickness and depth distributions of some epipelagic fish schools off southern California. Fishery Bulletin, 77 (2):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/thickness-and-
depth-distributions-some-epipelagic-fish-schools-southern-california;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/holliday.pdf.

Scotto:1979:LDCb

[2743] Lirerta E. Scotto. Larval development of the Cuban stone crab, Menippe nodifrons (Brachyura, Xanthidae), under laboratory conditions with notes on the status of the family Menippidae. Fishery Bulletin, 77(2):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/larval-development-
cuban-stone-crab-menlppe-nodifrons-brachyura-xanthidae-under-
laboratory;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1979/772/scotto.pdf.

Brill:1979:EBS

[2744] Richard W. Brill. The effect of body size on the standard metabolic rate of skipjack tuna, Katsuwonus pelamis. Fishery Bulletin, 77(2): ??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/effect-
body-size-standard-metabolic-rate-skipjack-tuna-katsuwonus-
pelamis;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/brill.pdf.

Siegel:1979:LPA

[2745] Robert A. Siegel, Joseph J. Mueller, and Brian J. Rothschild. A linear programming approach to determining harvesting capacity: a multiple species fishery. Fishery Bulletin, 77(2):??, 1979. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/linear-programming-
approach-determining-harvesting-capacity-multiple-species-fishery;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1979/772/siegel.pdf.

Hines:1979:ETD

[2746] Anson H. Hines. Effects of a thermal discharge on reproductive cycles in Mytilus edulis and Mytilus californianus (Mollusca, Bivalvia). Fishery Bulletin, 77(2):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
effects-thermal-discharge-reproductive-cycles-mytilus-edulis-
and-mytilus-californianus;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/hines.pdf.

Weinstein:1979:SMH

[2747] Michael P. Weinstein. Shallow marsh habitats as primary nurseries for fishes and shellfish, Cape Fear River, North Carolina. Fishery Bulletin, 77(2):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/shallow-
marsh-habitats-primary-nurseries-fishes-and-shellfish-cape-
fear-river-north;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1979/772/weinstein.pdf.

Sherburne:1979:IDP

[2748] Stuart W. Sherburne and Laurie L. Bean. Incidence and distribution of piscine erythrocytic necrosis and the microsporidian, Glugea hertwigi, in rainbow smelt, Osmerus mordax, from Massachusetts to the Canadian Maritimes. Fishery Bulletin, 77(2):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/incidence-and-
distribution-piscine-erythrocytic-necrosis-and-microsporidian-
glugea-hertwigi;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/sherburne%20(1).pdf.

Barham:1979:ACB

[2749] Eric G. Barham, Jay C. Sweeney, Stephen Leatherwood, Robert K. Beggs, and Cecilia L. Barham. Aerial census of the bottlenose dolphin, Tursiops truncatus, in a region of the Texas Coast. Fishery Bulletin, 77 (3):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/aerial-census-
bottlenose-dolphin-tursiops-truncatus-region-texas-coast;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1979/
773/barham.pdf.

Prezant:1979:AMP

[2750] Robert S. Prezant. An antipredation mechanism of the polychaete Phyllodoce mucosa with notes on similar mechanisms in other potential prey. Fishery Bulletin, 77(3):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/antipredation-mechanism-polychaete-phyllodoce-mucosa-
notes-similar-mechanisms-other;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1979/773/prezant.pdf.

Lancraft:1979:EPI

[2751] Thomas M. Lancraft and Bruce H. Robison. Evidence of postcapture ingestion by midwater fishes in trawl nets. Fishery Bulletin, 77 (3):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/evidence-
postcapture-ingestion-midwater-fishes-trawl-nets;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
lancraft.pdf.

Haynes:1979:LMP

[2752] Evan Haynes. Larval morphology of Pandalus tridens and a summary of the principal morphological characteristics of North Pacific pandalid shrimp larvae. Fishery Bulletin, 77(3):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/larval-morphology-pandalus-tridens-and-summary-
principal-morphological-characteristics;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1979/773/haynes.pdf.

Scarnecchia:1979:CWH

[2753] Dennis L. Scarnecchia and Harry H. Wagner. Contribution of wild and hatchery-reared Coho salmon, Oncorhynchus kisutch, to the Oregon ocean sport fishery. Fishery Bulletin, 77(3):??, 1979. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/contribution-wild-and-hatchery-
reared-coho-salmon-oncorhynchus-kisutch-oregon-ocean-sport;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1979/773/scarnecchia.pdf.

Newkirk:1979:IEA

[2754] G. F. Newkirk and D. L. Waugh. Inhibitory effect of the alga Pavlova lutherii on growth of mussel, Mytilus edulis, larvae. Fishery Bulletin, 77(3):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/inhibitory-effect-alga-pavlova-lutherii-growth-
mussel-mytilus-edulis-larvae;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/newkirk.pdf.

Healey:1979:UNR

[2755] M. C. Healey. Utilization of the Nanaimo River Estuary by juvenile Chinook salmon, Oncorhynchus tshawytscha. Fishery Bulletin, 77(3):653– 668, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/utilization-

nanaimo-river-estuary-juvenile-chinook-salmon-oncorhynchus-
tshawytscha;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1979/773/healey.pdf.

Weihs:1979:ESC

[2756] Daniel Weihs. Energetic significance of changes in swimming modes during growth of larval anchovy, Engraulis mordax. Fishery Bulletin, 77 (3):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/energetic-
significance-changes-swimming-modes-during-growth-larval-anchovy-
engraulis-mordax;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1979/773/weihs.pdf.

Hubbs:1979:RSPa

[2757] Carl L. Hubbs and Robert L. Wisner. Revision of the sauries (Pisces, Scomberesocidae) with descriptions of two new genera and one new species. Fishery Bulletin, 77(3):521–566, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/revision-savries-pisces-scomberesocidae-descriptions-
two-new-genera-and-one-new-species;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1979/773/hubbs.pdf.

Hubbs:1979:RSPb

[2758] Carl L. Hubbs and Robert L. Wisner. Revision of the sauries (Pisces, Scomberesocidae) with descriptions of two new genera and one new species. Fishery Bulletin, 77(3):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/revision-sauries-pisces-scomberesocidae-descriptions-
two-new-genera-and-one-new-species;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/galt_0.pdf.

Hunter:1979:SIB

[2759] J. Roe Hunter and Stephen R. Goldberg. Spawning incidence and batch fecundity in northern anchovy, Engraulis mordax. Fishery Bulletin, 77 (3):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/spawning-
incidence-and-batch-fecundity-northern-anchovy-engraulis-mordax;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1979/773/hunter.pdf.

Schlieder:1979:EDA

[2760] Rodric A. Schlieder. Effects of desiccation and autospasy on egg hatching success in stone crab, Menippe mercenaria. Fishery Bulletin, 77(3):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/effects-desiccation-and-autospasy-egg-hatching-
success-stone-crab-menippe-mercenaria;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/schlieder.pdf.

Johnson:1979:ETS

[2761] John Kenneth Johnson. Effects of temperature and salinity on production and hatching of dormant eggs of Acartia californiensis (Copepoda) in an Oregon Estuary. Fishery Bulletin, 77(3):??, 1979. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/effects-temperature-and-salinity-
production-and-hatching-dormant-eggs-acartia;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1979/773/johnson.
pdf.

Howe:1979:FRO

[2762] Kevin M. Howe, David L. Stein, and Carl E. Bond. First records off Oregon of the pelagic fishes Paralepis atlantica, Gonostoma atlanticum, and Aphanopus carbo, with notes on the anatomy of Aphanopus carbo. Fishery Bulletin, 77(3):700–703, 1979. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/first-records-oregon-
pelagic-fishes-paralepis-atlantica-gonostoma-atlanticum-and-
aphanopus;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/howe.pdf.

Judkins:1979:CAD

[2763] David C. Judkins, Creighton D. Wirick, and Wayne E. Esaias. Composition, abundance, and distribution of zooplankton in the New York Bight, September 1974–September 1975. Fishery Bulletin, 77(3):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/composition-
abundance-and-distribution-zooplankton-new-york-bight-september-
1974-september;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1979/773/judkins.pdf.

Buckley:1979:CRA

[2764] L. J. Buckley. Changes in ribonucleic acid, deoxyribonucleic acid, and protein content during ontogenesis in winter flounder, Pseudopleuronectes americanus, and effect of starvation. Fishery Bulletin, 77 (3):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/changes-
ribonucleic-acid-deoxyribonucleic-acid-and-protein-content-
during-ontogenesis;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/buckley.pdf.

Krouse:1979:DCC

[2765] Jay S. Krouse. Distribution and catch composition of Jonah crab, Cancer borealis, and rock crab, Cancer irroratus, near Boothbay Harbor, Maine. Fishery Bulletin, 77(3):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/distribution-and-catch-composition-jonah-crab-cancer-
borealis-and-rock-crab-cancer;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1979/773/krouse.pdf.

Scherer:1979:EEL

[2766] Michael D. Scherer and Donald W. Bourne. Eggs and early larvae of small mouth flounder, Etropus microstomus. Fishery Bulletin, 77(3):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/eggs-and-early-larvae-small-mouth-flounder-etropus-
microstomus;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/scherer.pdf.

Badcock:1979:RSD

[2767] Julian Badcock and Ronald C. Baird. Remarks on systematics, development, and distribution of the hatchetfish genus Sternoptyx (Pisces, Stomiatoidei). Fishery Bulletin, 77(4):803–820, 1979. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/remarks-systematics-development-
and-distribution-hatchetfish-genus-sternoptyx-pisces;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1979/774/
badcock.pdf.

Leatherwood:1979:DMR

[2768] Stephen Leatherwood, William F. Perrin, Vicky L. Kirby, Carl L. Hubbs, and Marilyn Dahlheim. Distribution and movements of Risso’s dolphin, Grampus griseus, in the Eastern North Pacific. Fishery Bulletin, 77(4): ??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/distribution-
and-movements-rissos-dolphin-grampus-griseus-eastern-north-
pacific;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1979/774/leatherwood.pdf.

Farfante:1979:RPS

[2769] Isabel Perez Farfante. Revision of the penaeid shrimp genus Penaeopsis (Crustacea: Decapoda). Fishery Bulletin, 77(4):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/revision-
penaeid-shrimp-genus-penaeopsis-crustacea-decapoda;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
perez_0.pdf.

Collins:1979:BRS

[2770] Jeff Collins, Kermit D. Reppond, and Fern A. Bullard. Black rockfish, Sebastes melanops: Changes in physical, chemical, and sensory properties when held in ice and in carbon dioxide modified refrigerated seawater. Fishery Bulletin, 77(4):??, 1979. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/black-rockfish-sebastes-melanops-
changes-physical-chemical-and-sensory-properties-when-held;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1979/774/collins.pdf.

Loeb:1979:VDD

[2771] Valerie J. Loeb. Vertical distribution and development of larval fishes in the North Pacific Central Gyre during summer. Fishery Bulletin, 77 (4):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/vertical-
distribution-and-development-larval-fishes-north-pacific-central-
gyre-during;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1979/774/loeb.pdf.

Caillouet:1979:TTD

[2772] Charles W. Caillouet, Frank J. Patella, and William B. Jackson. Trends toward decreasing size of brown shrimp, Penaeus aztecus, and white shrimp, Penaeus setiferus, in reported annual catches from Texas and Louisiana. Fishery Bulletin, 77(4):985–989, 1979. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/trends-toward-decreasing-
size-brown-shrimp-penaeus-aztecus-and-white-shrimp-penaeus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/caillouet.pdf.

Collins:1979:DLJ

[2773] L. Alan Collins, John H. Finucane, and Lyman E. Barger. Description of larval and juvenile red snapper, Lutjanus campechanus. Fishery Bulletin, 77(4):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
description-larval-and-juvenile-red-snapper-lutjanus-campechanus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1979/774/collins1.pdf.

Love:1979:IOR

[2774] Milton S. Love. Isolation of olive rockfish, Sebastes serranoides, populations off Southern California. Fishery Bulletin, 77(4):??, 1979. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/isolation-olive-rockfish-
sebastes-serranoides-populations-southern-california;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1979/774/
love.pdf.

Haynes:1979:SZC

[2775] Evan Haynes. Stage I zoeae of a crangonid shrimp, Crangon franciscorum angustimana, hatched from ovigerous females collected in Kachemak Bay, Alaska. Fishery Bulletin, 77(4):??, 1979. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/stage-i-zoeae-crangonid-shrimp-
crangon-franciscorum-angustimana-hatched-ovigerous-females;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/haynes_0.pdf.

Ferraro:1979:EDA

[2776] Steven P. Ferraro. Embryonic development of Atlantic menhaden, Brevoortia tyrannus, and a fish embryo age estimation method. Fishery Bulletin, 77(4):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/embryonic-development-atlantic-menhaden-brevoortia-
tyrannus-and-fish-embryo-age-estimation;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1979/774/ferraro.pdf.

Norris:1979:BHS

[2777] Kenneth S. Norris and Thomas P. Dohl. Behavior of the Hawaiian spinner dolphin, Stenella longirostris. Fishery Bulletin, 77(4): ??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/behavior-
hawaiian-spinner-dolphin-stenella-longirostris;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1979/774/norris.
pdf.

Baglin:1979:LWR

[2778] Raymond E. Baglin, Jr. Length–weight relationships of western Atlantic bluefin tuna, Thunnus thynnus. Fishery Bulletin, 77(4): ??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/length-
weight-relationships-western-atlantic-bluefin-tuna-thunnus-
thynnus;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/baglin_0.pdf.

Fogarty:1979:ETV

[2779] Michael J. Fogarty and David V. D. Borden. Effects of trap venting on gear selectivity in the inshore Rhode Island American lobster, Homarus americanus, fishery. Fishery Bulletin, 77(4):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/effects-trap-venting-
gear-selectivity-inshore-rhode-island-american-lobster-homarus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1979/774/fogarty.pdf.

Pieper:1979:AMM

[2780] R. E. Pieper and B. G. Bargo. Acoustic measurements of a migrating layer of the Mexican lampfish, Triphoturus mexicanus, at 102 kilohertz. Fishery Bulletin, 77(4):??, 1979. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/acoustic-measurements-migrating-
layer-mexican-lampfish-triphoturus-mexicanus-102-kilohertz;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1979/774/pieper.pdf.

Bulak:1979:DAI

[2781] James S. Bulak and Roy C. Heidinger. Developmental anatomy and inflation of the gas bladder in striped bass, Morone saxatilis. Fishery Bulletin, 77(4):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
developmental-anatomy-and-inflation-gas-bladder-striped-bass-
morone-saxatilis;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/bulak.pdf.

Kimura:1979:LMB

[2782] Daniel K. Kimura. Likelihood methods for the von Bertalanffy growth curve. Fishery Bulletin, 77(4):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/likelihood-methods-von-bertalanffy-growth-
curve;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1979/774/kimura.pdf.

Williams:1979:GSN

[2783] John G. Williams. Growth and survival in newly settled spat of the Manila clam, Tapes japonica. Fishery Bulletin, 77(4):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/growth-
and-survival-newly-settled-spat-manila-clam-tapes-japonica;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1979/774/williams.pdf.

Grabe:1979:FAA

[2784] Stephen A. Grabe. Food of age 1 and 2 Atlantic tomcod, Microgadus tomcod, from Haverstraw Bay, Hudson River, New York. Fishery Bulletin, 77(4):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/food-age-1-and-2-atlantic-tomcod-microgadus-tomcod-
haverstraw-bay-hudson-river-new-york;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/grabe.pdf.

Lang:1979:LDH

[2785] William H. Lang and Alan M. Young. Larval development of Hypoconcha sabulosa (Decapoda: Dromiidae). Fishery Bulletin, 77(4):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/larval-development-
hypoconcha-sabulosa-decapoda-dromiidae;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1979/774/lang.pdf.

Winter:1979:RDS

[2786] Gary W. Winter, Carl B. Schreck, and John D. McIntyre. Resistance of different stocks and transferrin genotypes of Coho salmon, Oncorhynchus kisutch, and steelhead trout, Salmo gairdneri, to bacterial kidney disease and vibriosis. Fishery Bulletin, 77(4):795–802, 1979. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/resistance-different-stocks-and-
transferrin-genotypes-coho-salmon-oncorhynchus-kisutch-and;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1979/774/winter.pdf.

Anonymous:1979:I

[2787] Anonymous. Index. Fishery Bulletin, 77(4):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/index-106;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/index_
6.pdf.

Laroche:1979:DOL

[2788] Wayne A. Laroche and Sally L. Richardson. Development and occurrence of larvae and juveniles of the rockfishes Sebastes flavidus and Sebastes melanops (Scorpaenidae) off Oregon. Fishery Bulletin, 77(4):??, 1979. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
development-and-occurrence-larvae-and-juveniles-rockfishes-
sebastes-flavidus-and-sebastes;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1979/774/laroche.pdf.

Wursig:1979:BED

[2789] Bernd Würsig and Melany Würsig. Behavior and ecology of the dusky dolphin, Lagenorhynchus obscurus, in the South Atlantic. Fishery Bulletin, 77(4):871–890, 1979. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
behavior-and-ecology-dusky-dolphin-lagenorhynchus-obscurus-
south-atlantic;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1979/774/wursig.pdf.

Hunter:1980:ELH

[2790] John R. Hunter and Carol A. Kimbrell. Early life history of Pacific mackerel, Scomber japonicus. Fishery Bulletin, 78(1):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic).  URL https://spo.nmfs.noaa.gov/content/early-
life-history-pacific-mackerel-scomber-japonicus;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1980/781/
hunter.pdf.

Hines:1980:OSO

[2791] Anson H. Hines and Thomas R. Loughlin. Observations of sea otters digging for clams at Monterey Harbor, California. Fishery Bulletin, 78(1):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/observations-sea-otters-digging-clams-monterey-
harbor-california;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/hines_0.pdf.

Morse:1980:MSF

[2792] Wallace W. Morse. Maturity, spawning, and fecundity of Atlantic croaker, Micropogonias undulatus, occurring north of Cape Hatteras, North Carolina. Fishery Bulletin, 78(1):??, 1980. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/maturity-spawning-and-fecundity-
atlantic-croaker-micropogonias-undulatus-occurring-north;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/morse1.pdf.

Mendelssohn:1980:UMD

[2793] Roy Mendelssohn. Using Markov decision models and related techniques for purposes other than simple optimization: Analyzing the consequences of policy alternatives on the management of salmon runs. Fishery Bulletin, 78(1):??, 1980. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/using-markov-decision-models-and-
related-techniques-purposes-other-simple-optimization;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1980/781/
mendelssohn.pdf.

Weis:1980:EZF

[2794] Peddrick Weis and Judith Shulman Weis. Effect of zinc on fin regeneration in the mummichog, Fundulus heteroclitus, and its interaction with methylmercury. Fishery Bulletin, 78(1):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/effect-zinc-
fin-regeneration-mummichog-fundulus-heteroclitus-and-its-interaction;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/weis_0.pdf.

Miller:1980:CSD

[2795] Robert E. Miller, Douglas W. Campbell, and Pamela J. Lunsford. Comparison of sampling devices for the juvenile blue crab, Callinectes sapidus. Fishery Bulletin, 78(1):196–198, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/comparison-sampling-devices-juvenile-
blue-crab-callinectes-sapidus;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/miller_1.pdf.

Morse:1980:SFA

[2796] Wallace W. Morse. Spawning and fecundity of Atlantic mackerel, Scomber scombrus, in the Middle Atlantic Bight. Fishery Bulletin, 78(1):

??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/spawning-and-
fecundity-atlantic-mackerel-scomber-scombrus-middle-atlantic-
bight;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1980/781/morse.pdf.

Testaverde:1980:SDA

[2797] Salvatore A. Testaverde and James G. Mead. Southern distribution of the Atlantic whitesided dolphin, Lagenorhynchus acutus, in the western North Atlantic. Fishery Bulletin, 78(1):??, 1980. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/southern-distribution-atlantic-
whitesided-dolphin-lagenorhynchus-acutus-western-north;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
weis_1.pdf.

Peterson:1980:OEL

[2798] R. H. Peterson, P. H. Johansen, and J. L. Metcalfe. Observations on early life stages of Atlantic tomcod, Microgadus tomcod. Fishery Bulletin, 78(1):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/observations-early-life-stages-atlantic-tomcod-
microgadus-tomcod;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1980/781/peterson.pdf.

Matarese:1980:ARS

[2799] Ann C. Matarese and David L. Stein. Additional records of the sculpin psychrolutes phrictus in the eastern Bering Sea and off Oregon. Fishery Bulletin, 78(1):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
additional-records-sculpin-psychrolutes-phrictus-eastern-bering-
sea-and-oregon;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/matarese.pdf.

Pietsch:1980:SDC

[2800] Theodore W. Pietsch and John P. Van Duzer. Systematics and distribution of ceratioid anglerfishes of the family Melanocetidae with the description of a new species from the Eastern North Pacific Ocean. Fishery Bulletin, 78(1):59–87, 1980. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
systematics-and-distribution-ceratioid-anglerfishes-family-
melanocetidae-description-new;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1980/781/pietsch.pdf.

Odell:1980:RMS

[2801] Daniel K. Odell, Edward D. Asper, Joe Baucom, and Lanny H. Cornell. A recurrent mass stranding of the false killer whale, Pseudorca crassidens, in Florida. Fishery Bulletin, 78(1):??, 1980. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/recurrent-mass-stranding-
false-killer-whale-pseudorca-crassidens-florida;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
odell.pdf.

Adams:1980:LHP

[2802] Peter B. Adams. Life history patterns in marine fishes and their consequences for fisheries management. Fishery Bulletin,

78(1):??, 1980.  CODEN FSYBAY.  ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
life-history-patterns-marine-fishes-and-their-consequences-
fisheries-management;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1980/781/adams.pdf.

Powles:1980:DLS

[2803] Howard Powles. Descriptions of larval silver perch, Bairdiella chrysoura, banded drum, Larimus fasciatus, and star drum, Stellifer lanceolatus (Sciaenidae). Fishery Bulletin, 78(1):??, 1980. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/descriptions-larval-silver-perch-
bairdiella-chrysoura-banded-drum-larimus-fasciatus-and;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1980/781/
powles.pdf.

Branstetter:1980:OFS

[2804] Steven Branstetter and Robert L. Shipp. Occurrence of the finetooth shark, Carcharhinus isodon, off Dauphin Island, Alabama. Fishery Bulletin, 78(1):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/occurrence-finetooth-shark-carcharhinus-isodon-
dauphin-island-alabama;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/branstetter.pdf.

Ambler:1980:SMC

[2805] Julie W. Ambler. Species of Munidopsis (Crustacea, Galatheidae) occurring off Oregon and in adjacent waters. Fishery Bulletin, 78(1):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/species-munidopsis-
crustacea-galatheidae-occurring-oregon-and-adjacent-waters;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1980/781/ambler.pdf.

Stout:1980:ORF

[2806] Virginia F. Stout. Organochlorine residues in fishes from the Northwest Atlantic Ocean and Gulf of Mexico. Fishery Bulletin, 78(1):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
organochlorine-residues-fishes-northwest-atlantic-ocean-and-
gulf-mexico;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1980/781/stout.pdf.

Baglin:1980:SRP

[2807] Raymond E. Baglin, Jr., Mark I. Farber, William H. Lenarz, and John M. Mason, Jr. Shedding rates of plastic and metal dart tags from Atlantic bluefin tuna, Thunnus thynnus. Fishery Bulletin, 78(1):179– 185, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/shedding-
rates-plastic-and-metal-dart-tags-atlantic-bluefin-tuna-thunnus-
thynnus;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/baglin%20(1).pdf.

Grimes:1980:RBV

[2808] Churchill B. Grimes and Gene R. Huntsman. Reproductive biology of the vermilion snapper, Rhomboplites aurorubens, from North Carolina and South Carolina. Fishery Bulletin, 78(1):??, 1980. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/reproductive-biology-vermilion-
snapper-rhomboplites-aurorubens-north-carolina-and-south;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1980/
781/grimes.pdf.

Weihs:1980:RDC

[2809] Daniel Weihs. Respiration and depth control as possible reasons for swimming of northern anchovy, Engraulis mordax, yolk-sac larvae. Fishery Bulletin, 78(1):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/respiration-and-depth-control-possible-reasons-
swimming-northern-anchovy-engraulis-mordax;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1980/781/weihs.pdf.

Haynes:1980:ILG

[2810] James M. Haynes and Robert H. Gray. Influence of little goose dam on upstream movements of adult Chinook salmon, Oncorhynchus tshawytscha. Fishery Bulletin, 78(1):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/influence-little-goose-dam-upstream-movements-
adult-chinook-salmon-oncorhynchus-tshawytscha.

Benirschke:1980:ODS

[2811] K. Benirschke, Mary L. Johnson, and Rolf J. Benirschke. Is ovulation in dolphins, Stenella longirostris and Stenella attenuata, always copulation-induced? Fishery Bulletin, 78(2):507–528, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/ovulation-dolphins-
stenella-longirostris-and-stenella-attenuata-always-copulation-
induced;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1980/782/benirschke.pdf.

Pietsch:1980:CAPa

[2812] Theodore W. Pietsch and Jeffrey A. Seigel. Ceratioid anglerfishes of the Philippine Archipelago, with descriptions of five new species. Fishery Bulletin, 78(2):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
ceratioid-anglerfishes-philippine-archipelago-descriptions-
five-new-species;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1980/782/pietsch.pdf.

Knight:1980:LDE

[2813] Margaret D. Knight. Larval development of Euphausia eximia (Crustacea: Euphausiacea) with notes on its vertical distribution and morphological divergence between populations. Fishery Bulletin, 78(2):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/larval-development-
euphausia-eximia-crustacea-euphausiacea-notes-its-vertical-
distribution;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/knight.pdf.

Kraeuter:1980:ELP

[2814] John N. Kraeuter and Michael Castagna. Effects of large predators on the field culture of the hard clam, Mercenaria mercenaria. Fishery Bulletin, 78(2):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
effects-large-predators-field-culture-hard-clam-mercenaria-
mercenaria;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/kraeuter.pdf.

Bullard:1980:IMA

[2815] Fern A. Bullard and Jeff Collins. An improved method to analyze trimethylamine in fish and the interference of ammonia and dimethylamine. Fishery Bulletin, 78(2):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/improved-method-analyze-trimethylamine-fish-and-
interference-ammonia-and-dimethylamine;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1980/782/bullard.pdf.

Randall:1980:SCE

[2816] John E. Randall. A survey of ciguatera at Enewetak and Bikini, Marshall Islands, with notes on the systematics and food habits of ciguatoxic fishes. Fishery Bulletin, 78(2):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/survey-ciguatera-enewetak-and-bikini-
marshall-islands-notes-systematics-and-food-habits;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1980/782/
randall.pdf.

Stoner:1980:FELa

[2817] Allan W. Stoner. Feeding ecology of Lagodon rhomboides (Pisces: Sparidae): variation and functional responses. Fishery Bulletin, 78 (2):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/feeding-
ecology-lagodon-rhomboides-pisces-sparidae-variation-and-functional-
responses-0;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/stoner_1.pdf.

Parker:1980:DME

[2818] Keith Parker. A direct method for estimating northern anchovy, Engraulis mordax, spawning biomass. Fishery Bulletin, 78(2):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/direct-method-estimating-
northern-anchovy-engraulis-mordax-spawning-biomass;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
parker.pdf.

Ebel:1980:TCS

[2819] Wesley J. Ebel. Transportation of Chinook salmon, Oncorhynchus tshawytscha, and steelhead, Salmo gairdneri, smolts in the Columbia River and effects on adult returns. Fishery Bulletin, 78(2):??, 1980. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/transportation-chinook-salmon-
oncorhynchus-tshawytscha-and-steelhead-salmo-gairdneri-smolts;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1980/782/ebel.pdf.

Richardson:1980:ELBa

[2820] Sally L. Richardson, Jean R. Dunn, and Nancy Anne Naplin. Eggs and larvae of butter sole, Isopsetta isolepis (Pleuronectidae), off Oregon and Washington. Fishery Bulletin, 78(2):401– 417, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/eggs-and-
larvae-butter-sole-isopsetta-isolepis-pleuronectidae-oregon-
and-washington;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1980/782/richardson.pdf.

Mead:1980:OMSa

[2821] James G. Mead, Daniel K. Odell, Randall S. Wells, and Michael D. Scott. Observations on a mass stranding of spinner dolphin, Stenella longirostris, from the west coast of Florida. Fishery Bulletin, 78(2): ??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/observations-
mass-stranding-spinner-dolphin-stenella-longirostris-west-coast-
florida-0;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/mead.pdf.

Pitcher:1980:FHS

[2822] Kenneth W. Pitcher. Food of the harbor seal, Phoca vitulina richardsi, in the Gulf of Alaska. Fishery Bulletin, 78(2):??, 1980. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/food-harbor-seal-phoca-vitulina-
richardsi-gulf-alaska;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/pitcher.pdf.

Ebeling:1980:AVRa

[2823] Alfred W. Ebeling, Ralph J. Larson, William S. Alevizon, and Richard N. Bray. Annual variability of reef-fish assemblages in kelp forests off Santa Barbara, California. Fishery Bulletin, 78(2):??, 1980. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/annual-variability-
reef-fish-assemblages-kelp-forests-santa-barbara-california;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1980/782/ebeling.pdf.

Smyth:1980:CDP

[2824] Peter O. Smyth. Callinectes (Decapoda: Portunidae) larvae in the Middle Atlantic Bight, 1975–77. Fishery Bulletin, 78(2):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/callinectes-decapoda-
portunidae-larvae-middle-atlantic-bight-1975-77;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1980/782/
smyth.pdf.

Ebeling:1980:AVRb

[2825] Alfred W. Ebeling, Ralph J. Larson, William S. Alevizon, and Richard N. Bray. Annual variability of reef-fish assemblages in kelp forests off Santa Barbara, California. Fishery Bulletin, 78(2):??, 1980. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/annual-variability-
reef-fish-assemblages-kelp-forests-santa-barbara-california-
0;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/ebeling.pdf.

Johnson:1980:PGS

[2826] James H. Johnson. Production and growth of subyearling Coho salmon, Oncorhynchus kisutch, Chinook salmon, Oncorhynchus tshawytscha, and steelhead, Salmo gairdneri, in Orwell Brook, tributary of Salmon River, New York. Fishery Bulletin, 78(2):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/production-and-growth-subyearling-coho-salmon-
oncorhynchus-kisutch-chinook-salmon;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/johnson_1.pdf.

Ferraro:1980:DTS

[2827] Steven P. Ferraro. Daily time of spawning of 12 fishes in the Peconic Bays, New York. Fishery Bulletin, 78(2):??, 1980. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/daily-time-spawning-
12-fishes-peconic-bays-new-york;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1980/782/ferraro.pdf.

Stoner:1980:FELb

[2828] Allan W. Stoner. Feeding ecology of Lagodon rhomboides (Pisces: Sparidae): Variation and functional responses. Fishery Bulletin, 78 (2):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/feeding-
ecology-lagodon-rhomboides-pisces-sparidae-variation-and-functional-
responses;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1980/782/stoner.pdf.

Pietsch:1980:CAPb

[2829] Theodore W. Pietsch and Jeffrey A. Seigel. Ceratioid anglerfishes of the Philippine Archipelago, with descriptions of five new species. Fishery Bulletin, 78(2):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
ceratioid-anglerfishes-philippine-archipelago-descriptions-
five-new-species-0;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/pietsch.pdf.

Mead:1980:OMSb

[2830] James G. Mead, Daniel K. Odell, Randall S. Wells, and Michael D. Scott. Observations on a mass stranding of spinner dolphin, Stenella longirostris, from the West Coast of Florida. Fishery Bulletin, 78(2): ??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/observations-
mass-stranding-spinner-dolphin-stenella-longirostris-west-coast-
florida;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1980/782/mead.pdf.

Weinstein:1980:RTT

[2831] Michael P. Weinstein, Sidney L. Weiss, Ronald G. Hodson, and Lawrence R. Gerry. Retention of three taxa of postlarval fishes in an intensively flushed tidal estuary, Cape Fear River, North Carolina. Fishery Bulletin, 78(2):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
retention-three-taxa-postlarval-fishes-intensively-flushed-
tidal-estuary-cape-fear-river;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1980/782/weinstein.pdf.

Richardson:1980:ELBb

[2832] Sally L. Richardson, Jean R. Dunn, and Nancy Anne Naplin. Eggs and larvae of butter sole, Isopsetta isolepis (Pleuronectidae), off Oregon and Washington. Fishery Bulletin, 78(2):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/eggs-and-
larvae-butter-sole-isopsetta-isolepis-pleuronectidae-oregon-
and-washington-0;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/richardson_0.pdf.

OConnell:1980:PSN

[2833] Charles P. O’Connell. Percentage of starving northern anchovy, Engraulis mordax, larvae in the sea as estimated by histological methods. Fishery Bulletin, 78(2):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/percentage-starving-northern-anchovy-engraulis-
mordax-larvae-sea-estimated-histological;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1980/782/oconnell.
pdf.

Huppert:1980:AUS

[2834] D. D. Huppert. An analysis of the United States demand for fish meal. Fishery Bulletin, 78(2):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/analysis-united-states-demand-fish-meal;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
huppert.pdf.

Pearcy:1980:LOC

[2835] William G. Pearcy. A large, opening-closing midwater trawl for sampling oceanic nekton, and comparison of catches with an Isaacs–Kidd midwater trawl. Fishery Bulletin, 78(2):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/large-opening-closing-midwater-trawl-
sampling-oceanic-nekton-and-comparison-catches-isaacs;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
pearcy_2.pdf.

Oliver:1980:RBW

[2836] John S. Oliver, Peter N. Slattery, Larry W. Hulberg, and James W. Nybakken. Relationships between wave disturbance and zonation of benthic invertebrate communities along a subtidal high-energy beach in Monterey Bay, California. Fishery Bulletin, 78(2):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/relationships-
between-wave-disturbance-and-zonation-benthic-invertebrate-

communities-along;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1980/782/oliver.pdf.

Potthoff:1980:DSF

[2837] Thomas Potthoff. Development and structure of fins and fin supports in dolphin fishes Coryphaena hippurus and Coryphaena equiselis (Coryphaenidae). Fishery Bulletin, 78(2):??, 1980. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/development-and-structure-fins-
and-fin-supports-dolphin-fishes-coryphaena-hippurus-and;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
potthoff_0.pdf.

Coe:1980:PBS

[2838] James M. Coe and Warren E. Stuntz. Passive behavior by the spotted dolphin, Stenella attenuata, in tuna purse seine nets. Fishery Bulletin, 78(2):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/passive-behavior-spotted-dolphin-stenella-
attenuata-tuna-purse-seine-nets;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/coe.pdf.

Bailey:1980:ESD

[2839] Jack E. Bailey, Stanley D. Rice, Jerome J. Pella, and Sidney G. Taylor. Effects of seeding density of pink salmon, Oncorhynchus gorbuscha, eggs on water chemistry, fry characteristics, and fry survival in gravel incubators. Fishery Bulletin, 78(3):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/effects-seeding-density-pink-salmon-oncorhynchus-
gorbuscha-eggs-water-chemistry-fry;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1980/783/bailey.pdf.

Laroche:1980:RNA

[2840] Joanne Lyczkowski Laroche and Sally L. Richardson. Reproduction of northern anchovy, Engraulis mordax, off Oregon and Washington. Fishery Bulletin, 78(3):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/reproduction-northern-anchovy-engraulis-mordax-
oregon-and-washington;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1980/783/laroche.pdf.

Fogarty:1980:MTA

[2841] Michael J. Fogarty, David V. D. Borden, and Howard J. Russell. Movements of tagged American lobster, Homarus americanus, off Rhode Island. Fishery Bulletin, 78(3):771–780, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/movements-tagged-american-lobster-homarus-
americanus-rhode-island;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fogarty.pdf.

Hunter:1980:ECN

[2842] John R. Hunter and Carol A. Kimbrell. Egg cannibalism in the northern anchovy, Engraulis mordax. Fishery Bulletin, 78(3):??, 1980. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/egg-cannibalism-northern-
anchovy-engraulis-mordax;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/hunter_0.pdf.

Bishop:1980:OCH

[2843] James M. Bishop, James G. Gosselink, and James H. Stone. Oxygen consumption and hemolymph osmolality of brown shrimp, Penaeus aztecus. Fishery Bulletin, 78(3):??, 1980. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/oxygen-consumption-and-hemolymph-
osmolality-brown-shrimp-penaeus-aztecus;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1980/783/bishop.pdf.

Lenarz:1980:SSC

[2844] William H. Lenarz and Peter B. Adams. Some statistical considerations of the design of trawl surveys for rockfish (Scorpaenidae). Fishery Bulletin, 78(3):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/some-
statistical-considerations-design-trawl-surveys-rockfish-scorpaenidae;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1980/783/lenarz.pdf.

Rensel:1980:FCG

[2845] John E. Rensel and Earl F. Prentice. Factors controlling growth and survival of cultured spot prawn, Pandalus platyceras, in Puget Sound, Washington. Fishery Bulletin, 78(3):??, 1980. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/factors-controlling-growth-and-
survival-cultured-spot-prawn-pandalus-platyceras-puget-sound;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/rensel_0.pdf.

Quinn:1980:DDS

[2846] Thomas P. Quinn, Bruce S. Miller, and R. Craig Wingert. Depth distribution and seasonal and diel movements of ratfish, Hydrolagus colliei, in Puget Sound, Washington. Fishery Bulletin, 78(3):816–821, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/depth-distribution-
and-seasonal-and-diel-movements-ratfish-hydrolagus-colliei-
puget-sound;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/quinn.pdf.

Cheng:1980:DAH

[2847] Lanna Cheng and Eric Shulenberger. Distribution and abundance of Halobates species (Insecta: Heteroptera) in the Eastern Tropical Pacific. Fishery Bulletin, 78(3):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/distribution-and-abundance-halobates-
species-insecta-heteroptera-eastern-tropical-pacific;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1980/783/
cheng.pdf.

Powell:1980:ELD

[2848] Allyn B. Powell and Herbert R. Gordy. Egg and larval development of the spot, Leiostomus xanthurus (Sciaenidae). Fishery Bulletin, 78 (3):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/egg-and-
larval-development-spot-leiostomus-xanthurus-sciaenidae;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1980/783/
powell.pdf.

Theilacker:1980:RCS

[2849] Gail H. Theilacker. Rearing container size affects morphology and nutritional condition of larval jack mackerel, Trachurus symmetricus. Fishery Bulletin, 78(3):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/rearing-container-size-affects-morphology-
and-nutritional-condition-larval-jack-mackerel;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/theilacker1.
pdf.

Pearson:1980:DPH

[2850] Walter H. Pearson, Peter C. Sugarman, Dana L. Woodruff, J. W. Blaylock, and Bori L. Olla. Detection of petroleum hydrocarbons by the Dungeness crab, Cancer magister. Fishery Bulletin, 78(3):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/detection-petroleum-
hydrocarbons-dungeness-crab-cancer-magister;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/pearson.
pdf.

Clarke:1980:DFS

[2851] Thomas A. Clarke. Diets of fourteen species of vertically migrating mesopelagic fishes in Hawaiian waters. Fishery Bulletin, 78 (3):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/diets-
fourteen-species-vertically-migrating-mesopelagic-fishes-hawaiian-
waters;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1980/783/clarke.pdf.

Rice:1980:SSE

[2852] Stanley D. Rice and Jack E. Bailey. Survival, size, and emergence of pink salmon, Oncorhynchus gorbuscha, alevins after short- and long-term exposures to ammonia. Fishery Bulletin, 78(3):??, 1980. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/survival-size-and-emergence-pink-
salmon-oncorhynchus-gorbuscha-alevins-after-short-and-long;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1980/783/rice.pdf.

Carney:1980:EMW

[2853] Robert S. Carney and Andrew G. Carey, Jr. Effectiveness of metering wheels for measurement of area sampled by beam trawls. Fishery Bulletin, 78(3):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/effectiveness-metering-wheels-measurement-area-
sampled-beam-trawls;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/carney.pdf.

Cressey:1980:BCP

[2854] Roger Cressey and Hillary Boyle Cressey. Bomolochid copepods parasitic on the eyes of Indo-West Pacific clupeid fishes. Fishery Bulletin, 78(3):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/bomolochid-copepods-
parasitic-eyes-indo-west-pacific-clupeid-fishes;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1980/783/
cressey.pdf.

Shane:1980:OMD

[2855] Susan H. Shane. Occurrence, movements, and distribution of bottlenose dolphin, Tursiops truncatus, in Southern Texas. Fishery Bulletin, 78(3): ??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/occurrence-
movements-and-distribution-bottlenose-dolphin-tursiops-truncatus-
southern-texas;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1980/783/shane.pdf.

Pitcher:1980:SCF

[2856] Kenneth W. Pitcher. Stomach contents and feces as indicators of harbor seal, Phoca vitulina, foods in the Gulf of Alaska. Fishery Bulletin, 78 (3):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/stomach-
contents-and-feces-indicators-harbor-seal-phoca-vitulina-foods-
gulf-alaska;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/pitcher%20(1).pdf.

Hankin:1980:MRP

[2857] David G. Hankin. A multistage recruitment process in laboratory fish populations: Implications for models of fish population dynamics. Fishery Bulletin, 78(3):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
multistage-recruitment-process-laboratory-fish-populations-
implications-models-fish;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1980/783/hankin.pdf.

Theilacker:1980:CBM

[2858] Gail H. Theilacker. Changes in body measurements of larval northern anchovy, Engraulis mordax, and other fishes due to handling and preservation. Fishery Bulletin, 78(3):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/changes-body-measurements-larval-northern-
anchovy-engraulis-mordax-and-other-fishes-due;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1980/783/theilacker.
pdf.

Christensen:1980:SRC

[2859] Darryl J. Christensen and Walter J. Clifford. The 1978 spring recreational catch of Atlantic mackerel, Scomber scombrus, off the Middle Atlantic region. Fishery Bulletin, 78(3):??, 1980. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/1978-spring-recreational-catch-
atlantic-mackerel-scomber-scombrus-middle-atlantic-region;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/christensen.pdf.

Horn:1980:DSV

[2860] Michael H. Horn. Diel and seasonal variation in abundance and diversity of shallow-water fish populations in Morro Bay, California. Fishery Bulletin, 78(3):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/diel-and-seasonal-variation-abundance-and-
diversity-shallow-water-fish-populations-morro;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1980/783/horn.
pdf.

Rice:1980:ECE

[2861] D. W. Rice, Jr., F. L. Harrison, and A. Jearld, Jr. Effects of copper on early life history stages of northern anchovy, Engraulis mordax. Fishery Bulletin, 78(3):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
effects-copper-early-life-history-stages-northern-anchovy-engraulis-
mordax;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/rice1.pdf.

Beardsley:1980:SPO

[2862] Grant L. Beardsley. Size and possible origin of sailfish, Istiophorus platypterus, from the eastern Atlantic Ocean. Fishery Bulletin, 78(3): 805–808, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/size-and-
possible-origin-sailfish-istiophorus-platypterus-eastern-atlantic-
ocean;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/beardsley.pdf.

Howell:1980:TEG

[2863] W. Huntting Howell. Temperature effects on growth and yolk utilization in yellowtail flounder, Limanda ferruginea, yolk-sac larvae. Fishery Bulletin, 78(3):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
temperature-effects-growth-and-yolk-utilization-yellowtail-
flounder-limanda-ferruginea-yolk;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1980/783/howell.pdf.

Morgan:1980:ALE

[2864] Steven G. Morgan. Aspects of larval ecology of Squilla empusa (Crustacea, Stomatopoda) in Chesapeake Bay. Fishery Bulletin, 78(3): ??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/aspects-
larval-ecology-squilla-empusa-crustacea-stomatopoda-chesapeake-
bay;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/morgan.pdf.

Rice:1980:ACP

[2865] Stanley D. Rice and Jack E. Bailey. Ammonia concentrations in pink salmon, Oncorhynchus gorbuscha, redds of sashin creek, southeastern Alaska. Fishery Bulletin, 78(3):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/ammonia-concentrations-pink-salmon-oncorhynchus-
gorbuscha-redds-sashin-creek-southeastern;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/rice2.pdf.

Mendelssohn:1980:UBJ

[2866] Roy Mendelssohn. Using Box–Jenkins models to forecast fishery dynamics: Identification, estimation, and checking. Fishery Bulletin, 78(4):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/using-box-jenkins-models-forecast-fishery-
dynamics-identification-estimation-and-checking;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1980/784/
mendelssohn.pdf.

McClellan:1980:RME

[2867] G. L. McClellan and J. K. Leong. A radiologic method for examination of the gastrointestinal tract in the Atlantic ridley, Lepidochelys kempi, and loggerhead, Caretta caretta, marine turtles. Fishery Bulletin, 78(4):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/radiologic-method-examination-gastrointestinal-
tract-atlantic-ridley-lepidochelys-kempi-and;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/mclellan.
pdf.

Modde:1980:SFO

[2868] Timothy Modde and Stephen T. Ross. Seasonality of fishes occupying a surf zone habitat in the Northern Gulf of Mexico. Fishery Bulletin, 78(4):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/seasonality-fishes-occupying-surf-zone-habitat-
northern-gulf-mexico;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1980/784/modde.pdf.

Clausen:1980:SFP

[2869] David M. Clausen. Summer food of Pacific cod, Gadus macrocephalus, in coastal waters of southeastern Alaska. Fishery Bulletin, 78(4):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/summer-food-pacific-
cod-gadus-macrocephalus-coastal-waters-southeastern-alaska;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/clausen.pdf.

Richardson:1980:SBE

[2870] Sally L. Richardson. Spawning biomass and early life of northern anchovy, Engraulis mordax, in the northern subpopulation off Oregon and Washington. Fishery Bulletin, 78(4):??, 1980. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/spawning-biomass-and-early-life-
northern-anchovy-engraulis-mordax-northern-subpopulation;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1980/
784/richardson.pdf.

Khilnani:1980:UGY

[2871] Arvind Khilnani. Use of Griffin’s yield model for the Gulf of Mexico shrimp fishery. Fishery Bulletin, 78(4):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/use-griffins-yield-model-gulf-mexico-
shrimp-fishery;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/khilnani.pdf.

Bray:1980:IWC

[2872] Richard N. Bray. Influence of water currents and zooplankton densities on daily foraging movements of blacksmith, Chromis punctipinnis, a planktivorous reef fish. Fishery Bulletin, 78(4):??, 1980. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/influence-water-currents-
and-zooplankton-densities-daily-foraging-movements-blacksmith;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1980/784/bray.pdf.

Ainley:1980:PSP

[2873] David G. Ainley, Craig S. Strong, Harriet R. Huber, T. James Lwis, and Stephen H. Morrell. Predation by sharks on pinnipeds at the Farallon Islands. Fishery Bulletin, 78(4):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/predation-sharks-pinnipeds-farallon-islands;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
ainley.pdf.

Goldberg:1980:SSC

[2874] Stephen R. Goldberg. Seasonal spawning cycle of the Pacific butterfish, Peprilus simillimus (Stromateidae). Fishery Bulletin, 78(4): ??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/seasonal-
spawning-cycle-pacific-butterfish-peprilus-simillimus-stromateidae;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/goldberg_0.pdf.

Breiwick:1980:EIP

[2875] Jeffrey M. Breiwick, Edward D. Mitchell, and Douglas G. Chapman. Estimated initial population size of the Bering Sea stock of bowhead whale, Balaena mysticetus: an iterative method. Fishery Bulletin, 78(4):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
estimated-initial-population-size-bering-sea-stock-bowhead-
whale-balaena-mysticetus;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1980/784/breiwick.pdf.

Griswold:1980:SOR

[2876] Carolyn A. Griswold and Jerome Prezioso. In situ observations on reproductive behavior of the long-finned squid, Loligo pealei. Fishery Bulletin, 78(4):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/situ-
observations-reproductive-behavior-long-finned-squid-loligo-
pealei;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/griswold.pdf.

Davis:1980:EIS

[2877] Gary E. Davis. Effects of injuries on spiny lobster, Panulirus argus, and implications for fishery management. Fishery Bulletin, 78 (4):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/effects-
injuries-spiny-lobster-panulirus-argus-and-implications-fishery-
management;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/davis_4.pdf.

Durbin:1980:VSS

[2878] Ann G. Durbin, Edward G. Durbin, Peter G. Verity, and Thomas J. Smayda. Voluntary swimming speeds and respiration rates of a filterfeeding planktivore, the Atlantic menhaden, Brevoortia tyrannus (Pisces: Clupeidae). Fishery Bulletin, 78(4):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/voluntary-swimming-speeds-and-respiration-
rates-filter-feeding-planktivore-atlantic;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1980/784/durbin.pdf.

Lewis:1980:SSM

[2879] Robert M. Lewis and Charles M. Roithmayr. Spawning and sexual maturity of Gulf menhaden, Brevoortia patronus. Fishery Bulletin, 78(4): ??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/spawning-and-
sexual-maturity-gulf-menhaden-brevoortia-patronus;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
lewis_0.pdf.

Anonymous:1980:I

[2880] Anonymous. Index. Fishery Bulletin, 78(4):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/index-107;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/index%
20(1)_0.pdf.

Laroche:1980:DLS

[2881] Wayne A. Laroche. Development of larval smooth flounder, Liopsetta putnami, with a redescription of development of winter flounder, Pseudopleuronectes americanus (family Pleuronectidae). Fishery Bulletin, 78 (4):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/development-
larval-smooth-flounder-liopsetta-putnami-redescription-development-

winter;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1980/784/laroche.pdf.

Stroud:1980:FPW

[2882] Richard K. Stroud, Clifford H. Fiscus, and Hiroshi Kajimura. Food of the Pacific white-sided dolphin, Lagenorhynchus obliquidens, Dall’s porpoise, Phocoenoides dalli, and northern fur seal, Callorhinus ursinus, off California and Washington. Fishery Bulletin, 78(4):951–959, 1980. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/food-pacific-white-sided-
dolphin-lagenorhynchus-obliquidens-dalls-porpoise-phocoenoides;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/stroud.pdf.

Matarese:1980:LDP

[2883] ANN C. Matarese, Sally L. Richardson, and Jean R. Dunn. Larval development of Pacific tomcod, Microgadus proximus, in the Northeast Pacific Ocean with comparative notes on larvae of walleye pollock, Theragra chalcogramma, and Pacific cod, Gadus macrocephalus (Gadidae). Fishery Bulletin, 78(4):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
larval-development-pacific-tomcod-microgadus-proximus-northeast-
pacific-ocean-comparative;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1980/784/matarese.pdf.

Marliave:1980:SLP

[2884] Jeffrey B. Marliave. Spawn and larvae of the Pacific sandfish, Trichodon trichodon. Fishery Bulletin, 78(4):??, 1980. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/spawn-and-larvae-pacific-
sandfish-trichodon-trichodon;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/marliave.pdf.

Koslow:1981:FSS

[2885] J. Anthony Koslow. Feeding selectivity of schools of northern anchovy, Engraulis mordax, in the Southern California Bight. Fishery Bulletin, 79(1):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/feeding-
selectivity-schools-northern-anchovy-engraulis-mordax-southern-
california-bight;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1981/791/koslow.pdf.

Haynes:1981:DSI

[2886] Evan Haynes. Description of Stage II zoeae of snow crab, Chionoecetes bairdi, (Oxyrhyncha, Majidae) from plankton of lower Cook Inlet, Alaska. Fishery Bulletin, 79(1):177–182, 1981. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/description-stage-ii-zoeae-snow-
crab-chionoecetes-bairdi-oxyrhyncha-majidae-plankton-lower;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/haynes%20(2).pdf.

Al-Judaimi:1981:PCN

[2887] Manal M. Al-Judaimi, A. K. Jafri, and K. A. George. Proximate composition and nutritive value of some important food fishes from the Arabian Gulf. Fishery Bulletin, 79(1):211–212, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic).  URL https://spo.nmfs.noaa.gov/content/proximate-
composition-and-nutritive-value-some-important-food-fishes-
arabian-gulf;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/al-judaimi.pdf.

Richardson:1981:CKL

[2888] Sally L. Richardson. Current knowledge of larvae of sculpins (Pisces: Cottidae and allies) in northeast Pacific genera with notes on intergeneric relationships. Fishery Bulletin, 79(1):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/current-knowledge-larvae-
sculpins-pisces-cottidae-and-allies-northeast-pacific-genera-
notes;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1981/791/richardson.pdf.

Spotte:1981:FRC

[2889] Stephen Spotte and Gary Adams. Feeding rate of captive adult female northern fur seals, Callorhinus ursinus. Fishery Bulletin, 79(1):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/feeding-rate-captive-
adult-female-northern-fur-seals-callorhinus-ursinus;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
spotte.pdf.

Richardson:1981:PEL

[2890] Sally L. Richardson. Pelagic eggs and larvae of the deepsea sole, Embassichthys bathybius (Pisces: Pleuronectidae), with comments on generic affinities. Fishery Bulletin, 79(1):163–170, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/pelagic-eggs-and-larvae-deepsea-sole-
embassichthys-bathybius-pisces-pleuronectidae-comments;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1981/791/
richardson1.pdf.

Kaya:1981:IST

[2891] Calvin M. Kaya, Andrew E. Dizon, and Sharon D. Hendrix. Induced spawning of a tuna, Euthynnus affinus. Fishery Bulletin, 79(1):185– 187, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/induced-
spawning-tuna-euthynnus-affinus;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/kaya.pdf.

Beardsley:1981:ACE

[2892] Grant L. Beardsley and Ramon J. Conser. An analysis of catch and effort data from the U.S. recreational fishery for billfishes (Istiophoridae) in the Western North Atlantic Ocean and Gulf of Mexico, 1971–78. Fishery Bulletin, 79(1):??, 1981. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/analysis-catch-and-effort-data-
us-recreational-fishery-billfishes-istiophoridae-western;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1981/
791/beardsley.pdf.

Townsend:1981:GAS

[2893] David W. Townsend and Joseph J. Graham. Growth and age structure of larval Atlantic herring, Clupea harengus harengus, in the Sheepscot River Estuaky, Maine, as determined by daily growth increments in otoliths. Fishery Bulletin, 79(1):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/growth-and-age-structure-
larval-atlantic-herring-clupea-harengus-harengus-sheepscot-
river;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1981/791/townsend.pdf.

Frost:1981:TIS

[2894] Kathryn J. Frost and Lloyd F. Lowry. Trophic importance of some marine gadids in northern Alaska and their body-otolith size relationships. Fishery Bulletin, 79(1):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/trophic-importance-some-marine-gadids-northern-

alaska-and-their-body-otolith-size;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/frost_0.pdf.

Blackburn:1981:ODL

[2895] Maurice Blackburn and D. L. Serventy. Observations on distribution and life history of skipjack tuna, Katsuwonus pelamis, in Australian waters. Fishery Bulletin, 79(1):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/observations-distribution-and-life-history-skipjack-
tuna-katsuwonus-pelamis-australian;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1981/791/blackburn.pdf.

Uchiyama:1981:AGS

[2896] James H. Uchiyama and Paul Struhsaker. Age and growth of skipjack tuna, Katsuwonus pelamis, and yellowfin tuna, Thunnus albacares, as indicated by daily growth increments of Sagittae. Fishery Bulletin, 79 (1):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/age-and-
growth-skipjack-tuna-katsuwonus-pelamis-and-yellowfin-tuna-
thunnus-albacares;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1981/791/uchiyama.pdf.

Williams:1981:CRA

[2897] Austin B. Williams and David McN. Williams. Carolinian records for American lobster, Homarus americanus, and tropical swimming crab, Callinectes bocourti. Postulated means of dispersal. Fishery Bulletin, 79(1):192–198, 1981. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
carolinian-records-american-lobster-homarus-americanus-and-
tropical-swimming-crab;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/williams_1.pdf.

Gooding:1981:RRL

[2898] Reginald M. Gooding, William H. Neill, and Andrew E. Dizon. Respiration rates and low-oxygen tolerance limits in skipjack tuna, Katsuwonus pelamis. Fishery Bulletin, 79(1):??, 1981. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/respiration-rates-and-low-oxygen-
tolerance-limits-skipjack-tuna-katsuwonus-pelamis;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1981/791/
gooding.pdf.

Fernholm:1981:WAH

[2899] Bo Fernholm and Carl L. Hubbs. Western Atlantic hagfishes of the genus Eptatretus (Myxinidae) with description of two new species. Fishery Bulletin, 79(1):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/western-atlantic-hagfishes-genus-
eptatretus-myxinidae-description-two-new-species;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
fernholm.pdf.

Barker:1981:MAH

[2900] Seth L. Barker, David W. Townsend, and John S. Hacunda. Mortalities of Atlantic herring, Clupea h. harengus, smooth flounder, Liopsetta putnami, and rainbow smelt, Osmerus mordax, larvae exposed to acute thermal shock. Fishery Bulletin, 79(1):198–200, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/mortalities-atlantic-
herring-clupea-h-harengus-smooth-flounder-liopsetta-putnami-
and;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/barker.pdf.

Hobson:1981:CNA

[2901] Edmund S. Hobson, William N. McFarland, and James R. Chess. Crepuscular and nocturnal activities of Californian nearshore fishes, with consideration of their scotopic visual pigments and the photic environment. Fishery Bulletin, 79(1):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/crepuscular-and-nocturnal-activities-
californian-nearshore-fishes-consideration-their;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1981/791/
hobson.pdf.

Webb:1981:BSP

[2902] P. W. Webb and R. T. Corolla. Burst swimming performance of northern anchovy, Engraulis mordax, larvae. Fishery Bulletin, 79 (1):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/burst-
swimming-performance-northern-anchovy-engraulis-mordax-larvae;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/webb.pdf.

Bowman:1981:FSN

[2903] Ray E. Bowman. Food of 10 species of northwest Atlantic juvenile groundfish. Fishery Bulletin, 79(1):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/food-10-species-northwest-atlantic-juvenile-
groundfish;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/bowman.pdf.

Kappenman:1981:MGC

[2904] Russell F. Kappenman. A method for growth curve comparisons. Fishery Bulletin, 79(1):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/method-growth-curve-comparisons;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1981/791/kappenman.
pdf.

Weihs:1981:ESP

[2905] Daniel Weihs. Effects of swimming path curvature on the energetics of fish motion. Fishery Bulletin, 79(1):??, 1981. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/effects-swimming-path-curvature-
energetics-fish-motion;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/weihs.pdf.

Libby:1981:DSR

[2906] David A. Libby. Difference in sex ratios of the anadromous alewife, Alosa pseudoharengus, between the top and bottom of a fishway at Damariscotta Lake, Maine. Fishery Bulletin, 79(1):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/difference-
sex-ratios-anadromous-alewife-alosa-pseudoharengus-between-
top-and-bottom;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/libby.pdf.

Hunter:1981:SEF

[2907] J. Roe Hunter and Roderick Leong. The spawning energetics of female northern anchovy, Engraulis mordax. Fishery Bulletin, 79(2):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/spawning-energetics-
female-northern-anchovy-engraulis-mordax;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1981/792/hunter.pdf.

Laroche:1981:DLJ

[2908] Wayne A. Laroche and Sally L. Richardson. Development of larvae and juveniles of the rockfishes Sebastes entomelas and S. zacentrus (family Scorpaenidae) and occurrence off Oregon, with notes on head spines of S. Mystinus, S. Flavidus, and S. melanops. Fishery Bulletin, 79(2):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/development-larvae-
and-juveniles-rockfishes-sebastes-entomelas-and-s-zacentrus-
family;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1981/792/laroche.pdf.

Laurs:1981:GRN

[2909] R. Michael Laurs and Jerry A. Wetherall. Growth rates of North Pacific albacore, Thunnus alalunga, based on tag returns. Fishery Bulletin, 79(2):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
growth-rates-north-pacific-albacore-thunnus-alalunga-based-
tag-returns;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1981/792/laurs.pdf.

Lynde:1981:EFD

[2910] C. M. Lynde. Economic feasibility of domestic groundfish harvest from Western Alaska waters: a comparison of vessel types, fishing strategies, and processor locations. Fishery Bulletin, 79(2):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/economic-feasibility-
domestic-groundfish-harvest-western-alaska-waters-comparison-
vessel;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1981/792/lynde.pdf.

Somerton:1981:CLH

[2911] David A. Somerton. Contribution to the life history of the deep-sea king crab, Lithodes couesi, in the Gulf of Alaska. Fishery Bulletin, 79(2): ??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/contribution-
life-history-deep-sea-king-crab-lithodes-couesi-gulf-alaska;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1981/792/somerton.pdf.

Ahrenholz:1981:REG

[2912] Dean W. Ahrenholz. Recruitment and exploitation of Gulf menhaden, Brevoortia patronus. Fishery Bulletin, 79(2):??, 1981. CODEN FSYBAY.

ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/recruitment-and-exploitation-gulf-menhaden-
brevoortia-patronus;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1981/792/ahrenholz.pdf.

Anderson:1981:SMS

[2913] James Jay Anderson. A stochastic model for the size of fish schools. Fishery Bulletin, 79(2):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/stochastic-model-size-fish-schools;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1981/
792/anderson.pdf.

Carey:1981:DPA

[2914] Francis G. Carey and Bruce H. Robison. Daily patterns in the activities of swordfish, Xiphias gladius, observed by acoustic telemetry. Fishery Bulletin, 79(2):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/daily-
patterns-activities-swordfish-xiphias-gladius-observed-acoustic-
telemetry;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1981/792/carey.pdf.

Andryszak:1981:CLD

[2915] Bryan L. Andryszak and Robert H. Gore. The complete larval development in the laboratory of Micropanope sculptipes (Crustacea, Decapoda, Xanthidae) with a comparison of larval characters in Western Atlantic xanthid genera. Fishery Bulletin, 79(3): ??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/complete-
larval-development-laboratory-micropanope-sculptipes-crustacea-
decapoda-xanthidae;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1981/793/andryszak.pdf.

Tanaka:1981:EPF

[2916] Kuniaki Tanaka, Yasuo Mugiya, and Juro Yamada. Effects of photoperiod and feeding on daily growth patterns in otoliths of juvenile Tilapia nilotica. Fishery Bulletin, 79(3):??, 1981. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/effects-photoperiod-and-feeding-
daily-growth-patterns-otoliths-juvenile-tilapia-nilotica;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1981/
793/tanaka.pdf.

Demartini:1981:OCF

[2917] E. E. Demartini and Robert K. Fountain. Ovarian cycling frequency and batch fecundity in the queenfish, Seriphus politus: Attributes representative of serial spawning fishes. Fishery Bulletin, 79 (3):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/ovarian-
cycling-frequency-and-batch-fecundity-queenfish-seriphus-politus-
attributes;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1981/793/demartini.pdf.

Wahle:1981:ENR

[2918] Roy J. Wahle and Ed Chaney. Establishment of nonindigenous runs of spring Chinook salmon, Oncorhynchus tshawytscha, in the Wind River Drainage of the Columbia River, 1955–63. Fishery Bulletin, 79(3):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/establishment-
nonindigenous-runs-spring-chinook-salmon-oncorhynchus-tshawytscha-
wind-river;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1981/793/wahle.pdf.

Love:1981:GRF

[2919] Milton S. Love and William V. Westphal. Growth, reproduction, and food habits of olive rockfish, Sebastes serranoides, off Central California. Fishery Bulletin, 79(3):533–545, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/growth-reproduction-and-food-habits-olive-
rockfish-sebastes-serranoides-central-california;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1981/793/
love.pdf.

Wankowski:1981:EGS

[2920] J. W. J. Wankowski. Estimated growth of surface-schooling skipjack tuna, Katsuwonus pelamis, and yellowfin tuna, Thunnus albacares, from the Papua New Guinea Region. Fishery Bulletin, 79(3):??, 1981. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/estimated-growth-surface-
schooling-skipjack-tuna-katsuwonus-pelamis-and-yellowfin-tuna;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1981/793/wankowski.pdf.

Medved:1981:FBB

[2921] Robert J. Medved and Joseph A. Marshall. Feeding behavior and biology of young sandbar sharks, Carcharhinus plumbeus (Pisces, Carcharhinidae), in Chincoteague Bay, Virginia. Fishery Bulletin, 79(3):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/feeding-behavior-and-
biology-young-sandbar-sharks-carcharhinus-plumbeus-pisces;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1981/
793/medved.pdf.

Haynes:1981:EZS

[2922] Evan Haynes. Early zoeal stages of Lebbeus polaris, Eualus suckleyi, E. fabricii, Spirontocaris arcuata, S. ochotensis, and Heptacarpus camtschaticus (Crustacea, Decapoda, Caridea, Hippolytidae) and morphological characterization of zoeae of Spirontocaris and related genera. Fishery Bulletin, 79(3):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/early-zoeal-stages-lebbeus-polaris-eualus-
suckleyi-e-fabricii-spirontocaris-arcuata-s;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/haynes%
20(3).pdf.

Neves:1981:ODA

[2923] Richard J. Neves. Offshore distribution of alewife, Alosa pseudoharengus, and blueback herring, Alosa aestivalis, along the Atlantic Coast. Fishery Bulletin, 79(3):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/offshore-distribution-alewife-alosa-pseudoharengus-
and-blueback-herring-alosa-aestivalis;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1981/793/neves.pdf.

Goldberg:1981:SSC

[2924] Stephen R. Goldberg. Seasonal spawning cycle of the black croaker, Cheilotrema saturnum (Sciaenidae). Fishery Bulletin, 79(3):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/seasonal-
spawning-cycle-black-croaker-cheilotrema-saturnum-sciaenidae;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/goldberg%20(1)_0.pdf.

Pietsch:1981:ORA

[2925] Theodore W. Pietsch. The osteology and relationships of the anglerfish genus Tetrabrachium with comments on lophiiform classification. Fishery Bulletin, 79(3):387–419, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/osteology-and-relationships-anglerfish-genus-tetrabrachium-

comments-lophiiform;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1981/793/pietsch.pdf.

Antonelis:1981:PGC

[2926] George A. Antonelis, Jr., Stephen Leatherwood, and Daniel K. Odell. Population growth and censuses of the northern elephant seal, Mirounga angustirostris, on the California Channel Islands, 1958–78. Fishery Bulletin, 79(3):562–567, 1981. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/population-growth-and-censuses-
northern-elephant-seal-mirounga-angustirostris-california;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/antonelis.pdf.

Pitcher:1981:PSS

[2927] Kenneth W. Pitcher. Prey of the Steller sea lion, Eumetopias jubatus, in the Gulf of Alaska. Fishery Bulletin, 79(3):??, 1981. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/prey-steller-sea-lion-eumetopias-
jubatus-gulf-alaska;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1981/793/pitcher.pdf.

Brege:1981:GCY

[2928] Dean A. Brege. Growth characteristics of young-of-the-year walleye, Stizostedion vitreum vitreum, in John Day Reservoir on the Columbia River, 1979. Fishery Bulletin, 79(3):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/growth-characteristics-young-year-walleye-
stizostedion-vitreum-vitreum-john-day-reservoir;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
brege.pdf.

Robinson:1981:SCS

[2929] William E. Robinson, William E. Wehling, M. Patricia Morse, and Guy C. McLeod. Seasonal changes in soft-body component indices and energy reserves in the Atlantic deep-sea scallop, Placopecten magellanicus. Fishery Bulletin, 79(3):449–458, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/seasonal-changes-soft-body-component-
indices-and-energy-reserves-atlantic-deep-sea-scallop;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1981/793/
robinson.pdf.

Holt:1981:ETS

[2930] Joan Holt, Robert Godbout, and C. R. Arnold. Effects of temperature and salinity on egg hatching and larval survival of red drum, Sciaenops ocellata. Fishery Bulletin, 79(3):569–573, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/effects-temperature-and-salinity-egg-
hatching-and-larval-survival-red-drum-sciaenops;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
holt.pdf.

Durbin:1981:AEN

[2931] Edward G. Durbin and Ann G. Durbin. Assimilation efficiency and nitrogen excretion of a filter-feeding planktivore, the Atlantic menhaden, Brevoortia tyrannus (Pisces: Clupeidae). Fishery Bulletin, 79(4):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/assimilation-
efficiency-and-nitrogen-excretion-filter-feeding-planktivore-
atlantic-menhaden;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1981/794/durbin.pdf.

Pearson:1981:ICA

[2932] Walter H. Pearson, Peter C. Sugarman, Dana L. Woodruff, and Bori L. Olla. Impairment of the chemosensory antennular flicking response in the Dungeness crab, Cancer magister, by petroleum hydrocarbons. Fishery Bulletin, 79(4):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
impairment-chemosensory-antennular-flicking-response-dungeness-
crab-cancer-magister;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1981/794/pearson.pdf.

Ainley:1981:MSH

[2933] David G. Ainley, Anthony R. DeGange, Linda L. Jones, and Richard J. Beach. Mortality of seabirds in high-seas salmon gill nets. Fishery Bulletin, 79(4):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
mortality-seabirds-high-seas-salmon-gill-nets;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/ainley%
20(1).pdf.

Flint:1981:GMS

[2934] R. Warren Flint and Nancy N. Rabalais. Gulf of Mexico shrimp production: a food web hypothesis. Fishery Bulletin, 79(4):??, 1981.

CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/gulf-mexico-shrimp-
production-food-web-hypothesis;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1981/794/flint.pdf.

Shlossman:1981:RMP

[2935] Philip A. Shlossman and Mark E. Chittenden, Jr. Reproduction, movements, and population dynamics of the sand seatrout, Cynoscion arenarius. Fishery Bulletin, 79(4):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/reproduction-movements-and-population-dynamics-
sand-seatrout-cynoscion-arenarius;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1981/794/shlossman.pdf.

OConnell:1981:HIL

[2936] Charles P. O’Connell and Pedro A. Paloma. Histochemical indications of liver glycogen in samples of emaciated and robust larvae of the northern anchovy, Engraulis mordax. Fishery Bulletin, 79(4):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
histochemical-indications-liver-glycogen-samples-emaciated-
and-robust-larvae-northern;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/oconnell_0.pdf.

Gabriel:1981:FSD

[2937] Wendy L. Gabriel and William G. Pearcy. Feeding selectivity of Dover sole, Microstomus pacificus, off Oregon. Fishery Bulletin, 79(4):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/feeding-selectivity-
dover-sole-microstomus-pacificus-oregon;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1981/794/gabriel.pdf.

Stevenson:1981:MYE

[2938] David K. Stevenson and Francisco Carranza. Maximum yield estimates for the Pacific thread herring, Opisthonema spp., fishery in Costa Rica. Fishery Bulletin, 79(4):??, 1981. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/maximum-yield-estimates-pacific-
thread-herring-opisthonema-spp-fishery-costa-rica;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1981/794/
stevenson.pdf.

Anonymous:1981:I

[2939] Anonymous. Index. Fishery Bulletin, 79(4):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/index-108;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/index%
20(2)_1.pdf.

Gruber:1981:TSB

[2940] S. H. Gruber and L. J. V. Compagno. Taxonomic status and biology of the bigeye thresher, Alopias superciliosus. Fishery Bulletin, 79(4): 617–640, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/taxonomic-
status-and-biology-bigeye-thresher-alopias-superciliosus;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1981/
794/gruber.pdf.

Toll:1981:CDS

[2941] Ronald B. Toll and Steven C. Hess. Cephalopods in the diet of the swordfish, Xiphias gladius, from the Florida Straits. Fishery Bulletin, 79(4):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
cephalopods-diet-swordfish-xiphias-gladius-florida-straits;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1981/794/toll.pdf.

Hacunda:1981:TRA

[2942] John S. Hacunda. Trophic relationships among demersal fishes in a coastal area of the Gulf of Maine. Fishery Bulletin, 79(4): ??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/trophic-
relationships-among-demersal-fishes-coastal-area-gulf-maine;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1981/794/hacunda.pdf.

Webb:1981:RNA

[2943] P. W. Webb. Responses of northern anchovy, Engraulis mordax, larvae to predation by a biting planktivore, Amphiprion percula. Fishery Bulletin, 79(4):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/responses-northern-anchovy-engraulis-mordax-larvae-
predation-biting-planktivore-amphiprion;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/webb%20(
1).pdf.

Irvine:1981:MAA

[2944] A. Blair Irvine, Michael D. Scott, Randall S. Wells, and John H. Kaufmann. Movements and activities of the Atlantic bottlenose dolphin, Tursiops truncatus, near Sarasota, Florida. Fishery Bulletin, 79(4):671– 688, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/movements-and-
activities-atlantic-bottlenose-dolphin-tursiops-truncatus-near-
sarasota;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1981/794/irvine.pdf.

Boehlert:1981:EPT

[2945] George W. Boehlert. The effects of photoperiod and temperature on laboratory growth of juvenile Sebastes diploproa and a comparison with growth in the field. Fishery Bulletin, 79(4):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/effects-photoperiod-
and-temperature-laboratory-growth-juvenile-sebastes-diploproa-
and;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/boehlert_0.pdf.

Kendall:1981:DDD

[2946] Arthur W. Kendall, Jr. and N. A. Naplin. Diel-depth distribution of summer ichthyoplankton in the Middle Atlantic Bight. Fishery Bulletin, 79(4):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/diel-depth-distribution-summer-ichthyoplankton-
middle-atlantic-bight;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1981/794/kendall.pdf.

Love:1981:CBA

[2947] Milton S. Love and William V. Westphal. A correlation between annual catches of Dungeness crab, Cancer magister, along the west coast of North America and mean annual sunspot number. Fishery Bulletin, 79 (4):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/correlation-
between-annual-catches-dungeness-crab-cancer-magister-along-
west-coast-north;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/love_1.pdf.

McHugh:1981:MFD

[2948] J. L. McHugh. Marine fisheries of Delaware. Fishery Bulletin, 79 (4):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/marine-
fisheries-delaware;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1981/794/mchugh.pdf.

Ennis:1981:FAL

[2949] G. P. Ennis. Fecundity of the American lobster, Homarus americanus, in Newfoundland waters. Fishery Bulletin, 79(4):??, 1981. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/fecundity-american-
lobster-homarus-americanus-newfoundland-waters;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/ennis_
0.pdf.

Baglin:1982:RBW

[2950] Raymond E. Baglin, Jr. Reproductive biology of Western Atlantic bluefin tuna. Fishery Bulletin, 80(1):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/reproductive-biology-western-atlantic-bluefin-tuna;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1982/801/baglin.pdf.

Conover:1982:OWM

[2951] David O. Conover and Steven A. Murawski. Offshore winter migration of the Atlantic silverside, Menidia menidia. Fishery Bulletin, 80 (1):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/offshore-
winter-migration-atlantic-silverside-menidia-menidia;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
conover.pdf.

Hjort:1982:PDA

[2952] R. C. Hjort and C. B. Schreck. Phenotypic differences among stocks of hatchery and wild Coho salmon, Oncorhynchus kisutch, in Oregon, Washington, and California. Fishery Bulletin, 80(1):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/phenotypic-
differences-among-stocks-hatchery-and-wild-coho-salmon-oncorhynchus-
kisutch;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1982/801/hjort.pdf.

Rosenberg:1982:GDM

[2953] Andrew A. Rosenberg and Joanne Lyczkowski Laroche. Growth during metamorphosis of English sole, Parophrys vetulus. Fishery Bulletin, 80(1):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
growth-during-metamorphosis-english-sole-parophrys-vetulus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/rosenberg.pdf.

Irvine:1982:ETT

[2954] A. B. Irvine, R. S. Wells, and M. D. Scott. An evaluation of techniques for tagging small odontocete cetaceans. Fishery Bulletin, 80(1):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/evaluation-techniques-tagging-small-odontocete-
cetaceans;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1982/801/irvine.pdf.

Pratt:1982:OLW

[2955] Harold L. Pratt, Jr., John G. Casey, and Robert B. Conklin. Observations on large white sharks, Carcharodon carcharias, off Long Island, New York. Fishery Bulletin, 80(1):??, 1982. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/observations-large-white-sharks-
carcharodon-carcharias-long-island-new-york;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/pratt.
pdf.

Laroche:1982:AGP

[2956] Joanne Lyczkowski Laroche, Sally L. Richardson, and Andrew A. Rosenberg. Age and growth of a pleuronectid, Parophrys vetulus, during the pelagic larval period in Oregon coastal waters. Fishery Bulletin, 80 (1):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/age-and-
growth-pleuronectid-parophrys-vetulus-during-pelagic-larval-
period-oregon-coastal;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1982/801/laroche.pdf.

Gibson:1982:NET

[2957] D. M. Gibson. A note on the estimation of trimethylamine in fish muscle. Fishery Bulletin, 80(1):??, 1982. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/note-estimation-trimethylamine-
fish-muscle;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/gibson.pdf.

Manooch:1982:DAA

[2958] Charles S. Manooch III and Charles A. Barans. Distribution, abundance, and age and growth of the tomtate, Haemulon aurolineatum, along the Southeastern United States Coast. Fishery Bulletin, 80(1):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/distribution-abundance-and-age-and-growth-
tomtate-haemulon-aurolineatum-along-southeastern;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1982/801/
manooch.pdf.

Crass:1982:SDW

[2959] Dennis W. Crass and Robert H. Gray. Snout dimorphism in white sturgeon, Acipenser transmontanus, from the Columbia River at Hanford, Washington. Fishery Bulletin, 80(1):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/snout-dimorphism-white-sturgeon-acipenser-
transmontanus-columbia-river-hanford-washington;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
crass.pdf.

Murawski:1982:GOQ

[2960] Steven A. Murawski, John W. Ropes, and Fredric M. Serchuki. Growth of the ocean quahog, Arctica islandica, in the Middle Atlantic Bight. Fishery Bulletin, 80(1):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
growth-ocean-quahog-arctica-islandica-middle-atlantic-bight;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1982/801/murawski.pdf.

Tucker:1982:LDC

[2961] John W. Tucker, Jr. Larval development of Citharichthys cornutus, C. gymnorhinus, C. spilopterus, and Etropus crossotus (Bothidae), with notes on larval occurrence. Fishery Bulletin, 80(1):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/larval-development-
citharichthys-cornutus-c-gymnorhinus-c-spilopterus-and-etropus-
crossotus;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1982/801/tucker.pdf.

Wiebe:1982:ATN

[2962] P. H. Wiebe, S. H. Boyd, B. M. Davis, and J. L. Cox. Avoidance of towed nets by the euphausiid Nematoscelis megalops. Fishery Bulletin, 80(1):

75–91, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/avoidance-
towed-nets-euphausiid-nematoscelis-megalops;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1982/801/wiebe.
pdf.

Clarke:1982:FHS

[2963] Thomas A. Clarke. Feeding habits of stomiatoid fishes from Hawaiian waters. Fishery Bulletin, 80(2):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/feeding-habits-stomiatoid-fishes-
hawaiian-waters;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1982/802/clarke.pdf.

Potthoff:1982:DVCa

[2964] Thomas Potthoff and Sharon Kelley. Development of the vertebral column, fins and fin supports, branchiostegal rays, and squamation in the swordfish, Xiphias gladius. Fishery Bulletin, 80(2):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/development-
vertebral-column-fins-and-fin-supports-branchiostegal-rays-
and-squamation;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1982/802/potthoff.pdf.

Dawson:1982:ELT

[2965] Margaret A. Dawson. Effects of long-term mercury exposure on hematology of striped bass, Morone saxatilis. Fishery Bulletin,

80(2):??, 1982.  CODEN FSYBAY.  ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
effects-long-term-mercury-exposure-hematology-striped-bass-
morone-saxatilis;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/dawson.pdf.

Hain:1982:FBH

[2966] James H. W. Hain, Gary R. Carter, Scott D. Kraus, Charles A. Mayo, and Howard E. Winn. Feeding behavior of the humpback whale, Megaptera novaeangliae, in the Western North Atlantic. Fishery Bulletin, 80(2):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/feeding-behavior-humpback-whale-megaptera-
novaeangliae-western-north-atlantic;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1982/802/hain.pdf.

Radtke:1982:IFO

[2967] R. L. Radtke and J. M. Dean. Increment formation in the otoliths of embryos, larvae, and juveniles of the mummichog, Fundulus heteroclitus. Fishery Bulletin, 80(2):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/increment-formation-otoliths-embryos-
larvae-and-juveniles-mummichog-fundulus-heteroclitus;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1982/802/
radtke.pdf.

Kaya:1982:RSM

[2968] Calvin M. Kaya, Andrew E. Dizon, Sharon D. Hendricks, Thomas K. Kazama, and Martina K. K. Queenth. Rapid and spontaneous maturation, ovulation, and spawning of ova by newly captured skipjack tuna, Katsuwonus pelamis. Fishery Bulletin, 80(2):??, 1982. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/rapid-and-spontaneous-maturation-
ovulation-and-spawning-ova-newly-captured-skipjack-tuna;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
kaya%20(1).pdf.

Haynes:1982:DLG

[2969] Evan Haynes. Description of larvae of the golden king crab, Lithodes aequispina, reared in the laboratory. Fishery Bulletin, 80(2):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/description-larvae-
golden-king-crab-lithodes-aequispina-reared-laboratory;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1982/802/
haynes.pdf.

Rosenberg:1982:GJE

[2970] Andrew A. Rosenberg. Growth of juvenile English sole, Parophrys vetulus, in estuarine and open coastal nursery grounds. Fishery Bulletin, 80 (2):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/growth-
juvenile-english-sole-parophrys-vetulus-estuarine-and-open-
coastal-nursery-grounds;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1982/802/rosenberg.pdf.

Lo:1982:EMI

[2971] Nancy C. H. Lo, Joseph E. Powers, and Bruce E. Wahlen. Estimating and monitoring incidental dolphin mortality in the eastern tropical Pacific tuna purse seine fishery. Fishery Bulletin, 80(2):396–401, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/estimating-
and-monitoring-incidental-dolphin-mortality-eastern-tropical-
pacific-tuna-purse;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/lo.pdf.

Knight:1982:LDS

[2972] Margaret Knight and Makoto Omori. The larval development of Sergestes similis Hansen (Crustacea, Decapoda, Sergestidae) reared in the laboratory. Fishery Bulletin, 80(2):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/larval-development-sergestes-similis-
hansen-crustacea-decapoda-sergestidae-reared;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1982/802/knight.
pdf.

Skinner:1982:IWQ

[2973] Renate H. Skinner. The interrelation of water quality, gill parasites, and gill pathology of some fishes from South Biscayne Bay, Florida. Fishery Bulletin, 80(2):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/interrelation-water-quality-gill-parasites-and-
gill-pathology-some-fishes-south-biscayne;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1982/802/skinner.pdf.

Joyce:1982:WDP

[2974] Gerald G. Joyce, John V. Rosapepe, and Junroku Ogasawara. White Dall’s porpoise sighted in the North Pacific. Fishery Bulletin, 80(2):401– 402, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/white-dalls-
porpoise-sighted-north-pacific;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/joyce.pdf.

Lough:1982:AGL

[2975] R. Gregory Lough, Michael Pennington, George R. Bolz, and Andrew A. Rosenberg. Age and growth of larval Atlantic herring, Clupea harengus L., in the Gulf of Maine–Georges Bank region based on otolith growth increments. Fishery Bulletin, 80(2):187–199, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/age-and-growth-larval-atlantic-
herring-clupea-harengus-l-gulf-maine-georges-bank-region;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1982/
802/lough.pdf.

Trumble:1982:SGH

[2976] Robert J. Trumble, Richard E. Thorne, and Norman A. Lemberg. The Strait of Georgia herring fishery: a case history of timely management aided by hydroacoustic surveys. Fishery Bulletin, 80(2):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/strait-georgia-herring-fishery-case-history-timely-
management-aided-hydroacoustic-surveys;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1982/802/trumble.pdf.

Ainley:1982:PFC

[2977] David G. Ainley, Harriet R. Huber, and Kevin M. Bailey. Population fluctuations of California sea lions and the Pacific whiting fishery off Central California. Fishery Bulletin, 80(2):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/population-fluctuations-california-sea-lions-
and-pacific-whiting-fishery-central-california;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1982/802/ainley.
pdf.

Mann:1982:SCG

[2978] Roger Mann. The seasonal cycle of gonadal development in Arctica islandica from the Southern New England Shelf. Fishery Bulletin, 80 (2):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/seasonal-
cycle-gonadal-development-arctica-islandica-southern-new-england-
shelf;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1982/802/mann.pdf.

Whitledge:1982:RNN

[2979] Terry E. Whitledge. Regeneration of nitrogen by the nekton and its significance in the Northwest Africa Upwelling ecosystem. Fishery Bulletin, 80(2):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
regeneration-nitrogen-nekton-and-its-significance-northwest-
africa-upwelling-ecosystem;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1982/802/whitledge.pdf.

Au:1982:MSD

[2980] D. Au and W. Perryman. Movement and speed of dolphin schools responding to an approaching ship. Fishery Bulletin, 80(2):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/movement-and-speed-
dolphin-schools-responding-approaching-ship;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1982/802/au.pdf.

Matarese:1982:LDL

[2981] Ann C. Matarese and Jeffrey B. Marliave. Larval development of laboratory-reared rosylip sculpin, Ascelichthys rhodorus (Cottidae). Fishery Bulletin, 80(2):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/larval-development-laboratory-reared-
rosylip-sculpin-ascelichthys-rhodorus-cottidae;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1982/802/matarese.
pdf.

Wolff:1982:BKE

[2982] Gary A. Wolff. A beak key for eight Eastern Tropical Pacific cephalopod species with relationships between their beak dimensions and size. Fishery Bulletin, 80(2):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
beak-key-eight-eastern-tropical-pacific-cephalopod-species-
relationships-between-their-beak;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1982/802/wolff.pdf.

Brundage:1982:ASA

[2983] Harold M. Brundage III and Robert E. Meadows. The Atlantic sturgeon, Acipenser oxyrhynchus, in the Delaware River estuary. Fishery Bulletin, 80(2):337–343, 1982. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
atlantic-sturgeon-acipenser-oxyrhynchus-delaware-river-estuary;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1982/802/brundage.pdf.

Miller:1982:EPI

[2984] Ruth Miller and John Spinelli. The effect of protease inhibitors on proteolysis in parasitized Pacific whiting, Merluccius productus, muscle. Fishery Bulletin, 80(2):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
effect-protease-inhibitors-proteolysis-parasitized-pacific-
whiting-merluccius-productus;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1982/802/miller.pdf.

Potthoff:1982:DVCb

[2985] Thomas Potthoff and Sharon Kelley. Development of the vertebral column, fins and fin supports, branchiostegal rays, and squamation in the swordfish, Xiphias gladius. Fishery Bulletin, 80(2):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/development-
ofthe-vertebral-column-fins-and-fin-supports-branchiostegal-
rays-and-squamation;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/potthoff%20(1).pdf.

Devries:1982:SAD

[2986] Douglas A. Devries and Mark E. Chittenden, Jr. Spawning, age determination, longevity, and mortality of the silver seatrout, Cynoscion nothus, in the Gulf of Mexico. Fishery Bulletin, 80(3):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/spawning-age-
determination-longevity-and-mortality-silver-seatrout-cynoscion-
nothus-gulf;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1982/803/devries.pdf.

Ulanowicz:1982:ICF

[2987] Robert E. Ulanowicz, Mohammed Liaquat Ali, Alice Vivian, Donald R. Heinle, William A. Richkus, and J. Kevin Summers. Identifying climatic factors influencing commercial fish and shellfish landings in Maryland. Fishery Bulletin, 80(3):611–619, 1982. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/identifying-climatic-
factors-influencing-commercial-fish-and-shellfish-landings-
maryland;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1982/803/ulanowicz.pdf.

Eldridge:1982:BGS

[2988] Maxwell B. Eldridge, Jeannetie A. Whipple, and Michael J. Bowers. Bioenergetics and growth of striped bass, Morone saxatilis, embryos and larvae. Fishery Bulletin, 80(3):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/bioenergetics-and-growth-
striped-bass-morone-saxatilis-embryos-and-larvae;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1982/803/
eldridge.pdf.

Watson:1982:DEL

[2989] William Watson. Development of eggs and larvae of the white croaker, Genyonemus lineatus Ayres (Pisces: Sciaenidae), off the southern California coast. Fishery Bulletin, 80(3):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/development-eggs-and-larvae-white-croaker-
genyonemus-lineatus-ayres-pisces-sciaenidae;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1982/803/watson.
pdf.

Geoghegan:1982:RMP

[2990] Paul Geoghegan and Mark E. Chittenden, Jr. Reproduction, movements, and population dynamics of the longspine porgy, Stenotomus caprinus. Fishery Bulletin, 80(3):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/reproduction-movements-and-population-dynamics-
longspine-porgy-stenotomus-caprinus;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1982/803/geoghegan.pdf.

Ralston:1982:MAC

[2991] Stephen Ralston and Jeffrey J. Polovina. A multispecies analysis of the commercial deep-sea handline fishery in Hawaii. Fishery Bulletin, 80(3):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/multispecies-analysis-commercial-deep-sea-
handline-fishery-hawaii;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/ralston_0.pdf.

Anger:1982:ECC

[2992] Klaus Anger and Ralph R. Dawirs. Elemental composition (C, N, H) and energy in growing and starving larvae of Hyas araneus (Decapoda, Majidae). Fishery Bulletin, 80(3):??, 1982. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/elemental-composition-c-n-h-and-
energy-growing-and-starving-larvae-hyas-araneus-decapoda;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1982/
803/anger.pdf.

Hogue:1982:FEA

[2993] E. W. Hogue and A. G. Carey, Jr. Feeding ecology of 0-age flatfishes at a nursery ground on the Oregon Coast. Fishery Bulletin, 80 (3):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/feeding-
ecology-o-age-flatfishes-nursery-ground-oregon-coast;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1982/803/
hogue.pdf.

Kane:1982:ESL

[2994] Joseph Kane. Effect of season and location on the relationship between zooplankton displacement volume and dry weight in the northwest Atlantic. Fishery Bulletin, 80(3):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/effect-season-and-location-relationship-between-
zooplankton-displacement-volume-and-dry;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/kane.pdf.

Bailey:1982:ELH

[2995] Kevin M. Bailey. The early life history of the Pacific hake, Merluccius productus. Fishery Bulletin, 80(3):??, 1982. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/early-life-history-
pacific-hake-merluccius-productus;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1982/803/bailey.pdf.

Irvine:1982:ASM

[2996] A. Blair Irvine, John E. Caffin, and Howard I. Kochman. Aerial surveys for manatees and dolphins in Western Peninsular Florida. Fishery Bulletin, 80(3):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
aerial-surveys-manatees-and-dolphins-western-peninsular-florida;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1982/803/irvine.pdf.

Brousseau:1982:EES

[2997] Diane J. Brousseau, Jenny A. Baglivo, and George E. Lang, Jr. Estimation of equilibrium settlement rates for benthic marine invertebrates: Its application to Mya arenaria (Mollusca: Pelecypoda). Fishery Bulletin, 80(3):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
estimation-equilibrium-settlement-rates-benthic-marine-invertebrates-
its-application-mya;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/brousseau.pdf.

Bath:1982:BWP

[2998] D. W. Bath and J. M. O’Connor. THE biology of the white perch, Morone americana, in the Hudson River Estuary. Fishery Bulletin, 80(3): ??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/biology-white-
perch-morone-americana-hudson-river-estuary;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1982/803/bath.
pdf.

Macy:1982:DAO

[2999] William K. Macy III. Development and application of an objective method for classifying long-finned squid, Loligo pealei, into sexual maturity stages. Fishery Bulletin, 80(3):449–459, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/development-and-application-objective-
method-classifying-long-finned-squid-loligo-pealei;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1982/803/
macy.pdf.

Holt:1982:GJR

[3000] Scott A. Holt and Connie R. Arnold. Growth of juvenile red snapper, Lutjanus campechanus, in the northwestern Gulf of Mexico. Fishery Bulletin, 80(3):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/growth-
juvenile-red-snapper-lutjanus-campechanus-northwestern-gulf-
mexico;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/holt%20(1).pdf.

Colby:1982:PSA

[3001] David R. Colby, Donald E. Hoss, and J. H. S. Blaxter. Pressure sensitivity of Atlantic herring, Clupea harengus L., larvae. Fishery Bulletin, 80(3):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/pressure-
sensitivity-atlantic-herring-clupea-harengus-l-larvae;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1982/803/
colby.pdf.

Morgan:1982:SRR

[3002] G. R. Morgan, B. F. Phillips, and L. M. Joll. Stock and recruitment relationships in Panulirus cygnus, the commercial rock (spiny) lobster of Western Australia. Fishery Bulletin, 80(3):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic).  URL https://spo.nmfs.noaa.gov/content/stock-and-
recruitment-relationships-panulirus-cygnus-commercial-rock-
spiny-lobster-western;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1982/803/morgan.pdf.

Banas:1982:ABP

[3003] P. T. Banas, D. E. Smith, and D. C. Biggs. An association between a pelagic octopod, Argonauta sp. Linnaeus 1758, and aggregate salps. Fishery Bulletin, 80(3):648–650, 1982. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/association-between-
pelagic-octopod-argonauta-sp-linnaeus-1758-and-aggregate-salps;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/banas.pdf.

Daniels:1982:FES

[3004] Robert A. Daniels. Feeding ecology of some fishes of the Antarctic Peninsula. Fishery Bulletin, 80(3):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/feeding-ecology-some-fishes-antarctic-peninsula;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1982/803/daniels.pdf.

Rothlisberg:1982:VME

[3005] Peter C. Rothlisberg. Vertical migration and its effect on dispersal of penaeid shrimp larvae in the Gulf of Carpentaria, Australia. Fishery Bulletin, 80(3):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
vertical-migration-and-its-effect-dispersal-penaeid-shrimp-
larvae-gulf-carpentaria;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1982/803/rothlisberg.pdf.

Miller:1982:MJW

[3006] David R. Miller. Migration of a juvenile wolf eel, Anarrhichthys ocellatus, from Port Hardy, British Columbia, to Willapa Bay, Washington. Fishery Bulletin, 80(3):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/migration-juvenile-wolf-eel-anarrhichthys-
ocellatus-port-hardy-british-columbia-willapa-bay;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
miller%20(1).pdf.

Allen:1982:SAC

[3007] Larry G. Allen. Seasonal abundance, composition, and productivity of the littoral fish assemblage in Upper Newport Bay, California. Fishery Bulletin, 80(4):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
seasonal-abundance-composition-and-productivity-littoral-fish-
assemblage-upper-newport-bay;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1982/804/allen.pdf.

Laroche:1982:TPA

[3008] Joanne Lyczkowski Laroche. Trophic patterns among larvae of five species of sculpins (family: Cottidae) in a Maine estuary. Fishery Bulletin, 80(4):827–840, 1982. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
trophic-patterns-among-larvae-five-species-sculpins-family-
cottidae-maine-estuary;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1982/804/laroche.pdf.

Botsford:1982:CCCa

[3009] Louis W. Botsford, Richard D. Methot, Jr., and James E. Wilen. Cycle covariation in the California king salmon, Oncorhynchus tshawytscha, silver salmon, O. kisutch, and Dungeness crab, Cancer magister, fisheries. Fishery Bulletin, 80(4):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/cycle-covariation-california-king-salmon-oncorhynchus-
tshawytscha-silver-salmon-o-kisutch;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/botsford.pdf.

Anonymous:1982:I

[3010] Anonymous. Index. Fishery Bulletin, 80(4):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/index-109;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/index%
20(3)_0.pdf.

Beacham:1982:PBC

[3011] Terry D. Beacham and Paul Starr. Population biology of chum salmon, Oncorhynchus keta, from the Fraser River, British Columbia. Fishery Bulletin, 80(4):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
population-biology-chum-salmon-oncorhynchus-keta-fraser-river-
british-columbia;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1982/804/beacham.pdf.

Millikin:1982:QQN

[3012] Mark R. Millikin. Qualitative and quantitative nutrient requirements of fishes: a review. Fishery Bulletin, 80(4):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/qualitative-and-quantitative-nutrient-
requirements-fishes-review;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1982/804/millikin.pdf.

Boehlert:1982:FWR

[3013] George W. Boehlert, W. H. Barss, and P. B. Lamberson. Fecundity of the widow rockfish, Sebastes entomelas, off the coast of Oregon. Fishery Bulletin, 80(4):881–884, 1982. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fecundity-widow-rockfish-sebastes-entomelas-coast-oregon;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/boehlert%20(1).pdf.

Botsford:1982:CCCb

[3014] Louis W. Botsford, Richard D. Methot, Jr., and James E. Wilen. Cyclic covariation in the California king salmon, Oncorhynchus tshawytscha, silver salmon, O. kisutch, and Dungeness crab, Cancer magister, fisheries. Fishery Bulletin, 80(4):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/cyclic-covariation-california-king-salmon-
oncorhynchus-tshawytscha-silver-salmon-o-kisutch;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1982/804/
botsford.pdf.

Peterson:1982:FHJ

[3015] William T. Peterson, Richard D. Brodeur, and William G. Pearcy. Food habits of juvenile salmon in the Oregon Coastal Zone, June 1979. Fishery Bulletin, 80(4):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
food-habits-juvenile-salmon-oregon-coastal-zone-june-1979;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1982/
804/peterson.pdf.

Babinchak:1982:CSA

[3016] John A. Babinchak, Daniel Goldmintz, and Gary P. Richards. A comparative study of autochthonous bacterial flora on the gills of the blue crab, Callinectes sapidus, and its environment. Fishery Bulletin, 80(4):884–890, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/comparative-study-autochthonous-bacterial-
flora-gills-blue-crab-callinectes-sapidus-and-its;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
babinchak.pdf.

Burgess:1982:FNS

[3017] Lourdes Alvina Burgess. Four new species of squid (Oegopsida: Enoploteuthis) from the Central Pacific and a description of adult Enoploteuthis reticulata. Fishery Bulletin, 80(4):??, 1982. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/four-new-species-squid-oegopsida-
enoploteuthis-central-pacific-and-description-adult;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1982/804/
burgess.pdf.

Waltz:1982:BWP

[3018] C. Wayne Waltz, William A. Roumillat, and Charles A. Wenner. Biology of the whitebone porgy, Calamus leucosteus, in the South Atlantic Bight. Fishery Bulletin, 80(4):??, 1982. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/biology-whitebone-porgy-calamus-
leucosteus-south-atlantic-bight;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1982/804/waltz.pdf.

LeBoeuf:1982:WSP

[3019] Burney J. Le Boeuf, Marianne Riedman, and Raymond S. Keys. White shark predation on pinnipeds in California coastal waters. Fishery Bulletin, 80(4):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/white-shark-predation-pinnipeds-california-coastal-
waters;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/leboeuf.pdf.

Colin:1982:SLD

[3020] Patrick L. Colin. Spawning and larval development of the hogfish, Lachnolaimus maximus (Pisces: Labridae). Fishery Bulletin, 80(4):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/spawning-and-larval-
development-hogfish-lachnolaimus-maximus-pisces-labridae;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1982/
804/colin.pdf.

Watkins:1982:ORW

[3021] William A. Watkins and William E. Schevill. Observations of right whales, Eubalaena glacialis, in Cape Cod waters. Fishery Bulletin, 80(4): ??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/observations-
right-whales-eubalaena-glacialis-cape-cod-waters;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1982/804/
watkins.pdf.

Schlotterbeck:1982:VST

[3022] Robert E. Schlotterbeck and David W. Conally. Vertical stratification of three nearshore southern California larval fishes (Engraulis mordax, Genyonemus lineatus, and Seriphus politus). Fishery Bulletin, 80 (4):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/vertical-
stratification-three-nearshore-southern-california-larval-fishes-
engraulis-mordax;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/schlotterbeck.pdf.

Creaser:1982:LHS

[3023] Edwin P. Creaser and David A. Clifford. Life history studies of the sandworm, Nereis virens Sars, in the Sheepscot Estuary, Maine. Fishery Bulletin, 80(4):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
life-history-studies-sandworm-nereis-virens-sars-sheepscot-
estuary-maine;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1982/804/creaser.pdf.

Webb:1982:SKS

[3024] P. W. Webb and Raymond S. Keyes. Swimming kinematics of sharks. Fishery Bulletin, 80(4):803–812, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/swimming-kinematics-sharks;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1982/804/webb.pdf.

Libby:1982:DLP

[3025] David A. Libby. Decrease in length at predominant ages during a spawning migration of the alewife, Alosa pseudoharengus. Fishery Bulletin, 80(4):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
decrease-length-predominant-ages-during-spawning-migration-
alewife-alosa-pseudoharengus;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/libby%20(1).pdf.

Hettler:1982:RWT

[3026] William F. Hettler and Alexander J. Chester. The relationship of winter temperature and spring landings of pink shrimp, Penaeus duorarum, in North Carolina. Fishery Bulletin, 80(4):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/relationship-
winter-temperature-and-spring-landings-pink-shrimp-penaeus-
duorarum-north;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1982/804/hettler.pdf.

Wetherall:1982:ADT

[3027] Jerry A. Wetherall. Analysis of double-tagging experiments. Fishery Bulletin, 80(4):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/analysis-double-tagging-experiments;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1982/804/
wetherall.pdf.

Goldberg:1982:SSC

[3028] Stephen R. Goldberg. Seasonal spawning cycle of the longfin sanddab, Citharichthys xanthostigma (Bothidae). Fishery Bulletin, 80(4): ??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/seasonal-
spawning-cycle-longfin-sanddab-citharichthys-xanthostigma-bothidae;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/goldberg%20(2).pdf.

Langton:1982:DOB

[3029] Richard W. Langton. Diet overlap between Atlantic cod, Gadus morhua, silver hake, Merluccius bilinearis, and fifteen other Northwest Atlantic finfish. Fishery Bulletin, 80(4):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/diet-overlap-between-atlantic-cod-gadus-morhua-
silver-hake-merluccius-bilinearis-and;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1982/804/langton.pdf.

Anonymous:1982:MTM

[3030] Anonymous. In memoriam — Thomas A. Manar. Fishery Bulletin, 80 (4):??, 1982. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/memoriam-
thomas-manar;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/inmemoriam_0.pdf.

Robinson:1982:OTS

[3031] Gary R. Robinson. Otter trawl sampling bias of the gill parasite, Lironeca vulgaris (Isopoda, Cymothoidae), from sanddab hosts, Citharichthys spp. Fishery Bulletin, 80(4):??, 1982. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/otter-trawl-sampling-bias-gill-
parasite-lironeca-vulgaris-isopoda-cymothoidae-sanddab-hosts;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/robinson.pdf.

Appeldoorn:1983:VGR

[3032] Richard S. Appeldoorn. Variation in the growth rate of Mya arenaria and its relationship to the environment as analyzed through principal components analysis and the ω parameter of the von Bertalanffy equation. Fishery Bulletin, 81(1):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
variation-growth-rate-mya-arenaria-and-its-relationship-environment-
analyzed-through;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1983/811/appeldoorn.pdf.

Swartzman:1983:IBS

[3033] Gordon L. Swartzman and Robert T. Haar. Interactions between fur seal populations and fisheries in the Bering Sea. Fishery Bulletin, 81(1):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/interactions-between-fur-seal-populations-
and-fisheries-bering-sea;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1983/811/swartzman.pdf.

Shenker:1983:DSR

[3034] Jonathan M. Shenker. Distribution, size relationships, and food habits of juvenile king-of-the-salmon, Trachipterus altivelis, caught off the Oregon coast. Fishery Bulletin, 81(1):??, 1983. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/distribution-size-relationships-
and-food-habits-juvenile-king-salmon-trachipterus-altivelis;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/shenker.pdf.

Bartoo:1983:SAF

[3035] Norman W. Bartoo and Keith R. Parker. Stochastic age-frequency estimation using the von Bertalanffy growth equation. Fishery Bulletin, 81 (1):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/stochastic-
age-frequency-estimation-using-von-bertalanffy-growth-equation;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1983/811/bartoo.pdf.

Dunn:1983:DDY

[3036] Jean R. Dunn. Development and distribution of the young of northern smoothtongue, Leuroglossus schmidti (Bathylagidae), in the northeast Pacific, with comments on the systematics of the genus Leuroglossus Gilbert. Fishery Bulletin, 81(1):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/development-and-distribution-young-northern-smoothtongue-
leuroglossus-schmidti-bathylagidae;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/dunn.pdf.

Bond:1983:NML

[3037] Carl E. Bond, Ting T. Kan, and Katherine W. Myers. Notes on the marine life of the river lamprey, Lampetra ayresi, in Yaquina Bay, Oregon, and the Columbia River estuary. Fishery Bulletin, 81(1):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/notes-marine-
life-river-lamprey-lampetra-ayresi-yaquina-bay-oregon-and-columbia-
river;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/bond.pdf.

Hanan:1983:RAB

[3038] Doyle A. Hanan. Review and analysis of the bluefin tuna, Thunnus thynnus, fishery in the Eastern North Pacific Ocean. Fishery Bulletin, 81(1): ??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/review-and-
analysis-bluefin-tuna-thunnus-thynnus-fishery-eastern-north-
pacific-ocean;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1983/811/hanan.pdf.

Shaklee:1983:BGP

[3039] James B. Shaklee, Richard W. Brill, and Robin Acerra. Biochemical genetics of Pacific blue marlin, Makaira nigricans, from Hawaiian waters. Fishery Bulletin, 81(1):85–90, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/biochemical-genetics-pacific-blue-marlin-makaira-nigricans-
hawaiian-waters;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/shaklee.pdf.

Menz:1983:EES

[3040] Fredric C. Menz and Donald P. Wilton. An economic evaluation of the St. Lawrence River–eastern Lake Ontario bass fishery. Fishery Bulletin, 81(1):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
economic-evaluation-st-lawrence-river-eastern-lake-ontario-
bass-fishery;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/menz.pdf.

Katz:1983:DTL

[3041] S. J. Katz, C. B. Grimes, and K. W. Able. Delineation of tilefish, Lopholatilus chamaeleonticeps, stocks along the United States East Coast and in the Gulf of Mexico. Fishery Bulletin, 81(1):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/delineation-
tilefish-lopholatilus-chamaeleonticeps-stocks-along-united-
states-east-coast;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1983/811/katz.pdf.

Johnson:1983:AGM

[3042] Allyn G. Johnson, William A. Fable, Jr., Mark L. Williams, and Lyman E. Barger. Age, growth, and mortality of king mackerel, Scomberomorus cavalla, from the southeastern United States. Fishery Bulletin, 81 (1):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/age-growth-
and-mortality-king-mackerel-scomberomorus-cavalla-southeastern-
united-states;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/johnson%20(1)_0.pdf.

Langton:1983:FHY

[3043] Richard W. Langton. Food habits of yellowtail flounder, Limanda ferruginea (Storer), from off the Northeastern United States. Fishery Bulletin, 81(1):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/food-habits-yellowtail-flounder-limanda-ferruginea-
storer-northeastern-united-states;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1983/811/langton.pdf.

Durbin:1983:ASG

[3044] Ann Gail Durbin, Edward G. Durbin, Thomas J. Smayda, and Peter G. Verity. Age, size, growth, and chemical composition of Atlantic menhaden, Brevoortia tyrannus, from Narragansett Bay, Rhode Island. Fishery Bulletin, 81(1):133–141, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/age-size-growth-and-chemical-composition-atlantic-menhaden-
brevoortia-tyrannus-narragansett;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/durbin.pdf.

Parsons:1983:RBA

[3045] Glenn R. Parsons. The reproductive biology of the Atlantic sharpnose shark, Rhizoprionodon terraenovae (Richardson). Fishery Bulletin, 81 (1):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/reproductive-
biology-atlantic-sharpnose-shark-rhizoprionodon-terraenovae-
richardson;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1983/811/parsons.pdf.

Vreeland:1983:HFC

[3046] Robert R. Vreeland and Roy J. Wahle. Homing and fisheries contribution of marked Coho salmon, Oncorhynchus kisutch, released at two Columbia River locations. Fishery Bulletin, 81(1): ??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/homing-
and-fisheries-contribution-marked-coho-salmon-oncorhynchus-
kisutch-released-two;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/vreeland.pdf.

Richards:1983:EBI

[3047] R. Anne Richards, J. Stanley Cobb, and Michael J. Fogarty. Effects of behavioral interactions on the catchability of American lobster, Homarus americanus, and two species of Cancer crab. Fishery Bulletin, 81(1):51–60, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/effects-behavioral-interactions-
catchability-american-lobster-homarus-americanus-and-two;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1983/
811/richards.pdf.

Colton:1983:MPB

[3048] Douglas E. Colton and William S. Alevizon. Movement patterns of bonefish, Albula vulpes, in Bahamian waters. Fishery Bulletin, 81 (1):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/movement-
patterns-bonefish-albula-vulpes-bahamian-waters;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
colton_0.pdf.

Smith:1983:CST

[3049] Tim D. Smith. Changes in size of three dolphin (Stenella spp.) populations in the Eastern Tropical Pacific. Fishery Bulletin, 81(1):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/changes-size-three-
dolphin-stenella-spp-populations-eastern-tropical-pacific;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1983/
811/smith.pdf.

Kleppel:1983:AFT

[3050] G. S. Kleppel and E. Manzanilla. Analyses of feeding in two marine copepods from Santa Monica Bay, California. Fishery Bulletin, 81(1):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/analyses-feeding-two-
marine-copepods-santa-monica-bay-california;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/kleppel.
pdf.

Cressey:1983:CSF

[3051] Roger F. Cressey, Bruce B. Collette, and Joseph L. Russo. Copepods and scombrid fishes: a study in host–parasite relationships. Fishery Bulletin, 81(2):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/copepods-
and-scombrid-fishes-study-host-parasite-relationships;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1983/812/
cressey.pdf.

Robison:1983:MFE

[3052] Bruce H. Robison and James E. Craddock. Mesopelagic fishes eaten by Fraser’s dolphin, Lagenodelphis hosei. Fishery Bulletin, 81(2):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/mesopelagic-
fishes-eaten-frasers-dolphin-lagenodelphis-hosei;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
robison.pdf.

Moffitt:1983:HLM

[3053] Robert B. Moffitt. Heterocarpus longirostris MacGilchrist from the Northern Mariana Islands. Fishery Bulletin, 81(2):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/heterocarpus-longirostris-macgilchrist-
northern-mariana-islands;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/moffitt.pdf.

Durbin:1983:ENB

[3054] Edward G. Durbin and Ann G. Durbin. Energy and nitrogen budgets for the Atlantic menhaden, Brevoortia tyrannus (Pisces: Clupeidae), a filter-feeding planktivore. Fishery Bulletin, 81(2): ??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/energy-
and-nitrogen-budgets-atlantic-menhaden-brevoortia-tyrannus-
pisces-clupeidae-filter;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1983/812/durbin.pdf.

Gadbois:1983:SPB

[3055] Donald F. Gadbois and Richard S. Maney. Survey of polychlorinated biphenyls in selected finfish species from United States coastal waters. Fishery Bulletin, 81(2):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/survey-polychlorinated-biphenyls-selected-
finfish-species-united-states-coastal-waters;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/gadbois_
0.pdf.

Gilmore:1983:RED

[3056] R. Grant Gilmore, Jon W. Dodrill, and Patricia A. Linley. Reproduction and embryonic development of the sand tiger shark, Odontaspis taurus (Rafinesque). Fishery Bulletin, 81(2):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/reproduction-and-embryonic-development-sand-
tiger-shark-odontaspis-taurus-rafinesque;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1983/812/gilmore.pdf.

Divita:1983:FCF

[3057] Regina Divita, Mischelle Creel, and Peter F. Sheridan. Foods of coastal fishes during brown shrimp, Penaeus aztecus, migration from Texas estuaries (June–July 1981). Fishery Bulletin, 81(2):396–404, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/foods-coastal-
fishes-during-brown-shrimp-penaeus-aztecus-migration-texas-
estuaries-june;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/divita.pdf.

Howell:1983:SCO

[3058] W. Huntiing Howell. Seasonal changes in the ovaries of adult yellowtail flounder, Limanda ferruginea. Fishery Bulletin, 81(2):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/seasonal-changes-ovaries-adult-
yellowtail-flounder-limanda-ferruginea;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1983/812/howell.pdf.

Lewis:1983:OSL

[3059] Robert M. Lewis and Mayo H. Judy. The occurrence of spot, Leiostomus xanthurus, and Atlantic croaker, Micropogonias undulatus, larvae in Onslow Bay and Newport River estuary, North Carolina. Fishery Bulletin, 81(2):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/occurrence-spot-leiostomus-xanthurus-
and-atlantic-croaker-micropogonias-undulatus-larvae;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
lewis%20(1).pdf.

Johnson:1983:MDH

[3060] James R. Johnson and Joseph G. Loesch. Morphology and development of hatchery-cultured American shad, Alosa sapidissima (Wilson). Fishery Bulletin, 81(2):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
morphology-and-development-hatchery-cultured-american-shad-
alosa-sapidissima-wilson;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1983/812/johnson.pdf.

Wespestad:1983:MSA

[3061] Vidar G. Wespestad, Kenneth Thorsen, and Sally A. Mizroch. Movement of sablefish, Anoplopoma fimbria, in the northeastern Pacific Ocean as determined by tagging experiments (1971–80). Fishery Bulletin, 81(2):415–420, 1983. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
movement-sablefish-anoplopoma-fimbria-northeastern-pacific-
ocean-determined-tagging;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/wespestad.pdf.

Beacham:1983:VMS

[3062] Terry D. Beacham. Variability in median size and age at sexual maturity of Atlantic cod, Gadus morhua, on the Scotian Shelf in the Northwest Atlantic Ocean. Fishery Bulletin, 81(2):??, 1983. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/variability-median-size-and-age-
sexual-maturity-atlantic-cod-gadus-morhua-scotian-shelf;   https:

//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1983/812/
beacham.pdf.

Quinn:1983:LTV

[3063] William H. Quinn and Victor T. Neal. Long-term variations in the Southern Oscillation, El Niño, and Chilean subtropical rainfall. Fishery Bulletin, 81(2):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/long-term-variations-southern-oscillation-
el-nino-and-chilean-subtropical-rainfall;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1983/812/quinn.pdf.

Dudley:1983:WAS

[3064] Richard G. Dudley and T. Glenn McGahee. Winter and altered spring movements of striped bass in the Savannah River, Georgia. Fishery Bulletin, 81(2):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
winter-and-altered-spring-movements-striped-bass-savannah-river-
georgia;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/dudley%20(1).pdf.

Brown:1983:AMF

[3065] Robin F. Brown and Bruce R. Mate. Abundance, movements, and feeding habits of harbor seals, Phoca vitulina, at Netarts and Tillamook Bays, Oregon. Fishery Bulletin, 81(2):??, 1983. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/abundance-movements-
and-feeding-habits-harbor-seals-phoca-vitulina-netarts-and-
tillamook;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1983/812/brown.pdf.

Reilly:1983:PAG

[3066] Stephen B. Reilly, Dale W. Rice, and Allen A. Wolman. Population assessment of the gray whale, Eschrichtius robustus, from California Shore Censuses, 1967–80. Fishery Bulletin, 81(2):??, 1983. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/population-assessment-gray-whale-
eschrichtius-robustus-california-shore-censuses-1967-80;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1983/812/
reilly.pdf.

Whoriskey:1983:IFR

[3067] F. G. Whoriskey, Jr. Intertidal feeding and refuging by cunners, Tautogolabrus adspersus (Labridae). Fishery Bulletin, 81(2):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/intertidal-feeding-and-refuging-
cunners-tautogolabrus-adspersus-labridae;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/whoriskey%
20(1).pdf.

Campbell:1983:SMF

[3068] A. Campbell and M. D. Eagles. Size at maturity and fecundity of rock crabs, Cancer irroratus, from the Bay of Fundy and Southwestern Nova Scotia. Fishery Bulletin, 81(2):??, 1983. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/size-maturity-and-fecundity-rock-
crabs-cancer-irroratus-bay-fundy-and-southwestern-nova;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1983/812/
campell.pdf.

Venrick:1983:PSP

[3069] E. L. Venrick. Percent similarity: the prediction of bias. Fishery Bulletin, 81(2):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/percent-
similarity-prediction-bias;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1983/812/venrick.pdf.

Peterson:1983:RET

[3070] Charles H. Peterson, Henry C. Summerson, and Stephen R. Fegley. Relative efficiency of two clam rakes and their contrasting impacts on seagrass biomass. Fishery Bulletin, 81(2):429–434, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/relative-efficiency-
two-clam-rakes-and-their-contrasting-impacts-seagrass-biomass;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/peterson_0.pdf.

Ralston:1983:AWD

[3071] Stephen Ralston and Garret T. Miyamoto. Analyzing the width of daily otolith increments to age the Hawaiian snapper, Pristipomoides filamentosus. Fishery Bulletin, 81(3):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/analyzing-width-daily-otolith-increments-age-hawaiian-

snapper-pristipomoides-filamentosus;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1983/813/ralston.pdf.

MacCall:1983:PPI

[3072] Alec D. MacCall, Keith R. Parker, Ronald Leithiser, and Bill Jessee. Power plant impact assessment: a simple fishery production model approach. Fishery Bulletin, 81(3):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/power-plant-impact-assessment-simple-
fishery-production-model-approach;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/maccall.pdf.

Grimes:1983:TTD

[3073] C. B. Grimes, S. C. Turner, and K. W. Able. A technique for tagging deepwater fish. Fishery Bulletin, 81(3):663–666, 1983. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/technique-tagging-deepwater-fish;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/grimes.pdf.

Ross:1983:RBB

[3074] Jeffrey L. Ross and John V. Merriner. Reproductive biology of the blueline tilefish, Caulolatilus microps, off North Carolina and South Carolina. Fishery Bulletin, 81(3):??, 1983. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/reproductive-biology-
blueline-tilefish-caulolatilus-microps-north-carolina-and-south;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1983/813/ross.pdf.

Somerton:1983:SSM

[3075] David A. Somerton and Richard A. MacIntosh. The size at sexual maturity of blue king crab, Paralithodes platypus, in Alaska. Fishery Bulletin, 81(3):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/size-
sexual-maturity-blue-king-crab-paralithodes-platypus-alaska;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/somerton_2.pdf.

Rothlisberg:1983:FAD

[3076] Peter C. Rothlisberg and Charles B. Miller. Factors affecting the distribution, abundance, and survival of Pandalus jordani (Decapoda, Pandalidae) larvae off the Oregon coast. Fishery Bulletin, 81(3):

??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/factors-
affecting-distribution-abundance-and-survival-pandalus-jordani-
decapoda-pandalidae;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1983/813/rothlisberg.pdf.

Livingston:1983:FHP

[3077] Patricia A. Livingston. Food habits of Pacific whiting, Merluccius productus, off the west coast of North America, 1967 and 1980. Fishery Bulletin, 81(3):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/food-
habits-pacific-whiting-merluccius-productus-west-coast-north-
america-1967-and-1980;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/livingston.pdf.

Allen:1983:TSP

[3078] Larry G. Allen and Edward E. Demartini. Temporal and spatial patterns of nearshore distribution and abundance of the pelagic fishes off San Onofre–Oceanside, California. Fishery Bulletin, 81(3):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/temporal-and-spatial-
patterns-nearshore-distribution-and-abundance-pelagic-fishes-
san;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1983/813/allen.pdf.

Wenner:1983:SAD

[3079] Charles A. Wenner. Species associations and day–night variability of trawl-caught fishes from the inshore sponge-coral habitat, South Atlantic Bight. Fishery Bulletin, 81(3):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/species-associations-and-day-night-variability-
trawl-caught-fishes-inshore-sponge-coral;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1983/813/wenner.pdf.

Clausen:1983:FWP

[3080] David M. Clausen. Food of walleye pollock, Theragra chalcogramma, in an embayment of southeastern Alaska. Fishery Bulletin, 81(3):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/food-walleye-pollock-
theragra-chalcogramma-embayment-southeastern-alaska;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
clausen%20(1).pdf.

Durbin:1983:SCS

[3081] E. G. Durbin, A. G. Durbin, R. W. Langton, and R. E. Bowman. Stomach contents of silver hake, Merluccius bilinearis, and Atlantic cod, Gadus morhua, and estimation of their daily rations. Fishery Bulletin, 81(3):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/stomach-contents-silver-hake-merluccius-bilinearis-
and-atlantic-cod-gadus-morhua-and;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1983/813/durbin.pdf.

Williams:1983:RGO

[3082] Vern R. Williams and Thomas A. Clarke. Reproduction, growth, and other aspects of the biology of the gold spot herring, Herklotsichthys quadrimaculatus (Clupeidae), a recent introduction to Hawaii. Fishery Bulletin, 81(3):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/reproduction-growth-and-other-aspects-biology-gold-
spot-herring-herklotsichthys;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1983/813/williams.pdf.

Sheridan:1983:SFT

[3083] Peter F. Sheridan and David L. Trimm. Summer foods of Texas coastal fishes relative to age and habitat. Fishery Bulletin, 81(3):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/summer-foods-texas-
coastal-fishes-relative-age-and-habitat;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/sheridan.
pdf.

Oliver:1983:WOR

[3084] John S. Oliver, Peter N. Slattery, Edmund F. O’Connor, and Lloyd F. Lowry. Walrus, Odobenus rosmarus, feeding in the Bering Sea: a benthic perspective. Fishery Bulletin, 81(3):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/walrus-odobenus-rosmarus-
feeding-bering-sea-benthic-perspective;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1983/813/oliver.pdf.

Winkler:1983:EBP

[3085] Delaine L. Winkler, Keith L. Duncan, Jo Ellen Hose, and Harold W. Puffer. Effects of benzo(a)pyrene on the early development of California grunion, Leuresthes tenuis (Pisces, Atherinidae). Fishery Bulletin, 81 (3):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/effects-
benzoapyrene-early-development-california-grunion-leuresthes-
tenuis-pisces;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1983/813/winkler.pdf.

Daniels:1983:LHS

[3086] Robert A. Daniels and Peter B. Moyle. Life history of splittail (Cyprinidae: Pogonichthys macrolepidotus) in the Sacramento–San Joaquin estuary. Fishery Bulletin, 81(3):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/life-history-splittail-cyprinidae-pogonichthys-
macrolepidotus-sacramento-san-joaquin;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/daniels.pdf.

Oliver:1983:CGW

[3087] John S. Oliver, Peter N. Slaitery, Mark A. Silberstein, and Edmund F. O’Connor. A comparison of gray whale, Eschrichtius robustus, feeding in the Bering Sea and Baja California. Fishery Bulletin, 81(3):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
comparison-gray-whale-eschrichtius-robustus-feeding-bering-
sea-and-baja-california;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1983/813/oliver1.pdf.

McCabe:1983:SHJ

[3088] George T. McCabe, Jr., Clifford W. Long, and Steve L. Leek. Survival and homing of juvenile Coho salmon, Oncorhynchus kisutch, transported by barge. Fishery Bulletin, 81(3):412–415, 1983. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/survival-and-homing-
juvenile-coho-salmon-oncorhynchus-kisutch-transported-barge;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/mccabe.pdf.

Hunte:1983:LHE

[3089] Wayne Hunte and Robin Mahon. Life history and exploitation of Macrobrachium faustinum in a tropical high-gradient river. Fishery Bulletin, 81(3):654–660, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/life-
history-and-exploitation-macrobrachium-faustinum-tropical-high-
gradient-river;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/hunte.pdf.

Power:1983:SNA

[3090] James H. Power and James D. McCleave. Simulation of the North Atlantic Ocean Drift of Anguilla leptocephali. Fishery Bulletin, 81(3):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/simulation-
north-atlantic-ocean-drift-anguilla-leptocephali;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1983/813/
power.pdf.

Bowering:1983:AGS

[3091] W. R. Bowering. Age, growth, and sexual maturity of Greenland halibut, Reinhardtius hippoglossoides (Walbaum), in the Canadian Northwest Atlantic. Fishery Bulletin, 81(3):??, 1983. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/age-growth-and-sexual-maturity-
greenland-halibut-reinhardtius-hippoglossoides-walbaum;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
bowering.pdf.

Smith:1983:ICH

[3092] G. J. D. Smith, A. J. Read, and D. E. Gaskins. Incidental catch of harbor porpoise, Phocoena phocoena (L.), in herring weirs in Charlotte County, New Brunswick, Canada. Fishery Bulletin, 81(3):660–662, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/incidental-catch-harbor-porpoise-phocoena-phocoena-
l-herring-weirs-charlotte-county-new;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/smith_0.pdf.

Bolz:1983:GLA

[3093] George R. Bolz and R. Gregory Lough. Growth of larval Atlantic cod, Gadus morhua, and haddock, Melanogrammus aeglefinus, on Georges Bank, spring 1981. Fishery Bulletin, 81(4):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/growth-
larval-atlantic-cod-gadus-morhua-and-haddock-melanogrammus-
aeglefinus-georges-bank;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1983/814/bolz.pdf.

Sherman:1983:CZL

[3094] K. Sherman, J. R. Green, J. R. Goulet, and L. Ejsymont. Coherence in zooplankton of a large Northwest Atlantic ecosystem. Fishery Bulletin, 81(4):855–862, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/coherence-zooplankton-large-northwest-atlantic-
ecosystem;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1983/814/sherman.pdf.

Watson:1983:RLP

[3095] William Watson. Redescription of larvae of the pigfish, Orthopristis chrysoptera Linnaeus (Pisces, Haemulidae). Fishery Bulletin, 81(4):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/redescription-larvae-
pigfish-orthopristis-chrysoptera-linnaeus-pisces-haemulidae;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/watson_0.pdf.

Oxenford:1983:AGD

[3096] Hazel A. Oxenford and Wayne Hunte. Age and growth of dolphin, Coryphaena hippurus, as determined by growth rings in otoliths. Fishery Bulletin, 81(4):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
age-and-growth-dolphin-coryphaena-hippurus-determined-growth-
rings-otoliths;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/oxenford.pdf.

Brousseau:1983:ARB

[3097] Diane J. Brousseau. Aspects of reproduction of the blue mussel, Mytilus edulis (Pelecypoda: Mytilidae) in Long Island Sound. Fishery Bulletin, 81(4):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
aspects-reproduction-blue-mussel-mytilus-edulis-pelecypoda-
mytilidae-long-island-sound;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1983/814/brousseau.pdf.

Trent:1983:SSR

[3098] Lee Trent, Roy O. Williams, Ronald G. Taylor, Carl H. Saloman, and Charles H. Manooch III. Size, sex ratio, and recruitment in various fisheries of king mackerel, Scomberomorus cavalla, in the Southeastern United States. Fishery Bulletin, 81(4):709–721, 1983. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/size-sex-ratio-and-recruitment-
various-fisheries-king-mackerel-scomberomorus-cavalla;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1983/814/
trent.pdf.

Williams:1983:MCP

[3099] Austin B. Williams. The mud crab, Panopeus herbstii, s.l. partition into six species (Decapoda: Xanthidae). Fishery Bulletin, 81 (4):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/mud-crab-
panopeus-herbstii-sl-partition-six-species-decapoda-xanthidae;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/williams%20(1)_0.pdf.

Kraus:1983:CAS

[3100] Scott D. Kraus, James R. Gilbert, and John H. Prescott. A comparison of aerial, shipboard, and land-based survey methodology for the harbor porpoise, Phocoena phocoena. Fishery Bulletin, 81(4):910–913, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/comparison-
aerial-shipboard-and-land-based-survey-methodology-harbor-porpoise-
phocoena;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/kraus.pdf.

Fritz:1983:HCM

[3101] Lowell W. Fritz and Dexter S. Haven. Hard clam, Mercenaria mercenaria: Shell growth patterns in Chesapeake Bay. Fishery Bulletin, 81 (4):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/hard-clam-
mercenaria-mercenaria-shell-growth-patterns-chesapeake-bay;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1983/814/fritz.pdf.

Turner:1983:GMA

[3102] S. C. Turner, C. B. Grimes, and K. W. Able. Growth, mortality, and age/ size structure of the fisheries for tilefish, Lopholatilus chamaelonticeps, in the Middle Atlantic–southern New England Region. Fishery Bulletin, 81 (4):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/growth-
mortality-and-agesize-structure-fisheries-tilefish-lopholatilus-
chamaelonticeps;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1983/814/turner.pdf.

Sullivan:1983:EAH

[3103] Bolling Sullivan, Katie Miller, Kathleen Singleton, Anthony G. Scheer, and Austin B. Williams. Electrophoretic analyses of hemocyanins from four species of mud crabs, genus Panopeus, with observations on the ecology of P. obesus. Fishery Bulletin, 81(4):??, 1983. CODEN FSYBAY.

ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/electrophoretic-analyses-hemocyanins-
four-species-mud-crabs-genus-panopeus-observations;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
sullivan.pdf.

Itzkowitz:1983:TFD

[3104] Norman Itzkowitz and J. R. Schubel. Tolerance of five-day-old winter flounder, Pseudopleuronectus americanus, larvae to thermal shock. Fishery Bulletin, 81(4):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
tolerance-five-day-old-winter-flounder-pseudopleuronectus-americanus-
larvae-thermal-shock;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/itzkowitz.pdf.

Fry:1983:FSM

[3105] Brian Fry. Fish and shrimp migrations in the Northern Gulf of Mexico analyzed using stable C, N, and S isotope ratios. Fishery Bulletin, 81(4):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fish-and-
shrimp-migrations-northern-gulf-mexico-analyzed-using-stable-
c-n-and-s-isotope;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1983/814/fry.pdf.

Peterson:1983:MRT

[3106] C. H. Peterson, P. B. Duncan, H. C. Summerson, and G. W. Safrit, Jr. A mark-recapture test of annual periodicity of internal growth band deposition in shells of hard clams, Mercenaria mercenaria, from a population along the southeastern United States. Fishery Bulletin, 81(4):765–779, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/mark-recapture-test-
annual-periodicity-internal-growth-band-deposition-shells-hard-
clams;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/peterson%20(1).pdf.

Reames:1983:MCP

[3107] Robert C. Reames and Austin B. Williams. Mud crabs of the Panopeus herbstii H. M. Edw., s.l., complex in Alabama, U.S.A. Fishery Bulletin, 81(4):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
mud-crabs-panopeus-herbstii-h-m-edw-sl-complex-alabama-usa;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/reames.pdf.

Mathews:1983:MRS

[3108] Stephen B. Mathews and Morris W. Barker. Movements of rockfish (Sebastes) tagged in northern Puget Sound, Washington. Fishery Bulletin, 81(4):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/movements-
rockfish-sebastes-tagged-northern-puget-sound-washington;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/mathews.pdf.

Grant:1983:BGP

[3109] W. Stewart Grant, Richard Bakkala, Fred M. Utter, David J. Teel, and Tokimasa Kobayashi. Biochemical genetic population structure of yellowfin sole, Limanda aspera, of the North Pacific Ocean and Bering Sea. Fishery Bulletin, 81(4):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/biochemical-genetic-population-structure-yellowfin-
sole-limanda-aspera-north-pacific-ocean;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1983/814/grant.pdf.

Walsh:1983:EDL

[3110] William A. Walsh and William A. Lund, Jr. Early development of the longhorn sculpin, Myoxocephalus octodecemspinosus. Fishery Bulletin, 81(4):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/early-
development-longhorn-sculpin-myoxocephalus-octodecemspinosus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/walsh.pdf.

Haynes:1983:ETR

[3111] Evan B. Haynes and Steve E. Ignell. Effect of temperature on rate of embryonic development of walleye pollock, Theragra chalcogramma. Fishery Bulletin, 81(4):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
effect-temperature-rate-embryonic-development-walleye-pollock-
theragra-chalcogramma;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/haynes%20(4).pdf.

Anonymous:1983:I

[3112] Anonymous. Index. Fishery Bulletin, 81(4):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/index-110;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/index%
20(4)_0.pdf.

Huntsman:1983:YPR

[3113] Gene R. Huntsman, Charles S. Manooch III, and Churchill B. Grimes. Yield per recruit models of some reef fishes of the U.S. South Atlantic Bight. Fishery Bulletin, 81(4):679–695, 1983. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/yield-recruit-models-some-reef-
fishes-us-south-atlantic-bight;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1983/814/huntsman.pdf.

Shepherd:1983:GHV

[3114] Gary Shepherd and Churchill B. Grimes. Geographic and historic variations in growth of weakfish, Cynoscion regalis, in the Middle Atlantic Bight. Fishery Bulletin, 81(4):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/geographic-and-historic-variations-growth-
weakfish-cynoscion-regalis-middle-atlantic-bight;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
shepherd.pdf.

Govoni:1983:HPT

[3115] John J. Govoni. Helminth parasitism of three larval fishes in the northern Gulf of Mexico. Fishery Bulletin, 81(4):??, 1983. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/helminth-parasitism-three-larval-
fishes-northern-gulf-mexico;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/govoni.pdf.

Majkowski:1983:ECL

[3116] J. Majkowski and J. Hampton. The estimation of a catch level which stabilizes the parental biomass of an exploited fish stock. Fishery Bulletin, 81(4):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
estimation-catch-level-which-stabilizes-parental-biomass-exploited-
fish-stock;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1983/814/majkowski.pdf.

McCabe:1983:IBJ

[3117] George T. McCabe, Jr., William D. Muir, Robert L. Emmett, and Joseph T. Durkin. Interrelationships between juvenile salmonids and nonsalmonid fish in the Columbia River estuary. Fishery Bulletin, 81(4):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
interrelationships-between-juvenile-salmonids-and-nonsalmonid-
fish-columbia-river-estuary;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/mccabe%20(1).pdf.

Hoenig:1983:EUL

[3118] John M. Hoenig. Empirical use of longevity data to estimate mortality rates. Fishery Bulletin, 81(4):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/empirical-use-longevity-data-estimate-mortality-
rates;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/hoenig.pdf.

Methot:1983:SVS

[3119] Richard D. Methot, Jr. Seasonal variation in survival of larval northern anchovy, Engraulis mordax, estimated from the age distribution of juveniles. Fishery Bulletin, 81(4):??, 1983. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/seasonal-variation-
survival-larval-northern-anchovy-engraulis-mordax-estimated-
age;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1983/814/methot.pdf.

Stoner:1983:DFS

[3120] Allan W. Stoner. Distribution of fishes in seagrass meadows: Role of macrophyte biomass and species composition. Fishery Bulletin, 81(4):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/distribution-fishes-
seagrass-meadows-role-macrophyte-biomass-and-species-composition;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/stoner%20(1).pdf.

VanHeukelem:1983:GGQ

[3121] W. Van Heukelem, M. C. Christman, C. E. Epifanio, and S. D. Sulkin. Growth of Geryon quinquedens (Brachyura: Geryonidae) juveniles in the laboratory. Fishery Bulletin, 81(4):??, 1983. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/growth-geryon-quinquedens-brachyura-geryonidae-
juveniles-laboratory;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/vanheukelem.pdf.

Barnett:1984:DIS

[3122] Arthur M. Barnett, Andrew E. Jahn, Peter D. Sertic, and William Watson. Distribution of ichthyoplankton off San Onofre, California, and methods for sampling very shallow coastal waters. Fishery Bulletin, 82(1):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/distribution-ichthyoplankton-san-onofre-california-
and-methods-sampling-very-shallow;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/barnett.pdf.

Ropes:1984:DAG

[3123] John W. Ropes, Douglas S. Jones, Steven A. Murawski, Fredric M. Serchuk, and Ambrose Jearld, Jr. Documentation of annual growth lines in ocean quahogs, Arctica islandica Linne. Fishery Bulletin, 82(1):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/documentation-annual-
growth-lines-ocean-quahogs-arctica-islandica-linne;   https:
//spo.nmfs.noaa.gov/sites/default/files/ropes.pdf.

Bowman:1984:FSH

[3124] Ray E. Bowman. Food of silver hake, Merluccius bilinearis. Fishery Bulletin, 82(1):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
food-silver-hake-merluccius-bilinearis;   https://spo.nmfs.
noaa.gov/sites/default/files/bowman.pdf.

Larson:1984:AVD

[3125] Ralph J. Larson and Edward E. DeMartini. Abundance and vertical distribution of fishes in a cobble-bottom kelp forest off San Onofre, California. Fishery Bulletin, 82(1):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/abundance-and-vertical-distribution-fishes-cobble-bottom-
kelp-forest-san-onofre-california;   https://spo.nmfs.noaa.
gov/sites/default/files/larson.pdf.

Coyer:1984:IAA

[3126] James A. Coyer. The invertebrate assemblage associated with the giant kelp, Macrocystis pyrifera, at Santa Catalina Island, California: a general description with emphasis on amphipods, copepods, mysids, and shrimps. Fishery Bulletin, 82(1):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/invertebrate-assemblage-associated-giant-kelp-macrocystis-

pyrifera-santa-catalina-island;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1984/821/coyer.pdf.

Antonelis:1984:SSP

[3127] George A. Antonelis, Jr., Clifford H. Fiscus, and Robert L. Delong. Spring and summer prey of California sea lions, Zalophus caufornianus, at San Miguel Island, California, 1978–79. Fishery Bulletin, 82(1):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/spring-and-
summer-prey-california-sea-lions-zalophus-caufornianus-san-
miguel-island;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1984/821/antonelis.pdf.

Griswold:1984:LDS

[3128] Carolyn A. Griswold and Thomas W. McKenney. Larval development of the scup, Stenotomus chrysops (Pisces: Sparidae). Fishery Bulletin, 82(1):77–84, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/larval-development-scup-stenotomus-chrysops-pisces-
sparidae;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1984/821/griswold.pdf.

Hettler:1984:DEL

[3129] William F. Hettler. Description of eggs, larvae, and early juveniles of Gulf menhaden, Breuoortia patronus, and comparisons with Atlantic menhaden, B. tryannus, and yellowfin menhaden, B. smithi. Fishery Bulletin, 82(1):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
description-eggs-larvae-and-early-juveniles-gulf-menhaden-breuoortia-
patronus-and;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/hettler.pdf.

McGurk:1984:RDO

[3130] Michael D. McGurk. Ring deposition in the otoliths of larval Pacific herring, Clupea harengus pallasi. Fishery Bulletin, 82(1): ??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/ring-
deposition-otoliths-larval-pacific-herring-clupea-harengus-
pallasi;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/mcgurk.pdf.

Macdonald:1984:FFA

[3131] J. Stevenson Macdonald, Michael J. Dadswell, Ralph G. Appy, Gary D. Melvin, and David A. Methven. Fishes, fish assemblages, and their seasonal movements in the Lower Bay of Fundy and Passamaquoddy Bay, Canada. Fishery Bulletin, 82(1):121–139, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/fishes-fish-
assemblages-and-their-seasonal-movements-lower-bay-fundy-and-
passamaquoddy-bay;   https://spo.nmfs.noaa.gov/sites/default/
files/macdonald.pdf.

Tilseth:1984:DDL

[3132] S. Tilseth and B. Ellertsen. The detection and distribution of larval Arcto–Norwegian cod, Gadus morhua food organisms by an in situ particle counter. Fishery Bulletin, 82(1):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/detection-and-distribution-larval-arcto-norwegian-
cod-gadus-morhua-food-organisms-situ;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1984/821/tilseth.pdf.

Ewing:1984:EST

[3133] R. D. Ewing, C. E. Hart, C. A. Fustich, and Greg Concannon. Effects of size and time of release on seaward migrations of spring Chinook salmon, Oncorhynchus tshawytscha. Fishery Bulletin, 82(1):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/effects-size-
and-time-release-seaward-migrations-spring-chinook-salmon-oncorhynchus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/ewing.pdf.

Campana:1984:IEA

[3134] Steven E. Campana. Interactive effects of age and environmental modifiers on the production of daily growth increments in otoliths of plainfin midshipman, Porichthys notatus. Fishery Bulletin, 82(1):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/interactive-effects-
age-and-environmental-modifiers-production-daily-growth-increments;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/campana.pdf.

Love:1984:ALH

[3135] Milton S. Love, Gerald E. McGowen, William Westphal, Robert J. Lavenberg, and Linda Martin. Aspects of the life history and fishery of the white croaker, Genyonemus lineatus (Sciaenidae), off California. Fishery Bulletin, 82(1):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/aspects-life-history-and-fishery-white-croaker-
genyonemus-lineatus-sciaenidae-california.

Morris:1984:FHB

[3136] Pamela A. Morris. Feeding habits of blacksmith, Chromis punctipinnis, associated with a thermal outfall. Fishery Bulletin, 82(1): ??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/feeding-
habits-blacksmith-chromis-punctipinnis-associated-thermal-outfall;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/morris.pdf.

Myrick:1984:CDL

[3137] Albert C. Myrick, Jr., Edward W. Shallenberger, Ingrid Kang, and David B. Mackay. Calibration of dental layers in seven captive Hawaiian spinner dolphins, Stenella longirostris, based on tetracycline labeling. Fishery Bulletin, 82(1):??, 1984. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/calibration-dental-layers-seven-
captive-hawaiian-spinner-dolphins-stenella-longirostris;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1984/821/
love.pdf.

Schmitt:1984:MGI

[3138] P. D. Schmitt. Marking growth increments in otoliths of larval and juvenile fish by immersion in tetracycline to examine the rate of increment formation. Fishery Bulletin, 82(1):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/marking-growth-increments-otoliths-larval-
and-juvenile-fish-immersion-tetracycline-examine;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
schmitt.pdf.

Ross:1984:RBD

[3139] Steve W. Ross. Reproduction of the banded drum, Larimus fasciatus, in North Carolina. Fishery Bulletin, 82(1):??, 1984. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/reproduction-banded-drum-larimus-
fasciatus-north-carolina;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1984/821/mcgurk.pdf.

Ennis:1984:TRV

[3140] G. P. Ennis. Tag-recapture validation of molt and egg extrusion predictions based upon pleopod examination in the American lobster, Homarus americanus. Fishery Bulletin, 82(1):??, 1984. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/tag-recapture-validation-
molt-and-egg-extrusion-predictions-based-upon-pleopod-examination;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/ennis.pdf.

Ennis:1984:CPF

[3141] G. P. Ennis. Comparison of physiological and functional sizematurity relationships in two Newfoundland populations of lobsters Homarus americanus. Fishery Bulletin, 82(1):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/comparison-physiological-
and-functional-size-maturity-relationships-two-newfoundland;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/ennis2.pdf.

Echeverria:1984:CBT

[3142] Tina Echeverria and William H. Lenarz. Conversions between total, fork, and standard lengths in 35 species of Sebastes from California. Fishery Bulletin, 82(1):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
conversions-between-total-fork-and-standard-lengths-35-species-
sebastes-california;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/echeverria.pdf.

Pearcy:1984:DVF

[3143] W. Pearcy, T. Nishiyama, T. Fujii, and K. Masuda. Diel variations in the feeding habits of Pacific salmon caught in gill nets during a 24-hour period in the Gulf of Alaska. Fishery Bulletin, 82(2): 391–399, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/diel-
variations-feeding-habits-pacific-salmon-caught-gill-nets-during-
24-hour-period-gulf;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/pearcy.pdf.

Haynes:1984:EZS

[3144] Evan B. Haynes. Early zoeal stages of Placetron wosnessenskii and Rhinolithodes wosnessenskii (Decapoda, Anomura, Lithodidae) and review of lithodid larvae of the northern North Pacific Ocean. Fishery Bulletin, 82(2):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/early-zoeal-stages-placetron-wosnessenskii-and-
rhinolithodes-wosnessenskii-decapoda-anomura;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/haynes.
pdf.

Goldberg:1984:PFH

[3145] Stephen R. Goldberg, Victor Hugo Alarcon, and Juergen Alheit. Postovulatory follicle histology of the Pacific sardine, Sardinops sagax from Peru. Fishery Bulletin, 82(2):443–445, 1984. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/postovulatory-follicle-histology-
pacific-sardine-sardinops-sagax-peru;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/goldberg%20(3)
.pdf.

Ropes:1984:SAS

[3146] John W. Ropes, Steven A. Murawski, and Fredric M. Serchuk. Size, age, sexual maturity; and sex ratio in ocean quahogs, Arctica islandica Linne, off Long Island, New York. Fishery Bulletin, 82(2):??, 1984. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/size-age-sexual-maturity-
and-sex-ratio-ocean-quahogs-arctica-islandica-linne-long-island;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/ropes.pdf.

Buck:1984:QQB

[3147] John D. Buck. Quantitative and qualitative bacteriology of elasmobranch fish from the Gulf of Mexico. Fishery Bulletin, 82(2):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/quantitative-
and-qualitative-bacteriology-elasmobranch-fish-gulf-mexico;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/buck.pdf.

Shimek:1984:NSP

[3148] Ronald L. Shimek, David Fyfe, Leah Ramsey, Anne Bergey, Joel Elliott, and Stewart Guy. A note on spawning of the Pacific market squid, Loligo opalescens (Berry, 1911), in the Barkley Sound region, Vancouver Island, Canada. Fishery Bulletin, 82 (2):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/note-
spawning-pacific-market-squid-loligo-opalescens-berry-1911-
barkley-sound-region;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/shimek.pdf.

Ruggerone:1984:ACP

[3149] Gregory T. Ruggerone and Donald E. Rogers. Arctic char predation on sockeye salmon smolts at Little Togiak River, Alaska. Fishery Bulletin, 82 (2):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/arctic-char-
predation-sockeye-salmon-smolts-little-togiak-river-alaska;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/ruggerone.pdf.

Botton:1984:DFH

[3150] Mark L. Botton and Harold H. Haskin. Distribution and feeding of the horseshoe crab, Limulus polyphemus, on the continental shelf off New Jersey. Fishery Bulletin, 82(2):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/distribution-and-feeding-
horseshoe-crab-limulus-polyphemus-continental-shelf-new-jersey;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/botton.pdf.

Epperly:1984:AVE

[3151] Sheryan P. Epperly and Walter R. Nelson. Arithmetic versus exponential calculation of mean biomass. Fishery Bulletin, 82(2):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/arithmetic-
versus-exponential-calculation-mean-biomass;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/epperly.
pdf.

Standard:1984:RMP

[3152] Gary W. Standard and Mark E. Chittenden, Jr. Reproduction, movements, and population dynamics of the banded drum, Larimus fasciatus, in the Gulf of Mexico. Fishery Bulletin, 82(2):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/reproduction-
movements-and-population-dynamics-banded-drum-larimus-fasciatus-
gulf-mexico;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/standard.pdf.

Maule:1984:FEW

[3153] Alec G. Maule and Howard F. Horton. Feeding ecology of walleye, Stizostedion vitreum vitreum, in the mid-Columbia River, with emphasis on the interactions between walleye and juvenile anadromous fishes. Fishery Bulletin, 82(2):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/feeding-ecology-walleye-stizostedion-
vitreum-vitreum-mid-columbia-river-emphasis;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/maule.
pdf.

Tettey:1984:IIU

[3154] Ernest Tettey, Christopher Pardy, Wade Griffin, and A. Nelson Swartz. Implications of investing under different economic conditions on the profitability of Gulf of Mexico shrimp vessels operating out of Texas. Fishery Bulletin, 82(2):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
implications-investing-under-different-economic-conditions-
profitability-gulf-mexico-shrimp;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/tettey.pdf.

Rockette:1984:BDS

[3155] Mark D. Rockette, Gary W. Standard, and Mark E. Chittenden, Jr. Bathymetric distribution, spawning periodicity, sex ratios, and size compositions of the mantis shrimp, Squilla empusa, in the northwestern Gulf of Mexico. Fishery Bulletin, 82(2):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/bathymetric-distribution-
spawning-periodicity-sex-ratios-and-size-compositions-mantis;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/rockett.pdf.

Minello:1984:SVH

[3156] Thomas J. Minello, Roger J. Zimmerman, and Gilbert Zamora, Jr. Selection of vegetated habitat by brown shrimp, Penaeus aztecus, in a Galveston Bay salt marsh. Fishery Bulletin, 82(2):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/selection-vegetated-
habitat-brown-shrimp-penaeus-aztecus-galveston-bay-salt-marsh;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/zimmerman.pdf.

Crowe:1984:DLW

[3157] Barbara J. Crowe. Distribution, length–weight relationship, and length-frequency data of southern kingfish, Menticirrhus americanus, in Mississippi. Fishery Bulletin, 82(2):??, 1984. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/distribution-length-
weight-relationship-and-length-frequency-data-southern-kingfish;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/crowe.pdf.

Brodeur:1984:FHD

[3158] Richard D. Brodeur and William G. Pearcy. Food habits and dietary overlap of some shelf rockfishes (genus Sebastes) from the northeastern Pacific Ocean. Fishery Bulletin, 82(2):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/food-habits-and-dietary-overlap-some-
shelf-rockfishes-genus-sebastes-northeastern-pacific;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
brodeur.pdf.

Radtke:1984:SEM

[3159] Richard Radtke. Scanning electron microscope evidence for yearly growth zones in giant bluefin tuna, Thunnus thynnus, otoliths from daily increments. Fishery Bulletin, 82(2):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/scanning-electron-microscope-evidence-yearly-growth-
zones-giant-bluefin-tuna-thunnus;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/radtke.pdf.

Colvocoresses:1984:SAC

[3160] J. A. Colvocoresses and J. A. Musick. Species associations and community composition of Middle Atlantic Bight continental shelf demersal fishes. Fishery Bulletin, 82(2):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/species-associations-and-community-composition-
middle-atlantic-bight-continental-shelf;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/colvocoresses.
pdf.

Payne:1984:YCA

[3161] P. Michael Payne and David C. Schneider. Yearly changes in abundance of harbor seals, Phoca vitulina, at a winter haul-out site in Massachusetts. Fishery Bulletin, 82(2):??, 1984. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/yearly-changes-abundance-harbor-
seals-phoca-vitulina-winter-haul-out-site-massachusetts;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
payne.pdf.

Shepherd:1984:RWC

[3162] Gary R. Shepherd and Churchill B. Grimes. Reproduction of weakfish, Cynoscion regalis, in the New York Bight and evidence for geographically specific life history characteristics. Fishery Bulletin, 82(3):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
reproduction-weakfish-cynoscion-regalis-new-york-bight-and-
evidence-geographically-specific;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1984/823/shepherd.pdf.

Brousseau:1984:SPG

[3163] Diane J. Brousseau and Jenny A. Baglivo. Sensitivity of the population growth rate to changes in single life history parameters: its application to Mya arenaria (Mollusca: Pelecypoda). Fishery Bulletin, 82(3):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/sensitivity-
population-growth-rate-changes-single-life-history-parameters-
its-application;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/brousseau%20(1).pdf.

Smith:1984:CTO

[3164] Stephen M. Smith, James G. Hoff, Steven P. O’Neil, and Michael P. Weinstein. Community and trophic organization of nekton utilizing shallow marsh habitats, York River, Virginia. Fishery Bulletin, 82(3):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/community-and-
trophic-organization-nekton-utilizing-shallow-marsh-habitats-
york-river;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1984/823/smith.pdf.

Sherburne:1984:OPE

[3165] Stuart W. Sherburne. The occurrence of piscine erythrocytic necrosis (PEN) in the sea lamprey, Petromyzon marinus, from several Maine localities. Fishery Bulletin, 82(3):??, 1984. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/occurrence-piscine-erythrocytic-

necrosis-pen-sea-lamprey-petromyzon-marinus-several-maine;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/sherburne_0.pdf.

Stevens:1984:DAG

[3166] Bradley G. Stevens and David A. Armstrong. Distribution, abundance, and growth of juvenile Dungeness crabs, Cancer magister, in Grays Harbor Estuary, Washington. Fishery Bulletin,

82(3):??, 1984.  CODEN FSYBAY.  ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
distribution-abundance-and-growth-juvenile-dungeness-crabs-
cancer-magister-grays-harbor;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1984/823/stevens.pdf.

Allen:1984:EDH

[3167] Sarah G. Allen, David G. Ainsley, Gary W. Page, and Christine A. Ribic. The effect of distribution on harbor seal haul out patterns at Bolinas Lagoon, California. Fishery Bulletin, 82(3):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/effect-distribution-harbor-seal-haul-
out-patterns-bolinas-lagoon-california;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/allen_0.
pdf.

Helfman:1984:CAE

[3168] Gene S. Helfman, Earl L. Bozeman, and Edward B. Brothers. Comparison of American eel growth rates from tag returns and length-age analyses. Fishery Bulletin, 82(3):519–522, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/comparison-american-eel-growth-rates-
tag-returns-and-length-age-analyses;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/helfman.pdf.

Condrey:1984:DDS

[3169] Richard E. Condrey. Density-dependent searching time: Implications in surplus-production models. Fishery Bulletin, 82(3):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/density-
dependent-searching-time-implications-surplus-production-models;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1984/823/condrey.pdf.

Haynes:1984:DES

[3170] Evan B. Haynes. Description of early stage zoeae of Spirontocaris murdochi (Decapoda, Hippolytidae) reared in the laboratory. Fishery Bulletin, 82(3):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
description-early-stage-zoeae-spirontocaris-murdochi-decapoda-
hippolytidae-reared;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/haynes%20(5).pdf.

Hoss:1984:FLO

[3171] D. E. Hoss and G. Phonlor. Field and laboratory observations on diurnal swim bladder inflation–deflation in larvae of Gulf menhaden, Brevoortia patronus. Fishery Bulletin, 82(3):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/field-and-laboratory-observations-diurnal-
swim-bladder-inflation-deflation-larvae-gulf;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1984/823/hoss.
pdf.

Ennis:1984:IMS

[3172] G. P. Ennis. Incidence of molting and spawning in the same season in female lobsters, Homarus americanus. Fishery Bulletin, 82(3):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
incidence-molting-and-spawning-same-season-female-lobsters-
homarus-americanus;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/ennis%20(1).pdf.

Johnson:1984:AGM

[3173] Allyn G. Johnson and Carl H. Saloman. Age, growth, and mortality of gray triggerfish, Balistes capriscus, from the Northeastern Gulf of Mexico. Fishery Bulletin, 82(3):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/age-growth-and-mortality-gray-triggerfish-
balistes-capriscus-northeastern-gulf-mexico;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1984/823/johnson.
pdf.

Love:1984:POR

[3174] Milton S. Love, Kimberly Shriner, and Pamela Morris. Parasites of olive rockfish, Sebastes serranoides (Scorpaenidae), off central California. Fishery Bulletin, 82(3):530–537, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/parasites-olive-rockfish-
sebastes-serranoides-scorpaenidae-central-california;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
love%20(1).pdf.

Shaklee:1984:GVPa

[3175] James B. Shaklee and Paul B. Samollow. Genetic variation and population structure in a spiny lobster, Panulirus marginatus, in the Hawaiian Archipelago. Fishery Bulletin, 82(4):??, 1984. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/genetic-variation-and-population-
structure-spiny-lobster-panulirus-marginatus-hawaiian;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1984/824/
shaklee.pdf.

Shaklee:1984:GVPb

[3176] James B. Shaklee and Paul B. Samollow. Genetic variation and population structure in a deepwater snapper, Pristipomoides filamentosus, in the Hawaiian Archipelago. Fishery Bulletin, 82(4): ??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/genetic-
variation-and-population-structure-deepwater-snapper-pristipomoides-
filamentosus;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1984/824/shaklee1.pdf.

Collette:1984:MSB

[3177] Bruce B. Collette and Joseph L. Russo. Morphology, systematics, and biology of the Spanish mackerels (Scomberomorus, Scombridae). Fishery Bulletin, 82(4):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
morphology-systematics-and-biology-spanish-mackerels-scomberomorus-
scombridae;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/collette%20(1).pdf.

Hendrickx:1984:DAS

[3178] M. E. Hendrickx. Distribution and abundance of Sicyonia penicillata Lockington, 1879 in the Gulf of California, with some notes on its biology. Fishery Bulletin, 82(4):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/distribution-and-abundance-sicyonia-penicillata-lockington-
1879-gulf-california-some-notes;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/hendrickx.pdf.

Anonymous:1984:I

[3179] Anonymous. Index. Fishery Bulletin, 82(4):??, 1984. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/index-111;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/index%
20(5).pdf.

Neilson:1985:VDS

[3180] John D. Neilson, Glen H. Geen, and Brian Chan. Variability in dimensions of salmonid otolith nuclei: Implications for stock identification and microstructure interpretation. Fishery Bulletin, 83(1):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/variability-dimensions-salmonid-
otolith-nuclei-implications-stock-identification-and;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1985/831/
neilson.pdf.

Neilson:1985:EFR

[3181] John D. Neilson and Glen H. Geen. Effects of feeding regimes and diel temperature cycles on otolith increment formation in juvenile Chinook salmon, Oncorhynchus tshawytscha. Fishery Bulletin, 83(1):91– 101, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/effects-
feeding-regimes-and-diel-temperature-cycles-otolith-increment-
formation-juvenile;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1985/831/neilson2.pdf.

Farfante:1985:RSG

[3182] Isabel Perez Farfante. The rock shrimp genus Sicyonia (Crustacea: Decapoda: Penaeoidea) in the Eastern Pacific. Fishery Bulletin, 83(1):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/rock-shrimp-
genus-sicyonia-crustacea-decapoda-penaeoidea-eastern-pacific;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1985/831/perez.pdf.

Boehlert:1985:UOC

[3183] George W. Boehlert. Using objective criteria and multiple regression models for age determination in fishes. Fishery Bulletin, 83(2):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/using-objective-criteria-
and-multiple-regression-models-age-determination-fishes;   https:

//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1985/832/
boehlert.pdf.

Cross:1985:FEA

[3184] Jeffrey N. Cross. Fin erosion among fishes collected near a Southern California municipal wastewater outfall (1971–82). Fishery Bulletin, 83(2): ??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/fin-erosion-
among-fishes-collected-near-southern-california-municipal-wastewater-
outfall;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1985/832/cross.pdf.

Demartini:1985:DDV

[3185] Edward E. Demartini, Larry G. Allen, Robert K. Fountain, and Dale Roberts. Diel and depth variations in the sex-specific abundance, size composition, and food habits of queenfish, Seriphus politus (Sciaenidae). Fishery Bulletin, 83(2):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/diel-and-depth-variations-sex-specific-abundance-size-
composition-and-food-habits-queenfish;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1985/832/demartini.pdf.

Hewitt:1985:RDS

[3186] Roger P. Hewitt. Reaction of dolphins to a survey vessel: Effects on census data. Fishery Bulletin, 83(2):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/reaction-dolphins-survey-vessel-effects-census-
data;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1985/832/hewitt.pdf.

Hunter:1985:RAO

[3187] J. Roe Hunter and Beverly J. Macewicz. Rates of atresia in the ovary of captive and wild northern anchovy; Engraulis mordax. Fishery Bulletin, 83(2):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/rates-
atresia-ovary-captive-and-wild-northern-anchovy-engraulis-mordax;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1985/832/hunter.pdf.

Lo:1985:EPC

[3188] Nancy C. H. Lo. Egg production of the central stock of northern anchovy, Engraulis mordax, 1951–82. Fishery Bulletin, 83(2):

137–150, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/egg-
production-central-stock-northern-anchovy-engraulis-mordax-
1951-82;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1985/832/lo.pdf.

Mullin:1985:VSN

[3189] M. M. Mullin, E. R. Brooks, F. M. H. Reid, J. Napp, and E. F. Stewart. Vertical structure of nearshore plankton off Southern California: a storm and a larval fish food web. Fishery Bulletin, 83 (2):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/vertical-
structure-nearshore-plankton-southern-california-storm-and-
larval-fish-food-web;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1985/832/mullin.pdf.

Guillemot:1985:SCF

[3190] Patrick J. Guillemot, Ralph J. Larson, and William H. Lenarz. Seasonal cycles of fat and gonad volume in five species of Northern California rockfish (Scorpaenidae: Sebastes). Fishery Bulletin, 83(3):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/seasonal-cycles-fat-
and-gonad-volume-five-species-northern-california-rockfish-
scorpaenidae;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1985/833/guillemot.pdf.

Rosenblum:1985:SCS

[3191] Shelly E. Rosenblum and Thomas M. Niesen. The spawning cycle of soft-shell clam, Mya arenaria, in San Francisco Bay. Fishery Bulletin, 83(3):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/spawning-
cycle-soft-shell-clam-mya-arenaria-san-francisco-bay;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1985/833/
rosenblum.pdf.

Hui:1985:UTC

[3192] Clifford A. Hui. Undersea topography and the comparative distributions of two pelagic cetaceans. Fishery Bulletin, 83(3):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/undersea-topography-
and-comparative-distributions-two-pelagic-cetaceans;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
hui_0.pdf.

Haynes:1985:MDI

[3193] Evan B. Haynes. Morphological development, identification, and biology of larvae of Pandalidae, Hippolytidae, and Crangonidae (Crustacea, Decapoda) of the Northern North Pacific Ocean. Fishery Bulletin, 83(3): ??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/morphological-
development-identification-and-biology-larvae-pandalidae-hippolytidae-
and;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1985/833/haynes.pdf.

Stevenson:1985:LHC

[3194] David K. Stevenson and Fran Pierce. Life history characteristics of Pandalus montagui and Dichelopandalus leptocerus in Penobscot Bay, Maine. Fishery Bulletin, 83(3):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/life-history-characteristics-pandalus-montagui-and-
dichelopandalus-leptocerus-penobscot-bay;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1985/833/stevenson.
pdf.

Boehlert:1985:LJG

[3195] George W. Boehlert and Mary M. Yoklavich. Larval and juvenile growth of sablefish, Anoplopoma fimbria, as determined from otolith increments. Fishery Bulletin, 83(3):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/larval-and-juvenile-growth-sablefish-
anoplopoma-fimbria-determined-otolith-increments;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
boehlert%20(2).pdf.

Hunter:1985:VTS

[3196] John Hunter and Ragan Nicholl. Visual threshold for schooling in northern anchovy Engraulis mordax. Fishery Bulletin, 83(3): ??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/visual-
threshold-schooling-northern-anchovy-engraulis-mordax;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1985/833/
hunter.pdf.

Wursig:1985:BBW

[3197] Bernd Würsig, Eleanor M. Dorsey, Mark A. Fraker, Roger S. Payne, and W. John Richardson. Behavior of bowhead whales, Balaena mysticetus, summering in the Beaufort Sea: a description. Fishery Bulletin, 83(3):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/behavior-
bowhead-whales-balaena-mysticetus-summering-beaufort-sea-description;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1985/833/wursig.pdf.

Ankenbrandt:1985:FHB

[3198] Lisa Ankenbrandt. Food habits of bait-caught skipjack tuna, Katsuwonus pelamis, from the Southwestern Atlantic Ocean. Fishery Bulletin, 83(3): ??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/food-habits-
bait-caught-skipjack-tuna-katsuwonus-pelamis-southwestern-atlantic-
ocean;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1985/833/ankenbrandt.pdf.

Jones:1985:WSD

[3199] Cynthia Jones. Within-season differences in growth of larval Atlantic herring, Clupea harengus harengus. Fishery Bulletin, 83(3):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/within-season-
differences-growth-larval-atlantic-herring-clupea-harengus-
harengus;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1985/833/jones.pdf.

Grossman:1985:RBT

[3200] Gary N. Grossman, Michael J. Harris, and Joseph E. Hightower. The relationship between tilefish, Lopholatilus chamaeleonticeps, abundance and sediment composition off Georgia. Fishery Bulletin, 83(3):443–447, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/relationship-between-
tilefish-lopholatilus-chamaeleonticeps-abundance-and-sediment;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/grossman.pdf.

Conover:1985:FLA

[3201] David O. Conover. Field and laboratory assessment of patterns in fecundity of a multiple spawning fish: the Atlantic silverside Menidia menidia. Fishery Bulletin, 83(3):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/field-and-laboratory-assessment-patterns-fecundity-
multiple-spawning-fish-atlantic;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1985/833/conover.pdf.

Lester:1985:PST

[3202] R. J. G. Lester, A. Barnes, and G. Habib. Parasites of skipjack tuna, Katsuwonus pelamis: Fishery implications. Fishery Bulletin, 83(3):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/parasites-skipjack-tuna-katsuwonus-pelamis-fishery-
implications;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1985/833/lester.pdf.

Matarese:1985:DOL

[3203] Ann C. Matarese and Beverly M. Vinter. The development and occurrence of larvae of the longfin Irish lord, Hemilepidotus zapus (Cottidae). Fishery Bulletin, 83(3):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/development-and-occurrence-larvae-longfin-
irish-lord-hemilepidotus-zapus-cottidae;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/matarese%
20(1).pdf.

Flierl:1985:PIW

[3204] G. R. Flierl and J. S. Wroblewski. The possible influence of warm core Gulf Stream rings upon shelf water larval fish distribution. Fishery Bulletin, 83(3):313–330, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/possible-influence-warm-core-gulf-stream-rings-
upon-shelf-water-larval-fish-distribution;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1985/833/flierl.pdf.

Love:1985:DPF

[3205] Milton S. Love, William Westphal, and Robson A. Collins. Distributional patterns of fishes captured aboard commercial passenger fishing vessels along the Northern Channel Islands, California. Fishery Bulletin, 83(3):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/distributional-patterns-fishes-captured-aboard-
commercial-passenger-fishing-vessels-along;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1985/833/love.pdf.

Polovina:1985:AEE

[3206] Jeffrey J. Polovina and Mark D. Ow. An approach to estimating an ecosystem box model. Fishery Bulletin, 83(3):??, 1985. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/approach-estimating-
ecosystem-box-model;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/polovina.pdf.

Gaskin:1985:HPP

[3207] David E. Gaskin and Alan P. Watson. The harbor porpoise, Phocoena phocoena, in Fish Harbour, New Brunswick, Canada: Occupancy, distribution, and movements. Fishery Bulletin, 83(3):??, 1985. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/harbor-porpoise-phocoena-
phocoena-fish-harbour-new-brunswick-canada-occupancy-distribution;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1985/833/gaskin.pdf.

McFarland:1985:RPY

[3208] W. N. McFarland, E. B. Brothers, J. C. Ogden, M. J. Shulman, E. L. Bermingham, and N. M. Kotchian-Prentiss. Recruitment patterns in young French grunts, Haemulon flavolineatum (family Haemulidae), at St. Croix, Virgin Islands. Fishery Bulletin, 83(3):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/recruitment-
patterns-young-french-grunts-haemulon-flavolineatum-family-
haemulidae-st-croix;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1985/833/mcfarland.pdf.

Sedberry:1985:FFT

[3209] George R. Sedberry. Food and feeding of the torptate, Haemulon aurolineatum (Pisces, Haemulidae), in the South Atlantic Bight. Fishery Bulletin, 83(3):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/food-
and-feeding-torptate-haemulon-aurolineatum-pisces-haemulidae-
south-atlantic-bight;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/sedberry.pdf.

Gerrodette:1985:CLP

[3210] Tim Gerrodette, Daniel Goodman, and Jay Barlow. Confidence limits for population projections when vital rates vary randomly. Fishery Bulletin, 83(3):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
confidence-limits-population-projections-when-vital-rates-vary-
randomly;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1985/833/gerrodette.pdf.

Medved:1985:SCY

[3211] Robert J. Medved, Charles E. Stillwell, and John J. Casey. Stomach contents of young sandbar sharks, Carcharhinus plumbeus, in Chincoteague Bay, Virginia. Fishery Bulletin, 83(3):395–402, 1985. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/stomach-contents-young-sandbar-
sharks-carcharhinus-plumbeus-chincoteague-bay-virginia;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1985/833/
medved.pdf.

Hines:1985:SRC

[3212] Anson H. Hines, Kenric E. Osgood, and Joseph J. Miklas. Semilunar reproductive cycles in Fundulus heteroclitus (Pisces: Cyprinodontidae) in an area without lunar tidal cycles. Fishery Bulletin, 83 (3):467–472, 1985. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
semilunar-reproductive-cycles-fundulus-heteroclitus-pisces-
cyprinodontidae-area-without;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/hines%20(1).pdf.

Boehlert:1985:VDI

[3213] George W. Boehlert, Dena M. Gadomski, and Bruce C. Mundy. Vertical distribution of ichthyoplankton off the Oregon Coast in spring and summer months. Fishery Bulletin, 83(4):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/vertical-distribution-ichthyoplanklun-oregon-
coast-spring-and-summer-months;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1985/834/boehlert.pdf.

Singer:1985:FHJ

[3214] Michael M. Singer. Food habits of juvenile rockfishes (Sebastes) in a Central California kelp forest. Fishery Bulletin, 83(4):531–541, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/food-habits-juvenile-
rockfishes-sebastes-central-california-kelp-forest;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1985/834/
singer.pdf.

Libby:1985:CSO

[3215] David A. Libby. A comparison of scale and otolith aging methods for the alewife, Alosa pseudoharengus. Fishery Bulletin, 83(4):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/comparison-scale-and-
otolith-aging-methods-alewife-alosa-pseudoharengus.

Brown:1985:FAG

[3216] R. S. Brown and N. Caputi. Factors affecting the growth of undersize western rock lobster, Panulirus cygnus George, returned by fishermen to the sea. Fishery Bulletin, 83(4):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/factors-affecting-growth-
undersize-western-rock-lobster-panulirus-cygnus-george-returned;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1985/834/brown.pdf.

Wahlen:1985:OEI

[3217] Bruce E. Wahlen and Tim D. Smith. Observer effect on incidental dolphin mortality in the Eastern Tropical Pacific tuna fishery. Fishery Bulletin, 83(4):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/observer-
effect-incidental-dolphin-mortality-eastern-tropical-pacific-
tuna-fishery;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1985/834/wahlen.pdf.

Maule:1985:PCR

[3218] Alec G. Maule and Howard F. Horton. Probable causes of the rapid growth and high fecundity of walleye, Stizostedion vitreum vitreum, in the mid-Columbia River. Fishery Bulletin, 83(4):??, 1985. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/probable-causes-rapid-growth-and-
high-fecundity-walleye-stizostedion-vitreum-vitreum-mid;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
maule%20(1).pdf.

Freeman:1985:ALH

[3219] Mary C. Freeman, Nate Neally, and Gary D. Grossman. Aspects of the life history of the fluffy sculpin, Oligocottus snyderi. Fishery Bulletin, 83 (4):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/aspects-
life-history-fluffy-sculpin-oligocottus-snyderi;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1985/834/
freeman.pdf.

Read:1985:RTM

[3220] Andrew J. Read and David E. Gaskin. Radio tracking the movements and activities of harbor porpoises, Phocoena phocoena (L.), in the Bay of Fundy, Canada. Fishery Bulletin, 83(4):??, 1985. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/radio-tracking-movements-
and-activities-harbor-porpoises-phocoena-phocoena-l-bay-fundy;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/read%20(1).pdf.

Taylor:1985:BAS

[3221] D. M. Taylor, R. G. Hooper, and G. P. Ennis. Biological aspects of the spring breeding migration of snow crabs, Chionoecetes opilio, in Bonne Bay, Newfoundland (Canada). Fishery Bulletin, 83(4):707–711, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/biological-aspects-
spring-breeding-migration-snow-crabs-chionoecetes-opilio-bonne-
bay;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/taylor_3.pdf.

Hohn:1985:EPG

[3222] Aleta A. Hohn and P. S. Hammond. Early postnatal growth of the spotted dolphin, Stenella attenuata, in the offsore Eastern Tropical Pacific. Fishery Bulletin, 83(4):??, 1985. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/early-postnatal-growth-spotted-
dolphin-stenella-attenuata-offsore-eastern-tropical-pacific;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1985/834/hohn.pdf.

Warlen:1985:AGD

[3223] Stanley M. Warlen and Alexander J. Chester. Age, growth, and distribution of larval spot, Leiostomus xanthurus, off North Carolina. Fishery Bulletin, 83(4):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/age-growth-and-distribution-larval-spot-leiostomus-
xanthurus-north-carolina;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/warlen.pdf.

Creed:1985:FDR

[3224] Robert P. Creed, Jr. Feeding, diet, and repeat spawning of blueblack herring, Alosa aestivalis, from the Chowan River, North Carolina.

Fishery Bulletin, 83(4):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/feeding-diet-and-repeat-spawning-blueblack-herring-
alosa-aestivalis-chowan-river-north;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/creed%20(1).pdf.

Jamieson:1985:SSF

[3225] G. S. Jamieson and A. Campbell. Sea scallop fishing impact on American lobsters in the Gulf of St. Lawrence. Fishery Bulletin, 83(4):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/sea-scallop-
fishing-impact-american-lobsters-gulf-st-lawrence;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1985/834/
jamieson.pdf.

Peterson:1985:ABD

[3226] Charles H. Peterson, P. Bruce Duncan, Henry C. Summerson, and Brian F. Beal. Annual band deposition within shells of the hard clam, Mercenaria mercenaria: consistency across habitat near Cape Lookout, North Carolina. Fishery Bulletin, 83(4):671–677, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/annual-band-
deposition-within-shells-hard-clam-mercenaria-mercenaria-consistency-
across;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/peterson%20(2).pdf.

Anonymous:1985:I

[3227] Anonymous. Index. Fishery Bulletin, 83(4):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/index-112;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/index%
20(6).pdf.

Albers:1985:DPC

[3228] W. D. Albers and P. J. Anderson. Diet of Pacific cod, Gadus macrocephalus, and predation on the northern pink shrimp, Pandalus borealis, in Pavlof Bay, Alaska. Fishery Bulletin, 83(4):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/diet-pacific-
cod-gadus-macrocephalus-and-predation-northern-pink-shrimp-
pandalus-borealis;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1985/834/albers.pdf.

Johnson:1985:PBA

[3229] Phyllis T. Johnson. Parasites of benthic amphipods: Microsporidans of Ampelisca agassizi (Judd) and some other gammarideans. Fishery Bulletin, 83(4):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
parasites-benthic-amphipods-microsporidans-ampelisca-agassizi-
judd-and-some-other;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1985/834/johnson.pdf.

Sullivan:1985:SSJ

[3230] Loretta F. Sullivan, Dennis A. Emiliani, and K. Neal Baxter. Standing stock of juvenile brown shrimp, Penaeus aztecus, in Texas coastal ponds. Fishery Bulletin, 83(4):677–682, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/standing-stock-juvenile-brown-shrimp-
penaeus-aztecus-texas-coastal-ponds;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/sullivan%20(1)
.pdf.

Au:1985:DHE

[3231] David W. K. Au and Wayne L. Perryman. Dolphin habitats in the Eastern Tropical Pacific. Fishery Bulletin, 83(4):??, 1985. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/dolphin-habitats-eastern-
tropical-pacific;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1985/834/au.pdf.

Overholtz:1985:LTR

[3232] William J. Overholtz and Albert V. Tyler. Long-term responses of the demersal fish assemblages of Georges Bank. Fishery Bulletin, 83(4):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
long-term-responses-demersal-fish-assemblages-georges-bank;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1985/834/overholtz.pdf.

Thomas:1985:PLB

[3233] David H. Thomas. A possible link between Coho (silver) salmon enhancement and a decline in central California Dungeness crab abundance. Fishery Bulletin, 83(4):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
possible-link-between-coho-silver-salmon-enhancement-and-decline-

central-california;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/thomas_1.pdf.

Barlow:1985:VTB

[3234] Jay Barlow. Variability, trends, and biases in reproductive rates of spotted dolphins, Stenella attenuata. Fishery Bulletin, 83(4):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/variability-trends-
and-biases-reproductive-rates-spotted-dolphins-stenella-attenuata;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1985/834/barlow.pdf.

Parrish:1985:RVG

[3235] R. H. Parrish, D. L. Mallicoate, and K. F. Mais. Regional variations in the growth and age composition of northern anchovy, Engraulis mordax. Fishery Bulletin, 83(4):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/regional-variations-growth-and-age-composition-
northern-anchovy-engraulis-mordax;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1985/834/parrish.pdf.

Parsons:1985:NLH

[3236] Glenn R. Parsons. Notes on the life history of the catshark, Scyliorhinus meadi. Fishery Bulletin, 83(4):??, 1985. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/notes-life-history-catshark-scyliorhinus-meadi;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/parsons.pdf.

Keirans:1986:DPC

[3237] Walter J. Keirans, Sidney S. Herman, and R. G. Malsberger. Differentiation of Prionotus carolinus and Prionotus evolans eggs in Hereford Inlet Estuary, Southern New Jersey, using immunodiffusion. Fishery Bulletin, 84(1):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
differentiation-prionotus-carolinus-and-prionotus-evolans-eggs-
hereford-inlet-estuary;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1986/841/keirans.pdf.

Theilacker:1986:SIM

[3238] Gail H. Theilacker. Starvation-induced mortality of young seacaught jack mackerel, Trachurus symmetricus, determined with histological and morphological methods. Fishery Bulletin, 84(1):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/starvation-
induced-mortality-young-sea-caught-jack-mackerel-trachurus-
symmetricus;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/theilacker.pdf.

Johnson:1986:PBAa

[3239] Phyllis T. Johnson. Parasites of benthid amphipods: ciliates. Fishery Bulletin, 84(1):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/parasites-benthid-amphipods-ciliates;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
johnson2.pdf.

Lo:1986:IMD

[3240] N. C. H. Lo and T. D. Smith. Incidental mortality of dolphins in the Eastern Tropical Pacific, 1959–72. Fishery Bulletin, 84(1):27–34, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/incidental-mortality-dolphins-eastern-tropical-
pacific-1959-72;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1986/841/lo.pdf.

Hunt:1986:EEC

[3241] John H. Hunt, William C. Lyons, and Frank S. Kennedy, Jr. Effects of exposure and confinement on spiny lobsters, Panulirus argus, used as attractants in the Florida fishery. Fishery Bulletin, 84(1):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/effects-exposure-and-
confinement-spiny-lobsters-panulirus-argus-used-attractants-
florida;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/hunt.pdf.

Mason:1986:FPH

[3242] J. C. Mason. Fecundity of the Pacific hake, Merluccius productus, spawning in Canadian waters. Fishery Bulletin, 84(1): ??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fecundity-
pacific-hake-merluccius-productus-spawning-canadian-waters;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/mason.pdf.

Mlddleton:1986:ADF

[3243] Robert W. Mlddleton and John A. Musick. The abundance and distribution of the family Macrouridae (Pisces: Gadiformes) in the Norfolk Canyon area. Fishery Bulletin, 84(1):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/abundance-and-distribution-family-macrouridae-
pisces-gadiformes-norfolk-canyon-area;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1986/841/middleton.pdf.

Krygier:1986:REO

[3244] E. E. Krygier and W. G. Pearcy. The role of estuarine and offshore nursery areas for young English sole, Parophrys vetulus Girard, of Oregon. Fishery Bulletin, 84(1):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/role-estuarine-and-offshore-nursery-areas-
young-english-sole-parophrys-vetulus-girard;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/krygier_
0.pdf.

Selzer:1986:SAI

[3245] Lawrence A. Selzer, Greg Early, Patricia M. Fiorelli, P. Michael Payne, and Robert Prescott. Stranded animals as indicators of prey utilization by harbor seals, Phoca vitulina concolor, in southern New England. Fishery Bulletin, 84(1):??, 1986. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/stranded-animals-indicators-prey-
utilization-harbor-seals-phoca-vitulina-concolor-southern;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/selzer.pdf.

Moyle:1986:PDA

[3246] Peter B. Moyle, Robert A. Daniels, Bruce Herbold, and Donald M. Baltz. Patterns in distribution and abundance of a noncoevolved assemblage of estuarine fishes in California. Fishery Bulletin, 84(1):105– 117, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/patterns-
distribution-and-abundance-noncoevolved-assemblage-estuarine-
fishes-california;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1986/841/moyle.pdf.

Weber:1986:SRG

[3247] Earl C. Weber and Stephen R. Goldberg. The sex ratio and gonad indices of swordfish, Xiphias gladius, caught off the coast of southern California in 1978. Fishery Bulletin, 84(1):185–186, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/sex-ratio-and-gonad-
indices-swordfish-xiphias-gladius-caught-coast-southern-california-
1978;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/weber%20(1).pdf.

Ranck:1986:GCS

[3248] Carol L. Ranck, Fred M. Utter, George B. Milner, and Gary B. Smith. Genetic confinnation of specific distinction of arrowtooth flounder, Atheresthes stomius, and Kamchatka flounder, A. evermanni. Fishery Bulletin, 84(1):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/genetic-confinnation-specific-distinction-
arrowtooth-flounder-atheresthes-stomius-and;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/ranck.
pdf.

Ralston:1986:ESC

[3249] Stephen Ralston, Reginald M. Gooding, and Gerald M. Ludwig. An ecological survey and comparison of bottom fish resource assessments (submersible versus handline fishing) at Johnston Atoll. Fishery Bulletin, 84 (1):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/ecological-
survey-and-comparison-bottom-fish-resource-assessments-submersible-
versus;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1986/841/ralston.pdf.

Uchiyama:1986:GDC

[3250] James H. Uchiyama, Raymond K. Burch, and Syd A. Kraul, Jr. Growth of dolphins, Coryphaena hippurus and C. equiselis, in Hawaiian waters as determined by daily increments on otoliths. Fishery Bulletin, 84(1):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/growth-dolphins-coryphaena-hippurus-and-c-
equiselis-hawaiian-waters-determined-daily;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/uchiyama.
pdf.

Beacham:1986:TQS

[3251] Terry D. Beacham. Type, quantity, and size of food of Pacific salmon (Oncorhynchus) in the Strait of Juan de Fuca, British Columbia. Fishery Bulletin, 84(1):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
type-quantity-and-size-food-pacific-salmon-oncorhynchus-strait-
juan-de-fuca-british;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1986/841/beacham.pdf.

Renaud:1986:HLC

[3252] Maurice L. Renaud. Hypoxia in Louisiana coastal waters during 1983: Implications for fisheries. Fishery Bulletin, 84(1):??, 1986. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/hypoxia-louisiana-coastal-waters-
during-1983-implications-fisheries;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1986/841/renaud.pdf.

Frost:1986:SWP

[3253] Kathryn J. Frost and Lloyd F. Lowry. Sizes of walleye pollock, Theragra chalcogramma, consumed by marine mammals in the Bering Sea. Fishery Bulletin, 84(1):??, 1986. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/sizes-walleye-pollock-theragra-
chalcogramma-consumed-marine-mammals-bering-sea;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
frost%20(1).pdf.

Johnson:1986:RIB

[3254] P. T. Johnson, R. A. MacIntosh, and D. A. Somerton. Rhizocephalan infection in blue king crabs, Paralithodes platypus, from Olga Bay, Kodiak Island, Alaska. Fishery Bulletin, 84(1):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/rhizocephalan-
infection-blue-king-crabs-paralithodes-platypus-olga-bay-kodiak-
island-alaska;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1986/841/johnson.pdf.

Steimle:1986:OTM

[3255] Frank W. Steimle, Paul D. Boehm, Vincent S. Zdanowicz, and Ralph A. Brun. Organic and trace metal levels in ocean quahog, Arctica islandica Linne, from the Northwestern Atlantic. Fishery Bulletin, 84(1): ??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/organic-and-
trace-metal-levels-ocean-quahog-arctica-islandica-linne-northwestern-
atlantic;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1986/841/steimle.pdf.

VanEngel:1986:OSP

[3256] W. A. Van Engel, R. E. Harris, Jr., and D. E. Zwerner. Occurrence of some parasites and a commensal in the American lobster, Homarus americanus, from the Mid-Atlantic Bight. Fishery Bulletin, 84(1):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/occurrence-
some-parasites-and-commensal-american-lobster-homarus-americanus-
mid-atlantic;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/vanengel.pdf.

Jones:1986:DAL

[3257] Cynthia Jones. Determining age of larval fish with the otolith increment technique. Fishery Bulletin, 84(1):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/determining-age-larval-fish-otolith-increment-
technique;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1986/841/jones.pdf.

Willason:1986:PNC

[3258] Stewart W. Willason, John Favuzzi, and James L. Cox. Patchiness and nutritional condition of zooplankton in the California Current. Fishery Bulletin, 84(1):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/patchiness-and-nutritional-condition-
zooplankton-california-current;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1986/841/willason.pdf.

Collette:1986:RFA

[3259] Bruce B. Collette. Resilience of the fish assemblage in New England tidepools. Fishery Bulletin, 84(1):??, 1986. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/resilience-fish-assemblage-new-
england-tidepools;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/collette%20(2).pdf.

Humes:1986:CAL

[3260] Arthur G. Humes. Copepodids and adults of Leptinogaster major (Williams, 1907), a poecilostomatoid copepod living in Mya arenaria L. and other marine bivalve mollusks. Fishery Bulletin, 84(2):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/copepodids-
and-adults-leptinogaster-major-williams-1907-poecilostomatoid-

copepod-living-mya;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1986/842/humes.pdf.

Polovina:1986:VCV

[3261] Jeffrey J. Polovina. A variable catchability version of the Leslie model with application to an intensive fishing experiment on a multispecies stock. Fishery Bulletin, 84(2):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/variable-catchability-version-leslie-
model-application-intensive-fishing-experiment;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1986/842/polovina.
pdf.

Graham:1986:MPS

[3262] Jeffrey B. Graham, Richard H. Rosenblatt, and Darcy L. Gibson. Morphology and possible swimming mode of a yellowfin tuna, Thunnus albacares, lacking one pectoral fin. Fishery Bulletin, 84(2):463–468, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/morphology-
and-possible-swimming-mode-yellowfin-tuna-thunnus-albacares-
lacking-one-pectoral;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/graham_0.pdf.

Kenney:1986:CHU

[3263] Robert D. Kenney and Howard E. Winn. Cetacean high-use habitats of the Northeast United States Continental Shelf. Fishery Bulletin, 84(2):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/cetacean-high-use-habitats-northeast-united-states-
continental-shelf;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1986/842/kenney.pdf.

Sen:1986:MPS

[3264] A. R. Sen. Methodological problems in sampling commercial rockfish landings. Fishery Bulletin, 84(2):??, 1986. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/methodological-problems-sampling-
commercial-rockfish-landings;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1986/842/sen.pdf.

Ratty:1986:CAA

[3265] F. J. Ratty, Y. C. Song, and R. M. Laurs. Chromosomal analysis of albacore, Thunnus alalunga, and yellowfin, Thunnus albacares, and skipjack, Katsuwonus pelamis, tuna. Fishery Bulletin, 84(2):469–476, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/chromosomal-analysis-albacore-thunnus-alalunga-and-
yellowfin-thunnus-albacares-and-skipjack;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/ratty.pdf.

Lo:1986:MLS

[3266] Nancy C. H. Lo. Modeling life-stage-specific instantaneous mortality rates, an application to northern anchovy Engraulis mordax, eggs and larvae. Fishery Bulletin, 84(2):??, 1986. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/modeling-life-stage-
specific-instantaneous-mortality-rates-application-northern-
anchovy;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1986/842/lo.pdf.

Squires:1986:EVP

[3267] Dale Squires. Ex-vessel price linkages in the New England fishing industry. Fishery Bulletin, 84(2):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/ex-vessel-price-linkages-new-england-
fishing-industry;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1986/842/squires.pdf.

Stier:1986:ASS

[3268] Kathleen Stier and Boyd Kynard. Abundance, size, and sex ratio of adult sea-run sea lamprey, Petromyzon marinus, in the Connecticut River. Fishery Bulletin, 84(2):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/abundance-size-and-sex-ratio-adult-sea-run-sea-
lamprey-petromyzon-marinus-connecticut-river;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/stier.
pdf.

Bayer:1986:SNO

[3269] Range D. Bayer. Seabirds near an Oregon estuarine salmon hatchery in 1982 and during the 1983 El Niño. Fishery Bulletin, 84(2): ??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/seabirds-
near-oregon-estuarine-salmon-hatchery-1982-and-during-1983-
el-nino;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1986/842/bayer.pdf.

Matsuura:1986:EDL

[3270] Yasunobu Matsuura and Nelson Takumi Yoneda. Early development of the lophiid anglerfish, Lophius gastrophysus. Fishery Bulletin, 84 (2):429–436, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/early-
development-lophiid-anglerfish-lophius-gastrophysus;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1986/842/
matsuura.pdf.

Sykes:1986:CSO

[3271] Stephen D. Sykes and Louis W. Botsford. Chinook salmon, Oncorhynchus tshawytscha, spawning escapement based on multiple mark– recapture of carcasses. Fishery Bulletin, 84(2):??, 1986. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/chinook-salmon-oncorhynchus-tshawytscha-
spawning-escapement-based-multiple-mark-recapture;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1986/842/
sykes.pdf.

Mason:1986:IOS

[3272] J. C. Mason and A. C. Phillips. An improved otter surface sampler. Fishery Bulletin, 84(2):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/improved-otter-surface-sampler;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/mason%20(
1).pdf.

Bigg:1986:ANS

[3273] Michael A. Bigg. Arrival of northern fur seals, Callorhinus ursinus, on St Paul Island, Alaska. Fishery Bulletin, 84(2):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/arrival-northern-fur-
seals-callorhinus-ursinus-st-paul-island-alaska;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1986/842/
bigg.pdf.

Myrick:1986:RBF

[3274] A. C. Myrick, Jr., A. A. Hohn, J. Barlow, and P. A. Sloan. Reproductive biology of female spotted dolphins, Stenella attenuata, from the Eastern Tropical Pacific. Fishery Bulletin, 84(2):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/reproductive-

biology-female-spotted-dolphins-stenella-attenuata-eastern-
tropical-pacific;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1986/842/myrick.pdf.

Wilkins:1986:DEM

[3275] Mark E. Wilkins. Development and evaluation of methodologies for assessing and monitoring the abundance of widow rockfish, Sebastes entomelas. Fishery Bulletin, 84(2):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/development-and-evaluation-methodologies-assessing-
and-monitoring-abundance-widow-rockfish;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1986/842/wilkins.pdf.

Grover:1986:MES

[3276] Jill J. Grover and Bori L. Olla. Morphological evidence for starvation and prey size selection of sea-caught larval sablefish, Anoplopoma fimbria. Fishery Bulletin, 84(2):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/morphological-evidence-starvation-and-prey-size-selection-
sea-caught-larval-sablefish;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/grover_1.pdf.

Cummings:1986:SBB

[3277] William C. Cummings, Paul O. Thompson, and Samuel J. Ha. Sounds from Bryde, Balaenoptera edeni, and finback, B. Physalus, whales in the Gulf of California. Fishery Bulletin, 84(2):??, 1986. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/sounds-bryde-balaenoptera-edeni-
and-finback-b-physalus-whales-gulf-california;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1986/842/cummings.
pdf.

Nelson:1986:PFC

[3278] Walter R. Nelson and Dean W. Ahrenholz. Population and fishery characteristics of Gulf menhaden, Brevoortia patronus. Fishery Bulletin, 84 (2):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/population-
and-fishery-characteristics-gulf-menhaden-brevoortia-patronus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1986/842/nelson.pdf.

Leber:1986:CSS

[3279] Kenneth M. Leber and Holly S. Greening. Community studies in seagrass meadows: a comparison of two methods for sampling macroinvertebrates and fishes. Fishery Bulletin, 84(2):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic).  URL https://spo.nmfs.noaa.gov/content/community-
studies-seagrass-meadows-comparison-two-methods-sampling-macroinvertebrates-
and;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/leber.pdf.

Dohl:1986:DAC

[3280] Thomas P. Dohl, Michael L. Bonnell, and R. Glenn Ford. Distribution and abundance of common dolphin, Delphinus delphis, in the Southern California Bight: a quantitative assessment based upon aerial transect data. Fishery Bulletin, 84(2):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/distribution-and-abundance-common-dolphin-
delphinus-delphis-southern-california-bight;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1986/842/dohl.
pdf.

Payne:1986:DHW

[3281] P. Michael Payne, John R. Nicolas, Loretta O’Brien, and Kevin D. Powers. The distribution of the humpback whale, Megaptera novaeangliae, on Georges Bank and in the Gulf of Maine in relation to densities of the sand eel, Ammodytes americanus. Fishery Bulletin, 84(2):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
distribution-humpback-whale-megaptera-novaeangliae-georges-
bank-and-gulf-maine-relation-t0;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1986/842/payne.pdf.

Oxenford:1986:PIS

[3282] Hazel A. Oxenford and Wayne Hunte. A preliminary investigation of the stock structure of the dolphin, Coryphaena hippurus, in the western central Atlantic. Fishery Bulletin, 84(2):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/preliminary-investigation-stock-structure-
dolphin-coryphaena-hippurus-western-central;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/oxenford%
20(1).pdf.

Hawkes:1986:LWR

[3283] Clayton R. Hawkes, Theodore R. Meyers, and Thomas C. Shirley. Length–weight relationships of blue, Paralithodes platypus, and golden, Lithodes aequispina, king crabs parasitized by the rhizocephalan, Briarosaccus callosus Boschma. Fishery Bulletin, 84(2):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/length-weight-relationships-blue-
paralithodes-platypus-and-golden-lithodes-aequispina-king;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1986/
842/hawkes.pdf.

Perez:1986:IFE

[3284] Michael A. Perez and Elizabeth E. Mooney. Increased food and energy consumption of lactating northern fur seals, Callorhinus ursinus. Fishery Bulletin, 84(2):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
increased-food-and-energy-consumption-lactating-northern-fur-
seals-callorhinus-ursinus;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1986/842/perez.pdf.

Forward:1986:ETS

[3285] Richard B. Forward, Jr., Blanca Rojas De Mendiola, and Richard T. Barber. Effects of temperature on swimming speed of the dinoflagellate Gymnodinium splendens. Fishery Bulletin, 84(2):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/effects-temperature-
swimming-speed-dinoflagellate-gymnodinium-splendens;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
forward.pdf.

Pennington:1986:SST

[3286] Michael Pennington. Some statistical techniques for estimating abundance indices from trawl surveys. Fishery Bulletin, 84(3): ??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14292;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/pennington.pdf.

Yang:1986:FHD

[3287] M. S. Yang and P. A. Livingston. Food habits and diet overlap of two congeneric species, Atheresthes stomias and Atheresthes evermanni, in the eastern Bering Sea. Fishery Bulletin, 84(3):??, 1986. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14300;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/yang.pdf.

Lenarz:1986:CVF

[3288] William H. Lenarz and Tina Wyllie Echeverria. Comparison of visceral fat and gonadal fat volumes of yellowtail rockfish, Sebastes flavidus, during a normal year and a year of El Niño conditions. Fishery Bulletin, 84(3):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14308;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/lenarz_0.pdf.

Reilly:1986:RID

[3289] Stephen B. Reilly and Jay Barlow. Rates of increase in dolphin population size. Fishery Bulletin, 84(3):??, 1986. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14293;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/reilly.pdf.

Shepard:1986:ECC

[3290] Andrew N. Shepard, Roger R. Theroux, Richard A. Cooper, and Joseph R. Uzmann. Ecology of Ceriantharia (Coelenterata, Anthozoa) of the northwest Atlantic from Cape Hatteras to Nova Scotia. Fishery Bulletin, 84(3):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14301;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/shepard.pdf.

Sorensen:1986:DFA

[3291] Peter W. Sorensen, Marco L. Bianchini, and Howard E. Winn. Diel foraging activity of American eels, Anguilla rostrata (Lesueur), in a Rhode Island estuary. Fishery Bulletin, 84(3):746–747, 1986. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14309;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/sorensen.pdf.

Atkinson:1986:DTD

[3292] C. Allen Atkinson. Discrete-time difference model for simulating interacting fish population dynamics. Fishery Bulletin, 84(3):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14294;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/atkinson.pdf.

Potthoff:1986:CBD

[3293] Thomas Potthoff, Sharon Kelley, and Joaquin C. Javech. Cartilage and bone development in scombroid fishes. Fishery Bulletin, 84(3): 647–678, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14302;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/potthoff_1.pdf.

Killam:1986:FRL

[3294] Kristie Killam and Glenn Parsons. First record of the longfin mako, Isurus paucus, in the Gulf of Mexico. Fishery Bulletin, 84 (3):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14310;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/killam.pdf.

Parsons:1986:FNS

[3295] O. G. Parsons and G. E. Tucker. Fecundity of northern shrimp, Pandulus borealis, (Crustacea, Decapoda) in areas of the Northwest Atlantic. Fishery Bulletin, 84(3):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/node/
14295;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/parsons%20(1).pdf.

Reis:1986:AGM

[3296] Enir Girondi Reis. Age and growth of the marine catfish, Netuma barba (Siluriformes, Ariidae), in the estuary of the Patos Lagoon (Brasil). Fishery Bulletin, 84(3):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/node/
14303;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/reis.pdf.

Stier:1986:MSR

[3297] Kathleen Stier and Boyd Kynard. Movement of sea-run sea lampreys, Petromyzon marinus, during the spawning migration in the Connecticut River. Fishery Bulletin, 84(3):??, 1986. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14311;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/stier%20(1).pdf.

Wahlen:1986:IDM

[3298] Bruce E. Wahlen. Incidental dolphin mortality in the eastern tropical Pacific tuna fishery, 1973 through 1978. Fishery Bulletin, 84 (3):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14296;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/wahlen.pdf.

Dandonneau:1986:MSS

[3299] Yves Dandonneau. Monitoring the sea surface chlorophyll concentration in the tropical Pacific: consequences of the 1982–83 El Niño. Fishery Bulletin, 84(3):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14304;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/dandonneau.pdf.

Hogans:1986:VMF

[3300] W. E. Hogans and P. C. F. Hurley. Variations in the morphology of Fistulicola plicatus Rudolphi (1802) (Cestoda:Pseudophyllidea) from the swordfish, Xiphias gladius L., in the Northwest Atlantic Ocean. Fishery Bulletin, 84(3):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/node/
14312;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/hogans.pdf.

Prince:1986:LAV

[3301] Eric D. Prince, Dennis W. Lee, Charles A. Wilson, and John M. Dean. Longevity and age validation of a tag-recaptured Atlantic sailfish, Istiophorus platypterus, using dorsal spines and otoliths. Fishery Bulletin, 84(3):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14289;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/prince.pdf.

Somerton:1986:DRB

[3302] David A. Somerton and Robert S. Otto. Distribution and reproductive biology of the golden king crab, Lithodes aequispina, in the eastern Bering Sea. Fishery Bulletin, 84(3):??, 1986. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14297;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/somerton_3.pdf.

Rogers:1986:ATS

[3303] S. Gordon Rogers, Hiram T. Langston, and Timothy E. Targett. Anatomical trauma to sponge-coral reef fishes captured by trawling and angling. Fishery Bulletin, 84(3):697–704, 1986. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14305;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/rogers_0.pdf.

Parrish:1986:ADF

[3304] Richard H. Parrish, Donna L. Mallicoate, and Richard A. Klingbeil. Age dependent fecundity, number of spawnings per year, sex ratio, and maturation stages in northern anchovy, Engraulis mordax. Fishery Bulletin, 84(3):503–517, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/node/14290;   https://spo.nmfs.noaa.gov/node/
14291;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/parrish.pdf;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/parrish_0.pdf.

Power:1986:MDN

[3305] James H. Power. A model of the drift of northern anchovy, Engraulis mordax, larvae in the California Current. Fishery Bulletin, 84(3):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14298;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/power.pdf.

Quast:1986:APE

[3306] Jay C. Quast. Annual production of eviscerated body weight, fat, and gonads by Pacific herring, Clupea harengus pallasi, near Auke Bay, southeastern Alaska. Fishery Bulletin, 84(3):??, 1986. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14306;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/quast_0.pdf.

Johnson:1986:PBAb

[3307] Phyllis T. Johnson. Parasites of benthic amphipods: dinoflagellates (Duboscquodinida: Syndinidae). Fishery Bulletin, 84(3):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14299;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/johnson_2.pdf.

Wenner:1986:CLH

[3308] Charles A. Wenner, William A. Roumillat, and C. Wayne Waltz. Contributions to the life history of black sea bass, Centropristis striata, off the southeastern United States. Fishery Bulletin, 84(3): 723–741, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14307;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/wenner_0.pdf.

Bodkin:1986:FAM

[3309] James Lee Bodkin. Fish assemblages in macrocystis and nereocystis kelp forests off Central California. Fishery Bulletin, 84 (4):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/fish-
assemblages-macrocystis-and-nereocystis-kelp-forests-central-
california;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1986/844/bodkin.pdf.

McGowan:1986:NAE

[3310] Michael F. McGowan. Northern anchovy, Engraulis mordax, spawning in San Francisco Bay, California, 1978–79, relative to hydrography and zooplankton prey of adults and larvae. Fishery Bulletin, 84(4):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/northern-anchovy-engraulis-mordax-spawning-san-
francisco-bay-california-1978-79-relative;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1986/844/mcgowan.pdf.

Tester:1986:IIL

[3311] Patricia A. Tester and Andrew G. Carey, Jr. Instar identification and life history aspects of juvenile deepwater spider crabs, Chionoecetes tannerl Rathbun. Fishery Bulletin, 84(4):??, 1986. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/instar-identification-and-life-
history-aspects-juvenile-deepwater-spider-crabs-chionoecetes;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1986/844/tester.pdf.

Dailey:1986:ARB

[3312] Murray D. Dailey and Stephen Ralston. Aspects of the reproductive biology, spatial distribution, growth, and mortality of the deepwater caridean shrimp, Heterocarpus laevigatus, in Hawaii. Fishery Bulletin, 84(4):

915–925, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/aspects-
reproductive-biology-spatial-distribution-growth-and-mortality-
deepwater-caridean;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1986/844/dailey.pdf.

Pauly:1986:SME

[3313] Daniel Pauly. A simple method for estimating the food consumption of fish populations from growth data and food conversion experiments. Fishery Bulletin, 84(4):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/simple-method-estimating-food-consumption-fish-
populations-growth-data-and-food-conversion;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1986/844/pauly.
pdf.

Yang:1986:GBS

[3314] W. T. Yang, R. F. Hixon, P. E. Turk, M. E. Krejci, W. H. Hulet, and R. T. Hanlon. Growth, behavior, and sexual maturation of the market squid, Loligo opalescens, cultured through the life cycle. Fishery Bulletin, 84(4):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/growth-behavior-and-sexual-maturation-market-squid-
loligo-opalescens-cultured-through-life;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1986/844/yang.pdf.

Ditty:1986:INW

[3315] James G. Ditty. Ichthyoplankton in neritic waters of the Northern Gulf of Mexico off Louisiana: Composition, relative abundance, and seasonality. Fishery Bulletin, 84(4):??, 1986. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/ichthyoplankton-neritic-waters-
northern-gulf-mexico-louisiana-composition-relative;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1986/844/
ditty.pdf.

Perez:1986:DNS

[3316] Michael A. Perez and Michael A. Bigg. Diet of northern fur seals, Callorhinus ursinus, off Western North America. Fishery Bulletin, 84(4): ??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/diet-northern-
fur-seals-callorhinus-ursinus-western-north-america;   https:

//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1986/844/
perez.pdf.

Cody:1986:CCM

[3317] Terry J. Cody and Billy E. Fuls. Comparison of catches in 4.3 m and 12.2 m shrimp trawls in the Gulf of Mexico. Fishery Bulletin, 84(4):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14313;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/cody.pdf.

Farley:1986:NOE

[3318] C. A. Farley, S. V. Otto, and C. L. Reinisch. New occurrence of epizootic sarcoma in Chesapeake Bay soft shell clams, Mya arenaria. Fishery Bulletin, 84(4):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
new-occurrence-epizootic-sarcoma-chesapeake-bay-soft-shell-
clams-mya-arenaria;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1986/844/farley.pdf.

Polovina:1986:AYA

[3319] Jeffrey J. Polovina and Stephen Ralston. An approach to yield assessment for unexploited resources with application to the deep swpe fishes of the Marianas. Fishery Bulletin, 84(4):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/approach-yield-assessment-unexploited-resources-
application-deep-swpe-fishes-marianas;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1986/844/polovina.pdf.

Powell:1986:ELH

[3320] Allyn B. Powell and Germano Phonlor. Early life history of Atlantic menhaden, Brevoortia tyrannus, and Gulf menhaden, B. patronus. Fishery Bulletin, 84(4):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/node/
14314;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/powell_0.pdf.

Finucane:1986:RBK

[3321] John H. Finucane, L. Alan Collins, Harold A. Brusher, and Carl H. Saloman. Reproductive biology of king mackerel, Scomberomorus cavalla, from the Southeastern United States. Fishery Bulletin, 84(4):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/reproductive-biology-
king-mackerel-scomberomorus-cavalla-southeastern-united-states;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1986/844/finucane.pdf.

Ralston:1986:IFE

[3322] Stephen Ralston. An intensive fishing experiment for the caridean shrimp, Heterocarpus laevigatus, at Alamagan Island in the Mariana Archipelago. Fishery Bulletin, 84(4):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/intensive-fishing-experiment-caridean-shrimp-
heterocarpus-laevigatus-alamagan-island;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1986/844/ralston.pdf.

Podniesinski:1986:SBM

[3323] Greg S. Podniesinski and Bernard J. McAlice. Seasonality of blue mussel, Mytilus edulis L., larvae in the Damariscotta River estuary, Maine, 1969–77. Fishery Bulletin, 84(4):??, 1986. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14315;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/PODNIESINSKI.pdf.

Folkvord:1986:SSV

[3324] Arild Folkvord and John R. Hunter. Size-specific vulnerability of northern anchovy, Engraulis mordax, larvae to predation by fishes. Fishery Bulletin, 84(4):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/size-specific-vulnerability-northern-
anchovy-engraulis-mordax-larvae-predation-fishes;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1986/844/
folkvord.pdf.

Rexstad:1986:SCF

[3325] Eric A. Rexstad and Ellen K. Pikitch. Stomach contents and food consumption estimates of Pacific hake, Merluccius productus. Fishery Bulletin, 84(4):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
stomach-contents-and-food-consumption-estimates-pacific-hake-
merluccius-productus;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1986/844/rexstad.pdf.

Anonymous:1986:I

[3326] Anonymous. Index. Fishery Bulletin, 84(4):??, 1986. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14316;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/index_7.pdf.

Houde:1986:SGS

[3327] Edward D. Houde and Lawrence Lubbers III. Survival and growth of striped bass, Morone saxatilis, and Morone hybrid larvae: Laboratory and pond enclosure experiments. Fishery Bulletin, 84(4): ??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/survival-
and-growth-striped-bass-morone-saxatilis-and-morone-hybrid-
larvae-laboratory-and;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1986/844/houde.pdf.

Smith:1986:ORB

[3328] Joseph W. Smith and John V. Merriner. Observations on the reproductive biology of the cownose ray, Rhinoptera bonasus, in Chesapeake Bay. Fishery Bulletin, 84(4):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/observations-reproductive-biology-cownose-
ray-rhinoptera-bonasus-chesapeake-bay;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1986/844/smith.pdf.

Hunter:1986:SFS

[3329] J. Roe Hunter, Beverly J. Macewicz, and John R. Sibert. The spawning frequency of skipjack tuna, Katsuwonus pelamis, from the South Pacific. Fishery Bulletin, 84(4):??, 1986. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/spawning-frequency-skipjack-tuna-
katsuwonus-pelamis-south-pacific;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1986/844/hunter.pdf.

Stepien:1986:LHL

[3330] Carol A. Stepien. Life history and larvel development of the giant kelpfish, Heterostichus rostratus Girard, 1854. Fishery Bulletin, 84(4):??, 1986. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/life-history-
and-larvel-development-giant-kelpfish-heterostichus-rostratus-
girard-1854;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1986/844/stepien.pdf.

Grover:1987:ENN

[3331] Jill J. Grover and Bori L. Olla. Effects of an El Niño event on the food habits of larval sablefish, Anoplopoma fimbria, off Oregon and Washington. Fishery Bulletin, 85(1):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/effects-el-nino-event-food-habits-larval-sablefish-
anoplopoma-fimbria-oregon-and-washington;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1987/851/grover.pdf.

Warner:1987:SFS

[3332] John Warner and Boyd Kynard. Scavenger feeding by subadult striped bass, Morone saxatilis, below a low-head hydroelectric dam. Fishery Bulletin, 85(1):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/scavenger-feeding-subadult-striped-bass-morone-
saxatilis-below-low-head-hydroelectric-dam;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/warner%20(
1).pdf.

Hanlon:1987:LRS

[3333] Roger T. Hanlon, Philip E. Turk, Phillip G. Lee, and Won Tack Yang. Laboratory rearing of the squid Loligo pealei to the juvenile stage: growth comparisons with fishery data. Fishery Bulletin, 85(1):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14324;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/hanlon.pdf.

Love:1987:LHF

[3334] Milton S. Love, Brita Axell, Pamela Morris, Robson Collins, and Andrew Brooks. Life history and fishery of the California scorpionfish, Scorpaena guttata, within the Southern California Bight. Fishery Bulletin, 85(1):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
life-hisiory-and-fishery-california-scorpionfish-scorpaena-
guttata-within-southern;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1987/851/love.pdf.

Horwood:1987:BVA

[3335] J. W. Horwood. Bias and variance in Allen’s recruitment rate method. Fishery Bulletin, 85(1):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14317;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/horwood.pdf.

Chapman:1987:CPS

[3336] Robert W. Chapman. Changes in the population structure of male striped bass, Morone saxatilis, spawning in the three areas of the Chesapeake Bay from 1984 to 1986. Fishery Bulletin, 85(1):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14325;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/chapman_0.pdf.

Lutz:1987:VBC

[3337] Peter L. Lutz and Ann Dunbar-Cooper. Variations in the blood chemistry of the loggerhead sea turtle, Caretta caretta. Fishery Bulletin, 85(1):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/variations-blood-chemistry-loggerhead-sea-
turtle-caretta-caretta;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1987/851/lutz.pdf.

Silvert:1987:CNE

[3338] W. Silvert and D. Pauly. On the compatibility of a new expression for gross conversion efficiency with the von Bertalanffy growth equation. Fishery Bulletin, 85(1):??, 1987. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14318;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/silvert.pdf.

Minello:1987:FPJ

[3339] Thomas J. Minello, Roger J. Zimmerman, and Eduardo X. Martinez. Fish predation on juvenile brown shrimp, Penaeus aztecus Ives: Effects of turbidity and substratum on predation rates. Fishery Bulletin, 85(1):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/fish-
predation-juvenile-brown-shrimp-penaeus-aztecus-ives-effects-
turbidity-and-substratum;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1987/851/minello.pdf.

McCabe:1987:ERD

[3340] George T. McCabe, Jr., Robert L. Emmett, Travis C. Coley, and Robert J. McConnell. Effect of a river-dominated estuary on the prevalence of carcino nemertes errans, an egg predator of the Dungeness crab, Cancer magister. Fishery Bulletin, 85(1):??, 1987. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14319;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/mccabe_0.pdf.

Utter:1987:GES

[3341] Fred Utter, David Teel, George Milner, and Donald McIsaac. Genetic estimates of stock compositions of 1983 Chinook salmon, Oncorhynchus tshawytscha, harvests off the Washington Coast and the Columbia River. Fishery Bulletin, 85(1):13–23, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/genetic-estimates-stock-compositions-1983-
chinook-salmon-oncorhynchus-tshawytscha-harvests;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1987/851/
utter.pdf.

Kynard:1987:SSM

[3342] Boyd Kynard and John P. Warner. Spring and summer movements of subadult striped bass, Morone saxatilis, in the Connecticut River. Fishery Bulletin, 85(1):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/node/
14320;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/kynard.pdf.

Brill:1987:SMR

[3343] Richard W. Brill. On the standard metabolic rates of tropical tunas, including the effect of body size and acute temperature change. Fishery Bulletin, 85(1):??, 1987. CODEN FSY-

BAY.  ISSN 0090-0656 (print), 1937-4518 (electronic).  URL

https://spo.nmfs.noaa.gov/content/standard-metabolic-rates-
tropical-tunas-including-effect-body-size-and-acute-temperature;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1987/851/brill.pdf.

Vermeer:1987:EAE

[3344] Gregory K. Vermeer. Effects of air exposure on desiccation rate, hemolymph chemistry, and escape behavior of the spiny lobster, Panulirus argus. Fishery Bulletin, 85(1):??, 1987. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/effects-air-exposure-desiccation-
rate-hemolymph-chemistry-and-escape-behavior-spiny-lobster;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1987/851/vermeer.pdf.

Walsh:1987:HPS

[3345] S. J. Walsh. Habitat partitioning by size in witch flounder, Glyptocephalus cynoglossus: a reevaluation with additional data and adjustments for gear selectivity. Fishery Bulletin, 85(1):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14321;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/walsh_1.pdf.

Clarke:1987:FSF

[3346] Thomas A. Clarke. Fecundity and spawning frequency of the Hawaiian anchovy or nehu, Encrasicholina purpurea. Fishery Bulletin, 85(1): ??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/fecundity-and-
spawning-frequency-hawaiian-anchovy-or-nehu-encrasicholina-
purpurea;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1987/851/clarke.pdf.

Waples:1987:PLD

[3347] Robin S. Waples and Richard H. Rosenblatt. Patterns of larval drift in Southern California marine shore fishes inferred from allozyme data. Fishery Bulletin, 85(1):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/patterns-larval-drift-southern-california-
marine-shore-fishes-inferred-allozyme-data;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1987/851/waples.pdf.

Mathews:1987:MTL

[3348] S. R. Mathews and M. LaRiviere. Movement of tagged lingcod, Ophiodon elongatus, in the Pacific northwest. Fishery Bulletin, 85 (1):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14322;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/mathews_0.pdf.

Gartner:1987:LPM

[3349] John V. Gartner, Jr., Thomas L. Hopkins, Ronald C. Baird, and Dean M. Milliken. The lanternfishes (Pisces: Myctophidae) of the Eastern Gulf of Mexico. Fishery Bulletin, 85(1):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/lanternfishes-
pisces-myctophidae-eastern-gulf-mexico;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1987/851/gartner.pdf.

Wilson:1987:JBC

[3350] K. A. Wilson, K. L. Heck, Jr., and K. W. Able. Juvenile blue crab, Callinectes sapidus, survival: an evaluation of eelgrass, Zostera marina, as refuge. Fishery Bulletin, 85(1):??, 1987. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/juvenile-blue-crab-callinectes-
sapidus-survival-evaluation-eelgrass-zostera-marina-refuge;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1987/851/wilson.pdf.

Vondracek:1987:DRG

[3351] Bruce Vondracek. Digestion rates and gastric evacuation times in relation to temperature of the Sacramento squawfish, Ptychocheilus grandis. Fishery Bulletin, 85(1):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14323;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/vondracek.pdf.

Hales:1987:DAR

[3352] L. Stanton Hales, Jr. Distribution, abundance, reproduction, food habits, age, and growth of round scad, Decapterus punctatus, in the South Atlantic Bight. Fishery Bulletin, 85(2):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/distribution-abundance-reproduction-food-habits-
age-and-growth-round-scad-decapterus;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1987/852/hales.pdf.

York:1987:ENN

[3353] Anne E. York and Patrick Kozloff. On the estimation of the numbers of northern fur seal, Callorhinus ursinus, pups born on St. Paul Island, 1980–86. Fishery Bulletin, 85(2):??, 1987. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14328;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/york.pdf.

Iversen:1987:BDB

[3354] Edwin S. Iversen, Edward S. Rutherford, Scott P. Bannerot, and Darryl E. Jory. Biological data on Berry Islands (Bahamas) queen conchs, Strombus gigas, with mariculture and fisheries management implications. Fishery Bulletin, 85(2):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/biological-data-berry-islands-bahamas-queen-conchs-
strombus-gigas-mariculture-and-fisheries;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1987/852/iversen.pdf.

Borowsky:1987:LSP

[3355] Betty Borowsky. Laboratory studies of the pattern of reproduction of the isopod crustacean idotea baltica. Fishery Bulletin, 85 (2):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14329;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/borowsky.pdf.

Lai:1987:OAE

[3356] Han-Lin Lai. Optimum allocation for estimating age composition using age–length key. Fishery Bulletin, 85(2):??, 1987. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/optimum-allocation-
estimating-age-composition-using-age-length-key;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1987/852/
lai.pdf.

OLeary:1987:OFF

[3357] John O’Leary and Douglas G. Smith. Occurrence of the first freshwater migration of the gizzard shad, Dorosoma cepedianum, in the Connecticut River, Massachusetts. Fishery Bulletin, 85(2):??, 1987. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14330;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/oleary.pdf.

Branstetter:1987:CAG

[3358] Steven Branstetter, J. A. Musick, and J. A. Colvocoresses. A comparison of the age and growth of the tiger shark, Galeocerdo cuvieri, from off Virginia and from the Northwestern Gulf of Mexico. Fishery Bulletin, 85(2): ??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/comparison-
age-and-growth-tiger-shark-galeocerdo-cuvieri-virginia-and-
northwestern-gulf;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1987/852/branstetter.pdf.

Merrick:1987:DAN

[3359] Richard L. Merrick, Thomas R. Loughlin, and Donald G. Calkins. Decline in abundance of the northern sea lion, Eumetopias jubatus, in Alaska, 1956–86. Fishery Bulletin, 85(2):??, 1987. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/decline-abundance-northern-sea-
lion-eumetopias-jubatus-alaska-1956-86;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1987/852/merrick.pdf.

Echeverria:1987:ROL

[3360] Tina Wyllie Echeverria. Relationship of otolith length to total length in rockfishes from northern and central California. Fishery Bulletin, 85(2):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14331;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/echeverria2.pdf.

Echeverria:1987:TFS

[3361] Tina Wyllie Echeverria. Thirty-four species of California rockfishes: Maturity and seasonality of reproduction. Fishery Bulletin, 85(2):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/thirty-four-species-
california-rockfishes-maturity-and-seasonality-reproduction;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1987/852/echeverria.pdf.

Moffitt:1987:DYD

[3362] Robert B. Moffitt and Jeffrey J. Polovina. Distribution and yield of the deepwater shrimp Heterocarpus resource in the Marianas. Fishery Bulletin, 85(2):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
distribution-and-yield-deepwater-shrimp-heterocarpus-resource-
marianas;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1987/852/moffitt.pdf.

LeBoeuf:1987:CWN

[3363] Burney J. Le Boeuf, John E. McCosker, and John Hewitt. Crater wounds on northern elephant seals: the cookiecutter shark strikes again. Fishery Bulletin, 85(2):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/node/
14332;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/leboeuf_0.pdf.

Farfante:1987:RGP

[3364] Isabel Perez Farfante. Revision of the Gamba prawn genus Pseudaristeus, with description of two new species (Crustacea: Decapoda: Penaeoidea). Fishery Bulletin, 85(2):??, 1987. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/revision-gamba-prawn-
genus-pseudaristeus-description-two-new-species-crustacea-decapoda;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1987/852/farfante.pdf.

Peterson:1987:ECM

[3365] Charles H. Peterson, Henry C. Summerson, and Stephen R. Fegley. Ecological consequences of mechanical harvesting of clams. Fishery Bulletin, 85(2):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
ecological-consequences-mechanical-harvesting-clams;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1987/852/
peterson.pdf.

Feeney:1987:DEL

[3366] Richard F. Feeney. Development of the eggs and larvae of the yellowchin sculpin, Icelinus quadriseriatus (Pisces: Cottidae). Fishery Bulletin, 85 (2):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/development-
eggs-and-larvae-yellowchin-sculpin-icelinus-quadriseriatus-
pisces-cottidae;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1987/852/feeney.pdf.

Jones:1987:VOI

[3367] Cynthia Jones and Edward B. Brothers. Validation of the otolith increment aging technique for striped bass, Morone saxatilis, larvae reared under suboptimal feeding conditions. Fishery Bulletin, 85(2):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14326;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/jones_1.pdf.

Greeley:1987:OGD

[3368] Mark S. Greeley, Jr., Daniel R. Calder, and Robin A. Wallace. Oocyte growth and development in the striped mullet, Mugil cephalus, during seasonal ovarian recrudescence: Relationship to fecundity and size at maturity. Fishery Bulletin, 85(2):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/oocyte-growth-
and-development-striped-mullet-mugil-cephalus-during-seasonal-
ovarian;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1987/852/greeley.pdf.

Theilacker:1987:FEG

[3369] Gail H. Theilacker. Feeding ecology and growth energetics of larval northern anchovy, Engraulis mordax. Fishery Bulletin, 85(2): ??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14327;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/theilacker%20(1).pdf.

Groot:1987:HMS

[3370] C. Groot and T. P. Qulnn. Homing migration of sockeye salmon, Oncorhynchus nerka, to the Fraser River. Fishery Bulletin, 85(3):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/homing-migration-sockeye-
salmon-oncorhynchus-nerka-fraser-river;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1987/853/groot.pdf.

Wenner:1987:EGC

[3371] Elizabeth L. Wenner, Glenn F. Ulrich, and John B. Wise. Exploration for golden crab, Geryon fennerl, in the South Atlantic Bight: Distribution, population structure, and gear assessment. Fishery Bulletin, 85(3):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/exploration-golden-
crab-geryon-fennerl-south-atlantic-bight-distribution-population;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1987/853/wenner.pdf.

Collins:1987:CLH

[3372] Mark R. Collins, C. Wayne Waltz, William A. Roumillat, and Daryl L. Stubbs. Contribution to the life history and reproductive biology of gag, Mycteroperca microlepis (Serranidae), in the South Atlantic Bight. Fishery Bulletin, 85(3):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/node/
14340;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/collins_2.pdf.

Hinckley:1987:RBW

[3373] Sarah Hinckley. The reproductive biology of walleye pollock, Theragra chalcogramma, in the Bering Sea, with reference to spawning stock structure. Fishery Bulletin, 85(3):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/reproductive-biology-walleye-pollock-
theragra-chalcogramma-bering-sea-reference-spawning;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1987/853/
hinckley.pdf.

Best:1987:ELC

[3374] Peter B. Best. Estimates of the landed catch of right (and other whalebone) whales in the American fishery, 1805–1909. Fishery Bulletin, 85(3):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14333;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/best.pdf.

Ropes:1987:AGR

[3375] John Ropes. Age and growth, reproductive cycle, and histochemical tests for heavy metals in hard clams, Mercenaria mercenaria, from raritan bay, 1974–75. Fishery Bulletin, 85(3):??, 1987. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14341;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/ropes_1.pdf.

Holt:1987:EDD

[3376] Rennie S. Holt. Estimating density of dolphin schools in the Eastern Tropical Pacific Ocean by line transect methods. Fishery Bulletin, 85(3): ??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/estimating-
density-dolphin-schools-eastern-tropical-pacific-ocean-line-
transect-methods;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1987/853/holt.pdf.

Armetta:1987:ABH

[3377] Therese M. Armetta and Bradley G. Stevens. Aspects of the biology of the hair crab, Erimacrus isenbeckii, in the eastern Bering Sea. Fishery Bulletin, 85(3):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/node/
14334;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/armetta.pdf.

Holt:1987:RVS

[3378] Rennie S. Holt, Tim Gerrodette, and John B. Cologne. Research vessel survey design for monitoring dolphin abundance in the Eastern Tropical Pacific. Fishery Bulletin, 85(3):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/research-vessel-survey-design-monitoring-dolphin-
abundance-eastern-tropical-pacific;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1987/853/holt2.pdf.

Ahrenholz:1987:PFC

[3379] Dean W. Ahrenholz, Walter R. Nelson, and Sheryan P. Epperly. Population and fishery characteristics of Atlantic menhaden, Brevoortia tyrannus. Fishery Bulletin, 85(3):??, 1987. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14335;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/ahrenholz.pdf.

Kendall:1987:DFG

[3380] A. W. Kendall, Jr., M. E. Clarke, M. M. Yoklavich, and G. W. Boehlert. Distribution, feeding, and growth of larval walleye pollock, Theragra chalcogramma, from Shelikof Strait, Gulf of Alaska. Fishery Bulletin, 85(3):499–521, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/distribution-feeding-and-growth-larval-walleye-
pollock-theragra-chalcogramma-shelikof;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1987/853/kendall.pdf.

Bowering:1987:DWF

[3381] W. R. Bowering. Distribution of witch flounder, Glyptocephalus cynoglossus, in the southern Labrador and eastern Newfoundland area and changes in certain biological parameters after 20 years of exploitation. Fishery Bulletin, 85(3):??, 1987. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14336;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/bowering_0.pdf.

Becker:1987:PCS

[3382] D. Scott Becker and Kenneth K. Chew. Predation on Capitella spp. by small-mouthed pleuronectids in Puget Sound, Washington. Fishery Bulletin, 85(3):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
predation-capitella-spp-small-mouthed-pleuronectids-puget-sound-
washington;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1987/853/becker.pdf.

Kornfield:1987:DMD

[3383] I. Kornfield and S. M. Bogdanowicz. Differentiation of mitochondrial DNA in Atlantic herring, Clupea harengus. Fishery Bulletin, 85(3):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/differentiation-mitochondrial-dna-atlantic-herring-
clupea-harengus;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1987/853/kornfield.pdf.

Mercaldo-Allen:1987:HGV

[3384] Renee Mercaldo-Allen and Frederick P. Thurberg. Heart and gill ventilatory activity in the lobster, Homarus americanus, at various temperatures. Fishery Bulletin, 85(3):??, 1987. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14337;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/mercaldo-allen.pdf.

Butler:1987:TSM

[3385] Richard W. Butler, Walter A. Nelson, and Tyrrell A. Henwood. A trawl survey method for estimating loggerhead turtle, Caretta caretta, abundance in five Eastern Florida channels and inlets. Fishery Bulletin, 85(3):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/trawl-
survey-method-estimating-loggerhead-turtle-caretta-caretta-
abundance-five-eastern;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1987/853/butler.pdf.

Mugiya:1987:PDB

[3386] Yasuo Mugiya. Phase difference between calcification and organic matrix formation in the diurnal growth of otoliths in the rainbow trout, Salmo gairdneri. Fishery Bulletin, 85(3):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/phase-difference-between-
calcification-and-organic-matrix-formation-diurnal-growth-otoliths;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1987/853/mugiya.pdf.

Foreman:1987:MST

[3387] Terry Foreman. A method of simultaneously tagging large oceanic fish and injecting them with tetracycline. Fishery Bulletin, 85 (3):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14338;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/foreman.pdf.

Dalley:1987:ELH

[3388] E. L. Dalley and G. H. Winters. Early life history of sand lance (Ammodytes), with evidence for spawning of A. dubius in Fortune Bay, Newfoundland. Fishery Bulletin, 85(3):??, 1987. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/early-life-history-
sand-lance-ammodytes-evidence-spawning-dubius-fortune-bay-newfoundland;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1987/853/dalley.pdf.

Sogard:1987:DDD

[3389] Susan M. Sogard, Donald E. Hoss, and John J. Govoni. Density and depth distribution of larval Gulf menhaden, Brevoortia patronus, Atlantic croaker, Micropogonias undulatus, and spot, Leiostomus xanthurus, in the Northern Gulf of Mexico. Fishery Bulletin, 85(3):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/density-and-
depth-distribution-larval-gulf-menhaden-brevoortia-patronus-
atlantic-croaker;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1987/853/sogard.pdf.

Schaefer:1987:SRK

[3390] Kurt M. Schaefer. Second record of the kawakawa, Euthynnus affinis, from the eastern Pacific Ocean. Fishery Bulletin, 85(3):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14339;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/schaefer.pdf.

Able:1987:SST

[3391] K. W. Able, D. C. Twichell, C. B. Grimes, and R. S. Jones. Sidescan sonar as a tool for detection of demersal fish habitats. Fishery Bulletin, 85 (4):725–736, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/sidescan-
sonar-tool-detection-demersal-fish-habitats;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1987/854/able.
pdf.

ONeil:1987:FHS

[3392] Steven P. O’Neil and Michael P. Weinstein. Feeding habitats of spot, Leiostomus xanthurus, in polyhaline versus meso-oligohaline tidal creeks and shoals. Fishery Bulletin, 85(4):??, 1987. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/feeding-habitats-spot-leiostomus-
xanthurus-polyhaline-versus-meso-oligohaline-tidal-creeks;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1987/
854/oneil.pdf.

Yoklavich:1987:DGI

[3393] Mary M. Yoklavich and George W. Boehlert. Daily growth increments in otoliths of juvenile black rockfish, Sebastes melanops: an evaluation of autoradiography as a new method of validation. Fishery Bulletin, 85(4):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14346;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/yoklavich.pdf.

Appeldoorn:1987:AMO

[3394] Richard S. Appeldoorn. Assessment of mortality in an offshore population of queen conch, Strombus gigas L., in southwest Puerto Rico. Fishery Bulletin, 85(4):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/assessment-mortality-offshore-population-queen-
conch-strombus-gigas-l-southwest-puerto-rico;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1987/854/appeldoorn.
pdf.

Quast:1987:MVP

[3395] Jay C. Quast. Morphometric variation of Pacific Ocean perch, Sebastes alutus, off Western North America. Fishery Bulletin,

85(4):??, 1987.  CODEN FSYBAY.  ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
morphometric-variation-pacific-ocean-perch-sebastes-alutus-
western-north-america;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1987/854/quast.pdf.

Anonymous:1987:I

[3396] Anonymous. Index. Fishery Bulletin, 85(4):??, 1987. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14347;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/index%20(1)_1.pdf.

Brill:1987:PEI

[3397] Richard W. Brill, Robert Bourke, James A. Brock, and Murray D. Dailey. Prevalence and effects of infection of tile dorsal aorta in yellowfin tuna, Thunnus albacares, by the larval cestode, Dasyrhynchus talismani. Fishery Bulletin, 85(4):767–776, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/prevalence-and-effects-infection-tile-dorsal-aorta-
yellowfin-tuna-thunnus-albacares-larval;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1987/854/brill.pdf.

Reisenbichler:1987:GVC

[3398] R. R. Reisenbichler and S. R. Phelps. Genetic variation in Chinook, Oncorhynchus tshawytscha, and Coho, O. kisutch, salmon from the north coast of Washington. Fishery Bulletin, 85(4): ??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/genetic-
variation-chinook-oncorhynchus-tshawytscha-and-coho-0-kisutch-
salmon-north-coast;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1987/854/reisenbichler.pdf.

Fable:1987:AGS

[3399] William A. Fable, Jr., Allyn G. Johnson, and Lyman E. Barger. Age and growth of Spanish mackerel, Scomberomorus maculatus, from Florida and the Gulf of Mexico. Fishery Bulletin, 85(4):??, 1987. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/age-and-growth-spanish-mackerel-
scomberomorus-maculatus-florida-and-gulf-mexico;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1987/854/
fable.pdf.

Smith:1987:EBT

[3400] Eric M. Smith and Penelope T. Howell. The effects of bottom trawling on American lobsters, Homarus americanus, in Long Island Sound. Fishery Bulletin, 85(4):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/effects-bottom-trawling-american-lobsters-
homarus-americanus-long-island-sound;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1987/854/smith.pdf.

Kleiber:1987:AIB

[3401] P. Kleiber and B. Baker. Assessment of interaction between North Pacific albacore, Thunnus alalunga, fisheries by use of a simulation model. Fishery Bulletin, 85(4):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/assessment-interaction-between-north-
pacific-albacore-thunnus-alalunga-fisheries-use;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1987/854/
kleiber.pdf.

Botton:1987:PHC

[3402] Mark L. Botton and John W. Ropes. Populations of horseshoe crabs, Limulus polyphemus, on the northwestern Atlantic continental shelf. Fishery Bulletin, 85(4):??, 1987. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14342;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/botton_0.pdf.

Kudo:1987:FAC

[3403] George Kudo, Harold J. Barnett, and Richard W. Nelson. Factors affecting cooked texture quality of Pacific whiting, Merluccius productus, fillets with particular emphasis on the effects of infection by the myxosporeans Kudoa paniformis and K. thyrsitis. Fishery Bulletin, 85(4):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/factors-affecting-cooked-
texture-quality-pacific-whiting-merluccius-productus-fillets;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1987/854/kudo.pdf.

Henwood:1987:AST

[3404] Tyrrell A. Henwood and Warren E. Stuntz. Analysis of sea turtle captures and mortalities during commercial shrimp trawling. Fishery Bulletin, 85(4):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14343;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/henwood.pdf.

Lai:1987:ADP

[3405] Han-Lin Lai, Donald R. Gunderson, and Loh Lee Low. Age determination of Pacific cod, Gadus macrocephalus, using five ageing methods. Fishery Bulletin, 85(4):713–723, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/age-determination-pacific-cod-gadus-macrocephalus-
using-five-ageing-methods;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1987/854/lai.pdf.

Render:1987:RBL

[3406] Jeffrey H. Render and Robert L. Allen. The relationship between lunar phase and Gulf butterfish, Peprilus burti catch rate. Fishery Bulletin, 85(4):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14344;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/render.pdf.

Leis:1987:DAB

[3407] Jeffrey M. Leis, Barry Goldman, and Shoji Ueyanagi. Distribution and abundance of billfish larvae (Pisces: Istiophoridae) in the Great Barrier Reef Lagoon and coral sea near Lizard Island, Australia. Fishery Bulletin, 85(4):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
distribution-and-abundance-billfish-larvae-pisces-istiophoridae-
great-barrier-reef-lagoon;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1987/854/leis.pdf.

Fisher:1987:MCO

[3408] J. P. Fisher and W. C. Pearcy. Movements of Coho, Oncorhynchus kisutch, and Chinook, O. tshawytscha, salmon tagged at sea off Oregon, Washington, and Vancouver Island during the summers 1982–85. Fishery Bulletin, 85(4):??, 1987. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/node/
14345;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fisher_0.pdf.

Loeb:1988:IVI

[3409] Valerie J. Loeb and Omar Rojas. Interannual variation of ichthyoplankton composition and abundance relations off Northern Chile, 1964–83. Fishery Bulletin, 86(1):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/interannual-variation-ichthyoplankton-composition-and-
abundance-relations-northern-chile;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1988/861/loeb.pdf.

Shirley:1988:AID

[3410] Susan M. Shirley and Thomas C. Shirley. Appendage injury in Dungeness crabs, Cancer magister, in southeastern Alaska. Fishery Bulletin, 86(1):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14350;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/shirley.pdf.

McEachron:1988:RBC

[3411] Lawrence W. McEachron, Jeff F. Doerzbacher, Gary C. Matlock, Albert W. Green, and Gary E. Saul. Reducing the bycatch in a commercial trotline fishery. Fishery Bulletin, 86(1):109– 117, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/reducing-
bycatch-commercial-trotline-fishery;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1988/861/mceachron.pdf.

Currens:1988:RUO

[3412] Kenneth P. Currens, Carl B. Schreck, and Hiram W. Li. Reexamination of the use of otolith nuclear dimensions to identify juvenile anadromous and nonanadromous rainbow trout, Salmo gairdneri. Fishery Bulletin, 86(1):160–163, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14351;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/currens.pdf.

Notarbartolo-Di-Sciara:1988:NHR

[3413] Giuseppe Notarbartolo-Di-Sciara. Natural history of the rays of the genus Mobula in the Gulf of California. Fishery Bulletin, 86(1):45–66, 1988. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
natural-history-rays-genus-mobula-gulf-california;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1988/861/
notar.pdf.

Butler:1988:ASV

[3414] John L. Butler and Darlene Pickett. Age-specific vulnerability of Pacific sardine, Sardillops sagax, larvae to predation by northern anchovy, Engraulis mordax. Fishery Bulletin, 86(1):??, 1988. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14352;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/currens_0.pdf.

OBrien:1988:SVC

[3415] Loretia O’Brien and Ralph K. Mayo. Sources of variation in catch per unit effort of yellowtail flounder, Llmanda ferruginea (Storer), harvested off the coast of New England. Fishery Bulletin, 86(1): ??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/sources-
variation-catch-unit-effort-yellowtail-flounder-llmanda-ferruginea-
storer-harvested;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1988/861/obrien.pdf.

Epifanio:1988:ISW

[3416] Charles E. Epifanio, David Goshorn, and Timothy E. Targett. Induction of spawning in the weakfish, Cynoscion regalis. Fishery Bulletin, 86 (1):168–171, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14353;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/epifanio.pdf.

Warlen:1988:AGL

[3417] Stanley M. Warlen. Age and growth of larval Gulf menhaden, Brevoortia patronus, in the Northern Gulf of Mexico. Fishery Bulletin, 86(1):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/age-and-growth-larval-gulf-menhaden-brevoortia-
patronus-northern-gulf-mexico;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1988/861/warlen.pdf.

Bruce:1988:LDB

[3418] B. D. Bruce. Larval development of blue grenadier, Macruronus novaezelandiae (Hector), in Tasmanian waters. Fishery Bulletin, 86(1):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/larval-development-
blue-grenadier-macruronus-novaezelandiae-hector-tasmanian-waters;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1988/861/bruce.pdf.

Williams:1988:NDE

[3419] Austin B. Williams. Notes on decapod and euphausiid crustaceans, Continental Margin, Western Atlantic, Georges Bank to Western Florida, USA. Fishery Bulletin, 86(1):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/notes-decapod-and-euphausiid-crustaceans-continental-
margin-western-atlantic-georges-bank;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1988/861/williams.pdf.

Cowan:1988:DAT

[3420] James H. Cowan, Jr. and Richard F. Shaw. The distribution, abundance, and transport of larval sciaenids collected during winter and early spring from the Continental Shelf waters off West Louisianaa. Fishery Bulletin, 86(1):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/distribution-abundance-and-transport-
larval-sciaenids-collected-during-winter-and-early;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1988/861/
cowan.pdf.

Vetter:1988:ENM

[3421] E. F. Vetter. Estimation of natural mortality in fish stocks: a review. Fishery Bulletin, 86(1):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14348;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/vetter.pdf.

Hamner:1988:BSR

[3422] William M. Hamner, Gregory S. Stone, and Bryan S. Obst. Behavior of southern right whales, Eubalaena australis, feeding on the Antarctic krill, Euphausia superba. Fishery Bulletin, 86(1):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/behavior-southern-right-whales-
eubalaena-australis-feeding-antarctic-krill-euphausia;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1988/861/
hamner.pdf.

Penson:1988:EAN

[3423] John B. Penson, Jr., Ernest O. Tetty, and Wade L. Griffin. An econometric analysis of net investment in Gulf shrimp fishing vessels. Fishery Bulletin, 86(1):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14349;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/penson.pdf.

Dutil:1988:EML

[3424] J.-D. Dutil and J.-M. Coutu. Early marine life of Atlantic salmon, Salmo salar, postsmolts in the Northern Gulf of St. Lawrence. Fishery Bulletin, 86(2):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/early-
marine-life-atlantic-salmon-salmo-salar-postsmolts-northern-
gulf-st-lawrence;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1988/862/dutil.pdf.

Polacheck:1988:ARB

[3425] Tom Polacheck. Analyses of the relationship between the distribution of searching effort, tuna catches, and dolphin sightings within individual purse seine cruises. Fishery Bulletin, 86(2):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/analyses-
relationship-between-distribution-searching-effort-tuna-catches-
and-dolphin;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1988/862/polacheck.pdf.

Mullin:1988:ELL

[3426] M. M. Mullin and E. R. Brooks. Extractable lipofuscin in larval marine fish. Fishery Bulletin, 86(2):??, 1988. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14359;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/mullin.pdf.

Herrnkind:1988:ESR

[3427] William F. Herrnkind, Mark J. Butler IV, and Richard A. Tankersley. The effects of siltation on recruitment of spiny lobsters, Panulirus argus. Fishery Bulletin, 86(2):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/effects-siltation-recruitment-spiny-lobsters-panulirus-
argus;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1988/862/herrnkind.pdf.

Shenker:1988:OAN

[3428] Jonathan M. Shenker. Oceanographic associations of neustonic larval and juvenile fishes and Dungeness crab Megalopae off Oregon. Fishery Bulletin, 86(2):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
oceanographic-associations-neustonic-larval-and-juvenile-fishes-
and-dungeness-crab;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1988/862/shenker.pdf.

Jahn:1988:RFS

[3429] A. E. Jahn, D. M. Gadomski, and M. L. Sowby. On the role of food-seeking in the suprabenthic habit of larval white croaker, Genyonemus uneatus (Pisces: Sciaenidae). Fishery Bulletin, 86(2): ??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/role-food-
seeking-suprabenthic-habit-larval-white-croaker-genyonemus-
uneatus-pisces;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1988/862/jahn.pdf.

Watson:1988:CTE

[3430] Cheryl Watson, Robert E. Bourke, and Richard W. Brill. A comprehensive theory on the etiology of burnt tuna. Fishery Bulletin, 86(2):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/comprehensive-theory-etiology-burnt-tuna;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1988/862/
watson.pdf.

Kirkley:1988:LIA

[3431] James E. Kirkley and Dale E. Squires. A limited information approach for determining capital stock and investment in a fishery. Fishery Bulletin, 86(2):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
limited-information-approach-determining-capital-stock-and-
investment-fishery;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1988/862/kirkley.pdf.

Williams:1988:NMD

[3432] Austin B. Williams. New marine decapod crustaceans from waters influenced by hydrothermal discharge, brine, and hydrocarbon seepage. Fishery Bulletin, 86(2):??, 1988. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/new-marine-decapod-crustaceans-
waters-influenced-hydrothermal-discharge-brine-and;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1988/862/
williams.pdf.

Martin:1988:MSG

[3433] Joel W. Martin, Frank M. Truesdale, and Darryl L. Felder. The megalopa stage of the Gulf stonecrab, Menippe adina Williams and Felder, 1986, with a comparison of megalopae in the genus Menippe. Fishery Bulletin, 86(2):289–297, 1988. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
megalopa-stage-gulf-stonecrab-menippe-adina-williams-and-felder-
1986-comparison-megalopae;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1988/862/martin.pdf.

Chen:1988:NRS

[3434] Che-Tsung Chen, Tzyh-Chang Leu, and Shoou-Jeng Joung. Notes on reproduction in the scalloped hammerhead, Sphyrna lewini, in northeastern Taiwan waters. Fishery Bulletin, 86(2):389–393, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14355;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/chen.pdf.

Bolz:1988:GTF

[3435] George R. Bolz and R. Gregory Lough. Growth through the first six months of Atlantic cod, Gadus morhua, and haddock, Melanogrammus aeglefinus, based on daily otolith increments. Fishery Bulletin, 86(2):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
growth-through-first-six-montiis-atlantic-cod-gadus-morhua-
and-haddock-melanogrammus;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1988/862/bolz.pdf.

Murphy:1988:SDJ

[3436] Michael L. Murphy, John F. Thedinga, and K. V. Koski. Size and diet of juvenile Pacific salmon during seaward migration through a small estuary in Southeastern Alaska. Fishery Bulletin, 86(2):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/size-and-diet-juvenile-pacific-
salmon-during-seaward-migration-through-small-estuary;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1988/862/
murphy.pdf.

Collins:1988:OYY

[3437] Mark R. Collins and Charles A. Wenner. Occurrence of youngof-the-year king Scomberomorus cavalla, and Spanish, S. maculatus, materials in commercial-type shrimp trawls along the Atlantic coast of the southeast United States. Fishery Bulletin, 86 (2):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14356;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/collins%20(1)_0.pdf.

Brown-Peterson:1988:RBS

[3438] Nancy Brown-Peterson, Peter Thomas, and Connie R. Arnold. Reproductive biology of the spotted seatrout, Cynoscion nebulosus, in South Texas. Fishery Bulletin, 86(2):??, 1988. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/reproductive-biology-
spotted-seatrout-cynoscion-nebulosus-south-texas;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1988/862/
brown.pdf.

Nyman:1988:RBS

[3439] Robert M. Nyman and David O. Conover. The relation between spawning season and the recruitment of young-of-the-year bluefish, Pomatomus saltatrix, to New York. Fishery Bulletin, 86(2):??, 1988. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/relation-between-spawning-season-
and-recruitment-young-year-bluefish-pomatomus-saltatrix;   https:

//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1988/862/
nyman.pdf.

Dew:1988:SCC

[3440] C. Braxton Dew. Stomach contents of commercially caught Hudson River striped bass, Morone saxatilis, 1973–1975. Fishery Bulletin, 86(2):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14357;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/dew.pdf.

Dagg:1988:WTD

[3441] M. J. Dagg, P. B. Ortner, and F. Al-Yamani. Winter-time distribution and abundance of copepod Nauplii in the Northern Gulf of Mexico. Fishery Bulletin, 86(2):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/winter-time-distribution-and-abundance-copepod-
nauplii-northern-gulf-mexico;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1988/862/dagg.pdf.

Pearcy:1988:MCS

[3442] William G. Pearcy and Joseph P. Fisher. Migrations of Coho salmon, Oncorhynchus kisutch, during their first summer in the ocean. Fishery Bulletin, 86(2):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/migrations-coho-salmon-oncorhynchus-kisutch-during-
their-first-summer-ocean;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1988/862/pearcy.pdf.

Vernet:1988:AAP

[3443] Maria Vernet, John R. Hunter, and Russell D. Vetter. Accumulation of age pigments (lipofuscin) in two cold-water fishes. Fishery Bulletin, 86 (2):401–407, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14358;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/vernet.pdf.

Barlow:1988:HPPa

[3444] Jay Barlow. Harbor porpoise, Phocoena phocoena, abundance estimation for California, Oregon, and Washington: I. Ship surveys. Fishery Bulletin, 86(3):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
harbor-porpoise-phocoena-phocoena-abundance-estimation-california-

oregon-and-washington-i;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1988/863/barlow.pdf.

Sigler:1988:EVS

[3445] Michael F. Sigler and Jeffrey T. Fujioka. Evaluation of variability in sablefish, Anoplopoma fimbria, abundance indices in the Gulf of Alaska using the bootstrap method. Fishery Bulletin, 86(3):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/evaluation-
variability-sablefish-anoplopoma-fimbria-abundance-indices-
gulf-alaska-using;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1988/863/sigler.pdf.

McIntyre:1988:PKR

[3446] John N. McIntyre, Reginald R. Reisenbichler, John M. Emlen, Richard L. Wilmot, and James E. Finn. Predation of Karluk River sockeye salmon by Coho salmon and char. Fishery Bulletin, 86(3):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14364;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/mcintyre.pdf.

Barlow:1988:HPPb

[3447] Jay Barlow, Charles W. Oliver, Terry D. Jackson, and Barbara L. Taylor. Harbor porpoise, Phocoena phocoena, abundance estimation for California, Oregon, and Washington: II. Aerial surveys. Fishery Bulletin, 86(3): 433–444, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/harbor-
porpoise-phocoena-phocoena-abundance-estimation-california-
oregon-and-washington-ii;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1988/863/barlow2.pdf.

Stoner:1988:FPA

[3448] Allan W. Stoner and Roger J. Zimmerman. Food pathways associated with penaeid shrimps in a mangrove-fringed estuary. Fishery Bulletin, 86 (3):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/food-pathways-
associated-penaeid-shrimps-mangrove-fringed-estuary;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1988/863/
stoner.pdf.

Brousseau:1988:LTT

[3449] Diane J. Brousseau and Jenny A. Baglivo. Life tables for two field populations of soft-shell clam, Mya arenarla, (Mollusca: Pelecypoda) from Long Island Sound. Fishery Bulletin, 86(3):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/life-tables-
two-field-populations-soft-shell-clam-mya-arenarla-mollusca-
pelecypoda-long;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1988/863/brousseau.pdf.

Watanabe:1988:GPS

[3450] Yoshiro Watanabe, John L. Butler, and Tsukasa Mori. Growth of Pacific saury, Cowlahls saira, in the Northeastern and Northwestern Pacific Ocean. Fishery Bulletin, 86(3):489–498, 1988. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/growth-pacific-saury-
cowlahls-salra-northeastern-and-northwestern-pacific-ocean;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1988/863/watanabe.pdf.

Davies:1988:OUS

[3451] N. M. Davies, R. W. Gauldie, S. A. Crane, and R. K. Thompson. Otolith ultrastructure of smooth oreo, Pseudocyttus maculatus, and black oreo, Allocyttus sp., species. Fishery Bulletin, 86(3):499– 515, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/otolith-
ultrastructure-smooth-oreo-pseudocyttus-maculatus-and-black-
oreo-allocyttus-sp;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1988/863/davies.pdf.

Wilson:1988:EEP

[3452] Elizabeth A. Wilson, Eric N. Powell, and Sammy M. Ray. The effect of the ectoparasitic pyramidellid snail, Boonea impressa, on the growth and health of oysters, Crassostrea virginica, under field conditions. Fishery Bulletin, 86(3):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/effect-ectoparasitic-pyramidellid-snail-boonea-
impressa-growth-and-health-oysters;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1988/863/wilson.pdf.

Farley:1988:LTS

[3453] C. Austin Farley, Peter H. Wolf, and Ralph A. Elston. A longterm study of ‘microcell’ disease in oysters with a description of a new genus, Mikrocytos (g. n.), and two new species, Mikrocytos mackini (sp. n.) and Mikrocytos roughleyi (sp. n.). Fishery Bulletin, 86(3):581–593, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/long-term-study-microcell-disease-oysters-
description-new-genus-mikrocytos-g-n-and-two-new;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1988/863/
farley.pdf.

Boehlert:1988:PBA

[3454] George W. Boehlert and Takashi Sasaki. Pelagic biogeography of the armorhead, Pseudopentaceros wheeleri, and recruitment to isolated seamounts in the North Pacific Ocean. Fishery Bulletin, 86 (3):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14360;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/boehlert_1.pdf.

Hoss:1988:MRS

[3455] Donald E. Hoss, Linda Coston-Clements, David S. Peters, and Patricia A. Tester. Metabolic responses of spot, Leiostomus xanthurus, and Atlantic croaker, Micropogonias undulatus, larvae to cold temperatures encountered following recruitment to estuaries. Fishery Bulletin, 86(3):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
metabolic-responses-spot-leiostomus-xanthurus-and-atlantic-
croaker-micropogonias-undulatus;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1988/863/hoss.pdf.

Williams:1988:CTA

[3456] Austin R. Williams. Cojoined twin adult shrimp (Decapoda: Penaeidae). Fishery Bulletin, 86(3):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14361;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/williams_2.pdf.

Jamieson:1988:OCC

[3457] Glen S. Jamieson and Antan C. Phillips. Occurrence of cancer crab (C. magister and C. oregonensis) megalopae off the West Coast of Vancouver Island, British Columbia. Fishery Bulletin, 86(3):??, 1988. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/occurrence-cancer-crab-c-
magister-and-c-oregonensis-megalopae-west-coast-vancouver-island;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1988/863/jamieson.pdf.

Gould:1988:NMG

[3458] Edith Gould, Diane Rusanowsky, and Donna A. Luedke. Note on muscle glycogen as an indicator of spawning potential in the sea scallop, Placopecten magellanicus. Fishery Bulletin, 86(3):597–601, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14362;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/gould.pdf.

Savoy:1988:TSD

[3459] Thomas F. Savoy and Victor A. Crecco. The timing and significance of density-dependent and density-independent mortality of American shad, Alosa sapidissima. Fishery Bulletin, 86(3): ??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/timing-
and-significance-density-dependent-and-density-independent-
mortality-american-shad;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1988/863/savoy.pdf.

Jamieson:1988:VDM

[3460] Glen S. Jamieson and Ellen K. Pikitch. Vertical distribution and mass mortality of prawns, Pandalus platyceras, in Saanich Inlet, British Columbia. Fishery Bulletin, 86(3):??, 1988. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14363;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/jamieson2.pdf.

Lin:1988:SFF

[3461] Biing-Hwan Lin and Nancy A. Williams. Specifying a functional form for the influence of hatchery smolt release on adult salmon production. Fishery Bulletin, 86(4):655–662, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/specifying-functional-form-influence-
hatchery-smolt-release-adult-salmon-production;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1988/864/biing-
hwan.pdf.

Hobson:1988:TRB

[3462] Edmund S. Hobson and James R. Chess. Trophic relations of the blue rockfish, Sebastes mystinus, in a coastal upwelling system off northern California. Fishery Bulletin, 86(4):??, 1988. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14365;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/hobson_0.pdf.

Graves:1988:EIE

[3463] John E. Graves, Marie A. Simovich, and Kurt M. Schaefer. Electrophoretic identification of early juvenile yellowfin tuna, Thunnus albacares. Fishery Bulletin, 86(4):835–838, 1988. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14373;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/graves_0.pdf.

Blaylock:1988:DAB

[3464] Robert A. Blaylock. Distribution and abundance of the bottlenose dolphin, Tursiops truncatus (Montagu, 1821), in Virginia. Fishery Bulletin, 86(4):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
distribution-and-abundance-bottlenose-dolphin-tursiops-truncatus-
montagu-1821-virginia;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1988/864/blaylock.pdf.

Grimes:1988:RBT

[3465] Churchill R. Grimes, Charles F. Idelberger, Kenneth W. Able, and Stephen C. Turner. The reproductive biology of tilefish, Lopkolatilus chamaeleonticeps Goode and Bean, from the United States Mid-Atlantic Bight, and the effects of fishing on the breeding system. Fishery Bulletin, 86(4):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14366;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/grimes_0.pdf.

Cahoon:1988:CDZ

[3466] Lawrence B. Cahoon and Craig R. Tronzo. A comparison of demersal zooplankton collected at Alligator Reef, Florida, using emergence and reentry traps. Fishery Bulletin, 86(4):??, 1988. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14374;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/cahoon.pdf.

Castro:1988:CTL

[3467] Margarida Castro and Karim Erzini. Comparison of two lengthfrequency based packages for estimating growth and mortality parameters using simulated samples with varying recruitment patterns. Fishery Bulletin, 86(4):645–653, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/comparison-two-length-frequency-based-packages-

estimating-growth-and-mortality-parameters;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1988/864/castro.pdf.

Hargreaves:1988:FMD

[3468] N. R. Hargreaves. A field method for determining prey preferences of predators. Fishery Bulletin, 86(4):??, 1988. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14367;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/hargreaves.pdf.

Anonymous:1988:I

[3469] Anonymous. Index. Fishery Bulletin, 86(4):??, 1988. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14375;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/index%20(2)_2.pdf.

Cross:1988:ABT

[3470] Jeffrey N. Cross. Aspects of the biology of two scyliorhinid sharks, Apristurus brunneus and Parmaturus xaniurus, from the Upper Continental Slope off Southern California. Fishery Bulletin, 86(4):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/aspects-biology-two-
scyliorhinid-sharks-apristurus-brunneus-and-parmaturus-xaniurus-
upper;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1988/864/cross.pdf.

Seki:1988:FHT

[3471] Michael P. Seki and Michael W. Callahan. The feeding habits of two deep slope snappers, Pristipomoides zonatus and P. auricilla, at Pathfinder Reef, Mariana Archipelago. Fishery Bulletin, 86(4):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14368;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/seki.pdf.

Ford:1988:EMP

[3472] Richard F. Ford, Bruce F. Phillips, and Lindsay M. Joll. Experimental manipulation of population density and its effects on growth and mortality of juvenile western rock lobsters, Panulirus cygnus George. Fishery Bulletin, 86(4):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/experimental-manipulation-population-
density-and-its-effects-growth-and-mortality-juvenile;   https:

//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1988/864/
ford.pdf.

Ditty:1988:SDD

[3473] James G. Ditty, Glen G. Zieske, and Richard F. Shaw. Seasonality and depth distribution of larval fishes in the northern Gulf of Mexico above latitude 26^{o}00’N. Fishery Bulletin, 86(4):811–823, 1988. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14369;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/ditty.pdf.

Barshaw:1988:LTS

[3474] Diana E. Barshaw and Donald R. Bryant-Rich. A long-term study on the behavior and survival of early juvenile American lobster, Homarus americanus, in three naturalistic substrates: Eelgrass, mud, and rocks. Fishery Bulletin, 86(4):789–796, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/long-term-study-behavior-
and-survival-early-juvenile-american-lobster-homarus-americanus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1988/864/barshaw.pdf.

Mendelssohn:1988:SPE

[3475] Roy Mendelssohn. Some problems in estimating population sizes from catch-at-age data. Fishery Bulletin, 86(4):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/some-problems-estimating-
population-sizes-catch-age-data;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1988/864/mendelssohn.pdf.

Rogers:1988:UWE

[3476] Christopher W. Rogers, Donald R. Gunderson, and David A. Armstrong. Utilization of a Washington estuary by juvenile English sole, Parophrys vetulus. Fishery Bulletin, 86(4):823–831, 1988. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14370;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/rogers%20(1).pdf.

Bayliff:1988:ISS

[3477] William H. Bayliff. Integrity of schools of skipjack tuna, Katsuwonus pelamis, in the Eastern Pacific Ocean, as determined from tagging data. Fishery Bulletin, 86(4):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/integrity-schools-skipjack-tuna-katsuwonus-pelamis-
eastern-pacific-ocean-determined-tagging;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1988/864/bayliff.pdf.

Shanks:1988:FSH

[3478] Alan L. Shanks. Further support for the hypothesis that internal waves can cause shoreward transport of larval invertebrates and fish. Fishery Bulletin, 86(4):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/further-support-hypothesis-internal-waves-
can-cause-shoreward-transport-larval;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1988/864/shanks.pdf.

Safrit:1988:LWR

[3479] Glen W. Safrit and Frank J. Schwartz. Length–weight relationships for Gulf flounder, Paralichthys albigutta, from North Carolina. Fishery Bulletin, 86(4):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14371;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/safrit.pdf.

Beacham:1988:AMD

[3480] Terry D. Beacham, Clyde B. Murray, and Ruth E. Withler. Age, morphology, developmental biology, and biochemical genetic variation of Yukon River fall chum salmon, Oncorhynchus keta, and comparisons with British Columbia populations. Fishery Bulletin, 86 (4):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/age-
morphology-developmental-biology-and-biochemical-genetic-variation-
yukon-river-fall;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1988/864/beacham.pdf.

Yang:1988:FHD

[3481] M. S. Yang and P. A. Livingston. Food habits and daily ration of Greenland halibut, Reinhardtius hippoglossoides, in the Eastern Bering Sea. Fishery Bulletin, 86(4):??, 1988. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/food-habits-and-daily-ration-greenland-halibut-
reinhardtius-hippoglossoides-eastern-bering;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1988/864/yang.
pdf.

Wiebe:1988:FRE

[3482] Peter H. Wiebe. Functional regression equations for zooplankton displacement volume, wet weight, dry weight, and carbon: a correction. Fishery Bulletin, 86(4):??, 1988. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14372;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/wiebe_0.pdf.

Agnello:1989:EVF

[3483] Richard J. Agnello. The economic value of fishing success: an application of socioeconomic survey data. Fishery Bulletin, 87(1):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/economic-value-fishing-success-
application-socioeconomic-survey-data;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1989/871/agnello.pdf.

Polacheck:1989:YTT

[3484] Tom Polacheck. Yellowfin tuna, Thunnus albacares, catch rates in the Western Pacific. Fishery Bulletin, 87(1):??, 1989. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/yellowfin-tuna-thunnus-albacares-
catch-rates-western-pacific;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1989/871/polacheck.pdf.

Thresher:1989:DAG

[3485] R. E. Thresher, B. D. Bruce, D. M. Furlani, and J. S. Gunn. Distribution, advection, and growth of larvae of the southern temperate gadoid, Macruronus novaezelandiae (Teleostei: Merlucciidae), in Australian coastal waters. Fishery Bulletin, 87(1):??, 1989. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14378;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/thresher.pdf.

Collins:1989:AGK

[3486] Mark R. Collins, David J. Schmidt, C. Wayne Waltz, and James L. Pickney. Age and growth of king mackerel, Scomberomorus cavalla, from the Atlantic Coast of the United States. Fishery Bulletin, 87(1):49–61, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/age-and-growth-king-mackerel-scomberomorus-cavalla-
atlantic-coast-united-states;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1989/871/collins.pdf.

Ropes:1989:FHF

[3487] John W. Ropes. The food habits of five crab species at Pettaquamscutt River, Rhode Island. Fishery Bulletin, 87(1):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/food-habits-five-crab-species-pettaquamscutt-
river-rhode-island;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1989/871/ropes.pdf.

Crawford:1989:SIW

[3488] Maurice K. Crawford, Churchill B. Grimes, and Norman E. Buroker. Stock identification of weakfish, Cynoscion regalis, in the Middle Atlantic region. Fishery Bulletin, 87(1):??, 1989. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14379;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/crawford.pdf.

Emmett:1989:ETN

[3489] B. Emmett and G. S. Jamieson. An experimental transplant of northern abalone, Haliotis kamtschatkana, in Barkley Sound, British Columbia. Fishery Bulletin, 87(1):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/experimental-transplant-northern-
abalone-haliotis-kamtschatkana-barkley-sound-british;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1989/871/
emmett.pdf.

Smith:1989:HVN

[3490] Ingrid Smith, Mark S. Fonseca, Jose A. Rivera, and Keith A. Rittmaster. Habitat value of natural versus recently transplanted eelgrass, Zostera marina, for the bay scallop, Argopecten irradians. Fishery Bulletin, 87(1):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/habitat-value-natural-versus-recently-transplanted-
eelgrass-zostera-marina-bay-scallop;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1989/871/smith.pdf.

Moring:1989:FHA

[3491] John R. Moring. Food habits and algal associations of juvenile lumpfish, Cyclopterus lumpus L., in intertidal waters. Fishery Bulletin, 87(1):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14380;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/moring.pdf.

Feldkamp:1989:EBE

[3492] Steven D. Feldkamp, Daniel P. Costa, and Gregory K. Dekrey. Energetic and behavioral effects of net entanglement on juvenile northern fur seals, Callorhinus ursinus. Fishery Bulletin, 87(1):??, 1989. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/energetic-and-behavioral-effects-
net-entanglement-juvenile-northern-fur-seals-callorhinus;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1989/
871/feldkamp.pdf.

Thorrold:1989:ESE

[3493] Simon R. Thorrold. Estimating some early life history parameters in a tropical clupeid, Herklotsichthys castelnaui, from daily growth increments in otoliths. Fishery Bulletin, 87(1):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/estimating-some-early-life-history-parameters-
tropical-clupeid-herklotsichthys-castelnaui;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1989/871/thorrold.
pdf.

Gartner:1989:EFM

[3494] John V. Gartner, Jr., Walter J. Conley, and Thomas L. Hopkins. Escapement by fishes from midwater trawls: a case study using lanternfishes (Pisces: Myctophidae). Fishery Bulletin, 87(1):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/escapement-
fishes-midwater-trawls-case-study-using-lanternfishes-pisces-
myctophidae;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1989/871/gartner.pdf.

Wenner:1989:SCA

[3495] Elizabeth L. Wenner and Charles A. Wenner. Seasonal composition and abundance of decapod and stomatopod crustaceans from coastal habitats, Southeastern United States. Fishery Bulletin, 87(1):155–176, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/seasonal-composition-and-abundance-decapod-and-
stomatopod-crustaceans-coastal-habitats;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1989/871/wenner.pdf.

Hightower:1989:STL

[3496] Joseph E. Hightower and Gary D. Grossman. Status of the tilefish, Lopholatilus chamaeleonticeps, fishery off South Carolina and Georgia and recommendations for management. Fishery Bulletin, 87(1): ??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/status-
tilefish-lopholatilus-chamaeleonticeps-fishery-south-carolina-
and-georgia-and;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1989/871/hightower.pdf.

Williams:1989:CSG

[3497] Austin B. Williams and C. Allan Child. Comparison of some genera and species of box crabs (Brachyura: Calappidae), Southwestern North Atlantic, with description of a new genus and species. Fishery Bulletin, 87 (1):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/comparison-
some-genera-and-species-box-crabs-brachyura-calappidae-southwestern-
north;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1989/871/williams.pdf.

Lauth:1989:SSC

[3498] Robert R. Lauth. Seasonal spawning cycle, spawning frequency and batch fecundity of the cabezon, Scorpaenichthys marmoratus, in Puget Sound, Washington. Fishery Bulletin, 87(1):??, 1989. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/seasonal-spawning-cycle-spawning-
frequency-and-batch-fecundity-cabezon-scorpaenichthys;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1989/871/
lauth.pdf.

Ralston:1989:NID

[3499] Stephen Ralston and Happy A. Williams. Numerical integration of daily growth increments: an efficient means of ageing tropical fishes for stock assessment. Fishery Bulletin, 87(1):??, 1989. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14376;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/ralston_1.pdf.

Margulies:1989:EFC

[3500] Daniel Margulies. Effects of food concentration and temperature on development, growth, and survival of white perch, Morone americana, eggs and larvae. Fishery Bulletin, 87(1):??, 1989. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/effects-food-concentration-and-
temperature-development-growth-and-survival-white-perch;   https:

//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1989/871/
margulies.pdf.

Beckman:1989:AGR

[3501] Daniel W. Beckman, Charles A. Wilson, and A. Louise Stanley. Age and growth of red drum, Sciaenops ocellatus, from offshore waters of the northern Gulf of Mexico. Fishery Bulletin, 87(1):17–28, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14377;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/beckman.pdf.

Bailey:1989:DSD

[3502] Kevin Bailey. Description and surface distribution of juvenile Peruvian jack mackerel, Trachurus murphyi, Nichols from the Subtropical Convergence Zone of the Central South Pacific. Fishery Bulletin, 87(2):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/description-
and-surface-distribution-juvenile-peruvian-jack-mackerel-trachurus-
murphyi;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1989/872/bailey.pdf.

Szelistowski:1989:MMS

[3503] William A. Szelistowski and Juan Garita. Mass mortality of sciaenid fishes in the Gulf of Nicoya, Costa Rica. Fishery Bulletin, 87(2):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14381;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/szelistowski.pdf.

Jackson:1989:USM

[3504] George David Jackson. The use of statolith microstructures to analyze life history events in the small tropical cephalopod Idiosepius pygmaeus. Fishery Bulletin, 87(2):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/use-statolith-microstructures-analyze-life-history-
events-small-tropical-cephalopod;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1989/872/jackson.pdf.

Barshaw:1989:GSE

[3505] Diana E. Barshaw. Growth and survival of early juvenile American lobsters, Homarus americanus, on a diet of plankton. Fishery Bulletin, 87(2):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14382;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/barshaw.pdf.

Mullen:1989:AFT

[3506] Ashley J. Mullen. Aggregation of fish through variable diffusivity. Fishery Bulletin, 87(2):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
aggregation-fish-through-variable-diffusivity;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1989/872/mullen.
pdf.

Vidal:1989:BOF

[3507] Omar Vidal and Gary Pechter. Behavioral observations on fin whale, Balaenoptera physalus, in the presence of killer whale, Orcinus orca. Fishery Bulletin, 87(2):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/node/
14383;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/vidal.pdf.

Sheridan:1989:FIR

[3508] Peter F. Sheridan, Refugio G. Castro M., Frank J. Patella, Jr., and Gilbert Zamora, Jr. Factors influencing recapture patterns of tagged penaeid shrimp in the Western Gulf of Mexico. Fishery Bulletin, 87(2):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/factors-influencing-
recapture-patterns-tagged-penaeid-shrimp-western-gulf-mexico;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1989/872/sheridan.pdf.

Squires:1989:IJU

[3509] Dale Squires, Samuel F. Herrick, Jr., and James Hastie. Integration of Japanese and United States sablefish markets. Fishery Bulletin, 87(2):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
integration-japanese-and-united-states-sablefish-markets;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1989/
872/squires.pdf.

Tegner:1989:PBR

[3510] Mia J. Tegner, Paul A. Breen, and Cleridy E. Lennert. Population biology of red abalones, Haliotis rufescens, in Southern California and management of the red and pink, H. corrugata, abalone fisheries. Fishery Bulletin, 87(2):313–339, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/population-biology-red-abalones-haliotis-rufescens-
southern-california-and-management-red;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1989/872/tegner.pdf.

Tucker:1989:EUB

[3511] John W. Tucker, Jr. Energy utilization in bay anchovy, Anchoa mitchilli, and black sea bass, Centropristis striata striata, eggs and larvae. Fishery Bulletin, 87(2):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/energy-utilization-bay-anchov-anchoa-mitchilli-and-
black-sea-bass-centropristis-striata;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1989/872/tucker.pdf.

Utter:1989:GPS

[3512] F. Utter, G. Milner, G. Ståhl, and D. Teel. Genetic population structure of Chinook salmon, Oncorhynchus tshawytscha, in the Pacific Northwest. Fishery Bulletin, 87(2):239–264, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/genetic-population-structure-chinook-
salmon-oncorhynchus-tshawytscha-pacific-northwest;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1989/872/
utter.pdf.

Chambers:1989:ESF

[3513] R. Christopher Chambers, William C. Leggett, and Joseph A. Brown. Egg size, female effects, and the correlations between early life history traits of capelin, Mallotus villosus: an appraisal at the individual level. Fishery Bulletin, 87(3):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/egg-size-female-effects-and-correlations-between-early-
life-history-traits-capelin-mallotus;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1989/873/chambers.pdf.

Moksness:1989:ABC

[3514] Erlend Moksness and Vidar Wespestad. Ageing and back-calculating growth rates of Pacific herring, Clupea pallasii, larvae by reading daily otolith increments. Fishery Bulletin, 87(3):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/ageing-and-back-calculating-growth-rates-
pacific-herring-clupea-pallasii-larvae-reading;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1989/873/moksness.
pdf.

Watanabe:1989:LPM

[3515] Yoshiro Watanabe and Nancy C. H. Lo. Larval production and mortality of Pacific saury, Cololabis sairn, in the Northwestern Pacific Ocean. Fishery Bulletin, 87(3):??, 1989. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/larval-production-and-mortality-
pacific-saury-cololabis-sairn-northwestern-pacific-ocean;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1989/
873/watanabe.pdf.

Jackson:1989:KGP

[3516] Christopher J. Jackson, Peter C. Rothlisberg, Robert C. Pendrey, and Mair T. Beamish. A key to genera of the penaeid larvae and early postlarvae of the Indo-west Pacific region, with descriptions of the larval development of Atypopenaeus formosus Dall and Metapenaeopsis palmensis Haswell (Decapoda: Penaeoidea: Penaeidae) reared in the laboratory. Fishery Bulletin, 87(3):??, 1989. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14388;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/jackson_1.pdf.

Clarke:1989:SDS

[3517] Thomas A. Clarke. Seasonal differences in spawning, egg size, and early development time of the Hawaiian anchovy or nehu, Encrasicholina purpurea. Fishery Bulletin, 87(3):??, 1989. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/seasonal-differences-
spawning-egg-size-and-early-development-time-hawaiian-anchovy-
or-nehu;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1989/873/clarke.pdf.

Morse:1989:CGM

[3518] Wallace W. Morse. Catchability, growth, and mortality of larval fishes. Fishery Bulletin, 87(3):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/catchability-growth-and-mortality-larval-fishes;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1989/873/
morse.pdf.

Watson:1989:LFD

[3519] William Watson and Raymond L. Davis, Jr. Larval fish diets in shallow coastal waters off San Onofre, California. Fishery Bulletin, 87(3):

??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/larval-fish-
diets-shallow-coastal-waters-san-onofre-california;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1989/873/
watson.pdf.

Marliave:1989:LLF

[3520] Jeffrey B. Marliave and Alex E. Peden. Larvae of Liparis fucensis and Liparis callyodon: Is the ‘cottoid bubblemorph’ phylogenetically significant? Fishery Bulletin, 87(3):??, 1989. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14389;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/marliave_0.pdf.

Hedgecock:1989:GMV

[3521] Dennis Hedgecock, Elmarie S. Hutchinson, Gang Li, Frederic L. Sly, and Keith Nelson. Genetic and morphometric variation in the Pacific sardine, Sardinops sagax caerulea: Comparisons and contrasts with historical data and with variability in the northern anchovy, Engraulis mordax. Fishery Bulletin, 87(3):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/genetic-and-morphometric-variation-pacific-sardine-
sardinops-sagax-caerulea-comparisons-and;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1989/873/hedgecock.
pdf.

Mullin:1989:BES

[3522] M. M. Mullin and A. Conversi. Biomasses of euphausiids and smaller zooplankton in the California Current-geographic and interannual comparisons relative to the Pacific whiting, Merluccius productus, fishery. Fishery Bulletin, 87(3):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
biomasses-euphausiids-and-smaller-zooplankton-california-current-
geographic-and-interannual;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1989/873/mullin.pdf.

Wroblewski:1989:OWC

[3523] J. S. Wroblewski, James G. Richman, and George L. Mellor. Optimal wind conditions for the survival of larval northern anchovy, Engraulis mordax: a modeling investigation. Fishery Bulletin, 87(3): ??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/optimal-
wind-conditions-survival-larval-northern-anchovy-engraulis-

mordax-modeling;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1989/873/wroblewski.pdf.

Houde:1989:CGM

[3524] Edward D. Houde. Comparative growth, mortality, and energetics of marine fish larvae: Temperature and implied latitudinal effects. Fishery Bulletin, 87(3):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/comparative-growth-mortality-and-energetics-marine-
fish-larvae-temperature-and-implied;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1989/873/houde.pdf.

Owen:1989:SSP

[3525] R. W. Owen, N. C. H. Lo, J. L. Butler, G. H. Theilacker, A. Alvarino, J. R. Hunter, and Y. Watanabe. Spawning and survival patterns of larval northern anchovy, Engraulis mordax, in contrasting environments — a site-intensive study. Fishery Bulletin, 87(3):673–688, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/spawning-and-
survival-patterns-larval-northern-anchovy-engraulis-mordax-
contrasting;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1989/873/owen.pdf.

Yamashita:1989:LSB

[3526] Yoh Yamashita and Kevin M. Bailey. A laboratory study of the bioenergetics of larval walleye pollock, Theragra chalcogramma. Fishery Bulletin, 87(3):525–536, 1989. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
laboratory-study-bioenergetics-larval-walleye-pollock-theragra-
chalcogramma;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1989/873/yamashita.pdf.

Leighton:1989:AGH

[3527] David L. Leighton. Abalone (genus Haliotis) mariculture on the North American Pacific Coast. Fishery Bulletin, 87(3):??, 1989. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/abalone-genus-haliotis-mariculture-
north-american-pacific-coast;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1989/873/leighton.pdf.

Pearcy:1989:RTD

[3528] William G. Pearcy, Christopher D. Wilson, Alton W. Chung, and John W. Chapman. Residence times, distribution, and production of juvenile chum salmon, Oncorhynchus keta, in Netarts Bay, Oregon. Fishery Bulletin, 87(3):553–568, 1989. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/residence-times-distribution-and-
production-juvenile-chum-salmon-oncorhynchus-keta-netarts;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1989/
873/pearcy.pdf.

Anonymous:1989:TP

[3529] Anonymous. Title page. Fishery Bulletin, 87(3):??, 1989. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14384;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/title%20page.pdf.

Lo:1989:PBE

[3530] Nancy C. H. Lo, John R. Hunter, and Roger P. Hewitt. Precision and bias of estimates of larval mortality. Fishery Bulletin, 87(3):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/precision-and-bias-
estimates-larval-mortality;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1989/873/lo.pdf.

Smith:1989:CMR

[3531] Paul E. Smith, Haydee Santander, and Juergen Alheit. Comparison of the mortality rates of Pacific sardine, Sardinops sagax, and Peruvian anchovy, Engraulis ringens, eggs off Peru. Fishery Bulletin, 87(3):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/comparison-mortality-
rates-pacific-sardine-sardinops-sagax-and-peruvian-anchovy-
engraulis;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1989/873/smith.pdf.

Vlymen:1989:RLR

[3532] Lillian Vlymen. Reuben lasker: a remembrance. Fishery Bulletin, 87(3):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14385;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/vlynem.pdf.

Margulies:1989:SSV

[3533] Daniel Margulies. Size-specific vulnerability to predation and sensory system development of white seabass, Atractoscion nobilis, larvae. Fishery Bulletin, 87(3):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/size-specific-vulnerability-predation-and-
sensory-system-development-white-seabass;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1989/873/margulies.
pdf.

Stevens:1989:DDL

[3534] Elizabeth G. Stevens, William Watson, and H. Geoffrey Moser. Development and distribution of larvae and pelagic juveniles of three kyphosid fishes (Girella nigricans, Medialuna californiensis, and Hermosilla azurea) off California and Baja California. Fishery Bulletin, 87 (3):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/development-
and-distribution-larvae-and-pelagic-juveniles-three-kyphosid-
fishes-girella;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1989/873/stevens.pdf.

Pauly:1989:ERL

[3535] Daniel Pauly. An eponym for reuben lasker. Fishery Bulletin, 87 (3):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14386;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/pauly.pdf.

Butler:1989:GDL

[3536] John L. Butler. Growth during the larval and juvenile stages of the northern anchovy, Engraulis mordax, in the California Current during 1980–84. Fishery Bulletin, 87(3):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/growth-during-larval-and-
juvenile-stages-northern-anchovy-engraulis-mordax-california;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1989/873/butler.pdf.

McGowan:1989:BTT

[3537] Michael F. McGowan and William J. Richards. Bluefin tuna, Thunnus thynnus, larvae in the Gulf Stream off the Southeastern United States: Satellite and shipboard observations of their environment. Fishery Bulletin, 87(3):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
bluefin-tuna-thunnus-thynnus-larvae-gulf-stream-southeastern-
united-states-satellite-and;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1989/873/mcgowan.pdf.

Theilacker:1989:MCH

[3538] Gail H. Theilacker and Yoshiro Watanabe. Midgut cell height defines nutritional status of laboratory raised larval northern anchovy, Engraulis mordax. Fishery Bulletin, 87(3):457–469, 1989. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/midgut-cell-height-
defines-nutritional-status-laboratory-raised-larval-northern-
anchovy;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1989/873/theilacker.pdf.

Somerton:1989:MCC

[3539] David A. Somerton and Donald R. Kobayashi. A method for correcting catches of fish larvae for the size selection of plankton nets. Fishery Bulletin, 87(3):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14387;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/somerton%20(1).pdf.

Boehlert:1989:TSG

[3540] George W. Boehlert, Mary M. Yoklavich, and Dudley B. Chelton. Time series of growth in the genus Sebastes from the Northeast Pacific Ocean. Fishery Bulletin, 87(4):??, 1989. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/time-series-growth-
genus-sebastes-northeast-pacific-ocean;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1989/874/boehlert.
pdf.

Mulligan:1989:OFN

[3541] K. P. Mulligan, R. W. Gauldie, and R. Thomson. Otoconia from four New Zealand Chimaeriformes. Fishery Bulletin, 87(4):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/otoconia-four-
new-zealand-chimaeriformes;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1989/874/mulligan.pdf.

Davis:1989:TLA

[3542] Tim L. O. Davis, Greg P. Jenkins, Mori Yukinawa, and Yasuo Nishikawa. Tuna larvae abundance: Comparative estimates from concurrent Japanese and Australian sampling programs. Fishery Bulletin, 87(4):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14392;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/davis_5.pdf.

Charles-Dominique:1989:CEP

[3543] Emmanuel Charles-Dominique. Catch efficiencies of purse and beach seines in Ivory Coast lagoons. Fishery Bulletin, 87(4):??, 1989. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/catch-efficiencies-purse-
and-beach-seines-ivory-coast-lagoons;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1989/874/charles-dominique.
pdf.

Neira:1989:LDA

[3544] Francisco J. Neira. Larval development of the Australian devilfish, Gymnapistes marmoratus (Teleostei: Scorpaenidae). Fishery Bulletin, 87 (4):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/larval-
development-australian-devilfish-gymnapistes-marmoratus-teleostei-
scorpaenidae;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1989/874/neira.pdf.

Witzell:1989:LST

[3545] W. N. Witzell. Longbill spearfish, Tetrapturus pfluegeri, incidentally caught by recreational billfishermen in the western North Atlantic Ocean, 1974–86. Fishery Bulletin, 87(4):??, 1989. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14393;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/witzell.pdf.

Dauble:1989:SDJ

[3546] Dennis D. Dauble, Thomas L. Page, and R. William Hanf, Jr. Spatial distribution of juvenile salmonids in the Hanford Reach, Columbia River. Fishery Bulletin, 87(4):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/spatial-distribution-juvenile-salmonids-hanford-reach-
columbia-river;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1989/874/dauble.pdf.

Pearcy:1989:SOD

[3547] W. G. Pearcy, D. L. Stein, M. A. Hixon, E. K. Pikitch, W. H. Barss, and R. M. Starr. Submersible observations of deep-reef fishes of Heceta Bank, Oregon. Fishery Bulletin, 87(4):955–965, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/submersible-
observations-deep-reef-fishes-heceta-bank-oregon;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1989/874/
pearcy.pdf.

Parsons:1989:ADL

[3548] Glenn R. Parsons and Kevin M. Peters. Age determination in larval and juvenile sheepshead, Archosargus probatocephalus. Fishery Bulletin, 87(4):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14394;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/parsons%20(2).pdf.

Edwards:1989:ENL

[3549] Elizabeth F. Edwards and Pierre M. Kleiber. Effects of nonrandomness on line transect estimates of dolphin school abundance. Fishery Bulletin, 87(4):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/effects-
nonrandomness-line-transect-estimates-dolphin-school-abundance;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1989/874/edwards.pdf.

Quinn:1989:SAI

[3550] T. P. Quinn, E. L. Brannon, and A. H. Dittman. Spatial aspects of imprinting and homing in Coho salmon, Oncorhynchus kisutch. Fishery Bulletin, 87(4):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
spatial-aspects-imprinting-and-homing-coho-salmon-oncorhynchus-
kisutch;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1989/874/quinn.pdf.

Yoklavich:1989:RSD

[3551] Mary M. Yoklavich and Ellen K. Pikitch. Reproductive status of Dover sole, Microstomus pacificus, off northern Oregon. Fishery Bulletin, 87(4):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14395;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/yoklavich%20(1).pdf.

Everson:1989:MRT

[3552] Alan R. Everson, Happy A. Williams, and Bernard M. Ito. Maturation and reproduction in two Hawaiian eteline snappers, uku, Aprion virescens, and onaga, Etelis coruscans. Fishery Bulletin, 87(4):877–888, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/maturation-
and-reproduction-two-hawaiian-eteline-snappers-uku-aprion-virescens-
and-onaga;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1989/874/everson.pdf.

Sedberry:1989:FHW

[3553] George R. Sedberry. Feeding habits of whitebone porgy, Calamus leucosteus (Teleostei: Sparidae), associated with hard bottom reefs off the Southeastern United States. Fishery Bulletin, 87(4):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/feeding-habits-whitebone-porgy-
calamus-leucosteus-teleostei-sparidae-associated-hard-bottom;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1989/874/sedberry.pdf.

Vecchione:1989:DGH

[3554] Michael Vecchione and Victoria A. Hand. Digestive-gland histology in paralarval squids (Cephalopoda: Loliginidae). Fishery Bulletin, 87(4):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14396;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/vecchione.pdf.

Hill:1989:CAG

[3555] Kevin T. Hill, Gregor M. Cailliet, and Richard L. Radtke. A comparative analysis of growth zones in four calcified structures of Pacific blue marlin, Makaim nigricans. Fishery Bulletin, 87(4):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/comparative-analysis-
growth-zones-four-calcified-structures-pacific-blue-marlin-
makaim;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1989/874/hill.pdf.

Williams:1989:NSL

[3556] Austin B. Williams and Keiji Baba. New squat lobsters (Galatheidae) from the Pacific Ocean: Mariana Back Arc Basin, East Pacific Rise, and Cascadia Basin. Fishery Bulletin, 87(4):??, 1989. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/new-squat-lobsters-galatheidae-
pacific-ocean-mariana-back-arc-basin-east-pacific-rise-and;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1989/874/williams.pdf.

Parrish:1989:IHJ

[3557] Frank A. Parrish. Identification of habitat of juvenile snappers in Hawaii. Fishery Bulletin, 87(4):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14397;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/parrish%20(1).pdf.

Langton:1989:PSM

[3558] Richard W. Langton and Joseph R. Uzmann. A photographic survey of the megafauna of the central and eastern Gulf of Maine. Fishery Bulletin, 87(4):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/photographic-survey-megafauna-central-and-
eastern-gulf-maine;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1989/874/langton.pdf.

Killam:1989:AGB

[3559] Kristie A. Killam and Glenn R. Parsons. Age and growth of the blacktip shark, Carcharhinus limbatus, near Tampa Bay, Florida. Fishery Bulletin, 87(4):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14390;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/killam%20(1).pdf.

Anonymous:1989:I

[3560] Anonymous. Index. Fishery Bulletin, 87(4):??, 1989. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14398;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/index%20(3)_1.pdf.

Livingston:1989:ITP

[3561] Patricia A. Livingston. Interannual trends in Pacific cod, Gadus macrocephalus, predation on three commercially important crab species in the eastern Bering Sea. Fishery Bulletin, 87(4):??, 1989. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/interannual-trends-pacific-cod-
gadus-macrocephalus-predation-three-commercially-important;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1989/874/livingston.pdf.

Ohman:1989:CSS

[3562] M. D. Ohman and J. R. Wilkinson. Comparative standing stocks of mesozooplankton and macrozooplankton in the southern sector of the California Current system. Fishery Bulletin, 87(4):??, 1989. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14391;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/ohman.pdf.

Able:1990:PSF

[3563] Kenneth W. Able, R. Edmond Matheson, Wallace W. Morse, Michael P. Fahay, and Gary Shepherd. Patterns of summer flounder Paralichthys dentatus early life history in the Mid-Atlantic Bight and New Jersey estuaries. Fishery Bulletin, 88(1):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/patterns-summer-flounder-paralichthys-dentatus-
early-life-history-mid-atlantic-bight-and;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1990/881/able.pdf.

Herrick:1990:MFF

[3564] Samuel F. Herrick, Jr. and Dale Squires. On measuring fishing fleet productivity: Development and demonstration of an analytical framework. Fishery Bulletin, 88(1):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/measuring-fishing-fleet-productivity-
development-and-demonstration-analytical-framework;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1990/881/
herrick.pdf.

Ratty:1990:GMH

[3565] Frank J. Ratty, R. Michael Laurs, and Raymond M. Kelly. Gonad morphology, histology, and spermatogenesis in South Pacific albacore tuna Thunnus alalunga (Scombridae). Fishery Bulletin, 88 (1):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/gonad-
morphology-histology-and-spermatogenesis-south-pacific-albacore-
tuna-thunnus-alalunga;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1990/881/ratty.pdf.

Barry:1990:IDP

[3566] James P. Barry and Mia J. Tegner. Inferring demographic processes from size-frequency distributions: Simple models indicate specific patterns of growth and mortality. Fishery Bulletin, 88(1):??, 1990. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/inferring-demographic-processes-
size-frequency-distributions-simple-models-indicate;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1990/881/
barry.pdf.

Ho:1990:PBH

[3567] Ju shey Ho. Phylogeny and biogeography of hakes (Merluccius; Teleostei): a cladistic analysis. Fishery Bulletin, 88(1):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/phylogeny-and-
biogeography-hakes-merluccius-teleostei-cladistic-analysis;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1990/881/ho.pdf.

Steimle:1990:PDG

[3568] Frank W. Steimle. Population dynamics, growth, and production estimates for the sand dollar Echinarachnius parma. Fishery Bulletin, 88(1):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/population-dynamics-growth-and-production-
estimates-sand-dollar-echinarachnius-parma;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1990/881/steimle.pdf.

Benfield:1990:GEP

[3569] Mark C. Benfield, Jaap R. Bosschieter, and Anthony T. Forbes. Growth and emigration of Penaeus indicus H. Milne-Edwards (Crustacea:Decapoda:Penaeidae) in the St. Lucia Estuary, Southern Africa. Fishery Bulletin, 88(1):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/growth-and-emigration-penaeus-indicus-h-milne-edwards-
crustaceadecapodapenaeidae-st-lucia;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1990/881/benfield.pdf.

Holt:1990:MTD

[3570] Rennie S. Holt and Stephanie N. Sexton. Monitoring trends in dolphin abundance in the Eastern Tropical Pacific using research vessels over a long sampling period: Analyses of 1986 data, the first year. Fishery Bulletin, 88(1):105–111, 1990. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
monitoring-trends-dolphin-abundance-eastern-tropical-pacific-

using-research-vessels-over;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1990/881/holt.pdf.

Uchiyama:1990:AEC

[3571] James H. Uchiyama and Jeffrey D. Sampaga. Age estimation and composition of pelagic armorhead Pseudopentaceros wheeleri from the Hancock Seamounts. Fishery Bulletin, 88(1):??, 1990. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/age-estimation-and-composition-
pelagic-armorhead-pseudopentaceros-wheeleri-hancock;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1990/881/
uchiyama.pdf.

Carls:1990:ADG

[3572] Mark G. Carls and Stanley D. Rice. Abnormal development and growth reductions of pollock Theragra chalcogramma embryos exposed to watersoluble fractions of oil. Fishery Bulletin, 88(1):??, 1990. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/abnormal-development-and-growth-
reductions-pollock-theragra-chalcogramma-embryos-exposed;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1990/
881/carls.pdf.

Jackson:1990:AGT

[3573] George David Jackson. Age and growth of the tropical nearshore loliginid squid Sepioteuthis lessoniana determined from statolith growth-ring analysis. Fishery Bulletin, 88(1):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/age-and-growth-tropical-nearshore-loliginid-squid-
sepioteuthis-lessoniana-determined;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1990/881/jackson.pdf.

Wespestad:1990:OGS

[3574] Vidar G. Wespestad and Erlend Moksness. Observations on growth and survival during the early life history of Pacific herring Clupea pallasii from Bristol Bay, Alaska, in a marine mesocosm. Fishery Bulletin, 88(1):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
observations-growth-and-survival-during-early-life-history-
pacific-herring-clupea-pallasii;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1990/881/wespestad.pdf.

DeBrosse:1990:CSM

[3575] Gregory A. DeBrosse, Adam J. Baldinger, and Patsy A. McLaughlin. A comparative study of the megalopal stages of Cancer oregonensis Dana and C. productus Randall (Decapoda: Brachyura: Cancridae) from the Northeastern Pacific. Fishery Bulletin, 88(1):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/comparative-
study-megalopal-stages-cancer-oregonensis-dana-and-c-productus-
randall-decapoda;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1990/881/debrosse.pdf.

Jefferson:1990:SDD

[3576] Thomas A. Jefferson. Sexual dimorphism and development of external features in Dall’s porpoise Phocoenoides dalli. Fishery Bulletin, 88 (1):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/sexual-
dimorphism-and-development-external-features-dalls-porpoise-
phocoenoides-dalli;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1990/881/jefferson.pdf.

Hayse:1990:FHA

[3577] John W. Hayse. Feeding habits, age, growth, and reproduction of Atlantic spadefish Chaetodipterus faber (Pisces: Ephippidae) in South Carolina. Fishery Bulletin, 88(1):??, 1990. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14399;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/hayse.pdf.

Fisher:1990:DRT

[3578] Joseph P. Fisher and William G. Pearcy. Distribution and residence times of juvenile fall and spring Chinook salmon in Coos Bay, Oregon. Fishery Bulletin, 88(1):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/distribution-and-residence-times-juvenile-fall-and-
spring-chinook-salmon-coos-bay-oregon;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1990/881/fisher.pdf.

Kendall:1990:ELD

[3579] Arthur W. Kendall, Jr. and Susan J. Picquelle. Egg and larval distributions of walleye pollock Theragra chalcogramma in Shelikof Strait, Gulf of Alaska. Fishery Bulletin, 88(1):133–154, 1990. CODEN FSYBAY.

ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/egg-and-larval-distributions-walleye-
pollock-theragra-chalcogramma-shelikof-strait-gulf;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1990/881/
kendall.pdf.

Gauldie:1990:HBN

[3580] Robert W. Gauldie and Zophie Czochanska. Hyperostosic bones from the New Zealand snapper Chrysophrys auratus (Sparidae). Fishery Bulletin, 88(1):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
hyperostosic-bones-new-zealand-snapper-chrysophrys-auratus-
sparidae;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1990/881/gauldie.pdf.

Maillet:1990:ESF

[3581] Gary L. Maillet and David M. Checkley, Jr. Effects of starvation on the frequency of formation and width of growth increments in Sagittae of laboratory-reared Atlantic menhaden Brevoortia tyrannus larvae. Fishery Bulletin, 88(1):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
effects-starvation-frequency-formation-and-width-growth-increments-
sagittae-laboratory;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1990/881/maillet.pdf.

Graves:1990:BGS

[3582] John E. Graves, Michelle J. Curtis, Paul A. Oeth, and Robin S. Waples. Biochemical genetics of southern California basses of the genus Paralabrax: Specific identification of fresh and ethanol-preserved individual eggs and early larvae. Fishery Bulletin, 88(1):59–66, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/biochemical-genetics-
southern-california-basses-genus-paralabrax-specific-identification;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1990/881/graves.pdf.

Podesta:1990:MPA

[3583] Guillermo P. Podesta. Migratory pattern of Argentine hake Merluccius hubbsi and oceanic processes in the southwestern Atlantic Ocean. Fishery Bulletin, 88(1):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/migratory-pattern-argentine-hake-merluccius-hubbsi-

and-oceanic-processes-southwestern;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1990/881/podesta.pdf.

Iversen:1990:ESV

[3584] Edwin S. Iversen, Scott P. Bannerot, and Darryl E. Jory. Evidence of survival value related to burying behavior in queen conch Strombus gigas. Fishery Bulletin, 88(2):??, 1990. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/evidence-survival-value-related-
burying-behavior-queen-conch-strombus-gigas;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1990/882/iversen.
pdf.

Silber:1990:ODV

[3585] Gregory K. Silber. Occurrence and distribution of the vaquita Phocoena sinus in the Northern Gulf of California. Fishery Bulletin, 88(2):339–346, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/occurrence-and-distribution-vaquita-phocoena-sinus-
northern-gulf-california;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1990/882/silber.pdf.

Kim:1990:ODL

[3586] Suam Kim and Bohyun Bang. Oceanic dispersion of larval fish and its implication for mortality estimates: Case study of walleye pollock larvae in Shelikof Strait, Alaska. Fishery Bulletin, 88(2):303–311, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/oceanic-dispersion-larval-fish-
and-its-implication-mortality-estimates-case-study-walleye;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1990/882/kim.pdf.

Smith:1990:LST

[3587] Susan E. Smith and Norman J. Abramson. Leopard shark Triakis semifasciata distribution, mortality rate, yield, and stock replenishment estimates based on a tagging study in San Francisco Bay. Fishery Bulletin, 88(2):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/leopard-
shark-triakis-semifasciata-distribution-mortality-rate-yield-
and-stock;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1990/882/smith.pdf.

Kope:1990:DFA

[3588] Robert G. Kope and Louis W. Botsford. Determination of factors affecting recruitment of Chinook salmon Oncorhynchus tshawytscha in central California. Fishery Bulletin, 88(2):??, 1990. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/determination-factors-affecting-
recruitment-chinook-salmon-oncorhynchus-tshawytscha-central;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1990/882/kope.pdf.

Stergiou:1990:SAM

[3589] Konstantinos I. Stergiou. A seasonal autoregressive model of the anchovy Engraulis encrasicolus fishery in the Eastern Mediterranean. Fishery Bulletin, 88(2):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
seasonal-autoregressive-model-anchovy-engraulis-encrasicolus-
fishery-eastern-mediterranean;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1990/882/stergiou.pdf.

Barshaw:1990:TTA

[3590] Diana E. Barshaw and Kenneth W. Able. Tethering as a technique for assessing predation rates in different habitats: an evaluation using juvenile lobsters Homarus americanus. Fishery Bulletin, 88(2):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/tethering-technique-
assessing-predation-rates-different-habitats-evaluation-using-
juvenille;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1990/882/barshaw.pdf.

Matthews:1990:CSH

[3591] Kathleen R. Matthews. A comparative study of habitat use by young-of-the-year, subadult, and adult rockfishes on four habitat types in central Puget Sound. Fishery Bulletin, 88(2):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/comparative-
study-habitat-use-young-year-subadult-and-adult-rockfishes-
four-habitat-types;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1990/882/matthews.pdf.

Sturm:1990:AGR

[3592] Maxwell G. de L. Sturm and Premila Salter. Age, growth, and reproduction of the king mackerel Scomberomorus cavalla (Cuvier) in Trinidad waters. Fishery Bulletin, 88(2):361–370, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/age-growth-and-reproduction-king-mackerel-
scomberomorus-cavalla-cuvier-trinidad-waters;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1990/882/sturm.
pdf.

Cockcroft:1990:AGR

[3593] Victor Gavin Cockcroft and Graham James Berry Ross. Age, growth, and reproduction of bottlenose dolphins Tursiops truncatus from the East Coast of Southern Africa. Fishery Bulletin, 88(2):289–302, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/age-growth-
and-reproduction-bottlenose-dolphins-tursiops-truncatus-east-
coast-southern;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1990/882/cockcroft.pdf.

Nizinski:1990:STS

[3594] Martha S. Nizinski, Bruce B. Collette, and Betsy B. Washington. Separation of two species of sand lances, Ammodytes americanus and A. dubius, in the western North Atlantic. Fishery Bulletin, 88(2):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/separation-
two-species-sand-lances-ammodytes-americanus-and-dubius-western-
north-atlantic;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1990/882/nizinski.pdf.

Waring:1990:ITM

[3595] Gordon T. Waring, Patricia Gerrior, P. Michael Payne, Betsy L. Parry, and John R. Nicolas. Incidental take of marine mammals in foreign fishery activities off the northeast United States, 1977–88. Fishery Bulletin, 88(2):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/incidental-take-marine-mammals-foreign-fishery-
activities-northeast-united-states-1977-88;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1990/882/waring.pdf.

Eggleston:1990:SCP

[3596] David B. Eggleston and Eleanor A. Bochenek. Stomach contents and parasite infestation of school bluefin tuna Thunnus thynnus collected from the Middle Atlantic Bight, Virginia. Fishery Bulletin, 88(2):389–395, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/stomach-contents-and-parasite-infestation-
school-bluefin-tuna-thunus-thynnus-collected;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1990/882/eggleston.
pdf.

Ojeda:1990:DAS

[3597] F. Patricio Ojeda and John H. Dearborn. Diversity, abundance, and spatial distribution of fishes and crustaceans in the rocky subtidal zone of the Gulf of Maine. Fishery Bulletin, 88(2):??, 1990. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/diversity-abundance-and-spatial-
distribution-fishes-and-crustaceans-rocky-subtidal-zone;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1990/882/
ojeda.pdf.

Seipt:1990:PCI

[3598] Irene E. Seipt, Phillip J. Clapham, Charles A. Mayo, and Mary P. Hawvermale. Population characteristics of individually identified fin whales Balaenoptera physalus in Massachusetts Bay. Fishery Bulletin, 88(2):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14400;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/seipt.pdf.

Hampton:1990:TSS

[3599] John Hampton and Geoffrey P. Kirkwood. Tag shedding by southern bluefin tuna Thunnus maccoyii. Fishery Bulletin, 88(2): ??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/tag-
shedding-southern-bluefin-tuna-thunnus-maccoyii;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1990/882/
hampton.pdf.

Shelton:1990:OSA

[3600] Peter A. Shelton and Larry Hutchings. Ocean stability and anchovy spawning in the Southern Benguela Current Region. Fishery Bulletin, 88(2):323–338, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/ocean-
stability-and-anchovy-spawning-southern-benguela-current-region;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1990/882/shelton.pdf.

Holland:1990:HVMa

[3601] Kim Holland, Richard Brill, and Randolph K. C. Chang. Horizontal and vertical movements of Pacific blue marlin captured and released using sportfishing gear. Fishery Bulletin, 88(2):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/horizontal-
and-vertical-movements-pacific-blue-marlin-captured-and-released-
using;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1990/882/holland.pdf.

Shields:1990:CWS

[3602] Jeffrey D. Shields and Armand M. Kurls. Carcinonemertes wickhami n. sp. (Nemertea), a symbiotic egg predator from the spiny lobster Panulirus interruptus in Southern California, with remarks on symbiont–host adaptations. Fishery Bulletin, 88(2):??, 1990. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/carcinonemertes-wickhami-
n-sp-nemertea-symbiotic-egg-predator-spiny-lobster-panulirus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1990/882/shields.pdf.

Buckley:1990:EWT

[3603] Lawrence J. Buckley, Alphonse S. Smigielski, Thomas A. Halavik, and Geoffrey C. Laurence. Effects of water temperature on size and biochemical composition of winter flounder Pseudopleuronectes americanus at hatching and feeding initiation. Fishery Bulletin, 88 (3):419–428, 1990. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
effects-water-temperature-size-and-biochemical-composition-
winter-flounder;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1990/883/buckley.pdf.

Manickchand-Heileman:1990:RAG

[3604] Sherry C. Manickchand-Heileman and Julien S. Kenny. Reproduction, age, and growth of the whitemouth croaker Micropogonias furnieri (Desmarest 1823) in Trinidad waters. Fishery Bulletin, 88(3):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/reproduction-age-and-
growth-whitemouth-croaker-micropogonias-furnieri-desmarest-
1823;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1990/883/manickchand-heileman.pdf.

Wilson:1990:IMI

[3605] Raymond R. Wilson, Jr. and Michael D. Tringali. Improved methods for isolation of fish mtDNA by ultracentrifugation and visualization of restriction fragments using fluorochrome dye: Results from Gulf of Mexico clupeids. Fishery Bulletin, 88(3):611–615, 1990. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/improved-methods-isolation-fish-
mtdna-ultracentrifugation-and-visualization-restriction;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1990/883/
wilson.pdf.

Dawson:1990:BCW

[3606] Margaret A. Dawson. Blood chemistry of the windowpane flounder scophthalmus aquosus in Long Island Sound: Geographical, seasonal, and experimental variations. Fishery Bulletin, 88(3): ??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/blood-
chemistry-windowpane-flounder-scophthalmus-aquosus-long-island-
sound-geographical;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1990/883/dawson.pdf.

Murphy:1990:RGM

[3607] Michael D. Murphy and Ronald G. Taylor. Reproduction, growth, and mortality of red drum Sciaenops acellatus in Florida waters. Fishery Bulletin, 88(3):531–542, 1990. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
reproduction-growth-and-mortality-red-drum-sciaenops-acellatus-
florida-waters;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1990/883/murphy.pdf.

Hinckley:1990:VES

[3608] Sarah Hinckley. Variation of egg size of walleye pollock Theragra chalcogramma with a preliminary examination of the effect of egg size on larval size. Fishery Bulletin, 88(3):??, 1990. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14401;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/hinckley.pdf.

Dentzau:1990:RMA

[3609] Michael W. Dentzau and Mark E. Chittenden, Jr. Reproduction, movements, and apparent population dynamics of the Atlantic threadfin Polydactylus octonemus in the Gulf of Mexico. Fishery Bulletin, 88(3):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/reproduction-movements-and-apparent-population-
dynamics-atlantic-threadfin-polydactylus;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1990/883/dentzau.pdf.

Okiyama:1990:CRS

[3610] Muneo Okiyama. Contrast in reproductive style between two species of sandfishes (family Trichodontidae). Fishery Bulletin,

88(3):??, 1990.  CODEN FSYBAY.  ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
contrast-reproductive-style-between-two-species-sandfishes-
family-trichodontidae;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1990/883/okiyama.pdf.

Grover:1990:FEL

[3611] Jill J. Grover. Feeding ecology of late-larval and early-juvenile walleye pollock Theragra chalcogramma from the Gulf of Alaska in 1987. Fishery Bulletin, 88(3):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/feeding-ecology-late-larval-and-early-juvenile-
walleye-pollock-theragra-chalcogramma-gulf;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1990/883/grover.pdf.

Ostrande:1990:DPR

[3612] Gary K. Ostrande, James J. Anderson, Jeffrey P. Fisher, Marsha L. Landolt, and Richard M. Kocan. Decreased performance of rainbow trout Oncorhynchus mykiss emergence behaviors following embryonic exposure to benzo[a]pyrene. Fishery Bulletin, 88(3): ??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/decreased-
performance-rainbow-trout-oncorhynchus-mykiss-emergence-behaviors-
following;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1990/883/ostrander.pdf.

Hoenig:1990:ERS

[3613] John M. Hoenig, Pierre Pepin, and William D. Lawing. Estimating relative survival rate for two groups of larval fishes from field data: Do older larvae survive better than young? Fishery Bulletin, 88(3):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/estimating-
relative-survival-rate-two-groups-larval-fishes-field-data-

do-older-larvae;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1990/883/hoenig.pdf.

Radtke:1990:AGA

[3614] Richard L. Radtke and Thomas F. Hourigan. Age and growth of the Antarctic fish nototheniops nudifrons. Fishery Bulletin, 88 (3):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/age-
and-growth-antarctic-fish-nototheniops-nudifrons;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1990/883/
radtke.pdf.

Holland:1990:HVMb

[3615] Kim N. Holland, Richard W. Brill, and Randolph K. C. Chang. Horizontal and vertical movements of yellowfin and bigeye tuna associated with fish aggregating devices. Fishery Bulletin, 88(3):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/horizontal-
and-vertical-movements-yellowfin-and-bigeye-tuna-associated-
fish-aggregating;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1990/883/holland.pdf.

Richards:1990:PIT

[3616] William J. Richards, Thomas Potthoff, and Jong man Kim. Problems identifying tuna larvae species (Pisces: Scombridae: Thunnus) from the Gulf of Mexico. Fishery Bulletin, 88(3):??, 1990. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/problems-identifying-tuna-larvae-
species-pisces-scombridae-thunnus-gulf-mexico;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1990/883/richards.
pdf.

Love:1990:CFT

[3617] Milton S. Love and William Westphal. Comparison of fishes taken by a sportfishing party vessel around oil platforms and adjacent natural reefs near Santa Barbara, California. Fishery Bulletin, 88(3):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/comparison-
fishes-taken-sportfishing-party-vessel-around-oil-platforms-
and-adjacent-natural;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1990/883/love.pdf.

Stoner:1990:DBQ

[3618] Allan W. Stoner and Janice M. Waite. Distribution and behavior of queen conch Strombus gigas relative to seagrass standing crop. Fishery Bulletin, 88(3):573–585, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/distribution-and-behavior-queen-conch-strombus-
gigas-relative-seagrass-standing-crop;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1990/883/stoner.pdf.

Lowry:1990:FHC

[3619] Mark S. Lowry, Charles W. Oliver, Carolyn Macky, and Jeannie B. Wexler. Food habits of California sea lions Zalophus californianus at San Clemente Island, California, 1981–86. Fishery Bulletin, 88(3):509–521, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/food-habits-california-sea-lions-zalophus-
californianus-san-clemente-island-california-1981;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1990/883/
lowry.pdf.

Wicksten:1990:KHS

[3620] Mary K. Wicksten. Key to the hippolytid shrimp of the eastern Pacific Ocean. Fishery Bulletin, 88(3):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/key-hippolytid-shrimp-eastern-pacific-ocean;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1990/883/
wicksten.pdf.

Kalish:1990:UOM

[3621] John M. Kalish. Use of otolith microchemistry to distinguish the progeny of sympatric anadromous and non-anadromous salmonids. Fishery Bulletin, 88(4):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/use-otolith-microchemistry-distinguish-progeny-
sympatric-anadromous-and-non-anadromous;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1990/884/kalish.pdf.

Taylor:1990:GPM

[3622] David M. Taylor and John M. Hoenig. Growth per molt of male snow crab Chionoecetes opilio from Conception and Bonavista Bays, Newfoundland. Fishery Bulletin, 88(4):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/growth-molt-male-snow-crab-chionoecetes-opilio-
conception-and-bonavista-bays-newfoundland;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1990/884/taylor.pdf.

Anonymous:1990:I

[3623] Anonymous. Index. Fishery Bulletin, 88(4):??, 1990. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/index-97;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
index_1.pdf.

Kulbicki:1990:CBF

[3624] Michel Kulbicki and Laurent Wantlez. Comparison between fish bycatch from shrimp trawlnet and visual censuses in St. Vincent Bay, New Caledonia. Fishery Bulletin, 88(4):??, 1990. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/comparison-between-fish-bycatch-
shrimp-trawlnet-and-visual-censuses-st-vincent-bay-new;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1990/884/
kulbicki.pdf.

Tercelro:1990:PAP

[3625] Mark Tercelro and Josef S. Idoine. A practical assessment of the performance of Shepherd’s Length Composition Analysis (SRLCA): Application to Gulf of Maine northern shrimp Pandalus borealis survey data. Fishery Bulletin, 88(4):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/practical-assessment-performance-
shepherds-length-composition-analysis-srlca-application;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1990/884/
terceiro.pdf.

Anonymous:1990:PAM

[3626] Anonymous. Publications awards, 1986–88 memorial to John Alan Gulland erratum. Fishery Bulletin, 88(4):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/publications-awards-1986-88-memorial-john-
alan-gulland-erratum;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/awards.pdf.

Page:1990:FAL

[3627] Henry M. Page and Yann O. Ricard. Food availability as a limiting factor to mussel Mytilus edulis growth in California Coastal waters. Fishery Bulletin, 88(4):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
food-availability-limiting-factor-mussel-mytilus-edulis-growth-
california-coastal-waters;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1990/884/page.pdf.

Wilk:1990:ACG

[3628] Stuart J. Wilk, Wallace W. Morse, and Linda L. Stehlik. Annual cycles of gonad-somatic indices as indicators of spawning activity for selected species of finfish collected from the New York Bight. Fishery Bulletin, 88(4):775–786, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/annual-cycles-gonad-somatic-indices-indicators-
spawning-activity-selected-species-finfish;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1990/884/wilk.pdf.

Barger:1990:AGB

[3629] Lyman E. Barger. Age and growth of bluefish Pomatomus saltatrix from the Northern Gulf of Mexico and U.S. South Atlantic Coast. Fishery Bulletin, 88(4):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/age-
and-growth-bluefish-pomatomus-saltatrix-northern-gulf-mexico-
and-us-south-atlantic;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1990/884/barger.pdf.

Payne:1990:RFA

[3630] P. Michael Payne, David N. Wiley, Sharon B. Young, Sharon Pittman, Phillip J. Clapham, and Jack W. Jossi. Recent fluctuations in the abundance of baleen whales in the southern Gulf of Maine in relation to changes in selected prey. Fishery Bulletin, 88(4):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/recent-fluctuations-
abundance-baleen-whales-southern-gulf-maine-relation-changes-
selected;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1990/884/payne.pdf.

Wisner:1990:TNS

[3631] Robert L. Wisner and Charmion B. McMillan. Three new species of hagfishes, genus Eptatretus (Cyclostomata, Myxinidae), from the Pacific Coast of North America, with new data on E. deani and E. stoutii. Fishery Bulletin, 88(4):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
three-new-species-hagfishes-genus-eptatretus-cyclostomata-myxinidae-

pacific-coast-north;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1990/884/wisner.pdf.

Fisher:1990:SSC

[3632] Joseph P. Fisher and William G. Pearcy. Spacing of scale circuli versus growth rate in young Coho salmon. Fishery Bulletin, 88(4):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
spacing-scale-circuli-versus-growth-rate-young-coho-salmon;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1990/884/fisher.pdf.

Schwartz:1990:LWA

[3633] Frank J. Schwartz. Length–weight, age and growth, and landings observations for sheepshead Archosargus probatocephalus from North Carolina. Fishery Bulletin, 88(4):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/length-weight-age-and-growth-and-landings-
observations-sheepshead-archosargus;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1990/884/schwartz.pdf.

Brodeur:1990:TRJ

[3634] Richard D. Brodeur and William G. Pearcy. Trophic relations of juvenile Pacific salmon off the Oregon and Washington coast. Fishery Bulletin, 88(4):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/trophic-relations-juvenile-pacific-salmon-
oregon-and-washington-coast;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/brodeur_0.pdf.

Grover:1990:FHL

[3635] Jill J. Grover and Bori L. Olla. Food habits of larval sablefish Anoplopoma fimbria from the Bering Sea. Fishery Bulletin,

88(4):??, 1990.  CODEN FSYBAY.  ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
food-habits-larval-sablefish-anoplopoma-fimbria-bering-sea;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1990/884/grover.pdf.

Seeb:1990:GVH

[3636] Lisa W. Seeb, James E. Seeb, and Jeffrey J. Polovina. Genetic variation in highly exploited spiny lobster Panulirus marginatus populations from the Hawaiian Archipelago. Fishery Bulletin, 88(4):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
genetic-variation-highly-exploited-spiny-lobster-panulirus-
marginatus-populations-hawaiian;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1990/884/seeb.pdf.

Pearcy:1990:DBJ

[3637] William G. Pearcy, Richard D. Brodeur, and Joseph P. Fisher. Distribution and biology of juvenile cutthroat trout Oncorhynchus clarki clarki and steelhead O. mykiss in coastal waters off Oregon and Washington. Fishery Bulletin, 88(4):697–711, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/distribution-and-biology-
juvenile-cutthroat-trout-onchorhynchus-clarki-clarki-and-steelhead;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/pearcy_0.pdf.

Hightower:1990:MHP

[3638] Joseph E. Hightower. Multispecies harvesting policies for Washington– Oregon–California rockfish trawl fisheries. Fishery Bulletin, 88(4):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/multispecies-
harvesting-policies-washington-oregon-california-rockfish-trawl-
fisheries;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1990/884/hightower.pdf.

Stanley:1990:FDB

[3639] David R. Stanley and Charles A. Wilson. A fishery-dependent based study of fish species composition and associated catch rates around oil and gas structures off Louisiana. Fishery Bulletin,

88(4):??, 1990.  CODEN FSYBAY.  ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishery-dependent-based-study-fish-species-composition-and-
associated-catch-rates-around;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1990/884/stanley.pdf.

Stevens:1990:SKT

[3640] Bradley G. Stevens. Survival of king and Tanner crabs captured by commercial sole trawls. Fishery Bulletin, 88(4):??, 1990. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/survival-king-and-tanner-crabs-
captured-commercial-sole-trawls;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/stevens.pdf.

Jagielo:1990:MTL

[3641] Thomas H. Jagielo. Movement of tagged lingcod Ophiodon elongatus at Neah Bay, Washington. Fishery Bulletin, 88(4):??, 1990. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/movement-tagged-lingcod-ophiodon-
elongatus-neah-bay-washington;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1990/884/jagielo.pdf.

Szedlmayer:1990:DGA

[3642] Stephen T. Szedlmayer, Michael P. Weinstein, and John A. Musick. Differential growth among cohorts of age-0 weakfish Cynoscion regalis in Chesapeake Bay. Fishery Bulletin, 88(4):??, 1990. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/differential-growth-
among-cohorts-age-o-weakfish-cynoscion-regalis-chesapeake-bay;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1990/884/szedlmayer.pdf.

Olson:1990:CTT

[3643] Robert J. Olson and Vernon P. Scholey. Captive tunas in a tropical marine research laboratory: Growth of late-larval and early juvenile black skipjack euthynnus lineatus. Fishery Bulletin, 88(4):??, 1990. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/captive-tunas-
tropical-marine-research-laboratory-growth-late-larval-and-
early-juvenile;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/olson.pdf.

Beckman:1991:AGR

[3644] Daniel W. Beckman, A. Louise Stanley, Jeffrey H. Render, and Charles A. Wilson. Age and growth-rate estimation of sheepshead Archosargus probatocephalus in Louisiana waters using otoliths. Fishery Bulletin, 89(1):1–8, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/age-and-growth-rate-estimation-sheepshead-
archosargus-probatocephalus-louisiana-waters;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1991/891/beckman.
pdf.

Lozano-Alvarez:1991:FCG

[3645] Enrique Lozano-Alvarez, Patricia Briones-Fourzan, and Bruce F. Phillips. Fishery characteristics, growth, and movements of the spiny lobster Panulirus argus in Bahia de la Ascension, Mexico. Fishery Bulletin, 89(1):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishery-characteristics-growth-and-movements-spiny-lobster-
panulirus-argus-bahia-de-la;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1991/891/lozano-alvarez.pdf.

DeMartini:1991:AVF

[3646] Edward E. DeMartini. Annual variations in fecundity, egg size, and the gonadal and somatic conditions of queenfish Seriphus politus (Sciaenidae). Fishery Bulletin, 89(1):??, 1991. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/annual-variations-fecundity-egg-
size-and-gonadal-and-somatic-conditions-queenfish-seriphus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1991/891/demartini.pdf.

Matlock:1991:CTT

[3647] Gary C. Matlock, Walter R. Nelson, Robert S. Jones, Albert W. Green, Terry J. Cody, Elmer Gutherz, and Jeff Doerzbacher. Comparison of two techniques for estimating tilefish, yellowedge grouper, and other deepwater fish populations. Fishery Bulletin, 89(1):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/comparison-two-techniques-estimating-
tilefish-yellowedge-grouper-and-other-deepwater-fish;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1991/891/
matlock.pdf.

Ferreira:1991:AGS

[3648] Beatrice Padovanl Ferreira and Carolus Maria Vooren. Age, growth, and structure of vertebra in the school shark Galeorhinus galeus (Linnaeus, 1758) from Southern Brazil. Fishery Bulletin, 89(1):??, 1991. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/age-growth-and-structure-
vertebra-school-shark-galeorhinus-galeus-linnaeus-1758-southern;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1991/891/ferreira.pdf.

Murphy:1991:RAG

[3649] Michael D. Murphy and Mark E. Chittenden, Jr. Reproduction, age and growth, and movements of the Gulf butterfish Peprilus burti. Fishery Bulletin, 89(1):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/reproduction-age-and-growth-and-movements-gulf-
butterfish-peprilus-burti;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1991/891/murphy.pdf.

Grant:1991:CCS

[3650] George C. Grant. Chaetognatha from the Central and Southern Middle Atlantic Bight: Species composition, temperature-salinity relationships, and interspecific associations. Fishery Bulletin, 89(1):??, 1991. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/chaetognatha-central-and-
southern-middle-atlantic-bight-species-composition-temperature;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1991/891/grant.pdf.

Overholtz:1991:ICR

[3651] William J. Overholtz, Steven A. Murawski, and William L. Michaels. Impact of compensatory responses on assessment advice for the Northwest Atlantic mackerel stock. Fishery Bulletin, 89(1):??, 1991. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/impact-compensatory-
responses-assessment-advice-northwest-atlantic-mackerel-stock;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1991/891/overholtz.pdf.

Hannah:1991:FIC

[3652] Robert W. Hannah and Stephen A. Jones. Fishery-induced changes in the population structure of pink shrimp Pandalus jordani. Fishery Bulletin, 89(1):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
induced-changes-population-structure-pink-shrimp-pandalus-jordani;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1991/891/hannah.pdf.

Polovlna:1991:EHR

[3653] Jeffrey J. Polovlna. Evaluation of hatchery releases of juveniles to enhance rockfish stocks, with application to Pacific Ocean perch Sebastes alutus. Fishery Bulletin, 89(1):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/evaluation-hatchery-releases-juveniles-enhance-
rockfish-stocks-application-pacific-ocean;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1991/891/polovina.
pdf.

Kimura:1991:BRB

[3654] Daniel K. Kimura and Julaine J. Lyons. Between-reader bias and variability in the age-determination process. Fishery Bulletin, 89(1):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/between-reader-bias-
and-variability-age-determination-process;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1991/891/kimura.pdf.

Somerton:1991:DDF

[3655] David A. Somerton. Detecting differences in fish diets. Fishery Bulletin, 89(1):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/detecting-differences-fish-diets;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1991/891/
somerton.pdf.

Lavalli:1991:SGE

[3656] Kari L. Lavalli. Survival and growth of early-juvenile American lobsters Homarus americanus through their first season while fed diets of mesoplankton, microplankton, and frozen brine shrimp. Fishery Bulletin, 89 (1):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/survival-and-
growth-early-juvenile-american-lobsters-homarus-americanus-
through-their-first;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1991/891/lavalli.pdf.

Stanley:1991:FAA

[3657] David R. Stanley and Charles A. Wilson. Factors affecting the abundance of selected fishes near oil and gas platforms in the Northern Gulf of Mexico. Fishery Bulletin, 89(1):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/factors-affecting-abundance-selected-
fishes-near-oil-and-gas-platforms-northern-gulf-mexico;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1991/891/
stanley.pdf.

Ledgerwood:1991:DSM

[3658] Richard D. Ledgerwood, Frank. P. Thrower, and Earl M. Dawley. Diel sampling of migratory juvenile salmonids in the Columbia River Estuary. Fishery Bulletin, 89(1):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/diel-sampling-migratory-juvenile-salmonids-columba-

river-estuary;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1991/891/ledgerwood.pdf.

Temte:1991:USP

[3659] Jonathan L. Temte. Use of serum progesterone and testosterone to estimate sexual maturity in Dall’s porpoise Phocoenoides dalli. Fishery Bulletin, 89(1):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/use-
serum-progesterone-and-testosterone-estimate-sexual-maturity-
dalls-porpoise;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1991/891/temte.pdf.

Goebel:1991:DPF

[3660] Michael E. Goebel, John L. Bengtson, Robert L. Delong, Roger L. Gentry, and Thomas R. Loughlin. Diving patterns and foraging locations of female northern fur seals. Fishery Bulletin, 89 (2):171–179, 1991. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
diving-patterns-and-foraging-locations-female-northern-fur-
seals;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1991/892/goebel.pdf.

Munroe:1991:WAT

[3661] Thomas A. Munroe. Western Atlantic tonguefishes of the Symphurus plagusia complex (Cynoglossidae: Pleuronectiformes), with descriptions of two new species. Fishery Bulletin, 89(2):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/western-atlantic-tonguefishes-symphurus-
plagusia-complex-cynoglossidae-pleuronectiformes;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1991/892/
munroe.pdf.

Govoni:1991:PPF

[3662] John J. Govoni and John E. Olney. Potential predation on fish eggs by the lobate ctenophore Mnemiopsis leidyi within and outside the Chesapeake Bay plume. Fishery Bulletin, 89(2):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/potential-predation-fish-eggs-lobate-
ctenophore-mnemiopsis-leidyi-within-and-outside;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1991/892/
govoni.pdf.

Schaefer:1991:GVM

[3663] Kurt M. Schaefer. Geographic variation in morphometric characters and gill-raker counts of yellowfin tuna Thunnus albacares from the Pacific Ocean. Fishery Bulletin, 89(2):??, 1991. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/geographic-variation-
morphometric-characters-and-gill-raker-counts-yellowfin-tuna-
thunnus;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1991/892/schaefer.pdf.

Grant:1991:DSB

[3664] George C. Grant and John E. Olney. Distribution of striped bass Marone saxatilis (Walbaum) eggs and larvae in major Virginia rivers. Fishery Bulletin, 89(2):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/distribution-striped-bass-marone-saxatilis-walbaum-
eggs-and-larvae-major-virginia-rivers;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1991/892/grant.pdf.

Shields:1991:FRP

[3665] Jeffrey D. Shields, Robert K. Okazaki, and Armand M. Kurls. Fecundity and the reproductive potential of the yellow rock crab Cancer anthonyi. Fishery Bulletin, 89(2):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/fecundity-and-reproductive-potential-
yellow-rock-crab-cancer-anthonyi;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1991/892/shields.pdf.

Kramer:1991:GMM

[3666] Sharon Hendrix Kramer. Growth, mortality, and movements of juvenile California halibut Paralichthys californicus in shallow coastal and bay habitats of San Diego County, California. Fishery Bulletin, 89(2):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
growth-mortality-and-movements-juvenile-california-halibut-
paralichthys-californicus;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1991/892/kramer.pdf.

Somerton:1991:RWL

[3667] David A. Somerton and Donald R. Kobayashi. Robustness of the Wetherall length-based method to population disequilibria. Fishery Bulletin, 89 (2):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/robustness-
wetherall-length-based-method-population-disequilibria;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1991/892/
somerton.pdf.

Lowry:1991:SAV

[3668] Mark S. Lowry, Brent S. Stewart, Carolyn B. Heath, Pamela K. Yochem, and John M. Francis. Seasonal and annual variability in the diet of California sea lions Zalophus californianus at San Nicolas Island, California, 1981–86. Fishery Bulletin, 89(2):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/seasonal-and-annual-variability-diet-
california-sea-lions-zalophus-californianus-san;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1991/892/
lowry.pdf.

Sutter:1991:MPS

[3669] Frederick C. Sutter III, Roy O. Williams, and Mark F. Godcharles. Movement patterns and stock affinities of king mackerel in the southeastern United States. Fishery Bulletin, 89(2):315– 324, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/movement-
patterns-and-stock-affinities-king-mackerel-southeastern-united-
states;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1991/892/sutter.pdf.

Matsui:1991:DYM

[3670] Tetsuo Matsui. Description of young of the mesopelagic Platytroctids holtbyrnia latifrons and Sagamichthys abei (Pisces, Alepocephaloidea) from the Northeastern Pacific Ocean. Fishery Bulletin, 89(2):209–219, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/description-
young-mesopelagic-platytroctids-holtbyrnia-latifrons-and-sagamichthys-
abei;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1991/892/matsui.pdf.

Szedlmayer:1991:AEC

[3671] Stephen T. Szedlmayer, Margaret M. Szedlmayer, and Michael E. Sieracki. Automated enumeration by computer digitization of age-0 weakfish Cynoscion regalis scale circuli. Fishery Bulletin, 89(2): ??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/automated-
enumeration-computer-digitization-age-o-weakfish-cynoscion-
regalis-scale-circuli;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1991/892/szedlmayer.pdf.

Milton:1991:MSS

[3672] David A. Milton and Stephen J. M. Blaber. Maturation, spawning seasonality, and proximate spawning stimuli of six species of tuna baitfish in the Solomon Islands. Fishery Bulletin, 89(2):??, 1991. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/maturation-spawning-seasonality-
and-proximate-spawning-stimuli-six-species-tuna-baitfish;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1991/
892/milton.pdf.

Taylor:1991:ETA

[3673] David M. Taylor and John M. Hoenig. Effect of tag anchor location on retention/survival through molt in male snow crabs Chionoecetes opilio. Fishery Bulletin, 89(2):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/effect-tag-anchor-location-retentionsurvival-through-
molt-male-snow-crabs-chionoecetes;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1991/892/taylor.pdf.

Mugiya:1991:BRB

[3674] Yasuo Mugiya and Hirotaka Oka. Biochemical relationship between otolith and somatic growth in the rainbow trout Oncorhynchus mykiss: Consequence of starvation, resumed feeding, and diel variations. Fishery Bulletin, 89(2):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/biochemical-relationship-between-
otolith-and-somatic-growth-rainbow-trout-oncorhynchus;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1991/892/
mugiya.pdf.

Baldwin:1991:LPK

[3675] Carole C. Baldwin and G. David Johnson. A larva of the poorly known serranid fish Jeboehlkia gladifer (Teleostei: Serranidae: Epinephelinae). Fishery Bulletin, 89(3):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/larva-poorly-known-serranid-fish-jeboehlkia-gladifer-
teleostei-serranidae-epinephelinae;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1991/893/baldwin.pdf.

Power:1991:SOS

[3676] James H. Power and L. Nelson May, Jr. Satellite observed seasurface temperatures and yellowfin tuna catch and effort in the Gulf of Mexico. Fishery Bulletin, 89(3):??, 1991. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/satellite-observed-sea-surface-
temperatures-and-yellowfin-tuna-catch-and-effort-gulf-mexico;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1991/893/power.pdf.

Dailey:1991:PFT

[3677] Murray D. Dailey and Wolfgang K. Vogelbein. Parasite fauna of three species of Antarctic whales with reference to their use as potential stock indicators. Fishery Bulletin, 89(3):355–365, 1991. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/parasite-fauna-three-
species-antarctic-whales-reference-their-use-potential-stock;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1991/893/dailey.pdf.

Prince:1991:EAG

[3678] Eric D. Prince, Dennis W. Lee, James R. Zweifel, and Edward B. Brothers. Estimating age and growth of young Atlantic blue marlin makalra nigricans from otolith microstructure. Fishery Bulletin, 89(3):441–459, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/estimating-age-and-
growth-young-atlantic-blue-marlin-makalra-nigricans-otolith;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1991/893/prince.pdf.

Forney:1991:DTH

[3679] Karin A. Forney, Doyle A. Hanan, and Jay Barlow. Detecting trends in harbor porpoise abundance from aerial surveys using analysis of covariance. Fishery Bulletin, 89(3):??, 1991. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/detecting-trends-harbor-porpoise-
abundance-aerial-surveys-using-analysis-covariance;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1991/893/
forney.pdf.

Smith:1991:EAG

[3680] M. Kimberly Smith and Eric Kostlan. Estimates of age and growth of ehu Etelis carbunculus in four regions of the Pacific from density of daily increments in otoliths. Fishery Bulletin, 89(3):461– 472, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/estimates-
age-and-growth-ehu-etelis-carbunculus-four-regions-pacific-
density-daily;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1991/893/smith.pdf.

Kleiber:1991:CPE

[3681] Pierre Kleiber and Christina Perrin. Catch-per-effort and stock status in the U.S. North Pacific albacore fishery: Reappraisal of both. Fishery Bulletin, 89(3):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/catch-effort-and-stock-status-us-north-pacific-
albacore-fishery-reappraisal-both;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1991/893/kleiber.pdf.

Stehlik:1991:DAF

[3682] Linda L. Stehlik, Clyde L. MacKenzie, Jr., and Wallace W. Morse. Distribution and abundance of four brachyuran crabs on the Northwest Atlantic shelf. Fishery Bulletin, 89(3):??, 1991. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/distribution-and-abundance-four-
brachyuran-crabs-northwest-atlantic-shelf;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1991/893/stehlik.pdf.

Long:1991:APW

[3683] Douglas J. Long. Apparent predation by a white shark Carcharodon carcharias on a pygmy sperm whale Kogia breviceps. Fishery Bulletin, 89(3):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/apparent-
predation-white-shark-carcharodon-carcharias-pygmy-sperm-whale-
kogia-breviceps;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1991/893/long.pdf.

Sudekum:1991:LHE

[3684] Anthony E. Sudekum, James D. Parrish, Richard L. Radtke, and Stephen Ralston. Life history and ecology of large jacks in undisturbed, shallow, oceanic communities. Fishery Bulletin, 89(3):493–513, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/life-history-and-ecology-large-
jacks-undisturbed-shallow-oceanic-communities;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1991/893/sudekum.
pdf.

Oda:1991:DEL

[3685] Debra Oda. Development of eggs and larvae of California halibut Paralichthys californicus and fantail sole Xystreurys liolepis (Pisces: Paralichthyidae). Fishery Bulletin, 89(3):??, 1991. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/development-eggs-and-
larvae-california-halibut-paralichthys-californicus-and-fantail-
sole;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1991/893/oda.pdf.

Vecchione:1991:OPE

[3686] Michael Vecchione. Observations on the paralarval ecology of a euryhaline squid Lolliguncula brevis (Cephalopoda: Loliginidae). Fishery Bulletin, 89(3):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
observations-paralarval-ecology-euryhaline-squid-lolliguncula-
brevis-cephalopoda;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1991/893/vecchione.pdf.

Pearson:1991:AGP

[3687] Donald E. Pearson, Joseph E. Hightower, and Jacqueline T. H. Chan. Age, growth, and potential yield for shortbelly rockfish Sebastes jordani. Fishery Bulletin, 89(3):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/age-growth-and-potential-yield-shortbelly-rockfish-
sebastes-jordani;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1991/893/pearson.pdf.

Woodbury:1991:IVG

[3688] David Woodbury and Stephen Ralston. Interannual variation in growth rates and back-calculated birthdate distributions of pelagic juvenile rockfishes (Sebastes spp.) off the Central California Coast. Fishery Bulletin, 89(3):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
interannual-variation-growth-rates-and-back-calculated-birthdate-
distributions-pelagic;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1991/893/woodbury.pdf.

Au:1991:PNT

[3689] David W. Au. Polyspecific nature of tuna schools: Shark, dolphin, and seabird associates. Fishery Bulletin, 89(3):??, 1991. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/polyspecific-nature-
tuna-schools-shark-dolphin-and-seabird-associates;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1991/893/
au.pdf.

Perrin:1991:GVE

[3690] William F. Perrin, Priscilla A. Akin, and Jerry V. Kashiwada. Geographic variation in external morphology of the spinner dolphin Stenella longirostris in the eastern Pacific and implications for conservation. Fishery Bulletin, 89(3):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
geographic-variation-external-morphology-spinner-dolphin-stenella-
longirostris-eastern;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1991/893/perrin.pdf.

Boggs:1991:BGN

[3691] Christofer H. Boggs. Bioenergetics and growth of northern anchovy Engraulis mordax. Fishery Bulletin, 89(4):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/bioenergetics-and-growth-
northern-anchovy-engraulis-mordax;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1991/894/boggs.pdf.

Laidig:1991:DGE

[3692] Thomas E. Laidig, Stephen Ralston, and James R. Bence. Dynamics of growth in the early life history of shortbelly rockfish Sebastes jordani. Fishery Bulletin, 89(4):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/dynamics-growth-early-life-history-shortbelly-rockfish-
sebastes-jordani;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1991/894/laidig.pdf.

Schiel:1991:PSA

[3693] David R. Schiel and Paul A. Breen. Population structure, ageing, and fishing mortality of the New Zealand abalone Haliotis iris. Fishery Bulletin, 89(4):681–691, 1991. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/population-structure-ageing-and-
fishing-mortality-new-zealand-abalone-haliotis-iris;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1991/894/
schiel.pdf.

Booman:1991:RSN

[3694] Clelia Booman, Arild Folkvord, and John R. Hunter. Responsiveness of starved northern anchovy Engraulis mordax larvae to predatory attacks by adult anchovy. Fishery Bulletin, 89(4):??, 1991. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/responsiveness-starved-northern-
anchovy-engraulis-mordax-larvae-predatory-attacks-adult;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1991/894/
booman.pdf.

Lemaitre:1991:LSN

[3695] Rafael Lemaitre and Sérgio de Almeida Rodrigues. Lepidophthalmus sinuensis: a new species of ghost shrimp (Decapoda: Thalassinidea: Callianassidae) of importance to the commercial culture of penaeid shrimps on the Caribbean Coast of Colombia, with observations on its ecology. Fishery Bulletin, 89(4):

623–630, 1991.  CODEN FSYBAY.  ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
lepidophthalmus-sinuensis-new-species-ghost-shrimp-decapoda-
thalassinidea-callianassidae;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1991/894/lemaitre.pdf.

Stoner:1991:EAH

[3696] Allan W. Stoner and Veronique J. Sandt. Experimental analysis of habitat quality for juvenile queen conch in seagrass meadows. Fishery Bulletin, 89(4):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
experimental-analysis-habitat-quality-juvenile-queen-conch-
seagrass-meadows;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1991/894/stoner.pdf.

Caillouet:1991:STS

[3697] Charles W. Caillouet, Jr., Marcel J. Duronslet, Andre M. Landry, Jr., Dickie B. Revera, Donna J. Shaver, Kerry M. Stanley, Robert W. Heinly, and Erich K. Stabenau. Sea turtle strandings and shrimp fishing effort in the Northwestern Gulf of Mexico, 1986–89. Fishery Bulletin, 89(4):712– 718, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/sea-turtle-
strandings-and-shrimp-fishing-effort-northwestern-gulf-mexico-
1986-89;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1991/894/calliouet.pdf.

Lyczkowski-Shultz:1991:DVD

[3698] Joanne Lyczkowski-Shultz and John P. Steen, Jr. Diel vertical distribution of red drum Sciaenops acellatus larvae in the Northcentral Gulf of Mexico. Fishery Bulletin, 89(4):??, 1991. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/diel-vertical-distribution-red-
drum-sciaenops-acellatus-larvae-northcentral-gulf-mexico;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1991/
894/lyczkowski-shultz.pdf.

Sutter:1991:GMK

[3699] Frederick C. Sutter III, Roy O. Williams, and Mark F. Godcharles. Growth and mortality of king mackerel Scomberomorus cavalla tagged in the Southeastern United States. Fishery Bulletin, 89(4):733–737, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/growth-and-mortality-
king-mackerel-scomberomorus-cavalla-tagged-southeastern-united-
states;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1991/894/sutter.pdf.

Gadomski:1991:ETE

[3700] Dena M. Gadomski and Steven M. Caddell. Effects of temperature on early-life-history stages of California halibut Paralichthys californicus. Fishery Bulletin, 89(4):??, 1991. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/effects-temperature-early-life-
history-stages-california-halibut-paralichthys-californicus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1991/894/gadomski.pdf.

McKenna:1991:TRW

[3701] James E. McKenna, Jr. Trophic relationships within the Antarctic demersal fish community of South Georgia Island. Fishery Bulletin, 89 (4):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/trophic-
relationships-within-antarctic-demersal-fish-community-south-
georgia-island;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1991/894/mckenna.pdf.

Gamboa:1991:OSV

[3702] David Aurloles Gamboa. Otolith size versus weight and body-length relationships for eleven fish species of Baja California, Mexico. Fishery Bulletin, 89(4):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
otolith-size-versus-weight-and-body-length-relationships-eleven-
fish-species-baja;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1991/894/gamboa.pdf.

Overholtz:1991:DCP

[3703] William J. Overholtz and Gordon T. Waring. Diet composition of pilot whales Globicephala sp. and common dolphins Delphinus delphis in the Mid-Atlantic Bight during spring 1989. Fishery Bulletin, 89(4):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/diet-
composition-pilot-whales-globicephala-sp-and-common-dolphins-
delphinus-delphis-mid;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1991/894/overholtz.pdf.

Grover:1991:TRA

[3704] Jill J. Grover. Trophic relationship of age-0 and age-1 walleye pollock Theragra chalcogramma collected together in the eastern Bering Sea. Fishery Bulletin, 89(4):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/trophic-relationship-age-0-and-age-1-
walleye-pollock-theragra-chalcogramma-collected;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1991/894/
grover.pdf.

Peres:1991:SDR

[3705] Mónica Brick Peres and Carolus Maria Vooren. Sexual development, reproductive cycle, and fecundity of the school shark Galeorhinus galeus off southern Brazil. Fishery Bulletin, 89(4):??, 1991. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/sexual-development-reproductive-
cycle-and-fecundity-school-shark-galeorhinus-galeus;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1991/894/
peres.pdf.

Hampton:1991:EOS

[3706] John Hampton. Estimation ot southern bluefin tuna Thunnus maccoyii growth parameters from tagging data, using von bertalantfy models incorporating individual variation. Fishery Bulletin, 89(4):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/estimation-ot-
southern-bluefin-tuna-thunnus-maccoyii-growth-parameters-tagging-

data-using;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1991/894/hampton1.pdf.

Rey:1991:RBH

[3707] Juan Carlos Rey and Ramon Munoz-Chapuli. Relation between hook depth and fishing efficiency in surface longline gear. Fishery Bulletin, 89(4):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/relation-
between-hook-depth-and-fishing-efficiency-surface-longline-
gear;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1991/894/rey.pdf.

Anderson:1991:AGM

[3708] Paul J. Anderson. Age, growth, and mortality of the northern shrimp Pandalus borealis Kröyer in Pavlof Bay, Alaska. Fishery Bulletin, 89(4):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
age-growth-and-mortality-northern-shrimp-pandalus-borealis-
kroyer-pavlof-bay-alaska;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1991/894/anderson.pdf.

Hampton:1991:ESB

[3709] John Hampton. Estimation of southern bluefin tuna Thunnus maccoyii natural mortality and movement rates from tagging experiments. Fishery Bulletin, 89(4):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/estimation-southern-bluefin-tuna-
thunnus-maccoyii-natural-mortality-and-movement-rates;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1991/894/
hampton2.pdf.

Robichaud:1991:DSC

[3710] David A. Robichaud, Robert W. Elner, and Richard F. J. Bailey. Differential selection of crab Chionoecetes opilio and hyas spp. as prey by sympatric cod Gadus morhua and thorny skate Raja radiata. Fishery Bulletin, 89(4):??, 1991. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
differential-selection-crab-chionoecetes-opilio-and-hyas-spp-
prey-sympatric-cod-gadus;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1991/894/robichaud.pdf.

Buckland:1992:ETA

[3711] Stephen T. Buckland, Karen L. Cattanach, and Alejandro A. Anganuzzi. Estimating trends in abundance of dolphins associated with tuna in the eastern tropical Pacific Ocean, using sightings data collected on commercial tuna vessels. Fishery Bulletin, 90(1):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/estimating-
trends-abundance-dolphins-associated-tuna-eastern-tropical-
pacific-ocean-using;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1992/901/buckland.pdf.

Safford:1992:LBG

[3712] Susan E. Safford and Henry Booke. Lack of biochemical genetic and morphometric evidence for discrete stocks of Northwest Atlantic herring Clupea harengus harengus. Fishery Bulletin, 90(1):??, 1992. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/lack-biochemical-genetic-
and-morphometric-evidence-discrete-stocks-northwest-atlantic;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1992/901/safford.pdf.

Chen:1992:CFG

[3713] Welhzong Chen, John J. Govoni, and Stanley M. Warlen. Comparison of feeding and growth of larval round herring Etrumeus teres and Gulf menhaden Brevoortia patronus. Fishery Bulletin, 90(1):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/comparison-
feeding-and-growth-larval-round-herring-etrumeus-teres-and-
gulf-menhaden;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1992/901/chen.pdf.

Sissenwine:1992:INZ

[3714] Michael P. Sissenwine and Pamela M. Mace. ITOs in New Zealand: The era of fixed quota in perpetuity. Fishery Bulletin, 90(1):147– 160, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/itos-
new-zealand-era-fixed-quota-perpetuity;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1992/901/sissenwine.
pdf.

DAmours:1992:ACO

[3715] Denis D’Amours and François Gregoire. Analytical correction for oversampled Atlantic mackerel Scomber scombrus eggs collected with oblique plankton tows. Fishery Bulletin, 90(1):190–196, 1992. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/analytical-correction-oversampled-
atlantic-mackerel-scomber-scombrus-eggs-collected-oblique;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1992/
901/d%27amours.pdf.

Sterglou:1992:VMC

[3716] Konstantinos I. Sterglou. Variability of monthly catches of anchovy Engraulis encrasicolus in the Aegean Sea. Fishery Bulletin, 90(1):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/variability-monthly-
catches-anchovy-engraulis-encrasicolus-aegean-sea;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1992/901/
sterglou.pdf.

Douglas:1992:GVC

[3717] Michael E. Douglas, Gary D. Schnell, Daniel J. Hough, and William F. Perrin. Geographic variation in cranial morphology of spinner dolphins Stenella longirostris in the eastern tropical Pacific Ocean. Fishery Bulletin, 90(1):54–76, 1992. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
geographic-variation-cranial-morphology-spinner-dolphins-stenella-
longirostris-eastern;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1992/901/douglas.pdf.

Stoner:1992:SRA

[3718] Allan W. Stoner, Veronique J. Sandt, and Isabelle F. Boldron-Metalron. Seasonality in reproductive activity and larval abundance of queen conch Strombus gigas. Fishery Bulletin, 90(1):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/seasonality-reproductive-
activity-and-larval-abundance-queen-conch-strombus-gigas;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1992/
901/stoner.pdf.

Gall:1992:GVP

[3719] Graham A. E. Gall, Devin Bartley, Boyd Bentley, Jon Brodziak, Richard Gomulkiewicz, and Marc Mangel. Geographic variation in population genetic structure of Chinook salmon from California and Oregon. Fishery Bulletin, 90(1):??, 1992. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/geographic-variation-population-

genetic-structure-chinook-salmon-california-and-oregon;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1992/901/
gall.pdf.

Wainwright:1992:PED

[3720] Thomas C. Wainwright, David A. Armstrong, Paul A. Dinnel, Jose M. Orensanz, and Katherine A. McGraw. Predicting effects of dredging on a crab population: an equivalent adult loss approach. Fishery Bulletin, 90(1):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/predicting-effects-dredging-crab-
population-equivalent-adult-loss-approach;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1992/901/wainwright.
pdf.

Hunter:1992:FSM

[3721] J. Roe Hunter, Beverly J. Macewicz, N. Chyan huei Lo, and Carol A. Kimbrell. Fecundity, spawning, and maturity of female Dover sole, Microstomus pacificus, with an evaluation of assumptions and precision. Fishery Bulletin, 90(1):101–128, 1992. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/fecundity-spawning-and-maturity-
female-dover-sole-microstomus-pacificus-evaluation;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1992/901/
hunter.pdf.

Nakatani:1992:CEL

[3722] Toshikuni Nakatani and Arthur W. Kendall, Jr. Comparisons of earlylife-history characteristics of walleye pollock Theragra chalcogramma in Shelikof Strait, Gulf of Alaska, and Funka Bay, Hokkaido, Japan. Fishery Bulletin, 90(1):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/comparisons-early-life-history-characteristics-
walleye-pollock-theragra-chalcogramma;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1992/901/kendall.pdf.

Rajaguru:1992:ABS

[3723] Arjuna Rajaguru and Gopalsamy Shantha. Association between the sessile barnacle Xenobalanus globicipitis (Coronulidae) and the bottlenose dolphin Tursiops truncatus (Delphinidae) from the Bay of Bengal, India, with a summary of previous records from cetaceans. Fishery Bulletin, 90 (1):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/association-

between-sessile-barnacle-xenobalanus-globicipitis-coronulidae-
and-bottlenose;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1992/901/rajaguru.pdf.

Armstrong:1992:AGR

[3724] Michael P. Armstrong, John A. Musick, and James A. Colvocoresses. Age, growth, and reproduction of the goosefish Lophius americanus (Pisces: Lophiiformes). Fishery Bulletin, 90(2):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/age-growth-
and-reproduction-goosefish-lophius-americanus-pisces-lophiiformes;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1992/902/armstrong.pdf.

Markle:1992:MOE

[3725] Douglas F. Markle, Phillip M. Harris, and Christopher L. Toole. Metamorphosis and an overview of early-life-history stages in Dover sole Microstomus pacificus. Fishery Bulletin, 90(2):??, 1992. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/metamorphosis-and-overview-early-
life-history-stages-dover-sole-microstomus-pacificus;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1992/902/
markle.pdf.

Toole:1992:EMP

[3726] Christopher L. Toole and Roger L. Nielsen. Effects of microprobe precision on hypotheses related to otolith Sr:Ca ratios. Fishery Bulletin, 90 (2):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/effects-
microprobe-precision-hypotheses-related-otolith-srca-ratios;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1992/902/toole.pdf.

Bowers:1992:ARC

[3727] Michael J. Bowers. Annual reproductive cycle of oocytes and embryos of yellowtail rockfish Sebastes flavidus (family Scorpaenidae). Fishery Bulletin, 90(2):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
annual-reproductive-cycle-oocytes-and-embryos-yellowtail-rockfish-
sebastes-flavidus-family;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1992/902/bowers.pdf.

Matlock:1992:GFF

[3728] Gary C. Matlock. Growth of five fishes in Texas bays in the 1960s. Fishery Bulletin, 90(2):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/growth-five-fishes-texas-bays-1960s;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1992/902/
matlock.pdf.

Bullock:1992:AGR

[3729] Lewis H. Bullock, Michael D. Murphy, Mark F. Godcharles, and Michael E. Mitchell. Age, growth, and reproduction of jewfish Epinephelus itajara in the eastern Gulf of Mexico. Fishery Bulletin, 90(2):243–249, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/age-
growth-and-reproduction-jewfish-epinephelus-itajara-eastern-
gulf-mexico;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1992/902/bullock.pdf.

Parrack:1992:ESA

[3730] Michael L. Parrack. Estimating stock abundance from size data. Fishery Bulletin, 90(2):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/estimating-stock-abundance-size-data;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1992/902/
parrack.pdf.

Bumguardner:1992:LTC

[3731] Britt W. Bumguardner, Robert L. Colura, and Gary C. Matlock. Longterm coded wire tag retention in juvenile Sciaenops ocellatus. Fishery Bulletin, 90(2):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
long-term-coded-wire-tag-retention-juvenile-sciaenops-ocellatus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1992/902/bumguardner.pdf.

Rajaguru:1992:BTC

[3732] Arjuna Rajaguru. Biology of two co-occurring tonguefishes, Cynoglossus arel and C. lida (Pleuronectiformes: Cynoglossidae), from Indian waters. Fishery Bulletin, 90(2):??, 1992. CODEN FSY-

BAY.  ISSN 0090-0656 (print), 1937-4518 (electronic).  URL

https://spo.nmfs.noaa.gov/content/biology-two-co-occurring-
tonguefishes-cynoglossus-arel-and-c-lida-pleuronectiformes;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1992/902/rajaguru.pdf.

Campton:1992:GPA

[3733] Donald E. Campton, Carl J. Berg, Jr., Lynn M. Robison, and Robert A. Glazer. Genetic patchiness among populations of queen conch Strombus gigas in the Florida keys and bimini. Fishery Bulletin, 90(2):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/genetic-patchiness-among-
populations-queen-conch-strombus-gigas-florida-keys-and-bimini;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1992/902/campton.pdf.

Restrepo:1992:MMP

[3734] Victor R. Restrepo. A mortality model for a population in which harvested individuals do not necessarily die: The stone crab. Fishery Bulletin, 90(2):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
mortality-model-population-which-harvested-individuals-do-not-
necessarily-die-stone-crab;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1992/902/restrepo.pdf.

Davis:1992:GML

[3735] Tim L. O. Davis and Grant J. West. Growth and mortality of Lutjanus vittus (Quoy and Gaimard) from the North West Shelf of Australia. Fishery Bulletin, 90(2):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/growth-and-mortality-lutjanus-vittus-quoy-
and-gaimard-north-west-shelf-australia;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1992/902/davis.pdf.

Salvado:1992:OCD

[3736] Carlos A. M. Salvado, Pierre Kleiber, and Andrew E. Dizon. Optimal course by dolphins for detection avoidance. Fishery Bulletin, 90(2):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
optimal-course-dolphins-detection-avoidance;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1992/902/salvado.
pdf.

Hettler:1992:CWT

[3737] William F. Hettler. Correlation of winter temperature and landings of pink shrimp Penaeus duorarum in North Carolina. Fishery Bulletin, 90 (2):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/correlation-
winter-temperature-and-landings-pink-shrimp-penaeus-duorarum-
north-carolina;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1992/902/hettler.pdf.

Somerton:1992:IMM

[3738] David A. Somerton and Donald R. Kobayashi. Inverse method for mortality and growth estimation: a new method for larval fishes. Fishery Bulletin, 90(2):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/inverse-method-mortality-and-growth-estimation-new-
method-larval-fishes;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1992/902/somerton.pdf.

Hyndes:1992:ISO

[3739] Glenn A. Hyndes, Nell R. Loneragan, and Ian C. Potter. Influence of sectioning otoliths on marginal increment trends and age and growth estimates for the flathead Platycephalus speculator. Fishery Bulletin, 90 (2):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/influence-
sectioning-otoliths-marginal-increment-trends-and-age-and-growth-
estimates;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1992/902/hyndes.pdf.

Stone:1992:SDR

[3740] Heath H. Stone and Brian M. Jessup. Seasonal distribution of river herring (Hosa pseudaharengus) and A. aestivalis off the Atlantic coast of Nova Scotia. Fishery Bulletin, 90(2):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/seasonal-distribution-river-herring-hosa-pseudaharengus-
and-aestivalis-atlantic-coast-nova;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1992/902/stone.pdf.

Campbell:1992:ETR

[3741] R. Page Campbell, Terry J. Cody, C. E. Bryan, Gary C. Matlock, Maury F. Osborn, and Albert W. Green. An estimate of tagreporting rate of commercial shrimpers in two Texas bays. Fishery Bulletin, 90(3):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
estimate-tag-reporting-rate-commercial-shrimpers-two-texas-
bays;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1992/903/campbell.pdf.

Sadovy:1992:AGR

[3742] Yvonne Sadovy, Miguel Figuerola, and Ana Roman. Age and growth of red hind Epinephelus guttatus in Puerto Rico and St. Thomas. Fishery Bulletin, 90(3):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
age-and-growth-red-hind-epinephelus-guttatus-puerto-rico-and-
st-thomas;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1992/903/sadovy.pdf.

Zuniga:1992:LDT

[3743] Humberto N. Zuñiga and Enzo S. Acuña. Larval development of two sympatric flounders, Paralichthys adspersus (Steindachner, 1867) and Paralichthys microps (Günther, 1881) from the Bay of Coquimbo, Chile. Fishery Bulletin, 90(3):607–620, 1992. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/larval-development-two-sympatric-
flounders-paralichthys-adspersus-steindachner-1867-and;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1992/903/
zuniga.pdf.

Edwards:1992:PDL

[3744] Elizabeth F. Edwards and Peter C. Perkins. Power to detect linear trends in dolphin abundance: Estimates from tuna-vessel observer data, 1975–89. Fishery Bulletin, 90(3):??, 1992. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/power-detect-linear-
trends-dolphin-abundance-estimates-tuna-vessel-observer-data-
1975-89;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1992/903/edwards.pdf.

Sogard:1992:ELH

[3745] Susan M. Sogard, Kenneth W. Able, and Michael P. Fahay. Early life history of the tautog Tautoga onitis in the Mid-Atlantic Bight. Fishery Bulletin, 90(3):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/early-
life-history-tautog-tautoga-onitis-mid-atlantic-bight;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1992/903/
sogard.pdf.

Feeney:1992:PYL

[3746] Richard F. Feeney. Post-yolksac larval development of two southern California sculpins, Clinocottus analyis and Orthonopias triacis (Pisces: Cottidae). Fishery Bulletin, 90(3):??, 1992. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/post-yolksac-larval-development-
two-southern-california-sculpins-clinocottus-analyis-and;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1992/
903/feeney.pdf.

Stein:1992:FHA

[3747] David L. Stein, Brian N. Tissot, Mark A. Hixon, and William Barss. Fish-habitat associations on a deep reef at the edge of the Oregon continental shelf. Fishery Bulletin, 90(3):540–551, 1992. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/fish-habitat-associations-deep-
reef-edge-oregon-continental-shelf;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1992/903/stein.pdf.

Graves:1992:GAW

[3748] John E. Graves, Jan R. McDowell, and M. Lisa Jones. A genetic analysis of weakfish Cynoscion regalis stock structure along the mid-Atlantic coast. Fishery Bulletin, 90(3):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/genetic-analysis-weakfish-cynoscion-regalis-stock-
structure-along-mid-atlantic-coast;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1992/903/graves.pdf.

Anonymous:1992:MAS

[3749] Anonymous. Management advice from a simple dynamic pool model. Fishery Bulletin, 90(3):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/management-advice-simple-dynamic-
pool-model;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1992/903/thompson1.pdf.

Moffitt:1992:AEB

[3750] Robert B. Moffitt and Frank A. Parrish. An assessment of the exploitable biomass of Heterocarpus laevigatus in the main Hawaiian Islands. Part 2: Observations from a submersible. Fishery Bulletin, 90(3): ??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/assessment-
exploitable-biomass-heterocarpus-laevigatus-main-hawaiian-islands-
part-2;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1992/903/moffitt.pdf.

Thompson:1992:BAM

[3751] Grant G. Thompson. A Bayesian approach to management advice when stock-recruitment parameters are uncertain. Fishery Bulletin, 90(3):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
bayesian-approach-management-advice-when-stock-recruitment-
parameters-are-uncertain;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1992/903/thompson2.pdf.

Polovina:1992:VSL

[3752] Jeffrey J. Polovina and Gary T. Mitchum. Variability in spiny lobster Panulirus marginatus recruitment and sea level in the Northwestern Hawaiian Islands. Fishery Bulletin, 90(3):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/variability-spiny-lobster-panulirus-marginatus-
recruitment-and-sea-level-northwestern;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1992/903/polovina.
pdf.

Walker:1992:EPS

[3753] H. J. Walker, Jr. and Keith W. Radford. Eastern Pacific species of the genus Umbrina (Pisces: Sciaenidae) with a description of a new species. Fishery Bulletin, 90(3):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/eastern-pacific-species-genus-umbrina-pisces-sciaenidae-
description-new-species;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1992/903/walker.pdf.

Ralston:1992:AEB

[3754] Stephen Ralston and Darryl T. Tagami. An assessment of the exploitable biomass of Heterocarpus laevigatus in the main Hawaiian Islands. Part 1: Trapping surveys, depletion experiment, and length structure. Fishery Bulletin, 90(3):??, 1992. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/assessment-exploitable-biomass-
heterocarpus-laevigatus-main-hawaiian-islands-part-1;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1992/903/
ralston.pdf.

Weinrich:1992:BRH

[3755] Mason T. Weinrich, Richard H. Lambertson, Cynthia R. Belt, Mark R. Schilling, Heidi J. Iken, and Stephen E. Syrjala. Behavioral reactions of humpback whales Megaptera novaeangliae to biopsy procedures. Fishery Bulletin, 90(3):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/behavioral-reactions-humpback-whales-megaptera-
novaeangliae-biopsy-procedures;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1992/903/weinrich.pdf.

Baker:1992:PCI

[3756] C. Scott Baker, Janice M. Straley, and Anjanette Perry. Population characteristics of individually identified humpback whales in southeastern Alaska: Summer and fall 1986. Fishery Bulletin, 90(3):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/population-
characteristics-individually-identified-humpback-whales-southeastern-
alaska;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1992/903/baker.pdf.

Reilly:1992:IVO

[3757] Carol A. Reilly, Tina Wyllie Echeverria, and Stephen Ralston. Interannual variation and overlap in the diets of pelagic juvenile rockfish (genus: Sebastes) off central California. Fishery Bulletin, 90(3):505–515, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/interannual-
variation-and-overlap-diets-pelagic-juvenile-rockfish-genus-
sebastes-central;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1992/903/reilly.pdf.

Wigley:1992:STD

[3758] Susan E. Wigley and Fredric M. Serchuk. Spatial and temporal distribution of juvenile Atlantic cod Gadus morhua in the Georges Bank– Southern New England region. Fishery Bulletin, 90(3):599–606, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/spatial-and-temporal-
distribution-juvenile-atlantic-cod-gadus-morhua-georges-bank-
southern;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1992/903/wigley.pdf.

Allen:1992:ADS

[3759] Larry G. Allen and Michael P. Franklin. Abundance, distribution, and settlement of young-of-the-year white seabass Atractoscion nobilis in the Southern California Bight, 1988–89. Fishery Bulletin, 90 (4):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/abundance-
distribution-and-settlement-young-year-white-seabass-atractoscion-
nobilis;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1992/904/allen.pdf.

Leffler:1992:AVG

[3760] Deborah L. Leffler and Richard F. Shaw. Age validation, growth, and mortality of larval Atlantic bumper (Carangidae: Chloroscombrus chrysurus) in the northern Gulf of Mexico. Fishery Bulletin, 90(4):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
age-validation-growth-and-mortality-larval-atlantic-bumper-
carangidae-chloroscombrus;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1992/904/leffler.pdf.

Skillman:1992:LTC

[3761] Robert A. Skillman and George H. Balazs. Leatherback turtle captured by ingestion of squid bait on swordfish longline. Fishery Bulletin, 90(4):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
leatherback-turtle-captured-ingestion-squid-bait-swordfish-
longline;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1992/904/skillman.pdf.

Boggs:1992:DCT

[3762] Christofer H. Boggs. Depth, capture time, and hooked longevity of longline-caught pelagic fish: Timing bites of fish with chips. Fishery Bulletin, 90(4):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/depth-
capture-time-and-hooked-longevity-longline-caught-pelagic-fish-
timing-bites-fish;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1992/904/boggs.pdf.

Matthews:1992:TSH

[3763] Kathleen R. Matthews. A telemetric study of the home ranges and homing routes of lingcod Ophiodon elongatus on shallow rocky reefs off Vancouver Island, British Columbia. Fishery Bulletin, 90(4):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/telemetric-
study-home-ranges-and-homing-routes-lingcod-ophiodon-elongatus-
shallow-rocky;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1992/904/matthews.pdf.

Somerton:1992:PDP

[3764] David A. Somerton and Bert S. Kikkawa. Population dynamics of pelagic armorhead Pseudopentaceros wheeleri on Southeast Hancock Seamount. Fishery Bulletin, 90(4):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/population-dynamics-pelagic-armorhead-pseudopentaceros-
wheeleri-southeast-hancock-seamount;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1992/904/somerton.pdf.

Boothroyd:1992:RAL

[3765] Frank A. Boothroyd and Gerald P. Ennis. Reproduction in American lobsters Homarus americanus transplanted northward to St. Michael’s Bay, Labrador. Fishery Bulletin, 90(4):??, 1992. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/reproduction-american-lobsters-
homarus-americanus-transplanted-northward-st-michaels-bay;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1992/
904/boothroyd.pdf.

Miller:1992:IBD

[3766] W. George Miller. An investigation of bottlenose dolphin Tursiops truncatus deaths in East Matagorda Bay, Texas, January 1990. Fishery Bulletin, 90(4):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/investigation-bottlenose-dolphin-tursiops-
truncatus-deaths-east-matagorda-bay-texas-january;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1992/904/
miller.pdf.

Taylor:1992:RBS

[3767] Ronald G. Taylor and Michael D. Murphy. Reproductive biology of the swordfish Xiphias gladius in the Straits of Florida and adjacent waters. Fishery Bulletin, 90(4):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/reproductive-biology-swordfish-xiphias-gladius-
straits-florida-and-adjacent-waters;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1992/904/taylor.pdf.

Ditty:1992:LDD

[3768] James G. Ditty and Richard F. Shaw. Larval development, distribution, and ecology of cobia Rachycentron canadum (family: Rachycentridae) in the northern Gulf of Mexico. Fishery Bulletin, 90(4):668–677, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/larval-development-
distribution-and-ecology-cobia-rachycentron-canadum-family;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1992/904/ditty.pdf.

Parrish:1992:EGF

[3769] Frank A. Parrish and Thomas K. Kasama. Evaluation of ghost fishing in the Hawaiian lobster fishery. Fishery Bulletin, 90(4):720–725, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/evaluation-
ghost-fishing-hawaiian-lobster-fishery;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1992/904/parrish.pdf.

Utter:1992:GIP

[3770] Fred M. Utter, Robin S. Waples, and David J. Teel. Genetic isolation of previously indistinguishable Chinook salmon populations of the Snake and Klamath Rivers: Limitations of negative data. Fishery Bulletin, 90(4):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/genetic-isolation-previously-indistinguishable-
chinook-salmon-populations-snake-and-klamath;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1992/904/utter.
pdf.

Edwards:1992:EAT

[3771] Elizabeth F. Edwards. Energetics of associated tunas and dolphins in the eastern tropical Pacific Ocean: a basis for the bond. Fishery Bulletin, 90 (4):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/energetics-
associated-tunas-and-dolphins-eastern-tropical-pacific-ocean-
basis-bond;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1992/904/edwards.pdf.

Pearcy:1992:MAT

[3772] William G. Pearcy. Movements of acoustically-tagged yellowtail rockfish Sebastes flavidus on Heceta Bank, Oregon. Fishery Bulletin, 90(4):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
movements-acoustically-tagged-yellowtail-rockfish-sebastes-
flavidus-heceta-bank-oregon;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1992/904/pearcy.pdf.

Eggleston:1992:ASS

[3773] David B. Eggleston, Romuald N. Lipcius, and David L. Miller. Artificial shelters and survival of juvenile Caribbean spiny lobster Panulirus argus: Spatial, habitat, and lobster size effects. Fishery Bulletin, 90(4):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/artificial-shelters-and-survival-juvenile-
caribbean-spiny-lobster-panulirus-argus-spatial;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1992/904/
eggleston.pdf.

Restrepo:1992:SSA

[3774] Victor R. Restrepo, John M. Hoenig, Joseph E. Powers, James W. Baird, and Stephen C. Turner. A simple simulation approach to risk and cost analysis, with applications to swordfish and cod fisheries. Fishery Bulletin, 90(4):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/simple-simulation-approach-risk-
and-cost-analysis-applications-swordfish-and-cod-fisheries;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1992/904/restrepo.pdf.

Graves:1992:SSB

[3775] John E. Graves, Jan R. McDowell, Ana M. Beardsley, and Daniel R. Scoles. Stock structure of the bluefish Poatomus saltatrix along the mid-Atlantic coast. Fishery Bulletin, 90(4):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/stock-structure-bluefish-
poatomus-saltatrix-along-mid-atlantic-coast;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1992/904/graves.
pdf.

Schilling:1992:BII

[3776] Mark R. Schilling, Irene Selpt, Mason T. Weinrich, Steven E. Frohock, Anne E. Kuhlberg, and Phillip J. Clapham. Behavior of individuallyidentified sei whales Balaenoptera borealis during an episodic influx into the southern Gulf of Maine in 1986. Fishery Bulletin, 90(4): ??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/behavior-
individually-identified-sei-whales-balaenoptera-borealis-during-
episodic-influx;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1992/904/schilling.pdf.

Jensen:1992:DPL

[3777] Gregory C. Jensen, Helle B. Anderson, and David A. Armstrong. Differentiating paralithodes larvae using telson spines: a tail of two species. Fishery Bulletin, 90(4):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/differentiating-paralithodes-larvae-using-
telson-spines-tail-two-species;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1992/904/jensen.pdf.

Secor:1992:AOM

[3778] David H. Secor. Application of otolith microchemistry analysis to investigate anadromy in Chesapeake Bay striped bass Morone saxatilis. Fishery Bulletin, 90(4):??, 1992. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
application-otolith-microchemistry-analysis-investigate-anadromy-
chesapeake-bay-striped;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1992/904/secor.pdf.

DeMartini:1993:CSL

[3779] Edward E. DeMartini, Denise M. Ellis, and Victor A. Honda. Comparisons of spiny lobster Panulirus marginatus fecundity, egg size, and spawning frequency before and after exploitation. Fishery Bulletin, 91(1):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/comparisons-spiny-lobster-panulirus-marginatus-
fecundity-egg-size-and-spawning-frequency;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1993/911/demartini.
pdf.

Lowerre-Barbieri:1993:NMO

[3780] Susan K. Lowerre-Barbieri and Luiz R. Barbieri. A new method of oocyte separation and preservation for fish reproduction studies. Fishery Bulletin, 91(1):165–170, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/new-method-oocyte-separation-and-
preservation-fish-reproduction-studies;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1993/911/lowerre-barbieri.
pdf.

Flores-Coto:1993:STG

[3781] César Flores-Coto and Stanley M. Warlen. Spawning time, growth, and recruitment of larval spot Leiostomus xanthurus into a North Carolina estuary. Fishery Bulletin, 91(1):8–22, 1993. CODEN FSYBAY.

ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/spawning-time-growth-and-recruitment-
larval-spot-leiostomus-xanthurus-north-carolina;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1993/911/
flores-coto.pdf.

Marks:1993:OSD

[3782] Rick E. Marks and David O. Conover. Ontogenetic shift in the diet of young-of-year bluefish Pomatomus saltatrix during the oceanic phase of the early life history. Fishery Bulletin, 91(1):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/ontogenetic-
shift-diet-young-year-bluefish-pomatomus-saltatrix-during-oceanic-
phase-early;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1993/911/marks.pdf.

Francis:1993:UOS

[3783] Malcolm P. Francis, Maryann W. Williams, Andrea C. Pryce, Susan Pollard, and Stephen G. Scott. Uncoupling of otolith and somatic growth in Pagrus auratus (Sparidae). Fishery Bulletin, 91(1):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/uncoupling-otolith-and-somatic-
growth-pagrus-auratus-sparidae;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1993/911/francis.pdf.

Olson:1993:VHM

[3784] Alan F. Olson and Thomas P. Quinn. Vertical and horizontal movements of adult Chinook salmon Oncorhynchus tshawytscha in the Columbia River estuary. Fishery Bulletin, 91(1):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/vertical-and-horizontal-movements-adult-chinook-
salmon-oncorhynchus-tshawytscha-columbia;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1993/911/olson.pdf.

Fuiman:1993:TES

[3785] Lee A. Fuiman and David R. Ottey. Temperature effects on spontaneous behavior of larval and juvenile red drum Sciaenops ocellatus, and implications for foraging. Fishery Bulletin, 91(1):??, 1993. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/temperature-effects-spontaneous-
behavior-larval-and-juvenile-red-drum-sciaenops-ocellatus;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1993/
911/fuiman.pdf.

Powell:1993:CEL

[3786] Allyn B. Powell. A comparison of early-life-history traits in Atlantic menhaden Brevoortia tyrannus and Gulf menhaden B. patronus. Fishery Bulletin, 91(1):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/comparison-early-life-history-traits-
atlantic-menhaden-brevoortia-tyrannus-and-gulf;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1993/911/powell.
pdf.

Haldorson:1993:APM

[3787] Lewis Haldorson, Marc Pritchett, David Sterrit, and John Watts. Abundance patterns of marine fish larvae during spring in a southeastern Alaskan bay. Fishery Bulletin, 91(1):36–44, 1993. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/abundance-patterns-marine-fish-
larvae-during-spring-southeastern-alaskan-bay;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1993/911/haldorson.
pdf.

Renaud:1993:LST

[3788] Maurice Renaud, Gregg Gitschlag, Edward Klima, Arvind Shah, Dennis Koi, and James Nance. Loss of shrimp by turtle excluder devices (TEDs) in coastal waters of the United States, North Carolina to Texas: March 1988–August 1990. Fishery Bulletin, 91(1):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/loss-shrimp-turtle-
excluder-devices-teds-coastal-waters-united-states-north-carolina-
texas;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1993/911/renaud.pdf.

Jearld:1993:EGB

[3789] Ambrose Jearld, Jr., Sherry L. Sass, and Melinda F. Davis. Early growth, behavior, and otolith development of the winter flounder Pleuronectes americanus. Fishery Bulletin, 91(1):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/early-growth-behavior-and-otolith-development-
winter-flounder-pleuronectes-americanus;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1993/911/jearld.pdf.

Singhas:1993:OEP

[3790] Lynda S. Singhas, Terry L. West, and William G. Ambrose, Jr. Occurrence of Echeneibothrium (Platyhelminthes, Cestoda) in the calico scallop Argopecten gibbus from North Carolina. Fishery Bulletin, 91(1):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/occurrence-echeneibothrium-platyhelminthes-cestoda-
calico-scallop-argopecten-gibbus-north;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1993/911/singhas.pdf.

Jordan:1993:LDT

[3791] Alan R. Jordan and Barry D. Bruce. Larval development of three roughy species complexes (Pisces: Trachichthyidae) from southern Australian waters, with comments on the occurrence of orange roughy Hoplostethus atlanticus. Fishery Bulletin, 91(1):??, 1993. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/larval-development-three-roughy-
species-complexes-pisces-trachichthyidae-southern;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1993/911/
jordan.pdf.

Stillwell:1993:FHS

[3792] Charles E. Stillwell and Nancy E. Kohler. Food habits of the sandbar shark Carcharhinus plumbeus off the U.S. northeast coast, with estimates of daily ration. Fishery Bulletin, 91(1):??, 1993. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/food-habits-sandbar-
shark-carcharhinus-plumbeus-us-northeast-coast-estimates-daily-
ration;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1993/911/stillwell.pdf.

Chittenden:1993:STO

[3793] Mark E. Chittenden, Jr., Luiz R. Barbieri, and Cynthia M. Jones. Spatial and temporal occurrence of Spanish mackerel Scomberomorus maculatus in Chesapeake Bay. Fishery Bulletin, 91(1):??, 1993. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/spatial-and-temporal-occurrence-
spanish-mackerel-scomberomorus-maculatus-chesapeake-bay;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1993/911/
chittenden.pdf.

Krieger:1993:DAR

[3794] Kenneth J. Krieger. Distribution and abundance of rockfish determined from a submersible and by bottom trawling. Fishery Bulletin, 91(1):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
distribution-and-abundance-rockfish-determined-submersible-
and-bottom-trawling;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1993/911/krieger.pdf.

Forward:1993:SIA

[3795] Richard B. Forward, Leslie M. McKelvey, William F. Hettler, and Donald E. Hoss. Swimbladder inflation of the Atlantic menhaden, Brevoortia tyrannus. Fishery Bulletin, 91(2):??, 1993. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/swimbladder-inflation-atlantic-
menhaden-brevoortia-tyrannus;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1993/912/forward.pdf.

Prager:1993:DCC

[3796] Michael H. Prager and Alec D. MacCall. Detection of contaminant and climate effects on spawning success of three pelagic fish stocks off southern California: Northern anchovy Engraulis mordax, Pacific sardine Sardinops sagax, and chub mackerel Scomber japonicus. Fishery Bulletin, 91(2):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/detection-contaminant-and-climate-effects-spawning-
success-three-pelagic-fish-stocks;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1993/912/prager.pdf.

Anganuzzi:1993:CTD

[3797] Alejandro A. Anganuzzi. A comparison of tests for detecting trends in abundance indices of dolphins. Fishery Bulletin, 91(2):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/comparison-
tests-detecting-trends-abundance-indices-dolphins;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1993/912/
anganuzzi.pdf.

Haynes:1993:SZL

[3798] Evan B. Haynes. Stage-I zoeae of laboratory-hatched Lopholithodes mandtii (Decapoda, Anomura, Lithodidae). Fishery Bulletin, 91(2): ??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/stage-i-zoeae-
laboratory-hatched-lopholithodes-mandtii-decapoda-anomura-lithodidae;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1993/912/haynes.pdf.

Shepherd:1993:LBA

[3799] Gary R. Shepherd and Josef S. Idoine. Length-based analyses of yield and spawning biomass per recruit for black sea bass Centropristis striata, a protogynous hermaphrodite. Fishery Bulletin, 91(2):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/length-based-
analyses-yield-and-spawning-biomass-recruit-black-sea-bass-
centropristis;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1993/912/shepherd.pdf.

Bartoo:1993:EDG

[3800] Norman Bartoo and David Holts. Estimated drift gillnet selectivity for albacore Thunnus alalunga. Fishery Bulletin, 91(2):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic).  URL https://spo.nmfs.noaa.gov/content/estimated-
drift-gillnet-selectivity-albacore-thunnus-alalunga;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1993/912/
bartoo.pdf.

Jackson:1993:SVR

[3801] George D. Jackson. Seasonal variation in reproductive investment in the tropical loliginid squid Loligo chinensis and the small tropical sepioid Idiosepius pygmaeus. Fishery Bulletin, 91(2):??, 1993. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/seasonal-variation-reproductive-
investment-tropical-loliginid-squid-loligo-chinensis-and;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1993/
912/jackson.pdf.

Smith:1993:LDP

[3802] Wallace G. Smith and Wallace W. Morse. Larval distribution patterns: Early signals for the collapse/recovery of Atlantic herring Clupea harengus in the Georges Bank area. Fishery Bulletin, 91(2):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/larval-distribution-
patterns-early-signals-collapserecovery-atlantic-herring-clupea;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1993/912/smith.pdf.

Choat:1993:CTN

[3803] John H. Choat, Peter J. Doherty, Brigid A. Kerrigan, and Jeffrey M. Leis. A comparison of towed nets, purse seine, and light-aggregation devices for sampling larvae and pelagic juveniles of coral reef fishes. Fishery Bulletin, 91(2):195–209, 1993. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
comparison-towed-nets-purse-seine-and-light-aggregation-devices-
sampling-larvae-and-pelagic;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1993/912/choat.pdf.

Kimura:1993:EBG

[3804] Daniel K. Kimura, Allen M. Shimada, and Sandra A. Lowe. Estimating von Bertalanffy growth parameters of sablefish Anoplopoma fimbria and Pacific cod Gadus macrocephalus using tag-recapture data. Fishery Bulletin, 91(2):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
estimating-von-bertalanffy-growth-parameters-sablefish-anoplopoma-
fimbria-and-pacific-cod;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1993/912/kimura.pdf.

Squire:1993:RAP

[3805] James L. Squire, Jr. Relative abundance of pelagic resources utilized by the California purse-seine fishery: Results of an airborne monitoring program, 1962–90. Fishery Bulletin, 91(2):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/relative-abundance-pelagic-resources-
utilized-california-purse-seine-fishery-results;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1993/912/
squire.pdf.

Comyns:1993:IDU

[3806] Bruce H. Comyns and George C. Grant. Identification and distribution of Urophycis and Phycis (Pisces, Gadidae) larvae and pelagic juveniles in the U.S. Middle Atlantic Bight. Fishery Bulletin, 91(2):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/identification-and-distribution-
urophycis-and-phycis-pisces-gadidae-larvae-and-pelagic;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1993/912/
comyns.pdf.

Lai:1993:OSD

[3807] Han-Lin Lai. Optimal sampling design for using the age–length key to estimate age composition of a fish population. Fishery Bulletin, 91 (2):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/optimal-
sampling-design-using-age-length-key-estimate-age-composition-

fish-population;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1993/912/lai.pdf.

Tringali:1993:DHF

[3808] Michael D. Tringali and Raymond R. Wilson, Jr. Differences in haplotype frequencies of mtDNA of the Spanish sardine Sardinella aurita between specimens from the eastern Gulf of Mexico and southern Brazil. Fishery Bulletin, 91(2):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
differences-haplotype-frequencies-mtdna-spanish-sardine-sardinella-
aurita-between-specimens;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1993/912/tringali.pdf.

Davis:1993:MRS

[3809] Tim L. O. Davis and Grant J. West. Maturation, reproductive seasonality, fecundity, and spawning frequency in Lutjanus vittus (Quoy and Gaimard) from the North West Shelf of Australia. Fishery Bulletin, 91(2):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
maturation-reproductive-seasonality-fecundity-and-spawning-
frequency-lutjanus-vittus-quoy;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1993/912/davis.pdf.

Lough:1993:VDP

[3810] R. Gregory Lough and David C. Potter. Vertical distribution patterns and diel migrations of larval and juvenile haddock Melanogrammus aeglefinus and Atlantic cod Gadus morhua on Georges Bank. Fishery Bulletin, 91(2):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/vertical-distribution-patterns-and-diel-migrations-
larval-and-juvenile-haddock;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1993/912/lough.pdf.

Ebert:1993:IDP

[3811] Thomas A. Ebert, Stephen C. Schroeter, and John D. Dixon. Inferring demographic processes from size-frequency distributions: Effect of pulsed recruitment on simple models. Fishery Bulletin, 91(2):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/inferring-demographic-processes-
size-frequency-distributions-effect-pulsed-recruitment;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1993/912/
ebert.pdf.

McBride:1993:RBP

[3812] Richard S. McBride, Jeffrey L. Ross, and David O. Conover. Recruitment of bluefish Pomatomus saltatrix to estuaries of the U.S. South Atlantic Bight. Fishery Bulletin, 91(2):??, 1993. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/recruitment-bluefish-pomatomus-
saltatrix-estuaries-us-south-atlantic-bight;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1993/912/mcbride.
pdf.

Fitzhugh:1993:ODF

[3813] Gary R. Fitzhugh, Bruce A. Thompson, and Theron G. Snider III. Ovarian development, fecundity, and spawning frequency of black drum Pogonias cromis in Louisiana. Fishery Bulletin, 91 (2):244–253, 1993. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
ovarian-development-fecundity-and-spawning-frequency-black-
drum-pogonias-cromis-louisiana;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1993/912/fitzhugh.pdf.

OConnell:1993:HSD

[3814] Victoria M. O’Connell and David W. Carlile. Habitat-specific density of adult yelloweye rockfish Sebastes ruberrimus in the eastern Gulf of Alaska. Fishery Bulletin, 91(2):??, 1993. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/habitat-specific-density-adult-
yelloweye-rockfish-sebastes-ruberrimus-eastern-gulf-alaska;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1993/912/oconnell.pdf.

Matlock:1993:DVB

[3815] Gary C. Matlock, Robert L. Colura, and Lawrence W. McEachron. Direct validation of black drum (Pogonias cromis) ages determined from scales. Fishery Bulletin, 91(3):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/direct-validation-black-drum-pogonias-cromis-ages-
determined-scales;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1993/913/matlock.pdf.

Yamanaka:1993:MTL

[3816] K. Lynne Yamanaka and Laura J. Richards. Movements of transplanted lingcod, Ophiodon elongatus, determined by ultrasonic telemetry. Fishery Bulletin, 91(3):582–587, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/movements-transplanted-lingcod-ophiodon-elongatus-determined-
ultrasonic-telemetry;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1993/913/yamanaka.pdf.

Busby:1993:DLE

[3817] Morgan S. Busby and David A. Ambrose. Development of larval and early juvenile pygmy poacher, Odontopyxis trispinosa, and blacktip poacher, Xeneretmus latifrons (Scorpaeniformes: Agonidae). Fishery Bulletin, 91 (3):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/development-
larval-and-early-juvenile-pygmy-poacher-odontopyxis-trispinosa-
and-blacktip;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1993/913/busby.pdf.

Millar:1993:IBH

[3818] Russell B. Millar. Incorporation of between-haul variation using bootstrapping and nonparametric estimation of selection curves. Fishery Bulletin, 91(3):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
incorporation-between-haul-variation-using-bootstrapping-and-
nonparametric-estimation;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1993/913/millar.pdf.

DeMartini:1993:MPF

[3819] Edward E. DeMartini. Modeling the potential of fishery reserves for managing Pacific coral reef fishes. Fishery Bulletin, 91(3):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/modeling-potential-
fishery-reserves-managing-pacific-coral-reef-fishes;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1993/913/
demartini.pdf.

Olesiuk:1993:APC

[3820] Peter F. Olesiuk. Annual prey consumption by harbor seals (Phoca vitulina) in the Strait of Georgia, British Columbia. Fishery Bulletin, 91(3):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
annual-prey-consumption-harbor-seals-phoca-vitulina-strait-
georgia-british-columbia;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1993/913/olesiuk.pdf.

Edwards:1993:AEP

[3821] Elizabeth F. Edwards. Allometry of energetics parameters in spotted dolphin (Stenella attenuata) from the eastern tropical Pacific Ocean. Fishery Bulletin, 91(3):??, 1993. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/allometry-energetics-parameters-
spotted-dolphin-stenella-attenuata-eastern-tropical-pacific;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1993/913/edwards.pdf.

Pearson:1993:BOC

[3822] Donald E. Pearson, David A. Douglas, and Bill Barss. Biological observations from the Cobb Seamount rockfish fishery. Fishery Bulletin, 91(3): ??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/biological-
observations-cobb-seamount-rockfish-fishery;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1993/913/pearson.
pdf.

Gaskin:1993:HPP

[3823] David E. Gaskin, Satoru Yamamoto, and Akito Kawamura. Harbor porpoise, Phocoena phocoena (L.), in the coastal waters of northern Japan. Fishery Bulletin, 91(3):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/harbor-porpoise-phocoena-phocoena-l-coastal-waters-
northern-japan;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1993/913/gaskin.pdf.

Sandt:1993:OSH

[3824] Veronique J. Sandt and Allan W. Stoner. Ontogenetic shift in habitat by early juvenile queen conch, Strombus gigas, patterns and potential mechanisms. Fishery Bulletin, 91(3):516–525, 1993. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/ontogenetic-shift-habitat-early-
juvenile-queen-conch-strombus-gigas-patterns-and-potential;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1993/913/sandt.pdf.

Hayes:1993:SME

[3825] Daniel B. Hayes. A statistical method for evaluating differences between age–length keys with application to Georges Bank haddock, Melanogrammus aeglefinus. Fishery Bulletin, 91(3):??, 1993. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/statistical-method-
evaluating-differences-between-age-length-keys-application-
georges-bank;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1993/913/hayes.pdf.

Schmidt:1993:AGM

[3826] David J. Schmidt, Mark R. Collins, and David M. Wyanski. Age, growth, maturity, and spawning of Spanish mackerel, Scomberomorus maculatus (Mitchill), from the Atlantic Coast of the southeastern United States. Fishery Bulletin, 91(3):??, 1993. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/age-growth-maturity-and-spawning-
spanish-mackerel-scomberomorus-maculatus-mitchill-atlantic;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1993/913/schmidt.pdf.

Humphreys:1993:UMG

[3827] Robert L. Humphreys, Jr., Mark A. Crossler, and Craig M. Rowland. Use of a monogenean gill parasite and feasibility of condition indices for identifying new recruits to a seamount population of armorhead Pseudopentaceros wheeleri (Pentacerotidae). Fishery Bulletin, 91(3):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/use-monogenean-gill-
parasite-and-feasibility-condition-indices-identifying-new-
recruits;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1993/913/humphreys.pdf.

Terceiro:1993:CAM

[3828] Mark Terceiro and Jeffrey L. Ross. A comparison of alternative methods for the estimation of age from length data for Atlantic coast bluefish (Pomatomus saltatrix). Fishery Bulletin, 91(3):534–549, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/comparison-alternative-methods-
estimation-age-length-data-atlantic-coast-bluefish-pomatomus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1993/913/terceiro.pdf.

Kane:1993:VZB

[3829] Joseph Kane. Variability of zooplankton biomass and dominant species abundance on Georges Bank, 1977–1986. Fishery Bulletin, 91(3):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/variability-
zooplankton-biomass-and-dominant-species-abundance-georges-

bank-1977-1986;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1993/913/kane.pdf.

Witting:1993:EBS

[3830] David A. Witting and Kenneth W. Able. Effects of body size on probability of predation for juvenile summer and winter flounder based on laboratory experiements. Fishery Bulletin, 91(3): ??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/effects-
body-size-probability-predation-juvenile-summer-and-winter-
flounder-based;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1993/913/witting.pdf.

Brill:1993:EEE

[3831] Richard W. Brill and David B. Holts. Effects of entanglement and escape from high-seas driftnets on rates of natural mortality of North Pacific albacore, Thunnus alalunga. Fishery Bulletin, 91(4):798–803, 1993. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/effects-entanglement-and-
escape-high-seas-driftnets-rates-natural-mortality-north-pacific;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1993/914/brill.pdf.

Lovrich:1993:RBF

[3832] Gustavo A. Lovrich and Julio H. Vinuesa. Reproductive biology of the false southern king crab (Paralomis granulosa, Lithodidae) in the Beagle Channel, Argentina. Fishery Bulletin, 91(4):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/reproductive-biology-
false-southern-king-crab-paralomis-granulosa-lithodidae-beagle;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1993/914/lovrich.pdf.

VanWaerebeek:1993:GVS

[3833] Koen Van Waerebeek. Geographic variation and sexual dimorphism in the skull of the dusky dolphin, Lagenorhynchus obscurus (Gray, 1828). Fishery Bulletin, 91(4):??, 1993. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/geographic-variation-and-sexual-
dimorphism-skull-dusky-dolphin-lagenorhynchus-obscurus-gray;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1993/914/waerebeek.pdf.

Chang:1993:AFR

[3834] Sukwoo Chang. Analysis of fishery resources: potential risk from sewage sludge dumping at the deepwater dumpsite off New Jersey. Fishery Bulletin, 91(4):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/analysis-fishery-resources-potential-risk-
sewage-sludge-dumping-deepwater-dumpsite-new;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1993/914/chang.
pdf.

Lozano-Alvarez:1993:OSV

[3835] Enrique Lozano-Alvarez, Patricia Briones-Fourzan, and Fernando Negrete-Soto. Occurrence and seasonal variations of spiny lobsters, Panulirus argus (Latreille), on the shelf outside Bahia de la Ascension, Mexico. Fishery Bulletin, 91(4):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/occurrence-and-seasonal-variations-spiny-lobsters-panulirus-
argus-latreille-shelf-outside;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1993/914/lozano-alvarez.pdf.

Weisberg:1993:SDA

[3836] Stephen B. Weisberg and William H. Burton. Spring distribution and abundance of ichthyoplankton in the tidal Delaware River. Fishery Bulletin, 91(4):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
spring-distribution-and-abundance-ichthyoplankton-tidal-delaware-
river;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1993/914/weisberg.pdf.

Chivers:1993:CAS

[3837] Susan J. Chivers and Al C. Myrick, Jr. Comparison of age at sexual maturity and other reproductive parameters for two stocks of spotted dolphin, Stenella attenuata. Fishery Bulletin, 91(4):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/comparison-age-sexual-maturity-
and-other-reproductive-parameters-two-stocks-spotted-dolphin;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1993/914/chivers.pdf.

Sampson:1993:ACS

[3838] David B. Sampson. The assumption of constant selectivity and the stock assessment for widow rockfish, Sebastes entomelas. Fishery Bulletin, 91(4):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
assumption-constant-selectivity-and-stock-assessment-widow-
rockfish-sebastes-entomelas;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1993/914/sampson.pdf.

Witherell:1993:GMW

[3839] David B. Witherell and Jay Burnett. Growth and maturation of winter flounder, Pleuronectes americanus, in Massachusetts. Fishery Bulletin, 91(4):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/growth-and-maturation-winter-flounder-pleuronectes-
americanus-massachusetts;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1993/914/witherell.pdf.

Chow:1993:PRA

[3840] Seinen Chow, Elizabeth Clarke, and Patrick J. Walsh. PCR-RFLP analysis on thirteen western Atlantic snappers (subfamily Lutjaninae): a simple method for species and stock identification. Fishery Bulletin, 91 (4):619–627, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/pcr-
rflp-analysis-thirteen-western-atlantic-snappers-subfamily-
lutjaninae-simple-method;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1993/914/chow.pdf.

Scoles:1993:GAP

[3841] Daniel R. Scoles and John E. Graves. Genetic analysis of the population structure of yellowfin tuna, Thunnus albacares, from the Pacific Ocean. Fishery Bulletin, 91(4):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/genetic-analysis-population-structure-yellowfin-tuna-
thunnus-albacares-pacific-ocean;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1993/914/scoles.pdf.

Edwards:1993:EDG

[3842] Elizabeth F. Edwards and Christina Perrin. Effects of dolphin group type, percent coverage, and fleet size on estimates of annual dolphin mortality derived from 1987 U.S. tuna vessel observer data. Fishery Bulletin, 91(4):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
effects-dolphin-group-type-percent-coverage-and-fleet-size-
estimates-annual-dolphin;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1993/914/edwards.pdf.

Sedberry:1993:PBF

[3843] George R. Sedberry and Nicole Cuellar. Planktonic and benthic feeding by the reef-associated vermilion snapper, Rhomboplites aurorubens (Teleostei, Lutjanidae). Fishery Bulletin, 91(4):699–709, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/planktonic-and-benthic-feeding-
reef-associated-vermilion-snapper-rhomboplites-aurorubens;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1993/
914/sedberry.pdf.

Gilpatrick:1993:MPE

[3844] James W. Gilpatrick, Jr. Method and precision in estimation of dolphin school size with vertical aerial photography. Fishery Bulletin, 91(4):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/method-and-precision-
estimation-dolphin-school-size-vertical-aerial-photography;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1993/914/gilpatrick.pdf.

Temte:1993:LVB

[3845] Jonathan L. Temte. Latitudinal variation in the birth timing of captive California sea lions and other captive North Pacific pinnipeds. Fishery Bulletin, 91(4):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
latitudinal-variation-birth-timing-captive-california-sea-lions-
and-other-captive-north;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1993/914/temte.pdf.

Hannah:1993:WPS

[3846] Robert W. Hannah and Neil T. Richmond. Weight of pink shrimp, Pandalus jordani, after commercial harvest and handling. Fishery Bulletin, 91(4):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/weight-pink-shrimp-pandalus-jordani-after-
commercial-harvest-and-handling;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1993/914/hannah.pdf.

Thompson:1993:VSD

[3847] Grant G. Thompson. Variations on a simple dynamic pool model. Fishery Bulletin, 91(4):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/variations-simple-dynamic-pool-model;   https:/

/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1993/914/
thompson.pdf.

Alados:1993:DII

[3848] Consuela L. Alados, Juan Escos, and John M. Emlen. Developmental instability as an indicator of environmental stress in the Pacific hake (Merluccius productus). Fishery Bulletin, 91(4):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/developmental-
instability-indicator-environmental-stress-pacific-hake-merluccius-
productus;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1993/914/alados.pdf.

Labelle:1993:DAG

[3849] Marc Labelle, John Hampton, Kevin Bailey, Talbot Murray, David A. Fournier, and John R. Sibert. Determination of age and growth of South Pacific albacore (Thunnus alalunga) using three methodologies. Fishery Bulletin, 91(4):??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
determination-age-and-growth-south-pacific-albacore-thunnus-
alalunga-using-three;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1993/914/labelle.pdf.

Wade:1993:EHP

[3850] Paul R. Wade. Estimation of historical population size of the eastern spinner dolphin (Stenella longirostris orientallis). Fishery Bulletin, 91(4): ??, 1993. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/estimation-
historical-population-size-eastern-spinner-dolphin-stenella-
longirostris;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1993/914/wade.pdf.

Incze:1994:DAC

[3851] Lewis S. Incze and Terri Ainaire. Distribution and abundance of copepod nauplii and other small (40–300 μ m) zooplankton during spring in Shelikof Strait, Alaska. Fishery Bulletin, 92(1):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/distribution-
and-abundance-copepod-nauplii-and-other-small-40-300-%C2%B5m-
zooplankton-during;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1994/921/incze.pdf.

Stoner:1994:QCS

[3852] Allan W. Stoner and Kirsten C. Schwarte. Queen conch, Strombus gigas, reproductive stocks in the central Bahamas: distribution and probable sources. Fishery Bulletin, 92(1):??, 1994. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/queen-conch-strombus-
gigas-reproductive-stocks-central-bahamas-distribution-and-
probable;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1994/921/stoner.pdf.

Barbieri:1994:AGM

[3853] Luiz R. Barbieri, Mark E. Chittenden, Jr., and Cynthia M. Jones. Age, growth, and mortality of Atlantic croaker, Micropogonias undulatus, in the Chesapeake Bay region, with a discussion of apparent geographic changes in population dynamics. Fishery Bulletin, 92(1):??, 1994. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/age-growth-and-mortality-
atlantic-croaker-micropogonias-undulatus-chesapeake-bay-region;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1994/921/barbieri.pdf.

Johnson:1994:EDS

[3854] Allyn G. Johnson, William A. Fable, Jr., Churchill B. Grimes, Lee Trent, and Javier Vasconcelos Perez. Evidence for distinct stocks of king mackerel, Scomberomorus cavalla, in the Gulf of Mexico. Fishery Bulletin, 92(1):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/evidence-
distinct-stocks-king-mackerel-scomberomorus-cavalla-gulf-mexico;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1994/921/johnson.pdf.

Wilber:1994:IAR

[3855] Dara H. Wilber. The influence of Apalachicola River flows on blue crab, Callinectes sapidus, in north Florida. Fishery Bulletin, 92(1): ??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/influence-
apalachicola-river-flows-blue-crab-callinectes-sapidus-north-
florida;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1994/921/wilber.pdf.

Bigelow:1994:AGO

[3856] Keith A. Bigelow. Age and growth of the oceanic squid Onychoteuthis borealijaponica in the North Pacific. Fishery Bulletin, 92(1):13–25, 1994. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
age-and-growth-oceanic-squid-onychoteuthis-borealijaponica-
north-pacific;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1994/921/bigelow.pdf.

Lenarz:1994:EWL

[3857] William H. Lenarz. Estimation of weight–length relationships from group measurements. Fishery Bulletin, 92(1):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/estimation-weight-length-
relationships-group-measurements;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1994/921/lenarz.pdf.

Burke:1994:DKR

[3858] Vincent J. Burke, Stephen J. Morreale, and Edward A. Standora. Diet of the Kemp’s ridley sea turtle, Lepidochelys kempii, in New York waters. Fishery Bulletin, 92(1):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/diet-kemps-ridley-sea-turtle-lepidochelys-
kempii-new-york-waters;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1994/921/burke.pdf.

Milton:1994:RBE

[3859] David A. Milton, Stephen J. M. Blaber, and Nicholas J. F. Rawlinson. Reproductive biology and egg production of three species of Clupeidae from Kiribati, tropical central Pacific. Fishery Bulletin, 92(1):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/reproductive-biology-and-
egg-production-three-species-clupeidae-kiribati-tropical-central;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1994/921/milton.pdf.

Ditty:1994:LDT

[3860] James G. Ditty and Richard F. Shaw. Larval development of tripletail, Lobotes surinamensis (Pisces: Lobotidae), and their spatial and temporal distribution in the northern Gulf of Mexico. Fishery Bulletin, 92(1):33–45, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/larval-development-tripletail-lobotes-surinamensis-
pisces-lobotidae-and-their-spatial-and;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1994/921/ditty.pdf.

Perryman:1994:ESS

[3861] Wayne L. Perryman and Morgan S. Lynn. Examination of stock and school structure of striped dolphin (Stenella coeruleoalba) in the eastern Pacific from aerial photogrammetry. Fishery Bulletin, 92(1):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/examination-
stock-and-school-structure-striped-dolphin-stenella-coeruleoalba-
eastern;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1994/921/perryman.pdf.

Fargo:1994:OMH

[3862] Jeff Fargo and Albert V. Tyler. Oocyte maturation in Hecate Strait English sole (Pleuronectes vetulus). Fishery Bulletin, 92(1): ??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/oocyte-
maturation-hecate-strait-english-sole-pleuronectes-vetulus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1994/921/fargo.pdf.

Polovina:1994:SLR

[3863] Jeffrey J. Polovina and Gary T. Mitchum. Spiny lobster recruitment and sea level: results of a 1990 forecast. Fishery Bulletin, 92(1):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/spiny-lobster-
recruitment-and-sea-level-results-1990-forecast;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1994/921/
polovina.pdf.

Ferreira:1994:AVE

[3864] Beatrice Padovani Ferreira and Garry Russ. Age validation and estimation of growth rate of the coral trout, Plectropomus leopardus, (Lacépède 1802) from Lizard Island, Northern Great Barrier Reef. Fishery Bulletin, 92(1):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/age-validation-and-estimation-growth-rate-
coral-trout-plectropomus-leopardus-lacepede-1802;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1994/921/
ferreira.pdf.

Punsly:1994:PTC

[3865] Richard G. Punsly, Patrick K. Tomlinson, and Ashley J. Mullen. Potential tuna catches in the eastern Pacific Ocean from schools not associated with dolphins. Fishery Bulletin, 92(1):??, 1994. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/potential-tuna-catches-eastern-
pacific-ocean-schools-not-associated-dolphins;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1994/921/punsly.
pdf.

Gold:1994:GDR

[3866] John R. Gold and Linda R. Richardson. Genetic distinctness of red drum (Sciaenaps acellatus) from Mosquito Lagoon, east-central Florida. Fishery Bulletin, 92(1):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
genetic-distinctness-red-drum-sciaenaps-acellatus-mosquito-
lagoon-east-central-florida;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1994/921/gold.pdf.

Sinclair:1994:PSN

[3867] Elizabeth Sinclair, Thomas Loughlin, and William Pearcy. Prey selection by northern fur seals (Callorhinus ursinus) in the eastern Bering Sea. Fishery Bulletin, 92(1):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/prey-selection-northern-fur-seals-callorhinus-ursinus-
eastern-bering-sea;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1994/921/sinclair.pdf.

Aurioles-Gamboa:1994:AMS

[3868] David Aurioles-Gamboa, Maria Isabel Castro-Gonzalez, and Ricardo Perez-Flores. Annual mass strandings of pelagic red crabs, Pleuroncodes planipes (Crustacea: Anomura: Galatheidae), in Bahja Magdalena, Baja California Sur, Mexico. Fishery Bulletin, 92(2):??, 1994. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/annual-mass-strandings-pelagic-
red-crabs-pleuroncodes-planipes-crustacea-anomura;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1994/922/
aurioles-gamboa.pdf.

Fiedler:1994:IVD

[3869] Paul C. Fiedler and Stephen B. Reilly. Interannual variability of dolphin habitats in the eastern tropical Pacific. II: Effects on abundances estimated from tuna vessel sightings, 1975– 1990. Fishery Bulletin, 92(2):451–463, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/interannual-variability-dolphin-

habitats-eastern-tropical-pacific-ii-effects-abundances;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1994/922/
fiedler.pdf.

Taylor:1994:SCC

[3870] David M. Taylor, Paul G. O’Keefe, and Charles Fitzpatrick. A snow crab, Chionoecetes opilio (Decapoda, Majidae), fishery collapse in Newfoundland. Fishery Bulletin, 92(2):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/snow-crab-chionoecetes-opilio-decapoda-majidae-
fishery-collapse-newfoundland;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1994/922/taylor.pdf.

Blood:1994:EDW

[3871] Deborah M. Blood, Ann C. Matarese, and Mary M. Yoklavich. Embryonic development of walleye pollock, Theragra chalcogramma, from Shelikof Strait, Gulf of Alaska. Fishery Bulletin, 92(2):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic).  URL https://spo.nmfs.noaa.gov/content/embryonic-
development-walleye-pollock-theragra-chalcogramma-shelikof-
strait-gulf-alaska;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1994/922/blood.pdf.

Kastelle:1994:UPR

[3872] Craig R. Kastelle, Daniel K. Kimura, Ahmad E. Nevissi, and Donald R. Gunderson. Using Pb-21 O/Ra-226 disequilibria for sablefish, Anoplopoma fimbria, age validation. Fishery Bulletin, 92(2):292–301, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/using-pb-21-ora-226-
disequilibria-sablefish-anoplopoma-fimbria-age-validation;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1994/
922/kastelle.pdf.

Warlen:1994:STR

[3873] Stanley M. Warlen. Spawning time and recruitment dynamics of larval Atlantic menhaden, Brevoortia tyrannus, into a North Carolina estuary. Fishery Bulletin, 92(2):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/spawning-time-and-recruitment-dynamics-larval-atlantic-
menhaden-brevoortia-tyrannus-north;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1994/922/warlen.pdf.

Brodeur:1994:DVD

[3874] Richard D. Brodeur and William C. Rugen. Diel vertical distribution of ichthyoplankton in the northern Gulf of Alaska. Fishery Bulletin, 92(2):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/diel-
vertical-distribution-ichthyoplankton-northern-gulf-alaska;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1994/922/brodeur.pdf.

Moltschaniwskyj:1994:DAT

[3875] Natalie A. Moltschaniwskyj and Peter Doherty. Distribution and abundance of two juvenile tropical Photololigo species (Cephalopoda: Loliginidae) in the central Great Barrier Reef Lagoon. Fishery Bulletin, 92(2): ??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/distribution-
and-abundance-two-juvenile-tropical-photololigo-species-cephalopoda;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1994/922/moltschaniwskyj.pdf.

Clark:1994:CPO

[3876] Malcolm R. Clark and Dianne M. Tracey. Changes in a population of orange roughy, Hoplostethus atlanticus, with commercial exploitation on the Challenger Plateau, New Zealand. Fishery Bulletin, 92(2):236–253, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/changes-population-
orange-roughy-hoplostethus-atlanticus-commercial-exploitation-
challenger;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1994/922/clark.pdf.

Murie:1994:ODA

[3877] Debra J. Murie, Daryl C. Parkyn, Bruce G. Clapp, and Geoffrey G. Krause. Observations on the distribution and activities of rockfish, Sebastes spp., in Saanich Inlet, British Columbia, from the Pisces IV submersible. Fishery Bulletin, 92(2):313–323, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/observations-
distribution-and-activities-rockfish-sebastes-spp-saanich-inlet-
british;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1994/922/murie.pdf.

Daniel:1994:MGI

[3878] Louis B. Daniel III and John E. Graves. Morphometric and genetic identification of eggs of spring-spawning sciaenids in lower Chesapeake Bay. Fishery Bulletin, 92(2):254–261, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/morphometric-and-genetic-
identification-eggs-spring-spawning-sciaenids-lower-chesapeake-
bay;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1994/922/daniel.pdf.

Perrin:1994:RGV

[3879] William F. Perrin, Gary D. Schnell, Daniel J. Hough, James W. Gilpatrick, Jr., and Jerry V. Kashiwada. Reexamination of geographic variation in cranial morphology of the pantropical spotted dolphin, Stenella attenuata, in the eastern Pacific. Fishery Bulletin, 92(2):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/reexamination-
geographic-variation-cranial-morphology-pantropical-spotted-
dolphin-stenella;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1994/922/perrin.pdf.

Ditty:1994:RDA

[3880] James G. Ditty, Richard F. Shaw, and Joseph S. Cope. A redescription of Atlantic spadefish larvae, Chaetodipterus faber (family: Ephippidae), and their distribution, abundance, and seasonal occurrence in the northern Gulf of Mexico. Fishery Bulletin, 92 (2):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/re-
description-atlantic-spadefish-larvae-chaetodipterus-faber-
family-ephippidae-and-their;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1994/922/ditty1.pdf.

Powell:1994:MOP

[3881] Eric N. Powell, John M. Klinck, Eileen E. Hofmann, and Sammy M. Ray. Modeling oyster populations. IV: Rates of mortality, population crashes, and management. Fishery Bulletin, 92(2):347–373, 1994. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/modeling-oyster-populations-iv-
rates-mortality-population-crashes-and-management;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1994/922/
powell.pdf.

Ditty:1994:LDD

[3882] James G. Ditty, Richard F. Shaw, Churchill B. Grimes, and Joseph S. Cope. Larval development, distribution, and abundance of common dolphin, Coryphaena hippurus, and pompano dolphin, C. equiselis (family: Coryphaenidae), in the northern Gulf of Mexico. Fishery Bulletin, 92(2):275–291, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/larval-development-distribution-
and-abundance-common-dolphin-coryphaena-hippurus-and;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1994/922/
ditty2.pdf.

Reilly:1994:IVD

[3883] Stephen B. Reilly and Paul C. Fiedler. Interannual variability of dolphin habitats in the eastern tropical Pacific. I: Research vessel surveys, 1986–1990. Fishery Bulletin, 92(2):434–450, 1994. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/interannual-variability-dolphin-
habitats-eastern-tropical-pacific-i-research-vessel-surveys;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1994/922/reilly.pdf.

Ennevor:1994:MMC

[3884] Bridget C. Ennevor. Mass marking Coho salmon, Oncorhynchus kisutch, fry with lanthanum and cerium. Fishery Bulletin, 92 (2):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/mass-
marking-coho-salmon-oncorhynchus-kisutch-fry-lanthanum-and-
cerium;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1994/922/ennevor.pdf.

Stoner:1994:EOJ

[3885] Allan W. Stoner and Megan Davis. Experimental outplanting of juvenile queen conch, Strombus gigas: comparison of wild and hatcheryreared stocks. Fishery Bulletin, 92(2):390–411, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/experimental-outplanting-
juvenile-queen-conch-strombus-gigas-comparison-wild-and-hatchery;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1994/922/stoner.pdf.

DiGiacomo:1994:RBC

[3886] Edgardo E. Di Giácomo and Maria Raquel Perier. Reproductive biology of the cockfish, Callorhynchus callorhynchus (Holocephali: Callorhynchidae), in Patagonian waters (Argentina). Fishery Bulletin, 92(3):531–539, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/reproductive-biology-
cockfish-callorhynchus-callorhynchus-holocephali-callorhynchidae;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1994/923/digiacomo.pdf.

Sarda:1994:STS

[3887] Francisco Sarda, Joan E. Cartes, and Walter Norbis. Spatiotemporal structure of the deep-water shrimp Aristeus antennatus (Decapoda: Aristeidae) population in the western Mediterranean. Fishery Bulletin, 92(3):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/spatio-temporal-structure-deep-water-
shrimp-aristeus-antennatus-decapoda-aristeidae;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1994/923/sarda.
pdf.

Forward:1994:SDA

[3888] Richard B. Forward, Jr., William F. Hettler, and Donald E. Hoss. Swimbladder deflation in the Atlantic menhaden, Brevoortia tyrannus. Fishery Bulletin, 92(3):641–646, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/swimbladder-deflation-atlantic-menhaden-brevoortia-
tyrannus;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1994/923/forward.pdf.

Stanley:1994:MTA

[3889] Richard D. Stanley, Bruce M. Leaman, Lewis Haldorson, and Victoria M. O’Connell. Movements of tagged adult yellowtail rockfish, Sebastes flavidus, off the west coast of North America. Fishery Bulletin, 92(3): 655–663, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/movements-
tagged-adult-yellowtail-rockfish-sebastes-flavidus-west-coast-
north-america;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1994/923/stanley.pdf.

Gilman:1994:EBN

[3890] Sharon L. Gilman. An energy budget for northern sand lance, Ammodytes dubius, on Georges Bank, 1977–1986. Fishery Bulletin, 92(3):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
energy-budget-northern-sand-lance-ammodytes-dubius-georges-
bank-1977-1986;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1994/923/gilman.pdf.

Steimle:1994:PPR

[3891] Frank W. Steimle, Dorothy Jeffress, Stephen A. Fromm, Robert N. Reid, Joseph J. Vitaliano, and Ann Frame. Predator–prey relationships of winter flounder, Pleuronectes americanus, in the New York Bight apex. Fishery Bulletin, 92(3):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
predator-prey-relationships-winter-flounder-pleuronectes-americanus-
new-york-bight-apex;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1994/923/steimle.pdf.

Castillo:1994:EIR

[3892] Gonzalo C. Castillo, Hiram W. Li, and James T. Golden. Environmentally induced recruitment variation in petrale sole, Eopsetta jordani. Fishery Bulletin, 92(3):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/environmentally-induced-recruitment-variation-petrale-
sole-eopsetta-jordani;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1994/923/castillo.pdf.

Kendall:1994:VDE

[3893] Arthur W. Kendall, Jr., Lewis S. Incze, Peter B. Ortner, Shailer R. Cummings, and Patricia K. Brown. The vertical distribution of eggs and larvae of walleye pollock, Theragra chalcogramma, in Shelikof Strait, Gulf of Alaska. Fishery Bulletin, 92(3):??, 1994. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/vertical-distribution-eggs-and-
larvae-walleye-pollock-theragra-chalcogramma-shelikof-strait;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1994/923/kendall.pdf.

Weinberg:1994:RAM

[3894] Kenneth L. Weinberg. Rockfish assemblages of the middle shelf and upper slope off Oregon and Washington. Fishery Bulletin, 92(3):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/rockfish-assemblages-
middle-shelf-and-upper-slope-oregon-and-washington;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1994/923/
weinberg.pdf.

Creaser:1994:DFA

[3895] Edwin P. Creaser and Herbert C. Perkins. The distribution, food, and age of juvenile bluefish, Pomatomus saltatrix, in Maine. Fishery Bulletin, 92(3):494–508, 1994. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
distribution-food-and-age-juvenile-bluefish-pomatomus-saltatrix-
maine;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1994/923/creaser.pdf.

Lowerre-Barbieri:1994:CVO

[3896] Susan K. Lowerre-Barbieri, Mark E. Chittenden, Jr., and Cynthia M. Jones. A comparison of a validated otolith method to age weakfish, Cynoscion regalis, with the traditional scale method. Fishery Bulletin, 92(3):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
comparison-validated-otolith-method-age-weakfish-cynoscion-
regalis-traditional-scale-method;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1994/923/lowerre-barbieri.pdf.

Xiao:1994:SGM

[3897] Yongshun Xiao and David C. Ramm. A simple generalized model of allometry, with examples of length and weight relationships for 14 species of groundfish. Fishery Bulletin, 92(3):664–670, 1994. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/simple-generalized-model-
allometry-examples-length-and-weight-relationships-14-species;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1994/923/xiao.pdf.

David:1994:DBS

[3898] Andrew W. David, J. Jeffery Isley, and Churchill B. Grimes. Differences between the sagitta, lapillus, and asteriscus in estimating age and growth in juvenile red drum, Sciaenaps ocellatus. Fishery Bulletin, 92(3): ??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/differences-
between-sagitta-lapillus-and-asteriscus-estimating-age-and-
growth-juvenile-red;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1994/923/david.pdf.

McKinnon:1994:FHD

[3899] Jeff McKinnon. Feeding habits of the dusky dolphin, Lagenorhynchus obscurus, in the coastal waters of central Peru. Fishery Bulletin, 92(3):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/feeding-habits-dusky-
dolphin-lagenorhynchus-obscurus-coastal-waters-central-peru;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1994/923/mckinnon.pdf.

Dee:1994:RTE

[3900] Anderson J. Dee and James D. Parish. Reproductive and trophic ecology of the soldierfish myripristis amaena in tropical fisheries. Fishery Bulletin, 92(3):516–530, 1994. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
reproductive-and-trophic-ecology-soldierfish-myripristis-amaena-
tropical-fisheries;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1994/923/dee.pdf.

Muter:1994:DAG

[3901] Franz-Josef Muter and Brenda L. Norcross. Distribution, abundance, and growth of larval walleye pollock, Theragra chalcogramma, in an Alaskan fjord. Fishery Bulletin, 92(3):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/distribution-abundance-and-growth-larval-walleye-
pollock-theragra-chalcogramma-alaskan;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1994/923/muter.pdf.

Delancey:1994:RLT

[3902] Lawrence B. Delancey, James E. Jenkins, and J. David Whitaker. Results of long-term, seasonal sampling for Penaeus postlarvae at Breach Inlet, South Carolina. Fishery Bulletin, 92(3):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/results-long-term-seasonal-sampling-penaeus-
postlarvae-breach-inlet-south-carolina;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1994/923/delancey.
pdf.

Norcross:1994:IVR

[3903] Brenda L. Norcross and David M. Wyanski. Interannual variation in the recruitment pattern and abundance of age-0 summer flounder Paralichthys dentatus, in Virginia estuaries. Fishery Bulletin, 92(3):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/interannual-variation-recruitment-pattern-
and-abundance-age-o-summer-flounder-paralichthys;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1994/923/
norcross.pdf.

Kendall:1994:SEL

[3904] Arthur W. Kendall, Jr. and Ann C. Matarese. Status of early life history descriptions of marine teleosts. Fishery Bulletin, 92(4):

??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/status-
early-life-history-descriptions-marine-teleosts;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1994/924/
kendall.pdf.

Shimada:1994:SMP

[3905] Allen M. Shimada and Daniel K. Kimura. Seasonal movements of Pacific cod, Gadus macrocephalus, in the eastern Bering Sea and adjacent waters based on tag-recapture data. Fishery Bulletin, 92(4):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/seasonal-movements-pacific-cod-
gadus-macrocephalus-eastern-bering-sea-and-adjacent-waters;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1994/924/shimada.pdf.

Koslow:1994:CEA

[3906] Julian A. Koslow, Karl Aiken, Stephanie Auil, and Antoinette Clementson. Catch and effort analysis of the reef fisheries of Jamaica and Belize. Fishery Bulletin, 92(4):737–747, 1994. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/catch-and-effort-analysis-reef-
fisheries-jamaica-and-belize;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1994/924/koslow.pdf.

Thresher:1994:EEP

[3907] Ronald E. Thresher, Craig H. Proctor, John S. Gunn, and Ian R. Harrowfield. An evaluation of electron-probe microanalysis of otoliths for stock delineation and identification of nursery areas in a southern temperate groundfish, Nemadactylus macropterus (Cheilodactylidae). Fishery Bulletin, 92(4):814–840, 1994. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
evaluation-electron-probe-microanalysis-otoliths-stock-delineation-
and-identification;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1994/924/thresher.pdf.

Laurenson:1994:CBG

[3908] Laurie J. B. Laurenson, Ian C. Potter, and Norm G. Hall. Comparisons between generalized growth curves for two estuarine populations of the eel tailed catfish, Cnidoglanis macrocephalus. Fishery Bulletin, 92(4):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/comparisons-between-
generalized-growth-curves-two-estuarine-populations-eel-tailed-

catfish;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1994/924/laurenson.pdf.

Wilson:1994:RBR

[3909] Charles A. Wilson and David L. Nieland. Reproductive biology of red drum, Sciaenaps acellatus, from the neritic waters of the northern Gulf of Mexico. Fishery Bulletin, 92(4):841–850, 1994. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/reproductive-biology-red-
drum-sciaenaps-acellatus-neritic-waters-northern-gulf-mexico;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1994/924/wilson1.pdf.

Barbieri:1994:MSO

[3910] Luiz R. Barbieri, Mark E. Chittenden, Jr., and Susan K. Lowerre-Barbieri. Maturity, spawning, and ovarian cycle of Atlantic croaker, Micropogonias undulatus, in the Chesapeake Bay and adjacent coastal waters. Fishery Bulletin, 92(4):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/maturity-spawning-and-ovarian-cycle-atlantic-croaker-
micropogonias-undulatus-chesapeake-bay;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1994/924/barbieri.
pdf.

Lehodey:1994:MDA

[3911] Patrick Lehodey, Paul Marchal, and Rene Grandperrin. Modelling the distribution of alfonsino, Beryx splendens, over the seamounts of New Caledonia. Fishery Bulletin, 92(4):??, 1994. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/modelling-distribution-alfonsino-
beryx-splendens-over-seamounts-new-caledonia;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1994/924/lehodey.
pdf.

Anonymous:1994:BPC

[3912] Anonymous. Biology and population characteristics of Squalus mitsukurii from a seamount in the central North Pacific Ocean. Fishery Bulletin, 92(4):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/biology-
and-population-characteristics-squalus-mitsukurii-seamount-
central-north-pacific;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1994/924/wilson2.pdf.

Biesiot:1994:BHC

[3913] Patricia M. Biesiot, Robert E. Caylor, and James S. Franks. Biochemical and histological changes during ovarian development of cobia, Rachycentron canadum, from the northern Gulf of Mexico. Fishery Bulletin, 92(4): ??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/biochemical-
and-histological-changes-during-ovarian-development-cobia-rachycentron-
canadum;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1994/924/biesiot.pdf.

McCabe:1994:SEL

[3914] George T. McCabe, Jr. and Charles A. Tracy. Spawning and early life history of white sturgeon, Acipenser transmontanus, in the lower Columbia River. Fishery Bulletin, 92(4):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/spawning-and-early-life-history-white-sturgeon-
acipenser-transmontanus-lower-columbia-river;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1994/924/mccabe.
pdf.

Erzini:1994:MDC

[3915] Karim Erzini and Margarida Castro. Measures of dispersion as constraints for length-frequency analysis. Fishery Bulletin, 92(4): ??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/measures-
dispersion-constraints-length-frequency-analysis;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1994/924/
erzini.pdf.

Mullin:1994:CUC

[3916] Keith D. Mullin, Wayne Hoggard, Carol L. Roden, Ren R. Lohoefener, Carolyn M. Rogers, and Brian Taggart. Cetaceans on the upper continental slope in the north-central Gulf of Mexico. Fishery Bulletin, 92 (4):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/cetaceans-
upper-continental-slope-north-central-gulf-mexico;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1994/924/
mullin.pdf.

Giorgi:1994:MCJ

[3917] Albert E. Giorgi, David R. Miller, and Benjamin P. Sandford. Migratory characteristics of juvenile ocean-type Chinook salmon, Oncorhynchus tshawytscha, in John Day Reservoir on the Columbia River. Fishery Bulletin, 92(4):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
migratory-characteristics-juvenile-ocean-type-chinook-salmon-
oncorhynchus-tshawytscha-john;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1994/924/giorgi.pdf.

Parker:1994:VTM

[3918] Richard O. Parker, Alexander J. Chester, and Russell S. Nelson. A video transect method for estimating reef fish abundance, composition, and habitat utilization at Gray’s Reef National Marine Sanctuary, Georgia. Fishery Bulletin, 92(4):??, 1994. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/video-transect-method-estimating-
reef-fish-abundance-composition-and-habitat-utilization;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1994/924/
parker.pdf.

Jacobson:1994:BBA

[3919] Larry D. Jacobson, Nancy C. H. Lo, and J. Thomas Barnes. A biomassbased assessment model for northern anchovy, Engraulis mordax. Fishery Bulletin, 92(4):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
biomass-based-assessment-model-northern-anchovy-engraulis-mordax;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1994/924/jacobson.pdf.

Raynie:1994:CLP

[3920] Richard C. Raynie and Richard F. Shaw. A comparison of larval and postlarval Gulf menhaden, Brevoortia patronus, growth rates between an offshore spawning ground and an estuarine nursery. Fishery Bulletin, 92(4):??, 1994. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/comparison-larval-and-postlarval-gulf-menhaden-
brevoortia-patronus-growth-rates-between;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1994/924/raynie.pdf.

Forney:1995:ACC

[3921] Karin A. Forney, Jay Barlow, and James V. Carretta. The abundance of cetaceans in California waters. Part II: Aerial surveys in winter and spring of 1991 and 1992. Fishery Bulletin, 93(1):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/abundance-cetaceans-california-

waters-part-ii-aerial-surveys-winter-and-spring-1991-and;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1995/
931/forney.pdf.

Schmid:1995:MTP

[3922] Jeffrey R. Schmid. Marine turtle populations on the east-central coast of Florida: results of tagging studies at Cape Canaveral, Florida, 1986–1991. Fishery Bulletin, 93(1):??, 1995. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/marine-turtle-populations-east-
central-coast-florida-results-tagging-studies-cape-canaveral;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1995/931/schmid.pdf.

Barlow:1995:ACC

[3923] Jay Barlow. The abundance of cetaceans in California waters. Part I: Ship surveys in summer and fall of 1991. Fishery Bulletin, 93(1):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/abundance-cetaceans-
california-waters-part-i-ship-surveys-summer-and-fall-1991;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1995/931/barlow.pdf.

Kornfield:1995:AHC

[3924] Irv Kornfield, Austin B. Williams, and Robert S. Steneck. Assignment of Homarus capensis (Herbst, 1972), the Cape lobster of South Africa, to the new genus Homarinus (Decapoda: Nephropidae). Fishery Bulletin, 93(1):97–102, 1995. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
assignment-homarus-capensis-herbst-1972-cape-lobster-south-
africa-new-genus-homarinus;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1995/931/kornfield.pdf.

Secor:1995:VOB

[3925] David H. Secor, T. Mark Trice, and Harry T. Hornick. Validation of otolith-based ageing and a comparison of otolith and scale-based ageing in mark-recaptured Chesapeake Bay striped bass, Marone saxatilis. Fishery Bulletin, 93(1):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
validation-otolith-based-ageing-and-comparison-otolith-and-
scale-based-ageing-mark;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1995/931/secor.pdf.

Benetti:1995:GRC

[3926] Daniel D. Benetti, Edwin S. Iversen, and Anthony C. Ostrowski. Growth rates of captive dolphin, Coryphaena hippurus, in Hawaii. Fishery Bulletin, 93(1):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
growth-rates-captive-dolphin-coryphaena-hippurus-hawaii;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1995/931/
benetti.pdf.

Laidig:1995:PUO

[3927] Thomas E. Laidig and Stephen Ralston. The potential use of otolith characters in identifying larval rockfish (Sebastes spp.). Fishery Bulletin, 93 (1):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/potential-use-
otolith-characters-identifying-larval-rockfish-sebastes-spp;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1995/931/laidig.pdf.

Szedlmayer:1995:ESM

[3928] Stephen T. Szedlmayer and Jeffrey C. Howe. An evaluation of six marking methods for age-0 red drum, Sciaenops ocellatus. Fishery Bulletin, 93(1):191–195, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/evaluation-six-marking-methods-age-o-
red-drum-sciaenops-ocellatus;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1995/931/szedlmayer.pdf.

Bruce:1995:LDK

[3929] Barry D. Bruce. Larval development of King George whiting, Sillaginodes punctata, school whiting, Sillago bassensis, and yellow fin whiting, Sillago schomburgkii (Percoidei: Sillaginidae), from South Australian waters. Fishery Bulletin, 93(1):??, 1995. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/larval-development-king-george-
whiting-sillaginodes-punctata-school-whiting-sillago;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1995/931/
bruce.pdf.

Matsuura:1995:CSP

[3930] Yasunobu Matsuura and Roger Hewitt. Changes in the spatial patchiness of Pacific mackerel, Scomberjaponicus, larvae with increasing age and size. Fishery Bulletin, 93(1):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/changes-spatial-patchiness-pacific-mackerel-scomberjaponicus-
larvae-increasing-age-and-size;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1995/931/matsuura.pdf.

Wiley:1995:SMH

[3931] David N. Wiley, Regina A. Asmutis, Thomas D. Pitchford, and Damon P. Gannon. Stranding and mortality of humpback whales, Megaptera novaeangliae, in the mid-Atlantic and southeast United States, 1985–1992. Fishery Bulletin, 93(1):196–205, 1995. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/stranding-and-mortality-humpback-
whales-megaptera-novaeangliae-mid-atlantic-and-southeast;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1995/
931/wiley.pdf.

Carls:1995:RTC

[3932] Mark G. Carls and Charles E. O’Clair. Responses of Tanner crabs, Chionoecetes bairdi, exposed to cold air. Fishery Bulletin, 93(1): ??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/responses-
tanner-crabs-chionoecetes-bairdi-exposed-cold-air;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1995/931/
carls.pdf.

Milton:1995:ATS

[3933] David A. Milton, Steven A. Short, Michael F. O’Neill, and Stephen J. M. Blaber. Ageing of three species of tropical snapper (Lutjanidae) from the Gulf of Carpentaria, Australia, using radiometry and otolith ring counts. Fishery Bulletin, 93(1):103–115, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/ageing-
three-species-tropical-snapper-lutjanidae-gulf-carpentaria-
australia-using;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1995/931/milton.pdf.

Canino:1995:MCT

[3934] Michael F. Canino and Elaine M. Caldarone. Modification and comparison of two fluorometric techniques for determining nucleic acid contents of fish larvae. Fishery Bulletin, 93(1):??, 1995. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14402;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/canino.pdf.

Cortes:1995:DAA

[3935] Enric Cortes. Demographic analysis of the Atlantic sharpnose shark, Rhizoprionodon terraenovae, in the Gulf of Mexico. Fishery Bulletin, 93(1): ??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/demographic-
analysis-atlantic-sharpnose-shark-rhizoprionodon-terraenovae-
gulf-mexico;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1995/931/cortes.pdf.

Natanson:1995:AGE

[3936] Lisa J. Natanson, John G. Casey, and Nancy E. Kohler. Age and growth estimates for the dusky shark, Carcharhinus obscurus, in the western North Atlantic Ocean. Fishery Bulletin, 93(1):??, 1995. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/age-and-growth-estimates-dusky-
shark-carcharhinus-obscurus-western-north-atlantic-ocean;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1995/
931/natanson.pdf.

Ellis:1995:EVC

[3937] Denise M. Ellis and Edward E. DeMartini. Evaluation of a video camera technique for indexing abundances of juvenile pink snapper, Pristipomoides filamentosus, and other Hawaiian insular shelf fishes. Fishery Bulletin, 93(1):67–77, 1995. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
evaluation-video-camera-technique-indexing-abundances-juvenile-
pink-snapper-pristipomoides;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1995/931/ellis.pdf.

Rickey:1995:MSS

[3938] Martha H. Rickey. Maturity, spawning, and seasonal movement of arrowtooth flounder, Atheresthes stomias, off Washington. Fishery Bulletin, 93(1):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/maturity-spawning-and-seasonal-movement-arrowtooth-
flounder-atheresthes-stomias-washington;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1995/931/rickey.pdf.

Finnerty:1995:ECS

[3939] John R. Finnerty and Barbara A. Block. Evolution of cytochrome b in the Scombroidei (Teleostei): molecular insights into billfish (Istiophoridae and Xiphiidae) relationships. Fishery Bulletin, 93(1):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/evolution-cytochrome-
b-scombroidei-teleostei-molecular-insights-billfish-istiophoridae-
and;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1995/931/finnerty.pdf.

Riley:1995:GML

[3940] Cecilia M. Riley, G. Joan Holt, and Connie R. Arnold. Growth and morphology of larval and juvenile captive bred yellowtail snapper, Ocyurus chrysurus. Fishery Bulletin, 93(1):??, 1995. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/growth-and-morphology-larval-and-
juvenile-captive-bred-yellowtail-snapper-ocyurus-chrysurus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1995/931/riley.pdf.

Felley:1995:AHU

[3941] James D. Felley and Michael Vecchione. Assessing habitat use by nekton on the continental slope using archived videotapes from submersibles. Fishery Bulletin, 93(2):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/assessing-habitat-use-nekton-continental-slope-
using-archived-videotapes-submersibles;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1995/932/felley.pdf.

Pepin:1995:ALW

[3942] Pierre Pepin. An analysis of the length–weight relationship of larval fish: limitations of the general allometric model. Fishery Bulletin, 93 (2):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/analysis-
length-weight-relationship-larval-fish-limitations-general-
allometric-model;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1995/932/pepin.pdf.

Fenton:1995:RAB

[3943] Gwen E. Fenton and Stephen A. Short. Radiometric analysis of blue grenadier, Macruronus novaezelandiae, otolith cores. Fishery Bulletin, 93(2):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/radiometric-analysis-blue-grenadier-macruronus-
novaezelandiae-otolith-cores;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1995/932/fenton.pdf.

Restrepo:1995:ASC

[3944] Victor R. Restrepo and Christopher M. Legault. Approximations for solving the catch equation when it involves a ‘plus group’. Fishery Bulletin, 93(2):308–314, 1995. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
approximations-solving-catch-equation-when-it-involves-plus-
group;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1995/932/restrepo.pdf.

Fisher:1995:DMG

[3945] Joseph P. Fisher and William G. Pearcy. Distribution, migration, and growth of juvenile Chinook salmon, Oncorhynchus tshawytscha, off Oregon and Washington. Fishery Bulletin, 93(2):??, 1995. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/distribution-migration-and-growth-
juvenile-chinook-salmon-oncorhynchus-tshawytscha-oregon;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1995/932/
fisher.pdf.

Rutherford:1995:ITC

[3946] Edward S. Rutherford and Edward D. Houde. The influence of temperature on cohort-specific growth, survival, and recruitment of striped bass, Morone saxatilis, larvae in Chesapeake Bay. Fishery Bulletin, 93 (2):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/influence-
temperature-cohort-specific-growth-survival-and-recruitment-
striped-bass-morone;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1995/932/rutherford.pdf.

Ahrenholz:1995:COA

[3947] Dean W. Ahrenholz, Gary R. Fitzhugh, James A. Rice, Stephen W. Nixon, and Wilson C. Pritchard. Confidence of otolith ageing through the juvenile stage for Atlantic menhaden, Brevoortia tyrannus. Fishery Bulletin, 93(2):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
confidence-otolith-ageing-through-juvenile-stage-atlantic-menhaden-
brevoortia-tyrannus;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1995/932/ahrenholz.pdf.

Garduno-Argueta:1995:SDD

[3948] Hector Garduno-Argueta and Jose A. Calderon-Perez. Seasonal depth distribution of the crystal shrimp, Penaeus brevirostris (Crustacea: Decapoda, Penaeidae), and its possible relation to temperature and oxygen concentration off southern Sinaloa, Mexico. Fishery Bulletin, 93(2):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/seasonal-depth-distribution-crystal-shrimp-penaeus-
brevirostris-crustacea-decapoda;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1995/932/garduno-argueta.pdf.

Theilacker:1995:CLW

[3949] Gail H. Theilacker and Stephen M. Porter. Condition of larval walleye pollock, Theragra chalcogramma, in the western Gulf of Alaska assessed with histological and shrinkage indices. Fishery Bulletin, 93(2):333–344, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/condition-larval-walleye-pollock-theragra-
chalcogramma-western-gulf-alaska-assessed;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1995/932/theilacker.
pdf.

Bisbal:1995:DSC

[3950] Gustavo A. Bisbal and David A. Bengtson. Description of the starving condition in summer flounder, Paralichthys dentatus, early life history stages. Fishery Bulletin, 93(2):??, 1995. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/description-starving-condition-
summer-flounder-paralichthys-dentatus-early-life-history;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1995/
932/bisbal.pdf.

Helser:1995:PQM

[3951] Thomas E. Helser and Daniel B. Hayes. Providing quantitative management advice from stock abundance indices based on research surveys. Fishery Bulletin, 93(2):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/providing-quantitative-management-advice-stock-
abundance-indices-based-research-surveys;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1995/932/helser.pdf.

Wade:1995:REI

[3952] Paul R. Wade. Revised estimates of incidental kill of dolphins (Delphinidae) by the purse-seine tuna fishery in the eastern tropical Pacific, 1959–1972. Fishery Bulletin, 93(2):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/revised-estimates-incidental-kill-dolphins-
delphinidae-purse-seine-tuna-fishery-eastern;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1995/932/wade.
pdf.

Carlson:1995:CYA

[3953] H. Richard Carlson. Consistent yearly appearance of age-0 walleye pollock, Theragra chalcogramma, at a coastal site in southeastern Alaska, 1973–1994. Fishery Bulletin, 93(2):??, 1995. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/consistent-yearly-appearance-age-
0-walleye-pollock-theragra-chalcogramma-coastal-site;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1995/932/
carlson.pdf.

Hernandez-Garcia:1995:DSX

[3954] Vicente Hernandez-Garcia. The diet of the swordfish Xiphias gladius Linnaeus, 1758, in the central east Atlantic, with emphasis on the role of cephalopods. Fishery Bulletin, 93(2):??, 1995. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/diet-swordfish-xiphias-gladius-
linnaeus-1758-central-east-atlantic-emphasis-role;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1995/932/
hernandez-garcia.pdf.

Yano:1995:KWO

[3955] Kazunari Yano and Marilyn E. Dahlheim. Killer whale, Orcinus orca, depredation on longline catches of bottomfish in the southeastern Bering Sea and adjacent waters. Fishery Bulletin, 93(2):355– 372, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/killer-
whale-orcinus-orca-depredation-longline-catches-bottomfish-
southeastern-bering-sea;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1995/932/yano.pdf.

Doyle:1995:NIW

[3956] Miriam J. Doyle, William C. Rugen, and Richard D. Brodeur. Neustonic ichthyoplankton in the western Gulf of Alaska during spring. Fishery Bulletin, 93(2):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
neustonic-ichthyoplankton-western-gulf-alaska-during-spring;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1995/932/doyle.pdf.

Kohler:1995:LWR

[3957] Nancy E. Kohler, John G. Casey, and Patricia A. Turner. Length– weight relationships for 13 species of sharks from the western North Atlantic. Fishery Bulletin, 93(2):??, 1995. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/length-weight-relationships-13-
species-sharks-western-north-atlantic;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1995/932/kohler.pdf.

Zeldis:1995:ARV

[3958] John R. Zeldis, Paul J. Grimes, and Jonathan K. V. Ingerson. Ascent rates, vertical distribution, and a thermal history model of development of orange roughy, Hoplostethus atlanticus, eggs in the water column. Fishery Bulletin, 93(2):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/ascent-rates-vertical-distribution-and-thermal-
history-model-development-orange-roughy;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1995/932/zeldis.pdf.

Epperly:1995:ASS

[3959] Sheryan P. Epperly, Joanne Braun, and Alexander J. Chester. Aerial surveys for sea turtles in North Carolina inshore waters. Fishery Bulletin, 93(2):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/aerial-
surveys-sea-turtles-north-carolina-inshore-waters;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1995/932/
epperly.pdf.

Norton:1995:NDR

[3960] Elizabeth C. Norton and R. Bruce MacFarlane. Nutritional dynamics of reproduction in viviparous yellowtail rockfish, Sebastes flavidus. Fishery Bulletin, 93(2):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
nutritional-dynamics-reproduction-viviparous-yellowtail-rockfish-
sebastes-flavidus;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1995/932/norton.pdf.

Adams:1995:PEP

[3961] Peter B. Adams, John L. Butler, Charles H. Baxter, Thomas E. Laidig, Katherine A. Dahlin, and W. Waldo Wakefield. Population estimates of Pacific coast groundfishes from video transects and swept-area trawls. Fishery Bulletin, 93(3):??, 1995. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/population-estimates-
pacific-coast-groundfishes-video-transects-and-swept-area-trawls;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1995/933/adams.pdf.

Renaud:1995:AJG

[3962] Maurice L. Renaud, James A. Carpenter, Jo A. Williams, and Sharon A. Manzella-Tirpak. Activities of juvenile green turtles, Chelonia mydas, at a jettied pass in South Texas. Fishery Bulletin, 93(3):586–593, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/activities-juvenile-
green-turtles-chelonia-mydas-jettied-pass-south-texas;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1995/933/
renaud.pdf.

Armstrong:1995:FHE

[3963] Janet L. Armstrong, David A. Armstrong, and Stephen B. Mathews. Food habits of estuarine staghorn sculpin, Leptocottus armatus, with focus on consumption of juvenile Dungeness crab, Cancer magister. Fishery Bulletin, 93(3):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
food-habits-estuarine-staghorn-sculpin-leptocottus-armatus-
focus-consumption-juvenile;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1995/933/armstrong.pdf.

Sakuma:1995:DPM

[3964] Keith M. Sakuma and Ralph J. Larson. Distribution of pelagic metamorphic-stage sanddabs Citharichthys sordidus and c. stigmaeus within areas of upwelling off central California. Fishery Bulletin, 93(3):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/distribution-pelagic-
metamorphic-stage-sanddabs-citharichthys-sordidus-and-c-stigmaeus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1995/933/sakuma.pdf.

Criales:1995:LDT

[3965] Maria M. Criales and Thomas N. Lee. Larval distribution and transport of penaeoid shrimps during the presence of the Tortugas Gyre in May–June 1991. Fishery Bulletin, 93(3):471–482, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/larval-distribution-
and-transport-penaeoid-shrimps-during-presence-tortugas-gyre-
may-june;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1995/933/criales.pdf.

Tserpes:1995:DAG

[3966] George Tserpes and Nikolaos Tsimenides. Determination of age and growth of swordfish, Xiphias gladius L., 1758, in the eastern Mediterranean using anal-fin spines. Fishery Bulletin, 93(3):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/determination-
age-and-growth-swordfish-xiphias-gladius-l-1758-eastern-mediterranean-
using;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1995/933/tserpes.pdf.

Ferrero:1995:GRC

[3967] Richard C. Ferrero and William A. Walker. Growth and reproduction of the common dolphin, Delphinus delphis Linnaeus, in the offshore waters of the North Pacific Ocean. Fishery Bulletin, 93(3):483– 494, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/growth-
and-reproduction-common-dolphin-delphinus-delphis-linnaeus-
offshore-waters-north;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1995/933/ferrero.pdf.

Wisner:1995:RNW

[3968] Robert L. Wisner and Charmion B. McMillan. Review of new world hagfishes of the genus Myxine (Agnatha, Myxinidae) with descriptions of nine new species. Fishery Bulletin, 93(3):??, 1995. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/review-new-world-hagfishes-genus-
myxine-agnatha-myxinidae-descriptions-nine-new-species;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1995/933/
wisner.pdf.

Fitzhugh:1995:TIP

[3969] Gary R. Fitzhugh and William F. Hettler. Temperature influence on postovulatory follicle degeneration in Atlantic menhaden, Brevoortia tyrannus. Fishery Bulletin, 93(3):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/temperature-influence-postovulatory-follicle-degeneration-
atlantic-menhaden-brevoortia;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1995/933/fitzhugh.pdf.

Worthington:1995:CBG

[3970] Duncan G. Worthington, Neil L. Andrew, and Gary Hamer. Covariation between growth and morphology suggests alternative size limits for the blacklip abalone, Haliotis rubra, in New South Wales, Australia. Fishery Bulletin, 93(3):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/covariation-between-growth-and-morphology-
suggests-alternative-size-limits-blacklip-abalone;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1995/933/
worthington.pdf.

Lutcavage:1995:FDP

[3971] Molly Lutcavage and Scott Kraus. The feasibility of direct photographic assessment of giant bluefin tuna, Thunnus thynnus, in New England waters. Fishery Bulletin, 93(3):??, 1995. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/feasibility-direct-
photographic-assessment-giant-bluefin-tuna-thunnus-thynnus-
new-england;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1995/933/lutcavage.pdf.

Atran:1995:AWF

[3972] Steven M. Atran and Joseph G. Loesch. An analysis of weekly fluctuations in catchability coefficients. Fishery Bulletin, 93(3): ??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14403;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/atran.pdf.

Murphy:1995:OGC

[3973] Margaret A. Murphy. Occurrence and group characteristics of minke whales, Balaenoptera acutorostrata, in Massachusetts Bay and Cape Cod Bay. Fishery Bulletin, 93(3):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/occurrence-and-group-characteristics-
minke-whales-balaenoptera-acutorostrata-massachusetts;   https:

//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1995/933/
murphy.pdf.

Mueller:1995:SSM

[3974] Karl W. Mueller. Size structure of mutton snapper, Lutjanus analis, associated with unexploited artificial patch reefs in the central Bahamas. Fishery Bulletin, 93(3):??, 1995. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14404;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/mueller.pdf.

Able:1995:ELH

[3975] Kenneth W. Able, Michael P. Fahay, and Gary R. Shepherd. Early life history of black sea bass, Centropristis striata, in the Mid-Atlantic Bight and a New Jersey estuary. Fishery Bulletin, 93(3):??, 1995. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/early-life-history-black-
sea-bass-centropristis-striata-mid-atlantic-bight-and-new-jersey;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1995/933/able.pdf.

Powell:1995:ELD

[3976] Allyn B. Powell and Theresa Henley. Egg and larval development of laboratory-reared Gulf flounder, Paralichthys albigutta, and southern flounder, P. lethostigma (Pisces, Paralichthyidae). Fishery Bulletin, 93(3):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
egg-and-larval-development-laboratory-reared-gulf-flounder-
paralichthys-albigutta-and;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1995/933/powell.pdf.

Brodeur:1995:SDE

[3977] Richard D. Brodeur, Morgan S. Busby, and Matthew T. Wilson. Summer distribution of early life stages of walleye pollock, Theragra chalcogramma, and associated species in the western Gulf of Alaska. Fishery Bulletin, 93(4):??, 1995. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/summer-distribution-early-life-
stages-walleye-pollock-theragra-chalcogramma-and-associated;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1995/934/brodeur.pdf.

Mojica:1995:RBA

[3978] Raymond Mojica, Jr., Jonathan M. Shenker, Christopher W. Harnden, and Daniel E. Wagner. Recruitment of bonefish, Albula vulpes, around Lee Stocking Island, Bahamas. Fishery Bulletin, 93(4):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/recruitment-bonefish-
albula-vulpes-around-lee-stocking-island-bahamas;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1995/934/
mojica.pdf.

Crabtree:1995:AGT

[3979] Roy E. Crabtree, Edward C. Cyr, and John M. Dean. Age and growth of tarpon, Megalops atlanticus, from South Florida waters. Fishery Bulletin, 93(4):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/age-
and-growth-tarpon-megalops-atlanticus-south-florida-waters;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1995/934/crabtree.pdf.

Morse:1995:DLH

[3980] Wallace W. Morse and Kenneth W. Able. Distribution and life history of windowpanne, Scophthalmus aquosus, off the northeastern United States. Fishery Bulletin, 93(4):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/distribution-and-life-history-windowpanne-
scophthalmus-aquosus-northeastern-united-states;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1995/934/
morse.pdf.

Ferreira:1995:PSL

[3981] Beatrice P. Ferreira and Garry R. Russ. Population structure of the leopard coralgrouper, Plectropomus leopardus, on fished and unfished reefs off Townsville, Central Great Barrier Reef, Australia. Fishery Bulletin, 93(4):629–642, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/population-structure-leopard-coralgrouper-
plectropomus-leopardus-fished-and-unfished-reefs;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1995/934/
ferreira.pdf.

Porch:1995:TBA

[3982] Clay E. Porch. Trajectory-based approaches to estimating velocity and diffusion from tagging data. Fishery Bulletin, 93(4):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/trajectory-
based-approaches-estimating-velocity-and-diffusion-tagging-
data;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1995/934/porch.pdf.

Forney:1995:DAH

[3983] Karin A. Forney. A decline in the abundance of harbor porpoise, Phocoena phocoena, in nearshore waters off California, 1986–93. Fishery Bulletin, 93(4):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
decline-abundance-harbor-porpoise-phocoena-phocoena-nearshore-
waters-california-1986-93;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1995/934/forney.pdf.

Ralston:1995:DYC

[3984] Stephen Ralston and Daniel F. Howard. On the development of year-class strength and cohort variability in two northern California rockfishes. Fishery Bulletin, 93(4):??, 1995. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/development-year-class-strength-
and-cohort-variability-two-northern-california-rockfishes;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1995/
934/ralston.pdf.

Jackson:1995:SIS

[3985] George D. Jackson. Seasonal influences on statolith growth in the tropical nearshore loliginid squid Loligo chinensis (Cephalopoda: Loliginidae) off Townsville, North Queensland, Australia. Fishery Bulletin, 93(4):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/seasonal-influences-
statolith-growth-tropical-nearshore-loliginid-squid-loligo-
chinensis;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1995/934/jackson.pdf.

Sainte-Marie:1995:FSC

[3986] Bernard Sainte-Marie and Chantal Carriere. Fertilization of the second clutch of eggs of snow crab, Chionoecetes opilio, from females mated once or twice after their molt to maturity. Fishery Bulletin, 93(4):759–764, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/fertilization-second-
clutch-eggs-snow-crab-chionoecetes-opilio-females-mated-once-

or-twice;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1995/934/sainte-marie.pdf.

Lowerre-Barbieri:1995:AGW

[3987] Susan K. Lowerre-Barbieri, Mark E. Chittenden, Jr., and Luiz R. Barbieri. Age and growth of weakfish, Cynoscion regalis, in the Chesapeake Bay region with a discussion of historical changes in maximum size. Fishery Bulletin, 93(4):643–656, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/age-and-growth-weakfish-cynoscion-regalis-
chesapeake-bay-region-discussion-historical;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1995/934/lowerre-
barbieri.pdf.

Sakuma:1995:DLP

[3988] Keith M. Sakuma and Thomas E. Laidig. Description of larval and pelagic juvenile chilipepper, Sebastes goodei (family Scorpaenidae), with an examination of larval growth. Fishery Bulletin, 93(4):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/description-larval-
and-pelagic-juvenile-chilipepper-sebastes-goodei-family-scorpaenidae;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1995/934/sakuma.pdf.

Merrick:1995:CSS

[3989] Richard L. Merrick, Robin Brown, Donald G. Calkins, and Thomas R. Loughlin. A comparison of Stellar sea lion, Eumetopias jubatus, pup masses between rookeries with increasing and decreasing populations. Fishery Bulletin, 93(4):753–758, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/comparison-stellar-sea-lion-eumetopias-jubatus-
pup-masses-between-rookeries-increasing-and;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1995/934/merrick.
pdf.

Williams:1995:RFS

[3990] John G. Williams and Gene M. Matthews. A review of flow and survival relationships for spring and summer Chinook salmon, Oncorhynchus tshawytscha, from the Snake River Basin. Fishery Bulletin, 93(4):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/review-flow-
and-survival-relationships-spring-and-summer-chinook-salmon-

oncorhynchus;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1995/934/williams.pdf.

Mertz:1995:EPR

[3991] Gordon Mertz and Ransom A. Myers. Estimating the predictability of recruitment. Fishery Bulletin, 93(4):??, 1995. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/estimating-predictability-recruitment;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1995/934/
mertz.pdf.

Anonymous:1995:I

[3992] Anonymous. Index. Fishery Bulletin, 93(4):??, 1995. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14405;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/index%20(4)_1.pdf.

Adams:1996:NMB

[3993] Peter B. Adams and Daniel F. Howard. Natural mortality of blue rockfish, Sebastes mystinus, during their first year in nearshore benthic habitats. Fishery Bulletin, 94(1):??, 1996. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/natural-mortality-blue-rockfish-
sebastes-mystinus-during-their-first-year-nearshore-benthic;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1996/941/adams.pdf.

Lowry:1996:CNE

[3994] Mark S. Lowry, Wayne L. Perryman, Morgan S. Lynn, Robin L. Westlake, and Fred Julian. Counts of northern elephant seals, Mirounga angustirostris, from large-format aerial photographs taken at rookeries in southern California during the breeding season. Fishery Bulletin, 94 (1):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/counts-
northern-elephant-seals-mirounga-angustirostris-large-format-
aerial-photographs;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1996/941/lowry.pdf.

Wintner:1996:AGD

[3995] Sabine P. Wintner and Geremy Cliff. Age and growth determination of the blacktip shark, Carcharhincus limbatus, from the east coast of South Africa. Fishery Bulletin, 94(1):135–144, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/age-and-growth-determination-blacktip-shark-
carcharhincus-limbatus-east-coast-south-africa;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1996/941/wintner.
pdf.

Anderl:1996:FAO

[3996] Delsa M. Anderl, Akira Nishimura, and Sandra A. Lowe. Is the first annulus on the otolith of the Atka mackerel, Pleurogrammus monopterygius, missing? Fishery Bulletin, 94(1):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/first-annulus-
otolith-atka-mackerel-pleurogrammus-monopterygius-missing;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1996/
941/anderl.pdf.

McGurk:1996:AMM

[3997] Michael D. McGurk. Allometry of marine mortality of Pacific salmon. Fishery Bulletin, 94(1):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/allometry-marine-mortality-pacific-salmon;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1996/941/mcgurk.pdf.

Xu:1996:AAE

[3998] Xucai Xu and Hussain M. A. Mohammed. An alternative approach to estimating growth parameters from length-frequency data, with application to green tiger prawns. Fishery Bulletin, 94(1):145–155, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/alternative-approach-
estimating-growth-parameters-length-frequency-data-application-
green;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1996/941/xu.pdf.

Arkhipkin:1996:DSS

[3999] Alexander I. Arkhipkin, Vyacheslav A. Bizikov, Vasily V. Krylov, and Kir N. Nesis. Distribution, stock structure, and growth of the squid Berryteuthis magister (Berry, 1913) (Cephalopoda, Gonatidae) during summer and fall in the western Bering Sea. Fishery Bulletin, 94(1):1–30, 1996. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
distribution-stock-structure-and-growth-squid-berryteuthis-

magister-berry-1913-cephalopoda;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1996/941/arkhipkin.pdf.

Oliveira:1996:FVA

[4000] Kenneth Oliveira. Field validation of annular growth rings in the American eel, Anguilla rostrata, using tetracycline-marked otoliths. Fishery Bulletin, 94(1):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/field-validation-annular-growth-rings-american-eel-
anguilla-rostrata-using-tetracycline;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1996/941/oliveira.pdf.

Gadomski:1996:ETD

[4001] Dena M. Gadomski and Steven M. Caddell. Effects of temperature on the development and survival of eggs of four coastal California fishes. Fishery Bulletin, 94(1):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/node/
14407;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/gadomski.pdf.

Bembo:1996:SDA

[4002] David G. Bembo, Gary R. Carvalho, Michael Snow, Nando Cingolani, and Tony J. Pitcher. Stock discrimination among European anchovies, Engraulis encrasicolus, by means of PCR-amplified mitochondrial DNA analysis. Fishery Bulletin, 94(1):??, 1996. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/stock-discrimination-
among-european-anchovies-engraulis-encrasicolus-means-pcr-amplified;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1996/941/bembo.pdf.

Pearson:1996:THZ

[4003] Donald E. Pearson. Timing of hyaline-zone formation as related to sex, location, and year of capture in otoliths of the widow rockfish, Sebastes entomelas. Fishery Bulletin, 94(1):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/timing-hyaline-zone-formation-related-sex-
location-and-year-capture-otoliths-widow-rockfish;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1996/941/
pearson.pdf.

Ikeda:1996:MBC

[4004] Tsutomu Ikeda. Metabolism, body composition, and energy budget of the mesopelagic fish Maurolicus muelleri in the Sea of Japan. Fishery Bulletin, 94(1):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/metabolism-body-composition-and-
energy-budget-mesopelagic-fish-maurolicus-muelleri-sea;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1996/941/
ikeda.pdf.

Ralston:1996:AAE

[4005] Stephen Ralston, Edward B. Brothers, Dale A. Roberts, and Keith M. Sakuma. Accuracy of age estimates for larval Sebastes jordani. Fishery Bulletin, 94(1):89–97, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/accuracy-age-estimates-larval-sebastes-jordani;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1996/941/ralston.pdf.

Jackson:1996:VSA

[4006] George D. Jackson and Julia Yeatman. Variation in size and age at maturity in Photololigo (Mollusca: Cephalopoda) from the northwest shelf of Australia. Fishery Bulletin, 94(1):??, 1996. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/variation-size-and-age-maturity-
photololigo-mollusca-cephalopoda-northwest-shelf-australia;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1996/941/jackson.pdf.

Schaefer:1996:STF

[4007] Kurt M. Schaefer. Spawning time, frequency, and batch fecundity of yellowfin tuna, Thunnus albacares, near Clipperton Atoll in the eastern Pacific Ocean. Fishery Bulletin, 94(1):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/spawning-time-frequency-and-batch-fecundity-
yellowfin-tuna-thunnus-albacares-near;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1996/941/schaefer.pdf.

Karlou-Riga:1996:OAR

[4008] Constantina Karlou-Riga and Panos S. Economidis. Ovarian atretic rates and sexual maturity of European horse mackerel, Trachurus trachurus (L.), in the Saronikos Gulf (Greece). Fishery Bulletin, 94 (1):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/ovarian-
atretic-rates-and-sexual-maturity-eurpoean-horse-mackerel-trachurus-
trachurus-l;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1996/941/karlou-riga.pdf.

Starr:1996:CSS

[4009] Richard M. Starr, David S. Fox, Mark. A. Hixon, Brian N. Tissot, Gary E. Johnson, and William H. Barss. Comparison of submersiblesurvey and hydroacoustic-survey estimates of fish density on a rocky bank. Fishery Bulletin, 94(1):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/comparison-submersible-survey-and-hydroacoustic-
survey-estimates-fish-density-rocky-bank;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1996/941/starr.pdf.

Lang:1996:UOM

[4010] Kathy L. Lang, Frank P. Almeida, George R. Bolz, and Michael P. Fahay. The use of otolith microstructure in resolving issues of first year growth and spawning seasonality of white hake, Urophycis tenuis, in the Gulf of Maine–Georges Bank region. Fishery Bulletin, 94(1):170–175, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/use-otolith-
microstructure-resolving-issues-first-year-growth-and-spawning-
seasonality;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1996/941/lang.pdf.

Stequert:1996:AGY

[4011] Bernard Stéquert, Jacques Panfili, and John Mark Dean. Age and growth of yellowfin tuna, Thunnus albacares, from the western Indian Ocean, based on otolith microstructure. Fishery Bulletin, 94(1): 124–134, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/age-
and-growth-yellowfin-tuna-thunnus-albacares-western-indian-
ocean-based-otolith;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1996/941/stequert.pdf.

Fortier:1996:CPF

[4012] Louis Fortier and Ann Villeneuve. Cannibalism and predation on fish larvae by larvae of Atlantic mackerel, Scomber scombrus: trophodynamics and potential impact on recruitment. Fishery Bulletin, 94(2):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/cannibalism-

and-predation-fish-larvae-larvae-atlantic-mackerel-scomber-
scombrus;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1996/942/fortier.pdf.

Sminkey:1996:DAS

[4013] Thomas R. Sminkey and John A. Musick. Demographic analysis of the sandbar shark, Carcharhinus plumbeus, in the western North Atlantic. Fishery Bulletin, 94(2):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/demographic-analysis-sandbar-shark-carcharhinus-plumbeus-
western-north-atlantic;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1996/942/sminkey.pdf.

Franks:1996:SCJ

[4014] James S. Franks, Nikola M. Garber, and James R. Warren. Stomach contents of juvenile cobia, Rachycentron canadum, from the northern Gulf of Mexico. Fishery Bulletin, 94(2):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/stomach-contents-ofjuvenile-cobia-rachycentron-
canadum-northern-gulf-mexico;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1996/942/franks.pdf.

Somerton:1996:CBG

[4015] David A. Somerton and William Donaldson. Contribution to the biology of the grooved and triangle Tanner crabs, Chionoecetes tanneri and c. angulatus, in the eastern Bering Sea. Fishery Bulletin, 94(2):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/contribution-biology-grooved-and-triangle-
tanner-crabs-chionoecetes-tanneri-and-c-angulatus;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1996/942/
somerton.pdf.

Krieger:1996:CCR

[4016] Kenneth J. Krieger and Michael F. Sigler. Catchability coefficient for rockfish estimated from trawl and submersible surveys. Fishery Bulletin, 94(2):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
catchability-coefficient-rockfish-estimated-trawl-and-submersible-
surveys;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1996/942/krieger.pdf.

Zimmermann:1996:BDA

[4017] Mark Zimmermann and Pamela Goddard. Biology and distribution of arrowtooth, Atheresthes stomias, and Kamchatka, A. evermanni, flounders in Alaskan waters. Fishery Bulletin, 94(2):358–370, 1996. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/biology-and-distribution-
arrowtooth-atheresthes-stomias-and-kamchatka-evermanni-flounders;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1996/942/zimmermann.pdf.

Baelde:1996:BDR

[4018] Pascale Baelde. Biology and dynamics of the reproduction of blueeye trevalla, Hyperoglyphe antarctica (Centrolophidae), off Tasmania, southern Australia. Fishery Bulletin, 94(2):??, 1996. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/biology-and-dynamics-
reproduction-blue-eye-trevalla-hyperoglyphe-antarctica-centrolophidae;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1996/942/baelde.pdf.

Laidig:1996:DPL

[4019] Thomas E. Laidig, Keith M. Sakuma, and Mary M. Nishimoto. Description of pelagic larval and juvenile stripetail rockfish, Sebastes saxicola (family Scorpaenidae), with an examination of larval growth. Fishery Bulletin, 94(2):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/description-pelagic-larval-and-juvenile-stripetail-
rockfish-sebastes-saxicola-family;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1996/942/laidig.pdf.

Caillouet:1996:RBS

[4020] Charles W. Caillouet, Jr., Donna J. Shaver, Wendy G. Teas, James M. Nance, Dickie B. Revera, and Andrea C. Cannon. Relationship between sea turtle stranding rates and shrimp fishing intensities in the northwestern Gulf of Mexico: 1986–1989 versus 1990–1993. Fishery Bulletin, 94(2):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14408;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/caillouet_0.pdf.

Bertram:1996:SDP

[4021] Douglas F. Bertram. Size-dependent predation risk in larval fishes: mechanistic inferences and levels of analysis. Fishery Bulletin, 94(2):??, 1996.

CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/size-dependent-predation-
risk-larval-fishes-mechanistic-inferences-and-levels-analysis;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1996/942/bertram.pdf.

Lavaniegos:1996:VDE

[4022] Bertha E. Lavaniegos. Vertical distribution of euphausiid life stages in waters adjacent to Baja California. Fishery Bulletin, 94(2): ??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/vertical-
distribution-euphausiid-life-stages-waters-adjacent-baja-california;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1996/942/lavaniegos.pdf.

Anonymous:1996:E

[4023] Anonymous. Erratum. Fishery Bulletin, 94(2):??, 1996. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14409;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/erratum.pdf.

Brodziak:1996:GLF

[4024] Jon K. T. Brodziak and William K. Macy III. Growth of long-finned squid, Loligo pealei, in the northwest Atlantic. Fishery Bulletin, 94(2):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/growth-long-finned-
squid-loligo-pealei-northwest-atlantic;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1996/942/brodziak.
pdf.

Newman:1996:VPS

[4025] Stephen J. Newman, David McB. Williams, and Garry R. Russ. Variabilty in the population structure of Lutjanus adetii (Castelnau, 1873) and L. quinquelineatus (Bloch, 1790) among reefs in the central Great Barrier Reef, Australia. Fishery Bulletin, 94(2):313–329, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/variabilty-population-structure-
lutjanus-adetii-castelnau-1873-and-l-quinquelineatus-bloch;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1996/942/newman.pdf.

DeMartini:1996:BAR

[4026] Edward E. DeMartini, Frank A. Parrish, and Denise M. Ellis. Barotrauma-associated regurgitation of food: implications for diet studies of Hawaiian pink snapper, Pristipomoides filamentosus (family Lutjanidae). Fishery Bulletin, 94(2):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/barotrauma-associated-regurgitation-food-implications-
diet-studies-hawaiian-pink-snapper;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1996/942/demartini.pdf.

Perkins:1996:MME

[4027] Peter C. Perkins and Elizabeth F. Edwards. A mixture model for estimating discarded bycatch from data with many zero observations: tuna discards in the eastern tropical Pacific Ocean. Fishery Bulletin, 94(2):330– 340, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/mixture-model-
estimating-discarded-bycatch-data-many-zero-observations-tuna-
discards;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1996/942/perkins.pdf.

Dorsey:1996:CAD

[4028] Suzanne E. Dorsey, Edward D. Houde, and John C. Gamble. Cohort abundances and daily variability in mortality of eggs and yolk-sac larvae of bay anchovy, Anchoa mitchilli, in Chesapeake Bay. Fishery Bulletin, 94(2):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
cohort-abundances-and-daily-variability-mortality-eggs-and-
yolk-sac-larvae-bay-anchovy;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1996/942/dorsey.pdf.

Read:1996:WCH

[4029] Andrew J. Read, John R. Nicolas, and James E. Craddock. Winter capture of a harbor porpoise in a pelagic drift net off North Carolina. Fishery Bulletin, 94(2):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
winter-capture-harbor-porpoise-pelagic-drift-net-north-carolina;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1996/942/read.pdf.

Love:1996:ALH

[4030] Milton S. Love, Andrew Brooks, Diego Busatta, John Stephens, and Paul A. Gregory. Aspects of the life histories of the kelp bass, Paralabrax clathratus, and barred sand bass, P. nebulifer; from the Southern California Bight. Fishery Bulletin, 94(3):??, 1996. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/aspects-life-histories-kelp-bass-
paralabrax-clathratus-and-barred-sand-bass-p-nebulifer;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1996/943/
love.pdf.

Pennington:1996:EMV

[4031] Michael Pennington. Estimating the mean and variance from highly skewed marine data. Fishery Bulletin, 94(3):??, 1996. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14410;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/pennington%20(1).
pdf.

Bolz:1996:AGL

[4032] George R. Bolz and Bruce R. Burns. Age and growth of larval Atlantic herring, Clupea harengus: a comparative study. Fishery Bulletin, 94 (3):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/age-and-
growth-larval-atlantic-herring-clupea-harengus-comparative-
study;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1996/943/bolz.pdf.

Orsi:1996:MDO

[4033] Joseph A. Orsi and Herbert W. Jaenicke. Marine distribution and origin of prerecruit Chinook salmon, Oncorhynchus tshawytscha, in southeastern Alaska. Fishery Bulletin, 94(3):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/marine-distribution-and-origin-prerecruit-
chinook-salmon-oncorhynchus-tshawytscha;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1996/943/orsi.pdf.

Bower:1996:EPD

[4034] John R. Bower. Estimated paralarval drift and inferred hatching sites for Ommastrephes bartramii (Cephalopoda: Ommastrephidae) near the Hawaiian Archipelago. Fishery Bulletin, 94(3):??, 1996. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/estimated-paralarval-drift-and-
inferred-hatching-sites-ommastrephes-bartramii-cephalopoda;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1996/943/bower.pdf.

Quinn:1996:OGS

[4035] Thomas P. Quinn, Jennifer L. Nielsen, Christina Gan, Martin J. Unwin, Richard Wilmot, Charles Guthrie, and Fred M. Utter. Origin and genetic structure of Chinook salmon, Oncorhynchus tshawytscha, transplanted from California to New Zealand: allozyme and mtDNA evidence. Fishery Bulletin, 94(3):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/origin-and-genetic-structure-chinook-
salmon-oncorhynchus-tshawytscha-transplanted;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1996/943/quinn.
pdf.

Broadhurst:1996:ESM

[4036] Matt K. Broadhurst, Steven J. Kennelly, and Gerard O’Doherty. Effects of square-mesh panels in codends and of haulback delay on bycatch reduction in the oceanic prawn-trawl fishery of New South Wales, Australia. Fishery Bulletin, 94(3):??, 1996. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/effects-square-mesh-
panels-codends-and-haulback-delay-bycatch-reduction-oceanic-
prawn-trawl;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1996/943/broadhurst.pdf.

Rountree:1996:SAG

[4037] Rodney A. Rountree and Kenneth W. Able. Seasonal abundance, growth, and foraging habits of juvenile smooth dogfish, Mustelus canis, in a New Jersey estuary. Fishery Bulletin, 94(3):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/seasonal-abundance-
growth-and-foraging-habits-juvenile-smooth-dogfish-mustelus-
canis-new;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1996/943/rountree.pdf.

Collette:1996:RFT

[4038] Bruce B. Collette and Christopher R. Aadland. Revision of the frigate tunas (Scombridae, Auxis), with descriptions of two new subspecies from the eastern Pacific. Fishery Bulletin, 94(3):423– 441, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/revision-
frigate-tunas-scombridae-auxis-descriptions-two-new-subspecies-
eastern-pacific;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1996/943/collette.pdf.

Sanchez:1996:MJS

[4039] Guillermo Sanchez, Harriet M. Perry, Christine B. Trigg, Michael Vecchione, Clyde Roper, and Patricia M. Biesiot. Morphometry of juvenile and subadult Loligo pealei and l. plei from the northern Gulf of Mexico. Fishery Bulletin, 94(3):??, 1996. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/morphometry-juvenile-
and-subadult-loligo-pealei-and-l-plei-northern-gulf-mexico;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1996/943/sanchez.pdf.

Crabtree:1996:AGM

[4040] Roy E. Crabtree, Christopher W. Harnden, Derke Snodgrass, and Connie Stevens. Age, growth, and mortality of bonefish, Albula vulpes, from the waters of the Florida keys. Fishery Bulletin, 94(3):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/age-growth-
and-mortality-bonefish-albula-vulpes-waters-florida-keys;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1996/
943/crabtree.pdf.

Stoner:1996:QCS

[4041] Allan W. Stoner and Melody Ray. Queen conch, Strombus gigas, in fished and unfished locations of the Bahamas: effects of a marine fishery reserve on adults, juveniles, and larval production. Fishery Bulletin, 94(3):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/queen-conch-strombus-
gigas-fished-and-unfished-locations-bahamas-effects-marine-
fishery;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1996/943/stoner.pdf.

Grijalva-Chon:1996:AVT

[4042] Jose M. Grijalva-Chon, Jorge De la Rosa-Velez, and Oscar Sosa-Nishizaki. Allozyme variability in two samples of swordfish, Xiphias gladius L., in the North Pacific Ocean. Fishery Bulletin, 94(3):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/allozyme-variability-two-
samples-swordfish-xiphias-gladius-l-north-pacific-ocean;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1996/943/
grijalva-chon.pdf.

VanDoornik:1996:TPC

[4043] Donald M. Van Doornik, George B. Milner, and Gary A. Winans. Transferrin polymorphism in Coho salmon,Oncorhynchus kisutch, and its application to genetic stock identification. Fishery Bulletin, 94(3):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/transferrin-polymorphism-
coho-salmononcorhynchus-kisutch-and-its-application-genetic-
stock;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1996/943/van%20doornik.pdf.

Leber:1996:MSE

[4044] Kenneth M. Leber, Steve M. Arce, David A. Sterritt, and Nathan P. Brennan. Marine stock-enhancement potential in nursery habitats of striped mullet, Mugil cephalus, in Hawaii. Fishery Bulletin, 94(3):452– 471, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/marine-stock-
enhancement-potential-nursery-habitats-striped-mullet-mugil-
cephalus-hawaii;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1996/943/leber.pdf.

Zheng:1996:CLA

[4045] Jie Zheng, Margaret C. Murphy, and Gordon H. Kruse. A catchlength analysis for crab populations. Fishery Bulletin, 94(3):??, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/catch-length-
analysis-crab-populations;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1996/943/zheng.pdf.

Friedland:1996:PSG

[4046] Kevin D. Friedland, Ruth E. Haas, and Tim F. Sheehan. Post-smolt growth, maturation, and survival of two stocks of Atlantic salmon. Fishery Bulletin, 94(4):654–663, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/post-smolt-growth-maturation-and-survival-
two-stocks-atlantic-salmon;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1996/944/friedland.pdf.

Sheridan:1996:FFP

[4047] Pete Sheridan. Forecasting the fishery for pink shrimp, Penaeus duorarum, on the Tortugas Grounds, Florida. Fishery Bulletin, 94(4):743– 755, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/forecasting-

fishery-pink-shrimp-penaeus-duorarum-tortugas-grounds-florida;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1996/944/sheridan.pdf.

Heist:1996:MDD

[4048] Edward J. Heist, John A. Musick, and John E. Graves. Mitochondrial DNA diversity and divergence among sharpnose sharks, Rhizoprionodon terraenovae, from the Gulf of Mexico and Mid-Atlantic Bight. Fishery Bulletin, 94(4):664–668, 1996. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/mitochondrial-dna-diversity-and-
divergence-among-sharpnose-sharks-rhizoprionodon;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/1996/944/
heist.pdf.

Torres:1996:ELR

[4049] Joseph J. Torres, Ross I. Brightman, Joe Donnelly, and John Harvey. Energetics of larval red drum, Sciaenops ocellatus. Part I: Oxygen consumption, specific dynamic action, and nitrogen excretion. Fishery Bulletin, 94(4):756–765, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/energetics-larval-red-drum-sciaenops-ocellatus-
part-i-oxygen-consumption-specific-dynamic;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1996/944/torres.pdf.

Hickford:1996:GSN

[4050] Michael J. H. Hickford and David R. Schiel. Gillnetting in southern New Zealand: duration effects of sets and entanglement modes of fish. Fishery Bulletin, 94(4):669–677, 1996. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/gillnetting-southern-new-zealand-
duration-effects-sets-and-entanglement-modes-fish;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1996/944/
hickford.pdf.

Yoklavich:1996:NAL

[4051] Mary M. Yoklavich, Valerie J. Loeb, Mary Nishimoto, and Brendan Daly. Nearshore assemblages of larval rockfishes and their physical environment off central California during an extended El Niño event, 1991–1993. Fishery Bulletin, 94(4):766–782, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/nearshore-assemblages-larval-rockfishes-and-their-physical-
environment-central-california;   https://spo.nmfs.noaa.gov/

node/14413;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1996/944/yoklavich.pdf;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/yoklavich%20(
2).pdf.

Allen:1996:DVS

[4052] M. James Allen, Patricia V. Velez, Dario W. Deihl, Sarah E. McFadden, and Michael Kelsh. Demographic variability in seafood consumption rates among recreational anglers of Santa Monica Bay, California, in 1991–1992. Fishery Bulletin, 94(4):597–610, 1996. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/demographic-variability-seafood-
consumption-rates-among-recreational-anglers-santa-monica;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1996/
944/allen.pdf.

Kimura:1996:KFD

[4053] Daniel K. Kimura, James W. Balsiger, and Daniel H. Ito. Kalman filtering the delay–difference equation: practical approaches and simulations. Fishery Bulletin, 94(4):678–691, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/kalman-filtering-delay-difference-equation-practical-
approaches-and-simulations;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1996/944/kimura.pdf.

Beacham:1996:UMD

[4054] Terry D. Beacham. The use of minisatellite DNA variation for stock identification of chum salmon, Oncorhynchus keta. Fishery Bulletin, 94 (4):611–627, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/use-
minisatellite-dna-variation-stock-identification-chum-salmon-
oncorhynchus-keta;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1996/944/beacham.pdf.

Lowry:1996:GVP

[4055] Patricia S. Lowry, Nicholas G. Elliott, Gordon K. Yearsley, and Robert D. Ward. Genetic variation and phylogenetic relationships of seven oreo species (Teleostei, Oreosomatidae) inferred from allozyme analysis. Fishery Bulletin, 94(4):792–706, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/genetic-variation-and-phylogenetic-relationships-seven-
oreo-species-teleostei-oreosomatidae;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1996/944/lowry.pdf.

Anonymous:1996:A

[4056] Anonymous. Announcement. Fishery Bulletin, 94(4):787, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14415;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/announcement_0.pdf.

Chapman:1996:RCW

[4057] Frank A. Chapman, J. P. Van Eenennaam, and S. I. Doroshov. The reproductive condition of white sturgeon, Acipenser transmentanus, in San Francisco Bay, California. Fishery Bulletin, 94(4):628–634, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/reproductive-
condition-white-sturgeon-acipenser-transmentanus-san-fransisco-
bay-california;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1996/944/chapman.pdf.

Myers:1996:FRR

[4058] Ransom A. Myers and Nicholas J. Barrowman. Is fish recruitment related to spawner abundance? Fishery Bulletin, 94(4):707–724, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/fish-recruitment-related-
spawner-abundance;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1996/944/myers.pdf.

Anonymous:1996:FBI

[4059] Anonymous. Fishery Bulletin index: Volume 94 (1–4), 1996. Fishery Bulletin, 94(4):788–796, 1996. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/node/
14416;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/index%20(7).pdf.

Cuellar:1996:RSM

[4060] Nicole Cuellar, George R. Sedberry, and David W. Wyanski. Reproductive seasonality, maturation, fecundity, and spawning frequency of the vermilion snapper, Rhomboplites aurorubens, off the southeastern United States. Fishery Bulletin, 94(4):635–653, 1996. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/reproductive-seasonality-
maturation-fecundity-and-spawning-frequency-vermilion-snapper;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1996/944/cuellar.pdf.

Ramon:1996:SSA

[4061] Darlene Ramon and Kevin Bailey. Spawning seasonality of albacore, Thunnus alalunga, in the South Pacific Ocean. Fishery Bulletin, 94(4):725–733, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/spawning-seasonality-albacore-thunnus-alalunga-
south-pacific-ocean;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1996/944/ramon.pdf.

Francis:1996:DHG

[4062] R. I. C. Chris Francis. Do herring grow faster than orange roughy? Fishery Bulletin, 94(4):783–786, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/do-herring-grow-faster-orange-roughy;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1996/944/francis.pdf.

Schultz:1996:IMS

[4063] Duane R. Schultz, Patricia I. Arnold, Thomas R. Capo, Claire B. Paris-Umouzy, Joseph E. Serafy, and William J. Richards. Immunologic methods for species identification of early life stages of lutjanid fishes from the western central Atlantic. Part I: Characterization of an interspecies protein. Fishery Bulletin, 94(4):734–742, 1996. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/immunologic-methods-
species-identification-early-life-stages-lutjanid-fishes-western;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1996/944/schultz.pdf.

Hinton:1997:UGI

[4064] Michael G. Hinton, Ronald G. Taylor, and Michael D. Murphy. Use of gonad indices to estimate the status of reproductive activity of female swordfish, Xiphias gladius: a validated classification method. Fishery Bulletin, 95(1):80–84, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/use-gonad-indices-estimate-status-reproductive-
activity-female-swordfish-xiphias-gladius;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1997/951/hinton.pdf.

Shi:1997:GSE

[4065] Yunbing Shi, Donald R. Gunderson, and Patrick J. Sullivan. Growth and survival of 0+ English sole, Pleuronectes vetulus, in estuaries and adjacent nearshore waters off Washington. Fishery Bulletin, 95(1):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/growth-and-survival-
0-english-sole-pleuronectes-vetulus-estuaries-and-adjacent-
nearshore;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1997/951/shi.pdf.

Kane:1997:PST

[4066] Joseph Kane. Persistent spatial and temporal abundance patterns for late-stage copepodites of Centropages hamatus (Copepoda: Calanoida) in the U.S. northeast continental shelf ecosystem. Fishery Bulletin, 95(1):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/persistent-spatial-and-temporal-abundance-
patterns-late-stage-copepodites-centropages;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1997/951/kane.
pdf.

Zavala-Gonzalez:1997:ECS

[4067] Alfredo Zavala-González and Eric Mellink. Entanglement of California sea lions, Zalophus californianus californianus, in fishing gear in the central-northern part of the Gulf of California, Mexico. Fishery Bulletin, 95(1):180–184, 1997. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
entanglement-california-sea-lions-zalophus-californianus-californianus-
fishing-gear-central;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1997/951/zavala-gonzalez.pdf.

Livingston:1997:EAP

[4068] Mary E. Livingston, Marianne Vignaux, and Kathy A. Schofield. Estimating the annual proportion of nonspawning adults in New Zealand hoki, Macruronus novaezelandiae. Fishery Bulletin, 95(1):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/estimating-
annual-proportion-nonspawning-adults-new-zealand-hoki-macruronus-
novaezelandiae;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1997/951/livingston.pdf.

Fisher:1997:DOJ

[4069] Joseph P. Fisher and William G. Pearcy. Dietary overlap of juvenile fall- and spring-run Chinook salmon, Oncorhynchus tshawytscha, in Coos Bay, Oregon. Fishery Bulletin, 95(1):??, 1997. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14419;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/fisher%20(1)_0.pdf.

Bertram:1997:IVG

[4070] Douglas F. Bertram, Thomas J. Miller, and William C. Leggett. Individual variation in growth and development during the early life stages of winter flounder, Pleuronectes americanus. Fishery Bulletin, 95(1):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/individual-variation-growth-and-development-during-
early-life-stages-winter-flounder;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1997/951/bertram.pdf.

McKenna:1997:SDF

[4071] James E. McKenna, Jr. Structure and dynamics of the fishery harvest in Broward County, Florida, during 1989. Fishery Bulletin, 95 (1):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/structure-
and-dynamics-fishery-harvest-broward-county-florida-during-
1989;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1997/951/mckenna.pdf.

Anonymous:1997:Ea

[4072] Anonymous. Erratum. Fishery Bulletin, 95(1):??, 1997. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14421;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/erratum%20(1).pdf.

Crockford:1997:AEL

[4073] Susan Janet Crockford. Archeological evidence of large northern bluefin tuna, Thunnus thynnus, in coastal waters of British Columbia and northern Washington. Fishery Bulletin, 95(1):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/archeological-
evidence-large-northern-bluefin-tuna-thunnus-thynnus-coastal-
waters-british;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1997/951/crockford.pdf.

McQuinn:1997:YCT

[4074] Ian H. McQuinn. Year-class twinning in sympatric seasonal spawning populations of Atlantic herring, Clupea harengus. Fishery Bulletin, 95(1): ??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/year-class-

twinning-sympatric-seasonal-spawning-populations-atlantic-herring-
clupea;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1997/951/mcquinn.pdf.

Goodyear:1997:FAD

[4075] C. Phillip Goodyear. Fish age determined from length: an evaluation of three methods using simulated red snapper data. Fishery Bulletin, 95(1):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fish-age-
determined-length-evaluation-three-methods-using-simulated-
red-snapper-data;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1997/951/goodyear.pdf.

Parrish:1997:NHR

[4076] Frank A. Parrish, Edward E. DeMartini, and Denise M. Ellis. Nursery habitat in relation to production of juvenile pink snapper, Pristipomoides filamentosus, in the Hawaiian Archipelago. Fishery Bulletin, 95 (1):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/nursery-
habitat-relation-production-juvenile-pink-snapper-pristipomoides-
filamentosus;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1997/951/parrish.pdf.

Griffiths:1997:LHS

[4077] Marc H. Griffiths. The life history and stock separation of silver kob, Argyrosomus inodorus, in South African waters. Fishery Bulletin, 95(1): ??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/life-history-
and-stock-separation-silver-kob-argyrosomus-inodorus-south-
african-waters;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1997/951/griffiths.pdf.

Restrepo:1997:AHB

[4078] Victor R. Restrepo and Joseph E. Powers. Application of highbreakdown robust regression to tuned stock assessment models. Fishery Bulletin, 95(1):149–160, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/application-high-breakdown-robust-regression-tuned-
stock-assessment-models;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1997/951/restrepo.pdf.

Hampton:1997:ETR

[4079] John Hampton. Estimates of tag-reporting and tag-shedding rates in a large-scale tuna tagging experiment in the western tropical Pacific Ocean. Fishery Bulletin, 95(1):??, 1997. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/estimates-tag-reporting-and-tag-
shedding-rates-large-scale-tuna-tagging-experiment-western;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1997/951/hampton.pdf.

Richardson:1997:MDD

[4080] Linda R. Richardson and John R. Gold. Mitochondrial DNA diversity in and population structure of red grouper, Epinephelus morio, from the Gulf of Mexico. Fishery Bulletin, 95(1):??, 1997. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/mitochondrial-dna-diversity-and-
population-structure-red-grouper-epinephelus-morio-gulf;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1997/951/
richardson.pdf.

Alvarez:1997:BAA

[4081] Federico Álvarez and Francisco Alemany. Birthdate analysis and its application to the study of recruitment of the Atlanto–Iberian sardine Sardina pilchardus. Fishery Bulletin, 95(2):187–194, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/birthdate-analysis-
and-its-application-study-recruitment-atlanto-iberian-sardine-
sardina;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1997/952/alvarez.pdf.

Leber:1997:IRS

[4082] Kenneth M. Leber, H. Lee Blankenship, Steve M. Arce, and Nathan P. Brennan. Influence of release season on size- dependent survival of cultured striped mullet, Mugli cephalus, in a Hawaiian estuary. Fishery Bulletin, 95(2):267–279, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/influence-release-season-
size-dependent-survival-cultured-striped-mullet-mugli-cephalus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1997/952/leber.pdf.

Timmons:1997:AGS

[4083] Maryellen Timmons and Richard N. Bray. Age, growth, and sexual maturity of shovelnose guitarfish, Rhinobatos productus (Ayres). Fishery Bulletin, 95(2):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/age-growth-and-sexual-maturity-shovelnose-
guitarfish-rhinobatos-productus-ayres;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1997/952/timmons.pdf.

Barber:1997:DFA

[4084] Willard E. Barber, Ronald L. Smith, Mark Vallarino, and Robert M. Meyer. Demersal fish assemblages of the notheastern Chukchi Sea, Alaska. Fishery Bulletin, 95(2):195–209, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/demersal-fish-assemblages-notheastern-chukchi-
sea-alaska;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1997/952/barber.pdf.

Legault:1997:CAC

[4085] Christopher M. Legault and Nelson M. Ehrhardt. Correcting annual catches from seasonal fisheries for use in virtual population analysis. Fishery Bulletin, 95(2):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
correcting-annual-catches-seasonal-fisheries-use-virtual-population-
analysis;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1997/952/legault.pdf.

Tucker:1997:DLR

[4086] John W. Tucker and Sabine R. Alshuth. Development of laboratoryreared sheepshead, Archosargus probatocephalus (Pisces: Sparidae). Fishery Bulletin, 95(2):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/development-laboratory-reared-sheepshead-archosargus-
probatocephalus-pisces-sparidae;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1997/952/tucker.pdf.

Biggs:1997:NPC

[4087] Douglas C. Biggs, Robert A. Zimmerman, Rebeca Gasca, Eduardo Suarez-Morales, Ivan Castellanos, and Robert R. Leben. Note on plankton and cold-core rings in the Gulf of Mexico. Fishery Bulletin, 95(2): ??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/note-plankton-

and-cold-core-rings-gulf-mexico;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1997/952/biggs.pdf.

Lenarz:1997:ETL

[4088] William H. Lenarz and Franklin R. Shaw. Estimates of tag loss from double-tagged sablefish, Anoplopoma fimbria. Fishery Bulletin, 95(2): ??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/estimates-tag-
loss-double-tagged-sablefish-anoplopoma-fimbria;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/1997/952/
lenarz.pdf.

Abitia-Cardenas:1997:FHE

[4089] Leonardo A. Abitia-Cardenas, Felipe Galvan-Magaña, and Jesus Rodriguez-Romero. Food habits and energy values of prey of striped marlin, Tetrapturus audax, off the coast of Mexico. Fishery Bulletin, 95 (2):360–368, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14422;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/abitia-cardenas.pdf.

Broadhurst:1997:ENG

[4090] Matt K. Broadhurst, Steven J. Kennelly, John W. Watson, and Ian K. Workman. Evaluations of the Nordmore grid and secondary bycatch-reducing devices (BRD’s) in the Hunter River prawn-trawl fishery, Australia. Fishery Bulletin, 95(2):209–218, 1997. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/evaluations-nordmore-
grid-and-secondary-bycatch-reducing-devices-brds-hunter-river-
prawn;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1997/952/broadhurst.pdf.

Lutcavage:1997:ASG

[4091] Molly Lutcavage, Scott Kraus, and Wayne Hoggard. Aerial survey of giant bluefin tuna, Thunnus thynnus, in the Great Bahama Bank, Straits of Florida, 1995. Fishery Bulletin, 95(2):??, 1997. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/aerial-survey-giant-bluefin-tuna-
thunnus-thynnus-great-bahama-bank-straits-florida-1995;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1997/952/
lutcavage.pdf.

Gunderson:1997:SPD

[4092] Donald R. Gunderson. Spatial patterns in the dynamics of slope rockfish stocks and their implications for management. Fishery Bulletin, 95 (2):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/spatial-
patterns-dynamics-slope-rockfish-stocks-and-their-implications-
management;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1997/952/gunderson.pdf.

Martini:1997:PPA

[4093] Frederic Martini, John B. Heiser, and Michael P. Lesser. A population profile for Atlantic hagfish, Myxine glutinosa (L.), in the Gulf of Maine. Part I: Morphometries and reproductive state. Fishery Bulletin, 95(2):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/population-profile-
atlantic-hagfish-myxine-glutinosa-l-gulf-maine-part-i-morphometries-
and;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1997/952/martini.pdf.

Johnson:1997:FEW

[4094] Lyndal L. Johnson, Sean Y. Sol, Daniel P. Lomax, Gregory M. Nelson, Catherine A. Sloan, and Edmundo Casillas. Fecundity and egg weight in English sole, Pleuronectes vetulus, from Puget Sound, Washington: influence of nutritional status and chemical contaminants. Fishery Bulletin, 95(2):??, 1997. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/fecundity-and-egg-weight-english-
sole-pleuronectes-vetulus-puget-sound-washington-influence;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1997/952/johnson.pdf.

McDermott:1997:RCS

[4095] Susanne F. McDermott and Sandra A. Lowe. The reproductive cycle and sexual maturity of Atka mackerel, Pleurogrammus monopterygius, in Alaska waters. Fishery Bulletin, 95(2):??, 1997. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/reproductive-cycle-
and-sexual-maturity-atka-mackerel-pleurogrammus-monopterygius-
alaska;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1997/952/mcdermott.pdf.

Lamkin:1997:LCA

[4096] John Lamkin. The Loop Current and the abundance of larval Cubiceps pauciradiatus (Pisces: Nomeidae) in the Gulf of Mexico: evidence for physical and biological interaction. Fishery Bulletin, 95 (2):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/loop-
current-and-abundance-larval-cubiceps-pauciradiatus-pisces-
nomeidae-gulf-mexico;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1997/952/lamkin.pdf.

Robertson:1997:POP

[4097] Kelly M. Robertson and Susan J. Chivers. Prey occurrence in pantropical spotted dolphins, Stenella attenuata, from the eastern tropical Pacific. Fishery Bulletin, 95(2):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/prey-occurrence-pantropical-spotted-dolphins-stenella-
attenuata-eastern-tropical-pacific;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/1997/952/robertson.pdf.

Lazzari:1997:PER

[4098] Mark A. Lazzari, David K. Stevenson, and Stephen M. Ezzy. Physical environment and recruitment variabiliity of Atlantic herring, Clupea harengus, in the Gulf of Maine. Fishery Bulletin,

95(2):??, 1997.  CODEN FSYBAY.  ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
physical-environment-and-recruitment-variabiliity-atlantic-
herring-clupea-harengus-gulf;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1997/952/lazzari.pdf.

Sekiguchi:1997:VDS

[4099] Keiko Sekiguchi and Peter B. Best. In vitro digestibility of some prey species of dolphins. Fishery Bulletin, 95(2):??, 1997. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/vitro-digestibility-some-
prey-species-dolphins;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1997/952/sekiguchi.pdf.

Armstrong:1997:SOC

[4100] Michael P. Armstrong. Seasonal and ontogenetic changes in distribution and abundance of smooth flounder, Pleuronectes putnami, and winter flounder, Pleuronectes americanus, along estuarine depth and salinity gradients. Fishery Bulletin, 95(3):??, 1997. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/seasonal-and-ontogenetic-changes-
distribution-and-abundance-smooth-flounder-pleuronectes;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1997/953/
armstrong.pdf.

Paperno:1997:DGI

[4101] Richard Paperno, Timothy E. Targett, and Paul A. Grecay. Daily growth increments in otoliths of juvenile weakfish, Cynoscion regalis: experimental assessment of changes in increment width with changes in feeding rate, growth rate, and condition factor. Fishery Bulletin, 95(3):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/daily-growth-increments-otoliths-juvenile-weakfish-
cynoscion-regalis-experimental;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1997/953/paperno.pdf.

Zimmermann:1997:MFA

[4102] Mark Zimmermann. Maturity and fecundity of arrowtooth flounder, Atheresthes stomias, from the Gulf of Alaska. Fishery Bulletin, 95 (3):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/maturity-
and-fecundity-arrowtooth-flounder-atheresthes-stomias-gulf-
alaska;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1997/953/zimmermann.pdf.

Brightman:1997:ELR

[4103] Ross I. Brightman, Joseph J. Torres, Joseph Donnelly, and M. Elizabeth Clarke. Energetics of larval red drum, Sciaenops ocellatus. Part II: Growth and biochemical indicators. Fishery Bulletin, 95(3):431–444, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/energetics-larval-red-drum-sciaenops-ocellatus-
part-ii-growth-and-biochemical-indicators;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1997/953/brightman.
pdf.

Peters:1997:DAG

[4104] John S. Peters and David J. Schmidt. Daily age and growth of larval and early juvenile Spanish mackerel, Scomberomorus maculatus, from the South Atlantic Bight. Fishery Bulletin, 95(3):??, 1997. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/daily-age-and-growth-larval-and-
early-juvenile-spanish-mackerel-scomberomorus-maculatus;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1997/953/
peters.pdf.

Norcross:1997:HMJ

[4105] Brenda L. Norcross, Franz-Josef Müter, and Brenda A. Holladay. Habitat models for juvenile pleuronectids around Kodiak Island, Alaska. Fishery Bulletin, 95(3):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/node/
14423;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/norcross.pdf.

Cadrin:1997:RAV

[4106] Steven X. Cadrin and Douglas S. Vaughan. Retrospective analysis of virtual population estimates for Atlantic menhaden stock assessment. Fishery Bulletin, 95(3):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/retrospective-analysis-virtual-population-
estimates-atlantic-menhaden-stock-assessment;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1997/953/cadrin.
pdf.

Roelke:1997:USI

[4107] Lynn A. Roelke and Luis A. Cifuentes. Use of stable isotopes to assess groups of king mackerel, Scomberomorus cavalla, in the Gulf of Mexico and southeastern Florida. Fishery Bulletin, 95(3):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/use-stable-
isotopes-assess-groups-king-mackerel-scomberomorus-cavalla-
gulf-mexico-and;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1997/953/roelke.pdf.

Anonymous:1997:Eb

[4108] Anonymous. Erratum. Fishery Bulletin, 95(3):??, 1997. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14424;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/erratum%20(2).pdf.

Crabtree:1997:MRS

[4109] Roy E. Crabtree, Derke Snodgrass, and Christopher W. Harnden. Maturation and reproductive seasonality in bonefish, Albula vulpes, from the waters of the Florida keys. Fishery Bulletin, 95(3):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/maturation-
and-reproductive-seasonality-bonefish-albula-vulpes-waters-
florida-keys;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1997/953/crabtree.pdf.

Rogers:1997:EFI

[4110] Donna R. Rogers, Barton D. Rogers, Janaka A. de Silva, and Vernon L. Wright. Effectiveness of four industry-developed bycatch reduction devices in Louisiana’s inshore waters. Fishery Bulletin, 95(3):552–565, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/effectiveness-four-
industry-developed-bycatch-reduction-devices-louisianas-inshore-
waters;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/1997/953/rogers.pdf.

Estrella:1997:SMO

[4111] Bruce T. Estrella and Thomas D. Morrissey. Seasonal movement of offshore American lobster, Homarus americanus, tagged along the eastern shore of Cape Cod, Massachusetts. Fishery Bulletin, 95(3): ??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/seasonal-
movement-offshore-american-lobster-homarus-americanus-tagged-
along-eastern-shore;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1997/953/estrella.pdf.

Rodriguez:1997:AGT

[4112] Martha J. Román Rodriguez and M. Gregory Hammann. Age and growth of totoaba, Totoaba macdonaldi (Sciaenidae), in the upper Gulf of California. Fishery Bulletin, 95(3):620–628, 1997. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/age-and-growth-totoaba-totoaba-
macdonald-sciaenidae-upper-gulf-california;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1997/953/roman-rodriguez.
pdf.

Hettler:1997:DVA

[4113] William F. Hettler, Jr., David S. Peters, David R. Colby, and Elisabeth H. Laban. Daily variability in abundance of larval fishes inside Beaufort Inlet. Fishery Bulletin, 95(3):477–493, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/daily-variability-abundance-larval-fishes-

inside-beaufort-inlet;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1997/953/hettler.pdf.

Thedinga:1997:MRG

[4114] John F. Thedinga, Adam Moles, and Jeffrey T. Fujioka. Mark retention and growth of jet-injected juvenile marine fish. Fishery Bulletin, 95(3):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/mark-retention-and-
growth-jet-injected-juvenile-marine-fish;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1997/953/thedinga.
pdf.

Munehara:1997:RBE

[4115] Hiroyuki Munehara. The reproductive biology and early life stages of Podothecus sachi (Pisces: Agonidae). Fishery Bulletin, 95(3):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/reproductive-biology-and-
early-life-stages-podothecus-sachi-pisces-agonidae;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1997/953/
munehara.pdf.

Ward:1997:GPS

[4116] Robert D. Ward, Nicholas G. Elliott, Bronwyn H. Innes, Adam J. Smolenski, and Peter M. Grewe. Global population structure of yellowfin tuna, Thunnus albacares, inferred from allozyme and mitochondrial DNA variation. Fishery Bulletin, 95(3):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/global-population-structure-yellowfin-tuna-
thunnus-albacares-inferred-allozyme-and;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1997/953/ward.pdf.

Andrew:1997:OSE

[4117] Neil L. Andrew and Yong Chen. Optimal sampling for estimating the size structure and mean size of abalone caught in a New South Wales fishery. Fishery Bulletin, 95(3):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/optimal-sampling-estimating-size-structure-and-
mean-size-abalone-caught-new-south-wales;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1997/953/andrew.pdf.

Nichol:1997:EGB

[4118] Daniel G. Nichol. Effects of geography and bathymetry on growth and maturity of yellowfin soul, Pleuronectes asper, in the eastern Bering Sea. Fishery Bulletin, 95(3):??, 1997. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/effects-geography-and-bathymetry-
growth-and-maturity-yellowfin-soul-pleuronectes-asper;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/1997/953/
nichol.pdf.

Zeldis:1997:EOR

[4119] John R. Zeldis, R. I. Chris Francis, Malcolm R. Clark, Jonathan K. V. Ingerson, Paul J. Grimes, and Marianne Vignaux. An estimate of orange roughy, Hoplostethus atlanticus, biomass using the daily fecundity reduction method. Fishery Bulletin, 95(3):576– 597, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/estimate-
orange-roughy-hoplostethus-atlanticus-biomass-using-daily-fecundity-
reduction;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1997/953/zeldis.pdf.

Barbieri:1997:YPR

[4120] Luiz R. Barbieri, Mark E. Chittenden, Jr., and Cynthia M. Jones. Yieldper-recruit analysis and management strategies for Atlantic croaker, Micropogonias undulatus, in the Middle Atlantic Bight. Fishery Bulletin, 95 (4):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/yield-recruit-
analysis-and-management-strategies-atlantic-croaker-micropogonias-
undulatus;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1997/954/barbieri.pdf.

Gaughan:1997:ADF

[4121] Daniel J. Gaughan and Ian C. Potter. Analysis of diet and feeding strategies within an assemblage of estuarine larval fish and an objective assessment of dietary niche overlap. Fishery Bulletin, 95(4):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/analysis-diet-
and-feeding-strategies-within-assemblage-estuarine-larval-fish-
and-objective;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1997/954/gaughan.pdf.

Rotunno:1997:TSS

[4122] Teresa Rotunno and Robert K. Cowen. Temporal and spatial spawning patterns of the Atlantic butterfish, Peprilus triacanthus, in the South and Middle Atlantic Bights. Fishery Bulletin, 95(4):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/temporal-and-spatial-
spawning-patterns-atlantic-butterfish-peprilus-triacanthus-
south-and;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1997/954/rotunno.pdf.

Bradbury:1997:DSA

[4123] Carole Bradbury, John M. Green, and Michael Bruce-Lockhart. Daily and seasonal activity patterns of female cunnet, Tautogolabrus adspersus (Labridae), in Newfoundland. Fishery Bulletin, 95 (4):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/daily-
and-seasonal-activity-patterns-female-cunnet-tautogolabrus-
adspersus-labridae;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1997/954/bradbury.pdf.

Grana-Raffucci:1997:ADL

[4124] Felix A. Grana-Raffucci and Richard S. Appledoorn. Age determination of larval strombid gastropods by means of growth increment counts in statoliths. Fishery Bulletin, 95(4):857–862, 1997. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/age-determination-larval-
strombid-gastropods-means-growth-increment-counts-statoliths;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1997/954/grana-raffucci.pdf.

Smith:1997:CTG

[4125] Peter J. Smith, Peter G. Benson, and S. Margaret McVeagh. A comparison of three genetic methods used for stock discrimination of orange roughy, Hoplostethus atlanticus: allozymes, mitochondrial DNA, and random amplified polymorphic DNA. Fishery Bulletin, 95(4):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
comparison-three-genetic-methods-used-stock-discrimination-
orange-roughy-hoplostethus;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1997/954/smith.pdf.

Broadhurst:1997:CSM

[4126] Matt K. Broadhurst and Steven J. Kennelly. The composite squaremesh panel: a modification to codends for reducing unwanted bycatch and increasing catches of prawns throughout the New South Wales oceanic prawn-trawl fishery. Fishery Bulletin, 95(4):653– 664, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/composite-
square-mesh-panel-modification-codends-reducing-unwanted-bycatch-
and-increasing;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1997/954/broadhurst.pdf.

Harris:1997:CLH

[4127] Patrick J. Harris and John C. McGovern. Changes in the life history of red porgy, Pagrus pagrus, from the southeastern United States, 1972–1994. Fishery Bulletin, 95(4):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/changes-life-history-red-
porgy-pagrus-pagrus-southeastern-united-states-1972-1994;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/1997/
954/harris.pdf.

Unwin:1997:SCS

[4128] Martin J. Unwin. Survival of Chinook salmon, Oncorhynchus tshawytscha, from a spawning tributary of the Rakaia River, New Zealand, in relation to spring and summer mainstem flows. Fishery Bulletin, 95(4):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
survival-chinook-salmon-oncorhynchus-tshawytscha-spawning-tributary-
rakaia-river-new;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1997/954/unwin.pdf.

Buckel:1997:MFP

[4129] Jeffrey A. Buckel and David O. Conover. Movements, feeding periods, and daily ration of piscivorous young-of-the-year bluefish, Pomatomus saltatrix, in the Hudson River estuary. Fishery Bulletin, 95 (4):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/movements-
feeding-periods-and-daily-ration-piscivorous-young-year-bluefish-
pomatomus;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1997/954/buckel.pdf.

Johnson:1997:RSP

[4130] Scott W. Johnson, Mark G. Carls, Robert P. Stone, Christine C. Brodersen, and Stanley D. Rice. Reproductive success of Pacific herring, Clupea pallasi, in Prince William Sound, Alaska, six years after the Exxon Valdez oil spill. Fishery Bulletin, 95(4):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/reproductive-success-
pacific-herring-clupea-pallasi-prince-william-sound-alaska-

six-years;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/1997/954/johnson.pdf.

Wallin:1997:STR

[4131] Julie E. Wallin, John M. Ransier, Sondra Fox, and Robert H. McMichael, Jr. Short-term retention of coded wire and internal anchor tags in juvenile common snook, Centropomus undecimalis. Fishery Bulletin, 95(4):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/short-term-retention-coded-wire-
and-internal-anchor-tags-juvenile-common-snook-centropomus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1997/954/wallin.pdf.

Chaloupka:1997:PGF

[4132] Milani Chaloupka and George R. Zug. A polyphasic growth function for the endangered kemps ridley sea turtle, Lepidochelys kempii. Fishery Bulletin, 95(4):849–856, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/polyphasic-growth-function-endangered-kemps-ridley-
sea-turtle-lepidochelys-kempii;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/1997/954/chaloupka.pdf.

Murphy:1997:BCR

[4133] Michael D. Murphy. Bias in Chapman–Robson and least-squares estimators of mortality rates for steady-state populations. Fishery Bulletin, 95 (4):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/bias-chapman-
robson-and-least-squares-estimators-mortality-rates-steady-
state-populations;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1997/954/murphy.pdf.

Weinrich:1997:SDH

[4134] Mason Weinrich, Malcolm Martin, Rachel Griffiths, Jennifer Bove, and Mark Schilling. A shift in distribution of humpback whales, Megaptera novaeangliae, in response to prey in the southern Gulf of Maine. Fishery Bulletin, 95(4):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
shift-distribution-humpback-whales-megaptera-novaeangliae-response-
prey-southern-gulf-maine;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1997/954/weinrich.pdf.

Craig:1997:PBH

[4135] Peter C. Craig, J. Howard Choat, Lynda M. Axe, and Suesan Saucerman. Population biology and harvest of the coral reef surgeonfish Acanthurus lineatus in American Samoa. Fishery Bulletin, 95(4):680–693, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/population-biology-and-harvest-coral-reef-
surgeonfish-acanthurus-lineatus-american-samoa;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/1997/954/craig.
pdf.

Myers:1997:MPG

[4136] Ransom A. Myers, Gordon Mertz, and P. Stacey Fowlow. Maximum population growth rates and recovery times for Atlantic cod, Gadus morhua. Fishery Bulletin, 95(4):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/maximum-population-growth-rates-and-recovery-times-
atlantic-cod-gadus-morhua;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/1997/954/myers.pdf.

Zhao:1997:TVS

[4137] Boxian Zhao and John C. McGovern. Temporal variation in sexual maturity and gear-specific sex ratio of the vermilion snapper, Rhomboplites aurorubens, in the South Atlantic Bight. Fishery Bulletin, 95(4):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/temporal-variation-
sexual-maturity-and-gear-specific-sex-ratio-vermilion-snapper;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1997/954/zhao.pdf.

DeVries:1997:STV

[4138] Douglas A. DeVries and Churchill B. Grimes. Spatial and temporal variation in age and growth of king mackerel, Scomberomorus cavalla, 1977–1992. Fishery Bulletin, 95(4):??, 1997. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/spatial-and-temporal-variation-
age-and-growth-king-mackerel-scomberomorus-cavalla-1977-1992;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1997/954/devries.pdf.

Nixon:1997:AGL

[4139] Stephen W. Nixon and Cynthia M. Jones. Age and growth of larval and juvenile Atlantic croaker, Micropogonias undulatus, from the Middle Atlantic Bight and estuarine waters of Virginia. Fishery Bulletin, 95 (4):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/age-and-
growth-larval-and-juvenile-atlantic-croaker-micropogonias-undulatus-
middle-atlantic;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/1997/954/nixon.pdf.

Anonymous:1997:R

[4140] Anonymous. 1997 referees. Fishery Bulletin, 95(4):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14425;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/referees.pdf.

Donohoe:1997:AGD

[4141] Christopher J. Donohoe. Age, growth, distribution, and food habits of recently settled white seabass, Atractoscion nobilis, off San Diego County, California. Fishery Bulletin, 95(4):??, 1997. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/age-growth-distribution-and-food-
habits-recently-settled-white-seabass-atractoscion-nobilis;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
1997/954/donohoe.pdf.

Ramon:1997:EFF

[4142] Darlene Ramon and Norm Bartoo. The effects of formalin and freezing on ovaries of albacore, Thunnus alalunga. Fishery Bulletin, 95(4):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/effects-formalin-and-
freezing-ovaries-albacore-thunnus-alalunga;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/1997/954/ramon.pdf.

Anonymous:1997:AI

[4143] Anonymous. Annual index. Fishery Bulletin, 95(4):??, 1997. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14426;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/index%20(8).pdf.

Munroe:1998:SET

[4144] Thomas A. Munroe. Systematics and ecology of tonguefishes of the genus Symphurus (Cynoglossidae: Pleuronectiformes) from the western Atlantic Ocean. Fishery Bulletin, 96(1):??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/systematics-and-ecology-tonguefishes-

genus-symphurus-cynoglossidae-pleuronectiformes;   https://
spo.nmfs.noaa.gov/content/systematics-and-ecology-tonguefishes-
genus-symphurus-cynoglossidae-pleuronectiformes-0;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
munro_0_0.pdf.

Campana:1998:RNT

[4145] Steven E. Campana and Cynthia M. Jones. Radiocarbon from nuclear testing applied to age validation of black drum, Pogonias cromis. Fishery Bulletin, 96(2):??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/radiocarbon-nuclear-testing-applied-age-validation-
black-drum-pogonias-cromis;   https://spo.nmfs.noaa.gov/content/
radiocarbon-nuclear-testing-applied-age-validation-black-drum-
pogonias-cromis-0;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/campana_0_0.pdf.

deSilva:1998:DPP

[4146] Janaka A. de Silva and Richard E. Condrey. Discerning patterns in patchy data: a categorical approach using Gulf menhaden, Brevoortia patronus, bycatch. Fishery Bulletin, 96(2):??, 1998. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/discerning-patterns-patchy-data-
categorical-approach-using-gulf-menhaden-brevoortia;   https:
//spo.nmfs.noaa.gov/content/discerning-patterns-patchy-data-
categorical-approach-using-gulf-menhaden-brevoortia-0;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
deSilva_0.pdf.

Edwards:1998:ETD

[4147] Elizabeth F. Edwards and Peter C. Perkins. Estimated tuna discard from dolphin, school, and log sets in the eastern tropical Pacific Ocean, 1989–1992. Fishery Bulletin, 96(2):??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/estimated-tuna-discard-dolphin-school-and-
log-sets-eastern-tropical-pacific-ocean-1989-0;   https://spo.
nmfs.noaa.gov/content/estimated-tuna-discard-dolphin-school-
and-log-sets-eastern-tropical-pacific-ocean-1989-1992;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
edwards_0.pdf.

Farley:1998:RDS

[4148] Jessica H. Farley and Tim L. O. Davis. Reproductive dynamics of southern bluefin tuna, Thunnus maccoyii. Fishery Bulletin, 96(2):??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/reproductive-
dynamics-southern-bluefin-tuna-thunnus-maccoyii;   https://
spo.nmfs.noaa.gov/content/reproductive-dynamics-southern-bluefin-
tuna-thunnus-maccoyii-0;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/farley_0.pdf.

Fiedler:1998:DPA

[4149] Paul C. Fiedler, Jay Barlow, and Tim Gerrodette. Dolphin prey abundance determined from acoustic backscatter data in eastern Pacific surveys. Fishery Bulletin, 96(2):237–247, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/dolphin-prey-abundance-determined-acoustic-
backscatter-data-eastern-pacific-surveys;   https://spo.nmfs.
noaa.gov/content/dolphin-prey-abundance-determined-acoustic-
backscatter-data-eastern-pacific-surveys-0;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/fiedler_
0.pdf.

Grover:1998:FHP

[4150] Jill J. Grover. Feeding habits of pelagic summer flounder, Paralichthys dentatus, larvae in oceanic and estuarine habitats. Fishery Bulletin, 96 (2):??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/feeding-
habits-pelagic-summer-flounder-paralichthys-dentatus-larvae-
oceanic-and-estuarine;   https://spo.nmfs.noaa.gov/content/
feeding-habits-pelagic-summer-flounder-paralichthys-dentatus-
larvae-oceanic-and-estuarine-0;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/grover_0.pdf.

Hyndes:1998:ACG

[4151] Glenn A. Hyndes, Margaret E. Platell, Ian C. Potter, and Rodney C. J. Lenanton. Age composition, growth, reproductive biology, and recruitment of King George whiting, Sillaginodes punctata, in coastal waters of southwestern Australia. Fishery Bulletin, 96 (2):??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/age-
composition-growth-reproductive-biology-and-recruitment-king-
george-whiting;   https://spo.nmfs.noaa.gov/content/age-composition-

growth-reproductive-biology-and-recruitment-king-george-whiting-
0;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/hyndes_0.pdf.

Julian:1998:EMM

[4152] Fred Julian and Marilyn Beeson. Estimates of marine mammal, turtle, and seabird mortality for two California gillnet fisheries: 1990–1995. Fishery Bulletin, 96(2):??, 1998. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/estimates-marine-mammal-turtle-
and-seabird-mortality-two-california-gillnet-fisheries-0;
https://spo.nmfs.noaa.gov/content/estimates-marine-mammal-turtle-
and-seabird-mortality-two-california-gillnet-fisheries-1990;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/julian_0.pdf.

Landingham:1998:FHJ

[4153] Joyce H. Landingham, Molly V. Sturdevant, and Richard D. Brodeur. Feeding habits of juvenile Pacific salmon in marine waters of southeastern Alaska and northern British Columbia. Fishery Bulletin, 96(2): 285–302, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/feeding-
habits-juvenile-pacific-salmon-marine-waters-southeastern-alaska-
and-northern;   https://spo.nmfs.noaa.gov/content/feeding-habits-
juvenile-pacific-salmon-marine-waters-southeastern-alaska-and-
northern-0;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/landingham_0.pdf.

McBride:1998:ICS

[4154] Richard S. McBride, Joseph B. O’Gorman, and Kenneth W. Able. Interspecific comparisons of searobin (Prionotus spp.) movements, size structure, and abundance in the temperate western North Atlantic. Fishery Bulletin, 96(2):303–314, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/interspecific-comparisons-searobin-prionotus-spp-movements-
size-structure-and-abundance;   https://spo.nmfs.noaa.gov/content/
interspecific-comparisons-searobin-prionotus-spp-movements-
size-structure-and-abundance-0;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/mcbride_0_0.pdf.

Nelson:1998:AGM

[4155] Gary A. Nelson. Abundance, growth, and mortality of young-of-the-year pinfish, Lagodon rhomboides, in three estuaries along the Gulf Coast of Florida. Fishery Bulletin, 96(2):315–328, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/abundance-growth-and-mortality-young-year-
pinfish-lagodon-rhomboids-three-estuaries-along;   https://
spo.nmfs.noaa.gov/content/abundance-growth-and-mortality-young-
year-pinfish-lagodon-rhomboids-three-estuaries-along-0;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
nelson_0_0.pdf.

Nero:1998:LFA

[4156] Redwood W. Nero, Charles H. Thompson, and Richard H. Love. Lowfrequency acoustic measurements of Pacific hake, Merluccius productus, off the west coast of the United States. Fishery Bulletin, 96(2): ??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/low-frequency-
acoustic-measurements-pacific-hake-merluccius-productus-west-
coast-united;   https://spo.nmfs.noaa.gov/content/low-frequency-
acoustic-measurements-pacific-hake-merluccius-productus-west-
coast-united-0;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/nero_0.pdf.

Sakurai:1998:PDS

[4157] Izumi Sakurai, Takashi Horii, Osamu Murakami, and Shigeru Nakao. Population dynamics and stock size prediction for the sunray surfclam, Mactra chinensis, at Tomakomai, southwest Hokkaido, Japan. Fishery Bulletin, 96(2):??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
population-dynamics-and-stock-size-prediction-sunray-surfclam-
mactra-chinensis-tomakomai;   https://spo.nmfs.noaa.gov/content/
population-dynamics-and-stock-size-prediction-sunray-surfclam-
mactra-chinensis-tomakomai-0;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/sakurai_0.pdf.

Beckmann:1998:RBG

[4158] Andreas T. Beckmann, Donald R. Gunderson, Bruce S. Miller, Raymond M. Buckley, and Betty Goetz. Reproductive biology, growth, and natural mortality of Puget Sound rockfish, Sebastes emphaeus (Starks, 1911). Fishery Bulletin, 96(2):??, 1998. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/reproductive-biology-
growth-and-natural-mortality-puget-sound-rockfish-sebastes-
emphaeus;   https://spo.nmfs.noaa.gov/content/reproductive-
biology-growth-and-natural-mortality-puget-sound-rockfish-sebastes-

emphaeus-0;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/beckmann_0.pdf.

Fernandez:1998:AGC

[4159] Stephanie Fernandez and Aleta A. Hohn. Age, growth, and calving season of bottlenose dolphins, Tursiops truncatus, off coastal Texas. Fishery Bulletin, 96(2):??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
age-growth-and-calving-season-bottlenose-dolphins-tursiops-
truncatus-coastal-texas;   https://spo.nmfs.noaa.gov/content/
age-growth-and-calving-season-bottlenose-dolphins-tursiops-
truncatus-coastal-texas-0;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fernandez_0.pdf.

Alonso:1998:DDD

[4160] Mariano Koen Alonso, Enrique Alberto Crespo, Néstor Aníbal García, Susana Noemí Pedraza, and Mariano Alberto Coscarella. Diet of dusky dolphins, Lagenorhynchus obscurus, in waters off Patagonia, Argentina. Fishery Bulletin, 96(2):??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/diet-dusky-dolphins-lagenorhynchus-obscurus-waters-
patagonia-argentina;   https://spo.nmfs.noaa.gov/content/diet-
dusky-dolphins-lagenorhynchus-obscurus-waters-patagonia-argentina-
0;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/alonso_0.pdf.

Macchi:1998:PES

[4161] Gustavo J. Macchi. Preliminary estimate of spawning frequency and batch fecundity of striped weakfish, Cynoscion striatus, in coastal waters off Buenos Aires province. Fishery Bulletin, 96(2):??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/preliminary-estimate-
spawning-frequency-and-batch-fecundity-striped-weakfish-cynoscion;
 https://spo.nmfs.noaa.gov/content/preliminary-estimate-spawning-
frequency-and-batch-fecundity-striped-weakfish-cynoscion-0;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/macchi_1_0.pdf.

Murphy:1998:DVA

[4162] Michael D. Murphy, Douglas H. Adams, Derek M. Tremain, and Brent L. Winner. Direct validation of ages determined for adult black drum, Pogonias cromis, in east-central Florida, with notes on black drum migration. Fishery Bulletin, 96(2):??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/direct-validation-ages-determined-adult-black-drum-
pogonias-cromis-east-central-florida;   https://spo.nmfs.noaa.
gov/content/direct-validation-ages-determined-adult-black-drum-
pogonias-cromis-east-central-florida-0;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/murphy_0.
pdf.

Sluka:1998:ISF

[4163] Robert D. Sluka and Kathleen M. Sullivan. The influence of spear fishing on species composition and size of groupers on patch reefs in the upper Florida keys. Fishery Bulletin, 96(2):??, 1998. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/influence-spear-fishing-species-
composition-and-size-groupers-patch-reefs-upper-florida;   https:
//spo.nmfs.noaa.gov/content/influence-spear-fishing-species-
composition-and-size-groupers-patch-reefs-upper-florida-0;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/sluka_0.pdf.

Ault:1998:RMA

[4164] Jerald S. Ault, James A. Bohnsack, and Geoffrey A. Meester. A retrospective (1979–1996) multispecies assessment of coral reef fish stocks in the Florida keys. Fishery Bulletin, 96(3):395–414, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/retrospective-1979-
1996-multispecies-assessment-coral-reef-fish-stocks-florida-
keys;   https://spo.nmfs.noaa.gov/content/retrospective-1979-
1996-multispecies-assessment-coral-reef-fish-stocks-florida-
keys-0;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/ault_0.pdf.

Collins:1998:RPS

[4165] L. Alan Collins, Allyn G. Johnson, Christopher C. Koenig, and M. Scott Baker, Jr. Reproductive patterns, sex ratio, and fecundity in gag, Mycteroperca microlepis (Serranidae), a protogynous grouper from the northeastern Gulf of Mexico. Fishery Bulletin, 96(3):??, 1998. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/reproductive-patterns-sex-ratio-
and-fecundity-gag-mycteroperca-microlepis-serranidae;   https:
//spo.nmfs.noaa.gov/content/reproductive-patterns-sex-ratio-
and-fecundity-gag-mycteroperca-microlepis-serranidae-0;   https:

//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
collins_0.pdf.

Gannon:1998:AFH

[4166] Damon P. Gannon, James E. Craddock, and Andrew J. Read. Autumn food habits of harbor porpoises, Phocoena phocoena, in the Gulf of Maine. Fishery Bulletin, 96(3):428–437, 1998. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/autumn-food-habits-
harbor-porpoises-phocoena-phocoena-gulf-maine;   https://spo.
nmfs.noaa.gov/content/autumn-food-habits-harbor-porpoises-phocoena-
phocoena-gulf-maine-0;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/gannon_0.pdf.

Gorfine:1998:CTU

[4167] Harry K. Gorfine, David A. Forbes, and Anne S. Gason. A comparison of two underwater census methods for estimating the abundance of the commercially important blacklip abalone, Haliotis rubra. Fishery Bulletin, 96(3):438–450, 1998. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/comparison-two-underwater-census-
methods-estimating-abundance-commercially-important;   https:
//spo.nmfs.noaa.gov/content/comparison-two-underwater-census-
methods-estimating-abundance-commercially-important-0;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
gorfine_0.pdf.

Jones:1998:AGM

[4168] Cynthia M. Jones and Brian Wells. Age, growth, and mortality of black drum, Pogonias cromis, in the Chesapeake Bay region. Fishery Bulletin, 96(3):??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/age-
growth-and-mortality-black-drum-pogonias-cromis-chesapeake-
bay-region;   https://spo.nmfs.noaa.gov/content/age-growth-
and-mortality-black-drum-pogonias-cromis-chesapeake-bay-region-
0;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/jones_0.pdf.

Kimura:1998:SSM

[4169] Daniel K. Kimura, Allen M. Shimada, and Franklin R. Shaw. Stock structure and movement of tagged sablefish, Anoplopoma fimbria, in offshore northeast Pacific waters and the effects of El Niño–Southern Oscillation on migration and growth. Fishery Bulletin, 96(3):462– 481, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/stock-
structure-and-movement-tagged-sablefish-anoplopoma-fimbria-
offshore-northeast-0;   https://spo.nmfs.noaa.gov/content/stock-
structure-and-movement-tagged-sablefish-anoplopoma-fimbria-
offshore-northeast-pacific;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/kimura_0.pdf.

Liu:1998:AGE

[4170] Kwang-Ming Liu, Po-Jen Chiang, and Che-Tsung Chen. Age and growth estimates of the bigeye thresher shark, Alopias superciliosus, in northeastern Taiwan waters. Fishery Bulletin, 96(3):482–491, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/age-and-growth-estimates-bigeye-
thresher-shark-alopias-superciliosus-northeastern-taiwan;
https://spo.nmfs.noaa.gov/content/age-and-growth-estimates-
bigeye-thresher-shark-alopias-superciliosus-northeastern-taiwan-
0;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/liu_0.pdf.

Love:1998:DNR

[4171] Milton S. Love, Jennifer E. Caselle, and Kevin Herbinson. Declines in nearshore rockfish recruitment and populations in the southern California Bight as measured by impingement rates in coastal electrical power generating stations. Fishery Bulletin, 96(3):492–501, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/declines-nearshore-
rockfish-recruitment-and-populations-southern-california-bight-
0;   https://spo.nmfs.noaa.gov/content/declines-nearshore-rockfish-
recruitment-and-populations-southern-california-bight-measured;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/love_0.pdf.

Lowe:1998:GVG

[4172] Sandra A. Lowe, Donald M. Van Doornik, and Gary A. Winans. Geographic variation in genetic and growth patterns of Atka mackerel, Pleurogrammus monopterygius (Hexagrammidae), in the Aleutian archipelago. Fishery Bulletin, 96(3):502–515, 1998. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/geographic-variation-genetic-and-
growth-patterns-atka-mackerel-pleurogrammus-0;   https://spo.
nmfs.noaa.gov/content/geographic-variation-genetic-and-growth-

patterns-atka-mackerel-pleurogrammus-monopterygius;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
lowe_0.pdf.

Martini:1998:PPH

[4173] Frederic H. Martini, Michael P. Lesser, and John B. Heiser. A population profile for hagfish, Myxine glutinosa, in the Gulf of Maine. Part 2: Morphological variation in populations of Myxine in the North Atlantic Ocean. Fishery Bulletin, 96(3):516–524, 1998. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/population-profile-
hagfish-myxine-glutinosa-gulf-maine-part-2-morphological-variation;
 https://spo.nmfs.noaa.gov/content/population-profile-hagfish-
myxine-glutinosa-gulf-maine-part-2-morphological-variation-
0;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/martini_0.pdf.

Merkouris:1998:LLG

[4174] Susan E. Merkouris, Lisa W. Seeb, and Margaret C. Murphy. Low levels of genetic diversity in highly exploited populations of Alaskan Tanner crabs, Chionoecetes bairdi, and Alaskan and Atlantic snow crabs, C. opilio. Fishery Bulletin, 96(3):525–537, 1998. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/low-levels-genetic-
diversity-highly-exploited-populations-alaskan-tanner-crabs-
0;   https://spo.nmfs.noaa.gov/content/low-levels-genetic-diversity-
highly-exploited-populations-alaskan-tanner-crabs-chionoecetes;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/merkouris_0.pdf.

Munro:1998:DRB

[4175] Peter T. Munro. A decision rule based on the mean square error for correcting relative fishing power differences in trawl survey data. Fishery Bulletin, 96(3):??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/decision-rule-based-mean-square-
error-correcting-relative-fishing-power-differences-trawl;
https://spo.nmfs.noaa.gov/content/decision-rule-based-mean-
square-error-correcting-relative-fishing-power-differences-
trawl-0;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/munro_1_0.pdf.

Nichol:1998:ABS

[4176] Daniel G. Nichol. Annual and between-sex variability of yellowfin sole, Pleuronectes aspe, spring–summer distributions in the eastern Bering Sea. Fishery Bulletin, 96(3):??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/annual-and-between-sex-variability-yellowfin-sole-pleuronectes-
aspe-spring-summer;   https://spo.nmfs.noaa.gov/content/annual-
and-between-sex-variability-yellowfin-sole-pleuronectes-aspe-
spring-summer-0;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/nichol_0.pdf.

Rocha-Olivares:1998:AGM

[4177] Axayacatl Rocha-Olivares. Age, growth, mortality, and population characteristics of the Pacific red snapper, Lutjanus peru, off the southeast coast of Baja California, Mexico. Fishery Bulletin, 96(3):??, 1998. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/age-growth-mortality-and-
population-characteristics-pacific-red-snapper-lutjanus-peru;
 https://spo.nmfs.noaa.gov/content/age-growth-mortality-and-
population-characteristics-pacific-red-snapper-lutjanus-peru-
0;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/rocha-olivares_0.pdf.

Scharf:1998:EDA

[4178] Frederick S. Scharf, Richard M. Yetter, Adam P. Summers, and Francis Juanes. Enhancing diet analyses of piscivorous fishes in the Northwest Atlantic by identification and reconstruction of original prey sizes from ingested remains. Fishery Bulletin, 96(3): ??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/enhancing-
diet-analyses-piscivorous-fishes-northwest-atlantic-identification-
and;   https://spo.nmfs.noaa.gov/content/enhancing-diet-analyses-
piscivorous-fishes-northwest-atlantic-identification-and-0;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/scharf_0.pdf.

Schmid:1998:MTP

[4179] Jeffrey R. Schmid. Marine turtle populations on the west-central coast of Florida: results of tagging studies at Cedar Keys, Florida, 1986–1995. Fishery Bulletin, 96(3):??, 1998. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/marine-turtle-populations-west-
central-coast-florida-results-tagging-studies-cedar-keys;

https://spo.nmfs.noaa.gov/content/marine-turtle-populations-
west-central-coast-florida-results-tagging-studies-cedar-keys-
0;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/schmid_0.pdf.

Secor:1998:EHT

[4180] David H. Secor and Troy E. Gunderson. Effects of hypoxia and temperature on survival, growth, and respiration of juvenile Atlantic sturgeon, Acipenser oxyrinchus. Fishery Bulletin, 96(3):??, 1998. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/effects-hypoxia-and-temperature-
survival-growth-and-respiration-juvenile-atlantic-0;   https:
//spo.nmfs.noaa.gov/content/effects-hypoxia-and-temperature-
survival-growth-and-respiration-juvenile-atlantic-sturgeon;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/secor_0.pdf.

Waldman:1998:MPB

[4181] John R. Waldman, Reese E. Bender, and Isaac I. Wirgin. Multiple population bottlenecks and DNA diversity in populations of wild striped bass, Morone saxatilis. Fishery Bulletin, 96(3):614–620, 1998. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/multiple-population-bottlenecks-
and-dna-diversity-populations-wild-striped-bass-morone;   https:
//spo.nmfs.noaa.gov/content/multiple-population-bottlenecks-
and-dna-diversity-populations-wild-striped-bass-morone-0;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/waldman_0.pdf.

Bejda:1998:CSU

[4182] Allen J. Bejda and Beth A. Phelan. Can scales be used to sex winter flounder, Pleuronectes americanus? Fishery Bulletin, 96(3):??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/can-scales-be-used-
sex-winter-flounder-pleuronectes-americanus;   https://spo.
nmfs.noaa.gov/content/can-scales-be-used-sex-winter-flounder-
pleuronectes-americanus-0;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/bejda_0.pdf.

Gelsleichter:1998:ETO

[4183] James Gelsleichter, Enric Cortés, Charles A. Manire, Robert E. Hueter, and John A. Musick. Evaluation of toxicity of oxytetracycline on growth of captive nurse sharks, Ginglymostoma cirratum. Fishery Bulletin, 96(3):??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/evaluation-toxicity-oxytetracycline-growth-captive-
nurse-sharks-ginglymostoma-cirratum;   https://spo.nmfs.noaa.
gov/content/evaluation-toxicity-oxytetracycline-growth-captive-
nurse-sharks-ginglymostoma-cirratum-0;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/gelsleichter_
0.pdf.

Koeller:1998:EAL

[4184] Peter Koeller and Gregory Crowell. Electrotaxis in American lobsters, Homarus americanus, and its potential use in sampling early benthic-phase animals. Fishery Bulletin, 96(3):??, 1998. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/electrotaxis-american-lobsters-
homarus-americanus-and-its-potential-use-sampling-early;   https:
//spo.nmfs.noaa.gov/content/electrotaxis-american-lobsters-
homarus-americanus-and-its-potential-use-sampling-early-0;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/koeller_0.pdf.

Pepin:1998:CPD

[4185] Pierre Pepin, John F. Dower, and William C. Leggett. Changes in the probability density function of larval fish body length following preservation. Fishery Bulletin, 96(3):633–640, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/changes-probability-density-function-
larval-fish-body-length-following-preservation;   https://spo.
nmfs.noaa.gov/content/changes-probability-density-function-
larval-fish-body-length-following-preservation-0;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
pepin_0.pdf.

Tanabe:1998:SJS

[4186] Toshiyuki Tanabe and Kodo Niu. Sampling juvenile skipjack tuna, Katsuwonus pelamis, and other tunas, Thunnus spp., using midwater trawls in the tropical western Pacific. Fishery Bulletin, 96(3):??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/sampling-juvenile-skipjack-tuna-katsuwonus-pelamis-
and-other-tunas-thunnus-spp-using;   https://spo.nmfs.noaa.
gov/content/sampling-juvenile-skipjack-tuna-katsuwonus-pelamis-

and-other-tunas-thunnus-spp-using-0;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/tanabe_0.pdf.

Wells:1998:EMB

[4187] Randall S. Wells, Suzanne Hofmann, and Tristen L. Moors. Entanglement and mortality of bottlenose dolphins, Tursiops runcatus, in recreational fishing gear in Florida. Fishery Bulletin, 96(3):647–650, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/entanglement-
and-mortality-bottlenose-dolphins-tursiops-runcatus-recreational-
fishing-0;   https://spo.nmfs.noaa.gov/content/entanglement-
and-mortality-bottlenose-dolphins-tursiops-runcatus-recreational-
fishing-gear;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/wells_0_0.pdf.

Anonymous:1998:R

[4188] Anonymous. 1998 reviewers. Fishery Bulletin, 96(4):??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14427;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/reviewers.pdf.

Anonymous:1998:AI

[4189] Anonymous. Annual index. Fishery Bulletin, 96(4):??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14428;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/index%20(9).pdf.

Begg:1998:SDS

[4190] Gavin A. Begg, Mike Cappo, Darren S. Cameron, Steve Boyle, and Michelle J. Sellin. Stock discrimination of school mackerel, Scomberomorus queenslandicus, and spotted mackerel, Scomberomorus munroi, in coastal waters of eastern Australia by analysis of minor and trace elements in whole otoliths. Fishery Bulletin, 96(4):653–666, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/stock-discrimination-school-mackerel-scomberomorus-
queenslandicus-and-spotted-mackerel;   https://spo.nmfs.noaa.
gov/content/stock-discrimination-school-mackerel-scomberomorus-
queenslandicus-and-spotted-mackerel-0;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/begg_0.pdf.

Buencuerpo:1998:PSA

[4191] Valentin Buencuerpo, Santiago Rios, and Julio Morón. Pelagic sharks associated with the swordfish, Xiphias gladius, fishery in the eastern North Atlantic Ocean and the Strait of Gibraltar. Fishery Bulletin, 96 (4):667–685, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/pelagic-
sharks-associated-swordfish-xiphias-gladius-fishery-eastern-
north-atlantic-ocean;   https://spo.nmfs.noaa.gov/content/pelagic-
sharks-associated-swordfish-xiphias-gladius-fishery-eastern-
north-atlantic-ocean-0;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/buencuerpo_0.pdf.

Cooper:1998:MIF

[4192] J. Andrew Cooper and François Chapleau. Monophyly and intrarelationships of the family Pleuronectidae (Pleuronectiformes), with a revised classification. Fishery Bulletin, 96(4):??, 1998. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/monophyly-and-intrarelationships-
family-pleuronectidae-pleuronectiformes-revised;   https://
spo.nmfs.noaa.gov/content/monophyly-and-intrarelationships-
family-pleuronectidae-pleuronectiformes-revised-0;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
cooper_0.pdf.

Cox:1998:DBH

[4193] Tara M. Cox, Andrew J. Read, Susan Barco, Joyce Evans, Damon P. Gannon, Heather N. Koopman, William A. McLellan, Kimberly Murray, John Nicolas, D. Ann Pabst, Charles W. Potter, W. Mark Swingle, Victoria G. Thayer, Kathleen M. Touhey, and Andrew J. Westgate. Documenting the bycatch of harbor porpoises, Phocoena phocoena, in coastal gillnet fisheries from stranded carcasses. Fishery Bulletin, 96(4):727–734, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/documenting-
bycatch-harbor-porpoises-phocoena-phocoena-coastal-gillnet-
fisheries-stranded;   https://spo.nmfs.noaa.gov/content/documenting-
bycatch-harbor-porpoises-phocoena-phocoena-coastal-gillnet-
fisheries-stranded-0;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/cox_0.pdf.

Crabtree:1998:AGR

[4194] Roy E. Crabtree and Lewis H. Bullock. Age, growth, and reproduction of black grouper, Mycteroperca bonaci, in Florida waters. Fishery Bulletin, 96(4):??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/age-growth-and-reproduction-black-grouper-
mycteroperca-bonaci-florida-waters;   https://spo.nmfs.noaa.
gov/content/age-growth-and-reproduction-black-grouper-mycteroperca-
bonaci-florida-waters-0;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/crabtree1_0.pdf.

Crabtree:1998:FHB

[4195] Roy E. Crabtree, Connie Stevens, Derke Snodgrass, and Fredrik J. Stengard. Feeding habits of bonefish, Albula vulpes, from the waters of the Florida keys. Fishery Bulletin, 96(4):??, 1998. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/feeding-habits-bonefish-albula-
vulpes-waters-florida-keys;   https://spo.nmfs.noaa.gov/content/
feeding-habits-bonefish-albula-vulpes-waters-florida-keys-0;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/crabtree2_0.pdf.

Gold:1998:PSG

[4196] John R. Gold and Linda R. Richardson. Population structure in greater amberjack, Seriola dumerili, from the Gulf of Mexico and the western Atlantic Ocean. Fishery Bulletin, 96(4):??, 1998. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/population-structure-
greater-amberjack-seriola-dumerili-gulf-mexico-and-western-
atlantic;   https://spo.nmfs.noaa.gov/content/population-structure-
greater-amberjack-seriola-dumerili-gulf-mexico-and-western-
atlantic-0;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/gold_0.pdf.

Jackson:1998:AGR

[4197] George D. Jackson and Vicki A. Wadley. Age, growth, and reproduction of the tropical squid Nototodarus hawaiiensis (Cephalopoda: Ommastrephidae) off the North West Slope of Australia. Fishery Bulletin, 96(4): ??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/age-growth-
and-reproduction-tropical-squid-nototodarus-hawaiiensis-cephalopoda;
 https://spo.nmfs.noaa.gov/content/age-growth-and-reproduction-
tropical-squid-nototodarus-hawaiiensis-cephalopoda-0;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
jackson_0.pdf.

Laidig:1998:DPL

[4198] Thomas E. Laidig and Keith M. Sakuma. Description of pelagic larval and juvenile grass rockfish, Sebastes rastrefliger (family Scorpaenidae), with an examination of age and growth. Fishery Bulletin, 96(4):??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/description-pelagic-
larval-and-juvenile-grass-rockfish-sebastes-rastrefliger-family;
 https://spo.nmfs.noaa.gov/content/description-pelagic-larval-
and-juvenile-grass-rockfish-sebastes-rastrefliger-family-0;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/laidig_0_0.pdf.

McGovern:1998:CSR

[4199] John C. McGovern, David M. Wyanski, Oleg Pashuk, Charles S. Manooch II, and George R. Sedberry. Changes in the sex ratio and size at maturity of gag, Mycteroperca microlepis, from the Atlantic coast of the southeastern United States during 1976–1995. Fishery Bulletin, 96(4): 797–807, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/changes-
sex-ratio-and-size-maturity-gag-mycteroperca-microlepis-atlantic-
coast-0;   https://spo.nmfs.noaa.gov/content/changes-sex-ratio-
and-size-maturity-gag-mycteroperca-microlepis-atlantic-coast-
southeastern;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/mcgovern_0.pdf.

Lomond:1998:TPB

[4200] Tammy M. Lomond, David C. Schneider, and David A. Methven. Transition from pelagic to benthic prey for age group 0–1 Atlantic cod, Gadus morhua. Fishery Bulletin, 96(4):908–911, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/transition-pelagic-benthic-prey-age-group-
0-1-atlantic-cod-gadus-morhua;   https://spo.nmfs.noaa.gov/
content/transition-pelagic-benthic-prey-age-group-0-1-atlantic-
cod-gadus-morhua-0;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/lomond_0.pdf.

Raum-Suryan:1998:DAH

[4201] Kimberly L. Raum-Suryan and James T. Harvey. Distribution and abundance of and habitat use by harbor porpoise, Phocoena phocoena, off the northern San Juan Islands, Washington. Fishery Bulletin, 96(4):??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/

distribution-and-abundance-and-habitat-use-harbor-porpoise-
phocoena-phocoena-northern-san;   https://spo.nmfs.noaa.gov/
content/distribution-and-abundance-and-habitat-use-harbor-porpoise-
phocoena-phocoena-northern-san-0;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/raum-suryan_0.pdf.

Scoles:1998:GPM

[4202] Daniel R. Scoles, Bruce B. Collette, and John E. Graves. Global phylogeography of mackerels of the genus Scomber. Fishery Bulletin, 96(4): 823–842, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/global-
phylogeography-mackerels-genus-scomber;   https://spo.nmfs.
noaa.gov/content/global-phylogeography-mackerels-genus-scomber-
0.

Small:1998:PSS

[4203] Maureen P. Small, Ruth Withler, and Terry D. Beacham. Population structure and stock identification of British Columbia Coho salmon, Oncorhynchus kisutch, based on microsatellite DNA variation. Fishery Bulletin, 96(4):??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
population-structure-and-stock-identification-british-columbia-
coho-salmon-oncorhynchus;   https://spo.nmfs.noaa.gov/content/
population-structure-and-stock-identification-british-columbia-
coho-salmon-oncorhynchus-0;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/small_0.pdf.

Sol:1998:GDA

[4204] Sean Y. Sol, O. Paul Olson, Daniel P. Lomax, and Lyndal L. Johnson. Gonadal development and associated changes in plasma reproductive steroids in English sole, Pleuronectes vetulus, from Puget Sound, Washington. Fishery Bulletin, 96(4):??, 1998. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/gonadal-development-
and-associated-changes-plasma-reproductive-steroids-english-
sole;   https://spo.nmfs.noaa.gov/content/gonadal-development-
and-associated-changes-plasma-reproductive-steroids-english-
sole-0;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/sol_0.pdf.

Somerton:1998:PGK

[4205] David A. Somerton and William Donaldson. Parasitism of the golden king crab, Lithodes aequispinus, by two species of snailfish, genus Careproctus. Fishery Bulletin, 96(4):??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/parasitism-golden-king-crab-lithodes-
aequispinus-two-species-snailfish-genus-careproctus;   https:
//spo.nmfs.noaa.gov/content/parasitism-golden-king-crab-lithodes-
aequispinus-two-species-snailfish-genus-careproctus-0;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
somerton_1_0.pdf.

Stoner:1998:SRQ

[4206] Allan W. Stoner, Melody Ray-Culp, and Sheila M. O’Connell. Settlement and recruitment of queen conch, Strombus gigas, in seagrass meadows: associations with habitat and micropredators. Fishery Bulletin, 96(4): ??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/settlement-
and-recruitment-queen-conch-strombus-gigas-seagrass-meadows-
associations-0;   https://spo.nmfs.noaa.gov/content/settlement-
and-recruitment-queen-conch-strombus-gigas-seagrass-meadows-
associations-habitat;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/stoner_0.pdf.

Watanabe:1998:GTL

[4207] Yoshiro Watanabe and Motohiko Nakamura. Growth trajectory of the larval Japanese sardine, Sardinops melanostictus, transported into the Pacific coastal waters off central Japan. Fishery Bulletin, 96(4):??, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/growth-trajectory-larval-japanese-sardine-
sardinops-melanostictus-transported-pacific;   https://spo.
nmfs.noaa.gov/content/growth-trajectory-larval-japanese-sardine-
sardinops-melanostictus-transported-pacific-0;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/watanabe_
0.pdf.

Moles:1998:MPP

[4208] Adam Moles, Jonathan Heifetz, and David C. Love. Metazoan parasites as potential markers for selected Gulf of Alaska rockfishes. Fishery Bulletin, 96(4):912–916, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/metazoan-parasites-potential-markers-selected-gulf-
alaska-rockfishes;   https://spo.nmfs.noaa.gov/content/metazoan-
parasites-potential-markers-selected-gulf-alaska-rockfishes-

0;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/moles_0.pdf.

Somarakis:1998:CRL

[4209] Stylianos Somarakis, Barbara Catalano, and Nikolaos Tsimenides. Catchability and retention of larval European anchovy, Engraulis enerasicolus, with bongo nets. Fishery Bulletin, 96(4):917–923, 1998. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/catchability-
and-retention-larval-european-anchovy-engraulis-enerasicolus-
bongo-nets;   https://spo.nmfs.noaa.gov/content/catchability-
and-retention-larval-european-anchovy-engraulis-enerasicolus-
bongo-nets-0;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/somarakis_0.pdf.

Broadhurst:1999:FRE

[4210] Matt K. Broadhurst, Steven J. Kennelly, and Steve Eayrs. Flowrelated effects in prawn-trawl codends: potential for increasing the escape of unwanted fish through square-mesh panels. Fishery Bulletin, 97(1):1–8, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/flow-related-effects-prawn-trawl-codends-potential-
increasing-escape-unwanted-fish-through;   https://spo.nmfs.
noaa.gov/sites/default/files/01broadh.pdf.

Brodziak:1999:AEE

[4211] Jon Brodziak and Lisa Hendrickson. An analysis of environmental effects on survey catches of squids Loligo pealei and Illex illecebrosus in the northwest Atlantic. Fishery Bulletin, 97(1):??, 1999. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/analysis-environmental-effects-
survey-catches-squids-loligo-pealei-and-illex-illecebrosus;
 https://spo.nmfs.noaa.gov/sites/default/files/02brodzi.pdf.

Chen:1999:MDE

[4212] Yong Chen and Steve S. Montgomery. Modeling the dynamics of eastern rock lobster, Jasus verreauxi, stock in New South Wales, Australia. Fishery Bulletin, 97(1):??, 1999. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/modeling-dynamics-eastern-rock-
lobster-jasus-verreauxi-stock-new-south-wales-australia;   https:
//spo.nmfs.noaa.gov/sites/default/files/03chenfi.pdf.

Deree:1999:AGD

[4213] Heather L. Deree. Age and growth, dietary habits, and parasitism of the fourbeard rockling, Enchelyopus cimbrius, from the Gulf of Maine. Fishery Bulletin, 97(1):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
age-and-growth-dietary-habits-and-parasitism-fourbeard-rockling-
enchelyopus-cimbrius-gulf;   https://spo.nmfs.noaa.gov/sites/
default/files/04dereef.pdf.

Heist:1999:GIS

[4214] Edward J. Heist and John R. Gold. Genetic identification of sharks in the U.S. Atlantic large coastal shark fishery. Fishery Bulletin, 97(1):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/genetic-identification-sharks-us-atlantic-
large-coastal-shark-fishery;   https://spo.nmfs.noaa.gov/sites/
default/files/05heistf.pdf.

Koeller:1999:ITE

[4215] Peter Koeller. Influence of temperature and effort on lobster catches at different temporal and spatial scales and the implications for stock assessments. Fishery Bulletin, 97(1):??, 1999. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/influence-temperature-and-effort-
lobster-catches-different-temporal-and-spatial-scales-and;
https://spo.nmfs.noaa.gov/sites/default/files/06koelle.pdf.

Lantry:1999:ETB

[4216] Brian F. Lantry, Donald J. Stewart, Peter S. Rand, and Edward L. Mills. Evaluation of total-body electrical conductivity to estimate wholebody water content of yellow perch, Perca flavescens, and alewife, Alosa pseudoharengus. Fishery Bulletin, 97(1):??, 1999. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/evaluation-total-body-electrical-
conductivity-estimate-whole-body-water-content-yellow;   https:
//spo.nmfs.noaa.gov/sites/default/files/07lantry.pdf.

Limburg:1999:GMR

[4217] Karin E. Limburg, Michael L. Pace, and Kristin K. Arend. Growth, mortality, and recruitment of larval Morone spp. in relation to food availability and temperature in the Hudson River. Fishery Bulletin, 97(1):80–91, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/growth-mortality-and-recruitment-larval-morone-spp-
relation-food-availability-and;   https://spo.nmfs.noaa.gov/
sites/default/files/08limbur.pdf.

Lorenzo:1999:BDB

[4218] José M. Lorenzo and José G. Pajuelo. Biology of a deep benthopelagic fish, roudi escolar Promethichthys prometheus (Gempylidae), off the Canary Islands. Fishery Bulletin, 97(1):??, 1999. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/biology-deep-benthopelagic-fish-
roudi-escolar-promethichthys-prometheus-gempylidae-canary;
https://spo.nmfs.noaa.gov/sites/default/files/09lorenz.pdf.

Love:1999:ALH

[4219] Milton S. Love and Korie Johnson. Aspects of the life histories of grass rockfish, Sebastes rastrelliger, and brown rockfish, S. auriculatus, from southern California. Fishery Bulletin, 97(1):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/aspects-life-
histories-grass-rockfish-sebastes-rastrelliger-and-brown-rockfish-
s;   https://spo.nmfs.noaa.gov/sites/default/files/10lovefi.
pdf.

McMillan:1999:TNS

[4220] Charmion B. McMillan. Three new species of hagfish (Myxinidae, Eptatretus) from the Galápagos Islands. Fishery Bulletin, 97(1): ??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/three-new-
species-hagfish-myxinidae-eptatretus-gal%C3%A1pagos-islands;
 https://spo.nmfs.noaa.gov/sites/default/files/11mcmill.pdf.

Pollard:1999:CMJ

[4221] Morgan J. Pollard, Michael J. Kingsford, and Stephen C. Battaglene. Chemical marking of juvenile snapper, Pagrus auratus (Sparidae), by incorporation of strontium into dorsal spines. Fishery Bulletin, 97(1): 118–131, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/chemical-
marking-juvenile-snapper-pagrus-auratus-sparidae-incorporation-
strontium-dorsal;   https://spo.nmfs.noaa.gov/sites/default/
files/12pollar.pdf.

Polovina:1999:ATP

[4222] Jeffrey J. Polovina, Pierre Kleiber, and Donald R. Kobayashi. Application of TOPEX-POSEIDON satellite altimetry to simulate transport dynamics of larvae of spiny lobster, Panulirus marginatus, in the Northwestern Hawaiian Islands, 1993–1996. Fishery Bulletin, 97(1):132–143, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/application-topex-poseidon-satellite-altimetry-
simulate-transport-dynamics-larvae-spiny;   https://spo.nmfs.
noaa.gov/sites/default/files/13polovi.pdf.

Poole:1999:PSA

[4223] David Poole, Geof H. Givens, and Adrian E. Raftery. A proposed stock assessment method and its application to bowhead whales, Balaena mysticetus. Fishery Bulletin, 97(1):144–152, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/proposed-stock-assessment-method-and-
its-application-bowhead-whales-balaena-mysticetus;   https:
//spo.nmfs.noaa.gov/sites/default/files/14poolef.pdf.

Wintner:1999:AGD

[4224] Sabine P. Wintner and Geremy Cliff. Age and growth determination of the white shark, Carcharodon carcharias, from the east coast of South Africa. Fishery Bulletin, 97(1):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/age-and-growth-determination-white-shark-carcharodon-
carcharias-east-coast-south-africa;   https://spo.nmfs.noaa.
gov/sites/default/files/15winter.pdf.

Xiao:1999:EIR

[4225] Yongshun Xiao, Lauren P. Brown, Terence I. Walker, and André E. Punt. Estimation of instantaneous rates of tag shedding for school shark, Galeorhinus galeus, and gummy shark, Mustelus antarcticus, by conditional likelihood. Fishery Bulletin, 97(1):170–184, 1999. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/estimation-instantaneous-rates-
tag-shedding-school-shark-galeorhinus-galeus-and-gummy-shark;
 https://spo.nmfs.noaa.gov/sites/default/files/16xiaofi.pdf.

Kennelly:1999:ADT

[4226] Steven J. Kennelly. Areas, depths, and times of high rates of scup, Stenotomus chrysops, discard by demersal fish trawling off the northeastern United States. Fishery Bulletin, 97(1):

185–192, 1999.  CODEN FSYBAY.  ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
areas-depths-and-times-high-rates-scup-stenotomus-chrysops-
discard-dermsal-fish-trawling;   https://spo.nmfs.noaa.gov/
sites/default/files/17kennel.pdf.

Meyer:1999:ELB

[4227] David L. Meyer, Mark S. Fonseca, Patricia L. Murphey, Robert H. McMichael, Jr., Michael M. Byerly, Michael W. LaCroix, Paula E. Whitfield, and Gordon W. Thayer. Effects of live-bait shrimp trawling on seagrass beds and fish bycatch in Tampa Bay, Florida. Fishery Bulletin, 97 (1):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/effects-live-
bait-shrimp-trawling-seagrass-beds-and-fish-bycatch-tampa-bay-
florida;   https://spo.nmfs.noaa.gov/sites/default/files/18meyerf.
pdf.

Witzell:1999:DRA

[4228] Wayne N. Witzell. Distribution and relative abundance of sea turtles caught incidentally by the U.S. pelagic longline fleet in the western North Atlantic Ocean, 1992–1995. Fishery Bulletin, 97(1):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/distribution-
and-relative-abundance-sea-turtles-caught-incidentally-us-pelagic-
longline;   https://spo.nmfs.noaa.gov/sites/default/files/19witzel.
pdf.

Cooper:1999:DIM

[4229] Andrew B. Cooper and Marc Mangel. The dangers of ignoring metapopulation structure for the conservation of salmonids. Fishery Bulletin, 97 (2):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/dangers-
ignoring-metapopulation-structure-conservation-salmonids;
https://spo.nmfs.noaa.gov/sites/default/files/01cooper.pdf.

Francis:1999:AGE

[4230] Malcolm P. Francis, Kevin P. Mulligan, Nick M. Davies, and Michael P. Beentjes. Age and growth estimates for New Zealand hapuku, Polyprion oxygeneios. Fishery Bulletin, 97(2):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/age-and-growth-estimates-new-zealand-
hapuku-polyprion-oxygeneios;   https://spo.nmfs.noaa.gov/sites/
default/files/02franci.pdf.

Goodyear:1999:APU

[4231] C. Phillip Goodyear. An analysis of the possible utility of time-area closures to minimize billfish bycatch by U.S. pelagic longlines. Fishery Bulletin, 97(2):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
analysis-possible-utility-time-area-closures-minimize-billfish-
bycatch-us-pelagic-longlines;   https://spo.nmfs.noaa.gov/sites/
default/files/03goodye.pdf.

Heifetz:1999:AVA

[4232] Jonathan Heifetz, Delsa Anderl, Nancy E. Maloney, and Thomas L. Rutecki. Age validation and analysis of ageing error from marked and recaptured sablefish, Anoplopoma fimbria. Fishery Bulletin, 97 (2):256–263, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/age-
validation-and-analysis-ageing-error-marked-and-recaptured-
sablefish-anoplopoma-fimbria;   https://spo.nmfs.noaa.gov/sites/
default/files/04heifet.pdf.

Krieger:1999:DAS

[4233] Kenneth J. Krieger and Daniel H. Ito. Distribution and abundance of shortraker rockfish, Sebastes borealis, and rougheye rockfish, S. aleutianus, determined from a manned submersible. Fishery Bulletin, 97(2):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/distribution-and-abundance-shortraker-rockfish-
sebastes-borealis-and-rougheye-rockfish-s;   https://spo.nmfs.
noaa.gov/sites/default/files/05kriege.pdf.

MacFarlane:1999:NDD

[4234] R. Bruce MacFarlane and Elizabeth C. Norton. Nutritional dynamics during embryonic development in the viviparous genus Sebastes and their application to the assessment of reproductive success. Fishery Bulletin, 97(2):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
nutritional-dynamics-during-embryonic-development-viviparous-
genus-sebastes-and-their;   https://spo.nmfs.noaa.gov/sites/
default/files/06macfar.pdf.

Macy:1999:MRR

[4235] William K. Macy III, Ann G. Durbin, and Edward G. Durbin. Metabolic rate in relation to temperature and swimming speed, and the cost of filter feeding in Atlantic menhaden, Brevoortia tyrannus. Fishery Bulletin, 97(2):282–293, 1999. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
metabolic-rate-relation-temperature-and-swimming-speed-and-
cost-filter-feeding-atlantic;   https://spo.nmfs.noaa.gov/sites/
default/files/07macyfi.pdf.

Miller:1999:RBO

[4236] Thomas J. Miller, Tomasz Herra, and William C. Leggett. The relation between otolith size and larval size at hatching for Atlantic cod, Gadus morhua. Fishery Bulletin, 97(2):204–305, 1999. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/relation-between-otolith-
size-and-larval-size-hatching-atlantic-cod-gadus-morhua;   https:
//spo.nmfs.noaa.gov/sites/default/files/08miller.pdf.

Quinonez-Velazquez:1999:AVG

[4237] Casimiro Quiñonez-Velázquez. Age validation and growth of larval and juvenile haddock, Melanogrammus aeglefinus, and pollock, Pollachius virens, on the Scotian Shelf. Fishery Bulletin, 97(2):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/age-validation-and-
growth-larval-and-juvenile-haddock-melanogrammus-aeglefinus-
and-pollock;   https://spo.nmfs.noaa.gov/sites/default/files/
09quione.pdf.

Stanley:1999:DVM

[4238] Richard D. Stanley, Robert Keiser, Bruce M. Leaman, and Ken D. Cooke. Diel vertical migration by yellowtail rockfish, Sebastes flavidus, and its impact on acoustic biomass estimation. Fishery Bulletin, 97(2): ??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/diel-vertical-
migration-yellowtail-rockfish-sebastes-flavidus-and-its-impact-
acoustic;   https://spo.nmfs.noaa.gov/sites/default/files/10stanle.
pdf.

Suryan:1999:VRP

[4239] Robert M. Suryan and James T. Harvey. Variability in reactions of Pacific harbor seals, Phoca vitulinarichardsi, to disturbance. Fishery Bulletin, 97(2):332–339, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/variability-reactions-pacific-harbor-

seals-phoca-vitulinarichardsi-disturbance;   https://spo.nmfs.
noaa.gov/sites/default/files/11suryan.pdf.

Terwilliger:1999:AGL

[4240] Mark. R. Terwilliger and Thomas A. Munroe. Age, growth, longevity, and mortality of blackcheek tonguefish, Symphurus plagiusa (Cynoglossidae: Pleuronectiformes), in Chesapeake Bay, Virginia. Fishery Bulletin, 97(2):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/age-
growth-longevity-and-mortality-blackcheek-tonguefish-symphurus-
plagiusa-cynoglossidae;   https://spo.nmfs.noaa.gov/sites/default/
files/12terwil.pdf.

Thompson:1999:ADG

[4241] Bruce A. Thompson, Marty Beasley, and Charles A. Wilson. Age distribution and growth of greater amberjack, Seriola dumerili, from the north-central Gulf of Mexico. Fishery Bulletin, 97(2):362– 371, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/age-
distribution-and-growth-greater-amberjack-seriola-dumerili-
north-central-gulf-mexico;   https://spo.nmfs.noaa.gov/sites/
default/files/13thomps.pdf.

Adams:1999:MLF

[4242] Douglas H. Adams and Robert H. McMichael, Jr. Mercury levels in four species of sharks from the Atlantic coast of Florida. Fishery Bulletin, 97 (2):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/mercury-
levels-four-species-sharks-atlantic-coast-florida;   https:
//spo.nmfs.noaa.gov/sites/default/files/14adamsf.pdf.

Beal:1999:USM

[4243] Brian F. Beal, Robert Bayer, M. Gayle Kraus, and Samuel R. Chapman. A unique shell marker in juvenile, hatchery-reared individuals of the softshell clam, Mya arenaria L. Fishery Bulletin, 97(2):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/unique-shell-
marker-juvenile-hatchery-reared-individuals-softshell-clam-
mya-arenaria-l;   https://spo.nmfs.noaa.gov/sites/default/files/
15bealfi.pdf.

Carlson:1999:ONJ

[4244] John K. Carlson. Occurrence of neonate and juvenile sandbar sharks, Carcharhinus plumbeus, in the northeastern Gulf of Mexico. Fishery Bulletin, 97(2):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/occurrence-neonate-and-juvenile-sandbar-sharks-
carcharhinus-plumbeus-northeastern-gulf;   https://spo.nmfs.
noaa.gov/sites/default/files/16carlso.pdf.

Holland:1999:DRT

[4245] Kim N. Holland, Pierre Kleiber, and Stephen M. Kajiura. Different residence times of yellowfin tuna, Thunnus albacares, and bigeye tuna, T. obesus, found in mixed aggregations over a seamount. Fishery Bulletin, 97(2):392–395, 1999. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
different-residence-times-yellowfin-tuna-thunnus-albacares-
and-bigeye-tuna-t-obesus-found;   https://spo.nmfs.noaa.gov/
sites/default/files/17hollan.pdf.

ONeill:1999:ASC

[4246] Michael F. O’Neill, David J. Die, Brian R. Taylor, and Malcolm J. Faddy. Accuracy of at-sea commercial size grading of tiger prawns (Penaeuse sculentus and P. semisulcatus) in the Australian northern prawn fishery. Fishery Bulletin, 97(2):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/accuracy-sea-commercial-size-grading-
tiger-prawns-penaeuse-sculentus-and-p-semisulcatus;   https:
//spo.nmfs.noaa.gov/sites/default/files/18oneill.pdf.

Vidal:1999:DTP

[4247] Erica A. G. Vidal and Manuel Haimovici. Digestive tract parasites in rhynchoteuthion squid paralarvae, particularly in Illex argentinus, (cephalopoda: Ommastrephidae). Fishery Bulletin,

97(2):??, 1999.  CODEN FSYBAY.  ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
digestive-tract-parasites-rhynchoteuthion-squid-paralarvae-
particularly-illex-argentinus;   https://spo.nmfs.noaa.gov/
sites/default/files/19vidalf.pdf.

Yang:1999:DPS

[4248] Mei-Sun Yang and Benjamin N. Page. Diet of Pacific sleeper shark, Somniosus pacificus, in the the Gulf of Alaska. Fishery Bulletin, 97(2):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
diet-pacific-sleeper-shark-somniosus-pacificus-gulf-alaska;
 https://spo.nmfs.noaa.gov/sites/default/files/20yangfi.pdf.

Austin:1999:MSA

[4249] Herbert M. Austin, Daniel Scoles, and Allison J. Abell. Morphometric separation of annual cohorts within mid-Atlantic bluefish, Pomatomus saltatrix, using discriminant function analysis. Fishery Bulletin, 97(3):411–420, 1999. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
morphometric-separation-annual-cohorts-within-mid-atlantic-
bluefish-pomatomus-saltatrix;   https://spo.nmfs.noaa.gov/sites/
default/files/01austin.pdf.

Bertignac:1999:EER

[4250] Michel Bertignac, John Hampton, and Atilio L. Coan, Jr. Estimates of exploitation rates for north Pacific albacore, Thunnus alalunga, from tagging data. Fishery Bulletin, 97(3):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/estimates-exploitation-rates-north-pacific-
albacore-thunnus-alalunga-tagging-data;   https://spo.nmfs.
noaa.gov/sites/default/files/02bertig.pdf.

Broadhurst:1999:USC

[4251] Matt K. Broadhurst, Roger B. Larsen, Steven J. Kennelly, and Paul E. McShane. Use and success of composite square-mesh codends in reducing bycatch and in improving size-selectivity of prawns in Gulf St. Vincent, South Australia. Fishery Bulletin, 97(3):??, 1999. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/use-and-success-composite-square-
mesh-codends-reducing-bycatch-and-improving-size;   https:/
/spo.nmfs.noaa.gov/sites/default/files/03broadh.pdf.

DeMartini:1999:MCE

[4252] Edward E. DeMartini and Boulderson B. Lau. Morphometric criteria for estimating sexual maturity in two snappers, Etelis carbunculus and Pristipomoides sieboldii. Fishery Bulletin, 97(3):??, 1999. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/morphometric-criteria-estimating-
sexual-maturity-two-snappers-etelis-carbunculus-and;   https:
//spo.nmfs.noaa.gov/sites/default/files/04demart.pdf.

Franks:1999:AGC

[4253] James S. Franks, James R. Warren, and Michael V. Buchanan. Age and growth of cobia, Rachycentron canadum, from the northeastern Gulf of Mexico. Fishery Bulletin, 97(3):459–471, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/age-and-growth-cobia-rachycentron-canadum-
northeastern-gulf-mexico;   https://spo.nmfs.noaa.gov/sites/
default/files/05franks.pdf.

Friedland:1999:GPP

[4254] Kevin D. Friedland, Jean-Denis Dutil, and Teresa Sadusky. Growth patterns in postsmolts and the nature of the marine juvenile nursery for Atlantic salmon, Salmo salar. Fishery Bulletin, 97(3):472– 481, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/growth-
patterns-postsmolts-and-nature-marine-juvenile-nursery-atlantic-
salmon-salmo-salar;   https://spo.nmfs.noaa.gov/sites/default/
files/06friedl.pdf.

Hannah:1999:NMI

[4255] Robert W. Hannah. A new method for indexing spawning stock and recruitment in ocean shrimp, Pandalus jordani, and preliminary evidence for a stock-recruitment relationship. Fishery Bulletin, 97(3):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/new-method-indexing-
spawning-stock-and-recruitment-ocean-shrimp-pandalus-jordani-
and;   https://spo.nmfs.noaa.gov/sites/default/files/07hannah.
pdf.

Labropoulou:1999:HSD

[4256] Mary Labropoulou, Athanasios Machias, and Nikolaos Tsimenides. Habitat selection and diet of juvenile red porgy, Pagrus pagrus (Linnaeus, 1758). Fishery Bulletin, 97(3):495–507, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/habitat-selection-and-diet-juvenile-red-porgy-
pagrus-pagrus-linnaeus-1758;   https://spo.nmfs.noaa.gov/sites/
default/files/08labrop.pdf.

Lindeman:1999:NHF

[4257] Kenyon C. Lindeman and David B. Snyder. Nearshore hardbottom fishes of southeast Florida and effects of habitat burial caused by dredging. Fishery Bulletin, 97(3):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/nearshore-hardbottom-fishes-southeast-florida-and-effects-
habitat-burial-caused-dredging;   https://spo.nmfs.noaa.gov/
sites/default/files/09lindem.pdf.

Nates:1999:GMG

[4258] Sergio F. Nates and Darryl L. Felder. Growth and maturation of the ghost shrimp Lepidophthalmus sinuensis Lemaitre and Rodrigues, 1991 (Crustacea, Decapoda, Callianassidae), a burrowing pest in penaeid shrimp culture ponds. Fishery Bulletin, 97(3):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/growth-and-
maturation-ghost-shrimp-lepidophthalmus-sinuensis-lemaitre-
and-rodrigues-1991;   https://spo.nmfs.noaa.gov/sites/default/
files/10natesf.pdf.

Perkins:1999:CRF

[4259] Peter C. Perkins and Elizabeth F. Edwards. Capture rate as a function of school size in pantropical spotted dolphins, Stenella attenuata, in the eastern tropical Pacific Ocean. Fishery Bulletin, 97(3):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/capture-rate-
function-school-size-pantropical-spotted-dolphins-stenella-
attenuata-eastern;   https://spo.nmfs.noaa.gov/sites/default/
files/11perkin.pdf.

Rilling:1999:RTV

[4260] Gene C. Rilling and Edward D. Houde. Regional and temporal variability in growth and mortality of bay anchovy, Anchoa mitchilli, larvae in Chesapeake Bay. Fishery Bulletin, 97(3):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/regional-and-
temporal-variability-growth-and-mortality-bay-anchovy-anchoa-
mitchilli-larvae;   https://spo.nmfs.noaa.gov/sites/default/
files/12rillin.pdf.

Roa:1999:ESS

[4261] Rubén Roa, Billy Ernst, and Fabián Tapia. Estimation of size at sexual maturity: an evaluation of analytical and resampling procedures. Fishery Bulletin, 97(3):570–580, 1999. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/estimation-size-sexual-maturity-

evaluation-analytical-and-resampling-procedures;   https://
spo.nmfs.noaa.gov/sites/default/files/13roafis.pdf.

Rooker:1999:STV

[4262] Jay R. Rooker, Scott A. Holt, G. Joan Holt, and Lea A. Fuiman. Spatial and temporal variability in growth, mortality, and recruitment potential of postsettlement red drum, Sciaenops ocellatus, in a subtropical estuary. Fishery Bulletin, 97(3):??, 1999. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/spatial-and-temporal-variability-
growth-mortality-and-recruitment-potential-postsettlement;
https://spo.nmfs.noaa.gov/sites/default/files/14rooker.pdf.

Sigler:1999:ESA

[4263] Michael F. Sigler. Estimation of sablefish, Anoplopoma fimbria, abundance off Alaska with an age-structured population model. Fishery Bulletin, 97(3):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
estimation-sablefish-anoplopoma-fimbria-abundance-alaska-age-
structured-population-model;   https://spo.nmfs.noaa.gov/sites/
default/files/15sigler.pdf.

Nowlis:1999:FBO

[4264] Joshua Sladek Nowlis and Callum M. Roberts. Fisheries benefits and optimal design of marine reserves. Fishery Bulletin, 97(3): ??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fisheries-
benefits-and-optimal-design-marine-reserves;   https://spo.
nmfs.noaa.gov/sites/default/files/16sladek.pdf.

Somerton:1999:NES

[4265] David A. Somerton and Robert S. Otto. Net efficiency of a survey trawl for snow crab, Chionoecetes opilio, and Tanner crab, C. bairdi. Fishery Bulletin, 97(3):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/net-efficiency-survey-trawl-snow-
crab-chionoecetes-opilio-and-tanner-crab-c-bairdi;   https:
//spo.nmfs.noaa.gov/sites/default/files/17somert.pdf.

Szedlmayer:1999:NHG

[4266] Stephen T. Szedlmayer and Joseph Conti. Nursery habitats, growth rates, and seasonality of age-0 red snapper, Lutjanus campechanus, in the northeast Gulf of Mexico. Fishery Bulletin, 97(3):??, 1999. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/nursery-habitats-growth-rates-and-
seasonality-age-0-red-snapper-lutjanus-campechanus;   https:
//spo.nmfs.noaa.gov/sites/default/files/18szedlm.pdf.

Thomas:1999:OCS

[4267] Ross Thomas and Natalie A. Moltschaniwskyj. Ontogenetic changes in size and shape of statoliths: implications for age and growth of the short-lived tropical squid Sepioteuthis lessoniana (Cephalopoda: Loliginidae). Fishery Bulletin, 97(3):??, 1999. CODEN FSY-

BAY.  ISSN 0090-0656 (print), 1937-4518 (electronic).  URL

https://spo.nmfs.noaa.gov/content/ontogenetic-changes-size-
and-shape-statoliths-implications-age-and-growth-short-lived;
 https://spo.nmfs.noaa.gov/sites/default/files/19thomas.pdf.

Valle:1999:DHU

[4268] Charles F. Valle, John W. O’Brien, and Kris B. Wiese. Differential habitat use by California halibut, Paralichthys californicus, barred sand bass, Paralabrax nebulifer, and other juvenile fishes in Alamitos Bay, California. Fishery Bulletin, 97(3):646–660, 1999. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/differential-habitat-use-
california-halibut-paralichthys-californicus-barred-sand-bass;
 https://spo.nmfs.noaa.gov/sites/default/files/20vallef.pdf.

VanEeckhaute:1999:MHM

[4269] Lutgarde A. M. Van Eeckhaute, Stratis Gavaris, and Edward A. Trippel. Movements of haddock, Melanogrammus aeglefinus, on eastern Georges Bank determined from a population model incorporating temporal and spatial detail. Fishery Bulletin, 97 (3):661–679, 1999. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
movements-haddock-melanogrammus-aeglefinus-eastern-georges-
bank-determined-population-model;   https://spo.nmfs.noaa.gov/
sites/default/files/21vaneec.pdf.

Woodbury:1999:RGO

[4270] David Woodbury. Reduction of growth in otoliths of widow and yellowtail rockfish (Sebastes entomelas and S. flavidus) during the 1983 El Niño. Fishery Bulletin, 97(3):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/reduction-growth-otoliths-widow-and-yellowtail-

rockfish-sebastes-entomelas-and-s-flavidus;   https://spo.nmfs.
noaa.gov/sites/default/files/22woodbu.pdf.

Xiao:1999:GAT

[4271] Yongshun Xiao. General age- and time-dependent growth models for animals. Fishery Bulletin, 97(3):??, 1999. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/general-age-and-time-dependent-
growth-models-animals;   https://spo.nmfs.noaa.gov/sites/default/
files/23xiaofi.pdf.

Ye:1999:AVC

[4272] Yimin Ye and Hussain M. A. Mohammed. An analysis of variation in catchability of green tiger prawn, Penaeus semisulcatus, in waters off Kuwait. Fishery Bulletin, 97(3):??, 1999. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/analysis-variation-
catchability-green-tiger-prawn-penaeus-semisulcatus-waters-
kuwait;   https://spo.nmfs.noaa.gov/sites/default/files/24yefish.
pdf.

Griffith:1999:SAS

[4273] Jason N. Griffith, Andrew P. Hendry, and Thomas P. Quinn. Straying of adult sockeye salmon, Oncorhynchus nerka, entering a non-natal hatchery. Fishery Bulletin, 97(3):713–716, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/straying-adult-sockeye-salmon-oncorhynchus-
nerka-entering-non-natal-hatchery;   https://spo.nmfs.noaa.
gov/sites/default/files/25griffi.pdf.

Hunt:1999:OFH

[4274] Sharon L. Hunt, Timothy J. Mulligan, and Kenichiro Komori. Oceanic feeding habits of Chinook salmon, Oncorhynchus tshawytscha, off northern California. Fishery Bulletin, 97(3):717–721, 1999. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/oceanic-feeding-habits-chinook-
salmon-oncorhynchus-tshawytscha-northern-california;   https:
//spo.nmfs.noaa.gov/sites/default/files/26huntfi.pdf.

Porter:1999:DDT

[4275] Steven M. Porter and Gail H. Theilacker. The development of the digestive tract and eye in larval walleye pollock, Theragra chalcogramma. Fishery Bulletin, 97(3):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/development-digestive-tract-and-eye-larval-walleye-
pollock-theragra-chalcogramma;   https://spo.nmfs.noaa.gov/
sites/default/files/27porter.pdf.

Winner:1999:ESA

[4276] Brent L. Winner, Robert H. McMichael, Jr., and Laurel L. Brant. Evaluation of small T-anchor and dart tags for use in marking hatchery-reared juvenile red drum, Sciaenops ocellatus. Fishery Bulletin, 97(3):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/evaluation-small-t-
anchor-and-dart-tags-use-marking-hatchery-reared-juvenile-red-
drum;   https://spo.nmfs.noaa.gov/sites/default/files/28winner.
pdf.

Anonymous:1999:A

[4277] Anonymous. Announcement. Fishery Bulletin, 97(4):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14429;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/28announc.pdf.

Anonymous:1999:I

[4278] Anonymous. Index. Fishery Bulletin, 97(4):??, 1999. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14430;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/29index.pdf.

Anonymous:1999:SF

[4279] Anonymous. Subscription form. Fishery Bulletin, 97(4):??, 1999. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/node/14431;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/subform.pdf.

Aceves-Medina:1999:LDS

[4280] Gerardo Aceves-Medina, Enrique A. González, and Ricardo J. Saldierna. Larval development of Symphurus williamsi (Cynoglossidae: Pleuronectiformes) from the Gulf of California. Fishery Bulletin, 97(4):738–745, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/larval-development-
symphurus-williamsi-cynoglossidae-pleuronectiformes-gulf-california;
 https://spo.nmfs.noaa.gov/content/larval-development-symphurus-
williamsi-cynoglossidae-pleuronectiformes-gulf-california-0;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/01aceves_0.pdf.

Arreguin-Sanchez:1999:CET

[4281] Francisco Arreguín-Sanchez and Tony J. Pitcher. Catchability estimates and their application to the red grouper (Epinephelus morio) fishery of the Campeche Bank, Mexico. Fishery Bulletin, 97(4):746–757, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/catchability-
estimates-and-their-application-red-grouper-epinephelus-morio-
fishery-0;   https://spo.nmfs.noaa.gov/content/catchability-
estimates-and-their-application-red-grouper-epinephelus-morio-
fishery-campeche;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/02arregu_0.pdf.

Buckel:1999:FHB

[4282] Jeffrey A. Buckel, Michael J. Fogarty, and David O. Conover. Foraging habits of bluefish, Pomatomus saltatrix, on the U.S. east coast continental shelf. Fishery Bulletin, 97(4):758–775, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/foraging-habits-bluefish-
pomatomus-saltatrix-us-east-coast-continental-shelf;   https:
//spo.nmfs.noaa.gov/content/foraging-habits-bluefish-pomatomus-
saltatrix-us-east-coast-continental-shelf-0;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/03buckel_
0.pdf.

Buckel:1999:MPF

[4283] Jeffrey A. Buckel, Michael J. Fogarty, and David O. Conover. Mutual prey of fish and humans: a comparison of biomass consumed by bluefish, Pomatomus saltatrix, with that harvested by fisheries. Fishery Bulletin, 97(4):776–785, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/mutual-prey-fish-and-humans-comparison-biomass-
consumed-bluefish-pomatomus-saltatrix;   https://spo.nmfs.noaa.
gov/content/mutual-prey-fish-and-humans-comparison-biomass-
consumed-bluefish-pomatomus-saltatrix-0;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/04buckel_
0.pdf.

Friedlander:1999:SSM

[4284] Alan M. Friedlander, George W. Boehlert, Michael E. Field, Janet E. Mason, James V. Gardner, and Peter Dartnell. Sidescan-sonar mapping of benthic trawl marks on the shelf and slope off Eureka, California. Fishery Bulletin, 97(4):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
sidescan-sonar-mapping-benthic-trawl-marks-shelf-and-slope-
eureka-california;   https://spo.nmfs.noaa.gov/content/sidescan-
sonar-mapping-benthic-trawl-marks-shelf-and-slope-eureka-california-
0;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/05friedl_0.pdf.

Gabr:1999:RVS

[4285] Howaida R. Gabr, Roger T. Hanlon, Salah G. El-Etreby, and Mahmoud H. Hanafy. Reproductive versus somatic tissue growth during the life cycle of the cuttlefish Sepia pharaonis Ehrenberg, 1831. Fishery Bulletin, 97(4):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/reproductive-versus-somatic-tissue-growth-
during-life-cycle-cuttlefish-sepia-pharaonis;   https://spo.
nmfs.noaa.gov/content/reproductive-versus-somatic-tissue-growth-
during-life-cycle-cuttlefish-sepia-pharaonis-0;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/06gabrfi_
0.pdf.

Gillanders:1999:AMY

[4286] Bronwyn M. Gillanders, Douglas J. Ferrell, and Neil L. Andrew. Aging methods for yellowtail kingfish, Seriola lalandi, and results from age- and size-based growth models. Fishery Bulletin, 97(4):812–827, 1999. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/aging-methods-yellowtail-
kingfish-seriola-lalandi-and-results-age-and-size-based-growth;
 https://spo.nmfs.noaa.gov/content/aging-methods-yellowtail-
kingfish-seriola-lalandi-and-results-age-and-size-based-growth-
0;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/07gillan_0.pdf.

Hood:1999:AGM

[4287] Peter B. Hood and Andrea K. Johnson. Age, growth, mortality, and reproduction of vermilion snapper, Rhomboplites aurorubens, from the eastern Gulf of Mexico. Fishery Bulletin, 97(4):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/age-growth-mortality-
and-reproduction-vermilion-snapper-rhomboplites-aurorubens-
eastern;   https://spo.nmfs.noaa.gov/content/age-growth-mortality-
and-reproduction-vermilion-snapper-rhomboplites-aurorubens-

eastern-0;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/08hoodfi_0.pdf.

Hunt:1999:MAC

[4288] Joseph J. Hunt, Wayne T. Stobo, and Frank Almeida. Movement of Atlantic cod, Gadus morhua, tagged in the Gulf of Maine area. Fishery Bulletin, 97(4):842–861, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/movement-atlantic-cod-gadus-morhua-tagged-gulf-
maine-area;   https://spo.nmfs.noaa.gov/content/movement-atlantic-
cod-gadus-morhua-tagged-gulf-maine-area-0;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/09huntfi_
0.pdf.

IbanezAguirre:1999:GAS

[4289] Ana L. Ibañez Aguirre, Manuel Gallardo-Cabello, and Xavier Chiappa Carrara. Growth analysis of striped mullet, Mugil cephalus, and white mullet, M. curema (Pisces: Mugilidae), in the Gulf of Mexico. Fishery Bulletin, 97(4):861–872, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/growth-analysis-striped-mullet-mugil-
cephalus-and-white-mullet-m-curema-pisces-mugilidae;   https:
//spo.nmfs.noaa.gov/content/growth-analysis-striped-mullet-
mugil-cephalus-and-white-mullet-m-curema-pisces-mugilidae-0;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/10ibanez_0.pdf.

Kaimmer:1999:DOH

[4290] Stephen M. Kaimmer. Direct observations on the hooking behavior of Pacific halibut, Hippoglossus stenolepis. Fishery Bulletin, 97(4):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/direct-observations-
hooking-behavior-pacific-halibut-hippoglossus-stenolepis;
https://spo.nmfs.noaa.gov/content/direct-observations-hooking-
behavior-pacific-halibut-hippoglossus-stenolepis-0;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
11kaimme_0.pdf.

Lankford:1999:MDA

[4291] Thomas E. Lankford, Jr., Timothy E. Targett, and Patrick M. Gaffney. Mitochondrial DNA analysis of population structure in the Atlantic croaker, Micropogonias undulatus (Perciformes: Sciaenidae). Fishery Bulletin, 97(4):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
mitochondrial-dna-analysis-population-structure-atlantic-croaker-
micropogonias-undulatus;   https://spo.nmfs.noaa.gov/content/
mitochondrial-dna-analysis-population-structure-atlantic-croaker-
micropogonias-undulatus-0;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/12lankfo_0.pdf.

Massuti:1999:OMA

[4292] Enric Massutí, Beatriz Morales-Nin, and Joan Moranta. Otolith microstructure, age, and growth patterns of dolphin, Coryphaena hippurus, in the western Mediterranean. Fishery Bulletin, 97(4):891– 899, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/otolith-
microstructure-age-and-growth-patterns-dolphin-coryphaena-hippurus-
western;   https://spo.nmfs.noaa.gov/content/otolith-microstructure-
age-and-growth-patterns-dolphin-coryphaena-hippurus-western-
0;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/13massut_0.pdf.

Matkin:1999:API

[4293] Craig O. Matkin, Graeme Ellis, Peter Olesiuk, and Eva Saulitis. Association patterns and inferred genealogies of resident killer whales, Orcinus orca, in Prince William Sound, Alaska. Fishery Bulletin, 97(4):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/association-patterns-and-
inferred-genealogies-resident-killer-whales-orcinus-orca-prince;
 https://spo.nmfs.noaa.gov/content/association-patterns-and-
inferred-genealogies-resident-killer-whales-orcinus-orca-prince-
0;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/14matkin_0.pdf.

Moser:1999:VDE

[4294] H. Geoffrey Moser and Tilman Pommeranz. Vertical distribution of eggs and larvae of northern anchovy, Engraulis mordax, and of the larvae of associated fishes at two sites in the Southern California Bight. Fishery Bulletin, 97(4):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/vertical-distribution-eggs-and-larvae-northern-
anchovy-engraulis-mordax-and-larvae;   https://spo.nmfs.noaa.
gov/content/vertical-distribution-eggs-and-larvae-northern-
anchovy-engraulis-mordax-and-larvae-0;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/15moserf_0.pdf.

Natanson:1999:GTS

[4295] Lisa J. Natanson, John G. Casey, Nancy E. Kohler, and Tristram Colket IV. Growth of the tiger shark, Galeocerdo cuvier, in the western North Atlantic based on tag returns and length frequencies; and a note on the effects of tagging. Fishery Bulletin, 97(4):944–953, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/growth-tiger-shark-
galeocerdo-cuvier-western-north-atlantic-based-tag-returns-
and-length;   https://spo.nmfs.noaa.gov/content/growth-tiger-
shark-galeocerdo-cuvier-western-north-atlantic-based-tag-returns-
and-length-0;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/16natans_0.pdf.

Petrik:1999:RAC

[4296] Rachel Petrik, Phillip S. Levin, Gregory W. Stunz, and John Malone. Recruitment of Atlantic croaker, Micropogonias undulatus: Do postsettlement processes disrupt or reinforce initial patterns of settlement? Fishery Bulletin, 97(4):954–961, 1999. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/recruitment-atlantic-
croaker-micropogonias-undulatus-do-postsettlement-processes-
disrupt-0;   https://spo.nmfs.noaa.gov/content/recruitment-
atlantic-croaker-micropogonias-undulatus-do-postsettlement-
processes-disrupt-or;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/17petrik_0.pdf.

Robards:1999:TGV

[4297] Martin D. Robards, John F. Piatt, Arthur B. Kettle, and Alisa A. Abookire. Temporal and geographic variation in fish communities of lower Cook Inlet, Alaska. Fishery Bulletin, 97(4):962– 977, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/temporal-
and-geographic-variation-fish-communities-lower-cook-inlet-
alaska;   https://spo.nmfs.noaa.gov/content/temporal-and-geographic-
variation-fish-communities-lower-cook-inlet-alaska-0;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
18robard_0.pdf.

Simpfendorfer:1999:MED

[4298] Colin A. Simpfendorfer. Mortality estimates and demographic analysis for the Australian sharpnose shark, Rhizoprionodon taylori, from northern Australia. Fishery Bulletin, 97(4):??, 1999. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/mortality-estimates-
and-demographic-analysis-australian-sharpnose-shark-rhizoprionodon;
 https://spo.nmfs.noaa.gov/content/mortality-estimates-and-
demographic-analysis-australian-sharpnose-shark-rhizoprionodon-
0;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/19simpfe_0.pdf.

Smale:1999:RFS

[4299] Malcolm J. Smale and André J. J. Goosen. Reproduction and feeding of spotted gully shark, Triakis megalopterus, off the Eastern Cape, South Africa. Fishery Bulletin, 97(4):??, 1999. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/reproduction-and-feeding-spotted-
gully-shark-triakis-megalopterus-eastern-cape-south-0;   https:
//spo.nmfs.noaa.gov/content/reproduction-and-feeding-spotted-
gully-shark-triakis-megalopterus-eastern-cape-south-africa;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/20smalef_0.pdf.

Stoner:1999:BWF

[4300] Allan W. Stoner, Allen J. Bejda, John P. Manderson, Beth A. Phelan, Linda L. Stehlik, and Jeffrey P. Pessutti. Behavior of winter flounder, Pseudopleuronectes americanus, during the reproductive season: laboratory and field observations on spawning, feeding, and locomotion. Fishery Bulletin, 97(4):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/behavior-winter-flounder-
pseudopleuronectes-americanus-during-reproductive-season;
https://spo.nmfs.noaa.gov/content/behavior-winter-flounder-
pseudopleuronectes-americanus-during-reproductive-season-0;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/21stoner_0.pdf.

UrbanR:1999:PSH

[4301] Jorge Urbán R., Carlos Alvarez F., Mario Salinas Z., Jeff Jacobsen, Kenneth C. Balcomb III, Armando Jaramillo L., Paloma Ladrón de Guevara P., and Anelio Aguayo L. Population size of humpback whale, Megaptera novaeangliae, in waters off the Pacific coast of Mexico. Fishery Bulletin, 97(4):1017–1024, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/population-size-humpback-
whale-megaptera-novaeangliae-waters-pacific-coast-mexico;
https://spo.nmfs.noaa.gov/content/population-size-humpback-

whale-megaptera-novaeangliae-waters-pacific-coast-mexico-0;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/22urbanf_0.pdf.

Vecchione:1999:EAS

[4302] Michael Vecchione. Extraordinary abundance of squid paralarvae in the tropical eastern Pacific Ocean during El Niño of 1987. Fishery Bulletin, 97(4):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
extraordinary-abundance-squid-paralarvae-tropical-eastern-pacific-
ocean-during-el-ni%C3%B1o-0;   https://spo.nmfs.noaa.gov/content/
extraordinary-abundance-squid-paralarvae-tropical-eastern-pacific-
ocean-during-el-ni%C3%B1o-1987;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/23vecchi_0.pdf.

Watson:1999:ELH

[4303] William Watson. Early life history stages of the whitetip flyingfish, Cheilopogon xenopterus (Gilbert, 1890) (Pisces: Exocoetidae). Fishery Bulletin, 97(4):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
early-life-history-stages-whitetip-flyingfish-cheilopogon-xenopterus-
gilbert-1890-pisces;   https://spo.nmfs.noaa.gov/content/early-
life-history-stages-whitetip-flyingfish-cheilopogon-xenopterus-
gilbert-1890-pisces-0;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/24watson_0.pdf.

Garlich-Miller:1999:EHP

[4304] Joel L. Garlich-Miller and Douglas M. Burn. Estimating the harvest of Pacific walrus, Odobenus rosmarus divergens, in Alaska. Fishery Bulletin, 97(4):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
estimating-harvest-pacific-walrus-odobenus-rosmarus-divergens-
alaska;   https://spo.nmfs.noaa.gov/content/estimating-harvest-
pacific-walrus-odobenus-rosmarus-divergens-alaska-0;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
25garlic_0.pdf.

Yang:1999:TRAa

[4305] Mei-Sun Yang. The trophic role of Atka mackerel, Pleurogrammus monopterygius, in the Aleutian Islands area primary tabs. Fishery Bulletin, 97(4):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
trophic-role-atka-mackerel-pleurogrammus-monopterygius-aleutian-

islands-area-primary-tabs;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/26yangfi_0.pdf.

Yang:1999:TRAb

[4306] Mei-Sun Yang. The trophic role of Atka mackerel, Pleurogrammus monopterygius, in the Aleutian Islands area. Fishery Bulletin, 97(4): ??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/trophic-role-
atka-mackerel-pleurogrammus-monopterygius-aleutian-islands-
area.

Zeller:1999:UTA

[4307] Dirk C. Zeller. Ultrasonic telemetry: its application to coral reef fisheries research. Fishery Bulletin, 97(4):??, 1999. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/ultrasonic-telemetry-its-application-
coral-reef-fisheries-research;   https://spo.nmfs.noaa.gov/
content/ultrasonic-telemetry-its-application-coral-reef-fisheries-
research-0;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/27zeller_0.pdf.

Anonymous:2000:SFa

[4308] Anonymous. Subscription form. Fishery Bulletin, 98(1):??, 2000. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/node/14432;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/subform%20(1)
.pdf.

Arkhipkin:2000:ISW

[4309] Alexander I. Arkhipkin. Intrapopulation structure of winter-spawned Argentine shortfin squid, Illex argentinus (Cephalopoda, Ommastrephidae), during its feeding period over the Patagonian Shelf. Fishery Bulletin, 98(1):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/intrapopulation-structure-winter-spawned-argentine-
shortfin-squid-illex-argentinus;   https://spo.nmfs.noaa.gov/
sites/default/files/01_2.pdf.

Beacham:2000:MDV

[4310] Terry D. Beacham, Khai D. Le, Monique R. Raap, Kim Hyatt, Wilf Luedke, and Ruth E. Withler. Microsatellite DNA variation and estimation of stock composition of sockeye salmon, Oncorhynchus nerka, in Barkley Sound, British Columbia. Fishery Bulletin, 98(1):??, 2000.

CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/microsatellite-dna-
variation-and-estimation-stock-composition-sockeye-salmon-oncorhynchus;
 https://spo.nmfs.noaa.gov/sites/default/files/02_2.pdf.

Clear:2000:DVA

[4311] Naomi P. Clear, John S. Gunn, and Anthony J. Rees. Direct validation of annual increments in the otoliths of juvenile southern bluefin tuna, Thunnus maccoyii, by means of a large-scale mark-recapture experiment with strontium chloride. Fishery Bulletin, 98(1):25–40, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/direct-validation-annual-
increments-otoliths-juvenile-southern-bluefin-tuna-thunnus;
 https://spo.nmfs.noaa.gov/sites/default/files/03_2.pdf.

Francis:2000:RED

[4312] Malcolm P. Francis and John D. Stevens. Reproduction, embryonic development, and growth of the porbeagle shark, Lamna nasus, in the southwest Pacific Ocean. Fishery Bulletin, 98(1):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/reproduction-
embryonic-development-and-growth-porbeagle-shark-lamna-nasus-
southwest-pacific;   https://spo.nmfs.noaa.gov/sites/default/
files/04_2.pdf.

Govoni:2000:DLS

[4313] John Jeffrey Govoni, Bruce W. Stender, and Oleg Pashuk. Distribution of larval swordfish, Xiphias gladius, and probable spawning off the southeastern United States. Fishery Bulletin, 98(1):64–74, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/distribution-larval-
swordfish-xiphias-gladius-and-probable-spawning-southeastern-
united;   https://spo.nmfs.noaa.gov/sites/default/files/05_
2.pdf.

Hayes:2000:BRP

[4314] Daniel B. Hayes. A biological reference point based on the Leslie matrix. Fishery Bulletin, 98(1):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/biological-reference-point-based-leslie-matrix;
 https://spo.nmfs.noaa.gov/sites/default/files/06_2.pdf.

Lenihan:2000:BES

[4315] Hunter S. Lenihan and Fiorenza Micheli. Biological effects of shellfish harvesting on oyster reefs: resolving a fishery conflict by ecological experimentation. Fishery Bulletin, 98(1):??, 2000. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/biological-effects-
shellfish-harvesting-oyster-reefs-resolving-fishery-conflict-
ecological;   https://spo.nmfs.noaa.gov/sites/default/files/
07_2.pdf.

Love:2000:FAA

[4316] Milton S. Love, Jennifer E. Caselle, and Linda Snook. Fish assemblages around seven oil platforms in the Santa Barbara Channel area. Fishery Bulletin, 98(1):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fish-assemblages-around-seven-oil-platforms-santa-barbara-channel-
area;   https://spo.nmfs.noaa.gov/sites/default/files/08_2.
pdf.

Nemerson:2000:SSF

[4317] David Nemerson, Steven Berkeley, and Carl Safina. Spawning site fidelity in Atlantic bluefin tuna, Thunnus thynnus: the use of size-frequency analysis to test for the presence of migrant east Atlantic bluefin tuna on Gulf of Mexico spawning grounds. Fishery Bulletin, 98(1):118–126, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/spawning-site-
fidelity-atlantic-bluefin-tuna-thunnus-thynnus-use-size-frequency-
analysis;   https://spo.nmfs.noaa.gov/sites/default/files/09_
2.pdf.

Seyoum:2000:AGP

[4318] Seifu Seyoum, Michael D. Tringali, Theresa M. Bert, Doug McElroy, and Rod Stokes. An analysis of genetic population structure in red drum, Sciaenops ocellatus, based on mtDNA control region sequences. Fishery Bulletin, 98(1):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/analysis-genetic-population-structure-red-drum-
sciaenops-ocellatus-based-mtdna-control;   https://spo.nmfs.
noaa.gov/sites/default/files/10_3.pdf.

Shields:2000:MHB

[4319] Jeffrey D. Shields and Christopher M. Squyars. Mortality and hematology of blue crabs, Callinectes sapidus, experimentally infected with the parasitic dinoflagellate Hematodinium perezi. Fishery Bulletin, 98(1): ??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/mortality-and-
hematology-blue-crabs-callinectes-sapidus-experimentally-infected-
parasitic;   https://spo.nmfs.noaa.gov/sites/default/files/
11_2.pdf.

Stevenson:2000:ADG

[4320] Jill T. Stevenson and David H. Secor. Age determination and growth of Hudson River Atlantic sturgeon, Acipenser oxyrinchus. Fishery Bulletin, 98(1):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/age-determination-and-growth-hudson-river-atlantic-
sturgeon-acipenser-oxyrinchus;   https://spo.nmfs.noaa.gov/
sites/default/files/12_2.pdf.

Steves:2000:SNH

[4321] Brian P. Steves, Robert K. Cowen, and Mark H. Malchoff. Settlement and nursery habitats for demersal fishes on the continental shelf of the New York Bight. Fishery Bulletin, 98(1):167–188, 2000. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/settlement-and-nursery-habitats-
demersal-fishes-continental-shelf-new-york-bight;   https:/
/spo.nmfs.noaa.gov/sites/default/files/13_3.pdf.

Wetherbee:2000:ADS

[4322] Bradley M. Wetherbee. Assemblage of deep-sea sharks on Chatham Rise, New Zealand. Fishery Bulletin, 98(1):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/assemblage-deep-sea-sharks-chatham-rise-new-
zealand;   https://spo.nmfs.noaa.gov/sites/default/files/14_
2.pdf.

Wyanski:2000:GPA

[4323] David M. Wyanski, D. Byron White, and Charles A. Barans. Growth, population age structure, and aspects of the reproductive biology of snowy grouper, Epinephelus niveatus, off North Carolina and South Carolina. Fishery Bulletin, 98(1):199–218, 2000. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/growth-population-age-structure-
and-aspects-reproductive-biology-snowy-grouper-epinephelus;
 https://spo.nmfs.noaa.gov/sites/default/files/15_2.pdf.

Balart:2000:NBC

[4324] Eduardo F. Balart, Jeanette González-García, and Carlos Villavicencio-Garayzar. Notes on the biology of Cephalurus cephalus and Parmaturus xaniurus (Chondrichthyes: Scyliorhinidae) from the west coast of Baja California Sur, México. Fishery Bulletin, 98(1):219–221, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/notes-biology-
cephalurus-cephalus-and-parmaturus-xaniurus-chondrichthyes-
scyliorhinidae;   https://spo.nmfs.noaa.gov/sites/default/files/
16_1.pdf.

Galindo-Bect:2000:PSL

[4325] Manuel S. Galindo-Bect, Edward P. Glenn, Henry M. Page, Kevin Fitzsimmons, Luis A. Galindo-Bect, Jose M. Hernandez-Ayon, Robert L. Petty, Jaqueline Garcia-Hernandez, and David Moore. Penaeid shrimp landings in the upper Gulf of California in relation to Colorado River freshwater discharge. Fishery Bulletin, 98(1):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/penaeid-
shrimp-landings-upper-gulf-california-relation-colorado-river-
freshwater-discharge;   https://spo.nmfs.noaa.gov/sites/default/
files/17_1.pdf.

Anonymous:2000:SFb

[4326] Anonymous. Subscription form. Fishery Bulletin, 98(2):??, 2000. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/node/14435;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/subform%20(2)
.pdf.

Acha:2000:SBM

[4327] Eduardo M. Acha and Gustavo J. Macchi. Spawning of Brazilian menhaden, Brevoortia aurea, in the Río de la Plata estuary off Argentina and Uruguay. Fishery Bulletin, 98(2):227–235, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/spawning-brazilian-
menhaden-brevoortia-aurea-r%C3%ADo-de-la-plata-estuary-argentina-
and-uruguay;   https://spo.nmfs.noaa.gov/node/14433;   https:
//spo.nmfs.noaa.gov/sites/default/files/01_1.pdf;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
01.pdf.

Brewster-Geisz:2000:MSS

[4328] Karyl K. Brewster-Geisz and Thomas J. Miller. Management of the sandbar shark, Carcharhinus plumbeus: implications of a stage-based model. Fishery Bulletin, 98(2):??, 2000. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/management-sandbar-
shark-carcharhinus-plumbeus-implications-stage-based-model;
 https://spo.nmfs.noaa.gov/sites/default/files/02_1.pdf.

Alonso:2000:FHS

[4329] Mariano Keon Alonso, Enrique A. Crespo, Susana N. Pedraza, Néstor A. García, and Mariano A. Coscarella. Food habits of the South American sea lion, Otaria flavescens, off Patagonia, Argentina. Fishery Bulletin, 98(2):250–263, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/food-
habits-south-american-sea-lion-otaria-flavescens-patagonia-
argentina;   https://spo.nmfs.noaa.gov/sites/default/files/
03_1.pdf.

Lo:2000:MSP

[4330] Nancy C. H. Lo, John R. Hunter, and James H. Churnside. Modeling statistical performance of an airborne lidar survey system for anchovy. Fishery Bulletin, 98(2):264–282, 2000. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/modeling-statistical-
performance-airborne-lidar-survey-system-anchovy;   https:/
/spo.nmfs.noaa.gov/sites/default/files/04_1.pdf.

Macchi:2000:SFB

[4331] Gustavo J. Macchi and Eduardo M. Acha. Spawning frequency and batch fecundity of Brazilian menhaden, Brevoortia aurea, in the Río de la Plata estuary off Argentina and Uruguay. Fishery Bulletin, 98(2):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/spawning-frequency-
and-batch-fecundity-brazilian-menhaden-brevoortia-aurea-r%C3%
ADo-de-la-plata;   https://spo.nmfs.noaa.gov/sites/default/
files/05_1.pdf.

Manickchand-Heileman:2000:AGY

[4332] Sherry C. Manickchand-Heileman and Dawn A. T. Phillip. Age and growth of the yellowedge grouper, Epinephelus flavolimbatus, and the yellowmouth grouper, Mycteroperca interstitialis, off Trinidad and Tobago.

Fishery Bulletin, 98(2):290–298, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/age-and-growth-yellowedge-grouper-epinephelus-flavolimbatus-
and-yellowmouth-grouper;   https://spo.nmfs.noaa.gov/sites/
default/files/06_1.pdf.

Mollet:2000:RBF

[4333] Henry F. Mollet, Geremy Cliff, Harold L. Pratt, Jr., and John D. Stevens. Reproductive biology of the female shortfin mako, Isurus oxyrinchus Rafinesque, 1810, with comments on the embryonic development of lamnoids. Fishery Bulletin, 98(2):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/reproductive-biology-female-shortfin-
mako-isurus-oxyrinchus-rafinesque-1810-comments;   https://
spo.nmfs.noaa.gov/sites/default/files/07_1.pdf.

Mortensen:2000:RBE

[4334] Donald Mortensen, Alex Wertheimer, Sidney Taylor, and Joyce Landingham. The relation between early marine growth of pink salmon, Oncorhynchus gorbuscha, and marine temperature, secondary production, and survival to adulthood. Fishery Bulletin, 98(2): ??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/relation-
between-early-marine-growth-pink-salmon-oncorhynchus-gorbuscha-
and-marine;   https://spo.nmfs.noaa.gov/sites/default/files/
08_1.pdf.

Orbacz:2000:GST

[4335] Elizabeth A. Orbacz and Patrick M. Gaffney. Genetic structure of tautog (Tautoga onitis) populations assayed by RFLP and DGGE analysis of mitochondrial and nuclear genes. Fishery Bulletin, 98(2):336–344, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/genetic-structure-tautog-
tautoga-onitis-populations-assayed-rflp-and-dgge-analysis;
https://spo.nmfs.noaa.gov/node/14434;   https://spo.nmfs.noaa.
gov/sites/default/files/09_1.pdf;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/09.pdf.

Rivera:2000:AGD

[4336] Glauco A. Rivera and Richard S. Appeldoorn. Age and growth of dolphinfish, Coryphaena hippurus, off Puerto Rico. Fishery Bulletin, 98(2):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
age-and-growth-dolphinfish-coryphaena-hippurus-puerto-rico;
 https://spo.nmfs.noaa.gov/sites/default/files/10_2.pdf.

Rocha-Olivares:2000:MID

[4337] Axayácatl Rocha-Olivares, H. Geoffrey Moser, and Jason Stannard. Molecular identification and description of pelagic young of the rockfishes Sebastes constellatus and Sebastes ensifer. Fishery Bulletin, 98 (2):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/molecular-
identification-and-description-pelagic-young-rockfishes-sebastes-
constellatus-and;   https://spo.nmfs.noaa.gov/sites/default/
files/11_1.pdf.

Schaefer:2000:SVR

[4338] Kurt M. Schaefer and Charles W. Oliver. Shape, volume, and resonance frequency of the swimbladder of yellowfin tuna, Thunnus albacares. Fishery Bulletin, 98(2):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
shape-volume-and-resonance-frequency-swimbladder-yellowfin-
tuna-thunnus-albacares;   https://spo.nmfs.noaa.gov/sites/default/
files/12_1.pdf.

Sevigny:2000:IDL

[4339] Jean-Marie Sévigny, Patrice Gagné, Yves de Lafontaine, and Julian Dodson. Identification and distribution of larvae of redfish (Sebastes fasciatus and S. mentella: Scorpaenidae) in the Gulf of St. Lawrence. Fishery Bulletin, 98(2):375–388, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/identification-and-distribution-larvae-redfish-
sebastes-fasciatus-and-s-mentella;   https://spo.nmfs.noaa.
gov/sites/default/files/13_2.pdf.

Stevens:2000:GFT

[4340] Bardley G. Stevens, Ivan Vining, Susie Byersdorfer, and William Donaldson. Ghost fishing by Tanner crab (Chionoecetes bairdi) pots off Kodiak, Alaska: pot density and catch per trap as determined from sidescan sonar and pot recovery data. Fishery Bulletin, 98 (2):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/ghost-
fishing-tanner-crab-chionoecetes-bairdi-pots-kodiak-alaska-
pot-density-and-catch-trap;   https://spo.nmfs.noaa.gov/sites/
default/files/14_1.pdf.

Stokesbury:2000:SDP

[4341] Kevin D. E. Stokesbury, Jay Kirsch, Evelyn D. Brown, Gary L. Thomas, and Brenda L. Norcross. Spatial distributions of Pacific herring, Clupea pallasi, and walleye pollock, Theragra chalcogramma, in Prince William Sound, Alaska. Fishery Bulletin, 98(2):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/spatial-
distributions-pacific-herring-clupea-pallasi-and-walleye-pollock-
theragra;   https://spo.nmfs.noaa.gov/sites/default/files/15_
1.pdf.

Volstad:2000:MED

[4342] Jon H. Vølstad, Alexei F. Sharov, Glenn Davis, and Brenda Davis. A method for estimating dredge catching efficiency for blue crabs, Callinectes sapidus, in Chesapeake Bay. Fishery Bulletin, 98(2):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/method-estimating-dredge-
catching-efficiency-blue-crabs-callinectes-sapidus-chesapeake-
bay;   https://spo.nmfs.noaa.gov/sites/default/files/16_0.pdf.

Ahrenholz:2000:PIF

[4343] Dean W. Ahrenholz, Deborah D. Squires, James A. Rice, Stephen W. Nixon, and Gary R. Fitzhugh. Periodicity of increment formation in otoliths of overwintering postlarval and prejuvenile Atlantic menhaden, Brevoortia tyrannus. Fishery Bulletin, 98(2):??, 2000. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/periodicity-increment-formation-
otoliths-overwintering-postlarval-and-prejuvenile-atlantic;
 https://spo.nmfs.noaa.gov/sites/default/files/17_0.pdf.

Powell:2000:LPJ

[4344] Allyn B. Powell, David G. Lindquist, and Jonathan A. Hare. Larval and pelagic juvenile fishes collected with three types of gear in Gulf Stream and shelf waters in Onslow Bay, North Carolina, and comments on ichthyoplankton distribution and hydrography. Fishery Bulletin, 98(2):427–438, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/larval-and-pelagic-juvenile-fishes-collected-three-
types-gear-gulf-stream-and-shelf-waters;   https://spo.nmfs.
noaa.gov/sites/default/files/18_0.pdf.

Anonymous:2000:SFc

[4345] Anonymous. Subscription form. Fishery Bulletin, 98(3):??, 2000. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/node/14436;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/subform%20(3)
.pdf.

Bolles:2000:DBS

[4346] Karen L. Bolles and Gavin A. Begg. Distinction between silver hake (Merluccius bilinearis) stocks in U.S. waters of the northwest Atlantic based on whole otolith morphometrics. Fishery Bulletin, 98(3):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/distinction-
between-silver-hake-merluccius-bilinearis-stocks-us-waters-
northwest-atlantic;   https://spo.nmfs.noaa.gov/sites/default/
files/01_0.pdf.

Broadhurst:2000:ETD

[4347] Matt K. Broadhurst, Paul E. McShane, and Roger B. Larsen. Effects of twine diameter and mesh size in the body of prawn trawls on bycatch in Gulf St. Vincent, Australia. Fishery Bulletin, 98(3):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/effects-twine-diameter-and-mesh-
size-body-prawn-trawls-bycatch-gulf-st-vincent-australia;
https://spo.nmfs.noaa.gov/sites/default/files/02_0.pdf.

Cappo:2000:NAV

[4348] Mike Cappo, Peter Eden, Stephen J. Newman, and Simon Robertson. A new approach to validation of periodicity and timing of opaque zone formation in the otoliths of eleven species of Lutjanus from the central Great Barrier Reef. Fishery Bulletin,

98(3):??, 2000.  CODEN FSYBAY.  ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
new-approach-validation-periodicity-and-timing-opaque-zone-
formation-otoliths-eleven;   https://spo.nmfs.noaa.gov/sites/
default/files/03_0.pdf.

DeMartini:2000:SMS

[4349] Edward E. DeMartini, James H. Uchiyama, and Happy A. Williams. Sexual maturity, sex ratio, and size composition of swordfish, Xiphias gladius, caught by the Hawaii-based pelagic longline fishery. Fishery Bulletin, 98(3):489–506, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/sexual-maturity-sex-ratio-and-size-composition-
swordfish-xiphias-gladius-caught-hawaii;   https://spo.nmfs.
noaa.gov/sites/default/files/04_0.pdf.

Drass:2000:LDR

[4350] Denice M. Drass, Kevin L. Bootes, Joanne Lyczkowski-Shultz, Bruce H. Comyns, G. Joan Holt, Cecilia M. Riley, and Ronald P. Phelps. Larval development of red snapper, Lutjanus campechanus, and comparisons with co-occurring snapper species. Fishery Bulletin, 98(3):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/larval-development-red-snapper-
lutjanus-campechanus-and-comparisons-co-occurring-snapper;
https://spo.nmfs.noaa.gov/sites/default/files/05_0.pdf.

McBride:2000:CDS

[4351] Richard S. McBride and Kim A. McKown. Consequences of dispersal of subtropically spawned crevalle jacks, Caranx hippos, to temperate estuaries. Fishery Bulletin, 98(3):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/consequences-dispersal-subtropically-
spawned-crevalle-jacks-caranx-hippos-temperate;   https://spo.
nmfs.noaa.gov/sites/default/files/06_0.pdf.

Orr:2000:RGL

[4352] James W. Orr and Ann C. Matarese. Revision of the genus Lepidopsetta gill, 1862 (teleostei: Pleuronectidae) based on larval and adult morphology, with a description of a new species from the North Pacific Ocean and Bering Sea. Fishery Bulletin, 98(3):??, 2000. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/revision-genus-lepidopsetta-gill-
1862-teleostei-pleuronectidae-based-larval-and-adult;   https:
//spo.nmfs.noaa.gov/sites/default/files/07_0.pdf.

Ortiz:2000:AME

[4353] Mauricio Ortiz, Christopher M. Legault, and Nelson M. Ehrhardt. An alternative method for estimating bycatch from the U.S. shrimp trawl fishery in the Gulf of Mexico, 1972—1995. Fishery Bulletin, 98(3):583–599, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/alternative-method-
estimating-bycatch-us-shrimp-trawl-fishery-gulf-mexico-1972%
E2%80%931995;   https://spo.nmfs.noaa.gov/sites/default/files/
08_0.pdf.

Pilling:2000:VAG

[4354] Graham M. Pilling, Richard S. Millner, Michael W. Easey, Christopher C. Mees, Shyama Rathacharen, and Roland Azemia. Validation of annual growth increments in the otoliths of the lethrinid Lethrinus mahsena and the lutjanid Aprion virescens from sites in the tropical Indian Ocean, with notes on the nature of growth increments in Pristipomoides filamentosus. Fishery Bulletin, 98(3):??, 2000. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/validation-annual-growth-
increments-otoliths-lethrinid-lethrinus-mahsena-and-lutjanid;
 https://spo.nmfs.noaa.gov/sites/default/files/09_0.pdf.

Taylor:2000:AGM

[4355] Ronald G. Taylor, James A. Whittington, Harry J. Grier, and Roy E. Crabtree. Age, growth, maturation, and protandric sex reversal in common snook, Centropomus undecimalis, from the east and west coasts of South Florida. Fishery Bulletin, 98(3): ??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/age-
growth-maturation-and-protandric-sex-reversal-common-snook-
centropomus-undecimalis-east;   https://spo.nmfs.noaa.gov/sites/
default/files/10_1.pdf.

Yoklavich:2000:HAD

[4356] Mary M. Yoklavich, H. Gary Greene, Gregor M. Cailliet, Deidre E. Sullivan, Robert N. Lea, and Milton S. Love. Habitat associations of deep-water rockfishes in a submarine canyon: an example of a natural refuge. Fishery Bulletin, 98(3):??, 2000. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/habitat-associations-deep-water-
rockfishes-submarine-canyon-example-natural-refuge;   https:
//spo.nmfs.noaa.gov/sites/default/files/11_0.pdf.

Bolden:2000:LDM

[4357] Stephania K. Bolden. Long-distance movement of a Nassau grouper (Epinephelus striatus) to a spawning aggregation in the central Bahamas. Fishery Bulletin, 98(3):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/long-distance-movement-nassau-grouper-epinephelus-striatus-
spawning-aggregation-central;   https://spo.nmfs.noaa.gov/sites/
default/files/12_0.pdf.

Dudley:2000:CSW

[4358] Sheldon F. J. Dudley, Michael D. Anderson-Reade, Greg S. Thompson, and Paul B. McMullen. Concurrent scavenging off a whale carcass by great white sharks, Carcharodon carcharias, and tiger sharks, Galeocerdo cuvier. Fishery Bulletin, 98(3):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/concurrent-scavenging-whale-carcass-great-
white-sharks-carcharodon-carcharias-and-tiger;   https://spo.
nmfs.noaa.gov/sites/default/files/13_1.pdf.

Powell:2000:VAE

[4359] Allyn B. Powell, Elisabeth H. Laban, Scott A. Holt, and G. Joan Holt. Validation of age estimates from otoliths of larval and juvenile spotted seatrout, Cynoscion nebulosus. Fishery Bulletin, 98(3):??, 2000. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/validation-age-estimates-
otoliths-larval-and-juvenile-spotted-seatrout-cynoscion-nebulosus;
 https://spo.nmfs.noaa.gov/sites/default/files/14_0.pdf.

Rooper:2000:CPH

[4360] Christopher N. Rooper and Lewis J. Haldorson. Consumption of Pacific herring (Clupea pallasi) eggs by greenling (Hexagrammidae) in Prince William Sound, Alaska. Fishery Bulletin, 98(3):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/consumption-pacific-
herring-clupea-pallasi-eggs-greenling-hexagrammidae-prince-
william;   https://spo.nmfs.noaa.gov/sites/default/files/15_
0.pdf.

Anonymous:2000:SFd

[4361] Anonymous. Subscription form. Fishery Bulletin, 98(4):??, 2000. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/node/14439;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/subform%20(4)
.pdf.

Anonymous:2000:A

[4362] Anonymous. Announcement. Fishery Bulletin, 98(4):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14437;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/announce.pdf.

Anonymous:2000:I

[4363] Anonymous. Index. Fishery Bulletin, 98(4):??, 2000. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14438;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/index%20(10).pdf.

Brodziak:2000:VLH

[4364] Jon Brodziak and Robert Mikus. Variation in life history parameters of Dover sole, Microstomus pacificus, off the coasts of Washington, Oregon, and northern California. Fishery Bulletin, 98(4): ??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/variation-
life-history-parameters-dover-sole-microstomus-pacificus-coasts-
washington-oregon;   https://spo.nmfs.noaa.gov/sites/default/
files/01.pdf.

delaRosa-Velez:2000:GST

[4365] Jorge de la Rosa-Vélez, Roberto Escobar-Fernández, Francisco Correa, Mercedes Maqueda-Cornejo, and Javier de la Torre-Cueto. Genetic structure of two commercial penaeids (Penaeus californiensis and P. stylirostris) from the Gulf of California, as revealed by allozyme variation. Fishery Bulletin, 98(4):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/genetic-structure-two-commercial-penaeids-
penaeus-californiensis-and-p-stylirostris-gulf;   https://spo.
nmfs.noaa.gov/sites/default/files/02.pdf.

Gledhill:2000:ILK

[4366] Christopher T. Gledhill and Joanne Lyczkowski-Shultz. Indices of larval king mackerel (Scomberomorus cavalla) abundance in the Gulf of Mexico for use in population assessments. Fishery Bulletin, 98(4):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/indices-larval-king-
mackerel-scomberomorus-cavalla-abundance-gulf-mexico-use-population;
 https://spo.nmfs.noaa.gov/sites/default/files/03.pdf.

Heckel:2000:EBS

[4367] Gisela Heckel, Kim E. Murphy, and Guillermo A. Compeán Jiménez. Evasive behavior of spotted and spinner dolphins (Stenella attenuata and S. longirostris) during fishing for yellowfin tuna (Thunnus albacares) in the eastern Pacific Ocean. Fishery Bulletin, 98(4):692–703, 2000. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/evasive-behavior-spotted-
and-spinner-dolphins-stenella-attenuata-and-s-longirostris-
during;   https://spo.nmfs.noaa.gov/sites/default/files/04.
pdf.

Hernandez:2000:LGA

[4368] Eduardo H. Hernandez and Leonardo R. Castro. Larval growth of the anchoveta Engraulis ringens during the winter spawning season off central Chile. Fishery Bulletin, 98(4):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/larval-growth-anchoveta-engraulis-ringens-
during-winter-spawning-season-central-chile;   https://spo.
nmfs.noaa.gov/sites/default/files/05.pdf.

Hoff:2000:BET

[4369] Gerald R. Hoff. Biology and ecology of threaded sculpin, Gymnocanthus pistilliger, in the eastern Bering Sea. Fishery Bulletin, 98(4):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/biology-and-ecology-
threaded-sculpin-gymnocanthus-pistilliger-eastern-bering-sea;
 https://spo.nmfs.noaa.gov/sites/default/files/06.pdf.

Hood:2000:AGM

[4370] Peter B. Hood and Andrea K. Johnson. Age, growth, mortality, and reproduction of red porgy, Pagrus pagrus, from the eastern Gulf of Mexico. Fishery Bulletin, 98(4):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/age-growth-mortality-and-
reproduction-red-porgy-pagrus-pagrus-eastern-gulf-mexico;
https://spo.nmfs.noaa.gov/sites/default/files/07.pdf.

Kingsley:2000:NDB

[4371] Michael C. S. Kingsley. Numbers and distribution of beluga whales, Delphinapterus leucas, in James Bay, eastern Hudson Bay, and Ungava Bay in Canada during the summer of 1993. Fishery Bulletin, 98(4):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/numbers-and-
distribution-beluga-whales-delphinapterus-leucas-james-bay-
eastern-hudson-bay;   https://spo.nmfs.noaa.gov/sites/default/
files/08.pdf.

Kwok:2000:AGC

[4372] Kai Yin Kwok and I-Hsun Ni. Age and growth of cutlassfishes, Trichiurus spp., from the South China Sea. Fishery Bulletin, 98(4): ??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/age-and-
growth-cutlassfishes-trichiurus-spp-south-china-sea;   https:
//spo.nmfs.noaa.gov/sites/default/files/09.pdf.

Lindley:2000:MPS

[4373] Steven T. Lindley, Michael S. Mohr, and Michael H. Prager. Monitoring protocol for Sacramento River winter Chinook salmon, Oncorhynchus tshawytscha: application of statistical power analysis to recovery of an endangered species. Fishery Bulletin, 98(4):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/monitoring-protocol-sacramento-river-
winter-chinook-salmon-oncorhynchus-tshawytscha;   https://spo.
nmfs.noaa.gov/sites/default/files/10_0.pdf.

Nonaka:2000:SVL

[4374] Ricardo H. Nonaka, Yasunobu Matsuura, and Kazuko Suzuki. Seasonal variation in larval fish assemblages in relation to oceanographic conditions in the abrolhos bank region off eastern Brazil. Fishery Bulletin, 98 (4):767–784, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/seasonal-
variation-larval-fish-assemblages-relation-oceanographic-conditions-
abrolhos-bank;   https://spo.nmfs.noaa.gov/sites/default/files/
11.pdf.

Sarre:2000:VAC

[4375] Gavin A. Sarre and Ian C. Potter. Variation in age compositions and growth rates of Acanthopagrus butcheri (Sparidae) among estuaries: some possible contributing factors. Fishery Bulletin, 98(4):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/variation-age-
compositions-and-growth-rates-acanthopagrus-butcheri-sparidae-
among-estuaries;   https://spo.nmfs.noaa.gov/sites/default/
files/12.pdf.

Secor:2000:DGY

[4376] David H. Secor, Edwin J. Niklitschek, Jill T. Stevenson, Troy E. Gunderson, Steven P. Minkkinen, Brian Richardson, Benjamin Florence, Michael Mangold, Jorgen Skjeveland, and Anne Henderson-Arzapalo. Dispersal and growth of yearling Atlantic sturgeon, Acipenser oxyrinchus, released into cheasapeake bay. Fishery Bulletin, 98(4): ??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/dispersal-
and-growth-yearling-atlantic-sturgeon-acipenser-oxyrinchus-
released-cheasapeake;   https://spo.nmfs.noaa.gov/sites/default/
files/13_0.pdf.

Simpfendorfer:2000:GRJ

[4377] Colin A. Simpfendorfer. Growth rates of juvenile dusky sharks, Carcharhinus obscurus (Lesueur, 1818), from southwestern Australia estimated from tag-recapture data. Fishery Bulletin, 98(4):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/growth-rates-
juvenile-dusky-sharks-carcharhinus-obscurus-lesueur-1818-southwestern;
 https://spo.nmfs.noaa.gov/sites/default/files/14.pdf.

Wilson:2000:EYR

[4378] Matthew T. Wilson. Effects of year and region on the abundance and size of age-0 walleye pollock, Theragra chalcogramma, in the western Gulf of Alaska, 1985—1988. Fishery Bulletin, 98(4):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/effects-year-
and-region-abundance-and-size-age-0-walleye-pollock-theragra-
chalcogramma;   https://spo.nmfs.noaa.gov/sites/default/files/
15.pdf.

Xiao:2000:UOB

[4379] Yongshun Xiao. Use of the original von Bertalanffy growth model to describe the growth of barramundi, Lates calcarifer (Bloch). Fishery Bulletin, 98(4):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/use-original-von-bertalanffy-growth-model-
describe-growth-barramundi-lates-calcarifer-bloch;   https:
//spo.nmfs.noaa.gov/sites/default/files/16.pdf.

Zenitani:2000:RBS

[4380] Hiromu Zenitani and Sakutaro Yamada. The relation between spawning area and biomass of Japanese pilchard, Sardinops melanostictus, along the Pacific coast of Japan. Fishery Bulletin, 98(4):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/relation-between-spawning-area-
and-biomass-japanese-pilchard-sardinops-melanostictus-along;
 https://spo.nmfs.noaa.gov/sites/default/files/17.pdf.

Ignell:2000:GPG

[4381] Steve Ignell and Evan Haynes. Geographic patterns in growth of the giant Pacific sea scallop,Patinopecten caurinus. Fishery Bulletin, 98(4):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/geographic-patterns-
growth-giant-pacific-sea-scalloppatinopecten-caurinus;   https:
//spo.nmfs.noaa.gov/sites/default/files/18.pdf.

Johnson:2000:PEM

[4382] Donna L. Johnson. Preliminary examination of the match-mismatch hypothesis and recruitment variability of yellowtail flounder,Limanda ferruginea. Fishery Bulletin, 98(4):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/preliminary-examination-match-mismatch-hypothesis-
and-recruitment-variability-yellowtail;   https://spo.nmfs.
noaa.gov/sites/default/files/19.pdf.

Masuda:2000:OAB

[4383] Reiji Masuda and Katsumi Tsukamoto. Onset of association behavior in striped jack, Pseudocaranx dentex, in relation to floating objects. Fishery Bulletin, 98(4):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
onset-association-behavior-striped-jack-pseudocaranx-dentex-
relation-floating-objects;   https://spo.nmfs.noaa.gov/sites/
default/files/20.pdf.

Sarver:2000:OPB

[4384] Shane K. Sarver, D. Wilson Freshwater, and Patrick J. Walsh. The occurrence of the provisional Brazilian subspecies of spiny lobster (Panulirus argus westonii) in Florida waters. Fishery Bulletin, 98(4):??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/occurrence-
provisional-brazilian-subspecies-spiny-lobster-panulirus-argus-
westonii-florida;   https://spo.nmfs.noaa.gov/sites/default/
files/21.pdf.

Wang:2000:CGC

[4385] You-Gan Wang and David A. Milton. On comparison of growth curves: How do we test whether growth rates differ? Fishery Bulletin, 98(4): ??, 2000. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/comparison-
growth-curves-how-do-we-test-whether-growth-rates-differ;
https://spo.nmfs.noaa.gov/sites/default/files/22.pdf.

Wing:2000:OPS

[4386] Bruce L. Wing, James M. Murphy, and Thomas L. Rutecki. Occurrence of Pacific sardine, Sardinops sagax, off southeastern Alaska. Fishery Bulletin, 98(4):881–883, 2000. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
occurrence-pacific-sardine-sardinops-sagax-southeastern-alaska;
 https://spo.nmfs.noaa.gov/sites/default/files/23.pdf.

Anonymous:2001:SFa

[4387] Anonymous. Subscription form. Fishery Bulletin, 99(1):??, 2001. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/node/14440;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/subform%20(5)
.pdf.

Begg:2001:UIO

[4388] Gavin A. Begg, William J. Overholtz, and Nancy J. Munroe. The use of internal otolith morphometrics for identification of haddock (Melanogrammus aeglefinus) stocks on Georges Bank. Fishery Bulletin, 99(1):1–14, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/use-
internal-otolith-morphometrics-identification-haddock-melanogrammus-
aeglefinus-stocks;   https://spo.nmfs.noaa.gov/sites/default/
files/1.pdf.

Brown-Peterson:2001:RBC

[4389] Nancy J. Brown-Peterson, Robin M. Overstreet, Jeffrey M. Lotz, James S. Franks, and Karen M. Burns. Reproductive biology of cobia, Rachycentron canadum, from coastal waters of the southern United States. Fishery Bulletin, 99(1):??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/reproductive-biology-cobia-rachycentron-canadum-
coastal-waters-southern-united-states;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2001/991/Brown-Peterson.
pdf.

Caretta:2001:ADD

[4390] James V. Caretta, Barbara L. Taylor, and Susan J. Chivers. Abundance and depth distribution of harbor porpoise (Phocoena phocoena) in northern California determined from a 1995 ship survey. Fishery Bulletin, 99(1):29–39, 2001. CODEN FSYBAY.

ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/abundance-and-depth-distribution-
harbor-porpoise-phocoena-phocoena-northern-california;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2001/991/
Carretta.pdf.

Dagorn:2001:AYT

[4391] Laurent Dagorn, Erwan Josse, and Pascal Bach. Association of yellowfin tuna (Thunnus albacares) with tracking vessels during ultrasonic telemetry experiments. Fishery Bulletin, 99(1):40–48, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/association-
yellowfin-tuna-thunnus-albacares-tracking-vessels-during-ultrasonic-
telemetry;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2001/991/Dagorn.pdf.

Gharrett:2001:IRS

[4392] Anthony J. Gharrett, Andrew K. Gray, and Jonathan Heifetz. Identification of rockfish (Sebastes spp.) by restriction site analysis of the mitochondrial ND-3/ND-4 and 12S/16S rRNA gene region. Fishery Bulletin, 99(1):??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/identification-rockfish-sebastes-
spp-restriction-site-analysis-mitochondrial-nd-3nd-4-and;
https://spo.nmfs.noaa.gov/sites/default/files/5.pdf.

Harvey:2001:CPA

[4393] Euan Harvey, David Fletcher, and Mark Shortis. A comparison of the precision and accuracy of estimates of reef-fish lengths determined visually by divers with estimates produced by a stereovideo system. Fishery Bulletin, 99(1):63–71, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/comparison-precision-and-
accuracy-estimates-reef-fish-lengths-determined-visually-divers;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2001/991/Harvey.pdf.

Harvey:2001:ISP

[4394] Euan Harvey, David Fletcher, and Mark Shortis. Improving the statistical power of length estimates of reef fish: a comparison of estimates determined visually by divers with estimates produced by a stereo-video system. Fishery Bulletin, 99(1):72–80, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic).  URL https://spo.nmfs.noaa.gov/content/improving-
statistical-power-length-estimates-reef-fish-comparison-estimates-
determined;   https://spo.nmfs.noaa.gov/sites/default/files/
7.pdf.

Lazzari:2001:DLF

[4395] Mark A. Lazzari. Dynamics of larval fish abundance in Penobscot Bay, Maine. Fishery Bulletin, 99(1):??, 2001. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/dynamics-larval-fish-abundance-
penobscot-bay-maine;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2001/991/Lazzari.pdf.

Nelson:2001:PSF

[4396] R. John Nelson, Maureen P. Small, Terry D. Beacham, and K. Janine Supernault. Population structure of Fraser River Chinook salmon (Oncorhynchus tshawytscha): an analysis using microsatellite DNA markers. Fishery Bulletin, 99(1):94–107, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/population-structure-fraser-river-chinook-
salmon-oncorhynchus-tshawytscha-analysis-using;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/2001/991/Nelson.
pdf.

Nichol:2001:ABF

[4397] Daniel G. Nichol and Erika I. Acuna. Annual and batch fecundities of yellowfin sole, Limanda aspera, in the eastern Bering Sea. Fishery Bulletin, 99(1):108–122, 2001. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
annual-and-batch-fecundities-yellowfin-sole-limanda-aspera-
eastern-bering-sea;   https://spo.nmfs.noaa.gov/sites/default/
files/10.pdf.

Noll:2001:ACG

[4398] Claire Noll, Natalia V. Varnavskaya, Evgeny A. Matzak, Sharon L. Hawkins, Victoria V. Midanaya, Oleg N. Katugin, Charles Russell, Natalya M. Kinas, Charles M. Guthrie III, Hiroshi Mayama, Fumio Yamazaki, Bruce P. Finney, and Anthony J. Gharrett. Analysis of contemporary genetic structure of even-broodyear populations of Asian and western Alaskan pink salmon, Oncorhynchus gorbuscha. Fishery Bulletin, 99(1):123–138, 2001. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
analysis-contemporary-genetic-structure-even-broodyear-populations-
asian-and-western;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2001/991/Noll.pdf.

Panfili:2001:VAE

[4399] Jacques Panfili and Javier Tomás. Validation of age estimation and back-calculation of fish length based on otolith microstructures in tilapias (Pisces, Cichlidae). Fishery Bulletin, 99(1):??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/validation-
age-estimation-and-back-calculation-fish-length-based-otolith-
microstructures;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2001/991/Panfi.pdf.

Pella:2001:BMA

[4400] Jerome Pella and Michele Masuda. Bayesian methods for analysis of stock mixtures from genetic characters. Fishery Bulletin, 99(1):151– 167, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/bayesian-
methods-analysis-stock-mixtures-genetic-characters;   https:
//spo.nmfs.noaa.gov/sites/default/files/13.pdf.

Soh:2001:PRM

[4401] SungKwon Soh, Donald R. Gunderson, and Daniel H. Ito. The potential role of marine reserves in the management of shortraker rockfish (Sebastes borealis) and rougheye rockfish (S. aleutianus) in the Gulf of Alaska. Fishery Bulletin, 99(1):168–179, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/potential-role-marine-reserves-management-
shortraker-rockfish-sebastes-borealis-and;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2001/991/Soh.pdf.

StJohn:2001:DLC

[4402] Jill St John, Garry R. Russ, Ian W. Brown, and Lyle C. Squire. The diet of the large coral reef serranid Plectropomus leopardus in two fishing zones on the Great Barrier Reef, Australia. Fishery Bulletin, 99(1): ??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/diet-large-
coral-reef-serranid-plectropomus-leopardus-two-fishing-zones-
great-barrier-reef;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2001/991/John.pdf.

Adam:2001:ETS

[4403] M. Shiham Adam and Geoffrey P. Kirkwood. Estimating tag-shedding rates for skipjack tuna, Katsuwonus pelamis, off the Maldives. Fishery Bulletin, 99(1):??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
estimating-tag-shedding-rates-skipjack-tuna-katsuwonus-pelamis-
maldives;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2001/991/Adam.pdf.

DeMaster:2001:ESS

[4404] Douglas P. DeMaster, Lloyd F. Lowry, Kathryn J. Frost, and Rebecca A. Bengston. The effect of sea state on estimates of abundance for beluga whales (Delphinapterus leucas) in Norton Sound, Alaska. Fishery Bulletin, 99(1):??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/effect-sea-state-estimates-abundance-
beluga-whales-delphinapterus-leucas-norton-sound;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/2001/991/
DeMaster.pdf.

Pierce:2001:AGS

[4405] Daryl J. Pierce, Behzad Mahmoudi, and Raymond R. Wilson, Jr. Age and growth of the scaled herring, Harengula jaguana, from Florida waters, as indicated by microstructure of the sagittae. Fishery Bulletin, 99 (1):202–209, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/age-
and-growth-scaled-herring-harengula-jaguana-florida-waters-
indicated-microstructure;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2001/991/Pierce.pdf.

Stone:2001:CCS

[4406] Heath H. Stone and Langille K. Dixon. A comparison of catches of swordfish, Xiphias gladius, and other pelagic species from Canadian longline gear configured with alternating monofilament and multifilament nylon gangions. Fishery Bulletin, 99(1):??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/comparison-catches-swordfish-xiphias-
gladius-and-other-pelagic-species-canadian-longline;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2001/991/
Stone.pdf.

Zimmermann:2001:DPN

[4407] Mark Zimmermann, Robin C. Harrison, and Anthony F. Jones. Differential parasitism by Naobranchia occidentalis (Copepoda: Naobranchiidae) and Nectobrachia indivisa (Copepoda: Lernaeopodidae) on northern rock sole (Lepidopsetta polyxystra Orr and Matarese, 2000) and southern rock sole (L. bilineata Ayres, 1855) in Alaskan waters. Fishery Bulletin, 99(2):371–380, 2001. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
differential-parasitism-naobranchia-occidentalis-copepoda-naobranchiidae-
and-nectobrachia;   https://spo.nmfs.noaa.gov/node/14441;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/zim.pdf;   https://spo.nmfs.noaa.gov/sites/default/files/
zim.pdf.

Anonymous:2001:SFb

[4408] Anonymous. Subscription form. Fishery Bulletin, 99(2):??, 2001. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/node/14442;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/subform%20(6)
.pdf.

Baumgartner:2001:CHN

[4409] Mark F. Baumgartner, Keith D. Mullin, L. Nelson May, and Thomas D. Leming. Cetacean habitats in the northern Gulf of Mexico. Fishery Bulletin, 99(2):??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
cetacean-habitats-northern-gulf-mexico;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2001/992/bau.pdf.

Bjorndal:2001:SGF

[4410] Karen A. Bjorndal, Alan B. Bolten, Bruce Koike, Barbara A. Schroeder, Donna J. Shaver, Wendy G. Teas, and Wayne N. Witzell. Somatic growth function for immature loggerhead sea turtles, Caretta caretta, in southeastern U.S. waters. Fishery Bulletin, 99(2):??, 2001. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/somatic-growth-function-immature-
loggerhead-sea-turtles-caretta-caretta-southeastern-us.

Bulllimore:2001:SCF

[4411] Blaise A. Bulllimore, Philip B. Newman, Michel J. Kaiser, Susanne E. Gilbert, and Kate M. Lock. A study of catches in a fleet of “ghost-fishing” pots. Fishery Bulletin, 99(2):??, 2001. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/study-catches-fleet-%
E2%80%9Cghost-fishing%E2%80%9D-pots;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2001/992/bul.pdf.

Burton:2001:AGM

[4412] Michael L. Burton. Age, growth, and mortality of gray snapper, Lutjanus griseus, from the east coast of Florida. Fishery Bulletin, 99(2):??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/age-growth-and-mortality-
gray-snapper-lutjanus-griseus-east-coast-florida.

Carrasson:2001:FEM

[4413] Maite Carrassón and Jesús Matallanas. Feeding ecology of the Mediterranean spiderfish, Bathypterois mediterraneus (Pisces: Chlorophthalmidae), on the western Mediterranean slope. Fishery Bulletin, 99(2): ??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/feeding-
ecology-mediterranean-spiderfish-bathypterois-mediterraneus-
pisces.

Choi:2001:LDK

[4414] Jung H. Choi and Sung Y. Hong. Larval development of the kishi velvet shrimp, Metapenaeopsis dalei (Rathbun) (Decapoda: Penaeidae), reared in the laboratory. Fishery Bulletin, 99(2):??, 2001. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/larval-development-kishi-velvet-
shrimp-metapenaeopsis-dalei-rathbun-decapoda-penaeidae;   https:
//spo.nmfs.noaa.gov/sites/default/files/cho.pdf.

Hobson:2001:IVP

[4415] Edmund S. Hobson, James R. Chess, and Daniel F. Howard. Interannual variation in predation on first-year Sebastes spp. by three northern California predators. Fishery Bulletin, 99(2):292–302, 2001. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/interannual-variation-predation-
first-year-sebastes-spp-three-northern-california-predators.

Hooker:2001:BRN

[4416] Sascha K. Hooker, Robin W. Baird, Sa’ad Al-Omari, Shannon Gowans, and Hal Whitehead. Behavioral reactions of northern bottlenose whales (Hyperoodon ampullatus) to biopsy darting and tag attachment procedures. Fishery Bulletin, 99(2):??, 2001. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/behavioral-reactions-
northern-bottlenose-whales-hyperoodon-ampullatus-biopsy-darting-
and.

Jacobson:2001:DDT

[4417] Larry D. Jacobson, Jon Brodziak, and Jean Rogers. Depth distributions and time-varying bottom trawl selectivities for Dover sole (Microstomus pacificus), sablefish (Anoplopoma fimbria), and thornyheads (Sebastolobus alascanus and S. altivelis) in a commercial fishery. Fishery Bulletin, 99(2):309–327, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/depth-distributions-and-time-varying-
bottom-trawl-selectivities-dover-sole-microstomus;   https:
//spo.nmfs.noaa.gov/sites/default/files/jac.pdf.

Jones:2001:YPR

[4418] Cynthia M. Jones and Brian K. Wells. Yield-per-recruit analysis for black drum, Pogonias cromis, along the east coast of the United States and management strategies for Chesapeake Bay. Fishery Bulletin, 99(2): ??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/yield-recruit-
analysis-black-drum-pogonias-cromis-along-east-coast-united-
states-and;   https://spo.nmfs.noaa.gov/sites/default/files/
jon.pdf.

Salthaug:2001:ACT

[4419] Are Salthaug. Adjustment of commercial trawling effort for Atlantic cod, Gadus morhua, due to increasing catching efficiency. Fishery Bulletin, 99(2):??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/adjustment-commercial-trawling-effort-atlantic-cod-
gadus-morhua-due-increasing-catching;   https://spo.nmfs.noaa.
gov/sites/default/files/sal.pdf.

Schaefer:2001:AST

[4420] Kurt M. Schaefer. Assessment of skipjack tuna (Katsuwonus pelamis) spawning activity in the eastern Pacific Ocean. Fishery Bulletin, 99(2): ??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/assessment-
skipjack-tuna-katsuwonus-pelamis-spawning-activity-eastern-
pacific-ocean;   https://spo.nmfs.noaa.gov/sites/default/files/
sch.pdf.

Smith:2001:BIS

[4421] Peter J. Smith and Peter G. Benson. Biochemical identification of shark fins and fillets from the coastal fisheries in New Zealand. Fishery Bulletin, 99(2):??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/biochemical-identification-shark-fins-and-
fillets-coastal-fisheries-new-zealand.

Yoneda:2001:RCF

[4422] Michio Yoneda, Muneharu Tokimura, Hitoshi Fujita, Naohiko Takeshita, Koji Takeshita, Michiya Matsuyama, and Shuhei Matsuura. Reproductive cycle, fecundity, and seasonal distribution of the anglerfish Lophius litulon in the East China and Yellow seas. Fishery Bulletin, 99(2):??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/reproductive-cycle-fecundity-and-seasonal-
distribution-anglerfish-lophius-litulon-east;   https://spo.
nmfs.noaa.gov/sites/default/files/yon.pdf.

Davis:2001:SDS

[4423] Tim L. O. Davis and Jessica H. Farley. Size distribution of southern bluefin tuna (Thunnus maccoyii) by depth on their spawning ground. Fishery Bulletin, 99(2):??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/size-distribution-southern-bluefin-tuna-thunnus-maccoyii-
depth-their-spawning-ground.

Anonymous:2001:SFc

[4424] Anonymous. Subscription form. Fishery Bulletin, 99(3):??, 2001. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/node/14443;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/subform%20(7)
.pdf.

Andrews:2001:RAV

[4425] Allen H. Andrews, Erica J. Burton, Kenneth H. Coale, Gregor M. Cailliet, and Roy E. Crabtree. Radiometric age validation of Atlantic tarpon, Megalops atlanticus. Fishery Bulletin, 99(3):??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/radiometric-
age-validation-atlantic-tarpon-megalops-atlanticus;   https:
//spo.nmfs.noaa.gov/sites/default/files/and.pdf.

Colura:2001:ESB

[4426] Robert L. Colura and Britt W. Bumguardner. Effect of the salt-box catch-bycatch separation procedure, as used by the Texas shrimp industry, on short-term survival of bycatch. Fishery Bulletin, 99(3):??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/effect-salt-
box-catch-bycatch-separation-procedure-used-texas-shrimp-industry-
short-term;   https://spo.nmfs.noaa.gov/sites/default/files/
col.pdf.

Gillanders:2001:TMF

[4427] Bronwyn M. Gillanders. Trace metals in four structures of fish and their use for estimates of stock structure. Fishery Bulletin, 99(3):??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/trace-metals-
four-structures-fish-and-their-use-estimates-stock-structure;
 https://spo.nmfs.noaa.gov/sites/default/files/gil.pdf.

Hanrahan:2001:ENF

[4428] Brian Hanrahan and Francis Juanes. Estimating the number of fish in Atlantic bluefin tuna (Thunnus thynnus thynnus) schools using models derived from captive school observations. Fishery Bulletin, 99(3):??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/estimating-
number-fish-atlantic-bluefin-tuna-thunnus-thynnus-thynnus-schools-
using-models;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2001/993/han.pdf.

Lucena:2001:EGS

[4429] Flávia M. Lucena and Carl M. O’Brien. Effects of gear selectivity and different calculation methods on estimating growth parameters of bluefish, Pomatomus saltatrix (Pisces: Pomatomidae), from southern Brazil. Fishery Bulletin, 99(3):??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/effects-gear-selectivity-and-different-calculation-
methods-estimating-growth-parameters;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2001/993/luc.pdf.

McBride:2001:NHL

[4430] Richard S. McBride, Timothy C. MacDonald, Richard E. Matheson, Jr., David A. Rydene, and Peter B. Hood. Nursery habitats for ladyfish, Elops saurus, along salinity gradients in two Florida estuaries. Fishery Bulletin, 99(3):??, 2001. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/nursery-habitats-ladyfish-elops-
saurus-along-salinity-gradients-two-florida-estuaries;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2001/993/
mcb.pdf.

McFarlane:2001:VFR

[4431] Gordon A. McFarlane and Jacquelynne R. King. The validity of the finray method of age determination for lingcod (Ophiodon elongatus). Fishery Bulletin, 99(3):??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
validity-fin-ray-method-age-determination-lingcod-ophiodon-
elongatus;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2001/993/mcf.pdf.

Rabe:2001:BGS

[4432] Jessica Rabe and Joseph A. Brown. The behavior, growth, and survival of witch flounder (Glyptocephalus cynoglossus) larvae in relation to prey availability: adaptations to an extended larval period. Fishery Bulletin, 99(3):??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
behavior-growth-and-survival-witch-flounder-glyptocephalus-
cynoglossus-larvae-relation-prey;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2001/993/rab.pdf.

Schubart:2001:LDB

[4433] Christoph D. Schubart, Jesús E. Conde, Carlos Carmona-Suárez, Rafael Robles, and Darryl L. Felder. Lack of divergence between 16S mtDNA sequences of the swimming crabs Callinectes bocourti and C. maracaiboensis (Brachyura: Portunidae) from Venezuela. Fishery Bulletin, 99(3):??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
lack-divergence-between-16s-mtdna-sequences-swimming-crabs-
callinectes-bocourti-and-c;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2001/993/sch.pdf.

Sturdevant:2001:FHP

[4434] Molly V. Sturdevant, Audra L. J. Brase, and Leland B. Hulbert. Feeding habits, prey fields, and potential competition of young-of-the-year walleye pollock (Theragra chalcogramma) and Pacific herring (Clupea pallasi) in Prince William@ Sound, Alaska, 1994–1995. Fishery Bulletin, 99(3):??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
feeding-habits-prey-fields-and-potential-competition-young-
year-walleye-pollock-theragra;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2001/993/stu.pdf.

Sun:2001:AGB

[4435] Chi-Lu Sun, Chien-Lung Huang, and Su-Zan Yeh. Age and growth of the bigeye tuna, Thunnus obesus, in the western Pacific Ocean. Fishery Bulletin, 99(3):502–509, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/age-and-growth-bigeye-tuna-thunnus-obesus-
western-pacific-ocean;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2001/993/sun.pdf.

Takahashi:2001:NOS

[4436] Kazutaka Takahashi and Kouichi Kawaguchi. Nocturnal occurrence of the swimming crab Ovalipes punctatus in the swash zone of a sandy beach in northeastern Japan. Fishery Bulletin, 99(3):??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/nocturnal-occurrence-
swimming-crab-ovalipes-punctatus-swash-zone-sandy-beach-northeastern;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2001/993/tak.pdf.

Anonymous:2001:AR

[4437] Anonymous. Acknowledgment of reviewers. Fishery Bulletin, 99 (4):??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14444;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/ack.pdf.

Anonymous:2001:I

[4438] Anonymous. Index. Fishery Bulletin, 99(4):??, 2001. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14445;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/ind.pdf.

Anonymous:2001:E

[4439] Anonymous. Erratum. Fishery Bulletin, 99(4):??, 2001. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14446;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/err.pdf.

Anonymous:2001:SFd

[4440] Anonymous. Subscription form. Fishery Bulletin, 99(4):??, 2001. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/node/14447;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/subform%20(8)
.pdf.

Arendt:2001:SOS

[4441] Michael D. Arendt, Jon A. Lucy, and Thomas A. Munroe. Seasonal occurrence and site-utilization patterns of adult tautog, Tautoga onitis (Labridae), at manmade and natural structures in lower Chesapeake Bay. Fishery Bulletin, 99(4):519–527, 2001. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/seasonal-occurrence-
and-site-utilization-patterns-adult-tautog-tautoga-onitis-labridae;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2001/994/are.pdf.

Gharrett:2001:PAM

[4442] Anthony J. Gharrett, Andrew K. Gray, and Vladimir Brykov. Phylogeographic analysis of mitochondrial DNA variation in Alaskan Coho salmon, Oncorhynchus kisutch. Fishery Bulletin, 99(4):529–544, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/phylogeographic-analysis-
mitochondrial-dna-variation-alaskan-coho-salmon-oncorhynchus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2001/994/gha.pdf.

Hurst:2001:DCR

[4443] Thomas P. Hurst and David O. Conover. Diet and consumption rates of overwintering YOY striped bass, Morone saxatilis, in the Hudson River. Fishery Bulletin, 99(4):??, 2001. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/diet-and-consumption-
rates-overwintering-yoy-striped-bass-morone-saxatilis-hudson-
river;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2001/994/hur.pdf.

Lo:2001:UCE

[4444] Nancy C. H. Lo, John R. Hunter, and Richard Charter. Use of a continuous egg sampler for ichthyoplankton surveys: application to the estimation of daily egg production of Pacific sardine (Sardinops sagax) off California. Fishery Bulletin, 99(4):554– 571, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/use-
continuous-egg-sampler-ichthyoplankton-surveys-application-
estimation-daily-egg;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2001/994/lo.pdf.

Loher:2001:GJR

[4445] Timothy Loher, David A. Armstrong, and Bradley G. Stevens. Growth of juvenile red king crab (Paralithodes camtschaticus) in Bristol Bay (Alaska) elucidated from field sampling and analysis of trawl-survey data. Fishery Bulletin, 99(4):572–587, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/growth-juvenile-red-king-
crab-paralithodes-camtschaticus-bristol-bay-alaska-elucidated;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2001/994/loh.pdf.

MacNair:2001:AGM

[4446] Leslie S. MacNair, Michael L. Domeier, and Calvin S. Y. Chun. Age, growth, and mortality of California halibut, Paralichthys californicus, along southern and central California. Fishery Bulletin, 99(4):588–600, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/age-growth-
and-mortality-california-halibut-paralichthys-californicus-
along-southern-and;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2001/994/mac.pdf.

Miyashita:2001:MDG

[4447] Shigeru Miyashita, Yoshifumi Sawada, Tokihiko Okada, Osamu Murata, and Hidemi Kumai. Morphological development and growth of laboratory-reared larval and juvenile Thunnus thynnus (Pisces: Scombridae). Fishery Bulletin, 99(4):??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/morphological-development-and-growth-
laboratory-reared-larval-and-juvenile-thunnus-thynnus;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2001/994/
miy.pdf.

Patterson:2001:AGR

[4448] William F. Patterson III, James H. Cowan, Jr., Charles A. Wilson, and Robert L. Shipp. Age and growth of red snapper, Lutjanus campechanus, from an artificial reef area off Alabama in the northern Gulf of Mexico. Fishery Bulletin, 99(4):617–627, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/age-and-growth-red-
snapper-lutjanus-campechanus-artificial-reef-area-alabama-northern-
gulf;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2001/994/pat.pdf.

Sipe:2001:CCS

[4449] Ann M. Sipe and Mark E. Chittenden, Jr. A comparison of calcified structures for aging summer flounder, Paralichthys dentatus. Fishery Bulletin, 99(4):??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
comparison-calcified-structures-aging-summer-flounder-paralichthys-
dentatus;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2001/994/sip.pdf.

Somerton:2001:BES

[4450] David A. Somerton and Peter Munro. Bridle efficiency of a survey trawl for flatfish. Fishery Bulletin, 99(4):??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/bridle-efficiency-survey-trawl-flatfish;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/2001/
994/som.pdf.

Wilson:2001:AGR

[4451] Charles Wilson and David L. Nieland. Age and growth of red snapper, Lutjanus campechanus, from the northern Gulf of Mexico off Louisiana. Fishery Bulletin, 99(4):??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/age-and-growth-red-snapper-lutjanus-campechanus-northern-
gulf-mexico-louisiana;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2001/994/wil.pdf.

Arendt:2001:SCA

[4452] Michael D. Arendt, John E. Olney, and Jon A. Lucy. Stomach content analysis of cobia, Rachycentron canadum, from lower Chesapeake Bay. Fishery Bulletin, 99(4):665–670, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/stomach-content-analysis-
cobia-rachycentron-canadum-lower-chesapeake-bay;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/2001/994/
are2.pdf.

Blandon:2001:PGP

[4453] Ivonne R. Blandon, Rocky Ward, Tim L. King, William J. Karel, and James P. Monaghan, Jr. Preliminary genetic population structure of southern flounder, Paralichthys lethostigma, along the Atlantic Coast and Gulf of Mexico. Fishery Bulletin, 99(4):??, 2001. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/preliminary-genetic-population-
structure-southern-flounder-paralichthys-lethostigma-along;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2001/994/bla.pdf.

Hernandez-Lopez:2001:ADD

[4454] Jose L. Hernández-López, Jose J. Castro-Hernández, and Vicente Hernández-García. Age determined from the daily deposition of concentric rings on common octopus (Octopus vulgaris) beaks. Fishery Bulletin, 99(4):679–684, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/age-
determined-daily-deposition-concentric-rings-common-octopus-
octopus-vulgaris-beaks;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2001/994/her.pdf.

Jones:2001:RFS

[4455] Thomas S. Jones and Karl I. Ugland. Reproduction of female spiny dogfish, Squalus acanthias, in the Oslofjord. Fishery Bulletin, 99(4): ??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/reproduction-
female-spiny-dogfish-squalus-acanthias-oslofjord;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/2001/994/
jon.pdf.

Porter:2001:ELS

[4456] Steven M. Porter, Annette L. Brown, and Kevin M. Bailey. Estimating live standard length of net-caught walleye pollock (Theragra chalcogramma) larvae using measurements in addition to standard length. Fishery Bulletin, 99(4):691–696, 2001. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/estimating-live-standard-length-
net-caught-walleye-pollock-theragra-chalcogramma-larvae;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2001/994/
por.pdf.

Takagi:2001:PSA

[4457] Motohiro Takagi, Tetsuro Okamura, Seinen Chow, and Nobuhiko Taniguchi. Preliminary study of albacore (Thunnus alalunga) stock differentiation inferred from microsatellite DNA analysis. Fishery Bulletin, 99(4):??, 2001. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
preliminary-study-albacore-thunnus-alalunga-stock-differentiation-
inferred-microsatellite;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2001/994/tak.pdf.

Anonymous:2002:SFa

[4458] Anonymous. Subscription form. Fishery Bulletin, 100(1):??, 2002. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/node/14448;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/subform%20(9)
.pdf.

Blick:2002:UAM

[4459] D. James Blick and Peter T. Hagen. The use of agreement measures and latent class models to assess the reliability of classifying thermally marked otoliths. Fishery Bulletin, 100(1):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/use-agreement-
measures-and-latent-class-models-assess-reliability-classifying-
thermally;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2002/1001/bli.pdf.

Carmona-Suarez:2002:LDA

[4460] Carlos A. Carmona-Suarez and Jesus E. Conde. Local distribution and abundance of swimming crabs (Callinectes spp. and Arenaeus cribrarius) on a tropical arid beach. Fishery Bulletin, 100(1):??, 2002. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/local-distribution-and-abundance-
swimming-crabs-callinectes-spp-and-arenaeus-cribrarius;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2002/1001/
car.pdf.

Crabtree:2002:AGR

[4461] Roy E. Crabtree, Peter B. Hood, and Derke Snodgrass. Age, growth, and reproduction of permit (Trachinotus falcatus) in Florida waters. Fishery Bulletin, 100(1):26–34, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/age-growth-and-reproduction-permit-trachinotus-
falcatus-florida-waters;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2002/1001/cra.pdf.

Denson:2002:TRL

[4462] Michael R. Denson, Wallace E. Jenkins, Arnold G. Woodward, and Theodore I. J. Smith. Tag-reporting levels for red drum (Sciaenops ocellatus) caught by anglers in South Carolina and Georgia estuaries. Fishery Bulletin, 100(1):??, 2002. CODEN FSY-

BAY.  ISSN 0090-0656 (print), 1937-4518 (electronic).  URL

https://spo.nmfs.noaa.gov/content/tag-reporting-levels-red-
drum-sciaenops-ocellatus-caught-anglers-south-carolina-and-
georgia;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2002/1001/den.pdf.

Faunce:2002:AGM

[4463] Craig H. Faunce, Heather M. Patterson, and Jerome J. Lorenz. Age, growth, and mortality of the Mayan cichlid (Cichlasoma urophthalmus) from the southeastern Everglades. Fishery Bulletin, 100(1): 42–50, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/age-
growth-and-mortality-mayan-cichlid-cichlasoma-urophthalmus-
southeastern-everglades;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2002/1001/fau.pdf.

Hastings:2002:PSS

[4464] Kelly K. Hastings and William J. Sydeman. Population status, seasonal variation in abundance, and long-term population trends of Steller sea lions (Eumetopias jubatus) at the South Farallon Islands, California. Fishery Bulletin, 100(1):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/population-status-seasonal-variation-abundance-
and-long-term-population-trends-steller-sea;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/2002/1001/has.
pdf.

McBride:2002:LSP

[4465] Richard S. McBride, Michael P. Fahay, and Kenneth W. Able. Larval and settlement periods of the northern searobin (Prionotus carolinus) and the striped searobin (P. evolans). Fishery Bulletin, 100(1):63–73, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/larval-and-
settlement-periods-northern-searobin-prionotus-carolinus-and-

striped-searobin-p;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2002/1001/mcb.pdf.

Pennington:2002:APF

[4466] Michael Pennington, Liza-Maré Burmeister, and Vidar Hjellvik. Assessing the precision of frequency distributions estimated from trawl-survey samples. Fishery Bulletin, 100(1):74–80, 2002. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/assessing-precision-
frequency-distributions-estimated-trawl-survey-samples;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2002/1001/
pen.pdf.

Potts:2002:EAR

[4467] Jennifer C. Potts and Charles S. Manooch III. Estimated ages of red porgy (Pagrus pagrus) from fishery-dependent and fishery-independent data and a comparison of growth parameters. Fishery Bulletin, 100(1): 81–89, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/estimated-
ages-red-porgy-pagrus-pagrus-fishery-dependent-and-fishery-
independent-data-and;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2002/1001/pot.pdf.

Romanov:2002:BTT

[4468] Evgeny V. Romanov. Bycatch taken by the tuna purse-seine fishery from the Indian Ocean. Fishery Bulletin, 100(1):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/bycatch-taken-tuna-purse-seine-
fishery-indian-ocean;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2002/1001/rom.pdf.

Sainte-Marie:2002:OSN

[4469] Bernard Sainte-Marie and Denis Chabot. Ontogenetic shifts in natural diet during benthic stages of American lobster (Homarus americanus), off the Magdalen Islands. Fishery Bulletin, 100(1):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/ontogenetic-
shifts-natural-diet-during-benthic-stages-american-lobster-
homarus-americanus;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2002/1001/sai.pdf.

DiNardo:2002:ELH

[4470] Gerard T. DiNardo, Edward E. DeMartini, and Wayne R. Haight. Estimates of lobster-handling mortality associated with the Northwestern Hawaiian Islands lobster-trap fishery. Fishery Bulletin, 100(1): 128–133, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/estimates-
lobster-handling-mortality-associated-northwestern-hawaiian-
islands-lobster-trap;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2002/1001/din.pdf.

Zug:2002:AGH

[4471] George R. Zug, George H. Balazs, Jerry A. Wetherall, Denise M. Parker, and Shawn K. K. Murakawa. Age and growth of Hawaiian green seaturtles (Chelonia mydas): an analysis based on skeletochronology. Fishery Bulletin, 100(1):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
age-and-growth-hawaiian-green-seaturtles-chelonia-mydas-analysis-
based-skeletochronology;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2002/1001/zug.pdf.

Graves:2002:EPS

[4472] John E. Graves, Brian E. Luckhurst, and Eric D. Prince. An evaluation of pop-up satellite tags for estimating postrelease survival of blue marlin (Makaira nigricans) from a recreational fishery. Fishery Bulletin, 100(1):134–142, 2002. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
evaluation-pop-satellite-tags-estimating-postrelease-survival-
blue-marlin-makaira-nigricans;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2002/1001/gra.pdf.

Hazin:2002:RBS

[4473] Fabio H. V. Hazin, Paulo G. Oliveira, and Matt K. Broadhurst. Reproduction of the blacknose shark (Carcharhinus acronotus) in coastal waters off northeastern Brazil. Fishery Bulletin, 100(1):143–148, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/reproduction-
blacknose-shark-carcharhinus-acronotus-coastal-waters-northeastern-
brazil;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2002/1001/haz.pdf.

Porch:2002:NGM

[4474] Clay E. Porch, Charles A. Wilson, and David L. Nieland. A new growth model for red drum (Sciaenops ocellatus) that accommodates seasonal and ontogenic changes in growth rates. Fishery Bulletin, 100(1):149–152, 2002. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
new-growth-model-red-drum-sciaenops-ocellatus-accommodates-
seasonal-and-ontogenic-changes;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2002/1001/por.pdf.

Anonymous:2002:SFb

[4475] Anonymous. Subscription form. Fishery Bulletin, 100(2):??, 2002. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/node/14449;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/subform%20(12)
.pdf.

Litwack:2002:AT

[4476] Avi Litwack. Avi test. Fishery Bulletin, 100(2):??, 2002. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/avi-test.

Brill:2002:HVM

[4477] Richard Brill, Molly Lutcavage, Greg Metzger, Peter Bushnell, Michael Arendt, Jon Lucy, Cheryl Watson, and David Foley. Horizontal and vertical movements of juvenile bluefin tuna (Thunnus thynnus), in relation to oceanographic conditions of the western North Atlantic, determined with ultrasonic telemetry. Fishery Bulletin, 100(2):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/horizontal-and-vertical-movements-juvenile-bluefin-
tuna-thunnus-thynnus-relation;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2002/1002/brillf.pdf.

Chase:2002:DDA

[4478] Bradford C. Chase. Differences in diet of Atlantic bluefin tuna (Thunnus thynnus) at five seasonal feeding grounds on the New England continental shelf. Fishery Bulletin, 100(2):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/differences-diet-atlantic-bluefin-tuna-thunnus-
thynnus-five-seasonal-feeding-grounds-new;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2002/1002/chasef.pdf.

Comeau:2002:MAL

[4479] Michel Comeau and Fernand Savoie. Movement of American lobster (Homarus americanus) in the southwestern Gulf of St. Lawrence. Fishery Bulletin, 100(2):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
movement-american-lobster-homarus-americanus-southwestern-gulf-
st-lawrence;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2002/1002/comeau.pdf.

Else:2002:STS

[4480] Page Else, Lewis Haldorson, and Kenneth Krieger. Shortspine thornyhead (Sebastolobus alascanus) abundance and habitat associations in the Gulf of Alaska. Fishery Bulletin, 100(2):193–199, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/shortspine-
thornyhead-sebastolobus-alascanus-abundance-and-habitat-associations-
gulf-alaska;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2002/1002/elsefi.pdf.

Hatfield:2002:GTP

[4481] Emma M. C. Hatfield and Steven X. Cadrin. Geographic and temporal patterns in size and maturity of the longfin inshore squid (Loligo pealeii) off the northeastern United States. Fishery Bulletin, 100(2):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
geographic-and-temporal-patterns-size-and-maturity-longfin-
inshore-squid-loligo-pealeii;   https://spo.nmfs.noaa.gov/sites/
default/files/05hatfie.pdf.

Hesp:2002:ASC

[4482] Sybrand A. Hesp, Ian C. Potter, and Norman G. Hall. Age and size composition, growth rate, reproductive biology, and habitats of the West Australian dhufish (Glaucosoma hebraicum) and their relevance to the management of this species. Fishery Bulletin, 100(2):214–227, 2002. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/age-and-size-composition-
growth-rate-reproductive-biology-and-habitats-west-australian;
 https://spo.nmfs.noaa.gov/sites/default/files/06hespfi.pdf.

Kitada:2002:LLS

[4483] Shuichi Kitada and Kiyoshi Tezuka. Longitudinal logbook survey designs for estimating recreational fishery catch, with application to ayu (Plecoglossus altivelis). Fishery Bulletin, 100(2):??, 2002. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/longitudinal-logbook-
survey-designs-estimating-recreational-fishery-catch-application-

ayu;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2002/1002/kitada.pdf.

MacFarlane:2002:PEJ

[4484] R. Bruce MacFarlane and Elizabeth C. Norton. Physiological ecology of juvenile Chinook salmon (Oncorhynchus tshawytscha) at the southern end of their distribution, the San Francisco Estuary and Gulf of the Farallones, California. Fishery Bulletin, 100(2):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/physiological-ecology-juvenile-
chinook-salmon-oncorhynchus-tshawytscha-southern-end-their;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2002/1002/macfar.pdf.

McFee:2002:BDT

[4485] Wayne E. McFee and Sally R. Hopkins-Murphy. Bottlenose dolphin (Tursiops truncatus) strandings in South Carolina, 1992–1996. Fishery Bulletin, 100(2):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
bottlenose-dolphin-tursiops-truncatus-strandings-south-carolina-
19921996;   https://spo.nmfs.noaa.gov/sites/default/files/09mcfeef.
pdf.

Natanson:2002:VAG

[4486] Lisa J. Natanson, Joseph J. Mello, and Steven E. Campana. Validated age and growth of the porbeagle shark (Lamna nasus) in the western North Atlantic Ocean. Fishery Bulletin, 100(2):266–278, 2002. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/validated-age-and-growth-
porbeagle-shark-lamna-nasus-western-north-atlantic-ocean;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/2002/
1002/natans.pdf.

Olson:2002:FHC

[4487] Robert J. Olson and Felipe Galván-Magaña. Food habits and consumption rates of common dolphinfish (Coryphaena hippurus) in the eastern Pacific Ocean. Fishery Bulletin, 100(2):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/food-habits-and-consumption-rates-common-
dolphinfish-coryphaena-hippurus-eastern-pacific;   https://
spo.nmfs.noaa.gov/sites/default/files/11olsonf.pdf.

Powles:2002:RSS

[4488] Perce M. Powles and Stanley M. Warlen. Recruitment season, size, and age of young American eels (Anguilla rostrata) entering an estuary near Beaufort, North Carolina. Fishery Bulletin, 100(2):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/recruitment-season-size-and-age-
young-american-eels-anguilla-rostrata-entering-estuary-near;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2002/1002/powles.pdf.

Shima:2002:CTS

[4489] Michiyo Shima, Anne Babcock Hollowed, and Glenn R. VanBlaricom. Changes over time in the spatial distribution of walleye pollock (Theragra chalcogramma) in the Gulf of Alaska, 1984–1996. Fishery Bulletin, 100(2):307–323, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/changes-over-time-spatial-distribution-walleye-
pollock-theragra-chalcogramma-gulf-alaska;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2002/1002/shimaf.pdf.

Starr:2002:MBS

[4490] Richard M. Starr, John N. Heine, Jason M. Felton, and Gregor M. Cailliet. Movements of bocaccio (Sebastes paucispinis) and greenspotted (S. chlorostictus) rockfishes in a Monterey submarine canyon: implications for the design of marine reserves. Fishery Bulletin, 100(2):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
movements-bocaccio-sebastes-paucispinis-and-greenspotted-s-
chlorostictus-rockfishes;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2002/1002/starrf.pdf.

Steele:2002:EBR

[4491] Philip Steele, Theresa M. Bert, Kristine H. Johnston, and Sandra Levett. Efficiency of bycatch reduction devices in small otter trawls used in the Florida shrimp fishery. Fishery Bulletin, 100(2):338–350, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/efficiency-bycatch-
reduction-devices-small-otter-trawls-used-florida-shrimp-fishery;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2002/1002/steele.pdf.

Vaughan:2002:SDA

[4492] Douglas S. Vaughan and Michael H. Prager. Severe decline in abundance of the red porgy (Pagrus pagrus) population off the southeastern United States. Fishery Bulletin, 100(2):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/severe-decline-abundance-red-porgy-pagrus-
pagrus-population-southeastern-united-states;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/2002/1002/vaugha.
pdf.

Abookire:2002:NDO

[4493] Alisa A. Abookire, John F. Piatt, and Suzann G. Speckman. A nearsurface, daytime occurrence of two mesopelagic fish species (Stenobrachius leucopsarus and Leuroglossus schmidti) in a glacial fjord. Fishery Bulletin, 100(2):376–380, 2002. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/nearsurface-daytime-occurrence-
two-mesopelagic-fish-species-stenobrachius-leucopsarus-and;
 https://spo.nmfs.noaa.gov/sites/default/files/17abooki.pdf.

Link:2002:OFL

[4494] Jason S. Link and Frank P. Almeida. Opportunistic feeding of longhorn sculpin (Myoxocephalus octodecemspinosus): Are scallop fishery discards an important food subsidy for scavengers on Georges Bank? Fishery Bulletin, 100(2):??, 2002. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/opportunistic-feeding-longhorn-
sculpin-myoxocephalus-octodecemspinosus-are-scallop-fishery;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2002/1002/linkfi.pdf.

Romanov:2002:FRY

[4495] Evgeny V. Romanov and Veniamin V. Zamorov. First record of a yellowfin tuna (Thunnus albacares) from the stomach of a longnose lancetfish (Alepisaurus ferox). Fishery Bulletin, 100(2):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/first-record-
yellowfin-tuna-thunnus-albacares-stomach-longnose-lancetfish-
alepisaurus-ferox;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2002/1002/romano.pdf.

Anonymous:2002:E

[4496] Anonymous. Errata. Fishery Bulletin, 100(3):??, 2002. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14450;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/errata%20(1).pdf.

Anonymous:2002:SFc

[4497] Anonymous. Subscription form. Fishery Bulletin, 100(3):??, 2002. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/node/14451;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/subform%20(13)
.pdf.

Allman:2002:TSD

[4498] Robert J. Allman and Churchill B. Grimes. Temporal and spatial dynamics of spawning, settlement, and growth of gray snapper (Lutjanus griseus) from the west Florida shelf as determined from otolith microstructures. Fishery Bulletin, 100(3):??, 2002. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/temporal-and-spatial-
dynamics-spawning-settlement-and-growth-gray-snapper-lutjanus-
griseus;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2002/1003/allman.pdf.

Anderson:2002:CVC

[4499] Todd W. Anderson, Claudine T. Bartels, Mark A. Hixon, Erich Bartels, Mark H. Carr, and Jonathan M. Shenker. Current velocity and catch efficiency in sampling settlement-stage larvae of coral-reef fishes. Fishery Bulletin, 100(3):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/current-velocity-and-catch-efficiency-sampling-
settlement-stage-larvae-coral-reef-fishes;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2002/1003/anders.pdf.

Boje:2002:ISM

[4500] Jesper Boje. Intermingling and seasonal migrations of Greenland halibut (Reinhardtius hippoglossoides) populations determined from tagging studies. Fishery Bulletin, 100(3):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/intermingling-and-seasonal-migrations-greenland-
halibut-reinhardtius-hippoglossoides;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2002/1003/bojefi.pdf.

Browne:2002:IPD

[4501] Patience Browne, Jeffrey L. Laake, and Robert L. DeLong. Improving pinniped diet analyses through identification of multiple skeletal structures in fecal samples. Fishery Bulletin, 100(3):423–433, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/improving-pinniped-
diet-analyses-through-identification-multiple-skeletal-structures-
fecal;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2002/1003/browne.pdf.

Laake:2002:PDC

[4502] Jeffrey L. Laake, Patience Browne, Robert L. DeLong, and Harriet R. Huber. Pinniped diet composition: a comparison of estimation models. Fishery Bulletin, 100(3):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/pinniped-diet-composition-comparison-
estimation-models;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2002/1003/laakef.pdf.

Davis:2002:VHM

[4503] Tim L. O. Davis and Clive A. Stanley. Vertical and horizontal movements of southern bluefin tuna (Thunnus maccoyii) in the Great Australian Bight observed with ultrasonic telemetry. Fishery Bulletin, 100(3):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/vertical-and-horizontal-movements-southern-bluefin-
tuna-thunnus-maccoyii-great-australian;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2002/1003/davisf.pdf.

Epperly:2002:TED

[4504] Sheryan P. Epperly and Wendy G. Teas. Turtle excluder devices — are the escape openings large enough? Fishery Bulletin, 100(3): ??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/turtle-
excluder-devices-are-escape-openings-large-enough;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2002/1003/
epperl.pdf.

Gibbons:2002:HUD

[4505] Mark J. Gibbons, André J. J. Goosen, and Patti A. Wickens. Habitat use by demersal nekton on the continental shelf in the Benguela ecosystem, southern Africa. Fishery Bulletin, 100(3):475–490, 2002.

CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/habitat-use-demersal-
nekton-continental-shelf-benguela-ecosystem-southern-africa;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2002/1003/gibbon.pdf.

Gold:2002:PSK

[4506] John R. Gold, Elena Pak, and Doug A. DeVries. Population structure of king mackerel (Scomberomorus cavalla) around peninsular Florida, as revealed by microsatellite DNA. Fishery Bulletin, 100(3):491–509, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/population-structure-
king-mackerel-scomberomorus-cavalla-around-peninsular-florida-
revealed;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2002/1003/goldfi.pdf.

Hazell:2002:DFS

[4507] R. W. Anthony Hazell, David S. Schoeman, and Mark N. Noffke. Do fluctuations in the somatic growth rate of rock lobster (Jasus lalandii) encompass all size classes? A re-assessment of juvenile growth. Fishery Bulletin, 100(3):510–518, 2002. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/do-fluctuations-somatic-growth-
rate-rock-lobster-jasus-lalandii-encompass-all-size-classes;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2002/1003/hazell.pdf.

Jacobson:2002:SRT

[4508] Larry D. Jacobson and Steven X. Cadrin. Stock-rebuilding time isopleths and constant-F stock-rebuilding plans for overfished stocks. Fishery Bulletin, 100(3):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/stock-rebuilding-time-isopleths-and-constant-f-
stock-rebuilding-plans-overfished-stocks;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2002/1003/jacobs.pdf.

McDowell:2002:NMD

[4509] Jan R. McDowell and John E. Graves. Nuclear and mitochondrial DNA markers for specific identification of istiophorid and xiphiid billfishes. Fishery Bulletin, 100(3):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/nuclear-and-mitochondrial-dna-markers-specific-

identification-istiophorid-and-xiphiid;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2002/1003/mcdowe.pdf.

McGarvey:2002:SGK

[4510] Richard McGarvey and Anthony J. Fowler. Seasonal growth of King George whiting (Sillaginodes punctata) estimated from lengthat-age samples of the legal-size harvest. Fishery Bulletin, 100(3): ??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/seasonal-
growth-king-george-whiting-sillaginodes-punctata-estimated-
length-age-samples;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2002/1003/mcgarv.pdf.

Mueter:2002:STP

[4511] Franz J. Mueter and Brenda L. Norcross. Spatial and temporal patterns in the demersal fish community on the shelf and upper slope regions of the Gulf of Alaska. Fishery Bulletin, 100(3):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/spatial-and-temporal-
patterns-demersal-fish-community-shelf-and-upper-slope-regions-
gulf;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2002/1003/mueter.pdf.

Nelson:2002:AGM

[4512] Gary A. Nelson. Age, growth, mortality, and distribution of pinfish (Lagodon rhomboides) in Tampa Bay and adjacent Gulf of Mexico waters. Fishery Bulletin, 100(3):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/age-growth-mortality-and-
distribution-pinfish-lagodon-rhomboides-tampa-bay-and-adjacent;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2002/1003/nelson.pdf.

Overholtz:2002:RGM

[4513] William J. Overholtz and Kevin D. Friedland. Recovery of the Gulf of Maine–Georges Bank Atlantic herring (Clupea harengus) complex: perspectives based on bottom trawl survey data. Fishery Bulletin, 100(3):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/recovery-gulf-maine-georges-bank-atlantic-
herring-clupea-harengus-complex-perspectives;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/2002/1003/overho.
pdf.

Warlen:2002:RLA

[4514] Stanley M. Warlen, Kenneth W. Able, and Elisabeth Laban. Recruitment of larval Atlantic menhaden (Brevoortia tyrannus) to North Carolina and New Jersey estuaries: evidence for larval transport northward along the east coast of the United States. Fishery Bulletin, 100(3):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/recruitment-larval-
atlantic-menhaden-brevoortia-tyrannus-north-carolina-and-new-
jersey;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2002/1003/warlen.pdf.

Witzell:2002:OIL

[4515] Wayne N. Witzell, Anna L. Bass, Michael J. Bresette, David A. Singewald, and Jonathan C. Gorham. Origin of immature loggerhead sea turtles (Caretta caretta) at Hutchinson Island, Florida: evidence from mtDNA markers. Fishery Bulletin, 100(3):??, 2002. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/origin-immature-loggerhead-sea-
turtles-caretta-caretta-hutchinson-island-florida-evidence;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2002/1003/witzel.pdf.

Bilkovic:2002:SAS

[4516] Donna Marie Bilkovic, John E. Olney, and Carl H. Hershner. Spawning of American shad (Alosa sapidissima) and striped bass (Morone saxatilis) in the Mattaponi and Pamunkey Rivers, Virginia. Fishery Bulletin, 100 (3):632–640, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/spawning-
american-shad-alosa-sapidissima-and-striped-bass-morone-saxatilis-
mattaponi-and;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2002/1003/bikovi.pdf.

McBride:2002:SGO

[4517] Richard S. McBride. Spawning, growth, and overwintering size of searobins (Triglidae: Prionotus carolinus and P. evolans). Fishery Bulletin, 100(3):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
spawning-growth-and-overwintering-size-searobins-triglidae-
prionotus-carolinus-and-p;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2002/1003/mcbrid.pdf.

Anonymous:2002:SFd

[4518] Anonymous. Subscription form. Fishery Bulletin, 100(4):??, 2002. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/node/14454;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/subform%20(14)
.pdf.

Anonymous:2002:AR

[4519] Anonymous. Acknowledgment of reviewers. Fishery Bulletin, 100 (4):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/node/14452;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/reviewers%20(1).pdf.

Anonymous:2002:I

[4520] Anonymous. Index. Fishery Bulletin, 100(4):??, 2002. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/node/14453;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/index%20(11).pdf.

Bostrom:2002:HBT

[4521] Meredith A. Bostrom, Bruce B. Collette, Brian E. Luckhurst, Kimberly S. Reece, and John E. Graves. Hybridization between two serranids, the coney (Cephalopholis fulva) and the creole-fish (Paranthias furcifer), at Bermuda. Fishery Bulletin, 100(4):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/hybridization-
between-two-serranids-coney-cephalopholis-fulva-and-creole-
fish-paranthias;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2002/1004/bostro.pdf.

Carter:2002:EEH

[4522] Sarah K. Carter and Glenn R. VanBlaricom. Effects of experimental harvest on red sea urchins (Strongylocentrotus franciscanus) in northern Washington. Fishery Bulletin, 100(4):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/effects-experimental-harvest-red-sea-
urchins-strongylocentrotus-franciscanus-northern;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/2002/1004/
carter.pdf.

Conrath:2002:AGS

[4523] Christina L. Conrath, James Gelsleichter, and John A. Musick. Age and growth of the smooth dogfish (Mustelus canis) in the northwest Atlantic Ocean. Fishery Bulletin, 100(4):674–682, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/age-and-growth-smooth-dogfish-mustelus-canis-
northwest-atlantic-ocean;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2002/1004/conrat.pdf.

Gaertner:2002:BBE

[4524] Daniel Gaertner, Frederic Menard, Carol Develter, and Javier Ariz. Bycatch of billfishes by the European tuna purse-seine fishery in the Atlantic Ocean. Fishery Bulletin, 100(4):??, 2002. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/bycatch-billfishes-
european-tuna-purse-seine-fishery-atlantic-ocean;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/2002/1004/
gaertn.pdf.

Griffiths:2002:LHS

[4525] Marc H. Griffiths. Life history of South African snoek, Thyrsites atun (Pisces: Gempylidae): a pelagic predator of the Benguela Ecosystem. Fishery Bulletin, 100(4):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/life-history-south-african-snoek-
thyrsites-atun-pisces-gempylidae-pelagic-predator-benguela;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2002/1004/griffi.pdf.

Hannah:2002:LAM

[4526] Robert W. Hannah, Steven J. Parker, and Erica L. Fruh. Length and age at maturity of female petrale sole (Eopsetta jordani) determined from samples collected prior to spawning aggregation. Fishery Bulletin, 100 (4):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/length-and-
age-maturity-female-petrale-sole-eopsetta-jordani-determined-
samples-collected;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2002/1004/hannah.pdf.

Hjellvik:2002:MEM

[4527] Vidar Hjellvik, Olav Rune Godø, and Dag Tjøstheim. The measurement error in marine survey catches: the bottom trawl case. Fishery Bulletin, 100(4):720–726, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/measurement-error-marine-survey-catches-bottom-
trawl-case;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2002/1004/hjellv.pdf.

Jensen:2002:RBP

[4528] Christopher F. Jensen, Lisa J. Natanson, Harold L. Pratt, Jr., Nancy Kohler, and Steven E. Campana. The reproductive biology of the porbeagle shark (Lamna nasus) in the western North Atlantic Ocean. Fishery Bulletin, 100(4):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/reproductive-biology-porbeagle-shark-
lamna-nasus-western-north-atlantic-ocean;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2002/1004/jensen.pdf.

Nasby-Lucas:2002:IST

[4529] Nicole M. Nasby-Lucas, Bob W. Embley, Mark A. Hixon, Susan G. Merle, Brian N. Tissot, and Dawn J. Wright. Integration of submersible transect data and high-resolution multibeam sonar imagery for a habitat-based groundfish assessment of Heceta Bank, Oregon. Fishery Bulletin, 100(4):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/integration-submersible-transect-data-and-
high-resolution-multibeam-sonar-imagery-habitat;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/2002/1004/
nasby-.pdf.

Olsen:2002:EST

[4530] Jeffrey B. Olsen, Susan E. Merkouris, and James E. Seeb. An examination of spatial and temporal genetic variation in walleye pollock (Theragra chalcogramma) using allozyme, mitochondrial DNA, and microsatellite data. Fishery Bulletin, 100(4):

752–764, 2002.  CODEN FSYBAY.  ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
examination-spatial-and-temporal-genetic-variation-walleye-
pollock-theragra-chalcogramma;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2002/1004/olsenf.pdf.

Schaefer:2002:MBH

[4531] Kurt M. Schaefer and Daniel W. Fuller. Movements, behavior, and habitat selection of bigeye tuna (Thunnus obesus) in the eastern equatorial Pacific, ascertained through archival tags. Fishery Bulletin, 100(4):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/movements-behavior-
and-habitat-selection-bigeye-tuna-thunnus-obesus-eastern-equatorial;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2002/1004/schaef.pdf.

Schirripa:2002:EBC

[4532] Michael J. Schirripa. An evaluation of back-calculation methodology using simulated otolith data. Fishery Bulletin, 100(4):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/evaluation-
back-calculation-methodology-using-simulated-otolith-data;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/2002/
1004/schirr.pdf.

Stobutzki:2002:SEC

[4533] Ilona C. Stobutzki, Margaret J. Miller, Don S. Heales, and David T. Brewer. Sustainability of elasmobranchs caught as bycatch in a tropical prawn (shrimp) trawl fishery. Fishery Bulletin, 100 (4):800–821, 2002. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
sustainability-elasmobranchs-caught-bycatch-tropical-prawn-
shrimp-trawl-fishery;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2002/1004/stobut.pdf.

Sun:2002:AGS

[4534] Chi-Lu Sun, Sheng-Ping Wang, and Su-Zan Yeh. Age and growth of the swordfish (Xiphias gladius L.) in the waters around Taiwan determined from anal-fin rays. Fishery Bulletin, 100(4):822–835, 2002. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/age-and-growth-swordfish-
xiphias-gladius-l-waters-around-taiwan-determined-anal-fin-
rays;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2002/1004/sunfis.pdf.

Williams:2002:DCO

[4535] Erik H. Williams and Stephen Ralston. Distribution and cooccurrence of rockfishes (family: Sebastidae) over trawlable shelf and slope habitats of California and southern Oregon. Fishery Bulletin, 100(4):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/distribution-and-co-occurrence-rockfishes-
family-sebastidae-over-trawlable-shelf-and-slope;   https:/

/spo.nmfs.noaa.gov/sites/default/files/pdf-content/2002/1004/
willia.pdf.

Latour:2002:PRT

[4536] Robert J. Latour, John M. Hoenig, and Kenneth H. Pollock. Properties of the residuals from two tag-recovery models. Fishery Bulletin, 100(4):856– 860, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/properties-
residuals-two-tag-recovery-models;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2002/1004/latour.pdf.

McFarlane:2002:PSU

[4537] Gordon A. McFarlane, Jacquelynne R. King, and Mark W. Saunders. Preliminary study on the use of neural arches in the age determination of bluntnose sixgill sharks (Hexanchus griseus). Fishery Bulletin, 100(4): ??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/preliminary-
study-use-neural-arches-age-determination-bluntnose-sixgill-
sharks-hexanchus;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2002/1004/mcfarl.pdf.

Mollet:2002:RIL

[4538] Henry F. Mollet, Antonio D. Testi, Leonard J. V. Compagno, and Malcolm P. Francis. Re-identification of a lamnid shark embryo. Fishery Bulletin, 100(4):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/re-identification-lamnid-shark-embryo;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2002/1004/
mollet.pdf.

Work:2002:NFS

[4539] Thierry M. Work and George H. Balazs. Necropsy findings in sea turtles taken as bycatch in the North Pacific longline fishery. Fishery Bulletin, 100(4):??, 2002. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/necropsy-
findings-sea-turtles-taken-bycatch-north-pacific-longline-fishery;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2002/1004/workfi.pdf.

Blaylock:2003:UPD

[4540] Reginald B. Blaylock, Leo Margolis (deceased), and John C. Holmes. The use of parasites in discriminating stocks of Pacific halibut (Hippoglossus stenolepis) in the northeast Pacific. Fishery Bulletin, 101(1):

1–9, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/use-parasites-
discriminating-stocks-pacific-halibut-hippoglossus-stenolepis-
northeast;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2003/1011/blaylo.pdf.

Comyns:2003:SSS

[4541] Bruce H. Comyns, Richard F. Shaw, and Joanne Lyczkowski-Shultz. Small-scale spatial and temporal variability in growth and mortality of fish larvae in the subtropical northcentral Gulf of Mexico: implications for assessing recruitment success. Fishery Bulletin, 101(1):10–21, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/small-scale-
spatial-and-temporal-variability-growth-and-mortality-fish-
larvae-subtropical;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2003/1011/comyns.pdf.

DeMartini:2003:TCP

[4542] Edward E. DeMartini, Gerard T. DiNardo, and Happy A. Williams. Temporal changes in population density, fecundity, and egg size of the Hawaiian spiny lobster (Panulirus marginatus) at Necker Bank, Northwestern Hawaiian Islands. Fishery Bulletin, 101(1):22– 31, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/temporal-
changes-population-density-fecundity-and-egg-size-hawaiian-
spiny-lobster-panulirus;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2003/1011/demart.pdf.

Friedlander:2003:IHR

[4543] Alan M. Friedlander and David A. Ziemann. Impact of hatchery releases on the recreational fishery for Pacific threadfin (Polydactylus sexfilis) in Hawaii. Fishery Bulletin, 101(1):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/impact-hatchery-releases-
recreational-fishery-pacific-threadfin-polydactylus-sexfilis;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2003/1011/friedl.pdf.

Hart:2003:YBP

[4544] Deborah R. Hart. Yield- and biomass-per-recruit analysis for rotational fisheries, with an application to the Atlantic sea scallop (Placopecten magellanicus). Fishery Bulletin, 101(1):??, 2003. CODEN FSYBAY.

ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/yield-and-biomass-recruit-analysis-rotational-
fisheries-application-atlantic-sea-scallop;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2003/1011/hartfi.pdf.

Hearn:2003:ELT

[4545] William S. Hearn and Thomas Polacheck. Estimating long-term growthrate changes of southern bluefin tuna (Thunnus maccoyii) from two periods of tag-return data. Fishery Bulletin, 101(1):??, 2003. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/estimating-long-term-growth-rate-
changes-southern-bluefin-tuna-thunnus-maccoyii-two-periods;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2003/1011/hearnf.pdf.

Loefer:2003:LHA

[4546] Joshua K. Loefer and George R. Sedberry. Life history of the Atlantic sharpnose shark (Rhizoprionodon terraenovae) (Richardson, 1836) off the southeastern United States. Fishery Bulletin, 101(1):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/life-history-atlantic-sharpnose-
shark-rhizoprionodon-terraenovae-richardson-1836;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/2003/1011/
loefer.pdf.

Maunder:2003:GFI

[4547] Mark N. Maunder and George M. Watters. A general framework for integrating environmental time series into stock assessment models: model description, simulation testing, and example. Fishery Bulletin, 101 (1):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/general-
framework-integrating-environmental-time-series-stock-assessment-
models-model;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2003/1011/maunde.pdf.

Miller:2003:SDA

[4548] Michael J. Miller, David M. Nemerson, and Kenneth W. Able. Seasonal distribution, abundance, and growth of young-of-the-year Atlantic croaker (Micropogonias undulatus) in Delaware Bay and adjacent marshes. Fishery Bulletin, 101(1):100–115, 2003. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/seasonal-distribution-abundance-
and-growth-young-year-atlantic-croaker-micropogonias;   https:

//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2003/1011/
miller.pdf.

Newman:2003:AVG

[4549] Stephen J. Newman and Iain J. Dunk. Age validation, growth, mortality, and additional population parameters of the goldband snapper (Pristipomoides multidens) off the Kimberley coast of northwestern Australia. Fishery Bulletin, 101(1):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/age-validation-growth-mortality-and-additional-
population-parameters-goldband-snapper;   https://spo.nmfs.
noaa.gov/sites/default/files/10newman.pdf.

Ralston:2003:AER

[4550] Stephen Ralston, James R. Bence, Maxwell B. Eldridge, and William H. Lenarz. An approach to estimating rockfish biomass based on larval production, with application to Sebastes jordani. Fishery Bulletin, 101(1): 129–146, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/approach-
estimating-rockfish-biomass-based-larval-production-application-
sebastes-jordani;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2003/1011/ralsto.pdf.

Winship:2003:PCS

[4551] Arliss J. Winship and Andrew W. Trites. Prey consumption of Steller sea lions (Eumetopias jubatus) off Alaska: How much prey do they require? Fishery Bulletin, 101(1):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/prey-consumption-steller-
sea-lions-eumetopias-jubatus-alaska-how-much-prey-do-they-require;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2003/1011/winshi.pdf.

Beerkircher:2003:MCD

[4552] Lawrence Beerkircher, Mahmood Shivji, and Enric Cortés. A Monte Carlo demographic analysis of the silky shark (Carcharhinus falciformis): implications of gear selectivity. Fishery Bulletin, 101(1):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/monte-carlo-
demographic-analysis-silky-shark-carcharhinus-falciformis-implications-
gear;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2003/1011/beerki.pdf.

Gunderson:2003:IEN

[4553] Donald R. Gunderson, Mark Zimmermann, Daniel G. Nichol, and Katherine Pearson. Indirect estimates of natural mortality rate for arrowtooth flounder (Atheresthes stomias) and darkblotched rockfish (Sebastes crameri). Fishery Bulletin, 101(1):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/indirect-estimates-natural-mortality-rate-
arrowtooth-flounder-atheresthes-stomias-and;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/2003/1011/gunder.
pdf.

Marcogliese:2003:UPS

[4554] David J. Marcogliese, Elaine Albert, Pierre Gagnon, and Jean-Marie Sevigny. Use of parasites in stock identification of the deepwater redfish (Sebastes mentella) in the Northwest Atlantic. Fishery Bulletin, 101(1):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/use-parasites-stock-
identification-deepwater-redfish-sebastes-mentella-northwest-
atlantic;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2003/1011/marcog.pdf.

Polovina:2003:DDD

[4555] Jeffrey J. Polovina, Evan Howell, Denise M. Parker, and George H. Balazs. Dive-depth distribution of loggerhead (Carretta carretta) and olive ridley (Lepidochelys olivacea) sea turtles in the central North Pacific: Might deep longline sets catch fewer turtles? Fishery Bulletin, 101 (1):189–193, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/dive-
depth-distribution-loggerhead-carretta-carretta-and-olive-ridley-
lepidochelys-olivacea;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2003/1011/polovi.pdf.

Smith:2003:AVL

[4556] Susan E. Smith, Robert A. Mitchell, and Dan Fuller. Agevalidation of a leopard shark (Triakis semifasciata) recaptured after 20 years. Fishery Bulletin, 101(1):194, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/age-validation-leopard-shark-triakis-
semifasciata-recaptured-after-20-years;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2003/1011/smithf.pdf.

Able:2003:UOE

[4557] Kenneth W. Able, Peter Rowe, Mark Burlas, and Don Byrne. Use of ocean and estuarine habitats by young-of-year bluefish (Pomatomus saltatrix) in the New York Bight. Fishery Bulletin, 101(2):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/use-ocean-and-
estuarine-habitats-young-year-bluefish-pomatomus-saltatrix-
new-york-bight;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2003/1012/ablefi.pdf.

Adam:2003:DBT

[4558] M. Shiham Adam, John Sibert, David Itano, and Kim Holland. Dynamics of bigeye (Thunnus obesus) and yellowfin (T. albacares) tuna in Hawaii’s pelagic fisheries: analysis of tagging data with a bulk transfer model incorporating size-specific attrition. Fishery Bulletin, 101(2):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/dynamics-bigeye-thunnus-obesus-and-yellowfin-t-
albacares-tuna-hawaiis-pelagic-fisheries;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2003/1012/adamfi.pdf.

Beacham:2003:GBP

[4559] Terry D. Beacham, K. Janine Supernault, Michael Wetklo, Bruce Deagle, Karen Labaree, James R. Irvine, John R. Candy, Kristina M. Miller, R. John Nelson, and Ruth E. Withler. The geographic basis for population structure in Fraser River Chinook salmon (Oncorhynchus tshawytscha). Fishery Bulletin, 101(2):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/geographic-basis-population-structure-
fraser-river-chinook-salmon-oncorhynchus-tshawytscha;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2003/1012/
beacha.pdf.

Beacham:2003:EAM

[4560] Terry D. Beacham, John R. Candy, K. Janine Supernault, Michael Wetklo, Bruce Deagle, Karen Labaree, James R. Irvine, Kristina M. Miller, R. John Nelson, and Ruth E. Withler. Evaluation and application of microsatellites for population identification of Fraser River Chinook salmon (Oncorhynchus tshawytscha). Fishery Bulletin, 101(2):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/evaluation-
and-application-microsatellites-population-identification-fraser-

river-chinook;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2003/1012/beacha2.pdf.

Butler:2003:BPD

[4561] John L. Butler, Larry D. Jacobson, J. Thomas Barnes, and H. Geoffrey Moser. Biology and population dynamics of cowcod (Sebastes levis) in the southern California Bight. Fishery Bulletin, 101(2):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/biology-and-population-dynamics-
cowcod-sebastes-levis-southern-california-bight;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/2003/1012/
butler.pdf.

Carlson:2003:LHP

[4562] John K. Carlson, Enric Cortés, and Dana M. Bethea. Life history and population dynamics of the finetooth shark (Carcharhinus isodon) in the northeastern Gulf of Mexico. Fishery Bulletin, 101(2):281–292, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/life-history-and-population-dynamics-finetooth-
shark-carcharhinus-isodon-northeastern-gulf;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/2003/1012/carlso.
pdf.

Francis:2003:QAV

[4563] Chris R. I. C. Francis, Rosemary J. Hurst, and James A. Renwick. Quantifying annual variation in catchability for commercial and research fishing. Fishery Bulletin, 101(2):293–304, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/quantifying-annual-variation-catchability-
commercial-and-research-fishing;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2003/1012/franci.pdf.

Grandcourt:2003:EIL

[4564] Edwin M. Grandcourt. The effect of intensive line fishing on the virgin biomass of a tropical deepwater snapper, the crimson jobfish (Pristipomoides filamentosus). Fishery Bulletin, 101(2):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/effect-
intensive-line-fishing-virgin-biomass-tropical-deepwater-snapper-
crimson-jobfish;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2003/1012/grandc.pdf.

Ju:2003:DAB

[4565] Se-Jong Ju, David H. Secor, and H. Roger Harvey. Demographic assessment of the blue crab (Callinectes sapidus) in Chesapeake Bay using extractable lipofuscins as age markers. Fishery Bulletin, 101(2):312–320, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/demographic-
assessment-blue-crab-callinectes-sapidus-chesapeake-bay-using-
extractable;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2003/1012/jufish.pdf.

Lindley:2003:MES

[4566] Steven T. Lindley and Michael S. Mohr. Modeling the effect of striped bass (Morone saxatilis) on the population viability of Sacramento River winter-run Chinook salmon (Oncorhynchus tshawytscha). Fishery Bulletin, 101(2):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
modeling-effect-striped-bass-morone-saxatilis-population-viability-
sacramento-river-winter;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2003/1012/lindle.pdf.

Macchi:2003:SEP

[4567] Gustavo J. Macchi, Eduardo M. Acha, and María I. Militelli. Seasonal egg production of whitemouth croaker (Micropogonias furnieri) in the Rio de la Plata estuary, Argentina–Uruguay. Fishery Bulletin, 101(2): ??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/seasonal-egg-
production-whitemouth-croaker-micropogonias-furnieri-rio-de-
la-plata-estuary;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2003/1012/macchi.pdf.

McDonough:2003:GRA

[4568] Christopher J. McDonough and Charles A. Wenner. Growth, recruitment, and abundance of juvenile striped mullet (Mugil cephalus) in South Carolina estuaries. Fishery Bulletin, 101(2):??, 2003. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/growth-recruitment-
and-abundance-juvenile-striped-mullet-mugil-cephalus-south-
carolina;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2003/1012/mcdono.pdf.

McFarlane:2003:MPS

[4569] Gordon A. McFarlane and Jacquelynne R. King. Migration patterns of spiny dogfish (Squalus acanthias) in the North Pacific Ocean. Fishery Bulletin, 101(2):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/migration-patterns-spiny-dogfish-squalus-acanthias-
north-pacific-ocean;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2003/1012/mcfarl.pdf.

Noell:2003:LDS

[4570] Craig J. Noell. Larval development of the southern sea garfish (Hyporhamphus melanochir) and the river garfish (H. regularis) (Beloniformes: Hemiramphidae) from South Australian waters. Fishery Bulletin, 101(2):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
larval-development-southern-sea-garfish-hyporhamphus-melanochir-
and-river-garfish-h;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2003/1012/noellf.pdf.

WeberRosas:2003:AGE

[4571] Fernando César Weber Rosas, André Silva Barreto, and Emygdio Leite de Araujo Monteiro-Filho. Age and growth of the estuarine dolphin (Sotalia guianensis) (Cetacea, Delphinidae) on the Parana coast, southern Brazil. Fishery Bulletin, 101(2):377–383, 2003. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/age-and-growth-estuarine-dolphin-
sotalia-guianensis-cetacea-delphinidae-parana-coast;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2003/1012/
rosasf.pdf.

Ross:2003:RVD

[4572] Steve W. Ross. The relative value of different estuarine nursery areas in North Carolina for transient juvenile marine fishes. Fishery Bulletin, 101(2):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/relative-
value-different-estuarine-nursery-areas-north-carolina-transient-
juvenile-marine;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2003/1012/rossfi.pdf.

Sulikowski:2003:AGE

[4573] James A. Sulikowski, Michael D. Morin, Seung H. Suk, and W. Huntting Howell. Age and growth estimates of the winter skate (Leucoraja ocellata) in the western Gulf of Maine. Fishery Bulletin, 101(2):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/age-and-growth-estimates-winter-
skate-leucoraja-ocellata-western-gulf-maine;   https://spo.

nmfs.noaa.gov/sites/default/files/pdf-content/2003/1012/suliko.
pdf.

Walter:2003:DCL

[4574] John F. Walter III and Herbert M. Austin. Diet composition of large striped bass (Morone saxatilis) in Chesapeake Bay. Fishery Bulletin, 101(2):414–423, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/diet-
composition-large-striped-bass-morone-saxatilis-chesapeake-
bay;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2003/1012/walter.pdf.

White:2003:RSF

[4575] Geoffrey G. White, Thomas A. Munroe, and Herbert M. Austin. Reproductive seasonality, fecundity, and spawning frequency of tautog (Tautoga onitis) in the lower Chesapeake Bay and coastal waters of Virginia. Fishery Bulletin, 101(2):424–442, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/reproductive-seasonality-
fecundity-and-spawning-frequency-tautog-tautoga-onitis-lower;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2003/1012/whitef.pdf.

Cordes:2003:ICH

[4576] Jan F. Cordes and John E. Graves. Investigation of congeneric hybridization in and stock structure of weakfish (Cynoscion regalis) inferred from analyses of nuclear and mitochondrial DNA loci. Fishery Bulletin, 101(2):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
investigation-congeneric-hybridization-and-stock-structure-
weakfish-cynoscion-regalis;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2003/1012/cordes.pdf.

Salthaug:2003:DAL

[4577] Are Salthaug. Dynamic age–length keys. Fishery Bulletin, 101(2): ??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/dynamic-
age-length-keys;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2003/1012/saltha.pdf.

Walls:2003:EBE

[4578] Elizabeth A. Walls and Jim Berkson. Effects of blood extraction on horseshoe crabs (Limulus polyphemus). Fishery Bulletin, 101(2):??, 2003.

CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/effects-blood-extraction-
horseshoe-crabs-limulus-polyphemus;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2003/1012/wallsf.pdf.

Brule:2003:RPB

[4579] Thierry Brulé, Ximena Renán, Teresa Colás-Marrufo, Yazmin Hauyon, Armin N. Tuz-Sulub, and Christian Déniel. Reproduction in the protogynous black grouper (Mycteroperca bonaci (Poey)) from the southern Gulf of Mexico. Fishery Bulletin, 101(3):??, 2003. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/reproduction-protogynous-black-
grouper-mycteroperca-bonaci-poey-southern-gulf-mexico;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2003/1013/
brulef.pdf.

Comeau:2003:ETT

[4580] Michel Comeau and Manon Mallet. The effect of timing of tagging on streamer-tag recapture rates for American lobster (Homarus americanus). Fishery Bulletin, 101(3):??, 2003. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/effect-timing-tagging-streamer-
tag-recapture-rates-american-lobster-homarus-americanus;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2003/1013/
comeau.pdf.

Diamond:2003:EBS

[4581] Sandra L. Diamond. Estimation of bycatch in shrimp trawl fisheries: a comparison of estimation methods using field data and simulated data. Fishery Bulletin, 101(3):??, 2003. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/estimation-bycatch-shrimp-trawl-
fisheries-comparison-estimation-methods-using-field-data;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/2003/
1013/diamon.pdf.

Hanselman:2003:AAC

[4582] Dana H. Hanselman, Terrance J. Quinn II, Chris Lunsford, Jonathan Heifetz, and David Clausen. Applications in adaptive cluster sampling of Gulf of Alaska rockfish. Fishery Bulletin, 101(3):501–513, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/applications-
adaptive-cluster-sampling-gulf-alaska-rockfish;   https://spo.

nmfs.noaa.gov/sites/default/files/pdf-content/2003/1013/hansel.
pdf.

Itoh:2003:MPY

[4583] Tomoyuki Itoh, Sachiko Tsuji, and Akira Nitta. Migration patterns of young Pacific bluefin tuna (Thunnus orientalis) determined with archival tags. Fishery Bulletin, 101(3):514–534, 2003. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/migration-patterns-young-pacific-
bluefin-tuna-thunnus-orientalis-determined-archival-tags;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/2003/
1013/itohfi.pdf.

Itoh:2003:SDA

[4584] Tomoyuki Itoh, Sachiko Tsuji, and Akira Nitta. Swimming depth, ambient water temperature preference, and feeding frequency of young Pacific bluefin tuna (Thunnus orientalis) determined with archival tags. Fishery Bulletin, 101(3):535–544, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/swimming-depth-ambient-water-temperature-preference-
and-feeding-frequency-young-pacific;   https://spo.nmfs.noaa.
gov/sites/default/files/06itohf2.pdf.

Jagielo:2003:DGD

[4585] Thomas Jagielo, Annette Hoffmann, Jack Tagart, and Mark Zimmermann. Demersal groundfish densities in trawlable and untrawlable habitats off Washington: implications for the estimation of habitat bias in trawl surveys. Fishery Bulletin, 101(3):545–565, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/demersal-groundfish-
densities-trawlable-and-untrawlable-habitats-washington-implications;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2003/1013/jagiel.pdf.

Loughlin:2003:DBI

[4586] Thomas R. Loughlin, Jeremy T. Sterling, Richard L. Merrick, John L. Sease, and Anne E. York. Diving behavior of immature Steller sea lions (Eumetopias jubatus). Fishery Bulletin, 101(3):??, 2003. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/diving-behavior-immature-steller-
sea-lions-eumetopias-jubatus;   https://spo.nmfs.noaa.gov/sites/
default/files/08loughl.pdf.

McBride:2003:SCH

[4587] Richard S. McBride, Justin R. Styer, and Rob Hudson. Spawning cycles and habitats for ballyhoo (Hemiramphus brasiliensis) and balao (H. balao) in south Florida. Fishery Bulletin, 101(3):583–589, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/spawning-cycles-and-
habitats-ballyhoo-hemiramphus-brasiliensis-and-balao-h-balao-
south;   https://spo.nmfs.noaa.gov/sites/default/files/09mcbrid.
pdf.

Morato:2003:DTR

[4588] Telmo Morato, Encarnacion Solà, Maria P. Grós, and Gui Menezes. Diets of thornback ray (Raja clavata) and tope shark (Galeorhinus galeus) in the bottom longline fishery of the Azores, northeastern Atlantic. Fishery Bulletin, 101(3):590–602, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/diets-thornback-ray-raja-clavata-and-tope-
shark-galeorhinus-galeus-bottom-longline-fishery;   https:/
/spo.nmfs.noaa.gov/sites/default/files/10morato.pdf.

Mullin:2003:ACS

[4589] Keith D. Mullin and Gregory L. Fulling. Abundance of cetaceans in the southern U.S. North Atlantic Ocean during summer 1998. Fishery Bulletin, 101(3):603–613, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/abundance-cetaceans-southern-us-north-atlantic-
ocean-during-summer-1998;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2003/1013/mullin.pdf.

Rogers-Bennett:2003:MRS

[4590] Laura Rogers-Bennett, Donald W. Rogers, William A. Bennett, and Thomas A. Ebert. Modeling red sea urchin (Strongylocentrotus franciscanus) growth using six growth functions. Fishery Bulletin, 101(3): ??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/modeling-red-
sea-urchin-strongylocentrotus-franciscanus-growth-using-six-
growth-functions;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2003/1013/rogers.pdf.

Skomal:2003:AGB

[4591] Gregory B. Skomal and Lisa J. Natanson. Age and growth of the blue shark (Prionace glauca) in the North Atlantic Ocean. Fishery Bulletin, 101(3):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/age-
and-growth-blue-shark-prionace-glauca-north-atlantic-ocean;
 https://spo.nmfs.noaa.gov/sites/default/files/13skomal.pdf.

Teel:2003:GAJ

[4592] David J. Teel, Donald M. Van Doornik, David R. Kuligowski, and W. Stewart Grant. Genetic analysis of juvenile Coho salmon (Oncorhynchus kisutch) off Oregon and Washington reveals few Columbia River wild fish. Fishery Bulletin, 101(3):??, 2003. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/genetic-analysis-juvenile-coho-
salmon-oncorhynchus-kisutch-oregon-and-washington-reveals;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/2003/
1013/teelfi.pdf.

Terceiro:2003:SPR

[4593] Mark Terceiro. The statistical properties of recreational catch rate data for some fish stocks off the northeast U.S. coast. Fishery Bulletin, 101(3):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/statistical-properties-recreational-catch-rate-
data-some-fish-stocks-northeast-us-coast;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2003/1013/tercei.pdf.

Williams:2003:SSV

[4594] Ashley J. Williams, Campbell R. Davies, Bruce D. Mapstone, and Garry R. Russ. Scales of spatial variation in demography of a large coral-reef fish — an exception to the typical model? Fishery Bulletin, 101(3):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/scales-spatial-variation-demography-large-
coral-reef-fish%E2%80%94-exception-typical-model;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/2003/1013/
willia.pdf.

Klimley:2003:OYT

[4595] A. Peter Klimley, Salvador J. Jorgensen, Arturo Muhlia-Melo, and Sallie C. Beavers. The occurrence of yellowfin tuna (Thunnus albacares) at espiritu santo seamount in the Gulf of California. Fishery Bulletin, 101(3):684–693, 2003. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
occurrence-yellowfin-tuna-thunnus-albacares-espiritu-santo-

seamount-gulf-california;   https://spo.nmfs.noaa.gov/sites/
default/files/17klimle.pdf.

Landaeta:2003:LDD

[4596] Mauricio F. Landaeta, Francisco J. Neira, and Leonardo R. Castro. Larvae of Dactylopsaron dimorphicum (Perciformes: Percophidae) from oceanic islands in the southeast Pacific. Fishery Bulletin, 101(3):693–697, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/larvae-dactylopsaron-
dimorphicum-perciformes-percophidae-oceanic-islands-southeast-
pacific;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2003/1013/landae.pdf.

Pooler:2003:ASM

[4597] Penelope S. Pooler, David R. Smith, Robert E. Loveland, Mark L. Botton, and Stewart F. Michels. Assessment of sampling methods to estimate horseshoe crab (Limulus polyphemus L.) egg density in Delaware Bay. Fishery Bulletin, 101(3):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/assessment-sampling-methods-estimate-
horseshoe-crab-limulus-polyphemus-l-egg-density;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/2003/1013/
pooler.pdf.

Powell:2003:LAD

[4598] Allyn B. Powell. Larval abundance, distribution, and spawning habits of spotted seatrout (Cynoscion nebulosus) in Florida Bay, Everglades National Park, Florida. Fishery Bulletin, 101(3): ??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/larval-
abundance-distribution-and-spawning-habits-spotted-seatrout-
cynoscion-nebulosus;   https://spo.nmfs.noaa.gov/sites/default/
files/20powell.pdf.

Zimmerman:2003:EAC

[4599] Christian E. Zimmerman and Roger L. Nielsen. Effect of analytical conditions in wavelength dispersive electron microprobe analysis on the measurement of strontium-to-calcium (Sr/Ca) ratios in otoliths of anadromous salmonids. Fishery Bulletin, 101(3): ??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/effect-
analytical-conditions-wavelength-dispersive-electron-microprobe-

analysis-measurement;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2003/1013/zimmer.pdf.

Baum:2003:BLS

[4600] Julia K. Baum, Jessica J. Meeuwig, and Amanda C. J. Vincent. Bycatch of lined seahorses (Hippocampus erectus) in a Gulf of Mexico shrimp trawl fishery. Fishery Bulletin, 101(4):721–731, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/bycatch-lined-
seahorses-hippocampus-erectus-gulf-mexico-shrimp-trawl-fishery;
 https://spo.nmfs.noaa.gov/sites/default/files/01baumfi.pdf.

Bjorndal:2003:ESP

[4601] Karen A. Bjorndal, Alan B. Bolten, and Helen R. Martins. Estimates of survival probabilities for oceanic-stage loggerhead sea turtles (Caretta caretta) in the North Atlantic. Fishery Bulletin, 101(4):732–736, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/estimates-survival-
probabilities-oceanic-stage-loggerhead-sea-turtles-caretta-
caretta-north;   https://spo.nmfs.noaa.gov/sites/default/files/
02bjornd.pdf.

Chen:2003:DGT

[4602] Yong Chen, Margaret Hunter, Robert Vadas, and Brian Beal. Developing a growth-transition matrix for the stock assessment of the green sea urchin (Strongylocentrotus droebachiensis) off Maine. Fishery Bulletin, 101(4):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
developing-growth-transition-matrix-stock-assessment-green-
sea-urchin-strongylocentrotus;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2003/1014/chenfi.pdf.

deLestang:2003:RBB

[4603] Simon de Lestang, Norman G. Hall, and Ian C. Potter. Reproductive biology of the blue swimmer crab (Portunus pelagicus, Decapoda: Portunidae) in five bodies of water on the west coast of Australia. Fishery Bulletin, 101(4):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
reproductive-biology-blue-swimmer-crab-portunus-pelagicus-decapoda-
portunidae-five-bodies;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2003/1014/delest.pdf.

Dew:2003:MAL

[4604] Jodi R. Dew, Jim Berkson, Eric M. Hallerman, and Standish K. Allen, Jr. A model for assessing the likelihood of self-sustaining populations resulting from commercial production of triploid suminoe oysters (Crassostrea ariakensis) in Chesapeake Bay. Fishery Bulletin, 101(4):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/model-
assessing-likelihood-self-sustaining-populations-resulting-
commercial-production;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2003/1014/dewfis.pdf.

Diaz-Jaimes:2003:ARV

[4605] Pindaro Diaz-Jaimes and Manuel Uribe-Alcocer. Allozyme and RAPD variation in the eastern Pacific yellowfin tuna (Thunnus albacares). Fishery Bulletin, 101(4):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
allozyme-and-rapd-variation-eastern-pacific-yellowfin-tuna-
thunnus-albacares;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2003/1014/diaz.pdf.

Govoni:2003:ELH

[4606] John Jeffrey Govoni, Elisabeth H. Laban, and Jonathan A. Hare. The early life history of swordfish (Xiphias gladius) in the western North Atlantic. Fishery Bulletin, 101(4):778–789, 2003. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/early-life-history-
swordfish-xiphias-gladius-western-north-atlantic;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/2003/1014/
govoni.pdf.

Heales:2003:DSS

[4607] Donald S. Heales, David T. Brewer, You-Gan Wang, and Peter N. Jones. Does the size of subsamples taken from multispecies trawl catches affect estimates of catch composition and abundance? Fishery Bulletin, 101(4):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/does-size-
subsamples-taken-multispecies-trawl-catches-affect-estimates-
catch-composition;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2003/1014/heales.pdf.

Laidig:2003:AGB

[4608] Thomas E. Laidig, Donald E. Pearson, and Lorraine L. Sinclair. Age and growth of blue rockfish (Sebastes mystinus) from central and northern California. Fishery Bulletin, 101(4):800–808, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/age-and-growth-blue-
rockfish-sebastes-mystinus-central-and-northern-california;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2003/1014/laidig.pdf.

MarinE:2003:RRW

[4609] Baumar J. Marin E., Antonio Quintero, Dany Bussière, and Julian J. Dodson. Reproduction and recruitment of white mullet (Mugil curema) to a tropical lagoon (Margarita Island, Venezuela) as revealed by otolith microstructure. Fishery Bulletin, 101(4):809–821, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/reproduction-
and-recruitment-white-mullet-mugil-curema-tropical-lagoon-margarita-
island;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2003/1014/marinf.pdf.

McDonough:2003:FSS

[4610] Christopher J. McDonough, William A. Roumillat, and Charles A. Wenner. Fecundity and spawning season of striped mullet (Mugil cephalus L.) in South Carolina estuaries. Fishery Bulletin, 101(4):822–834, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/fecundity-and-
spawning-season-striped-mullet-mugil-cephalus-l-south-carolina-
estuaries;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2003/1014/mcdono.pdf.

Nelson:2003:MFA

[4611] Peter A. Nelson. Marine fish assemblages associated with fish aggregating devices (FADs): effects of fish removal, FAD size, fouling communities, and prior recruits. Fishery Bulletin, 101(4):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/marine-fish-
assemblages-associated-fish-aggregating-devices-fads-effects-
fish-removal-fad;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2003/1014/nelson.pdf.

Pajuelo:2003:AGB

[4612] José G. Pajuelo, José M. Lorenzo, and Muriel Gregoire. Age and growth of the bastard grunt (Pomadasys incisus: Haemulidae) inhabiting the Canarian archipelago, Northwest Africa. Fishery Bulletin, 101(4):851–859, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/age-and-growth-bastard-grunt-pomadasys-incisus-
haemulidae-inhabiting-canarian-archipelago;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2003/1014/pajuel.pdf.

Punt:2003:EEM

[4613] Andre E. Punt. Evaluating the efficacy of managing west coast groundfish resources through simulations. Fishery Bulletin, 101(4):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/evaluating-efficacy-
managing-west-coast-groundfish-resources-through-simulations;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2003/1014/puntfi.pdf.

Purves:2003:DDD

[4614] Martin G. Purves, David J. Agnew, Guillermo Moreno, Tim Daw, Cynthia Yau, and Graham Pilling. Distribution, demography, and discard mortality of crabs caught as bycatch in an experimental pot fishery for toothfish (Dissostichus eleginoides) in the South Atlantic. Fishery Bulletin, 101(4):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
distribution-demography-and-discard-mortality-crabs-caught-
bycatch-experimental-pot-fishery;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2003/1014/purves.pdf.

Stabenau:2003:PEM

[4615] Erich K. Stabenau and Kimberly R. N. Vietti. The physiological effects of multiple forced submergences in loggerhead sea turtles (Caretta caretta). Fishery Bulletin, 101(4):??, 2003. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/physiological-effects-multiple-
forced-submergences-loggerhead-sea-turtles-caretta-caretta;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2003/1014/staben.pdf.

Stephenson:2003:QDT

[4616] Peter C. Stephenson and Norm G. Hall. Quantitative determination of the timing of otolith ring formation from marginal increments in four marine teleost species from northwestern Australia. Fishery Bulletin, 101(4):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/quantitative-determination-timing-otolith-ring-

formation-marginal-increments-four-marine;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2003/1014/stephe.pdf.

Chan:2003:ADB

[4617] Ricky W. K. Chan, Patricia I. Dixon, Julian G. Pepperell, and Dennis D. Reid. Application of DNA-based techniques for the identification of whaler sharks (Carcharhinus spp.) caught in protective beach meshing and by recreational fisheries off the coast of New South Wales. Fishery Bulletin, 101(4):??, 2003. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/application-dna-based-techniques-
identification-whaler-sharks-carcharhinus-spp-caught;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2003/1014/
chanfi.pdf.

Ebert:2003:RSU

[4618] Thomas A. Ebert and John R. Southon. Red sea urchins (Strongylocentrotus franciscanus) can live over 100 years: confirmation with A-bomb^{14} carbon. Fishery Bulletin, 101(4):915–922, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/red-sea-urchins-strongylocentrotus-franciscanus-
can-live-over-100-years-confirmation-bomb;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2003/1014/ebertf.pdf.

Fulling:2003:ADC

[4619] Gregory L. Fulling, Keith D. Mullin, and Carrie W. Hubard. Abundance and distribution of cetaceans in outer continental shelf waters of the U.S. Gulf of Mexico. Fishery Bulletin, 101(4):923–932, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/abundance-and-
distribution-cetaceans-outer-continental-shelf-waters-us-gulf-
mexico;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2003/1014/fullin.pdf.

Hata:2003:AHC

[4620] David Hata and Jim Berkson. Abundance of horseshoe crabs (Limulus polyphemus) in the Delaware Bay area. Fishery Bulletin, 101(4):??, 2003. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/abundance-horseshoe-
crabs-limulus-polyphemus-delaware-bay-area;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2003/1014/hatafi.pdf.

Kerstetter:2003:UPS

[4621] David W. Kerstetter, Brian E. Luckhurst, Eric D. Prince, and John E. Graves. Use of pop-up satellite archival tags to demonstrate survival of blue marlin (Makaira nigricans) released from pelagic longline gear. Fishery Bulletin, 101(4):??, 2003. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/use-pop-satellite-archival-tags-
demonstrate-survival-blue-marlin-makaira-nigricans-released;
 https://spo.nmfs.noaa.gov/sites/default/files/22kerste.pdf.

Alonzo:2004:ESS

[4622] Suzanne H. Alonzo and Marc Mangel. The effects of size-selective fisheries on the stock dynamics of and sperm limitation in sexchanging fish. Fishery Bulletin, 102(1):??, 2004. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/effects-size-selective-fisheries-
stock-dynamics-and-sperm-limitation-sex-changing-fish;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2004/1021/
alonzo.pdf.

Baba:2004:EBG

[4623] Katsuhisa Baba, Toshifumi Kawajiri, Yasuhiro Kuwahara, and Shigeru Nakao. An environmentally based growth model that uses finite difference calculus with maximum likelihood method: its application to the brackish water bivalve Corbicula japonica in Lake Abashiri, Japan. Fishery Bulletin, 102(1):14–24, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/environmentally-based-growth-model-uses-finite-
difference-calculus-maximum-likelihood;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2004/1021/baba.pdf.

Brodeur:2004:JSD

[4624] Rick D. Brodeur, Joseph P. Fisher, David J. Teel, Robert L. Emmett, Edmundo Casillas, and Todd W. Miller. Juvenile salmonid distribution, growth, condition, origin, and environmental and species associations in the Northern California Current. Fishery Bulletin, 102(1):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
juvenile-salmonid-distribution-growth-condition-origin-and-
environmental-and-species;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2004/1021/brodeur.pdf.

Garcia-Rodriguez:2004:STV

[4625] Francisco J. Garcia-Rodriguez and David Aurioles-Gamboa. Spatial and temporal variation in the diet of the California sea lion (Zalophus californianus) in the Gulf of California, Mexico. Fishery Bulletin, 102(1):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/spatial-and-temporal-variation-diet-california-sea-
lion-zalophus-californianus-gulf;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2004/1021/garcia.pdf.

Jung:2004:RSS

[4626] Sukgeun Jung and Edward D. Houde. Recruitment and spawningstock biomass distribution of bay anchovy (Anchoa mitchilli) in Chesapeake Bay. Fishery Bulletin, 102(1):??, 2004. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/recruitment-and-spawning-stock-
biomass-distribution-bay-anchovy-anchoa-mitchilli-chesapeake;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2004/1021/jung.pdf.

Kellison:2004:CEE

[4627] Todd G. Kellison and David B. Eggleston. Coupling ecology and economy: modeling optimal release scenarios for summer flounder (Paralichthys dentatus) stock enhancement. Fishery Bulletin, 102(1):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/coupling-ecology-and-
economy-modeling-optimal-release-scenarios-summer-flounder;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2004/1021/kellison.pdf.

Kritzer:2004:SSG

[4628] Jacob P. Kritzer. Sex-specific growth and mortality, spawning season, and female maturation of the stripey bass (Lutjanus carponotatus) on the great barrrier reef. Fishery Bulletin, 102(1):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/sex-specific-
growth-and-mortality-spawning-season-and-female-maturation-
stripey-bass;   https://spo.nmfs.noaa.gov/sites/default/files/
kritzer.pdf.

Orr:2004:EFH

[4629] Anthony J. Orr, Adria S. Banks, Steve Mellman, Harriet R. Huber, Robert L. DeLong, and Robin F. Brown. Examination of the foraging habits of Pacific harbor seal (Phoca vitulina richardsi) to describe their use of the Umpqua River, Oregon, and their predation on salmonids. Fishery Bulletin, 102(1):??, 2004. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/examination-foraging-
habits-pacific-harbor-seal-phoca-vitulina-richardsi-describe-
their-use;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2004/1021/orr.pdf.

Park:2004:LDS

[4630] Wongyu Park, R. Ian Perry, and Sung Yun Hong. Larval development of the sidestriped shrimp (Pandalopsis dispar Rathbun) (Crustacea, Decapoda, Pandalidae) reared in the laboratory. Fishery Bulletin, 102(1):118–126, 2004. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/larval-development-sidestriped-
shrimp-pandalopsis-dispar-rathbun-crustacea-decapoda;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2004/1021/
park.pdf.

Pearson:2004:SAD

[4631] Donald E. Pearson and Franklin R. Shaw. Sources of age determination errors for sablefish (Anoplopoma fimbria). Fishery Bulletin, 102(1): ??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/sources-age-
determination-errors-sablefish-anoplopoma-fimbria;   https:
//spo.nmfs.noaa.gov/sites/default/files/pearson.pdf.

Powell:2004:GMH

[4632] Allyn B. Powell, Robin T. Cheshire, Elisabeth H. Laban, James Colvocoresses, Patrick O’Donnell, and Marie Davidian. Growth, mortality, and hatchdate distributions of larval and juvenile spotted seatrout (Cynoscion nebulosus) in Florida Bay, Everglades National Park. Fishery Bulletin, 102(1):??, 2004. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/growth-mortality-and-hatchdate-
distributions-larval-and-juvenile-spotted-seatrout-cynoscion;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2004/1021/powell.pdf.

Santana:2004:ADG

[4633] Francisco M. Santana and Rosangela Lessa. Age determination and growth of the night shark (Carcharhinus signatus) off the northeastern Brazilian coast. Fishery Bulletin, 102(1):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/age-determination-and-growth-night-shark-
carcharhinus-signatus-northeastern-brazilian-coast;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2004/1021/
santana.pdf.

Smith:2004:DBP

[4634] Keith R. Smith, David A. Somerton, Mei-Sun Yang, and Daniel G. Nichol. Distribution and biology of prowfish (Zaprora silenus) in the northeast Pacific. Fishery Bulletin, 102(1):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/distribution-and-biology-
prowfish-zaprora-silenus-northeast-pacific;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2004/1021/smith.pdf.

Ward:2004:FLS

[4635] Peter Ward, Ransom A. Myers, and Wade Blanchard. Fish lost at sea: the effect of soak time on pelagic longline catches. Fishery Bulletin, 102 (1):179–195, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fish-lost-
sea-effect-soak-time-pelagic-longline-catches;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/2004/1021/ward.
pdf.

Watanabe:2004:EDD

[4636] Chikako Watanabe and Akihiko Yatsu. Effects of density-dependence and sea surface temperature on interannual variation in length-at-age of chub mackerel (Scomber japonicus) in the Kuroshio–Oyashio area during 1970–1997. Fishery Bulletin, 102(1):??, 2004. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/effects-density-dependence-and-
sea-surface-temperature-interannual-variation-length-age;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/2004/
1021/watanabe.pdf.

Llanos-Rivera:2004:LSE

[4637] Alejandra Llanos-Rivera and Leonardo R. Castro. Latitudinal and seasonal egg-size variation of the anchoveta (Engraulis ringens) off the Chilean coast. Fishery Bulletin, 102(1):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/latitudinal-and-seasonal-
egg-size-variation-anchoveta-engraulis-ringens-chilean-coast;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2004/1021/llanos.pdf.

Purcell:2004:MMG

[4638] Maureen Purcell, Greg Mackey, Eric LaHood, Harriet Huber, and Linda Park. Molecular methods for the genetic identification of salmonid prey from Pacific harbor seal (Phoca vitulina richardsi) scat. Fishery Bulletin, 102(1):213–220, 2004. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
molecular-methods-genetic-identification-salmonid-prey-pacific-
harbor-seal-phoca-vitulina;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2004/1021/purcell.pdf.

Weng:2004:DVM

[4639] Kevin C. Weng and Barbara A. Block. Diel vertical migration of the bigeye thresher shark (Alopias superciliosus), a species possessing orbital retia mirabilia. Fishery Bulletin, 102(1):221–229, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/diel-vertical-migration-bigeye-
thresher-shark-alopias-superciliosus-species-possessing;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2004/1021/
weng.pdf.

Archer:2004:AEU

[4640] Frederick Archer, Tim Gerrodette, Susan Chivers, and Alan Jackson. Annual estimates of the unobserved incidental kill of pantropical spotted dolphin (Stenella attenuata attenuata) calves in the tuna purseseine fishery of the eastern tropical Pacific. Fishery Bulletin, 102 (2):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/annual-
estimates-unobserved-incidental-kill-pantropical-spotted-dolphin-
stenella-attenuata;   https://spo.nmfs.noaa.gov/sites/default/
files/archer.pdf.

Chernova:2004:RNS

[4641] Natalia V. Chernova and David L. Stein. A remarkable new species of Psednos (Teleostei: Liparidae) from the western North Atlantic Ocean. Fishery Bulletin, 102(2):??, 2004. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/remarkable-new-species-psednos-
teleostei-liparidae-western-north-atlantic-ocean;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/2004/1022/
chernova.pdf.

Chiang:2004:AGS

[4642] Wei-Chuan Chiang, Chi-Lu Sun, Su-Zan Yeh, and Wei-Cheng Su. Age and growth of sailfish (Istiophorus platypterus) in waters off eastern Taiwan. Fishery Bulletin, 102(2):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/age-and-growth-sailfish-istiophorus-platypterus-
waters-eastern-taiwan;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2004/1022/chiang.pdf.

Clark:2004:HUM

[4643] Randall D. Clark, John D. Christensen, Mark E. Monaco, Philip A. Caldwell, Geoffrey A. Matthews, and Thomas J. Minello. A habitatuse model to determine essential fish habitat for juvenile brown shrimp (Farfantepenaeus aztecus) in Galveston Bay, Texas. Fishery Bulletin, 102(2):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/habitat-use-model-determine-essential-fish-habitat-
juvenile-brown-shrimp-farfantepenaeus;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2004/1022/clark.pdf.

Delgado:2004:TSR

[4644] Gabriel A. Delgado, Claudine T. Bartels, Robert A. Glazer, Nancy J. Brown-Peterson, and Kevin J. McCarthy. Translocation as a strategy to rehabilitate the queen conch (Strombus gigas) population in the Florida keys. Fishery Bulletin, 102(2):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/translocation-strategy-rehabilitate-queen-
conch-strombus-gigas-population-florida-keys;   https://spo.
nmfs.noaa.gov/sites/default/files/delgado.pdf.

Lage:2004:GDA

[4645] Christopher Lage, Kristen Kuhn, and Irv Kornfield. Genetic differentiation among Atlantic cod (Gadus morhua) from Browns Bank, Georges Bank, and Nantucket Shoals. Fishery Bulletin, 102(2):289– 297, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/genetic-
differentiation-among-atlantic-cod-gadus-morhua-browns-bank-
georges-bank-and;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2004/1022/lage.pdf.

Lenihan:2004:COR

[4646] Hunter S. Lenihan and Charles H. Peterson. Conserving oyster reef habitat by switching from dredging and tonging to diverharvesting. Fishery Bulletin, 102(2):298–305, 2004. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/conserving-oyster-reef-habitat-
switching-dredging-and-tonging-diver-harvesting;   https://
spo.nmfs.noaa.gov/sites/default/files/lenihan.pdf.

Macewicz:2004:FED

[4647] Beverly J. Macewicz, John R. Hunter, Nancy C. H. Lo, and Erin L. LaCasella. Fecundity, egg deposition, and mortality of market squid (Loligo opalescens). Fishery Bulletin, 102(2):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/fecundity-egg-deposition-
and-mortality-market-squid-loligo-opalescens;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/2004/1022/macewicz.
pdf.

Orr:2004:DRT

[4648] James W. Orr and James E. Blackburn. The dusky rockfishes (Teleostei: Scorpaeniformes) of the North Pacific Ocean: resurrection of Sebastes variabilis (Pallas, 1814) and a redescription of Sebastes ciliatus (Tilesius, 1813). Fishery Bulletin, 102(2):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic).  URL https://spo.nmfs.noaa.gov/content/dusky-
rockfishes-teleostei-scorpaeniformes-north-pacific-ocean-resurrection-
sebastes;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2004/1022/orr.pdf.

Powers:2004:RER

[4649] Joseph E. Powers. Recruitment as an evolving random process of aggregation and mortality. Fishery Bulletin, 102(2):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/recruitment-
evolving-random-process-aggregation-and-mortality;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2004/1022/
powers.pdf.

Szedlmayer:2004:DSJ

[4650] Stephen T. Szedlmayer and Jason D. Lee. Diet shifts of juvenile red snapper (Lutjanus campechanus) with changes in habitat and fish size. Fishery Bulletin, 102(2):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/diet-shifts-juvenile-red-

snapper-lutjanus-campechanus-changes-habitat-and-fish-size;
 https://spo.nmfs.noaa.gov/sites/default/files/szedlmayer.pdf.

Webb:2004:IGR

[4651] Stacey Webb and Ronald T. Kneib. Individual growth rates and movement of juvenile white shrimp (Litopenaeus setiferus) in a tidal marsh nursery. Fishery Bulletin, 102(2):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/individual-growth-rates-and-movement-
juvenile-white-shrimp-litopenaeus-setiferus-tidal;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2004/1022/
webb.pdf.

Forsythe:2004:DCM

[4652] John Forsythe, Nuutti Kangas, and Roger T. Hanlon. Does the California market squid (Loligo opalescens) spawn naturally during the day or at night? A note on the successful use of ROVs to obtain basic fisheries biology data. Fishery Bulletin, 102(2):389–392, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/does-california-market-squid-loligo-opalescens-
spawn-naturally-during-day-or-night-note;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2004/1022/forsythe.
pdf.

Kotas:2004:ICL

[4653] Jorge E. Kotas, Silvio dos Santos, Venâncio G. de Azevedo, Berenice M. G. Gallo, and Paulo C. R. Barata. Incidental capture of loggerhead (Caretta caretta) and leatherback (Dermochelys coriacea) sea turtles by the pelagic longline fishery off southern Brazil. Fishery Bulletin, 102(2): ??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/incidental-
capture-loggerhead-caretta-caretta-and-leatherback-dermochelys-
coriacea-sea;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2004/1022/kotas.pdf.

Yang:2004:DCP

[4654] Mei-Sun Yang. Diet changes of Pacific cod (Gadus macrocephalus) in Pavlof Bay associated with climate changes in the Gulf of Alaska between 1980 and 1995. Fishery Bulletin, 102(2):??, 2004. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/diet-changes-pacific-cod-gadus-
macrocephalus-pavlof-bay-associated-climate-changes-gulf;

https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/2004/
1022/yang.pdf.

Abascal:2004:TDM

[4655] Francisco J. Abascal, César Megina, and Antonio Medina. Testicular development in migrant and spawning bluefin tuna (Thunnus thynnus (L.)) from the eastern Atlantic and Mediterranean. Fishery Bulletin, 102(3):407–417, 2004. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
testicular-development-migrant-and-spawning-bluefin-tuna-thunnus-
thynnus-l-eastern-atlantic;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2004/1023/abascal.pdf.

Bobko:2004:MOC

[4656] Stephen J. Bobko and Steven A. Berkeley. Maturity, ovarian cycle, fecundity, and age-specific parturition of black rockfish (Sebastes melanops). Fishery Bulletin, 102(3):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/maturity-ovarian-cycle-fecundity-and-
age-specific-parturition-black-rockfish-sebastes;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/2004/1023/
bobko.pdf.

Brock:2004:MOO

[4657] Daniel J. Brock and Timothy M. Ward. Maori octopus (Octopus maorum) bycatch and southern rock lobster (Jasus edwardsii) mortality in the South Australian rock lobster fishery. Fishery Bulletin, 102(3):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/maori-octopus-
octopus-maorum-bycatch-and-southern-rock-lobster-jasus-edwardsii-
mortality;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2004/1023/brock.pdf.

Dawson:2004:SBS

[4658] Stephen Dawson, Elisabeth Slooten, Sam DuFresne, Paul Wade, and Deanna Clement. Small-boat surveys for coastal dolphins: line-transect surveys for Hector’s dolphins (Cephalorhynchus hectori). Fishery Bulletin, 102(3):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/small-
boat-surveys-coastal-dolphins-line-transect-surveys-hectors-
dolphins-cephalorhynchus;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2004/1023/dawson.pdf.

Laidig:2004:DGL

[4659] Thomas E. Laidig, Keith M. Sakuma, and Jason A. Stannard. Description and growth of larval and pelagic juvenile pygmy rockfish (Sebastes wilsoni) (family Sebastidae). Fishery Bulletin, 102(3):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/description-
and-growth-larval-and-pelagic-juvenile-pygmy-rockfish-sebastes-
wilsoni-family;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2004/1023/laidig.pdf.

McGarvey:2004:EER

[4660] Richard McGarvey. Estimating the emigration rate of fish stocks from marine sanctuaries using tag-recovery data. Fishery Bulletin, 102(3):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/estimating-
emigration-rate-fish-stocks-marine-sanctuaries-using-tag-recovery-
data;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2004/1023/mcgarvey.pdf.

Roumillat:2004:RDF

[4661] William A. Roumillat and Myra C. Brouwer. Reproductive dynamics of female spotted seatrout (Cynoscion nebulosus) in South Carolina. Fishery Bulletin, 102(3):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/reproductive-dynamics-female-spotted-
seatrout-cynoscion-nebulosus-south-carolina;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/2004/1023/roumillat.
pdf.

Taggart:2004:EDC

[4662] S. James Taggart, Charles E. O’Clair, Thomas C. Shirley, and Jennifer Mondragon. Estimating Dungeness crab (Cancer magister) abundance: crab pots and dive transects compared. Fishery Bulletin, 102(3):488–497, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/estimating-dungeness-crab-cancer-magister-
abundance-crab-pots-and-dive-transects-compared;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/2004/1023/
taggart.pdf.

Tollit:2004:MIS

[4663] Dominic J. Tollit, Susan G. Heaslip, Tonya K. Zeppelin, Ruth Joy, Katherine A. Call, and Andrew W. Trites. A method to improve size estimates of walleye pollock (Theragra chalcogramma) and Atka mackerel (Pleurogrammus monopterygius) consumed by pinnipeds: digestion correction factors applied to bones and otoliths recovered in scats. Fishery Bulletin, 102(3):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
method-improve-size-estimates-walleye-pollock-theragra-chalcogramma-
and-atka-mackerel;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2004/1023/tollit1.pdf.

Zeppelin:2004:SWP

[4664] Tonya K. Zeppelin, Dominic J. Tollit, Katherine A. Call, Trevor J. Orchard, and Carolyn J. Gudmundson. Sizes of walleye pollock (Theragra chalcogramma) and Atka mackerel (Pleurogrammus monopterygius) consumed by the western stock of Steller sea lions (Eumetopias jubatus) in Alaska from 1998 to 2000. Fishery Bulletin, 102(3):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/sizes-walleye-pollock-theragra-
chalcogramma-and-atka-mackerel-pleurogrammus-monopterygius;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2004/1023/zeppelin.pdf.

Tollit:2004:SWP

[4665] Dominic J. Tollit, Susan G. Heaslip, and Andrew Trites. Sizes of walleye pollock (Theragra chalcogramma) consumed by the eastern stock of Steller sea lions (Eumetopias jubatus) in Southeast Alaska from 1994 to 1999. Fishery Bulletin, 102(3):522–532, 2004. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/sizes-walleye-pollock-theragra-
chalcogramma-consumed-eastern-stock-steller-sea-lions;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2004/1023/
tollit2.pdf.

Tremain:2004:MMS

[4666] Derek M. Tremain, Christopher W. Harnden, and Douglas H. Adams. Multidirectional movements of sportfish species between an estuarine no-take zone and surrounding waters of the Indian River Lagoon, Florida. Fishery Bulletin, 102(3):533–544, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/multidirectional-movements-sportfish-
species-between-estuarine-no-take-zone-and-surrounding;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2004/1023/
tremain.pdf.

Wells:2004:DAG

[4667] R. J. David Wells and Jay R. Rooker. Distribution, age, and growth of young-of-the-year greater amberjack (Seriola dumerili) associated with pelagic Sargassum. Fishery Bulletin, 102(3):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/distribution-age-and-
growth-young-year-greater-amberjack-seriola-dumerili-associated;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2004/1023/wells.pdf.

Hiroishi:2004:IFP

[4668] Shingo Hiroishi, Yasutaka Yuki, Eriko Yuruzume, Yosuke Onishi, Tomoji Ikeda, Hironobu Komaki, and Muneo Okiyama. Identification of formalin-preserved eggs of red sea bream (Pagrus major) (Pisces: Sparidae) using monoclonal antibodies. Fishery Bulletin, 102(3):555–560, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/identification-formalin-preserved-eggs-red-sea-
bream-pagrus-major-pisces-sparidae-using;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2004/1023/hiroishi.
pdf.

Calambokidis:2004:DAH

[4669] John Calambokidis, Gretchen H. Steiger, David K. Ellifrit, Barry L. Troutman, and C. Edward Bowlby. Distribution and abundance of humpback whales (Megaptera novaeangliae) and other marine mammals off the northern Washington coast. Fishery Bulletin, 102(4):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/distribution-
and-abundance-humpback-whales-megaptera-novaeangliae-and-other-
marine-mammals;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2004/1024/calam.pdf.

Danilewicz:2004:RBM

[4670] Daniel Danilewicz, Juan A. Claver, Alejo L. Pérez Carrera, Eduardo R. Secchi, and Nelson F. Fontoura. Reproductive biology of male franciscanas (Pontoporia blainvillei) (Mammalia: Cetacea) from Rio Grande do Sul, southern Brazil. Fishery Bulletin, 102(4):581–592, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/reproductive-biology-
male-franciscanas-pontoporia-blainvillei-mammalia-cetacea-rio-
grande;   https://spo.nmfs.noaa.gov/sites/default/files/danil.
pdf.

Fischer:2004:RSL

[4671] Andrew J. Fischer, M. Scott Baker, Jr., and Charles A. Wilson. Red snapper (Lutjanus campechanus) demographic structure in the northern Gulf of Mexico based on spatial patterns in growth rates and morphometrics. Fishery Bulletin, 102(4):??, 2004. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/red-snapper-lutjanus-campechanus-
demographic-structure-northern-gulf-mexico-based-spatial;
https://spo.nmfs.noaa.gov/sites/default/files/fisch.pdf.

FitzGerald:2004:ESO

[4672] Jennifer L. FitzGerald, Simon R. Thorrold, Kevin M. Bailey, Annette L. Brown, and Kenneth P. Severin. Elemental signatures in otoliths of larval walleye pollock (Theragra chalcogramma) from the northeast Pacific Ocean. Fishery Bulletin, 102(4):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/elemental-signatures-otoliths-larval-
walleye-pollock-theragra-chalcogramma-northeast;   https://
spo.nmfs.noaa.gov/sites/default/files/fitzg.pdf.

Gaughan:2004:SCD

[4673] Daniel J. Gaughan, Timothy I. Leary, Ronald W. Mitchell, and Ian W. Wright. A sudden collapse in distribution of Pacific sardine (Sardinops sagax) off southwestern Australia enables an objective re-assessment of biomass estimates. Fishery Bulletin, 102(4):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/sudden-collapse-distribution-pacific-
sardine-sardinops-sagax-southwestern-australia-enables;   https:
//spo.nmfs.noaa.gov/sites/default/files/gaugh.pdf.

Griffiths:2004:FRT

[4674] Shane P. Griffiths, Ron J. West, Andy R. Davis, and Ken G. Russell. Fish recolonization in temperate Australian rockpools: a quantitative experimental approach. Fishery Bulletin, 102(4): ??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/fish-
recolonization-temperate-australian-rockpools-quantitative-
experimental-approach;   https://spo.nmfs.noaa.gov/content/
fish-recolonization-temperate-australian-rockpools-quantitative-
experimental-approach-0;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2004/1024/griff.pdf.

Hesp:2004:FIT

[4675] S. Alexander Hesp, Ian C. Potter, and Sonja R. M. Schubert. Factors influencing the timing and frequency of spawning and fecundity of the goldlined seabream (Rhabdosargus sarba) (Sparidae) in the lower reaches of an estuary. Fishery Bulletin, 102(4):648–660, 2004. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/factors-influencing-timing-and-
frequency-spawning-and-fecundity-goldlined-seabream;   https:
//spo.nmfs.noaa.gov/sites/default/files/hesp.pdf.

Maxwell:2004:FDC

[4676] Michael R. Maxwell, Annette Henry, Christopher D. Elvidge, Jeffrey Safran, Vinita R. Hobson, Ingrid Nelson, Benjamin T. Tuttle, John B. Dietz, and John R. Hunter. Fishery dynamics of the California market squid (Loligo opalescens), as measured by satellite remote sensing. Fishery Bulletin, 102(4):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishery-dynamics-california-market-squid-loligo-opalescens-
measured-satellite-remote;   https://spo.nmfs.noaa.gov/sites/
default/files/maxwe.pdf.

Murray:2004:MDS

[4677] Kimberly T. Murray. Magnitude and distribution of sea turtle bycatch in the sea scallop (Placopecten magellanicus) dredge fishery in two areas of the northwestern Atlantic Ocean, 2001–2002. Fishery Bulletin, 102(4): ??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/magnitude-and-
distribution-sea-turtle-bycatch-sea-scallop-placopecten-magellanicus-
dredge;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2004/1024/murra.pdf.

Snover:2004:VIA

[4678] Melissa L. Snover and Aleta A. Hohn. Validation and interpretation of annual skeletal marks in loggerhead (Caretta caretta) and Kemp’s ridley (Lepidochelys kempii) sea turtles. Fishery Bulletin, 102(4):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/validation-
and-interpretation-annual-skeletal-marks-loggerhead-caretta-
caretta-and-kemps;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2004/1024/snove.pdf.

Stehlik:2004:HRE

[4679] Linda L. Stehlik, Robert A. Pikanowski, and Donald G. McMillan. The Hudson–Raritan Estuary as a crossroads for distribution of blue (Callinectes sapidus), lady (Ovalipes ocellatus), and Atlantic rock (Cancer irroratus) crabs. Fishery Bulletin, 102(4):693– 710, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/hudson-
raritan-estuary-crossroads-distribution-blue-callinectes-sapidus-
lady-ovalipes;   https://spo.nmfs.noaa.gov/sites/default/files/
stehl.pdf.

Stevens:2004:RVA

[4680] Melissa M. Stevens, Allen H. Andrews, Gregor M. Cailliet, Kenneth H. Coale, and Craig C. Lundstrom. Radiometric validation of age, growth, and longevity for the blackgill rockfish (Sebastes melanostomus). Fishery Bulletin, 102(4):711–722, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/lundstrom-radiometric-validation-age-
growth-and-longevity-blackgill-rockfish-sebastes;   https:/
/spo.nmfs.noaa.gov/content/radiometric-validation-age-growth-
and-longevity-blackgill-rockfish-sebastes-melanostomus;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2004/1024/
steve.pdf.

Tolan:2004:DLP

[4681] James M. Tolan and David A. Newstead. Descriptions of larval, prejuvenile, and juvenile finescale menhaden (Brevoortia gunteri) (family Clupeidae), and comparisons to Gulf menhaden (B. patronus). Fishery Bulletin, 102(4):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/descriptions-larval-prejuvenile-
and-juvenile-finescale-menhaden-brevoortia-gunteri-family;
https://spo.nmfs.noaa.gov/content/descriptions-larval-prejuvenile-
and-juvenile-finescale-menhaden-brevoortia-gunteri-family-0;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2004/1024/tolan.pdf.

Uchikawa:2004:DMA

[4682] Kazuhisa Uchikawa, John R. Bower, Yasuko Sato, and Yasunori Sakurai. Diet of the minimal armhook squid (Berryteuthis anonychus) (cephalopoda: Gonatidae) in the northeast Pacific during spring. Fishery Bulletin, 102(4):733–739, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/diet-minimal-armhook-squid-berryteuthis-anonychus-cephalopoda-
gonatidae-northeast-pacific;   https://spo.nmfs.noaa.gov/content/
diet-minimal-armhook-squid-berryteuthis-anonychus-cephalopoda-
gonatidae-northeast-pacific-0;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2004/1024/uchik.pdf.

Weinberg:2004:CPS

[4683] Kenneth L. Weinberg, Robert S. Otto, and David A. Somerton. Capture probability of a survey trawl for red king crab (Paralithodes camtschaticus). Fishery Bulletin, 102(4):740–749, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/capture-probability-survey-trawl-red-king-
crab-paralithodes-camtschaticus.

Kerstetter:2004:ESP

[4684] David W. Kerstetter, Jeffery J. Polovina, and John E. Graves. Evidence of shark predation and scavenging on fishes equipped with pop-up satellite archival tags. Fishery Bulletin, 102(4):750–756, 2004. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/evidence-shark-predation-
and-scavenging-fishes-equipped-pop-satellite-archival-tags;
 https://spo.nmfs.noaa.gov/content/evidence-shark-predation-
and-scavenging-fishes-equipped-pop-satellite-archival-tags-
0;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2004/1024/kerst.pdf;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/kerst.pdf.

Laptikhovsky:2004:SRR

[4685] Vladimir V. Laptikhovsky. Survival rates of rays discarded by the bottom trawl squid fishery off the Falkland Islands. Fishery Bulletin, 102(4):??, 2004. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/survival-rates-rays-
discarded-bottom-trawl-squid-fishery-falkland-islands;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
lapti.pdf.

Bochenek:2005:ASS

[4686] Eleanor A. Bochenek, Eric N. Powell, Allison J. Bonner, and Sarah E. Banta. An assessment of scup (Stenotomus chrysops) and black sea bass (Centropristas striata) discards in the directed otter trawl fisheries in the Mid-Atlantic Bight. Fishery Bulletin, 103(1):1–14, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/assessment-
scup-stenotomus-chrysops-and-black-sea-bass-centropristas-striata-
discards;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2005/1031/boch.pdf.

Cooper:2005:FST

[4687] Daniel W. Cooper, Katherine E. Pearson, and Donald R. Gunderson. Fecundity of shortspine thornyhead (Sebastolobus alascanus) and longspine thornyhead (S. altivelis) (Scorpaenidae) from the northeastern Pacific Ocean, determined by stereological and gravimetric techniques. Fishery Bulletin, 103(1):15–22, 2005. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/fecundity-shortspine-
thornyhead-sebastolobus-alascanus-and-longspine-thornyhead-
s-altivelis;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2005/1031/coop.pdf.

DeMartini:2005:RPL

[4688] Edward E. DeMartini, Marti L. McCracken, Robert B. Moffitt, and Jerry A. Wetherall. Relative pleopod length as an indicator of size at sexual maturity in slipper (Scyllarides squammosus) and spiny Hawaiian (Panulirus marginatus) lobsters. Fishery Bulletin, 103(1): ??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/relative-
pleopod-length-indicator-size-sexual-maturity-slipper-scyllarides-
squammosus-and;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2005/1031/dema.pdf.

Fisher:2005:SCG

[4689] Joseph P. Fisher and William G. Pearcy. Seasonal changes in growth of Coho salmon (Oncorhynchus kisutch) off Oregon and Washington and concurrent changes in the spacing of scale circuli. Fishery Bulletin, 103 (1):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/seasonal-
changes-growth-coho-salmon-oncorhynchus-kisutch-oregon-and-
washington-and;   https://spo.nmfs.noaa.gov/sites/default/files/
fish.pdf.

Groeneveld:2005:ECR

[4690] Johan C. Groeneveld, Jimmy P. Khanyile, and David S. Schoeman. Escapement of the Cape rock lobster (Jasus lalandii) through the mesh and entrance of commercial traps. Fishery Bulletin, 103(1):52–62, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/escapement-
cape-rock-lobster-jasus-lalandii-through-mesh-and-entrance-
commercial-traps;   https://spo.nmfs.noaa.gov/sites/default/
files/groe.pdf.

Grusha:2005:QDL

[4691] Donna S. Grusha and Mark R. Patterson. Quantification of drag and lift imposed by pop-up satellite archival tags and estimation of the metabolic cost to cownose rays (Rhinoptera bonasus). Fishery Bulletin, 103(1):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/quantification-drag-and-lift-imposed-pop-satellite-
archival-tags-and-estimation-metabolic;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2005/1031/grus.pdf.

Harvey:2005:ENN

[4692] Chris J. Harvey. Effects of El Niño events on energy demand and egg production of rockfish (Scorpaenidae: Sebastes): a bioenergetics approach. Fishery Bulletin, 103(1):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/effects-el-nino-events-energy-demand-
and-egg-production-rockfish-scorpaenidae-sebastes;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2005/1031/
harv.pdf.

Horodysky:2005:APS

[4693] Andrij Z. Horodysky and John Graves. Application of pop-up satellite archival tag technology to estimate postrelease survival of white marlin (Tetrapturus albidus) caught on circle and straight-shank (“J”) hooks in the western North Atlantic recreational fishery. Fishery Bulletin, 103(1): ??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/application-
pop-satellite-archival-tag-technology-estimate-postrelease-
survival-white;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2005/1031/horo.pdf.

Kerr:2005:AVQ

[4694] Lisa A. Kerr, Allen H. Andrews, Kristen Munk, Kenneth H. Coale, Brian R. Frantz, Gregor M. Cailliet, and Thomas A. Brown. Age validation of quillback rockfish (Sebastes maliger) using bomb radiocarbon. Fishery Bulletin, 103(1):??, 2005. CODEN FSY-

BAY.  ISSN 0090-0656 (print), 1937-4518 (electronic).  URL

https://spo.nmfs.noaa.gov/content/age-validation-quillback-
rockfish-sebastes-maliger-using-bomb-radiocarbon;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/2005/1031/
kerr.pdf.

Marancik:2005:CSS

[4695] Katrin E. Marancik, Lisa M. Clough, and Jonathan A. Hare. Crossshelf and seasonal variation in larval fish assemblages on the southeast United States continental shelf off the coast of Georgia. Fishery Bulletin, 103(1):108–129, 2005. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
cross-shelf-and-seasonal-variation-larval-fish-assemblages-
southeast-united-states;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2005/1031/mara.pdf.

OFarrell:2005:YCF

[4696] Michael R. O’Farrell and Ralph J. Larson. Year-class formation in Pacific herring (Clupea pallasi) estimated from spawning-date distributions of juveniles in San Francisco Bay, California. Fishery Bulletin, 103(1):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/year-class-formation-
pacific-herring-clupea-pallasi-estimated-spawning-date-distributions;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2005/1031/ofar.pdf.

Parker:2005:DOL

[4697] Denise M. Parker, William J. Cooke, and George H. Balazs. Diet of oceanic loggerhead sea turtles (Caretta caretta) in the central North Pacific. Fishery Bulletin, 103(1):142–152, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/diet-oceanic-loggerhead-sea-turtles-caretta-
caretta-central-north-pacific;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2005/1031/park.pdf.

Roberson:2005:IVA

[4698] Nancy E. Roberson, Daniel K. Kimura, Donald R. Gunderson, and Allen M. Shimada. Indirect validation of the age-reading method for Pacific cod (Gadus macrocephalus) using otoliths from marked and recaptured fish. Fishery Bulletin, 103(1):??, 2005. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/indirect-validation-age-reading-
method-pacific-cod-gadus-macrocephalus-using-otoliths;   https:

//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2005/1031/
robe.pdf.

Sulikowski:2005:AGE

[4699] James A. Sulikowski, Jeff Kneebone, Scott Elzey, Joe Jurek, Patrick D. Danley, W. Huntting Howell, and Paul C. W. Tsang. Age and growth estimates of the thorny skate (Amblyraja radiata) in the western Gulf of Maine. Fishery Bulletin, 103(1):161–168, 2005. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/age-and-growth-estimates-
thorny-skate-amblyraja-radiata-western-gulf-maine;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2005/1031/
suli.pdf.

Tracey:2005:AVG

[4700] Sean R. Tracey and Jeremy M. Lyle. Age validation, growth modeling, and mortality estimates for striped trumpeter (Latris lineata) from southeastern Australia: making the most of patchy data. Fishery Bulletin, 103(1):169–182, 2005. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/age-validation-growth-modeling-
and-mortality-estimates-striped-trumpeter-latris-lineata;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/2005/
1031/trac.pdf.

Trnski:2005:LDE

[4701] Thomas Trnski, Amanda C. Hay, and D. Stewart Fielder. Larval development of estuary perch (Macquaria colonorum) and Australian bass (M. novemaculeata) (Perciformes: Percichthyidae), and comments on their life history. Fishery Bulletin, 103(1):183–194, 2005. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/larval-development-estuary-perch-
macquaria-colonorum-and-australian-bass-m-novemaculeata;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2005/1031/
trns.pdf.

Venerus:2005:ELH

[4702] Leonardo A. Venerus, Laura Machinandiarena, Martin D. Ehrlich, and Ana M. Parma. Early life history of the Argentine sandperch Pseudopercis semifasciata (Pinguipedidae) off northern Patagonia. Fishery Bulletin, 103(1):195–206, 2005. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
early-life-history-argentine-sandperch-pseudopercis-semifasciata-

pinguipedidae-northern;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2005/1031/vene.pdf.

Wilson:2005:GVA

[4703] Matthew T. Wilson, Annette L. Brown, and Kathryn L. Mier. Geographic variation among age-0 walleye pollock (Theragra chalcogramma): evidence of mesoscale variation in nursery quality? Fishery Bulletin, 103(1):207–218, 2005. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
geographic-variation-among-age-0-walleye-pollock-theragra-chalcogramma-
evidence-mesoscale;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2005/1031/wils.pdf.

Markaida:2005:TSJ

[4704] Unai Markaida, Joshua J. C. Rosenthal, and William F. Gilly. Tagging studies on the jumbo squid (Dosidicus gigas) in the Gulf of California, Mexico. Fishery Bulletin, 103(1):219–226, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/tagging-studies-jumbo-squid-dosidicus-
gigas-gulf-california-mexico;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2005/1031/mark.pdf.

Alonzo:2005:SCR

[4705] Suzanne H. Alonzo and Marc Mangel. Sex-change rules, stock dynamics, and the performance of spawning-per-recruit measures in protogynous stocks. Fishery Bulletin, 103(2):??, 2005. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/sex-change-rules-stock-dynamics-
and-performance-spawning-recruit-measures-protogynous;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2005/1032/
alonzo.pdf.

Brandon:2005:NGS

[4706] Elisif A. A. Brandon, Donald G. Calkins, Thomas R. Loughlin, and Randall W. Davis. Neonatal growth of Steller sea lion (Eumetopias jubatus) pups in Alaska. Fishery Bulletin, 103(2):??, 2005. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/neonatal-growth-steller-sea-lion-
eumetopias-jubatus-pups-alaska;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2005/1032/brandon.pdf.

Brouwer:2005:RBC

[4707] Stephen L. Brouwer and Marc H. Griffiths. Reproductive biology of carpenter seabream (Argyrozona argyrozona) (Pisces: Sparidae) in a marine protected area. Fishery Bulletin, 103(2):??, 2005. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/reproductive-biology-carpenter-
seabream-argyrozona-argyrozona-pisces-sparidae-marine;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2005/1032/
brouwer.pdf.

Burn:2005:DSO

[4708] Douglas M. Burn and Angela M. Doroff. Decline in sea otter (Enhydra lutris) populations along the Alaska Peninsula, 1986–2001. Fishery Bulletin, 103(2):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/decline-sea-otter-enhydra-lutris-populations-along-
alaska-peninsula-1986-2001;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2005/1032/burn.pdf.

Carlson:2005:GDS

[4709] John K. Carlson and Ivy E. Baremore. Growth dynamics of the spinner shark (Carcharhinus brevipinna) off the United States southeast and Gulf of Mexico coasts: a comparison of methods. Fishery Bulletin, 103 (2):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/growth-
dynamics-spinner-shark-carcharhinus-brevipinna-united-states-
southeast-and-gulf;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2005/1032/carlson.pdf.

Domeier:2005:TPB

[4710] Michael L. Domeier, Dale Kiefer, Nicole Nasby-Lucas, Adam Wagschal, and Frank O’Brien. Tracking Pacific bluefin tuna (Thunnus thynnus orientalis) in the northeastern Pacific with an automated algorithm that estimates latitude by matching sea-surface-temperature data from satellites with temperature data from tags on fish. Fishery Bulletin, 103 (2):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/tracking-
pacific-bluefin-tuna-thunnus-thynnus-orientalis-northeastern-
pacific-automated;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2005/1032/domeier.pdf.

Fischer:2005:AGM

[4711] Andrew J. Fischer, M. Scott Baker, Jr., Charles A. Wilson, and David L. Nieland. Age, growth, mortality, and radiometric age validation of gray snapper (Lutjanus griseus) from Louisiana. Fishery Bulletin, 103(2):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/age-growth-mortality-
and-radiometric-age-validation-gray-snapper-lutjanus-griseus-
louisiana;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2005/1032/fischer.pdf.

Grabowski:2005:EES

[4712] Robert C. Grabowski, Thomas Windholz, and Yong Chen. Estimating exploitable stock biomass for the Maine green sea urchin (Strongylocentrotus droebachiensis) fishery using a spatial statistics approach. Fishery Bulletin, 103(2):320–330, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/estimating-exploitable-stock-biomass-
maine-green-sea-urchin-strongylocentrotus;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2005/1032/grabowski.
pdf.

Lowry:2005:ADC

[4713] Mark S. Lowry and Karin A. Forney. Abundance and distribution of California sea lions (Zalophus californianus) in central and northern California during 1998 and summer 1999. Fishery Bulletin, 103(2):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/abundance-and-distribution-california-sea-
lions-zalophus-californianus-central-and-northern;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2005/1032/
lowry.pdf.

Mackie:2005:VSF

[4714] Michael C. Mackie, Paul D. Lewis, Daniel J. Gaughan, and Stephen J. Newman. Variability in spawning frequency and reproductive development of the narrow-barred Spanish mackerel (Scomberomorus commerson) along the west coast of Australia. Fishery Bulletin, 103(2):344–354, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/variability-spawning-
frequency-and-reproductive-development-narrow-barred-spanish-
mackerel;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2005/1032/mackie.pdf.

Ruggerone:2005:SMG

[4715] Gregory T. Ruggerone, Ed Farley, Jennifer Nielson, and Peter Hagen. Seasonal marine growth of Bristol Bay sockeye salmon (Oncorhynchus nerka) in relation to competition with Asian pink salmon (O. gorbuscha) and the 1977 ocean regime shift. Fishery Bulletin, 103 (2):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/seasonal-
marine-growth-bristol-bay-sockeye-salmon-oncorhynchus-nerka-
relation-competition;   https://spo.nmfs.noaa.gov/sites/default/
files/ruggerone.pdf.

Shoji:2005:DFC

[4716] Jun Shoji and Masaru Tanaka. Distribution, feeding condition, and growth of Japanese Spanish mackerel (Scomberomorus niphonius) larvae in the Seto Inland Sea. Fishery Bulletin, 103(2):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/distribution-feeding-condition-
and-growth-japanese-spanish-mackerel-scomberomorus-niphonius;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2005/1032/shoji.pdf.

Wang:2005:MLE

[4717] You-Gan Wang and Nick Ellis. Maximum likelihood estimation of mortality and growth with individual variability from multiple length-frequency data. Fishery Bulletin, 103(2):??, 2005. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/maximum-likelihood-
estimation-mortality-and-growth-individual-variability-multiple-
length;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2005/1032/wang.pdf.

Williams:2005:EFG

[4718] Erik H. Williams and Kyle W. Shertzer. Effects of fishing on growth traits: a simulation analysis. Fishery Bulletin, 103(2): ??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/effects-
fishing-growth-traits-simulation-analysis;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2005/1032/williams.
pdf.

Burton:2005:PEI

[4719] Michael L. Burton, Kenneth J. Brennan, Roldan C. Muñoz, and Richard O. Parker, Jr. Preliminary evidence of increased spawning aggregations of mutton snapper (Lutjanus analis) at Riley’s Hump two years after establishment of the Tortugas South Ecological Reserve. Fishery Bulletin, 103(2):??, 2005. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/preliminary-evidence-increased-
spawning-aggregations-mutton-snapper-lutjanus-analis-rileys;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2005/1032/burton.pdf.

Carpentieri:2005:FHE

[4720] Paolo Carpentieri, Francesco Colloca, Massimiliano Cardinale, Andrea Belluscio, and Giandomenico D. Ardizzone. Feeding habits of European hake (Merluccius merluccius) in the central Mediterranean Sea. Fishery Bulletin, 103(2):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
feeding-habits-european-hake-merluccius-merluccius-central-
mediterranean-sea;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2005/1032/carpentieri.pdf.

Gobert:2005:BQS

[4721] Bertrand Gobert, Alain Guillou, Peter Murray, Patrick Berthou, Maria D. Oqueli Turcios, Ester Lopez, Pascal Lorance, Jérôme Huet, Nicolas Diaz, and Paul Gervain. Biology of queen snapper (Etelis oculatus: Lutjanidae) in the Caribbean. Fishery Bulletin, 103(2):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
biology-queen-snapper-etelis-oculatus-lutjanidae-caribbean;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2005/1032/gobert.pdf.

Graham:2005:CSB

[4722] Rachel T. Graham and Daniel W. Castellanos. Courtship and spawning behaviors of carangid species in Belize. Fishery Bulletin, 103(2):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
courtship-and-spawning-behaviors-carangid-species-belize;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/2005/
1032/graham.pdf.

Hewitt:2005:CTA

[4723] David A. Hewitt and John M. Hoenig. Comparison of two approaches for estimating natural mortality based on longevity. Fishery Bulletin, 103(2):

??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/comparison-
two-approaches-estimating-natural-mortality-based-longevity;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2005/1032/hewitt.pdf.

Lindquist:2005:ECS

[4724] David C. Lindquist and Richard F. Shaw. Effects of current speed and turbidity on stationary light-trap catches of larval and juvenile fishes. Fishery Bulletin, 103(2):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/effects-current-speed-and-turbidity-stationary-
light-trap-catches-larval-and-juvenile;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2005/1032/lindquist.
pdf.

Macchi:2005:CCS

[4725] Gustavo J. Macchi, Marcelo Pájaro, and Adrián Madirolas. Can a change in the spawning pattern of Argentine hake (Merluccius hubbsi) affect its recruitment? Fishery Bulletin, 103(2):445–452, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/can-change-spawning-
pattern-argentine-hake-merluccius-hubbsi-affect-its-recruitment;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2005/1032/macchi.pdf.

Raymundo-Huizar:2005:FHD

[4726] Alma R. Raymundo-Huizar, Horacio Pérez-España, Maite Mascaró, and Xavier Chiappa-Carrara. Feeding habits of the dwarf weakfish (Cynoscion nannus) off the coasts of Jalisco and Colima, Mexico. Fishery Bulletin, 103(2):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/feeding-habits-dwarf-weakfish-cynoscion-nannus-
coasts-jalisco-and-colima-mexico;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2005/1032/huizar.pdf.

Wood:2005:UBM

[4727] Anthony D. Wood. Using bone measurements to estimate the original sizes of bluefish (Pomatomus saltatrix) from digested remains. Fishery Bulletin, 103(2):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/using-bone-measurements-estimate-original-
sizes-bluefish-pomatomus-saltatrix-digested;   https://spo.

nmfs.noaa.gov/sites/default/files/pdf-content/2005/1032/wood.
pdf.

Dressel:2005:UPI

[4728] Sherri C. Dressel and Brenda L. Norcross. Using poststratification to improve abundance estimates from multispecies surveys: a study of juvenile flatfishes. Fishery Bulletin, 103(3):??, 2005. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/using-poststratification-improve-
abundance-estimates-multispecies-surveys-study-juvenile;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2005/1033/
dressel.pdf.

Francis:2005:LMT

[4729] Malcolm P. Francis and Clinton Duffy. Length at maturity in three pelagic sharks (Lamna nasus, Isurus oxyrinchus, and Prionace glauca) from New Zealand. Fishery Bulletin, 103(3):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/length-maturity-three-pelagic-sharks-
lamna-nasus-isurus-oxyrinchus-and-prionace-glauca-new;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2005/1033/
francis.pdf.

Fritz:2005:SFD

[4730] Lowell W. Fritz and Eric S. Brown. Survey- and fishery-derived estimates of Pacific cod (Gadus macrocephalus) biomass: implications for strategies to reduce interactions between groundfish fisheries and Steller sea lions (Eumetopias jubatus). Fishery Bulletin, 103(3):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/survey-and-fishery-
derived-estimates-pacific-cod-gadus-macrocephalus-biomass-implications;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2005/1033/fritz.pdf.

Greig:2005:MGS

[4731] Thomas W. Greig, M. Katherine Moore, Cheryl M. Woodley, and Joseph M. Quattro. Mitochondrial gene sequences useful for species identification of western North Atlantic Ocean sharks. Fishery Bulletin, 103 (3):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/mitochondrial-
gene-sequences-useful-species-identification-western-north-
atlantic-ocean;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2005/1033/greig.pdf.

Hawkins:2005:GVR

[4732] Sharon L. Hawkins, Jonathan Heifetz, Christine M. Kondzela, John E. Pohl, Richard L. Wilmot, Oleg N. Katugin, and Vladimir N. Tuponogov. Genetic variation of rougheye rockfish (Sebastes aleutianus) and shortraker rockfish (S. borealis) inferred from allozymes. Fishery Bulletin, 103(3):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/genetic-
variation-rougheye-rockfish-sebastes-aleutianus-and-shortraker-
rockfish-s-borealis;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2005/1033/hawkins.pdf.

Sulikowski:2005:RCT

[4733] James A. Sulikowski, Jeff Kneebone, Scott Elzey, Joe Jurek, Patrick D. Danley, W. Huntting Howell, and Paul C. W. Tsang. The reproductive cycle of the thorny skate (Amblyraja radiata) in the western Gulf of Maine. Fishery Bulletin, 103(3):??, 2005. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/reproductive-cycle-
thorny-skate-amblyraja-radiata-western-gulf-maine;   https:
//spo.nmfs.noaa.gov/sites/default/files/sulikowski.pdf.

Fey:2005:ETO

[4734] Dariusz P. Fey, Gretchen E. Bath Martin, James A. Morris, and Jonathan Hare. Effect of type of otolith and preparation technique on age estimation of larval and juvenile spot (Leiostomus xanthurus). Fishery Bulletin, 103(3):544–552, 2005. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
effect-type-otolith-and-preparation-technique-age-estimation-
larval-and-juvenile-spot;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2005/1033/fey.pdf.

Piner:2005:PUO

[4735] Kevin R. Piner, Melissa A. Haltuch, and John R. Wallace. Preliminary use of oxygen stable isotopes and the 1983 El Niño to assess the accuracy of aging black rockfish (Sebastes melanops). Fishery Bulletin, 103(3):553–558, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/preliminary-use-oxygen-stable-isotopes-and-1983-el-
nino-assess-accuracy-aging-black;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2005/1033/piner.pdf.

Kingsford:2005:PGM

[4736] Michael J. Kingsford and Julian M. Hughes. Patterns of growth, mortality, and size of the tropical damselfish Acanthochromis polyacanthus across the continental shelf of the Great Barrier Reef. Fishery Bulletin, 103(4):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/patterns-
growth-mortality-and-size-tropical-damselfish-acanthochromis-
polyacanthus-across;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2005/1034/king.pdf.

Kotwicki:2005:VDW

[4737] Stan Kotwicki, Troy W. Buckley, Taina Honkalehto, and Gary Walters. Variation in the distribution of walleye pollock (Theragra chalcogramma) with temperature and implications for seasonal migration. Fishery Bulletin, 103(4):??, 2005. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/variation-distribution-walleye-
pollock-theragra-chalcogramma-temperature-and-implications;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2005/1034/kotw.pdf.

Luthy:2005:TIL

[4738] Stacy A. Luthy, Robert K. Cowen, Joseph E. Serafy, and Jan R. Mc-Dowell. Toward identification of larval sailfish (Istiophorus platypterus), white marlin (Tetrapturus albidus), and blue marlin (Makaira nigricans) in the western North Atlantic Ocean. Fishery Bulletin, 103 (4):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/toward-
identification-larval-sailfish-istiophorus-platypterus-white-
marlin-tetrapturus;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2005/1034/luth.pdf.

McDonough:2005:SDG

[4739] Christopher J. McDonough, William A. Roumillat, and Charles A. Wenner. Sexual differentiation and gonad development in striped mullet (Mugil cephalus L.) from South Carolina estuaries. Fishery Bulletin, 103 (4):601–618, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/sexual-
differentiation-and-gonad-development-striped-mullet-mugil-
cephalus-l-south-carolina;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2005/1034/mcdo.pdf.

Megalofonou:2005:ICE

[4740] Persefoni Megalofonou, Constantinos Yannopoulos, Dimitrios Damalas, Gregorio De Metrio, Michele Deflorio, Jose M. de la Serna, and David Macias. Incidental catch and estimated discards of pelagic sharks from the swordfish and tuna fisheries in the Mediterranean Sea. Fishery Bulletin, 103(4):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
incidental-catch-and-estimated-discards-pelagic-sharks-swordfish-
and-tuna-fisheries;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2005/1034/mega.pdf.

Narimatsu:2005:RBF

[4741] Yoji Narimatsu, Daiji Kitagawa, Tsutomu Hattori, and Hirobumi Onodera. Reproductive biology of female Rikuzen sole (Dexistes rikuzenius). Fishery Bulletin, 103(4):635–647, 2005. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/reproductive-biology-
female-rikuzen-sole-dexistes-rikuzenius;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2005/1034/nari.pdf.

Porter:2005:TSD

[4742] Steven M. Porter. Temporal and spatial distribution and abundance of flathead sole (Hippoglossoides elassodon) eggs and larvae in the western Gulf of Alaska. Fishery Bulletin, 103(4):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/temporal-and-spatial-
distribution-and-abundance-flathead-sole-hippoglossoides-elassodon;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2005/1034/port.pdf.

Prince:2005:MSW

[4743] Eric D. Prince, Robert K. Cowen, Eric S. Orbesen, Stacy A. Luthy, Joel K. Llopiz, David E. Richardson, and Joseph E. Serafy. Movements and spawning of white marlin (Tetrapturus albidus) and blue marlin (Makaira nigricans) off Punta Cana, Dominican Republic. Fishery Bulletin, 103(4):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
movements-and-spawning-white-marlin-tetrapturus-albidus-and-
blue-marlin-makaira-nigricans;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2005/1034/prin.pdf.

Stanley:2005:LHC

[4744] Richard D. Stanley and Allen R. Kronlund. Life history characteristics for silvergray rockfish (Sebastes brevispinis) in British Columbia waters and the implications for stock assessmentand management. Fishery Bulletin, 103(4):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/life-
history-characteristics-silvergray-rockfish-sebastes-brevispinis-
british-columbia;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2005/1034/stan.pdf.

Weise:2005:ICS

[4745] Michael J. Weise and James T. Harvey. Impact of the California sea lion (Zalophus californianus) on salmon fisheries in Monterey Bay, California. Fishery Bulletin, 103(4):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/impact-california-sea-lion-zalophus-californianus-salmon-
fisheries-monterey-bay-california;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2005/1034/weis.pdf.

Welsford:2005:EGC

[4746] Dirk C. Welsford and Jeremy M. Lyle. Estimates of growth and comparisons of growth rates determined from length- and age-based models for populations of purple wrasse (Notolabrus fucicola). Fishery Bulletin, 103(4):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
estimates-growth-and-comparisonsof-growth-rates-determined-
length-and-age-based-models;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2005/1034/wels.pdf.

Bishop:2005:EHM

[4747] Melanie J. Bishop, Charles H. Peterson, Henry C. Summerson, and David Gaskill. Effects of harvesting methods on sustainability of a bay scallop fushery: dredging uproots seagrass and displaces recruits. Fishery Bulletin, 103(4):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/effects-harvesting-methods-sustainability-
bay-scallop-fushery-dredging-uproots-seagrass-and;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2005/1034/
bish.pdf.

Diaz:2005:LCB

[4748] Guillermo A. Diaz and Joseph E. Serafy. Longline-caught blue shark (Prionace glauca): factors affecting the numbers available for live release.

Fishery Bulletin, 103(4):??, 2005. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/longline-caught-blue-shark-prionace-glauca-factors-
affecting-numbers-available-live-release;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2005/1034/diaz.pdf.

Fey:2005:LCL

[4749] Dariusz P. Fey and Jonathan A. Hare. Length correction for larval and early-juvenile Atlantic menhaden (Brevoortia tyrannus) after preservation in alcohol. Fishery Bulletin, 103(4):??, 2005. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/length-correction-larval-
and-early-juvenile-atlantic-menhaden-brevoortia-tyrannus-after;
 https://spo.nmfs.noaa.gov/sites/default/files/fey.pdf.

Hare:2005:CAL

[4750] Jonathan A. Hare and John J. Govoni. Comparison of average larval fish vertical distributions among species exhibiting different transport pathways on the southeast United States continental shelf. Fishery Bulletin, 103(4):??, 2005. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/comparison-average-larval-fish-
vertical-distributions-among-species-exhibiting-different;
https://spo.nmfs.noaa.gov/sites/default/files/hare.pdf.

Rochet:2006:PAF

[4751] Marie-Joëlle Rochet, Jean-François Cadiou, and Verena M. Trenkel. Precision and accuracy of fish length measurements obtained with two visual underwater methods. Fishery Bulletin, 104(1):1–9, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/precision-and-
accuracy-fish-length-measurements-obtained-two-visual-underwater-
methods;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2006/1041/rochet.pdf.

Witteveen:2006:EPC

[4752] Briana H. Witteveen, Robert J. Foy, and Kate M. Wynne. The effect of predation (current and historical) by humpback whales (Megaptera novaeangliae) on fish abundance near Kodiak Island, Alaska. Fishery Bulletin, 104(1):10–20, 2006. CODEN FSY-

BAY.  ISSN 0090-0656 (print), 1937-4518 (electronic).  URL

https://spo.nmfs.noaa.gov/content/effect-predation-current-
and-historical-humpback-whales-megaptera-novaeangliae-fish;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2006/1041/witteveen.pdf.

Weinberg:2006:VTG

[4753] Kenneth L. Weinberg and David A. Somerton. Variation in trawl geometry due to unequal warp length. Fishery Bulletin, 104(1):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/variation-trawl-geometry-
due-unequal-warp-length;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2006/1041/weinberg.pdf.

Kotwicki:2006:EAS

[4754] Stan Kotwicki, Kenneth L. Weinberg, and David A. Somerton. The effect of autotrawl systems on the performance of a survey trawl. Fishery Bulletin, 104(1):35–45, 2006. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
effect-autotrawl-systems-performance-survey-trawl;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2006/1041/
kotwicki.pdf.

Zeidberg:2006:FCM

[4755] Louis D. Zeidberg, William M. Hamner, Nikolay P. Nezlin, and Annette Henry. The fishery for California market squid (Loligo opalescens) (cephalopoda: Myopsida), from 1981 through 2003. Fishery Bulletin, 104(1):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
california-market-squid-loligo-opalescens-cephalopoda-myopsida-
1981-through-2003;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2006/1041/zeidberg.pdf.

Criales:2006:VSC

[4756] Maria M. Criales, John D. Wang, Joan A. Browder, Michael B. Robblee, Thomas L. Jackson, and Clinton Hittle. Variability in supply and cross-shelf transport of pink shrimp (Farfantepenaeus duorarum) postlarvae into western Florida Bay. Fishery Bulletin, 104(1):60–74, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/variability-supply-
and-cross-shelf-transport-pink-shrimp-farfantepenaeus-duorarum;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2006/1041/criales.pdf.

Grandcourt:2006:BAP

[4757] Edwin M. Grandcourt, Thabit Z. Al Abdessalaam, Ahmed T. Al Shamsi, and Franklin Francis. Biology and assessment of the painted sweetlips (Diagramma pictum (Thunberg, 1792)) and the spangled emperor (Lethrinus nebulosus (Forsskål, 1775)) in the southern Arabian Gulf. Fishery Bulletin, 104(1):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/biology-and-assessment-painted-sweetlips-diagramma-
pictum-thunberg-1792-and-spangled;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2006/1041/grandcourt.pdf.

Porch:2006:CFS

[4758] Clay E. Porch, Anne-Marie Eklund, and Gerald P. Scott. A catchfree stock assessment model with application to goliath grouper (Epinephelus itajara) off southern Florida. Fishery Bulletin, 104(1):89– 101, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/catch-free-
stock-assessment-model-application-goliath-grouper-epinephelus-
itajara-southern;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2006/1041/porch.pdf.

Tuckey:2006:FAF

[4759] Troy D. Tuckey and Mark Dehaven. Fish assemblages found in tidal-creek and seagrass habitats in the Suwannee River estuary. Fishery Bulletin, 104(1):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/fish-
assemblages-found-tidal-creek-and-seagrass-habitats-suwannee-
river-estuary;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2006/1041/tuckey.pdf.

Clarke:2006:PIE

[4760] Lora M. Clarke, Alistair D. M. Dove, and David O. Conover. Prevalence, intensity, and effect of a nematode (Philometra saltatrix) in the ovaries of bluefish (Pomatomus saltatrix). Fishery Bulletin, 104(1):118–124, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/prevalence-
intensity-and-effect-nematode-philometra-saltatrix-ovaries-
bluefish-pomatomus;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2006/1041/clarke.pdf.

Edwards:2006:DUS

[4761] Elizabeth F. Edwards. Duration of unassisted swimming activity for spotted dolphin (Stenella attenuata) calves: implications for mother– calf separation during tuna purse-seine sets. Fishery Bulletin, 104(1):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/duration-unassisted-
swimming-activity-spotted-dolphin-stenella-attenuata-calves;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2006/1041/edwards.pdf.

Saillant:2006:PSV

[4762] Eric Saillant and John R. Gold. Population structure and variance effective size of red snapper (Lutjanus campechanus) in the northern Gulf of Mexico. Fishery Bulletin, 104(1):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/population-structure-and-
variance-effective-size-red-snapper-lutjanus-campechanus-northern;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2006/1041/saillant.pdf.

Friedland:2006:RBS

[4763] Kevin D. Friedland, Lora M. Clarke, Jean-Denis Dutil, and Matti Salminen. The relationship between smolt and postsmolt growth for Atlantic salmon (Salmo salar) in the Gulf of St. Lawrence. Fishery Bulletin, 104 (1):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/relationship-
between-smolt-and-postsmolt-growth-atlantic-salmon-salmo-salar-
gulf-st;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2006/1041/friedland.pdf.

Garcia-Diaz:2006:SSM

[4764] Mercedes García-Díaz, José A. González, María J. Lorente, and Victor M. Tuset. Spawning season, maturity sizes, and fecundity in blacktail comber (Serranus atricauda) (Serranidae) from the easterncentral Atlantic. Fishery Bulletin, 104(2):??, 2006. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/spawning-season-maturity-sizes-
and-fecundity-blacktail-comber-serranus-atricauda-serranidae;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2006/1042/garcia.pdf.

Tissot:2006:BIF

[4765] Brian N. Tissot, Mary M. Yoklavich, Milton S. Love, Keri York, and Mark Amend. Benthic invertebrates that form habitat on deep banks off southern California, with special reference to deep sea coral. Fishery Bulletin, 104(2):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
benthic-invertebrates-form-habitat-deep-banks-southern-california-
special-reference-deep;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2006/1042/tissot.pdf.

Li:2006:KSR

[4766] Zhuozhuo Li, Andrew K. Gray, Milton S. Love, Akira Goto, Takashi Asahida, and Anthony J. Gharrett. A key to selected rockfishes (Sebastes spp.) based on mitochondrial DNA restriction fragment analysis. Fishery Bulletin, 104(2):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/key-selected-rockfishes-sebastes-spp-
based-mitochondrial-dna-restriction-fragment-analysis;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2006/1042/
li.pdf.

Torres-Orozco:2006:VYT

[4767] Ernesto Torres-Orozco, Arturo Muhlia-Melo, Armando Trasviña, and Sofía Ortega-García. Variation in yellowfin tuna (Thunnus albacares) catches related to El Niño-Southern Oscillation events at the entrance to the Gulf of California. Fishery Bulletin, 104(2):??, 2006. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/variation-yellowfin-tuna-thunnus-
albacares-catches-related-el-nino-southern-oscillation;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2006/1042/
torres.pdf.

Pepin:2006:EER

[4768] Pierre Pepin. Estimating the encounter rate of Atlantic capelin (Mallotus villosus) with fish eggs, based on stomach content analysis. Fishery Bulletin, 104(2):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
estimating-encounter-rate-atlantic-capelin-mallotus-villosus-
fish-eggs-based-stomach;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2006/1042/pepin.pdf.

Davis:2006:PMA

[4769] Michelle L. Davis, Jim Berkson, and Marcella Kelly. A production modeling approach to the assessment of the horseshoe crab (Limulus polyphemus) population in Delaware Bay. Fishery Bulletin, 104(2):215–225, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/production-modeling-
approach-assessment-horseshoe-crab-limulus-polyphemus-population;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2006/1042/davis.pdf.

Hoff:2006:BIR

[4770] Gerald R. Hoff. Biodiversity as an index of regime shift in the eastern Bering Sea. Fishery Bulletin, 104(2):??, 2006. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/biodiversity-index-regime-shift-
eastern-bering-sea;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2006/1042/hoff.pdf.

Pietsch:2006:TDN

[4771] Theodore W. Pietsch and James W. Orr. Triglops dorothy, a new species of sculpin (Teleostei: Scorpaeniformes: Cottidae) from the southern Sea of Okhotsk. Fishery Bulletin, 104(2):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/triglops-dorothy-new-species-sculpin-
teleostei-scorpaeniformes-cottidae-southern-sea;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/2006/1042/
pietsch.pdf.

Chen:2006:CTF

[4772] Yong Chen, Sally Sherman, Carl Wilson, John Sowles, and Minoru Kanaiwa. A comparison of two fishery-independent survey programs used to define the population structure of American lobster (Homarus americanus) in the Gulf of Maine. Fishery Bulletin, 104(2):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/comparison-two-fishery-independent-survey-programs-
used-define-population-structure;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2006/1042/chen.pdf.

Walsh:2006:JFA

[4773] Harvey J. Walsh, Katrin E. Marancik, and Jonathan A. Hare. Juvenile fish assemblages collected on unconsolidated sediments of the southeast United States continental shelf. Fishery Bulletin, 104(2):256–277, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/juvenile-fish-
assemblages-collected-unconsolidated-sediments-southeast-united-
states;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2006/1042/walsh.pdf.

Goldman:2006:GMS

[4774] Kenneth J. Goldman and John A. Musick. Growth and maturity of salmon sharks (Lamna ditropis) in the eastern and western North Pacific, and comments on back-calculation methods. Fishery Bulletin, 104(2):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/growth-and-maturity-
salmon-sharks-lamna-ditropis-eastern-and-western-north-pacific-
and;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2006/1042/goldman.pdf.

Hurton:2006:PCM

[4775] Lenka Hurton and Jim Berkson. Potential causes of mortality for horseshoe crabs (Limulus polyphemus) during the biomedical bleeding process. Fishery Bulletin, 104(2):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/potential-causes-mortality-horseshoe-crabs-limulus-
polyphemus-during-biomedical-bleeding;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2006/1042/hurton.pdf.

Kirshenbaum:2006:STM

[4776] Sheril Kirshenbaum and Yong Chen Scott Feindel. A study of tagging methods for the sea cucumber Cucumaria frondosa in the waters off Maine. Fishery Bulletin, 104(2):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/study-tagging-methods-sea-cucumber-cucumaria-frondosa-
waters-maine;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2006/1042/kirshenbaum.pdf.

Kimura:2006:PPE

[4777] Daniel K. Kimura and Martin W. Dorn. Parameterizing probabilities for estimating age-composition distributions for mixture models. Fishery Bulletin, 104(2):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
parameterizing-probabilities-estimating-age-composition-distributions-
mixture-models;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2006/1042/kimura.pdf.

May-Ku:2006:MDS

[4778] Marco A. May-Kú, Uriel Ordóñez-López, and Omar Defeo. Morphometric differentiation in small juveniles of the pink spotted shrimp (Farfantepenaeus brasiliensis) and the southern pink shrimp (F. notialis) in the Yucatan Peninsula, Mexico. Fishery Bulletin, 104(2):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/morphometric-
differentiation-small-juveniles-pink-spotted-shrimp-farfantepenaeus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2006/1042/may.pdf.

Cope:2006:EIL

[4779] Jason M. Cope. Exploring intraspecific life history patterns in sharks. Fishery Bulletin, 104(2):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/exploring-intraspecific-life-history-
patterns-sharks;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2006/1042/cope.pdf.

Dowd:2006:SRM

[4780] W. Wesley Dowd, Richard W. Brill, Peter G. Bushnell, and John A. Musick. Standard and routine metabolic rates of juvenile sandbar sharks (Carcharhinus plumbeus), including the effects of body mass and acute temperature change. Fishery Bulletin, 104(3):??, 2006. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/standard-and-routine-metabolic-
rates-juvenile-sandbar-sharks-carcharhinus-plumbeus;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2006/1043/
dowd.pdf.

Dowd:2006:ECR

[4781] W. Wesley Dowd, Richard W. Brill, Peter G. Bushnell, and John A. Musick. Estimating consumption rates of juvenile sandbar sharks (Carcharhinus plumbeus) in Chesapeake Bay, Virginia, using a bioenergetics model. Fishery Bulletin, 104(3):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/estimating-consumption-rates-juvenile-sandbar-sharks-
carcharhinus-plumbeus-chesapeake-bay;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2006/1043/dowd2.pdf.

Zollett:2006:DCB

[4782] Erika A. Zollett and Andrew J. Read. Depredation of catch by bottlenose dolphins (Tursiops truncatus) in the Florida king mackerel (Scomberomorus cavalla) troll fishery. Fishery Bulletin, 104(3):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/depredation-
catch-bottlenose-dolphins-tursiops-truncatus-florida-king-mackerel;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2006/1043/zollett.pdf.

Abookire:2006:RBS

[4783] Alisa A. Abookire. Reproductive biology, spawning season, and growth of female rex sole (Glyptocephalus zachirus) in the Gulf of Alaska. Fishery Bulletin, 104(3):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
reproductive-biology-spawning-season-and-growth-female-rex-
sole-glyptocephalus-zachirus;   https://spo.nmfs.noaa.gov/sites/
default/files/abookire.pdf.

Abesamis:2006:HMD

[4784] Rene A. Abesamis, Angel C. Alcala, and Garry R. Russ. How much does the fishery at Apo Island benefit from spillover of adult fish from the adjacent marine reserve? Fishery Bulletin, 104(3):360–375, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/how-much-does-fishery-apo-island-benefit-spillover-
adult-fish-adjacent-marine-reserve;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2006/1043/abesamis.pdf.

Li:2006:CIS

[4785] Zhuozhuo Li, Mary M. Nishimoto, Milton S. Love, and Anthony J. Gharrett. Comparing the identification of southern California juvenile rockfishes (genus Sebastes spp.) by restriction site analysis of the mitochondrial ND3/ND4 region and by morphological characteristics. Fishery Bulletin, 104(3):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/comparing-identification-
southern-california-juvenile-rockfishes-genus-sebastes-spp;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2006/1043/li.pdf.

Love:2006:PUO

[4786] Milton S. Love, Donna M. Schroeder, William Lenarz, Alec Mac-Call, Ann Scarborough Bull, and Lyman Thorsteinson. Potential use of offshore marine structures in rebuilding an overfished rockfish species, bocaccio (Sebastes paucispinis). Fishery Bulletin, 104(3):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/potential-use-
offshore-marine-structures-rebuilding-overfished-rockfish-species-

bocaccio;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2006/1043/love.pdf.

Emery:2006:DOG

[4787] Brian M. Emery, Libe Washburn, Milton S. Love, Mary M. Nishimoto, and J. Carter Ohlmann. Do oil and gas platforms off California reduce recruitment of bocaccio (Sebastes paucispinis) to natural habitat? An analysis based on trajectories derived from high-frequency radar. Fishery Bulletin, 104(3):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/do-oil-and-gas-platforms-
california-reduce-recruitment-bocaccio-sebastes-paucispinis;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2006/1043/emery.pdf.

Leis:2006:BOL

[4788] Jeffrey M. Leis, Amanda C. Hay, Domine L. Clark, I-Shiung Chen, and Kwang-Tsao Shao. Behavioral ontogeny in larvae and early juveniles of the giant trevally (Caranx ignobilis) (Pisces: Carangidae). Fishery Bulletin, 104(3):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/behavioral-ontogeny-larvae-and-early-juveniles-
giant-trevally-caranx-ignobilis-pisces;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2006/1043/leis.pdf.

Reddin:2006:DNT

[4789] David G. Reddin, Peter Downton, and Kevin D. Friedland. Diurnal and nocturnal temperatures for Atlantic salmon postsmolts (Salmo salar L.) during their early marine life. Fishery Bulletin, 104(3):415–428, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/diurnal-and-nocturnal-temperatures-atlantic-salmon-
postsmolts-salmo-salar-l-during-their;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2006/1043/reddin.pdf.

He:2006:PSS

[4790] Xi He, Marc Mangel, and Alec MacCall. A prior for steepness in stockrecruitment relationships, based on an evolutionary persistence principle. Fishery Bulletin, 104(3):429, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/prior-steepness-stock-recruitment-relationships-based-
evolutionary-persistence-principle;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/he.pdf.

Kerstetter:2006:SWM

[4791] David W. Kerstetter and John E. Graves. Survival of white marlin (Tetrapturus albidus) released from commercial pelagic longline gear in the western North Atlantic. Fishery Bulletin, 104(3): ??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/survival-
white-marlin-tetrapturus-albidus-released-commercial-pelagic-
longline-gear-western;   https://spo.nmfs.noaa.gov/sites/default/
files/kerstetter.pdf.

Gudmundson:2006:ATM

[4792] Carolyn J. Gudmundson, Tonya K. Zeppelin, and Rolf R. Ream. Application of two methods for determining diet of northern fur seals (Callorhinus ursinus). Fishery Bulletin, 104(3):445–455, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/application-two-methods-determining-diet-
northern-fur-seals-callorhinus-ursinus;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2006/1043/gudmundson.
pdf.

Smith:2006:AAH

[4793] David R. Smith, Michael J. Millard, and Sheila Eyler. Abundance of adult horseshoe crabs (Limulus polylphemus) in Delaware Bay estimated from a bay-wide mark-recapture study. Fishery Bulletin, 104(3): 456–464, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/abundance-
adult-horseshoe-crabs-limulus-polylphemus-delaware-bay-estimated-
bay-wide-mark;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2006/1043/smith.pdf.

Clark:2006:ECL

[4794] William G. Clark and Stephen M. Kaimmer. Estimates of commercial longline selectivity for Pacific halibut (Hippoglossus stenolepis) from multiple marking experiments. Fishery Bulletin, 104(3):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/estimates-commercial-
longline-selectivity-pacific-halibut-hippoglossus-stenolepis-
multiple;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2006/1043/clark.pdf.

Hattori:2006:HDN

[4795] Tsutomu Hattori, Akira Nishimura, Yoji Narimatsu, and Daiji Kitagawa. Hatching date, nursery grounds, and early growth of juvenile walleye pollock (Theragra chalcogramma) off northern Japan. Fishery Bulletin, 104(3):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/hatching-date-nursery-grounds-and-early-growth-
juvenile-walleye-pollock-theragra;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2006/1043/hattori.pdf.

DeForest:2006:DLE

[4796] Lisa G. De Forest and Morgan S. Busby. Development of larval and early juvenile penpoint gunnel (Apodichthys flavidus) (family: Pholidae). Fishery Bulletin, 104(3):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/development-larval-and-early-juvenile-penpoint-
gunnel-apodichthys-flavidus-family-pholidae;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/2006/1043/deforest.
pdf.

Gandini:2006:STP

[4797] Patricia Gandini and Esteban Frere. Spatial and temporal patterns in the bycatch of seabirds in the Argentinian longline fishery. Fishery Bulletin, 104(3):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
spatial-and-temporal-patterns-bycatch-seabirds-argentinian-
longline-fishery;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2006/1043/gandini.pdf.

Barrientos:2006:GDF

[4798] Claudio A. Barrientos, M. Teresa González, and Carlos A. Moreno. Geographical differences in the feeding patterns of red rockfish (Sebastes capensis) along South American coasts. Fishery Bulletin, 104(4):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/geographical-differences-
feeding-patterns-red-rockfish-sebastes-capensis-along-south;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2006/1044/barrientos.pdf.

Bergenius:2006:UOM

[4799] Mikaela A. J. Bergenius, Gavin A. Begg, and Bruce D. Mapstone. The use of otolith morphology to indicate the stock structure of common coral trout (Plectropomus leopardus) on the Great Barrier Reef, Australia. Fishery Bulletin, 104(4):498–511, 2006. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/use-otolith-morphology-indicate-

stock-structure-common-coral-trout-plectropomus-leopardus;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/2006/
1044/bergenius.pdf.

Mant:2006:BCM

[4800] Jason C. Mant, Michael J. Moran, Stephen J. Newman, S. Alex Hesp, Norman G. Hall, and Ian C. Potter. Biological characteristics and mortality of western butterfish (Pentapodus vitta), an abundant bycatch species of prawn trawling and recreational fishing in a large subtropical embayment. Fishery Bulletin, 104(4):??, 2006. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/biological-characteristics-and-
mortality-western-butterfish-pentapodus-vitta-abundant;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2006/1044/
mant.pdf.

Butler:2006:UMS

[4801] John Butler, Melissa Neuman, Deanna Pinkard, Rikk Kvitek, and Guy Cochrane. The use of multibeam sonar mapping techniques to refine population estimates of the endangered white abalone (Haliotis sorenseni). Fishery Bulletin, 104(4):521–532, 2006. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/use-multibeam-sonar-
mapping-techniques-refine-population-estimates-endangered-white-
abalone;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2006/1044/butler.pdf.

Scandol:2006:EGC

[4802] James P. Scandol, Tony J. Underwood, and Matt K. Broadhurst. Experiments in gear configuration to reduce bycatch in an estuarine squidtrawl fishery. Fishery Bulletin, 104(4):533–541, 2006. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/experiments-gear-configuration-
reduce-bycatch-estuarine-squid-trawl-fishery;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/2006/1044/scandol.
pdf.

Love:2006:RBF

[4803] Milton S. Love and Anne York. The relationships between fish assemblages and the amount of bottom horizontal beam exposed at California oil platforms: fish habitat preferences at man-made platforms and (by inference) at natural reefs. Fishery Bulletin, 104(4):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/relationships-
between-fish-assemblages-and-amount-bottom-horizontal-beam-
exposed-california;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2006/1044/love.pdf.

Corander:2006:BMI

[4804] Jukka Corander, Pekka Marttinen, and Samu Mäntyniemi. A Bayesian method for identification of stock mixtures from molecular marker data. Fishery Bulletin, 104(4):550–558, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/bayesian-method-identification-stock-
mixtures-molecular-marker-data;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2006/1044/corander.pdf.

Shepherd:2006:MPB

[4805] Gary R. Shepherd, Joshua Moser, David Deuel, and Pam Carlsen. The migration patterns of bluefish (Pomatomus saltatrix) along the Atlantic coast determined from tag recoveries. Fishery Bulletin, 104 (4):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/migration-
patterns-bluefish-pomatomus-saltatrix-along-atlantic-coast-
determined-tag;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2006/1044/shepherd.pdf.

Seitz:2006:ELB

[4806] Andrew C. Seitz, Brenda L. Norcross, Derek Wilson, and Jennifer L. Nielsen. Evaluating light-based geolocation for estimating demersal fish movements in high latitudes. Fishery Bulletin, 104(4):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/evaluating-
light-based-geolocation-estimating-demersal-fish-movements-
high-latitudes;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2006/1044/seitz.pdf.

Burks:2006:ADW

[4807] Carolyn M. Burks, William B. Driggers III, and Keith D. Mullin. Abundance and distribution of whale sharks (Rhincodon typus) in the northern Gulf of Mexico. Fishery Bulletin, 104(4):579–584, 2006. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/abundance-and-distribution-whale-
sharks-rhincodon-typus-northern-gulf-mexico;   https://spo.

nmfs.noaa.gov/sites/default/files/pdf-content/2006/1044/burks.
pdf.

Davis:2006:ESF

[4808] Michelle L. Davis and Jim Berkson. Effects of a simulated fishing moratorium on the stock assessment of red porgy (Pagrus pagrus). Fishery Bulletin, 104(4):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/effects-simulated-fishing-moratorium-
stock-assessment-red-porgy-pagrus-pagrus;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2006/1044/davis.pdf.

Taylor:2006:HUI

[4809] David L. Taylor, Peter M. Rowe, and Kenneth W. Able. Habitat use of the inner continental shelf off southern New Jersey by summerspawned bluefish (Pomatomus saltatrix). Fishery Bulletin, 104(4):593– 604, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/habitat-use-
inner-continental-shelf-southern-new-jersey-summer-spawned-
bluefish-pomatomus;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2006/1044/taylor.pdf.

Staudinger:2006:SSB

[4810] Michelle D. Staudinger. Seasonal and size-based predation on two species of squid by four fish predators on the Northwest Atlantic continental shelf. Fishery Bulletin, 104(4):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/seasonal-and-size-based-predation-two-species-squid-
four-fish-predators-northwest-atlantic;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2006/1044/staudinger.
pdf.

Rooper:2006:ERP

[4811] Christopher N. Rooper, Donald R. Gunderson, and David A. Armstrong. Evidence for resource partitioning and competition in nursery estuaries by juvenile flatfish in Oregon and Washington. Fishery Bulletin, 104(4):616–622, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/evidence-resource-partitioning-and-competition-
nursery-estuaries-juvenile-flatfish-oregon;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2006/1044/rooper.pdf.

LaPlante:2006:LSB

[4812] Lori H. LaPlante. Length-specific brood size and winter parturition in pink seaperch (Zalembius rosaceus) (Perciformes: Embiotocidae). Fishery Bulletin, 104(4):??, 2006. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/length-specific-brood-size-and-
winter-parturition-pink-seaperch-zalembius-rosaceus;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2006/1044/
laplante.pdf.

Riemer:2006:AFO

[4813] Susan D. Riemer and Robert Mikus. Aging fish otoliths recovered from Pacific harbor seal (Phoca vitulina) fecal samples. Fishery Bulletin, 104(4):??, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/aging-
fish-otoliths-recovered-pacific-harbor-seal-phoca-vitulina-
fecal-samples;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2006/1044/riemer.pdf.

Thedinga:2006:HAD

[4814] John F. Thedinga, Scott W. Johnson, and Donald G. Mortensen. Habitat, age, and diet of a forage fish in southeastern Alaska: Pacific sandfish (Trichodon trichodon). Fishery Bulletin, 104(4):631–637, 2006. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/habitat-age-and-diet-
forage-fish-southeastern-alaska-pacific-sandfish-trichodon-
trichodon;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2006/1044/thedinga.pdf.

Upadhyay:2006:GDY

[4815] Satyendra K. Upadhyay, Wang Jun, Su Yong-Quan, Ding Shao-Xiong, and Sonal Chaturvedi. Genetic diversity of yellow grouper (Epinephelus awoara) determined by random amplified polymorphic DNA (RAPD) analysis. Fishery Bulletin, 104(4):??, 2006. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/genetic-diversity-yellow-grouper-
epinephelus-awoara-determined-random-amplified-polymorphic;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2006/1044/upadhyay.pdf.

Gregg:2006:IPO

[4816] Jacob L. Gregg, Delsa M. Anderl, and Daniel K. Kimura. Improving the precision of otolith-based age estimates for Greenland halibut (Reinhardtius hippoglossoides) with preparation methods adapted for fragile sagittae. Fishery Bulletin, 104(4):643–648, 2006. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/improving-precision-
otolith-based-age-estimates-greenland-halibut-reinhardtius;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2006/1044/gregg.pdf.

Wexler:2007:TVG

[4817] Jeanne B. Wexler, Seinen Chow, Toshie Wakabayashi, Kenji Nohara, and Daniel Margulies. Temporal variation in growth of yellowfin tuna (Thunnus albacares) larvae in the Panama Bight, 1990–97. Fishery Bulletin, 105(1):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
temporal-variation-growth-yellowfin-tuna-thunnus-albacares-
larvae-panama-bight-1990-97;   https://spo.nmfs.noaa.gov/sites/
default/files/wexler.pdf.

McDermott:2007:AFB

[4818] Susanne F. McDermott, Katherine P. Maslenikov, and Donald R. Gunderson. Annual fecundity, batch fecundity, and oocyte atresia of Atka mackerel (Pleurogrammus monopterygius) in Alaskan waters. Fishery Bulletin, 105(1):19–29, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/annual-fecundity-batch-fecundity-and-
oocyte-atresia-atka-mackerel-pleurogrammus;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2007/1051/mcdermott.
pdf.

Hobbs:2007:MBI

[4819] James A. Hobbs, William A. Bennett, Jessica E. Burton, and Bradd Baskerville-Bridges. Modification of the biological intercept model to account for ontogenetic effects in laboratory-reared delta smelt (Hypomesus transpacificus). Fishery Bulletin, 105(1):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/modification-
biological-intercept-model-account-ontogenetic-effects-laboratory-
reared-delta;   https://spo.nmfs.noaa.gov/sites/default/files/
hobbs.pdf.

Laidig:2007:RBA

[4820] Thomas E. Laidig, James R. Chess, and Daniel F. Howard. Relationship between abundance of juvenile rockfishes (Sebastes spp.) and environmental variables documented off northern California and potential mechanisms for the covariation. Fishery Bulletin, 105(1):39–48, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/relationship-between-abundance-
juvenile-rockfishes-sebastes-spp-and-environmental-variables;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2007/1051/laidig.pdf.

Harley:2007:PUT

[4821] Shelton J. Harley and Jenny M. Suter. The potential use of time-area closures to reduce catches of bigeye tuna (Thunnus obesus) in the purseseine fishery of the eastern Pacific Ocean. Fishery Bulletin, 105(1):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/potential-use-time-
area-closures-reduce-catches-bigeye-tuna-thunnus-obesus-purse-
seine;   https://spo.nmfs.noaa.gov/sites/default/files/harley.
pdf.

Secor:2007:OMR

[4822] David H. Secor and Philip M. Piccoli. Oceanic migration rates of upper Chesapeake Bay striped bass (Morone saxatilis), determined by otolith microchemical analysis. Fishery Bulletin, 105(1): ??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/oceanic-
migration-rates-upper-chesapeake-bay-striped-bass-morone-saxatilis-
determined;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2007/1051/secor.pdf.

Matkin:2007:EVP

[4823] Craig O. Matkin, Lance G. Barrett-Lennard, Harald Yurk, David Ellifrit, and Andrew W. Trites. Ecotypic variation and predatory behavior among killer whales (Orcinus orca) off the eastern Aleutian Islands, Alaska. Fishery Bulletin, 105(1):??, 2007. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/ecotypic-variation-and-predatory-
behavior-among-killer-whales-orcinus-orca-eastern-aleutian;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2007/1051/matkin.pdf.

Stockhausen:2007:RON

[4824] William T. Stockhausen and Michael Fogarty. Removing observational noise from fisheries-independent time series data using ARIMA models. Fishery Bulletin, 105(1):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/removing-observational-noise-fisheries-independent-
time-series-data-using-arima-models;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2007/1051/stockhausen.pdf.

Pitcher:2007:ADE

[4825] Kenneth W. Pitcher, Peter F. Olesiuk, Robin F. Brown, Mark S. Lowry, Steven J. Jeffries, John L. Sease, Wayne L. Perryman, Charles E. Stinchcomb, and Lloyd F. Lowry. Abundance and distribution of the eastern North Pacific Steller sea lion (Eumetopias jubatus) population. Fishery Bulletin, 105(1):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
abundance-and-distribution-eastern-north-pacific-steller-sea-
lion-eumetopias-jubatus;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2007/1051/pitcher.pdf.

Gerritsen:2007:PES

[4826] Hans D. Gerritsen and David McGrath. Precision estimates and suggested sample sizes for length-frequency data. Fishery Bulletin, 105 (1):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/precision-
estimates-and-suggested-sample-sizes-length-frequency-data;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2007/1051/gerritsen.pdf.

Farley:2007:EMG

[4827] Edward V. Farley, Jr., James M. Murphy, Milo D. Adkison, Lisa B. Eisner, John H. Helle, Jamal H. Moss, and Jennifer Nielsen. Early marine growth in relation to marine-stage survival rates for Alaska sockeye salmon (Oncorhynchus nerka). Fishery Bulletin, 105(1):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/early-marine-growth-
relation-marine-stage-survival-rates-alaska-sockeye-salmon-
oncorhynchus;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2007/1051/farley.pdf.

Ganias:2007:DPF

[4828] Konstantinos Ganias, Cristina Nunes, and Yorgos Stratoudakis. Degeneration of postovulatory follicles in the Iberian sardine Sardina pilchardus: structural changes and factors affecting resorption. Fishery Bulletin, 105(1):131–139, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/degeneration-postovulatory-follicles-

iberian-sardine-sardina-pilchardus-structural-changes;   https:
//spo.nmfs.noaa.gov/sites/default/files/ganias.pdf.

Arrizabalaga:2007:UGD

[4829] Haritz Arrizabalaga, Victoria López-Rodas, Eduardo Costas, and Alberto González-Garcés. Use of genetic data to assess the uncertainty in stock assessments due to the assumed stock structure: the case of albacore (Thunnus alalunga) from the Atlantic Ocean. Fishery Bulletin, 105(1):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/use-genetic-data-assess-uncertainty-stock-
assessments-due-assumed-stock-structure-case;   https://spo.
nmfs.noaa.gov/sites/default/files/arrizabalaga.pdf.

Cherel:2007:NIF

[4830] Yves Cherel, Richard Sabatié, Michel Potier, Francis Marsac, and Frédéric Ménard. New information from fish diets on the importance of glassy flying squid (Hyaloteuthis pelagica) (Teuthoidea: Ommastrephidae) in the epipelagic cephalopod community of the tropical Atlantic Ocean. Fishery Bulletin, 105(1):147–152, 2007. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/new-information-fish-
diets-importance-glassy-flying-squid-hyaloteuthis-pelagica-
teuthoidea;   https://spo.nmfs.noaa.gov/sites/default/files/
cherel.pdf.

Conti:2007:MTS

[4831] Stéphane G. Conti, Benjamin D. Maurer, Mark A. Drawbridge, and David A. Demer. Measurements of total scattering spectra from bocaccio (Sebastes paucispinis). Fishery Bulletin, 105(1):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/measurements-total-
scattering-spectra-bocaccio-sebastes-paucispinis;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/2007/1051/
conti.pdf.

Gentner:2007:SAB

[4832] Brad Gentner. Sensitivity of angler benefit estimates from a model of recreational demand to the definition of the substitute sites considered by the angler. Fishery Bulletin, 105(2):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/sensitivity-angler-benefit-estimates-
model-recreational-demand-definition-substitute-sites;   https:

//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2007/1052/
gentner.pdf.

Anderson:2007:MHA

[4833] Tara J. Anderson and Mary M. Yoklavich. Multiscale habitat associations of deepwater demersal fishes off central California. Fishery Bulletin, 105(2):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/multiscale-habitat-associations-deepwater-demersal-
fishes-central-california;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2007/1052/anderson.pdf.

McClelland:2007:UEH

[4834] Gary McClelland and Jason Melendy. Use of endoparasitic helminths as tags in delineating stocks of American plaice (Hippoglossoides platessoides) from the southern Gulf of St. Lawrence and Cape Breton Shelf. Fishery Bulletin, 105(2):??, 2007. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/use-endoparasitic-helminths-tags-
delineating-stocks-american-plaice-hippoglossoides;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2007/1052/
mcclelland.pdf.

Hamilton:2007:ISC

[4835] Judy Hamilton and Brenda Konar. Implications of substrate complexity and kelp variability for south-central Alaskan nearshore fish communities. Fishery Bulletin, 105(2):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/implications-substrate-complexity-and-kelp-variability-
south-central-alaskan-nearshore-fish;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2007/1052/hamilton.pdf.

Ward:2007:GVS

[4836] Rocky Ward, Kevin Bowers, Rebecca Hensley, Brandon Mobley, and Ed Belouski. Genetic variability in spotted seatrout (Cynoscion nebulosus), determined with microsatellite DNA markers. Fishery Bulletin, 105(2):197–206, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/genetic-variability-spotted-seatrout-cynoscion-
nebulosus-determined-microsatellite-dna;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2007/1052/ward.pdf.

Smith-Vaniz:2007:RCJ

[4837] William F. Smith-Vaniz and Kent E. Carpenter. Review of the crevalle jacks, Caranx hippos complex (Teleostei: Carangidae), with a description of a new species from West Africa. Fishery Bulletin, 105 (2):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/review-
crevalle-jacks-caranx-hippos-complex-teleostei-carangidae-description-
new-species;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2007/1052/smith.pdf.

Trites:2007:DSS

[4838] Andrew W. Trites, Donald G. Calkins, and Arliss J. Winship. Diets of Steller sea lions (Eumetopias jubatus) in Southeast Alaska, 1993–1999. Fishery Bulletin, 105(2):2334–248, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/diets-steller-sea-lions-eumetopias-jubatus-southeast-
alaska-1993-1999;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2007/1052/trites.pdf.

Margulies:2007:SED

[4839] Daniel Margulies, Jenny M. Suter, Sharon L. Hunt, Robert J. Olson, Vernon P. Scholey, Jeanne B. Wexler, and Akio Nakazawa. Spawning and early development of captive yellowfin tuna (Thunnus albacares). Fishery Bulletin, 105(2):??, 2007. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/spawning-and-early-development-
captive-yellowfin-tuna-thunnus-albacares;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2007/1052/margulies.
pdf.

Zeller:2007:RES

[4840] Dirk Zeller, Shawn Booth, Gerald Davis, and Daniel Pauly. Reestimation of small-scale fishery catches for U.S. flag-associated island areas in the western Pacific: the last 50 years. Fishery Bulletin, 105(2):266–277, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/re-estimation-small-scale-fishery-catches-
us-flag-associated-island-areas-western-pacific;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/2007/1052/
zeller.pdf.

Somerton:2007:WGE

[4841] David A. Somerton, Peter T. Munro, and Kenneth L. Weinberg. Whole-gear efficiency of a benthic survey trawl for flatfish. Fishery Bulletin, 105(2):278–291, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/whole-gear-efficiency-benthic-survey-
trawl-flatfish;   https://spo.nmfs.noaa.gov/sites/default/files/
somerton.pdf.

Strange:2007:YPF

[4842] Rex M. Strange and Carol A. Stepien. Yellow (Perca flavescens) and Eurasian (P. fluviatilis) perch distinguished in fried fish samples by DNA analysis. Fishery Bulletin, 105(2):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/yellow-perca-flavescens-and-eurasian-
p-fluviatilis-perch-distinguished-fried-fish-samples;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2007/1052/
strange.pdf.

Cooper:2007:NMR

[4843] Daniel W. Cooper, Katherine P. Maslenikov, and Donald R. Gunderson. Natural mortality rate, annual fecundity, and maturity at length for Greenland halibut (Reinhardtius hippoglossoides) from the northeastern Pacific Ocean. Fishery Bulletin, 105(2):296–304, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/natural-
mortality-rate-annual-fecundity-and-maturity-length-greenland-
halibut-reinhardtius;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2007/1052/cooper.pdf.

Roemer:2007:VBC

[4844] Marja E. Roemer and Kenneth Oliveira. Validation of backcalculation equations for juvenile bluefish (Pomatomus saltatrix) with the use of tetracycline-marked otoliths. Fishery Bulletin, 105(2):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/validation-
back-calculation-equations-juvenile-bluefish-pomatomus-saltatrix-
use;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2007/1052/roemer.pdf.

Auth:2007:DVV

[4845] Toby D. Auth, Richard D. Brodeur, and Kathleen M. Fisher. Diel variation in vertical distribution of an offshore ichthyoplankton community off the Oregon coast. Fishery Bulletin, 105(3): 313–326, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/diel-
variation-vertical-distribution-offshore-ichthyoplankton-community-
oregon-coast;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2007/1053/auth.pdf.

Curtis:2007:VME

[4846] Janelle M. R. Curtis. Validation of a method for estimating realized annual fecundity in a multiple spawner, the long-snouted seahorse (Hippocampus guttulatus), using underwater visual census. Fishery Bulletin, 105(3):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
validation-method-estimating-realized-annual-fecundity-multiple-
spawner-long-snouted;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2007/1053/curtis.pdf.

Benson:2007:ADH

[4847] Scott R. Benson, Karin A. Forney, James T. Harvey, James V. Carretta, and Peter H. Dutton. Abundance, distribution, and habitat of leatherback turtles (Dermochelys coriacea) off California, 1990–2003. Fishery Bulletin, 105(3):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/abundance-distribution-and-habitat-leatherback-
turtles-dermochelys-coriacea-california-1990;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/2007/1053/benson.
pdf.

Bengtson:2007:CSM

[4848] John L. Bengtson, Alana V. Phillips, Elizabeth A. Mathews, and Michael A. Simpkins. Comparison of survey methods for estimating abundance of harbor seals (Phoca vitulina) in glacial fjords. Fishery Bulletin, 105(3):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/comparison-survey-methods-estimating-abundance-
harbor-seals-phoca-vitulina-glacial-fjords;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2007/1053/bengtson.
pdf.

DeMartini:2007:AGS

[4849] Edward E. DeMartini, James H. Uchiyama, Robert L. Humphreys, Jr., Jeffrey D. Sampaga, and Happy A. Williams. Age and growth of swordfish (Xiphias gladius) caught by the Hawaii-based pelagic longline fishery. Fishery Bulletin, 105(3):??, 2007. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/age-and-growth-swordfish-xiphias-
gladius-caught-hawaii-based-pelagic-longline-fishery;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2007/1053/
demartini.pdf.

Anderson:2007:SNA

[4850] Joel D. Anderson. Systematics of the North American menhadens: molecular evolutionary reconstructions in the genus Brevoortia (Clupeiformes: Clupeidae). Fishery Bulletin, 105(3):??, 2007. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/systematics-north-american-menhadens-
molecular-evolutionary-reconstructions-genus;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/2007/1053/anderson.
pdf.

Schaefer:2007:VMP

[4851] Kurt M. Schaefer and Daniel W. Fuller. Vertical movement patterns of skipjack tuna (Katsuwonus pelamis) in the eastern equatorial Pacific Ocean, as revealed with archival tags. Fishery Bulletin, 105(3): ??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/vertical-
movement-patterns-skipjack-tuna-katsuwonus-pelamis-eastern-
equatorial-pacific;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2007/1053/schaefer.pdf.

Golet:2007:DCN

[4852] Walter J. Golet, Andrew B. Cooper, Robert Campbell, and Molly Lutcavage. Decline in condition of northern bluefin tuna (Thunnus thynnus) in the Gulf of Maine. Fishery Bulletin, 105(3): ??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/decline-
condition-northern-bluefin-tuna-thunnus-thynnus-gulf-maine;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2007/1053/golet.pdf.

Stark:2007:GSV

[4853] James W. Stark. Geographic and seasonal variations in maturation and growth of female Pacific cod (Gadus macrocephalus) in the Gulf of Alaska and Bering Sea. Fishery Bulletin, 105(3):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/geographic-

and-seasonal-variations-maturation-and-growth-female-pacific-
cod-gadus;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2007/1053/stark.pdf.

Garrison:2007:IBM

[4854] Lance P. Garrison. Interactions between marine mammals and pelagic longline fishing gear in the U.S. Atlantic Ocean between 1992 and 2004. Fishery Bulletin, 105(3):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/interactions-between-marine-mammals-and-pelagic-longline-
fishing-gear-us-atlantic-ocean;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2007/1053/garrison.pdf.

Melville-Smith:2007:CEP

[4855] Roy Melville-Smith and Simon de Lestang. Changes in egg production of the western rock lobster (Panulirus cygnus) associated with appendage damage. Fishery Bulletin, 105(3):??, 2007. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/changes-egg-production-western-
rock-lobster-panulirus-cygnus-associated-appendage-damage;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/2007/
1053/smith.pdf.

Able:2007:DEM

[4856] Kenneth W. Able and Thomas M. Grothues. Diversity of estuarine movements of striped bass (Morone saxatilis): a synoptic examination of an estuarine system in southern New Jersey. Fishery Bulletin, 105 (3):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/diversity-
estuarine-movements-striped-bass-morone-saxatilis-synoptic-
examination-estuarine;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2007/1053/able.pdf.

Renshaw:2007:MMP

[4857] Mark A. Renshaw, Eric Saillant, Siya Lem, Philip Berry, and John R. Gold. Microsatellite multiplex panels for genetic studies of gray snapper (Lutjanus griseus) and lane snapper (Lutjanus synagris). Fishery Bulletin, 105(3):??, 2007. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/microsatellite-multiplex-panels-
genetic-studies-gray-snapper-lutjanus-griseus-and-lane;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2007/1053/
renshaw.pdf.

Kane:2007:ETS

[4858] Joseph Kane and Jacquelyn L. Anderson. Effect of towing speed on retention of zooplankton in bongo nets. Fishery Bulletin, 105(3):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/effect-towing-
speed-retention-zooplankton-bongo-nets;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2007/1053/kane.pdf.

Katakura:2007:ASF

[4859] Seiji Katakura, Hisatoshi Ikeda, Akira Nishimura, Tsuneo Nishiyama, and Yasunori Sakurai. An allometric smoothing function to describe the relation between otolith and somatic growth over the lifespan of walleye pollock (Theragra chalcogramma). Fishery Bulletin, 105(4):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/allometric-smoothing-
function-describe-relation-between-otolith-and-somatic-growth-
over;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2007/1054/katakura.pdf.

Able:2007:TMF

[4860] Kenneth W. Able, Peter J. Clarke, R. Christopher Chambers, and David A. Witting. Transitions in the morphological features, habitat use, and diet of young-of-the-year goosefish (Lophius americanus). Fishery Bulletin, 105(4):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/transitions-morphological-features-habitat-use-and-
diet-young-year-goosefish-lophius;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2007/1054/able.pdf.

Stewart:2007:DPM

[4861] Ian J. Stewart. Defining plausible migration rates based on historical tagging data: a Bayesian mark-recapture model applied to English sole (Parophrys vetulus). Fishery Bulletin, 105(4):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/defining-plausible-
migration-rates-based-historical-tagging-data-bayesian-mark-
recapture;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2007/1054/stewart.pdf.

Magel:2007:APN

[4862] Christopher Magel, Kirstin Wakefield, Nancy Targett, and Richard Brill. Activity in the pallial nerve of knobbed (Busycon carica) and channeled (Busycotypus canaliculatum) whelks recorded during exposure of the osphradium to odorant solutions. Fishery Bulletin, 105 (4):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/activity-
pallial-nerve-knobbed-busycon-carica-and-channeled-busycotypus-
canaliculatum;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2007/1054/magel.pdf.

Eveson:2007:IFO

[4863] J. Paige Eveson, Tom Polacheck, and Geoff M. Laslett. Incorporating fishery observer data into an integrated catch-at-age and multiyear tagging model for estimating mortality rates and abundance. Fishery Bulletin, 105(4):493–508, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/incorporating-fishery-observer-data-integrated-
catch-age-and-multiyear-tagging-model;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2007/1054/eveson.pdf.

Barlow:2007:APD

[4864] Jay Barlow and Karin A. Forney. Abundance and population density of cetaceans in the California Current ecosystem. Fishery Bulletin, 105(4): ??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/abundance-and-
population-density-cetaceans-california-current-ecosystem;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/2007/
1054/barlow.pdf.

Tupper:2007:SCV

[4865] Mark H. Tupper. Spillover of commercially valuable reef fishes from marine protected areas in Guam, Micronesia. Fishery Bulletin, 105 (4):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/spillover-
commercially-valuable-reef-fishes-marine-protected-areas-guam-
micronesia;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2007/1054/tupper.pdf.

Klibansky:2007:SSE

[4866] Nikolai Klibansky and Francis Juanes. Species-specific effects of four preservative treatments on oocytes and ovarian material of Atlantic cod (Gadus morhua), haddock (Melanogrammus aeglefinus), and American plaice (Hippoglossoides platessoides). Fishery Bulletin, 105(4):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/species-specific-effects-
four-preservative-treatments-oocytes-and-ovarian-material-atlantic;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2007/1054/klibansky.pdf.

Miller:2007:DTR

[4867] Todd W. Miller and Richard D. Brodeur. Diets of and trophic relationships among dominant marine nekton within the northern California Current ecosystem. Fishery Bulletin, 105(4):??, 2007. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/diets-and-trophic-relationships-
among-dominant-marine-nekton-within-northern-california;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2007/1054/
miller.pdf.

Lauth:2007:TDM

[4868] Robert R. Lauth, Jared Guthridge, Dan Nichol, Scott W. McEntire, and Nicola Hillgruber. Timing and duration of mating and brooding periods of Atka mackerel (Pleurogrammus monopterygius) in the North Pacific Ocean. Fishery Bulletin, 105(4):??, 2007. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/timing-and-duration-mating-and-
brooding-periods-atka-mackerel-pleurogrammus-monopterygius;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2007/1054/lauth.pdf.

Lauth:2007:DED

[4869] Robert R. Lauth and Deborah M. Blood. Description of embryonic development of Atka mackerel (Pleurogrammus monopterygius). Fishery Bulletin, 105(4):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
description-embryonic-development-atka-mackerel-pleurogrammus-
monopterygius;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2007/1054/lauth2.pdf.

Kimura:2007:UEB

[4870] Daniel K. Kimura. Using the empirical Bayes method to estimate and evaluate bycatch rates of seabirds from individual fishing vessels. Fishery Bulletin, 105(4):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/using-empirical-bayes-method-estimate-and-evaluate-
bycatch-rates-seabirds-individual;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2007/1054/kimura.pdf.

DeVries:2007:NEB

[4871] Douglas A. DeVries. No evidence of bias from fish behavior in the selectivity of size and sex of the protogynous red porgy (Pagrus pagrus, Sparidae) by hook-and-line gear. Fishery Bulletin, 105(4):??, 2007. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/no-evidence-bias-fish-behavior-
selectivity-size-and-sex-protogynous-red-porgy-pagrus-pagrus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2007/1054/devries.pdf.

Rooper:2008:EAR

[4872] Christopher N. Rooper. An ecological analysis of rockfish (Sebastes spp.) assemblages in the North Pacific Ocean along broad-scale environmental gradients. Fishery Bulletin, 106(1):??, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/ecological-analysis-rockfish-sebastes-spp-
assemblages-north-pacific-ocean-along-broad-scale;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2008/1061/
rooper.pdf.

Brooks:2008:SAP

[4873] Elizabeth N. Brooks, Kyle W. Shertzer, Todd Gedamke, and Douglas S. Vaughan. Stock assessment of protogynous fish: evaluating measures of spawning biomass used to estimate biological reference points. Fishery Bulletin, 106(1):??, 2008. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
stock-assessment-protogynous-fish-evaluating-measures-spawning-
biomass-used-estimate;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2008/1061/brooks.pdf.

Stevenson:2008:EPS

[4874] Duane E. Stevenson, James W. Orr, Gerald R. Hoff, and John D. McEachran. Emerging patterns of species richness, diversity, population density, and distribution in the skates (Rajidae) of Alaska. Fishery Bulletin, 106(1):??, 2008. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
emerging-patterns-species-richness-diversity-population-density-
and-distribution-skates;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2008/1061/stevenson.pdf.

Stiansen:2008:ECD

[4875] Stian Stiansen, Anders Fernö, Dag Furevik, Terje Jørgensen, and Svein Løkkeborg. Efficiency and catch dynamics of collapsible square and conical crab pots used in the red king crab (Paralithodes camtschaticus) fishery. Fishery Bulletin, 106(1):??, 2008. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/efficiency-and-catch-
dynamics-collapsible-square-and-conical-crab-pots-used-red-
king-crab;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2008/1061/stiansen.pdf.

Latour:2008:TDS

[4876] Robert J. Latour, James Gartland, Christopher F. Bonzek, and Rae-Marie A. Johnson. The trophic dynamics of summer flounder (Paralichthys dentatus) in Chesapeake Bay. Fishery Bulletin, 106(1): ??, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/trophic-
dynamics-summer-flounder-paralichthys-dentatus-chesapeake-bay;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2008/1061/latour.pdf.

Haddon:2008:UIL

[4877] Malcolm Haddon, Craig Mundy, and David Tarbath. Using an inverse-logistic model to describe growth increments of blacklip abalone (Haliotis rubra) in Tasmania. Fishery Bulletin, 106(1):58–71, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/using-inverse-
logistic-model-describe-growth-increments-blacklip-abalone-
haliotis-rubra;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2008/1061/haddon.pdf.

Byrd:2008:ECF

[4878] Barbie L. Byrd, Aleta A. Hohn, Fentress H. Munden, Gretchen N. Lovewell, and Rachel E. Lo Piccolo. Effects of commercial fishing regulations on stranding rates of bottlenose dolphin (Tursiops truncatus). Fishery Bulletin, 106(1):??, 2008. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
effects-commercial-fishing-regulations-stranding-rates-bottlenose-
dolphin-tursiops;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2008/1061/byrd.pdf.

Schwenke:2008:AGR

[4879] Kara L. Schwenke and Jeffrey A. Buckel. Age, growth, and reproduction of dolphinfish (Coryphaena hippurus) caught off the coast of North Carolina. Fishery Bulletin, 106(1):??, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/age-growth-and-reproduction-dolphinfish-coryphaena-
hippurus-caught-coast-north-carolina;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2008/1061/schwenke.pdf.

Vasslides:2008:ISS

[4880] James M. Vasslides and Kenneth W. Able. Importance of shoreface sand ridges as habitat for fishes off the northeast coast of the United States. Fishery Bulletin, 106(1):??, 2008. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/importance-shoreface-sand-ridges-
habitat-fishes-northeast-coast-united-states;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/2008/1061/vasslides.
pdf.

Orr:2008:SRR

[4881] James W. Orr and Sharon Hawkins. Species of the rougheye rockfish complex: resurrection of Sebastes melanostictus (Matsubara, 1934) and a redescription of Sebastes aleutianus (Jordan and Evermann, 1898) (Teleostei: Scorpaeniformes). Fishery Bulletin, 106(2): ??, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/species-
rougheye-rockfish-complex-resurrection-sebastes-melanostictus-
matsubara-1934-and;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2008/1062/orr.pdf.

Kingsford:2008:IEC

[4882] Michael J. Kingsford, Heather M. Patterson, and Matthew J. Flood. The influence of elemental chemistry on the widths of otolith increments in the neon damselfish (Pomacentrus coelestis). Fishery Bulletin, 106(2): 135–142, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/influence-
elemental-chemistry-widths-otolith-increments-neon-damselfish-
pomacentrus;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2008/1062/kingsford.pdf.

Aalbers:2008:SDL

[4883] Scott A. Aalbers. Seasonal, diel, and lunar spawning periodicities and associated sound production of white seabass (Atractoscion nobilis). Fishery Bulletin, 106(2):??, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/seasonal-diel-and-lunar-spawning-periodicities-and-
associated-sound-production-white;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2008/1062/aalbers.pdf.

Clardy:2008:STV

[4884] Todd R. Clardy, William F. Patterson III, Douglas A. DeVries, and Christopher Palmer. Spatial and temporal variability in the relative contribution of king mackerel (Scomberomorus cavalla) stocks to winter mixed fisheries off South Florida. Fishery Bulletin, 106(2):152–160, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/spatial-and-temporal-
variability-relative-contribution-king-mackerel-scomberomorus-
cavalla;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2008/1062/clardy.pdf.

Parnel:2008:ICC

[4885] Maria M. Parnel, Robert L. Emmett, and Richard D. Brodeur. Ichthyoplankton community in the Columbia River plume off Oregon: effects of fluctuating oceanographic conditions. Fishery Bulletin, 106(2):161–173, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/ichthyoplankton-community-columbia-river-plume-
oregon-effects-fluctuating-oceanographic;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2008/1062/parnel.pdf.

Overton:2008:IBA

[4886] Anthony S. Overton, Charles S. Manooch III, Joseph W. Smith, and Kenneth Brennan. Interactions between adult migratory striped bass (Morone saxatilis) and their prey during winter off the Virginia and North Carolina Atlantic coast from 1994 through 2007. Fishery Bulletin, 106(2): ??, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/interactions-
between-adult-migratory-striped-bass-morone-saxatilis-and-their-
prey-during;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2008/1062/overton.pdf.

Vandersea:2008:ILS

[4887] Mark W. Vandersea, R. Wayne Litaker, Katrin E. Marancik, Jonathan A. Hare, Harvey J. Walsh, Siya Lem, Melissa A. West, David M. Wyanski, Elisabeth H. Laban, and Patricia A. Tester. Identification of larval sea basses (Centropristis spp.) using ribosomal DNA-specific molecular assays. Fishery Bulletin, 106(2):183–193, 2008. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/identification-larval-sea-basses-
centropristis-spp-using-ribosomal-dna-specific-molecular;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/2008/
1062/vandersea.pdf.

Rhodes:2008:VRA

[4888] Kevin L. Rhodes and Mark. H. Tupper. The vulnerability of reproductively active squaretail coralgrouper (Plectropomus areolatus) to fishing. Fishery Bulletin, 106(2):??, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/vulnerability-reproductively-active-squaretail-
coralgrouper-plectropomus-areolatus-fishing;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/2008/1062/rhodes.
pdf.

Garcia-Rodriguez:2008:MDM

[4889] Francisco J. García-Rodríguez, Germán Ponce-Díaz, Isabel Muñoz-García, Rogelio González-Armas, and Ricardo Perez-Enriquez. Mitochondrial DNA markers to identify commercial spiny lobster species (Panulirus spp.) from the Pacific coast of Mexico: an application on phyllosoma larvae. Fishery Bulletin, 106(2):??, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/mitochondrial-dna-markers-identify-commercial-
spiny-lobster-species-panulirus-spp-pacific;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/2008/1062/garcia.
pdf.

Nolan:2008:AUC

[4890] Cormac J. Nolan and Bret S. Danilowicz. Advantages of using crest nets to sample presettlement larvae of reef fishes in the Caribbean sea. Fishery Bulletin, 106(2):213–221, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/advantages-using-crest-nets-sample-presettlement-larvae-
reef-fishes-caribbean-sea;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2008/1062/nolan.pdf.

Shertzer:2008:PBA

[4891] Kyle W. Shertzer, Michael H. Prager, and Erik H. Williams. A probability-based approach to setting annual catch levels. Fishery Bulletin, 106(3):225–232, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/probability-based-approach-setting-annual-
catch-levels;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2008/1063/shertzer.pdf.

Hoff:2008:NSA

[4892] Gerald R. Hoff. A nursery site of the Alaska skate (Bathyraja parmifera) in the eastern Bering Sea. Fishery Bulletin, 106(3):

??, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/nursery-
site-alaska-skate-bathyraja-parmifera-eastern-bering-sea;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/2008/
1063/hoff.pdf.

Beacham:2008:DPS

[4893] Terry D. Beacham, Nataly V. Varnavskaya, Khai D. Le, and Michael H. Wetklo. Determination of population structure and stock composition of chum salmon (Oncorhynchus keta) in Russia determined with microsatellites. Fishery Bulletin, 106(3):??, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/determination-population-
structure-and-stock-composition-chum-salmon-oncorhynchus-keta;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2008/1063/beacham.pdf.

Shertzer:2008:FAI

[4894] Kyle W. Shertzer and Erik H. Williams. Fish assemblages and indicator species: reef fishes off the southeastern United States. Fishery Bulletin, 106(3):??, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/fish-
assemblages-and-indicator-species-reef-fishes-southeastern-
united-states;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2008/1063/shertzer2.pdf.

Nichols:2008:STD

[4895] Owen C. Nichols, Robert D. Kenney, and Moira W. Brown. Spatial and temporal distribution of North Atlantic right whales (Eubalaena glacialis) in Cape Cod Bay, and implications for management. Fishery Bulletin, 106(3):270–280, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/spatial-and-temporal-distribution-north-atlantic-right-
whales-eubalaena-glacialis-cape-cod;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2008/1063/nichols.pdf.

Conners:2008:ECF

[4896] M. Elizabeth Conners and Peter Munro. Effects of commercial fishing on local abundance of Pacific cod (Gadus macrocephalus) in the Bering Sea. Fishery Bulletin, 106(3):??, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/effects-commercial-fishing-local-abundance-pacific-

cod-gadus-macrocephalus-bering-sea;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2008/1063/conners.pdf.

Zimmermann:2008:CEM

[4897] Mark Zimmermann and Christopher Rooper. Comparison of echogram measurements against data expectations and assumptions for distinguishing seafloor substrates. Fishery Bulletin, 106(3):??, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/comparison-echogram-
measurements-against-data-expectations-and-assumptions-distinguishing;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2008/1063/zimmermann.pdf.

Maloney:2008:ASM

[4898] Nancy E. Maloney and Michael F. Sigler. Age-specific movement patterns of sablefish (Anoplopoma fimbria) in Alaska. Fishery Bulletin, 106(3): ??, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/age-specific-
movement-patterns-sablefish-anoplopoma-fimbria-alaska;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2008/1063/
maloney.pdf.

Kilgour:2008:DRD

[4899] Morgan J. Kilgour and Thomas C. Shirley. Distribution of red deepsea crab (Chaceon quinquedens) by size and sex in the Gulf of Mexico. Fishery Bulletin, 106(3):??, 2008. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
distribution-red-deepsea-crab-chaceon-quinquedens-size-and-
sex-gulf-mexico;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2008/1063/kilgour.pdf.

Erzini:2008:CGF

[4900] Karim Erzini, Luís Bentes, Rui Coelho, Pedro G. Lino, Pedro Monteiro, Joaquim Ribeiro, and Jorge M. S. Conçalves. Catches in ghost-fishing octopus and fish traps in the northeastern Atlantic Ocean (Algarve, Portugal). Fishery Bulletin, 106(3):??, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/catches-ghost-
fishing-octopus-and-fish-traps-northeastern-atlantic-ocean-
algarve-portugal;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2008/1063/erzini.pdf.

Stark:2008:ALM

[4901] James W. Stark. Age- and length-at-maturity of female arrowtooth flounder (Atheresthes stomias) in the Gulf of Alaska. Fishery Bulletin, 106(3):??, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/age-
and-length-maturity-female-arrowtooth-flounder-atheresthes-
stomias-gulf-alaska;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2008/1063/stark.pdf.

Stoner:2008:ADM

[4902] Allan W. Stoner, Craig S. Rose, J. Eric Munk, Carwyn F. Hammond, and Michael W. Davis. An assessment of discard mortality for two Alaskan crab species, Tanner crab (Chionoecetes bairdi) and snow crab (C. opilio), based on reflex impairment. Fishery Bulletin, 106(4):??, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/assessment-discard-
mortality-two-alaskan-crab-species-tanner-crab-chionoecetes-
bairdi-and;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2008/1064/stoner.pdf.

Casazza:2008:FAP

[4903] Tara L. Casazza and Steve W. Ross. Fishes associated with pelagic Sargassum and open water lacking Sargassum in the Gulf Stream off North Carolina. Fishery Bulletin, 106(4):??, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/fishes-associated-pelagic-sargassum-and-open-
water-lacking-sargassum-gulf-stream-north;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2008/1064/casazza.
pdf.

Rodgveller:2008:EHC

[4904] Cara J. Rodgveller, Chris R. Lunsford, and Jeffrey T. Fujioka. Evidence of hook competition in longline surveys. Fishery Bulletin, 106(4): 364–374, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/evidence-
hook-competition-longline-surveys;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2008/1064/rodgveller.pdf.

Kastelle:2008:AVD

[4905] Craig R. Kastelle, Delsa M. Anderl, Daniel K. Kimura, and Chris G. Johnston. Age validation of Dover sole (Microstomus pacificus) by means of bomb radiocarbon. Fishery Bulletin, 106(4):??, 2008.

CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/age-validation-dover-
sole-microstomus-pacificus-means-bomb-radiocarbon;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2008/1064/
kastelle.pdf.

Maschner:2008:YTS

[4906] Herbert D. G. Maschner, Matthew W. Betts, Katherine L. Reedy-Maschner, and Andrew W. Trites. A 4500-year time series of Pacific cod (Gadus macrocephalus) size and abundance: archaeology, oceanic regime shifts, and sustainable fisheries. Fishery Bulletin, 106(4):??, 2008. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
4500-year-time-series-pacific-cod-gadus-macrocephalus-size-
and-abundance-archaeology;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2008/1064/maschner.pdf.

Kane:2008:PCB

[4907] Emily A. Kane, Paula A. Olson, Tim Gerrodette, and Paul C. Fiedler. Prevalence of the commensal barnacle Xenobalanus globicipitis on cetacean species in the eastern tropical Pacific Ocean, and a review of global occurrence. Fishery Bulletin, 106(4):??, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/prevalence-commensal-
barnacle-xenobalanus-globicipitis-cetacean-species-eastern-
tropical;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2008/1064/kane.pdf.

Fowler:2008:OOD

[4908] Ashley M. Fowler, Jeffrey M. Leis, and Iain M. Suthers. Onshore–offshore distribution and abundance of tuna larvae (Pisces: Scombridae: Thunnini) in near-reef waters of the Coral Sea. Fishery Bulletin, 106(4): 405–416, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/onshore-
offshore-distribution-and-abundance-tuna-larvae-pisces-scombridae-
thunnini-near;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2008/1064/fowler.pdf.

Jeffers:2008:HBE

[4909] Sarah A. Jeffers, William F. Patterson III, and James H. Cowan, Jr. Habitat and bycatch effects on population parameters of inshore lizardfish (Synodus foetens) in the north central Gulf of Mexico. Fishery Bulletin, 106(4):??, 2008. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
habitat-and-bycatch-effects-population-parameters-inshore-lizardfish-
synodus-foetens-north;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2008/1064/jeffers.pdf.

Brill:2008:ERD

[4910] Richard Brill, Christopher Magel, Michael Davis, Robert Hannah, and Polly Rankin. Effects of rapid decompression and exposure to bright light on visual function in black rockfish (Sebastes melanops) and Pacific halibut (Hippoglossus stenolepis). Fishery Bulletin, 106(4):??, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/effects-rapid-
decompression-and-exposure-bright-light-visual-function-black-
rockfish;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2008/1064/brill.pdf.

Asch:2008:CBM

[4911] Rebecca G. Asch and Jeremy S. Collie. Changes in a benthic megafaunal community due to disturbance from bottom fishing and the establishment of a fishery closure. Fishery Bulletin, 106(4):??, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/changes-benthic-megafaunal-community-
due-disturbance-bottom-fishing-and-establishment;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/2008/1064/
asch.pdf.

Vollen:2008:PBA

[4912] Tone Vollen and Ole T. Albert. Pelagic behavior of adult Greenland halibut (Reinhardtius hippoglossoides). Fishery Bulletin, 106(4): ??, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/pelagic-
behavior-adult-greenland-halibut-reinhardtius-hippoglossoides;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2008/1064/vollen.pdf.

Love:2008:ATE

[4913] Milton S. Love, Donna M. Schroeder, Linda Snook, Anne York, and Guy Cochrane. All their eggs in one basket: a rocky reef nursery for the longnose skate (Raja rhina Jordan & Gilbert, 1880) in the Southern California Bight. Fishery Bulletin, 106(4):??, 2008. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/all-their-eggs-one-basket-rocky-
reef-nursery-longnose-skate-raja-rhina-jordan-gilbert-1880;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2008/1064/love.pdf.

Karlsson:2008:MPR

[4914] Sten Karlsson, Mark A. Renshaw, Caird E. Rexroad III, and John R. Gold. Microsatellite primers for red drum (Sciaenops ocellatus). Fishery Bulletin, 106(4):476–482, 2008. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/microsatellite-primers-red-drum-sciaenops-
ocellatus;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2008/1064/karlsson.pdf.

Stehlik:2009:ESC

[4915] Linda L. Stehlik. Effects of seasonal change on activity rhythms and swimming behavior of age-0 bluefish (Pomatomus saltatrix) and a description of gliding behavior. Fishery Bulletin, 107(1): ??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/effects-
seasonal-change-activity-rhythms-and-swimming-behavior-age-
0-bluefish-pomotomus;   https://spo.nmfs.noaa.gov/sites/default/
files/stehlik.pdf.

Anderson:2009:EAB

[4916] Joel D. Anderson, Dusty L. McDonald, Glen R. Sutton, and William J. Karel. Evolutionary associations between sand seatrout (Cynoscion arenarius) and silver seatrout (C. nothus) inferred from morphological characters, mitochondrial DNA, and microsatellite markers. Fishery Bulletin, 107(1):??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
evolutionary-associations-between-sand-seatrout-cynoscion-arenarius-
and-silver-seatrout-c;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2009/1071/anderson.pdf.

McDonald:2009:SSA

[4917] Dusty L. McDonald, Joel D. Anderson, Britt W. Bumguardner, Fernando Martinez-Andrade, and Josh O. Harper. Spatial and seasonal abundance of sand seatrout (Cynoscion arenarius) and silver seatrout (C. nothus) off the coast of Texas, determined with twenty years of data (1987–2006). Fishery Bulletin, 107(1):??, 2009. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/spatial-and-seasonal-abundance-
sand-seatrout-cynoscion-arenarius-and-silver-seatrout-c;   https:

//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2009/1071/
mcdonald.pdf.

Tolan:2009:BRC

[4918] James M. Tolan and Mark Fisher. Biological response to changes in climate patterns: population increases of gray snapper (Lutjanus griseus) in Texas bays and estuaries. Fishery Bulletin, 107(1):??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/biological-response-
changes-climate-patterns-population-increases-gray-snapper-
lutjanus;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2009/1071/tolan.pdf.

Candy:2009:DPG

[4919] John R. Candy, R. Gregory Bonnell, Terry D. Beacham, Colin G. Wallace, and Ruth E. Withler. Dividing population genetic distance data with the software Partitioning Optimization with Restricted Growth Strings (PORGS): an application for Chinook salmon (Oncorhynchus tshawytscha), Vancouver Island, British Columbia. Fishery Bulletin, 107(1):??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/dividing-population-genetic-distance-data-software-
partitioning-optimization-restricted;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2009/1071/candy.pdf.

Coulson:2009:WBG

[4920] Peter G. Coulson, S. Alex Hesp, Norman G. Hall, and Ian C. Potter. The western blue groper (Achoerodus gouldii), a protogynous hermaphroditic labrid with exceptional longevity, late maturity, slow growth, and both late maturation and sex change. Fishery Bulletin, 107(1):??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/western-blue-groper-achoerodus-gouldii-protogynous-
hermaphroditic-labrid-exceptional;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2009/1071/coulson.pdf.

Wood:2009:RDD

[4921] Anthony D. Wood, Bradley M. Wetherbee, Francis Juanes, Nancy E. Kohler, and Cheryl Wilga. Recalculated diet and daily ration of the shortfin mako (Isurus oxyrinchus), with a focus on quantifying predation on bluefish (Pomatomus saltatrix) in the northwest Atlantic Ocean. Fishery Bulletin, 107(1):??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
recalculated-diet-and-daily-ration-shortfin-mako-isurus-oxyrinchus-
focus-quantifying;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2009/1071/wood.pdf.

Wood:2009:CBW

[4922] Abby Jane M. Wood, Jeremy S. Collie, and Jonathan A. Hare. A comparison between warm-water fish assemblages of Narragansett Bay and those of Long Island Sound waters. Fishery Bulletin, 107(1):89–100, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/comparison-between-
warm-water-fish-assemblages-narragansett-bay-and-those-long-
island-sound;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2009/1071/wood2.pdf.

Staudinger:2009:ROB

[4923] Michelle D. Staudinger, Francis Juanes, and Suzanne Carlson. Reconstruction of original body size and estimation of allometric relationships for the longfin inshore squid (Loligo pealeii) and northern shortfin squid (Illex illecebrosus). Fishery Bulletin, 107(1):

101–105, 2009.  CODEN FSYBAY.  ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
reconstruction-original-body-size-and-estimation-allometric-
relationships-longfin-inshore;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2009/1071/staudinger.pdf.

Powell:2009:MSRa

[4924] Eric N. Powell, John M. Klinck, Kathryn A. Ashton-Alcox, and John N. Kraeuter. Multiple stable reference points in oyster populations: biological relationships for the eastern oyster (Crassostrea virginica) in Delaware Bay. Fishery Bulletin, 107(2):??, 2009. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/multiple-stable-reference-points-
oyster-populations-biological-relationships-eastern-oyster;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2009/1072/powell.pdf.

Powell:2009:MSRb

[4925] Eric N. Powell, John M. Klinck, Kathryn A. Ashton-Alcox, and John N. Kraeuter. Multiple stable reference points in oyster populations: implications for reference point-based management. Fishery Bulletin, 107 (2):??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/multiple-
stable-reference-points-oyster-populations-implications-reference-
point-based;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2009/1072/powell2.pdf.

Witthames:2009:AMD

[4926] Peter R. Witthames, Anders Thorsen, Hilario Murua, Francisco Saborido-Rey, Lorraine N. Greenwood, Rosario Dominguez, Maria Korta, and Olav S. Kjesbu. Advances in methods for determining fecundity: application of the new methods to some marine fishes. Fishery Bulletin, 107(2):??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/advances-
methods-determining-fecundity-application-new-methods-some-
marine-fishes;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2009/1072/witthames.pdf.

MacAvoy:2009:AFM

[4927] Stephen E. MacAvoy, Greg C. Garman, and Stephen A. Macko. Anadromous fish as marine nutrient vectors. Fishery Bulletin, 107(2):??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/anadromous-
fish-marine-nutrient-vectors;   https://spo.nmfs.noaa.gov/sites/
default/files/macavoy.pdf.

Cartwright:2009:DEL

[4928] Rachael L. Cartwright. Description of early life history stages of the northern sculpin (Icelinus borealis Gilbert) (Teleostei: Cottidae). Fishery Bulletin, 107(2):??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/description-early-life-history-stages-northern-
sculpin-icelinus-borealis-gilbert;   https://spo.nmfs.noaa.
gov/sites/default/files/cartwright.pdf.

Ehrhardt:2009:MFC

[4929] Nelson M. Ehrhardt and Vallierre K. W. Deleveaux. Management of fishing capacity in a spiny lobster (Panulirus argus) fishery: analysis of trap performance under the Florida spiny lobster trap certificate program. Fishery Bulletin, 107(2):??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/management-fishing-capacity-spiny-lobster-
panulirus-argus-fishery-analysis-trap-performance;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2009/1072/
ehrhardt.pdf.

Harter:2009:AFP

[4930] Stacey L. Harter, Marta M. Ribera, Andrew N. Shepard, and John K. Reed. Assessment of fish populations and habitat on Oculina Bank, a deep-sea coral marine protected area off eastern Florida. Fishery Bulletin, 107(2):??, 2009. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/assessment-fish-populations-and-
habitat-oculina-bank-deep-sea-coral-marine-protected-area;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/2009/
1072/harter.pdf.

Sewall:2009:CBC

[4931] Fletcher F. Sewall and Cara J. Rodgveller. Changes in body composition and fatty acid profile during embryogenesis of quillback rockfish (Sebastes maliger). Fishery Bulletin, 107(2):??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/changes-body-composition-
and-fatty-acid-profile-during-embryogenesis-quillback-rockfish;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2009/1072/sewall.pdf.

Nichol:2009:EST

[4932] Daniel G. Nichol and David A. Somerton. Evidence of the selection of tidal streams by northern rock sole (Lepidopsetta polyxystra) for transport in the eastern Bering Sea. Fishery Bulletin, 107(2):??, 2009. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/evidence-selection-tidal-
streams-northern-rock-sole-lepidopsetta-polyxystra-transport;
 https://spo.nmfs.noaa.gov/sites/default/files/nichol_0.pdf.

Graham:2009:PWW

[4933] Larissa J. Graham, Mark L. Botton, David Hata, Robert E. Loveland, and Brian R. Murphy. Prosomal-width-to-weight relationships in American horseshoe crabs (Limulus polyphemus): examining conversion factors used to estimate landings. Fishery Bulletin, 107(2): ??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/prosomal-
width-weight-relationships-american-horseshoe-crabs-limulus-
polyphemus-examining;   https://spo.nmfs.noaa.gov/sites/default/
files/graham.pdf.

Beacham:2009:PSC

[4934] Terry D. Beacham, John R. Candy, Khai D. Le, and Michael Wetklo. Population structure of chum salmon (Oncorhynchus keta) across the Pacific rim, determined from microsatellite analysis. Fishery Bulletin, 107(2):??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/population-structure-chum-salmon-oncorhynchus-keta-
across-pacific-rim-determined;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2009/1072/beacham.pdf.

Chiang:2009:ASS

[4935] Wei-Chuan Chiang, Chi-Lu Sun, Sheng-Ping Wang, Su-Zan Yeh, Yong Chen, Wei-Cheng Su, Don-Chung Liu, and Wen-Yie Chen. Analysis of sex-specific spawning biomass per recruit of the sailfish (Istiophorus platypterus) in the waters off eastern Taiwan. Fishery Bulletin, 107 (3):??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/analysis-
sex-specific-spawning-biomass-recruit-sailfish-istiophorus-
platypterus-waters;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2009/1073/chiang.pdf.

Boldt:2009:ACD

[4936] Jennifer L. Boldt and Christopher N. Rooper. Abundance, condition, and diet of juvenile Pacific Ocean perch (Sebastes alutus) in the Aleutian Islands. Fishery Bulletin, 107(3):??, 2009. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/abundance-condition-
and-diet-juvenile-pacific-ocean-perch-sebastes-alutus-aleutian-
islands;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2009/1073/boldt.pdf.

Broadhurst:2009:MST

[4937] Matt K. Broadhurst, Russell B. Millar, Craig P. Brand, and Sebastan S. Uhlmann. Modified sorting technique to mitigate the collateral mortality of trawled school prawns (Metapenaeus macleayi). Fishery Bulletin, 107(3):??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/modified-sorting-technique-mitigate-collateral-
mortality-trawled-school-prawns-metapenaeus;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/2009/1073/broadhurst.
pdf.

Brill:2009:RFD

[4938] Richard Brill, Peter Bushnell, Leonie Smith, Coley Speaks, Rumya Sundaram, Eric Stroud, and John Wang. The repulsive and feedingdeterrent effects of electropositive metals on juvenile sandbar sharks (Carcharhinus plumbeus). Fishery Bulletin, 107(3):??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/repulsive-and-feeding-deterrent-
effects-electropositive-metals-juvenile-sandbar-sharks;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2009/1073/
brill.pdf.

Grizzle:2009:ELF

[4939] Raymond E. Grizzle, Larry G. Ward, Larry A. Mayer, Mashkoor A. Malik, Andrew B. Cooper, Holly A. Abeels, Jennifer K. Greene, Melissa A. Brodeur, and Andrew A. Rosenberg. Effects of a large fishing closure on benthic communities in the western Gulf of Maine: recovery from the effects of gillnets and otter trawls. Fishery Bulletin, 107(3): ??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/effects-large-
fishing-closure-benthic-communities-western-gulf-maine-recovery-
effects;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2009/1073/grizzle.pdf.

Riley:2009:DGH

[4940] Kenneth L. Riley, Charles R. Weirich, and David Cerino. Development and growth of hatchery-reared larval Florida pompano (Trachinotus carolinus). Fishery Bulletin, 107(3):318–328, 2009. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/development-and-growth-hatchery-
reared-larval-florida-pompano-trachinotus-carolinus;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2009/1073/
riley.pdf.

Mather:2009:UNN

[4941] Martha E. Mather, John T. Finn, Kristen H. Ferry, Linda A. Deegan, and Gary A. Nelson. Use of non-natal estuaries by migratory striped bass (Morone saxatilis) in summer. Fishery Bulletin, 107(3):??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/use-non-natal-estuaries-migratory-striped-bass-
morone-saxatilis-summer;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2009/1073/mather.pdf.

Hoolihan:2009:SMS

[4942] John P. Hoolihan, Nerida F. Perez, Ronald M. Faugue, Andrea M. Bernard, Rebekah L. Horn, Derke Snodgrass, and Duane R. Schultz. Surface mucous as a source of genomic DNA from Atlantic billfishes (Istiophoridae) and swordfish (Xiphiidae). Fishery Bulletin, 107(3): ??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/surface-
mucous-source-genomic-dna-atlantic-billfishes-istiophoridae-
and-swordfish-xiphiidae;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2009/1073/hoolihan.pdf.

Lopez-Rasgado:2009:AHQ

[4943] Francisco Javier López-Rasgado and Sharon Z. Herzka. Assessment of habitat quality for juvenile California halibut (Paralichthys californicus) in a seasonally arid estuary. Fishery Bulletin, 107(3):??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/assessment-habitat-
quality-juvenile-california-halibut-paralichthys-californicus-
seasonally;   https://spo.nmfs.noaa.gov/sites/default/files/
lopez_0.pdf.

Grigg:2009:SSR

[4944] Emma K. Grigg, A. Peter Klimley, Sarah G. Allen, Deborah E. Green, Deborah L. Elliott-Fisk, and Hal Markowitz. Spatial and seasonal relationships between Pacific harbor seals (Phoca vitulina richardii) and their prey, at multiple scales. Fishery Bulletin, 107(3):??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/spatial-and-seasonal-
relationships-between-pacific-harbor-seals-phoca-vitulina-richardii;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2009/1073/grigg.pdf.

Graves:2009:UPS

[4945] John E. Graves, Andrij Z. Horodysky, and Robert J. Latour. Use of pop-up satellite archival tag technology to study postrelease survival of and habitat use by estuarine and coastal fishes: an application to striped bass (Morone saxatilis). Fishery Bulletin, 107(3):373–383, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/use-pop-satellite-archival-tag-technology-
study-postrelease-survival-and-habitat-use;   https://spo.nmfs.
noaa.gov/sites/default/files/graves.pdf.

Craddock:2009:FHA

[4946] James E. Craddock, Pamela T. Polloni, Brett Hayward, and Frederick Wenzel. Food habits of Atlantic white-sided dolphins (Lagenorhynchus acutus) off the coast of New England. Fishery Bulletin, 107(3): ??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/food-
habits-atlantic-white-sided-dolphins-lagenorhynchus-acutus-
coast-new-england;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2009/1073/craddock.pdf.

Humborstad:2009:RIM

[4947] Odd-Børre Humborstad, Michael W. Davis, and Svein Løkkeborg. Reflex impairment as a measure of vitality and survival potential of Atlantic cod (Gadus morhua). Fishery Bulletin, 107(3):395– 402, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/reflex-
impairment-measure-vitality-and-survival-potential-atlantic-
cod-gadus-morhua;   https://spo.nmfs.noaa.gov/sites/default/
files/humborstad.pdf.

Bacheler:2009:VMP

[4948] Nathan M. Bacheler, Lee M. Paramore, Summer M. Burdick, Jeffrey A. Buckel, and Joseph E. Hightower. Variation in movement patterns of red drum (Sciaenops ocellatus) inferred from conventional tagging and ultrasonic telemetry. Fishery Bulletin, 107(4):??, 2009. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/variation-movement-patterns-red-
drum-sciaenops-ocellatus-inferred-conventional-tagging-and;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2009/1074/bacheler.pdf.

Sun:2009:RBB

[4949] Chi-Lu Sun, Yi-Jay Chang, Chien-Chung Tszeng, Su-Zan Yeh, and Nan-Jay Su. Reproductive biology of blue marlin (Makaira nigricans) in the western Pacific Ocean. Fishery Bulletin, 107 (4):420–432, 2009. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
reproductive-biology-blue-marlin-makaira-nigricans-western-
pacific-ocean;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2009/1074/sun.pdf.

Reese:2009:STD

[4950] Carl Reese, Nicola Hillgruber, Molly Sturdevant, Alex Wertheimer, William Smoker, and Rick Focht. Spatial and temporal distribution and the potential for estuarine interactions between wild and hatchery chum salmon (Oncorhynchus keta) in Taku Inlet, Alaska. Fishery Bulletin, 107 (4):??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/spatial-and-
temporal-distribution-and-potential-estuarine-interactions-
between-wild-and;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2009/1074/reese.pdf.

Stoner:2009:PDM

[4951] Allan W. Stoner. Prediction of discard mortality for Alaskan crabs after exposure to freezing temperatures, based on a reflex impairment index. Fishery Bulletin, 107(4):??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/prediction-discard-mortality-alaskan-
crabs-after-exposure-freezing-temperatures-based;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/2009/1074/
stoner.pdf.

Dierking:2009:DCP

[4952] Jan Dierking, Ivor D. Williams, and William J. Walsh. Diet composition and prey selection of the introduced grouper species peacock hind (Cephalopholis argus) in Hawaii. Fishery Bulletin, 107(4):??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/diet-composition-and-
prey-selection-introduced-grouper-species-peacock-hind-cephalopholis;
 https://spo.nmfs.noaa.gov/sites/default/files/dierking.pdf.

Cox:2009:EPA

[4953] M. Keith Cox and Ron Heintz. Electrical phase angle as a new method to measure fish condition. Fishery Bulletin, 107(4):??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/electrical-
phase-angle-new-method-measure-fish-condition;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/2009/1074/cox.
pdf.

Martinson:2009:GSS

[4954] Ellen C. Martinson, John H. Helle, Dennis L. Scarnecchia, and Houston H. Stokes. Growth and survival of sockeye salmon (Oncorhynchus nerka) from Karluk Lake and River, Alaska, in relation to climatic and oceanic regimes and indices, 1922–2000. Fishery Bulletin, 107(4):488– 500, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/growth-and-
survival-sockeye-salmon-oncorhynchus-nerka-karluk-lake-and-
river-alaska-relation;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2009/1074/martinson.pdf.

Shulzitski:2009:PCA

[4955] Kathryn Shulzitski, Michael A. McCartney, and Michael L. Burton. Population connectivity among Dry Tortugas, Florida, and Caribbean populations of mutton snapper (Lutjanus analis), inferred from multiple microsatellite loci. Fishery Bulletin, 107(4):??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/population-
connectivity-among-dry-tortugas-florida-and-caribbean-populations-
mutton-snapper;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2009/1074/shulzitski.pdf.

Lee:2009:DVT

[4956] Yong-Woo Lee and David B. Sampson. Dietary variations in three co-occurring rockfish species off the Pacific northwest during anomalous oceanographic events in 1998 and 1999. Fishery Bulletin, 107(4):??, 2009. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/dietary-variations-three-co-occurring-rockfish-
species-pacific-northwest-during-anomalous;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2009/1074/lee.pdf.

Polovina:2009:IRA

[4957] Jeffrey J. Polovina, Melanie Abecassis, Evan A. Howell, and Phoebe Woodworth. Increases in the relative abundance of mid-trophic level fishes concurrent with declines in apex predators in the subtropical North Pacific, 1996–2006. Fishery Bulletin, 107(4):??, 2009. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/increases-relative-abundance-mid-
trophic-level-fishes-concurrent-declines-apex-predators;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2009/1074/
polovina.pdf.

Chen:2010:MAI

[4958] Xinjun Chen, Siquan Tian, Yong Chen, and Bilin Liu. A modeling approach to identify optimal habitat and suitable fishing grounds for neon flying squid (Ommastrephes bartramii) in the Northwest Pacific Ocean. Fishery Bulletin, 108(1):1–14, 2010. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/modeling-approach-identify-optimal-habitat-and-
suitable-fishing-grounds-neon-flying-squid;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2010/1081/chen.pdf.

Cadigan:2010:SIA

[4959] Noel G. Cadigan and Jeff J. Dowden. Statistical inference about the relative efficiency of a new survey protocol, based on paired-tow survey calibration data. Fishery Bulletin, 108(1):??, 2010. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/statistical-inference-about-relative-
efficiency-new-survey-protocol-based-paired-tow-survey;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2010/1081/
cadigan.pdf.

Hannah:2010:ETO

[4960] Robert W. Hannah, Stephen A. Jones, William Miller, and Jayme S. Knight. Effects of trawling for ocean shrimp (Pandalus jordani) on macroinvertebrate abundance and diversity at four sites near Nehalem Bank, Oregon. Fishery Bulletin, 108(1):??, 2010. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/effects-trawling-ocean-shrimp-pandalus-
jordani-macroinvertebrate-abundance-and-diversity;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2010/1081/
hannah.pdf.

Barry:2010:ECE

[4961] Patrick D. Barry, Sherry L. Tamone, and David A. Tallmon. Evaluation of the capture efficiency and size selectivity of four pot types in the prospective fishery for North Pacific giant octopus (Enteroctopus dofleini). Fishery Bulletin, 108(1):39–44, 2010. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/evaluation-capture-
efficiency-and-size-selectivity-four-pot-types-prospective-
fishery-north;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2010/1081/barry.pdf.

Saldierna-Martinez:2010:LDS

[4962] Ricardo J. Saldierna-Martínez, Gerardo Aceves-Medina, and Enrique A. González-Navarro. Larval development of the spotfin tonguefish (Symphurus oligomerus Mahadeva and Munroe, 1990) (Pleuronectiformes:

Cynoglossidae) from the Gulf of California, Mexico. Fishery Bulletin, 108(1):??, 2010. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
larval-development-spotfin-tonguefish-symphurus-oligomerus-
mahadeva-and-munroe-1990;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2010/1081/martinez.pdf.

Butler:2010:FEA

[4963] Christopher M. Butler, Paul J. Rudershausen, and Jeffrey A. Buckel. Feeding ecology of Atlantic bluefin tuna (Thunnus thynnus) in North Carolina: diet, daily ration, and consumption of Atlantic menhaden (Brevoortia tyrannus). Fishery Bulletin, 108(1):??, 2010. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/feeding-ecology-atlantic-bluefin-
tuna-thunnus-thynnus-north-carolina-diet-daily-ration-and;
https://spo.nmfs.noaa.gov/sites/default/files/butler.pdf.

Chilton:2010:MGF

[4964] Elizabeth A. Chilton. Maturity and growth of female dusky rockfish (Sebastes variabilis) in the central Gulf of Alaska. Fishery Bulletin, 108(1):??, 2010. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/maturity-and-growth-female-dusky-rockfish-sebastes-
variabilis-central-gulf-alaska;   https://spo.nmfs.noaa.gov/
sites/default/files/chilton.pdf.

Miller:2010:ECS

[4965] Cara E. Miller and Donald M. Baltz. Environmental characterization of seasonal trends and foraging habitat of bottlenose dolphins (Tursiops truncatus) in northern Gulf of Mexico bays. Fishery Bulletin, 108(1):??, 2010. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/environmental-characterization-seasonal-trends-and-
foraging-habitat-bottlenose-dolphins;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2010/1081/miller.pdf.

Lynch:2010:MGI

[4966] Abigail J. Lynch, Jan R. McDowell, and John E. Graves. A molecular genetic investigation of the population structure of Atlantic menhaden (Brevoortia tyrannus). Fishery Bulletin, 108(1):??, 2010. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/molecular-genetic-investigation-
population-structure-atlantic-menhaden-brevoortia-tyrannus;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2010/1081/lynch.pdf.

Gerard:2010:VIS

[4967] Trika Gerard and Barbara Muhling. Variation in the isotopic signatures of juvenile gray snapper (Lutjanus griseus) from five southern Florida regions. Fishery Bulletin, 108(1):??, 2010. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/variation-isotopic-
signatures-juvenile-gray-snapper-lutjanus-griseus-five-southern-
florida;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2010/1081/gerard.pdf.

Gannon:2010:ATD

[4968] Damon P. Gannon and Janet G. Gannon. Assessing trends in the density of Atlantic croaker (Micropogonias undulatus): a comparison of passive acoustic and trawl methods. Fishery Bulletin, 108(1):??, 2010. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/assessing-trends-density-
atlantic-croaker-micropogonias-undulatus-comparison-passive;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2010/1081/gannon.pdf.

Tribuzio:2010:AGS

[4969] Cindy A. Tribuzio, Gordon H. Kruse, and Jeffrey T. Fujioka. Age and growth of spiny dogfish (Squalus acanthias) in the Gulf of Alaska: analysis of alternative growth models. Fishery Bulletin, 108(2):119–135, 2010. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/age-and-growth-spiny-
dogfish-squalus-acanthias-gulf-alaska-analysis-alternative-
growth;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2010/1082/tribuzio.pdf.

Rose:2010:EHF

[4970] Craig S. Rose, John R. Gauvin, and Carwyn F. Hammond. Effective herding of flatfish by cables with minimal seafloor contact. Fishery Bulletin, 108(2):136–144, 2010. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
effective-herding-flatfish-cables-minimal-seafloor-contact;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2010/1082/rose.pdf.

Ryer:2010:FHB

[4971] Clifford H. Ryer, Craig S. Rose, and Paul J. Iseri. Flatfish herding behavior in response to trawl sweeps: a comparison of diel responses to conventional sweeps and elevated sweeps. Fishery Bulletin, 108(2):145–154, 2010. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
flatfish-herding-behavior-response-trawl-sweeps-comparison-
diel-responses-conventional;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2010/1082/ryer.pdf.

Mateo:2010:STV

[4972] Ivan Mateo, Edward G. Durbin, David A. Bengtson, Richard Kingsley, Peter K. Swart, and Daisy Durant. Spatial and temporal variation in otolith chemistry for tautog (Tautoga onitis) in Narragansett Bay and Rhode Island coastal ponds. Fishery Bulletin, 108(2):??, 2010. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/spatial-and-
temporal-variation-otolith-chemistry-tautog-tautoga-onitis-
narragansett-bay-and;   https://spo.nmfs.noaa.gov/sites/default/
files/mateo.pdf.

Masuda:2010:FAA

[4973] Reiji Masuda, Masami Shiba, Yoh Yamashita, Masahiro Ueno, Yoshiaki Kai, Asami Nakanishi, Masaru Torikoshi, and Masaru Tanaka. Fish assemblages associated with three types of artificial reefs: density of assemblages and possible impacts on adjacent fish abundance. Fishery Bulletin, 108(2):162–173, 2010. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fish-assemblages-associated-three-types-artificial-
reefs-density-assemblages-and-possible;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2010/1082/masuda.pdf.

Lo:2010:BRP

[4974] Nancy C. H. Lo, Beverly J. Macewicz, and David A. Griffiths. Biomass and reproduction of Pacific sardine (Sardinops sagax) off the Pacific northwestern United States, 2003–2005. Fishery Bulletin, 108(2):174–192, 2010. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/biomass-and-reproduction-pacific-sardine-sardinops-
sagax-pacific-northwestern-united-states;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2010/1082/lo.pdf.

Hernandez:2010:SVI

[4975] Frank J. Hernandez, Jr., Sean P. Powers, and William M. Graham. Seasonal variability in ichthyoplankton abundance and assemblage composition in the northern Gulf of Mexico off Alabama. Fishery Bulletin, 108(2):??, 2010. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/seasonal-variability-ichthyoplankton-abundance-and-
assemblage-composition-northern-gulf;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2010/1082/hernandez.pdf.

Stevenson:2010:ORS

[4976] Duane E. Stevenson and Kristy A. Lewis. Observer-reported skate bycatch in the commercial groundfish fisheries of Alaska. Fishery Bulletin, 108(2):??, 2010. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/observer-
reported-skate-bycatch-commercial-groundfish-fisheries-alaska;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2010/1082/stevenson.pdf.

Fergusson:2010:ESE

[4977] Emily A. Fergusson, Molly V. Sturdevant, and Joseph A. Orsi. Effects of starvation on energy density of juvenile chum salmon (Oncorhynchus keta) captured in marine waters of Southeastern Alaska. Fishery Bulletin, 108(2):218–225, 2010. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
effects-starvation-energy-density-juvenile-chum-salmon-oncorhynchus-
keta-captured-marine;   https://spo.nmfs.noaa.gov/sites/default/
files/fergusson.pdf.

Fruh:2010:ASD

[4978] Erica L. Fruh, Aimee Keller, Jessica Trantham, and Victor Simon. Accuracy of sex determination for northeastern Pacific Ocean thornyheads (Sebastolobus altivelis and s. alascanus). Fishery Bulletin, 108(2):??, 2010. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/accuracy-sex-
determination-northeastern-pacific-ocean-thornyheads-sebastolobus-
altivelis;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2010/1082/fruh.pdf.

Jacobson:2010:MEB

[4979] Larry D. Jacobson, Kevin D. E. Stokesbury, Melissa A. Allard, Antonie Chute, Bradley P. Harris, Deborah Hart, Tom Jaffarian, Michael C. Marino II, Jacob I. Nogueira, and Paul Rago. Measurement errors in body size of sea scallops (Placopecten magellanicus) and their effect on stock assessment models. Fishery Bulletin, 108(2):233–247, 2010. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/measurement-errors-body-size-sea-scallops-
placopecten-magellanicus-and-their-effect-stock;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/2010/1082/
jacobson.pdf.

Hobbs:2010:AHP

[4980] Roderick C. Hobbs and Janice M. Waite. Abundance of harbor porpoise (Phocoena phocoena) in three Alaskan regions, corrected for observer errors due to perception bias and species misidentification, and corrected for animals submerged from view. Fishery Bulletin, 108 (3):??, 2010. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/abundance-
harbor-porpoise-phocoena-phocoena-three-alaskan-regions-corrected-
observer-errors;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2010/1083/hobbs.pdf.

Poisson:2010:ELC

[4981] François Poisson, Jean-Claude Gaertner, Marc Taquet, Jean-Pierre Durbec, and Keith Bigelow. Effects of lunar cycle and fishing operations on longline-caught pelagic fish: fishing performance, capture time, and survival of fish. Fishery Bulletin, 108(3):??, 2010. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/effects-
lunar-cycle-and-fishing-operations-longline-caught-pelagic-
fish-fishing-performance;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2010/1083/poisson.pdf.

McAllister:2010:UEE

[4982] Murdoch K. McAllister, Richard D. Stanley, and Paul Starr. Using experiments and expert judgment to model catchability of Pacific rockfishes in trawl surveys, with application to bocaccio (Sebastes paucispinis) off British Columbia. Fishery Bulletin, 108 (3):??, 2010. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/using-
experiments-and-expert-judgment-model-catchability-pacific-
rockfishes-trawl-surveys;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2010/1083/mcallister.pdf.

Patrick:2010:UPS

[4983] Wesley S. Patrick, Paul Spencer, Jason Link, Jason Cope, John Field, Donald Kobayashi, Peter Lawson, Todd Gedamke, Enric Cortés, Olav Ormseth, Keith Bigelow, and William Overholtz. Using productivity and susceptibility indices to assess the vulnerability of United States fish stocks to overfishing. Fishery Bulletin, 108(3):??, 2010. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/using-productivity-
and-susceptibility-indices-assess-vulnerability-united-states-
fish;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2010/1083/patrick.pdf.

Munroe:2010:DLH

[4984] Thomas A. Munroe and Steve W. Ross. Distribution and life history of two diminutive flatfishes, Citharichthys gymnorhinus and C. cornutus (Pleuronectiformes: Paralichthyidae), in the western North Atlantic. Fishery Bulletin, 108(3):??, 2010. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/distribution-and-life-history-two-diminutive-flatfishes-
citharichthys-gymnorhinus-and-c;   https://spo.nmfs.noaa.gov/
sites/default/files/munroe.pdf.

Passerotti:2010:AVG

[4985] Michelle S. Passerotti, John K. Carlson, Andrew N. Piercy, and Steven E. Campana. Age validation of great hammerhead shark (Sphyrna mokarran), determined by bomb radiocarbon analysis. Fishery Bulletin, 108(3):??, 2010. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/age-
validation-great-hammerhead-shark-sphyrna-mokarran-determined-
bomb-radiocarbon-analysis;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2010/1083/passerotti.pdf.

Williams:2010:USC

[4986] Kresimir Williams, Christopher N. Rooper, and Rick Towler. Use of stereo camera systems for assessment of rockfish abundance in untrawlable areas and for recording pollock behavior during midwater trawls. Fishery Bulletin, 108(3):352–362, 2010. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/use-stereo-camera-systems-assessment-
rockfish-abundance-untrawlable-areas-and-recording;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2010/1083/
williams.pdf.

Jones:2010:SCE

[4987] Ashlee A. Jones, Norman G. Hall, and Ian C. Potter. Species compositions of elasmobranchs caught by three different commercial fishing methods off southwestern Australia, and biological data for four abundant bycatch species. Fishery Bulletin, 108(4):365–381, 2010. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/species-compositions-
elasmobranchs-caught-three-different-commercial-fishing-methods;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2010/1084/jones.pdf.

Hurst:2010:OPT

[4988] Thomas P. Hurst, Benjamin J. Laurel, and Lorenzo Ciannelli. Ontogenetic patterns and temperature-dependent growth rates in early life stages of Pacific cod (Gadus macrocephalus). Fishery Bulletin, 108(4):382–392, 2010. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/ontogenetic-patterns-and-temperature-dependent-
growth-rates-early-life-stages-pacific-cod;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2010/1084/hurst.pdf.

Bollens:2010:FEJ

[4989] Stephen M. Bollens, Rian vanden Hooff, Mari Butler, Jeffery R. Cordell, and Bruce W. Frost. Feeding ecology of juvenile Pacific salmon (Oncorhynchus spp.) in a northeast Pacific fjord: diet, availability of zooplankton, selectivity for prey, and potential competition for prey resources. Fishery Bulletin, 108(4):??, 2010. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/feeding-ecology-juvenile-pacific-
salmon-oncorhynchus-spp-northeast-pacific-fjord-diet;   https:
//spo.nmfs.noaa.gov/sites/default/files/bollens.pdf.

DeCelles:2010:MPW

[4990] Gregory R. DeCelles and Steven X. Cadrin. Movement patterns of winter flounder (Pseudopleuronectes americanus) in the southern Gulf of Maine: observations with the use of passive acoustic telemetry. Fishery Bulletin, 108(4):??, 2010. CODEN FSY-

BAY.  ISSN 0090-0656 (print), 1937-4518 (electronic).  URL

https://spo.nmfs.noaa.gov/content/movement-patterns-winter-
flounder-pseudopleuronectes-americanus-southern-gulf-maine;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2010/1084/decelles.pdf.

Heupel:2010:DCE

[4991] Michelle R. Heupel, Ashley J. Williams, David J. Welch, Campbell R. Davies, Ann Penny, Jacob P. Kritzer, Ross J. Marriott, and Bruce D. Mapstone. Demographic characteristics of exploited tropical lutjanids: a comparative analysis. Fishery Bulletin, 108(4):??, 2010. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/demographic-
characteristics-exploited-tropical-lutjanids-comparative-analysis;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2010/1084/heupel.pdf.

Graves:2010:ACB

[4992] John E. Graves and Andrij Z. Horodysky. Asymmetric conservation benefits of circle hooks in multispecies billfish recreational fisheries: a synthesis of hook performance and analysis of blue marlin (Makaira nigricans) postrelease survival. Fishery Bulletin, 108(4):??, 2010. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
asymmetric-conservation-benefits-circle-hooks-multispecies-
billfish-recreational-fisheries;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2010/1084/graves.pdf.

Laidig:2010:IOC

[4993] Thomas E. Laidig. Influence of ocean conditions on the timing of early life history events for blue rockfish (Sebastes mystinus) off California. Fishery Bulletin, 108(4):??, 2010. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
influence-ocean-conditions-timing-early-life-history-events-
blue-rockfish-sebastes-mystinus;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2010/1084/laidig.pdf.

Dunton:2010:ADA

[4994] Keith J. Dunton, Adrian Jordaan, Kim A. McKown, David O. Conover, and Michael G. Frisk. Abundance and distribution of Atlantic sturgeon (Acipenser oxyrinchus) within the Northwest Atlantic Ocean, determined from five fishery-independent surveys. Fishery Bulletin, 108(4): ??, 2010. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/abundance-and-
distribution-atlantic-sturgeon-acipenser-oxyrinchus-within-
northwest-atlantic;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2010/1084/dunton.pdf.

Hart:2010:DNA

[4995] Ted D. Hart, Julia E. R. Clemons, W. Waldo Wakefield, and Selina S. Heppell. Day and night abundance, distribution, and activity patterns of demersal fishes on Heceta Bank, Oregon. Fishery Bulletin, 108(4):??, 2010. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
day-and-night-abundance-distribution-and-activity-patterns-
demersal-fishes-heceta-bank;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2010/1084/hart.pdf.

Simms:2010:DGM

[4996] Jeffrey R. Simms, Jay R. Rooker, Scott A. Holt, G. Joan Holt, and Jessica Bangma. Distribution, growth, and mortality of sailfish (Istiophorus platypterus) larvae in the northern Gulf of Mexico. Fishery Bulletin, 108 (4):??, 2010. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/distribution-
growth-and-mortality-sailfish-istiophorus-platypterus-larvae-
northern-gulf;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2010/1084/simms.pdf.

Campfield:2011:ICS

[4997] Patrick A. Campfield and Edward D. Houde. Ichthyoplankton community structure and comparative trophodynamics in an estuarine transition zone. Fishery Bulletin, 109(1):1–19, 2011. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/ichthyoplankton-community-structure-and-comparative-
trophodynamics-estuarine-transition;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2011/1091/campfield.pdf.

Friess:2011:AGS

[4998] Claudia Friess and George R. Sedberry. Age, growth, and spawning season of red bream (Beryx decadactylus) off the southeastern United States. Fishery Bulletin, 109(1):??, 2011. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/age-growth-and-spawning-season-red-bream-beryx-
decadactylus-southeastern-united-states;   https://spo.nmfs.
noaa.gov/sites/default/files/friess.pdf.

Cox:2011:MRR

[4999] M. Keith Cox, Ron Heintz, and Kyle Hartman. Measurements of resistance and reactance in fish with the use of bioelectrical impedance analysis: sources of error. Fishery Bulletin, 109(1):34–47, 2011. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/measurements-
resistance-and-reactance-fish-use-bioelectrical-impedance-analysis-
sources;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2011/1091/cox.pdf.

Sturdivant:2011:EEB

[5000] S. Kersey Sturdivant and Kelton L. Clark. An evaluation of the effects of blue crab (Callinectes sapidus) behavior on the efficacy of crab pots as a tool for estimating population abundance. Fishery Bulletin, 109(1):??, 2011. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/evaluation-effects-
blue-crab-callinectes-sapidus-behavior-efficacy-crab-pots-tool;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2011/1091/sturdivant.pdf.

Merritt:2011:BBC

[5001] Daniel Merritt, Mary K. Donovan, Christopher Kelley, Lynn Waterhouse, Michael Parke, Kevin Wong, and Jeffrey C. Drazen. BotCam: a baited camera system for nonextractive monitoring of bottomfish species. Fishery Bulletin, 109(1):??, 2011. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/botcam-baited-camera-system-nonextractive-
monitoring-bottomfish-species;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2011/1091/merritt.pdf.

Able:2011:LAS

[5002] Kenneth W. Able, Mark C. Sullivan, Jonathan A. Hare, Gretchen Bath-Martin, J. Christopher Taylor, and Roland Hagan. Larval abundance of summer flounder (Paralichthys dentatus) as a measure of recruitment and stock status. Fishery Bulletin, 109(1):??, 2011. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/larval-abundance-summer-flounder-
paralichthys-dentatus-measure-recruitment-and-stock-status;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2011/1091/able.pdf.

Kurita:2011:ETD

[5003] Yutaka Kurita, Yuichiro Fujinami, and Masafumi Amano. The effect of temperature on the duration of spawning markers — migratory-nucleus and hydrated oocytes and postovulatory follicles

— in the multiple-batch spawner Japanese flounder (Paralichthys olivaceus). Fishery Bulletin, 109(1):79–89, 2011. CODEN FSYBAY.

ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/effect-temperature-duration-spawning-
markers%E2%80%94migratory-nucleus-and-hydrated-oocytes-and;
 https://spo.nmfs.noaa.gov/sites/default/files/kurita.pdf.

Ballagh:2011:IMD

[5004] Aaron C. Ballagh, David Welch, Ashley J. Williams, Amos Mapleston, Andrew Tobin, and Nicholas Marton. Integrating methods for determining length-at-age to improve growth estimates for two large scombrids. Fishery Bulletin, 109(1):??, 2011. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/integrating-methods-determining-length-age-improve-
growth-estimates-two-large-scombrids;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2011/1091/ballagh.pdf.

Booth:2011:AVG

[5005] Anthony J. Booth, Alan J. Foulis, and Malcolm J. Smale. Age validation, growth, mortality, and demographic modeling of spotted gully shark (Triakis megalopterus) from the southeast coast of South Africa. Fishery Bulletin, 109(1):101–112, 2011. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/age-validation-growth-mortality-
and-demographic-modeling-spotted-gully-shark-triakis;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2011/1091/
booth.pdf.

Lane:2011:OTV

[5006] Hillary A. Lane, Andrew J. Westgate, and Heather N. Koopman. Ontogenetic and temporal variability in the fat content and fatty acid composition of Atlantic herring (Clupea harengus) from the Bay of Fundy, Canada. Fishery Bulletin, 109(1):113–122, 2011. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/ontogenetic-and-temporal-variability-
fat-content-and-fatty-acid-composition-atlantic;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/2011/1091/
lane.pdf.

DeMartini:2011:EBS

[5007] Edward E. DeMartini, Alan R. Everson, and Ryan S. Nichols. Estimates of body sizes at maturation and at sex change, and the spawning seasonality and sex ratio of the endemic Hawaiian grouper (Hyporthodus quernus, F. Epinephelidae). Fishery Bulletin, 109(1):123–134, 2011. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/estimates-body-sizes-
maturation-and-sex-change-and-spawning-seasonality-and-sex-
ratio;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2011/1091/dimartini.pdf.

Sanchez-Rubio:2011:CRH

[5008] Guillermo Sanchez-Rubio, Harriet M. Perry, Patricia M. Biesiot, Donald R. Johnson, and Romuald N. Lipcius. Climate-related hydrological regimes and their effects on abundance of juvenile blue crabs (Callinectes sapidus) in the northcentral Gulf of Mexico. Fishery Bulletin, 109(2):139–146, 2011. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/climate-related-hydrological-regimes-
and-their-effects-abundance-juvenile-blue-crabs;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/2011/1092/
rubio.pdf.

Fry:2011:MRS

[5009] Brian Fry. Mississippi River sustenance of brown shrimp (Farfantepenaeus aztecus) in Louisiana coastal waters. Fishery Bulletin, 109(2):??, 2011. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/mississippi-
river-sustenance-brown-shrimp-farfantepenaeus-aztecus-louisiana-
coastal-waters;   https://spo.nmfs.noaa.gov/sites/default/files/
fry_0.pdf.

Duffy-Anderson:2011:SSV

[5010] Janet T. Duffy-Anderson, Deborah M. Blood, and Kathryn L. Mier. Stage-specific vertical distribution of Alaska plaice (Pleuronectes quadrituberculatus) eggs in the eastern Bering Sea. Fishery Bulletin, 109(2):162–169, 2011. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/stage-specific-vertical-distribution-alaska-plaice-
pleuronectes-quadrituberculatus-eggs;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2011/1092/anderson.pdf.

Prista:2011:USM

[5011] Nuno Prista, Norou Diawara, Maria José Costa, and Cynthia Jones. Use of SARIMA models to assess data-poor fisheries: a case study with a sciaenid fishery off Portugal. Fishery Bulletin, 109(2):??, 2011. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/use-sarima-models-assess-
data-poor-fisheries-case-study-sciaenid-fishery-portugal;

https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/2011/
1092/prista.pdf.

Reum:2011:SDD

[5012] Jonathan C. P. Reum and Timothy E. Essington. Season- and depthdependent variability of a demersal fish assemblage in a large fjord estuary (Puget Sound, Washington). Fishery Bulletin, 109(2):??, 2011. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/season-and-
depth-dependent-variability-demersal-fish-assemblage-large-
fjord-estuary-puget;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2011/1092/reum.pdf.

He:2011:IAD

[5013] Xi He, Stephen Ralston, and Alec D. MacCall. Interactions of agedependent mortality and selectivity functions in age-based stock assessment models. Fishery Bulletin, 109(2):198–216, 2011. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/interactions-age-dependent-mortality-
and-selectivity-functions-age-based-stock-assessment;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2011/1092/
he.pdf.

Ralston:2011:MAA

[5014] Stephen Ralston, André E. Punt, Owen S. Hamel, John D. De-Vore, and Ramon J. Conser. A meta-analytic approach to quantifying scientific uncertainty in stock assessments. Fishery Bulletin, 109(2):??, 2011. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/meta-
analytic-approach-quantifying-scientific-uncertainty-stock-
assessments;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2011/1092/ralston.pdf.

Collins:2011:DDS

[5015] Angela B. Collins and Richard S. McBride. Demographics by depth: spatially explicit life-history dynamics of a protogynous reef fish. Fishery Bulletin, 109(2):??, 2011. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/demographics-depth-spatially-explicit-life-history-
dynamics-protogynous-reef-fish;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2011/1092/collins.pdf.

Chuwen:2011:CCR

[5016] Benjamin M. Chuwen, Ian C. Potter, Norman G. Hall, Steeg D. Hoeksema, and Laurie J. B. Laurenson. Changes in catch rates and length and age at maturity, but not growth, of an estuarine plotosid (Cnidoglanis macrocephalus) after heavy fishing. Fishery Bulletin, 109(3):??, 2011. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
changes-catch-rates-and-length-and-age-maturity-not-growth-
estuarine-plotosid-cnidoglanis;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2011/1093/chuwen.pdf.

Harding:2011:RSP

[5017] Jeffrey A. Harding, Arnold J. Ammann, and R. Bruce MacFarlane. Regional and seasonal patterns of epipelagic fish assemblages from the central California Current. Fishery Bulletin, 109 (3):261–281, 2011. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
regional-and-seasonal-patterns-epipelagic-fish-assemblages-
central-california-current;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2011/1093/harding.pdf.

Laurel:2011:ETH

[5018] Benjamin J. Laurel and Deborah M. Blood. The effects of temperature on hatching and survival of northern rock sole larvae (Lepidopsetta polyxystra). Fishery Bulletin, 109(3):??, 2011. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/effects-temperature-hatching-and-
survival-northern-rock-sole-larvae-lepidopsetta-polyxystra;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2011/1093/laurel.pdf.

Stephen:2011:CLH

[5019] Jessica A. Stephen, Patrick J. Harris, and Marcel J. M. Reichert. Comparison of life history parameters for landed and discarded fish captured off the southeastern United States. Fishery Bulletin, 109(3):292–304, 2011. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/comparison-life-history-parameters-landed-
and-discarded-fish-captured-southeastern-united;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/2011/1093/
stephen.pdf.

Macieira:2011:DPT

[5020] Raphael M. Macieira and Jean-Christophe Joyeux. Distribution patterns of tidepool fishes on a tropical flat reef. Fishery Bulletin, 109(3): ??, 2011. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/distribution-
patterns-tidepool-fishes-tropical-flat-reef;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/2011/1093/macieira.
pdf.

Wilson:2011:RVA

[5021] Matthew T. Wilson, Andre Buchheister, and Christina Jump. Regional variation in the annual feeding cycle of juvenile walleye pollock (Theragra chalcogramma) in the western Gulf of Alaska. Fishery Bulletin, 109(3): 316–326, 2011. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/regional-
variation-annual-feeding-cycle-juvenile-walleye-pollock-theragra-
chalcogramma;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2011/1093/wilson.pdf.

Locascio:2011:DST

[5022] James V. Locascio and David A. Mann. Diel and seasonal timing of sound production by black drum (Pogonias cromis). Fishery Bulletin, 109(3):??, 2011. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/diel-and-seasonal-timing-sound-production-
black-drum-pogonias-cromis;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2011/1093/locascio.pdf.

Musyl:2011:PSV

[5023] Michael K. Musyl, Richard W. Brill, Daniel S. Curran, Nuno M. Fragoso, Lianne M. McNaughton, Anders Nielsen, Bert S. Kikkawa, and Christopher D. Moyes. Postrelease survival, vertical and horizontal movements, and thermal habitats of five species of pelagic sharks in the central Pacific Ocean. Fishery Bulletin, 109(4):??, 2011. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/postrelease-survival-
vertical-and-horizontal-movements-and-thermal-habitats-five-
species;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2011/1094/1094musyl.pdf.

Riemer:2011:FHS

[5024] Susan D. Riemer, Bryan E. Wright, and Robin F. Brown. Food habits of Steller sea lions (Eumetopias jubatus) off Oregon and northern California, 1986–2007. Fishery Bulletin, 109(4):369–381, 2011. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/food-habits-steller-
sea-lions-eumetopias-jubatus-oregon-and-northern-california-
1986-2007;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2011/1094/1094riemer.pdf.

Wildes:2011:GVB

[5025] Sharon L. Wildes, Johanna J. Vollenweider, Hanhvan T. Nguyen, and Jeffrey R. Guyon. Genetic variation between outer-coastal and fjord populations of Pacific herring (Clupea pallasii) in the eastern Gulf of Alaska. Fishery Bulletin, 109(4):??, 2011. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/genetic-variation-between-outer-
coastal-and-fjord-populations-pacific-herring-clupea;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2011/1094/
1094wildes.pdf.

OConnell:2011:APM

[5026] Craig P. O’Connell, Daniel C. Abel, Eric M. Stroud, and Patrick H. Rice. Analysis of permanent magnets as elasmobranch bycatch reduction devices in hook-and-line and longline trials. Fishery Bulletin, 109(4):??, 2011. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/analysis-permanent-
magnets-elasmobranch-bycatch-reduction-devices-hook-and-line-
and;   https://spo.nmfs.noaa.gov/sites/default/files/1094oconnell.
pdf.

Javor:2011:OMP

[5027] Barbara Javor, Nancy Lo, and Russ Vetter. Otolith morphometrics and population structure of Pacific sardine (Sardinops sagax) along the west coast of North America. Fishery Bulletin, 109(4):402– 415, 2011. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/otolith-
morphometrics-and-population-structure-pacific-sardine-sardinops-
sagax-along-west;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2011/1094/1094javor.pdf.

Carson:2011:PSL

[5028] Evan W. Carson, Eric Saillant, Mark A. Renshaw, Nancie J. Cummings, and John R. Gold. Population structure, long-term connectivity, and effective size of mutton snapper (Lutjanus analis) in the Caribbean Sea and Florida Keys. Fishery Bulletin, 109(4):??, 2011. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/population-
structure-long-term-connectivity-and-effective-size-mutton-
snapper-lutjanus;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2011/1094/1094carson.pdf.

Muto:2011:GMD

[5029] Nozomu Muto, Yoshiaki Kai, and Tetsuji Nakabo. Genetic and morphological differences between Sebastes vulpes and S. zonatus (Teleostei: Scorpaeniformes: Scorpaenidae). Fishery Bulletin, 109(4): ??, 2011. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/genetic-and-
morphological-differences-between-sebastes-vulpes-and-s-zonatus-
teleostei;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2011/1094/1094muto.pdf.

Pearcy:2011:ODA

[5030] William G. Pearcy and Joseph P. Fisher. Ocean distribution of the American shad (Alosa sapidissima) along the Pacific coast of North America. Fishery Bulletin, 109(4):??, 2011. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/ocean-distribution-american-shad-alosa-sapidissima-
along-pacific-coast-north-america;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2011/1094/1094pearcy.pdf.

Andrews:2011:BRT

[5031] Allen H. Andrews, Lisa J. Natanson, Lisa A. Kerr, George H. Burgess, and Gregor M. Cailliet. Bomb radiocarbon and tag-recapture dating of sandbar shark (Carcharhinus plumbeus). Fishery Bulletin, 109(4):??, 2011. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/bomb-radiocarbon-and-
tag-recapture-dating-sandbar-shark-carcharhinus-plumbeus;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/2011/
1094/1094andrews.pdf.

Marancik:2012:STD

[5032] Katrin E. Marancik, David E. Richardson, Joanne Lyczkowski-Schultz, Robert K. Cowen, and Małgorzata Konieczna. Spatial and temporal distribution of grouper larvae (Serranidae: Epinephelinae: Epinephelini) in the Gulf of Mexico and Straits of Florida. Fishery Bulletin, 110(1):1–20, 2012. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/spatial-and-temporal-distribution-grouper-

larvae-serranidae-epinephelinae-epinephelini-gulf;   https:
//spo.nmfs.noaa.gov/sites/default/files/marancik.pdf.

Rooper:2012:CHB

[5033] Christopher N. Rooper and Michael H. Martin. Comparison of habitatbased indices of abundance with fishery-independent biomass estimates from bottom trawl surveys. Fishery Bulletin, 110(1):??, 2012. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/comparison-habitat-based-
indices-abundance-fishery-independent-biomass-estimates-bottom;
 https://spo.nmfs.noaa.gov/sites/default/files/rooper.pdf.

Rodgveller:2012:EMA

[5034] Cara J. Rodgveller, Chris R. Lunsford, and Jeffrey T. Fujioka. Effects of maternal age and size on embryonic energy reserves, developmental timing, and fecundity in quillback rockfish (Sebastes maliger). Fishery Bulletin, 110(1):35–45, 2012. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/effects-maternal-age-and-size-embryonic-energy-
reserves-developmental-timing-and-fecundity;   https://spo.
nmfs.noaa.gov/sites/default/files/rodgveller_0.pdf.

Loher:2012:LSE

[5035] Timothy Loher and Jessica C. Hobden. Length and sex effects on the spatial structure of catches of Pacific halibut (Hippoglossus stenolepis) on longline gear. Fishery Bulletin, 110(1):??, 2012. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/length-and-sex-effects-spatial-
structure-catches-pacific-halibut-hippoglossus-stenolepis;
https://spo.nmfs.noaa.gov/sites/default/files/loher.pdf.

Demer:2012:PCS

[5036] David A. Demer, Juan P. Zwolinski, Kyle A. Byers, George R. Cutter, Josiah S. Renfree, Thomas S. Sessions, and Beverly J. Macewicz. Prediction and confirmation of seasonal migration of Pacific sardine (Sardinops sagax) in the California Current ecosystem. Fishery Bulletin, 110(1):52–70, 2012. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/prediction-and-confirmation-seasonal-migration-
pacific-sardine-sardinops-sagax-california;   https://spo.nmfs.
noaa.gov/sites/default/files/demer.pdf.

Martinho:2012:CFE

[5037] Filipe Martinho, Cátia Sá, Joana Falcão, Henrique Nogueira Cabral, and Miguel Ângelo Pardal. Comparative feeding ecology of two elasmobranch species, Squalus blainville and Scyliorhinus canicula, off the coast of Portugal. Fishery Bulletin, 110(1):71–84, 2012. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/comparative-feeding-ecology-two-
elasmobranch-species-squalus-blainville-and-scyliorhinus;
https://spo.nmfs.noaa.gov/sites/default/files/martinho.pdf.

Weber:2012:EEC

[5038] Edward D. Weber and Sam McClatchie. Effect of environmental conditions on the distribution of Pacific mackerel (Scomber japonicus) larvae in the California Current system. Fishery Bulletin, 110(1):??, 2012. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/effect-environmental-
conditions-distribution-pacific-mackerel-scomber-japonicus-
larvae;   https://spo.nmfs.noaa.gov/sites/default/files/weber.
pdf.

Arendt:2012:CRD

[5039] Michael D. Arendt, Jeffrey A. Schwenter, Albert L. Segars, Julia I. Byrd, Philip P. Maier, J. David Whitaker, David W. Owens, Gaëlle Blanvillain, Joseph M. Quattro, and Mark A. Roberts. Catch rates and demographics of loggerhead sea turtles (Caretta caretta) captured from the Charleston, South Carolina, shipping channel during the period of mandatory use of turtle excluder devices (TEDs). Fishery Bulletin, 110 (1):??, 2012. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/catch-rates-
and-demographics-loggerhead-sea-turtles-caretta-caretta-captured-
charleston;   https://spo.nmfs.noaa.gov/sites/default/files/
arendt.pdf.

Zwolinski:2012:DAP

[5040] Juan P. Zwolinski, David A. Demer, Kyle A. Byers, George R. Cutter, Josiah S. Renfree, Thomas S. Sessions, and Beverly J. Macewicz. Distributions and abundances of Pacific sardine (Sardinops sagax) and other pelagic fishes in the California Current ecosystem during spring 2006, 2008, and 2010, estimated from acoustic—trawl surveys. Fishery Bulletin, 110(1):??, 2012. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/distributions-and-abundances-pacific-

sardine-sardinops-sagax-and-other-pelagic-fishes;   https:/
/spo.nmfs.noaa.gov/sites/default/files/zwolinski.pdf.

Lough:2012:LSR

[5041] R. Gregory Lough and Loretta O’Brien. Life-stage recruitment models for Atlantic cod (Gadus morhua) and haddock (Melanogrammus aeglefinus) on Georges Bank. Fishery Bulletin, 110(1):??, 2012. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/life-stage-
recruitment-models-atlantic-cod-gadus-morhua-and-haddock-melanogrammus;
 https://spo.nmfs.noaa.gov/sites/default/files/lough.pdf.

Matkin:2012:CAR

[5042] Craig O. Matkin, John W. Durban, Eva L. Saulitis, Russel D. Andrews, Janice M. Straley, Dena R. Matkin, and Graeme M. Ellis. Contrasting abundance and residency patterns of two sympatric populations of transient killer whales (Orcinus orca) in the northern Gulf of Alaska. Fishery Bulletin, 110(2):??, 2012. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
contrasting-abundance-and-residency-patterns-two-sympatric-
populations-transient-killer;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2012/1102/matkin.pdf.

Rudershausen:2012:CBC

[5043] Paul J. Rudershausen, Jeffrey A. Buckel, Greg E. Bolton, Randy W. Gregory, Tyler W. Averett, and Paul B. Conn. A comparison between circle hook and J hook performance in the dolphinfish, yellowfin tuna, and wahoo troll fishery off the coast of North Carolina. Fishery Bulletin, 110(2):??, 2012. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/comparison-between-circle-hook-and-j-hook-
performance-dolphinfish-yellowfin-tuna-and-wahoo;   https:/
/spo.nmfs.noaa.gov/sites/default/files/rudershausen.pdf.

Able:2012:DMH

[5044] Kenneth W. Able, Thomas M. Grothues, Jason T. Turnure, Donald M. Byrne, and Paul Clerkin. Distribution, movements, and habitat use of small striped bass (Morone saxatilis) across multiple spatial scales. Fishery Bulletin, 110(2):??, 2012. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
distribution-movements-and-habitat-use-small-striped-bass-morone-
saxatilis-across-0;   https://spo.nmfs.noaa.gov/sites/default/
files/able.pdf.

Tinus:2012:PPL

[5045] Craig A. Tinus. Prey preference of lingcod (Ophiodon elongatus), a top marine predator: implications for ecosystem-based fisheries management. Fishery Bulletin, 110(2):??, 2012. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/prey-preference-lingcod-ophiodon-
elongatus-top-marine-predator-implications-ecosystem-based;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2012/1102/tinus.pdf.

Keller:2012:VEN

[5046] Aimee A. Keller, John R. Wallace, Beth H. Horness, Owen S. Hamel, and Ian J. Stewart. Variations in eastern North Pacific demersal fish biomass based on the U.S. west coast groundfish bottom trawl survey (2003– 2010). Fishery Bulletin, 110(2):??, 2012. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/variations-eastern-north-pacific-demersal-fish-
biomass-based-us-west-coast-groundfish;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2012/1102/keller.pdf.

Ohnishi:2012:NBG

[5047] Shuhei Ohnishi, Takashi Yamakawa, Hiroshi Okamura, and Tatsuro Akamine. A note on the von Bertalanffy growth function concerning the allocation of surplus energy to reproduction. Fishery Bulletin, 110(2):223–229, 2012. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
note-von-bertalanffy-growth-function-concerning-allocation-
surplus-energy-reproduction;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2012/1102/ohnishi.pdf.

Baremore:2012:GSJ

[5048] Ivy E. Baremore, Dana M. Bethea, and Kate I. Andrews. Gillnet selectivity for juvenile blacktip sharks (Carcharhinus limbatus). Fishery Bulletin, 110(2):230–241, 2012. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
gillnet-selectivity-juvenile-blacktip-sharks-carcharhinus-limbatus;
 https://spo.nmfs.noaa.gov/sites/default/files/baremore.pdf.

Beacham:2012:PSP

[5049] Terry D. Beacham, Brenda McIntosh, Cathy MacConnachie, Brian Spilsted, and Bruce A. White. Population structure of pink salmon (Oncorhynchus gorbuscha) in British Columbia and Washington, determined with microsatellites. Fishery Bulletin, 110(2):??, 2012. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/population-structure-pink-salmon-
oncorhynchus-gorbuscha-british-columbia-and-washington;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2012/1102/
beacham.pdf.

Caldarone:2012:EBI

[5050] Elaine M. Caldarone, Sharon A. MacLean, and Beth Sharack. Evaluation of bioelectrical impedance analysis and Fulton’s condition factor as nonlethal techniques for estimating short-term responses in postsmolt Atlantic salmon (Salmo salar) to food availability. Fishery Bulletin, 110(2):257–270, 2012. CODEN FSY-

BAY.  ISSN 0090-0656 (print), 1937-4518 (electronic).  URL

https://spo.nmfs.noaa.gov/content/evaluation-bioelectrical-
impedance-analysis-and-fultons-condition-factor-nonlethal;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/2012/
1102/caldarone.pdf.

Stachura:2012:EDM

[5051] Megan M. Stachura, Chris R. Lunsford, Cara J. Rodgveller, and Jonathan Heifetz. Estimation of discard mortality of sablefish (Anoplopoma fimbria ) in Alaska longline fisheries. Fishery Bulletin, 110(2):271–279, 2012. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
estimation-discard-mortality-sablefish-anoplopoma-fimbria-alaska-
longline-fisheries;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2012/1102/stachura.pdf.

Powers:2012:ACD

[5052] Sean P. Powers, Crystal L. Hightower, J. Marcus Drymon, and Matthew W. Johnson. Age composition and distribution of red drum (Sciaenops ocellatus) in offshore waters of the north central Gulf of Mexico: an evaluation of a stock under a federal harvest moratorium. Fishery Bulletin, 110(3):283–292, 2012. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/age-composition-and-distribution-
red-drum-sciaenops-ocellatus-offshore-waters-north-central;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2012/1103/powers.pdf.

Boldt:2012:FIC

[5053] Jennifer L. Boldt, Troy W. Buckley, Christopher N. Rooper, and Kerim Aydin. Factors influencing cannibalism and abundance of walleye pollock (Theragra chalcogramma) on the eastern Bering Sea shelf, 1982–2006. Fishery Bulletin, 110(3):??, 2012. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/factors-influencing-cannibalism-
and-abundance-walleye-pollock-theragra-chalcogramma-eastern;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2012/1103/boldt.pdf.

Tagliafico:2012:ERS

[5054] Alejandro Tagliafico, Néstor Rago, Salomé Rangel, and Jeremy Mendoza. Exploitation and reproduction of the spotted eagle ray (Aetobatus narinari) in the Los Frailes Archipelago, Venezuela. Fishery Bulletin, 110(3): ??, 2012. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/exploitation-
and-reproduction-spotted-eagle-ray-aetobatus-narinari-los-frailes-
archipelago;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2012/1103/tagliafico.pdf.

Rooper:2012:ESS

[5055] Christopher N. Rooper, Michael H. Martin, John L. Butler, Darin T. Jones, and Mark Zimmermann. Estimating species and size composition of rockfishes to verify targets in acoustic surveys of untrawlable areas. Fishery Bulletin, 110(3):??, 2012. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/estimating-species-and-size-composition-rockfishes-
verify-targets-acoustic-surveys;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2012/1103/rooper.pdf.

Jones:2012:ERA

[5056] Darin T. Jones, Christopher D. Wilson, Alex De Robertis, Christopher N. Rooper, Thomas C. Weber, and John L. Butler. Evaluation of rockfish abundance in untrawlable habitat: combining acoustic and complementary sampling tools. Fishery Bulletin, 110(3):??, 2012. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/evaluation-rockfish-
abundance-untrawlable-habitat-combining-acoustic-and-complementary;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2012/1103/jones.pdf.

Barlow:2012:EME

[5057] Paige F. Barlow and Jim Berkson. Evaluating methods for estimating rare events with zero-heavy data: a simulation model estimating sea turtle bycatch in the pelagic longline fishery. Fishery Bulletin, 110(3):??, 2012. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/evaluating-methods-estimating-rare-events-
zero-heavy-data-simulation-model-estimating-sea;   https://
spo.nmfs.noaa.gov/sites/default/files/barlow.pdf.

Echave:2012:ICG

[5058] Katy B. Echave, Dana H. Hanselman, Milo D. Adkison, and Michael F. Sigler. Interdecadal change in growth of sablefish (Anoplopoma fimbria) in the northeast Pacific Ocean. Fishery Bulletin, 110(3):??, 2012. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/interdecadal-change-
growth-sablefish-anoplopoma-fimbria-northeast-pacific-ocean;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2012/1103/echave.pdf.

Hanselman:2012:AAT

[5059] Dana H. Hanselman, Paul D. Spencer, Denise R. McKelvey, and Michael H. Martin. Application of an acoustic-trawl survey design to improve estimates of rockfish biomass. Fishery Bulletin, 110(4):379–396, 2012. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
application-acoustic-trawl-survey-design-improve-estimates-
rockfish-biomass;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2012/1104/hanselman.pdf.

Lefebvre:2012:ISC

[5060] Lyndsey S. Lefebvre and Michael R. Denson. Inshore spawning of cobia (Rachycentron canadum) in South Carolina. Fishery Bulletin, 110(4):??, 2012. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
inshore-spawning-cobia-rachycentron-canadum-south-carolina;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2012/1104/lefebvre.pdf.

Fitzhugh:2012:RSA

[5061] Gary R. Fitzhugh, Kyle W. Shertzer, G. Todd Kellison, and David M. Wyanski. Review of size- and age-dependence in batch spawning: implications for stock assessment of fish species exhibiting indeterminate fecundity. Fishery Bulletin, 110(4):??, 2012. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/review-size-and-age-dependence-batch-
spawning-implications-stock-assessment-fish-species;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2012/1104/
fitzhugh.pdf.

Weitkamp:2012:SIV

[5062] Laurie A. Weitkamp, Paul J. Bentley, and Marisa N. C. Litz. Seasonal and interannual variation in juvenile salmonids and associated fish assemblage in open waters of the lower Columbia River estuary. Fishery Bulletin, 110(4):??, 2012. CODEN FSY-

BAY.  ISSN 0090-0656 (print), 1937-4518 (electronic).  URL

https://spo.nmfs.noaa.gov/content/seasonal-and-interannual-
variation-juvenile-salmonids-and-associated-fish-assemblage-
open;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2012/1104/weitkamp.pdf.

Ogburn:2012:ESI

[5063] Matthew B. Ogburn and Richard B. Forward, Jr. Effect of sampling interval on estimates of larval supply. Fishery Bulletin, 110(4):??, 2012. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/effect-sampling-interval-
estimates-larval-supply;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2012/1104/ogburn.pdf.

Syc:2012:CSA

[5064] Tara S. Syc and Stephen T. Szedlmayer. A comparison of size and age of red snapper (Lutjanus campechanus) with the age of artificial reefs in the northern Gulf of Mexico. Fishery Bulletin, 110(4):??, 2012. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/comparison-size-and-
age-red-snapper-lutjanus-campechanus-age-artificial-reefs-northern-
gulf;   https://spo.nmfs.noaa.gov/sites/default/files/syc.pdf.

Arendt:2012:TTI

[5065] Michael D. Arendt, Jeffrey A. Schwenter, Jessica Boynton, Albert L. Segars, Julia I. Byrd, J. David Whitaker, and Lindsey Parker. Temporal trends (2000–2011) and influences on fishery-independent catch rates for loggerhead sea turtles (Caretta caretta) at an important coastal foraging region in the southeastern United States. Fishery Bulletin, 110(4):??, 2012. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/temporal-trends-2000-
2011-and-influences-fishery-independent-catch-rates-loggerhead-
sea;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2012/1104/arendt.pdf.

Cowen:2013:ESI

[5066] Robert K. Cowen, Adam T. Greer, Cedric M. Guigand, Jonathan A. Hare, David E. Richardson, and Harvey J. Walsh. Evaluation of the In Situ Ichthyoplankton Imaging System (ISIIS): comparison with the traditional (bongo net) sampler. Fishery Bulletin, 111(1):1–12, 2013. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/evaluation-situ-ichthyoplankton-
imaging-system-isiis-comparison-traditional-bongo-net;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2013/cowen.
pdf.

Bromaghin:2013:NID

[5067] Jeffrey F. Bromaghin, Monique M. Lance, Elizabeth W. Elliott, Steven J. Jeffries, Alejandro Acevedo-Gutiérrez, and John M. Kennish. New insights into the diets of harbor seals (Phoca vitulina) in the Salish Sea revealed by analysis of fatty acid signatures. Fishery Bulletin, 111(1):13–26, 2013. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/new-insights-diets-harbor-seals-phoca-vitulina-
salish-sea-revealed-analysis-fatty-acid;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2013/bromaghin.pdf.

Howard:2013:FCH

[5068] Sarah M. S. Howard, Monique M. Lance, Steven J. Jeffries, and Alejandro Acevedo-Gutiérrez. Fish consumption by harbor seals (Phoca vitulina) in the San Juan Islands, Washington. Fishery Bulletin, 111 (1):27–41, 2013. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/fish-
consumption-harbor-seals-phoca-vitulina-san-juan-islands-washington;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2013/howard.pdf.

Rose:2013:QRU

[5069] Craig S. Rose, Carwyn F. Hammond, Allan W. Stoner, J. Eric Munk, and John R. Gauvin. Quantification and reduction of unobserved mortality rates for snow, southern Tanner, and red king crabs (Chionoecetes opilio, C. bairdi, and Paralithodes camtschaticus) after encounters with trawls on the seafloor. Fishery Bulletin, 111(1):42–53, 2013.

CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/quantification-and-
reduction-unobserved-mortality-rates-snow-southern-tanner-and-
red-king;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2013/rose.pdf.

Laidig:2013:RFT

[5070] Thomas E. Laidig, Lisa M. Krigsman, and Mary M. Yoklavich. Reactions of fishes to two underwater survey tools, a manned submersible and a remotely operated vehicle. Fishery Bulletin, 111(1):54–67, 2013. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/reactions-fishes-two-underwater-
survey-tools-manned-submersible-and-remotely-operated;   https:
//spo.nmfs.noaa.gov/sites/default/files/laidig.pdf.

Weber:2013:SCT

[5071] Thomas C. Weber, Christopher Rooper, John Butler, Darin Jones, and Chris Wilson. Seabed classification for trawlability determined with a multibeam echo sounder on Snakehead Bank in the Gulf of Alaska. Fishery Bulletin, 111(1):68–77, 2013. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/seabed-classification-trawlability-determined-multibeam-
echo-sounder-snakehead-bank-gulf;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2013/weber.pdf.

Staaf:2013:DOP

[5072] Danna J. Staaf, Jessica V. Redfern, William F. Gilly, William Watson, and Lisa T. Ballance. Distribution of ommastrephid paralarvae in the eastern tropical Pacific. Fishery Bulletin, 111(1):78–89, 2013. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/distribution-
ommastrephid-paralarvae-eastern-tropical-pacific;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/2013/staaf.
pdf.

Burchard:2013:SOH

[5073] Katie A. Burchard, Francis Juanes, Rodney A. Rountree, and William A. Roumillat. Staging ovaries of haddock (Melanogrammus aeglefinus): implications for maturity indices and field sampling practices. Fishery Bulletin, 111(1):90–106, 2013. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
staging-ovaries-haddock-melanogrammus-aeglefinus-implications-

maturity-indices-and-field;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2013/burchard.pdf.

Fulford:2013:QIV

[5074] Richard S. Fulford and Kevin Dillon. Quantifying intrapopulation variability instable isotope data for spotted seatrout (Cynoscion nebulosus). Fishery Bulletin, 111(2):111–121, 2013. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/quantifying-intrapopulation-variability-instable-
isotope-data-spotted-seatrout-cynoscion;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2013/fulford.pdf.

Frable:2013:NSW

[5075] Benjamin W. Frable, Carole C. Baldwin, Brendan M. Luther, and Lee A. Weigt. A new species of western Atlantic lizardfish (Teleostei: Synodontidae: Synodus) and resurrection of Synodus bondi Fowler, 1939, as a valid species from the Caribbean with redescriptions of S. bondi, S. foetens (Linnaeus, 1766), and S. intermedius (Agassiz). Fishery Bulletin, 111 (2):122–146, 2013. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/new-
species-western-atlantic-lizardfish-teleostei-synodontidae-
synodus-and-resurrection;   https://spo.nmfs.noaa.gov/sites/
default/files/frable.pdf.

Wells:2013:AVJ

[5076] R. J. David Wells, Susan E. Smith, Suzanne Kohin, Ellen Freund, Natalie Spear, and Darlene A. Ramon. Age validation of juvenile shortfin mako (Isurus oxyrinchus) tagged and marked with oxytetracycline off southern California. Fishery Bulletin, 111(2):147–160, 2013. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/age-validation-juvenile-shortfin-
mako-isurus-oxyrinchus-tagged-and-marked-oxytetracycline;
https://spo.nmfs.noaa.gov/sites/default/files/wells.pdf.

McBride:2013:CSA

[5077] Richard S. McBride, Tiffany E. Vidal, and Steven X. Cadrin. Changes in size and age at maturity of the northern stock of Tilefish (Lopholatilus chamaeleonticeps) after a period of overfishing. Fishery Bulletin, 111(2):161–174, 2013. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/changes-size-and-age-maturity-thenorthern-
stock-tilefish-lopholatiluschamaeleonticeps-after;   https:

//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2013/mcbride.
pdf.

Syamsuddin:2013:ENN

[5078] Mega L. Syamsuddin, Sei-Ichi Saitoh, Toru Hirawake, Samsul Bachri, and Agung B. Harto. Effects of El Niño–Southern Oscillation events on catches of bigeye tuna (Thunnus obesus) in the eastern Indian Ocean off Java. Fishery Bulletin, 111(2):175–188, 2013. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/effects-el-ni%C3%B1o%
E2%80%93southern-oscillation-events-catches-bigeye-tuna-thunnus-
obesus-eastern;   https://spo.nmfs.noaa.gov/sites/default/files/
syamsuddin.pdf.

Capossela:2013:MWE

[5079] Karen M. Capossela, Mary C. Fabrizio, and Richard W. Brill. Migratory and within-estuary behaviors of adult summer flounder (Paralichthys dentatus) in a lagoon system of the southern Mid-Atlantic Bight. Fishery Bulletin, 111(2):189–201, 2013. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/migratory-and-within-
estuary-behaviors-adult-summer-flounder-paralichthys-dentatus-
lagoon;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2013/capossela.pdf.

Essington:2013:SED

[5080] Timothy E. Essington, Kathryn Dodd, and Thomas P. Quinn. Shifts in the estuarine demersal fish community after a fishery closure in Puget Sound, Washington. Fishery Bulletin, 111(3):205–217, 2013. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/shifts-estuarine-demersal-fish-
community-after-fishery-closure-puget-sound-washington;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2013/essington.
pdf.

Bacheler:2013:IST

[5081] Nathan M. Bacheler, Valerio Bartolino, and Marcel J. M. Reichert. Influence of soak time and fish accumulation on catches of reef fishes in a multispecies trap survey. Fishery Bulletin, 111(3):218–232, 2013. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/influence-soak-time-
and-fish-accumulation-catches-reef-fishes-multispecies-trap-

survey;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2013/bacheler.pdf.

Moiseev:2013:EPF

[5082] Sergey I. Moiseev, Svetlana A. Moiseeva, Tatyana V. Ryazanova, and Anna M. Lapteva. Effects of pot fishing on the physical conditionof snow crab (Chionoecetes opilio) and southern Tanner crab (Chionoecetes bairdi). Fishery Bulletin, 111(3):233–251, 2013. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/effects-pot-
fishing-physical-conditionof-snow-crab-chionoecetes-opilio-
and-southern-tanner;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2013/moiseev.pdf.

Zudaire:2013:RPY

[5083] Iker Zudaire, Hilario Murua, Maitane Grande, and Nathalie Bodin. Reproductive potential of yellowfin tuna (Thunnus albacares) in the western Indian Ocean. Fishery Bulletin, 111(3):252–264, 2013. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/reproductive-potential-yellowfin-
tuna-thunnus-albacares-western-indian-ocean;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/2013/zudaire.
pdf.

Peemoeller:2013:ASS

[5084] Bhae-Jin Peemoeller and Bradley G. Stevens. Age, size, and sexual maturity of channeled whelk (Busycotypus canaliculatus) in Buzzards Bay, Massachusetts. Fishery Bulletin, 111(3):265–278, 2013. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/age-size-and-sexual-
maturity-channeledwhelk-busycotypus-canaliculatus-buzzards-
bay;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2013/peemoeller.pdf.

Wood:2013:MMY

[5085] Anthony D. Wood and Steven X. Cadrin. Mortality and movement of yellowtail flounder (Limanda ferruginea) tagged off New England. Fishery Bulletin, 111(3):279–287, 2013. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/mortality-and-movement-yellowtail-flounder-limanda-
ferruginea-tagged-new-england;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/2013/wood.pdf.

Misa:2013:ESH

[5086] William F. X. E. Misa, Jeffrey C. Drazen, Christopher D. Kelley, and Virginia N. Moriwake. Establishing species-habitat associations for 4 eteline snappers with the use of a baited stereo-video camera system. Fishery Bulletin, 111(4):293–308, 2013. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/establishing-species-habitat-associations-
4-eteline-snappers-use-baited-stereo-video-camera;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2013/misa.
pdf.

Cappo:2013:JRC

[5087] Mike Cappo, Ross J. Marriott, and Stephen J. Newman. James’s rule and causes and consequences of a latitudinal cline in the demography of John’s snapper (Lutjanus johnii) in coastal waters of Australia. Fishery Bulletin, 111(4):309–324, 2013. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/jamess-rule-and-causes-and-consequences-latitudinal-
cline-demography-johns-snapper-lutjanus;   https://spo.nmfs.
noaa.gov/sites/default/files/cappo.pdf.

Macchi:2013:MEF

[5088] Gustavo J. Macchi, Ezequiel Leonarduzzi, Marina V. Diaz, Marta Renzi, and Karina Rodrigues. Maternal effects on fecundity and egg quality of the Patagonian stock of Argentine hake (Merluccius hubbsi). Fishery Bulletin, 111(4):325–336, 2013. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/maternal-effects-fecundity-and-
egg-qualityof-patagonian-stock-argentine-hake-merluccius;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/2013/
macchi.pdf.

Porter:2013:UMM

[5089] Steven M. Porter and Kevin M. Bailey. Using measurements of muscle cell nuclear RNA with flow cytometry to improve assessment of larval condition of walleye pollock (Gadus chalcogrammus). Fishery Bulletin, 111(4):337–351, 2013. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/using-measurements-muscle-cell-
nuclear-rna-flow-cytometry-improve-assessment-larval;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2013/porter.
pdf.

Wuenschel:2013:HDO

[5090] Mark J. Wuenschel, Kenneth W. Able, James M. Vasslides, and Donald M. Byrne. Habitat and diet overlap of 4 piscivorous fishes: variation on the inner continental shelf off New Jersey. Fishery Bulletin, 111(4): 352–369, 2013. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/habitat-
and-diet-overlap-4-piscivorous-fishes-variation-inner-continental-
shelf-new-jersey;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2013/wuenschel.pdf.

Drymon:2013:MAF

[5091] J. Marcus Drymon, Laure Carassou, Sean P. Powers, Mark Grace, John Dindo, and Brian Dzwonkowski. Multiscale analysis of factors that affect the distribution of sharks throughout the northern Gulf of Mexico. Fishery Bulletin, 111(4):370–380, 2013. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/multiscale-analysis-factors-affect-thedistribution-
sharks-throughout-northern-gulf-mexico;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2013/drymon.pdf.

Wenzel:2013:FHS

[5092] Frederick W. Wenzel, Pamela T. Polloni, James E. Craddock, Damon P. Gannon, John R. Nicolas, Andrew J. Read, and Patricia E. Rosel. Food habits of Sowerby’s beaked whales (Mesoplodon bidens) taken in the pelagic drift gillnet fishery of the western North Atlantic. Fishery Bulletin, 111(4):381–389, 2013. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
food-habits-sowerbys-beaked-whales-mesoplodon-bidens-taken-
pelagic-drift-gillnet-fishery;   https://spo.nmfs.noaa.gov/
sites/default/files/wenzel.pdf.

Colmenero:2013:RBB

[5093] Ana I. Colmenero, Víctor M. Tuset, Laura Recasens, and Pilar Sánchez. Reproductive biology of black anglerfish (Lophius budegassa) in the northwestern Mediterranean Sea. Fishery Bulletin, 111(4):390–401, 2013. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/reproductive-biology-
black-anglerfish-lophius-budegassa-northwestern-mediterranean-
sea;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2013/colmenero.pdf.

Byrd:2014:SIM

[5094] Barbie L. Byrd, Aleta A. Hohn, Gretchen N. Lovewell, Karen M. Altman, Susan G. Barco, Ari Friedlaender, Craig A. Harms, William A. McLellan, Kathleen T. Moore, Patricia E. Rosel, and Victoria G. Thayer. Strandings as indicators of marine mammal biodiversity and human interactions off the coast of North Carolina. Fishery Bulletin, 112(1):1–23, 2014. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
strandings-indicators-marine-mammal-biodiversity-and-human-
interactions-coast-north;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/2014/byrd.pdf.

Darden:2014:PGC

[5095] Tanya L. Darden, Matthew J. Walker, Karl Brenkert, Justin R. Yost, and Michael R. Denson. Population genetics of cobia (Rachycentron canadum): implications for fishery management along the coast of the southeastern United States. Fishery Bulletin, 112(1):24–35, 2014. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/population-genetics-
cobia-rachycentron-canadum-implications-fishery-management-
along-coast;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2014/darden.pdf.

Heery:2014:COB

[5096] Eliza Heery and Jason M. Cope. Co-occurrence of bycatch and target species in the groundfish demersal trawl fishery of the U.S. west coast; with special consideration of rebuilding stocks. Fishery Bulletin, 112(1): 36–48, 2014. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/co-occurrence-
bycatch-and-target-species-groundfish-demersal-trawl-fishery-
us-west-coast;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2014/heery.pdf.

Munyandorero:2014:SCE

[5097] Joseph Munyandorero. In search of climate effects on Atlantic croaker (Micropogonias undulatus) stock off the U.S. Atlantic coast with Bayesian state-space biomass dynamic models. Fishery Bulletin, 112 (1):49–70, 2014. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/search-
climate-effects-atlantic-croaker-micropogonias-undulatus-stock-
us-atlantic-coast;   https://spo.nmfs.noaa.gov/sites/default/
files/munyandorero_0.pdf.

Anderson:2014:LGS

[5098] Joel D. Anderson and William J. Karel. Limited genetic structure of Gulf menhaden (Brevoortia patronus), as revealed by microsatellite markers developed for the genus Brevoortia (Clupeidae). Fishery Bulletin, 112(1):71–81, 2014. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/limited-genetic-structure-gulf-menhaden-brevoortia-
patronus-revealed-microsatellite-markers;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2014/anderson.pdf.

Fabrizio:2014:HRS

[5099] Mary C. Fabrizio, John P. Manderson, and Jeffrey P. Pessutti. Home range and seasonal movements of black sea bass (Centropristis striata) during their inshore residency at a reef in the Mid-Atlantic Bight. Fishery Bulletin, 112(1):82–97, 2014. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
home-range-and-seasonal-movements-black-sea-bass-centropristis-
striata-during-their-inshore;   https://spo.nmfs.noaa.gov/sites/
default/files/fabrizio_0.pdf.

Long:2014:EGF

[5100] W. Christopher Long, Peter A. Cummiskey, and J. Eric Munk. Effects of ghost fishing on the population of red king crab (Paralithodes camtschaticus) in Womens Bay, Kodiak Island, Alaska. Fishery Bulletin, 112(2–3):101–111, 2014. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/effects-ghost-fishing-population-red-king-crab-
paralithodes-camtschaticus-womens-bay-0;   https://spo.nmfs.
noaa.gov/content/effects-ghost-fishing-population-red-king-
crab-paralithodes-camtschaticus-womens-bay-kodiak;   https:
//spo.nmfs.noaa.gov/sites/default/files/long_0.pdf;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2014/long.
pdf.

Clemento:2014:ESN

[5101] Anthony J. Clemento, Eric D. Crandall, John Carlos Garza, and Eric C. Anderson. Evaluation of a single nucleotide polymorphism baseline for genetic stock identification of Chinook salmon (Oncorhynchus tshawytscha) in the California Current large marine ecosystem. Fishery Bulletin, 112(2–3):112–130, 2014. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/evaluation-single-nucleotide-polymorphism-baseline-

genetic-stock-identification-chinook;   https://spo.nmfs.noaa.
gov/content/evaluation-single-nucleotide-polymorphism-baseline-
genetic-stock-identification-chinook-0;   https://spo.nmfs.
noaa.gov/sites/default/files/clemento.pdf;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2014/clemento.pdf.

Callihan:2014:EDS

[5102] Jody L. Callihan, Charlton H. Godwin, and Jeffrey A. Buckel. Effect of demography on spatial distribution: movement patterns of the Albemarle Sound—Roanoke River stock of striped bass (Morone saxatilis) in relation to their recovery. Fishery Bulletin, 112(2–3):131– 143, 2014. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/effect-
demography-spatial-distribution-movement-patterns-albemarle-
sound%E2%80%93roanoke-river;   https://spo.nmfs.noaa.gov/sites/
default/files/callihan.pdf.

Callihan:2014:RRS

[5103] Jody L. Callihan, Charlton H. Godwin, and Jeffrey A. Buckel. Roanoke River stock of striped bass (Morone saxatilis) in relation to their recovery. Fishery Bulletin, 112(2–3):131–143, 2014. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/roanoke-river-stock-striped-bass-
morone-saxatilis-relation-their-recovery;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2014/callihan.pdf.

Manderson:2014:RTH

[5104] John P. Manderson, Linda L. Stehlik, Jeff Pessutti, John Rosendale, and Beth Phelan. Residence time and habitat duration for predators in a small mid-Atlantic estuary. Fishery Bulletin, 112(2–3):144–158, 2014. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/residence-time-and-
habitat-duration-predators-small-mid-atlantic-estuary;   https:
//spo.nmfs.noaa.gov/content/residence-time-and-habitat-duration-
predators-small-mid-atlantic-estuary-0;   https://spo.nmfs.
noaa.gov/sites/default/files/manderson.pdf;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2014/manderson.pdf.

Henderson:2014:EFS

[5105] E. Elizabeth Henderson, Karin A. Forney, Jay P. Barlow, John A. Hildebrand, Annie B. Douglas, John Calambokidis, and William J. Sydeman. Effects of fluctuations in sea-surface temperature on the occurrence of small cetaceans off southern California. Fishery Bulletin, 112(2–3):

159–177, 2014. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/effects-
fluctuations-sea-surface-temperature-occurrence-small-cetaceans-
southern-0;   https://spo.nmfs.noaa.gov/content/effects-fluctuations-
sea-surface-temperature-occurrence-small-cetaceans-southern-
california;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2014/henderson.pdf.

Clardy:2014:AGR

[5106] Samuel D. Clardy, Nancy J. Brown-Peterson, Mark S. Peterson, and Robert T. Leaf. Age, growth, and reproduction of southern kingfish (Menticirrhus americanus): a multivariate comparison with life history patterns in other sciaenids. Fishery Bulletin, 112(2–3):178–197, 2014. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/age-growth-
and-reproduction-southern-kingfish-menticirrhus-americanus-
multivariate;   https://spo.nmfs.noaa.gov/content/age-growth-
and-reproduction-southern-kingfish-menticirrhus-americanus-
multivariate-0;   https://spo.nmfs.noaa.gov/sites/default/files/
clardy.pdf;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2014/clardy.pdf.

Douglas:2014:SDA

[5107] Annie B. Douglas, John Calambokidis, Lisa M. Munger, Melissa S. Soldevilla, Megan C. Ferguson, Andrea M. Havron, Dominique L. Camacho, Greg S. Campbell, and John A. Hildebrand. Seasonal distribution and abundance of cetaceans off Southern California estimated from CalCOFI cruise data from 2004 to 2008. Fishery Bulletin, 112(2–3): 198–220, 2014. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/seasonal-
distribution-and-abundance-cetaceans-southern-california-estimated-
calcofi-0;   https://spo.nmfs.noaa.gov/content/seasonal-distribution-
and-abundance-cetaceans-southern-california-estimated-calcofi-
cruise;   https://spo.nmfs.noaa.gov/sites/default/files/douglas.
pdf;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2014/douglas.pdf.

Rulifson:2014:PDH

[5108] Roger A. Rulifson and Christopher F. Batsavage. Population demographics of Hickory Shad (Alosa mediocris) during a period of population growth. Fishery Bulletin, 112(2–3):221–236, 2014. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/population-demographics-hickory-shad-

alosa-mediocris-during-period-population-growth;   https://
spo.nmfs.noaa.gov/content/population-demographics-hickory-shad-
alosa-mediocris-during-period-population-growth-0;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/2014/rulifson.
pdf;   https://spo.nmfs.noaa.gov/sites/default/files/rulifson.
pdf.

Lovell:2014:USW

[5109] Sabrina J. Lovell and David W. Carter. The use of sampling weights in regression models of recreational fishing-site choices. Fishery Bulletin, 112(4):243–252, 2014. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/use-
sampling-weights-regression-models-recreational-fishing-site-
choices;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2014/lovell.pdf.

Conrath:2014:ARB

[5110] Christina L. Conrath and M. Elizabeth Conners. Aspects of the reproductive biology of the North Pacific giant octopus (Enteroctopus dofleini) in the Gulf of Alaska. Fishery Bulletin, 112(4):253–260, 2014. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/aspects-reproductive-
biology-north-pacific-giant-octopus-enteroctopus-dofleini-gulf-
alaska;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/2014/conrath.pdf.

Saari:2014:RDA

[5111] Courtney R. Saari, James H. Cowan, Jr., and Kevin M. Boswell. Regional differences in the age and growth of red snapper (Lutjanus campechanus) in the U.S. Gulf of Mexico. Fishery Bulletin, 112(4):261–273, 2014. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/regional-differences-age-and-growth-ofred-
snapper-lutjanus-campechanus-us-gulf-mexico;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/2014/saari.pdf.

Baremore:2014:VMD

[5112] Ivy E. Baremore and J. Drew Rosati. A validated, minimally deleterious method for aging sturgeon. Fishery Bulletin, 112(4):274–282, 2014. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/validated-minimally-
deleterious-method-aging-sturgeon;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/2014/baremore.pdf.

Campbell:2014:RMR

[5113] Matthew D. Campbell, William B. Driggers III, Beverly Sauls, and John F. Walter. Release mortality in the red snapper (Lutjanus campechanus) fishery: a meta-analysis of 3 decades of research. Fishery Bulletin, 112(4):283–296, 2014. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/release-mortality-red-snapper-lutjanus-campechanus-
fishery-meta-analysis-3-decades-research;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/2014/campbell.pdf.

Horning:2014:CBE

[5114] Markus Horning and Jo-Ann E. Mellish. In cold blood: evidence of Pacific sleeper shark (Somniosus pacificus) predation on Steller sea lions (Eumetopias jubatus) in the Gulf of Alaska. Fishery Bulletin, 112(4):297–310, 2014. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/cold-blood-evidence-pacific-sleeper-shark-
somniosus-pacificus-predation-steller-sea-lions;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/2014/horning.
pdf.

Condini:2014:AGD

[5115] Mario V. Condini, Cristiano Q. Albuquerque, and Alexandre M. Garcia. Age and growth of dusky grouper (Epinephelus marginatus) (Perciformes: Epinephelidae) in the southwestern Atlantic, with a size comparison of offshore and littoral habitats. Fishery Bulletin, 112(4):311–321, 2014. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/age-and-growth-dusky-
grouper-epinephelus-marginatus-perciformes-epinephelidae-southwestern;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
2014/condini.pdf.

Antoni:2014:DCM

[5116] Luca Antoni, Patricia L. Luque, Kaylie Naghshpour, Lionel Reynal, and Eric A. Saillant. Development and characterization of microsatellite markers for blackfin tuna (Thunnus atlanticus) with the use of Illumina paired-end sequencing. Fishery Bulletin, 112(4):322–325, 2014. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/development-
and-characterization-microsatellite-markers-blackfin-tuna-thunnus-
atlanticus;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/2014/antoni.pdf.

Midway:2014:SFP

[5117] Stephen R. Midway, Steven X. Cadrin, and Frederick S. Scharf. Southern flounder (Paralichthys lethostigma) stock structure inferred from otolith shape analysis. Fishery Bulletin, 112(4):326– 338, 2014. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/southern-
flounder-paralichthys-lethostigma-stock-structure-inferred-
otolith-shape-analysis;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/2014/midway.pdf.

Aalbers:2015:SMP

[5118] Scott A. Aalbers and Chugey A. Sepulveda. Seasonal movement patterns and temperature profiles of adult white seabass (Atractoscion nobilis) off California. Fishery Bulletin, 113(1):1–14, 2015. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/seasonal-movement-patterns-and-
temperature-profiles-adult-white-seabass-atractoscion;   https:
//spo.nmfs.noaa.gov/sites/default/files/aalbers.pdf.

Bacheler:2015:ERA

[5119] Nathan M. Bacheler and Kyle W. Shertzer. Estimating relative abundance and species richness from video surveys of reef fishes. Fishery Bulletin, 113(1):15–26, 2015. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/estimating-relative-abundance-and-species-richness-
video-surveys-reef-fishes;   https://spo.nmfs.noaa.gov/sites/
default/files/bacheler.pdf.

Burton:2015:AGM

[5120] Michael L. Burton, Jennifer C. Potts, Daniel R. Carr, Michael Cooper, and Jessica Lewis. Age, growth, and mortality of gray triggerfish (Balistes capriscus) from the southeastern United States. Fishery Bulletin, 113(1):27–39, 2015. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/age-
growth-and-mortality-gray-triggerfishbalistes-capriscus-southeastern-
united-states;   https://spo.nmfs.noaa.gov/sites/default/files/
burton.pdf.

Niamaimandi:2015:EVS

[5121] Nassir Niamaimandi and Gholam-Abbas Zarshenas. Economic valuation of stock enhancement of banana prawn (Fenneropenaeus merguiensis) in the Strait of Hormuz. Fishery Bulletin, 113(1):40–46, 2015. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/economic-valuation-stock-
enhancement-banana-prawn-fenneropenaeus-merguiensis-strait-
hormuz;   https://spo.nmfs.noaa.gov/sites/default/files/niamaimandi.
pdf.

Elzey:2015:CAS

[5122] Scott P. Elzey, Katie A. Rogers, and Kimberly J. Trull. Comparison of 4 aging structures in the American shad (Alosa sapidissima). Fishery Bulletin, 113(1):47–54, 2015. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
comparison-4-aging-structures-american-shad-alosa-sapidissima;
 https://spo.nmfs.noaa.gov/sites/default/files/elzey.pdf.

deAlmeidaDuarte:2015:SLS

[5123] Luis Felipe de Almeida Duarte, Evandro Severino-Rodrigues, Marcelo A. A. Pinheiro, and Maria A. Gasalla. Slipper lobsters (Scyllaridae) off the southeastern coast of Brazil: relative growth, population structure, and reproductive biology. Fishery Bulletin, 113(1): 55–68, 2015. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/slipper-
lobsters-scyllaridae-southeastern-coast-brazil-relative-growth-
population-structure;   https://spo.nmfs.noaa.gov/sites/default/
files/duarte.pdf.

Ligas:2015:MGR

[5124] Alessandro Ligas, Francesco Colloca, Mathieu G. Lundy, Alessandro Mannini, Paolo Sartor, Mario Sbrana, Alessandro Voliani, and Paola Belcari. Modeling the growth of recruits of European hake (Merluccius merluccius) in the northwestern Mediterranean Sea with generalized additive models. Fishery Bulletin, 113(1):69–81, 2015. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/modeling-growth-recruits-
european-hake-merluccius-merluccius-northwestern-mediterranean-
sea;   https://spo.nmfs.noaa.gov/sites/default/files/ligas.
pdf.

Lindholm:2015:EEB

[5125] James Lindholm, Mary Gleason, Donna Kline, Larissa Clary, Steve Rienecke, Alli Cramer, and Marc Los Huertos. Ecological effects of bottom trawling on the structural attributes of fish habitat in unconsolidated sediments along the central California outer continental shelf. Fishery Bulletin, 113(1):82–96, 2015. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/ecological-effects-
bottom-trawling-structural-attributes-fish-habitat-unconsolidated;
 https://spo.nmfs.noaa.gov/sites/default/files/lindholm.pdf.

Sagarese:2015:SOS

[5126] Skyler R. Sagarese, Michael G. Frisk, Robert M. Cerrato, Kathy A. Sosebee, John A. Musick, and Paul J. Rago. Spatiotemporal overlap of spiny dogfish (Squalus acanthias) and commercial fisheries in the northeast U.S. shelf large marine ecosystem. Fishery Bulletin, 113(2):101–120, 2015. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
spatiotemporal-overlap-spiny-dogfish-squalus-acanthias-and-
commercial-fisheries-northeast;   https://spo.nmfs.noaa.gov/
sites/default/files/sagarese.pdf.

Nakano:2015:DSS

[5127] Hikaru Nakano, Koji Hirakawa, and Jun Shoji. Development of swimming speed and schoolingbehavior of juvenile white rockfish (Sebastes cheni) in relation to ambient light intensity. Fishery Bulletin, 113(2):121– 128, 2015. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/development-
swimming-speed-and-schoolingbehavior-juvenile-white-rockfish-
sebastes-cheni;   https://spo.nmfs.noaa.gov/sites/default/files/
nakano.pdf.

Orr:2015:SNP

[5128] James W. Orr, Sharon Wildes, Yoshiaki Kai, Nate Raring, T. Nakabo, Oleg Katugin, and Jeff Guyon. Systematics of North Pacific sand lances of the genus Ammodytes based on molecular and morphological evidence, with the description of a new species from Japan. Fishery Bulletin, 113(2):129–156, 2015. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/systematics-north-pacific-sand-
lances-genus-ammodytes-based-molecular-and-morphological;
https://spo.nmfs.noaa.gov/sites/default/files/orr.pdf.

Garner:2015:DOF

[5129] Steven B. Garner and William F. Patterson. Direct observation of fishing effort, catch, and discard rates of charter boats targeting reef fishes in the northern Gulf of Mexico. Fishery Bulletin, 113(2):157– 166, 2015. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/direct-
observation-fishing-effort-catch-and-discard-rates-charter-

boats-targeting-reef;   https://spo.nmfs.noaa.gov/sites/default/
files/garner.pdf.

Turnure:2015:PIE

[5130] Jason T. Turnure, Kenneth W. Able, and Thomas M. Grothues. Patterns of intra-estuarine movement of adult weakfish (Cynoscion regalis): evidence of site affinity at seasonal and diel scales. Fishery Bulletin, 113(2):167–179, 2015. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
patterns-intra-estuarine-movement-adult-weakfish-cynoscion-
regalis-evidence-site-affinity;   https://spo.nmfs.noaa.gov/
sites/default/files/turnure.pdf.

Weinberg:2015:RVB

[5131] Kenneth L. Weinberg and Stan Kotwicki. Reducing variability in bottom contact and net width of a survey trawl by restraining door movement and applying a constant ratio of warp length to depth. Fishery Bulletin, 113(2):180–190, 2015. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/reducing-variability-
bottom-contact-and-net-width-survey-trawl-restraining-door-
movement;   https://spo.nmfs.noaa.gov/sites/default/files/weinberg_
0.pdf.

Fairchild:2015:SFA

[5132] Elizabeth A. Fairchild, Shelly Tallack, Scott P. Elzey, and Michael P. Armstrong. Spring feeding of Atlantic wolffish (Anarhichas lupus) on Stellwagen Bank, Massachusetts. Fishery Bulletin, 113(2):191– 201, 2015. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/spring-
feeding-atlantic-wolffish-anarhichas-lupus-stellwagen-bank-
massachusetts;   https://spo.nmfs.noaa.gov/sites/default/files/
fairchild.pdf.

Chang:2015:AOI

[5133] Ke-Yang Chang, Chih-Shin Chen, Hui-Yu Wang, Chin-Lau Kuo, and Tai-Sheng Chiu. The Antarctic Oscillation index as an environmental parameter for predicting catches of the Argentine shortfin squid (Illex argentinus) (cephalopoda: Ommastrephidae) in southwest Atlantic waters. Fishery Bulletin, 113(2):202–212, 2015. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/antarctic-oscillation-index-environmental-

parameter-predicting-catches-argentine-shortfin;   https://
spo.nmfs.noaa.gov/sites/default/files/chang.pdf.

Weitkamp:2015:SAS

[5134] Laurie A. Weitkamp, Susan A. Hinton, and Paul J. Bentley. Seasonal abundance, size, and host selection of western river (Lampetra ayresii) and Pacific (Entosphenus tridentatus) lampreys in the Columbia River estuary. Fishery Bulletin, 113(2):213–226, 2015. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/seasonal-abundance-
size-and-host-selection-western-river-lampetra-ayresii-and-
pacific;   https://spo.nmfs.noaa.gov/sites/default/files/weitkamp.
pdf.

Parrish:2015:ATM

[5135] Frank A. Parrish, Nicholas T. Hayman, Christopher Kelley, and Raymond C. Boland. Acoustic tagging and monitoring of cultured and wild juvenile crimson jobfish (Pristipomoides filamentosus) in a nursery habitat. Fishery Bulletin, 113(3):231–241, 2015. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/acoustic-tagging-and-monitoring-
cultured-and-wild-juvenile-crimson-jobfish-pristipomoides-0;
 https://spo.nmfs.noaa.gov/sites/default/files/parrish.pdf.

Dahlheim:2015:TCA

[5136] Marilyn E. Dahlheim, Alexandre N. Zerbini, Janice M. Waite, and Amy S. Kennedy. Temporal changes in abundance of harbor porpoise (Phocoena phocoena) inhabiting the inland waters of Southeast Alaska. Fishery Bulletin, 113(3):242–255, 2015. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/temporal-changes-abundance-harbor-porpoise-phocoena-
phocoena-inhabiting-inland-waters-0;   https://spo.nmfs.noaa.
gov/sites/default/files/dahlheim.pdf.

Ryer:2015:DDH

[5137] Clifford H. Ryer, William C. Long, Mara L. Spencer, and Paul Iseri. Depth distribution, habitat associations, and differential growth of newly settled southern Tanner crab (Chionoecetes bairdi) in embayments around Kodiak Island, Alaska. Fishery Bulletin, 113(3):256–269, 2015. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/depth-distribution-
habitat-associations-and-differential-growth-newly-settled-

southern-0;   https://spo.nmfs.noaa.gov/sites/default/files/
ryer.pdf.

Laman:2015:CEB

[5138] Edward A. Laman, Stan Kotwicki, and Christopher N. Rooper. Correlating environmental and biogenic factors with abundance and distribution of Pacific Ocean perch (Sebastes alutus) in the Aleutian Islands, Alaska. Fishery Bulletin, 113(3):270–289, 2015. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/correlating-environmental-and-biogenic-factors-
abundance-and-distribution-pacific-ocean;   https://spo.nmfs.
noaa.gov/sites/default/files/laman.pdf.

Willis:2015:DCF

[5139] C. Michelle Willis, Jonathan Richardson, Tracey Smart, Joseph Cowan, and Patrick Biondo. Diet composition, feeding strategy, and diet overlap of 3 sciaenids along the southeastern United States. Fishery Bulletin, 113(3):290–301, 2015. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/diet-composition-feeding-strategy-and-diet-overlap-
3-sciaenids-along-southeastern-united;   https://spo.nmfs.noaa.
gov/sites/default/files/willis.pdf.

Lin:2015:SYP

[5140] Yu-Jia Lin, Chi-Lu Sun, Yi-Jay Chang, and Wann-Nian Tzeng. Sensitivity of yield-per-recruit and spawning-biomass-per-recruit models to bias and imprecision in life history parameters: an example based on life history parameters of Japanese eel (Anguilla japonica). Fishery Bulletin, 113(3):302–312, 2015. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/sensitivity-yield-recruit-and-spawning-biomass-
recruit-models-bias-and-imprecision-life;   https://spo.nmfs.
noaa.gov/sites/default/files/lin.pdf.

Nichol:2015:SMM

[5141] Daniel G. Nichol and David A. Somerton. Seasonal migrations of morphometrically mature male snow crab (Chionoecetes opilio) in the eastern Bering Sea in relation to mating dynamics. Fishery Bulletin, 113(3):313–326, 2015. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
seasonal-migrations-morphometrically-mature-male-snow-crab-
chionoecetes-opilio-eastern;   https://spo.nmfs.noaa.gov/sites/
default/files/nichol.pdf.

Beck:2015:EOH

[5142] Steve Beck and Megan K. La Peyre. Effects of oyster harvest activities on Louisiana reef habitat and resident nekton communities. Fishery Bulletin, 113(3):327–340, 2015. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
effects-oyster-harvest-activities-louisiana-reef-habitat-and-
resident-nekton-communities;   https://spo.nmfs.noaa.gov/sites/
default/files/beck.pdf.

OFarrell:2015:IHF

[5143] Michael R. O’Farrell and William H. Satterthwaite. Inferred historical fishing mortality rates for an endangered population of Chinook salmon (Oncorhynchus tshawytscha). Fishery Bulletin, 113(3):341–351, 2015. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/inferred-historical-
fishing-mortality-rates-endangered-population-chinook-salmon;
 https://spo.nmfs.noaa.gov/sites/default/files/ofarrell.pdf.

Frable:2015:NSS

[5144] Benjamin W. Frable, D. Wolfe Wagman, Taylor N. Frierson, Andres Aguilar, and Brian L. Sidlauskas. A new species of Sebastes (Scorpaeniformes: Sebastidae) from the northeastern Pacific, with a redescription of the blue rockfish, S. mystinus (Jordan and Gilbert, 1881). Fishery Bulletin, 113(4):355–377, 2015. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/new-species-sebastes-scorpaeniformes-
sebastidae-northeastern-pacific-redescription-blue;   https:
//spo.nmfs.noaa.gov/sites/default/files/frable_0.pdf.

Voss:2015:FDD

[5145] Christine M. Voss, Joan A. Browder, Andrew Wood, and Adriane Michaelis. Factors driving the density of derelict crab pots and their associated bycatch in North Carolina waters. Fishery Bulletin, 113(4):378–390, 2015. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/factors-driving-density-derelict-crab-pots-and-
their-associated-bycatch-north-carolina;   https://spo.nmfs.
noaa.gov/sites/default/files/voss.pdf.

Sanchez-Rubio:2015:CRM

[5146] Guillermo Sanchez-Rubio and Harriet Perry. Climate-related meteorological and hydrological regimes and their influence on recruitment of Gulf menhaden (Brevoortia patronus) in the northern Gulf of Mexico. Fishery Bulletin, 113(4):391–406, 2015. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/climate-related-meteorological-
and-hydrological-regimes-and-their-influence-recruitment;
https://spo.nmfs.noaa.gov/sites/default/files/sanchez-rubio.
pdf.

Zischke:2015:PRS

[5147] Mitchell T. Zischke and Shane P. Griffiths. Per-recruit stock assessment of wahoo (Acanthocybium solandri) in the southwest Pacific Ocean. Fishery Bulletin, 113(4):407–418, 2015. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/recruit-stock-assessment-wahoo-
acanthocybium-solandri-southwest-pacific-ocean;   https://spo.
nmfs.noaa.gov/sites/default/files/zischke.pdf.

Merten:2015:GSD

[5148] Wessley B. Merten, Nikolaos V. Schizas, Matthew T. Craig, Richard S. Appeldoorn, and Donald L. Hammond. Genetic structure and dispersal capabilities of dolphinfish (Coryphaena hippurus) in the western central Atlantic. Fishery Bulletin, 113(4):419–429, 2015. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/genetic-structure-and-dispersal-
capabilities-dolphinfish-coryphaena-hippurus-western;   https:
//spo.nmfs.noaa.gov/sites/default/files/merten.pdf.

Underwood:2015:BDS

[5149] Melanie J. Underwood, Paul D. Winger, Anders Fernö, and Arill Engås. Behavior-dependent selectivity of yellowtail flounder (Limanda ferruginea) in the mouth of a commercial bottom trawl. Fishery Bulletin, 113(4):430–441, 2015. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
behavior-dependent-selectivity-yellowtail-flounder-limanda-
ferruginea-mouth-commercial;   https://spo.nmfs.noaa.gov/sites/
default/files/underwood.pdf.

Seyoum:2015:GDP

[5150] Seifu Seyoum, Angela B. Collins, Cecilia Puchulutegui, Richard S. McBride, and Michael D. Tringali. Genetically determined population structure of hogfish (Labridae: Lachnolaimus maximus) in the southeastern United States. Fishery Bulletin, 113(4):442–455, 2015. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/genetically-

determined-population-structure-hogfish-labridae-lachnolaimus-
maximus;   https://spo.nmfs.noaa.gov/sites/default/files/seyoum.
pdf.

Gomila:2015:ELH

[5151] Lujan Villanueva Gomila, Martín D. Ehrlich, and Leonardo A. Venerus. Early life history of the Argentine sea bass (Acanthistius patachonicus) (Pisces: Serranidae). Fishery Bulletin, 113(4):456–467, 2015. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/early-life-history-
argentine-sea-bass-acanthistius-patachonicus-pisces-serranidae;
 https://spo.nmfs.noaa.gov/sites/default/files/gomila.pdf.

Fernandez-Carvalho:2015:MAG

[5152] Joana Fernandez-Carvalho, Rui Coelho, Karim Erzini, and Miguel N. Santos. Modeling age and growth of the bigeye thresher (Alopias superciliosus) in the Atlantic Ocean. Fishery Bulletin, 113(4):468–481, 2015. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/modeling-age-
and-growth-bigeye-thresher-alopias-superciliosus-atlantic-ocean;
 https://spo.nmfs.noaa.gov/sites/default/files/fernandez-carvalho.
pdf.

Clarke:2016:EBA

[5153] Tayler M. Clarke, Mario Espinoza, Robert Ahrens, and Ingo S. Wehrtmann. Elasmobranch bycatch associated with the shrimp trawl fishery off the Pacific coast of Costa Rica, Central America. Fishery Bulletin, 114(1):1–17, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/elasmobranch-bycatch-associated-
shrimp-trawl-fishery-pacific-coast-costa-rica-central;   https:
//spo.nmfs.noaa.gov/sites/default/files/clarke.pdf.

Kupchik:2016:AGR

[5154] Matthew J. Kupchik and Richard F. Shaw. Age, growth, and recruitment of larval and early juvenile Atlantic croaker (Micropogonias undulatus), determined from analysis of otolith microstructure. Fishery Bulletin, 114(1):18–33, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/age-
growth-and-recruitment-larval-and-early-juvenile-atlantic-croaker-
micropogonias;   https://spo.nmfs.noaa.gov/sites/default/files/
kupchik.pdf.

Fang:2016:AGP

[5155] Zhou Fang, Jianhua Li, Katherine Thompson, Feifei Hu, Xinjun Chen, Bilin Liu, and Yong Chen. Age, growth, and population structure of the red flying squid (Ommastrephes bartramii) in the North Pacific Ocean, determined from beak microstructure. Fishery Bulletin, 114(1):34–44, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/age-growth-and-population-structure-red-flying-
squid-ommastrephes-bartramii-north-pacific;   https://spo.nmfs.
noaa.gov/sites/default/files/fang.pdf.

Ross:2016:FSA

[5156] Steve W. Ross, Mike Rhode, Stephen T. Viada, and Rod Mather. Fish species associated with shipwreck and natural hard-bottom habitats from the middle to outer continental shelf of the Middle Atlantic Bight near Norfolk Canyon. Fishery Bulletin, 114(1):45–57, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/fish-species-
associated-shipwreck-and-natural-hard-bottom-habitats-middle-
outer-continental;   https://spo.nmfs.noaa.gov/sites/default/
files/ross.pdf.

Long:2016:NQM

[5157] W. Christopher Long. A new quantitative model of multiple transitions between discrete stages, applied to the development of crustacean larvae. Fishery Bulletin, 114(1):58–66, 2016. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/new-quantitative-model-multiple-
transitions-between-discrete-stages-applied-development;   https:
//spo.nmfs.noaa.gov/sites/default/files/long.pdf.

Long:2016:EDG

[5158] W. Christopher Long and Scott B. Van Sant. Embryo development in golden king crab (Lithodes aequispinus). Fishery Bulletin, 114(1): 67–76, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/embryo-
development-golden-king-crab-lithodes-aequispinus;   https:
//spo.nmfs.noaa.gov/sites/default/files/long2.pdf.

Lytton:2016:AVN

[5159] Adam R. Lytton, Joseph C. Ballenger, Marcel J. M. Reichert, and Tracey I. Smart. Age validation of the North Atlantic stock of wreckfish (Polyprion americanus), based on bomb radiocarbon (^{14}C), and new estimates of life history parameters. Fishery Bulletin, 114(1):77–88, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/age-validation-north-
atlantic-stock-wreckfish-polyprion-americanus-based-bomb-radiocarbon;
 https://spo.nmfs.noaa.gov/sites/default/files/lytton.pdf.

Rodgveller:2016:AMS

[5160] Cara J. Rodgveller, James W. Stark, Katy B. Echave, and Peter-John F. Hulson. Age at maturity, skipped spawning, and fecundity of female sablefish (Anoplopoma fimbria) during the spawning season. Fishery Bulletin, 114(1):89–102, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/age-maturity-skipped-spawning-and-fecundity-female-
sablefish-anoplopoma-fimbria-during;   https://spo.nmfs.noaa.
gov/sites/default/files/rodgveller.pdf.

Locascio:2016:PAS

[5161] James V. Locascio and Michael L. Burton. A passive acoustic survey of fish sound production at Riley’s Hump within Tortugas South Ecological Reserve: implications regarding spawning and habitat use. Fishery Bulletin, 114(1):103–116, 2016. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
passive-acoustic-survey-fish-sound-production-riley%E2%80%99s-
hump-within-tortugas-south-ecological;   https://spo.nmfs.noaa.
gov/sites/default/files/locascio.pdf.

Smith:2016:UCT

[5162] Nicola S. Smith and Dirk Zeller. Unreported catch and tourist demand on local fisheries of small island states: the case of The Bahamas, 1950— 2010. Fishery Bulletin, 114(1):117–131, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/unreported-catch-and-tourist-demand-local-
fisheries-small-island-states-case-bahamas-1950;   https://
spo.nmfs.noaa.gov/sites/default/files/smith.pdf.

Watson:2016:ELS

[5163] William Watson, Sharon R. Charter, and Cynthia A. Taylor Lawley. Early larvae of the swordspine rockfish (Sebastes ensifer) identified by molecular methods. Fishery Bulletin, 114(2):135–143, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/early-larvae-
swordspine-rockfish-sebastes-ensifer-identified-molecular-methods;
 https://spo.nmfs.noaa.gov/sites/default/files/watson.pdf.

Tagliafico:2016:ERB

[5164] Alejandro Tagliafico, Nicolás Ehemann, María Salomé Rangel, and Néstor Rago. Exploitation and reproduction of the bullnose ray (Myliobatis freminvillei) caught in an artisanal fishery in La Pared, Margarita Island, Venezuela. Fishery Bulletin, 114(2):144–152, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/exploitation-
and-reproduction-bullnose-ray-myliobatis-freminvillei-caught-
artisanal-fishery;   https://spo.nmfs.noaa.gov/sites/default/
files/tagliafico.pdf.

Lewis:2016:IDB

[5165] Leah A. Lewis, David E. Richardson, Evgeny V. Zakharov, and Robert Hanner. Integrating DNA barcoding of fish eggs into ichthyoplankton monitoring programs. Fishery Bulletin, 114(2):153–165, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/integrating-dna-barcoding-fish-
eggs-ichthyoplankton-monitoring-programs;   https://spo.nmfs.
noaa.gov/sites/default/files/lewis.pdf.

Lopez:2016:MBD

[5166] Jon Lopez, Gala Moreno, Guillermo Boyra, and Laurent Dagorn. A model based on data from echosounder buoys to estimate biomass of fish species associated with fish aggregating devices. Fishery Bulletin, 114 (2):166–178, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/model-
based-data-echosounder-buoys-estimate-biomass-fish-species-
associated-fish;   https://spo.nmfs.noaa.gov/sites/default/
files/lopez.pdf.

Demetras:2016:DUR

[5167] Nicholas J. Demetras, David D. Huff, Cyril J. Michel, Joseph M. Smith, George R. Cutter, Sean A. Hayes, and Steven T. Lindley. Development of underwater recorders to quantify predation of juvenile Chinook salmon (Oncorhynchus tshawytscha) in a river environment. Fishery Bulletin, 114(2):179–185, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/development-underwater-recorders-quantify-
predation-juvenile-chinook-salmon-oncorhynchus;   https://spo.
nmfs.noaa.gov/sites/default/files/demetras.pdf.

Mathews:2016:HPE

[5168] Elizabeth A. Mathews, Lauri A. Jemison, Grey W. Pendleton, Karen M. Blejwas, Kevin E. Hood, and Kimberly L. Raum-Suryan. Haul-out patterns and effects of vessel disturbance on harbor seals (Phoca vitulina) on glacial ice in Tracy Arm, Alaska. Fishery Bulletin, 114(2): 186–202, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/haul-out-
patterns-and-effects-vessel-disturbance-harbor-seals-phoca-
vitulina-glacial-ice;   https://spo.nmfs.noaa.gov/sites/default/
files/Mathews.pdf.

Yasumiishi:2016:EPA

[5169] Ellen M. Yasumiishi, Keith R. Criddle, John H. Helle, Nicola Hillgruber, and Franz J. Mueter. Effect of population abundance and climate on the growth of 2 populations of chum salmon (Oncorhynchus keta) in the eastern North Pacific Ocean. Fishery Bulletin, 114(2):203–219, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/effect-population-abundance-and-climate-growth-
2-populations-chum-salmon-oncorhynchus-keta;   https://spo.
nmfs.noaa.gov/sites/default/files/Yasumiishi.pdf.

Contente:2016:FAS

[5170] Riguel F. Contente and Carmen L. D. B. Rossi-Wongtschowski. Fish assemblages on the southeastern Brazilian Bight, sampled by midwater trawl during spring and summer seasons: species composition, abundance, and environmental drivers. Fishery Bulletin, 114 (2):220–236, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/fish-
assemblages-southeastern-brazilian-bight-sampled-midwater-trawl-
during-spring-and;   https://spo.nmfs.noaa.gov/sites/default/
files/contente.pdf.

Weijerman:2016:TBC

[5171] Mariska Weijerman, Ivor Williams, Jay Gutierrez, Shanna Grafeld, Brent Tibbatts, and Gerry Davis. Trends in biomass of coral reef fishes, derived from shore-based creel surveys in Guam. Fishery Bulletin, 114(2):237–256, 2016. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
trends-biomass-coral-reef-fishes-derived-shore-based-creel-
surveys-guam;   https://spo.nmfs.noaa.gov/sites/default/files/
weijerman.pdf.

Walker:2016:AUC

[5172] Justin H. Walker, Arthur C. Trembanis, and Douglas C. Miller. Assessing the use of a camera system within an autonomous underwater vehicle for monitoring the distribution and density of sea scallops (Placopecten magellanicus) in the Mid-Atlantic Bight. Fishery Bulletin, 114(3):261–273, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/assessing-use-camera-
system-within-autonomous-underwater-vehicle-monitoring-distribution;
 https://spo.nmfs.noaa.gov/sites/default/files/walker.pdf.

VanNoord:2016:OID

[5173] Joel E. Van Noord, Robert J. Olson, Jessica V. Redfern, Leanne M. Duffy, and Ronald S. Kaufmann. Oceanographic influences on the diet of 3 surface-migrating myctophids in the eastern tropical Pacific Ocean. Fishery Bulletin, 114(3):274–287, 2016. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/oceanographic-influences-diet-3-
surface-migrating-myctophids-eastern-tropical-pacific-ocean;
 https://spo.nmfs.noaa.gov/sites/default/files/vannoord.pdf.

Caldarone:2016:ENA

[5174] Elaine M. Caldarone, Sharon A. MacLean, and Brian R. Beckman. Evaluation of nucleic acids and plasma IGF1 levels for estimating short-term responses of postsmolt Atlantic salmon (Salmo salar) to food availability. Fishery Bulletin, 114(3):288–301, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/evaluation-nucleic-acids-
and-plasma-igf1-levels-estimating-short-term-responses-postsmolt;
 https://spo.nmfs.noaa.gov/sites/default/files/caldarone.pdf.

LeClair:2016:SCA

[5175] Larry L. LeClair, Ocean Eveningsong, and Jesse M. Schultz. Seasonal changes in abundance and compelling evidence of migration for 2 rockfish species (Sebastes auriculatus and S. caurinus) inhabiting a nearshore, temperate-water artificial reef. Fishery Bulletin, 114(3):302–316, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/seasonal-changes-abundance-and-compelling-evidence-
migration-2-rockfish-species-sebastes;   https://spo.nmfs.noaa.
gov/sites/default/files/leclair.pdf.

Goldman:2016:FHR

[5176] Sarah F. Goldman, Dawn M. Glasgow, and Michelle M. Falk. Feeding habits of 2 reef-associated fishes, red porgy (Pagrus pagrus) and gray triggerfish (Balistes capriscus), off the southeastern United States. Fishery Bulletin, 114(3):317–329, 2016. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/feeding-habits-2-reef-associated-
fishes-red-porgy-pagrus-pagrus-and-gray-triggerfish;   https:
//spo.nmfs.noaa.gov/sites/default/files/goldman.pdf.

Syah:2016:PPF

[5177] Achmad F. Syah, Sei-Ichi Saitoh, Irene D. Alabia, and Toru Hirawake. Predicting potential fishing zones for Pacific saury (Cololabis saira) with maximum entropy models and remotely sensed data. Fishery Bulletin, 114(3):330–342, 2016. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
predicting-potential-fishing-zones-pacific-saury-cololabis-
saira-maximum-entropy-models-and;   https://spo.nmfs.noaa.gov/
sites/default/files/syah.pdf.

Stevens:2016:DTD

[5178] Bradley G. Stevens and Vincent Guida. Depth and temperature distribution, morphometrics, and sex ratios of red deepsea crab (Chaceon quinquedens) at 4 sampling sites in the Mid-Atlantic Bight. Fishery Bulletin, 114(3):343–359, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-

4518 (electronic). URL https://spo.nmfs.noaa.gov/content/depth-
and-temperature-distribution-morphometrics-and-sex-ratios-red-
deepsea-crab-chaceon;   https://spo.nmfs.noaa.gov/sites/default/
files/stevens.pdf.

Weinberg:2016:SSC

[5179] Kenneth L. Weinberg, Cynthia Yeung, David A. Somerton, Grant G. Thompson, and Patrick H. Ressler. Is the survey selectivity curve for Pacific cod (Gadus macrocephalus) dome-shaped? Direct evidence from trawl studies. Fishery Bulletin, 114(3):360–369, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/survey-selectivity-curve-pacific-cod-
gadus-macrocephalus-dome-shaped-direct-evidence-trawl;   https:
//spo.nmfs.noaa.gov/sites/default/files/weinberg.pdf.

Willey:2016:CCH

[5180] Angel L. Willey, Linda S. Barker, and Mark Sampson. A comparison of circle hook and J hook performance in the recreational shark fishery off Maryland. Fishery Bulletin, 114(3):370–372, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/comparison-circle-hook-and-j-hook-performance-
recreational-shark-fishery-maryland;   https://spo.nmfs.noaa.
gov/sites/default/files/willey.pdf.

Elzey:2016:INM

[5181] Scott P. Elzey and Kimberly J. Trull. Identification of a nonlethal method for aging tautog (Tautoga onitis). Fishery Bulletin, 114(4):377–385, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/identification-nonlethal-
method-aging-tautog-tautoga-onitis;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/elzey.pdf.

Laidig:2016:CDL

[5182] Thomas E. Laidig and Mary M. Yoklavich. A comparison of density and length of Pacific groundfishes observed from 2 survey vehicles: a manned submersible and a remotely operated vehicle. Fishery Bulletin, 114(4):386–396, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/comparison-density-and-length-pacific-groundfishes-
observed-2-survey-vehicles-manned;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/laidig.pdf.

Macchi:2016:SSP

[5183] Gustavo J. Macchi, Marina V. Diaz, Ezequiel Leonarduzzi, María Inés Militelli, and Karina Rodrigues. Skipped spawning in the Patagonian stock of Argentine hake (Merluccius hubbsi). Fishery Bulletin, 114(4): 397–408, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/skipped-
spawning-patagonian-stock-argentine-hake-merluccius-hubbsi;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/macchi.pdf.

Trianni:2016:LHC

[5184] Michael S. Trianni. Life history characteristics and stock status of the thumbprint emperor (Lethrinus harak) in Saipan Lagoon. Fishery Bulletin, 114(4):409–425, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/life-history-characteristics-and-
stock-status-thumbprint-emperor-lethrinus-harak-saipan;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
trianni.pdf.

Whitney:2016:SDD

[5185] Nicholas M. Whitney, Marc Taquet, Richard W. Brill, Charlotte Girard, Gail D. Schwieterman, Laurent Dagorn, and Kim N. Holland. Swimming depth of dolphinfish (Coryphaena hippurus) associated and unassociated with fish aggregating devices. Fishery Bulletin, 114(4):426–434, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/swimming-depth-dolphinfish-coryphaena-hippurus-
associated-and-unassociated-fish-aggregating;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/whitney.
pdf.

Lopez:2016:FAF

[5186] Miranda Lopez, Gavino Puggioni, and David A. Bengtson. First assessment of the field ecology of larval Atlantic silverside (Menidia menidia). Fishery Bulletin, 114(4):435–444, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/first-assessment-field-ecology-larval-atlantic-
silverside-menidia-menidia;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/lopez.pdf.

Bizzarro:2016:BBH

[5187] Joseph J. Bizzarro, Ashley N. Peterson, Jennifer M. Blaine, Jordan P. Balaban, H. Gary Greene, and Adam P. Summers. Burrowing behavior, habitat, and functional morphology of the Pacific sand lance (Ammodytes personatus). Fishery Bulletin, 114(4):445–460, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/burrowing-behavior-habitat-and-
functional-morphology-pacific-sand-lance-ammodytes;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
bizzarro.pdf.

Poulakis:2016:LTS

[5188] Gregg R. Poulakis, Philip W. Stevens, Amy A. Timmers, Christopher J. Stafford, Demian D. Chapman, Kevin A. Feldheim, Michelle R. Heupel, and Caitlin Curtis. Long-term site fidelity of endangered small-tooth sawfish (Pristis pectinata) from different mothers. Fishery Bulletin, 114(4):461–475, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/long-term-site-fidelity-endangered-small-tooth-
sawfish-pristis-pectinata-different-mothers;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/poulakis.
pdf.

Blaylock:2016:EVA

[5189] Jessica Blaylock and Gary R. Shepherd. Evaluating the vulnerability of an atypical protogynous hermaphrodite to fishery exploitation: results from a population model for black sea bass (Centropristis striata). Fishery Bulletin, 114(4):476–489, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/evaluating-vulnerability-
atypical-protogynous-hermaphrodite-fishery-exploitation-results;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/blaylock.pdf.

McIlwain:2016:DPO

[5190] Jennifer L. McIlwain, Aisha Ambu-ali, Nasr Al Jardani, Andrew. R. Halford, Hamed S. Al-Oufi, and David A. Feary. Demographic profile of an overexploited serranid, the orange-spotted grouper (Epinephelus coioides), from northern Oman. Fishery Bulletin, 114(4):490–502, 2016. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/demographic-profile-
overexploited-serranid-orange-spotted-grouper-epinephelus-coioides;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/mcilwain.pdf.

Mickle:2016:ECR

[5191] Paul F. Mickle, Jacob F. Schaefer, Susan B. Adams, Brian R. Kreiser, and William T. Slack. Environmental conditions of 2 river drainages into the northern Gulf of Mexico during successful hatching of Alabama shad (Alosa alabamae). Fishery Bulletin, 114(4):503–512, 2016. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/environmental-conditions-2-river-
drainages-northern-gulf-mexico-during-successful-hatching;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/mickle.pdf.

Steinhorst:2017:AEC

[5192] Kirk Steinhorst, Timothy Copeland, Michael W. Ackerman, William C. Schrader, and Eric C. Anderson. Abundance estimates and confidence intervals for the run composition of returning salmonids. Fishery Bulletin, 115(1):1–12, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/abundance-estimates-and-confidence-intervals-run-composition-
returning;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/steinhorst.pdf.

LopezQuintero:2017:FBA

[5193] Freddy Omar López Quintero, Javier E. Contreras-Reyes, Rodrigo Wiff, and Reinaldo B. Arellano-Valle. Flexible Bayesian analysis of the von Bertalanffy growth function with the use of a log-skew-t distribution. Fishery Bulletin, 115(1):13–26, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fishery-bulletin/flexible-bayesian-analysis-von-bertalanffy-
growth-function-use-log-skew-t;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/lopezquintero.pdf.

Dapp:2017:ULD

[5194] Derek R. Dapp, Charlie Huveneers, Terence I. Walker, John Mandelman, David W. Kerstetter, and Richard D. Reina. Using logbook data to determine the immediate mortality of blue sharks (Prionace glauca) and tiger sharks (Galeocerdo cuvier) caught in the commercial U.S. pelagic longline fishery. Fishery Bulletin, 115(1):27–41, 2017. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/fishery-bulletin/using-logbook-
data-determine-immediate-mortality-blue-sharks-prionace;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
dapp.pdf.

Kilada:2017:PAD

[5195] Raouf Kilada, Joel B. Webb, Kevin W. McNeel, Laura M. Slater, Quinn Smith, and Jayde Ferguson. Preliminary assessment of a direct age determination method for 3 commercially important crustaceans from Alaska. Fishery Bulletin, 115(1):42–49, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fishery-bulletin/preliminary-assessment-direct-age-
determination-method-3-commercially;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/kilada.pdf.

Driggers:2017:IBT

[5196] William B. Driggers III, Matthew D. Campbell, Kristin M. Hannan, Eric R. Hoffmayer, Christian M. Jones, Lisa M. Jones, and Adam G. Pollack. Influence of bait type on catch rates of predatory fish species on bottom longline gear in the northern Gulf of Mexico. Fishery Bulletin, 115(1):50–59, 2017. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishery-bulletin/influence-bait-type-catch-rates-predatory-
fish-species-bottom-longline-gear;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/driggers.pdf.

Colmenero:2017:RSW

[5197] Ana I. Colmenero, Víctor M. Tuset, and Pilar Sánchez. Reproductive strategy of white anglerfish (Lophius piscatorius) in Mediterranean waters: implications for management. Fishery Bulletin, 115(1):60– 73, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/reproductive-strategy-white-anglerfish-lophius-piscatorius-
mediterranean;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/colmenero.pdf.

Mace:2017:PDS

[5198] Marvin M. Mace III and Lawrence P. Rozas. Population dynamics and secondary production of juvenile white shrimp (Litopenaeus setiferus) along an estuarine salinity gradient. Fishery Bulletin, 115(1):74–88, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fishery-bulletin/population-dynamics-and-secondary-
production-juvenile-white-shrimp;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/mace.pdf.

DellApa:2017:MHD

[5199] Andrea Dell’Apa, Maria Grazia Pennino, and Christopher Bonzek. Modeling the habitat distribution of spiny dogfish (Squalus acanthias), by sex, in coastal waters of the northeastern United States. Fishery Bulletin, 115(1):89–100, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fishery-bulletin/modeling-habitat-distribution-
spiny-dogfish-squalus-acanthias-sex-coastal;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/dellapa.
pdf.

Hueter:2017:HVM

[5200] Robert E. Hueter, John P. Tyminski, John J. Morris, Alexei Ruiz Abierno, and Jorge Angulo Valdes. Horizontal and vertical movements of longfin makos (Isurus paucus) tracked with satellite-linked tags in the northwestern Atlantic Ocean. Fishery Bulletin, 115 (1):101–116, 2017. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishery-bulletin/horizontal-and-vertical-movements-longfin-
makos-isurus-paucus-tracked;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/hueter.pdf.

Ortega-Garcia:2017:AGL

[5201] Sofia Ortega-Garcia, Chugey Sepulveda, Scott Aalbers, Ulianov Jakes-Cota, and Ruben Rodriguez-Sanchez. Age, growth, and length—weight relationship of roosterfish (Nematistius pectoralis) in the eastern Pacific Ocean. Fishery Bulletin, 115(1):117–124, 2017. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/fishery-bulletin/age-growth-and-
length%25E2%2580%2593weight-relationship-roosterfish-nematistius;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/ortegagarcia.pdf.

Bradford:2017:AEC

[5202] Amanda L. Bradford, Karin A. Forney, Erin M. Oleson, and Jay Barlow. Abundance estimates of cetaceans from a line-transect survey within the U.S. Hawaiian Islands Exclusive Economic Zone. Fishery Bulletin, 115(2):129–142, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fishery-bulletin/abundance-estimates-cetaceans-
line-transect-survey-within-us-hawaiian;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/bradford.
pdf.

Stehlik:2017:UGN

[5203] Linda L. Stehlik, John P. Manderson, and Jeffrey Pessutti. Use of gill nets and telemetry in tracking movements and feeding of striped bass (Morone saxatilis), bluefish (Pomatomus saltatrix), and weakfish (Cynoscion regalis) at a salinity front in a small estuary. Fishery Bulletin, 115(2):143–154, 2017. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/fishery-bulletin/use-gill-nets-
and-telemetry-tracking-movements-and-feeding-striped-bass;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/stehlik.pdf.

Winner:2017:AGS

[5204] Brent L. Winner, Timothy C. MacDonald, and Kimberly B. Amendola. Age and growth of sheepshead (Archosargus probatocephalus) in Tampa Bay, Florida. Fishery Bulletin, 115(2):155–166, 2017. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/fishery-bulletin/age-and-growth-
sheepshead-archosargus-probatocephalus-tampa-bay-florida;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/winner.pdf.

Taylor:2017:FHD

[5205] David L. Taylor and Carissa L. Gervasi. Feeding habits and dietary overlap of age-0 winter flounder (Pseudopleuronectes americanus) and summer flounder (Paralichthys dentatus) in southern New England tidal rivers. Fishery Bulletin, 115(2):167–185, 2017. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/fishery-bulletin/feeding-habits-
and-dietary-overlap-age-0-winter-flounder-pseudopleuronectes;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/taylor.pdf.

Sanchez:2017:PCA

[5206] Phillip J. Sanchez, Richard S. Appeldoorn, Michelle T. Schärer-Umpierre, and James V. Locascio. Patterns of courtship acoustics and geophysical features at spawning sites of black grouper (Mycteroperca bonaci). Fishery Bulletin, 115(2):186–195, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/fishery-bulletin/patterns-courtship-acoustics-
and-geophysical-features-spawning-sites-black;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/sanchez.
pdf.

Balazik:2017:FVO

[5207] Matthew Balazik. First verified occurrence of the shortnose sturgeon (Acipenser brevirostrum) in the James River, Virginia. Fishery Bulletin, 115(2):186–195, 2017. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishery-bulletin/first-verified-occurrence-shortnose-sturgeon-
acipenser-brevirostrum-james;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/balazik.pdf.

Lemos:2017:SIT

[5208] Valéria M. Lemos, Cassiano Monteiro-Neto, Henrique Cabral, and João P. Vieira. Stock identification of tainha (Mugil liza) by analyzing stable carbon and oxygen isotopes in otoliths. Fishery Bulletin, 115(2):201–205, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fishery-bulletin/stock-identification-tainha-mugil-
liza-analyzing-stable-carbon-and-oxygen;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/lemos.pdf.

Kupchik:2017:ERT

[5209] Matthew J. Kupchik and Richard F. Shaw. Effects of recruitment through a coastal hydrodynamic boundary layer on growth and otolith microstructure of larval Gulf menhaden (Brevoortia patronus). Fishery Bulletin, 115(2):206–218, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/fishery-bulletin/effects-recruitment-
through-coastal-hydrodynamic-boundary-layer-growth-and;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
kupchik.pdf.

Santos:2017:ELM

[5210] Miguel N. Santos, Pedro G. Lino, and Rui Coelho. Effects of leader material on catches of shallow pelagic longline fisheries in the southwest Indian Ocean. Fishery Bulletin, 115(2):219–232, 2017. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/fishery-bulletin/effects-
leader-material-catches-shallow-pelagic-longline-fisheries;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/santos.pdf.

Sogard:2017:PMG

[5211] Susan M. Sogard and Steven A. Berkeley. Patterns of movement, growth, and survival of adult sablefish (Anoplopoma fimbria) at contrasting depths in slope waters off Oregon. Fishery Bulletin, 115(2):233–251, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/fishery-bulletin/patterns-movement-growth-
and-survival-adult-sablefish-anoplopoma-fimbria;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
sogard.pdf.

Alatorre-Ramirez:2017:TSM

[5212] Vanessa G. Alatorre-Ramirez, Felipe Galván-Magaña, Yassir E. Torres-Rojas, and Robert J. Olson. Trophic segregation of mixed schools of yellowfin tuna (Thunnus albacares) and skipjack tuna (Katsuwonus pelamis) caught in the eastern tropical Pacific Ocean. Fishery Bulletin, 115(2):252–268, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fishery-bulletin/trophic-segregation-mixed-schools-
yellowfin-tuna-thunnus-albacares-and;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/alatorre.pdf.

Walden:2017:MCP

[5213] John B. Walden, Rolf Färe, and Shawna Grosskopf. Measuring change in productivity of a fishery with the Bennet—Bowley indicator. Fishery Bulletin, 115(3):273–283, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fishery-bulletin/measuring-change-productivity-
fishery-bennet%25E2%2580%2593bowley-indicator;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/walden.
pdf.

Brown-Peterson:2017:RDG

[5214] Nancy J. Brown-Peterson, Robert T. Leaf, Amy M. Schueller, and Michael J. Andres. Reproductive dynamics of Gulf menhaden (Brevoortia patronus) in the northern Gulf of Mexico: effects on stock assessments. Fishery Bulletin, 115(3):284–299, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fishery-bulletin/reproductive-dynamics-gulf-menhaden-
brevoortia-patronus-northern-gulf;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/brown_peterson.
pdf.

Kuykendall:2017:MSE

[5215] Kelsey M. Kuykendall, Eric N. Powell, John M. Klinck, Paula T. Moreno, and Robert T. Leaf. Management strategy evaluation for the Atlantic surfclam (Spisula solidissima) using a spatially explicit, vessel-based fisheries model. Fishery Bulletin, 115(3):300–325, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/fishery-bulletin/management-strategy-
evaluation-atlantic-surfclam-spisula-solidissima-using;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
kuykendall.pdf.

Hulson:2017:IDS

[5216] Peter-John F. Hulson, Dana H. Hanselman, and S. Kalei Shotwell. Investigations into the distribution of sample sizes for determining age composition of multiple species. Fishery Bulletin, 115(3):326– 342, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/investigations-distribution-sample-sizes-determining-
age-composition;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/hulson.pdf.

Somerton:2017:QBF

[5217] David A. Somerton, Kresimir Williams, and Matthew D. Campbell. Quantifying the behavior of fish in response to a moving camera vehicle by using benthic stereo cameras and target tracking. Fishery Bulletin, 115(3):343–354, 2017. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishery-bulletin/quantifying-behavior-fish-response-moving-
camera-vehicle-using-benthic;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/somerton.pdf.

LopezQuintero:2017:IUL

[5218] Freddy O. López Quintero, Javier E. Contreras-Reyes, and Rodrigo Wiff. Incorporating uncertainty into a length-based estimator of natural mortality in fish populations. Fishery Bulletin, 115(3):355– 364, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/incorporating-uncertainty-length-based-estimator-natural-
mortality-fish;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/lopez_quintero.pdf.

Molina:2017:AGH

[5219] Juan M. Molina, Gabriela E. Blasina, and Andrea C. Lopez Cazorla. Age and growth of the highly exploited narrownose smooth-hound (Mustelus schmitti) (Pisces: Elasmobranchii). Fishery Bulletin, 115(3): 365–379, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/age-and-growth-highly-exploited-narrownose-smooth-
hound-mustelus-schmitti;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/molina.pdf.

Tolotti:2017:FSV

[5220] Mariana Tolotti, Robert Bauer, Fabien Forget, Pascal Bach, Laurent Dagorn, and Paulo Travassos. Fine-scale vertical movements of oceanic whitetip sharks (Carcharhinus longimanus). Fishery Bulletin, 115(3):380–395, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fishery-bulletin/fine-scale-vertical-movements-
oceanic-whitetip-sharks-carcharhinus;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/tolotti.pdf.

Macchi:2017:ESS

[5221] Gustavo J. Macchi, Karina Rodrigues, Marina V. Diaz, and María I. Militelli. Effects of skipped spawning on the reproductive potential of Argentine hake (Merluccius hubbsi). Fishery Bulletin, 115(3):396– 407, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/effects-skipped-spawning-reproductive-potential-argentine-
hake-merluccius;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/macchi_0.pdf.

Cullen:2017:UUV

[5222] Daniel W. Cullen and Bradley G. Stevens. Use of an underwater video system to record observations of black sea bass (Centropristis striata) in waters off the coast of Maryland. Fishery Bulletin, 115(3):408–418, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/fishery-bulletin/use-
underwater-video-system-record-observations-black-sea-bass;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/cullen.pdf.

Harter:2017:FAA

[5223] Stacey L. Harter, Heather Moe, John K. Reed, and Andrew W. David. Fish assemblages associated with red grouper pits at Pulley Ridge, a mesophotic reef in the Gulf of Mexico. Fishery Bulletin, 115(3):419– 432, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/fish-assemblages-associated-red-grouper-pits-pulley-
ridge-mesophotic-reef;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/harter_0.pdf.

Walsh:2017:SOF

[5224] Harvey J. Walsh and Vince G. Guida. Spring occurrence of fish and macro-invertebrate assemblages near designated wind energy areas on the northeast U.S. continental shelf. Fishery Bulletin, 115(4):437– 450, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/spring-occurrence-fish-and-macro-invertebrate-assemblages-
near-designated;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/walsh_0.pdf.

Gonzalez-Pestana:2017:TES

[5225] Adriana Gonzalez-Pestana, Nicolas Acuña-Perales, Javier Coasaca-Cespedes, Francisco Cordova-Zavaleta, Joanna Alfaro-Shigueto, Jeffrey C. Mangel, and Pepe Espinoza. Trophic ecology of the smooth hammerhead shark (Sphyrna zygaena) off the coast of northern Peru. Fishery Bulletin, 115(4):451–459, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fishery-bulletin/trophic-ecology-smooth-hammerhead-
shark-sphyrna-zygaena-coast-northern-peru;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/gonzalez-
pestana.pdf.

Dippold:2017:GMM

[5226] David A. Dippold, Robert T. Leaf, James S. Franks, and J. Read Hendon. Growth, mortality, and movement of cobia (Rachycentron canadum). Fishery Bulletin, 115(4):460–472, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fishery-bulletin/growth-mortality-and-movement-
cobia-rachycentron-canadum;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/dippold.pdf.

Natanson:2017:GPP

[5227] Lisa J. Natanson, Camilla T. McCandless, Kelsey James, and John Hoey. Gestation period and pupping seasonality of female spiny dogfish (Squalus acanthias) off southern New England. Fishery Bulletin, 115(4): 473–483, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/gestation-period-and-pupping-seasonality-female-spiny-
dogfish-squalus;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/natanson.pdf.

Aucoin:2017:NUS

[5228] Serge Aucoin, Steve Weege, Max Toebe, Jeff Guertin, Jonathan Gorham, and Michael Bresette. A new underwater shark capture method used by divers to catch and release nurse sharks (Ginglymostoma cirratum). Fishery Bulletin, 115(4):484–495, 2017. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/fishery-bulletin/new-underwater-
shark-capture-method-used-divers-catch-and-release-nurse;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/aucoin.pdf.

vonSzalay:2017:MPT

[5229] Paul G. von Szalay and David A. Somerton. A method for predicting trawlability in the Gulf of Alaska with the use of calibrated, split-beam, echosounder backscatter. Fishery Bulletin, 115(4):496– 503, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/method-predicting-trawlability-gulf-alaska-use-calibrated-

split-beam;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/szalay.pdf.

Peters:2017:STP

[5230] Rebecca Peters and Paulinus Chigbu. Spatial and temporal patterns of abundance of juvenile black sea bass (Centropristis striata) in Maryland coastal bays. Fishery Bulletin, 115(4):504–516, 2017. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/fishery-bulletin/spatial-
and-temporal-patterns-abundance-juvenile-black-sea-bass;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
peters.pdf.

delFavero:2017:SRV

[5231] Jana M. del Favero, Mario Katsuragawa, Fabricio S. C. Oliveira, Nayara C. Tannure, M. L. Zani-Teixeira, and Jefferson T. Turner. Seasonal and regional variation in egg size of the Argentine anchoita (Engraulis anchoita) in the southeastern Brazilian Bight. Fishery Bulletin, 115(4): 517–531, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/seasonal-and-regional-variation-egg-size-argentine-
anchoita-engraulis;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/delfavero.pdf.

Whitney:2017:PSR

[5232] Nicholas M. Whitney, Connor F. White, Paul A. Anderson, Robert E. Hueter, and Gregory B. Skomal. The physiological stress response, postrelease behavior, and mortality of blacktip sharks (Carcharhinus limbatus) caught on circle and J-hooks in the Florida recreational fishery. Fishery Bulletin, 115(4):532–543, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fishery-bulletin/physiological-stress-response-
postrelease-behavior-and-mortality-blacktip;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/whitney_
0.pdf.

Kagley:2017:RPM

[5233] Anna N. Kagley, Joseph M. Smith, Kurt L. Fresh, Kinsey E. Frick, and Thomas P. Quinn. Residency, partial migration, and late egress of subadult Chinook salmon (Oncorhynchus tshawytscha) and Coho salmon (O. kisutch) in Puget Sound, Washington. Fishery Bulletin, 115(4): 544–555, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/residency-partial-migration-and-late-egress-subadult-
chinook-salmon;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/kagley.pdf.

Torre:2017:DDA

[5234] Michael P. Torre, Danielle M. Lifavi, and Timothy E. Targett. Diel differences in abundance and diversity of fish species and blue crab (Callinectes sapidus) in the sandy beach shore zone of lower Delaware Bay. Fishery Bulletin, 115(4):556–565, 2017. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fishery-bulletin/diel-differences-abundance-and-diversity-
fish-species-and-blue-crab;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/torre.pdf.

Magel:2017:RVF

[5235] Christopher Magel, Clifford Ryer, and Richard Brill. Recovery of visual function in Pacific halibut (Hippoglossus stenolepis) after exposure to bright light. Fishery Bulletin, 115(4):566–575, 2017. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/fishery-bulletin/recovery-visual-
function-pacific-halibut-hippoglossus-stenolepis-after;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
magel.pdf.

McBride:2018:RFY

[5236] Richard S. McBride, Matthew K. Tweedie, and Kenneth Oliveira. Reproduction, first-year growth, and expansion of spawning and nursery grounds of black sea bass (Centropristis striata) into a warming Gulf of Maine. Fishery Bulletin, 116(3–4):323–336, September 2018. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

Driggers:2018:DAS

[5237] William B. Driggers III, Matthew D. Campbell, David S. Hanisko, Kristin M. Hannan, Eric R. Hoffmayer, Christian M. Jones, Adam G. Pollack, and David S. Portnoy. Distribution of angel sharks (Squatinidae) in United States waters of the western North Atlantic Ocean. Fishery Bulletin, 116(3–4):337–347, October 2018. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

Karlou-Riga:2018:SCO

[5238] Constantina Karlou-Riga, Dimitra Petza, and Konstantinos Ganias. Sex change and oscillating growth pattern of the picarel (Spicara smaris) in the Saronikos Gulf (Greece). Fishery Bulletin, 116(3–4):348–359, October 2018. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

Olaya-Restrepo:2019:EGP

[5239] Julián Olaya-Restrepo, Karim Erzini, and Mercedes González-Wangüemert. Estimation of growth parameters for the exploited sea cucumberHolothuria arguinensis from South Portugal. Fishery Bulletin, 116 (1):1–8, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/estimation-growth-parameters-exploited-sea-cucumber-
holothuria-arguinensis;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/olaya_restrepo.pdf.

Marin-Enriquez:2019:ESP

[5240] Emigdio Marín-Enríquez and Arturo Muhlia-Melo. Environmental and spatial preferences of dolphinfish (Coryphaena spp.) in the eastern Pacific Ocean off the coast of Mexico. Fishery Bulletin, 116(1): 9–20, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/environmental-and-spatial-preferences-dolphinfish-
coryphaena-spp-eastern;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/marin_enriquez.pdf.

Somerton:2019:EWI

[5241] David Somerton, Kenneth Weinberg, Peter Munro, Louis Rugolo, and Thomas Wilderbuer. The effects of wave-induced vessel motion on the geometry of a bottom survey trawl and the herding of yellowfin sole (Limanda aspera). Fishery Bulletin, 116(1):21–33, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/effects-wave-induced-
vessel-motion-geometry-bottom-survey-trawl-and-herding-yellowfin-
sole;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/somerton_0.pdf.

Rincon:2019:EEE

[5242] Margarita M. Rincón, Ignacio A. Catalán, Samu Mäntyniemi, Diego Macías, and Javier Ruiz. Embedding the effect of environmental conditions on recruitment and survival of the European anchovy (Engraulis encrasicolus): a Bayesian model with dual-time resolution. Fishery Bulletin, 116(1):34–49, 2019. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishery-bulletin/embedding-effect-environmental-conditions-

recruitment-and-survival-european;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/rincon1.pdf.

Gong:2019:GVJ

[5243] Yi Gong, Xinjun Chen, Yunkai Li, and Zhou Fang. Geographic variations of jumbo squid (Dosidicus gigas) based on gladius morphology. Fishery Bulletin, 116(1):50–59, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fishery-bulletin/geographic-variations-jumbo-squid-
dosidicus-gigas-based-gladius-morphology;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/gong.pdf.

Sme:2019:DSC

[5244] Noel Sme, Sarah Lyon, Michael Canino, Natalia Chernova, Jason O’Bryhim, Stacey Lance, Kenneth Jones, Franz Mueter, and Anthony Gharrett. Distinction of saffron cod (Eleginus gracilis) from several other gadid species by using microsatellite markers. Fishery Bulletin, 116(1):60–68, 2019. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishery-bulletin/distinction-saffron-cod-eleginus-gracilis-
several-other-gadid-species-using;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/sme1_0.pdf.

Lee:2019:FNS

[5245] Laura M. Lee and Jason E. Rock. The forgotten need for spatial persistence in catch data from fixed-station surveys. Fishery Bulletin, 116(1): 69–74, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/forgotten-need-spatial-persistence-catch-data-fixed-
station-surveys;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/lee.pdf.

Piner:2019:BEG

[5246] Kevin R. Piner, Hui-Hua Lee, and Lennon R. Thomas. Bias in estimates of growth when selectivity in models includes effects of gear and availability of fish. Fishery Bulletin, 116(1):75–80, 2019. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/fishery-bulletin/bias-estimates-
growth-when-selectivity-models-includes-effects-gear-and;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/piner.pdf.

Vasconcelos:2019:IPB

[5247] Joana Vasconcelos, Ana R. Vieira, Vera Sequeira, José A. González, Manfred Kaufmann, and Leonel Serrano Gordo. Identifying populations of the blue jack mackerel (Trachurus picturatus) in the northeast Atlantic by using geometric morphometrics and otolith shape analysis. Fishery Bulletin, 116(1):81–92, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fishery-bulletin/identifying-populations-blue-jack-
mackerel-trachurus-picturatus-northeast;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/vasconcelos.
pdf.

Sanchez-Rubio:2019:OPS

[5248] Guillermo Sanchez-Rubio, Harriet Perry, James S. Franks, and Donald R. Johnson. Occurrence of pelagic Sargassum in waters of the U.S. Gulf of Mexico in response to weather-related hydrographic regimes associated with decadal and interannual variability in global climate. Fishery Bulletin, 116(1):93–106, 2019. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishery-bulletin/occurrence-pelagic-sargassum-waters-us-gulf-
mexico-response-weather-related;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/sanchez_rubio.pdf.

French:2019:SRB

[5249] Kiyomi J. French, Nancy L. Shackell, and Cornelia E. den Heyer. Strong relationship between commercial catch of adult Atlantic halibut (Hippoglossus hippoglossus) and availability of suitable habitat for juveniles in the Northwest Atlantic Ocean. Fishery Bulletin, 116(2):111– 125, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/strong-relationship-between-commercial-catch-adult-
atlantic-halibut;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/french_0.pdf.

Yochum:2019:CLH

[5250] Noëlle Yochum, Allan W. Stoner, David B. Sampson, and Craig Rose. A comparison of laboratory-holding and tag-return methods for evaluating delayed mortality of Dungeness crab (Cancer magister) discarded in Oregon fisheries. Fishery Bulletin, 116(2):126– 141, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/comparison-laboratory-holding-and-tag-return-methods-

evaluating-delayed;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/yochum_0.pdf.

Weaver:2019:INC

[5251] Daniel M. Weaver, Stephen M. Coghlan, Jr., and Joseph Zydlewski. The influence of nutrients from carcasses of sea lamprey (Petromyzon marinus) on larval growth and spawner abundance. Fishery Bulletin, 116(2):142–152, 2019. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishery-bulletin/influence-nutrients-carcasses-sea-lamprey-
petromyzon-marinus-larval-growth;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/weaver.pdf.

Zhukova:2019:GMF

[5252] Kristina Zhukova, Olga A. Maznikova, and Alexei M. Orlov. Gonadal maturation of the female smooth lumpsucker (Aptocyclus ventricosus). Fishery Bulletin, 116(2):153–160, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/fishery-bulletin/gonadal-
maturation-female-smooth-lumpsucker-aptocyclus-ventricosus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/zhukova.pdf.

Clark:2019:UVS

[5253] Abigail S. Clark, Steven H. Jury, Jason S. Goldstein, Thomas G. Langley, and Winsor H. Watson III. Underwater video surveillance of American lobsters (Homarus americanus) to understand saturation levels in lobster traps. Fishery Bulletin, 116(2):161– 170, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/underwater-video-surveillance-american-lobsters-homarus-
americanus;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/clark.pdf.

Wang:2019:APV

[5254] Hui-Yu Wang and Mikko Heino. Adaptive and plastic variation in growth and maturation of the cutlassfishTrichiurus japonicus in the subtropical Pacific Ocean. Fishery Bulletin, 116(2):

171–182, 2019.  CODEN FSYBAY.  ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishery-bulletin/adaptive-and-plastic-variation-growth-and-
maturation-cutlassfish-trichiurus;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/wang.pdf.

Ayyildiz:2019:AGR

[5255] Hakan Ayyildiz and Aytac Altin. Age and growth rates at the early life stages of common pandora (Pagellus erythrinus) based on analysis of otolith microstructure. Fishery Bulletin, 116(2):183–189, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/fishery-bulletin/age-and-growth-
rates-early-life-stages-common-pandora-pagellus-erythrinus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/ayyildiz.pdf.

Wetzel:2019:ERD

[5256] Chantel R. Wetzel, André E. Punt, and Jason M. Cope. The effect of reduced data on the ability to monitor rebuilding of overfished fish stocks. Fishery Bulletin, 116(2):190–206, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/fishery-bulletin/effect-reduced-data-
ability-monitor-rebuilding-overfished-fish-stocks;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
wetzel.pdf.

Crow:2019:VDD

[5257] Gerald L. Crow, Bradley M. Wetherbee, Robert L. Humphreys, Jr., and Richard Young. Vertical distribution, diet, and reproduction of the velvet dogfish (Zameus squamulosus) in waters off Hawaii. Fishery Bulletin, 116 (2):207–214, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/vertical-distribution-diet-and-reproduction-velvet-
dogfish-zameus;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/crow.pdf.

Fox:2019:SOA

[5258] Adam G. Fox, Edward S. Stowe, Keith J. Dunton, and Douglas L. Peterson. Seasonal occurrence of Atlantic sturgeon (Acipenser oxyrinchus oxyrinchus) in the St. Johns River, Florida. Fishery Bulletin, 116(3–4): 219–227, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/seasonal-occurrence-atlantic-sturgeon-acipenser-oxyrinchus-
oxyrinchus-st;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/fox_0.pdf.

Woodworth-Jefcoats:2019:SAO

[5259] Phoebe A. Woodworth-Jefcoats, Jeffrey J. Polovina, and Jeffrey C. Drazen. Synergy among oceanographic variability, fishery expansion, and longline catch composition in the central North Pacific Ocean. Fishery Bulletin, 116(3–4):228–239, 2019. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/fishery-bulletin/synergy-among-
oceanographic-variability-fishery-expansion-and-longline;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/woodworth-jefcoats_1.pdf.

Schobernd:2019:EFL

[5260] Christina M. Schobernd, M. Conor McManus, Joanne Lyczkowski-Shultz, Nathan M. Bacheler, and Denice M. Drass. Extrusion of fish larvae from SEAMAP plankton sampling nets: a comparison between 0.333-mm and 0.202-mm mesh nets. Fishery Bulletin, 116(3–4):240– 253, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/extrusion-fish-larvae-seamap-plankton-sampling-nets-
comparison-between-0333;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/schobernd_0.pdf.

Oldemeyer:2019:MHB

[5261] Bryce N. Oldemeyer, Timothy Copeland, and Brian P. Kennedy. A multiyear hierarchical Bayesian mark-recapture model incorporating data on recurring salmonid behavior to account for sparse or missing data. Fishery Bulletin, 116(3–4):254–265, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fishery-bulletin/multiyear-hierarchical-bayesian-
mark-recapture-model-incorporating-data;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/oldemeyer_
0.pdf.

Johansson:2019:SDL

[5262] Mattias L. Johansson, Marisa N. C. Litz, Richard D. Brodeur, Tristan A. Britt, Camilo A. Vanegas, John R. Hyde, and Michael A. Banks. Seasonal distribution of late larval and juvenile rockfish (Sebastes spp.) and associated environmental conditions off Oregon and Washington: new insights based on genetics. Fishery Bulletin, 116(3–4):266–280, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fishery-bulletin/seasonal-distribution-late-larval-
and-juvenile-rockfish-sebastes-spp-and;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/johansson.
pdf.

Maynard:2019:MMA

[5263] George A. Maynard, Lisa K. Izzo, and Joseph D. Zydlewski. Movement and mortality of Atlantic salmon kelts (Salmo salar) released into the Penobscot River, Maine. Fishery Bulletin, 116(3–4):281– 290, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/movement-and-mortality-atlantic-salmon-kelts-salmo-
salar-released-penobscot;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/maynard.pdf.

Watters:2019:CLD

[5264] Diana L. Watters and E. J. Dick. A comparison of length distributions of rockfishes (Sebastes spp.) from submersible and trawl surveys off central California. Fishery Bulletin, 116(3–4):291–301, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/comparison-length-distributions-rockfishes-sebastes-
spp-submersible-and;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/watters.pdf.

Fey:2019:DDG

[5265] Dariusz P. Fey, Adam M. Lejk, and Martyna Greszkiewicz. Daily deposition of growth increments in sagittae and lapilli of laboratoryreared larval northern pike (Esox lucius). Fishery Bulletin, 116(3–4): 302–309, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/daily-deposition-growth-increments-sagittae-and-lapilli-
laboratory-reared;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/fey_0.pdf.

Cordova-Zavaleta:2019:FHB

[5266] Francisco Córdova-Zavaleta, Jaime Mendo, Sergio A. Briones-Hernández, Nicolas Acuña-Perales, Adriana Gonzalez-Pestana, Joanna Alfaro-Shigueto, and Jeffrey C. Mangel. Food habits of the blue shark, Prionace glauca (Linnaeus, 1758), in waters off northern Peru. Fishery Bulletin, 116(3–4):310–322, 2019. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishery-bulletin/food-habits-blue-shark-prionace-glauca-linnaeus-
1758-waters-northern-peru;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/cordova-zavaleta.pdf.

Powers:2019:CRF

[5267] Sean P. Powers and Kevin Anson. Compression and relaxation of fishing effort in response to changes in length of fishing season for red snapper (Lutjanus campechanus) in the northern Gulf of Mexico. Fishery Bulletin, 117(1–2):1–7, 2019. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishery-bulletin/compression-and-relaxation-fishing-effort-
response-changes-length-fishing;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/powers_0_0.pdf.

Stevenson:2019:IBF

[5268] Duane E. Stevenson, Gerald R. Hoff, James W. Orr, Ingrid Spies, and Christopher N. Rooper. Interactions between fisheries and early life stages of skates in nursery areas of the eastern Bering Sea. Fishery Bulletin, 117(1–2):8–14, 2019. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishery-bulletin/interactions-between-fisheries-and-early-life-
stages-skates-nursery-areas;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/stevenson_0.pdf.

Orbesen:2019:VPM

[5269] Eric S. Orbesen, Craig A. Brown, Derke Snodgrass, Joseph E. Serafy, and John F. Walter III. At-vessel and postrelease mortality rates of bluefin tuna (Thunnus thynnus) associated with pelagic longline gear in the northern Gulf of Mexico. Fishery Bulletin, 117(1–2):15– 23, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/vessel-and-postrelease-mortality-rates-bluefin-tuna-
thunnus-thynnus;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/orbesen.pdf.

Misawa:2019:PSO

[5270] Ryo Misawa, Yoji Narimatsu, Hiromitsu Endo, and Yoshiaki Kai. Population structure of the ocellate spot skate (Okamejei kenojei) inferred from variations in mitochondrial DNA (mtDNA) sequences and from morphological characters of regional populations. Fishery Bulletin, 117 (1–2):24–36, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/population-structure-ocellate-spot-skate-okamejei-
kenojei-inferred;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/misawa.pdf.

Brown-Peterson:2019:MMA

[5271] Nancy J. Brown-Peterson, Christopher R. Peterson, and Gary R. Fitzhugh. Multidecadal meta-analysis of reproductive parameters of female red snapper (Lutjanus campechanus) in the northern Gulf of Mexico. Fishery Bulletin, 117(1–2):37–49, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/fishery-bulletin/multidecadal-meta-analysis-
reproductive-parameters-female-red-snapper;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/brown-peterson.
pdf.

Natanson:2019:UOV

[5272] Lisa J. Natanson and Bethany M. Deacy. Using oxytetracycline validation for confirmation of changes in vertebral band-pair deposition rates with ontogeny in sandbar sharks (Carcharhinus plumbeus) in the western North Atlantic Ocean. Fishery Bulletin, 117(1–2):50– 58, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/using-oxytetracycline-validation-confirmation-changes-
vertebral-band-pair;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/natanson_0.pdf.

Schlick:2019:GAR

[5273] C. J. Carroll Schlick and Kim de Mutsert. Growth of adult river herring that spawn in tributaries of the Potomac River in northern Virginia. Fishery Bulletin, 117(1–2):59–69, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fishery-bulletin/growth-adult-river-herring-spawn-tributaries-
potomac-river-northern;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/schick_0.pdf.

Natanson:2019:PRP

[5274] Lisa J. Natanson, Bethany M. Deacy, Warren Joyce, and James Sulikowski. Presence of a resting population of female porbeagles (Lamna nasus), indicating a biennial reproductive cycle, in the western North Atlantic Ocean. Fishery Bulletin, 117(1–2):70– 77, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/presence-resting-population-female-porbeagles-lamna-
nasus-indicating;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/natanson2_0.pdf.

Matsumoto:2019:SMM

[5275] Rui Matsumoto, Yosuke Matsumoto, Keiichi Ueda, Miwa Suzuki, Kiyoshi Asahina, and Keiichi Sato. Sexual maturation in a male whale shark (Rhincodon typus) based on observations made over 20 years of captivity. Fishery Bulletin, 117(1–2):78–86, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fishery-bulletin/sexual-maturation-male-whale-shark-
rhincodon-typus-based-observations-made;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/matsumoto_
1_0.pdf.

Munoz:2019:CVD

[5276] Roldan C. Muñoz and Michael L. Burton. Comparison of video and diver observations of sharks from a fishery-independent trapvideo survey off east-central Florida, including utility of an alternative method of video analysis. Fishery Bulletin, 117(1–2):87– 96, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/comparison-video-and-diver-observations-sharks-fishery-
independent-trap;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/munoz_0.pdf.

VanDoornik:2019:IGA

[5277] Donald M. Van Doornik, David R. Kuligowski, Cheryl A. Morgan, Brian J. Burke, and Todd R. Seamons. Insights, from genetic analyses, into stock-specific distribution of juvenile steelhead (Oncorhynchus mykiss) from the Columbia River during early marine migration. Fishery Bulletin, 117(1–2):97–106, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fishery-bulletin/insights-genetic-analyses-stock-specific-
distribution-juvenile-steelhead;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/vandoornik.pdf.

Fujinami:2019:ADG

[5278] Yuki Fujinami, Yasuko Semba, and Sho Tanaka. Age determination and growth of the blue shark (Prionace glauca) in the western North Pacific Ocean. Fishery Bulletin, 117(1–2):107–120, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/fishery-bulletin/age-
determination-and-growth-blue-shark-prionace-glauca-western-
north;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/fujinami_0_0.pdf.

Fogg:2019:CAG

[5279] Alexander Q. Fogg, Joseph T. Evans, Mark S. Peterson, Nancy J. Brown-Peterson, Eric R. Hoffmayer, and G. Walter Ingram, Jr. Comparison of age and growth parameters of invasive red lionfish (Pterois volitans) across the northern Gulf of Mexico. Fishery Bulletin, 117(3): 1–15, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/comparison-age-and-growth-parameters-invasive-red-
lionfish-pterois-volitans;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/fogg_0.pdf.

Conrath:2019:RPL

[5280] Christine L. Conrath. Reproductive potential of light dusky rockfish (Sebastes variabilis) and northern rockfish (S. polyspinis) in the Gulf of Alaska. Fishery Bulletin, 117(3):16–26, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/fishery-bulletin/reproductive-potential-light-
dusky-rockfish-sebastes-variabilis-and;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/conrath_
0_0.pdf.

Chase:2019:FRL

[5281] Bradford C. Chase, Scott Elzey, Sara M. Turner, and Matt H. Ayer. Fecundity and reproductive life history of anadromous rainbow smelt (Osmerus mordax) in coastal waters of Massachusetts. Fishery Bulletin, 117 (3):27–44, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/fecundity-and-reproductive-life-history-anadromous-
rainbow-smelt-osmerus;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/chase_0_0.pdf.

Torres-Palacios:2019:AGP

[5282] Katherine Torres-Palacios, Paola A. Mejía-Falla, Andrés F. Navia, Víctor H. Cruz-Escalona, Roberto Félix-Uraga, and Casimiro Quiñonez-Velázquez. Age and growth parameters of the Panamic stingray (Urotrygon aspidura). Fishery Bulletin, 117(3):45–55, 2019. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/fishery-bulletin/age-and-growth-
parameters-panamic-stingray-urotrygon-aspidura;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/torres_
palacios_0.pdf.

Holder:2019:EFR

[5283] Anna M. Holder and John C. Field. An exploration of factors that relate to the occurrence of multiple brooding in rockfishes (Sebastes spp.). Fishery Bulletin, 117(3):56–64, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/fishery-bulletin/exploration-factors-
relate-occurrence-multiple-brooding-rockfishes-sebastes;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
holder_0_0.pdf.

Driggers:2019:ECB

[5284] William B. Driggers III and Kristin M. Hannan. Efficacy of 2 common bait types in reducing bycatch of coastal sharks on bottom longline gear in the absence of choice. Fishery Bulletin, 117(3):65– 71, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/efficacy-2-common-bait-types-reducing-bycatch-coastal-
sharks-bottom;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/driggers_1.pdf.

Jacobson:2019:UPC

[5285] Kym Jacobson, Rebecca Baldwin, Michael Banks, and Robert Emmett. Use of parasites to clarify residency and migration patterns of Pacific sardine (Sardinops sagax) in the California Current primary tabs. Fishery Bulletin, 117(3):196–210, 2019. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/fishery-bulletin/use-parasites-
clarify-residency-and-migration-patterns-pacific-sardine;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/jacobson_0.pdf.

Watson:2019:EAM

[5286] Winsor H. Watson III, Elizabeth Morrissey, Stephanie Sykes, Abigail S. Clark, and Steven Jury. Experimental analyses of the mechanisms leading to American lobster (Homarus americanus) ventless trap saturation. Fishery Bulletin, 117(3):211–219, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fishery-bulletin/experimental-analyses-mechanisms-
leading-american-lobster-homarus;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/watson.pdf.

Perkinson:2019:ESS

[5287] Matt Perkinson, Tanya Darden, Maggie Jamison, Matt J. Walker, Michael R. Denson, James Franks, Read Hendon, Susanna Musick, and Eric S. Orbesen. Evaluation of the stock structure of cobia (Rachycentron canadum) in the southeastern United States by using dart-tag and genetics data. Fishery Bulletin, 117(3):220– 233, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/evaluation-stock-structure-cobia-rachycentron-canadum-
southeastern-united;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/Perkinson_0.pdf.

Harris:2019:UOM

[5288] Jeremy P. Harris, Charles Hutchinson, and Sharon Wildes. Using otolith morphometric analysis to improve species discrimination of blackspotted rockfish (Sebastes melanostictus) and rougheye rockfish (S. aleutianus). Fishery Bulletin, 117(3):234–244, 2019. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/fishery-bulletin/using-otolith-
morphometric-analysis-improve-species-discrimination;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
harris.pdf.

Mendez-Macias:2019:TEO

[5289] José S. Méndez-Macías, Vanessa M. Velázquez-Chiquito, Colombo Estupiñán-Montaño, and Felipe Galván-Magaña. Trophic ecology and ontogenetic shift in the diet of the sicklefin smoothhound (Mustelus lunulatus) in the southeastern Pacific Ocean. Fishery Bulletin, 117(3): 245–257, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/trophic-ecology-and-ontogenetic-shift-diet-sicklefin-
smoothhound-mustelus;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/mendez_macias_0.pdf.

Goetz:2019:BTA

[5290] Frederick A. Goetz and Thomas P. Quinn. Behavioral thermoregulation by adult Chinook salmon (Oncorhynchus tshawytscha) in estuary and freshwater habitats prior to spawning. Fishery Bulletin, 117(3):258–271, 2019. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishery-bulletin/behavioral-thermoregulation-adult-chinook-
salmon-oncorhynchus-tshawytscha;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/goetz_0.pdf.

Schmitt:2019:MPD

[5291] Joseph D. Schmitt, Brandon K. Peoples, Aaron J. Bunch, Leandro Castello, and Donald J. Orth. Modeling the predation dynamics of invasive blue catfish (Ictalurus furcatus) in Chesapeake Bay. Fishery Bulletin, 117(4):277–290, 2019. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishery-bulletin/modeling-predation-dynamics-invasive-blue-
catfish-ictalurus-furcatus;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/schmitt_1_0.pdf.

Lefebvre:2019:RES

[5292] Lyndsey S. Lefebvre, Cherisa L. Friedlander, and John C. Field. Reproductive ecology and size-dependent fecundity in the petrale sole (Eopsetta jordani) in waters of California, Oregon, and Washington. Fishery Bulletin, 117(4):291–302, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fishery-bulletin/reproductive-ecology-and-size-
dependent-fecundity-petrale-sole-eopsetta;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/lefebvre_
0_0.pdf.

DAlessandro:2019:SPI

[5293] Dalin N. D’Alessandro and Deborah A. Duffield. Salmonid passive integrated transponder tags and coded wire tags found in the forestomach of a harbor porpoise (Phocoena phocoena) in southwestern Washington primary tabs. Fishery Bulletin, 117(4):303– 307, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/salmonid-passive-integrated-transponder-tags-and-coded-
wire-tags-found;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/d_alessandro_0.pdf.

Kolmos:2019:TCL

[5294] Kevin J. Kolmos, David M. Wyanski, D. Byron White, and Paulette P. Mikell. Temporal changes in the life history of snowy grouper (Hyporthodus niveatus) off North and South Carolina, and factors that influence spawning dynamics. Fishery Bulletin, 117(4):308– 321, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/temporal-changes-life-history-snowy-grouper-hyporthodus-
niveatus-north-and;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/Kolmos_0.pdf.

OMalley:2019:EEE

[5295] Joseph M. O’Malley, Corey B. Wakefield, Zack S. Oyafuso, Ryan S. Nichols, Brett Taylor, Ashley J. Williams, Maria Sapatu, and Michael Marsik. Effects of exploitation evident in age-based demography of 2 deepwater snappers, the goldeneye jobfish (Pristipomoides flavipinnis) in the Samoa Archipelago and the goldflag jobfish (P. auricilla) in the Mariana Archipelago. Fishery Bulletin, 117(4):322– 336, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/effects-exploitation-evident-age-based-demography-
2-deepwater-snappers;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/omalley_0_0.pdf.

Anweiler:2019:ETH

[5296] Katie V. Anweiler, Karl Brenkert, Tanya L. Darden, Eric J. McElroy, and Michael R. Denson. Effects of temperature and hypoxia on the metabolic performance of juvenile striped bass (Morone saxatilis). Fishery Bulletin, 117(4):337–347, 2019. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishery-bulletin/effects-temperature-and-hypoxia-metabolic-
performance-juvenile-striped-bass;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/anweiler_0.pdf.

Lear:2019:FSB

[5297] Karissa O. Lear, Gregg R. Poulakis, Rachel M. Scharer, Adrian C. Gleiss, and Nicholas M. Whitney. Fine-scale behavior and habitat use of the endangered smalltooth sawfish (Pristis pectinata): insights from accelerometry. Fishery Bulletin, 117(4):348–359, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/fishery-bulletin/fine-scale-behavior-
and-habitat-use-endangered-smalltooth-sawfish-pristis;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
lear_0.pdf.

May:2019:AMF

[5298] Rebecca K. May, S. Gregory Tolley, Rachel M. Scharer, Bass Dye, Felix Jose, and Gregg R. Poulakis. Automated monitoring of fine-scale movements of the endangered smalltooth sawfish (Pristis pectinata). Fishery Bulletin, 117(4):360–371, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fishery-bulletin/automated-monitoring-fine-scale-movements-
endangered-smalltooth-sawfish;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/may_0.pdf.

Brown:2019:SGM

[5299] H. Kevin Brown, Blake Price, Laura Lee, M. Scott Baker, Jr., and Sara E. Mirabilio. Simple gear modifications for achieving greater than 40% bycatch reduction in an estuarine shrimp trawl fishery. Fishery Bulletin, 117 (4):372–385, 2019. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/simple-gear-modifications-achieving-greater-40-bycatch-
reduction-estuarine;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/brown_0.pdf.

Murakumo:2020:PUO

[5300] Kiyomi Murakumo, Rui Matsumoto, Taketeru Tomita, Yousuke Matsumoto, and Keiichi Ueda. The power of ultrasound: observation of nearly the entire gestation and embryonic developmental process of captive reef manta rays (Mobula alfredi). Fishery Bulletin, 118(1): 1–8, 2020. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/power-ultrasound-observation-nearly-entire-gestation-
and-embryonic;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/murakumo_0.pdf.

Bank:2020:AVG

[5301] Cristiana M. Bank, Kenneth Oliveira, Sandra J. Sutherland, Michael P. Armstrong, Jorge Landa, and Steven X. Cadrin. Age validation for goosefish (Lophius americanus) in the northeastern United States. Fishery Bulletin, 118(1):8–20, 2020. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/fishery-bulletin/age-validation-goosefish-
lophius-americanus-northeastern-united-states;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/bank_
0_0.pdf.

Natanson:2020:URP

[5302] Lisa J. Natanson, Megan Winton, Heather Bowlby, Warren Joyce, Bethany Deacy, Rui Coelho, and Daniela Rosa. Updated reproductive parameters for the shortfin mako (Isurus oxyrinchus) in the North Atlantic Ocean with inferences of distribution by sex and reproductive stage. Fishery Bulletin, 118(1):21–36, 2020. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/fishery-bulletin/updated-
reproductive-parameters-shortfin-mako-isurus-oxyrinchus-north;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/natanson_2_0.pdf.

Bizzarro:2020:GIB

[5303] Joseph J. Bizzarro, Elizabeth A. Gilbert-Horvath, E. J. Dick, Aaron M. Berger, Katherine T. Schmidt, Don Pearson, Christine Petersen, Lisa A. Kautzi, Rebecca R. Miller, John C. Field, and John Carlos Garza. Genetic identification of blue rockfish (Sebastes mystinus) and deacon rockfish (S. diaconus) to enable life history analyses for stock assessment. Fishery Bulletin, 118(1):37–50, 2020. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/fishery-bulletin/genetic-
identification-blue-rockfish-sebastes-mystinus-and-deacon-rockfish;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/bizzarro_0_0.pdf.

Driggers:2020:SVF

[5304] William B. Driggers III, Matthew D. Campbell, Kristin M. Hannan, Eric R. Hoffmayer, Christian M. Jones, and James A. Sulikowski. Spatial variability in the fecundity of Atlantic sharpnose sharks (Rhizoprionodon terraenovae) in the northern Gulf of Mexico. Fishery Bulletin, 118(1):51–62, 2020. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fishery-bulletin/spatial-variability-fecundity-
atlantic-sharpnose-sharks-rhizoprionodon;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/driggers_
3_0.pdf.

Doukakis:2020:PSG

[5305] Phaedra Doukakis, Ethan A. Mora, Susan Wang, Paul Reilly, Russ Bellmer, Kristine Lesyna, Travis Tanaka, Natnael Hamda, Mary L. Moser, Daniel L. Erickson, Jason Vestre, Jon McVeigh, Kevin Stockmann, Kristen Duncan, and Steven T. Lindley. Postrelease survival of green sturgeon (Acipenser medirostris) encountered as bycatch in the trawl fishery that targets California halibut (Paralichthys californicus), estimated by using pop-up satellite archival tags. Fishery Bulletin, 118 (1):63–73, 2020. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/postrelease-survival-green-sturgeon-acipenser-medirostris-
encountered;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/doukakis_0.pdf.

Markin:2020:GJA

[5306] Erin L. Markin and David H. Secor. Growth of juvenile Atlantic sturgeon (Acipenser oxyrinchus oxyrinchus) in response to dual-season spawning and latitudinal thermal regimes. Fishery Bulletin, 118(1):74–86, 2020. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/fishery-bulletin/growth-juvenile-atlantic-
sturgeon-acipenser-oxyrinchus-oxyrinchus-response;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
markin_0.pdf.

Cahalan:2020:DIF

[5307] Jennifer Cahalan and Craig Faunce. Development and implementation of a fully randomized sampling design for a fishery monitoring program. Fishery Bulletin, 118(1):87–99, 2020. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/development-and-implementation-fully-
randomized-sampling-design-fishery-monitoring-program;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
cahalan_0.pdf.

Zemeckis:2020:DMB

[5308] Douglas R. Zemeckis, Jeff Kneebone, Connor W. Capizzano, Eleanor A. Bochenek, William S. Hoffman, Thomas M. Grothues, John W. Mandelman, and Olaf P. Jensen. Discard mortality of black sea bass (Centropristis striata) in a deepwater recreational fishery off New Jersey: role of swim bladder venting in reducing mortality. Fishery Bulletin, 118(2):105–119, 2020. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/discard-mortality-black-sea-bass-
centropristis-striata-deepwater-recreational-fishery-new;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/zemeckis_0.pdf.

Richardson:2020:CSD

[5309] David E. Richardson, Lauren Carter, Kiersten L. Curti, Katrin E. Marancik, and Martin Castonguay. Changes in the spawning distribution and biomass of Atlantic mackerel (Scomber scombrus) in the western Atlantic Ocean over 4 decades. Fishery Bulletin, 118(2):120–134, 2020. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/changes-spawning-distribution-and-biomass-atlantic-
mackerel-scomber-scombrus-western;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/richardson_1.
pdf.

Bolshakova:2020:ICN

[5310] Yana Yu. Bolshakova and Sergei A. Evseenko. Ichthyoplankton of the central North Atlantic Ocean: larval development of lanternfish species (Pisces: Myctophidae). Fishery Bulletin, 118(2):135– 144, 2020. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/ichthyoplankton-central-north-atlantic-ocean-larval-
development-lanternfish;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/bolshakova_1.pdf.

Langan:2020:ADA

[5311] Joseph A. Langan, M. Conor McManus, Douglas R. Zemeckis, and Jeremy S. Colli. Abundance and distribution of Atlantic cod (Gadus morhua) in a warming southern New England. Fishery Bulletin, 118(2): 145–156, 2020. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/abundance-
and-distribution-atlantic-cod-gadus-morhua-warming-southern-
new-england;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/langan_0.pdf.

Ben-Aderet:2020:RLH

[5312] Noah Ben-Aderet, Erin M. Johnston, Richard Cravey, and Stuart A. Sandin. Revisiting the life history of yellowtail jack (Seriola dorsalis) in the Southern California Bight: new evidence for ontogenetic habitat shifts and regional differences in a changing environment. Fishery Bulletin, 118(2):157–170, 2020. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
revisiting-life-history-yellowtail-jack-seriola-dorsalis-southern-
california-bight-new;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/ben_aderet.pdf.

Eayrs:2020:DTU

[5313] Stephen Eayrs, Michael Pol, Jon Knight, and Jim Ford. Development and testing of an ultra-low-opening trawl to improve utilization of quota of Atlantic cod (Gadus morhua) in the groundfish fishery of New England. Fishery Bulletin, 118(2):171–183, 2020. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/development-and-testing-ultra-low-opening-
trawl-improve-utilization-quota-atlantic-cod;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/eayrs_
0.pdf.

Orphanides:2020:DHP

[5314] Christopher D. Orphanides, Frederick W. Wenzel, and Jeremy S. Collie. Diet of harbor porpoises (Phocoena phocoena) on the continental shelf off southern New England. Fishery Bulletin, 118(2):184–197, 2020. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/diet-harbor-porpoises-phocoena-
phocoena-continental-shelf-southern-new-england;   https://
spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
orphanides_0.pdf.

Ingram:2020:AES

[5315] Evan C. Ingram, Douglas L. Peterson, and Adam G. Fox. Abundance of endangered shortnose sturgeon (Acipenser brevirostrum) in the Altamaha River in Georgia. Fishery Bulletin, 118(2):198–204, 2020. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/abundance-endangered-
shortnose-sturgeon-acipenser-brevirostrum-altamaha-river-georgia;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/ingram_0.pdf.

Scherrer:2020:EMO

[5316] Stephen R. Scherrer and Kevin C. Weng. Evaluating movements of opakapaka (Pristipomoides filamentosus) relative to a restricted fishing area by using acoustic telemetry and a depth-constrained estimator of linear home ranges. Fishery Bulletin, 118(3):209– 224, 2020. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/evaluating-movements-opakapaka-pristipomoides-filamentosus-
relative;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/1183scherrer.pdf.

Owings:2020:EBB

[5317] Meghan Owings, Christopher Chabot, and Winsor Watson III. Effects of the biomedical bleeding process on the behavior and hemocyanin levels of the American horseshoe crab (Limulus polyphemus). Fishery Bulletin, 118(3):225–239, 2020. CODEN FSY-

BAY.  ISSN 0090-0656 (print), 1937-4518 (electronic).  URL

https://spo.nmfs.noaa.gov/content/fishery-bulletin/effects-
biomedical-bleeding-process-behavior-and-hemocyanin-levels-
american;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-
content/fish-bull/1183owings.pdf.

Trujillo-Cordova:2020:UAS

[5318] Jorge A. Trujillo-Córdova, Emanuel Mimila-Herrera, Natalí Cárdenas-Palomo, and Jorge A. Herrera-Silveira. Use of aerial surveys for assessing abundance of the whale shark (Rhincodon typus) and the giant manta (Mobula birostris) in the northern Caribbean Sea off Mexico. Fishery Bulletin, 118(3):240–249, 2020. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/fishery-bulletin/use-
aerial-surveys-assessing-abundance-whale-shark-rhincodon-typus-
and;   https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/1183trujillo_cordova.pdf.

Kennedy:2020:IIV

[5319] James Kennedy and Halldór G. Ólafsson. Intra- and interannual variation in fecundity and egg size of lumpfish (Cyclopterus lumpus) in Iceland. Fishery Bulletin, 118(3):250–269, 2020. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fishery-bulletin/intra-and-interannual-variation-
fecundity-and-egg-size-lumpfish-cyclopterus;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/1183kennedy.
pdf.

Cook:2020:EFD

[5320] Melissa Cook, Jaymie L. Reneker, Redwood W. Nero, Brian A. Stacy, and David S. Hanisko. Effects of freezing on decomposition of sea turtle carcasses used for research studies. Fishery Bulletin, 118(3):268–274, 2020. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishery-bulletin/effects-freezing-decomposition-sea-turtle-
carcasses-used-research-studies;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/1183cook.pdf.

Mercaldo-Allen:2020:TRC

[5321] Renee Mercaldo-Allen, Paul Clark, Dylan Redman, Yuan Liu, Shannon Meseck, Lisa Milke, Julie M. Rose, and Ronald Goldberg. Temperature-related changes in species composition of juvenile finfish on a rock reef in Long Island Sound. Fishery Bulletin, 118(3):275– 283, 2020. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/temperature-related-changes-species-composition-juvenile-
finfish-rock-reef;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/1183mercaldo_allen.pdf.

Ardini:2020:CSM

[5322] Greg Ardini, Chad Demarest, and Katherine McArdle. Costs of atsea monitoring under government and private contracts in the groundfish fishery of the northeastern United States. Fishery Bulletin, 118(3): 284–296, 2020. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/costs-sea-monitoring-under-government-and-private-
contracts-groundfish;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/1183ardini.pdf.

Weber:2020:SRP

[5323] D. Nick Weber, Bryan S. Frazier, Nicholas M. Whitney, James Gelsleichter, and Gorka Sancho. Stress response and postrelease mortality of blacktip sharks (Carcharhinus limbatus) captured in shorebased and charter-boat-based recreational fisheries. Fishery Bulletin, 118(3):297–314, 2020. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishery-bulletin/stress-response-and-postrelease-mortality-
blacktip-sharks-carcharhinus;   https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-bull/1183weber.pdf.

Evseenko:2020:DLT

[5324] Sergei A. Evseenko (deceased) and Yana Yu. Bolshakova. A description of late and transformation stages of larvae of the Diaphus ostenfeldi in the western periphery of the South Pacific gyre. Fishery Bulletin, 118(4):319– 323, 2020. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/description-
late-and-transformation-stages-larvae-diaphus-ostenfeldi-western-
periphery;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/1184evseenko.pdf.

Jakes-Cota:2020:AGS

[5325] Ulianov Jakes-Cota, Arturo Tripp-Valdez, Marisol Arce-Acosta, and Francisco Omar López-Fuerte. Age and growth of stone scorpionfish (Scorpaena mystes) from the Gulf of California in Mexico. Fishery Bulletin, 118(4):324–328, 2020. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/fishery-bulletin/age-and-growth-
stone-scorpionfish-scorpaena-mystes-gulf-california-mexico;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/1184_jakes_cota.pdf.

Frazier:2020:GRB

[5326] Bryan S. Frazier, Dana M. Bethea, Robert E. Hueter, Camilla T. Mc-Candless, John P. Tyminski, and William B. Driggers III. Growth rates of bonnetheads (Sphyrna tiburo) estimated from tag-recapture data. Fishery Bulletin, 118(4):329–345, 2020. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/fishery-bulletin/growth-rates-bonnetheads-
sphyrna-tiburo-estimated-tag-recapture-data;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/1184frazier.
pdf.

Kemberling:2020:DRA

[5327] Adam A. Kemberling and M. Zachary Darnell. Distribution, relative abundance, and reproductive output of spawning female blue crab (Callinectes sapidus) in offshore waters of the Gulf of Mexico. Fishery Bulletin, 118(4):346–358, 2020. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishery-bulletin/distribution-relative-abundance-and-reproductive-
output-spawning-female;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/1184kemberling.pdf.

Marquez-Farias:2020:LMP

[5328] J. Fernando Márquez-Farías. Length at maturity of the Pacific angel shark (Squatina californica) in the artisanal elasmobranch fishery in the Gulf of California in Mexico. Fishery Bulletin, 118(4):359–364, 2020. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fishery-bulletin/length-maturity-pacific-angel-
shark-squatina-californica-artisanal;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/1184marquez_farias.
pdf.

Zamora-Garcia:2020:CRL

[5329] Oscar G. Zamora-García, J. Fernando Márquez-Farías, Aristóteles Stavrinaky-Suárez, Carlos Díaz-Avalos, Noemí I. Zamora-García, and Raúl E. Lara-Mendoza. Catch rate, length, and sex ratio of Pacific hake (Merluccius productus) in the northern Gulf of California. Fishery Bulletin, 118(4):365–379, 2020. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishery-bulletin/catch-rate-length-and-sex-ratio-pacific-hake-
merluccius-productus-northern;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/1184zamora_garcia.
pdf.

Siddeek:2020:DHC

[5330] Shareef M. Siddeek, Benjamin Daly, André E. Punt, Steve Martell, Jie Zheng, and Mark Stichert. Development of harvest control rules for hard-to-age crab stocks: the example of the golden king crab (Lithodes aequispinus) in the eastern Aleutian Islands in Alaska. Fishery Bulletin, 118(4):380–398, 2020. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fishery-bulletin/development-harvest-control-rules-
hard-age-crab-stocks-example-golden-king;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/1184siddeek.
pdf.

Kneebone:2020:SDH

[5331] Jeff Kneebone, Heather Bowlby, Joseph J. Mello, Camilla T. McCandless, Lisa J. Natanson, Brian Gervelis, Gregory B. Skomal, Nancy Kohler, and Diego Bernal. Seasonal distribution and habitat use of the common thresher shark (Alopias vulpinus) in the western North Atlantic Ocean inferred from fishery-dependent data. Fishery Bulletin, 118(4):399–412, 2020. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fishery-bulletin/seasonal-distribution-and-habitat-
use-common-thresher-shark-alopias;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/1184kneebone.
pdf.

Bohnsack:2021:UTH

[5332] James A. Bohnsack, Laura Jay W. Grove, and Joseph E. Serafy. Use of the term harvest when referring to wild stock exploitation. Fishery Bulletin, 119(1):1–2, 2021. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishery-bulletin/use-term-harvest-when-referring-wild-stock-
exploitation;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/1191bohnsack.pdf.

Castillo:2021:CAE

[5333] Cristopher G. Avalos Castillo and Omar Santana Morales. Characterization of the artisanal elasmobranch fisheries off the Pacific coast of Guatemala. Fishery Bulletin, 119(1):3–9, 2021. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/fishery-bulletin/characterization-artisanal-
elasmobranch-fisheries-pacific-coast-guatemala;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/1191avalos_
castillo.pdf.

Carrillo-Colin:2021:BEA

[5334] Luis D. Carrillo-Colín, J. Fernando Márquez-Farías, Raúl E. Lara-Mendoza, and Oscar G. Zamora-García. Bayesian estimation of the age and growth of the golden cownose ray (Rhinoptera steindachneri) in the southern Gulf of California in Mexico. Fishery Bulletin, 119(1): 10–20, 2021. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/bayesian-estimation-age-and-growth-golden-cownose-
ray-rhinoptera;   https://spo.nmfs.noaa.gov/sites/default/files/
pdf-content/fish-bull/1191carrillo_colin.pdf.

Schueller:2021:CAD

[5335] Amy M. Schueller, Amanda Rezek, Raymond M. Mroch III, Eric Fitzpatrick, and Alicia Cheripka. Comparison of ages determined by using an eberbach projector and a microscope to read scales from Atlantic menhaden (Brevoortia tyrannus) and Gulf menhaden (B. patronus). Fishery Bulletin, 119(1):21–32, 2021. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishery-bulletin/comparison-ages-determined-using-eberbach-
projector-and-microscope-read;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/1191schueller.pdf.

Gordeev:2021:FDD

[5336] Ilya Gordeev, Kristina Zhukova, and Svetlana Frenkel. The first data on the diet and reproduction of Fedorov’s lumpsucker (Eumicrotremus fedorovi). Fishery Bulletin, 119(1):33–40, 2021. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/fishery-bulletin/first-data-diet-and-
reproduction-fedorovs-lumpsucker-eumicrotremus-fedorovi;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
1191gordeev.pdf.

James:2021:VUV

[5337] Kelsey C. James, Lisa J. Natanson, Christopher Flight, Cindy Tribuzio, John Hoey, and Camilla McCandless. Validation of the use of vertebrae and dorsal-fin spines for age determination of spiny dogfish (Squalus acanthias) in the western North Atlantic Ocean. Fishery Bulletin, 119(1):41–49, 2021. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/fishery-bulletin/validation-use-vertebrae-
and-dorsal-fin-spines-age-determination-spiny;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/1191james.
pdf.

Garcia-Rodriguez:2021:OSF

[5338] Encarnación García-Rodríguez, Miguel Vivas, José M. Bellido, Antonio Esteban, and María Ángeles Torres. Ontogenetic shifts and feeding strategies of 7 key species of Gadiformes in the western Mediterranean Sea. Fishery Bulletin, 119(1):50–65, 2021. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/fishery-bulletin/ontogenetic-shifts-and-feeding-
strategies-7-key-species-gadiformes-western;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/1191garcia_
rodriguez.pdf.

Meseck:2021:EOA

[5339] Shannon L. Meseck, Renee Mercaldo-Allen, Paul Clark, Catherine Kuropat, Dylan Redman, David Veilleux, and Lisa Milke. Effects of ocean acidification on larval Atlantic surfclam (Spisula solidissima) from Long Island Sound in Connecticut. Fishery Bulletin, 119(1):66–76, 2021. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/fishery-bulletin/effects-
ocean-acidification-larval-atlantic-surfclam-spisula-solidissima;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/1191meseck.pdf.

Erickson:2021:ESS

[5340] Kenneth A. Erickson and Marc O. Nadon. An extension of the stepwise stochastic simulation approach for estimating distributions of missing life history parameter values for sharks, groupers, and other taxa. Fishery Bulletin, 119(1):77–92, 2021. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fishery-bulletin/extension-stepwise-stochastic-
simulation-approach-estimating-distributions;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/1191erickson.
pdf.

Kusaka:2021:GPS

[5341] Takahiro Kusaka, Kojiro Hara, Keisuke Furumitsu, Shinji Uehara, Yuta Yagi, Atsuko Yamaguchi, and Naoki Yagishita. Genetic population structure of the polkadot skate (Dipturus chinensis) around Japan, based on mitochondrial DNA sequences and polymorphic microsatellite loci. Fishery Bulletin, 119(2–3):97–111, 2021. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/genetic-population-
structure-polkadot-skate-dipturus-chinensis-around-japan-based;

 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/11923kusaka.pdf.

Bolshakova:2021:MLL

[5342] Yana Yu. Bolshakova, Sergei A. Evseenko, and Dmitry V. Bolshakov. Morphology of larvae of 6 Lampanyctus species in the central South Pacific Ocean with notes on their occurrence. Fishery Bulletin, 119(2–3): 112–122, 2021. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/morphology-larvae-6-lampanyctus-species-central-south-
pacific-ocean-notes;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/11923bolshakova.pdf.

Spanik:2021:UDB

[5343] Kevin R. Spanik, Tracey I. Smart, Marcel J. M. Reichert, and Tanya L. Darden. Using DNA barcoding to improve taxonomic resolution of the diet of red snapper (Lutjanus campechanus) along the Atlantic coast of the southeastern United States. Fishery Bulletin, 119(2–3):123–134, 2021. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/using-dna-barcoding-
improve-taxonomic-resolution-diet-red-snapper-lutjanus-campechanus;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/11923spanik.pdf.

Dennis:2021:UOC

[5344] Joshua D. Dennis, Gretchen Grammer, Tim Ward, Jonathan Smart, and Charlie Huveneers. Using otolith chronologies to identify extrinsic drivers of growth of 2 commercially targeted small pelagic fish species. Fishery Bulletin, 119(2–3):135–148, 2021. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/fishery-bulletin/using-otolith-
chronologies-identify-extrinsic-drivers-growth-2-commercially;
 https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/
fish-bull/11923dennis.pdf.

Li:2021:CPA

[5345] Bai Li, Kyle W. Shertzer, Patrick D. Lynch, James N. Ianelli, Christopher M. Legault, Erik H. Williams, Richard D. Methot, Jr., Elizabeth N. Brooks, Jonathan J. Deroba, Aaron M. Berger, Skyler R. Sagarese, Jon K. T. Brodziak, Ian G. Taylor, Melissa A. Karp, Chantel R. Wetzel, and Matthew Supernaw. A comparison of 4 primary age-structured stock assessment models used in the United States. Fishery Bulletin, 119(2–3):149–167, 2021. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
fishery-bulletin/comparison-4-primary-age-structured-stock-
assessment-models-used-united;   https://spo.nmfs.noaa.gov/
sites/default/files/pdf-content/fish-bull/11923li.pdf.

Coulson:2021:BCP

[5346] Peter G. Coulson and Jake A. Poad. Biological characteristics of the primitive flatfish Indian halibut (Psettodes erumei) from the tropical northeastern Indian Ocean, including implications of the use of incorrect aging methods on mortality estimates. Fishery Bulletin, 119(2–3): 168–183, 2021. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/biological-characteristics-primitive-flatfish-indian-
halibut-psettodes;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/11923coulson.pdf.

Stienessen:2021:CMT

[5347] Sarah C. Stienessen, Christopher N. Rooper, Thomas C. Weber, Darin T. Jones, Jodi L. Pirtle, and Christopher D. Wilson. Comparison of model types for prediction of seafloor trawlability in the Gulf of Alaska by using multibeam sonar data. Fishery Bulletin, 119(2–3):184–196, 2021. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fishery-bulletin/comparison-model-types-prediction-
seafloor-trawlability-gulf-alaska-using;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/11923stienessen.
pdf.

Berkman:2021:EEL

[5348] Stephanie A. Berkman, Trent M. Sutton, Franz J. Mueter, and Brian W. Elliott. Effects of early-life stage and environmental factors on the freshwater and marine survival of Chinook salmon (Oncorhynchus tshawytscha) in rivers of Southeast Alaska. Fishery Bulletin, 119(4):201–215, 2021. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/effects-early-life-stage-and-environmental-factors-
freshwater-and-marine-survival-chinook;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/1194_berkman.
pdf.

Campbell:2021:AAA

[5349] Matthew D. Campbell, Ariane Huddleston, David Somerton, M. Elizabeth Clarke, Waldo Wakefield, Steve Murawski, Chris Taylor, Hanumant Singh, Yogesh Girdhar, and Mary Yoklavich. Assessment of attraction and avoidance behaviors of fish in response to the proximity of transiting underwater vehicles. Fishery Bulletin, 119(4):216– 230, 2021. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/assessment-attraction-and-avoidance-behaviors-fish-
response-proximity;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/1194_campbell.pdf.

McElroy:2021:CBL

[5350] W. David McElroy, Jessica Blaylock, Gary R. Shepherd, Christopher M. Legault, Paul C. Nitschke, and Katherine A. Sosebee. Comparison of a bottom longline survey and a bottom trawl survey for 2 groundfish species in the Gulf of Maine to evaluate habitatrelated availability of large fish. Fishery Bulletin, 119(4):231–242, 2021. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/comparison-bottom-longline-survey-and-bottom-trawl-
survey-2-groundfish;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/1194_mcelroy.pdf.

Fox:2021:RSJ

[5351] Adam G. Fox, Nathaniel Q. Hancock, J. Andrew Marbury, Adam J. Kaeser, and Douglas L. Peterson. Recruitment and survival of juvenile Gulf sturgeon (Acipenser oxyrinchus desotoi) in the Apalachicola River in Florida. Fishery Bulletin, 119(4):243–254, 2021. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/fishery-bulletin/recruitment-and-
survival-juvenile-gulf-sturgeon-acipenser-oxyrinchus;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
1194_fox.pdf.

Bond:2021:PML

[5352] Emily P. Bond and Michael C. James. Postrelease movements of leatherback turtles (Dermochelys coriacea) following incidental capture in fishing gear in the Atlantic Ocean. Fishery Bulletin, 119(4):255–260, 2021. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fishery-bulletin/postrelease-movements-leatherback-
turtles-dermochelys-coriacea-following;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/1194_bond.
pdf.

Koob:2021:AVN

[5353] Elise R. Koob, Scott P. Elzey, John W. Mandelman, and Michael P. Armstrong. Age validation of the northern stock of black sea bass (Centropristis striata) in the Atlantic Ocean. Fishery Bulletin, 119(4):261–273, 2021. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.

gov/content/fishery-bulletin/age-validation-northern-stock-
black-sea-bass-centropristis-striata-atlantic;   https://spo.
nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/1194_
koob.pdf.

Poussard:2021:DBH

[5354] Leanne M. Poussard, Eric N. Powell, and Daniel R. Hennen. Discriminating between high- and low-quality field depletion experiments through simulation analysis. Fishery Bulletin, 119(4):274–293, 2021. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/fishery-bulletin/discriminating-
between-high-and-low-quality-field-depletion-experiments;
https://spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-
bull/1194_poussard.pdf.

Ferm:2022:FFH

[5355] Nissa C. Ferm, Janet Duffy-Anderson, and Thomas P. Hurst. Functional foraging habits and dietary overlap of yellowfin sole (Limanda aspera) and northern rock sole (Lepidopsetta polyxystra) in a coastal nursery of the Bering Sea. Fishery Bulletin, 120(1):1–12, 2022. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.

nmfs.noaa.gov/content/fishery-bulletin/functional-foraging-
habits-and-dietary-overlap-yellowfin-sole-limanda;   https:
//spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
1201_ferm.pdf.

Sutherland:2022:VMA

[5356] Sandra J. Sutherland and R. Anne Richards. Validation of methods for aging goosefish (Lophius americanus) based on lengthmode progression of a strong cohort. Fishery Bulletin, 120(1):13– 25, 2022. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/validation-methods-aging-goosefish-lophius-americanus-
based-length-mode;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/1201_sutherland.pdf.

Viducic:2022:RCB

[5357] Katie Viducic, Lisa J. Natanson, Megan V. Winton, and Austin Humphries. Reproductive characteristics for the blue shark (Prionace glauca) in the North Atlantic Ocean. Fishery Bulletin, 120(1):26– 38, 2022. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/reproductive-characteristics-blue-shark-prionace-glauca-
north-atlantic.

Scordino:2022:DNO

[5358] Jonathan J. Scordino, Adrianne M. Akmajian, and Stacy L. Edmondson. Dietary niche overlap and prey consumption for the Steller sea lion (Eumetopias jubatus) and California sea lion (Zalophus californianus) in northwest Washington during 2010–2013. Fishery Bulletin, 120(1): 39–54, 2022. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/fishery-
bulletin/dietary-niche-overlap-and-prey-consumption-steller-
sea-lion-eumetopias;   https://spo.nmfs.noaa.gov/sites/default/
files/pdf-content/fish-bull/120_scordino.pdf.

Siwicke:2022:CSB

[5359] Kevin A. Siwicke, Andrew C. Seitz, Cara J. Rodgveller, and Katy B. Echave. Characterizing spawning behavior of Greenland halibut (Reinhardtius hippoglossoides) in the eastern Bering Sea and Aleutian Islands. Fishery Bulletin, 120(1):55–67, 2022. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:/

/spo.nmfs.noaa.gov/content/fishery-bulletin/characterizing-
spawning-behavior-greenland-halibut-reinhardtius;   https:/
/spo.nmfs.noaa.gov/sites/default/files/pdf-content/fish-bull/
120_siwicke.pdf.

James:2022:OHU

[5360] Rebekah C. James, Tobey H. Curtis, Benjamin Galuardi, Gregory Metzger, Alisa Newton, Michael P. McCallister, G. Christopher Fischer, and Matthew J. Ajemian. Overwinter habitat use of young-of-the-year white sharks (Carcharodon carcharias) off the eastern United States. Fishery Bulletin, 120(1):68–73, 2022. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/fishery-bulletin/overwinter-habitat-use-young-year-
white-sharks-carcharodon-carcharias;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-content/fish-bull/120_james.pdf.

Wilborn:2022:FOU

[5361] Rachel E. Wilborn, Ingrid Spies, Pamela Goddard, Christopher N. Rooper, and James W. Orr. First observation of the use of coral habitat by larval northern rockfish (Sebastes polyspinis) in the western Gulf of Alaska. Fishery Bulletin, 120(1):74–78, 2022. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.

noaa.gov/content/fishery-bulletin/first-observation-use-coral-
habitat-larval-northern-rockfish-sebastes;   https://spo.nmfs.
noaa.gov/sites/default/files/pdf-content/fish-bull/120_wilborn.
pdf.

Pauly:2022:GRT

[5362] Daniel Pauly, Nicola Smith, and Mark Butler. Growth and related traits of the sheepswool sponge (Hippospongia lachne): practical and theoretical considerations. Fishery Bulletin, 120(2):??, ???? 2022. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/growth-and-related-
traits-sheepswool-sponge-hippospongia-lachne-practical-and-
theoretical.

Evseenko:2022:NDM

[5363] Sergei A. Evseenko, Dmitry V. Bolshakov, Yana Yu. Bolshakova, Yuri O. Kopylov-Guskov, Natalya V. Gordeeva, Evgeniia P. Karpova, and Alexander R. Boltachev. New data on the morphology of the Mekong blind sole (Typhlachirus elongatus) indicating the need for a revision of the genus. Fishery Bulletin, 120(2):??, ???? 2022. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/new-data-morphology-mekong-blind-
sole-typhlachirus-elongatus-indicating-need-revision-genus.

Rabaoui:2022:ELH

[5364] Lotfi Rabaoui, Lamia Yacoubi, Yu-Jia Lin, Thadickal V. Joydas, Mohammad A. Qurban, Panickan Premlal, Jinoy Gopalan, Zahid Nazeer, Manambrakat Vijayakumaran, Syed Ajmal Khan, and Rubén H. Roa-Ureta. Ecology, life history, and fisheries potential of the flathead lobster (Thenus orientalis) in the Arabian Gulf. Fishery Bulletin, 120(2):??, ???? 2022. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/ecology-life-history-
and-fisheries-potential-flathead-lobster-thenus-orientalis-
arabian.

Cooper:2022:SCP

[5365] Ross Cooper, Heidi Dewar, Barbara A. Muhling, Steven L. H. Teo, John Hyde, and Keith Bigelow. Spatiotemporal catch patterns and population distributions of bigeye Pacific opah (Lampris megalopsis) and smalleye Pacific opah (L. incognitus) in the eastern North Pacific Ocean. Fishery Bulletin, 120(2):??, ???? 2022. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/spatiotemporal-catch-
patterns-and-population-distributions-bigeye-pacific-opah-lampris.

Scordino:2022:CPS

[5366] Jonathan J. Scordino, Cydni Marshall, Adrianne M. Akmajian, Daniel Shay, and Randall James. Consumption of Pacific salmon (Oncorhynchus spp.) by California sea lions (Zalophus californianus) and Steller sea lions (Eumetopias jubatus) in northwest Washington during 2010–2013. Fishery Bulletin, 120(2):??, ???? 2022. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/consumption-pacific-
salmon-oncorhynchus-spp-california-sea-lions-zalophus-californianus-
and.

Hightower:2022:PDR

[5367] Crystal L. Hightower, J. Marcus Drymon, Amanda E. Jefferson, Matthew B. Jargowsky, Emily A. Seubert, Simon Dedman, John F. Mareska, and Sean P. Powers. Population dynamics, relative abundance, and habitat suitability for adult red drum (Sciaenops ocellatus) in nearshore waters of the north-central Gulf of Mexico. Fishery Bulletin, 120(2):??, ???? 2022. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/population-dynamics-relative-abundance-and-habitat-
suitability-adult-red-drum-sciaenops.

McKinzie:2022:MEG

[5368] Megan K. McKinzie and Stephen T. Szedlmayer. Mortality estimates for gray triggerfish (Balistes capriscus) in the northern Gulf of Mexico based on acoustic telemetry. Fishery Bulletin, 120 (2):176–186, ???? 2022. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/

content/population-dynamics-relative-abundance-and-habitat-
suitability-adult-red-drum-sciaenops.

Zokan:2022:LHS

[5369] Marcus A. Zokan, David M. Wyanski, D. Byron White, and George R. Sedberry. Life history of the spotted moray (Gymnothorax moringa) off the Atlantic coast of the southeastern United States: age, growth, and reproduction. Fishery Bulletin, 120(3–4):??, ???? 2022. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/life-history-spotted-
moray-gymnothorax-moringa-atlantic-coast-southeastern-united-
states.

Marquez-Farias:2022:BEA

[5370] J. Fernando Márquez-Farías, L. Daniel Carrillo-Colín, Allan Rosales-Valencia, Raúl E. Lara-Mendoza, and Oscar G. Zamora-García. Bayesian estimation of the age and growth of the round stingray (Urobatis halleri) in the Gulf of California in Mexico. Fishery Bulletin, 120 (3–4):??, ???? 2022. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
bayesian-estimation-age-and-growth-round-stingray-urobatis-
halleri-gulf-california-mexico.

Hightower:2022:EFC

[5371] Crystal L. Hightower, Kelly S. Boyle, and Sean P. Powers. Evaluation of factors contributing to postrelease mortality of greater amberjack (Seriola dumerili) in the northern Gulf of Mexico with depth and acceleration data from acoustic tags. Fishery Bulletin, 120(3–4):??, ???? 2022. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.gov/content/evaluation-factors-
contributing-postrelease-mortality-greater-amberjack-seriola-
dumerili.

Chin:2022:LHD

[5372] Andrew T. Chin, Morgan H. Bond, and Thomas P. Quinn. Life history differences between 2 anadromous populations of the northern form of the Dolly Varden (Salvelinus malma malma) in Bristol Bay in southwestern Alaska. Fishery Bulletin, 120(3–4):??, ???? 2022. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/life-history-differences-between-
2-anadromous-populations-northern-form-dolly-varden.

Krusic-Golub:2023:EOI

[5373] Kyne Krusic-Golub and Lisa Ailloud. Evaluating otolith increment deposition rates in bigeye tuna (Thunnus obesus) and yellowfin tuna (T. albacares) tagged in the Atlantic Ocean. Fishery Bulletin, 121(1–2):??, ???? 2023. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (elec-

tronic). URL https://spo.nmfs.noaa.gov/content/evaluating-
otolith-increment-deposition-rates-bigeye-tuna-thunnus-obesus-
and-yellowfin-tuna.

Nguyen:2023:JYA

[5374] Long V. Nguyen, Dat X. Mai, Quang M. Thai, and Tuan S. Vo. Juvenile yield and adult abundance, genetic diversity and structure, and linkages among marine habitats for goldlined spinefoot (Siganus guttatus) in the coastal waters of Vietnam. Fishery Bulletin, 121(1–2):??, ???? 2023. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL

https://spo.nmfs.noaa.gov/content/juvenile-yield-and-adult-
abundance-genetic-diversity-and-structure-and-linkages-among.

Buzzell:2023:EBR

[5375] Bobbie M. Buzzell, Adrianne M. Akmajian, and Alejandro Acevedo-Gutiérrez. Exploring the biotic resistance of the invasive green crab (Carcinus maenas) by examining the diet of North American river otters (Lontra canadensis). Fishery Bulletin, 121(1–2):??, ???? 2023. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518

(electronic). URL https://spo.nmfs.noaa.gov/content/exploring-
biotic-resistance-invasive-green-crab-carcinus-maenas-examining-
diet-north.

Cyrana:2023:OMR

[5376] Michael V. Cyrana, Alexander J. Ledet, and Henry L. Bart, Jr. Ovarian masculinization and reproductive impairment in 3 species of groundfish in and around the hypoxic zone in the Gulf of Mexico. Fishery Bulletin, 121(1–2):??, ???? 2023. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
ovarian-masculinization-and-reproductive-impairment-3-species-
groundfish-and-around-hypoxic.

Kneebone:2024:HSS

[5377] Jeff Kneebone, Connor F. White, Caroline Collatos, and Nicholas M. Whitney. High survivorship of sandbar sharks (Carcharhinus plumbeus) following catch and release in a growing land-based fishery in Massachusetts. Fishery Bulletin, 121(1–2):??, ???? 2024. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/high-survivorship-sandbar-sharks-
carcharhinus-plumbeus-following-catch-and-release-growing.

Lemos:2024:VOM

[5378] Valéria M. Lemos, Ana C. G. Mai, Ana C. Martins, José V. Martins, Sandra Andrade, Henrique Cabral, and João P. Vieira. Variation in otolith microchemistry for Lebranche mullet (Mugil liza) in southern Brazil and its potential use in identifying their nursery grounds. Fishery Bulletin, 121(1–2):??, ???? 2024. CODEN FSY-BAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https://

spo.nmfs.noaa.gov/content/variation-otolith-microchemistry-
lebranche-mullet-mugil-liza-southern-brazil-and-its.

Damiano:2024:SDD

[5379] Matthew D. Damiano, Mandy Karnauskas, Wessley Merten, and Jie Cao. Spatiotemporal dynamics of dolphinfish (Coryphaena hippurus) in the western Atlantic Ocean. Fishery Bulletin, 121(1– 2):??, ???? 2024. CODEN FSYBAY. ISSN 0090-0656 (print),

1937-4518 (electronic). URL https://spo.nmfs.noaa.gov/content/
spatiotemporal-dynamics-dolphinfish-coryphaena-hippurus-western-
atlantic-ocean.